A Complete Bibliography of Publications in the Journal of Theoretical Biology: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

18 March 2025 Version 1.12

Title word cross-reference

(ΔΨ) [LL17b]. (M, R) [PWG16]. −1 [Xie13a]. −2 [ZL18b]. 1 [GMMN18, STA15]. 2 [CLPZ14, CK17, DHK13, HBSF11, JO13, MS14d, Nos14, Shu13, TA16, WML^{+}17]. 3 [BPF^{+}19, CZJQ14, CK17, DS15b, DTP19, EBP15, ESGA15, GBC^{+}16, HGGCR13, KKGN16, KMM18, LHPF18, MPCTGJ^{+}15, Nos14, PHA^{+}17, RMSTG13, SK15, SKK18, TA16, ZKP15, ZSH^{+}16]. 4 [EBSW17, LZL^{+}19]. 5 [HKR^{+}19, NMZ19, SIK^{+}18]. ^{+} [AHW13, CFF12, MYC12, OCN10]. ^{1}5 [LDB^{+}14]. ^{2+} [CFF12, CN12a, DD13, EG10, Gri11, KCZ^{+}19, PRV^{+}14]. ^{6} [AH18, ZZN^{+}19]. _{2} [LDB^{+}14]. _{3} [CZT^{+}16, LL17b]. a [ASRM15]. α [BZJP18, EKvdKvFK13, HSiI^{+}19, KSK^{+}17, RAR19, SSZR17, WPH^{+}12, YCB16, dSMP^{+}11]. b [ASRM15]. β [BZJP18, CWP^{+}18, Goe15, Gün12, Gün13, KSKK15, KGM15b, MVGGB18, MB18b, MHMM11, SNS17, SWR11, SHLL11, SKK18, SAGC12, WGC13, ZZRZ11]. c [PBSM19]. C^{3} [Mic15b]. R_{0} [TDW16]. Δ [LL17b]. γ [OBK^{+}11, TRM^{+}14]. H_{∞} [QF10]. K [NMZ19, DLYW13, GAB14, HA15a, Hua16, LCL14, Tun13, WZY14, YZZ13, YW13a, YW13b]. κ [ÁTC^{+}14, GM13, HSiI^{+}19, KB19, TC11, WPH^{+}12, ZBA14]. N [SPSP12, FS15b, HY16, LZT12, LZT13, TP10a, vV11a]. N_{e} [ZGW16]. P_{2} [PIPB10]. R_{0} [CD16, DDSDW13]. σ [TLCZ12]. × [Sum13].

-2 [BMGC11a, BMGC11b]. -activated [CFF12]. -amyloid-neuron [WGC13]. -associated [HBSF11]. -based [JO13, ANMH11, Mil11]. -catenin [SWR11]. -catenin/TCF [SWR11]. -cells [MVGGB18, MB18b]. -circular [FS15b]. -clavaminic [RMRC^{+}16]. -clavulanic [RMRC^{+}16]. -dihydromethyltrisporate [EBSW17]. -dimensional [EBP15]. -grams [HY16]. -helical [dSMP^{+}11]. -Ig [LZL^{+}19]. -infected [STA15]. -interacting [HBSF11]. -lactamase [KSKK15, SKK18]. -LTR [CLPZ14, WML^{+}17]. -mer [HA15a, Hua16, WZY14]. -methyladenosine [AH18, ZZN^{+}19]. -methylcytosine [SIK^{+}18]. -nearest [NMZ19]. -Partite [LCL14]. -person [TP10a, SPSP12]. -phenylpyrimidine [TA16]. -player [DHK13, Shu13]. -protein [SAGC12]. -PseAAC [NMZ19]. -rays [TRM^{+}14]. -regular [CZJQ14]. -sheet [ZZRZ11]. -spaced [Tun13]. -state [MS14d]. -step [NMZ19]. -steps [HKR^{+}19]. -strand [SNS17, ZZRZ11]. -susceptible [TDKJ15]. -tier [SKK18]. -tubulin [KGM15b, SSZR17]. -turn [SHLL11]. -variable [DTP19]. -word [DLYW13, YW13a, YW13b]. -words [GAB14].

/chondrotin [SABB15]. /Cl [CZT^{+}16]. /K [OCN10].

09 [SY11].

1 [CH16, FH10a, FHG15, GKPB13, IGL^{+}12, ITR^{+}18, LLW15, LGK^{+}09, MR19, NTC^{+}11, SW11, SPG^{+}18, SH15, TXCW15, VBHM^{+}13]. 1-108 [Ano19-48]. 1-72 [Ano19-49]. 10 [MMFK10]. 118 [LHPF18]. 16S [WSTL16, WZ17]. 185 [BBM^{+}13a, BBM^{+}13b]. 1c [BAGG14].

2 [TTC19, BMGC11b, FH10b, GML10, Lab16]. 2’-O-methylation [TTC19]. 2009 [KLJ17, LJ15, LCJ18, TDE12]. 2013 [DHK13]. 2019 [Ano19b, Ano19a, Ano19-43, Ano19-38, Ano19-30, Ano19-34]. 21 [Ano19a, Ano19-30, Ano19-34]. 236 [SRAL12]. 256 [Lei10]. 257 [EPJ^{+}11]. 259 [ML12a]. 260 [DBBW11]. 261 [LGK^{+}12]. 262 [ZZR13]. 263 [CLG^{+}11]. 264 [dOGL13, SCÁBM11, ZG10a]. 266 [LBS^{+}11, PSS^{+}13]. 271 [FIS16a]. 277 [BMGC11a]. 278 [PCT19, Pan19]. 280 [Mul12]. 283 [CNG^{+}12]. 289 [JCLS^{+}21]. 293 [Mei13a]. 294 [LPvSP12, RTEKG19a]. 2A [DS15b]. 2b [VZB^{+}15]. 2D [NWB^{+}10, WXC10]. 2D-MH [WXC10]. 2L [JCG15]. 2LSAAC [AHJ18b]. 2methyl [TTC19].

3 [Sel12a, BZJP18]. 3’-to-5 [Sel12a]. 3’-to-5’-direction [Sel12a]. 307 [BPGS12a, Gün13, LZT13]. 310 [Mei13b, WL12a]. 311 [MLBA13]. 317 [TSM13a]. 321 [DHK13]. 329 [AEM14]. 330 [SC13a]. 335 [LHPF18]. 339 [GS14]. 340 [KDK14a]. 342 [PL14a]. 343 [KMLT14]. 344 [JRG14a, JRG14b]. 355 [JN14, NO14a]. 361 [GZ14b]. 365 [BPF15b, Van16b]. 372 [FS16a]. 380 [RTEKG19b]. 382 [SZ16]. 388 [ZSCL16a]. 3D [HB12a, aHLZ^{+}12, MDD13, PWC^{+}12, SGB12, SKH11, XM11]. 3D-geometric [SGB12]. 3R [RMRC^{+}16]. 3S [RMRC^{+}16]. 3ST [RLFS18].

413 [YCH^{+}17a]. 417 [Sch23]. 419 [JS18, SYY17a]. 426 [LPB18]. 431 [EJ19]. 437 [PP19]. 439 [SSKS20]. 447 [Lio24]. 454 [Sch21]. 456 [GHS20]. 464 [Mac11]. 483 [WLW20].

5 [Sel12a, SGD^{+}16]. 5’-UTRs [SGD^{+}16]. 5R [RMRC^{+}16]. 5S [RMRC^{+}16].

6-bisphosphate [Bar19]. 6DoF [GAL11b].

7 [Ano19b, Ano19-43, Ano19-38, OOY16]. 756 [Mac11].

Îto [OOY16].

A*0201 [BYY^{+}11]. A-peptide [WHL13]. A. [MHX^{+}14]. A/H1N1 [KLJ17, KJ19, LJ15, TXY^{+}12]. aberrant [PRN10]. Ability [HA15b, HN10, LDdA^{+}13, PDB^{+}15, PGH11, RT19, SMHB10]. Abiogenesis [Ser16, Tox18]. abiotic [LVLC18]. abnormal [CPF^{+}13]. Aboelkassem [SWPC^{+}16]. Abrupt [TBM^{+}13]. Absence [BWP10, BGLR^{+}19, DSLR18, HZG^{+}17]. absence-deriving [BGLR^{+}19]. absent [YZZ12, YZFY16]. absorption [DS18, Gal11a]. abstaining [MBC^{+}12a]. abundance [AF10, FVC15, GT11b, GBIR13, KFG15, KS14, KMM16, MG14b, SS15a, Su16, TKB^{+}18, TWW19, YSST13, YMZ18]. abundances [PC13]. abuse [CFS^{+}19]. abysses [Di 13a]. accelerate [Kor18]. Accelerated [CzST^{+}18, DZJR10]. accelerated-predator-satiety [DZJR10]. Accelerating [LK19]. accelerometric [WHHS15]. accessible [IMH15, TAI^{+}18]. accessory [SH15]. accompanied [iITIM15]. according [CC10, PVGV19, RMSTG13, SS15e, ZEJA11]. account [FD18, MP12, MA10, Ush16]. Accounting [Oko15, ZKTK10, LAS14, MKBE17a, MKBE17b, TSB10]. accounts [Wei12]. accumulate [HHMT19, KDK14a, KDK14b]. Accumulation [SK12b, AK15, Ber14, HL14, JCLS^{+}11, JCLS^{+}21, MYK^{+}11, MYC12, SW11, SAB10]. accuracies [RSD^{+}16]. accuracy [CC11b, KOS13, Kon12, LJ10, MSRW16, WYX^{+}17, ZLY14]. Accurate [KZL14, ZNCM15, SMZ^{+}17, USF^{+}18, ZCAB17, dEG11]. Accurately [DLL^{+}17, BMSEE14, IGL^{+}12, SHB^{+}17]. acetaminophen [BRN15, RWW^{+}15]. Acetolactate [EJ16]. acetylation [CCY^{+}19, SQS^{+}12, XWD^{+}10]. acetylcholinesterase [MR18, WB15]. achieve [CW11a, iHIM12]. achieves [XJ19]. achieving [WB10]. Acid [AH15, AHJ18b, BRK16, BMN16, BTF19, BB13, BWM^{+}19, CEdLSG^{+}16, CC10, CL13a, Cho11, DS19b, Di 14a, DWM15, Edg19, EBSW17, EMM10, FL12, FL13, FLW16, GT11a, GZ14a, GZ14b, HPB^{+}14, HYA14, HK11a, HYW11, HY13, JD16a, JD16b, JD17, JSF^{+}11, JCG15, JCG16, KHK15, KB11, KZL14, KHH^{+}17, KSKK15, KGM15b, LLG16, LL13a, LD11, LBS^{+}14, MJ11, MGGM10b, MFZ18, MZ18, MBE11, MP14b, NBL14, OFB10, PWZ^{+}19, PDC^{+}17, PWC^{+}15, QLC^{+}18, RBMS17, RMRC^{+}16, SRS^{+}15, SFV16, TP17, Tun13, WMK13, XNJ^{+}13, Xie15, XWW^{+}14, XSLZ16, YSH^{+}14, YG18, YGL^{+}10, ZPdFJ19, ZDG^{+}10, ZSZM14, ZZK14, ZNA^{+}16, ZLW^{+}19, dBJ11, dSMP^{+}11, vLFM^{+}19]. acid-mediated [MGGM10b]. acidic [KHK15]. acidification [MGGM10a, RTFP^{+}17]. acidosis [TDSM12]. acids [CL11, Cle10, FD18, FS15b, MP13b, RSD^{+}16, Rog19, Woh15, WXC10, YGMT12]. acinar [SMZ^{+}17]. Acinetobacter [RRG^{+}10, RRG^{+}12, SRP16]. acquaintance [CW15a]. acquired [FM14a, GJ12, PDW10, PRM14, TA10, WKB13]. acquisition [SI19]. across [AMM16, AHD^{+}18, ASL^{+}18, Bac15, BHR10, DFG^{+}18b, GABM12, LZG^{+}19, LL17b, WRC^{+}19, ZJS11]. actin [FKK14, HBW^{+}11, KSK^{+}11, KHNM16, MDEKH13, MS18, MED12, SW15]. acting [GN10, HG18b]. action [ACT12, Arc14, BZJP18, BHBH11, BCPM^{+}16, CF11, DBJ12, Gri15b, KMA10, LdLK11, MPNP12, PNL15, PN16, SNOV11, TWTA^{+}18, TSF^{+}19, WLD^{+}11]. actions [DL12a]. activated [CFF12, YST14, ZCA^{+}14]. activates [CSKZ19]. Activation [KL18, ABR11, BN12, BN15, BSC^{+}18, CS11a, EJ16, KHH10, KHS13, KPS17, LO15, MH11, MYC12, Mor10a, NZZ19, RG17, RG18, RG12, SH16, TSP15, UK17]. activations [SCF^{+}12]. activator [Das13]. activators [NZZ19]. Active [AMP12, RGA^{+}10, ASS19, CN10, CFGRB17, CT16, DK13c, EM11, FGH^{+}14, GTC19, GDPPSS^{+}11, MM12, NP15, Pat16, Pel18, PGF11, RW15b, SMJS14, SPL14, TC11, TSB10]. active-search [RW15b]. activities [FHG15, MFKS13, NF14b]. activity [BB11b, BPPC15, CP14, CW11c, DS10a, GVGC15, HB18, KSSM11, KIH16, KB10b, Kro10, LL13a, Mar12, NBW10, NBW11, OA15, PSG^{+}17, PVGV19, PPF17, RRC^{+}11, RG17, ZLL^{+}12]. activity-pCa [RG17]. actomyosin [ITN^{+}11]. acts [HTN14, MLBA12, MLBA13]. actually [RWH16]. actuation [MK18b]. acute [BFJ^{+}18, CMS16, LGK^{+}09, LGK^{+}12, LH14, MBBV14, NSH^{+}10, RZRSC19, SCLA10, SBM^{+}19, SW11, ZCT18]. acyl [PWH^{+}13]. Acylguanidine [SK19]. Acylguanidine-BACE1 [SK19]. adaptability [CMCS18, SM10a]. adaptable [SHW16]. Adaptation [GCB17, AT10, BPM^{+}12, BSH19, Don13, FWR19, GB13, GTSP16, Mil16, NS16, PW18a, SMG18, SSV^{+}15, SH16, TD17, TKTB18]. adaptations [DPvBvA12]. adapted [Jam16a]. adapting [ABIM10]. Adaptive [ADC19, BS15a, BEK10, DMO^{+}17, EL12, Ezo19, FGGT15, FH13b, GS13a, GS14, HD11, iHM17, OBE^{+}17, RPCW18, SSD13, SM17b, SHF16, WFZW13, ZPHS11, ZMT11, AR11, AK11, CJ12b, DGMY18, GK17, GSN^{+}11, HE16, IEN15, IMD16, KDG13, KMMS17, KR19, LS16a, LvBJ16, MZAI19, Mor13, NO15, NSH^{+}10, PNK16, RTEKG15, RTEKG19b, RAMS11, SDT17, STN19a, TM12, YCH^{+}17a, YCH^{+}17b, YC13, ZCXF13]. adaptively [COWA11, Zhu11]. ADCC [HGO^{+}18]. added [LS18]. Addendum [SZ16]. Adding [KA15, SHW16, FLM18]. addition [RD12]. Additional [SVA18]. Additive [Gri15b, Boc12, HK14, JSDEK14]. additive-by-additive [Boc12]. Additivity [MEJ18]. Addressing [NI11]. adherence [LS11a]. adherent [CCdS15]. adhesins [DMM^{+}14]. Adhesion [CCS18, CMS^{+}19, CCB11b, CML11, DTGC14, FPG11, GMMN18, HJLNZ11, HBW^{+}11, MFG14, MFMB12, MT15, NCLB16, PAS10, UAL10]. adhesions [ESGA15]. adhesive [IUK14, Sar10]. adipocytes [SJSG13]. adipose [PDL^{+}17, SGS15]. adjacent [CFF12]. Administration [HL18, BS15a, CHN^{+}15, RSD^{+}13, TTB^{+}18]. admissibility [MKE15]. adolescents [BSR14]. adopting [AH19]. ADP [CPS10, MS18]. ADPM [ZK18]. adrenal [Jam15, WZ12]. ADSVM [WMK16]. adult [DLHS11, GMM15, GM16b, ISZ18, NTED^{+}19, NNH^{+}16, RVA^{+}11, SSG^{+}19, SB19]. adults [CL13b, Gar10]. Advance [SL12]. advanced [SS17a, SCF^{+}12, SKD^{+}10]. advances [Wil11]. Advancing [WYX^{+}17]. advantage [BID15, Bul12b, Cro17, EHSR17, ESE15, HG18b, IUK14, KDK14a, KDK14b, Lar16, SG18]. advantageous [SL12]. advantages [BTK14]. advection [RLK10, VIBC12]. advection-diffusion [RLK10]. advective [LS15b]. advertisement [BR12a, FH13a]. Aedes [LBGW13, ASF^{+}15, HLH^{+}18]. aegypti [LBGW13, ASF^{+}15, MHX^{+}14]. Aerodynamic [DS12, MHSH12, FH14, LBS^{+}10, LBS^{+}11, Ush16]. aeruginosa [ALM^{+}19, MBC^{+}12b]. affect [AR11, BH13a, CV15, DLEMP19, GSF13, HGHL10, KLN^{+}12, RL17, SM16, SLL18, SJK18]. affected [GZ19, JSP^{+}16, LS14b, TB18]. affecting [LvdBP12, POW18]. affects [ATTN12, BS19b, Cut15, GSSBF18, HT13, KBL15, KRR14, MGS16, NIiT18b, Nak19, SN14, SHW16, XDC^{+}11]. affinity [TLT^{+}15, TCYY^{+}12, WFW19]. AFM [KMM18, SK15]. AFM-based [KMM18]. AFP [KCM^{+}11]. AFP-Pred [KCM^{+}11]. Africa [DE17, MS16]. African [CMMR13, GPD18, PDW11, STLJ18]. after [BPLM12, FS11a, HWM^{+}16, HHMT19, HRG11, JV14, LNRK11, SSME10, TTB^{+}18, TALC16, VLPH17]. AGAAAAGA [ZGY11, ZSW11]. against [BC19, DMM^{+}14, DL15a, Di 10, Di 17b, Dra19, GWCA14, GDPPSS^{+}11, HS15, JSF^{+}11, KJM17, KRDJ15, LZL^{+}19, MCL19, NGJ^{+}14, PGGvL^{+}19, ST17a, SRP16, TWR^{+}18, TXTW16, TAR16, YM16, Yam16b, YY18, ZCW13, vVE15]. agarwood [LCQ^{+}18]. Age [Kär11, MDE11, CHK16, DBB14a, Dol16, GGQ^{+}12, GMR16, JS15, JS17, JS18, KJ19, KMH19, KSPA17, LAS14, LKK13, LM11, LS16b, LSLL13, MDMG14, MKPVH16, OT13, Tak16, Yan10a, ZLT^{+}19]. age-dependent [KJ19, LAS14, LKK13, MKPVH16]. age-size [OT13]. Age-specific [MDE11, MDMG14]. age-stage [ZLT^{+}19]. age-structured [GMR16, KSPA17, LM11, LS16b, Tak16]. ageing [BM16, EdKM11, Gia13, KDK14a, KDK14b, PMD^{+}18, ZARK19]. Agent [CSR^{+}05, Gal10, GCB17, SRAL12, BRP^{+}18, BCS^{+}16, CS14a, CA17, Cro18, DJD10, FM15a, HWPL12, HM13, JSSZ12, KTJ19, LS16c, METC12, SGDL12, TCB13, TVA18, WFC^{+}14]. Agent-based [CSR^{+}05, Gal10, SRAL12, BCS^{+}16, CS14a, CA17, Cro18, DJD10, FM15a, HM13, JSSZ12, METC12, TCB13]. agent-environment [BRP^{+}18]. agents [ACCR11, GN10, HAuR^{+}18, IHNS16, JH11, LFZN11, TTB^{+}18, VW10, Wil13]. aggregate [MPLK14]. aggregates [GWW^{+}10, PAV10, YMW12]. Aggregation [BHP17, AGB^{+}14, DFMR19, GABM12, IOY15, KKS^{+}14, MCPF12, NPB13, RDD14, SM14, SS15a]. aggregations [PWE15, Tos11, Woo10]. aging [HH10, Hor13, Kit10, KS15, LC16, PS13, PVGV19, SC11b, Wu14]. ago [Kur14]. agonism [BMF^{+}18, HPML18]. agonist [HBSF11, PHC12]. agonists [DvDBD15, HP12]. AhR [ACM16]. AhR-gene [ACM16]. aid [SN18]. aided [SB15, WDH^{+}16]. AIDS [DCR18, HTK12, JO13, NM11, PT10, XTZ^{+}13, Yan10a]. Air [CXT19, IMW15, JG14, MAM16]. airborne [IMW15, RSD12, UD10]. airfoil [LS10b]. airway [AJD11, AJC12, AdSG14, Don13, Don17, EJK16, HGB^{+}13, PDT^{+}10, WTC10, WBB^{+}18]. airways [SCS15]. Akt [GC16, LCGMH12]. Akt/mTOR [LCGMH12]. al [JN14, JAHKH12, Mac11, NBW11]. al. [Di 14b, SWPC^{+}16]. alarms [SPG^{+}18]. albicans [HHS^{+}10]. albopictus [MHX^{+}14]. albumin [SCS10]. alcohols [KK15]. aldolase [Bar19]. aleurone [OFB10]. algae [JB18a, PKL12]. algal [AK15, BH11, STI13]. algebra [MFG^{+}13, MPT19]. Algebraic [NMS10, MPT19]. algorithm [AHAFA14, CPH^{+}14, DXW^{+}16, DV12, EHSR17, FCS18, Ish16, JW18, KKGN16, KF15, LT14, LJ10, LDWW14, LHH12, wLW10, MGM13b, NMZ19, PDNP16, PSAA13, QF10, SNS17, SM10a, SKH17, SMC17, WXmH10, WY13, WcW14, XXD^{+}17, YLC16, ZMNd^{+}10]. Algorithmic [BNG16]. Algorithmically [SWG^{+}15]. algorithms [Mor19, NNK^{+}15]. Alignment [HZLX11, YZZ12, YZZ13, GAB14, HLHY17, LWL^{+}11, LBS^{+}14, LPD^{+}16, MSND12, MPL16, MED12, PM11b, QQL10, SNS17, SVSP15, TP17, WDH^{+}16, ZZRZ11]. alignment-based [MSND12]. Alignment-free [HZLX11, GAB14, HLHY17, LWL^{+}11, TP17]. alkaline [KHS13, KHK15]. All-atom [Mas18]. all-or-nothing [DS10a]. Allee [AH17, BGM19, BBB18, BIMC17, HMM17, MG14b, MBP16, PBB10, PR13, SHvHB16, ZMW10, dSRM15]. allele [AM11, CJKR10, MKF^{+}14, SGW17, SL12, WL12a, WL12b]. alleles [ASRM15, GGM12, PKH11]. Allelic [KM18, MKE15]. allelochemicals [LFM11]. allelopathy [GW14]. allergen [HI17]. allergy [GMB11]. alleviate [CW14]. alleviation [KK18]. alliance [MGGGA11]. Allocation [TP14, BHBR12, DBB14a, DB19, GCZ^{+}12, KMHdlP10, MDMG14, MLT10, MK14c, MKJS13, MRS14, SCH^{+}19a, WKH16, YYST13]. Allometric [All11, BZN17, PRSC11, BS13, BDE^{+}14, Gla13, TW13]. allometries [GLZ^{+}13, Nij11]. Allometry [CW11a, SI10b, Bal13, MR10]. allostasis [RGBR17]. allosteric [PBBB10]. allostery [NGND12]. allow [Moo14, tBdR18]. allows [JP13]. almond [YLH12]. alone [GT12, Sch19b]. along [Bro13, CN10, Che14, CL10, PMD11, SSZR17, SPL14]. alopecia [DPC15]. Alpha [KA16, EG12, RGP13, VZB^{+}15]. alpha-2b [VZB^{+}15]. alphabet [HWHL15, dBJ11]. already [Di 14a]. also [GvVT14, SKS^{+}19]. alter [CK19, GWCA14, GBM18, MR15, SLD^{+}17]. alterations [DM11, KK18, SPMGR10]. altered [CML11, MPS16]. alternans [ESG^{+}16, MKMT14, VLCT19, WMT15]. alternate [CMGN17]. alternating [SM11b, ZLY14]. Alternative [MH18, NWWL18, QYO10, VA10, BWM^{+}19, GBRS19, HSR12, TCW14]. alters [Oku15, PW18b, WCPF15]. altruism [Gar10, KF11, LDF11b, SN16, SC12, YvBS18, ZZCZ17]. altruist [MV18]. altruistic [Sek12, dWV11]. Alveolar [HWMT17, HWM^{+}16, KKUM10]. always [BEK10]. Alzheimer [Ana11, DD13, GRR^{+}14, HAuR^{+}18, HRHAAA15, PS13, WB15]. ambient [YI18a]. ambiguous [DH16]. ambivalent [NC15]. ambushing [BR12a]. ameloblast [Cox10]. American [DB11]. Ames [HPB^{+}14]. amid [GSRR17]. amides [HAuR^{+}18]. Amino [AH15, CC10, YSH^{+}14, dSMP^{+}11, AHJ18b, BMN16, BB13, BWM^{+}19, CEdLSG^{+}16, CL11, CL13a, Cho11, Cle10, DS19b, Di 14a, DWM15, EMM10, FD18, FL12, FL13, FLW16, FS15b, GZ14a, GZ14b, HPB^{+}14, HYA14, HK11a, HYW11, HY13, JD16a, JD16b, JD17, JSF^{+}11, JCG15, JCG16, KHK15, KB11, KZL14, KSKK15, KGM15b, LLG16, LL13a, LD11, LBS^{+}14, MJ11, MFZ18, MZ18, MP13b, MBE11, MP14b, NBL14, PWZ^{+}19, PDC^{+}17, PWC^{+}15, QLC^{+}18, RBMS17, RSD^{+}16, Rog19, SRS^{+}15, SFV16, TP17, Tun13, WMK13, Woh15, WXC10, XNJ^{+}13, XWW^{+}14, XSLZ16, YGMT12, YG18, YGL^{+}10, ZPdFJ19, ZDG^{+}10, ZSZM14, ZZK14, ZNA^{+}16, ZLW^{+}19, dBJ11]. aminoacyl [Di 17a, Rog19]. aminoacyl-tRNA [Di 17a]. ammonite [MGC15]. amoeba [GF11, UI13]. amoeboid [AA12]. among [AN15, Arc14, Arc16, BKC^{+}16, CBBO16, Cha18, CL13b, DK13a, EAN14, Gar10, KGK10, LCCK12, Mal15, MHMM18, Mas14, MSBB13, MFP^{+}14, NF14b, NI11, NI19, Oka15, Pav19, PBSM19, SS15d, TA15, XSMF12]. amphibians [BKL14]. amplification [BM10, CMPS17, IDM15, JC10]. amplify [RS12, UT17]. amplitude [GZT15, NSK13, RL15, ZCSRG12, tBdR18]. amplitudes [SKK^{+}11]. amylase [KSK^{+}17]. amyloid [ADMR16, Mau15, SPG^{+}18, SMM^{+}13, SAGC12, WGC13, ZGY11, ZSW11]. amyloids [DD13]. amyotrophic [BGCB12, KKS^{+}14]. anabolic [BPGS12a, BPGS12b]. anaerobes [Mar17a]. anaerobic [SI19]. anal [MHMM18]. Analogue [BSMK11, IGG^{+}18]. analogues [RSD^{+}13, TA16]. analyse [MKPVH16]. analyses [KRDJ15, LPS^{+}14, LAH^{+}16, LZL^{+}19, MZWC10, Oko16, TT10a]. Analysing [LAJC19, BR12b, DSTD19, KS13, PBD13, WP17]. Analysis [BMN16, BF16, BBP13, CRLH^{+}19, CC11a, COG19, HNL^{+}11, JKS15, KB19, KMCJ17, KK18, LMM^{+}13, LW18, MFZ18, MZ18, OO11, OL19, PWH^{+}13, PWZ^{+}19, Pat19, SNC12, Str15, SBS17, YNY^{+}10, YWW^{+}14, YCR^{+}15, ZLZ^{+}16, ZLB13, ASK17, ABP^{+}11, AHD^{+}18, ABKS11, AAGO^{+}17, AP17, ALM^{+}19, ADR^{+}11b, BC19, BP18, BDID^{+}12, BUC14, BD10b, BLP10, BHO^{+}18, CRH^{+}15, CW11b, CMPS17, Che12, CW15b, CLL18, Che18, CCY^{+}19, CSS^{+}18, CMD^{+}10, CLW12, CGRRGM^{+}16, CMN^{+}11, DS10b, DQY^{+}15, DJD10, DB11, DG10a, DDLW10, DLYW13, DCS14, DOC17, DLHS11, DDSDW13, ESS19, FLA^{+}19, Fra19a, GAL11b, GZFX14, GP11a, GHBI16, GY19, GdGL17, Gou16, GMK18, GFS15, GZT15, GRRG16, GGG12, GRR^{+}14, HS15, HL10, HME12, HN10, HS14a, HYZS14, HDZ^{+}19, HYW11, Hua16, JTSG14, JB18a, JSZ12, JB19a, JWTW17, JPDP10, JG14, KMHdlP10, KFW12]. analysis [KMM18, KJSS10, KSKS13, KF12, KZ14, KLHS17, KTT^{+}19, KSPA17, KS10, Lab16, LKSM14, LAS14, LCJ16, LY14, LMYL15, LLC15, LZG^{+}19, LZY^{+}15, LCL14, LLJ18, LCQ^{+}18, LC10, LD18, Mac10, MK11, MMY^{+}12, ML09, ML12a, Mal18, MDB12a, MPP^{+}16a, MH12, MÇ16b, MN11, MDEKH13, MBBV14, MB18b, MAFK12, MPS^{+}11, MPT19, MNSZ16, MH14, NAK^{+}11, NMI18, pNMK^{+}10, Nos14, NSS16, OA15, OGE10, Oht12, OYY16, OBAI11, OABI12, PDB^{+}15, PNP^{+}16, PWC^{+}12, Pav14, PPF17, QLZQ11, QJR^{+}16, RLK10, RRG^{+}10, RAA^{+}16, RWW^{+}15, RSI11, RP18, RLBCJ13, RSV10, RLM^{+}14, RM19, RS14b, SAA10, ST17a, SKAG18, SS19, SBW11, SPMGR10, Sha14, She11a, SCF^{+}12, SG15a, SLW^{+}18a, SGD^{+}16, SGW^{+}18, SSKS18, SSKS20, SCS10, SCLC13, SAI14, SQZ^{+}16, TDHC^{+}18, TM16, TAR16, TZY18, TDE12, TAORS10, TF15, VLPH17, VZB^{+}15]. analysis [VM16, VH11b, WTQL10, WZJ^{+}13, WB15, Woo10, WCHC11, WLL^{+}14, WAC14, WGDZ18, XSLZ16, YISG14, YM14, YTGW16, YZFY16, YY15, ZCT18, ZYJL18, ZZL19, ZDY11, ZNC^{+}15, ZMC^{+}18, dBJ11, dLMV^{+}10, dlCGSA16, ISB^{+}11]. analysis-regression [HME12]. Analytic [JBSFB12, SPRF13, CM14, KFS^{+}13, Pie10]. Analytical [BHSB11, FVC15, MSC10, PM10a, SSFG15, VPFV13, AdlPLMZ13, Gal10, LCTG15, LPF11, LCGMH12, LB11, OYA14, WHWZ18]. analytically [SBJ13]. analyze [HZWH10, Kri18]. analyzed [CTPB10, vVLS14]. Analyzing [MS10b, Su16, AHJ^{+}18a, GLZ^{+}13, RSR11]. anastomosis [FRTP13]. Anatomical [RLSM17, DOC17]. anatomically [SMZ^{+}17]. ancestor [Cam15, Di 19c, Sch14]. ancestors [FS12]. Ancestral [ZSSM14, BJOS13, Ell15, GG16, Gri19, HC17, OGO19, RCD16, dF13]. ancestries [Mul11, Mul12]. ancestry [SS15e]. anchorage [CEP14]. Andean [FBU11]. androdioecy [VSLVB15]. androgen [HHSA15, HA15b, SSN^{+}14, YZY^{+}16]. androgens [Jam15]. Anelosimus [QJR^{+}16]. anemia [AGC18]. Anergy [ABD^{+}15, SI11a]. anesthetized [SKK^{+}11]. aneurysm [MAR^{+}17c, NF14a, ZKP15]. aneurysms [MTdS^{+}16]. angiogenesis [BJJR10, BZJP18, BBI14, CXWL11, FBFM12, FJC^{+}10, GBJ18, HGM^{+}16, SGGM11, SLD^{+}17]. angiogenesis-modulating [BZJP18]. angiogenic [HGM^{+}16, Kar16, SOF16, VDRL14]. angioplasty [ZLZ^{+}11]. angle [ARB13, HB16a, LL17a, Oka12]. angles [GP11a, MIH16, Rei12, TSdL14]. anguilliform [HFT15]. Animal [Che14, Fis19, WL12c, AZOLVH18, AHMA^{+}19, CB16, Eft13, EHBC10, Gau11, HFD17, IK15, JMZ13, KKWA18, KJ15, LWLM18, LXJ15, MJS11, NKL10, PWZ^{+}19, PCB14, PM17, PBD13, RD12, SS18, Sue12, TWP16, VLFF12, WP17]. Animals [SS15b, DS10a, Fue14, Fue15, Fue16, Jen10, LYK12, MROS15, NO18, Pon12, RT19, RW14b, SBCR10, SS15c]. anion [Kun16]. anions [RCL^{+}10]. anisogamy [ER12, Yan10b]. anisotropic [CFGRB17, Kro11, PH13]. anisotropy [KCJ^{+}11]. ankle [MS14b, RS14a, SNCM12, ZHAK14]. AnmK [DQY^{+}15]. AnmK-like [DQY^{+}15]. Annotating [LL12]. annotation [GTS15, YCR^{+}15]. annual [CMJD11, DP13b, HYN19, HN19, STLJ18]. annuals [BI12]. Anomalies [TK10b, Sat18]. ansatz [GS10]. Ant [KTO^{+}14, VPFV13, Amo15, AC12, CFR13, Cro18, Cza14, GKWG17, Möl12, NBW10, NBW11, PRMH14, RSC14, RRC^{+}11, SC11b, TD17, UPWK15, vTMW15]. antagonism [BP16, FHS^{+}14, MK14d]. Antagonistic [FE11, GL14, Gia13, LIPD12, MK16a]. Antarctic [SM17b]. antenna [IK15]. anterior [BKL14]. anthracis [KJM17]. anthrax [DFS11, GMS^{+}13, KJM17]. anthropocene [Fue18]. anthropogenic [HMM17]. Anti [RANO10, WGS10, AN13, BHR10, CJ12a, CLA^{+}16, EBE10, EN15, FM10, FM14a, HNR14, HSR12, HGM^{+}16, IHNS16, PYG^{+}19, SOF16, SH15, TDKJ15, YBC17]. anti-angiogenic [HGM^{+}16, SOF16]. anti-anti-dsDNA [AN13]. anti-cancer [CLA^{+}16, FM10, FM14a, IHNS16]. anti-dsDNA [AN13]. anti-inflammatory [PYG^{+}19]. anti-malarial [EN15]. anti-parallel [CJ12a]. anti-predator [BHR10, HSR12]. anti-resorptive [CLA^{+}16]. Anti-social [RANO10, WGS10, HNR14]. anti-tit-for-tat [YBC17]. anti-tumor [EBE10]. anti-viral [SH15]. anti-virulence [TDKJ15]. antiangiogenic [KK18, LBW^{+}13]. antibiotic [CL17a, LN13, PCL^{+}15, SBS17, TDKJ15, WSVT14]. antibiotic-resistant [TDKJ15]. antibiotics [JEA18]. antibodies [AN13, ADR^{+}11b, CDK11, FGH^{+}14, GWCA14, MNSZ16, Sch17, Sch23]. antibody [BJ18, HGO^{+}18, HTB^{+}13, JRR^{+}12, OBK^{+}11, RRG^{+}10, RRG^{+}12, STN19a, SRP16, TW12]. antibody-antigen [RRG^{+}12]. antibody-dependent [HGO^{+}18, HTB^{+}13, STN19a]. anticancer [HPB^{+}14, MBPS17, TTB^{+}18]. anticipatory [FHM18]. anticodons [Sel14a]. antidiabetic [BAGG14]. antifreeze [KCM^{+}11, MP14b]. antigen [MCL19, RRG^{+}12, RSI11]. antigenic [CST16, GAV15]. antimalarial [Kle14]. Antimicrobial [PDC^{+}17, PPF17, BC19, MB18a, ZSH^{+}16]. antimüllerian [MS13]. antioxidant [BRK19, DH18, FBAPMD13, LL13a]. antiparallel [ZDG^{+}10]. antipredator [BR16]. antiretroviral [WML^{+}17]. Antisense [Sel12b, Sel11]. antisocial [GT12]. antisymmetry [Kam11]. antitermination [Sel12b]. Antiviral [TGZF11, GN10, LCCC10, MSL^{+}16]. ants [CDD12, Cro19, GK17, KCMF11, MAW12, RRBM12, STN^{+}19b]. any [JS15, JS17, JS18]. aortic [ASC15, ASC16, SWPC^{+}16]. APC [SWR11]. ape [NH12]. aperture [ARB13]. Apis [EM11, MKJS13]. apnea [SEK^{+}10]. apoptosis [LTL^{+}15, LMM18, ML09, ML12a, PHTP^{+}12, YLWZ10, ZL18a]. apoptotic [CSKZ19, KKO^{+}18]. APP [HRHAAA15]. apparatus [TWH^{+}13]. apparent [CL10, PHC12]. appear [MAA18]. appearance [AIY16]. appendix [JEA18]. appetite [GCO^{+}11]. apple [vdSS12]. Apples [GP11a]. Application [BTG^{+}15, Cox10, DEK15, DS16, FGH^{+}14, HCS^{+}19, JS18, KP16, MJ11, MGB17, MAA18, PDC^{+}17, SDRA^{+}15, TSF^{+}19, VACGF17, WMPF^{+}15, ZLZ^{+}11, dLMV^{+}10, BGW15, CSM^{+}14, DLYZ11, DYS^{+}13, Dim10, GMR16, IMH15, MMG^{+}16, Mor11, OT13, OVKL14, OL19, PCB14, PRV^{+}14, RKMG12, SMD^{+}16, TTC10, TBA14, Tre19, VAT18, WWY12, WZJ^{+}13, YYH^{+}14, CMMR13, JS17]. Applications [BLP10, Gol16, Lan16, CZJQ14, EEHMH18, FR13b, FR14, GK16, LAS14, LLZ13, MH11, Pai19, PYG^{+}19, WZY14, XM11, XWW^{+}19, YZZ13, YCY14, ER18]. Applied [ZLW16, GSRR17, MB16, PRSC11]. Applying [DGW^{+}18, Mor19, SOIO10, CDM^{+}14, GK13]. apposition [PVGA12]. appraisal [LP17]. Approach [JSC^{+}16, KS10, ASK17, ACM16, ADS^{+}19, AMSSG16, ADCG14, AOR17, Amo15, ACvKA10, ARM18, ÁTC^{+}14, BBT^{+}15, BCGD10, BMdA17, BORA10, CFS^{+}19, CPF^{+}13, CL14, CVPCV^{+}15, CM14, CTL^{+}15, DB10, DGM15, DKP^{+}18, Dim17, DXW^{+}16, Dol19, DRPM17, EBE17, FKMG15, FKV19, GAGP^{+}14, GSYS10, Ghu18, GBC^{+}16, GGR11, GM16b, Gre15, GT15, Hal16, HXL16, HXL18a, HME12, HB10, HHF11, IHNS16, JAM18, JLX^{+}16, JSF^{+}11, JSB15, JPDP10, KCM^{+}11, KSK^{+}11, KSP18, KTH16, Kla10, Kon12, Kri16, KSM^{+}15, LvBJ16, LN13, cLCJ^{+}10, LKAJ18, LGPS17, LPvSP11, LPvSP12, MFN^{+}18, MPJH13, MSND12, MB17, MFM16, MT14, MM19, MKBE17a, MKBE17b, MTE15, MSRW16, MPS^{+}11, MPK^{+}12, MF15, MMCZM12, MHH13, MDD13, MJV16, NKM^{+}12, NGJ^{+}14, NHM10, NBW10, NBW11, OWB14, OBK^{+}11, PDM17, PM14, Pav19, PMKM10]. approach [PGF11, QW11, RLK10, RA10, RMRC^{+}16, Rev15, RRC^{+}11, RBKW19, RDP16, RSR11, RBMPP^{+}15, SMHB10, SL10, SBR16, SK11a, SCKL15, SLT^{+}18, SZ10, SDPC11, ŚKSRW16, SST19, TvMG16, TKB^{+}18, TKKE19, TTB^{+}18, TWC^{+}19, TT10a, TK10c, WHHS15, XNJ^{+}13, XSCS12, YS14, YBH^{+}19, YA14, YLH12, YS11, ZMAM19, ZGY11, ZJSC16, ZD18, dMP11]. Approached [ZLY14]. Approaches [JD16a, JD16b, AH16, BCPM^{+}16, CB15, EB15, Fer12, Gal10, GSCS11, HMMSRSD15, JAHKH12, Mic13, MMLK11, NRS^{+}16, OAJK10, QZ14, SPRF13, SD16, SYY17a, SYY17b, SSRA16]. appropriate [DZW10, Nak12]. Approximate [KBF18, Ou19, MKBE17b, PHK15, RSC14, Wal12, XGZ17]. Approximating [HBK12, OBHS19]. approximation [CCF^{+}14, FMLM12, GHBC14, HM11b, JRG14a, JRG14b, JRG14c, LLH^{+}13, MFG14, RB14, Wax11b]. approximations [MPBS17]. April [Ano10-39, Ano10-42, Ano11-42, Ano11z, Ano12-33, Ano13-47, Ano13-44, Ano14z, Ano14-35, Ano15-38, Ano15-34, Ano16-43, Ano16-42, Ano17-28, Ano17-42, Ano18-33, Ano18-36, Ano19-36, Ano19-39]. AQP1 [KH19]. aquatic [NTED^{+}19, QA15, Whi11]. Aquilaria [LCQ^{+}18]. Arabidopsis [BHH^{+}14, HNA15, MPP^{+}15, Mit16, MBK^{+}11, MBKB13, MLL^{+}16, OFT15, SCÁBM10, SCÁBM11, SSI^{+}16, SDRA^{+}15]. arachnid [FWSG17]. arbitrary [CM14, DNB12, KS15, SWG^{+}15]. arborescence [HB12b]. archaea [SGD^{+}16]. archaeal [LY14]. archetype [Hou13]. architecture [CRLH^{+}19, DKP^{+}18, KOF^{+}14, KSM^{+}15, MBKB13, PDNP16, RW14a, ST16]. area [BH13a, GABM12, IMH15, NdORC11, TAI^{+}18, TLM14, WCF12, ZGG^{+}19]. area-concentrated [BH13a]. areas [KG13, KMCJ17, KJ15, TM14, Tak16]. areata [DPC15]. Argentine [RRBM12]. arginine [AKR^{+}18]. Argulus [MNH^{+}12]. argument [ER12]. arguments [Di 17b, RW14b, ASH15]. arise [Ell15, LRA^{+}13]. arising [APW10, BPP^{+}16, DD13, ZW16]. Arkansas [GAY^{+}15]. armadillo [DS15b]. armed [Mor19]. Arms [IST11, CL10]. Armstrong [XF13]. army [GK17, GKWG17, TD17]. Arousal [PRK13]. arrangement [YM14]. array [QYO10]. arrays [WRH^{+}16]. arrivals [BBR12]. arsenic [SAKG13]. arsenicals [BTF19]. art [MDB12b]. artemisinin [CKZ^{+}17]. artemisinin-induced [CKZ^{+}17]. arterial [JSK16, Js12, LFD11, Ren13, ZLZ^{+}11, ZR16]. arteries [AV19, DLMK12, LZWK11, MH14, Van15, VSW10]. arterioles [LH13]. arteriovenous [FWR19]. artery [DPCM16, MAR^{+}17c]. arthritis [MF19]. arthropod [FMS^{+}10, YLY15]. article [JN14, gSxFH^{+}12]. articular [HRG11, SABB15, WGS15]. Artificial [GM16b, IL12, AJC12, CLG^{+}10, CLG^{+}11, GMK18, HL19, LCMC14, MFM16, Sal15, XSKA17]. artificially [Bry13]. artiodactyl [OB15]. aryl [TAR16]. Ascaris [CH18]. ascendency [BJ17]. ascorbic [BRK16, Edg19]. asexual [FOT^{+}15, JJ17, Jam16a, MO12]. Asia [TDM14]. Asia-pacific [TDM14]. aspects [BBJ^{+}10, BHR10, Cro17, Ger16, LDF^{+}11a]. Aspergillus [GT11a, GS13b, KHS13]. aspiration [MN11, TM12]. assay [DAA17, HYN19, IGL^{+}12, TLT^{+}15, TSMB14]. assays [HPM^{+}17, JSP^{+}16, JRB^{+}16]. assemblies [Kur10, KI15, ML14]. assembly [CC11a, CCB11a, CPH^{+}14, Dun11, EJ16, FWSG17, MH18, Mou12b, MKF^{+}14, MSIR10]. assess [FKB^{+}12, KCJ^{+}11, PGH11, RFME^{+}12]. assessed [SWPC^{+}16]. Assessing [BLV18, CDM^{+}14, GML10, HLH^{+}18, KCZ^{+}19, LDdA^{+}13, MHKA16, PZLF19, PRM12, TGZF11, ZJY19]. Assessment [Kro10, LG18, WHS^{+}13, CC11b, GT11a, HFD17, LFW^{+}18, NNK^{+}15, OSN18, Pan11, SN16, Sig12, US10]. Assessor [CCNT19]. assignment [dEG11]. assignments [DS19b]. assistance [KSBvL^{+}12]. assisted [BB13, KSK^{+}16, LSLL13, RMRC^{+}16]. associated [AD15, ADS^{+}19, AFM10, BRA15, Ber14, BB19, CRH^{+}15, CST^{+}12, HBSF11, KEHK17, Kri18, KKS^{+}14, LZG^{+}19, LMHF13, NBW10, NBW11, OABI12, PS13, RRG^{+}10, RRC^{+}11, RFdL15, SRP16, Voe18, WDL^{+}17, ZK18, ZDF^{+}12, ZXS^{+}19]. Association [RMM^{+}16b, AKR^{+}18, GZ19, HMWB13, LCG12, PS13, Sar10, Sha14, TDZ^{+}18, ZK10]. associations [PSAA13, Van16a]. associative [GJ10]. Assortativity [NR16, PPM12, PMP13]. Assortment [CAV16, AS12, BBD18, VMZK^{+}19]. assumption [LJW^{+}16, RR16, XGZ17]. assumption-free [XGZ17]. Assumptions [TWP16, MMB14, SWPC^{+}16]. asthma [Don16, SGDL12]. asthmatic [Don17, PDT^{+}10]. astragalus [PRSC11]. astrocyte [LL13b]. astrocytes [ABJ12, PRV^{+}14, ZMAM19]. astrocytic [KPD18]. Asymmetric [HSM19, HFT^{+}18, PGKZ17, ROF17, CLA17, DGJ14, LS15b, Lee16b, MG10a, yTRSC13, Mig16, MAA^{+}17, NdlPZA10, NTFK11, RPD14, ST16, Sel13, Sht17, TWW19, UI10]. Asymmetrical [FCC^{+}10, MDS16]. asymmetrically [RKMG12]. asymmetry [DB18, DZE11, HTT19, Hor13, KMZR18, KJ11a, MS14b]. asymptomatic [Gru18]. Asymptotic [MAFK12, BP18, DGW^{+}18, Rei12]. asymptotics [LK19]. asynchronism [GC11]. asynchronous [SAA10]. Athel [HS16]. atherosclerosis [Fok12, WBMM18]. atherosclerotic [BD12, FBM19, GCYL18, KDMD13]. atmospheric [AK11]. atom [Mas18, MIH16]. atomic [FTS15, Woh15, ZSW11]. atomic-resolution [ZSW11]. atoms [MPY14]. atopic [TDHC^{+}18]. ATP [BFGS10, CPS10, CW11c, MS14a, MS18, PMK15, STNT17]. ATP-binding [MS14a]. ATP-sensitive [CW11c]. atrial [LWB^{+}15, RLSM17]. attachment [FPG11]. attainment [DGJ15]. attempt [Mal18]. attenuate [RS12]. attenuation [JYY^{+}18]. attitude [SN11, SN14]. attitude-behavior [SN14]. attitudes [MJI14]. attract [Arc11]. attractant [RG10]. attraction [Str11]. attractive [KAN11]. Attractor [SAA10, HXL16, HXL18a, WCF12]. Attractors [SVB^{+}10, HRCA19, JZL13]. attribute [Cho11]. attrition [IK16]. atypical [Dol19]. AUC [BJ18]. auditory [Lug15, ZKMB19]. Aug [MR19]. August [Ano10-37, Ano10-45, Ano11y, Ano11-44, Ano12-28, Ano12-40, Ano13-43, Ano13-45, Ano14-32, Ano14-39, Ano15z, Ano15-46, Ano16-32, Ano16-40, Ano17-33, Ano17-45, Ano18y, Ano18z, Ano19-49, Ano19-29]. aurantia [AK11]. aureus [PDW10]. Australian [RSC14]. autapse [WSLC14]. autism [Jam14, Jam16b, VHM^{+}17]. Auto [BPG^{+}18, HGB^{+}13, ZL18a]. Auto-correlated [BPG^{+}18]. auto-cross [ZL18a]. auto-regulation [HGB^{+}13]. Autoantigen [PBP15]. Autocatalytic [GS17a, Hor17, GFH^{+}18, SHS13]. autocorrelation [HM11b, XLL^{+}10]. autoencoders [NK18]. autoimmune [AN10, AN13, BN15, JDSPK15, LHW14, PBP15, RB15, RBF15]. Autoimmunity [Coh15a, BN12, LPB17, LPB18, MBF^{+}15, MM15a, SI11a, Tau15]. automata [GMOP12, MMAS13, SZ18, WMD16, ZZS19]. automaton [BCPG18, CLCH10, PGHC12, SHB^{+}17]. autonomic [WHS^{+}13]. Autonomous [WZJ^{+}13]. autoparasitism [HYQ^{+}16]. autophagy [JL16]. autophosphorylation [DMP15, HKS15]. autoreactive [JDPK15]. autoregulatory [Das13]. autorotating [SIHO15]. autosomal [Sek12]. autotrophs [CG10]. auxin [FKB13, HTM15, MPP^{+}15, MBK^{+}11, MBKB13, SLC12]. availability [AR17, TRJD19]. available [ASF^{+}15, TT17]. Avascular [KHP^{+}12, BJO^{+}16, BAM^{+}11, LS15c, VDD^{+}17]. avenues [SSJK18]. average [FL12, FL13, GDF18, KTI18, LM15, NWZ15, VL11, ZCAB17]. average-level [NWZ15]. averaged [ZCS^{+}15]. averages [BCK19]. averse [MSBB13]. avert [AK15]. avian [BUL12a, BTW11, CW15b, DGD^{+}11, GBN14, JMS12, Oko16, PHH18, RC16, XSTW14]. avidity [JDPK15]. avium [AMSSG16, AHMA^{+}19, CRH^{+}15]. avoid [Car17, MB18c]. avoidance [CJRR11, Che12, TB18, Veg10]. Avoiding [LSS^{+}14, CW14, NTY16]. aware [RCL16, SMD^{+}16]. awareness [ABR19, Esc11, FGJ10, JKS15, ZN18]. axial [BBS18b, SA17]. axiomatic [MP16, OWB14]. axis [CMJD11, KL18, SSKK17]. axisymmetric [HPP10, POP12]. Axonal [LM15, DHHP14, ML12b, SL19]. axoneme [CTPB10, CL10]. axons [KK17b]. ayu [TIY^{+}11]. azar [MCCC^{+}10].

B [HSiI^{+}19, KB11, SQZ^{+}16, ZZR13, ABD^{+}15, AD15, ÁTC^{+}14, CKB18, GM13, GC18, HSiI^{+}19, Jam11, KB19, LZY^{+}15, MR11b, MG15, MSL^{+}16, NS11, Nak16a, PaCZ10, TC11, TCYY^{+}12, WPH^{+}12, ZBA14, ZZR10, ZRZ15]. B-cell [KB11, SQZ^{+}16]. B. [ZCSRG12]. B19 [KEHK17]. Babesiosis [ABB^{+}19]. Bac2A [BC19]. Bac2A-derived [BC19]. BACE1 [HRHAAA15, SK19]. Bacillus [KJM17, BMGC11a, BMGC11b, BLP10, RBSD10]. back [IIKT13, WCM15]. backbone [LL12, NSBL10]. Background [Abb10, BCKS13, DZW10, HTN14, HG18a, Lug15]. backgrounds [SBJ^{+}18]. backward [BSKV18, Ger16, UPWK15, ZTT18]. Bacon [RRC^{+}11]. BacPP [dEG11]. bacteria [AES^{+}13, BMN16, CFT11, Dai12, DHV19, FE11, HK17, Kee10, LW17, MCWF14, Mic15b, NS18, NGS^{+}16, OUMA10, Sht17, SGD^{+}16, SCLC13, TDKJ15, WST15, Wu14, Zuc14]. Bacterial [AABS16, JS17, JS18, Sch17, Sch23, TDKJ15, dEG11, AAJGCD15, BCKS13, BCPG18, BKPV15, CEdLSG^{+}16, CXC18, DBD15, FWMC14, FIJJ11, GSYS10, GCvWE^{+}14, HLS^{+}12, JS15, KDFM14, KAKK19, LDHD14, LG10a, LZ18, MBCM12, Mor11, MKPVH16, NvD17, SRD^{+}15, SC10b, SMS19, SPRF13, YSI18, YGL^{+}10, ZDF^{+}12]. bactericidal [XLG^{+}15]. bacteriophage [ESG10, SSSD17]. bacterium [GGGM16]. BAD [HBSF11, YLWZ10]. bagoti [RSC14]. balance [BUC14, CNDK15, HS14c, KTH16, MO12, PDRC16, STNT17, SSI^{+}16, SMdA16, Vaz10, VMZK^{+}19]. balanced [IIKH16]. Balancing [CB16, TED10, CXC18, KOS13]. Baldwin [SSF15]. balloon [ZLZ^{+}11]. banded [BS19b]. bandit [Mor19]. bands [HSG^{+}18]. bank [GCvWE^{+}14]. Bap [SRP16]. Barber [Ou19]. barcodes [LLC15]. barcoding [Bar18]. bare [BCF19]. bare-nosed [BCF19]. Bargaining [Ros13, GK10]. bark [KLB^{+}16]. barley [DKP^{+}18, OFB10]. barnacles [YYST13, YHI14]. Barr [HR12]. barrel [Pat19]. barrier [HPM^{+}17, MYN^{+}13, MAM16]. basal [RLM^{+}14]. Base [Tri10, LF10, PM14]. Based [GCB17, LPGTBMA19, LMJ^{+}18, MOSS15, ZKMB19, ACR^{+}17, AOM19, ACdlRMR^{+}11, AH19, AMSSG16, AA16, BRA15, BKR14, BSV^{+}10, BYY^{+}11, BWS10, BJ17, BPF^{+}19, BCS^{+}16, BRK19, CS14a, CSR^{+}05, CCG^{+}18, CLC11, CZW^{+}11, CPH^{+}14, CA17, CDD12, CZPC^{+}18, Cro18, DJD10, DHS^{+}15, DWL^{+}14, DQS^{+}15, DSPM14, Di 16, DYQ^{+}14, DXW^{+}16, DBG18, DZW10, DDB10, DZE11, Est10, FJS^{+}19, FP18, FHW^{+}10, FBAPMD13, FLCS^{+}15, FM15a, For10, FYZ15, FPR10, GK16, Gal10, GVGC15, GZ14a, GZ14b, GSYS10, GW13, GdGL17, GBC^{+}16, GJ15b, GGGM16, lGLLL14, GGX^{+}10, GSN^{+}11, GGCJ16, GZ19, HGM15, HB12a, HNO18, Hay16, aHLZ^{+}12, HFT^{+}18, HS12, HLS^{+}12, HXL18b, HZC^{+}10, HZWH10, HMM11, HYW11, HKR^{+}19, HM13, IMD16, Jab10, JM12, JAM18, JWS^{+}10, JZ17, JSSZ12, JR17a, JB11, JO13, JLH^{+}15b, KPK^{+}13, KKOM18, KAPR12, KBV11, KBV12, KHK^{+}17, KHNM16]. based [Kit10, KCJ^{+}11, Kon17, KSRN19, KMM18, KCD11a, KJSS10, KF15, KDK14a, KDK14b, KP16, KLHS17, KBF18, LJM15, LTSTD15, LFW^{+}18, LFM14, LX10, LLZ13, LLW^{+}18, LDWW14, LTL^{+}15, LPH19, LZMM15, LC16, LSSG10, LDH^{+}14, MS14a, MCL19, METC12, MJ11, MSND12, ML09, ML12a, MBE10, MMAS13, MV10, Mas18, Mas12, MA13, Mau15, MBBD13, MAFK12, MZAI19, Mei12a, Mei12b, Mei13a, Mei13b, Mei14, MFG14, MP12, MIH16, MGM13a, MGM13b, Möl12, MP14b, MKF^{+}14, MAFB18, MR18, MR19, NdMLLBLC13, NGND12, NBL10, NSO15, NKM^{+}12, NYSM12, Nos14, Oko16, OYY16, Opr10, PDM17, Pat16, PFJS15, PG18b, PWHW16, PRMH14, PMD^{+}18, PKH11, PSAA13, QW11, QLZQ11, QZ14, QW12, QLWC19, RRG^{+}12, RBMS17, RA10, RMRC^{+}16, RC11, Rev15, Rog19, Rou14, RBMPP^{+}15, RG12, SA17, SM17a, SL10, SK15, SK16a, SHB^{+}17]. based [SMG18, SMS19, SNY^{+}13, SSD12, SGDL12, SG15b, SOF16, SCS10, SH17, SDPC11, SRAL12, SAZ^{+}14, SQZ^{+}16, SNCM12, SA14, TTG19, TWR^{+}18, Tak17b, TF17, TCB13, UPWK15, Uit11, dlRVFK^{+}16, VG13, WL13, WMK15, WW12b, WY13, WMCL18, WZ17, Whi11, WK18b, WXC10, WGDZ18, XLL^{+}10, XM11, XNW17, XSCS12, YZMY18, YZW10, YZZ12, YZFY16, YYC19, YC13, YC15, YS11, YHY14, YZ19, ZMS11, ZOG14, ZJSC16, ZT16, ZZC^{+}17, ZDY11, ZC14, ZMS17, ZLDZ13, dF13, dSRM15, ANMH11, Mil11]. basement [DFC^{+}12]. bases [XM11]. Basic [CCAdS13, JD13, PLF18, BPP^{+}16, GBRS19, GHS18, GHS20, LLW15, MCN10, SS11, Wal10, WDL^{+}13]. Basis [IU18, Gri11, Kam11, KAZ13, MGC15, NS16, NHM10, PG18a, Sat17, Uit11, VBD10, YW13b]. bat [BLS^{+}12, KOS13, MSB16, VP18]. batch [Aya12]. bats [LLD^{+}17, LA15, MHSH12, VP18]. baumannii [RRG^{+}10, RRG^{+}12, SRP16]. Bayes [HTK14, MKBE17a, MKBE17b, SSP15]. Bayesian [Bac15, BSB^{+}13, CR14, CLW12, DGM15, DRPM17, GBIR13, GMR16, IS12, KEG17, KBF18, LTS^{+}16, MS14c, RM17b, SS15d, Ste13, VM12, Wal12, ZMC^{+}18, vLFM^{+}19]. Bayesness [MKE15]. BCG [BMGC11a, BMGC11b]. Bcl [HBSF11]. Bcl- [HBSF11]. BCR [TCYY^{+}12]. BDNF [LRA^{+}13]. Be [SYI17, Di 12, Di 18, GCM14, GvVT14, Gri15b, HC17, Har15, HKM12, Jen10, KNA^{+}18, LDA13, LPH19, Mar11, MN12, MBLC17, MALAN17, OO11, PDC^{+}17, PW18a, RDD14, RW14b, Sch19b, SWPC^{+}16, vVLS14]. beak [SYR11]. beam [She11a]. beard [Lai11]. beards [GvVT14]. beat [WMPF^{+}15]. beat-to-beat [WMPF^{+}15]. beautiful [BD10b]. because [Di 14a]. Becker [DFMR19]. become [BH13b, YSYI13]. bed [ADB^{+}13, NDZMA14, NMAZP16]. bed-net [ADB^{+}13, NDZMA14]. bed-nets [NMAZP16]. Beddington [GG12, GG14]. bednet [PRM14]. beds [BSMK11, SM16]. bee [BIMC17, MKJS13, NPS10]. beehives [STKE12]. been [BM14, Di 14b]. bees [SOIO10]. beetle [APS^{+}13, APBS15, KLB^{+}16, Whi11]. beetle-fungus [APS^{+}13, APBS15]. beetles [BBR12, FKMG15]. before [ARB13, CBC17]. behalf [LPH19]. behaving [CCG^{+}18]. behavior [ADCG14, AA12, Alv18, AdGM12, BPFR16, BBP13, CW11c, Che12, CSM^{+}14, CCdS15, Cox10, Déb17, FAFA17, FHM18, Fre10, GWCA14, HE16, JLQ^{+}12, KDG13, KMZR18, KOS13, KMM18, LDF^{+}11a, LRA^{+}13, MMAS13, MPP^{+}18, NBS^{+}13, OYIM10, OVKL14, PaCZ10, PJ18, SGGHS19, SN11, SN14, SPG^{+}18, SPH12, SH17, SSKK17, Tac10, TLK19, WSW10, WCHC11, WYL^{+}18, XLSF19]. behavior- [SPG^{+}18]. behavior-induced [MPP^{+}18]. Behavioral [Akç12, Kři13, Abr14, CA17, Cre10, HAH19, IDI11, ISB^{+}11, LXJ15, MKJS13, MMRCCC10, Ran12, RW15b]. behaviors [LLLV11, LLL13, QJR^{+}16, RN12, SHW16, TM15a, WHWZ18, ZG10a, ZG10b]. behaviour [ACH15, ADR^{+}11b, ASL^{+}18, BB10, BAR14, CJKR10, EME^{+}16, FD11b, GA16, Hor11a, JAB18, LDM16, PB18, PCN17, TB11, TCB13, Voh17, WPPD16, ZLB^{+}18]. behavioural [DBD12, FGJ10, NvD17, WL12c]. behaviours [Eft13, WWIG19]. behind [TF17]. belief [DWL^{+}14, Kur17, TI12a]. belief-based [DWL^{+}14]. believer [HLI15]. believers [SGA^{+}12]. bells [HM15]. benchmarking [PPF17]. bendability [Tri10]. Bending [CTPB10]. bends [HRG11]. beneath [DFC^{+}12]. Beneficial [LDF11b, HKM12]. benefit [VGZS18]. benefits [ARB13, HM15, JA13, Moo14, NT10, PBA12, RMB15]. benthic [JB18a]. benthic-pelagic [JB18a]. benthos [QA15]. Bernal [Pia19]. Bernoulli [Pia19]. Bertalanffy [OYA14, RRRTR10]. best [LN13, GDPPSS^{+}11]. Bet [MHF^{+}13, Uit11, YY18, MGT17]. Bet-hedging [MHF^{+}13, YY18, MGT17]. beta [CJ12a, HDZ^{+}19, JDSPK15, Mau15, She11b, ZDG^{+}10]. beta-cell [JDSPK15]. beta-helix [She11b]. beta-sheet [Mau15]. beta-strands [CJ12a, ZDG^{+}10]. better [Hou15, HXL18b, RAAS15]. between [ASC16, AGB^{+}14, ABKS11, ARG14, AH12a, AHW13, ARB13, BYJ17, Bal13, BTO14, BAGG14, BST14, BZJP18, BMN16, BLZ^{+}19, BMB^{+}18a, BL14, BYM^{+}18, CPGF^{+}16, CFZ14, CCS^{+}16, CK17, Das18, DBBW09, DBBW11, DI13, DD13, DAA17, Di 19a, DBG18, Dol16, DDS13, DPCM16, DHT16, ESW13, FD18, FB18a, FE10, Fil15, FVTS16, FZL18, GMZM15, GZT12, GCB17, GML10, GLZ^{+}11, GBM18, GZY12, GN10, Gün12, Gün13, Han10, Has14, HHSA15, IK15, IST11, IOY15, JC10, KTI18, KL13, KCMF11, KLM14, KMCJ17, KN11, KG12, Kro10, KSPA17, LL17a, LV17, LDH^{+}12, LKP^{+}12, LM14, LMCW18, LPF11, Lug15, MZWC10, MVGGB18, MGML10, MGO^{+}15, MK16a, MW14, MK19, MS18, MBLC17, Miy17, MD16, MSS10, MLL^{+}16, NGN13, NG11, NMAZP16, NTOI16, Ots11, PSJ15, PBB10, POP12]. between [Pel18, PB18, PL14a, PL14b, PB16, PBA12, Ram10, RPD14, RS19, RCD19, SW15, SGGM11, Sek12, SN11, Sel10, Sel15b, SS15a, SQS^{+}12, SZ15a, SZ16, SC12, SAB17, SOF16, SKK^{+}11, Str13, SK16b, SP14, SYY17a, SYY17b, TP14, TZY18, TK10c, UK17, VMZK^{+}19, Wal16, WSP14, WHT17, WZ18, WK18a, WZWY11, XCW^{+}18, XLSF19, YZW10, YZZ12, ZLZ^{+}15b, ZLZ^{+}16, ZLB13, dlEBRM15, vLBJ^{+}13, vNÁBG12]. between-host [CFZ14, LMCW18]. Beyond [AE17, CP14, KDST15, KGC18, HMW16, WF18, Pag19]. BFDT [GJ15b]. bFGF [TBMM19]. bi [ATRR10, DFMR19, HTK14, RSC14, SYI17, SLDP13, KHK^{+}17]. bi-directional [SYI17]. bi-gram [HTK14, SLDP13]. bi-modal [RSC14]. bi-monomeric [DFMR19]. Bi-PSSM [KHK^{+}17]. bi-stably [ATRR10]. biarticular [CSB15, Cle18a]. bias [GRRG16, IDM15, Jan15, KMCJ17, OL19, PCC15, ZLZ^{+}15b]. biased [BWB11, BMF^{+}18, Gar10, IH17, KHKI17, WL15]. biases [LY14]. biclustering [WY13]. bicyclic [SK10]. BID [HBSF11]. bidimensional [MR10]. Bidirectional [FZ14, HHJR11, HKS16, SS15f, WK12, YI18b]. Bidirectionality [MK17]. biennial [CE17]. bifunctional [SBR16]. Bifurcation [HS14a, LJ18, TDHC^{+}18, ZZ14b, AJM19, BA10, DLHS11, Ger16, KP16, MAR^{+}17c, TAORS10, UPWK15, VK10, ZTT18]. Bifurcations [GFS15, ZG10c]. bigram [DLL^{+}17]. Bihar [MCCC^{+}10]. bijection [LAPK14]. bijections [FDS13]. bijective [FGGS15]. Bilateral [WTQL10]. bilayer [KMZR18]. bilinear [Don16, MPCTGJ^{+}15]. Bimatrix [CK19, Oht10]. bimodal [FPD15, TF18]. bimodality [Mic13]. binary [GJ15b, Mul11, Mul12, QLWC19, TAI^{+}18, UI10]. bind [RBRB16]. binding [ÅCHS19, AFS19, ÁTC^{+}14, BMSEE14, BZJP18, BGCB12, BF16, BSM^{+}14, FLM18, HG18a, HKH18, JSZ12, JRR^{+}12, JB19a, JLQ^{+}12, KHP^{+}12, KBK16, KGM^{+}15a, KG12, KHH^{+}17, KC11b, LYM^{+}19, MS14a, MPY14, MSND12, NTC^{+}11, NHSX14, NHTS14, OBK^{+}11, PHÁBAI^{+}16, PBvdG10, QQJW13, RSS^{+}18, RG18, SGGHS19, TLT^{+}15, UT17, WB15, WHH17, ZGS^{+}10, ZLZ^{+}15b, pZWZ^{+}16, dS15a]. Binding-upon-folding [LYM^{+}19]. binomial [Mog15, WKH16]. binuclear [RS19]. bio [AANF16, DQS^{+}15, GKG^{+}18, MPCTGJ^{+}15, MS10b, NGN13, Sar10, ZZL^{+}11]. bio-adhesive [Sar10]. bio-computational [ZZL^{+}11]. bio-economic [GKG^{+}18]. bio-inspired [AANF16]. bio-macromolecular [MPCTGJ^{+}15]. bio-markers [DQS^{+}15]. bio-molecules [MS10b]. bioactivity [MR18, MR19]. biochemical [AdGM12, Chu18, FMS^{+}12, FVTS16, GM13, HKS16, KMD^{+}12, MK11, MC19, NBC16a, NBC16b, PP12, PP14, RLSN14, SOCF14, TP10c, TLK19]. biochemistry [OSF11]. biocidal [MW13]. bioconvective [ZG19]. Biodiversity [SDJ^{+}15, Bar18, DBD15, Edu16, IS10, MB12b, MBRRI19, PMSY17, SCF^{+}18, VY12, WF17]. bioenergetic [BBR10, CMGN17]. bioenergetics [GSL13]. biofabrication [YMW12]. biofilm [ALM^{+}19, FWB^{+}12, PWL^{+}11, RRG^{+}10, SRP16, SMS19]. biofilms [PBR17b, TGB^{+}18, ZSH^{+}16]. biofluids [Zad11]. biofuel [AK15, PBR17b]. biogeography [LPP^{+}19]. Bioinformatic [Sha14, BBT^{+}15]. bioinformatics [ADCG14]. Biokinetic [CMM19]. Biol [AEM14, BPF15b, BBM^{+}13a, BBM^{+}13b, BPGS12a, BMGC11a, CLG^{+}11, CNG^{+}12, DBBW11, EPJ^{+}11, FIS16a, FS16a, GZ14b, GS14, dOGL13, Gün13, JN14, JRG14a, JRG14b, JCLS^{+}21, KDK14a, KMLT14, LBS^{+}11, LGK^{+}12, Lei10, Lio24, LZT13, LHPF18, LPvSP12, ML12a, Mei13a, Mei13b, MLBA13, Mul12, NO14a, PSS^{+}13, PL14a, SCÁBM11, Sch21, Sch23, SZ16, SSKS20, SC13a, SRAL12, TSM13a, Van16b, WL12a, YCH^{+}17a, ZG10a, ZSCL16a, ZZR13]. biolarvicides [PNVN18]. bioliquid [AAJ15]. bioliquid-filled [AAJ15]. biologic [RAR19]. Biological [Di 10, Fue18, Gra15, MM15b, PGHC12, AAGO^{+}17, AKNP18, BPP^{+}16, BMI19, Bar11, BBWS^{+}13, BHP17, BHLH12, BE14, CCG^{+}18, CMBP19, CNG^{+}12, CG11, DB10, DS16, DG10a, Dra19, Est10, FPG11, FDM^{+}11, GI15, GE18, Gre16, GSN^{+}11, HRC^{+}12, HM10b, HZWH10, JPDP10, KSA16, Kit10, Kon17, KMM18, Kro10, LCSH14, LLS^{+}15, LBB^{+}13, LDH^{+}12, LLZ13, MK18a, MRPLAS15, MPK^{+}12, MdSPBL16, Mit14, MROS15, Mor16, OO18, OA12, PAK11, PBEI12, PR13, RGBR17, RHJ11, RKMG12, SSD13, SK15, SS19, SWTO15, SY11, SVA18, SLO10, TM16, TK19, TZY18, TF17, Tou14, Tsu19, VGPS18, Wal10, WTQL10, WST15, YZW10, YW13a, ZZ14b, ZGW16, vdSS12]. Biologically [DWG^{+}19, WMT10, APW10, Cle10, LDJW16]. biologist [Sza15b]. Biology [DHK13, EJ19, GHS20, JS18, LPB18, PCT19, Pan19, PP19, RTEKG19b, RTEKG19a, SYY17a, WLW20, AENK12, Ano19-27, DGW^{+}18, GMM^{+}13, Gri15a, KCE^{+}11, LA12, Mal18, MMLK11, PBRW11, PJL14, SS19, TT17, WSM12, dlEBRM15]. bioluminescent [FL13]. biomacromolecule [ZM10]. biomarkers [QZ14, ZZC14, ZC14]. biomass [MR10, PMKM10, RHT18]. biomaterials [Zad11]. biomathematical [RRL^{+}13, SGL10]. biomechanical [TF15, VF12, YWP13, dLMD^{+}11]. Biomechanics [CEP14, PWC^{+}12, ZTK12]. BioModels [BFS18]. Biomolecular [ZLY^{+}13, LCG12]. biomolecules [Cha17]. Biophysical [AH12a, BKSW16, FM15b, HRG11, KG15, PDG17, SSZR17, TALC16, ZCA^{+}14]. BioPlex [YWZ^{+}16]. biopoiesis [Pia19]. biopolymer [FK16]. bioreactor [CWB^{+}17, SSR^{+}19]. bioremediation [RAAS15]. biosensor [PDNP16]. biospheres [NWWL17]. biosynthesis [BP18, BSM^{+}14, BLP10, FD17, MB14, Pie12, PCS^{+}18]. biosynthesis/inactivation [Pie12]. biosynthetic [FD18]. biosystems [Dun11]. bipedal [SHW16, ZHAK14]. bipedalism [PC17]. bipedally [PC17]. bipolar [HGM15]. bird [ASL^{+}18, BORA10, GZT12, Ush16, vLSBBD11]. birds [AFD^{+}17, BH12, EzMMH18, GWCA14, Hur12, MSBB13, MHSH12, YY18, ZXH19]. Birth [CMMR13, LG10a, SGB19, CHK16, Hal12, Jam12, Jam13, JV14, KMH19, LL18, Par19, PMP13, RAD14, Sta10, SGW^{+}18, WZ15, ZWvG10]. Birth-and-Death [SGB19]. birth-death [Hal12, RAD14, Sta10, SGW^{+}18, WZ15]. birth-processes [Par19]. bisexual [AGM11]. bisigmoid [VLFF12]. bisphosphate [Bar19]. Bistability [Dil14, JCLS^{+}11, JCLS^{+}21, LAH^{+}16, YT12, BKL14, GGG19, HIT18, SC13a, SC13b]. bistable [AKS^{+}19, DD13, Gol10b, JZL13]. bite [SCF^{+}12]. bivalve [Alv18]. black [SJSK18]. bladder [BMGC11a, BMGC11b, MSA^{+}16]. BlaPred [SKK18]. BLAST [DYQ^{+}14]. bleaching [CMGN17]. bleb [WGO^{+}15]. blending [Gar11]. blight [CLK18]. blind [AGM11, GGM12]. blind-mating [GGM12]. blips [OMO13, STA15, WR14]. blister [LK15]. block [ZMNd^{+}10, ZCAB17, HHD^{+}16]. blockers [Ken19]. blocking [NTC^{+}11]. blood [AV19, AOM19, CXWL11, CPS19, CKZ^{+}17, DK13b, DLL^{+}18, EBP15, GDF17, GDF18, GMMN18, HPP10, HII^{+}19, HRHAAA15, HSLW16, KCD11a, KCD11b, LDB^{+}14, MAM16, OPS^{+}19, PIT12, PFJS15, POP12, QF10, SPMGR10, WFM^{+}13, WFC^{+}14, ZCT18]. blood-stage [CKZ^{+}17]. bloodstream [Voe18]. bloom [JSWY19]. bloom-forming [JSWY19]. blooms [Ric17, STI13]. blow [CP14, GGG12]. blow-up [CP14, GGG12]. Blue [TF16]. bluefin [TTW13]. Bluetongue [CSLE11]. blunt [Gha16]. BMP [vHHKB14, ÅCHS19]. boar [CR14]. Board [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o, Ano13p]. Board [Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano13x, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano14x, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h]. Board [Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z]. bodied [RW14b]. bodies [SSS17]. body [Bur10, CAG13, Che16, CLP11, Hor11a, LCHMP16, Mas18, MAFB18, Nij11, RGCML10, WSRG18, ZEJA11]. body-size [CAG13]. Bohr [Oko15, Oko16]. BOLD [DHRA10]. Boltzmann [MAR^{+}17c, WGH^{+}14]. bolus [dLMD^{+}11]. bond [GY19, GMMN18]. bonding [KJH^{+}19]. bonds [GZY12, LF10, MdFDM10, MBC^{+}12a]. Bone [RM17d, VH11b, ASH15, AOM19, BPGS12a, BPGS12b, Bue15, CGvG^{+}15, CLA^{+}16, CGRRGM^{+}16, FGAM12, JTSG14, KLN^{+}12, MH13, NHE^{+}16, PZS^{+}10, POLT12, PVGA12, RHJ11, SSS15, SAB10, TKTB18, Van16b]. bones [BFGAGA16, Mil16]. Boolean [ML12a, BJJR10, DHT19, DRC11, FK18, HXL16, HXL18a, HRCA19, LWH^{+}11, LT11, ML09, MBLV10, MPT19, Ric19, SAA10, SVB^{+}10, VC11, WDN^{+}19]. boosting [DFM^{+}14, LHFM16, RAF^{+}19, XXD^{+}17, ZZN^{+}19]. boosts [Cor18]. bootstrap [KK17b]. border [GM16b]. Bordered [Non10, SR10]. Bordetella [FGH^{+}14]. borehole [MCWF14]. borne [BPR15, BZ10, CBC^{+}18, CG11, DDSDW13, MCRG18, MRPH17, MS12b, MYOS14, MW13, PKZ12, ZLT^{+}19]. Borrelia [ÅCHS19]. both [BBS18b, BHBR12, CW14, Don17, DDS13, EG10, FD18, GSF13, HY13, Lei09, Lei10, Moo14, RTR^{+}13, TH17, XWC11]. bottleneck [DRPM17]. bottlenecks [CW15c, NT14]. bottom [KP12, QA15, SH17]. bottom-up [KP12]. Botts [JLQ^{+}12]. Boubaker [DZJR10, Mil11]. bound [SS17b]. boundaries [CF11, CKNB19, ESS19]. Boundary [WWIG19, YM16, BPS19, CFCM13, DHHP14, Fok12, HSG^{+}18]. Bounded [XWZ10, CA18]. bounding [Sac13, Ush16]. bounds [Lev14, PP12]. bouts [PR17]. bovine [ABB^{+}19, BVJE17, BSR^{+}11, POK^{+}12, SSS15, vLFM^{+}19]. Bowden [HS16]. bowel [WBRHE10, YXH^{+}14]. box [DCL18]. boxes [HHJR11]. BPES [Mil11]. brachiopod [SK11a]. Brain [GLOC10, RPGG^{+}19, AOT^{+}11, ÁTC^{+}14, BBJ^{+}10, BTO14, CCS15, CRGS18, DLHS11, HMM16, IBB^{+}15, KCJ^{+}11, MA11a, MAMEA15, MvAKR17, PJ13, Sat18, SSH^{+}19, SHM^{+}18, SPG^{+}18, ZLL^{+}12]. brainstem [PR17, ZKMB19, ZAL^{+}19]. brake [CST19]. branch [CS15a, SS12a]. branched [LMdL11]. branches [CLK18, MS12a]. branching [AGM11, CMMR13, DDHM12, DI10a, GL18b, GGM12, ID14, IS16, LSMC18, LLS^{+}15, LG19a, MB12b, NKOS11, OBE^{+}17, Par15, POW17, SCsMzX18, SK18, WL14, WK18b, ZCS^{+}15, vE11]. brassinosteroid [AP17]. brave [PP17b]. Brazil [DCR18]. Brazilian [PLF18]. BRCA1 [DM11]. BRCA2 [RMM^{+}16b]. Breadth [All11, CP11]. break [BKL^{+}15]. breakage [KDL16]. breakdown [ABG^{+}13, Zha15]. Breaking [yTRSC13, Len14, SWSLMJ19, SG18, TM15a]. breakup [GMNY14]. breast [BPF^{+}19, DM11, DLM^{+}19, HP16, HZL^{+}13, KPEK14, LLC15, NWB^{+}10, NJP18, OOY16, RTRRS^{+}17, RMM^{+}16b, SER^{+}12]. breath [HH10, KKU^{+}10, Tra16]. breath-hold [HH10, Tra16]. breathing [JB19b]. breeders [MW14]. Breeding [CRLH^{+}19, VDTF15, WK17]. breedings [VCF^{+}19]. Bregman [WcW14]. Bribery [VNS18]. bridge [MCL^{+}11, Sos18]. bridges [GKWG17]. Bridging [OPS^{+}19]. Bringing [Mor10b]. brittle [LMdL11]. Broad [XJ19, HS14c]. broadcast [MDB12b]. broader [WWB14]. broadly [CDK11]. brood [KTO^{+}14, STKE12]. brood/fungi [KTO^{+}14]. brood/fungi-warming [KTO^{+}14]. brown [KSKS13]. Brownian [EPP19]. browse [OSV^{+}16, YXH^{+}14]. Bruijn [WZ17]. bubble [Kue16]. bubble-driven [Kue16]. bubbles [GY18]. Buckling [She10, Van15, GF11, LH13]. bud [BCS^{+}16, LF16]. budding [WKS14]. Budget [TTB^{+}18, DB19, EBS11, Fil10, MP13a, AKNP18]. budgets [WSRG18]. buffer [AHKN10]. build [HRHAAA15]. Building [HSMCCR19, LFZN11, EJ17, EJ19, Pia19]. built [Lüt16]. built-in [Lüt16]. bulk [CEKM^{+}19]. bulk-surface [CEKM^{+}19]. bumble [SOIO10]. bumblebee [XS13]. buoyancy [HN10]. buoyant [Ios16]. burden [Cro19, GAY^{+}11]. burgdorferi [ÅCHS19]. Burkholderia [ADCG14, SSRA16]. burnetii [CMN^{+}11]. Burrows [YZW10]. burst [FR14]. bursting [MVGGB18, PAOMV17, TAORS10, WTZ^{+}11]. bursts [ARG17]. butterfly [LLK14, SFT14]. By-product [DH16, De 17]. Bystander [PMCS16, PSD^{+}18].

C [KB10a, LDB^{+}14, RS19, CM12, DK13a, GN10, KNA^{+}18, LMYL15, LL17b, MVH11, MLT10, Nak12, pNMK^{+}10, RDSD^{+}12, RCL16, SB15, dTKDV15, HWW^{+}14, RFS^{+}15]. C-element [pNMK^{+}10]. c-Met [KL18]. c-Myc [RFS^{+}15]. c-Src [HWW^{+}14]. C. [Sht17, ZLZ^{+}15b]. CA [NTC^{+}11, CFF12, CN12a, DD13, EG10, Gri11, KCZ^{+}19, PRV^{+}14]. CA-CTD [NTC^{+}11]. cadherin [WGS10]. Caenorhabditis [BLZ^{+}19, DI13, OYIM10]. caffeine [PFPR11, RRL^{+}13, RLW^{+}14]. Cajal [MS10a]. calanoid [RL17]. calcification [CL19]. calcifying [RTFP^{+}17]. calcium [AMP12, AHKN10, BTO14, CCF^{+}14, CSM^{+}14, FR13b, GDB^{+}19, IPS14, KLN^{+}12, KSSM11, KSK^{+}16, LSP13, LXZQ14, LO15, MS10a, PSP^{+}12, RG18, SGGHS19, SMZ^{+}17, ST17b, TK10a, WTC10, ZCA^{+}14]. calcium-activated [ZCA^{+}14]. calcium-assisted [KSK^{+}16]. calcium-binding [SGGHS19]. calcium-buffer [AHKN10]. calculate [MB17]. calculated [Mar12]. Calculating [KAKK19]. calculation [Ayd18, lDsWM^{+}15, HXL16, HXL18a, KFG^{+}14, TMS13]. calculations [ZZ18]. calf [SZ15a, SZ16]. Cali [BLV18]. calibrated [METC12]. Calibration [AP13a, HM13]. California [JMK^{+}17]. callus [MFM16]. Calmette [BMGC11a, BMGC11b]. calorimetry [RBB^{+}12]. calpain [NHTS14]. Calsequestrin [TK10a]. CAM [HBVP15]. CAMBIUM [DDB10]. Cambrian [GJ10]. Camelid [SRP16]. camellia [IST11, PWH^{+}13]. CaMKII [HKS15]. CaMKII-NMDAR [HKS15]. camouflage [Abb10, BJB18]. cAMP [Mor10a, SN12]. campaign [NM11]. campaigns [Hou16]. Campus [TXY^{+}12]. Can [GWCA14, Hou15, KDK14a, KDK14b, MN12, NdlPZA10, PFB10, RM17b, RC13, RS14c, Sch19b, SAB10, SWPC^{+}16, BT17, BH13b, BPG^{+}18, BMV12, BIMC17, BBD18, CV15, CLP11, CNM^{+}13, CW14, DZE11, Ezo19, FXML18, FR17, GvVT14, GF11, Gri15b, HBEW12, HMM17, Hor13, HKM12, Jen10, JA13, JEA18, Kis10, KNA^{+}18, Kor18, Kun16, LWLM18, LDA13, LPH19, MGML10, Mar11, MGT17, MYN^{+}15, MG17, MALAN17, MAA18, PDC^{+}17, PPBT11, PGPD19, PTB12, PSL^{+}10, RANO10, RN12, Ran12, RB13, RS12, RL15, Sal15, SWR11, TM15a, TIS10, TWH^{+}13, Uit11, VMZK^{+}19, WFW19, YSYI13, YI16, ZSSM14, ZSO16]. Canada [MPJH13, WDL^{+}13]. Canalicular [KTT^{+}19]. Canalization [LFM14, LP12a]. Canalization-based [LFM14]. Cancer [BST14, Har15, KB15, ZZL^{+}11, AMBH^{+}13, AMFR^{+}16, Ala15, Arc14, Arc16, ARM18, BS15a, BBS18a, BPF^{+}19, BMGC11a, BMGC11b, CWF^{+}16, CCS^{+}16, CLA^{+}16, CBGS18, CvBF18, DDHM12, DL12a, Dim10, DLL^{+}18, DBM^{+}18, DTGC14, DFM^{+}19, DDTL19, DLM^{+}19, EJ17, EJ19, FM10, FM14a, GXFF13, GY19, GLF^{+}15, HP16, HHSA15, HBA10, HA15b, HWPL12, HZL^{+}13, HM13, IHNS16, JMC^{+}10, Kar15, KRDJ15, KPEK14, KP16, LF17, LF19, LDA13, LMW^{+}10, LLC15, LX16, LTZ17, LTZ19, LMHF13, LSS^{+}14, LC16, LXLF14, LDH^{+}14, LD18, MB12a, MM19, MKMG^{+}14, MPZK16, NH19, NGJ^{+}14, NWZ15, NJP18, OOY16, PAS10, PH13, PYG^{+}19, PRN10, PGF11, RMM^{+}16b, SBM^{+}16, SMS17, SFMS16, SCKL15, SOF16, SSN^{+}14, ŚKSRW16, SER^{+}12, TDZ^{+}18, TK10c, VKKA12, VDRL14, WMN18, WCC14, YBH^{+}19, YCL^{+}17, ZWW14, ZMC^{+}18, ZXS^{+}19, dBE16, dSdS13]. cancer-specific [SFMS16]. cancerization [FLR14]. Cancerous [DS15b, RTRRS^{+}17]. cancers [BH13c, DQS^{+}15, LZG^{+}19, OPS^{+}19]. Cancitis [OPS^{+}19]. Candida [HHS^{+}10]. candidate [LTSTD15, PPC^{+}17]. Candidates [JSC^{+}16, CLC11, SPSM15]. canine [HHA17]. cannibalistic [LG19a, LG19b]. cannot [vVLS14]. Canonical [FHM18, Nem17, Sel16, ZGY11]. canopy [PAA11]. cap [WBMM18]. capabilities [ABGM11]. capability [SKSO12]. capacities [YiTS19]. capacity [ASGE14, CCNT19, DN18, KA17, RS14b, VZ19]. capelin [EBS11]. capillary [AMM16, ASGE14, BSMK11, LRA^{+}13, LDB^{+}14, MAM16, SSD11a]. Capitalizing [WMN18]. capripoxvirus [Cha18]. captive [HK17]. capture [DFMR19, GAY^{+}11, JZ17, KA15, Lab16, MNH^{+}12, RB13, VP18]. captures [OSF11]. capturing [ABKS11]. carbohydrate [KF12, MTLW^{+}10]. carbon [AK15, HWGT15a, HWGT15b, HWGT17, LRL^{+}16, LCMC14, MF15, TP14]. carbonate [HWGT15b, RTFP^{+}17]. carboxy [SSZR17]. carboxy-terminal [SSZR17]. carboxylase [WLD^{+}11]. carboxypeptidase [NRS^{+}16]. carcinogenesis [LTZ19, RBKW19, WCC14]. carcinoma [CM12, GCSP17, METC12, MGM13b, NWB^{+}10, SGGY10]. carcinomas [SS17a]. cardiac [CPGF^{+}16, CN12a, CF11, ESG^{+}16, GKMC10, GAPK10, HVL11, KCS^{+}15, LFW^{+}18, LdLK11, LW18, MKMT14, NAK^{+}11, OCN10, PW18b, ST17b, VLCT19, VM16, WMPF^{+}15, WMT16]. cardiomyocytes [HTK^{+}18, KSSM11]. cardiomyopathy [MCL^{+}11]. care [Ram10, Sal15, SRV11, SPRF13]. Cargo [LKZB15, ML12b, LM15, MK17, SL19]. Cargo-mooring [LKZB15]. Caribbean [BH11]. Carlo [KSK^{+}11]. Carnivora [TF15]. carnivorous [BJB18]. carotid [OB15]. carp [KCS^{+}15]. carrier [FKB13, LG18, PWH^{+}13]. carriers [BTK14, Jam11]. carrot [MR12]. carry [PPBD10]. Carrying [VZ19]. cartilage [LNRK11, SABB15, WGS15]. cascade [CCA17, KDG13, MBM11, SK10, ZG10c]. cascades [ANY14, ST16]. Cascading [SL10]. Case [HHA^{+}13, AHD^{+}18, Ant13, ARZ15, BLV18, BR12b, CBC^{+}18, Cle16, EJ17, EJ19, FMS^{+}10, GTC19, HWW^{+}14, HB18, Hou16, KKD18, LBMG16, LWLM18, LS16b, LS10c, MFM16, MRPH17, MPS^{+}11, OO11, OA11, Pav14, PSK10, RCL^{+}10, RP18, RSV10, SSG^{+}19, vV11b, SCS10]. case-control [KKD18]. casei [RZIR16]. casein [SG15a]. Cases [HC17, EAN14, KN11, NI11, Rao12]. Caste [LF13]. castrate [CBGS18]. castration [WMN18]. cat [PFB10]. Cataglyphis [MAW12]. catalysis [HS12]. catalysts [SP16]. catalytic [HS12, KIH16, SMJS14, SLT^{+}18, SK18, WHZ^{+}17]. catastrophe [FS11b, dASdC^{+}15]. catastrophes [PSV17, Sch18, Sch21]. catastrophic [SI10b, WS16]. categories [Men12, SNY^{+}13]. catenin/TCF [SWR11]. Caterpillar [Yam16b]. catheter [Voe18]. cation [KB10b]. cationic [PDC^{+}17]. cats [LZTD18]. cattle [CSLE11, GKG^{+}18, SSZB15]. caudal [Ios16]. caudate [GMK18]. caudate-putaminal [GMK18]. Caught [PBB10]. Caulobacter [SK12a]. causal [GCZ^{+}12, KEG17]. causation [CLG^{+}10, CLG^{+}11]. cause [BIMC17, HBEW12, Jam14, Jam15, MJI14, PSL^{+}10]. caused [BFGS10, CE17, YY18]. causes [CB17a, YSI18]. cava [NPMM15]. cavity [MBC^{+}12b]. CCAAT [HHJR11]. CD16b [OBK^{+}11]. CD200 [LBF13]. CD200-CD200R [LBF13]. CD200R [LBF13]. CD25 [BCPL10]. CD4 [BCPL10, EH17, GML10, LAH^{+}16, RAF^{+}14, RW12, SSRA16, TXWW12]. CD8 [BYY^{+}11, CTL^{+}15]. Cdk [GGG19]. CDK2 [TMS13]. CDPK5 [RM19]. Cell [JAK19, LRA^{+}13, PRN10, STN19a, SI12, Tre19, UI13, AAAV14, ASK17, AGB^{+}14, AUE15, AMBH^{+}13, AMFR^{+}16, AA12, ABV19, AP13a, AKdV^{+}14, AHAFA14, BTO14, BTO15, BJJR10, BGL^{+}19, BJO^{+}16, BYY^{+}11, BLNR15, BK19, BCPG18, BMB^{+}18b, BPGS12a, BPGS12b, CML10, CCS18, CMS^{+}19, CSR^{+}05, CN12a, CWB^{+}17, CCB11b, CLHB11, CNL14, CH16, CLCH10, CK17, CML11, CCdS15, Cox10, CTL^{+}15, CMS16, CKBCG12, CEKM^{+}19, DQY^{+}15, DBBW09, DBBW11, DK13b, DKR16, Dim10, DTGC14, DL12b, DL16, DBJ12, EdKM11, EBE14, EBE17, EUM^{+}16, ESGA15, FWB^{+}12, FP18, FZ14, FLM18, FLGGD^{+}14, FMLM12, FR17, For10, FZL18, GGQ^{+}12, GMOP12, GFM^{+}19, GGG19, GY19, Ghu18, Gol10a, Gol10b, Gol16, GMNY14, GM13, Gon13, GC18, GWW^{+}10, GBIR13, HII^{+}19, HPM^{+}17, HTM16, HHA17, HWPL12, HWGT15a, Hor13, HL19, HMJB15, HBW^{+}11]. cell [ITR^{+}18, JDPK15, JDSPK15, JD13, JB19a, JS15, JS17, JS18, JSB15, KBV12, KSP12, KBL15, KB11, KL11, KRR14, KTH16, KFG^{+}14, KS16, KCD11a, KCD11b, KKM12, KAKK19, Kro10, KSPA17, KBF18, LLTP19, LCTG15, LDF^{+}11a, LFZN11, LWH^{+}11, LAG^{+}14, LS15b, Lee16a, Lee16b, LMC^{+}13, LSMP14, LL17a, LFB^{+}16, LSY^{+}10, LL14, LXS15, LPTR14, LC16, LGPS17, LNRK11, MMJB17, MRF19, MH11, Mal10, Mal18, MOSS15, MMFK10, MEKK11, MBLV10, MSM^{+}14, MYC12, MPBS17, MBS19, MFG14, MGM13b, MPZK16, MKJS13, MDD13, MT15, MBK^{+}11, MLL^{+}16, MG15, NCLB16, Nic19, NTOI16, NKOS11, OKVN18, OS11, OOY16, OTGT10, PFJS15, PW19, Pie10, Pie12, PWE15, PGPD19, PGHC12, PAV10, RCH14, RJSC18, RAF^{+}14, RD12, RC13, SI11a, SDD15, SMM15, SA17, ST16, SKGM19, SDT17, SHB^{+}17]. cell [Sch19c, SP13a, SFMS16, SVB^{+}10, SAGAGB17, SLM17, SZ18, SMZ^{+}17, SK12b, SOF16, SSKS18, SSKS20, SZ10, SLC12, SLvdBMP10, SB19, SRAL12, SQZ^{+}16, SSRA16, SN17, TKK14, TIMI^{+}15, TGB^{+}18, TXWW12, TDZ^{+}18, TC11, TST^{+}13, TDSM12, TALC16, TT17, TSMB14, TCYY^{+}12, TK10c, VKKA12, VG13, VB19a, VBD10, WGS10, WMPF^{+}15, WRH^{+}16, Wal12, WML16, WGO^{+}15, WBB^{+}18, XJ19, YM14, YM16, YAK17, YCB16, YFB^{+}12, YLLL12, ZCT18, ZS12, ZMS17, ZWW14, ZZL^{+}11, Zhu11, ZBA18, ZZS19]. cell-adhesion [CCB11b]. cell-based [KBF18, LC16, PFJS15, VG13]. cell-controlled [Kro10]. cell-cycle [HWPL12, PGHC12]. cell-division [HTM16, LAG^{+}14]. cell-fate [LLTP19, MBK^{+}11]. cell-free [ITR^{+}18]. cell-matrix [ESGA15]. cell-mediated [ABV19]. Cell-to-cell [UI13, CH16, WMPF^{+}15]. Cells [LRA^{+}13, ABR11, Arc14, Arc16, BCPL10, BFH^{+}15, BTO15, BBS18a, BZJP18, BHBH11, BZL17, BCPM^{+}16, BPF^{+}19, BKPV15, CFF12, Di 18, Di 19c, DLL^{+}18, DBM^{+}18, DLM^{+}19, EG10, EH17, EME^{+}16, EUM^{+}16, Fer12, For10, FM15b, GML10, GMOP12, GHJ14, GMBK14, GLF^{+}15, HPM^{+}17, HVL11, HMJB15, JMC^{+}10, Kar15, KKO^{+}18, KHNM16, KPEK14, KAKK19, KP16, KC13, Lab16, LP17, LAH^{+}16, LCA^{+}15, LLDW14, LCGMH12, LSS^{+}14, MVGGB18, Mar17a, MB18b, MS10a, MSA^{+}16, MH14, NTOI16, NWZ15, RD14, RHS14a, RFdL15, Rou14, RW12, SI10a, SDD15, STNT17, SNY^{+}17, Sat17, STH18, SPMGR10, SSM^{+}18, SZ15b, ŚKSRW16, TvMG16, TWTA^{+}18, TRM^{+}14, WGS10, Wod18, ZM10, ZFWK17, Zhe16, ZMC^{+}18, d’O12, dSdS13]. Cellular [CAGM^{+}17, LL14, RSV10, ARG17, BK19, BCPG18, CCS15, CN10, CHL^{+}11, CLCH10, Di 18, Di 19c, FRP14, GMOP12, GGG19, GN10, HCK^{+}11, HGO^{+}18, HWGT15a, HZW^{+}14, JSDEK14, LS16a, LSP^{+}17, MSN14, MMAS13, MMLK11, NLB14, PAS10, POP^{+}19, PGHC12, RFME^{+}12, SS19, SHB^{+}17, SBSR13, SNC12, SZ18, TM15a, TvMG16, THC^{+}11, VLPH17, VBD10, Wod18, XMWC13, YCH^{+}17a, YCH^{+}17b, YMW12, ZZS19, dlCGSA16]. cellularity [SGS15]. censoring [NI11]. center [RS19]. centers [GA13, YS14]. centipede [RN12]. Central [Voe18, BRA15, FKB^{+}12]. Centralities [ER18]. centrality [EG12, Est10, FB12, KJSS10, Mal18, YBH^{+}19]. centric [VGPS18]. Centripetal [CFSB16]. centroid [GGCJ16]. centroid-based [GGCJ16]. cerebellar [ML11a]. cerebral [CPS19, EBP15, FHR13, KPD19, LDB^{+}14, MTdS^{+}16, MP19, NF14a, NGL^{+}10, ZKP15]. cerevisiae [AM12, BLP10, HZW^{+}14, LLCC12, LXC15, OB10, SSJK18, TK10b]. cerotegument [FWSG17]. certain [KDS18]. certification [Ala15]. CFD [JG14]. CFSBoost [RAF^{+}19]. CFTR [Mor10a]. cGAS [GSS19]. CGH [SS17a]. CGR [JLQ^{+}19]. Chad [LLN^{+}19]. Chagas [AZOLVH18, MMG^{+}16, RS14c]. chain [CJKR10, CHH10, CSB15, GK13, HB13a, KPG19, LYF^{+}15, MCL^{+}11, SV14, XDC^{+}11]. chains [HM11b, Jia16, LFZN11, MGML10, PBKR13, SI10b, UBP12, ZLW16]. chalcone [TAR16]. challenge [KMM16, Ser16, SLM17]. challenged [SBCR10]. challenges [GE18, KCE^{+}11, VFS^{+}15]. chance [Sch19b]. change [Abr14, AT10, CW11a, DFM^{+}14, DLSD15, DPvBvA12, JSZ10, KS13, KL11, KC17, MMC19b, PSV17, PW19, STJG12, SYI17, SGW17, Spe15, YSST13, Yam16a, YI18b]. Changes [PHTP^{+}12, AJC12, BVK10, CCB11b, EBSW17, FGGT15, HKH18, LXJ15, LCDH15, OVKL14, PDNP16, Rog19, STNT17, SLD^{+}17, STLJ18, TK10a, Wil11, vV11b]. changing [CMCS18, GW15, GJ12, GA16, GWCA14, MBRRI19, SBMH10, SRS18, WSVT14]. channel [BGCB12, GJ11, HZC^{+}10, KH19, LdLK11, MFZ18, MAA^{+}17, Mor10a, PDG17, XMWC13, ZCA^{+}14, ZZG^{+}16]. channel-drug [XMWC13]. channel-targeted [ZZG^{+}16]. channelopathy [PBS^{+}12]. channels [Bec14, CPGF^{+}16, CFF12, CW11c, CSM^{+}14, DM15, FD17, KB10b, LSP13, LD11]. Chaos [Kea16, BA19, GW13, GP11b, JLQ^{+}19, JC16, JSF^{+}11, PSL^{+}10, SCS10, SY12, TF16, WH11, XSLZ16, ZKHL16, ZG10c]. Chaotic [STG17, DGJ15, LDA13, MGM13a, MGM13b, STJG12, TLPH11]. chaperone [XKCG15]. character [CJ12a, Cut15, MC15b]. characterisation [DKP^{+}18]. Characteristic [ESS19, MTNM12, HgLL^{+}10]. characteristics [DS15b, DLYZ11, DDF^{+}14, HCK^{+}11, JYY^{+}18, Kon11a, LBS^{+}10, LBS^{+}11, LHD^{+}17, LYZ^{+}18, Rey13, ZLL17]. Characterization [ABN15, LL13a, OCN10, YWZ^{+}16, ASRM15, CA17, GRB^{+}13, HZLX11, LDH^{+}12, LCQ^{+}18, MdFDM10, MAMEA15, PDG17, PPC^{+}17, PPT^{+}16, QW12, RMM^{+}16b, WdLS17, WHL13, YLL^{+}14, pZWZ^{+}16]. characterize [JK12, RKJ^{+}11]. Characterizing [BPF15a, BPF15b, CWF^{+}16, SNY^{+}13]. characters [Cut15, GJ12, PRSC11, RCD16]. Charcot [AKR^{+}18]. Charge [DK13c, YS14]. charged [Smi11]. Chaube [Pan19]. Cheater [MV18, RD14]. Cheater-altruist [MV18]. cheaters [SKS^{+}19, SBS17]. cheating [UI13]. check [BBT^{+}15]. checkpoint [KBE^{+}13]. chelation [FE10]. Chemical [Wil11, Bau18, Bro13, Di 16, Di 17b, FL12, FL13, KS19, KF15, MF15, NGN13, PAK11, SHH15, SSV^{+}15, UI13, WLY^{+}17, ZS10]. chemicals [FSA15]. chemistry [RTFP^{+}17, Sza15b, dlEBRM15]. chemo [MDD13]. chemoactive [TVA18]. chemoimmunotherapy [RTEKG15, RTEKG19b]. chemoinformatic [HAuR^{+}18]. chemokinetic [SN17]. chemopreventive [TVA18]. chemoprophylaxis [EN15]. chemorepulsion [d’O12]. chemorepulsive [ANK10]. chemostat [FE11, LFM11, TWW19]. chemostat-like [LFM11]. chemotactic [CFT11, MOSS15, NDG14, RG10, SN17, WST15]. chemotaxis [Amo15, BPF^{+}19, FZ14, NvD17, NS18, PHH18, Pai19]. chemotaxis-based [BPF^{+}19]. chemotherapeutic [GSSBF18, HNP18, NSH^{+}10].

Chemotherapy [dG10b, BBJ^{+}10, BPLM12, Car17, CL18, CHN^{+}15, CvBF18, EJ17, EJ19, HL18, IBB^{+}15, LF19, LBW^{+}13, LIB^{+}17, PGHC12, RTEKG15, RTEKG19b, SGL10, SKD^{+}10, WHH17, ZLM12]. chemotherapy-induced [CL18]. chemoton [Gri15a]. chemotrophs [SI19]. Chi [Xie18]. chicken [SBSR13]. Chief [Kir11]. chikungunya [MHX^{+}14, MP13c]. child [FM18]. childhood [Nås12, NSH^{+}10]. children [BSR14]. chimpanzees [Sch14]. China [Kur14, ZZR13, GM17, HJR12, LZY^{+}15, TXY^{+}12, TXTW16, WXWL12, WZJ^{+}13, XTZ^{+}13, XSTW14, ZLZ15a, ZJY19, ZZR10, ZRZ15]. chiral [OA15]. chirality [OA15]. Chiroptera [AK11]. chitinase [KGP^{+}15]. Chlamydomonas [SLM17]. chloride [BHBH11, Mor10a, ZCA^{+}14]. chlorophenol [WPPD16]. chlorophenol-mineralising [WPPD16]. chlorophyll [GZT15, Laz13, MTE15]. chloroplasts [MB14, MSC10]. CHO [GHJ14]. choice [AGM11, CS14a, CC11b, Cro19, DB19, Gal10, GK10, HLI15, MI11b, MI11c, NPS10, TI12a, WSW10, YLH12]. choices [JRB^{+}16]. cholera [ERT13, HWGW18, LKO^{+}17]. cholerae [FBK11, HK14, SDK11]. cholesterol [BSM^{+}14, PCS^{+}18]. chondrocyte [CAGM^{+}17, CWP^{+}18]. chondrocytic [ABD^{+}11]. chondrogenesis [CWP^{+}18]. chondrotin [SABB15]. choosy [MG17]. Chou [Mei13a, Mei13b, AFS19, AH19, AH18, AH15, AHJ18b, BMN16, BRK19, CL13a, CCY^{+}19, CT18b, DHS^{+}15, EMM10, FL12, FL13, FLW16, HPB^{+}14, HYA14, HY13, HKR^{+}19, JD16b, JD17, JCG15, JCG16, KHK15, KHKI17, KZL14, Kri18, KSKK15, LZ18, Mei12a, Mei12b, MFZ18, MZ18, MBE11, MP14b, NBL14, NMZ19, PWZ^{+}19, QLC^{+}18, RBMS17, RSS^{+}18, SIK^{+}18, SRS^{+}15, SM18, STG19, SKK18, TTC19, TWC^{+}19, TP17, WMK13, WZM19, XNJ^{+}13, XSLZ16, YGL^{+}10, ZSZM14, ZZK14, ZK18, ZL18a, ZD18]. Chromatid [Fer12]. Chromatin [ASP13, Mic11, PDNP16, PSK10, SI12]. chromosomal [SP14, ZS12]. Chromosome [JS15, AA16, CL17b, DKP^{+}18, MZWC10, MK16a, PCB14, SK12a, SM11b, Unc11]. chromosomes [MYK17, SK10]. chronic [KEG17, LLDW14, LMYL15, NS11, OVKL14, PGPD19, RTR^{+}13, TBMM19]. chronological [Yan10a]. chrysomelid [FKMG15]. chytridiomycosis [JB11]. CICR [OTTF11]. cilia [KBL15]. ciliate [BKC^{+}16]. cilium [ABD^{+}11]. circadian [CLW12, DFT^{+}17, DB18, EBE14, EBE17, GM13, HBVP15, JC10, KTO^{+}14, LG13, LW12, OHWS18, RGG12, SCLA10, TAL19, VH11a, XCW^{+}18, Zha10b]. circRNA [Sos18]. circuit [FK18]. circuit-preserving [FK18]. circuitry [RM17a]. circuits [BPP^{+}16, DEN^{+}11, HMTA17, LPTR14, TK19]. Circular [EM16, GGR11, Sos18, AFM10, CC10, CM17, FGGS15, FS15b, FZL18, LLDW14, MP13b, Mic15b, MPP16b]. circulating [SMM^{+}13]. circulation [Hou15]. circumstances [HPvdB17]. cis [SDRA^{+}15]. cis-regulatory [SDRA^{+}15]. cisplatin [MB12a]. cities [LCCK12, ZZL19]. citric [GT11a]. city [MYOS14]. civilization [FBU11]. Cl [CZT^{+}16]. clade [ADR11a]. clades [HMM11]. clamp [HYZS14]. clamp/relaxation [KFG^{+}14]. clams [Alv18]. clarified [Hor17]. clarity [GKNT10]. class [BT17, BvLH14, Gra15, JSZ12, KHX^{+}19, KSKK15, Lio18, Lio24, LJ10, MI11c, NRKE18, NK18, RSD^{+}16, ZM10, ZLY14, ZZK14, ZKHL16]. class-structured [Gra15, Lio18, Lio24]. classes [CT18b, DYQ^{+}14, FDS13, HTK14, KZL14, MPCTGJ^{+}15, NBL14, OYY16, PT10, VK10, WCC13]. classic [ZLL17]. classical [Cle16, HS14a, Hua12]. Classification [AH15, BWS10, FLCS^{+}15, GM16b, AH19, AC15, BBP13, FBAPMD13, FLGGD^{+}14, GXFF13, GMM15, lGLLL14, GGCJ16, HRD14, JSZ12, JWS^{+}10, JS17, JS18, KHX^{+}19, MdFDM10, MP12, NBL10, NSBL10, RMSTG13, RKMG12, Tal12, VZB^{+}15, YGMT12]. classification-based [JWS^{+}10]. classifications [HYA14, XM11]. classifier [DSPM14, JLX^{+}15, WYX^{+}17, XWC11]. classifiers [BMSEE14, HRHAAA15, SM17a, XWD^{+}10]. Classify [Kri16, TP17]. classifying [SKK18]. clavaminic [RMRC^{+}16]. clavulanic [RMRC^{+}16]. cleansing [JG14]. clearance [KBV^{+}15, MG15, SBMH10]. cleavage [ATK10, Sat19]. click [TB11, WL10]. Climate [FBU11, AT10, BTK14, DPvBvA12, PSV17]. climatic [GS13a, GS14]. climb [GF11]. Climbing [GW12]. cline [IST11]. clinical [BC19, DH18, Dim10, FBFM12, METC12, PT10, PZLF19, PMCS16, SKD^{+}10, YD15]. cliques [LCL14]. clock [DFT^{+}17, EBE14, EBE17, LG13, MMB11, RGG12, TAL19, XCW^{+}18]. clock-cell [EBE14]. clocks [TCH14]. clonal [CMB^{+}12, FR17, HM10a, Ros15, RW12, ST17a]. Cloning [KGP^{+}15]. clonotypes [LCHMP16]. close [BWP10, Di 12]. close-facing [BWP10]. closed [CSB15, LLDW14, MH13]. closely [MMH^{+}12]. Clostridium [JEA18]. Closure [CLG^{+}10, CLG^{+}11, ABH^{+}11, FS16c, LCCB11, MM15b, PCBMM12, SYR11]. closures [PHK15]. clot [SHK14]. clouds [SW18]. clubs [TLW18]. clumped [BHKR14]. cluster [HLHY17, HYZ^{+}15]. Clustered [Don16, HK11b, MS12c, RBHK14, SBK16, TWTA^{+}18]. Clustering [HRC^{+}12, HY16, YFB^{+}12, GMK18, HASM17, KUV^{+}10, PWHW16, RD14, SAG19, SKH17, TP17, VZ19, WCSA12, WZ17, YCY14, ZYJL18, ZDY11]. clusters [DMSW10, GMNY14, KRR14, LL14]. CML [RCH14]. Co [Ant13, AMFL10, BB11b, Cam11, PBA12, SPS11, ADV^{+}10, CWP^{+}18, Dim17, DL16, EJK16, HNV^{+}16, HPM^{+}17, HAH19, HgLL^{+}10, LZG^{+}19, LDH^{+}14, MG14b, MBLC17, MALAN17, MH18, NZZ19, PDB^{+}15, PNP^{+}16, PDNP16, PGLZ14, PDW10, RANO10, SS18, SCKL15, SST19, SOBC12]. co-activators [NZZ19]. co-colonization [PDW10]. co-culture [HPM^{+}17]. co-cultures [CWP^{+}18]. co-dependence [DL16]. Co-dominance [Cam11]. Co-evolution [Ant13, AMFL10, PBA12, SPS11, MALAN17, MH18, RANO10]. co-existence [ADV^{+}10, MG14b]. Co-existent [BB11b]. co-existing [EJK16]. co-expression [LDH^{+}14]. co-flowering [MBLC17]. co-infection [HAH19]. co-localization [LZG^{+}19]. co-morbid [SST19]. co-occurrence [PDB^{+}15, PNP^{+}16, PDNP16]. co-optimization [SS18]. co-option [HNV^{+}16]. co-regulated [HgLL^{+}10]. co-regulation [PGLZ14, SCKL15]. co-transporter [Dim17]. Coaction [vDRT14]. coagulation [PFJS15, TSP15]. coalescent [ADR11a, CK15, GK16, Gre15, PP17a]. coalescent-based [GK16]. coalition [MGGGA11, NOS17]. coarse [KSK^{+}11, KG15, Men12]. coarse-grained [KG15]. coastal [SH17]. Coated [SPA18]. cocaine [SNY^{+}13]. coccolithophores [HWGT17]. code [AFM10, BWM^{+}19, Boz15, CC10, CL11, DS19b, Di 13a, Di 14a, DMCP14, Di 16, Di 17a, Di 17b, EM16, FD18, FDS13, aHLZ^{+}12, KA17, Len14, Mas16, MP13b, Mic15b, Nah14, Nem17, Rog19, Rou14, SM10a, Sel12b, Woh15]. coded [BBT^{+}15, Cle10, PBKR13, Sel12a, Sel17]. codes [BWM^{+}19, FGGS15, FS15b, FS16b, GGR11, Mas16, MP13b, MPP16b]. codified [Di 14a]. codifying [Di 14a]. coding [CK16, CJ12b, DHBS19, MZWC10, MWSM13, MP13b, Mic14, Mic15a, SI12, Sel13, SCLC13, Tak17a, TP10b, YCR^{+}15, ZTWL12, ZLB13]. Codon [Sel15a, DS19b, DCL18, KB10a, KK12b, LY14, LDWW14, NS13, PWH^{+}13, PMYHR12, QW11, RGP13, Woh15]. codon-amino [DS19b]. codons [KK12b, Rog19, Sel17]. coefficient [Flo11, LDB^{+}14, ZYJL18]. coefficients [AS16, ZYJL18]. coeruleus [PR17]. Coevolution [LG19a, Nak16b, TI13, TCB13, Wei17, ZLW18, AIB^{+}19, Di 16, Di 17a, Di 17b, DN18, Rou14, Saf13, SK11b, SA14, WFZW13, YYB^{+}19]. coevolutionary [MI11a, Van16a]. coexist [PFB10, YSYI13]. Coexistence [CG10, EI13, KK13, OAS10, PC10a, SvK12, SSC13, SM10b, AdlPLMZ13, BT17, BL15b, BLS^{+}17b, CH11, CMS16, GVSLG16, GCS12, GW19, HSM17, IC11, JCW13, Kar12, Kři13, KR19, LS15a, LRGA13, MGT17, NvHM13, NdlPZA10, PML18, SX12, SK16b, SKS^{+}19, TZ11, TSB^{+}17, VLCM10, VGL16, XF13, YLW^{+}14, ZLY^{+}17]. cofactors [RCL^{+}10]. cognition [GS17a]. cognitive [Gau11, Kuz19, RRL^{+}13, Wal18]. coherence [BDLR12, LV17, LCL14]. coherent [MDVT10, SKF16]. cohesion [CMBP19]. cohesive [Woo10]. cohesiveness [NTY16]. coil [MALAN17]. coiled [MALAN17]. coiling [Nos14]. coinfection [ITR^{+}18, TCW14, XSMF12]. coinfections [PGPD19]. cold [Kur14, SGAM12, SSV^{+}15]. Coleopteran [LBS^{+}10, LBS^{+}11]. coli [ABK^{+}12, ARZ15, BRG^{+}12, BLP10, HVSZ10, KF12, MBE10, MMH^{+}12, RG10, Tak17a, ZS12]. colimited [BBS10]. colitis [LMHF13]. colitis-associated [LMHF13]. collaborate [EM11]. collagen [HB12a, Kro10, KF10, MHH13, SDKF12]. collagenous [Kro10]. collapse [BIMC17, FBU11, MJI14, VWF19]. collateral [LH13]. collecting [KFG15]. Collective [CLA17, Str11, Arc14, BFW10, BBP13, CB16, Cox10, Cro19, DB10, Eft13, FD11b, FR17, Kea16, MPBS17, MPNP12, NCLB16, PNL15, PN16, RLK10, UPWK15, XLSF19, ZLB^{+}18]. collectively [Luk14]. collinearity [GG16]. colloidal [FWSG17]. Colombia [BLV18]. colon [LPP^{+}19, LMHF13]. colonic [FBC12, MFMB12, MTLW^{+}10, Wal12, WBRHE10, YAK17]. colonies [BIMC17, CFR13, CCdS15, LDHD14, MKJS13, SK12b, UPWK15].

Colonization [FR13a, PR13, BLNR15, PDW10, RL17, SPRF13, SDK11]. colonizing [NGS^{+}16]. colony [AC12, GK17, GS13b, GSF13, LDF^{+}11a, MSB16, RBSD10, STN^{+}19b]. colorectal [LF17, SS17a, SBM^{+}16]. Colored [YGMT12, JC10]. colour [FR13a]. columns [MYLK11]. COM [BAM^{+}15]. combat [Bos12, JMH^{+}19]. Combination [PDF18, YBC17, AH15, CvBF18, LF19, QW11, SYY17a, SYY17b]. combination-chemotherapy [CvBF18]. combinations [JSDEK14, XXD^{+}17]. combinatorial [SW16]. combinatorics [LG17, SFV16]. Combined [GK10, KJM17, AMM16, Ala15, BMGC11a, BMGC11b, DGM15, KSM^{+}15, LBW^{+}13, STI13, TDKJ15]. Combining [SSJK18, YLF^{+}17, HZL^{+}11, LTL^{+}15, NAK^{+}11, TP17, WYL^{+}19, ZLW^{+}19]. combs [MKJS13]. come [Ano19-27]. coming [TTN13]. comma [FS16b]. comma-free [FS16b]. Commensal [BL15b]. Comment [Ghu18, HS16, JMS12, Mac11, Mar11, Pan19, RRC^{+}11, Bal13, DHK13, NBW11, vV11b]. commercial [ZLZ^{+}11]. common [Cam15, De 17, FS12, GZ19, KP12, LJI17, LLY15, RBRB16, SK16a, SS15e, TLW18]. common-pool [LJI17, TLW18]. commons [MB18c, VSP11]. communication [CDD12, DRW14, EG10, Eft13, EBSW17, LLSO13, OO18, WK12]. communicative [SA13]. Communities [DWM15, Cor16, CG10, DSA^{+}16, FXML18, FB18a, JHE15a, JHE15b, KDMK16, MRPH17, MP13a, MH18, NTFK11, PSJ15, Pav19, PC13, PFB10, RK18, RR12, SL10, SM17b, SKS15, SRS18, SBS17, VA10, VGL16, WCPF15, ZADB15]. community [BZ10, CTA15, DBD15, DS19a, FGPR10, GL12a, Gri19, HL11, HLTW14, Jab10, KCS16, LJW^{+}16, MK14d, MI11c, MSIR10, PDW10, PDW11, Soz13, Spe15, TBM^{+}13, TLCZ12, WSTL16, XLSF19, YISG14, ZMT11]. community-acquired [PDW10]. community-structured [KCS16, TLCZ12]. commuters [MYOS14]. commuting [BLV18, LCJ18]. Compact [GLOC10, SBK16, WMK14, Wax11a]. Comparative [AMM16, GP11a, GRRG16, SPSM15, TAR16, ALH10, BPM^{+}12, DGD^{+}11, ESE15, GTS15, JSZ12, KKD18, KTT^{+}19, MH12, RSI11, SG15a, TF15]. compare [KCJ^{+}11]. Comparing [DGMY18, Dra19, LLS^{+}15, MFMPS10, MHX^{+}14, MSM^{+}14, WF17, DS10b, FHW^{+}10]. Comparison [AD16, CPGF^{+}16, HHSA15, HL18, LS16b, Wax11b, Zha11, ALH16, BMN16, Boc12, BL14, BWP10, DLYZ11, DGL12, DYS^{+}13, Di 19a, EB15, GAB14, HRD14, HZLX11, LAG^{+}14, LLL^{+}17, LGR^{+}12, LWL^{+}11, MHKA16, Pee12, RCD16, SMG18, SDRA^{+}15, TMD14, UK17, WP17, WZY14, YZZ12, YZZ13, YW13b, KA16]. comparisons [Di 13a, WS10]. Compartment [LO15, MEKK11, MHMM11, PBvdG10]. compartmental [BvLH14, LCJ16, LSSG10, MMFK10, MPC12, cSGFB17, TF18, TW12]. compartmentalisation [SLS17]. compartmentalization [Nak12, NCS^{+}13]. compartments [KS15]. compass [Jen10]. compatibility [LP14]. compensation [Das18]. compensatory [HWMT17, IIKT13, KIT^{+}16, KII19]. Competence [BPFR16]. competency [BF15]. competing [BWY^{+}17, Gol10b, IC11, LKP^{+}12, PP17b]. Competition [DDS13, GW14, JA13, Kři14, MZ17, NdlPZA10, NBA^{+}18, Pel18, PB16, RK18, SAB17, SK16b, ACMK12, BI12, BCF^{+}16a, CTB18, CKNB19, CP11, DHV19, DP13b, DKF17, FE10, GSRR17, GVSLG16, GBM18, HGM15, HP12, Han12, IGL^{+}12, ID19, KB15, Kar12, KSP12, KM19, LTHEK12, LSDD13, MBLC17, MSS10, MK14d, Nak16b, NT14, NTOI16, Oku15, PIPB10, PBB^{+}15b, PB18, PC10a, PDF18, RCH14, SK12b, Soz13, Str15, TCR13, US10, VY12, WZ18, WW19, WMN18, WHYMG17, YK11, ZADB15, dlCGSA16]. competition-antagonism-mutualism [MK14d]. competition-diffusion [KM19]. competition-facilitation [TCR13]. Competitive [Bar11, KKGN16, LS15a, SM14, XF13, CCA17, DZE11, FM14b, GL12a, Kar15, LJW^{+}16, MPP^{+}18, PBBB10, PDW10, RR12, SLvdBMP10, UT17, WMN18, YZY^{+}16]. competitor [EUM^{+}16, SGCD14]. competitors [PFB10]. complement [RG17]. complementarity [PBP15, SL10]. Complementary [MGS16, Mic15b, Sel12a]. complementarysynthesis [FK13]. complementation [SE10]. complete [JOAN14, MKBE17b, NIiT18a, OAP14, QW11]. completion [ESS19]. Complex [ABKS11, ASL^{+}18, KMD^{+}12, RPD14, TRJD19, VLCT19, ASK17, APBS15, AP19, AD13, Alv19, AN10, BPFR16, BEK10, BJ17, CRLH^{+}19, CL14, CW15a, Est10, EJ16, FWLW11, GZT12, GSCS11, HGM15, HKS15, HB10, IMA16, KOF^{+}14, KHB^{+}18, Kon11c, LJM15, LMC^{+}13, LYZ^{+}18, MK11, MTE15, MD16, MBP16, NW10, OA15, OYY16, Pła10, PW14, RCH14, RFME^{+}12, RDP16, SK19, SSH^{+}19, SAB17, TBM^{+}13, TK19, VL11, VNS18, WMCL18, Xie18, XSMF12, YLWZ10, YZ19, ZZC10, ZC14]. Complex-linear [KMD^{+}12]. complex-mediated [AN10]. complexation [BE14]. complexes [ACvKA10, DFM^{+}19, ER18, HNV^{+}16, HMMSRSD15, NGN13, NZZ19, PWHW16]. complexification [RAD14]. Complexity [AdGM12, PCC16, PDG12, AMFL10, BL15a, BORA10, CS15b, DS10b, FB18a, LXJ15, MGO^{+}15, MSS10, NBS^{+}13, PP12, TCH14, TAL19, THM10, VBHM^{+}13, YHY14, ZLY14, ZGW16, ZLL^{+}12]. Complexity-stability [PDG12]. Compliance [Sar10, ASC16, CPS19, FKK14]. Compliant [MS14b]. component [HME12, OUMA10, SG15a]. components [AH19, BGM19, GLR^{+}18, NMZ19, PWZ^{+}19, TTC19, TWC^{+}19, ZK18]. composite [FK16, HK11a, Rey13, SIK^{+}18, SHLL11]. Composition [AH15, AHJ18b, Bau18, BMN16, CL13a, Cho11, EUM^{+}16, EMM10, FL12, FL13, FLW16, GZ14a, GZ14b, Gri11, HPB^{+}14, HK11a, HWGT15b, HYW11, JSF^{+}11, KHK15, KHKI17, KZL14, KSKK15, LD11, MJ11, MFZ18, MZ18, MBE11, MP14b, NBL14, PDC^{+}17, QLC^{+}18, RBMS17, RW12, RRB10, SRS^{+}15, TP17, Tun13, WMK13, XNJ^{+}13, XSLZ16, YGL^{+}10, ZSZM14, ZZK14, ZNA^{+}16]. compositional [ML14]. compositions [HY13, JD16a, JD16b, JD17, PWZ^{+}19, ZLW^{+}19, ZLDZ13]. compounds [GDPPSS^{+}11]. Comprehensive [JSZ12, ASK17, KTJ19, OCN10, SLW^{+}18a]. compressibility [LL14]. compression [BNG16, FKDW15]. compressive [ZPHS11]. comprising [GI15]. compromised [CWM11, WCM15]. compulsive [RM17a]. computability [CLG^{+}10, CLG^{+}11]. Computation [BR15, ASP13, KBF18, PSK10, Wal12]. Computational [BZJP18, BLP10, DQY^{+}15, Deu18, EKvdKvFK13, FBK11, JB19a, KIH16, MMY^{+}12, MAM16, MAA^{+}17, PS13, PWC^{+}15, SBT^{+}18, SBMH10, TLT^{+}15, TSF^{+}19, VB19a, AKR^{+}18, AMBH^{+}13, ÁTC^{+}14, BTH^{+}13, BZ18, BFGAGA16, CCS15, CRGS18, CS11b, Fre10, GA13, GT15, HFT15, HN10, HHF11, IHNS16, IGL^{+}12, JAK19, KHK^{+}17, Kha10, KBL15, Kit10, KKG^{+}14, LL17a, LCQ^{+}18, LYZ^{+}18, LHL^{+}14, MFN^{+}18, MPJH13, MFMS^{+}18, MV10, MSRW16, MDD13, NGL^{+}10, OB15, OTTF11, PWC^{+}12, PCL^{+}15, QZ14, RPGG^{+}19, SSD11a, SSF15, SPG^{+}18, SZ10, SLC12, SLD^{+}17, Str13, SA13, SYY17a, SYY17b, SCT12, SST19, VPFV13, WHZ^{+}17, ZGY11, ZGW16, ZZL^{+}11, vHHKB14]. Computationally [Ros10b, RMRC^{+}16]. computationally-assisted [RMRC^{+}16]. compute [BMV12]. computed [CZPC^{+}18, MIH16]. Computer [LSP13, MSN14, MZ17, SRH19]. computers [Chu18]. Computing [RTFP^{+}17, CA17, FMS^{+}12, FVC15, ZSW^{+}14]. concentrated [BH13a]. concentration [FT12, GT14, HHA17, LS16c, MKRE18, Ou19, SWR11, SSKK17, VSW10, ZGS^{+}10, ZAL^{+}19]. concentrations [Bur10, FVTS16, LF17, Str17]. concentric [GAPK10]. concept [BB13, CKS15, EMM10, JSF^{+}11, MBE11, XNJ^{+}13, ZSZM14]. conceptual [HB18, O’M17]. concerning [RB15]. concerted [WLD^{+}11]. Concurrent [BCGD10, CHN^{+}15]. condition [GPT17, Kon11c, OBAI11, POW17, PTNKT11, Tox18]. Conditional [IIVR10, CW11b, GCS12, MBBD13, PMD11, RPCW18, RQB17, SP13b]. conditioned [WN19, ZLW14, ZLW16]. Conditions [BdABY13, EA17, FVTS16, RCL16, AK11, BDR10, CDO17, GT11a, GGM12, HSG^{+}18, HHT^{+}19, HS14c, KGM^{+}15a, LRL^{+}16, LDdA^{+}13, Lug15, LS11b, Nås12, NNvdB17, OFT15, OAS10, RGBR17, SI19, TL15, Tos11, ZG19]. conductance [BBS18b, For10, OTTF11]. Conducting [LSMP14]. conduction [Hal16, WMT15, WMT16]. conductivity [KFG^{+}14, LS15c, Tak17a]. cone [KTN^{+}10, RBB^{+}12]. confers [OTTF11]. confidence [EJ17, EJ19, KSA16, KK17b, TDW16]. Configuration [Den19, HBK12, RMRC^{+}16]. configurations [MAR^{+}17c, TG10]. confined [BP15, Lab16, RW14b]. Confinement [MH12]. confirm [Di 13a]. confirms [Sel17]. conflict [CHK16, For14, KM12, MJI14, Sak10]. conflicting [KOS13]. conflicts [Fis19, Sek12]. confluence [JSP^{+}16]. Confocal [KKO^{+}18]. Conformal [Mit16]. Conformation [She11b, DHB15]. Conformational [MdFDM10, BC19, CVPCV^{+}15, CT18b, SG15a, SQZ^{+}16, WHL13]. conformers [Mau15]. conformist [TSL12]. conformists [KL13]. Conformity [WL15, KL13, Van16a]. confusingly [HC17]. confusion [Hor17]. congestion [De 19]. Congo [OL19]. conifers [Cab18]. conjugated [MHD18]. conjugates [BJ18, YCL^{+}17]. conjugation [PCHM10]. conjunction [KHK15]. Conjunctivitis [Hur12]. connect [FRH15, TGL15]. connected [CM14]. Connecting [RC11]. connection [ASC16, AH12a]. connectionist [FIS11, FIS16a]. connections [DHT16, SS15a]. Connective [FAFA17]. connectivity [BPG^{+}18, DBB^{+}17, Hal16, HvLPdB15, KK17a, NMS10, RFME^{+}12, WMT10]. connexin [WMT16]. conotoxins [ZZG^{+}16]. conquer [JH11]. consanguinity [ANMH11]. consciousness [RBG^{+}17, Sat14]. consensus [MP16, SLM17, LLP^{+}19]. consequence [AGPK13, MDS16, NPB13]. Consequences [BAM^{+}15, CVO^{+}15, CP11, WF18, BD10b, CdPD11, DFM^{+}14, KG13, MA11b, RWSB11, SBSE14, VK10]. Consequential [VK10]. conservation [DZW10, KMN^{+}15, SAB17, SVA18, Tsu19]. conservative [Rog19]. conserved [ÅCHS19, AKR^{+}18, DS15b, FT12, JRR^{+}12, PEZ^{+}13, SRP16]. consideration [lDsWM^{+}15, KMM18]. Considerations [ABIM10, Fue16, Wil13]. considering [BBJ^{+}10, CDO17, Dol19, LJM15, MZAI19]. Consistency [Ste13]. consistent [BS15b, Gia13, JSK16, NRKE18, NPH^{+}12, SSS13]. consisting [LZTD18, TD18]. consolidation [WGS15, ZLL^{+}12]. consortia [SCH^{+}19a]. consortium [SK16b]. conspicuous [LS10a]. constant [BSMK11, CK18, KMM16, LLB^{+}18]. constants [VACGF17]. constituent [HHA17, WXC10]. Constitutive [FK16, LZWK11, PMD^{+}18, ZLZ^{+}11]. constraining [RT19]. constraint [NvHM13, RGCML10, SWTO15, ZGWL14]. Constraints [SKK^{+}12, BDS12, CBL14, Cam15, CCS18, CMBP19, GCM17a, HAK^{+}19, IS16, KHP^{+}12, LWB^{+}15, Lüt16, Mal17, MM15b, OABI12, PPW16, Ush16]. construct [ABIM10]. constructal [KTO^{+}14, LCBO^{+}12]. constructal-law [LCBO^{+}12]. Constructing [CN12b, Lev14, LT14]. Construction [TDM14, BRP^{+}18, GKWG17, JAM18, KGB13, KWO19, TM16, TTN13, dF13]. constructs [CLCH10]. consumed [WBM^{+}15]. consumer [AR11, Abr14, BV13, RB13, SSD11b, SSD12, WD11, tBdR18]. consumer-resource [AR11, RB13, SSD11b, WD11, tBdR18]. consumers [XF13]. consumption [MMK16, RZRSC19]. Contact [FPG11, AL17, AAJGCD15, BGW15, GAY^{+}15, HS15, HB16a, KG12, LL14, MvdDW13, MPLK14, NRKE18, PSL^{+}10, SZ18, SMC^{+}13, TC12, YSH^{+}14, YZMY18, YS11, ZCXF13]. contacts [CGW18, IU18]. contagion [NO15, NO18]. containing [LT14, MCL^{+}11, ZG19]. containment [AH12b]. contaminants [HPW^{+}13]. contamination [BSC19, MW13, WXWL12]. contemporary [KDMK16, SS19]. content [JS15, JS17, JS18, MBE10, MS12a, NHM10, Wax11b]. contest [MGS16]. context [CSR^{+}05, FRG^{+}13, KG13, LDB^{+}14, ON14, Sat17, SY11, SRAL12, WGS15]. contextual [GRRG16]. contingency [FRG^{+}13]. contingent [WFZW13]. continuation [JR17b]. Continuous [MT15, Sim14, TP10a, ATB14, BSKV18, BBS10, Cre10, DHK13, GZFX14, HMWB13, IOY15, KS14, LY11, ML10, RA10, RA18, Sch19c, SFS13, Shu13, WL12a, WL12b, WSH^{+}10, Yat14, vDRT14, vE11]. Continuous-time [Sim14, vDRT14]. continuous-valued [HMWB13]. Continuously [DI10a]. Continuum [JDPK15, MPBS17, MBS19, HGGCR13, KCZ^{+}19, Kro10, LHPF18, MFG14, SS14, SKH11]. Contract [Arc11]. contractile [LFW^{+}18, SH16]. contraction [HGGCR13, LHPF18, MAH12, SS14, SKH11, WGO^{+}15, YWP13]. contractions [DGNT17, Kro11]. contrast [LS16c, LPF11]. Contrasting [CB15, AK13]. contrasts [Ell15]. contribute [KEHK17, SAB10, VBHM^{+}13]. contributes [Laz13]. Contribution [DK13b, LLW15, MYK17, dS15a, BCBD19, CPW16, GGQ^{+}12, HFD17, JDSPK15, RT15]. contributions [ABJ12, Ano19-50, Bur19, KG12, PBP15]. Control [BS15a, DBM^{+}18, GK10, GCG14, HZG^{+}17, RJSC18, SSD13, SZ15b, TWTA^{+}18, ASK17, ABA11, ABIM10, BFJ^{+}18, BAGG14, BZL17, BCBD19, BAR14, CFCM13, CSLE11, CKN^{+}12, CH18, CDGV10, CMS16, CG11, CBGS18, DMS^{+}16, DROC11, DKR16, Don17, EHBC10, EB15, FE11, FD17, FH13b, Fra19a, Fra19b, GKMC10, GM19, GA16, GTC19, GM17, GP11b, GP12, GT15, GAY^{+}15, HS15, HCHI18, HTM16, HB16b, HJR12, HK11b, HCW18, JB19b, JWTW17, KKD18, Kee10, KAZ11, KAZ13, KEKB18, KSvdH18, KLJ17, KP16, KKYV18, KF11, KH19, LCCC10, LBGW13, LKK13, LCJ16, LC17, LSY^{+}10, LLZ13, LNH13, MIJ16, MA13, MAM16, MG10a, MZAI19, ML10, MS12b, MFKS13, NSS^{+}11, NMAZP16, NSH^{+}10, OA12, OA11, OCHHZ12, PVCEC18, PZLF19, PRM12, PRM14, QF10, ROF17, RB14, RS14a, SW18, SKK^{+}12, SSG^{+}19, SGW17, SSST19]. control [SBM^{+}19, SNCM12, TXTW16, Van17b, VCF^{+}19, Wal16, WPA17, YT12, YAK17, YD14, YLLL12, ZCT18, ZTT18, ZN18, ZLT^{+}19, Zhu11, ZZR10, ZZR13, dSKBS10]. controllability [IH17]. controlled [BFR14, GLOC10, Gri11, GRB^{+}13, Kro10, Lee16a, SK15, vLFM^{+}19]. Controlling [BVJE17, AHJ^{+}18a, HLI15, HRCA19, MB14, TI12a]. controls [MSB16, Sar10]. controversial [ACM16]. convalescent [HSLW16]. convection [DS16]. conventional [Gia13, HL18]. Convergence [Dal17, TF15, Gou16, NWZ15, WL12a, WL12b]. conversion [FBC12, KII19, MSC10, MFMB12, RRÚJ19, SAGC12]. Convex [TZY18, HHMT19]. convolution [TTC19]. convolutional [WYL^{+}19]. Cooperation [Arc16, ACCR11, Axe12, BP16, CSZT12, DBD12, Gha16, KCS16, Pła10, SSS17, TR12, Van16a, WK18b, XY14, Yan10b, ZZC10, AH17, ATB14, Arc11, BYJ17, BBD18, CS14a, COWA11, CWW18, CBP12, CAV16, DG12, Das12, De 19, DDRP13, GWNW15, GT12, GvVT14, HNR14, HKM12, Hou13, HCW18, ITO16, ID19, IIVR10, KNT10, Kon11c, KC11a, KI17, Lai18, LPH18, ML13, MSN14, Mas12, MJI14, MBBD13, MKMG^{+}14, MB19, NSKO18, NT10, NBA^{+}18, Now12, Oht12, PBR17a, PCN17, RANO10, RMB15, ROF17, RC11, Ros10a, Ros13, RRR15, SPS11, SPLP12, SN18, SP16, SLW18b, SK11b, SP13b, TM12, TGL15, TA15, Uit11, WSP14, Wol12, WFZW13, WFW19, YvBS18, YHZ14, YZX15, ZCW13, ZLY^{+}17, ZSS10, vDRT14, vVGA10]. Cooperation-based [WK18b]. Cooperative [CCAdS13, Kin18, KUV^{+}10, HVSZ10, HKB10, Kon11b, LDF^{+}11a, MW14, PBA12, QJR^{+}16, RN12, SHN12, SHH15, SJK18, TBQG14, WK17, ZZCZ17]. cooperativity [BSP18, CL10, CB13b, Mar17b, Mic11, ZBA18]. cooperators [FNH10, GCS12, LG10b, OO17b]. cooperi [KHB^{+}18, KFS^{+}13]. coordinate [Ben14, VC10]. coordinated [FW15, NS18]. coordinates [DGL12]. coordination [Cle18b, PEW18, RPGG^{+}19]. Cope [SSD12]. copepods [RL17]. copied [YCR^{+}15]. copper [BSC19]. copy [JOM16]. coral [BH11, CMGN17, KRSD11, MI11b, MI11c, RTFP^{+}17, TZ13]. coral-algal [BH11]. coral-Symbiodinium [CMGN17]. corals [RDTHSM13]. cord [ISZ18]. Core [WCC14, BFGS10, CKN^{+}12, CSKZ19, GXFF13, KBS^{+}19, KBS^{+}20, KF12, Sat14, VH11a]. core-shell [KBS^{+}19, KBS^{+}20]. coregulation [WM17]. cornerstones [BPFR16]. corneum [SSvdM10]. Cornish [HS16]. Cornish-Bowden [HS16]. coronary [DPCM16]. CoRR [All17]. correct [MA10]. corrected [OGO19]. correcting [Fra19b]. correction [HK16, ZTWL12]. Correlated [PPBD10, BPG^{+}18, BM16, Che10, Ell15, MKRE18, NdORC11, PNK16, Rey13, RAD14, WMCL18, ZZC14]. correlated/persistent [Che10]. Correlation [ESW13, WZWY11, YTGW16, AGB^{+}14, BE14, DCS14, DBG18, FSW^{+}16, HgLL^{+}10, JAB18, KDMD13, KN11, Kro10, LXS15, MC15b, RCD19, SB12, SCLC13, SS11, TMF17, ZYJL18, ZL18a]. Correlations [MK19, GCZ^{+}12, dOGL10, dOGL13, HH16, HHD^{+}16, HNA15, Sel10, TP10c, VDD12, WMHH10]. corresponding [DHT16, FS11a, FS15b, ZLZ^{+}16, ZLB13]. Corrigendum [AEM14, BBM^{+}13a, BBM^{+}13b, BPGS12a, BMGC11a, CLG^{+}11, CNG^{+}12, DBBW11, EJ19, EPJ^{+}11, FIS16a, FS16a, Fue15, GS14, GHS20, JS18, JRG14a, JCLS^{+}21, KBS^{+}20, KMLT14, LPB18, LBS^{+}11, LGK^{+}12, Lei10, Lio24, LHPF18, LPvSP12, ML12a, Mei13a, Mei13b, MLBA13, Mul12, NBC16a, PSS^{+}13, PCT19, PP19, RTEKG19b, RTEKG19a, SCÁBM11, Sch21, Sch23, SRAL12, SYY17a, TSM13a, Van16b, WL12a, WLW20, YCH^{+}17a, ZG10a, ZSCL16a, ZZR13]. corroborating [Di 17b]. corrugation [DS12]. corruption [HCW18, LSDI15, LJI17]. cortex [FHR13, NGL^{+}10, Pat19, PW19, TWTA^{+}18]. Cortical [DI13, AANF16, BTO14, KLN^{+}12, KPD18, KPD19, RNN15, SSD11a, WZ12]. Corticothalamic [HRDL14, FJR19, SLR17, ZKR15]. Corvus [GSV11]. Cost [ZKMB19, Bac15, BR13, Fue14, Fue15, HB12b, Lar10, LP14, SRV11, SS15b, SS18, Tra16, VDH^{+}15, WL19, ZZCZ17]. COStar [LDWW14]. Costly [Kis10, Akç12, BWL12, JR17b, MM18, Nak16c, SRV11]. costs [Fil15, GSL13, LLG16, PPBT11, vVE15]. Could [IK15, Dol16, FGAM12, OO11, PDL^{+}17, RDD14]. count [LM15, LS13b, Mog15, TXWW12]. count-guided [TXWW12]. counter [Ber14]. counter-strategy [Ber14]. counterpart [RLSN14]. counterposed [Di 16]. countries [DB10, RB14]. Country [MDMG14]. Country- [MDMG14]. counts [HS11, YW13b]. couple [AAJ15]. Coupled [CXWL11, CKS15, HI19, MROS15, BTO15, BMF^{+}18, CEKM^{+}19, DGJ15, For10, GM19, GL14, HPML18, KJ11b, MGC13, PBBB10, SW13a, SOCF14, SX12, TGLK19, WOB15]. Coupling [OSF11, WFM^{+}13, WFC^{+}14, ZS12, BRG^{+}12, BTO14, CFZ14, DvDBD15, DBD16, FD11a, FH14, GMMN18, Hor11b, JAK19, KHH10, KTH16, Kuz19, MDB12a, MT14, NHS^{+}16, SKAG18, SB19, SY12, TIS10, WMT16]. coupons [KFG15]. course [BBDB13, NS11]. CoV [SKPK17]. covalent [Dun11, SC13a, SC13b, Str17, ZJS11, ZZ10]. covalently [LLDW14]. covariance [FR13a, QZY17]. coverage [LLD^{+}17, YTGW16]. Coveting [WMD^{+}11]. cow [GKG^{+}18]. cow- [GKG^{+}18]. Cowan [CV19]. cows [BGF^{+}14, CFR^{+}14, CMR^{+}18]. Coxiella [CMN^{+}11]. CpG [HA15a]. Cracking [Nah14]. cranial [BFGAGA16, YKO^{+}16]. craniofacial [BDID^{+}12]. crassa [XCW^{+}18]. Craton [Kur14]. Crawling [WL13]. create [CZPC^{+}18, HA15b]. created [SI11b]. creating [KG13, PJ13]. credible [Di 16]. creeping [SAGAGB17]. crescentus [SK12a, Sht17]. crest [LFZN11, Sch19c]. criminology [SOIO10]. criteria [CJ12a, CZW^{+}11, Gri15a, Kro10, Opr10]. criterion [Di 19b]. Critical [HSiI^{+}19, LG13, APBS15, BCPM^{+}16, CGvG^{+}15, GXF13, HBT13, iHM17, MJV16, SCA13, SPH12, Van16b, Voh17, WMT10]. criticality [JRG14a, JRG14b, JRG14c]. critically [KKM12]. CRM107 [YCL^{+}17]. crop [ABR19, BSB14, SS11]. cropland [BCBD19]. cropping [RDMP11]. crops [AM14]. Cross [GCB17, HF17, MCL^{+}11, BJJR10, CMD^{+}10, FKK14, FZL18, HWW^{+}14, HTB^{+}13, KKOM18, LHFM16, MW13, MBK^{+}11, OF10, OAS10, SDD15, SCH^{+}19a, SCLC13, SJK18, SKS^{+}19, VGL16, WCF12, WA14, XJ19, YCH^{+}17a, YCH^{+}17b, YST14, ZL18a]. Cross-bridge [MCL^{+}11]. cross-contamination [MW13]. cross-correlation [SCLC13]. cross-enhancement [WA14]. cross-feeding [SCH^{+}19a, SJK18, SKS^{+}19, VGL16]. cross-immunity [CMD^{+}10, LHFM16]. Cross-jurisdictional [HF17]. cross-link [FKK14]. cross-linked [FKK14]. cross-reaction [OAS10]. cross-reactivity [HWW^{+}14, SDD15, XJ19, YCH^{+}17a, YCH^{+}17b]. cross-regulation [MBK^{+}11]. cross-specific [OF10]. cross-talk [BJJR10, FZL18]. cross-talking [YST14]. crossbridge [ST17b]. crossflow [KA15]. crosslink [DBJ12]. crossroads [dlEBRM15]. crosstalk [RA13, XCW^{+}18]. crouched [BBJDS11]. crowd [DS10a, ISB^{+}11]. Crowded [OO17a]. crowding [Cza14, DHB15, Vaz10]. crown [Cam11]. crows [LZTD18]. crucially [IH17]. cruzi [KZ14, PKZ12]. crypt [BZL17, FBC12, MFMB12]. cryptic [BC15, CLPZ14, PDC^{+}17]. cryptically [Sel13]. cryptosystem [Mor11]. crypts [YAK17]. crystalline [RCL^{+}10]. crystallization [JM12]. CSF [BPLM12, CHN^{+}15, SER^{+}12, ZLM12]. CSS [Cre10]. CTD [NTC^{+}11]. CTL [Lev14, SBR10]. CTLA [LZL^{+}19]. CTLA- [LZL^{+}19]. CTLs [ARM18]. CTPB [DK13c]. cuboctahedron [Den19]. cuckoldry [YY18]. cuckoos [LZTD18]. cue [Jen10]. cues [BJJR10, PDL^{+}17]. Culex [VSD^{+}17, YMZ18]. Culicidae [YMZ18]. culling [LSSG10]. cultivated [AM14]. Cultural [HLI15, Nak19, TI12a, CT14a, EL12, KS13, KLM14, ML10, Nak14, SC10a, SN11, WL12c]. culturally [WF18]. culturally-driven [WF18]. culture [GS17a, GZFX14, HPM^{+}17, KLM14, KWO15, KWO19, Nak16c, PMKM10, Saf13]. cultures [Aya12, BBS10, BPF^{+}19, CWP^{+}18, DBBW09, DBBW11, GMOP12, MMB18, PMKM10]. Culturomics [SD16]. cumin [MFM16]. Cuminum [MFM16]. Cumulative [CT14a, PC17, RAF^{+}19, EL12, KLM14, KWO15, KWO19, MBBV14, Nak14, Nak16c, Tal12]. curable [HHA^{+}13]. curbing [TXY^{+}12]. curing [PRM11]. current [AOM19, Dim17]. currents [BHBH11, STKE12]. curtailing [KM15]. Curvature [DSF11, CL10, JTSG14, KPD19, KPS17, Mac11, MPK^{+}12]. curve [DGL12, Opr10]. Curved [Mil16]. curves [Mar17b, PPT^{+}16, RG17, Str15, TCH14]. customize [NSH^{+}10]. customized [YD14]. cutter [KCMF11]. cutting [CLK18, RDMP11]. cyanobacteria [GWBL12, RSBB10]. cyanobacterium [iITIM15]. cycad [RBB^{+}12]. cycle [AAAV14, ASK17, ANGPB12, AN12, BTO14, BTO15, BSR^{+}11, CLHB11, DBBW09, DBBW11, EBE14, EBE17, FP18, GFM^{+}19, GGG19, GMNY14, HS14a, HWPL12, JAK19, JS15, JS17, JS18, KRR14, KCS^{+}15, KSPA17, LG12, MS13, MLT10, NIiT18a, OABI12, PRK13, PGHC12, POK^{+}12, RSD^{+}13, SB19, TDSM12, VHM^{+}17, WLD^{+}11, XSMF12, YFB^{+}12, vVN12]. cycle-regulated [FP18]. Cycles [KdSUS10, BI10, CE17, GK17, JOAN14, KGB13, RL17, RS10, Ric19, RBB16a, Sat10, SC13a, SC13b, Str17, SY15, YRD17, tBdR18]. Cyclic [LG19b, Bec14, CLA17, GDJC11, KSK^{+}11, LAJC19, LRGA13, SGD12, SGCD14, TBM^{+}13, TR12]. Cyclical [LM11, ZML19]. Cycling [HVSZ10, LG19a, MK14a, Nak16a, RRÚJ19]. cylinder [KA15]. cyminum [MFM16]. cyst [GTB10]. cysteine [TAR16]. cysteines [GGX^{+}10]. cystic [BTK14, KBV^{+}15]. cytochrome [RS19]. cytokine [CDM^{+}14, LHW14, NF16]. cytokines [RAR19]. cytokinesis [RD12, Wol14]. cytokinin [MBK^{+}11, MBKB13]. cytokinin-auxin [MBK^{+}11]. cytometry [OBK^{+}11]. cytoplasm [AAJ15, SA17, TM15a]. cytoplasmatic [Vaz10]. cytoplasmic [Rad16, SWR11, SS15f, YTK10, YYB^{+}19]. cytosine [KB10a]. cytosis [CKD11]. cytoskeletal [KHNM16, ONO15]. cytosolic [GDB^{+}19]. cytotoxic [LG12, TALC16]. cytotoxicity [HGO^{+}18].

D [LDWW14, LHPF18, RRC^{+}11, BPF^{+}19, CK17, DS15b, ESGA15, GBC^{+}16, GM16a, GMMN18, HGGCR13, KKGN16, KMM18, MPCTGJ^{+}15, Nos14, PHA^{+}17, RMSTG13, SK15, TA16, ZKP15, ZSH^{+}16]. D-motifs [RMSTG13]. D-QSAR [TA16]. D-star [LDWW14]. D1 [ASRM15]. daily [BLV18, DDB10, WTC16]. dairy [AMSSG16, BGF^{+}14, CFR^{+}14, CMR^{+}18, GKG^{+}18, LSSG10]. Daisyworld [BG12, BLW11]. Dalgarno [SGD^{+}16]. Damage [TALC16, ASK17, ADC19, CDM^{+}14, Eli17, GA14, JSDEK14, KBE^{+}13, KP16, LCTG15, MBBV14, VLPH17, ZSCL16a, ZSCL16b]. damages [MYK^{+}11]. Danio [Hal17]. Darby [HHA17]. Dark [Das12, LRL^{+}16]. Darwinian [CMM19, CS15b, GCM17b]. Darwinizing [Doo17]. Data [Ana11, JWTW17, LZA^{+}19, ZLB^{+}18, ASH15, ABKS11, AHR10, AD16, BCPL10, BMI19, BWP10, BMB^{+}18b, CR14, CPS19, CAG13, Dim10, EJ17, EJ19, GBIR13, GGCJ16, HYN19, HII^{+}19, HW13, HZL^{+}11, KKD18, KHX^{+}19, KA16, LJM15, LAPK14, LZMM15, MSP19, MAMEA15, Mor19, MS14c, Mun12, NNK^{+}15, NI11, Nos14, OGE10, PH13, PS13, PGLZ14, PBEI12, PBD13, QQL10, QYO10, RKMG12, RRRTR10, SS17a, SKH17, SWD16, SD16, SCKL15, SLW^{+}18a, SKK^{+}11, SGW^{+}18, TZ13, TWH^{+}13, TT17, VF12, WH14, XSKA17, YD15, ZMS11, ZLZ15a]. data-dependent [HW13]. Data-driven [Ana11, LZA^{+}19, ZLB^{+}18]. database [OSV^{+}16, YXH^{+}14]. dataset [CXC18, JZZK11]. datasets [SM17a]. day [VZB^{+}15]. days [KK18]. DBH [WZ17]. DC [Moo14, SKGM19]. DCIS [METC12, NWB^{+}10]. de-differentiation [Wod18]. Deadbeats [MG17]. deadly [KJM17]. DeAngelis [GG12, GG14]. Death [LNL^{+}11, SGB19, CPW16, CMMR13, FWB^{+}12, FLGGD^{+}14, GTSP16, Hal12, HBSF11, KFA12, LAS14, LG10a, RAD14, SZ18, Sta10, SGW^{+}18, TRM^{+}14, WZ15, ZLL17, ZZS19]. death-related [FLGGD^{+}14]. deaths [Zuc14]. DEB [EBS11, TTB^{+}18, LAPK14]. debate [Tjø10]. DEBkiss [JMZ13]. debt [SIVH17]. decay [JKS15, NDZMA14]. December [Ano10-38, Ano10-43, Ano11-27, Ano11-31, Ano12-42, Ano12-30, Ano13-31, Ano14-44, Ano15-37, Ano15-40, Ano16-35, Ano16y, Ano17-37, Ano18-39, Ano18-34, Ano19b, Ano19a]. decide [ZZL^{+}11]. Deciphering [AHMA^{+}19]. Decision [AOT^{+}11, GRR^{+}14, LTSTD15, LNL^{+}11, AHAFA14, BVJE17, CC11b, GJ15b, HIT18, RLK10, SM17a, SYI17, Sue12]. Decision-making [LTSTD15, RLK10, SYI17, Sue12]. decisions [AB16, Coo16, EM11, MPNP12, Mor19, SKPK17, vLSBBD11]. decline [Jus11]. declining [WWB14]. Decoding [PJ13]. decoherence [AOT^{+}11]. decomposition [BJO^{+}16, CZJQ14, PP14]. decompression [HRG11]. deconstruct [JTSG14]. Deconvolution [LB12, Lüt16]. decorated [ONO15]. decoys [MIH16]. decrease [KKO^{+}18, VMZK^{+}19]. decreases [HPvdB17]. dedicated [TK19]. deduce [GS10]. deduced [PKH11]. deep [IA17, MYK^{+}11, TLPH11, WYL^{+}19]. deeper [Di 19c]. deer [EEHMH18]. defeat [WS10]. defection [BYJ17, Bar11, CGD13, PTB12, ZLY^{+}17]. defectors [GCS12, Sch10, WB19]. defects [Don16, NAK^{+}11]. Defence [HK10, FB18b, HSR12, VB19b]. defences [BHR10, WCS19, WRC^{+}19]. defense [DL15a, IEN15, ST17a, Yam16b, vVE15]. defenses [LH18]. deficiency [CB13a]. deficit [BFGS10]. deficits [SHG16]. defines [JZL13]. Defining [KC11a]. Definition [KDST15, BGD^{+}13, Han17, NGND12, Pia19]. Definitions [LS16b]. deformable [BB15, CFGRB17, DV12, She10]. Deformation [MYK^{+}11, DS12, LCA^{+}15, MS14c, SW15, SAGAGB17, SGB12, THBM10]. deformations [SS14]. Degeneracy [WB10, LDH^{+}12, MZM^{+}16]. degenerate [LFW^{+}15, TMF17]. degenerated [KB10a]. degradation [MTLW^{+}10, TLL^{+}12]. degrading [SBS17]. Degree [CAG13, JSP^{+}16, Mar17b, SBK16, WL13, WMCL18, ZEJA11]. degree-correlated [WMCL18]. degrees [JCG15]. dehydrogenase [EBSW17, Pav14]. Dejerine [AD15]. delay [BT17, CV19, ESG^{+}16, FWLW11, HLH^{+}18, JP15, KCSB14, LS16a, LLL13, LBZ18, MSZW19, MS13, MS12b, MBP16, RCH14, SX12, SPA18, WYKT17]. delay-coupled [SX12]. delayed [BBI14, GT14, Iij12, JP13, KKWA18, LJ18, NKB16b, PJ13, UJLG14, WK17]. delayer [BB10]. delays [BH11, EKFV18, LZTD18, TAL19, WCHC11, ZZ14b]. Delbrück [Zhe15]. deleterious [AM11, CNDK15, JJ15, O’F11]. deleting [Sel17]. deletion [KDK14a, KDK14b, KGM15b, Ros15]. deletions [HTT19, Sel15a]. delineation [SS19]. delivered [BB19]. delivery [GSV16, KPS17, LLZ13, TBMM19]. Delosperma [KHB^{+}18, KFS^{+}13]. Delta [TSB^{+}17, KA16]. Delta/Notch [TSB^{+}17]. Delta/Notch-mediated [TSB^{+}17]. demand [LAPK14, SGGHS19, Wal16, dM10b]. deme [HI12, WL14]. deme-structured [WL14]. demineralisation [IvLP12]. demixing [ZWvG10]. Democratic [OL19]. Demographic [Oku15, YRD17, Ada16, DFM^{+}14, DB10, DCR18, LP14, LGBT15, MGCL13, MB10, PDSP13, PP17b, VDD12, ZL18b]. demographics [NTED^{+}19]. demography [BBR16, GAY^{+}11, NMAZP16, PSJ15, TRJD19, WKB13]. demography-driven [PSJ15]. demonstrable [MPC12]. demonstrate [AMFR^{+}16, TCH14]. denaturation [Kue16]. dendritic [LSP13, LLDW14]. Dengue [TXTW16, ZN18, ABKS11, BLV18, GTC19, Gru18, HB16b, HTB^{+}13, JWTW17, Lan16, LCC15, MHX^{+}14, MDMG14, yTRSC13, NBBCC18, PVCEC18, SDT17, STN19a, SKK^{+}12, SSG^{+}19, WA14]. denitrifying [GGGM16]. denoising [TWC^{+}19]. Dense [Thu15, YLY15, LSMC18, LIB^{+}17]. densities [LAG^{+}14, PKZ12, TGL15]. Density [GRBL13, HGHL10, MCPF12, NCN13, WCPF15, Bal10, BKR14, BH13a, CMS^{+}19, CK17, Das18, DK13c, DZE11, EPP19, GBC^{+}16, JSK16, KMZR18, KMHdlP10, KHRS10, KGC18, LS10c, MZWC10, MJ11, OGA16, PGH11, RWSB11, RB13, STA15, SSB^{+}15, SBSE14, TKTB18, Tos11, Vaz10, VSW10, dG10b]. density-dependence [OGA16]. Density-dependent [HGHL10, MCPF12, WCPF15, Bal10, EPP19, RWSB11, RB13, SBSE14]. dental [IvLP12]. dentate [GMM15, GM16b]. DENV [TWR^{+}18]. DENV-NS3 [TWR^{+}18]. departures [BBR12]. depend [CGHF14, SS11]. dependant [WKS14]. Dependence [DSLR18, Nij11, DL16, DZE11, EzMMH18, KHRS10, KG15, LS14b, LS10c, MBBV14, OGA16]. dependency [BB15, CSR^{+}05, PHÁBAI^{+}16, SRAL12]. dependent [Aya12, Bal10, BSB^{+}13, BDLR12, BFM10, EPP19, FAFA17, FR13b, GRCdL16, GSV11, GKW16, HMTA17, HGHL10, HGO^{+}18, Höh15, HW13, HTB^{+}13, HT10, ITN^{+}11, Jen10, JSK16, KMHdlP10, KJ19, KGC18, KH19, LAS14, LKK13, LH14, LLCC12, LLB^{+}18, MB10, MCPF12, MAS15b, MKPVH16, NG11, ØW16, PRN10, RLW^{+}14, RWSB11, RB13, RU16, RC16, SA17, STN19a, SSS15, SBSE14, SGD^{+}16, SW13b, TGLK19, TSP15, Tos11, THBM10, WCPF15, Wax11a, Wei17, YI18a, ZCA^{+}14, ZTT18, ZW16, vE11]. depending [HHA17, RS12]. depends [WGO^{+}15]. depletion [FS11a]. deployment [GA16, LvdBP12]. depolymerisation [ADMR16]. Depressing [JSZ10]. depression [BB11b, HKS16, KPD18, KPD19, Ros13, YK11]. deprivation [FPR10, RTR^{+}13, RRL^{+}13, RLW^{+}14]. depth [PLHL19]. Derivation [LAG^{+}14, DS13, GG12, MJS11, MKBE17a, MKBE17b]. derivatives [MR18, MR19, POW17, TAR16]. derived [ARD14, BC19, DBBW09, DBBW11, DTP19, FHG15, KCM^{+}11, KR15, LMM^{+}13, NTC^{+}11, SSM^{+}18, VHS17, WL12a, WL12b, WLY^{+}17, WSRG18, ZPdFJ19, ZNA^{+}16]. derives [MPNP12]. deriving [BGLR^{+}19]. dermal [MHMM11]. dermatitis [TDHC^{+}18]. dermis [ØW16]. desaturase [PWH^{+}13]. describe [BGD^{+}13, NKOS11, PLF19, WSH^{+}10, ZVMB10]. described [DD13]. describes [GB13, IGL^{+}12, PRC15]. describing [JH12]. description [ABK^{+}12, CJKR10, Dim17, HNO18, IDM15, IKHL16, MA10]. descriptions [MBS19, PLD^{+}17]. descriptor [YYH^{+}14]. descriptors [DHS^{+}15, FHG15, FLCS^{+}15, MPCTGJ^{+}15, NBL10, NSBL10, NBL14, NLB14]. desert [MAW12, RSC14]. Desertification [ZUFM17]. deserts [BS19b, CW15c]. Design [pNMK^{+}10, TM16, AKS^{+}19, BC19, CRLH^{+}19, Fra19a, Fra19b, GZ19, HAB^{+}17, JRB^{+}16, KTO^{+}14, KAN11, Mig16, PGGvdB12, SSJK18, SMJS14, VGPS18, VBV10, WB10]. designed [PPF17, TDT16]. Designing [HS15, III17, JB19a, KRDJ15]. designs [YNY^{+}10]. desired [Kon12]. desmoplastic [NF17, PM13]. despite [WWB14]. destabilization [GCYL18]. destabilize [Kis10, MGML10, TGC^{+}17]. destabilizing [HWPL12, RAMS11]. destination [PTC13]. destructive [ACCR11, MSP19, NF14a]. desynchronization [BTR18, FW15, SY12]. desynchronization/synchronization [BTR18]. detachment [SMS19, SOCF14]. detail [SWG^{+}15]. details [JG14]. Detecting [DGM15, For10, HBL10, NMZ19, WSK^{+}19, WKH16, LMW^{+}10, SD16, Tay16a, ZK10]. detection [AEZR^{+}16, AK11, CPV16, Cut15, Fer12, GNB^{+}13, JRR^{+}12, JRMS12, JP13, LJM15, LYZ^{+}18, LMJ^{+}18, Lug15, MSND12, Mar19, OO11, PDC^{+}17, PGGvdB12, PCC15, SVCS10, YYC19, YLF^{+}17, ZLY^{+}13]. detector [JSZ10, KL11, LJM^{+}17]. detects [LZG^{+}19]. deter [YLY15]. determinant [AOR17, CGvG^{+}15, CZ14, Cle18b, HWTN15, HH13, IM18, QZY17, Van16b]. Determinants [US12, Ada16, Bur10, HB13a, KI15, MDM16, Pav14, RS14b]. Determination [KCSB14, SSS15, ADCG14, CSKF10, Kam11, KFG^{+}14, LF13, MÇ15a, MBK^{+}11, TEY16, TC12, YHI14, YI18a]. determine [Bal13, CCS18, CS15a, Dol16, Don17, GSL13, LN13, MFKS13, RC13, TDHC^{+}18, VM16]. Determined [PP18, PP19, SSI^{+}16]. determines [KKM12, SCÁBM10, SCÁBM11, TRJD19, VY12, ZAL^{+}19]. Determining [ADR11a, NRKE18, SABB15, SN17, YAK17, ZFWK17, AGRR^{+}15, GLR^{+}18, GVCG12, KK17b, PKH11]. Deterministic [LDF^{+}11a, PJ18, IGL^{+}12, Pag19, PMW12, RBKW19, SAB10, ZZ10]. detrended [SCLC13]. detrending [DSTD19]. develop [CBGS18, MFG^{+}13, SWTO15]. developed [HKR^{+}19, MZAI19]. Developing [SCF^{+}12, WDL^{+}13, BCS^{+}16, GBN14, MK16b]. Development [FSA15, HBA10, ISZ18, KMA10, MBC^{+}12b, MBCM12, ŚKSRW16, ALD^{+}11, ASF^{+}15, BTO14, BSR^{+}11, BFGAGA16, CCNT19, CM12, CCB11b, CEP14, DTGD19, DM10a, DDB10, FWB^{+}12, GLZ^{+}11, GC18, Hal17, HTM15, iHIM12, KKOM18, LMC^{+}13, LTZ17, LH13, MM12, MFMS^{+}18, MBK^{+}11, NHE^{+}16, OCN10, PAS10, PHH18, PPC^{+}17, SBSR13, SGGY10, SWSLMJ19, SSRA16, TSB^{+}17, ZZL^{+}11]. Developmental [STNT17, BLNR15, BBDB13, HJWC11, LHD^{+}17, Vel17, WKH16, ZZS19]. developmental-dynamics [LHD^{+}17]. deviations [DPCM16]. devices [FPG11]. DEVS [ANMH11]. dForml [NMZ19]. DFT [DK13c]. Dharmendra [Pan19]. di- [Kři14]. diabetes [BSR14, BMM^{+}14, JDPK15, JDSPK15, KSU^{+}18, KCSB14, RAA^{+}16, WHS^{+}13]. diabetic [BPS15, LLB^{+}18]. diabetically [HBA17]. diabetics [CTW10]. diacylglycerol [CHD^{+}10]. diagnosed [Jam16b]. diagnosis [EAN14]. diagnostic [RRG^{+}12, SHSR15]. diagram [MMY^{+}12, Zho11]. diagrams [AJM19]. dialect [FM15a]. diameter [BSV^{+}10]. diapause [TI13]. diauxic [BRG^{+}12, Chu17]. diauxie [ABK^{+}12]. dichotomic [FDS13]. diclofenac [PYG^{+}19]. dictate [POP^{+}19]. dictator [SSM15]. Dictyostelium [ANK10, FIJJ11, MGT17, PWE15, SN12, SSS17, Tre19]. Dictyostelium/Pseudomonas [FIJJ11]. did [Di 14a]. didactic [CD16]. Diet [All11, Kři10, Rem15, TBA14]. dietary [POK^{+}12]. Difference [MYN^{+}13, MYN^{+}15, LL17b, Ots11, TBA14, ZZK14]. Differences [OGA16, ABKS11, Cam11, FS12, FBK11, KLM14, MS18, MM15a, Oko15, PVB^{+}10]. different [AK11, BMN16, BHR10, BKL14, CZN12, lDsWM^{+}15, DDF^{+}14, DHT16, Eft13, FtL18, HSG^{+}18, HY13, ITR^{+}18, JYZ18, KDG13, KMCJ17, LZ18, LDdA^{+}13, MHKA16, MNSZ16, MAR^{+}17c, NF14b, NGND12, NBL14, Pan11, PGGvL^{+}19, PBD11, RBMS17, SMM15, SS15e, SGW^{+}18, SAZ^{+}14, TIS10, TG10, UI10, VLFF12, WP17, jWlGQ19, WRC^{+}19, WLL^{+}14]. Differential [OAJK10, BMM^{+}14, ESG^{+}16, EAN14, FWLW11, GGM15, HJLNZ11, HLH^{+}18, JCLS^{+}11, JCLS^{+}21, MBCM12, NBC16a, NBC16b, OOY16, PCHM10, VH11a, WP17, Zad11]. differential-equation [GGM15]. differentially [ZZC14]. differentiate [ML12b]. differentiated [FM14b, HMJB15]. differentiation [BF13, CW15c, GMOP12, Gol10a, Gol10b, GZT15, iITIM15, KP12, KUV^{+}10, NKOS11, PWE15, RSBB10, SCÁBM10, SCÁBM11, ST16, SLL18, SB19, TZ11, TWB16, Wod18, YM14]. differs [Mal15]. difficile [JEA18]. diffuse [CL14]. diffusible [ON14, ZGS^{+}10]. diffusing [LDHD14, PB16, RS14b]. Diffusion [Flo11, MC16a, PH13, AGRR^{+}15, ATB14, BTG^{+}15, BF16, BMB^{+}18a, BI19, CWW18, CFGRB17, DK13b, DM10a, GHBC14, IGL^{+}12, KBS^{+}19, KBS^{+}20, KM19, KGM^{+}15a, MB19, PWL^{+}11, PD16, RLK10, RV16, RRRTR10, RRTR12, RRRRSPTR16, RSR11, RHT18, Sch18, Sch21, SAB^{+}19, SSvdM10, SKPK17, SSS15, SOBC12, TGB^{+}18, TWW19, TC11, Wax11b, ZM10]. diffusion-driven [KGM^{+}15a]. diffusion-reaction [ZM10]. diffusive [ZZ14b]. diffusively [Lab16]. diffusivity [TSMB14]. digesta [TF18]. digraph [cLCJ^{+}10]. dihydromethyltrisporate [EBSW17]. Dilemma [CZ14, GCS12, Lai18, MRPN12, ZLY^{+}17, BLS^{+}17b, CW14, FWW18, FNH10, JC16, Lai11, MN11, PCN17, TP10a, WLW19, WLW20, YA14, ZZC10, ZGWL14, ZZCZ17, MI19, PM10a]. dilemmas [AS12, CGD13, MR12, NO14a, NO14b, NR16, SSC13, WMD^{+}11]. Dilution [HvLPdB15, PPW16, CFR^{+}14, CMR^{+}18, NKL10, SKAG18]. Dimension [ZS10, BVK10, PHTP^{+}12, RBB16b, SKK^{+}11]. dimensional [AGPK13, ATB14, BTH^{+}13, BG12, DHHP14, EBP15, FZ14, FJC^{+}10, IMD16, IS16, MPY14, Mon18, MAA18, RMM^{+}16a, RBB16b, SMZ^{+}17, VBVD17, ZR16]. dimensionality [QW11, YYC19]. dimensions [Edu16, HB13b, RVA^{+}11]. dimer [FS11b, WLD^{+}11]. dimeric [GWX17]. dimers [SRS^{+}15]. dimethylarsinic [BTF19]. dimethylarsinous [BTF19]. dimidiata [MMG^{+}16]. diminishing [FOT^{+}15]. dimorphic [YHI14, ZAL^{+}19]. dimorphism [Mor13]. dimorphisms [AM10]. dinoflagellates [JSWY19]. dinosaur [PC17]. dinosaurs [BJB18]. Dinucleotide [FGGS15, FS16b, MPP16b, WZWY11, ZK18]. dinucleotides [GZY12]. dioxide [LCMC14]. dipeptide [KHKI17, LD11]. diploid [BID15, Fis19, HT16, Ots11, SLL18, TT10b]. diploids [Gra14]. Diptera [YMZ18]. Dipteran [FH10a, FH10b, TB11]. Direct [DLM^{+}19, HB13a, SNC12, CB16, GVSLG16, GSF13, GN10, KI17, PVGA12, ZLY^{+}17]. direct-acting [GN10]. directed [ABD^{+}15, SSZB15, SGW17, SMJS14, VSP11]. direction [Sel12a]. directional [BPG^{+}18, BHP17, FH10b, ID14, RCD19, SYI17, SG18, WD11]. disassembly [BZ10]. disaster [GA16]. discoideum [MGT17, PWE15, SSS17, Tre19]. discontinuous [GT15]. discordance [SPG^{+}18]. discount [DHP11]. discounted [LW15]. discounts [AK11]. discourses [SS19]. discovered [PGGvdB12]. Discovering [Ayd18, SFV16, SG15b]. discovery [CTSL11, GBC^{+}16, GDPPSS^{+}11, wLW10, SDRA^{+}15]. discrepancies [CK17]. Discrete [FS16c, GI15, Yat14, ANMH11, AEM13, AEM14, BSKV18, CL17b, Che10, Cor16, Cox10, GLR^{+}17, GW13, Gol10b, GPT17, Iij12, JMK^{+}17, KA11, LP12b, LC16, RA18, SKD^{+}10, TSS16, ZLDZ13]. discrete-cell [Cox10]. Discrete-time [FS16c, Cor16, GLR^{+}17, LP12b]. Discretionary [MG17]. discriminant [KF15, Pav14]. Discriminate [MIH16]. Discriminating [FL13]. Discrimination [KHK15, AH16, AHJ18b, Coh15b, CBP12, FHW^{+}10, IL12, Mei14, NNK^{+}15, PDM17, SC12, TCYY^{+}12]. Discriminatory [PDB^{+}15]. discs [Che19]. Disease [GA16, GLZ^{+}11, GGM15, MPLK14, AZOLVH18, AOM19, AKR^{+}18, ADS^{+}19, AB16, ABN15, BM12, BN15, BGLR^{+}19, BDR10, BZ10, CFZ14, Coh15a, CG14, DFM^{+}14, DGM15, DD13, DSTD19, EME^{+}16, ERT13, FGJ10, GGL^{+}16, GRR^{+}14, GMS^{+}13, GMR16, HAuR^{+}18, HYSG19, HRHAAA15, HSG^{+}18, HZL^{+}11, HvLPdB15, IMW15, JAB18, KKD18, KGJ12, KEG17, KST17, Lan16, LN19, LNH13, LSSG10, MV10, MPS16, ML10, MMG^{+}16, MYOS14, MvAKR17, MW13, NKM^{+}12, NG17, NC15, OA12, OA11, OVKL14, PS13, PBP15, PGKZ17, QZ14, Rao12, RSH19, RB14, RBF15, STJG12, SKPK17, SY11, SLJ^{+}10, TB15, WB15, WN19, WBRHE10, WDL^{+}13, YC15, ZLL17, ZJY19, ZZL19, ZLT^{+}19, Ana11]. disease-behaviour [JAB18]. disease-free [BDR10]. disease-induced [ZLL17]. diseases [AL17, BH13b, BB10, BPR15, CR14, CBC^{+}18, HHN14, LWRE14, MCRG18, MS12b, MS12c, OSV^{+}16, PS13, PKZ12, RSD12, RB15, Ros11, TXY^{+}12, TDW16, WW12a, YZMY18, ZC14, dM10b]. Disentangling [VZ19]. disequilibrium [LMW^{+}10]. disinclined [PC17]. disintegration [HBSF11, LT11]. dismutase [DWM15]. Disorder [RFS^{+}15, HGM15, JPBM17, RM17a, RCL^{+}10, RA17, VHM^{+}17]. disorders [PG18a, PG18b, PPC^{+}17, Wal10]. disparate [GP11a]. Dispersal [AdVG15, BBM^{+}97, HB12b, KA11, Kis16, LS10c, AWRD15, Alv19, AGGME18, BS19b, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, CVO^{+}15, DGJ14, DGJ15, EKFV18, FPD15, Gar10, GH11, GPR^{+}16, GPT17, HGHL10, HM11b, HASM17, Jab10, JHE15a, JHE15b, KK17a, Kis10, KJ11a, KW15, KJ15, LNH13, LK19, MSZW19, MCPF12, MAL^{+}11, NG11, NG17, NP11, NPS17, NPS18, PMD11, PGH11, PBA12, Ros10b, SCF^{+}18, SRCF11, TRG16, VZ19, Wan19, WK15, WK17, YLW^{+}14]. dispersal-limited [Jab10]. dispersers [NP15]. dispersing [HH16, HHD^{+}16]. dispersion [GMNY14, Js12, KKWA18, Kro11]. displacement [Che10, PIPB10]. disposition [Zho11]. Disrupting [OA11]. disruptive [LHD15]. dissect [CRLH^{+}19]. Dissecting [LCDH15]. Disseminators [CWW18]. dissimilar [ABP^{+}11]. dissimilarity [DLYZ11]. dissipative [MDVT10]. dissociation [FLW16, IIVR10, Xie13b]. dissolution [ZSW^{+}14]. dissolved [NR14]. distance [ABP^{+}11, BdABY13, BS19b, BHSB11, BvLH14, CLC11, CR16, DGMY18, Fan11, FPD15, Hua16, LTSTD15, LLL^{+}17, PGH11, QZY17, YW13a, YHY14]. distances [DGMY18, Mul11, Mul12, OGO19, SENS^{+}17]. distinct [PCT11, PCT19, Pan19, RMF^{+}18]. distinction [IL12]. distinguish [Mas18, SQS^{+}12, SY11]. distinguishability [LKO^{+}17]. Distinguishing [AGB^{+}14, MBM11, LLS^{+}15]. distorters [YTK10]. distortion [LLSO13]. Distributed [KS15, CCS15, HT19, ISZ18, LCG^{+}15, PDT^{+}10, RKMG12, RPCW18, TAL19]. Distribution [OBK^{+}11, POLT12, SS12a, ZLZ^{+}15b, AD16, BKR14, BH13a, Cam15, CPF^{+}13, CW11a, CW11b, CF11, DYS^{+}13, DK13c, EPP19, EBX17, GVGC15, Gol16, Kär11, LdLK11, MSK10, Meh17, MYLK11, MC15b, Nak12, NYSM12, Par15, PWHW16, Pia19, RTRRS^{+}17, RRB10, SSI^{+}16, Sch14, SJSG13, Su16, SM11b, SCT12, SSRA16, TLR14, Woh15, XGZ17, YZW10, YZZ12, Yin17, ZMS17, ZL18b]. Distributions [ZS10, ALH10, BR15, DLL^{+}18, Don13, DZW10, FVC15, GFM^{+}19, Gol10a, JL11, KS14, KS15, KdSUS10, LLLV11, LLG16, LSC^{+}18, MJS11, MTKY11, MPY14, MS16, MDM^{+}10, ML12b, MB16, Mul11, Mul12, Por13, SSH^{+}19, SY17, SBM^{+}12, TKKE19, TLM14, Wei12]. districts [LG18]. disturbance [HHD^{+}16]. disturbances [BZA16, GDF17, KSSM11, VAT18]. disulfide [KJH^{+}19, MdFDM10]. diurnal [HCS^{+}19]. divergence [AP13b, LZ18, Oka12, Rei12, SS15d]. Diverse [STN^{+}19b, AAGO^{+}17]. diversification [DI10b, GHBC14, HD11, IMD16, LS13a, NI19, PMR11, TGLK19, ZYD15]. diversified [IS17]. diversion [LRL^{+}16]. Diversity [Mal15, RT19, Sat10, CB17a, CdPD11, DLSD15, Edu16, Ezo19, FR13a, GWE11, Gre16, HDNB10, KDFM14, LS13a, MFMPS10, MP13a, MFKS13, Moo14, NdORC11, SC10a, SPLP12, SSI^{+}16, SAM17, SS12a, SPCM18, Su16, SLW18b, TD18, Tjø10, WL12c, WFZW13]. diversity-area [NdORC11]. diversity-stability [FR13a]. dives [IPMH12]. Divide [JH11]. divided [BCK19]. diving [Tra16]. division [AUE15, AKdV^{+}14, CL13b, Gon13, HHG14, HTM16, HMJB15, LCTG15, LAG^{+}14, LS15b, Lee16b, LL17a, LPTR14, MMJB17, NF14b, NSKO18, OKVN18, PW19, RC13, RSBB10, ST16, SLM17, Wal12, Zhu11]. divisions [STH18]. divorce [MB10]. Dkk3 [MPZK16]. Dll1 [BF13]. DMA [BTF19]. DNA [FS16a, dOGL13, JS18, ZSCL16a, ASK17, AFS19, ACH15, BBT^{+}15, BKL^{+}15, BWS10, BSB^{+}13, CK16, CL17b, CLL18, Den19, DDLW10, DLYW13, DHB15, Eli17, FHW^{+}10, FLM18, FS15a, GSS19, GY18, dOGL10, HRC^{+}12, HMMSRSD15, HA15a, HYZ^{+}15, HYZS14, HYW11, HY16, JS15, JS17, KKD18, KBE^{+}13, KRDJ15, KP16, KK12b, Kue16, KDL16, LZA^{+}19, LF10, LLDW14, LTP19, MPC12, Mor11, NHSX14, PRN10, QQL10, RSS^{+}18, RFM^{+}12, RVP19, Sel14b, Sos18, SCS10, SCLC13, TBB14, Tri10, TS17b, WZWY11, Xie11, XM11, Xie15, Xie18, YZZ12, YLC16, YCY14, Yin17, ZARK19, pZWZ^{+}16, ZSCL16b, Zol14]. DNA-binding [RSS^{+}18, pZWZ^{+}16]. DnaA [ZS12]. DNase [ZZC^{+}17]. Do [CS15a, GSL13, KLN^{+}12, STJG12, STH18, AM10, AD13, HHMT19, Hou15, LLW15, LRH15]. docking [BGCB12, FHG15, FCS18, GBC^{+}16, HAuR^{+}18, KJM17, KHH^{+}17, KGM15b, PYG^{+}19, RMM^{+}16b, RM19, SW13b, TWR^{+}18, TLT^{+}15, TAR16, TMS13, TA16, YNY^{+}10]. Does [BBJDS11, GF11, Oht12, RDMP11, RSC14, TMHM12, ZCW13, AR11, BEK10, HNP18, HNR14, KMH19, LCHMP16, POP12, RWH16, Sat18, SM16, SLL18, SS11]. DOF [HMMSRSD15]. dog [LLN^{+}19]. dolphin [SZ15a, SZ16]. domain [ADS^{+}19, CL14, DS15b, DAVS11, HWW^{+}14, HBSF11, LHL^{+}14, NZZ19, NHTS14, PSAA13, SCsMzX18, SRS18, WHZ^{+}17, ZJSC16, ZLW^{+}19]. domain-based [ZJSC16]. domain-domain [PSAA13]. domains [ASGE14, CL13a, DI13, Di 18, Di 19c, DV12, EKFV18, IK15, JSZ12, Kri18, MK18b, SGKCP19, SV17, Woh15]. domestic [HT13, NKL10]. dominance [BI10, Cam11, CLA17, FVC15, GDJC11, LBJ13, NBS^{+}13, PS12, SGD12, SGCD14, SLL18, TR12, WTZ^{+}11, YRD17]. dominance-subordinate [NBS^{+}13]. dominant [WRH^{+}16]. dominated [KH19, SCF^{+}18]. dominates [GJ12]. donation [HSLW16]. done [OO11]. donghaiense [JSWY19]. donor [MSRW16]. don’t [Pan11]. Dopamine [PJL14]. Döring [DFMR19]. Döring-type [DFMR19]. dormancy [Aya12, DBM^{+}18, Kar16]. dormant [MM12]. dorsal [ABH^{+}11]. dorso [SSKK17]. dorso-ventral [SSKK17]. dosage [Das18, HAB^{+}17, YD14]. Dose [RLW^{+}14, RKK^{+}19, BH13c, CHN^{+}15, GMS^{+}13, HPvdB17, LH14, LHD^{+}17, LIB^{+}17, SW19]. dose-dense [LIB^{+}17]. Dose-dependent [RLW^{+}14]. doses [RHJ11]. dosimetry [BHH^{+}14]. dosing [FM10, LN13, NSH^{+}10]. Double [PWHW16, BKL^{+}15, CL17b, KFL12, Kue16, KDL16, SNCM12]. Double-layer [PWHW16]. double-strand [BKL^{+}15, KDL16]. double-stranded [Kue16]. Dove [Ayd18, DN18]. down [ARB13, LG13, MPS^{+}11, SC10b, XNW17, YRD17]. downstream [ZZL19]. Downy [PS12]. doxorubicin [EPJ^{+}09, EPJ^{+}11]. DPC [KHKI17]. DPD [TAB^{+}13]. DPD-PDE [TAB^{+}13]. DPP [RSS^{+}18]. DPP-PseAAC [RSS^{+}18]. Dr [BD10b]. Drafting [SZ15a, SZ16]. Drag [LWY^{+}16, JLH^{+}15b]. dragonfly [LS10b, LS14a]. dramatic [VK10]. draw [cSGFB17]. drift [Bro13, FCS18, Gal11a, JJ15, NMI18, PMR11, RG10, SMdA16, Wax11b, ZW16]. DRIN [PJL14]. drive [EA17, HMM17, Kun16, MH12, OGA16, TR12, XKCG15]. driven [Ana11, ACCR11, AP13b, CFGRB17, Chu18, DHHP14, GPT17, Han12, HH16, HLH^{+}18, KB15, KFA12, Kin18, KGM^{+}15a, Kue16, LZA^{+}19, MC16a, MH13, MA11b, PSJ15, SK10, TS17a, TWH^{+}13, Urs14, US12, WF18, WDL^{+}13, YMZ18, ZLB^{+}18, dM10b]. driver [BST14, DBB14a, HRCA19, MGB17, Pol12]. drivers [AHP^{+}19, FR13a, LPP^{+}19]. drives [EA19, KEKB18, KMCH14, RM13]. driving [HM15, WDL^{+}17]. drones [LWY^{+}16]. Dronpa [HYZS14]. drop [BKR14]. droplet [ZWvG10]. Drosophila [Che19, DM10a, DM19, GAL11b, iHIM12, MBM11, RGG12, Sel12b]. drought [JMK^{+}17]. drowsy [BVK10]. drug [BS15a, BJ18, CWM11, CDO17, CTSL11, CL18, CGW18, DK13a, GY19, HNP18, HAB^{+}17, HZC^{+}10, JSDEK14, KHP^{+}12, Kle14, KS10, LP19, LLZ13, MPM10, MVH11, NLB14, OA15, PYG^{+}19, PRV^{+}14, RAF^{+}19, RDSD^{+}12, SBM^{+}16, SMJS14, TAR16, WCM15, XMWC13, XSMF12, XXD^{+}17, YD14]. drug-resistant [CL18, LP19]. drug-target [RAF^{+}19]. drugs [GSSBF18, GSV16, LF19, LBW^{+}13, LT14, MAFB18, PPC^{+}17, PBSM19, RCL16, WHH17]. dry [DDS13]. dsDNA [AN13]. DSSB [AD15]. DTI [PH13]. Dual [MAMEA15, EI13, For10, HZL^{+}13, MNSZ16, OGA16, TSF^{+}19, WMT16, ZGY11, ZCA^{+}14]. dual-cell [For10]. dual-lattice [EI13]. Dual-porosity [MAMEA15]. Duality [BYJ17, FVTS16]. ductal [GCSP17, METC12, NWB^{+}10, SGGY10]. ducts [FKMG15]. due [AJC12, ABR11, BSP18, CK17, Cza14, GN10, KKS^{+}14, MK14a, SMM15, SBM^{+}16, SRCF11, STKE12, SL19, VB19b, WZ18, YiTS19]. dung [BBR12]. duplex [LF10]. duplicated [Das13]. duplication [HNV^{+}16, KII19, Ots11, SAZ^{+}14]. duplications [bCRS17]. duration [CF11, LMCW18, MGS16, PGS10, YTGW16]. durations [BR16]. during [ASC15, AHAFA14, ALD^{+}11, BTO14, BAGG14, BRN15, BTK14, BBRT12, BKPV15, BFGAGA16, CLPZ14, Che12, CL18, Chu17, Cle18a, Cle18b, Cox10, CHN^{+}15, DPRSS11, DL15b, DAA17, DB18, DDF^{+}14, EJK16, FWB^{+}12, FM10, FPR10, GWBL12, HWMT17, HM10a, Hal17, HPM^{+}17, HMH14, HHE19, HRG11, IS12, ITR^{+}18, JV14, JMK^{+}17, KHB^{+}18, KCS^{+}15, KG12, Kuz19, KK18, LCTG15, LAH^{+}16, Lee16b, LMC^{+}13, LSM17, LO15, LRA^{+}13, MYK^{+}11, MMLK11, NPS17, Orr19, PHH18, PDRC16, PC10b, PVGA12, RRL^{+}13, RLW^{+}14, RJSC18, RS14a, SEK^{+}10, SK15, SKGM19, SHB^{+}17, SK10, SGS15, SAGC12, SSKK17, SP14, SNCM12, Tak17a, TEW12, TDE12, dlRVFK^{+}16, VF12, WDL^{+}17, Wil11, Xie13b, YM14, ZTWL12, ZWY15]. durotaxis [SVMN11]. dusty [BTK14]. Dwarf [YSYI13, YYST13]. dwell [Voe18]. dwell-time [Voe18]. dyad [GU13]. Dynamic [CS11a, CLA^{+}16, DFM^{+}19, FWB^{+}12, FRG^{+}13, FWR19, KV15, KAZ13, LS14a, LS19, MMJB17, NHS^{+}16, RSH19, SK15, TI12a, WOB15, dM10b, AJD11, BTG^{+}15, BB10, BHH^{+}14, BTW11, BC13, CTW10, CW15b, CLP11, Cre10, CMGN17, DAVS11, DB19, Don17, EBS11, FD17, GAL11b, GHBI16, Gol10b, HAuR^{+}18, HH16, HHD^{+}16, JSWY19, JRMS12, KMZR18, KK12a, KMM18, LCJ16, LM14, LDWW14, LBS^{+}14, LPD^{+}16, MHC16, MN11, MP13a, MPC12, MZAI19, MAS15a, Mun12, PDG17, PSS^{+}10, PSS^{+}13, PKL12, RRBM12, RS12, SNS17, Sat10, Sat14, SGA^{+}12, SVSP15, SHW16, SA14, SY15, TTG19, TDHC^{+}18, TB11, TXWW12, TXCW15, Tay16a, TLK19, WCJ11, WHH17, WYL^{+}18, XSKA17, XS13, YA14, ZMNd^{+}10, ZCSRG12, vV18, AKNP18, TTB^{+}18]. Dynamical [GCB17, LMYL15, LLL13, PaCZ10, VLCM10, AdGM12, CST^{+}12, DFM^{+}14, FLA^{+}19, GB13, GML10, KLB^{+}16, Lai18, NBA^{+}18, OA15, RWSB11, Sim14, Xie13a, ZJY19]. Dynamics [BBM^{+}97, CFF11, FTS15, FHS^{+}14, GL14, GM16a, HJR12, ITR^{+}18, KGJ12, LCT11, LLJ18, LPTR14, MMH^{+}12, MFKS13, MW13, Por13, RB14, SPSP12, SHK14, SC16, TF14, WSVT14, Xie13b, ZYJL18, ASC15, AR11, AGPK13, AOM19, AD15, ADS^{+}19, ABKS11, AHKN10, AH12a, AY15, Akç12, AKS^{+}19, AMBH^{+}13, AMSSG16, AMFR^{+}16, AGRR^{+}15, AABS16, AN12, AdlPLMZ13, AP13a, APW10, AM12, Arc14, Arg12, AWK15, ALD^{+}11, BYJ17, BC19, BPS19, BA10, BRG^{+}12, BCPL10, BM12, BDS12, BLV18, BPFR16, BZCS10, BBI14, BWY^{+}17, BHKR14, BN12, BN15, BMB^{+}18a, BJWL15, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, Bon10, BK19, BBRT12, BMF^{+}18, BPLM12, Bue15, BHO^{+}18, BJ18, CHD^{+}10, Cam11, CML10, CC11a, CMS^{+}19, CFZ14, CN12a, CS18, CZN12, CNL14, CFL^{+}15, CWF^{+}16, CVPCV^{+}15, CSM^{+}14, CS14b, CG14, CA18, CH18, CPPLAB18]. dynamics [CHN^{+}15, CSZT12, CGW18, DB10, DS19a, DMM^{+}14, DMO^{+}17, DM11, DSLR18, Del12, DL12a, DEN^{+}11, DG16, DGJ14, DGJ15, DHBS19, DRC11, DCN13, DCR18, DHB15, DFMR19, DP13b, DHRA10, EKFV18, EUM^{+}16, EN15, ECP^{+}16, FH10a, FH10b, FHLM10, FK18, FMS^{+}10, FtL18, FGGT15, FKV19, GLR^{+}18, GAL11b, GVGC15, GAGP^{+}14, GCSMMG^{+}12, GDS11, GMNY14, GXF13, GC18, GMMN18, GP12, GY18, GN10, GAV15, GPT17, GAY^{+}11, HBK12, HT19, HSD11, Han12, HI19, HII^{+}19, HBVP15, HNO18, Hau10, HNR14, HTM15, HE16, HKS15, HFT^{+}18, HB16a, HAH19, HBT13, HVL11, HR14, HSM17, HH16, HWPL12, HHE19, HB13a, HBW^{+}11, HTB^{+}13, HHT^{+}19, HPW^{+}13, HWL15, HHN14, HSLW16, Hur12, HR12, IMM16, Iij12, IGL^{+}12, IS12, IK16, Jab10, JB18b, JYY^{+}18, JSSZ12, JL16, JSB15, JMK^{+}17, KJM17, KHH10, KSP^{+}14]. dynamics [KHP^{+}12, Kea16, KMHdlP10, KGP^{+}15, KFA12, KB19, KGK10, KR15, KLJ17, KKU^{+}10, Kla10, KDMK16, KL10, KW10, KKYV18, KGC18, KGM15b, KSK^{+}17, LJM15, LG10a, LS10a, LCC15, LvBJ16, LP19, LHV16, LW15, LMCW18, LGR^{+}12, LYW11, LHD^{+}17, LPH18, LPH19, LP12b, LJS14, LH18, LJW^{+}16, LCDH15, LC16, Luk14, LW18, MTKY11, MPJH13, MZ17, MRPN12, MKKS14, MDVT10, MPP^{+}16a, MCC^{+}17, MBLV10, ML14, MMAS13, MFG^{+}13, MKMT14, MS10a, MS18, ML10, Mob10, MSE^{+}14, MS10b, MGM13a, MZMM^{+}14, MPNP12, MMRCCC10, MS14c, MA11b, MBP16, MI11a, Mou12a, MYOS14, MCCC^{+}10, MvAKR17, MSIR10, MG15, MSL^{+}16, NIiT19, Nak16a, NCLB16, NSO15, NTED^{+}19, NZZ19, NG17, NC15, NP13, OB15, Oht10, Oka15, OUMA10, OB10, OOY16, OAJK10, OBHS19, PAA11, PPBD10, PM11a, PBR17a, PR18, PT10, PNL15, PVB^{+}10, PSV17, PW18b]. dynamics [PCL^{+}15, PGKZ17, PZLF19, PM17, PMYHR12, PP17b, PKH11, PRM12, RW16, Rad16, RMF^{+}18, RN12, RFS^{+}15, RW10, Rem15, RB13, RBHK14, RBB16b, RSD12, RG17, Ros10b, RM19, RRÚJ19, SM14, SSPP15, SK19, SU11, SW15, SFT14, SBR10, SVB^{+}10, SI19, SW17, SBR16, SV17, SCF^{+}12, SKPK17, SMJS14, SW11, SKS15, SC11b, SAB17, SOCF14, SG15a, SP11, SY17, SSME10, SGS15, SLC12, SLO10, SLS17, SS11, SK11b, SSN^{+}14, SYY17a, SYY17b, SST19, SPL14, TED10, TBM^{+}13, TLCZ12, TCR13, TGC^{+}17, TDKJ15, TRM^{+}14, TK19, TMS13, TMD14, TRJD19, VSD^{+}17, Vel17, VABS18, WWY12, WCPF15, WCW15, WJL^{+}16, WML16, WYKT17, WML^{+}17, jWlGQ19, WW19, WMHH10, WGC13, Wod18, Xie11, Xie18, XYZ15, Yam16a, YI18a, YNY^{+}10, YBT^{+}17, YFB^{+}12, YST14, ZARK19, ZBA14, Zha10a, ZZ14a, ZZ18, ZN18]. dynamics [ZKTK10, ZWG10, ZSW^{+}14, ZLM12, ZSCL16a, ZSCL16b, ZZR10, ZZR13, ZRZ15, dTKDV15, vV11a, vdBD10]. dynein [CL10, SS15f]. dysfunction [LFW^{+}18, WHS^{+}13]. dysplasia [MGM13b]. dysregulated [HZL^{+}13]. Dytiscus [Whi11].

E-cadherin [WGS10]. E. [RG10]. each [RG17]. Ear [CZJQ14]. early [Bau18, CCB11b, CPV16, CTSL11, CEP14, DI13, DTGD19, DM10a, DCR18, Dun11, FD17, Fra16, JAB18, KG12, KC11a, Lab16, LJ15, Lev14, LJS14, MSBB13, PHTP^{+}12, PGGvdB12, PR17, SBB18, SCS14, Tre19, VDD12, WBMM18, YM14]. early-detection [CPV16]. early-stage [Lab16]. Earth [Di 18]. East [DB18, EAN14, LCJ16]. East-west [DB18]. eating [DDF^{+}14, JPBM17]. Eavesdropping [KL10]. Ebola [Ada16, BGW15, DE17, HSLW16, JB19a, LN19, LLD^{+}17, PZLF19, RPP16, TDW16]. EBV [CRLH^{+}19]. EC [RMSTG13]. EC-numbers [RMSTG13]. eccentric [GAPK10]. echinococcosis [KKUM10, WZJ^{+}13]. echinoid [CS11b]. echolocating [VP18]. echolocation [LA15]. eclectic [TT17]. EcmPred [KPK^{+}13]. Eco [HBT13, KK13, Urs14, AIB^{+}19, FtL18, HCS^{+}19, KGB13, KDMK16, LH18, SKAG18, YBT^{+}17]. eco-evo [FtL18]. Eco-evolutionary [HBT13, KK13, AIB^{+}19, KGB13, KDMK16, LH18, SKAG18, YBT^{+}17]. eco-genetic [HCS^{+}19]. Eco-hydrology [Urs14]. Ecological [BUL12a, GLR^{+}17, JMS12, KEKB18, SY15, TCW14, APBS15, All11, ALH16, BWB11, Bal10, FB18a, FR13a, GP12, Gri19, HT18, IS10, JHE15a, JHE15b, LS14b, MMB14, PM15, PSV17, RW15b, SRS18, SC16, VK10, VR17, WH11, WF18, vVE15]. ecologies [CN12b, DDP12, RDP16]. ecology [BCK19, HL11, KA15, Lio16, MRPH17, MW14, MN14, Pon12, TGL15, ZADB15]. econometrics [HYSG19]. Economic [DRW14, TDE12, CJRR11, ESE15, GKG^{+}18, KG13, MDM16, PT10, SK16a, TLW18]. economics [CT14c]. economy [Pon12]. ecosystem [AMRRI10, BSH19, BCF16b, BFM10, BJ17, Edu16, GL14, GK13, LCMC14, MK14a, MA13, MP13a, RM10]. ecosystem-based [MA13]. ecosystems [ABLH18, BDLR12, CC11a, CCB11a, FB12, GFH^{+}18, JB18a, MGML10, MPP^{+}18, PMSY17, QA15, SI11b, Sie18, THM10]. ectotherms [WSRG18]. Edge [WMCL18, HD17, YZMY18, ZZC14]. Edge-based [WMCL18, YZMY18]. edge-biomarkers [ZZC14]. EdgeMarker [ZZC14]. editing [DMCP14]. editor [JN14, gSxFH^{+}12, ARD12, JAHKH12]. Editorial [Sza15a, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o]. Editorial [Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano13x, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano14x, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g]. Editorial [Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z]. Editors [Kir11]. Editors-in-Chief [Kir11]. Edward [BCF16b]. EEG [BVK10, Ing16, SKK^{+}11]. EET [FD11a]. Effect [AP19, ABD^{+}11, BSC19, BBI14, EPP19, FMS^{+}10, GTC19, JJ15, LFD11, MLCH11, MGC13, MAR^{+}17c, SGAM12, SGCD14, TMM10, TI12b, TLR14, WHHS15, Wod18, YTK10, ZLL^{+}12, ARG19, AGRRTR11, AGRR^{+}15, Arc14, ACCR11, ABK^{+}12, Ayd18, BB15, BBB18, BCF^{+}16a, BG12, BHKR14, BH11, BD10b, BE14, Cam15, CMCS18, CWW18, CLA^{+}16, CMN^{+}11, DSA^{+}16, DZE11, Eli17, ECP^{+}16, GHBI16, GH11, GM13, Gon13, GJS^{+}10, HWMT17, HFT15, Has14, HB16a, HMM17, HTB^{+}13, JRB^{+}16, Kar15, KMZR18, KMCJ17, KMM18, Lai11, LS10a, LCCK12, LN13, LMJ^{+}16, LBJ13, LS11b, MPM10, MGT17, MB15, MFMB12, MBP16, MI11b, NKL10, Nak14, NSKO18, NdORC11, NI19, NC15, Oko15, Oko16, POW17, PSD^{+}18, PDW10, QA15, RGG12, RC16, RRTR12, RRRRSPTR16, RAR19, ST17a, SSF15, SBCR10, SXZY12, SNY^{+}13, SK16b, SM10b]. effect [TB16, Tal12, TIY^{+}11, UD10, VAT18, VSW10, WTC10, WMT16, WMHH10, WHH17, WCM15, WFM^{+}13, WFC^{+}14, XTZ^{+}13, YBH^{+}19, YLH12, ZMW10, dG10b, dSRM15]. Effective [AFS19, Gre16, SS18, AKS^{+}19, Bar16, CPH^{+}14, CG11, Ghu18, HI17, LCTG15, LGPS17, MA13, NF14a, Nak12, SSJK18, Sch14, Tak16, VCF^{+}19, XXD^{+}17]. Effectiveness [EN15, GAY^{+}15, SP13b]. effector [ME15]. effectors [d’O12]. Effects [AdSG14, ASGG17, BKSW16, CHL^{+}19, CF11, DGJ14, EBX17, GCZ^{+}12, GC11, HE16, HSR12, HY10, IIKT13, IEN15, Js12, KO18, KMHdlP10, KJ15, KII19, LdLK11, LG18, LFM11, NIiT19, NSO15, NNK10, NJP18, PMYHR12, RVA^{+}11, RHT18, STA15, SKPK17, TM14, TA13, VSD^{+}17, WK18a, YLW^{+}14, ZR16, ASC16, AHD^{+}18, AMBH^{+}13, AMFR^{+}16, AH17, AAJ15, ARM18, BGM19, BLV18, BGL10, BSV^{+}10, Boc12, BSS11, BIMC17, BMM^{+}14, BCPG18, BZ10, CLK18, Cle16, CSDR15, Cor16, CCAdS13, Cox10, DH16, DB18, DHB15, DS12, EBE14, FB12, FSW^{+}16, FGLS10, GMZM15, GVSLG16, GCvWE^{+}14, GS17b, GBRS19, HD17, HPP10, HT18, HHG14, HAH19, HH10, HWPL12, HWGT15b, JYY^{+}18, KSU^{+}18, KDG13, KA11, KPD19, KEKB18, KF11, LTHEK12, LL14, LLDW14, LWY^{+}16, LS14b, LLY15, LS11a, LDF11b, MG14b, MKBE17a, MKBE17b]. effects [MKRE18, MEJ18, MMRCCC10, NR14, OO17a, PDSP13, PSP^{+}12, PHÁBAI^{+}16, PBB10, PN18, PMK15, PRV^{+}14, PZLF19, PGHC12, PMCS16, PLF19, PML18, RRL^{+}13, RLW^{+}14, RHJ11, RTEKG15, RTEKG19b, RLSN14, Sac13, STI13, SHvHB16, SAB10, SJSK18, TSdL14, TDHC^{+}18, TCW14, UI10, VZ19, Vaz10, WXWL12, WML16, WTC16, WPA17, WVO^{+}15, YNY^{+}10, dM10b, dMP11, d’O12]. Efficacy [RMSTG13, GSSBF18, KMA10, KRR14, LBW^{+}13, LZL^{+}19, MDB12b, NDZMA14]. efficiency [BH13a, CLP11, DK13a, GMM^{+}13, GSF13, HLH^{+}18, JG14, LFB^{+}16, LJM^{+}17, LMJ^{+}16, Orr19, PBRW11, PVCEC18, SHW16, TP10b].

Efficient [NW10, SNS17, CLG^{+}10, CLG^{+}11, FD17, GRCdL16, HXL16, IMH15, IL13, Kha10, SW13a, TLY^{+}19, XGZ17, ZMNd^{+}10]. efflux [FKB13]. effort [KJM17]. efforts [HJR12, STKE12]. Eg5 [HNL^{+}11]. Egalitarian [Sch10, TT19]. EGFR [GY19, ZZ18, ZZL^{+}11]. EGFR-ErbB3 [GY19]. eggs [ICG16]. eggshells [BUL12a, JMS12]. egoism [SSC13]. EIAV [CS14b]. eIF [KSvdH18]. Eigen [Dal17]. eight [BRA15]. either [RS12]. EJC [ZLZ^{+}15b]. elaborated [CN12a]. elaboration [FHSG15]. Elastic [HBT13, KF10, RS12, SYR11, SDKF12, BB15, JH12, KFG^{+}14, MH14, ØW16, RA10, TI12b]. elasticity [CMJD11, MHH13, RS14a, RGCML10, SA17, SP13a, VH11b, ZHAK14]. elasto [PAV10]. elasto-visco-plastic [PAV10]. Elastosis [EJK16]. electric [HCK^{+}11, LL17b, PLHL19, SKGM19]. Electrical [Hal16, BZN17, CW11c, KSSM11]. electro [HGGCR13, LHPF18]. electro-mechanical [HGGCR13, LHPF18]. Electrodiffusion [GJ11]. Electroencephalographic [IPS14]. electrolyte [Smi11]. electromagnetic [Ghu18, LGPS17]. electromechanical [PW18b]. Electromigration [SKGM19]. Electron [RS19, MTE15]. electroosmotic [DFG^{+}18b]. electrophysiological [EME^{+}16, WMPF^{+}15]. electrophysiology [CRGS18, ESG^{+}16, RPGG^{+}19, RLSM17]. Electrostatic [ØW16, Smi14, DK13c]. electrotaxis [MDD13]. elegans [BLZ^{+}19, DI13, OYIM10, ZLZ^{+}15b]. Element [PWC^{+}12, BDID^{+}12, BSM^{+}14, CTPB10, FS11a, HB12a, JTSG14, KHB^{+}18, MH14, pNMK^{+}10, TF15, ZLZ^{+}11]. Elementary [DG10a, PLF19, FT12, MRS14]. elements [BJO^{+}16, SDRA^{+}15]. elephant [CMMR13, WHYMG17]. elevated [Kun16, LCMC14]. elevational [AGPK13]. elimination [CSS^{+}18, DSTD19, GM19, GGM15, LW17, LTL^{+}15, PBSM19, VKKA12]. Elman [MP12]. elongation [BS19a]. elucidates [HLTR19]. Elucidating [EJ16, dSKBS10]. Elucidation [OTGT10, VGPS18]. Embeddability [RLFS18]. embedded [AAJ15, AGGME18, MK18b]. ember [SVCS10]. embodied [RBG^{+}17]. Embolization [Gou15]. embryo [GLZ^{+}11, SCS14]. embryonic [ALD^{+}11, BF13, BJWL15, TRM^{+}14]. embryos [DI13, KKG^{+}14, vHHKB14]. emerge [CNM^{+}13]. Emergence [Mor10b, NF14b, RBB16a, SKS15, SA14, TR17, FCC^{+}10, FGPR10, GWNW15, GFH^{+}18, GGM15, HKS16, IDI11, MHX^{+}14, Mas16, OS13, OMO13, PR17, SS19, SCH^{+}19a, Tak17b, TT19, Unc11, WD12, WK17, Woo10, WSVT14]. Emergent [BFW10, WPPD16, BTG^{+}15, KL11, MT14, NvD17, TXY^{+}12, WdLS17]. emerges [SV14]. Emerging [CFZ14, ZKP15, AOM19, EPP19, LBGW13, TDW16, WDL^{+}13]. emerin [YX15]. EMG [Gol10a]. emigration [PGH11, Wei17]. Emiliania [HWGT15a, HWGT15b, HWGT17]. emission [BGF^{+}14, LDB^{+}14]. emissivity [BUL12a, JMS12]. emotional [NO15, NO18]. emphasis [JSWY19]. emphasising [JB18a]. Empirical [SSP15, BV13, Mor16, NBW10, NBW11, RRC^{+}11, RBMPP^{+}13, YSST13]. Empirically [Bec14, PRMH14]. Employing [TKK14, Pat16, RMB15, pZWZ^{+}16]. EMT6 [AP13a, BFGS10]. EMT6/Ro [AP13a, BFGS10]. enables [ADC19, GRCdL16, MC19, SSKK17]. enamel [BHLH12]. encephalitis [ZLCH18]. encode [RM17b, RBRB16]. encoded [KB11, ZP11]. Encoding [JP13, III17, LTL^{+}15, SIK^{+}18, Sel13]. encounter [RPCW18, WRC^{+}19]. encounter-conditional [RPCW18]. encrypted [BKC^{+}16]. end [Ano19-27, BCS^{+}16, Fan11, KMH19, LN19, LC11, WHH17]. end-binding [WHH17]. end-to-end [Fan11]. Endemic [FGJ10, ABN15, BH13b, BDR10, MHX^{+}14, MK19, RK19, RB14, TL15, TDM14, VABS18]. Endemics [LBGW13]. endocrine [BSR^{+}11, Hal17, TAORS10]. Endogenous [DHBS19, MÇ16b, HZW^{+}14]. endonucleases [KDL16]. endophytic [ADCG14]. endoreduplication [AKdV^{+}14]. endosymbiont [CEdLSG^{+}16, YT12]. endosymbionts [Unc11]. Endosymbiosis [KM17, AE17, Lan17, O’M17]. endosymbiotic [LP17, Sat17]. endothelial [ABR11, BZJP18, CMHM19, LFD11, SMM15, SNY^{+}17, VSW10]. endothelium [DvDBD15, PIT12]. endotherms [BS13, BW13, WSRG18]. endurance [GMMGVD12]. enduring [Gri15a]. enemies [De 17]. enemy [HJWC11, TTC10]. Energetic [DPRSS11, RW14b, HKH18, MC19, SS15b, SS18, SH17, SHW16]. Energetics [HH10, IPMH12, CEdLSG^{+}16, HFT15, RPGG^{+}19, RS14a, SW13a]. Energy [AKNP18, ACH15, Kuz19, SSH^{+}19, TTB^{+}18, Tra16, WSRG18, BRP^{+}18, EBS11, Fil10, FYZ15, GS13a, GS14, GLOC10, GTSP16, HMH14, IMH15, KF10, LWLM18, LAJC19, MPJH13, MTE15, MP13a, MSC10, Miy17, PKL12, Pel18, RZRSC19, RBMPP^{+}13, SW18, SW13a, SI19, Smi11, TvMG16, TTW13, TI12b, TKTB18, TMS13, YSH^{+}14, ZG10a, ZG10b, ZKMB19]. Energy-information [Kuz19]. energy-saving [TTW13]. enforce [MI19]. enfuvirtide [LS11a]. engineered [EA17, EA19, NHE^{+}16]. engineering [CLCH10, Ghu18, KSP18, LGPS17, SWTO15, TP10c]. engulfment [PGS10]. enhance [BT17, BPG^{+}18, BBJDS11, SC10b]. Enhanced [HYQ^{+}16, MW13, NK18, RW16, DS16, KSM^{+}15, RG17, SWSLMJ19, ZZC10]. enhancement [HTB^{+}13, yTRSC13, PSG^{+}17, STN19a, WA14]. Enhancer [PBBB10, FLM18, NZZ19]. enhancer-promoter [FLM18]. enhances [LP14, MK18a, PMSY17]. enhancing [NF16]. ennuplet [DMCP14]. Enquist [Bal13]. enriched [JB19a]. enriching [WDN^{+}19]. enrichment [KK15, Rem15, RM10, ZLW^{+}19]. ensemble [AH12b, Hua16, JLX^{+}15, JLX^{+}16, Mei14, Miy17, NS19, WYX^{+}17, WYL^{+}19, XWD^{+}10]. Ensembles [BK19]. ensure [FD17]. entanglement [KDL16]. enteric [HJLNZ11, LFZN11]. enterobacteria [dEG11]. enterocolitis [AES^{+}13, BTH^{+}13]. entities [Kro10]. entorhinal [TWTA^{+}18]. entrainable [TCH14]. entrainment [AUE15, DFT^{+}17]. entropic [AN17]. entropies [MGO^{+}15, XNJ^{+}13]. Entropy [RFME^{+}12, Chu18, DPRSS11, DZW10, GKNT10, LLG16, LXJ15, Mit14, Nem17, SSP15, Tsu19]. entropy-based [DZW10]. entropy-driven [Chu18]. entropy/clarity [GKNT10]. entry [DOC17, LCT11]. Enumerating [CM17]. envelope [MB14]. envelopes [KDMK16]. envious [GCM14]. Environ [FB12]. environment [AdVG15, BKR14, BRP^{+}18, CCW16, CNM^{+}13, DS18, GJ12, GCG14, HD11, KK13, Lab16, LLLV11, LFM11, LW10a, LW10b, LLL13, MG14b, MMFK10, NdlPZA10, RRBM12, Saf13, SV14, SM10b, UK17, VDD^{+}17, WST15]. Environmental [FR13a, HAK^{+}19, MP14a, RM13, Abr14, BJJR10, BDR10, DSA^{+}16, DHT19, EzMMH18, FGLS10, FGGT15, GGBW14, HBVP15, HSM19, HSM17, HHT^{+}19, HWL15, JWTW17, LRH15, Mal17, MLCH11, MJV16, NNvdB17, NWWL17, PMD11, PW14, RR12, gSxFH^{+}12, VBVD17, WXWL12, WD12]. Environmentally [CCW16, BCF19]. environments [BID15, CMCS18, DH16, De 19, FB18b, JRMS12, KL13, KM19, LX10, LNH13, LCDH15, Mar19, MHF^{+}13, NdORC11, OBE^{+}17, PEW18, PB16, RW14b, SI11b, SB17, WSVT14]. enzootic [RC16]. Enzymatic [MSSM18, CTJ10, KJ11b, OF10, Pie12, YI18a]. enzymatically [Pie12]. Enzyme [MRS14, Bur19, DK13c, DBJ12, ESS19, EBSW17, HAuR^{+}18, KHK15, NCS^{+}13, OFB10, RMSTG13, SC13a, SC13b, Str17]. enzymes [ONO15, PBKR13]. Enzymological [IDM15]. enzymology [XG12]. ephaptic [WMT16]. ephemeral [MBC^{+}12a]. Epidemic [CGW18, HS15, HK11b, MYOS14, ZLL17, ZCXF13, ABB^{+}19, Ada16, AB16, BBR16, BM10, BR15, BGLR^{+}19, DS19a, DGRB13, FJS^{+}19, FHM18, FS16c, GA16, GSCS11, HSD11, HR14, JKS15, KKYV18, LN19, LY11, LCT11, LLJ18, LJS14, MZAI19, MFP^{+}14, MPLK14, PSJ15, PM11a, PGGvdB12, PTC13, RBHK14, SSZB15, SKPK17, SMC^{+}13, cSGFB17, SS11, TA13, VABS18, WHT17, WMCL18, XTZ^{+}13, XLSF19, Yan10a, ZYJL18]. Epidemicity [MCRG18]. Epidemics [AGGME18, MDM^{+}10, MLMG^{+}15, NIiT18a, BSC19, BM12, BZCS10, BTK15, CJRR11, Che12, CE17, DCR18, EB15, GM17, HS15, HJR12, LAS14, LC17, LKL13, MG10a, PHK15, SBK16, SLJ^{+}10, US12, WCSA12, XSCS12, YS11, ZLCH18]. Epidemiological [MDM16, AHJ^{+}18a, CFF11, GA15, HASM17]. epidemiology [GGL^{+}16, Gru18, LDJW16, MR11b, Rao12, WA14]. epidermal [AMP12, GBC^{+}16, KSK^{+}16]. epididymal [Hir16]. Epigenetic [ASP13, Pee12, LCDH15, PG18a, PG18b, Sat10, Sha14]. epigenetic-based [PG18b]. epigenetics [Pee12]. epilepsies [MvAKR17]. epilepsy [HZG^{+}17]. epileptic [MGB17]. epimysium [XYN15]. epiphysis [CAA11]. episodes [PDW11]. episomes [Mor11]. epistasis [FSW^{+}16, Jam16a, KII19, PTNKT11, VDTF15]. epistatic [BST14, Gri15b, SLO10]. epistemic [KLHS17]. epithelia [AdSG14]. epithelial [BPS19, DFC^{+}12, PC10b, SHB^{+}17, TK10c, VAG16, WGS10, WBB^{+}18, vDVFH^{+}18]. epithelial-mesenchymal [TK10c]. epithelium [DFG^{+}18b, SOCF14, WTC10]. epitope [BBT^{+}15, BYY^{+}11, CRH^{+}15, CP11, KRDJ15, NGJ^{+}14, SQZ^{+}16, SSRA16, TWR^{+}18]. epitope-based [TWR^{+}18]. epitopes [JB19a, KB11]. Epstein [HR12]. eQTL [LZG^{+}19]. equal [Di 18, Gra14]. equals [BCK19]. Equation [Gra15, BMM^{+}14, GGM15, JCLS^{+}11, JCLS^{+}21, Lio18, Lio24, MBCM12, MB16, OYA14, Oko16, Pie10, RLK10, RG17, VACGF17, VM16, WP17, Zad11, ZS10, vVGSE12]. equations [CK18, DG11, ESG^{+}16, FWLW11, HLH^{+}18, Lio16, MROS15, NBC16a, NBC16b, OOY16, PCHM10, SF15, VLFF12]. equator [Kur14]. Equilibria [IKHL16]. Equilibrium [HT19, ABN15, BR13, CWF^{+}16, FXML18, FT12, HN10, JLQ^{+}12, JJ15, KS13, Kis10, KS16, Lio18, Lio24, Miy16, NWZ15, NP13, RRÚJ19, SC12, SABB15, Tsu19, UBP12, WdLS17, ZL18b]. Equisetum [KSRN19]. equitans [Di 14b]. equity [Tak17b]. equivalence [MPLK14]. Equivalent [FS11a, TZ13, vVGSE12]. eradicating [RCL16]. Eradication [vdBD10, DBM^{+}18, MDM16]. Eradication-resolution [vdBD10]. ErbB3 [GY19]. Erdös [LTP19]. Erection [ZWY15]. ergodic [KSPA17, SVB^{+}10]. Erikson [Pee12]. Erratum [BPF15b, dOGL13, Gün13, JRG14b, KDK14a, LZT13, NO14a, PL14a, SSKS20, SC13a]. Error [OF10, SE10, TH17, dASdC^{+}15, Fra19b, HK16, LCG^{+}15, Mas16, PP12, TMM10]. error-correcting [Fra19b]. errors [CV15, MI19, PW11, PW14, TH17]. erythroblastic [FKB^{+}12]. erythrocytes [CZT^{+}16]. erythropoiesis [CDGV10]. Escape [Coo16, Kar16, CPV16, Fre10, IMA16, Lev14, PVB^{+}10]. Escaping [VSP11]. Escherichia [ABK^{+}12, ARZ15, BRG^{+}12, BLP10, HVSZ10, KF12, MBE10, MMH^{+}12, Tak17a, ZS12]. Esml [JLX^{+}15]. esophageal [MGM13b]. ESS [Ayd18]. Essential [AFM10, LLL^{+}17, LLW^{+}18, LLP^{+}19, TM15a, VW10, ZRGW19, dASdC^{+}15]. establish [BGM19]. establishment [DDSDW13, Kle14, MJV16, NBW10, NBW11, RRC^{+}11, Tak16, Tox18]. Estimate [HN19, BMI19, Boc12, GHBC14, RBB^{+}12, VACGF17, ZKMB19]. estimated [KN11, CRLH^{+}19]. estimates [SENS^{+}17, TSMB14]. Estimating [AL17, AGRR^{+}15, AHR10, CPS19, CMCS18, HYN19, HS11, KD16, KFG15, NYSM12, PAA11, PGGvdB12, VGL16, BGF^{+}14, For10, HM11b, HRG11, MSS10, PM15, TDW16]. Estimation [AJC12, ADMR16, Lüt16, NI11, AEM13, AEM14, AP13a, DJD10, ERT13, KM18, LZA^{+}19, LSC^{+}18, MC19, NSS16, PP17a, Par19, RM17c, SQZZ17, SC10c, TKB^{+}18, UAL10, Wal12, WcW14, WZ15, ZDF^{+}12]. estimator [BS15b, Kon12, NPH^{+}12, YZFY16]. estimators [MKE15]. estrogen [ZP11]. estrous [BSR^{+}11]. ethylene [SABB15]. eubacteria [QW11, QLZQ11]. Eucalyptus [DDB10]. eukaryotes [CB17b, Dil14, Edg19, EM16, Hua12, Lan17, Mic15b, Rad16]. eukaryotic [FZ14, KSvdH18, KGM15b, LGEM17, MG10b, RFdL15, Sat17, SDRA^{+}15, SM11b, SN17]. Eumetopias [LLLV11]. euphotic [DD10]. Europe [MJV16]. eusocial [LF13]. evaluate [CPF^{+}13, FH14]. Evaluating [Bal13, MBPS17, NMI18, O’M17, RK19, HMTA17]. Evaluation [BGF^{+}14, KCD11a, MAFB18, VLFF12, XSKA17, ASH15, CRH^{+}15, GBRS19, IU10, LN13, LZL^{+}19, PNP^{+}16, RAAS15]. evasion [BBS18a, GL18a, Kar15, LBF13]. even [BC15]. evenness [DBD15]. event [ANMH11, BTR18, BJOS13, Lan17, MMH^{+}12, SKD^{+}10, ZLY^{+}13]. events [BZA16, GMMGVD12, Gol10b, Höh15, LCJ16, SP16, SPCM18]. Evidence [Di 17a, MPL16, RPGG^{+}19, Di 10, GDB^{+}19, Mor10b, PC17, RMRC^{+}16, She11b]. Evil [GvVT14]. evinces [BHLH12]. evo [FtL18]. evoked [LSB10, LB11, LB12]. Evolution [AK15, Arc13, AS16, AK16, ACvKA10, Axe12, BBB18, COWA11, CHK16, CKNB19, FZW17, FM10, FM14a, FPD15, GT12, GTSP16, Hou13, HCW18, Jam13, KK17a, KGB13, Kit10, KWO19, KNT10, KI17, Len14, LG12, Mas14, MA11b, MAL^{+}11, NH12, NF14b, NO18, Nak16c, Nei14, NOS17, NP13, ON14, OAV^{+}16, PBYB18, RC11, Rog19, Rou14, SSM15, SC12, SLW18b, TP10a, TM12, UI13, WK15, WK17, YHI14, YTK10, dWV11, Abb10, AdVG15, AFM10, AH12a, AIY16, Akç12, AOR17, AB11, Ant13, AMFL10, Arc11, Arg12, AK11, AT10, BGKL17, Bau18, BM11, BJ17, BD12, CT14a, Cha17, CSBK15, CW15c, Cle16, CBP12, CAV16, CS15b, CzST^{+}18, Cro17, CL13b, CKS15, Das13, DBD17, Déb17, DSLR18, DMCP14, Dun11, Edg19, ER12, Ell15, FB18b, FM15a, FH13a, FRG^{+}13, FM18]. evolution [FOT^{+}15, Gar10, GFH^{+}18, GJ12, GG14, GN14, GJ10, GJ15a, GKPB13, GCvWE^{+}14, GKW16, HNV^{+}16, HJWC11, HM10a, HAP^{+}16, HLI15, HT18, HTT19, HKM12, HMM11, HTK12, IIKT13, Ish16, ID14, ID19, IIVR10, Jan15, KG15, KMCH14, Kin18, Kor18, KW15, Kur14, KIT^{+}16, Lai11, LBB^{+}13, LGK^{+}09, LGK^{+}12, LvBJ16, LG19b, LSM17, LvdBP12, LGEM17, LSLL13, MTNM12, MB18a, MK16a, MYK17, Men12, MALAN17, MH18, MAA18, MH13, Mor16, MC15b, MM18, NF14a, NS11, Nak14, Nak19, NSKO18, NBS^{+}13, NMS10, NP11, NPS17, NPS18, Oht12, Ots11, Pan11, PIPB10, POW18, PBR17a, PMD11, PKZ12, PPBT11, PN16, PN18, Pie12, PGH11, Pol12, PBD11, PM10b, PBA12, Rad16, RANO10, RBR18, RBRB15, RBRB16, Ros13, RRR15, RCD19, RD17, SW16, SC10a, Sal15, SPS11]. evolution [SPLP12, SSF15, Sch19b, SHBT14, SN16, Sel10, SGW17, SSS17, SN18, SP16, SB17, SA13, SP13b, TGLK19, TI12a, TF17, TALC16, TB15, UI10, VSLVB15, Vel17, VB19b, WCS19, WKB13, WZLW15, Wil11, WK18a, Wod18, Wol12, WFW19, XY14, XW17, YK11, YT12, YvBS18, YSH^{+}14, Yan10b, YZX15, YG12, YX15, ZPdFJ19, ZCW13, ZADB15, ZLW16, ZMW10, ZMT11, tBdR18, vVE15, AM10]. evolution-based [HMM11]. Evolutionarily [BHBR12, KM12, BI12, DWL^{+}14]. Evolutionary [AY15, AN12, Arc14, BL14, BM16, Bur10, CZN12, DM11, Ezo12, Fis16b, Fra19a, Fra19b, GCM17a, HI12, HWTN15, HIT18, HHN14, IMA16, ID14, IS16, IK16, Kři10, LLZ13, LW15, LCQ^{+}18, MPNP12, NPB13, OGO19, PBR17a, PNL15, Rad16, RN12, RBB16b, Ros10a, SSPP15, Sek12, SGW17, SPCM18, SLO10, SG18, ŚKSRW16, TTN13, TEY16, WL12a, WL12b, WL14, WCW15, jWlGQ19, ZGWL14, ZWG10, Abr14, AH12a, AH12b, ATI12, ATTN12, AD13, AWRD15, AOR17, ATG12, AIB^{+}19, BS15b, BID15, BS11, BD10b, Bon10, BBDB13, BR13, Bul12b, CVO^{+}15, CSZT12, CK19, CB10c, DGMY18, DMO^{+}17, DHS^{+}15, DRCL12, DKR16, DI10a, DI10b, EHSR17, FL13, FLW16, Gal11a, GHBC14, GT11b, GC11, HBK12, HSM19, HFT^{+}18, HBT13, HK10, HY10, Höh15, IMM16, IMA14, Iij12, KGB13, KDMK16]. evolutionary [KK13, KZ14, Kri16, Kri18, KC17, KGC18, Lai18, Lar16, LG19a, LL12, LGBT15, Lio16, LPKH15a, LPKH15b, LPH18, LH18, LIPD12, Lun13, MP14a, MFMPS10, MRPN12, MDM16, Mon18, MB12b, MKPVH16, MSIR10, NMI18, Oht10, OBHS19, PBB^{+}15a, POW17, PLN14, PGG^{+}15, Pła10, RC11, RBB16a, RVP19, RSBB10, RM17d, SKAG18, SMS17, Sch17, Sch23, SMG18, SSSD17, STG19, Shi19, SKS15, SY17, SS15e, SM11a, SJK18, SC16, SS12b, TBM^{+}13, TA10, TLCZ12, Tay13, TZY18, Tre19, USF^{+}18, VGZS18, WSP14, WYX^{+}17, WLY^{+}17, WMN18, WCF12, Woo10, WK18b, WSVT14, XKCG15, YHZ14, YBT^{+}17, ZZK14, ZYD15]. evolutionary-based [DHS^{+}15]. evolutionary-dynamical [Lai18]. Evolvability [MH17, GJS^{+}10, Gri15b, IMA14, PMP13, VFS^{+}15, WB10]. evolvable [AIY16]. evolved [Bry13]. Evolving [MBRRI19, Now12, SAZ^{+}14, AH12b, CL14, CL19, De 17, EB15, FWW18, NR16, PSJ15, PW19, SB12, WLW19, WLW20]. EvoStruct [USF^{+}18]. EvoStruct-Sub [USF^{+}18]. Exact [KFW12, NSK13, PHK15, Pie10, SW17, ZLW14, Gri15a, Ros10b, ZM10]. examination [TF16]. Examining [BLS^{+}17b, VAT18, AHW13, ER12, Gla13, GS17b]. example [CR14, PKZ12, SGA^{+}12]. Examples [TvMG16, MF15]. excess [AHKN10]. excessiveness [VM16]. exchange [Bal13, BTG^{+}15, BP16, CC10, CZT^{+}16, GT15, LMJ^{+}16, PWC^{+}15, POP12, REC10, XYZ15]. Exchanges [Fer12, Sel13, Sel14b, Sel15b]. exchanging [Sel16, Sel17]. exchanging/deleting [Sel17]. excision [BKPV15]. Excitability [OMO13, Eli17, OTTF11]. excitable [BHBH11]. excitation [BZN17, CN12a, MTE15, SS14]. excitation-contraction [SS14]. excitation-energy [MTE15]. excitatory [CP14, KTH16]. excludability [De 19]. exclusion [CCA17, GP12, GA15, NvHM13, ØW16, PDW10, SLvdBMP10, XF13]. exclusion-principle [NvHM13]. exclusive [WKS14]. executive [RM17a]. exercise [BSR14, BS13, EJ17, EJ19, HLVR16, QF10, VM16]. exert [Cle18a]. exhaled [IMW15, KKU^{+}10]. exhibit [Sat18]. exhibiting [Bro13, JB19a]. exist [AM10]. Existence [Opr10, ADV^{+}10, FT12, Hay16, HXL18b, MG14b, WKS14]. existent [BB11b]. existing [EJK16, PVGA12]. exit [LCT11]. exit-entry [LCT11]. exocytosis [GU13, SWSLMJ19]. exogenous [CWP^{+}18]. exon [MSSM18, PBKR13, QYO10, ZLZ^{+}15b]. exon-coded [PBKR13]. exon-junctions [MSSM18]. exons [dOGL10, dOGL13, ZLZ^{+}15b]. exoskeleton [SKSO12]. Exosomal [LF17, Sha14]. expand [Dim17, RT19]. expanded [Sel17]. Expanding [CB10b, HL10, DBJ12]. expansion [DZJR10, FNH10, GTB10, HM10a, Mil11, Ou19, RSR11, Sel15a, WMD16]. expansions [CKNB19]. expectations [ALH10]. Expected [BCK19]. experiment [AHD^{+}18, LGR^{+}12, MBS19, PWL^{+}11, SYY17a, SYY17b, VBV10, WW12b, Zhe16]. Experimental [BPP^{+}16, CW11c, LFW^{+}18, WHWZ18, ASH15, AGC18, BMB^{+}18b, CDM^{+}14, DRW14, DDF^{+}14, FKMP10, GDPPSS^{+}11, GDB^{+}19, JRB^{+}16, LCSH14, LAG^{+}14, MAMEA15, MBF^{+}15, PMKS^{+}15, RMRC^{+}16, RGA^{+}10, YA14, ZBA14]. Experiments [JSP^{+}16, MAH12, QMJW15, EUM^{+}16, FS11a, HWPL12, JSK16, KP15, LSMP14, MSS10, PBB^{+}15b, PMCS16, Ran12, SGGM11, SBR10, SBMH10]. explain [BRB18, NMS10, PDL^{+}17, RN12, SGGHS19, TWH^{+}13, WR14, YY18]. explained [Fil15, Jen10, PVB^{+}10, Sch19b, SW15]. explaining [CMB^{+}12, GDJC11]. explains [Das18, DM10a, Eli17, EJK16, RG18, SSI^{+}16, SWD16, SP14, WKS14]. explanation [BPF^{+}19, FPdRdH10, MM15a, Sil16, UD10]. explicit [BLS^{+}17b, HDNB10, KF10, Nak16c, SSD12, SS11, TM15b, TKB^{+}18]. explicitly [MAL^{+}11]. Exploitation [Das12, CFR13, HS14c, Oku15, YvBS18]. exploiters [KR19]. Exploiting [NS19, RS14a, DAVS11]. Exploration [CFR13, HAuR^{+}18, LKAJ18, TF18, Gru18, HNR14, HAH19]. explore [AA12, MCL19, MJV16, Pav19, RC16, TvMG16]. explored [MDEKH13, XYZ15]. Exploring [CVPCV^{+}15, Edu16, HXL18b, LC17, MRPH17, TB16, WHH17, ZZ18]. explosion [GJ10]. Exponential [Hor11b, AEM13, AEM14, CAV16, MN12, SZ15b, Tak17a]. Exponentially [Gol10a, Gol16]. exposed [LCDH15]. Exposing [SAM17]. exposure [BHH^{+}14, DD10, KKGN16, RC16]. expressed [Sat19]. expression [BRG^{+}12, CB10b, Das18, Dil14, Eli17, Gol16, GMBK14, HVSZ10, HNA15, iHIM12, HgLL^{+}10, HZL^{+}13, IOK19, JP15, JOM16, LJM15, Lei09, Lei10, LXS15, LDH^{+}14, MPP^{+}15, PS13, PGLZ14, RLM^{+}14, RT15, SRS^{+}15, SLS17, Van17b, WOB15, WH14, WMT16, WC10, XGZ17, YXH^{+}14, ZS10]. expressions [BF13, DHT16]. Extended [JMS18, JG14, Kee10, Kur17, LDHD14, LLP^{+}19, MOSS15, Mic15a, OFT15, TI13, Zha10b]. Extending [HP12, KDMK16, AH18, AHJ18b, GNB^{+}13]. extension [BRB18, CHH10, CSB15, HHF11, SKF16]. Extensions [GW13, RGP13]. extensive [Meh17]. external [Cle18a, VB19a, ZGWL14]. externally [SGW17]. extinct [HMM17]. Extinction [GDS11, GGM12, HBEW12, HY10, Höh15, KDS18, LW10a, LW10b, LSS^{+}14, NdMLLBLC13, PDSP13, PD13, RT19, Sch18, Sch21, SIVH17, SSME10, SPCM18, SLJ^{+}10, TMM10, TLS^{+}13, TSS16, VB19b, WN19]. extinctions [CCA17, SL10]. Extra [POP^{+}19, TMS13]. Extra-cellular [POP^{+}19]. Extracellular [AH16, AdSG14, CRGS18, CPS10, EG10, KPK^{+}13, KSRN19, Lee16b, MÇ15a, SWSLMJ19, SBJ^{+}18, ZSZM14, ZAL^{+}19]. extract [LlLYB13]. extracted [HZL^{+}13, NSBL10, NK18]. Extracting [IA17]. Extraction [GAB14, NS16, AAJGCD15, CGRRGM^{+}16, DLL^{+}17, DM15, SLDP13, ZKHL16]. extraocular [GGC14]. extravagant [Bul12b]. extravascular [Kun16, MAM16]. extreme [IA17, ZZN^{+}19]. Extrinsic [OM10, KPS17, RLSN14]. eye [GGC14, Meh17, SSLB15].

fabp4 [BRA15]. face [ISB^{+}11]. face-to-face [ISB^{+}11]. facemasks [TDE12]. faces [CS14a, Mit14]. facilitate [UT17, WFW19, YI16]. facilitated [KPEK14, LWY^{+}16, PWC^{+}15, Rad16, TM12]. facilitates [Kin18, Mun12, PCN17]. Facilitation [MBLC17, CH11, DP13b, HvLPdB15, LTHEK12, MR15, MGCL13, RK18, TCR13]. facilities [GA16]. facing [BWP10, Coh15b]. factor [AOR17, BFR14, EKvdKvFK13, FLM18, GJ10, GBC^{+}16, HG18a, KHS13, KSvdH18, KPS17, MSND12, MHMM11, PSG^{+}17, TBB14, TSP15, dEG11]. factor- [EKvdKvFK13]. Factors [LvdBP12, DCR18, HT18, JWTW17, LP19, LVLC18, MFN^{+}18, MMFK10, MKBE17b, PHÁBAI^{+}16, SY11, WLL^{+}14, ZBA18]. factory [CEdLSG^{+}16]. facultative [BW13]. fade [BBR16]. fade-out [BBR16]. Fading [AJD11]. failed [DH18, Har15]. failure [BC15, GB13]. failures [YY18]. Fair [WF17]. fairness [AB11, NOS17, SSPP15, SSM15, WCW15]. faithful [MFKS13]. falciparum [CDO17, CKZ^{+}17, CMD^{+}10, GDPPSS^{+}11, KMLT14, KPLT14, RM19, SPMGR10, TEY10, TEW12, TAR16]. fallacy [BCK19]. False [LA12, KSP18, SN16]. Familial [LG18, MCL^{+}11]. families [CDC18, HWW^{+}14, Kin18]. family [Akç12, AEM13, AEM14, GHBC14, MKBE17b, Pav14, TT10a, Xie11]. farming [ABR19]. fascicle [KSM^{+}15, RW15a]. Fasciola [VPGDP^{+}14]. Fast [MB18b, SMC17, WZ15, Fue14, Fue15, Jia16, LM15, MSBB13, NdlPZA10, PB16, RKJ^{+}11, SL19, VWF19, WHL13]. fast- [PB16]. fast-folding [RKJ^{+}11]. Fast-slow [MB18b]. fasted [OABI12]. faster [RL15]. fastidiosa [AHD^{+}18]. fasting [BAGG14, HMH14]. fat [LK19]. fat-tailed [LK19]. fate [AHAFA14, CSKF10, LLTP19, MBK^{+}11, POP^{+}19]. fathead [HCHI18]. fatigue [BSS11, FPR10, KCZ^{+}19, PFPR11]. fatty [vLFM^{+}19]. favor [Di 17a]. favorable [VDTF15]. favoritism [KM12, Mas12, MJI14]. favour [TMHM12, YvBS18]. favoured [DMCP14]. favours [PP17b]. Fc [OBK^{+}11]. FDG [BDTR15]. feasibility [SRS18, VP18]. feather [PHH18]. Feature [AH15, AAJGCD15, KKD18, KHX^{+}19, AH19, AH16, DLL^{+}17, DDLW10, FP18, lGLLL14, HTK14, JSZ12, JD16b, JW18, KPK^{+}13, LWW14, LTL^{+}15, MS14a, MFZ18, NKM^{+}12, RAF^{+}19, SIK^{+}18, SM18, SLDP13, WYL^{+}19, ZKHL16]. feature-based [FP18]. features [ÅCHS19, AFS19, BCGD10, BRB18, BRK19, CCY^{+}19, lDsWM^{+}15, Dol19, HK11a, IA17, JR17a, KSA16, KHNM16, KFW12, LT14, LlLYB13, LXC15, MSK10, MZ18, NK18, NHSX14, PWH^{+}13, STJG12, SSSD17, SLT^{+}18, USF^{+}18, YLC16, YGL^{+}10, ZMS11, ZLY^{+}13, ZZG^{+}16, ZNC^{+}15, ZNA^{+}16]. February [Ano10-41, Ano10-46, Ano11-39, Ano11-43, Ano12-36, Ano12-38, Ano13-37, Ano13-32, Ano14y, Ano14-36, Ano15-28, Ano15-41, Ano16-31, Ano16z, Ano17-39, Ano17-32, Ano18-35, Ano18-40, Ano19-45, Ano19-37]. Fecal [JEA18]. fecundity [JB18b, POW17, POW18, TEW12, TEY16, WK15]. fecundity-affecting [POW18]. fecundity-dispersal [WK15]. fecundity-longevity [JB18b]. fed [OABI12]. feed [AGS18, RLM^{+}14]. feed-forward [RLM^{+}14]. Feedback [CMS16, GKMC10, UT17, ABIM10, AN13, BFJ^{+}18, BGL^{+}19, CMB^{+}12, CPS10, CDGV10, Cza14, Fra19b, GI15, GSS19, HSM19, iHIM12, JB19b, JRMS12, KHS13, KDFM14, LS16a, LKP^{+}12, LCGMH12, LWvB^{+}19, MK18a, MA13, MKMT14, MFKS13, Pag19, PRK13, Ram10, RW14a, RA18, SS15a, SXZY12, TWH^{+}13, Wod18, YFB^{+}12, YLLL12, Zab11, ZCSRG12, ZS10, dSKBS10]. Feedback-control [GKMC10]. Feedbacks [KMN16, NN18, AIB^{+}19, BWB11]. feedforward [LL13b, Pat19]. feedforward-loop [LL13b]. feeding [MA10, PGS10, RVA^{+}11, SCH^{+}19a, SJK18, SKS^{+}19, VGL16, ZWY15]. feedstuffs [TLL^{+}12]. feeling [KM17]. felid [CW11a]. female [FKMG15, HCHI18, MS13, Nak16b, YY18, ZXS^{+}19]. females [Gar14, Gra14, MG17, TT10b, YSST13]. femur [JTSG14]. Fengxiao [TXY^{+}12]. Fermentation [SSJK18, MÇ16b]. Fermi [LPKH15a, LPKH15b]. ferric [Ber14]. Fertility [DBD17, BCBD19, KMH19]. fertilization [TEY10, TEW12, YD14, YD15]. Fetal [JPBM17, NNH^{+}16]. fetus [BD10a]. Fever [LG18, PLF18, HTB^{+}13, JWTW17, SKK^{+}12, XSCS12]. fiber [ESGA15, Js12, KHH10, SP13a, SDKF12]. fibers [Giv10, Smi14]. fibre [CWB^{+}17, Kro10, SSR^{+}19]. fibril [SPG^{+}18, ZGY11]. Fibrillar [MHH13]. fibrillation [LWB^{+}15]. fibrillogenesis [SAGC12, VH11b]. fibrils [ZSW11]. fibrin [SNY^{+}17]. fibrinolytic [GGG12]. fibroblasts [CLW12, NJP18]. fibrosis [BTK14, KBV^{+}15, NPMM15, RCL16, WDL^{+}17]. fibrous [WBMM18]. ficus [TVA18]. ficus-indica [TVA18]. Fidelity [Déb17, XKCG15]. field [ASL^{+}18, BZ18, FLR14, FHR13, FR13b, FR14, GSCS11, HCK^{+}11, IPS14, KTJ19, LL13b, Mor10b, MvAKR17, Pat19, Rob11, SOIO10, TKB^{+}18, Tay16a, TAR16, YMW12, dlCGSA16]. fields [Pel18, PSD^{+}18, PW14, SKGM19, Tay16a]. Fifty [Kir11]. fighting [GVGC15]. filament [DLEMP19, Don13, KSK^{+}11, Kro11, MED12, MAH12, Por13, RG18, SK12a, Smi11, WGH^{+}14, MOSS15]. filament-mediated [SK12a]. filamentous [GS13b, LTS^{+}16]. filaments [KHNM16, Smi11, SH16]. Filippov [CLK18]. filled [AAJ15]. filling [CdPD11]. filter [DGM15, PJ13, SSS15]. filtered [AHR10, CRV16]. filtering [BF16, PP17a, PR18]. filters [NPMM15]. fimbriae [CB10a]. fin [Ios16]. final [BR15, NYSM12, Zhe16]. find [CLL18, KGB13]. finding [BBB18, GTC19, RKK^{+}19]. fine [MDB12b]. finger [HMMSRSD15]. finger-DNA [HMMSRSD15]. fingerprints [XMWC13]. fingers [LCBO^{+}12]. Finite [JTSG14, KHB^{+}18, PWC^{+}12, AY15, AM11, BTG^{+}15, BDID^{+}12, CTPB10, FZW17, FS11a, Gre15, HB12a, Han17, Has14, HI12, HHN14, KD16, MH14, RN12, Ros11, SS14, SKH11, Tay16b, TF15, WL12a, WL12b, WN19, YiTS19, ZLZ^{+}11]. finite-population [KD16]. finitely [IM18]. fire [ADV^{+}10, AC12, CP14, CLK18, DAVS11, HTK^{+}18, Urs14]. fire-blight [CLK18]. firing [AANF16, BHBH11, TWTA^{+}18]. firmness [vdSS12]. First [WCC13, CW14, Dol16, KMM18, KK18, Lai18, MJS11, Nah12, PGGvdB12, Rou14]. first-order [CW14]. First-spike [WCC13]. Fish [LYK12, BPG^{+}18, CT16, DHP11, FSA15, GS13a, GSL13, GS14, HPW^{+}13, Ios16, ICG16, NR14, NTY16, NBA^{+}18, OGA16, PLF19, gSxFH^{+}12, SPH12, Str13, TIY^{+}11, VDH^{+}15, WBM^{+}15]. Fisher [CS18, LPH19, PBYB18, SW17, TT10b, jWlGQ19, Wax11b, ZLW14, ZGW16, vV18]. Fisheries [LHD15, MA13, MNH^{+}12]. Fisheries-induced [LHD15]. fishery [TB16]. fishes [SYI17, Yam16a]. fishing [TM14]. Fission [KOS13, LSY^{+}10, SDPC11, Wu14]. Fission-fusion [KOS13]. Fit [PBRW11, Bal13, TMF17]. Fitness [CBBO16, CS18, WSW10, AH12b, AIY16, Arc13, CAV16, CMM19, CGB13, FSW^{+}16, FGLS10, FLA^{+}19, Gra15, Han17, HTN14, KS14, LSM17, LS16b, LCG^{+}15, LPH19, LPF11, Mar11, Miy17, MPNP12, NvD17, NSK13, OWB14, PNK16, PBB^{+}15b, PTNKT11, SY17, Tay13, TGL15, TDKJ15, WMD^{+}11, WSP14, vVGSE12, vVLS14]. fitness-based [LPH19]. fittest [SBJ13, ZZC10]. Fitting [RRRTR10, BPC15, TZ13]. Fitzhugh [CHL^{+}19]. five [KF15, ZNA^{+}16]. fix [DDRP13]. Fixation [AM11, CW15c, DS18, HT10, JJ17, JLH15a, LPKH15a, LPKH15b, BCF^{+}16a, Gre15, Jam16a, KII19, Mon18, SW17, Wax11a]. fixed [ARB13, MB18a, WSW10]. flagella [Ish16]. flagellar [NS18, OSF11]. flagellum [CL10]. flap [MHSH12, Sac15, Ush16]. flap-gliding [MHSH12, Sac15, Ush16]. flapping [DS12, FH14, LS18, SW13a, Ush16]. flare [vdBD10]. flare-ups [vdBD10]. Flat [BFGAGA16, HTM16, RW14b]. flatness [SW15]. flavin [LT14]. flavin-containing [LT14]. flavonoid [MR18, MR19]. flavor [dLMD^{+}11]. flavour [DDF^{+}14, dLMV^{+}10]. Fleeing [Coo16]. flexibility [Cza14, GRCdL16, TED10, TvMG16, WZWY11, YX15]. flexible [BUC14, FCS18, Ios16, WGH^{+}14]. flexural [THBM10]. flies [GWCA14]. Flight [SW18, FH10a, FH10b, FH14, GAL11b, LS14a, LS19, MHSH12, Sac13, Sac15, SW13a, TB11, Ush16, XS13]. flights [EPP19, Han12, RPCW18, TA15]. FLN [JAM18]. floating [RDD14, STI13]. floating-leaved [STI13]. flock [ZVMB10]. flora [ISM^{+}11, YSI18]. floral [CPPLAB18, FH13a, TGB^{+}18]. Florida [CBC^{+}18]. Florigen [SSI^{+}16]. Flow [FRTP13, LMdL11, STKE12, AV19, ABR11, Cab18, CPS19, CR19, CKD11, Don17, EBP15, GDF17, GDF18, GSF13, HWM^{+}16, HT13, JBSFB12, KSBvL^{+}12, KTT^{+}19, LMM^{+}13, LLK14, LCJ18, LDB^{+}14, MYK17, Mig16, OBK^{+}11, PIT12, SM16, Str13, TKTB18, TAB^{+}13, VF12, Zad11, ZG19]. flow-assistance [KSBvL^{+}12]. Flow-induced [LMdL11]. flower [Abb10, SCÁBM10, SCÁBM11]. flowering [MYN^{+}13, MYN^{+}15, MBLC17, PPT^{+}16, TIS10]. flows [MGML10, PCT11, PCT19, Pan19, SAGAGB17]. flu [LC17]. flu-like [LC17]. Fluctuating [RRÚJ19, DS18, Del12, FB18b, HD11, LCDH15, Mar19, NR14, PEW18, SMG18, SM10b]. Fluctuation [BGL10, AC12, Lei09, Lei10, Zhe16]. fluctuations [DDP12, HMM16, JRG14a, JRG14b, JRG14c, KMM16, MLCH11, MK17, NS18, PPBD10, STA15, SGD12, TP10c, TLR14, VDD12, WTC16, YST14]. Fluid [WO11, ASH15, AABS16, ABR11, BJWL15, DS16, DFG^{+}18b, GTB10, HWM^{+}16, KMZR18, KTT^{+}19, LS15c, LS16c, MPJH13, MN12, MH13, OB15, RS12, SCS15, SOCF14, TKTB18, WTC10]. fluid-biochemical [SOCF14]. fluid-tissue [BJWL15]. fluid-to-solid [MH13]. fluids [Bur10, PCT11, PCT19, Pan19, RL15]. fluke [BHO^{+}18]. fluorescence [GT14, GZT15, Laz13, MTE15, TKK14, TP10c]. fluorescens [FE10]. flux [BUC14, KJSS10, KdSUS10, LRL^{+}16, MDB12a, MT14, MRS14, Ou19, Vaz10, XW17]. fluxes [FMS^{+}12, FVTS16, HWGT15a, KLN^{+}12, LD18, MAA^{+}17]. fly [GAL11b]. flying [JLH^{+}15b, LA15, MPJH13]. flytrap [Leh18]. FMN [MPY14]. FMN-binding [MPY14]. Fn14 [KB19]. Fn14-NF- [KB19]. focused [HRHAAA15]. focusing [CFSB16]. fog [BFM10]. fog-dependent [BFM10]. fold [AC15, Bry13, IA17, KB10a, LBS^{+}14, LPD^{+}16, RSD^{+}16, SRS^{+}15, SLDP13, She11b]. foldability [Miy17]. Folding [XYZ15, BB13, CL17b, CdPD11, Gin10, HH13, HWHL15, Kin18, Kri18, LYM^{+}19, MWSM13, Miy17, NGL^{+}10, RA10, RKJ^{+}11, RBMPP^{+}13, RBMPP^{+}15, SV14, TI12b, TP17, Wal10, XLL^{+}10, ZGZ^{+}10]. folding-upon-binding [LYM^{+}19]. folds [APW10, GNB^{+}13, Mas18, SV14]. foliaceus [MNH^{+}12]. Follicle [BSR^{+}11, PVGV19]. follicular [LAH^{+}16]. follow [ZL18b]. followers [KW15]. following [BRG^{+}12, CDM^{+}14, CDK11, CS14b, FD11b, JKF19, VZB^{+}15]. follows [MH13]. Food [HAB11, ZEJA11, ACT12, AD13, AWRD15, ASF^{+}15, BV13, BEK10, CAG13, DDF^{+}14, DKF17, GK13, HSR12, HY10, HB13a, KDG13, KPG19, Kři14, cLCJ^{+}10, MGML10, MB15, MD16, MEJ18, MA10, MW13, PBRW11, PDG12, RPD14, RBB16a, RBB16b, RM13, SGS15, SVA18, SI10b, TBM^{+}13, TD17, TD18, VL11, WPPD16, WBM^{+}15, XDC^{+}11, ZG10a, ZG10b, dLMD^{+}11]. food-borne [MW13]. food-mediated [MA10]. food-web [MD16, PDG12, RPD14]. food-webs [MD16]. foodweb [Kla10]. foot [RB14, ZJY19]. foot-and-mouth [RB14]. forager [EM11, RPCW18, SP11]. Foragers [EM11]. Foraging [BC15, Mor19, ACT12, Amo15, AMFL10, BPFR16, BEK10, Cza14, GVCG12, GSL13, HH10, HS14c, Kři10, KR19, LTHEK12, MK14b, PRMH14, QJR^{+}16, RRBM12, SK16a, STN^{+}19b, SH17, TED10, TD17, UPWK15, VPFV13, ZMT11]. foraging-related [ZMT11]. Force [CHH10, HMM16, TS17a, ITN^{+}11, IS12, JLH^{+}15b, LL17b, RG18, SCF^{+}12, WC10]. Force-driven [TS17a]. Force-extension [CHH10]. force-length [RG18, WC10]. forced [GP12, GA15, Ros10b, TGZF11]. forces [AHP^{+}19, KG12, LS18, SCF^{+}12, Smi14, WBB^{+}18, ZLB^{+}18]. forcing [BM10, DFM^{+}14, HBVP15, LKK13, PM11a, RVP19, TWS15, VR17]. forecast [SIVH17]. forecasting [HYSG19, Jab15]. forecasts [PM15]. Forest [FBAPMD13, AEZR^{+}16, Cam11, Dra19, GDS11, JM12, JLX^{+}16, KCM^{+}11, KPK^{+}13, Kär11, LCMC14, MS14a, MZ17, Oka15, PAA11, SS15a, SAB17, ZL18b]. Forests [WZM19, GZ14a, GZ14b]. forgotten [TF17]. forked [SKSO12]. form [AH19, CS11b, CL13a, FL12, FL13, HYA14, JAK19, JD16b, KZL14, MGC15, SRS^{+}15, Uit11, WMK13, ZZK14, ZZCZ17]. formal [DWG^{+}19]. formalism [LN13]. Formalizing [WDN^{+}19]. Formation [GCB17, TCYY^{+}12, Amo15, BB11a, BFM10, BC13, BPF^{+}19, Bue15, BFGAGA16, CRC15, CLPZ14, CMB^{+}12, Che19, CGRRGM^{+}16, CPPLAB18, CAA11, DNB12, ESGA15, FKMP10, FGAM12, FK13, GKMC10, Gal10, GA13, GWW^{+}10, HTM16, HKS15, Hir16, iITIM15, KRR14, KS16, Kon17, KLHS17, Kun16, LFM14, LS15b, Lee16a, Lee16b, LCQ^{+}18, LRA^{+}13, MTdS^{+}16, MEKK11, MKMT14, MGGGA11, MBC^{+}12b, MN14, NOS17, PHH18, PHA^{+}17, PW19, SRK^{+}12, SNY^{+}17, SBP13, SHK14, SYSM11, SM16, SYSF17, SSKK17, SCS14, TGB^{+}18, WH11, WRH^{+}16, WZ18, WDL^{+}17, WBMM18, WMT15, YWP13, YG18, YKO^{+}16, ZW16, ZKTK10, vDVFH^{+}18]. formation/resorption [FGAM12]. formations [MN14, Mit17]. formed [HZW^{+}14, MSSM18, NHTS14]. forming [JSWY19, KB11, SPG^{+}18]. forms [BL15a, Jam15, TT10a]. formula [BGD^{+}13, CHH10, JHE15a, JHE15b]. formulae [FVC15]. Formulas [IU10]. formulate [AH18]. formulation [Mei12a, Mei12b, Mei13a, Mei13b, MKF^{+}14, RGA^{+}10]. formulations [CPGF^{+}16, KBV11]. formylation [NMZ19]. Fortune [PP17b]. forward [AGS18, BSKV18, RLM^{+}14, XS13]. fossil [DRCL12]. fossilized [SGW^{+}18]. fossils [Dai12]. fosters [Nik15]. foundation [BB15, Di 16, Voe18]. foundational [Sza15b]. founded [Bec14]. Founder [Sza15b]. Four [ZADB15, DBD15, Di 18, III17, JYZ18, LRGA13, Mas18, MSSM18, NHTS14, PGGvL^{+}19, SSS13]. four-body [Mas18]. four-species [DBD15]. four-taxon [SSS13]. Fourier [CL17b, GT14, HYZ^{+}15, YCY14, YY15, ZDY11]. fourth [Che10]. Fractal [Kea16, LC10, PNP^{+}16, SHM^{+}18, BVK10, DS10b, LHH12, NAK^{+}11, NHSX14, PHTP^{+}12, PDB^{+}15, RBSD10, SKK^{+}11, Su16, Wei12, ZKP15]. fraction [JSK16, RJSC18]. Fractional [KSKS13, DG11, HSD11, MKBE17b, PD16]. fractional-order [HSD11]. fracture [CGvG^{+}15, LBB^{+}13, SAB10, Van16b]. fractured [Zad11]. fragment [PBKR13, SRP16]. fragmentation [BJPP12]. fragmented [GWBG14]. fragments [VHS17]. frame [CS15b, CMJD11, Mic14, Mic15a, SSCJ16]. Frameshifted [Wil13, Sel12b]. frameshifting [Xie13a]. framework [ASH15, AHD^{+}18, ANK10, AC15, BW13, BR12b, CdPD11, CG11, DG10a, DHK13, FWLW11, GXFF13, GGL^{+}16, HAB11, HB18, HMM11, KFG15, LX10, LTS^{+}16, Luo14, MK16b, OGE10, PBB^{+}15a, PGT^{+}19, RBF15, SMS17, Shu13, SLT^{+}18, SH15, TT19, ZGZ^{+}10, dSKBS10]. Frankenhaeuser [MAS15b]. FRAP [HL10]. fraternity [SS12b, SSC13]. Free [BRP^{+}18, BPS19, BFJ^{+}18, BC15, BDR10, BMB^{+}18b, CFCM13, FS16b, Fok12, FYZ15, GJS^{+}10, GAB14, HLHY17, HZLX11, ITR^{+}18, KSKS13, LKSM14, LL13a, LKL13, LWL^{+}11, LAJC19, MPBS17, Miy17, RBMPP^{+}13, TMS13, TP17, XGZ17, XLSF19, YZZ12, YZZ13]. free-boundary [CFCM13]. Free-energy [BRP^{+}18, LAJC19]. freedom [WL13]. freezing [KSRN19]. frequencies [AIY16, AGM11, AEM13, AEM14, DLYZ11, MKE15, MKF^{+}14, SGW17, Sin15]. Frequency [OAJK10, CNDK15, HCK^{+}11, HM15, HSG^{+}18, HT10, JSZ10, JJ15, KM18, KGC18, SKK^{+}11, SM11a, WMK13, ZLY14, ZW16, ZAL^{+}19]. frequency-dependent [ZW16]. frequent [CDC18]. fresh [Kři11, MFM16, Ros15]. freshwater [PLF19]. Friction [WL13, HB16a, KC13, MLMG^{+}15]. Friction-based [WL13]. friend [TCW14]. frigatebirds [SW18]. frog [LO15, RP18]. frogs [JB11]. Front [Fra16, HSD11, Han12, dRFI15, ZUFM17]. fronts [MSM^{+}14, MLCH11]. FRS [Ano19-50]. fructose [Bar19]. fruit [AKdV^{+}14, GAL11b]. Fruiting [SSS17]. frustration [HKH18]. FtsZ [Sht17]. Fu [WZM19]. Fu-SulfPred [WZM19]. fuel [KK15]. Full [Kae17, TAORS10, BMM^{+}14, RN12]. function [Aya12, ASF^{+}15, BKR14, CST19, CLC11, DL15b, Don17, ESW13, FKB^{+}12, For14, GJ15b, GS17b, HRC^{+}12, IMH15, JK12, Kin18, LRA^{+}13, MKBE17a, MS14b, MIH16, OBK^{+}11, RMSTG13, RBMPP^{+}13, SS19, SS17b, TT17, TF15, VBV10, WTQL10, ZHC10, dBB14b]. function-driven [Kin18]. function-oriented [SS19]. function-specific [JK12]. Functional [ABJ12, CW11c, GVGC15, AF10, ASRM15, BRA15, CCY^{+}19, CNM^{+}13, DS13, DOC17, FB12, FWLW11, GTS15, GG12, GG14, GBC^{+}16, HH14, III17, JAM18, KDFM14, Kha10, LDH^{+}12, LCL14, LMJ^{+}18, Mor10b, PC17, PRV^{+}14, SL10, SFMS16, SK11a, SGDL12, SVA18, TWB16, WMK16, ZOG14, ZLW^{+}19, ZSS10, vLBJ^{+}13]. Functionality [LT11, PJL14]. functionally [SHN12]. functioning [AOT^{+}11]. functions [AGB^{+}14, AFM10, ARD14, DGMY18, DHT19, DBG18, MR15, MP16, PLF19, SKSO12, UK17, WSRG18, YSI17, ZSCL16a, ZSCL16b]. Fundamental [Gra18, HTT19, LLG16, LE19, OGE10, Di 18, Di 19c, Esc11, vV18]. fundamentals [TF16]. funds [SU11]. Fungal [NC15, Bos12, CWM11, CTB18, DEK15, DTP19, EBSW17, VBVD17, WCM15, YLY15]. fungi [BMN16, DQY^{+}15, LTS^{+}16]. fungi-warming [KTO^{+}14]. fungus [APS^{+}13, APBS15, GS13b, KCMF11]. Further [BD10b, HFD17, Jam11, FLM18, JV14]. Fusarium [WMD16]. Fusing [WZM19, HK11a, TWC^{+}19, YGL^{+}10]. Fusion [LCTG15, DFG^{+}18a, EBSW17, GPD18, KOS13, LS11a, Pel18, Rad16, SW13b, YMW12, YC15, ZMS11]. fusions [MK16a]. Fussy [Cro19]. futile [OABI12]. futile-cycle-associated [OABI12]. future [BM12, IS10, Kir11]. fuzzy [BKR14, GKNT10, GJ15b, HZWH10, JCG16, JW18, MGM13b].

G [BMF^{+}18, BPLM12, CHN^{+}15, HPML18, PBBB10, ZLM12, DHB15, Dol19, Ken19, KB10a]. G-CSF [BPLM12, CHN^{+}15, ZLM12]. G-Protein [Ken19]. G-protein-coupled [HPML18, PBBB10]. G-Protein/Receptor [Ken19]. G-quadruplex [DHB15, Dol19]. G1661A [ACM16]. G2 [KBE^{+}13]. G2019S [ADS^{+}19]. G4Catchall [Dol19]. G6PD [CB13a]. GA [DCL18]. GABA [MBE11]. Gaia [AN17, Doo17, NWWL18]. gain [ASH15, BCK19, LvdBP12, SW18]. Gains [PLN14]. gait [STJG12, WHHS15]. gaits [Ush16]. GAL [AM12]. galactolipid [MB14]. galactose [MMK16]. galectin [GBN14]. galls [Yam16b]. galvanotaxis [SKGM19]. Game [AS12, DB19, HHS^{+}10, TWB16, Abb10, AOR17, Ayd18, BBR12, BMdA17, BB10, CZN12, Che12, CWW18, CSZT12, CK19, CB10c, DH16, DMO^{+}17, DN18, HFT^{+}18, HS14b, IMM16, IRS11, IK16, JLQ^{+}19, JC16, KNT10, Kon11c, LLZ13, LPH18, LPvSP11, LPvSP12, MI19, MGS16, NIiT18b, NiTI18, NIiT19, OO17b, PEW18, QMJW15, Ram10, RN12, ROF17, RSH19, RC11, Ros10a, RM17d, SMS17, SSM15, Shi19, SGA^{+}12, SN18, SCS10, TT19, TGL15, TR17, jWlGQ19, WL19, WHYMG17, XSLZ16, ZMAM19, ZGWL14, ZKHL16, ZLY^{+}17, ZWG10, dMP11, PS12]. game-theoretic [DH16, MGS16, WHYMG17, ZMAM19]. game-theoretical [SMS17]. Games [AN15, Fis16b, KVN16, LSDI15, LJI17, ŚKSRW16, ATTN12, ATG12, Arc16, AS16, BR12b, CLA17, CZ14, CK19, DDRP13, DHK13, FTS15, FNH10, Gal11a, GCM17a, GT11b, GC11, Has14, Hau10, HI12, HSM19, HWTN15, IM18, IOY15, Jan15, JOAN14, JSF^{+}11, JR17b, KC17, KGC18, Kur19, Lai18, LGBT15, LPKH15a, LPKH15b, LRGA13, Mob10, MV18, MB19, NOS17, NCN13, Oht10, PIPB10, PLN14, PGG^{+}15, Pła10, SHN12, SPSP12, SU11, Shu13, SS12b, Tay13, TR12, UI10, VNS18, XWZ10, YHZ14, ZZC10, ZCW13, ZWC^{+}13, ZZCZ17, vVN12]. gametocyte [TEY10, TEW12, TEY16]. gamma [Gol16, HYN19]. gangliogenesis [HJLNZ11]. Gánti [HS16, CB15, Gri15a, Sza15b]. Gántian [GJ15a]. gap [For10, JCG15, KPD18, MAA^{+}17, PAA11, SGGM11]. garden [KCMF11]. gas [GT15, REC10]. gaster [MAW12]. gastric [OSV^{+}16, PRN10]. gastrocnemius [CSB15, ZPM^{+}15]. gastrointestinal [PPC^{+}17, PBS^{+}12]. gastropod [Nos14, NSS16]. gastropods [IUK14]. gastrulation [YM14]. gate [Zha10b]. gated [Bec14, PDG17]. gating [FIS11, FIS16a, LdLK11, ZCA^{+}14]. Gause [Kři11, KP15, LJ18]. Gaussian [BMI19, CKF17, Gol10a, LCG^{+}15, MKBE17a, MKRE18, PPT^{+}16, ZMS17]. Gaussian-distributed [LCG^{+}15]. GC [KB11]. GC-rich [KB11]. GCPII [NRS^{+}16]. gDNA [pZWZ^{+}16]. gDNA-Prot [pZWZ^{+}16]. GED [LJM15]. geese [MPJH13]. gel [FKMP10, HHA17, RVZ14]. gels [Kro10, SNY^{+}17, SABB15]. gender [DBBM11, FS12]. gender-specific [DBBM11]. Gene [Hua12, OHWS18, YLF^{+}17, ZT16, ACM16, AD15, AKR^{+}18, ABIM10, AGS18, BRA15, BRG^{+}12, BPP^{+}16, Bar18, BSB^{+}13, bCR19, DTGD19, Dil14, DTY12, EA17, EA19, FXML18, GG16, Gin10, GHBC14, Gol16, GMBK14, Gri15b, GGCJ16, HNV^{+}16, HTM16, HL14, HXL16, HXL18a, HG18a, HNA15, HgLL^{+}10, HZL^{+}11, HK16, HT13, IOK19, III17, JP15, JOM16, JYY^{+}18, JSF^{+}11, JK12, KSP18, KF11, KII19, LJM15, LWH^{+}11, Lei09, Lei10, LXS15, LLL^{+}17, LDH^{+}14, MTKY11, MC16a, MMC19a, MH11, MBE10, MPP^{+}16a, MH12, MYK17, MHD18, NMS10, Orr19, Ots11, PAOMV17, PPBD10, PWH^{+}13, PS13, PPM12, PMP13, PGLZ14, Pol12, PKH11, PSK10, RMM^{+}16b, RLM^{+}14, RT15, RRÚJ19, Sak10, SCÁBM10, SCÁBM11, SSS13, SMC17, SLHS13, SLS17, Van17b, WOB15, WMK13, WH14, WM17, XSKA17, XGZ17, YXH^{+}14, YGMT12, YBH^{+}19, YCR^{+}15, ZS10]. gene [ZTWL12]. gene-attenuation [JYY^{+}18]. gene-based [PKH11]. gene-regulatory [RLM^{+}14]. genealogical [LG18, Mul11, Mul12]. genealogies [PPTC19]. General [MA11a, WZM19, AFS19, AH19, BMN16, BDR10, BR12b, CL13a, CCY^{+}19, CXC18, CT18b, CdPD11, DHS^{+}15, FL12, FL13, FLW16, FM10, GML10, HN19, HYA14, HKR^{+}19, JLQ^{+}12, JYZ18, JD16b, JD17, JCG15, JCG16, KFB18, KHKI17, KZL14, Kri16, Kri18, LZ18, MFZ18, MZ18, MDM^{+}10, MK16b, NBL14, Non10, QLC^{+}18, RSS^{+}18, RPCW18, SRS^{+}15, SM17a, SM18, SR10, SRV11, STG19, SKK18, WMK13, XSLZ16, YG12, YFB^{+}12, ZCAB17, ZZK14, ZD18, ZNA^{+}16]. generalised [BSB14]. Generalising [BLS17a]. generalist [FKV19, SK16b]. generality [MN14, THM10]. generalization [AW12, BM11]. Generalized [Est10, KC14, CGD13, Hal12, HYW11, IEN15, KHX^{+}19, MvAKR17, PDG12, RRRTR10, WMT10, WGH^{+}14, ZWC^{+}13, vLBJ^{+}13]. generalizes [BKR14]. generate [AD13, TM15a, VDH^{+}15, ZSO16]. generated [Mul11, Mul12, Sat10, SWG^{+}15, YvBS18]. generates [OS11]. generating [LLC15, Ots11, WXC10]. Generation [LSM17, PKH11, Bue15, CRLH^{+}19, DHBS19, FJR19, ITN^{+}11, LWW14, MMFK10, MDM^{+}10, Pat19]. generations [CNL14, Oht12]. generative [BBP13, Hal16]. generator [PRK13]. generic [JMZ13]. generosity [FZW17]. genes [ADCG14, All17, AGM11, BRA15, BKC^{+}16, Di 12, Di 14b, FP18, GXFF13, GLZ^{+}11, GRR^{+}14, HZL^{+}11, KB10a, KB11, LG17, LZG^{+}19, Mic15b, Ots11, Sek12, Sel12b, Sel12a, Sel13, SL12, TDZ^{+}18, TR12, YLC16, YC15, YCR^{+}15, Zab11, ZLZ^{+}16, ZLB13, ZSW^{+}14, ZXS^{+}19]. genesis [PBYB18]. Genetic [DXW^{+}16, GWW^{+}18, HLTR19, LS13b, ZZX^{+}16, ABA11, AK16, BMdA17, BM11, BWM^{+}19, Boz15, CRLH^{+}19, CC10, COG19, CL11, CKF17, CVO^{+}15, DKP^{+}18, DS19b, Deu18, Di 13a, Di 14a, DMCP14, Di 16, Di 17a, Di 17b, FD18, FRP14, FDS13, FGLS10, Fra19b, GKNT10, GE18, Gon13, GJS^{+}10, Gri15b, HL14, HG18b, HCS^{+}19, Ish16, IU10, Kam11, KA17, LR16, Len14, LMW^{+}10, LT14, LPTR14, wLW10, Mas16, Mic14, Mic15a, Nem17, NMI18, NdORC11, pNMK^{+}10, OM10, OCHHZ12, PMYHR12, Rog19, Rou14, SM10a, Sel12b, SLL18, Tal12, UT17, Voh17, WPA17, Wax11b, WZY14, Woh15, ZW16]. genetic-metabolic [COG19]. genetically [DROC11, GRB^{+}13]. genetics [AM14, BSKV18, CM14, Mau15, NGS^{+}16, SL12, ZLW16]. Genome [ABP^{+}11, CW11b, HHJR11, MZWC10, TDZ^{+}18, ABG^{+}13, BYY^{+}11, BLZ^{+}19, BKC^{+}16, FMS^{+}12, HZLX11, HMM11, KEHK17, Li11, LZL^{+}17, MKBE17a, MKBE17b, MKRE18, MSIR10, QLZQ11, TLY^{+}19, WRWX19, YT12, YZFY16, YY15, YHY14, Zha15]. genome-scale [FMS^{+}12]. Genome-wide [TDZ^{+}18, MKBE17a, MKBE17b, MKRE18]. genomes [EM16, FS15a, FS16a, GRRG16, HTT19, KEHK17, LLW15, LY14, QW11, SM11b, YCR^{+}15]. genomewide [ZK10]. Genomic [Sel15b, SSCJ16, AD16, CL17b, DCS14, dOGL10, dOGL13, JYZ18, MPS^{+}11, SVSP15, TLY^{+}19]. genomics [GTS15, SPSM15]. genotype [AGM11, Bec14, IMA14, IMA16, LP12a, MHKA16, NvD17, NNK^{+}15, VPGDP^{+}14]. genotype-phenotype [Bec14, IMA14, IMA16]. genotype-to-fitness [NvD17]. genotypes [MKBE17a, MKBE17b, Mou12b, SS11, TTP17]. genus [NNK^{+}15, WSTL16]. geodesic [VPFV13]. Geographic [SSD11b, SOIO10]. Geographical [IST11]. geological [DLSD15, FGGT15]. Geometric [NSS16, Tou14, GZY12, Nos14, PWC^{+}12, PW11, RG18, SGB12, TKKE19, Ush16, WY13]. geometric-mechanistic [RG18]. Geometrical [KDMK16, FKK14, LWB^{+}15]. geometries [CL14, CL19]. Geometry [FJ10, WXmH10, AP19, CW11a, CF11, CGB13, DI13, GL18b, KBS^{+}19, Lee16a, LS10b, Mar12, MPT19, Sum13, TC11, TSMB14, KBS^{+}20]. germ [SFMS16, TIMI^{+}15]. germination [BI12, LHD^{+}17]. get [Cza14, vVGA10]. Giant [DDP12, TDT16]. Giardia [GDPPSS^{+}11]. gibberellic [OFB10]. gibberellin [AP17, BP18]. gills [Str13]. Ginzburg [MROS15]. giraffes [HN10]. give [Rey13]. gives [SAB17]. giving [lDsWM^{+}15]. GJB1 [AKR^{+}18]. glaciation [BTK14]. gland [CFR^{+}14]. Glazier [SMS19]. GlcNAcylation [ZNC^{+}15]. Glia [ZAL^{+}19, IBB^{+}15, Sat19]. Glia-mediated [ZAL^{+}19]. glia-neuron [IBB^{+}15]. glial [BTO15, EG10, PD16]. gliding [AABS16, MHSH12, Sac15, Ush16]. gliobastoma [CRC15]. glioblastoma [Mal10]. glioma [Mal18, PH13, SHSR15, THC^{+}11]. Global [BMI19, CW15b, KLHS17, LP12b, MBBV14, RBMPP^{+}13, SSZB15, UI10, WW19, WGO^{+}15, XLL^{+}10, ZCT18, Arg12, BBI14, CJKR10, CT18b, DLSD15, DDSDW13, JZL13, MGM13b, PM11b, Pie10, QQL10, VK10]. globins [IK15]. glomerular [AMM16]. glomerulonephritis [AN13]. glomerulus [ML11a]. glossal [ZWY15]. glucocorticoid [BSC^{+}18, MZMM^{+}14]. glucocorticoid-induced [MZMM^{+}14]. Glucose [BD10a, MKMG^{+}14, RBSD10, BAGG14, CTW10, GdGL17, HBA17, KSU^{+}18, KTI18, LPGTBMA19, LLB^{+}18, QF10, SW13b, SW19, WCHC11]. glucose-dependent [SW13b]. glucose-dose [SW19]. glucose-insulin [CTW10, GdGL17, HBA17, WCHC11]. Glucose-lactate [MKMG^{+}14]. glutamate [MPS^{+}11, ML11a, NRS^{+}16, Pav14]. glutamatergic [CS11a]. glutathione [BRN15, KIH16]. glutathionylation [ZNA^{+}16]. glycerol [GZFX14]. glycine [Sat19]. glycocalyx [LFD11, VSW10]. Glycocodon [Nah12]. glycocodons [Nah12]. GlycoGAIT [OSV^{+}16]. glycogenes [OSV^{+}16]. glycol [SABB15]. Glycolysis [Arc14, GHJ14, YLWZ10]. glycolytic [STNT17]. glycoprotein [DLM^{+}19, JB19a]. glycosaminoglycan [FS11a, XLLF17]. glycosylation [GHBI16]. GM [SER^{+}12]. GM-CSF [SER^{+}12]. Gneg [SC10b]. Gneg-mPLoc [SC10b]. GnRH [DLHS11, RSD^{+}13]. go [DS10a, RA17, vIK11, LWW14]. goals [MI11b]. GOASVM [WMK13]. goat [VCF^{+}19]. goldfish [CPS10]. Golgi [AH19, JD16a, JD16b, TWH^{+}13, ZLW^{+}19]. Golgi-resident [JD16a, JD16b]. Gompertz [AGRR^{+}15, BM16, JKF19, RRRRSPTR16, TMF17]. Gompertzian [AGRRTR11]. gondii [JSSZ12]. gonorrhea [MMRCCC10]. gonorrhoeae [JSC^{+}16]. good [LvBJ16, MBLC17, OAP14, SYI17]. goods [AS12, AS13, Arc16, AS16, CWW18, CSZT12, CB10c, DDRP13, Hau10, KNT10, MV18, ON14, OO17b, ROF17, SU11, WL19, XWZ10, ZCW13, ZWC^{+}13]. goodsgames [WH11]. gossip [LLSO13, SN16]. gossip-mediated [SN16]. govern [GL18b]. governance [Dra19]. governed [ASH15, BC13]. governing [BMB^{+}18a, ZSW^{+}14]. gp120 [VBHM^{+}13]. GPI [SMJS14]. graded [AMM16, MBM11]. graded-size [AMM16]. gradient [Che19, DNB12, FGGT15, KAN11, Lio18, Lio24, MR15, MFL^{+}18, NSS16, Pel18, PSD^{+}18, RG10, SA17, XXD^{+}17, ZZN^{+}19, Zhu11]. gradients [AGPK13, DM10a, GS13a, GS14, HWMT17, MBM11, Mal15, PMD11].

Graduated [IL13]. graft [GB13]. grained [KSK^{+}11, KG15]. gram [AAJGCD15, BC19, HTK14, LTL^{+}15, SLDP13, CXC18, DHS^{+}15, HLS^{+}12, LZ18, SRD^{+}15, SC10b, USF^{+}18]. Gram-negative [CXC18, DHS^{+}15, LZ18, SC10b]. gram-positive [BC19, DHS^{+}15, HLS^{+}12, USF^{+}18]. graminearum [WMD16]. grams [HY16]. Grandmothering [KMCH14, CHK16, Sek12, KMH19]. Graner [SMS19]. granularity [LlLYB13]. granules [SW13b]. Granuloma [SYSF17, MEKK11, WDL^{+}17]. granuloma-associated [WDL^{+}17]. granulomas [GSV16]. graph [DB11, DMSW10, FBAPMD13, FLCS^{+}15, HNO18, MP14a, Mac14, MKRE18, QLZQ11, TP17, WY13, WZ17]. graph-based [WZ17]. graph-structured [Mac14]. graph-theoretic [DMSW10]. graphene [ACH15]. graphic [WXC10]. Graphical [BFS18, KZ14, aHLZ^{+}12, KDMK16, XM11]. graphs [ATTN12, JOAN14, Mon18, NIiT18a, NRKE18, OBHS19, Tay13, WCW15, ZGWL14]. grass [ADV^{+}10, DAVS11]. grasses [DDS13]. grasslands [LJW^{+}16]. grave [Sat18]. gravitational [Thu15]. gravitropic [CMJD11]. gravity [TSdL14]. Gray [aHLZ^{+}12, PNP^{+}16, PDNP16]. grazing [AZOLVH18, KO18, Sie18]. greater [Di 14a, Di 18, NPB13]. Greedy [PNK16]. Green [Lai11, DM15, GvVT14, GPD18, JG14]. Green-beard [Lai11]. grey [PDB^{+}15]. grid [TWP16]. grooming [SDPC11]. Grounded [ARB13]. groundwater [BDLR12]. groundwater-dependent [BDLR12]. Group [PN18, Sal10, SP16, SDPC11, vVGSE12, BJOS13, CMBP19, CFF11, CDD12, CB10c, GSRR17, HB10, KM12, KW15, KF11, KI17, LTHEK12, LW15, LSSG10, MJS11, MSN14, Mas12, MJI14, MPNP12, NF14b, NO18, PN16, PRMH14, PTB12, SSPP15, SN18, Sim14, SRCF11, YK11, YBH^{+}19, ZGZ^{+}10, ZLW18, vVLS14]. Group-level [SP16]. group-living [NO18]. group-mass [CDD12]. group-structured [PTB12]. group-targeted [CFF11]. group-wise [KI17]. grouped [DS10a]. grouping [SL19, vVGA10]. groupings [dSMP^{+}11]. Groups [TP10a, Cro19, Eft13, HKB10, KDFM14, LA15, MBC^{+}12a, NTOI16, PCB14, Sue12, TZY18]. growing [CKD11, DFC^{+}12, GMBK14, Hir16, KBF18, LFM14, MSK10, RR16, SGKCP19, SK12b, ZTK12]. grown [PP18, PP19]. Growth [DCR18, ISZ18, PPW16, Ren13, ATI12, AGRRTR11, AR17, ABD^{+}11, ASGG17, BCGD10, BPC15, BSV^{+}10, BDE^{+}14, BP16, BAM^{+}11, CXWL11, CFCM13, CKZ^{+}17, CTJ10, CAGM^{+}17, CS11b, CWB^{+}17, CCB11b, CL14, CFL^{+}15, CL19, CL18, Chu17, CLCH10, CCAdS13, CCdS15, DFG^{+}18a, DHHP14, DTGC14, DTP19, DV12, EdKM11, EBE14, FHLM10, Fil10, Fra16, FJC^{+}10, GCSP17, Ghu18, GS13b, GBC^{+}16, GAPK10, GLF^{+}15, GPT17, GDB^{+}19, HWMT17, HB13b, IGL^{+}12, JAK19, JD13, JKF19, JMS18, Jus11, KHP^{+}12, KBV12, KBKK14, KTN^{+}10, LG10a, LL14, LTS^{+}16, LDdA^{+}13, LCGMH12, LXLF14, LGPS17, Mit16, MBCM12, Mol18, Mor13, MGC12, MP19, MHMM11, MJV16, NR14, NF17, NNH^{+}16, Nij11, NHS^{+}16, NHM10, NSS16, OBE^{+}17, OYA14, OT13, OB10, PH13, PBYB18, Pie10, PBR17b, PSG^{+}17, PLD^{+}17, PLF19, PM13, RRRTR10, RRTR12, RRRRSPTR16]. growth [Sal15, Sch18, Sch21, SZ15b, SAB17, SSLB15, SZ18, Spe15, SHSR15, SER^{+}12, Tak17a, TGB^{+}18, TS17a, TRM^{+}14, TTB^{+}18, TP14, TT10a, TAB^{+}13, VBV10, VZ19, Veg10, VBVD17, VDD^{+}17, WGS10, WSRG18, WBMH12, WGO^{+}15, WFM^{+}13, XY14, YZY^{+}16, Yat14, YBT^{+}17, ZS12, ZZS19]. growth-II [FJC^{+}10]. growths [VLFF12]. GRP78 [ÁTC^{+}14]. GRP78-NF- [ÁTC^{+}14]. Guangdong [HJR12, TXTW16]. Guangzhou [JWTW17]. guanidine [PRM11]. guanine [KB10a]. Guaranteed [PP12]. Guerin [BMGC11a, BMGC11b]. guidance [BHP17, KTN^{+}10]. guide [GM19]. guided [TXWW12, TXCW15]. guideline [SRS18]. guides [BSP18, CHN^{+}15, LSMC18]. guild [Ezo19, IEN15]. guises [CD16]. gut [EzMMH18, KDFM14]. GWAS [LZG^{+}19]. Gymnema [RAA^{+}16]. gymnemagenin [RAA^{+}16]. gynodioecy [PM10b, VSLVB15]. gyrotactic [CFSB16, ZG19].

H [AHW13, GTS15, LDB^{+}14]. H1 [KGM15b]. H1-S2 [KGM15b]. H1N1 [NI11, TDE12, KLJ17, KJ19, LJ15, LCJ18, PMKS^{+}15, SY11, TXY^{+}12]. H1N1-2009 [NI11]. H1N1/09 [SY11]. H37Rv [MG14a]. H3N2 [MAA18]. H7N9 [CW15b, XSTW14]. Habitat [BJPP12, HD17, HMWB13, HH16, YLH12, BORA10, CA18, GWE11, LLH^{+}13, PC10a, VZ19, YLW^{+}14]. habitat-preference [PC10a]. habitats [AGPK13, GPR^{+}16, Kis16, VIBC12]. habitual [Mil16]. haematopoiesis [MMJB17]. haematopoietic [EUM^{+}16]. Haemophilus [DMM^{+}14, KM15]. hair [ANGPB12]. hairpin [TI12b]. hairs [ZWY15]. hairy [PMKM10, PS12]. half [HLI15, KKGN16, TSB10]. half-believer [HLI15]. half-sarcomere [TSB10]. halide [RCL^{+}10]. hallmarks [PBP15]. hallucinations [HRDL14, HRD14]. halophilic [GZ14a, GZ14b]. halves [Di 12]. Hamilton [AW12, SFD12, Tay16b, WWB14, vVAH^{+}17]. hamster [TKK14]. hamstrings [CSB15, Cle18a]. Han [PP19, PP18]. hand [CT14b, Gra18, HXL18b, LE19, PVCEC18]. handed [She11b]. handedness [Kam11]. handicap [Chu10]. handling [GBM18, LLB^{+}18]. hands [HXL18b]. haplodiploidy [Gar14]. haploid [BID15, LS16b, TT10b]. haplotype [AEM13, AEM14, CPH^{+}14, Das13, Mou12b, MKF^{+}14, TDZ^{+}18, ZMNd^{+}10]. harbor [SH17]. Hard [GW15]. hardwood [dBB14b]. haremic [SYI17]. harmonic [ARR14, FJR19]. Harsh [De 19]. harsher [DH16]. Harvest [PBD11, TM15b]. Harvest-induced [PBD11]. harvested [KPG19]. harvesting [BI19, CK18, DHP11, GW15, KJ15, LJ18, LP12b, PRMH14, PBD11, SHF16, ZZ14b]. harveyi [FBK11, HK14]. hash [ZLY^{+}13]. Hatze [RG17]. Hausdorff [YHY14]. having [Di 12, SSZB15]. Hawk [Ayd18, DN18]. hawk-dove [DN18]. HbA [BAGG14]. HbA1c [KTI18]. HBV [Nak16a, GM17, LLDW14, LZY^{+}15]. HCO [CZT^{+}16]. HCV [PBSM19, Nak12, SSST19, VZB^{+}15]. HDV [GM17]. head [MB12a, MM19]. heal [RS14c]. Healing [FKDW15, CGvG^{+}15, CSDR15, FBFM12, GBJ18, MHMM11, NCLB16, PC10b, Van16b, VG13]. health [BYM^{+}18, EME^{+}16, Kea16, LCCK12, SPRF13, ZARK19]. healthier [ADC19]. healthy [BRK16, LSS^{+}14, SCT12]. Heart [MKMT14, PW18b, BZN17, CST19, FPdRdH10, KAZ11, KAZ13, KCS^{+}15]. heat [DWG^{+}19, MG10b, RZIR16]. heating [RBB^{+}12]. heavy [GW12]. Hebbian [RA13]. hedging [MGT17, MHF^{+}13, Uit11, YY18]. heel [JH12]. heel-region [JH12]. heels [ZPM^{+}15]. height [AdSG14, BNG16, GDS11, JCLS^{+}11, JCLS^{+}21]. Helical [KI15, EG12, dSMP^{+}11]. helicases [Xie15]. Helicobacter [LAH^{+}16]. Helix [Kur10, KC11b, LZMM15, She11b]. Helix-helix [Kur10, KC11b]. help [Ezo19, FXML18, PFB10, Ran12]. helper [GML10, LAH^{+}16, RW12]. helping [Bar11, CT14b, MW14, WK17, YG12]. hematological [MPC12]. hematopoiesis [LM11, ZCT18]. hemodialysis [dlRVFK^{+}16]. hemodynamic [ARD14, CPS19, DHRA10, SSD11a, SKF16]. hemodynamics [OB15, RPGG^{+}19, SKF16]. hemoglobin [Kri16]. hemoglobins [Oko15, Oko16]. hemorrhage [GCYL18]. hepatic [DLL^{+}18, SWG^{+}15, SBMH10]. hepatica [VPGDP^{+}14]. Hepatitis [DK13a, ZZR13, dTKDV15, CM12, CKB18, GM16a, GN10, Jam11, KNA^{+}18, LHW14, LMYL15, LZY^{+}15, MR11b, MVH11, MLT10, MG15, MSL^{+}16, NS11, Nak12, Nak16a, PaCZ10, PBSM19, RDSD^{+}12, RCL16, ZZR10, ZRZ15, GC18]. hepatocellular [CM12]. hepatocyte [GWW^{+}10]. hepatocyte-stellate [GWW^{+}10]. hepatocytes [CPS10, DL12a, SAKG13]. hepatotoxicity [CTSL11, SBMH10]. HER2 [GBC^{+}16]. herbicide [EJ16, GT14, RDMP11]. Herbivore [IC11, LH18]. Herbivore-induced [IC11]. herbivores [DAVS11, HSR12, Yam16b]. herbivores-fire [DAVS11]. herbivory [YLY15, vVE15]. herd [AMSSG16, ASS19, CMN^{+}11, GKG^{+}18, OO17a]. herd-specific [GKG^{+}18]. herds [GKG^{+}18, KO18, LSSG10]. hermaphrodite [KFB18]. hermaphrodites [BHBR12, XY14, YSYI13]. hermaphroditic [KMHdlP10]. herpes [BYY^{+}11, OA11]. Herpestes [TF15]. herpesviral [FS15a, FS16a]. Hes1 [BF13, STX^{+}11]. hetero [WLD^{+}11]. hetero-dimer [WLD^{+}11]. heterocyst [iITIM15]. heterodimer [GY19]. Heterogeneity [GWE11, HTN14, MCC^{+}17, ABV19, BPR15, CP11, CG14, CMN^{+}11, Don17, EBX17, FHSG15, GGG19, GLF^{+}15, HD17, HR14, IOK19, JYY^{+}18, KFG15, LVLC18, NiTI18, NJP18, POW18, PGHC12, PRM12, RD17, RAMS11, SS15d, SZ10, VZ19, Wei12, YSI18]. Heterogeneous [CPF^{+}13, GW19, KMM16, Moo14, AGPK13, AK16, BZCS10, BI19, CCW16, CNM^{+}13, CGW18, DGJ14, FR17, GA16, HMWB13, HH16, HHD^{+}16, HLTR19, IM10, IL13, JMS18, KD16, Kle14, LS10c, LW18, MKKS14, MBS19, MG10a, OBE^{+}17, SBK16, SR12, Soz13, SB17, VIBC12, WMT16, WLY^{+}17, ZSH^{+}16, dWV11, dlCGSA16]. heterologous [BLP10]. heteromorphic [BI10]. heteroscedasticity [Mal17]. heterosexuals [MHMM18]. heterospecific [SBSE14]. heterothermy [BW13]. heterotrophs [CG10]. heterozygosity [Gin10]. heterozygote [BTK14]. Heuristic [Gri15a, MMCZM12, WZ17]. heuristics [RPCW18]. hexameric [Xie15]. hexaploid [ASRM15]. hexapods [GW12]. HGF [KL18]. HGF/c [KL18]. HGF/c-Met [KL18]. Hidden [Syi14, TS17b, BPC15, SG15b, Woh15]. Hierarchical [JSC^{+}16, Mic11, OS11, HASM17, HCW18, JAM18, MKBE17a, MKBE17b, YCY14]. hierarchies [MP16]. Hierarchy [LMJ^{+}18, FS15b]. Hierarchy-Based [LMJ^{+}18]. High [HCK^{+}11, HMM17, Jam15, MAA18, NBL10, Wal16, ASL^{+}18, CT14b, CA17, CZPC^{+}18, DMM^{+}14, GP11b, GAB14, HWW^{+}14, HPvdB17, IKHL16, IGG^{+}18, IMD16, Jam14, JR17b, KSA16, KFA12, LSMP14, LG18, LJ10, LFD11, MHMM18, Mon18, NC15, PDW11, SKK^{+}11, Str17, ZPM^{+}15, ZEJA11]. high-accuracy [LJ10]. high-dimensional [IMD16, Mon18]. High-frequency [HCK^{+}11]. high-performance [CA17]. Higher [MBLV10, RBHK14, KB10a, KB11, THBM10, vV18, KB10a]. Higher-order [RBHK14]. highlighting [SS19]. highlights [CRGS18, PW19, SGD^{+}16]. Highly [BTW11, ZCAB17, AKR^{+}18, DS15b, FCS18, HII^{+}19, JRR^{+}12, KB11, NP15, TDM14]. Higuchi [SKK^{+}11]. Hill [AS16, GS10, LYW11, RG17]. hinder [GWNW15]. hinoeuma [HLI15, TI12a]. hints [IK15]. hip [BAM^{+}15, SNCM12]. hippocampal [CW11c]. hippocampus [OABB18, PDB^{+}15]. Hippos [BCF16b]. Hipposideridae [AK11]. Histone [JK12, LLY15]. histopathological [KCJ^{+}11]. Historical [SS19, TIY^{+}11, BCPM^{+}16, LP17]. histories [GS13a, GS14, SSD12]. history [BWB11, BJOS13, BZA16, BDE^{+}14, BHLH12, Chu10, DBB14a, Fil10, HIT18, JMZ13, Kři11, MP13a, NNK10, PBD11, SMG18, TRJD19, YYST13, dMP11, tBdR18]. hitting [JRR^{+}12]. HIV [LGK^{+}12, ARG14, ARG19, BFH^{+}15, CLPZ14, CH16, CDK11, DK13a, DBBM11, DCR18, DL16, FHG15, GKPB13, HVM13, HTK12, IGL^{+}12, ITR^{+}18, KK12a, LLW15, LGK^{+}09, LS11a, MR11a, MHMM18, MPM10, MFP^{+}14, MR19, NTC^{+}11, NM11, OMO13, PT10, PDW11, RVZ14, RW12, STA15, SW11, SPG^{+}18, SH15, TXWW12, TXCW15, TCB13, VGZS18, VBHM^{+}13, WW12b, WR14, WML^{+}17, WDH^{+}11, XTZ^{+}13, Yan10a]. HIV- [STA15]. HIV-1 [LGK^{+}12, CH16, FHG15, GKPB13, IGL^{+}12, ITR^{+}18, LLW15, LGK^{+}09, MR19, NTC^{+}11, SW11, SPG^{+}18, SH15, TXCW15, VBHM^{+}13]. HIV-specific [RW12]. HIV/AIDS [DCR18, NM11, PT10, XTZ^{+}13, Yan10a]. hive [EM11]. Hjort [SPH12]. HLA [BYY^{+}11, JB19a, SSRA16]. HLA-A*0201-restricted [BYY^{+}11]. HMM [LPD^{+}16, SBT^{+}18]. HMM-HMM [LPD^{+}16]. HMW [DMM^{+}14]. Hodgkin [CPGF^{+}16, GL12b, MZM^{+}16, WCC13, WSLC14]. Hogeweg [SMS19]. Hohenberg [MROS15]. hold [HH10, Tra16]. Holling [CNM^{+}13, DS13, Mor10b]. hollow [CWB^{+}17, SSR^{+}19]. holobionts [O’M17]. holometabolous [iHM17]. Holotestoid [CS11b]. home [HAH19, WW12a]. Homeostasis [AGS18, BGL^{+}19, BLW11, CN12a, CKN^{+}12, GT11a, KSK^{+}16, LKP^{+}12, LLB^{+}18, LC16, MPP^{+}15]. homeostatic [TM16]. homeothermy [Kur14]. homing [SBW11]. Homo [KHH^{+}17]. homochirality [Tox18]. homogeneous [AR11, CFT11, CK16, FHSG15, JAHKH12, Kis10, LX10, MB18a, NdlPZA10, Oht12, PB16, Ros11, SS12a, VDD^{+}17]. homogenization [ASH15, EBP15, NP15, SAD17]. homogenize [FB12]. homolog [Mei12b, Mei13b]. Homologous [NSS^{+}11, KB11, LC11, MG14a, SM11b]. Homology [GTS15, KHH^{+}17, SRP16, BMSEE14, GNB^{+}13, RMM^{+}16b]. homophily [RMB15]. homoplasy [Cut15]. homotopy [WcW14]. honest [SHBT14]. Honesty [RZ16, Bul12b]. honey [BIMC17, MKJS13, NPS10]. honeybee [Gal10, LTSTD15, LWY^{+}16, SK16a, ZWY15]. honeybees [STKE12]. Hong [HHE19, ZLCH18]. Hooke [RBMPP^{+}15]. Hopf [BA10, SGKCP19]. hopping [RS14a]. Horizon [OO17a]. Horizontal [FXML18, BCF^{+}16a]. Hormonal [Yam16a, YI18a]. hormone [CTJ10, CFF12, CLA^{+}16, HBA10, MS13, MLL^{+}16, NH19, PHC12]. hormone-agonist [PHC12]. hormone-induced [CFF12]. hormone-releasing [CTJ10]. hormones [ZP11]. horses [ACR^{+}17, CS14b]. hospital [CL17a, HLTW14, PDW10, WXWL12, YTGW16]. hospital-acquired [PDW10]. Host [CCW16, DMS^{+}16, JB18b, RFME^{+}12, WDH^{+}11, AIB^{+}19, Ber14, BPR15, Bon10, BYM^{+}18, CR14, CL17a, CFZ14, CSBK15, CB10a, CS14b, CP11, DMM^{+}14, DFS11, DL15a, DRPM17, FB18b, FIJJ11, GCG14, GRBL13, HBEW12, HYQ^{+}16, Hur12, JCW13, Kor18, LM14, LMCW18, LvdBP12, MRPH17, MBBV14, MFG^{+}13, MGC13, MB12b, NS11, NBBCC18, NGS^{+}16, OMO13, PRC15, PKZ12, PVB^{+}10, PBB^{+}15b, PCL^{+}15, PSL^{+}10, RMF^{+}18, RC16, SR12, TWS15, TB15, VDD12, VB19b, WW19, WBMH12, WK18a, YT12, YCH^{+}17a, YCH^{+}17b]. Host-jumping [CCW16]. host-macroparasite [TWS15]. host-parasite [AIB^{+}19, CR14, CSBK15, GCG14]. host-parasitoid [HYQ^{+}16, PSL^{+}10]. Host-pathogen [JB18b, Bon10, FIJJ11, HBEW12, OMO13]. host-phage [JCW13]. host-population [PVB^{+}10]. hosts [BPFR16, DMM^{+}14, HASM17, JB18b, JH11]. hot [CZW^{+}11]. hotspots [LLCC12, LXC15, NK18]. house [Cro19]. house-hunting [Cro19]. Household [Ada16]. households [KN11]. hover [FH10a, FH10b]. hoverflies [DS12]. hovering [DS12, GAL11b, LBS^{+}10, LBS^{+}11, LS14a, LS19, NC16, XS13]. Hox [GG16]. HRCT [LMM^{+}13]. HRCT-derived [LMM^{+}13]. Hsp70 [DWG^{+}19]. HSR [RZIR16]. HTLV [LM14]. HTLV-I [LM14]. hull [TZY18]. Hum [WMK15]. Human [CzST^{+}18, EME^{+}16, HW11, IGG^{+}18, LLSO13, MMB14, PJL14, PTC13, YLL^{+}14, ABG^{+}13, ACM16, ANGPB12, Ant13, BCPL10, BFGAGA16, CPF^{+}13, CW15b, CZT^{+}16, CVPCV^{+}15, DBBW09, DBBW11, DMM^{+}14, DGD^{+}11, DTY12, EBP15, EMM10, FLM18, FS15a, FS16a, GGC14, GSGV11, GBC^{+}16, GMM15, GM16b, HHJR11, HLI15, HII^{+}19, HWW^{+}14, HKB10, HXL18b, HKH18, Jam12, KSA16, KKUM10, KLM14, KDFM14, KCJ^{+}11, LPGTBMA19, Li11, LXZQ14, LZL^{+}17, LZG^{+}19, LTP19, LTZ19, LT14, LGR^{+}12, LDF11b, MIJ16, MV10, MS16, MA11a, Meh17, MPS^{+}11, MTLW^{+}10, NH12, NTED^{+}19, NNH^{+}16, NI19, PRSC11, PVGV19, PRM12, RLW^{+}14, RKMG12, RS14a, RSD^{+}13, Ros10a, SGL10, Sel13, Sel15b, SSFG15, SSB^{+}15, SCF^{+}18, SDK11, Sue12, SCT12, TKK14, TI12a, WMK15, Zha11, Zha15, ZJSC16]. human-dominated [SCF^{+}18]. human-protein [WMK15]. Humans [Flo11, AOR17, BD10b, BVK10, LWLM18, Sch14, TR17, WVO^{+}15]. humidity [AK11, KMM18, UD10]. humile [vTMW15]. hummingbirds [SYR11]. hunted [HF17]. Hunting [AH17, CR14, Cro19, HH10, RW15b]. Huxley [CPGF^{+}16, GL12b, MZM^{+}16, MAS15b, WSLC14]. huxleyi [HWGT15a, HWGT15b, HWGT17]. Hybrid [CLCH10, FKV19, GE18, TZ13, AV19, AMBH^{+}13, AH16, BCPG18, BCS^{+}16, FKB^{+}12, GZFX14, HTK14, IGL^{+}12, JL16, KMMS17, KRR14, Kri16, LHH12, MEKK11, MK14d, NHSX14, PHA^{+}17, PGHC12, RBKW19, ZZG^{+}16]. hybridization [BJOS13, Sel16]. hybridizing [HK12]. Hydra [Cor16, DKS15]. hydrating [MHH13]. Hydration [LCG12, ØW16]. hydration-dependent [ØW16]. Hydraulic [Che16, KFG^{+}14, KC14, LS15c]. hydraulics [RHS14b, RHS14a]. hydrocarbons [KK15]. hydrocephalic [MA11a]. hydrocephalus [MAMEA15]. hydrochloride [PRM11]. hydrochloride-induced [PRM11]. Hydrodynamic [CMBP19, GI10, Ish16, SKSO12, WWIG19]. Hydrodynamics [DS16, TTW13, ICG16, NAK^{+}11, OSF11, RVA^{+}11]. Hydrogen [GY19, GZY12, Kun16, LF10, WJL^{+}16, vLFM^{+}19]. hydrogen-methanogen [WJL^{+}16]. hydrogenotrophic [LWvB^{+}19]. hydrolase [KJM17]. hydrology [Urs14]. hydromechanical [KC14, MAMEA15]. hydrophobic [FWSG17]. hydrophobicity [BKR14]. hydrostatic [NHE^{+}16]. hydrothermal [Dai12]. hyemale [KSRN19]. hyper [AY15, AANF16, PDT^{+}10, QF10]. hyper- [QF10]. hyper-responsiveness [PDT^{+}10]. hyper-synchronized [AANF16]. hyperactivated [OSF11, SOCF14]. hyperactivation [CKBCG12]. hyperarcs [MT14]. Hyperchaos [BA19]. hypercycles [GFS15]. hyperemia [WHS^{+}13]. hyperglycemia [SAI14]. hyperglycinemia [Sat18]. hypergraph [CPH^{+}14]. hypernetwork [PCC16]. hyperpredation [BH12]. hypersecretion [Goe15]. hypersensitive [ZZC^{+}17]. hypertension [JSK16]. hypertrophic [MCL^{+}11]. hypertrophy [CAGM^{+}17]. hypervariability [VBHM^{+}13]. hyphae [DEK15, YLY15]. Hypodicranotus [SKSO12]. hypofractionated [MM19]. hypoglycemic [QF10]. Hypopredation [BH12]. hypostome [SKSO12]. Hypotheses [Jam12, CS11b, JZZK11]. Hypothesis [BC15, BTK14, Fre10, GK16, Jam15, Kur14, All17, GA13, GU13, HD17, Jam11, Jam14, Mal15, Mal10, PRH^{+}11, RGL^{+}17, Rou14, SAG19, SDJ^{+}15, SPH12, TIMI^{+}15, WZ12, ZP11]. hypoxia [BB19, HLVR16, KK18, SH17]. hypoxia-regulated [BB19]. hysteresis [PDG17, PHC12].

I30M [AD15]. I30T [AD15]. iCDI [XMWC13]. iCDI-PseFpt [XMWC13]. Icelandic [EBS11]. idea [AH18]. ideal [ZL18b]. Identical [Ell15, FS15a, FS16a, MAFK12]. Identifiability [AHR10, CK15, ERT13, FM18, EJ17, EJ19, LCSH14, RSV10]. identifiable [HSMCCR19]. Identification [AH18, FLW16, HA15a, IHNS16, JSC^{+}16, LFW^{+}15, LLY15, SBT^{+}18, SSSD17, STG19, SCKL15, SSRA16, THC^{+}11, WDH^{+}16, WZM19, Yin17, ZNA^{+}16, ZXS^{+}19, CMN^{+}11, DQS^{+}15, DDTL19, GZFX14, GRB^{+}13, HW11, HGM^{+}16, JSZ12, KHK^{+}17, KS10, LZL^{+}17, LLL^{+}17, MPZK16, PRSC11, RRG^{+}12, SPSM15, SBR16, VBV10, YWW^{+}14, YC15, YZ19, ZNCM15]. identified [MG14a]. identifies [MBBV14, TDZ^{+}18]. Identify [JLQ^{+}19, KHKI17, KF15, LZ18, TTC19, XMWC13, ZK18, ACdlRMR^{+}11, GGL^{+}16, NRS^{+}16, NKM^{+}12, YLC16]. Identifying [CA17, CG11, GYWJ10, GZ14a, GZ14b, HMMSRSD15, HPvdB17, LP19, LLW^{+}18, PGLZ14, PBEI12, SIK^{+}18, SFMS16, SLHS13, UPWK15, WDL^{+}17, YGL^{+}10, ZZC14, ZZN^{+}19, ZSO16, CLC11, Di 19b, HKR^{+}19, JLX^{+}15, JCG15, LTS^{+}16, NLB14, QZ14, WMK16, XWC11]. identity [SLW18b]. ideologies [DBD17]. idiotypic [LPB17, LPB18, MBF^{+}15]. if [Bal13, IL13, YSYI13, vV11b]. IFN [Gün12, Gün13, VZB^{+}15, ZXC^{+}10]. IFS [MS14a]. IG [Goe15, LZL^{+}19]. IgM [OBK^{+}11]. II [BPF15b, AOM19, Akç12, BPF15a, CCB11a, Das18, DS13, Fra19b, FJC^{+}10, HWMT17, HRD14, HH16, KGP^{+}15, KDL16, LW10b, MKBE17b, MTE15, MFKS13, NRS^{+}16, OAJK10, RHS14b, RBRB16, XYZ15]. III [BTF19, SAKG13, ACdlRMR^{+}11, CNM^{+}13, DS13, Mor10b, SVA18]. IL-10 [MMFK10]. IL-2 [GML10, Lab16]. illegal [LSDI15]. illness [EAN14, vNÁBG12]. illumination [DM15, PP18, PP19]. illumination-invariant [DM15]. illusion [SSF15]. illustrate [DBBW09, DBBW11]. illustrated [UI10]. iLM [JCG15]. iLM-2L [JCG15]. iLoc [XWC11]. iLoc-Virus [XWC11]. image [BMI19, CLL18, CZPC^{+}18, MAR^{+}17c, NSBL10, Nos14]. images [KKOM18, KSM^{+}15, LYK12, RMM^{+}16a, TKK14]. Imaging [PH13, KKO^{+}18]. imbalanced [SM17a]. iMem [AHJ18b]. iMem-2LSAAC [AHJ18b]. iMethyl [AH18]. iMethyl-STTNC [AH18]. imitate [CCNT19]. Imitation [Gal11a, WYKT17, Mas14, dMP11]. imitation-game [dMP11]. Immature [BBM^{+}97, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c]. immigration [EAN14, GKW16, Wei17]. immobile [KGM^{+}15a]. immortal [Wal12]. Immune [GBM18, Kar15, Kor18, yTRSC13, RDTHSM13, WBRHE10, AK13, ABGM11, ARG14, ABV19, AN10, AHAFA14, BBS18a, BAM^{+}11, CWM11, CdMB10, CKB18, CP11, CTL^{+}15, DFM^{+}14, FMLM12, FHS^{+}14, GLR^{+}18, GL18a, HI19, HL19, LHFM16, LW17, LMYL15, LBF13, wLW10, MCL19, ME15, MOEdP16, MRF19, MGM13a, NKB16b, OAS10, PVB^{+}10, PEZ^{+}13, RH19, RTEKG12, RTEKG19a, SBSR13, SR12, SMA11, SOF16, TXCW15, VAG16, WDH^{+}11, WCM15, YCH^{+}17a, YCH^{+}17b, d’O12]. immune-mediated [RH19]. Immunetworks [DEN^{+}11]. immunity [CST16, CMD^{+}10, DMM^{+}14, EBE10, FGH^{+}14, HTB^{+}13, LKO^{+}17, LHFM16, LM14, LG12, MMLK11, Moo14, PVB^{+}10, PCL^{+}15, PRM14, RTEKG15, RTEKG19b, SDT17, STN19a, WKB13]. Immunization [SMC^{+}13, CW15a, LDJW16, NW10, dMP11]. Immuno [BFH^{+}15]. Immuno-modulatory [BFH^{+}15]. immunodeficiency [HII^{+}19, SBCR10]. immunodominance [CP11]. immunodynamic [RKK^{+}19]. immunoediting [KB15]. immunogenic [BBT^{+}15, KB11, SSRA16]. immunogenicity [CRH^{+}15]. Immunoinformatics [TWR^{+}18]. immunological [MYC12]. immunopathology [VC15]. immunoprecipitation [FKMP10]. immunoprophylaxis [WW12b]. immunostimulation [RKK^{+}19]. immunostimulation/immunodynamic [RKK^{+}19]. immunosuppression [SSM^{+}18]. immunosuppressive [LHL^{+}14]. immunotherapies [EH17]. immunotherapy [BMGC11a, BMGC11b, GMB11, HI17, JO13, LX16]. immunotoxicants [BYM^{+}18]. Impact [AD15, BAR14, FZL18, GL12a, HLTW14, HvLPdB15, JWTW17, PIT12, PT10, PBvdG10, PTC13, RVZ14, SSG^{+}19, TD18, TC12, TRG16, VR17, AKR^{+}18, ADB^{+}13, BM12, BTG^{+}15, Bon10, BKPV15, CH18, DBD15, DBBM11, DKF17, HWGW18, HB16a, HAH19, HJR12, HPW^{+}13, HWL15, JP19, Kle14, KJ11a, KM15, Kur10, LTZ17, LZY^{+}15, LNH13, LYZ^{+}18, MHMM18, MPP^{+}16a, MB18a, MS12a, MBC^{+}12a, MS16, MB19, NDZMA14, NGS^{+}16, PAS10, PSJ15, PNVN18, Rao12, SEK^{+}10, SCC11, SW11, SH17, SB17, STLJ18, WHYMG17, YSH^{+}14, YTGW16, YCR^{+}15, dMP11]. Impacts [TB18, WCSA12, AB16, GK13, MSZW19, NIiT18b, ZADB15]. impaired [HBA17]. impairment [RTR^{+}13]. impairments [ML12b]. Impasses [VFS^{+}15]. imperfect [AM10, Gha16, MLMG^{+}15]. Imperialist [KKGN16]. implant [PVGA12]. implantation [NPMM15]. implementation [CA17]. Implications [BZJP18, BGCB12, BYM^{+}18, CWM11, DTGC14, GTC19, HMJB15, LG13, LMCW18, LH18, MKKS14, OO18, QA15, SSM^{+}18, TF15, ZARK19, ABKS11, BJWL15, BDE^{+}14, BAM^{+}11, Cro17, FBFM12, GL14, Gin10, Gol16, GK10, KSBvL^{+}12, LCGMH12, MGO^{+}15, MKMG^{+}14, NMAZP16, PG18b, PMCS16, RCH14, RGBR17, TEY10, TEY16, VH11b, WK17, dM10b, MFP^{+}14]. imply [RWH16]. import [ABKS11]. Importance [BRG^{+}12, EUM^{+}16, BUL12a, CB10b, FB12, JMS12, KBF18, LHW14, cLCJ^{+}10, LD18, MvdDW13, PW19, SSP15, SPL14, TZ11, WD12, vVE15]. important [Cle18a, EKvdKvFK13, IK15, SHN12, Sos18, TAR16]. importation [CBC^{+}18, NI11]. imported [EAN14, NI11]. improve [CR14, CLP11, IA17, MAW12, NS19, SS15c, WHL13]. Improved [HM13, IMH15, Mou12b, SVCS10, CW15a, MB12a, MSRW16, YY15]. improves [Mar19, MP14b]. Improving [FW15, MHC16, RSD^{+}16, WSTL16, ZLY14]. impulse [RAR19]. Impulsive [BRK16, Yan10a, LY11, LZT12, LLL13, LZT13, ML10]. imputation [MHKA16]. in-frame [SSCJ16]. in-group [KM12, MJI14]. in-host [CL17a]. in-mouth [DDF^{+}14]. In-silico [JSC^{+}16, MÇ15a, RAA^{+}16, GHJ14, LKAJ18, LZL^{+}19, MÇ16b]. In-situ [KFG^{+}14]. inactivation [KAKK19, Pie12]. inbreeding [Cam15, YK11]. Incentives [Akç12, Arc11, CSZT12, TLW18]. incidence [KMCJ17, PGGvdB12]. incisions [KHB^{+}18]. inclination [TSdL14]. inclined [PC17]. include [CK19, FHSG15, HP12]. including [AAJ15, CN10, FIS11, FIS16a, GCSMMG^{+}12, MP13a, MP13c]. inclusion [KDMD13, TWH^{+}13]. Inclusive [OWB14, Tay13, Arc13, Mar11, TGL15, vVGSE12, vVLS14]. incoherent [YS14]. incompatibility [Cut15, VSLVB15, YI16]. Incomplete [BLNR15, bCRS17, TEY10, TEW12]. incompletely [KSP12]. incompressible [LZWK11]. inconsistency [SN11, SN14]. incorporates [ZGW16]. Incorporating [CG14, OABI12, BRK19, CTJ10, CMS^{+}19, CL13a, CCY^{+}19, CT18b, DHS^{+}15, EG10, FL12, FL13, FLW16, HYA14, JLX^{+}15, JLQ^{+}19, JD17, JCG15, JCG16, KHKI17, KZL14, Kri16, LZ18, NBL14, QLC^{+}18, SM18, WMK13, WTC16, XDD^{+}15, XSLZ16, ZCAB17, ZK18, ZNA^{+}16]. Incorporation [BUC14, PFPR11, APW10]. incorrect [RMB15]. increase [OO11, RT19, TSB10, VMZK^{+}19, WMT10]. Increased [HBEW12, PVCEC18, ZZL19, Gri15b, LNH13, ZGW16]. increases [Bar11, CNDK15, Fra19b, GSSBF18, HPvdB17, yTRSC13, ML11a]. increasing [AK15]. increasingly [RW14b]. incubation [GMS^{+}13, KN11, NI11]. Independence [Gal10, TSB10]. Independent [FD17, Ell15, LZY^{+}15, Mun12]. indeterminate [PLD^{+}17]. Index [WF17, CXT19, PCB14, SNY^{+}13]. indexes [Su16]. India [MFP^{+}14]. Indian [MCCC^{+}10]. indica [TVA18]. indicate [ARB13]. indicates [SN17, YM14]. indicator [GT14]. indicators [DSTD19, Dol16]. indicaxanthin [TVA18]. indices [ACdlRMR^{+}11, FBAPMD13, WF17]. Indirect [BBD18, GN15, MJI14, NO14a, NO14b, TSM13a, TSM13b, CB16, DLM^{+}19, FB12, KDG13, MHC16, Mas12, OSN18, Pan11, Sig12, US10]. INDISIM [GGGM16]. INDISIM-Paracoccus [GGGM16]. indispensable [LPS^{+}14]. Individual [BBWS^{+}13, DBG18, GW13, HS14b, IMD16, STI13, SH17, AMSSG16, BGD^{+}13, BC13, BLS^{+}17b, DB19, Dim10, DDF^{+}14, EPP19, Fil10, FGLS10, FM15b, GCM17b, GGGM16, HFT^{+}18, HMM16, JB11, KKO^{+}18, KAKK19, LP14, MCL19, MFG14, MPNP12, NNH^{+}16, QMJW15, SMG18, Sek12, TKKE19, YS11, Zuc14, vLSBBD11]. Individual-based [GW13, IMD16, SH17, GGGM16, MCL19, SMG18, YS11]. individual-cell [MFG14]. individuality [KC13]. individualized [RD14]. individuals [CWM11, FS16c, HSMCCR19, HR12, Mas14, MP13a, PGH11, TA15, VHM^{+}17, WCM15]. individuated [Vel17]. indivisible [KNT10]. induce [ESG^{+}16, MGT17, TIS10, ZYD15]. induced [ASH15, AGC18, AP13a, BCKS13, BKL^{+}15, BGD^{+}13, BKPV15, CKZ^{+}17, CFF12, CXT19, CL18, FBK11, FPG11, GKMC10, GMNY14, GL12b, HKH18, IC11, Kar16, KKWA18, KTT^{+}19, KK18, LHD15, LWRE14, Lee16a, LXZQ14, LMM18, LMdL11, LL17b, MB17, MNH^{+}12, MPP^{+}18, MZMM^{+}14, MSA^{+}16, PRM11, PHTP^{+}12, PDNP16, PJ18, PRN10, PBD11, RBSD10, SSM^{+}18, SXZY12, Ste12, SLD^{+}17, TST^{+}13, ZLL17, ZMS17]. induces [Has14, MMC19b, NS18, SSS17, SY12]. Inducing [CWP^{+}18, NS19]. induction [CKB19, DWG^{+}19, GJ15b, LF16, MTE15, Mic13]. Industry [RFME^{+}12]. inertial [RS12, VDH^{+}15]. inevitable [TLPH11]. infancy [BFGAGA16, PR17]. infanticide [LPvSP11, LPvSP12]. infants [SEK^{+}10]. infected [CLK18, Gru18, HII^{+}19, LFB^{+}16, LLD^{+}17, MB12b, STA15, SPMGR10, VG13, ZLT^{+}19]. Infection [RW12, AWK15, CWM11, CKZ^{+}17, CL17a, CM12, CXT19, CH16, CDK11, CS14b, DJD10, DL16, FCC^{+}10, FGH^{+}14, FHS^{+}14, GKPB13, GM16a, GC18, GN10, HT19, HYN19, HN19, HHA^{+}13, HAH19, HVM13, HTK12, ITR^{+}18, JMH^{+}19, JEA18, JCW13, KMN^{+}15, KOS13, KNA^{+}18, LAH^{+}16, LGK^{+}09, LGK^{+}12, Lev14, LLDW14, LMYL15, LM14, LMCW18, LCT11, LBZ18, MRPH17, MK19, MMLK11, MSE^{+}14, MMP13, MG15, NBBCC18, OMO13, PDRC16, PRN10, PCC15, PSL^{+}10, RVZ14, RMF^{+}18, RM10, RW14a, RS14c, STN19a, SBCR10, SW11, SPRF13, SMA11, SAM17, SLJ^{+}10, TB18, VDD12, WW12b, WWY12, WDL^{+}17, WCM15, XSTW14, Yan10a, YCH^{+}17a, YCH^{+}17b, ZZL19]. infection-induced [PRN10]. Infections [TGC^{+}17, ABKS11, BN15, BMB^{+}18a, CMD^{+}10, CH18, Gru18, GKG^{+}18, HE16, Hur12, LH14, MBC^{+}12a, Nås12, Ric17, SDT17, SCC11, TB18, VAG16, Voe18, WXWL12, WBMH12]. infectious [AL17, BB10, CT18a, GGL^{+}16, GMR16, HHN14, IMW15, JH11, RSH19, RSD12, Ros11, SBK16, TXY^{+}12, TDW16, WW12a, ZZL19]. infectivity [HR14, KN11]. infer [BTF19]. Inference [AIY16, BSKV18, Bac15, BP18, BSB^{+}13, DN18, GK16, GMR16, IS12, JB11, LKO^{+}17, MS14c, NdMLLBLC13, OGE10, SMD^{+}16, SSP15, SS15d, Ste13, WS10, XSKA17, ZT16, ZSSM14]. inferences [Ell15, TF15]. inferior [NPMM15]. Inferring [AGRRTR11, AMRRI10, BMB^{+}18b, bCRS17, DRPM17, KDS18, LBB^{+}13, PHPM18, RRTR12, RNN15, TLS^{+}13, SLT^{+}18]. infiltration [KF10]. infiltrative [DTGC14]. infinitely [Hay16]. inflammasome [VLPH17]. inflammation [AN10, AES^{+}13, BFJ^{+}18, MBBV14, OPS^{+}19, SCA13, WBRHE10, ZFWK17]. inflammatory [CSDR15, LHW14, OSV^{+}16, PYG^{+}19, PMKS^{+}15, RZRSC19, REC10, SCLA10, WBRHE10, WLL^{+}14]. inflated [Mog15]. inflorescence [SSI^{+}16]. Influence [BGM19, CCB11b, CZT^{+}16, FKK14, Gau11, JOM16, Kro11, PCBMM12, SLJ^{+}10, WW12a, WSLC14, WML^{+}17, ZCSRG12, BKR14, BHBH11, CBBO16, CTB18, DI13, DCR18, DvDBD15, Edu16, GDF17, GPR^{+}16, GY18, IPS14, MBPS17, MBK^{+}11, NHE^{+}16, NDG14, NZZ19, PPM12, PMP13, PB18, RM10, RS14a, RRB10, SCLA10, SP13a, SS15d, SLS17, WGS10, ZW16, dLMD^{+}11]. Influences [SHvHB16, HR14, Ing16, MM18, PGH11]. influencing [DDF^{+}14, MJV16, ØW16]. influenza [BTW11, CW15b, DD18, DGD^{+}11, HHE19, KLJ17, KJ19, Lan16, LCCC10, LCCK12, LKK13, LJ15, LCJ18, LFB^{+}16, MDE11, MZAI19, MAA18, MMP13, NI11, OAS10, OS13, PGGvL^{+}19, PBB^{+}15b, PMKS^{+}15, SY11, TLT^{+}15, TGZF11, TC12, UD10, YCH^{+}17a, YCH^{+}17b, vNÁBG12, FHS^{+}14, XSTW14]. influenza-infected [LFB^{+}16]. influenza-like [vNÁBG12]. influenzae [DMM^{+}14, KM15]. influx [SOBC12]. Information [Bar18, MBE10, MWSM13, Mau15, PEW18, Wal16, ACT12, BORA10, BBD18, BP16, CC11b, CWW18, CDM^{+}14, CB16, FL13, FLW16, FRP14, GSF13, GGX^{+}10, KOS13, KZL14, Kri16, KRSD11, KST17, KA17, Kuz19, LLSO13, LTSTD15, LL12, LLW^{+}18, LJ10, LLJ18, LPH18, MK14b, MS12a, MRPLAS15, MP12, Mit14, Nem17, NDG14, NHM10, Pia19, RMB15, SN16, SSCJ16, STG19, SLO10, TP10b, WSH^{+}10, WYX^{+}17, WHT17, WZ12, WLY^{+}17, Wol11, XLL^{+}10, XNJ^{+}13, YZ19, ZP11, ZCAB17, ZZK14]. Information-theoretic [Bar18, MWSM13]. informational [DPRSS11, ZLW16]. informed [BGLR^{+}19, JWTW17]. ingredient [MSIR10]. ingredients [CCAdS13]. Ingroup [Mas12, Jan15]. inhalation [DFS11]. inhalational [GMS^{+}13]. inheritance [ASP13, BCF^{+}16a, Gar11, Sha14]. inhibit [KJM17, POP12]. inhibition [HAuR^{+}18, JP13, Kar15, Lee16a, LL14, MPZK16, PJ13, PMK15, SZ18, SBB18, SER^{+}12, TTB^{+}18, TSB^{+}17, VACGF17, ZXC^{+}10]. inhibition-induced [Lee16a]. Inhibitor [DS15b, GPD18, LF19, LS11a, RAR19]. inhibitors [ARG19, CCB11b, DGD^{+}11, GBC^{+}16, IHNS16, Ken19, KRR14, MFZ18, MR18, MR19, NRS^{+}16, SBR16, TMS13, TA16]. Inhibitory [KTH16, BB11b, FHG15, Hor11b, Pat16, PJ18]. inhibits [GDB^{+}19]. inhomogeneous [STA15]. Initial [GGC14, DBD15, EBX17, JSP^{+}16, KAKK19, Pan11, PL14a, PL14b, PW11, WDH^{+}11]. initiate [MYN^{+}15]. Initiation [LWB^{+}15, BST14, Coh15a, GYWJ10, GRRG16, KDMD13, KSvdH18, SLM17, SGD^{+}16, WKS14]. initiator [ZS12]. initio [Cle16, Kue16]. inject [PBSM19, RCL16]. injecting [DK13a, MVH11, RDSD^{+}12]. injection [ABN15, ARM18]. injections [SAI14]. injunctive [SAB17]. injuries [ÁTC^{+}14, MYK^{+}11]. injury [PNP^{+}16, POP^{+}19]. innate [ARG14, FHS^{+}14, LW17, SMA11]. inner [MB14]. innovate [CCNT19]. Innovation [PSK10]. inoculum [LH14]. inorganic [Bur10, HWGT15b, HWGT17]. input [LLL13, NK18, RA13]. Insect [DM15, DPRSS11, FH10a, FH10b, GT15, GSF13, IH17, IS17, KA11, LBS^{+}10, LBS^{+}11, LBGW13, LS18, NC15, WPA17]. insecticide [HPvdB17, HS14b, NMAZP16]. insecticide-treated [HS14b, NMAZP16]. Insects [BBM^{+}97, ABA11, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, iHM17, JLH^{+}15b, TT10b, WBMH12]. insert [Sel12b]. inserted [TKTB18]. insertions [HTT19]. Insight [NBA^{+}18, SMJS14, dBB14b, HP16, RZIR16]. Insights [ALM^{+}19, BCPL10, BTW11, BCPG18, CJRR11, Dai12, DHB15, GBJ^{+}11, LFZN11, LC10, MKMG^{+}14, MZMM^{+}14, NS16, PDF18, PMCS16, SK19, SG15a, SLD^{+}17, Tal12, VKKA12, ADS^{+}19, ASRM15, ÁTC^{+}14, DQY^{+}15, HNP18, HAuR^{+}18, KHH^{+}17, PBSM19, PGHC12, Zho11, AKNP18]. Insilico [MNSZ16]. inspired [AANF16, LA15, LDJW16]. Instabilities [GGG12, NvHM13]. instability [AK16, ACvKA10, CST^{+}12, EJK16, GF11, KGM^{+}15a, SA17, SAG19, SYR11, TIMI^{+}15, Van15, WHH17]. instantaneous [UJLG14]. institution [Gha16]. institutional [CSZT12]. institutions [Ant13, SHBT14]. instrument [Gha16]. insufficiency [Sim14]. insufficient [KI17]. Insulin [Goe15, BSR14, CTW10, GdGL17, GA14, GS17b, HWD^{+}14, HBA17, LLB^{+}18, SW13b, SW19, SAI14, WCHC11]. insulin-dependent [LLB^{+}18]. insulin-mediated [GS17b]. insurance [SDJ^{+}15, TLW18, ZCW13]. intake [LLK14, WSRG18]. integral [ABIM10, BGL^{+}19]. integrase [WHZ^{+}17]. Integrate [HTK^{+}18, CP14, LLW15]. Integrated [FTPN10, MB18b, SW19, ACH18, DKP^{+}18, FP18, HWPL12, IHNS16, MVGGB18, PGS10, RPGG^{+}19, SAKG13, TTC10, WBB^{+}18]. Integrating [APBS15, CLA^{+}16, YYST13, CRGS18, LLW^{+}18, PCL^{+}15, STG19, SZ10, YLWZ10, ZL18a]. Integration [BS13, KSA16, Ben14, Che14, CV15, GC18, Jen10, KJH^{+}19, LNL^{+}11, PWC^{+}12, PPM12, RB15, VC10, VM12]. integrative [Kuz19, SMHB10, SBR16, SLT^{+}18]. Integrin [CML11, BZJP18]. integrin-binding [BZJP18]. integrity [SW19]. integro [Zad11]. integro-partial [Zad11]. intelligence [HL19, MFM16]. Intelligent [EM11, KHK^{+}17]. intensification [CLPZ14, WML^{+}17]. intensities [jWlGQ19]. intensity [AM11, CWW18, HWGT15b, MM18, Tay16a, VM16]. intensive [BAGG14]. Inter [Don17, HH13, RG18, TCR13, CL17a, LP14, LF10, LLL^{+}17, STLJ18]. Inter-airway [Don17]. inter-annual [STLJ18]. inter-base-pair [LF10]. Inter-filament [RG18]. inter-host [CL17a]. inter-individual [LP14]. inter-nucleotide [LLL^{+}17]. Inter-residue [HH13]. Inter-species [TCR13]. interact [HPM^{+}17]. interacting [BGKL17, FS16c, HBSF11, JSB15, Lab16, MPY14, MK19, PDM17, Thu15, WCSA12]. Interaction [JC10, KC17, LS14b, PJL14, XLSF19, ARG14, AEZR^{+}16, ALM^{+}19, BZJP18, Boc12, BLS^{+}12, CZW^{+}11, CK19, DS15b, DK13b, DSPM14, EBE14, ER18, GI10, GCG14, HNL^{+}11, HVSZ10, HYZS14, HH13, JB19a, JR17a, KSK^{+}17, KS10, LS15a, LWB^{+}15, LX10, LL12, cLCJ^{+}10, LYZ^{+}18, LMJ^{+}18, LSLL13, MMY^{+}12, Mal18, MG14a, MR12, MGM13a, MMCZM12, MA11b, ML11b, Nak14, NLB14, NC15, PRC15, PL14a, PL14b, PM17, PSAA13, QW12, QZY17, QJR^{+}16, RAF^{+}19, RAR19, SW15, SAB^{+}19, SN14, Smi11, SF15, SAZ^{+}14, TIMI^{+}15, TG10, TRJD19, UI13, UPWK15, VL11, WD11, XMWC13, YWW^{+}14, ZMAM19, ZJSC16, ZT16, ZLB13, Zho11]. interaction-based [UPWK15]. Interactions [ADV^{+}10, BBM^{+}97, BYM^{+}18, KCMF11, Mal17, XSMF12, APBS15, AP17, AES^{+}13, ÁTC^{+}14, BPC15, BST14, BS19a, BH11, BSR^{+}11, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, Bon10, BR12b, CVH^{+}10, CW13, CCS18, CMCS18, CB10a, DBG18, FLM18, FGPR10, FIJJ11, GWBL12, GSRR17, GMZM15, GAGP^{+}14, GL14, GN15, GMB11, Gün12, Gün13, HMMSRSD15, HB13a, HYQ^{+}16, IBB^{+}15, Ing16, JB18a, JLX^{+}15, JLQ^{+}19, JA13, KA11, Kur10, Lai18, LDH^{+}12, LW15, LCL14, MCL19, MRF19, MGGM10b, MALAN17, MMB14, NKB16a, NTOI16, ØW16, OMO13, PCN17, RZ16, RTEKG12, RTEKG19a, RR12, RM13, SK19, SMS17, SFMS16, SLO10, TBQG14, Tay16b, TWC^{+}19, WRZ^{+}12, WYX^{+}17, WYL^{+}19, WLY^{+}17, WGC13, WHL13, WFW19, YYC19, YLW^{+}14, YC13, ZJS11, ZCAB17, ZPHS11, ZZ18, ZRGW19, ZLW18, ZK10, vLSBBD11]. interactions-in [Lai18]. interactive [SLW18b]. Interactome [LPS^{+}14, MGO^{+}15, KSA16]. interacts [Don17, ZSCL16a]. interbreeding [HT13]. intercellular [BSP18, EG10, PRV^{+}14, PC10b, TS17a]. interceptor [BE14]. interconnected [PGKZ17]. intercourse [MHMM18]. interdependence [Gal10]. interdependent [VNS18]. interface [DI13, JR17a, NTOI16, YG18]. interfaces [BBWS^{+}13, KC11b, MMY^{+}12, ZJS11]. Interference [ACMK12, GP11b, NKB16a, ZADB15]. interferon [HYN19]. interferon-gamma [HYN19]. intergenerational [CL13b, SN14]. intergroup [KM12, Mas12]. interior [dSMP^{+}11]. interleukin [BMGC11a, BMGC11b, JO13, MBBV14]. interleukin- [JO13]. interleukin-6 [MBBV14]. interlocked [UT17]. intermediate [DLEMP19, Por13, RKJ^{+}11, Wax11b, ZYD15]. Intermittent [DD18, SNCM12, DDTL19, HHSA15, HA15b, MSIR10, PP18, PP19, SSN^{+}14, WR14, YZY^{+}16]. intermolecular [DMP15, SK19]. Internal [HBW^{+}11, Cle18a, DTY12, Gal11a, MHC16, WZ18]. internal/external [Cle18a]. Internet [HASM17]. interneurons [GMK18]. interphase [AA16, Fer12]. Interplay [BMB^{+}18a, NMAZP16, SYY17a, SYY17b, WHT17, AHW13, AWRD15, BI19, DAA17, GML10, GN10, HB16b, KFL12, MS12a, MD16, MLL^{+}16]. interpolar [DL15b]. interpolation [RP18, SRD^{+}15]. interpretability [RG17]. interpretable [WMK15]. interpretation [KDST15, MR10, Mun12, SPMGR10]. interpretations [ACM16]. Interpreting [SBCR10]. Interrogating [WRZ^{+}12]. interrupt [ZXH19]. interscan [BR16]. intersecting [DEN^{+}11]. Interspecies [SB12, FSA15, MC15b]. interspecific [AS13, Cam11, Edu16, PS12, Shi19]. Interstitial [HWM^{+}16, LS15c, LCA^{+}15, LS16c, MS10a, WFM^{+}13, WFC^{+}14]. Interval [AH15, CLC11, CR16, DLYW13, MDM^{+}10, ZZ14b]. Interval-based [CLC11]. intervality [CAG13, ZEJA11]. intervals [KK17b, NSO15]. intervention [AB16, CJL15, DK13a, HS15, KdlRJ18, LJ15, LSSG10]. interventions [BSR14, HAH19, LCCK12, LCJ16, POK^{+}12]. intestinal [BZL17, DOC17, HI19, PCT11, PCT19, Pan19, SBSR13]. intestine [ISM^{+}11, SDK11, TLL^{+}12, YSI17, YSI18]. intimal [Fok12]. Intra [Sek12, CMJD11, Ezo19, FRP14, IEN15, PDF18]. intra-annual [CMJD11]. intra-cellular [FRP14]. intra-guild [Ezo19, IEN15]. Intra-individual [Sek12]. intra-tumoral [PDF18]. intracellular [AHW13, BTO14, CCF^{+}14, CKN^{+}12, DCN13, GN10, GDB^{+}19, LXZQ14, MDEKH13, MLT10, MK18b, NS11, Nak16a, Pel18, SBJ^{+}18, STX^{+}11, SPL14, TC11, WTC10]. intracellular-to-extracellular [SBJ^{+}18]. intragenomic [Sak10]. intraguild [AF10, SvK12, VA10]. intrahost [MSE^{+}14]. intraluminal [VF12]. intramammary [GKG^{+}18]. intramembranous [CGRRGM^{+}16]. Intransitivity [LRGA13, LS15a]. intraplaque [GCYL18]. intraspecific [Edu16]. intratumour [LVLC18]. intrauterine [Jam14]. intricate [MK16a].

Intrinsic [DCS14, LYW11, GLF^{+}15, HFT15, IU10, KFW12, Lei09, Lei10, MPM10, PR18, PMD^{+}18, RCL^{+}10, RSC14, THC^{+}11]. introduce [SPA18]. introduced [BGM19]. introducing [SGCD14]. Introduction [GKK13]. intron [NHTS14, ZGZ^{+}10]. intron-positions [NHTS14]. introns [BLZ^{+}19, ZLZ^{+}15b, ZLZ^{+}16, ZLB13]. Intruder [CK19]. intrusion [HK10]. inundative [BAR14]. invader [DG16]. invades [Bar11]. invading [ASGG17]. invadopodia [SRK^{+}12]. invagination [TIMI^{+}15]. invariable [CK15]. invariance [BP15]. invariant [DM15, GGL^{+}16]. invariants [KMD^{+}12]. invasibility [PC10a]. Invasion [AP19, BV12, BWY^{+}17, BSL^{+}17, BDR10, FNH10, MMG^{+}16, NP15, Ros11, ACMK12, BCF^{+}16a, BPR15, CC11a, CNL14, DE17, DTGC14, ESG10, FJC^{+}10, GCSP17, GFM^{+}19, GRBL13, GP12, HPM^{+}17, KPEK14, MKKS14, MGGM10b, MLCH11, PAS10, PH13, PM11a, SMS17, SSZB15, SP13a, THC^{+}11, WB19]. Invasions [Bal10, VIBC12, BBWS^{+}13, LK19, PR13]. invasive [AP19, BJPP12, CPV16, Dra19, MSP19, MJV16]. invented [KK12b]. Inverse [RBB^{+}12, Dol16]. Inversion [RMRC^{+}16, WcW14]. invertase [Sel12a]. inverted [BRB18, SNCM12]. investigate [APBS15, ACM16, LPP^{+}19, SH15]. investigated [KGM15b]. Investigating [AHD^{+}18, AGC18, KGM^{+}15a, MH14, PMK15, TK10c, HPB^{+}14, HWGT15a]. Investigation [AHJ^{+}18a, HL14, SCLC13, YIS^{+}17, EH17, FKMP10, HBA17, KMM18, KAN11, LBS^{+}10, LBS^{+}11, LBW^{+}13, MFMB12, PZS^{+}10, PDW11, SEK^{+}10, SVA18, SSRA16, TDT16]. investigations [GAY^{+}15]. investment [AOR17, BHR10, ER12, FPD15, LvBJ16, MG17, TD17, WRC^{+}19]. investments [VSP11]. invicta [AC12]. involution [MZMM^{+}14]. involved [ADCG14, CL10, GRR^{+}14, HBW^{+}11, ZXS^{+}19]. Involvement [KHS13]. involving [AN13, KSU^{+}18, KIT^{+}16, SVA18]. iodine [KSM^{+}15]. iodine-enhanced [KSM^{+}15]. ion [Bec14, CPGF^{+}16, CN12a, HZC^{+}10, LSP13, LD11, LD18, MFZ18, MSC10, PDG17, ZZG^{+}16]. ionizing [LHD^{+}17]. ions [Bur10, FM14b]. IP [HP12]. IPF [CPF^{+}13]. iPHLoc [SSSD17]. iPHLoc-ES [SSSD17]. iPPI [JLX^{+}15, JLQ^{+}19]. iPPI-Esml [JLX^{+}15]. iPPI-PseAAC [JLQ^{+}19]. iRNA [TTC19]. iRNA-PseKNC [TTC19]. iron [Ber14, CKN^{+}12, DWM15, FE10, FE11, MBC^{+}12b]. iron-starvation [MBC^{+}12b]. irradiated [TRM^{+}14]. irradiation [AP13a, KTO^{+}14]. irradiation-induced [AP13a]. irrespective [BMSEE14]. Irreversible [ML09, ML12a, Jia16]. irritable [YXH^{+}14]. IRS [HWD^{+}14]. iRSpot [ZK18]. iRSpot-ADPM [ZK18]. ischemic [SXZY12]. ISDTool [LHL^{+}14]. island [HA15a, TLY^{+}19]. Islands [SFT14, CM14, FKB^{+}12, LPS^{+}14]. islet [SW19]. isocline [KP15]. isoforms [SG15a]. isolate [SPG^{+}18]. isolated [BC19, FRH15, Kur14]. isolates [AHMA^{+}19]. isolating [MYN^{+}13]. isolation [BdABY13, HAH19, KK15, LCCC10, PMD11, SPCM18]. isometric [DGNT17]. isomorphic [BI10]. Isoniazid [LBMG16, LGMV14]. isoprene [KKU^{+}10]. isothermal [LDdA^{+}13]. isotonic [DGNT17]. isotope [CFR^{+}14, CMR^{+}18, HWGT17, Rem15]. Israeli [LG18]. ISSR [CMPS17]. issue [GKK13, TBB14, TK19]. issues [DBM^{+}18, KLHS17, RBF15].

Iterated [CZ14, GCS12, MN11, QMJW15]. iterative [OO18, QQJW13, ZNCM15]. Iturin [KSK^{+}17]. Ixodes [DB11, WDL^{+}13, DDSDW13].

J [AEM14, BPF15b, BBM^{+}13a, BBM^{+}13b, BPGS12a, BMGC11a, CLG^{+}11, CNG^{+}12, DBBW11, EPJ^{+}11, FIS16a, FS16a, GZ14b, GS14, dOGL13, Gün13, JN14, JRG14a, JRG14b, JCLS^{+}21, KDK14a, KMLT14, LBS^{+}11, LGK^{+}12, Lei10, Lio24, LZT13, LHPF18, LPvSP12, ML12a, Mei13a, Mei13b, MLBA13, Mul12, NO14a, PSS^{+}13, PL14a, RRC^{+}11, SCÁBM11, Sch21, Sch23, SZ16, SSKS20, SC13a, SRAL12, TSM13a, Van16b, WL12a, ZG10a, ZSCL16a, ZZR13]. J.Theor [YCH^{+}17a]. JAK [RG12]. JAK-STAT [RG12]. JAK2 [ZFWK17]. January [Ano10-28, Ano10-35, Ano11-35, Ano11-48, Ano12z, Ano12-43, Ano13-29, Ano13-41, Ano14-42, Ano14-31, Ano15-43, Ano15-44, Ano16-29, Ano16-45, Ano17-35, Ano17-44, Ano18-30, Ano18-44, Ano19-42, Ano19-44]. Japan [HYN19, HN19, SFT14, TI12a]. Japanese [ZLCH18]. jasmonate [BB11a]. javanicus [TF15]. jellyfish [HM11a, HM15]. jeopardize [EVRE14]. jet [DB18, HM11a, LG13]. jigsaw [SSP15]. Jinha [JAHKH12]. JNK [KHH10]. job [SMM^{+}13]. Johne [LSSG10]. Joining [RM17c, LC11]. Joint [BPC15, MKBE17a, MKBE17b, NP11, NPS18, SQZZ17, YK11, BRP^{+}18, Edu16, ID19, KS16, MSZW19, MFMS^{+}18, MPT19, SCF^{+}12, VSLVB15, ZLZ^{+}15b]. jointly [GLZ^{+}13]. joker [ACCR11]. Journal [DHK13, EJ19, GHS20, JS18, LPB18, PCT19, Pan19, PP19, RTEKG19b, RTEKG19a, SYY17a, WLW20]. journey [NGND12, TK19]. JTB [Kir11, TK19]. jubatus [LLLV11]. judging [PBRW11]. July [Ano10-30, Ano10-48, Ano11-40, Ano11-33, Ano12-45, Ano12-27, Ano13-28, Ano13y, Ano14-27, Ano14-34, Ano15-30, Ano15-27, Ano16-30, Ano16-27, Ano17z, Ano17-27, Ano18-32, Ano18-27, Ano19-33, Ano19-47]. Jumping [SMHB10, CCW16, Cle18a, Cle18b, RP18]. junction [For10, JSK16, MAA^{+}17]. junctions [KDMD13, KPD18, MSSM18, TS17a]. June [Ano10-47, Ano10-31, Ano11-32, Ano11-37, Ano12-47, Ano12-32, Ano13-42, Ano13-46, Ano14-29, Ano14-46, Ano15-36, Ano16-36, Ano16-34, Ano17-46, Ano17-43, Ano18-41, Ano18-38, Ano19-31, Ano19-32]. jurisdictional [HF17]. just [Di 12, JV14, WGO^{+}15]. juvenile [Gar14, RVA^{+}11]. juveniles [Gar10].

K-gap [JCG15]. k-spaced [JCG16]. K-tuple [LFW^{+}15]. Kala [MCCC^{+}10]. Kala-azar [MCCC^{+}10]. Kalman [SSS15]. Kanazawa [BD10b]. Kauffman [WMT10]. Kawasaki [YLH12]. KEGG [HgLL^{+}10]. keratinocyte [NCLB16, PSG^{+}17]. Kerkhoff [Bal13]. Kernel [KOF^{+}14, Kon17, Mei12a, Mei13a, XWW^{+}14]. kernel-based [Kon17]. kernelized [LBS^{+}14]. kernels [FPD15, LSDD13]. Key [LL13a, MPZK16, AFS19, ALM^{+}19, CRGS18, CMN^{+}11, EBSW17, GLR^{+}18, GRB^{+}13, HMMSRSD15, Höh15, LPP^{+}19, LP19, Luo14, MSIR10, RPGG^{+}19, SSRA16]. key-role [Höh15]. kidney [CRV16, HHA17, PNP^{+}16, WLL^{+}14, ZCS^{+}15]. kill [HNP18]. Killer [WF18, FM15a, SR14]. Killers [CT18a]. killing [NC15, Unc11, ZMS17]. Kimura [RLFS18]. Kin [Gar11, Soz13, LvBJ16, Sim14, Uit11]. kinase [ADS^{+}19, DQY^{+}15, HKH18, SG15a]. kinase-1 [SG15a]. kinds [Di 18, KF15]. kinematic [GP11a, KHH10]. kinematics [RP18, SW13a]. kinematics-energetics [SW13a]. kinesin [GWX17, HHF11, SSZR17]. kinesin-1 [SSZR17]. Kinetic [HPML18, JLQ^{+}12, LDB^{+}14, Mic13, PRM11, RA17, BHBH11, FTPN10, KKO^{+}18, KIH16, Lei09, Lei10, MC19, MRS14, SSJK18, VACGF17, YIS^{+}17]. kinetic-order [YIS^{+}17]. Kinetics [SAGC12, TSP15, AdSG14, BB13, Bur19, DGW^{+}18, Das18, Dim10, DFMR19, DBJ12, ESS19, FT12, HH13, JSWY19, KMN^{+}15, MBM11, MTE15, MCL^{+}11, Nij11, OBK^{+}11, PAK11, SGGHS19, Sar10, ST17b, TF18, UAL10, VBV10, Van17b, VBD10]. kinetics-on-demand [SGGHS19]. kinetochore [SK10]. king [WHHS15]. kingdom [STG17]. kinship [DBD17]. KISS [dSdS13]. kleptoparasitism [BBR12, BCFR10]. Kmer [LFW^{+}15]. knifefish [ISZ18]. KNN [NMZ19]. knock [MdSPBL16]. knock-outs [MdSPBL16]. knockout [MJC^{+}11]. know [ZSSM14]. Knowledge [BYY^{+}11, MV10, PM14, Lar16, LWLM18, Mas18, Mei12b, Mei13b, SNS17]. Knowledge-based [BYY^{+}11, MV10, Mas18]. known [KMA10, PMW12]. Kohonen [GMK18]. koi [OA11]. Kolkata [SSG^{+}19]. Kolokotrones [Mac11]. Kong [HHE19, ZLCH18]. Korea [AHJ^{+}18a, EN15, KLJ17, KJ19, LJ15, LCJ16, LCJ18, SKPK17, WCJ11]. KRAS [SBM^{+}16]. Kullback [LZ18]. Kuramoto [MGB17].

L [Edg19, KP15]. L-ascorbic [Edg19]. L-shaped [KP15]. L. [JRR^{+}12, MFM16, RMF^{+}18, RZIR16]. L1 [DSPM14]. L1-logreg [DSPM14]. label [Mei12b, Mei13b, WMK14, WMK16, WLY^{+}17, XWC11]. labelling [KCD11a]. labor [CL13b, HHG14, NF14b, NSKO18, Ran12, RC13, RSBB10]. laboratory [SYY17a, SYY17b]. labyrinth [BA19, OBE^{+}17]. labyrinths [VPFV13]. lack [Di 16]. lactamase [KSKK15, SKK18]. lactate [AHW13, MKMG^{+}14]. lactating [CFR^{+}14, CMR^{+}18]. Lactobacillus [LDdA^{+}13, RZIR16]. lactose [Mic13]. lacunarity [PHTP^{+}12]. laevis [SSKK17]. Lag [Ber14, Chu17, DB18, LG13, WYL^{+}18]. lag-phase [Chu17]. Lake [BCF16b]. lakes [SM17b]. Lambda [Voh17]. lamblia [GDPPSS^{+}11]. Lamellipodia [SW15]. Lamellipodium [MOSS15, Kun16]. Landau [MROS15]. landing [MS14b]. landing-take [MS14b]. Landmark [LKSM14]. Landmark-free [LKSM14]. landscape [AH12b, AP19, Alv18, CVPCV^{+}15, CA17, KS14, LWLM18, PNK16]. landscapes [CBBO16, CGB13, FSW^{+}16, FLA^{+}19, FRG^{+}13, GGBW14, GWBG14, HMWB13, HH16, HHD^{+}16, LCG^{+}15, LLH^{+}13, LS10c, LAJC19, NSK13, NP15, PTNKT11, XW17]. Langevin [APW10]. Langstroth [STKE12]. language [CB17a, KL10, Nak19, SFS13]. languages [PMR11]. lanuginosus [KGP^{+}15]. Laplace [AEM14, AEM13]. Large [CL17b, CDC18, Ing16, SN18, tBdR18, BJB18, Fue16, FOT^{+}15, HXL16, HMM17, ITO16, JJ17, Mei14, MAA^{+}17, MP13c, MH14, Ots11, RA10, Ram10, TZ11, WZ17]. Large-amplitude [tBdR18]. Large-scale [CL17b, CDC18, Ing16, HXL16, ITO16, Mei14, MP13c, Ots11, RA10, WZ17]. larger [CW11a, RW14b]. larger-bodied [RW14b]. larvae [BPG^{+}18, SPH12, YSI17, YSI18]. larval [YHI14]. Lasso [KHX^{+}19, WMK15]. LASSO-based [WMK15]. last [BTK14, CHK16, KMH19, NdMLLBLC13]. latency [ARG14, ARG19, VGZS18, WCC13]. latent [BL14, CK16, MM12, TC12, WR14]. Lateral [KDG13, Lee16a, XS13, BGCB12, KKS^{+}14, SSS13, SLHS13, TSB^{+}17, ZBA18, FH10b]. Lateral-directional [FH10b]. lateralization [MPL16]. laterally [MDS16]. latitudes [Kur14]. Lattice [MAR^{+}17c, BMB^{+}18b, EI13, FNH10, HM11b, Kar12, MPBS17, MFG14, PM10b, ST17a, Smi14, WGH^{+}14, YLW^{+}14]. lattice-Boltzmann [WGH^{+}14]. lattice-free [BMB^{+}18b, MPBS17]. lattice-spring [WGH^{+}14]. lattice-structured [YLW^{+}14]. Launching [Axe12]. Law [DPCM16, CdPD11, GVCG12, HB10, JKF19, KMN^{+}15, KHRS10, KK15, LCBO^{+}12, PCT11, PCT19, Pan19, PR17, PWHW16, RBMPP^{+}15, TLCZ12, ZL18b, KMM16]. law-based [RBMPP^{+}15]. laws [GMMGVD12, LLTP19, MC19]. LAX3 [MPP^{+}15]. layer [AABS16, AC15, CWB^{+}17, DD10, DK13b, FK13, HGB^{+}13, LCA^{+}15, LFD11, OFB10, PWHW16, RRB10, VSW10, WMK16]. layer-specific [LCA^{+}15]. Layered [PP14, ZHC10]. layers [EPJ^{+}09, EPJ^{+}11, PDB^{+}15]. laying [RRBM12, ZVMB10]. LCRs [VBHM^{+}13]. LDL [LFD11]. leaching [RAAS15]. lead [ANY14, BEK10, PGPD19, Sal15]. Leader [CDD12]. Leader-based [CDD12]. leaderless [CB16, SGD^{+}16]. leaderless-dependent [SGD^{+}16]. leaders [KW15]. leadership [HKB10]. leading [MFN^{+}18, NBW10, NBW11, RRC^{+}11]. leads [CCA17, Eft13, FPD15, ITO16, LCTG15, SK16a, STN^{+}19b, ZLY^{+}17]. Leaf [RHS14b, RHS14a, CdPD11, HTM16, KCMF11, LFM14, TSdL14, WCF12]. leaf-cutter [KCMF11]. leaking [SJK18]. Leaky [MGGM10a, JSK16, KDMD13]. learned [LDA13]. learners [CZN12, KMMS17, Mar19]. Learning [BSC^{+}18, FLCS^{+}15, AH16, AMFL10, BBP13, DHT19, EHSR17, EL12, Fer11, GJ10, IA17, Kea16, KMMS17, MN11, Mei12a, Mei12b, Mei13a, Mei13b, Mei14, MHKA16, Mor19, NS19, SMG18, SLT^{+}18, TM12, TBA14, XWC11, ZPHS11, ZNCM15]. leave [IIVR10]. leaved [STI13]. leaves [KHB^{+}18, LKSM14, LL17b, Mit16, TDT16]. Leaving [GT12, Jia16, PMW12]. lectins [OSV^{+}16]. left [DPCM16, Gra18, LE19, She11b, SWSLMJ19]. left-hand [LE19]. left-handed [She11b]. left-right [SWSLMJ19]. leg [BAM^{+}15, BBJDS11, SS15b, SS15c, SS18]. legacy [O’M17]. Legal [RFME^{+}12]. legged [MS14b]. legs [ARB13]. Leibler [LZ18]. Leishmania [DJD10, SMJS14, WWIG19]. leishmaniasis [SYSF17]. Lempel [YHY14, ZLY14]. lend [CT14b]. Length [SH16, AJC12, ABD^{+}11, Don13, EBX17, GSV11, IST11, Kee10, LSC^{+}18, MZWC10, MR10, Pon12, Por13, RG18, SCLC13, TSB10, THBM10, WC10]. length-biomass [MR10]. length-dependent [GSV11, THBM10]. lengths [CS15a, JB18a, POLT12, SS12a, TSB10, YW13a]. lens [SSLB15]. lentiviruses [VCF^{+}19]. leptin [CVPCV^{+}15, KSU^{+}18]. lesion [KDMD13]. less [HII^{+}19, MG17]. Lessons [TXY^{+}12, Tjø10, AK15, CEdLSG^{+}16, PGGvL^{+}19]. lethal [Bon10]. lethality [TMM10]. Letter [ARD12, JN14, JAHKH12, Kir11, gSxFH^{+}12]. leukaemia [CMS16, SBM^{+}19]. leukemia [KEG17, NSH^{+}10, ZCT18]. leukocyte [GMMN18]. level [AHJ18b, ANY14, BB13, GBRS19, HBW^{+}11, JCG15, KB10a, Kon12, MN11, NBBCC18, NWZ15, NBW10, NBW11, PDB^{+}15, PNP^{+}16, PDNP16, RRC^{+}11, SC10a, SL10, SCC11, SFD12, SP16, SBSE14, WSTL16, XGZ17, YG18, vLSBBD11]. levels [FGLS10, Jam14, Jam15, TM12, YST14]. Levy [EPP19, BC15, CHL^{+}19, Han12, HS14c, Rey13, RSC14, RPCW18, TA15]. LGT [SMC17]. L’Homme [LCCB11]. Lie [SFSJ12, Sum13]. Liebig [GVCG12]. Life [BWB11, BDE^{+}14, CBC17, DBB14a, SSD12, Tox19, Bau18, BZA16, BI10, BBDB13, BHLH12, CLG^{+}10, CLG^{+}11, Chu10, Coo16, Dai12, Di 19b, Di 19c, Dol16, DMSW10, Fil10, GS13a, GS14, Gri15a, Han10, HIT18, JMZ13, KMH19, LCCB11, LG12, MCN10, MP13a, Mau15, MLT10, NNH^{+}16, Nik15, NNK10, OT13, Pia19, PBD11, RBRB15, RVP19, RRR15, Sat10, SMG18, Tox17, TRJD19, VW10, WSK^{+}19, XSMF12, YYST13, tBdR18]. life-course [BBDB13]. life-cycle [LG12, MLT10, XSMF12]. Life-history [BDE^{+}14, PBD11, SMG18, TRJD19]. life-Symmetry [RVP19]. Lifelong [BCPL10, MS13]. lifespan [KTI18, KS15, Nei14, RRR15, SKK^{+}12]. lifespan-shortening [SKK^{+}12]. lifespans [HZW^{+}14, KBV12]. lifetime [PAA11]. lifetimes [PC13]. Lift [JLH^{+}15b]. ligament [SGAM12]. ligand [ÅCHS19, CML11, FCS18, HWW^{+}14, KG12, KC11b, QQJW13]. ligand-binding [ÅCHS19]. ligands [AH12b, GDPPSS^{+}11, WB15, YNY^{+}10]. ligase [PRH^{+}11]. Light [Jen10, ÅCHS19, CFL^{+}15, HWGT15b, KSP12, LRL^{+}16, LL17b, MZ17, MSC10, MCL^{+}11, OFT15, PKL12, SM17b, dBJ11]. Light-dependent [Jen10]. light-induced [LL17b]. light/dark [LRL^{+}16]. lignin [BMN16]. like [AES^{+}13, AOT^{+}11, ABK^{+}12, CHH10, DQY^{+}15, DZE11, FPG11, KFG15, KS14, LC17, LFM11, Mas14, RGP13, SSKK17, VPGDP^{+}14, vNÁBG12]. like-minded [Mas14]. likelihood [BFLS17, HC17, LSC^{+}18, MKF^{+}14, SENS^{+}17, SC10c, UAL10, WZ15]. likelihoods [HS11]. likely [Ber14, CHL^{+}19]. limb [CSB15, GBN14, Pon12]. limbs [GP11a]. limit [GF11, MGCL13, Uit11]. Limitations [GSCS11, BH12, FE11, GMMGVD12, LHD^{+}17, MP16]. Limited [Chu17, CNDK15, HKM12, IH17, BBD18, HWGW18, Jab10, KFA12, KST17, Lar16, LCCC10, MS14c, SSvdM10, SI19, SYSM11, TGB^{+}18, YSYI13]. Limiter [SSD13, FH13b]. Limiting [AGM11, LSDD13, MPS16, SY17, DK13a, RC11, Sch14]. limits [BL15a, Chu18, HM11a, Par19, Sch18, Sch21, TDW16, Zab11]. line [OOY16, cSGFB17, Voe18, ZBA18]. line-associated [Voe18]. lineage [bCRS17, CNL14, Di 14b, LWH^{+}11, SSKS18, SSKS20, WML16]. lineages [RJSC18, SS15d, YAK17, YLLL12]. Linear [CCF^{+}14, YW13a, ACdlRMR^{+}11, BCF16b, COG19, CZW^{+}11, Das18, JAHKH12, JRG14a, JRG14b, JRG14c, KMD^{+}12, KCJ^{+}11, LCSH14, LT14, MI19, MPT19, NSKO18, Pav14, PHPM18, SRD^{+}15, SKF16, SG15b, TGB^{+}18, Van17b, VLFF12, WRH^{+}16]. linearized [LZWK11]. Linepithema [vTMW15]. lines [DBBW09, DBBW11, DKP^{+}18, FLM18, TKK14, TW13]. linguistic [MPS^{+}11]. Link [DPCM16, EB15, BC13, CNL14, FKK14, RBRB15, RBRB16, TK10c]. linkage [Gar14, Has14, ITO16, JAM18, RRÚJ19, SKH17]. linked [AKR^{+}18, AGM11, BUC14, BTK14, FKK14, GGM12, Sek12]. linker [HHF11]. Linking [CML11, Mor16, PMD^{+}18, GHS18, GHS20, HS14a, LHD^{+}17, PVB^{+}10]. links [CZJQ14, WSP14]. lion [LLLV11]. lipid [BMSEE14, FBM19, She11a, ZWvG10]. lipid-structured [FBM19]. lipids [MK18b]. lipograms [LCF17]. lipoic [KHH^{+}17]. lipopeptide [KSK^{+}17]. lipophilicity [Mar12]. lipoprotein [JSK16, VSW10]. Liquid [LLK14, AdSG14, WBB^{+}18]. Liquid-intake [LLK14]. liquoring [SSST19]. Lispro [SAI14]. listeners [SN16]. Lite [LDWW14]. Lite-based [LDWW14]. Literal [TWP16]. literature [PBSM19]. litura [KSK^{+}17]. live [ZM10]. liver [BDTR15, BRN15, BHO^{+}18, CLHB11, CMR^{+}18, LHW14, LLB^{+}18, OBAI11, SSR^{+}19, ZXS^{+}19]. livers [OABI12]. livestock [Cha18]. living [BL15a, BCPM^{+}16, GKWG17, NO18, SA17]. load [MSP19, OMO13, VMZK^{+}19, WML^{+}17]. loaded [She10, SGB12]. loading [JBSFB12, KTT^{+}19, LCA^{+}15, MYK^{+}11, Mil16, RS12, TKTB18, WYL^{+}18]. loads [DPRSS11]. lobule [DLL^{+}18]. Loc [WMK14]. Local [AT10, CJKR10, MKJS13, NT14, OO17a, TGL15, ATRR10, BBM^{+}97, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, BJPP12, CBC^{+}18, CT18b, DL15a, FGJ10, GWBL12, Gal11a, GWW^{+}10, HB16a, IU18, LTP19, LPH18, LPH19, MH12, MGCL13, MMAS13, MDEKH13, PGLZ14, Soz13, Str11, TKK14, UI10, WSH^{+}10, WGO^{+}15, XLL^{+}10, YK11, YZFY16, dBJ11]. localised [ABR11]. localization [ABD^{+}11, CXC18, DHS^{+}15, DTY12, FKB13, iHIM12, HLS^{+}12, LLW^{+}18, LZG^{+}19, Mei12a, Mei12b, Mei13a, Mei13b, SRD^{+}15, SC10b, STG19, Sht17, SBJ^{+}18, TKK14, USF^{+}18, WMK14, WMK15, XWC11, ZL18a, ZD18]. localizations [AAJGCD15, YLL^{+}14]. localized [MYC12]. Locally [HH16, HHD^{+}16]. location [BSB^{+}13, BCPG18, GC16, LWW14, LTL^{+}15, WMK13]. locations [FL12, HY13, JD17, QLC^{+}18, SSSD17, SOIO10, ZMS11, ZLW^{+}19]. Loci [HG18b, LZG^{+}19, Mog15, YI16]. locked [Opr10]. Locking [XWW^{+}19, LV17, WSLC14]. Locomotion [Ios16, HFT15, IUK14, MS14b, OYIM10, SS15b, SS15c, WL13]. locomotor [Pon12]. locus [HT13, Wax11a, XW17, PR17]. lodging [BSB14]. Log [Gla13]. Log-transformation [Gla13]. logarithmic [Bal13]. logging [LSDI15]. logic [PPM12, SCS14]. logical [AJM19, TK19]. logistic [Fra16, GAGP^{+}14, GW13, JMS18, KKD18, LLL13, Sch18, Sch21]. logistics [KBKK14]. logreg [DSPM14]. loners [GT12]. Long [BS19b, BD12, GA14, MK18a, MBP16, MI11b, OYIM10, AGGME18, ABLH18, BHLH12, BHSB11, BvLH14, FLM18, FHM18, HKS15, HKS16, JP19, MPP^{+}16a, MS12a, MSL^{+}16, NS11, RRR15, SKSO12, SBB14, VZB^{+}15, WGS15]. Long-distance [BS19b, BHSB11, BvLH14]. long-range [AGGME18, JP19]. Long-tail [OYIM10]. Long-term [GA14, MBP16, MI11b, ABLH18, MPP^{+}16a, MSL^{+}16, SBB14, VZB^{+}15, WGS15]. longevity [CHK16, JB18b, KMCH14]. Longitudinal [JB18a, NNH^{+}16, FH10a]. look [AA12, Kři11, RVP19, Ros15, Zhe16]. look-and-run [AA12]. Looking [Bro13, NGND12, SLL18]. Loop [DB11, ACT12, AGS18, DD13, GI15, JB19b, KHS13, KGM15b, LL13b, MK18a, MB17, RLM^{+}14]. Loop5 [HNL^{+}11]. loops [AN13, iHIM12, MB14, RA18, Sel14a, UT17, ZCSRG12]. loser [SYI17]. losers [MG17]. loss [BV13, LS13a, LKO^{+}17, Mal10, RT19, SRH19, SPCM18, WGS10, ZPM^{+}15]. losses [bCRS17]. Lost [BLS17a]. Lotka [DG11, DZJR10, Mil11, THM10]. lottery [AR17]. Love [AW12]. Low [JSK16, LHD^{+}17, VBHM^{+}13, BTW11, DGJ15, DPCM16, GGBW14, GMS^{+}13, GAY^{+}11, MGO^{+}15, MAA18, OTTF11, PR13, SV14, SSV^{+}15, UBP12, VSW10, ZZK14]. low-complexity [MGO^{+}15]. Low-density [JSK16]. Low-dose [LHD^{+}17, GMS^{+}13]. low-similarity [ZZK14]. lower [CSB15, Lev14]. lower-bounds [Lev14]. lowers [HBSF11, IH17]. LRRK2 [ADS^{+}19]. LT2 [SPSM15]. LTB4 [PDRC16]. LTD [HKS16]. LTP [HKS16, SBB18]. LTR [CLPZ14, WML^{+}17]. lubrication [dLMV^{+}10]. LUCA [CBC17]. lumbricoides [CH18]. lumen [GTB10]. luminal [DvDBD15]. luminescence [FBK11]. lumped [BSMK11]. Lumping [IMM16]. lung [BJWL15, BSB^{+}13, CPF^{+}13, CvBF18, Don17, GL18b, GY19, HWMT17, HHMT19, HGB^{+}13, LFB^{+}16, LTZ17, LTZ19, MMFK10, RS14b, SCS15, SMA11, SCKL15, TDZ^{+}18, TLPH11]. lunge [PGS10]. lungs [Don17, KBV^{+}15, SCT12]. Lupus [LZL^{+}19]. Luria [Zhe15]. LVQ [YC13]. LXA4 [PDRC16]. Lyme [WDL^{+}13]. lymph [CZPC^{+}18, GMM^{+}13, MMFK10, MMLK11]. lymphangiogenesis [BPS15]. lymphatic [WFM^{+}13, WFC^{+}14]. lymphoblastic [NSH^{+}10]. lymphocyte [DP13a, LG12, PDNP16]. lymphocyte-mediated [LG12]. lymphocytes [ABD^{+}15, WMHH10]. lymphoguanylin [PPC^{+}17]. Lynn [CB17b, LP17, Sat17]. Lys [JLX^{+}16]. lysine [CCY^{+}19, JLX^{+}16, JCG15, JCG16, SQS^{+}12, XDD^{+}15, XWD^{+}10]. lysis [BL14]. LysRS [NTC^{+}11]. LZ [DYS^{+}13]. LZ-word [DYS^{+}13]. LZCC [Zho11].

M [Pan19]. M. [PCC15, SCC11]. M1 [dBE16]. M2 [dBE16]. macaque [HII^{+}19]. macaques [WVO^{+}15]. Machine [AH16, FLCS^{+}15, GYWJ10, IA17, JW18, KHX^{+}19, KSKK15, LYF^{+}15, MHKA16, MBE11, Nic19, NHSX14, PWG16, SMG18, SLT^{+}18, XWD^{+}10, pZWZ^{+}16, ZZC^{+}17, LCCB11]. machine-learning [SLT^{+}18]. machinery [GHS18, GHS20]. machines [LRA^{+}13, ZDG^{+}10]. MacKay [DSF11]. Macro [KPD18]. macroalgae [BI10]. macrocysts [SSS17]. macroecological [TKKE19]. macromolecular [DS16, MPCTGJ^{+}15, ØW16, Rou14, Vaz10]. macromolecules [SKGM19]. macroparasite [TWS15]. macrophage [CHD^{+}10, FKB^{+}12, FBM19, KPEK14]. macrophage-facilitated [KPEK14]. Macrophages [EH17, AHMA^{+}19, BB19, HHS^{+}10, NJP18, PL14a, PL14b, dBE16]. macrophytes [STI13]. macroscopic [GS10, GHS18, GHS20, KDMD13, METC12]. Madin [HHA17]. Magnesium [KBK16, CZT^{+}16]. magnetic [FR13b, MPL16, Tay16a]. Magnetohydrodynamic [ZG19]. Magnitude [PCC15, Oko15]. Magnus [WRWX19]. Mahalanobis [QZY17]. Main [DDF^{+}14]. Maini [Ano19-50, Bur19]. mainland [XTZ^{+}13, XSTW14, ZJY19]. maintain [Ezo19, FXML18, LCHMP16, XKCG15]. maintainable [HK16]. maintained [Bon12, DEG^{+}14]. maintaining [LHW14, Miy16]. maintains [Cza14]. Maintenance [Lar10, XDC^{+}11, DBD15, GWE11, NSH^{+}10, PCBMM12, PRM14, SBB18, SLvdBMP10, TR17, VSLVB15]. major [BKL^{+}15, DJD10, Jam15, MMFK10, SMJS14, SZ10]. Majority [MS14d]. make [HNP18, MG17, Pan11]. Making [BFW10, LLSO13, AOT^{+}11, LTSTD15, LNL^{+}11, RLK10, SYI17, Sue12]. malaria [ADB^{+}13, BAR14, CSS^{+}18, DROC11, EN15, GM19, GWCA14, KMCJ17, LP19, MS16, NKL10, NT14, NDZMA14, NMAZP16, NTED^{+}19, OL19, PNVN18, PEZ^{+}13, PRM12, PRM14, TEY16]. malariae [CMD^{+}10]. malarial [EN15]. Male [FKMG15, TEW12, BWB11, DEG^{+}14, HIT18, MG17, Nak16b, RQB17, Unc11, YSYI13, YYB^{+}19, YY18, ZLT^{+}19, ZXS^{+}19]. male-caused [YY18]. male-killing [Unc11]. male-male [Nak16b]. male-sterility [YYB^{+}19]. males [Gar14, Gra14, TT10b, YSYI13, YYST13]. malignancies [MPC12]. malignancy [SGGY10]. malignant [RTRRS^{+}17, SHvHB16]. Mallotus [EBS11]. mammal [Sil16]. Mammalia [TF15]. mammalian [ASK17, AA16, BZN17, Bec14, CLHB11, CKBCG12, DTGD19, GGG19, Jam13, LF16, LW12, VB19a, Xie11]. mammals [FPdRdH10, NNK10, PC17, PEZ^{+}13, SPO^{+}10, VAT18]. mammary [BCS^{+}16, CFR^{+}14]. manage [YTGW16]. management [ACH18, ABR19, GL14, HYSG19, HF17, IL13, KKUM10, MM19, MA13, NNvdB17, NBA^{+}18, PGF11, RDMP11, TB16, TTC10, TSL12, VCF^{+}19]. Managing [KPG19, MGB17]. manganese [KBK16]. mange [BCF19]. manifold [Bro13, BBP13, TMF17]. manipulation [AK13, Fer11, HXL18b, SK15]. manipulations [KMM18]. Many [BWM^{+}19, Bur19, CBBO16, CD16, Ezo12, GT11b, Hay16, HG18b, LCHMP16]. many-player [GT11b]. map [Gau11, HGM15, IMA14, MGM13a, NvD17, OABB18, TWP16, WDL^{+}13]. map-based [MGM13a]. MAP1B [KK17b]. MAPK [KFL12, ZG10c].

Mapping [BJJR10, DRC11, FK18, FWLW11, KSM^{+}15, Mog15, SQZZ17]. maps [Bec14, FH13b, GMK18, RNVP10]. MAPT [HRHAAA15]. March [Ano10z, Ano10-33, Ano11-41, Ano11-45, Ano12y, Ano12-37, Ano13-33, Ano13z, Ano14-45, Ano14-28, Ano15-31, Ano15y, Ano16-28, Ano16-41, Ano17-36, Ano17-29, Ano18-29, Ano18-48, Ano19-40, Ano19-35]. margin [IS17]. marginal [Di 17a]. marginalis [Whi11]. marginally [Bro13]. margins [HBT13]. Margulis [CB17b, LP17, Sat17]. Marie [AKR^{+}18]. Marie-tooth [AKR^{+}18]. marine [KG13, MP13a, MROS15, NNK10, TM14, Tak16, VAT18, YSYI13]. marital [MMC19b, ZSW^{+}14]. marker [MKBE17a, MKBE17b, MKRE18, TF18]. markers [DQS^{+}15, FMS^{+}10, Mor11, PHTP^{+}12, Tal12, WCC14]. market [Ant13]. markets [Bar11, MLMG^{+}15, Ran12]. Markov [BCFR10, CJKR10, CPGF^{+}16, CSM^{+}14, FLGGD^{+}14, HM11b, Jia16, LYF^{+}15, MC15b, RFME^{+}12, RLFS18, SG15b, SFSJ12, Sum13, TSS16, UBP12, ZLW16]. Markovian [LZT13, ACT12, BBR16, LZT12, PHK15, Veg10]. marriage [Arc13]. marrow [AOM19]. Marshall [vV11b]. Martingales [Mon18]. Maspin [AMFR^{+}16, AMBH^{+}13]. Mass [FT12, Mac11, CDD12, Cro18, DBD17, EN15, HPP10, Höh15, Hor11a, LLG16, LS18, MKBE17a, MK14c, PIPB10, POP12, SSME10, WSRG18, Woh15]. mass-extinction [Höh15]. mass-recruiting [Cro18]. mast [TI13]. master [ZS10]. masticated [DDF^{+}14]. masting [AM10]. Mastomys [GRBL13]. masu [HIT18]. matching [AdVG15, ESS19, MBBD13, ZLZ^{+}15b, ZLZ^{+}16]. Mate [MM18, WSW10, BBB18, CC11b, HLI15, NT14, SLL18, TI12a, YK11]. mate-finding [BBB18]. material [CW11a, Fil15, GSF13]. Maternal [FPD15, FGH^{+}14, Fil15, Jam14, Jam15]. mates [SBJ13]. Mathematical [ASC15, AOM19, AP17, AM12, ADR^{+}11b, BCPM^{+}16, BB19, BHO^{+}18, BJ18, CWM11, CDO17, CSBK15, CWB^{+}17, CKB18, CSS^{+}18, CSKZ19, CSKF10, CHN^{+}15, CDGV10, DGNT17, DBM^{+}18, DPC15, DTGC14, FBC12, Fok12, FM13, FIJJ11, GT11a, GE18, GBJ^{+}11, GSS19, GMB11, GCYL18, HTM15, Hir16, HHE19, HBA17, IBB^{+}15, IvLP12, iITIM15, KCMF11, KMZR18, KSSM11, KLJ17, KdlRJ18, KPEK14, KSK^{+}16, LF19, LBMG16, LSMP14, LMHF13, MOEdP16, MRF19, MVH11, MFL^{+}18, MMK16, MSA^{+}16, MTLW^{+}10, NKB16a, NPMM15, NTOI16, PH13, Pai19, PRC15, PBSM19, PVGV19, PLD^{+}17, RMM^{+}16a, RZRSC19, RKMG12, RGL^{+}17, RBB^{+}12, SRK^{+}12, SBM^{+}16, SNY^{+}17, SDT17, SBW11, SSM^{+}18, Sht17, SD16, SMA11, SSR^{+}19, SRH19, SW13b, SPL14, TLL^{+}12, TRM^{+}14, UK17, VHM^{+}17, WVO^{+}15, WCM15, YCL^{+}17, ZARK19, ZLZ15a, ATK10, AHW13, ANGPB12, ABR19]. mathematical [ANK10, ABH^{+}11, AHMA^{+}19, ABV19, Ano19-27, AHAFA14, ALH16, ARM18, BRK16, BTR18, BCPL10, BPP^{+}16, BBJ^{+}10, BTH^{+}13, BSM^{+}14, BPS15, BTF19, BGD^{+}13, BSR^{+}11, BCF16b, Bue15, BMGC11a, BMGC11b, Bur19, CFS^{+}19, CVH^{+}10, CPW16, CPF^{+}13, CCG^{+}18, Che12, CWP^{+}18, CKN^{+}12, CTB18, CCAdS13, DG11, DL12a, DMP15, DLHS11, DHT16, DFG^{+}18b, EPJ^{+}09, EPJ^{+}11, FBFM12, FKDW15, FBU11, FWLW11, GKMC10, GLR^{+}18, GA13, Ger16, GTC19, HL11, HI19, HHSA15, HLVR16, HBA10, HGO^{+}18, HGM^{+}16, JMH^{+}19, JSDEK14, KSU^{+}18, KSP^{+}14, KBE^{+}13, KJ19, KHNM16, KFW12, KCS^{+}15, KS16, KK18, LPP^{+}19, LF17, LSP^{+}17, LZA^{+}19, LA12, LF16, Leh18, LHW14, LSY^{+}10, LX16, LZY^{+}15, LLB^{+}18, LDdA^{+}13, LXLF14, LNRK11, MZ17, MM12, MT14, MGGM10b, MYC12, MBBV14, MKMG^{+}14, MBC^{+}12b, MS12b, MSE^{+}14, MF19, Mor10b]. mathematical [Mor16, NS11, NH19, NS13, OFT15, OO11, PCT11, PCT19, Pan19, PDRC16, PMK15, PL14a, PL14b, PDF18, PCS^{+}18, PC10b, PMKS^{+}15, POK^{+}12, QJR^{+}16, RM17a, RVZ14, RTR^{+}13, RRBM12, RWW^{+}15, REC10, RAF^{+}14, RH19, RTEKG12, RTEKG19a, RSD^{+}13, RB15, RBG^{+}17, SGAM12, SBSR13, SBP13, SHK14, SPO^{+}10, SK12a, SYSF17, SY11, SZ10, SW19, SHSR15, SH15, Tac10, TWTA^{+}18, TDHC^{+}18, TEY10, TAL19, TT17, TDM14, TK10c, VKKA12, WXWL12, Wod18, WBB^{+}18, XYN15, YKO^{+}16, ZCS^{+}15, dSKBS10]. MATHT [SLW^{+}18a]. Mating [LPvSP11, LPvSP12, Ant13, BWB11, BM14, BAR14, CC11b, FH13a, GGM12, HT16, LG18, LF13, MO12, NH12, ZKTK10]. matrices [RLFS18, Sum13]. Matrix [AH16, Bue15, DLL^{+}17, ESGA15, GCM17a, HSMCCR19, HHF11, KPK^{+}13, KHH10, KHK^{+}17, KC17, PDB^{+}15, PNP^{+}16, PDNP16, PLN14, POP^{+}19, PSAA13, QLC^{+}18, SLDP13, WYX^{+}17, WZY14, ZSZM14]. matters [Das12, Gau11, PCBMM12]. Maturation [BT17, BI12, LZTD18, PBD11, WSRG18]. mature [BLZ^{+}19, DKR16]. maturity [ZZS19]. max [BBRT12, MKF^{+}14]. max-plus [BBRT12]. Max-SAT [MKF^{+}14]. maxima [RG18]. Maximal [FS16b, HK16, MPP16b, CHL^{+}19, JD16b, LYF^{+}15, Mic15b, SSB^{+}15]. Maximization [BCBD19, Arc13, LPF11, OWB14]. maximize [SWR11]. Maximizing [ARZ15, KOS13]. Maximum [BH13c, LSC^{+}18, MKF^{+}14, SENS^{+}17, TM15b, UAL10, BR14, BFLS17, Fue16, GL12a, GK13, HC17, KPK^{+}13, PMKM10, TP14, WZ15]. Maximum-likelihood [LSC^{+}18]. May [Ano10-27, Ano10-34, Ano11-47, Ano11-30, Ano12-46, Ano12-35, Ano13-30, Ano13-27, Ano14-30, Ano14-33, Ano15-39, Ano15-29, Ano16-39, Ano16-38, Ano17-41, Ano17-34, Ano18-28, Ano18-37, Ano19-46, Ano19-41, CR14, CW11a, DI13, Di 12, EVRE14, GL18b, GBM18, HNP18, MJI14, POP^{+}19, SS15b, SS15c, TGC^{+}17, VGZS18, WR14, YY18]. MCF [OOY16]. MCF-7 [OOY16]. MCF7 [DLM^{+}19]. McGehee [XF13]. MDCK [HHA17]. Mdm2 [Eli17, OAJK10, STX^{+}11]. meal [BSR14]. mean [FLGGD^{+}14, GSCS11, KTJ19, PDSP13, STLJ18, dlCGSA16]. mean-field [GSCS11, KTJ19, dlCGSA16]. meaning [LDB^{+}14, ZDF^{+}12]. meaningful [WMT10]. means [AGRR^{+}15, JM12, QF10, RRRRSPTR16, SM10a, WMD^{+}11]. Measles [SGA^{+}12, KN11, Lan16]. measure [DLYZ11, FHW^{+}10, HY16, Hua16, LXJ15, MK11, Mar17b, PBD13, YW13b, YCY14]. measured [PW18a]. measurement [Han17, LLB^{+}18]. measurements [ADMR16, BPP^{+}16, KBF18, METC12, SPMGR10]. Measures [Orr19, ALH16, DZW10, Est10, Gia13, HMM11, WSK^{+}19]. Measuring [DPRSS11, FSW^{+}16, XTZ^{+}13]. Mechanical [LH13, MS18, MGC12, WHWZ18, AA12, BSP18, BZL17, CPF^{+}13, Cle18b, Fue14, Fue15, GHS18, GHS20, HGGCR13, IS17, Js12, KSSM11, Kue16, LHPF18, MBPS17, MGC15, PDL^{+}17, RHJ11, SGAM12, SAB10, TSdL14, TIMI^{+}15, WBB^{+}18, XWW^{+}19, ZR16, Ran12]. Mechanics [vDVFH^{+}18, ABD^{+}15, ATG12, BK19, DCL18, FPG11, Ing16, KBL15, Kro10, MN12, RS14a, SHB^{+}17, SAGAGB17, ST17b, WO11]. mechanism [ADC19, BZJP18, BL15b, CSKZ19, CW14, DGJ15, Di 16, DZE11, FD11b, GWW^{+}18, HTM15, HMM16, HDZ^{+}19, IK15, JLQ^{+}12, Kee10, KFS^{+}13, KPS17, LCG12, MSN14, Mor10a, PJ13, Pat16, RG12, Sht17, SWSLMJ19, TGB^{+}18, TB11, TWH^{+}13, TSB10, TCYY^{+}12, WK18b, XF13, XKCG15, YG18, ZWY15, Zho11]. Mechanisms [HK14, JCW13, SYSM11, AGB^{+}14, CCB11b, CMS16, Deu18, DTP19, Eft13, GZFX14, GT11a, GTSP16, HS14a, JLH^{+}15b, KMA10, KEG17, Kon17, LKO^{+}17, MMLK11, MZMM^{+}14, NWWL18, PHÁBAI^{+}16, Pat19, RDD14, SBMH10, SZ15a, SZ16, SGD^{+}16, SH17, THC^{+}11, TSP15, WDL^{+}17, WBRHE10]. Mechanistic [DDF^{+}14, AR17, BSR14, BGF^{+}14, CKZ^{+}17, EHBC10, GCSMMG^{+}12, GG12, JG14, KB19, NWB^{+}10, OCN10, RG18, SL10, SAKG13, TW12, WJL^{+}16, YCB16, vLFM^{+}19]. mechano [MDD13]. mechano-chemo-thermo-electrotaxis [MDD13]. mechanobiological [CAA11]. mechanobiology [ZTK12]. mechanochemical [HGB^{+}13, MROS15, MHMM11, MAH12, SKH11]. mechanomolecular [SK10]. mechanosensation [KBL15]. mechanosensitivity [KLN^{+}12]. Media [YTGW16, CFT11, CFGRB17, LLD^{+}17, MBPS17, MBCM12]. median [VACGF17]. mediate [DHBS19]. mediated [ATI12, ABV19, AN10, AES^{+}13, AC12, BWB11, BIMC17, CB10a, DHV19, EBSW17, FMLM12, GS17b, Gri11, HWD^{+}14, KR19, LW12, LG12, Mal10, MGGM10b, Mic11, MK17, MH18, MA10, NF14a, NKB16a, Pie12, RC11, RK18, RH19, SDT17, STN19a, SN16, SK12a, TSB^{+}17, UAL10, VM12, WFW19, ZAL^{+}19]. mediates [Gar10, RG18, TK10a]. mediating [LVLC18]. mediators [CSDR15, GU13]. medical [HWGW18, RCH14]. Mediterranean [LG18]. medium [FKMP10, Gou16, MDVT10, RTFP^{+}17]. medulla [PNP^{+}16]. medulloblastoma [KSP^{+}14]. meets [Ros10a]. MEG [KA16]. meiosis [SM11b, WKS14]. meiotic [TR12]. Melanoblast [ALD^{+}11]. melanogaster [GAL11b]. melanoma [CCB11b, EBE10, EH17, HPM^{+}17, HLTR19]. melioidosis [SSRA16]. mellifera [EM11, MKJS13]. Melophorus [RSC14]. Mem [WMK16]. Mem-ADSVM [WMK16]. members [ACdlRMR^{+}11, NF14b]. membrane [AH15, AHJ18b, BC19, CRH^{+}15, CN12a, CKD11, CL13a, DK13b, DTY12, DFC^{+}12, GTS15, HYA14, HK11a, HK12, HMM16, KHK^{+}17, KHKI17, KPS17, LSMP14, LL17b, MJ11, MB14, Mei14, MSA^{+}16, PGT^{+}19, PWC^{+}15, RD12, RFdL15, SM17a, SM18, Sar10, SW15, SHB^{+}17, SS17b, WMK16, WXmH10]. membrane-associated [RFdL15]. membrane-penetrating [BC19]. membrane-substrate [Sar10]. membranes [KMZR18, MPK^{+}12]. membranous [GA13]. MemHyb [HK12]. memorizing [Tac10]. Memory [DHT19, Gau11, ME15, RLSN14, AJD11, DN18, HKS15, HKM12, KLHS17, Nak16c, OABB18, PAA11, RLK10, RA17, SI12, SXZY12, TK19, UK17, WMT15, YCH^{+}17a, YCH^{+}17b]. men [MFP^{+}14]. meningitis [EKvdKvFK13]. menopause [CL13b, MS13, TR17]. menstrual [HS14a, RSD^{+}13]. Menten [LJ18, OM10, Van17a]. mer [HA15a, Hua16, WZY14]. merged [Di 14b]. merging [Di 12]. meristem [MLL^{+}16]. meristematic [GRB^{+}13]. merit [Tak17b]. merit-based [Tak17b]. Merkel [KC13]. MERS [EAN14, AHJ^{+}18a, SKPK17]. MERS-CoV [SKPK17]. mesenchymal [CMHM19, TK10c]. mesendoderm [BKL14]. meshes [MKKS14]. Meshwork [SCsMzX18, SNY^{+}17]. mesic [MGCL13]. mesoderm [BKL14, JAK19]. Mesoscopic [CSM^{+}14, SHG16]. messenger [FZL18, JWS^{+}10]. Messor [PRMH14]. meta [LWH^{+}11, TD18, XSCS12]. meta-Boolean [LWH^{+}11]. meta-food [TD18]. meta-population [XSCS12]. Metabolic [OABI12, BRG^{+}12, CPW16, COG19, DSF11, HHG14, HB10, IvdSW16, KB15, KJSS10, KdSUS10, LCCB11, Mac11, MDB12a, MT14, MKMG^{+}14, MJC^{+}11, MMK16, MF15, MRS14, NKM^{+}12, OA15, ONO15, PYG^{+}19, PMK15, PCBMM12, RS10, RBB^{+}12, SHN12, SI19, SSD12, SK16b, WOB15, Wal16, WcW14, YIS^{+}17]. Metabolically [CKS15]. metabolism [AEE19, BRK16, BDTR15, BRN15, CCS15, CRGS18, DQY^{+}15, FTPN10, GLOC10, KSU^{+}18, KK18, LRL^{+}16, LPGTBMA19, LWvB^{+}19, OBAI11, OCHHZ12, PCC16, RPGG^{+}19, RWW^{+}15, Tox18, ZXS^{+}19]. metabolism-associated [ZXS^{+}19]. metabolizable [LT14]. metabotropic [MPS^{+}11]. metacommunities [AR11, BWL12, FGPR10, HD17, HGHL10, MSZW19, MAL^{+}11, RHT18, RAMS11, SCF^{+}18]. metagenome [DCS14]. metagraph [LT11]. metamorphosis [DPRSS11, iHM17]. metaphase [DL15b]. metaplasticity [FR14]. Metapopulation [LCJ18, NIiT18b, NiTI18, NIiT19, YiTS19, BGL10, CV19, GB12, Kis10, KJ11a, KKYV18, LLN^{+}19, LS10a, MP13c, NIiT18a, NdlPZA10, NG17, NGS^{+}16, NPS18, TL15, Tak16, TSS16, WHT17]. metapopulations [Ayd18, DGJ14, HBEW12, HvLPdB15, JH11, NP11, SSD13, TRG16]. metastases [HM10a]. metastasis [CLA^{+}16, LD18, RH19, SS17a]. metastatic [BH13c, CBGS18, HP16, WMN18]. methane [BGF^{+}14, vLFM^{+}19]. methanogen [WJL^{+}16]. methanogens [LWvB^{+}19]. methicillin [PDW10]. methicillin-resistant [PDW10]. methionine [VHM^{+}17]. method [BPM^{+}12, BTK15, CJ12b, CTPB10, DLYW13, DYQ^{+}14, lDsWM^{+}15, FHW^{+}10, FYZ15, GJ15b, lGLLL14, GGCJ16, HYA14, HLHY17, HYZ^{+}15, HZWH10, JZZK11, Kha10, KGB13, Kon12, KK17b, LZL^{+}17, LMJ^{+}18, MMG^{+}16, MAR^{+}17c, Opr10, Ou19, PAK11, PNP^{+}16, PRSC11, PWHW16, QQL10, QQJW13, QLWC19, RBMS17, RP18, SDRA^{+}15, SMS19, SQS^{+}12, SQZ^{+}16, TLY^{+}19, TAB^{+}13, TP17, VACGF17, WTQL10, WYL^{+}19, WZ17, Wu14, WGH^{+}14, XGZ17, YG18, YC13, YC15, YDC^{+}13, ZTWL12, ZKHL16, ZZC^{+}17, ZDY11, ZK10]. methodological [JPDP10]. Methods [GLR^{+}18, OBHS19, Boc12, DGW^{+}18, DEK15, Jab15, KOF^{+}14, KCD11a, LYZ^{+}18, MSM^{+}14, MNH^{+}12, MHKA16, MSS10, PDB^{+}15, PWC^{+}12, Ros10b, RCD16, SCF^{+}12, SBB14, TMD14, WH14, WP17, XWW^{+}14, dlCGSA16]. methyladenosine [AH18, ZZN^{+}19]. methylation [BSB^{+}13, JCG15, KKD18, LZA^{+}19, MPC12, PRN10, SQS^{+}12, TTC19, ZARK19]. methylcytosine [SIK^{+}18]. metric [GKNT10]. metrics [CA17]. metronomic [BH13c, HL18]. mevalonate [DGW^{+}18, PCS^{+}18]. mexicana [WWIG19]. Mexico [OL19]. MFSC [AH19]. MGMT [BSB^{+}13]. mGneg [CXC18]. MH [WXC10]. MIA-QSAR [MR18]. mica [Han10]. mice [AGC18, BTF19, LBW^{+}13, SNY^{+}13, TTB^{+}18]. Michaelis [LJ18, OM10, Van17a]. Micro [PMSY17, GWE11, MAM16, SK15]. micro-habitat [GWE11]. Micro-scale [PMSY17, MAM16]. micro/nanoparticles [SK15]. microalgae [AK15, BBS10, FGLS10, GCSMMG^{+}12, MMB18, NHS^{+}16, PBR17b, PP18, PP19]. microarray [GGCJ16, KHX^{+}19, OGE10, PS13, XSKA17]. microbe [GW19]. microbes [BID15]. Microbial [Aya12, SKS^{+}19, Ber14, BSL^{+}17, BWL12, CG10, DG12, DDP12, ESE15, FXML18, FE11, ISM^{+}11, Jus11, KFA12, MB18a, SCH^{+}19a, SI11b, SYY17a, SYY17b, SBS17, VBV10, WPPD16, WSTL16, XDC^{+}11]. microbicides [DBBM11]. microbiome [HI19]. microbiota [LPP^{+}19, MTLW^{+}10]. microchannels [SP13a]. microcircuits [RNN15]. microcirculation [LHW14]. microcrack [POLT12]. microCT [KSM^{+}15]. microdamage [SAB10]. microenvironment [CMHM19, CLA^{+}16, KL18]. microenvironmental [CL14, CL19]. microevolutionary [SBJ^{+}18]. microfibril [HB12a]. microhabitats [HSR12]. Micromechanical [FAFA17, WYL^{+}18, dBB14b]. microorganisms [GI10, KK15, NNvdB17, ZG19]. microridges [LWY^{+}16]. MicroRNA [LW12, ATI12, HW11, LF17, LFW^{+}15]. MicroRNA-mediated [LW12, ATI12]. microscale [PMK15, Str13]. microscopic [CSM^{+}14, DTP19, GAV15, GHS18, GHS20, METC12]. microscopy [KKO^{+}18, TKK14]. Microstructural [NF14a, PMD^{+}18]. Microstructurally [PMD^{+}18]. Microstructurally-based [PMD^{+}18]. microtopographies [Alv19]. Microtubule [WdVMD15, BB15, HCK^{+}11, MAFK12, SSZR17, WHH17]. microtubule-based [MAFK12]. microtubules [AAJ15, CN10, DEK15, DL15b, SA17, She10, SPL14, THBM10]. Microvascular [CR19, EBP15, PA15]. microvasculature [MP19]. microvessels [KCJ^{+}11]. mid [KMH19]. mid-life [KMH19]. Middle [EAN14, LCJ16]. Mig6 [ZZ18]. might [DMCP14, Har15, SGGHS19]. migrant [WW12a]. migrating [AA12, MSBB13]. migration [ASL^{+}18, Bon12, BMB^{+}18b, CML10, CCS18, CNDK15, Cox10, DMO^{+}17, GMOP12, GFM^{+}19, HBEW12, HPM^{+}17, HYQ^{+}16, KCS16, LFZN11, LZT12, LZT13, Lun13, MPBS17, MPS16, MFG14, MA10, MK16b, MDD13, NIiT19, NCLB16, NP15, NdlPZA10, PGG^{+}15, RPD14, Sch19c, SLJ^{+}10, SN17, TD18, WRH^{+}16, WW12a, WCW15, YiTS19, vLSBBD11]. Mildvan [TMF17]. milieu [LHW14]. milk [Sil16]. mimetic [GWE11, SFT14]. mimic [BTR18]. Mimicking [SKD^{+}10, SABB15]. mimicry [LBJ13, PS12, Ros15, YLY15, Yam16b]. mimotope [SQZ^{+}16]. mind [EM11, Nik15, PW19, GDPPSS^{+}11]. MIND-BEST [GDPPSS^{+}11]. minded [Mas14]. mineral [CKS15, DDS13]. mineralising [WPPD16]. Mineralization [MH13, VH11b]. Mineralization-driven [MH13]. mineralized [HB12a]. Minimal [Jus11, RGBR17, SHS13, CEdLSG^{+}16, Chu17, DS19b, JD16b, LPB17, LPB18, MPP^{+}18, SK10, SM10b, YZFY16]. minimalistic [ABKS11, IDI11]. minimaxity [MKE15]. minimisation [Bac15]. minimise [CPV16, CHN^{+}15]. minimization [BRP^{+}18]. minimize [FR17]. minimized [Mas16]. minimizing [KOS13]. Minimum [Sat14, SWTO15, FYZ15, HB12b, KPK^{+}13, QZY17]. Mining [HZL^{+}11, CDC18]. minnows [HCHI18]. minor [MKF^{+}14, NYSM12]. minority [JZ17]. miRNA [DHBS19, FD17, HW13, HZL^{+}13, SFMS16, SCKL15]. miRNA-miRNA [SCKL15]. miRNA-mRNA [HZL^{+}13, SFMS16]. miRNAs [HZL^{+}13, WSH^{+}10]. Mirror [LYK12]. miscarriage [Jam15]. misleading [HC17]. mismatch [BCKS13, SEK^{+}10]. missense [MPP^{+}16a, RMM^{+}16b]. missing [RBRB15, RBRB16, SPH12]. misuse [HS14b]. mitigated [BTW11, Kar16]. mitigating [KdlRJ18, TXY^{+}12]. mitigation [MPLK14]. Mitochondria [HK17, Har15, HBSF11, Sel12b, Sel15b, VHS17]. Mitochondrial [KDK14a, KDK14b, Rad16, BBR10, GA14, OTTF11, Sel11, Sel12a, Sel13, Sel14a, Sel15a, Sel16, Sel17, SB15]. mitochondrion [YLWZ10]. mitomiRs [VHS17]. mitosing [LP17, Mar17a, Sat17]. mitosis [SK10, SP14, WKS14]. mitral [LCA^{+}15]. mixability [LPF11]. Mixed [ŚKSRW16, ZWC^{+}13, ASGE14, BSV^{+}10, CMD^{+}10, KSP12, MO12, NWWL17, Pag19, PCC15, RRB10, SCC11, SNCM12, TBB14, UI10, WL19, ZDF^{+}12, ZLW18, ZK10]. mixes [MFP^{+}14]. Mixing [YSI17, FHSG15, MPLK14, Rad16, TLPH11]. mixotrophs [CG10]. mixture [PPT^{+}16, VZB^{+}15]. mixtures [QLWC19]. mLASSO [WMK15]. mLASSO-Hum [WMK15]. MM [ZZ18]. MM-PBSA [ZZ18]. MMP [SMS17]. MMP-TIMP [SMS17]. MMPBSA [KJM17]. MMTV [Mic11]. mobile [RW16, TA15]. mobility [BV12, IH17, MAW12, MP13c, PTC13, SK11b, XTZ^{+}13]. mobilization [SW13b]. modal [RSC14]. mode [BI12, GL18b, KHH^{+}17, LD11, MZAI19, RW15b, WSLC14]. mode-locking [WSLC14]. Model [ACH18, All11, BS15a, Das13, DP13b, FHS^{+}14, GCB17, KBV^{+}15, LKO^{+}17, LL14, LSLL13, MB18b, Mor10a, Mun12, NAK^{+}11, NF17, OA15, PVGA12, RSI11, RG12, SBR10, SGS15, SS15f, TT17, WBRHE10, ZKMB19, ACR^{+}17, ABB^{+}19, AAAV14, AV19, ACT12, AHD^{+}18, ABKS11, ANMH11, ATK10, AHW13, AJD11, AMBH^{+}13, AMFR^{+}16, ANGPB12, ABR19, ABR11, AA12, AWRD15, ABH^{+}11, ABV19, ABJ12, AdlPLMZ13, AP13a, AM12, ABN15, AR17, AHAFA14, AHJ18b, AP13b, AN17, AOT^{+}11, ABK^{+}12, ASF^{+}15, Bac15, BTR18, BSB14, BPS19, Bal13, BBR16, BKR14, BP18, BB15, BFJ^{+}18, BBJ^{+}10, BTH^{+}13, BTG^{+}15, BBR12, BF13, BFR14, BF15, BGKL17, BSV^{+}10, BJJR10, BBS18a, BBR10, BCF19, BKL^{+}15, BGF^{+}14, BHBH11, BDID^{+}12, BSM^{+}14, BB10, BPS15, BHH^{+}14, BTF19, BM10, BSR^{+}11, BSS11, BCF16b, BSB^{+}13, BLS^{+}12]. model [BR14, BMM^{+}14, BCPG18, BBRT12, BDR10, BJ17, BZ10, BC13, BMB^{+}18b, BPGS12a, BPGS12b, Bue15, BHSB11, BMGC11a, BMGC11b, BFGAGA16, BCS^{+}16, CHL^{+}19, CJKR10, CS11a, CCS15, CRGS18, CFCM13, CVH^{+}10, CPW16, CN10, CTA15, CDO17, CKZ^{+}17, CC11a, CCB11a, CTJ10, CGvG^{+}15, CMS^{+}19, CCG^{+}18, CC11b, CAGM^{+}17, CL17a, CFZ14, CN12a, CS11b, CK16, CHH10, CKD11, CJRR11, CW15b, CWP^{+}18, CKB18, CH16, Che14, CKN^{+}12, CK15, CJL15, CTB18, CFF11, CMHM19, Chu10, Chu17, Cle16, CA17, CB16, CG14, CV19, CZPC^{+}18, CGRRGM^{+}16, CAA11, CMJD11, CFR^{+}14, CMR^{+}18, Cro18, CMS16, CMGN17, CEKM^{+}19, CKS15, DFM^{+}14, DS15b, Dal17, DJD10, DDHM12, DS19a, DG11, DBBW09, DBBW11, DLSD15, DH16, DAVS11, DGNT17, DB19, DL12a, DAA17, Di 13b, DROC11, DHHP14, DCN13, Dim10, DGRB13]. model [DPC15, DI10b, DDTL19, DDB10, DHRA10, DTP19, DLHS11, DWG^{+}19, DDSDW13, DFG^{+}18b, DBJ12, ESG^{+}16, EG10, EBS11, EHBC10, ESW13, ESE15, EPJ^{+}09, EPJ^{+}11, EI13, FD18, FIS11, FIS16a, FZ14, FS11a, FE10, FM15a, FKB^{+}12, FBFM12, FKDW15, FBU11, Fok12, FLR14, FBM19, FRG^{+}13, FM18, Fra16, Fre10, GSRR17, GLR^{+}18, GCSP17, GLR^{+}17, GCSMMG^{+}12, GJ11, GMOP12, GM19, GA13, GGG19, Ger16, GTC19, GI10, GW13, GTB10, GLOC10, GdGL17, GAPK10, Gol16, GCG14, GBJ^{+}11, GS17b, GGGM16, GC16, GMMN18, Gri15a, GT15, Gri11, GAV15, GDJC11, GSN^{+}11, GW12, GHS18, GHS20, GWX17, GKG^{+}18, GMR16, HGM15, HBK12, HMTA17, HT19, HB12a, HFT15, HD17, HP12, Han12, HNO18, HTK^{+}18, HKS16, HS14a, HLVR16, HGGCR13, HGB^{+}13, HM11b, HASM17, HBA10, HGO^{+}18, HWGT15a, HWGT17]. model [HS12, Hor11a, HDNB10, HB18, HT16, HPW^{+}13, HRG11, HKR^{+}19, HGM^{+}16, IBB^{+}15, ISZ18, IGL^{+}12, IS17, JAK19, JDPK15, Jab10, JMZ13, JB18a, JMH^{+}19, Jen10, JSK16, JWS^{+}10, JSSZ12, JL16, JSP^{+}16, JMS18, JB11, JSDEK14, JMK^{+}17, JCLS^{+}11, JCLS^{+}21, JKS15, KSU^{+}18, KS13, KA11, KSK^{+}11, KGJ12, KPG19, KCZ^{+}19, KB15, Kar12, KP12, KHH10, KSP^{+}14, KFB18, KBV11, KBE^{+}13, KMA10, KHK15, KHK^{+}17, KG15, KB19, KGK10, KRR14, KMCH14, KLJ17, KdlRJ18, KJ19, KCSB14, KNA^{+}18, Kit10, KPEK14, KTN^{+}10, KSK^{+}16, KEG17, KCS^{+}15, KS16, KKG^{+}14, Kon17, Kon11b, KFS^{+}13, KSRN19, KL18, KCD11a, KCD11b, KA16, KF12, Kři11, KLB^{+}16, KLHS17, KIT^{+}16, KK17b, KF10, KC14, LLN^{+}19, LPP^{+}19, LF17, LPB17, LPB18, LSP^{+}17, LWH^{+}11, LFW^{+}18, LF16, LP12a, LKK13, LJ15, LCJ16, LCJ18, LM11, LPGTBMA19, LGMV14]. model [LDA13, LHW14, LHFM16, LFB^{+}16, LF10, LMW^{+}10, LSY^{+}10, LWW14, LMYL15, LYF^{+}15, LX16, LL18, LLP^{+}19, LKL13, LS14a, LZY^{+}15, LLH^{+}13, LAPK14, LTS^{+}16, LMJ^{+}16, LGR^{+}12, LW10a, LW10b, LCT11, LZWK11, LWL^{+}11, LZT12, LZT13, LO15, LHD^{+}17, LJ18, LBZ18, LP12b, LMHF13, LLB^{+}18, LSS^{+}14, LHPF18, LDB^{+}14, LC16, LXLF14, LSSG10, LZTD18, LNRK11, LB11, LHL^{+}14, MCL19, MZ17, MOEdP16, Mal17, MTdS^{+}16, MBE10, Mal10, MS13, MVGGB18, MMFK10, MEKK11, MFMS^{+}18, MGGM10b, MYC12, Mar12, MMAS13, MBPS17, MN11, MDEKH13, MBBV14, MTE15, MP13a, MB10, MAM16, MKMG^{+}14, MPC12, MZAI19, MA11a, MSC10, Mic11, MFG14, MED12, MS12b, MMK16, MSE^{+}14, Mog15, MGM13a, MGM13b, MGB17, MF19, MBCM12, Möl12, MROS15, MAS15a, MAS15b, MS14d, MP13c, MKF^{+}14, MKPVH16, MLL^{+}16, MHMM11, MMB11, MR18, NIiT18a, NIiT18b, NiTI18]. model [NS11, Nak12, NH19, NBS^{+}13, NT14, NSO15, NKB16a, NKB16b, New11, NdlPZA10, NBA^{+}18, NTED^{+}19, NWWL17, NGL^{+}10, NHS^{+}16, NTOI16, NSH^{+}10, NWB^{+}10, NBW10, NBW11, OFT15, OCN10, OO11, OSF11, OOY16, OL19, OTTF11, OPS^{+}19, Ou19, PAA11, PHH18, PBYB18, PSS^{+}10, PSS^{+}13, PCT11, PCT19, PaCZ10, Pan19, Pat16, PR17, Pat19, PFJS15, PPW16, PRC15, PDRC16, PG18a, PG18b, PBvdG10, PWL^{+}11, PSV17, PMK15, PCBMM12, PL14a, PL14b, PMW12, PBBB10, Pie12, PC10a, PDT^{+}10, PBR17b, PCS^{+}18, PVGV19, PC10b, PW11, PW14, PGS10, PLHL19, PGHC12, PKH11, PP18, PP19, PP14, PAV10, POK^{+}12, PD16, PS12, PFPR11, RCH14, RM17a, RVZ14, RSS^{+}18, RTR^{+}13, RRL^{+}13, RLW^{+}14, RGA^{+}10, RRBM12, RPGG^{+}19, RWW^{+}15, REC10, RB13, RH19, RRC^{+}11, RLSM17, RA17, RB14, RPP16]. model [RTEKG12, RTEKG15, RTEKG19b, RTEKG19a, RC16, RSD^{+}13, RG18, RBB^{+}12, RRRTR10, RRR15, RW15b, RPCW18, RKJ^{+}11, RQB17, RAR19, RCL16, RD17, RBG^{+}17, Sac15, SMM15, SSD11a, SHG16, SRD^{+}15, ST17a, SC10a, SW13a, SSS13, SGAM12, SEK^{+}10, SGB19, SL10, SKF16, SI12, SSZR17, SYI17, SCA13, SBJ13, SAB^{+}19, SWD16, SBSR13, SGL10, Sch17, Sch23, SSvdM10, SN16, SFT14, SI11b, SW17, SS14, SCS15, SHK14, She10, SR10, She11a, SRV11, gSxFH^{+}12, SMS19, SCF^{+}12, Shi19, SGA^{+}12, SFS13, SPO^{+}10, SK10, SK12a, SGB12, SYSF17, SAB17, SSLB15, SGGY10, SY11, SOCF14, SZ18, SYR11, SMA11, SSME10, SOBC12, SSD11b, SSD12, SGDL12, SDK11, SRH19, SWPC^{+}16, SAD17, SKH11, SGW^{+}18, SKD^{+}10, SAKG13, SVMN11, SLC12, SH17, SLD^{+}17, SB19, Str13, SHSR15, SL12, SAZ^{+}14]. model [SNCM12, SA13, SCT12, ST17b, SM10b, Tac10, TTG19, TWTA^{+}18, TM15b, TDHC^{+}18, TXWW12, TZ13, TXCW15, Tay16a, TEY10, TDKJ15, TSS16, TR17, TW12, TT10a, TALC16, TGZF11, TDM14, Tre19, TSF^{+}19, TF15, TCB13, TT10b, UPWK15, dlRVFK^{+}16, Van16b, VK10, VH11a, VPGDP^{+}14, Vaz10, VLFF12, Veg10, VG13, VM12, VB19a, VDRL14, Wal12, WW12b, WXWL12, WWY12, WPH^{+}12, WJL^{+}16, WML16, WTC16, WBMM18, WL10, Wax11b, WMD16, Wei12, WZY14, WVO^{+}15, WMN18, WCF12, WCJ11, WK12, Wol14, WLD^{+}11, WCHC11, WHZ^{+}17, WBB^{+}18, WGDZ18, WYL^{+}18, WHYMG17, XYN15, Xie11, Xie13a, Xie15, Xie18, XS13, XGZ17, XNW17, XSCS12, Yan10a, YZMY18, YGMT12, YZZ13, YWP13, YW13a, YZY^{+}16, YG12, YD15, YCB16, YY15, YiTS19, YLH12, YKO^{+}16, YBT^{+}17, YLLL12, YMZ18, ZBA14, ZPdFJ19, ZCA^{+}14, ZYJL18, ZK18, ZTT18, ZJY19, ZLZ^{+}11, ZKR15, ZGW16]. model [ZMS17, ZZX^{+}16, ZSW^{+}14, ZWW14, ZMC^{+}18, ZSS10, ZPM^{+}15, ZVMB10, ZCS^{+}15, ZAL^{+}19, ZZS19, dLMD^{+}11, dTKDV15, vHHKB14, vVLS14, vdSS12, MOSS15]. Model-Based [ZKMB19, RG12, KBV11]. model-experiment [LGR^{+}12]. model-free [BFJ^{+}18]. Model-independent [Mun12]. modeled [KHRS10, MWSM13]. Modeling [ASC16, ARG19, ATI12, AAGO^{+}17, AMSSG16, ABA11, Alv18, Amo15, AN10, AES^{+}13, AGGME18, AC12, AFD^{+}17, BR16, BZCS10, BZL17, BVK10, BBS10, BGW15, BKPV15, CHD^{+}10, CFT11, CB13a, Cha18, CTW10, CBC^{+}18, CL18, CB10a, CSDR15, CT16, DE17, DNB12, DFS11, DFT^{+}17, DBBM11, EA19, EBE10, EJ17, EJ19, EEHMH18, EBE14, ESGA15, FGH^{+}14, FKB^{+}12, GWBL12, GML10, GSGV11, GKPB13, GBN14, GMBK14, GPD18, Gon13, GLF^{+}15, GSV16, GMS^{+}13, GAY^{+}11, HT18, HWGW18, HVM13, HH10, HBW^{+}11, HR12, ITN^{+}11, JO13, KBS^{+}19, Kon12, KB10b, KST17, LRL^{+}16, LCMC14, LGK^{+}09, LGK^{+}12, LW17, LC11, LLDW14, LH14, LXZQ14, LGBT15, LA15, LS10c, LSB10, LL17b, MB14, MKRE18, NCLB16, NJP18, OAP14, ONO15, OB10, OVKL14, PNVN18, PKL12, PRV^{+}14, Rem15, RHJ11, RS14a, RRRRSPTR16, SCLA10]. Modeling [SP13a, SPRF13, SMZ^{+}17, SER^{+}12, TGLK19, TAI^{+}18, TTB^{+}18, TWH^{+}13, TSB^{+}17, TLK19, VP18, WZJ^{+}13, WGC13, XCW^{+}18, YS14, YMW12, ZBA18, ZZR10, ZZR13, vTMW15, AMM16, ASC15, AOM19, AD15, ARG14, Ana11, BRK16, BCPL10, BPP^{+}16, BSV^{+}10, BZJP18, BTF19, BTW11, BBP13, BJ18, CWM11, CLPZ14, CSR^{+}05, CLCH10, CLA^{+}16, CSKF10, CCAdS13, CTL^{+}15, CvBF18, DLEMP19, DFG^{+}18a, DHT19, DBM^{+}18, DGD^{+}11, DFM^{+}19, EKvdKvFK13, FAFA17, FSA15, FBK11, FHM18, FBC12, FPR10, FIJJ11, FKV19, GKMC10, GTS15, GL18a, Giv10, GCO^{+}11, GBJ^{+}11, GSV11, GSS19, GMB11, GCYL18, Hal16, HB10, HCS^{+}19, HWD^{+}14, HM13, KCMF11, KDMD13, KMZR18, KSSM11, KAPR12, KBV12, KKO^{+}18, KIH16, KKU^{+}10, KTH16, Kla10, KHH^{+}17, KTJ19, KMLT14, KPLT14, KK18, LF19, LBMG16, LL17a, LDB^{+}14, MPJH13, MRF19]. modeling [MYK^{+}11, MH11, MFM16, MFL^{+}18, MBC^{+}12b, MZMM^{+}14, Mol18, MAFB18, MMP13, MAH12, NPMM15, PAOMV17, PDF18, PBSM19, PZLF19, QJR^{+}16, RMM^{+}16a, RZRSC19, RMM^{+}16b, RM19, SRK^{+}12, SBM^{+}16, SK15, SNY^{+}17, SDT17, SSJK18, SMG18, SRP16, SSM^{+}18, SKPK17, SMJS14, Sht17, SR14, SKK^{+}11, SW13b, SW19, SDPC11, SRAL12, SH15, ŚKSRW16, SPL14, TLL^{+}12, TI12a, TCH14, TRM^{+}14, TKTB18, TC12, TA16, VKKA12, VGPS18, VHM^{+}17, VZB^{+}15, WHWZ18, Whi11, WCM15, XWW^{+}19, YLWZ10, YCL^{+}17, Zad11, ZGY11, ZLZ15a, ZXC^{+}10, vLFM^{+}19, ALD^{+}11, KBS^{+}20]. modelled [RG17]. Modelling [AKS^{+}19, ASGE14, AKdV^{+}14, AK11, ASL^{+}18, BDTR15, BTO14, BTO15, BBWS^{+}13, BJO^{+}16, Bos12, BCPG18, BPPC15, BMF^{+}18, BPGS12a, BPGS12b, CM12, CST16, CLK18, CDD12, CMN^{+}11, CKBCG12, Cut15, DGW^{+}18, Don13, DS10a, DFC^{+}12, EH17, EME^{+}16, ECP^{+}16, FKK14, FGLS10, FKB13, GDF18, GS13b, GBJ18, Gün12, Gün13, HKS15, HKS16, HT13, IS10, IMW15, JYY^{+}18, JSB15, KHS13, KMHdlP10, KBKK14, KM15, LWH^{+}11, LWRE14, LBGW13, LMC^{+}13, LS11a, LWvB^{+}19, MR11b, MSP19, MCWF14, MNH^{+}12, MPP^{+}15, MS12c, Mor16, NNH^{+}16, NBC16a, NBC16b, NM11, PSP^{+}12, PCHM10, PWE15, PGHC12, PSL^{+}10, RGG12, RDSD^{+}12, RLM^{+}14, SW11, SR12, SJSG13, TT19, TTC10, TXTW16, TBMM19, TDSM12, TAB^{+}13, VAG16, VBVD17, WPA17, WBMH12, WBM^{+}15, YCH^{+}17a, YCH^{+}17b, ZLCH18, ZN18, ANK10, AP17, BVJE17, BB19, BCFR10, CXWL11]. modelling [CRC15, CR14, CSBK15, CWB^{+}17, CSKZ19, CHN^{+}15, DBG18, DTGC14, DDF^{+}14, EBE17, EJ16, FWB^{+}12, FK16, FGAM12, FTPN10, FM13, GT11a, HWPL12, HLTR19, HB13b, HHS^{+}10, IvLP12, JSP^{+}16, KPD18, KHB^{+}18, Kun16, Leh18, LTZ19, LMM18, MZWC10, METC12, MMC19a, MM19, MSM^{+}14, MVH11, MSBB13, MFP^{+}14, MG10b, Mor10b, MK16b, MSA^{+}16, MDD13, MTLW^{+}10, MP19, MJV16, NF14a, PH13, PGT^{+}19, PWC^{+}15, PGGvL^{+}19, PHA^{+}17, PMD^{+}18, PM13, RKK^{+}19, RKMG12, STA15, SSR^{+}19, STX^{+}11, TC11, TF16, VC10, WMPF^{+}15, ZLB^{+}18, dBB14b, dSKBS10, PW19]. Models [AK13, FD11a, IIKT13, LK15, MGGGA11, OFB10, PWC^{+}12, PSD^{+}18, SCC11, SLM17, dSRM15, APBS15, ACdlRMR^{+}11, AHJ^{+}18a, Ala15, AKNP18, AD13, AHR10, AHMA^{+}19, BSR14, BT17, BGL10, BSH19, BM11, BBI14, BCPM^{+}16, BKSW16, BPR15, BJPP12, BKL14, BvLH14, BSC^{+}18, CS14a, CAG13, COG19, Coh15b, CMMR13, CNG^{+}12, Cor16, Cre10, Di 19a, DM15, Dil14, DMP15, Dra19, DHT16, FJS^{+}19, FLCS^{+}15, FDM^{+}11, FS16c, Gal10, GLZ^{+}13, GFM^{+}19, GK13, GI15, GE18, GXF13, GA14, GWW^{+}10, GBIR13, GGM15, Gru18, HSD11, HHSA15, HSMCCR19, HVL11, HHE19, HPML18, IMD16, JBSFB12, Jus11, Kae17, KFW12, KM19, KW10, KKYV18, KA15, Kři10, KGC18, KJ11b, KSPA17, KBF18, LS13a, LCSH14, LZA^{+}19, LY11, LL13a, LJS14, LDdA^{+}13, LW18, MR11a, MHC16, MBLV10, MKBE17b, MKRE18, MPBS17, MFG^{+}13]. models [ML10, MGS16, MA10, MC15b, MT15, Non10, NKOS11, NPS18, OB15, OGE10, OHWS18, OT13, OO17a, OTGT10, Pai19, PRM11, PM11a, POW17, PCC16, PBRW11, PDG12, PPT^{+}16, RW15a, Rei12, RC11, RFME^{+}12, RGL^{+}17, RB15, Ros13, SAA10, SE10, SMD^{+}16, SBP13, SBK16, SFD12, Sim14, SG15b, SS12a, cSGFB17, SFSJ12, SBS17, Tal12, TA13, TZ13, TBA14, TF18, Tjø10, TAL19, TAORS10, TWP16, VC11, WL12a, WL12b, Wal18, Wod18, WA14, Yat14, ZARK19, ZLL17, Zha10b, ZLW14, ZLT^{+}19, dlCGSA16, vDRT14, vE11, vVGSE12]. Moderate [VMZK^{+}19]. Modern [Flo11, BCPM^{+}16, Dun11]. Modes [PGG^{+}15, BMN16, EJ16, FLA^{+}19, HY13, JYZ18, KZ14, LZ18, MVGGB18, MRS14, Opr10, PKZ12, RBMS17, SM17a, SK15, SGW^{+}18, TCW14]. modification [Abr14, Dun11, ESG10, JK12, LS14b, SI12, SC13a, SC13b, Str17, Wol14, XG12, XWW^{+}14, ZZ10]. modifications [DMCP14, FGLS10, LLY15]. modified [AGRR^{+}15, AAJ15, DROC11, Gol10a, Gol16, GZ19, PLF19, RRRRSPTR16, WdLS17, XSLZ16, ZGW16]. modifier [JLQ^{+}12]. modifies [RR12]. modular [AAGO^{+}17, WB19]. modularity [ML13]. modulated [CLCH10, DTGD19, GZT15, MPK^{+}12, Zhu11]. modulating [BZJP18, ZGS^{+}10]. Modulation [MYC12, Bac15, MMFK10, PKL12, PBBB10, VGZS18, ZAL^{+}19]. modulations [HKS16]. modulatory [BFH^{+}15, WBRHE10]. module [HSiI^{+}19, LMJ^{+}18, PEZ^{+}13, RPD14, YLF^{+}17]. modules [HZL^{+}11, Kři14, LDH^{+}12, LDH^{+}14, PML18]. modulus [KFG^{+}14]. Moffitt [BBDB13]. moieties [CM17]. mold [Tre19]. Molecular [ADS^{+}19, ASRM15, DHB15, FR17, KSK^{+}17, KMLT14, KPLT14, MTKY11, NZZ19, PYG^{+}19, RMM^{+}16b, Ros15, SMJS14, YLWZ10, ZZ14a, ASH15, AD15, AFM10, BWP10, BE14, CN10, CVPCV^{+}15, Dai12, DEK15, DMM^{+}14, DCL18, Dun11, EJ16, FHG15, FZ14, FLCS^{+}15, GHBI16, HAuR^{+}18, HBW^{+}11, HZW^{+}14, IIKT13, Ing16, JZL13, JL16, KJM17, Kee10, KGP^{+}15, KIT^{+}16, KF10, LCG12, LT14, LKZB15, MPP^{+}16a, Mar12, MAFK12, MS18, MFKS13, ML11b, NNK^{+}15, NS16, Par19, QZ14, RFME^{+}12, RM19, SN12, SS19, Sel10, SG15a, SP14, TWR^{+}18, TGLK19, TLT^{+}15, TAR16, TMS13, XMWC13, XYZ15, YNY^{+}10, YCL^{+}17, Zha10a, ZGY11, ZZ18]. molecular-cellular [HZW^{+}14]. molecule [Di 19a, LAJC19, TS17b, ZZC14]. molecules [AMM16, AIY16, JB19a, KIT^{+}16, LCG12, MS10b, MAA^{+}17, Zol14]. Moment [Lio16, Che10, FS16c, JSB15, PHK15]. moments [ADMR16, BRK19, Cle18a, DK13c, KJH^{+}19, Sch18, Sch21, vV18]. momentum [IPS14]. monedula [GSV11]. monitor [MFKS13]. Monitoring [CPV16, BGLR^{+}19, GLR^{+}17, PGGvdB12]. monkeys [GPD18]. mono [VLFF12]. monoclonal [ADR^{+}11b, FBC12, MFMB12]. monocots [TDT16]. monocytogenes [JRR^{+}12, RMF^{+}18]. Monoecious [AM10, HT16]. monolayer [DFC^{+}12, NTOI16]. monomeric [DFMR19, Xie15]. monooxygenases [LT14]. monophyletic [Di 13b]. monoploid [Ots11]. monotonic [BBR16]. Monte [KSK^{+}11]. moonlighting [Bar19]. mooring [LKZB15]. Mopeia [GRBL13]. Moral [Sig12]. Morales [JLQ^{+}12]. Moran [CAV16, Dal17, GH11, HT16, LPKH15a, LPKH15b, LPH19]. morbid [SST19]. MoRFPred [SBT^{+}18]. MoRFPred-plus [SBT^{+}18]. MoRFs [SBT^{+}18]. morpho [MGC15]. morpho-mechanical [MGC15]. morphodynamic [GBN14]. morphogen [DNB12, MBM11]. morphogenesis [IS12, KRSD11, LL17a, LC10, MGC12, ZCS^{+}15]. Morphogenetic [BJWL15, RV16]. Morphological [CCB11b, DKP^{+}18, AHR10, KHNM16, Sel10, SPMGR10]. morphologically [Wu14]. morphologies [LSMC18]. morphology [KHNM16, LFZN11, MBC^{+}12b, NWB^{+}10, PDL^{+}17, RMM^{+}16a, SMM15, SK11a]. Morphometric [KKOM18, Nij11, VF12]. morphometrics [PWC^{+}12]. mortality [AOR17, Aya12, DBB14a, Dol16, GMS^{+}13, MDE11, RU16, RC13, WKH16]. mosquito [BPR15, HS14b, HLH^{+}18, LBGW13, LL18, NMAZP16, SSG^{+}19, TGC^{+}17, VSD^{+}17, WTC16, ZLT^{+}19]. mosquito-borne [BPR15, ZLT^{+}19]. mosquitoes [BAR14, ECP^{+}16, LS19, PEZ^{+}13]. most [Cam15, FS12, Sil16, Zhe15]. mother [BD10a, SZ15a, SZ16]. motif [AGS18, KBK16, MZ18, SG15b, wLW10, SM11a]. motifs [BM11, EM16, FLM18, HA15a, HLS^{+}12, Kur10, LKP^{+}12, LTP19, LL13b, MD16, PR18, RMSTG13, Str15]. motile [DHV19, LFZN11]. motility [HBW^{+}11, ICG16, LL14, OSF11, SOCF14, VG13]. motion [BJB18, BFW10, CML10, EPP19, FH10a, FH10b, FH14, GWBL12, KO18, RLK10, RCL^{+}10, SK15, SNY^{+}13, Str11, SN17]. motivate [BP18]. motivates [RMB15]. motivational [PMW12]. motive [LL17b, Sch10]. Motor [DL15b, CST19, DLEMP19, DEK15, LM15, LAJC19, MK17, SL19, WdVMD15]. motor-brake [CST19]. motor-mediated [MK17]. motors [LKZB15, MAFK12, NS18, SP14]. mound [KTO^{+}14]. mountain [APS^{+}13, APBS15]. mountainous [ASL^{+}18]. mouse [ALD^{+}11, DLHS11, EUM^{+}16, HHJR11, MPS^{+}11, TRM^{+}14, Zha11]. mouth [DDF^{+}14, RB14, ZJY19]. mouthparts [LWY^{+}16]. moved [Kur14]. Movement [HCS^{+}19, PP17b, AR11, BLS^{+}17b, CB16, DL12b, EPP19, FD11b, Gro17, GWX17, JSB15, KSBvL^{+}12, LWLM18, LXJ15, PWE15, PBD13, PRM12, RLBCJ13, RAMS11, SK10, SK12a, SS15f, WP17, YLH12, vHHKB14]. movements [JP19, KHB^{+}18, MDS16, SBSE14]. Moving [HMW16, DHHP14, Gou16, KA15, MOSS15, WdVMD15]. moving-boundary [DHHP14]. mPLoc [SC10b]. MPZ [AD15]. MRI [LS16c]. mRMR [JW18, MS14a]. mRMR-IFS [MS14a]. mRNA [BBRT12, CB10b, Das18, Di 14a, DM10a, HZL^{+}13, RT15, SFMS16, SC11a, YST14, ZLZ^{+}16]. mRNAs [BLZ^{+}19, Di 14a, RBRB16, Sha14]. MSC [CWP^{+}18]. MSC/chondrocyte [CWP^{+}18]. MSM [MMRCCC10]. mTOR [DFM^{+}19, LCGMH12]. mucedo [ESW13]. much [AW12, KB10a, MK14a, OAP14, ZSSM14]. mucociliary [KBV^{+}15]. Mucor [ESW13]. mucus [IGG^{+}18, IUK14]. Muller [pNMK^{+}10, WL10]. Müllerian [GWE11, LBJ13]. Multi [AH19, Das18, De 19, EBP15, JSZ12, Mei12a, Mei13a, NKOS11, vVN12, ABKS11, ABN15, BPR15, BDR10, Cox10, GSRR17, GXFF13, Giv10, GVSLG16, HTK14, HHT^{+}19, IS16, JCW13, KHX^{+}19, KCZ^{+}19, KRDJ15, KTJ19, LSSG10, MRPH17, MEKK11, MPBS17, MPC12, Mei12b, Mei13b, Mor19, MLL^{+}16, NGJ^{+}14, PLN14, PTNKT11, PBD13, SSG^{+}19, Shi19, SFD12, SR12, SRS18, SH15, SL19, VBV10, WMK14, WMK16, XWC11, XW17, YCR^{+}15, YZ19, ZSH^{+}16, ZWW14, dlCGSA16]. multi-agent [KTJ19]. multi-armed [Mor19]. Multi-class [JSZ12, KHX^{+}19]. multi-compartment [MEKK11]. multi-compartmental [MPC12]. multi-copied [YCR^{+}15]. multi-dimensional [IS16]. multi-epitope [KRDJ15, NGJ^{+}14]. multi-functional [WMK16]. multi-group [GSRR17, LSSG10]. multi-host [BPR15, MRPH17]. Multi-kernel [Mei12a, Mei13a]. multi-label [Mei12b, Mei13b, WMK14, WMK16, XWC11]. multi-level [SFD12]. multi-localization [Mei12b, Mei13b]. multi-locus [XW17]. multi-motor [SL19]. multi-parameter [VBV10]. multi-patch [SSG^{+}19]. multi-peaked [PTNKT11]. multi-phenotypic [ZWW14]. Multi-player [De 19, vVN12, PLN14]. multi-profile [HTK14]. multi-regimes [HHT^{+}19]. Multi-scale [EBP15, Cox10, Giv10, KCZ^{+}19, MLL^{+}16, PBD13, SH15, dlCGSA16]. multi-source [YZ19]. multi-species [MPBS17, Shi19, ZSH^{+}16]. Multi-step [Das18]. multi-strain [ABKS11, ABN15, BDR10, GVSLG16, JCW13, SR12]. multi-task [GXFF13]. multi-trophic [SRS18]. Multi-type [NKOS11]. Multi-voting [AH19]. multiactivity [KG13]. Multiallelic [Bon12]. multibody [SCF^{+}12]. multicellular [DFG^{+}18a, Edg19, EPJ^{+}09, EPJ^{+}11, HNO18, HNA15, iITIM15, KUV^{+}10, LC16, MFL^{+}18, OKVN18, TDSM12, VH11a, Vel17]. multicellularity [KVN16, LSM17, TWB16]. Multiclass [lGLLL14]. Multicopy [FS11b]. Multidimensional [RW15a, GM16b, HB18]. Multidimensionality [LH18]. multifactor [YYC19]. Multifocality [FLR14]. multiforme [CRC15]. Multifractal [YISG14, SCLC13]. multifunctional [HAuR^{+}18]. Multigame [Has14]. multigroup [CFF11]. multihost [PGKZ17]. multilayered [PCT11, PCT19, Pan19]. multilevel [FK18, GCM14, Luo14, PGG^{+}15, SSM15, TGL15]. multilocus [VPGDP^{+}14, YYC19]. Multinomial [KBV11, CW11b]. multiparameter [MK11]. Multiparticle [Kon11a]. Multiphase [HB13b, PM13, MBPS17]. multiplayer [BR12b, GCS12, HWTN15, PBR17a, PGG^{+}15]. Multiple [Edu16, iHIM12, NP15, NWWL17, SHG16, AHD^{+}18, AH12b, AWK15, CZW^{+}11, ERT13, EJK16, GLR^{+}18, GLZ^{+}13, GT11b, HS14a, HB18, HY13, Hua16, JHE15a, JHE15b, JZZK11, JWTW17, KSA16, KTN^{+}10, KS16, KF11, KIT^{+}16, LM15, MIJ16, MGT17, MAFK12, MK19, MSB16, QJR^{+}16, SBK16, SVSP15, SAM17, SL12, SKS^{+}19, TM14, WSTL16, WW19, WCC14, XWC11, YLH12, ZNC^{+}15, vLBJ^{+}13]. multiple-allele [SL12]. multiple-segmented [Hua16]. multiplex [JAB18]. multiplicative [CKF17]. Multiplicity [RBR18]. Multiscale [CRC15, RLK10, BCPG18, CGvG^{+}15, CMHM19, EBE17, GM19, GAPK10, HM11b, HGM^{+}16, KNA^{+}18, LFW^{+}18, LMJ^{+}16, MPK^{+}12, MHH13, MMP13, PDT^{+}10, PGHC12, PMCS16, SKD^{+}10, SW19, ST17b, Van16b, ZPM^{+}15]. Multispecies [ML14, YLLL12, FJC^{+}10, KG13, KL18, NF17]. Multistability [LM11, OCHHZ12, SLR17, MPP^{+}18, SX12, SY12]. multistage [LTZ19]. multistationarity [BFS18, KS19]. multitarget [ZMS17]. multitemplate [IDM15]. multivalued [DRC11]. Multivariate [MKBE17b, BPM^{+}12, GMK18, GBIR13, Gri11, Gri15b, MKBE17a]. multiwhisker [Pat19]. multocida [GTS15]. Münch [JBSFB12]. murine [ABD^{+}11, MSE^{+}14]. Murray [DPCM16]. muscle [AJD11, AJC12, BRN15, DGNT17, Don13, GMOP12, Giv10, GHS18, GHS20, HGGCR13, KCZ^{+}19, Kro11, KSM^{+}15, LO15, LCGMH12, LHPF18, MAH12, MH14, NF16, RGA^{+}10, RW15a, RS12, RS14a, RG17, RG18, SS14, SCF^{+}12, SGB12, Smi14, SAD17, SKH11, SH16, ST17b, TvMG16, TSB10, WC10, ZPM^{+}15]. muscle-tendon [RS12]. muscle-tendons [RS14a]. muscles [ASGE14, BSS11, CSB15, CT16, GGC14, XYN15, ZPM^{+}15]. muscular [HFT15]. musculoskeletal [SCF^{+}12]. mushroom [MBC^{+}12b]. mussel [SM16]. must [LCBO^{+}12]. musth [WHYMG17]. mutagen [BE14]. mutagen-interceptor [BE14]. mutagenesis [BCKS13, BKPV15, MV10]. mutagenicity [HPB^{+}14]. mutant [AH12b, ABN15, GGM12, Gre15, TLT^{+}15, YCL^{+}17, ZFWK17]. mutant-alleles [GGM12]. mutants [ASGG17, CNDK15, HT10, KDK14a, KDK14b, Zha10a, dRFI15]. Mutation [MO12, NIiT18b, AY15, ATTN12, AM11, CJKR10, CNDK15, EA19, FM18, IIKT13, MCL^{+}11, NIiT19, NPH^{+}12, NZZ19, O’F11, SC10a, SBM^{+}16, SMdA16, SAZ^{+}14, WCM15, WZ15, XKCG15, YGMT12, ZTWL12]. mutation-drift [SMdA16]. Mutation-selection [MO12, CNDK15, SC10a]. mutational [KEHK17, LC16]. mutations [AP13b, BST14, BTK14, CCS^{+}16, FSW^{+}16, JJ15, KDK14a, KDK14b, KII19, LMW^{+}10, Mob10, SAB^{+}19, Sel11, SW11, SK12b, TH17, UI10]. mutator [Gou16, Jam16a]. mutators [JJ17]. mutual [HTM16, KM17, MP12, SSCJ16, SN18, UK17]. mutual-aid [SN18]. mutualism [APS^{+}13, AS13, DH16, De 17, KCMF11, LZTD18, MK14d, Shi19, Uit11]. mutualisms [Ezo12, GK10, LIPD12, Rev15, Wan19]. mutualistic [Ezo19, FB18a, GAGP^{+}14, GL14, JA13, MH18, Unc11]. mutualists [EI13, KR19, SJSK18]. mutually [Pat16, WKS14]. my [TCW14]. Myc [RFS^{+}15]. mycelia [Bos12]. mycobacterial [FL12, KHK^{+}17, KHKI17]. Mycobacterium [AMSSG16, AHMA^{+}19, CRH^{+}15, GDB^{+}19, LKAJ18, MM12, PDRC16, PL14a, PL14b, WDL^{+}17, MG14a, PRC15, KS10]. mycobacterium-host [PRC15]. mycoplasma [Hur12]. mycorrhiza [TP14]. myelin [MHD18]. myelinated [MAS15a]. myeloid [CMS16, KEG17, SSM^{+}18, SBM^{+}19, ZCT18]. myeloid-derived [SSM^{+}18]. myeloma [GLR^{+}18, HB18, KS16]. myeloproliferative [ZFWK17]. myocardial [CDM^{+}14, KSSM11]. myocarditis [RBF15]. myocyte [CN12a, LFW^{+}18]. myocyte-based [LFW^{+}18]. myocytes [GKMC10]. myofibrils [MCL^{+}11]. myofibroblasts [HHMT19]. myogenesis [FD17]. myogenic [LSB10, LB11, LB12]. myosin [ITN^{+}11, Kro11, MCL^{+}11, MED12]. myosin-dependent [ITN^{+}11]. Myr [Kur14].

N [HS14b, LDB^{+}14, TT19, AH18, ZZN^{+}19]. N-player [HS14b, TT19]. Nagumo [CHL^{+}19]. naïve [SLvdBMP10]. Nano [TT17]. Nano-QSAR [TT17]. Nanoarchaeota [Di 14b]. Nanoarchaeum [Di 14b]. nanofluid [ZG19]. nanomimic [KIH16]. nanoparticle [ALM^{+}19, GSV16]. nanoparticles [MHD18, SK15]. nanopatterned [XLG^{+}15]. nanopatterns [CML11]. nanopores [ACH15]. natalensis [GRBL13]. national [LCJ18]. native [Mas18, MIH16]. Natural [BWL12, Gra18, Hua16, LE19, MMH10, SFV16, Sel17, BWB11, BZ18, BBS18b, Cro17, Di 16, HJWC11, HM15, KGK10, KA17, Orr19, RM17b, RLSN14, SGW17, SFS13, TTC10, TMHM12, TVA18, YDC^{+}13, dMP11, vV18]. naturally [PRM14]. Nature [CEdLSG^{+}16, CTA15, Mac11, Esc11, KK12b, LG10b, XLG^{+}15, YDC^{+}13, ZGG^{+}19]. Nav1.6 [BGCB12]. navigation [Möl12, PW11, TWP16]. navigational [CB16, PW14, SBW11]. NCSTN [HRHAAA15]. N’Djamena [LLN^{+}19]. Neanderthals [Flo11]. near [IG14, ICG16, SL17]. near-polytomy [SL17]. nearest [ABP^{+}11, NMZ19]. Necessary [Sin15, Bal13, DDRP13, GPT17, JEA18, PTNKT11, Tox18]. necessity [vV18]. neck [HHF11, MB12a, MM19]. necrosis [EKvdKvFK13]. Necrotic [BFGS10]. necrotizing [AES^{+}13, BTH^{+}13]. nectar [EM11]. need [CW11a, TT10a, WS10]. needs [KOS13]. Negative [CMB^{+}12, CPS10, DZE11, Zab11, AAJGCD15, BBD18, CXC18, CB13b, Cza14, DHS^{+}15, GI15, KHS13, LS16a, LKP^{+}12, LZ18, MK18a, Mog15, NJP18, Ric19, SRD^{+}15, SS15a, SC10b, Tos11, ZS10, ZW16]. negative-feedback-regulated [ZS10]. Neighbor [RM17c, NMZ19, SJSK18]. Neighborhood [MR15, ZLY^{+}13]. neighbors [WMD^{+}11]. neighbour [BSB^{+}13, MBLC17]. neighbour-dependent [BSB^{+}13]. Neighbourhood [HBK12, YZX15]. Neisseria [JSC^{+}16]. neither [GCM14, Non10]. Neoadjuvant [MB12a]. neocortical [Ing16]. neoplasms [ZFWK17]. neoplastic [Mal10]. neostriate [GMK18]. neovascularization [GCYL18]. nephritis [LZL^{+}19]. nerve [BPPC15, GU13, MAS15a]. nervous [HJLNZ11]. Nest [AR17, Cro18, Gal10, LTSTD15, NPS10, SOIO10]. nest-site [Cro18, Gal10, NPS10]. Nested [TA10, AKNP18, JCW13]. nesting [NSO15]. nestmate [New11]. net [ADB^{+}13, HS14b, KA16, NDZMA14]. nets [HS14b, NMAZP16, YGMT12]. Network [BTG^{+}15, DQS^{+}15, DBB^{+}17, LMJ^{+}18, ML13, PJL14, WLL^{+}14, YLF^{+}17, ZC14, AENK12, AJC12, ABJ12, AANF16, ABIM10, BRA15, BMSEE14, BDS12, BJJR10, BBS18b, BC13, BKL14, bCR19, CW15a, CM14, CV19, DGM15, DG10a, DHBS19, DWM15, DL12b, DL16, FW15, For10, GSYS10, GB12, GBN14, GLZ^{+}11, GXF13, GSN^{+}11, HCK^{+}11, HRCA19, HYZS14, HZW^{+}14, HZL^{+}11, HWD^{+}14, ITN^{+}11, JAM18, JAB18, JR17a, KSK^{+}11, KHK15, KEG17, KSKS13, KS10, LPB17, LPB18, Lar16, LMM^{+}13, LNL^{+}11, LLW^{+}18, LKL13, LS14b, LL13b, LDJW16, LYZ^{+}18, LSLL13, MvdDW13, ML09, ML12a, MBE10, Mal18, MZM^{+}16, MG14a, MPT19, MBF^{+}15, MGC13, MF15, MGM13b, MK14d, MP13c, MYOS14, NvD17, NRKE18, NKM^{+}12, OHWS18, OABI12, OAJK10, PPBD10, PS13, PM14, PR18, PHK15, PHA^{+}17, PRV^{+}14, PMKM10, QYO10, RPP16, RDSD^{+}12, SN12, SCÁBM10, SCÁBM11, SSvdM10, SWTO15]. network [SLT^{+}18, SW19, Str15, TTC19, TK10b, VC11, WH14, WYL^{+}19, WLW19, WLW20, WLY^{+}17, WCC14, WAC14, XSKA17, XNW17, XSCS12, YZMY18, YWW^{+}14, YWZ^{+}16, YBH^{+}19, YC13, YZ19, ZMAM19, ZOG14, ZSO16].

Network-based [DQS^{+}15, ZC14, BRA15, GSYS10, ML09, ML12a, NKM^{+}12, XSCS12]. networked [MAF^{+}19]. networking [KAZ11, KAZ13, NLB14, XMWC13]. Networks [LS13b, AAGO^{+}17, All11, AD13, AdGM12, Arc16, AGGME18, BTO15, BFS18, BZCS10, BLS17a, BB11b, Bro13, BR12b, BFLS17, CP14, COWA11, CST^{+}12, CR19, DBD12, DTGD19, DG10a, EG10, EBP15, EB15, Est10, ER18, FKK14, FK16, FJS^{+}19, Fer11, FMS^{+}12, FVTS16, FRTP13, FM18, FWR19, GS17a, GSCS11, GE18, GXF13, GCZ^{+}12, GMM15, GM16b, GMK18, GJS^{+}10, HBK12, HS15, Hal17, Hal16, Hay16, HL14, HXL16, HXL18a, HB16b, HB10, HM10b, HBL10, HMW16, IMM16, IMA16, IS10, ISB^{+}11, IU10, IvdSW16, IM10, JC16, JCW13, JKS15, KMD^{+}12, KS19, KSP18, KR15, KTH16, Kon11c, KKM12, KJSS10, KKYV18, KdSUS10, KJ11b, LLSO13, LCMC14, LWH^{+}11, LR16, LS14b, LCL14, LLJ18, LMJ^{+}18, LC10, LT11, MC16a, MMC19a, MK11, MT14, MBLV10, MC19, MTE15, Mig16, MGC13]. networks [MJC^{+}11, MdSPBL16, MS10b, MFKS13, MF15, MS12c, MLBA12, MLBA13, MRS14, MBRRI19, ML11b, NSS^{+}11, NGN13, NW10, NMS10, OYY16, Opr10, PAOMV17, PSJ15, PPM12, PMP13, PGKZ17, PMYHR12, PP12, PP14, QZ14, RA10, Ric19, RFME^{+}12, RBHK14, RLSN14, SAA10, SSD11a, SW16, SHN12, SSZB15, SP13a, SBP13, SMC17, SVB^{+}10, SWTO15, SSH^{+}19, SNC12, SBK16, SHM^{+}18, SCKL15, SMC^{+}13, SHS13, SAZ^{+}14, TL15, TP10c, TM16, TLK19, UK17, VPGDP^{+}14, VNS18, WOB15, WcW14, WHT17, WMCL18, WM17, WB19, WDN^{+}19, WL15, WL12c, WMT10, XLSF19, YIS^{+}17, YSH^{+}14, YAK17, YA14, YS11, ZZC10, ZG10a, ZG10b, ZGWL14, ZYJL18, ZHC10, ZCXF13]. Neural [BMSEE14, FHR13, FR13b, FR14, GMM15, KAZ11, RFME^{+}12, Rob11, Sch19c, SST19, AJC12, AANF16, BF13, FIS11, FIS16a, GXF13, GM16b, GMK18, KAZ13, KHK15, KA16, LFZN11, LCMC14, MPL16, MGM13b, MvAKR17, Opr10, PM14, PMKM10, RS14a, SX12, TTC19, Tay16a, Wal16, WYL^{+}19, YM14, YC13]. Neuraminidase [DGD^{+}11, TLT^{+}15, YNY^{+}10]. neurite [vE11]. neurocranium [GA13, VDH^{+}15]. neurogenesis [BFR14, SB19, TSB^{+}17]. Neuroid [TTG19]. neurological [Sat18]. neuromorphological [GMK18]. Neuromyelitis [LSP^{+}17]. neuron [GMM15, GL12b, IBB^{+}15, RMM^{+}16a, RNN15, Thu15, WSLC14, WGC13]. neuronal [ABJ12, BB11b, CP14, CTJ10, DHT16, FW15, Fer11, KAZ11, KTN^{+}10, LL13b, Pat16, SHG16, ZMAM19]. Neurons [GMM15, ZMAM19, BPF15a, BPF15b, CW11c, DLHS11, GM16b, HMM16, Hor11b, KB10b, MZM^{+}16, PJ18, RKMG12, WCC13, ZKMB19]. neuroprotection [NRS^{+}16]. neuroprotective [ÁTC^{+}14]. neurosphere [SZ18, ZZS19]. Neurospora [XCW^{+}18]. neurotoxic [WB15]. neurotoxins [MZ18]. Neurotrophin [MHD18]. Neurovascular [DvDBD15, CS11a, DBD16, FD11a, WK12]. Neutral [AHP^{+}19, KS14, PMR11, BdABY13, DSA^{+}16, DZE11, Gre15, HL11, HSM17, Hor11a, HT13, KS13, Mas16, PC13, SK12b, TZ13, WCPF15, Wax11a, YISG14, ZPdFJ19]. Neutral-like [KS14, DZE11]. neutrality [SC16]. neutralization [MR11a]. neutralizing [CDK11]. neutrally [Ios16, KHRS10, SB12]. neutropenia [CHN^{+}15, LM11]. Neutrophil [BPLM12, CHN^{+}15, ZLM12]. neutrophils [OBK^{+}11]. newborn [LZY^{+}15]. Newman [MZM^{+}16]. newt [JMK^{+}17]. Newtonian [AABS16, ZG19]. next [MMFK10, PGGvL^{+}19]. next-generation [MMFK10]. NF [ÁTC^{+}14, KB19, GM13, HHJR11, HSiI^{+}19, TC11, WPH^{+}12, ZBA14]. NF- [GM13, HSiI^{+}19, TC11, WPH^{+}12, ZBA14]. NF-Y [HHJR11]. Ngn2 [BF13]. Niche [GFH^{+}18, AHP^{+}19, BGL^{+}19, BRP^{+}18, DB11, DZE11, HL11, KWO19, KDMK16, Mal15, PC13, RT19, TZ11, TZ13, ZZL^{+}11, dSdS13, tBdR18]. niche-based [DZE11]. niche-neutral [TZ13]. niche-stabilized [PC13]. niches [DEG^{+}14, LH18]. nidulans [GS13b, KHS13]. niger [GT11a]. Nile [MRPH17, RC16, ZXH19]. NIS [Cre10]. nitrate [LRL^{+}16]. nitric [DK13b, DBD16]. nitrogen [BBS10, DDS13, MHC16, TP14]. NL [GDPPSS^{+}11]. NLRP3 [VLPH17]. NLRP3-inflammasome [VLPH17]. NMDA [ML11a]. NMDAR [HKS15]. NMR [ZZ14a]. No [JEA18, WWB14, EHSR17]. nocturnally [MYN^{+}15, MSBB13]. node [CZPC^{+}18, GMM^{+}13, MMFK10]. nodes [Di 19c, HRCA19, MMLK11, MGB17, SSZB15]. Noise [GMNY14, HNA15, OO17a, Van17a, BPP^{+}16, BFW10, CHL^{+}19, CCF^{+}14, CKF17, DHT16, FR17, HSM17, HM10b, IU10, JZL13, JRG14a, JRG14b, JRG14c, KHRS10, Kon11a, Kon11b, Kon12, Lei09, Lei10, LXS15, LYW11, Lug15, LRH15, MB17, OM10, OS11, Pat16, PJ18, PR18, Pol12, RAD14, RLSN14, SVB^{+}10, TST^{+}13, Van17b, WRZ^{+}12, YRD17, ZCSRG12, ZLL^{+}12, dSdS13]. noise-based [Pat16]. Noise-induced [GMNY14, MB17, PJ18, TST^{+}13]. Noise-plasticity [HNA15]. noises [JC10]. Noisy [SWD16, BPF15a, BPF15b, JRMS12, LLSO13, PBEI12]. Non [ACdlRMR^{+}11, BBM^{+}97, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, GJ12, GWW^{+}10, GSF13, Meh17, SC11b, SY12, Tsu19, ZJS11, Akç12, AABS16, BBR16, BZ18, BCF16b, Bon10, CKF17, DKS15, Das18, DL15a, DHBS19, EI13, FT12, GY19, Giv10, HSMCCR19, HA15b, Hua12, Jam16a, JLQ^{+}12, JAHKH12, Kae17, KS13, Kar15, LS13a, LC11, LDdA^{+}13, LC16, MWSM13, MSP19, MV10, MG14a, MAF^{+}19, NP13, Oht12, PHK15, Pie12, PB16, RK19, Sel12a, Sel16, SCLC13, SS17b, TGB^{+}18, TDZ^{+}18, Van17b, VLFF12, Veg10, WKH16, YBT^{+}17]. Non-acquired [GJ12]. non-binomial [WKH16]. non-canonical [Sel16]. non-classical [Hua12]. non-coding [DHBS19, MWSM13, SCLC13]. non-complementary [Sel12a]. non-costly [Akç12]. Non-covalent [ZJS11]. non-destructive [MSP19]. non-endemic [RK19]. non-enzymatically [Pie12]. Non-equilibrium [Tsu19, FT12, JLQ^{+}12, KS13, NP13]. Non-extensive [Meh17]. non-Gaussian [CKF17]. non-homogeneous [PB16]. non-homologous [LC11]. non-identifiable [HSMCCR19]. non-isothermal [LDdA^{+}13]. non-lethal [Bon10]. Non-linear [ACdlRMR^{+}11, BCF16b, Das18, JAHKH12, TGB^{+}18, Van17b, VLFF12]. Non-local [BBM^{+}97, BBM^{+}13a, BBM^{+}13b, BBM^{+}13c, GWW^{+}10]. non-Markovian [PHK15, Veg10]. non-membrane-bound [SS17b]. non-monotonic [BBR16]. non-mutational [LC16]. non-mutator [Jam16a]. non-mutualists [EI13]. non-natural [BZ18]. non-Newtonian [AABS16]. non-normality [MAF^{+}19]. non-overlapping [Oht12]. Non-random [GSF13, SY12]. non-senescence [DKS15]. non-small [GY19, TDZ^{+}18]. non-spatial [YBT^{+}17]. non-specific [DL15a]. Non-stationary [SC11b, Kae17, LS13a]. non-stem [Kar15]. non-synonymous [MV10]. non-trivial [HA15b]. non-uniformities [Giv10]. nonequilibria [NN18]. Nonequilibrium [PDG17, FMS^{+}12, Kla10]. nonergodic [Wal18]. nonhuman [KLM14]. Nonlinear [Cle16, CA18, FJR19, HBVP15, SKF16, She11a, ARR14, AHJ^{+}18a, ABV19, AS16, BBP13, DFMR19, FJC^{+}10, GZFX14, MDEKH13, MBBV14, PBEI12, SF15, XF13, XLLF17, YZY^{+}16, YST14, ZLW18, Zol14, YiTS19]. nonlinearities [HFT15, NN18]. Nonlinearity [TM15a, SM10b, TST^{+}13]. Nonlocal [Sie18, Ayd18, CCS18, SA17, She10, She11a]. nonmutator [JJ17]. nonnative [Mas18]. nonparametric [Jab15]. Nonself [Coh15b]. nonsmooth [CXT19]. nontypeable [DMM^{+}14]. nor [GCM14, Non10]. norm [Tak17b]. normal [BPS15, LFD11]. normality [MAF^{+}19]. Normalized [PM11b, ZLY14]. norms [MJI14, SAB17]. north [DB18, Kur14, DB11]. north-south [DB18]. northeast [BLS^{+}12]. Northwest [VPGDP^{+}14]. nose [MSB16]. nosed [BCF19]. nosocomial [HLTW14]. Notch-mediated [TSB^{+}17]. note [CFGRB17, CR16, CD16, JV14, SK18, Ano19-51]. nothing [DS10a]. notice [BBM^{+}13c, GZ14b]. notifications [YTGW16]. notion [AHJ18b]. notional [AK11]. notions [MRPLAS15]. notum [iHIM12]. Nouvellet [RRC^{+}11]. Novel [Dai12, FJS^{+}19, MPCTGJ^{+}15, RVP19, ACMK12, BW13, CA17, DCS14, GBC^{+}16, HLHY17, KRDJ15, Kla10, LS10a, LZL^{+}17, LLP^{+}19, LCL14, MB12a, NGJ^{+}14, QLWC19, RA10, RP18, SM18, SKH17, SDRA^{+}15, SBB18, SQZ^{+}16, TAR16, TP17, VM16, WTQL10, Wu14, WGDZ18, XSTW14, YW13b, YYH^{+}14, YC13, YC15, ZMNd^{+}10, ZGY11, ZTWL12, ZCA^{+}14, pZWZ^{+}16, ZKHL16, ZLZ^{+}11, ZDY11, ZMS17]. Novelty [Mar19, Fue18]. November [Ano10y, Ano10-32, Ano11-28, Ano11-46, Ano12-34, Ano12-41, Ano13-38, Ano13-34, Ano14-43, Ano14-47, Ano15-32, Ano15-35, Ano16-44, Ano16-33, Ano17-30, Ano17y, Ano18-43, Ano18-42, Ano19-43]. NREM [PRK13]. NRTIs [KMA10]. NS3 [TWR^{+}18]. nsSNP [KKS^{+}14]. NTCP [KBV11]. Nuclear [KKO^{+}18, DAA17, KKG^{+}14, MP13b, SAGAGB17, SLS17, YYB^{+}19]. nuclei [HZG^{+}17]. nucleic [Xie15]. Nucleo [SWR11]. Nucleo-cytoplasmic [SWR11]. nucleoside [KJM17]. nucleosomal [WZWY11]. nucleosome [GZY12, JYZ18, Tri10]. nucleotide [ABP^{+}11, Bec14, BKPV15, FS15a, FS16a, HNL^{+}11, LLL^{+}17, LHH12, MV10, Sel13, Sin15, Yin17, ZDY11]. nucleotide-based [ZDY11]. nucleotide-gated [Bec14]. nucleotides [AdSG14, BM11, Sel14b, Sel15b, Sel16, Sel17]. nucleus [GMM15, GM16b, KTH16, KB10b, SP13a, VH11a]. NucPosPred [JYZ18]. null [BZ10, JZZK11, NBW10, NBW11, RRC^{+}11]. null-model [NBW10, NBW11, RRC^{+}11]. number [DDP12, Di 14a, Di 18, GBRS19, HM11a, HHA17, HRCA19, HK16, JOM16, KL13, KAKK19, LS10b, PIPB10, PCB14, PLF18, SI10a, STH18, SV14, WDL^{+}13, YI16, Zhe16]. numbers [EM11, Gre16, HHA^{+}13, KD16, RMSTG13]. Numeral [Woh15]. Numerical [DLYZ11, FWSG17, HWGT15a, KA15, LBS^{+}10, LBS^{+}11, MN11, PFJS15, Rev15, AABS16, AP13a, DAA17, FKMG15, GVCG12, HM11a, HWGT17, HM15, HZLX11, LSMP14, OA15, PAOMV17, SPRF13, SVMN11, SAI14, ZS10, pZWZ^{+}16, ZSH^{+}16]. Numericalization [CJ12b]. nurseries [CPV16]. nursery [LIPD12]. Nutrient [MGML10, TGB^{+}18, BID15, BWL12, CTB18, CLCH10, DMS^{+}16, GVCG12, Gro17, HAK^{+}19, KFA12, KK18, NNvdB17, Ou19, PWL^{+}11, RM10, WKS14, WST15]. Nutrient-limited [TGB^{+}18]. nutrient-poor [BID15]. nutrients [CCB11b, KSP12]. nutrition [HFD17]. Nutritional [KRSD11, EHBC10]. Nye [Ou19].

O [TTC19, ZNC^{+}15, LDB^{+}14, LDB^{+}14]. O-GlcNAcylation [ZNC^{+}15]. oases [CW15c]. obesity [BRA15]. objectives [BUC14]. Objects [MRPLAS15, RDD14]. obligate [KCMF11, O’F11]. observability [Gha16]. observable [Mar19]. Observation [ŚKSRW16, GDB^{+}19, MI19, OSN18]. observations [Di 13a, KSRN19, RB15, ZBA14]. observed [ARB13, ASL^{+}18, BM14, NRKE18, RN12, SGGHS19, SN17]. Observer [MZAI19]. Observer-based [MZAI19]. obsessive [RM17a]. Obstacle [NTY16]. obstructed [DLL^{+}18]. obtained [WF17]. occult [GC18]. occupancy [MR11a, PBvdG10, UBP12, Xie13b]. occurrence [AIY16, Cut15, DPCM16, MD16, PDB^{+}15, PNP^{+}16, PDNP16]. occurrences [AMRRI10]. occurs [Cle16, DLMK12, HHA17]. OCD [SST19]. ocean [Di 13a, RTFP^{+}17]. oceanic [Ric17]. octahedron [Den19]. October [Ano10-40, Ano10-44, Ano11-29, Ano11-36, Ano12-44, Ano12-31, Ano13-35, Ano13-40, Ano14-40, Ano14-38, Ano15-42, Ano15-33, Ano16-37, Ano16-47, Ano17-40, Ano17-47, Ano18-46, Ano18-45, Ano19-38, Ano19-30]. octopus [ZKMB19]. ODE [KW10]. ODEs [KNA^{+}18, TBMM19]. odor [Alv18]. odor-landscape [Alv18]. odour [New11]. oestrous [POK^{+}12]. off [CLK18, HMH14, JB18b, KZ14, MS14b, MFG14, PBB10, WK15]. off-lattice [MFG14]. offending [BBDB13]. offs [Chu18, FtL18, MLT10, PBD11]. offset [DZE11]. Offspring [AOR17, Cam15, CL13b, Déb17, Fil15, For14, Jam11, MHKA16, OGA16, Sak10]. oil [AK15, BKR14]. ok [Mar11]. old [Di 13a]. Older [YHZ14]. oligodendrocytes [WRH^{+}16]. oligomeric [IKHL16]. oligomers [TCYY^{+}12]. oligonucleotides [DCS14]. omnigenic [PG18a, PG18b]. omnivore [GVCG12]. oncogenesis [DKR16]. Oncolytic [JKF19, BB19, KW10, ME15, OAP14, SZ15b]. One [GHS18, GHS20, AGPK13, AS12, BG12, Di 14a, DHHP14, Ezo12, Jam14, JZ17, JZZK11, KK12b, Lai18, LFM11, MMH^{+}12, New11, NTED^{+}19, TW13, Toz15, WL13, WW19]. one-dimensional [BG12, DHHP14]. one-shot [AS12]. one-sided [JZ17]. one-to-many [Ezo12]. ones [BC15]. online [Ran12]. only [Di 17a, Hor11a, KK12b, TT10a, TF17]. onset [EAN14, EBX17, HDZ^{+}19, MFMS^{+}18, SPA18]. onto [DRC11]. ontogenetic [WSRG18, tBdR18]. Ontology [Hua12, YLF^{+}17, ZT16, DTY12, WMK13]. oocyte [HB16a, SOBC12]. oocytes [KG12]. oogenesis [RBMS17]. OOgenesis Pred [RBMS17]. open [MB17]. open-loop [MB17]. operates [WKH16]. Operating [Str17, LKZB15]. operation [BAM^{+}15]. Operational [ABK^{+}12, HPML18, Ram10, SKPK17]. operator [HVSZ10]. operators [LG10a]. operculum [XWW^{+}19]. operon [HVSZ10, Mic13, ZCSRG12]. OPG [PZS^{+}10]. opinion [SAB17]. Opisthorchis [BHO^{+}18]. opponents [YA14]. opportunities [AB11, SH15]. opportunity [PVCEC18, YSYI13]. opposing [DEG^{+}14, MJI14, SP14]. opt [Kur19]. opt-out [Kur19]. optica [LSP^{+}17]. optical [KCJ^{+}11]. Optimal [BI10, BHR10, CFCM13, CJL15, CBGS18, DHP11, EzMMH18, GVCG12, GA16, GKWG17, HS14c, Ish16, KKUM10, LCCC10, LKK13, LDM16, MIJ16, MK14c, Mit17, MI11c, MRS14, NNvdB17, NS13, NC15, OT13, PP17a, PRM14, SI10a, SDD15, ST16, SSB^{+}15, SSST19, SBM^{+}19, TBQG14, Vaz10, YD14, ZSW11, AS16, Cha18, CL11, CMM19, DB19, FD18, Fil10, KGJ12, KTO^{+}14, KLJ17, Kři10, KST17, KA17, LCC15, LC17, MDMG14, MW14, Mig16, O’F11, RW15b, STN^{+}19b, SL17, TEY10, TEW12, TBA14, TBMM19, VBV10, WCJ11, WSW10, YTGW16, ZCT18, ZZ14b, ZLT^{+}19]. Optimality [CKB19, LLP^{+}19, RFM^{+}12]. optimally [TDT16]. optimisation [CHN^{+}15, FCS18, SCF^{+}12, SK11a, Tra16]. optimise [EM11]. Optimization [Car17, JH12, MG10a, BMI19, JAHKH12, LDWW14, MDS16, MMAS13, MGM13b, NF16, PP18, PP19, RA10, SSJK18, SK16a, SS18, WOB15, YYC19, ZXC^{+}10]. optimization-based [YYC19]. optimize [DXW^{+}16, QQJW13]. optimized [HB18, TAI^{+}18, ZZN^{+}19]. optimizer [ZZN^{+}19]. Optimizing [HL19, IDM15, PD13]. Optimum [NC16, SK11a, TTC10]. option [HNV^{+}16, IIVR10, Kur19]. Optional [JOAN14, PCN17, GN15]. options [CBBO16]. Opuntia [TVA18]. OPV [Hou16]. ORAI [SAB^{+}19]. oral [KR15]. oranges [GP11a]. orb [Mac10]. orb-web [Mac10]. orbit [HA15b]. order [CW14, DS19b, Gou16, HSD11, JRG14a, JRG14b, JRG14c, KDST15, LYF^{+}15, MBLV10, PP14, PD16, RA17, RBHK14, THBM10, XG12, YIS^{+}17, ZRGW19]. order-disorder [RA17]. ordered [SKH17]. Ordering [Luk14, ZEJA11, ASH15, CzST^{+}18, RNVP10]. ordering-induced [ASH15]. ordinal [SQZZ17]. ordinary [BMM^{+}14]. ordination [Pav19]. organ [CPPLAB18, KKOM18, MS14c, SCÁBM10, SCÁBM11]. organelle [DTY12, RFdL15]. organelles [All17, SS17b]. organic [HWGT15b, HWGT17]. organisation [CCS18, KKG^{+}14, MMJB17, MM15b, Pai19]. Organising [KdSUS10]. organism [HNA15, SI11b]. organism-created [SI11b]. organismal [KMN16]. organisms [DLSD15, OKVN18, Ots11, Vel17]. Organization [SSH^{+}19, ABD^{+}11, AMRRI10, CML11, DPRSS11, Di 17a, LS15b, MDS16, NBS^{+}13, OS11, RC13, RDP16, SHM^{+}18, TED10, Zha10b]. organized [CDD12, IU18, VY12]. organizing [GMK18]. organogenesis [Cox10]. organs [CCG^{+}18, GRB^{+}13, Wei12]. Orientation [SPH12, ABD^{+}11, BZL17, Jen10, Js12, Kro11, Mac10, PW11, ZDG^{+}10]. oriented [CFCM13, CB17a, HAB^{+}17, SS19]. Orientia [KGK10]. Origin [Mau15, Bau18, Cle10, CB17b, Dai12, Di 13a, Di 13b, Di 16, Di 17a, Di 17b, Di 19a, GS17a, GG16, Han10, Jam14, Koe15, Lan17, LP17, LC16, Mar17a, NI11, Pel18, PC17, Rad16, Sat17, SHH15, SS17b, Tox17, VC15]. original [Sat17, TLT^{+}15]. originality [HMM11]. originate [Di 14a]. originated [Di 14b]. originates [Laz13]. originating [Miy17]. Origins [GJ12, KSvdH18, ZML19, BBDB13]. ornamentation [DEG^{+}14]. ornaments [MM18]. Ornstein [BGKL17]. orthogonal [MMG^{+}16]. oryzae [JSF^{+}11, JSF^{+}11]. Oscillating [KFA12, BF13, GH11, GTSP16, GFS15, RR16]. oscillation [AJC12, HSiI^{+}19, HL14, KKM12, SN12, SWD16, SP14, TSB^{+}17, Toz15, ZML19]. Oscillations [FWMC14, Pag19, ARR14, ARG17, BH13b, BKSW16, CCF^{+}14, CdMB10, CFF12, CLW12, Dil14, GGG19, GM13, HLVR16, Hor11b, HDZ^{+}19, HBA17, JAK19, LKP^{+}12, MC16a, MCWF14, MvAKR17, OCHHZ12, PSP^{+}12, PJ13, TST^{+}13, UT17, WTC10, ZKR15]. Oscillator [MB18b, CTJ10, COG19, JC10, MVGGB18, Zha10b, ZAL^{+}19]. oscillators [CHL^{+}11, Gon13, MK18a, Pag19, SX12, Tou14, Zha10b]. Oscillatory [Mob10, SK11b, DHBS19, DFMR19, GDF17, HSG^{+}18, IU10, MMAS13, RFS^{+}15, SW19, WCHC11]. osmolarity [SABB15]. Osmotic [DFG^{+}18b]. osseointegration [PVGA12]. ossification [CGRRGM^{+}16, FHLM10, GA13]. osteoarthritis [WGS15]. osteocyte [Bue15, KLN^{+}12]. Osteocytes [Bue15]. osteoporosis [TSF^{+}19]. other [Jam15, NT14, PS13, RG17, SSS17, SS12b, TDW16]. other-regarding [SS12b]. otherwise [TZ13]. OTUs [WZ17]. out-frame [SSCJ16]. out-group [KM12]. out-of-phase [DGJ15]. out-of-place [HY16]. Outbreak [PM11a, AHJ^{+}18a, BTW11, CW15b, CBC^{+}18, HWGW18, JWTW17, LCJ16, LLD^{+}17, NI11, SKPK17, WN19]. outbreaks [BGW15, DGM15, FHSG15, HYSG19, KLB^{+}16, LBGW13, MW13, NYSM12, OA11, TA13]. outcome [CK19, DKF17, GBM18, HBA10, MGS16, NdlPZA10, Oku15, POP^{+}19, TTB^{+}18, VZB^{+}15]. outcomes [AB16, KC17, POK^{+}12, WD11, dBE16]. outcompeting [Flo11]. outer [GTS15]. outflows [TF18]. outperform [BC15, RPCW18, TZ13]. outperforms [Kur17]. output [GMM^{+}13, VDH^{+}15, ZAL^{+}19]. outs [MdSPBL16]. ovarian [DDHM12]. over-infection [FCC^{+}10]. overall [Laz13, Mar17b]. overcomes [CGD13]. overdose [BRN15]. overfished [TM14]. overhead [Fil15]. overlap [MAH12]. Overlapping [Sel12a, GSRR17, LG17, Oht12, Sel12b, Sel13]. overlaps [APBS15]. overlay [LLJ18]. overloaded [KSSM11]. overlooked [RFdL15]. overproducing [SJK18]. Overpunishing [DDRP13]. oversampling [JZ17, ZD18]. overshoot [CWF^{+}16]. oversimplified [AKNP18]. overuse [LJI17]. Overview [Ano19-28, CKS15, Tau15]. ovigenic [DPvBvA12]. ovulatory [GS17b, Nak16b]. own [HFT^{+}18, RBRB16]. Owner [CK19]. oxidase [RS19]. oxidative [GA14, KMN16, STNT17]. oxide [DK13b, DBD16]. oxide-red [DK13b]. oxidopamine [PDNP16]. oxidopamine-induced [PDNP16]. Oxygen [CGvG^{+}15, Van16b, ASGE14, CRGS18, DLL^{+}18, GSGV11, JMH^{+}19, LMJ^{+}16, NR14, PSV17, RGL^{+}17, RS14b, SSFG15, SSB^{+}15]. oxygenation [SEK^{+}10]. Oxytricha [BKC^{+}16]. oyster [JCLS^{+}11, JCLS^{+}21, PKH11]. ozone [LCMC14].

P [RRC^{+}11, DLM^{+}19]. P-glycoprotein [DLM^{+}19]. P. [NKL10, RRC^{+}11]. P1 [KMLT14, KPLT14]. p300 [DK13c]. p53 [CSKZ19, DCN13, Eli17, HAB^{+}17, LNL^{+}11, LYW11, ZSCL16a, ZSCL16b, OAJK10, STX^{+}11]. p53-Mdm2 [OAJK10, STX^{+}11]. P56S [KKS^{+}14]. P UNSAT [HXL18a]. PacC [KHS13]. pacemaker [DB18, HVL11, LW18]. pacemakers [LWB^{+}15]. pacemaking [MS10a]. pacific [TDM14, GDJC11, SH17, TTW13]. package [ZLZ^{+}11]. Packard [Bal13]. packing [CK17, LSDD13, YM16]. Pages [Ano19-43, Ano19-38, Ano19-30, Ano19-34, Ano10-39, Ano10-42, Ano10-37, Ano10-45, Ano10-38, Ano10-43, Ano10-41, Ano10-46, Ano10-28, Ano10-35, Ano10-47, Ano10-31, Ano10-30, Ano10-48, Ano10z, Ano10-33, Ano10-27, Ano10-34, Ano10y, Ano10-32, Ano10-40, Ano10-44, Ano10-29, Ano10-36, Ano11-42, Ano11z, Ano11y, Ano11-44, Ano11-27, Ano11-31, Ano11-39, Ano11-43, Ano11-35, Ano11-48, Ano11-32, Ano11-37, Ano11-40, Ano11-33, Ano11-41, Ano11-45, Ano11-47, Ano11-30, Ano11-28, Ano11-46, Ano11-29, Ano11-36, Ano11-34, Ano11-38, Ano12-33, Ano12-28, Ano12-40, Ano12-42, Ano12-30, Ano12-36, Ano12-38, Ano12z, Ano12-43, Ano12-47, Ano12-32, Ano12-45, Ano12-27, Ano12y, Ano12-37, Ano12-46, Ano12-35, Ano12-34, Ano12-41, Ano12-44, Ano12-31, Ano12-29, Ano12-39, Ano13-47, Ano13-44, Ano13-43, Ano13-45, Ano13-31, Ano13-37, Ano13-32, Ano13-29, Ano13-41, Ano13-42, Ano13-46, Ano13-28, Ano13y]. Pages [Ano13-33, Ano13z, Ano13-30, Ano13-27, Ano13-38, Ano13-34, Ano13-35, Ano13-40, Ano13-39, Ano13-36, Ano14z, Ano14-35, Ano14-32, Ano14-39, Ano14-44, Ano14y, Ano14-36, Ano14-42, Ano14-31, Ano14-29, Ano14-46, Ano14-27, Ano14-34, Ano14-45, Ano14-28, Ano14-30, Ano14-33, Ano14-43, Ano14-40, Ano14-38, Ano14-47, Ano14-37, Ano14-41, Ano15-38, Ano15-34, Ano15z, Ano15-46, Ano15-37, Ano15-40, Ano15-28, Ano15-41, Ano15-43, Ano15-44, Ano15-36, Ano15-30, Ano15-27, Ano15-31, Ano15y, Ano15-39, Ano15-29, Ano15-32, Ano15-35, Ano15-42, Ano15-33, Ano15-45, Ano16-43, Ano16-42, Ano16-32, Ano16-40, Ano16-35, Ano16y, Ano16-31, Ano16z, Ano16-29, Ano16-45, Ano16-36, Ano16-34, Ano16-30, Ano16-27, Ano16-28, Ano16-41, Ano16-39, Ano16-38, Ano16-44, Ano16-33, Ano16-37, Ano16-47, Ano16-46, Ano17-28, Ano17-42, Ano17-33, Ano17-45, Ano17-37, Ano17-39, Ano17-32, Ano17-35, Ano17-44, Ano17-46]. Pages [Ano17-43, Ano17z, Ano17-27, Ano17-36, Ano17-29, Ano17-41, Ano17-34, Ano17-30, Ano17y, Ano17-40, Ano17-47, Ano17-31, Ano17-38, Ano18-33, Ano18-36, Ano18y, Ano18z, Ano18-39, Ano18-34, Ano18-35, Ano18-40, Ano18-30, Ano18-44, Ano18-41, Ano18-38, Ano18-32, Ano18-27, Ano18-29, Ano18-48, Ano18-28, Ano18-37, Ano18-43, Ano18-42, Ano18-46, Ano18-45, Ano18-47, Ano18-31, Ano19-49, Ano19-36, Ano19-39, Ano19-29, Ano19-45, Ano19-37, Ano19-42, Ano19-44, Ano19-31, Ano19-32, Ano19-33, Ano19-47, Ano19-40, Ano19-35, Ano19-46, Ano19-41, Ano19-48]. Pair [HM11b, AGB^{+}14, DBG18, For10, GZ19, LF10, LLH^{+}13, MBC^{+}12a, PJ18, RB14, Tri10, ZKTK10, dSMP^{+}11]. pair- [dSMP^{+}11]. pair-correlation [AGB^{+}14]. paired [Fan11]. pairing [GZT12, Koe15, ZDG^{+}10]. pairs [JCG15, JCG16, KDST15, Tun13, ZZC14]. pairwise [HS11, LJS14, MEJ18, SENS^{+}17, SDRA^{+}15, SBK16, WS10]. palate [WHWZ18]. palisade [CRC15]. PAM [WMD16]. pancreas [TSB^{+}17]. pancreatic [JDSPK15, LX16, LXLF14, MVGGB18, MB18b]. pandemic [BGL10, LCCC10, MDE11, PGGvL^{+}19, TGZF11, TDE12, YTGW16]. Pandey [Pan19]. panspermia [Di 10]. PAO1 [ALM^{+}19]. paper [BMdA17, BD10b, Mob10, NIiT18b, NiTI18, NIiT19, Pan19]. Papilio [SFT14]. papillomavirus [MIJ16, MMB14]. papillomaviruses [EMM10]. ParA [SK12a]. Paracoccus [GGGM16]. paracrine [VLPH17]. paradigm [Wal10]. paradigms [AOM19]. paradox [EH17, JTSG14, KMMS17, KWO15, MH17, NT10, VB19b]. Paradoxical [CDK11, SBB18, NIiT18b]. parallel [CJ12a, ZDG^{+}10]. parallel/antiparallel [ZDG^{+}10]. Paramecium [OS11]. Parameter [BP18, DJD10, DDTL19, JB11, LS10b, RD14, Wal12, WcW14, CCS^{+}16, DAVS11, DSLR18, Dol16, GZFX14, KLHS17, LCSH14, LZA^{+}19, LAPK14, MC19, SQZZ17, VBV10, vTMW15]. parameterising [KBF18]. parameterizing [KLM14]. parameters [Arg12, Bal10, BMI19, BMB^{+}18b, BE14, BM16, CL11, CMN^{+}11, EzMMH18, FKK14, JH12, KMM18, KK17b, Lei09, Lei10, Li11, MJV16, PHPM18, PDG12, PMD^{+}18, PP18, PP19, PMKM10, RAAS15, SSD11a, SSV^{+}15, SSS15, VZ19, WST15, ZZ14b]. parametric [CE17, Jab15]. parametrization [ISZ18]. parapatric [YI16, YI17]. Parasite [MB15, RFME^{+}12, VB19b, AIB^{+}19, BYM^{+}18, CR14, CDO17, CKZ^{+}17, CSBK15, CB10a, EBSW17, GCG14, MFG^{+}13, MNH^{+}12, NT14, PEZ^{+}13, TEY16, VMZK^{+}19, XSMF12]. parasite-host [CB10a]. parasite-induced [MNH^{+}12]. Parasitella [EBSW17]. parasites [BPFR16, BM14, CCW16, WW19, WKB13]. parasitic [EBSW17, YHI14]. parasitica [EBSW17]. Parasitism [PML18, BWB11, FB18b, LZTD18, WW19]. parasitism-competition [WW19]. parasitism-mutualism-predation [LZTD18]. parasitoid [HYQ^{+}16, PSL^{+}10]. parasitoids [DPvBvA12]. parathyroid [CLA^{+}16]. paratuberculosis [AMSSG16, AHMA^{+}19, BVJE17, CRH^{+}15]. parenchymal [HWMT17]. Parent [CB17a, BD10b, Déb17, For14, MHKA16, Sak10]. parent-offspring [Déb17, For14, MHKA16, Sak10]. parentage [KSKS13]. Parental [SRV11, ER12, MG17, Ram10]. Pareto [LLP^{+}19]. Parkinson [ADS^{+}19, HSG^{+}18, HDZ^{+}19, KKD18, MvAKR17, STJG12]. Parkinsonian [FW15]. parotid [SMZ^{+}17]. parsimony [BFLS17]. Parsing [FIS11, FIS16a]. Part [HWMT17, AOM19, Laz13, Mog15, FH10a, FH10b]. parthenogenesis [SSD11b]. Partial [Kři10, MT14, WP17, WA14, CMD^{+}10, DMO^{+}17, HP12, Lun13, MK14b, RNVP10, SC10c, Zad11]. partially [Mar19]. Participation [PPBT11]. Particle [WMD16, BF16, DGM15, FCS18, GSN^{+}11, KA15, MMAS13, MN12, PP18, PP19, SF15, YYC19, ZZN^{+}19]. particles [BHP17, KMM18, MAR^{+}17c, TLPH11]. particulate [HWGT15b, HWGT17]. Partite [LCL14]. partition [LLW^{+}18]. Partitioning [SM17b, CPH^{+}14, CFR^{+}14, CMR^{+}18, DB11, Gre16, GSF13, LPTR14, SSvdM10, SA14, ZMNd^{+}10]. partitioning-limited [SSvdM10]. Partner [CS14a, Tak17b, Ezo19, GK10, YHZ14]. partners [Arc11, ZGS^{+}10]. partnership [HHA^{+}13]. partnerships [KK12a]. parts [Nij11]. party [JR17b]. Parvovirus [KEHK17]. passenger [BST14]. passerine [GZT12, Hur12]. passing [OM10]. Passive [ST17b, CLP11, Don13, GAL11b]. Past [Kir11, BM12, MPNP12]. Pasteurella [GTS15]. pastoris [MÇ15a, MÇ16b]. Patch [PMW12, BLV18, DGJ14, GPT17, GKW16, HS14c, KA11, Kři10, LCC15, LBZ18, QW12, SSG^{+}19, SQZ^{+}16, TWW19, Wan19, Wei17]. patch-based [QW12]. patch-type [GKW16, Wei17]. patches [AR11, CT18a, FRH15, LTHEK12, NG11, RPD14, TD18, YiTS19]. patchily [RPCW18]. Patchy [SRCF11, LNH13, TD17]. patella [Cle18b]. paternal [AOR17, Sal15, YTK10]. Path [Ben14, VM12, Che14, CV15, CR16, NWZ15, RSBB10, VC10, vTMW15]. path-wise [NWZ15]. Pathogen [CT18a, MGC13, AH12b, Bon10, CTL^{+}15, FIJJ11, GVSLG16, HBEW12, JB18b, KGB13, LMCW18, MSP19, MB12b, OMO13, PM11a, PB18, RDTHSM13, ST17a, WMD16, WK18a]. pathogen-receptors [AH12b]. pathogen-specific [CTL^{+}15]. pathogenesis [Ana11, JO13]. Pathogenic [Sel11, BTW11, HII^{+}19, LLC15]. pathogenicity [LPS^{+}14]. pathogens [BL15b, CPV16, CG11, DB11, DL15a, DDSDW13, FE10, FE11, GBM18, JB18b, LvdBP12, O’F11, PSL^{+}10]. pathologic [SHK14]. pathological [BBI14, ML12b, RNN15, Wil13]. pathologies [Wal18]. pathology [GRR^{+}14]. paths [BMV12, Che14, NIiT19, Sch19b, VPFV13]. Pathway [MJC^{+}11, OBAI11, WAC14, ÁTC^{+}14, BP18, BB11a, BSM^{+}14, CCS^{+}16, DGW^{+}18, GSS19, lGLLL14, HKS16, LCGMH12, MHD18, NS18, PCS^{+}18, RFME^{+}12, TC11, VHM^{+}17, VLPH17, WPH^{+}12, YST14]. pathways [AP17, BKL^{+}15, Coh15a, EG10, ERT13, Gün12, Gün13, HAP^{+}16, HHG14, HgLL^{+}10, KFL12, PH13, PS13, RW10, STX^{+}11, ZXC^{+}10, ZXS^{+}19]. Patient [METC12, HM13, TDHC^{+}18]. Patient-calibrated [METC12]. patient-specific [HM13, TDHC^{+}18]. patients [Dim10, HA15b, LX16, PGF11, STA15, WR14, WML^{+}17, WHS^{+}13, WLL^{+}14, ZXS^{+}19]. pats [BBR12]. Pattern [FK13, GCB17, GRB^{+}13, WH11, ASP13, BFM10, CDC18, FKMP10, GKMC10, Gal10, HT19, Hir16, iITIM15, JB19b, Kon17, Lee16a, LWL^{+}11, MN14, Mit17, MKJS13, PHH18, PDNP16, RBSD10, SCsMzX18, SM16, SCS14, TGB^{+}18, Tri10, VY12, WZ18, YM16, YKO^{+}16, ZSO16]. patterned [Sie18]. patterning [Dun11, GBN14, HWMT17, WdVMD15, Wol11]. Patterns [MAF^{+}19, AL17, ASL^{+}18, Ayd18, BV12, BS19b, BB11b, BKC^{+}16, CFT11, CDM^{+}14, DBB14a, DTGC14, Don17, DZE11, EJK16, GZY12, HTM15, HDNB10, HgLL^{+}10, JB18a, KDST15, Kri18, LSMC18, LLLV11, LBB^{+}13, LG18, LH14, LTS^{+}16, LRA^{+}13, MSZW19, MPM10, MGCL13, MPS16, MCPF12, MALAN17, MK14c, MROS15, MPLK14, MBRRI19, ML11b, NTY16, Nos14, PS13, PM10a, QW11, Ram10, RV16, RRRTR10, SGKCP19, SWD16, SS15a, SBMH10, Sht17, SAM17, SRCF11, Syi14, TKK14, TKKE19, TA13, TC12, VLCT19, WZWY11, YvBS18, ZKP15, ZML19]. pause [RA17]. pause-and-go [RA17]. Pavlovian [PM10a]. payoff [Ayd18, MI19, MB16]. PBMDR [YYC19]. PBPK [MAFB18, BTF19]. PBSA [ZZ18]. pCa [RG17]. PCD [LJM15]. PCD-GED [LJM15]. PCOS [HS14a]. PCR [IDM15]. PCR-Shrinking [IDM15]. PDC [XXD^{+}17]. PDC-SGB [XXD^{+}17]. PDCD5 [ZSCL16a, ZSCL16b]. PDE [DTP19, KNA^{+}18, TAB^{+}13]. PDE4B [TA16]. PDGFR [YCB16]. pDHS [ZZC^{+}17]. pDHS-SVM [ZZC^{+}17]. peak [FLA^{+}19, LCG^{+}15, SPA18]. peaked [PTNKT11]. peaks [CGB13]. PECM [ZSZM14]. pectinate [SSS13]. peculiar [LLY15]. pediatric [BB10]. pedigree [CRLH^{+}19]. pedigrees [CMPS17]. PEG [VZB^{+}15]. PEG-IFN [VZB^{+}15]. pelagic [JB18a]. Pella [BR14]. pellucida [KG12]. pelotons [Tre19]. pendulum [BRB18, SNCM12]. penetrance [BLNR15]. penetrating [BC19]. penetration [EPJ^{+}09, EPJ^{+}11]. penguins [WHHS15]. penile [FKMG15]. pennation [RW15a]. pentacoded [Sel15a]. people [IL13, PBSM19, RCL16]. peptide [HWW^{+}14, NTC^{+}11, NF16, NGJ^{+}14, SV14, SAGC12, TWR^{+}18, WHL13]. peptides [BC19, BZJP18, BYY^{+}11, CL13a, HPB^{+}14, HME12, JB19a, LL13a, LYM^{+}19, MÇ15a, MÇ16b, PDC^{+}17, PPF17, Sel15a, Sel17, Tox17]. Percent [KB11]. Perception [MBBD13, ACT12, JP19, PLHL19, SP11, TCB13]. perception-action [ACT12]. perceptions [FWW18, MMH10]. Perceptual [BR12a]. Percolation [BGL10]. perennial [MK14c]. perfectly [ABIM10]. Performance [Chu18, AK11, CA17, GMMGVD12, HWTN15, Mar19, MHSH12, NBL10, OO18, PBYB18, RTR^{+}13, RRL^{+}13, Tal12]. perfused [OABI12]. perfusion [CXWL11, DAA17, MAM16, SEK^{+}10]. pergandei [PRMH14]. peri [PVGA12]. peri-implant [PVGA12]. pericarp [AKdV^{+}14, IST11]. period [BL14, BHLH12, CP14, GDF18, GMS^{+}13, HTK^{+}18, JMK^{+}17, KN11, MK18a, NI11, SPH12, Tac10, TC12]. period-average [GDF18]. period-memorizing [Tac10]. Periodic [LHFM16, MGM13a, DGJ15, GZY12, HA15b, Hou16, LBZ18, MB18a, Ros10b, WZWY11]. periodically [GP12]. Periodicity [RLBCJ13, CK16, GP11b, US12]. periods [BAGG14]. peripheral [AOM19, HII^{+}19, NAK^{+}11]. peristalsis [ISM^{+}11, VF12, dLMV^{+}10].

Peristaltic [PCT11, PCT19, Pan19, BJWL15]. perivascular [WO11]. permeability [ASH15, BBR10, DK13b, LDB^{+}14, OTTF11]. permeable [LSMP14]. permutations [GRB^{+}13]. permuted [MP13b]. peroxidase [KIH16]. peroxidases [BMN16]. peroxide [Kun16]. Persistence [LW10a, LW10b, Zha15, AF10, BTG^{+}15, BJPP12, GB12, GRBL13, JA13, KJ11a, Kur17, LLH^{+}13, LM14, LS11b, MBC^{+}12a, MB15, yTRSC13, PRC15, SBM^{+}12, TSS16, UD10, VAT18, ZSH^{+}16, JH11]. persistence-time [SBM^{+}12]. Persistent [CML10, LG10b, Zuc14, BBDB13, Che10]. persists [SN11]. person [NOS17, Pła10, SPSP12, TP10a]. personal [Tay13]. personality [Del12, Pła10]. Personalized [BS15a, BSR14, GdGL17, KCSB14]. perspective [ABGM11, ABD^{+}15, BK19, BRP^{+}18, CHL^{+}19, DKR16, DMSW10, KGM15b, LXS15, NN18, PJL14, RAF^{+}14, SPG^{+}18, SY15, WRZ^{+}12, WWB14, dlEBRM15]. perspectives [DMO^{+}17, KK13, LA15, Sek12, Wal16]. Perturbation [AABS16, GAL11b, MDEKH13, MZMM^{+}14, Ou19]. perturbations [ABLH18, KC14, MEJ18]. pertussis [FGH^{+}14]. pest [ACH18, ABR19, KGJ12, ROF17, TTC10]. pesticide [HT18, KGJ12, TTC10]. pests [IH17]. PET [Mun12]. petiole [GRCdL16]. Petri [YGMT12]. PGE2 [PDRC16]. PGlcS [ZNC^{+}15]. pH [LL17b, AHW13, GT11a, IvLP12, KHS13, KDFM14, LL17b, SLC12]. phage [ARZ15, JCW13, LW17, Voh17]. phagocyte [Kun16]. phagocytic [MMFK10]. phagocytosis [SZ18]. phantom [WTZ^{+}11]. pharmacodynamic [ADR^{+}11b, CvBF18, VZB^{+}15]. Pharmacokinetic [CvBF18, ADR^{+}11b, BSC^{+}18, MAFB18, VZB^{+}15]. pharmacokinetic-pharmacodynamic [ADR^{+}11b]. Pharmacokinetic/pharmacodynamic [CvBF18]. pharmacokinetics [ACR^{+}17, BPLM12]. pharmacological [OA15, POK^{+}12, RG17, SBB14]. pharmacology [LBMG16]. Pharmacophore [TA16, GBC^{+}16]. pharmacophore-based [GBC^{+}16]. pharyngeal [dLMV^{+}10]. Phase [DMM^{+}14, GP11b, JP13, PJ13, Zha10b, AUE15, Ber14, Chu17, CSDR15, DFG^{+}18a, DGJ15, GCSP17, KMM18, LG13, LV17, MH13, OFT15, Opr10, PM10a, SM14, SSH^{+}19, Tou14, VLCM10, WBMM18, WTZ^{+}11, YMW12]. phase-associated [Ber14]. Phase-delayed [JP13, PJ13]. phase-locked [Opr10]. phase-locking [LV17]. phasic [GK17, TD17]. phenological [APBS15]. phenomena [GGG12, ISM^{+}11, Pai19, PA15]. Phenomenological [BSV^{+}10, LPGTBMA19, BCGD10, BSS11, MAS15a, SKAG18, WMPF^{+}15]. Phenomenological-Based [LPGTBMA19]. phenomenology [Oka11]. phenomenon [DKS15]. Phenotype [LP12a, VGPS18, WFW19, BJJR10, BBS18a, Bec14, Das13, Fil15, GCM17b, Gri11, HGM^{+}16, IMA14, IMA16, IMD16, Kit10, LCDH15, SSP15, TK10c, WDN^{+}19]. Phenotype-centric [VGPS18]. Phenotype-genotype [LP12a]. phenotype-structured [LCDH15]. phenotypes [FBK11, FWLW11, HK14, NvD17, RBR18, TDHC^{+}18, WL12a, WL12b, WLL^{+}14]. Phenotypic [PVB^{+}10, SSF15, UJLG14, ACM16, ABV19, BCGD10, BLP10, CRLH^{+}19, CWF^{+}16, Fra19b, HLTR19, KDFM14, LVLC18, Mou12a, NWZ15, NI19, RD17, SA13, SA14, ZWW14, ZMC^{+}18, ZMW10]. phenotypic- [CRLH^{+}19]. phenotypical [DXW^{+}16]. phenotyping [DKP^{+}18, KOF^{+}14]. phenylalanine [AEE19, CFR^{+}14, CMR^{+}18]. phenylhydrazine [AGC18]. phenylhydrazine-induced [AGC18]. phenylpyrimidine [TA16]. pheromone [EBSW17]. pheromones [Amo15]. Phi [MIH16]. Philip [Ano19-50]. Philippines [KdlRJ18]. phloem [JBSFB12]. Phoca [SH17]. Phogly [XDD^{+}15]. Phogly-PseAAC [XDD^{+}15]. Phosphatase [DS15b]. phosphate [JSWY19, TP14]. phosphoglycerylation [JCG16, XDD^{+}15]. Phospholipid [ZWvG10]. phospholipidosis [NS19]. phosphoprotein [KMLT14, KPLT14]. phosphorelays [KFW12]. phosphoribosylamine [Koe15]. phosphoribosylnicotinate [Koe15]. Phosphoribosylphosphate [Koe15]. phosphorus [BBS10]. phosphorylated [ZZ18]. Phosphorylation [HKH18, FKB13, HSiI^{+}19, KFL12, RGG12, STNT17]. Phosphorylation-induced [HKH18]. photoacclimation [GCSMMG^{+}12, NHS^{+}16]. photodynamic [LMM18]. photoinhibition [NHS^{+}16]. photomorphogenesis [PRH^{+}11]. photoperiodic [DFT^{+}17]. photoprotection [IK15]. photoreceptor [CVH^{+}10, CW13, CPW16]. photoreceptors [Meh17]. photosynthesis [GCSMMG^{+}12, LJM^{+}17, YS14]. photosynthetic [DMS^{+}16, PKL12]. photosystem [Laz13, MTE15]. phreatophytic [TLR14]. phyllotaxis [GRB^{+}13, Oka11, Rei12]. Phylogenetic [Bar16, CTA15, HYW11, LAS14, SC10c, BPM^{+}12, BLS17a, BFLS17, Cut15, DDLW10, DLYW13, FM18, GK16, GD16, HS11, Hua16, HMW16, LS13a, MS12a, SMD^{+}16, SS12a, Ste13, SPCM18, TZY18, WS10, YY15]. phylogenetics [AHR10]. phylogenies [Kae17, Par15]. phylogeny [JPDP10, SL17, YHY14, dF13]. Physarum [BMV12, Tac10]. Physical [OB15, Oka11, SNY^{+}13, WST15, Zab11, dSRM15]. Physically [HB12a]. physico [Di 16, Di 17b, SSV^{+}15]. physico-chemical [Di 16, Di 17b, SSV^{+}15]. Physicochemical [SBT^{+}18, CL13a, CT18b, FD18, HYA14, JLX^{+}15, JD16a, JD17, RSD^{+}16, STG19, WXC10, XSLZ16]. physics [LCBO^{+}12]. Physiological [DPvBvA12, KKU^{+}10, MFN^{+}18, Ush16, WCHC11, CTW10, CB13b, DSLR18, DLMK12, DCN13, EzMMH18, Gri11, LCA^{+}15, LHD^{+}17, PRK13, RG17, SS17b]. Physiologically [KAPR12, ACR^{+}17, FPR10, MAFB18, NSO15, dlRVFK^{+}16]. physiologically-based [NSO15]. Physiology [Mar17a, ARD14, BLW11, PRC15]. phytobenthic [SM17b]. phytodesalination [RAAS15]. Phytoplankton [BSH19, CFSB16, CFL^{+}15, DD10, Gro17, KSP12, MHC16, MYLK11, RRB10]. PI [MZAI19]. Pichia [MÇ15a, MÇ16b]. Pick [FtL18]. picture [Goe15]. piece [SPH12]. Piecewise [TXWW12, TXCW15, COG19]. Piecing [SSP15]. Piezoelectricity [FGAM12]. piezophilic [NS16]. pig [LT14, SMM^{+}13]. pigeon [PW11]. pigeons [SBW11]. pigment [DFG^{+}18b]. pigmentosa [CW13, CPW16, RGL^{+}17]. pillars [Pia19]. pilot [SAI14]. Pimephales [HCHI18]. pin [HTM15]. Pina [ASRM15]. Pina-D1 [ASRM15]. pine [APS^{+}13, APBS15, CRLH^{+}19, LK15]. Pink [ZLL^{+}12, KHRS10, Toz15]. pinnatifida [MJV16]. pinning [KM19]. pipiens [YMZ18]. piscine [HH14]. pistillata [KRSD11]. Pisum [NNK^{+}15]. PKPD [ADR^{+}11b]. place [HY16]. placement [DI13]. placenta [SSFG15, SSB^{+}15]. placental [GSGV11, LMJ^{+}16, PWC^{+}15]. plain [GW12]. planar [YM16]. plane [MSSM18]. Planes [NHTS14]. planform [NC16]. planktic [DLSD15]. plankton [CMBP19, PSV17, PHPM18, PMSY17, RM10, Ric17]. plankton-oxygen [PSV17]. planktonic [CG10]. planning [KBV11]. plans [KCSB14]. Plant [KR19, LRH15, AEE19, AMRRI10, BDLR12, CMB^{+}12, CPV16, Che16, CEP14, CG11, DFT^{+}17, lDsWM^{+}15, DBJ12, FCC^{+}10, FM13, FM15b, GGBW14, GWBG14, GZT15, GRRG16, HAP^{+}16, HTM15, HB13a, IK15, IC11, KA11, KOF^{+}14, KHNM16, KFG^{+}14, KFS^{+}13, LKSM14, LWRE14, LvdBP12, LH18, Mei12b, Mei13b, ML10, MK14c, NKB16a, OBE^{+}17, OHWS18, PCB14, Pie10, PRH^{+}11, PPTC19, RK18, ST17a, Sat10, SLC12, SIHO15, TIS10, TAL19, Yam16b, ZZC^{+}17, vVE15]. plant-insect [KA11]. plant-pollinator [RK18]. plant-predator [HB13a]. plant-virus [NKB16a]. plantarum [LDdA^{+}13]. planting [CK18]. plants [AGGME18, BHSB11, BvLH14, CMB^{+}12, DHT19, DP13b, Kha10, KBKK14, LRH15, NKB16b, SI12, SNOV11, WZ18, YLY15]. plaque [GCYL18, IvLP12]. plaques [BD12, FBM19]. Plasma [RFdL15, BAGG14, CKD11, CRV16, SHB^{+}17]. plasmid [FS11b, GY18, ZKTK10, ZDF^{+}12]. plasmids [BBT^{+}15, Mic15b]. plasminogen [Sil16]. Plasmodium [CB13a, CDO17, CKZ^{+}17, CMD^{+}10, GDPPSS^{+}11, KMLT14, KPLT14, RM19, SPMGR10, TEY10, TEW12, TAR16, Tac10]. plastic [HD11, PAV10]. plasticity [CWF^{+}16, DAA17, FHR13, FR13b, HKS16, HNA15, KAZ13, Mou12a, Rob11, SSF15, SA13, SA14, UJLG14, ZWW14, ZMC^{+}18]. plastids [Sat17]. plate [ABD^{+}11, CAGM^{+}17, FHLM10]. platelets [SPA18]. plausible [FPdRdH10, SBMH10]. play [SPS11]. played [dOGL10, dOGL13]. player [De 19, DHK13, GT11b, HS14b, KC17, Kur19, Lai18, PLN14, Shu13, TT19, vVN12]. players [vV11a]. playground [SD16]. plays [Lai18, RPGG^{+}19]. pleiotropy [Gia13]. Plk1 [SBR16]. pLoc bal [CXC18]. pLoc bal-mGneg [CXC18]. ploidy [YG12]. plot [Mun12]. plus [BBRT12, SBT^{+}18]. pneumococcal [MSE^{+}14, SMA11]. pneumonectomy [HWM^{+}16, HHMT19]. pneumonia [MSE^{+}14]. Pocketknife [Sel14a]. Podar [Di 14b]. podosomes [HBW^{+}11]. point [AP13b, HT19, KDST15, MKE15, MS14c, PP17a, SSN^{+}14, TH17, YI17]. points [AIY16, DI10a, KMN16, SSJK18]. Poisson [GBIR13]. Poland [GMR16]. polar [DGL12]. polarisation [CEKM^{+}19, VSW10, dBE16]. polarised [JD13]. polarity [LS15b, Lee16b, YM16]. polarization [FZ14, SKGM19]. policies [KTJ19, LJ15, PGGvL^{+}19]. policing [OAV^{+}16]. policy [GK13, ZXH19]. polio [KR15]. poliovirus [Hou15]. polioviruses [KR15]. pollen [CKD11, YLH12]. Pollination [Wan19, EVRE14, LIPD12]. Pollination-mutualisms [Wan19]. Pollinator [FH13a, TIS10, EVRE14, RK18]. pollinator-flower [Abb10]. pollinators [YLH12]. polluted [LW10a, LW10b, LLL13]. poly [CDK11, SABB15]. poly-specific [CDK11]. polyadenylation [JWS^{+}10]. polydomous [CFR13]. polygamous [Arc13]. polygenic [HT18]. polyglycine [Di 14a]. polyhedra [Den19]. polyhedral [CZJQ14]. polyhydroxybutyrate [SBJ^{+}18]. polyketide [BLP10]. polymer [Fan11]. Polymerase [BS19a, Das18, IDM15, LYF^{+}15, SBJ^{+}18]. polymerase-template [IDM15]. polymerases [Xie11]. Polymerization [SGGHS19, Sel12a]. polymers [SG18]. Polymorphic [Fis16b, AY15, BRA15]. polymorphism [ACM16, BCF^{+}16a, Bon12, DEG^{+}14, IOY15, Kis16, LBJ13, OO17b, SMD^{+}16]. polymorphism-aware [SMD^{+}16]. Polymorphisms [Ram10, LR16, MV10, OGO19, ZW16]. polynomial [DZJR10]. polynomials [MMG^{+}16, Mil11]. polypeptide [PBKR13, RA10]. polypeptides [MNSZ16]. polyphyletic [Di 13b]. polyploid [PPTC19]. polysome [SC11a]. polytes [SFT14]. polytomy [SL17]. polytopic [EG12]. pool [LJI17, SHBT14, TLW18]. pooled [BR13]. pooled-equilibrium [BR13]. poor [BID15, CT14b, PT10, YSST13]. poorly [LS15c]. Population [CM14, Kam11, SI19, SBSE14, TWW19, WZLW15, AR11, AY15, AKS^{+}19, ADC19, Arg12, AR17, ASGG17, BTG^{+}15, BBR10, BBWS^{+}13, BSKV18, BHKR14, Bon10, BHBR12, BC13, BPGS12a, BPGS12b, CC11a, CMS^{+}19, CFT11, CA18, DB10, DMM^{+}14, EKFV18, EBE14, EPP19, EVRE14, ECP^{+}16, FGLS10, FR17, FH13b, GGQ^{+}12, GW15, GVGC15, GAGP^{+}14, GA16, GS13b, Gre15, GP12, GBRS19, GMR16, HN19, HSMCCR19, HR14, HM11b, Hor13, HTB^{+}13, HPW^{+}13, JJ17, Jam16a, JL16, JP19, JMK^{+}17, KMHdlP10, KFB18, KD16, KG15, KAKK19, KSPA17, LS15a, LR16, LG10a, LHV16, LFM11, LL18, LP12b, MZ17, ML14, MB18a, MSK10, MZAI19, MCPF12, MD16, MMRCCC10, MA11b, MBP16, NKL10, NSO15, NBBCC18, NGS^{+}16, NP13, PBR17a, PVB^{+}10, PGH11, PKH11, Ram10, RJSC18, RWSB11, RRR15, RR12, ST17a, Sal15]. population [SI12, Sch14, SFT14, SCC11, SV17, SFS13, SZ10, SL12, SY15, SYY17a, SYY17b, ŚKSRW16, TKB^{+}18, TLCZ12, TGC^{+}17, TDKJ15, TMD14, VSD^{+}17, VZ19, VK10, VAT18, VZB^{+}15, WR14, WML16, WKB13, WMHH10, XSCS12, YMZ18, ZKTK10, dTKDV15, vV11a]. Population-level [SBSE14, GBRS19, SCC11]. populations [AGPK13, ATRR10, BGM19, BV12, BGKL17, BI19, CZN12, CMMR13, DBBW09, DBBW11, Déb17, DKR16, DBG18, FHSG15, FZW17, Fis19, FBM19, FS16c, FNH10, FOT^{+}15, GvVT14, GSYS10, GGBW14, GWBG14, GH11, Gra15, GPT17, Han17, HAB11, Has14, HI12, HMWB13, HH16, HHD^{+}16, HMM17, HK11b, HHN14, ITR^{+}18, JMS18, JSB15, Jus11, KCS16, KD16, KFA12, KEHK17, KS15, KUV^{+}10, Lab16, LS16b, LW15, LGBT15, Lio18, Lio24, LCDH15, Mac14, MR11a, MKKS14, MH12, MKBE17a, MKBE17b, MBS19, MP13a, MBC^{+}12a, MG10a, MZMM^{+}14, MSS10, Moo14, MP13c, NI19, Oht12, Pat16, PNL15, PP17b, PKH11, PTB12, RN12, Ros11, SSD13, SPSP12, SMdA16, SGW17, SC12, SM11a, Tay16b, TB18, TRJD19, WL12a, WL12b, WL14, WGS10, WSP14, WL19, WBM^{+}15, YG12, ZADB15, ZDF^{+}12, ZLW18, dASdC^{+}15, dWV11, dlCGSA16]. porcine [SMM^{+}13]. pore [OTTF11]. pores [MAA^{+}17]. poroelastic [DHRA10, XLLF17]. porosity [MAMEA15]. porous [CFT11, MBPS17]. poroviscoelastic [MA11a]. poroviscoelasticity [MAMEA15]. Position [KHK^{+}17, AKR^{+}18, DLL^{+}17, GGC14, QW11, QLC^{+}18, SLDP13, WYX^{+}17, XDD^{+}15]. position-specific [WYX^{+}17, XDD^{+}15]. Positional [Wol11, Che14, JD16a, JD17, KRSD11]. positional-specific [JD17]. Positioning [BAM^{+}15, Lee16b, JYZ18, NDG14, Sht17]. positions [KB10a, LL13a, NHTS14]. Positive [Eli17, FGPR10, KF11, LCGMH12, Ric19, Sel10, SXZY12, AN13, BC19, DD13, DHS^{+}15, HLS^{+}12, HR12, JRMS12, LKP^{+}12, RA18, SRD^{+}15, TWH^{+}13, USF^{+}18]. Positively [RAD14]. positron [LDB^{+}14]. possessing [JRR^{+}12]. possibility [IOY15]. Possible [BTK14, GG16, GK13, Han10, MA11b, ÁTC^{+}14, BGCB12, Coh15a, DGJ15, HNR14, HKS15, Kro10, Mal10, SWSLMJ19, YLY15, d’O12]. Post [EBSW17, BLZ^{+}19, GRB^{+}13, MMC19b, RRR15, XG12, ZLB13]. post-marital [MMC19b]. post-meristematic [GRB^{+}13]. post-reproductive [RRR15]. post-spliced [BLZ^{+}19, ZLB13]. Post-translational [EBSW17, XG12]. postbuckling [She10]. posterior [HS11, ZBA18]. postganglionic [BPPC15]. postharvest [vdSS12]. postnatal [Mol18]. postsynaptic [MZ18]. posttranslational [XWW^{+}14]. posture [BBJDS11]. potassium [CRGS18, CW11c, FD11a, GJ11, ZAL^{+}19]. potency [ADR^{+}11b, ONO15, PDC^{+}17]. Potential [BLZ^{+}19, MG14a, PYG^{+}19, APW10, BHBH11, CF11, CVO^{+}15, DSTD19, HN10, HZC^{+}10, IHNS16, KS10, LdLK11, LLD^{+}17, LSB10, LB11, LB12, LL17b, MGGM10a, MV10, Mas18, MN14, NTC^{+}11, NS19, NRS^{+}16, NYSM12, PRK13, PPC^{+}17, RRG^{+}12, RL17, RGCML10, SZ18, SNOV11, TWR^{+}18, TWTA^{+}18, Thu15, TGZF11, Unc11, XSTW14, XW17, YLH12, JSC^{+}16]. potentially [BHBR12, PMSY17]. potentials [AAJGCD15, CS18, FMS^{+}12, HMM16]. potentiation [HKS15, HKS16, SBB14]. potentiators [Mor10a]. Potts [LL14]. Power [KMM16, HB10, HYZ^{+}15, KHRS10, PCT11, PCT19, Pan19, PR17, PWHW16, RS12, VDH^{+}15, ZDY11, ZL18b]. power-law [HB10, PCT11, PCT19, Pan19, PWHW16, ZL18b]. powerful [JZZK11, LWL^{+}11]. pp [JN14]. PPI [LJM15, LCL14, YZ19, ZHC10]. Practical [Ger16, Zhe15]. Practically [NH19]. Practopoiesis [Nik15]. pre [DLMK12, Mau15, Mol18, PVGA12, RT15, SPS11, SHSR15, VDTF15, WSH^{+}10]. pre-breeding [VDTF15]. pre-diagnostic [SHSR15]. pre-existing [PVGA12]. pre-miRNAs [WSH^{+}10]. pre-mRNA [RT15]. pre-play [SPS11]. pre-RNA [Mau15]. pre-stressed [DLMK12]. preadipocyte [CSKF10]. prebiotic [CKS15, ML14, RBRB16, SHH15]. precipitation [WTC16]. Precise [JRR^{+}12, RRG^{+}12, Zhu11]. Precision [NNK^{+}15, TMS13]. preconditioning [SXZY12]. precultivation [ABK^{+}12]. precursor [LFW^{+}15]. Pred [KCM^{+}11]. Predation [MB12b, PDSP13, AF10, HMH14, IEN15, KGB13, LZTD18, PBB10, PFB10, SP11, VA10, Woo10]. Predator [GSL13, SV17, Abb10, BT17, BH13b, BHR10, CLA17, CGHF14, CK18, CNM^{+}13, DS13, DZJR10, DKF17, EKFV18, FtL18, Fre10, GLR^{+}17, GVGC15, GWE11, GPR^{+}16, HAK^{+}19, HGHL10, HSR12, HB13a, HWL15, KGJ12, KB15, Kři11, Kři13, KP15, LHV16, LJ18, MSZW19, MMAS13, MFG^{+}13, MA11b, MI11a, Mou12a, NTOI16, PPW16, PHPM18, SWD16, SBSE14, SVA18, SY12, VWF19, WRC^{+}19, YiTS19, ZZ14b, ZMW10, ZMT11, ZYD15]. predator-2 [PHPM18]. predator-prey [HAK^{+}19]. predators [Alv18, AH17, BT17, BR12a, FKV19, LG19a, LG19b, SvK12, TI13, TB15, ZYD15, vLBJ^{+}13]. predatory [IPMH12]. Predict [CXC18, FL12, HZC^{+}10, JLX^{+}16, ZZK14, pZWZ^{+}16, BMSEE14, BJ18, FGAM12, GMZM15, HYA14, JSF^{+}11, LXC15, MJ11, PH13, PWHW16, RW14b, WYL^{+}18, XNJ^{+}13, ZZG^{+}16]. predictably [HD11]. predicted [HH14, Hal17, HCHI18, KZL14, LJ10, MALAN17, PDC^{+}17, RT19, ZKHL16]. Predicting [ÅCHS19, All11, ADR^{+}11b, CL13a, CT18b, CTL^{+}15, GSCS11, GMM^{+}13, GGX^{+}10, HPB^{+}14, HK11a, HK12, HN10, HS12, HLS^{+}12, HY13, JYZ18, JD16a, JD17, JCG16, LS13a, LD11, LJW^{+}16, MPCTGJ^{+}15, Mei12b, Mei13b, MSB16, NHSX14, OYY16, QLC^{+}18, SM17a, SM18, SCA13, SCF^{+}12, SKK18, TWC^{+}19, WGS10, WS16, XTZ^{+}13, ZL18a, ZRGW19, ZLW^{+}19, ZLDZ13, AAJGCD15, BW13, DLL^{+}17, DTP19, Hua12, KCM^{+}11, LLP^{+}19, LKL13, LHH12, LHL^{+}14, MV10, QLWC19, RBMS17, RW15a, SW13a, SC10b, SOIO10, WXmH10, WHZ^{+}17, YD14, YC13]. Prediction [ARR14, BRK19, CZW^{+}11, CCY^{+}19, GXFF13, HTK14, HWHL15, JD16b, JR17a, JW18, KPK^{+}13, KJH^{+}19, KSKK15, LYF^{+}15, LTL^{+}15, MPY14, MBE11, NBL14, Pav14, QW12, SS17a, SMS19, SHLL11, Tun13, VZB^{+}15, WYL^{+}19, WLY^{+}17, XWW^{+}14, XDD^{+}15, XXD^{+}17, ZMS11, ZDG^{+}10, ZSZM14, ZJSC16, ZD18, ZNC^{+}15, vdSS12, AFS19, CJ12b, CTSL11, Cho11, DSPM14, DYQ^{+}14, Dol19, DTY12, EMM10, FHG15, FYZ15, GMMGVD12, GJ15b, HA15a, HK11b, HW13, IMH15, JM12, JWS^{+}10, KKGN16, KGP^{+}15, KZL14, LWW14, LJ10, LLCC12, LDdA^{+}13, LZMM15, Mar12, MKBE17a, MKBE17b, MKRE18, MSRW16, MFZ18, MZ18, Möl12, MP14b, MMCZM12, MP19, MR18, MR19, NK18, NS19, PWZ^{+}19, PM14, PMKM10, PSAA13, QQJW13, RSS^{+}18, RSD^{+}16, RAF^{+}19, RCL^{+}10, SKK18, SQZ^{+}16, TLY^{+}19, WRZ^{+}12, WYX^{+}17, XLL^{+}10, XWD^{+}10, ZCAB17, ZPHS11, ZLY14, ZKHL16, ZZC^{+}17, ZHC10, dEG11]. Predictions [FKDW15, KHRS10, KSP18, METC12, QZY17, SPA18, TWP16]. predictive [ABB^{+}19, KBE^{+}13, MPC12, NSH^{+}10, Non10]. predictor [IIKH16, JZ17, JCG15, MS14a, USF^{+}18, WMK13, WMK14, WMK15, WMK16]. predictors [PPF17]. predicts [ASC16, Bac15, HBA10, Lai11, MD16, Tsu19]. predominant [KBK16]. predominantly [MR10]. prefer [SS15b]. preference [PC10a, PHPM18, SS12b, ZLZ^{+}16]. preferences [Ant13, BC13, Kři10, KR19, Sal10, ZDG^{+}10]. preferential [FHSG15]. preferentially [HHMT19]. Preferring [SBJ13, Jab15]. Pregnancy [For14, SPO^{+}10]. preliminary [BVK10, LZL^{+}19]. premature [TH17]. prenatal [BFGAGA16]. prenylation [HKR^{+}19]. PrEP [RVZ14]. prepared [PGGvL^{+}19]. preparedness [PGGvL^{+}19]. preprocessing [GZ14a, GZ14b]. Prerequisite [Tox19]. prescription [RD14]. presence [ASK17, Alv18, AWK15, BYJ17, BZA16, BAM^{+}11, CW15c, CDK11, GT12, KM18, KS14, KGM^{+}15a, MOSS15, MB18a, yTRSC13, NM11, RRRRSPTR16, SIVH17, Sil16, TB15, TH17, VIBC12, WXWL12, WdVMD15, ZN18]. present [Kir11, MPP^{+}16a]. Presenting [TTG19]. Preservation [Tox18]. preserve [MFMPS10]. preserving [AdGM12, FK18, WXmH10]. presomitic [JAK19]. pressure [AM11, BGM19, CMS^{+}19, CCG^{+}18, DLMK12, HWMT17, IKHL16, JSK16, KEHK17, KFG^{+}14, LFD11, LS15c, LS16c, NHE^{+}16, PR13, TSL12, WWB14, WFM^{+}13, WFC^{+}14]. pressure-based [CCG^{+}18]. pressure-clamp [KFG^{+}14]. pressure-clamp/relaxation [KFG^{+}14]. pressures [Rey13]. presynaptic [MZ18]. preterm [SEK^{+}10]. PREvaIL [SLT^{+}18]. prevalence [ADB^{+}13, KST17, LZY^{+}15, MHMM18, MS16, MM15a, PDW11, SSG^{+}19, SGD^{+}16, YHZ14]. prevent [RANO10]. preventing [FHSG15, LJI17]. prevention [CJL15, PBSM19]. Prey [AF10, BH13a, Abb10, AK11, BT17, BH13b, BC15, BBB18, BR12a, CLA17, CGHF14, CK18, DS13, DZJR10, EKFV18, FtL18, Fre10, GLR^{+}17, GG14, GSL13, GPR^{+}16, HAK^{+}19, HGHL10, HWL15, IEN15, KB15, Kři11, Kři13, KP15, LS10a, LG19a, LG19b, LHV16, LJ18, MSZW19, MMAS13, MFG^{+}13, MA11b, MI11a, Mou12a, NTOI16, PPW16, PHPM18, RPCW18, SWD16, SV17, SBSE14, SVA18, SY12, Tos11, VP18, VWF19, WRC^{+}19, YiTS19, ZZ14b, ZMW10, ZMT11, vLBJ^{+}13]. prey-predator [BT17, GLR^{+}17, YiTS19]. prey-resource [SY12]. prey-shared [KB15]. prey-subsidy [EKFV18, LHV16, SV17]. prey/pollinator [Abb10]. prey/pollinator-flower [Abb10]. Price [Gra15, vVGSE12, Lio18, Lio24]. prices [HMM17]. Primary [KBL15, ABKS11, ABD^{+}11, DBBW09, DBBW11, Di 18, Di 19c, FHLM10, FHW^{+}10, GGC14, GA13, HWM^{+}16, LRL^{+}16, PRV^{+}14, SDT17, XM11, YZZ12, ZPHS11]. Primate [BHLH12, RGP13, SCF^{+}12]. primates [LPvSP11, LPvSP12, Nak16b, SDPC11]. priming [MEKK11, SW13b]. primitively [LF13]. primordia [PHH18]. Primordial [VFS^{+}15, Mau15]. primordium [ZBA18]. principal [HME12, SG15a]. Principle [Pee12, CHH10, FMS^{+}12, HRC^{+}12, LKZB15, LIB^{+}17, Mig16, NvHM13, WB10]. principles [DRW14, KAN11, KSPA17, MJS11, TM16, VGPS18, dSRM15]. prion [DFMR19, PRM11, Wil13, Zha10a, Zha11, ZGY11, ZSW11, ZZ14a]. prior [SNS17]. priorities [SAB17]. prioritise [CSS^{+}18]. Prioritization [KJ19, FP18, LLW^{+}18, YC15]. Prioritizing [HZL^{+}13]. priors [KM18, MKBE17a, MKBE17b, RM17b]. Prisoner [CZ14, MI19, PM10a, BLS^{+}17b, JC16, TP10a, YA14, ZZC10, ZGWL14, FNH10, GCS12, Lai11, Lai18, MN11, ZLY^{+}17]. prisoners [PCN17]. private [OSN18]. pro [DPvBvA12]. pro-ovigenic [DPvBvA12]. Probabilistic [BDID^{+}12, EAN14, PPTC19, SRS^{+}15, Ala15, CMPS17, FM18, KHK15, KMCH14, OGE10, PDM17, ZGG^{+}19]. probabilities [ADR11a, CS15a, HS11, HT10, JLH15a, JR17b, LPKH15a, SLDP13]. Probability [Mul11, Mul12, AP13a, BBR16, BPR15, CPV16, Cut15, HTK14, Hou15, JJ17, Jam16a, LAG^{+}14, LN19, MPY14, MKBE17a, Mon18, OGE10, SW17, VGL16, Wax11a, YG18]. probable [Sil16, SS15e]. probands [Jam16b]. probed [Zha10a]. probes [KFG^{+}14]. Probing [GNB^{+}13]. Probiotics [RD14]. problem [AKNP18, Arc14, ASP13, BM16, CFS^{+}19, CB15, DSTD19, DZJR10, Fok12, GA15, HB12b, Mil11, MAR^{+}17c, Van16a, VLCM10]. problems [Coh15b, Mor19, MRS14, YBC17]. proboscis [LLK14]. procedure [MW13]. process [AGRR^{+}15, BGKL17, BCFR10, CJKR10, CCB11a, CAV16, DDHM12, DRCL12, DDB10, FCC^{+}10, GFH^{+}18, Gou16, Gri19, Hal12, Höh15, JSWY19, KB15, LPH19, Mal10, MFG^{+}13, NPS10, NPMM15, PP17a, RRRTR10, RRRRSPTR16, SW16, SCÁBM10, SCÁBM11, SGAM12, SSJK18, SBW11, SBJ^{+}18, TGLK19, TA10, TF17, jWlGQ19, WdLS17, WZ15, ZCT18, ZTWL12]. process-based [DDB10]. processed [CRV16]. processes [AGM11, BA10, BMI19, BSKV18, BMB^{+}18a, CK15, Edu16, GGM12, Ing16, JG14, Kuz19, LPKH15a, LPKH15b, LPH19, MRPLAS15, MH18, NYSM12, PAS10, PBB^{+}15a, Par19, RRTR12, RSR11, SS15e, SMC^{+}13, SK18]. processing [BZ18, CC11b, ESW13, HYSG19, RWH16]. processive [GWX17]. processivity [SSZR17]. produce [DBB14a, KK15, Sel13, Sel15a]. produced [LFM11, MDVT10, SK15]. produces [KL11, MOSS15]. producing [FE11, GT11a]. product [DH16, De 17, DDF^{+}14, ZK18]. production [BSL^{+}17, GSL13, GS17b, HWGT15b, JLH^{+}15b, LDF^{+}11a, LvBJ16, MÇ15a, MÇ16b, OFB10, PBR17b, Sak10, STNT17, SPA18, Tsu19, WBM^{+}15, YLY15, ZCT18, vLFM^{+}19]. productivity [FGLS10, SDJ^{+}15, SK16b, YLH12]. products [DDF^{+}14]. Professor [Ano19-50]. profile [AMP12, CW11b, DYQ^{+}14, HTK14, SC11a, SG15b, WDH^{+}16]. profile-profile [WDH^{+}16]. profiles [CCS18, IvLP12, JK12, Pła10, RU16, SBT^{+}18, TK10a, TBB14, VLFF12]. profiling [HZL^{+}13, SOIO10]. progeny [DBBW09, DBBW11]. progesterone [LXZQ14]. progesterone-induced [LXZQ14]. prognosis [LX16]. program [VDTF15]. programmed [HZW^{+}14, Xie13a]. programmes [dMP11]. programming [CZW^{+}11, JPBM17, LPD^{+}16, NS13, SNS17, ZMNd^{+}10]. programs [NKOS11]. progression [CW13, CLHB11, DD13, GL18a, GA14, HTK12, KEG17, KC11a, LTZ19, LCGMH12, METC12, NWB^{+}10, SB19]. Progressive [SVSP15, Len14, SCKL15]. projected [DTY12]. Projecting [MHMM18]. projection [PRSC11, WMK14]. projections [WXmH10]. prokaryote [CCY^{+}19]. prokaryotes [GYWJ10, Hua12, XKCG15]. prokaryotic [AAAV14, LXS15, YCR^{+}15]. proliferation [ALD^{+}11, BJO^{+}16, BMB^{+}18b, GMOP12, Gol10a, Gol10b, Gol16, HPM^{+}17, LD18, MFL^{+}18, MFMB12, MPZK16, TSMB14, WRH^{+}16]. proliferative [SZ18]. Prologue [TK19]. promelas [HCHI18]. promise [Ran12]. Promote [GWNW15, BBD18, GW19, HNR14, JA13, Lai18, PTB12, RRR15, SSS17, SJSK18, Tos11, WDH^{+}11]. promoter [BSB^{+}13, CBL14, FLM18, HA15a, HBSF11, RW16, SDRA^{+}15, dEG11]. promoters [HHJR11, MMH^{+}12, RW16, ZLDZ13]. promotes [BLW11, CS14a, CH11, ML13, MB12b, NCLB16, SN18, SKS^{+}19, WSP14, XY14, YHZ14, ZZCZ17]. promoting [BM14, BLS^{+}17b, LHW14, Mal10]. Promotion [LPH18]. proof [HRC^{+}12]. propagated [HS15]. propagating [Hor11b]. Propagation [NS18, Van17b, BZN17, BSB^{+}13, HM10b, HMM16, KKWA18, OTTF11, PRV^{+}14, ST17a, SNOV11, WML16, WHT17, WLY^{+}17, YM16, ZUFM17, dRFI15]. propagule [BGM19, PR13]. propel [RL15]. propensity [JM12, NPS18, PGH11, XDD^{+}15]. propeptide [NF16]. proper [KP16, SW19]. Properties [SBT^{+}18, ÅCHS19, CS18, CW11a, CL13a, CT18b, DLEMP19, FD18, FLM18, FLGGD^{+}14, GML10, HT19, HYA14, HAuR^{+}18, HZL^{+}11, JBSFB12, JLX^{+}15, JD16a, JD17, KCM^{+}11, LL13a, LCG^{+}15, Luo14, MBE10, MT14, MGC13, PM11a, PW19, PDW11, RSD^{+}16, RHS14a, SFV16, SG15a, SDKF12, WM17, WN19, WXC10, XSLZ16, YWW^{+}14, YWZ^{+}16, ZR16]. property [GRB^{+}13, Js12, Su16]. prophage [CKB19]. prophylaxis [RWSB11]. Proportion [WF17, Dol16, IU18, JMC^{+}10, PWE15]. proportional [Gla13, RWH16]. proportionality [KMM16]. proportions [HXL18b]. proposal [FRP14, SNCM12]. propulsion [FKMG15, HM11a]. propulsive [KG12]. Prorocentrum [JSWY19]. Prosperity [CST^{+}12]. prostate [CBGS18, Dim10, DDTL19, HHSA15, HBA10, HA15b, NH19, PGF11, SSN^{+}14, WMN18, YZY^{+}16]. prosthesis [JH12]. Prot [pZWZ^{+}16]. Protandry [Mor13]. protean [Fre10]. protease [FHG15, IHNS16, TAR16]. proteasome [SBB18]. protected [KG13, KJ15, TM14, Tak16]. protecting [BH12]. protective [HE16]. Protein [DS15b, KKS^{+}14, KS10, LJM15, LCF17, LBS^{+}14, LPD^{+}16, MALAN17, Nah14, NSBL10, QZY17, SV14, SBT^{+}18, SSV^{+}15, TKK14, Wal10, WZM19, XSLZ16, YDC^{+}13, YZ19, ZT16, AFS19, ACdlRMR^{+}11, AEZR^{+}16, AH15, AC15, AP13b, BBT^{+}15, BKR14, BB15, BSM^{+}14, BB13, BGD^{+}13, Boz15, BMF^{+}18, Bry13, CJ12a, CZW^{+}11, CL13a, CSKZ19, Cho11, DLEMP19, DI13, DHS^{+}15, DSPM14, DM10a, DM19, DYQ^{+}14, DZW10, EHBC10, EG12, ER18, FJ10, FLGGD^{+}14, FCS18, GTS15, Gin10, GJ15b, GNB^{+}13, HNV^{+}16, HMTA17, HYA14, HK11a, HK12, HTK14, aHLZ^{+}12, HLHY17, HMH14, HPML18, HLS^{+}12, HH13, HY13, HWHL15, IA17, JSZ12, JM12, JLQ^{+}19, JD16a, JD16b, JD17, JCG15, KSA16, KHP^{+}12, Kee10, KKGN16, KRDJ15, KHKI17, Kin18, KZL14, Kri18, KSK^{+}17, Kur10, KK17b, LL12, LWW14, LZ18, LJ10, LlLYB13, LCL14, LYZ^{+}18, LMJ^{+}18, LSLL13]. protein [MPY14, MJ11, MMY^{+}12, Mal18, MdFDM10, MPCTGJ^{+}15, MGO^{+}15, MÇ15a, MÇ16b, Mei12a, Mei12b, Mei13a, Mei13b, MG14a, MIH16, Miy16, Miy17, MPZK16, MP14b, NBL10, NBL14, NLB14, NSS^{+}11, NMZ19, NHTS14, OYY16, PDM17, PAOMV17, PWH^{+}13, PM14, Pav14, PBvdG10, PWHW16, PM11b, PBBB10, PSAA13, QW12, QQJW13, QLC^{+}18, RRG^{+}10, RSS^{+}18, RSD^{+}16, RCL^{+}10, RM19, RT15, RBMPP^{+}13, RBMPP^{+}15, SNS17, SRS^{+}15, SM17a, SM18, SFV16, SRP16, Sel13, SC11a, SLDP13, SSSD17, STG19, SHLL11, SQS^{+}12, Sht17, SBB18, SG15b, SAGC12, SAZ^{+}14, SPL14, TWR^{+}18, TP10b, TAI^{+}18, TWC^{+}19, THBM10, TP17, USF^{+}18, VBHM^{+}13, WMK14, WMK15, WXmH10, WRZ^{+}12, WYX^{+}17, WYL^{+}19, WLY^{+}17, WXC10, WGDZ18, XLL^{+}10, XNJ^{+}13, XGZ17, YSH^{+}14, YWW^{+}14, YG18, YYH^{+}14, YC13, YCR^{+}15, YLF^{+}17, ZMS11, ZCAB17, ZPHS11, Zha11, ZZRZ11, ZS12, ZTWL12, ZLY14, ZZK14, ZLZ^{+}16]. protein [pZWZ^{+}16, ZJSC16, ZKHL16, ZL18a, ZD18, ZRGW19, ZLB13, ZNCM15, ZNC^{+}15, ZLW^{+}19, Zho11, ZHC10, dBJ11, dS15a]. protein-coupled [BMF^{+}18]. protein-drug [NLB14]. protein-ligand [FCS18]. protein-ligand-binding [QQJW13]. Protein-protein [MALAN17, QZY17, ZT16, ZRGW19]. protein-RNA [LL12, PDM17]. Protein/Receptor [Ken19]. proteins [ÅCHS19, ARG14, AH19, AH16, ALM^{+}19, AAJGCD15, AHJ18b, BMSEE14, BRK19, CRH^{+}15, CXC18, CRV16, CT18b, DTY12, ESW13, FL12, FL13, FLW16, FBAPMD13, FLCS^{+}15, GZ14a, GZ14b, GDPPSS^{+}11, HZC^{+}10, Hua12, HKR^{+}19, IKHL16, III17, IK15, JLX^{+}15, JLX^{+}16, KCM^{+}11, KPK^{+}13, KHK^{+}17, KB11, KBK16, Kri16, Kri18, LL12, LLW^{+}18, LLP^{+}19, LTL^{+}15, LAJC19, Mei14, MP12, MBE11, NHSX14, RBMS17, RBRB16, RKJ^{+}11, Rou14, SRD^{+}15, SGGHS19, Sha14, SC10b, SSV^{+}15, Ste12, SH15, SSRA16, Tox18, WMK16, WdVMD15, WHH17, Wil13, XWC11, XDD^{+}15, XKCG15, YLL^{+}14, YGL^{+}10, ZJS11, ZGS^{+}10, Zha11, ZZ14a, ZSZM14, pZWZ^{+}16, ZRGW19, ZNA^{+}16, ZXC^{+}10, dSMP^{+}11]. proteolysis [Sel17]. proteome [LKAJ18, RNVP10]. proteomes [LLG16]. proteomics [WAC14]. proto [Di 14a]. proto-mRNAs [Di 14a]. protocell [ZSS10]. protocells [VMZK^{+}19]. protocol [AGC18, SSN^{+}14, Zhe15]. protogynous [YSST13]. proton [LL17b, Mar12]. prototype [GBJ^{+}11]. prototypic [ANGPB12]. protozoal [HFD17]. provide [IK15, TLW18, VGZS18]. provided [WZ12]. provides [PJ13, PVCEC18, TCYY^{+}12, Zho11]. Province [HJR12, TXTW16]. provinces [STG17]. proving [WSM12]. proxy [MHC16]. prune [BLS17a]. pruning [LMdL11, dG10b]. PseAAC [Mei13a, Mei13b, AFS19, AH19, AH18, BRK19, CCY^{+}19, CXC18, CT18b, DHS^{+}15, HYA14, HKR^{+}19, JLX^{+}15, JLX^{+}16, JLQ^{+}19, JCG15, JCG16, KHKI17, KJH^{+}19, Kri16, Kri18, LZ18, Mei12a, Mei12b, NBL10, NMZ19, RSS^{+}18, SM17a, SM18, STG19, SKK18, WZM19, XMWC13, XDD^{+}15, ZCAB17, ZL18a, ZD18]. PseFpt [XMWC13]. PseKNC [JYZ18, SIK^{+}18, TTC19]. Pseudo [AH15, AHJ18b, BMN16, CL13a, Cho11, EMM10, FL12, FL13, FLW16, GZ14a, GZ14b, HPB^{+}14, HK11a, HYW11, HY13, HGM^{+}16, JD16a, JD16b, JD17, JSF^{+}11, KHK15, KZL14, KSKK15, LD11, MJ11, MFZ18, MZ18, MBE11, MP14b, NBL14, NMZ19, PWZ^{+}19, QLC^{+}18, RBMS17, SRS^{+}15, TTC19, TWC^{+}19, TP17, WMK13, XNJ^{+}13, XSLZ16, YGL^{+}10, ZSZM14, ZZK14, ZK18, ZNA^{+}16, ZLW^{+}19, ZLDZ13]. pseudo-amino [GZ14a, GZ14b, HYW11, JD16b, JD17, KSKK15, MJ11, MBE11, QLC^{+}18, WMK13, YGL^{+}10, ZNA^{+}16, ZLW^{+}19]. pseudo-average [FL12]. pseudo-multiscale [HGM^{+}16]. pseudo-position [QLC^{+}18]. pseudo-trinucleotide [ZLDZ13]. pseudomallei [SSRA16]. Pseudomonas [ALM^{+}19, FE10, FIJJ11, MBC^{+}12b, SBJ^{+}18]. pseudopodium [GF11]. Psi [MIH16, DYQ^{+}14, PRM11]. PSI-BLAST [DYQ^{+}14]. PSO [FJS^{+}19]. PSO-based [FJS^{+}19]. psoriasis [RAR19]. pSSbond [KJH^{+}19]. pSSbond-PseAAC [KJH^{+}19]. PSSM [DLL^{+}17, HK12, KHK^{+}17, LZ18, LTL^{+}15, ZCAB17, ZL18a]. PSSM-Suc [DLL^{+}17]. pSuc [JLX^{+}16]. pSuc-Lys [JLX^{+}16]. psychiatric [PG18a, PG18b]. psychological [FWW18]. psychophysical [vTMW15]. PTH [CLA^{+}16, TSF^{+}19]. PTTQ18 [GY18]. Public [CJRR11, AS12, AS13, Arc16, AS16, Che12, CWW18, CSZT12, CB10c, DDRP13, Hau10, LCCK12, LvBJ16, MK14b, MV18, ON14, OO17b, ROF17, SU11, SHBT14, SP13b, WH11, WL19, XWZ10, ZCW13, ZWC^{+}13]. publications [SS19]. Publisher [Ano19-51]. pulmonary [GBJ^{+}11, NAK^{+}11, REC10]. pulsatile [GDF18]. pulse [ABLH18, BB11a, GZT15, LL18, ZTT18]. pulse-amplitude-modulated [GZT15]. pulses [HB18, LSMP14]. pump [Dim17, OCN10]. pumping [Alv18, YSI17]. Punctuated [WdLS17]. pungent [Di 17b]. punishers [OO17b]. punishing [Sch10]. Punishment [HNR14, PTB12, ZZCZ17, CW14, GWNW15, GT12, IL13, JR17b, KCS16, RANO10, SHBT14, SP13b, WL19, Wol12, ZCW13, dWV11]. pupylation [Tun13]. purely [TBQG14]. purifying [OGO19]. puroindoline [ASRM15]. puroindoline- [ASRM15]. Pursuit [ACT12, Gri15a, VP18]. putamen [RKMG12]. putaminal [GMK18]. Putative [Sel14a, AIY16, SPSM15, SBR16]. Pütter [OYA14]. Putting [JPDP10]. PUVCs [JSC^{+}16]. puzzle [DMO^{+}17, SSP15]. puzzles [TK19]. pv [JSF^{+}11]. pyemia [BC19]. pylori [LAH^{+}16].

QEPS [NNH^{+}16]. Qrr [HK14]. QSAR [FHG15, HME12, LL13a, LT14, MR18, MR19, TT17, TA16]. QSPR [VPGDP^{+}14]. QSPR-like [VPGDP^{+}14]. QTL [Boc12, SQZZ17]. quadratic [KF15]. quadruplex [DHB15, Dol19]. quails [ARB13]. Qualia [Esc11]. Qualitative [SAI14, CS18, Di 19b, KFW12, LvdBP12, SVA18]. qualitatively [FtL18]. quality [CXT19, GAB14, HMWB13, LTSTD15, LLH^{+}13, RFME^{+}12, SC10b, SB17]. Quantal [NHM10]. Quantification [HPP10, LCA^{+}15, LDH^{+}12, MZM^{+}16, Nos14, PBS^{+}12, DBG18, KCJ^{+}11, LDB^{+}14, PCB14, SMS19]. quantifies [CKZ^{+}17, LAPK14]. quantify [Mar17b, RTR^{+}13]. Quantifying [Boz15, CPW16, DP13a, GDF17, HPM^{+}17, HHG14, IOK19, JRB^{+}16, KSBvL^{+}12, KHNM16, LL17a, LTP19, cLCJ^{+}10, MB18a, NDZMA14, PBB^{+}15b, XW17, YBH^{+}19]. Quantitative [BE14, FPR10, Oko16, RCL^{+}10, SPMGR10, VH11b, Wil13, YM14, BKL^{+}15, FSA15, FLR14, HL10, HP16, HT13, KAPR12, LN13, LvdBP12, MAMEA15, Mog15, MHH13, PWC^{+}12, PFPR11, RKMG12, RCD16, SGGY10, SVA18]. Quantum [AOT^{+}11, ABK^{+}12, KDL16]. Quantum-like [AOT^{+}11, ABK^{+}12]. quarantine [DGRB13, TXY^{+}12]. quartets [GK16]. Quasi [AGPK13, ALH10, MK11, SGB19, BKSW16, CXC18, DEK15, JRG14a, JRG14b, JRG14c, Lio18, Lio24, PM10a, RRÚJ19, ZWG10]. quasi-balancing [CXC18]. quasi-equilibrium [Lio18, Lio24]. quasi-linkage [RRÚJ19]. Quasi-multiparameter [MK11]. Quasi-one-dimensional [AGPK13]. quasi-regular [PM10a]. Quasi-stationary [ALH10, ZWG10]. quasi-steady [DEK15, JRG14a, JRG14b, JRG14c]. quasi-steady-state [BKSW16]. Quasispecies [BDS12, Dal17, LCG^{+}15, MTNM12, SE10, TMM10]. quasistable [PJ18]. quaternion [RP18]. queen [SJSK18, STG17]. Queller [Mar11]. Quercus [AM10]. quest [JMZ13, Sal10]. queues [TF14]. quiet [SNCM12]. quirks [BFLS17]. quorum [BSL^{+}17, FBK11, LDHD14, MKPVH16, NBW10, NBW11, RAF^{+}14, RRC^{+}11]. quorum-sensing [RAF^{+}14].

r [LDB^{+}14, LDB^{+}14]. R.Br. [RAA^{+}16]. R1 [Ano14-47]. R2 [Ano14-47]. R2787H [RMM^{+}16b]. R3P [WMK14]. R3P-Loc [WMK14]. R58Q [MCL^{+}11]. rabbit [GMS^{+}13, Zha10a, Zha11, ZZ14a]. Rabbits [BH12]. rabies [BLS^{+}12, HJR12, LLN^{+}19]. Rac1 [ABR11]. race [Coo16, IST11]. radial [BTO15, BHP17, BBS18b, Che10, Sat19]. radially [She10]. radiation [BKL^{+}15, BGD^{+}13, BCPM^{+}16, HCK^{+}11, LTZ17, LHD^{+}17, PSD^{+}18, PMCS16, Sel14b, SC16, WVO^{+}15, ZG19, ZMS17]. radiation-induced [BKL^{+}15, BGD^{+}13, ZMS17]. radical [LL13a]. radioactive [BHH^{+}14]. radiobiological [BBJ^{+}10]. Radiosensitive [YBH^{+}19]. radiosensitivity [BGD^{+}13]. radiotherapy [BBJ^{+}10, KBV11, MM19, OOY16, YBH^{+}19]. radius [WGDZ18]. raffle [PIPB10]. raider [GSRR17]. rain [ADV^{+}10]. rainfall [LDM16, STLJ18, VSD^{+}17]. raising [MAW12]. raltegravir [CLPZ14, WML^{+}17]. ramp [TSB10]. Random [ATB14, FBAPMD13, FCS18, Gol10b, MSIR10, Par15, AEZR^{+}16, BP15, CML10, CC11b, Che10, GZ14a, GZ14b, GSF13, HMM16, HT16, HT10, JM12, JLX^{+}16, JKS15, Kam11, KCM^{+}11, KPK^{+}13, KO18, LMCW18, MS14a, MKKS14, Mit17, MAA18, NIiT18a, NRKE18, QMJW15, Rey13, SFV16, Sch18, Sch21, SGD12, SVB^{+}10, SLHS13, SPCM18, SK18, SY12, SN17, TGLK19, WMK14, Wax11b]. random-walk [CC11b]. randomization [NS13]. randomly [Fan11]. randomness [LTP19, MKBE17a, MKBE17b]. Range [PSD^{+}18, AGGME18, BBS18b, CKNB19, CVO^{+}15, FLM18, HBT13, HS14c, JP19, NG17, RN12, SGW^{+}18]. rangeland [BCBD19]. ranges [AT10]. ranging [Pon12]. RANK [PZS^{+}10]. RANK-RANKL-OPG [PZS^{+}10]. Ranking [Hua12, TLY^{+}19]. rankings [WF17]. RANKL [PZS^{+}10]. RAP [Kri18]. Rapid [BTK15, GU13, HMM16, SYR11, XJ19]. Rappaport [Wol14].

Rare [HA15a, GZ19, HMM17, YI18b]. rarely [BM14]. rat [Pat19, TSF^{+}19]. ratchet [WL10]. Rate [WS10, AY15, BBR16, BS15b, BS19a, BL14, CZT^{+}16, CK15, CKB19, CCdS15, DGJ15, DHP11, FAFA17, Gia13, Hor13, KAZ11, KAZ13, KG15, Kit10, LG18, LLB^{+}18, MZWC10, MC19, MB10, MKMT14, NSS16, PBYB18, PW18b, RU16, RRRRSPTR16, RBMPP^{+}15, SS15d, SL17, SLJ^{+}10, TP14, TSMB14, Tsu19, Van17a, Wal16, WST15, WZLW15]. rate-dependent [FAFA17]. Rate-variation [WS10]. rates [BGF^{+}14, EBE14, FGH^{+}14, FPdRdH10, GHBC14, HPM^{+}17, HS12, HWHL15, Jia16, KFA12, Lan16, LKK13, Lev14, yTRSC13, NiTI18, NdMLLBLC13, NPH^{+}12, NC15, O’F11, PBYB18, RDMP11, RBB^{+}12, Sel10, SAZ^{+}14, TTC10, WZ15, XLL^{+}10, XKCG15, ZDF^{+}12, vE11]. rather [Di 12]. ratio [AM10, Arg12, ALH10, Del12, IDM15, JV14, Jam16b, Kun16, LLW15, LXS15, LS14b, MCN10, MB12a, MS14d, NG11, PKZ12, Ram10, SS11, TEY10, UI13, YTK10, YK11]. Rational [BC19, LN13]. rationality [HS14b, XWZ10]. Rationalizing [Coh15a]. ratios [Jam11, Jam12, Jam13, MT14, NT14, NG11, TEW12, TEY16]. rats [LLB^{+}18, SKK^{+}11]. RAW [CHD^{+}10]. ray [CZPC^{+}18, ZZ14a]. rays [TRM^{+}14]. Rb [CCS^{+}16]. RBC [KTI18]. RBSURFpred [TAI^{+}18]. Re [ER12, ASH15, DOC17, GGM15, HLTR19, Nak16a, dBE16]. re-cycling [Nak16a]. re-emergence [GGM15]. re-entry [DOC17]. re-evaluation [ASH15]. Re-examining [ER12]. re-polarisation [dBE16]. re-sensitisation [HLTR19]. reactance [Don16]. Reaction [PWL^{+}11, SAB^{+}19, BFS18, BMB^{+}18a, Bro13, DGW^{+}18, HS12, KS19, KBS^{+}19, KBS^{+}20, Kon11a, Kon11b, Kon12, MTE15, MSSM18, MF15, NGN13, NBC16a, NBC16b, OM10, OAS10, PP14, RV16, RSR11, SI19, SCF^{+}12, SOBC12, TP10c, TWH^{+}13, WZ18, YS14, YIS^{+}17, YZX15, ZM10]. Reaction-diffusion [PWL^{+}11, SAB^{+}19, RV16, SOBC12]. reactions [FWMC14, KJ11b, LYF^{+}15, NGN13, SHN12, Van17a, YI18a]. Reactivated [KEHK17]. reactive [VP18, VR17, WHS^{+}13]. reactivity [ALH16, HWW^{+}14, SDD15, XJ19, YCH^{+}17a, YCH^{+}17b]. reactor [AIY16]. reading [Mic14, Mic15a]. real [IL12, RRRTR10, TAI^{+}18, ZKP15]. realism [ZGW16]. Realistic [XG12, GFM^{+}19, PDG12, vTMW15]. realistically [DNB12]. realities [Mor16]. realized [HTM16, YI18a]. realizing [YDC^{+}13]. really [LM15, Nic19]. rearrangement [ITN^{+}11, SENS^{+}17]. reasoning [LA12]. recall [DHT19]. recapitulate [DS19b]. recapitulates [SHB^{+}17]. RecBCD [Xie18]. receivers [EM11]. receiving [LX16]. recently [Di 12]. receptive [BZ18, Pat19]. Receptor [BJJR10, Ken19, AES^{+}13, BMF^{+}18, GBC^{+}16, GBIR13, Gri11, HP12, KG12, Kri18, LC16, ML11a, MBE11, NDG14, PBvdG10, RSI11, UAL10]. receptor-2 [GBC^{+}16]. receptor-ligand [KG12]. receptor-mediated [AES^{+}13, UAL10]. Receptors [PJL14, AH12b, HPML18, MPS^{+}11, OBK^{+}11, PBBB10]. recessive [HTM16]. Reciprocal [PTNKT11, SC13a, SC13b, VDTF15, Mou12a]. reciprocally [PJ18]. Reciprocity [TP10a, TSM13a, BBD18, DGC11, GN15, IM10, KV15, KI17, Kur17, Mas12, MJI14, NO14a, NO14b, OSN18, Pan11, PPBT11, Sig12, Syi14, TSM13b, US10, ZLY^{+}17, vDRT14]. Recognition [FLM18, AC15, Bry13, CSR^{+}05, GvVT14, HWW^{+}14, IA17, LCG12, LDH^{+}14, Lug15, LBS^{+}14, LPD^{+}16, MCL19, Nah14, New11, PDNP16, RSD^{+}16, SRS^{+}15, SLDP13, SPO^{+}10, SM11b, SRAL12, Xie18]. recognizing [WSH^{+}10]. recoil [RS12]. recombinant [DKP^{+}18, MÇ15a, MÇ16b]. recombination [FM18, GKPB13, HG18b, JJ15, KII19, LLCC12, LXC15, MZWC10, MYK17, NK18, VDTF15, ZK18, dASdC^{+}15]. recombination-mutation [FM18]. recommendation [HGM^{+}16]. Reconceiving [Tau15]. reconciled [MN12]. Reconciliation [bCR19, SMC17]. reconciliations [bCRS17]. reconfiguration [LMdL11]. reconsidered [CA17]. reconstruct [LAJC19]. reconstructed [DRCL12, Hal12, Höh15, KCJ^{+}11]. Reconstructing [CMPS17, NvD17]. Reconstruction [KSM^{+}15, Lan16, ASP13, DG10a, Ell15, HB12b, Kae17, PWC^{+}12, PPTC19, QYO10, RCD16, WH14, WSTL16, ZOG14]. record [Bue15, DRCL12]. recordings [For10]. records [TLS^{+}13]. recover [ABLH18]. recovered [Gru18]. recovery [AGC18, Gou15, LG13, RJSC18, SSME10]. recruiting [Cro18]. recruitment [AdlPLMZ13, CDD12, Ezo19, QA15]. rectification [For10]. Recurrence [Nås12, FHM18, FLR14, OYY16, SGB19]. recurrent [BV12, TA13, VABS18]. Recurring [BKC^{+}16]. recursions [RLBCJ13]. recursive [FIS11, FIS16a, LTL^{+}15]. recycling [BWL12, SCS15]. Red [Fre10, DK13b, KCD11a, KCD11b, SPMGR10, STG17]. redox [FWMC14, SI19]. reduce [Hor13, SI11a, SS15b]. Reduced [HVL11, KJ11b, TD17, HY16, OPS^{+}19, PWZ^{+}19, ZGW16]. reduces [LP14]. Reducing [CK17, TAL19, LZY^{+}15]. reductase [WLD^{+}11]. Reduction [AENK12, VC11, AdGM12, BP18, BFH^{+}15, BKSW16, GJ15b, KSP18, Kon11b, KLHS17, LWY^{+}16, PAK11, PP12, PP14, QW11, RAD14, VG13, YYC19, ZS10]. reductions [MNH^{+}12]. redundancy [JD16b, KPK^{+}13, MJC^{+}11]. redundant [HK14, YCR^{+}15]. reef [JCLS^{+}11, JCLS^{+}21, TZ13]. Reentrainment [DB18]. reentry [RLSM17]. Reference [BS15a, BGF^{+}14, Boz15, DLSD15]. refilling [dlRVFK^{+}16]. Refining [ZOG14, SCKL15]. reflection [NBC16a, NBC16b]. Reflections [Jam13, Ser16]. reflex [Bac15]. reformulation [MC19]. refractoriness [MAS15b]. Refractory [CP14, HTK^{+}18]. refuge [CGHF14, Coo16, Kři11, KP15]. refuges [Kři13]. regarding [JN14, JAHKH12, SS12b]. regenerating [CLHB11]. regeneration [LNRK11, MZMM^{+}14, PGPD19, RJSC18]. Regime [PSV17, SI11b, AHKN10, CCdS15, JAB18, JRG14a, JRG14b, JRG14c, LW10a, LW10b, SI10b, SY15, Urs14]. regimen [LN13]. regimens [HL18]. regimes [AHD^{+}18, HHT^{+}19, PJ18, PBD11, Str17]. Region [CzST^{+}18, WZJ^{+}13, CJ12b, CW15b, JRR^{+}12, JH12, PRM14, SRP16, YM14, ZLZ^{+}15b]. regional [BVJE17, DL15a]. regions [DPCM16, KGM15b, LZMM15, MGO^{+}15, RW16, RBRB16, SDRA^{+}15, TDM14, VH11a, VBHM^{+}13, WSTL16]. regraft [BLS17a]. regression [HME12, HZL^{+}11, KKD18, Mog15, TAI^{+}18, WMK14, YW13a]. regret [EHSR17]. regrets [EHSR17]. Regular [CE17, CZJQ14, JC16, JKS15, LF10, PM10a, SAI14]. regularity [KKM12]. regularization [WcW14]. regularized [TAI^{+}18]. regulate [HMJB15, ZBA18]. regulated [BB19, FP18, Hor11a, HgLL^{+}10, JOM16, KTN^{+}10, ZS10]. regulates [PEZ^{+}13, Zhu11]. regulating [YAK17]. Regulation [AHW13, CLHB11, JSWY19, KP12, AMP12, ATI12, BPP^{+}16, BHBH11, BK19, BHLH12, CFF12, CKD11, Das18, Deu18, Dil14, DM19, EdKM11, EBE17, EBSW17, FD17, FZL18, GHJ14, GTB10, GCO^{+}11, GWW^{+}18, GS17b, HTM16, HGB^{+}13, HG18a, HMJB15, HBA17, IvdSW16, IU18, JAK19, JL16, JK12, LW12, MH11, MBF^{+}15, Mic11, MMK16, MBK^{+}11, NCS^{+}13, NWWL17, PGLZ14, PWE15, PSK10, RZIR16, RW14a, Sat10, SK16a, SBR10, SCKL15, SC13a, SC13b, ST17b, UPWK15, VH11a, VLPH17, WM17, WD12, Wod18, WCHC11, WBB^{+}18, Xie18]. regulator [Ber14, RLM^{+}14, ZSCL16a, ZSCL16b]. Regulatory [LS13b, ML11b, OUMA10, SCS14, WML16, AEE19, AGS18, BCPL10, Bar19, BSM^{+}14, CKF17, CSKZ19, DTGD19, DG10a, Fra19a, Fra19b, GML10, GJS^{+}10, GGG12, Hal17, HCHI18, HL14, HXL16, HXL18a, JDPK15, JZL13, KSP18, LWH^{+}11, LAH^{+}16, LZL^{+}17, MTKY11, MC16a, MMC19a, MBE10, MB14, MCL^{+}11, MS10b, MFKS13, NS18, NMS10, OHWS18, OABI12, PAOMV17, PPBD10, PS13, PPM12, PMP13, QYO10, RLM^{+}14, SI10a, SDD15, SCÁBM10, SCÁBM11, SDRA^{+}15, SWTO15, Str15, TK10b, TLK19, Vel17, XSKA17]. regulon [AM12]. reinforcement [Mar19, MN11, Mor19, TM12]. reinforces [WFZW13]. rejection [EBE10]. RelA [HSiI^{+}19]. relapse [HP16, Mal10]. Relapses [CB13a]. related [ÅCHS19, BTR18, FLGGD^{+}14, GY19, Gol10a, Gol16, HKS15, HZL^{+}13, KLHS17, LCT11, LBZ18, MCRG18, RAA^{+}16, ZMT11]. Relatedness [TG10]. Relating [Pon12]. relation [Das18, Fil15, GS10, LWH^{+}11, LP12a, NGN13, TP14, WCF12, ZGG^{+}19, vNÁBG12]. relations [All11, BLZ^{+}19, Fre10, KFG^{+}14, MT14, PDG12, Sin15]. Relationship [CCS^{+}16, IOY15, BAGG14, DBBW09, DBBW11, FSA15, GZT12, GLZ^{+}11, GABM12, GZY12, KTI18, Kri18, KSPA17, LL17a, LV17, LL13a, MGO^{+}15, MK16a, MW14, SKK^{+}11, WC10]. Relationships [Miy17, FD18, FB18a, FR13a, Gla13, HZL^{+}13, MZWC10, MI19, NdORC11, PGLZ14, PHC12, TLM14, TZY18, THM10, dBB14b, dBJ11]. relative [AK11, BI10, HGHL10, KG12, PBYB18, PBB^{+}15b, SM10b]. relatives [AN15]. Relativity [BAM^{+}11]. relaxation [KFG^{+}14, WHWZ18]. relaxed [KHX^{+}19, Smi14, Wol14]. relay [HZG^{+}17]. release [BTO14, BAR14, CPS10, CSM^{+}14, DDF^{+}14, HYN19, LL18, TK10a, WMN18, dLMV^{+}10, dLMD^{+}11]. releases [IH17, TTC10, ZLT^{+}19]. releasing [CTJ10]. relevance [CDGV10, DS10b, GE18, Gol10a, JD16b, KPK^{+}13]. relevant [AL17, APW10]. reliability [ML11a, PM15, ZOG14]. reliable [DQS^{+}15, JP13]. reliably [ANY14]. religions [DI10b]. relocalization [BGD^{+}13]. REM [PRK13]. REM/NREM [PRK13]. remaining [OA15]. remarks [Cho11]. rematching [QMJW15]. remedies [YBC17]. remigration [NSO15]. remodeling [CPF^{+}13, EJK16, HHMT19, Mic11, PZS^{+}10, RM17d]. remodelling [FGAM12, NF14a]. remopleuridid [SKSO12]. removal [EB15, Jia16, UBP12]. Renaissance [DGRB13]. rendezvous [FH13a]. René [TK19]. renegade [Har15]. renewable [TSL12]. Renormalization [HB10]. reoccurring [VAT18]. reorganization [CMCS18]. reorientation [Kro10]. repair [BCKS13, BKL^{+}15, BKPV15, RFM^{+}12, SGAM12, TALC16]. repeated [IM18, JR17b, Kur19, KBF18, RZ16, TP10a, MI19]. repeats [Sel15b, WDH^{+}16, Yin17]. repercussions [TGZF11]. reperfusion [MYK^{+}11, PNP^{+}16]. repertoire [GZT12, KFG15, SLvdBMP10, XJ19]. repertoires [RSI11]. replacement [MMB14]. replacements [BWM^{+}19]. replanting [CLK18]. replica [XYZ15]. replica-exchange [XYZ15]. replicase [SHH15]. replicating [AIY16, Mau15, SG18]. replication [FLA^{+}19, GN10, JS15, KDK14a, KDK14b, NS11, Nak12, Nak16a, OF10, PBB^{+}15b, Sel14b, TH17, ZS12]. replicative [Hor13, HZW^{+}14]. Replicator [CKS15, Hau10, SU11, AY15, BYJ17, Gou16, KGC18, MB16, vV11a]. replicator-mutator [Gou16]. reply [DSF11, Di 14b, LE19, vV11b]. repolarization [PW18b]. Repopulation [EUM^{+}16]. reported [Jam11]. reporting [Rao12]. Representation [BB13, BWS10, DGL12, DDLW10, aHLZ^{+}12, JLQ^{+}19, JSF^{+}11, MM12, SCS10, WXC10, WRWX19, ZKHL16]. representations [Mal15, XM11]. representing [SBSR13]. repressilator [Dil14]. repressing [SLS17]. repression [UK17]. reproduce [Wu14]. reproduced [Nak16a]. reproduces [ZBA14]. Reproducibility [JSP^{+}16]. reproducing [OFT15]. Reproduction [LP14, AM10, CH11, Cro17, EHSR17, EBS11, EEHMH18, Fil15, Fis19, Fre10, GBRS19, HHA^{+}13, LRH15, MO12, NR14, PBB10, PLF18, QA15, SRV11, SSD11b]. Reproductive [Mac14, SR10, CH11, FS12, Gar14, Gra14, Hal17, Jam15, KD16, KSKS13, LLW15, MCN10, MS13, Non10, PMD11, RRR15, SRV11, SS11, WDL^{+}13, YK11, YY18]. reptiles [PC17]. Republic [EN15, KLJ17, KJ19, LCJ16, LCJ18, OL19]. repulsive [KAN11, TBQG14]. reputation [Hau10, Mas12]. Reputations [TSM13a, TSM13b]. required [Yam16a]. requirement [SBB18]. requirements [BFS18]. rerio [Hal17]. resampling [JZ17]. Rescorla [TMHM12]. rescue [Oht12, WSVT14]. rescuing [SBB14]. research [Ano19-50, KSBvL^{+}12]. reserve [MHC16]. reserve-utilization [MHC16]. reserves [NNvdB17]. reservoir [BFH^{+}15, VABS18, WR14]. reservoirs [KGK10]. reset [FW15]. resetting [AUE15, Opr10, WTZ^{+}11]. resided [YLWZ10]. Residence [DD10, LCC15, MMC19b]. residency [MGS16]. resident [DG16, JD16a, JD16b]. resident-invader [DG16]. residual [Ren13, ZR16]. Residue [EG12, HYZS14, AKR^{+}18, HH13, JR17a, QW12, QQJW13]. residues [KB11, KGM15b, LBS^{+}14, MS14a, MPY14, PDM17, SLT^{+}18, Zho11]. Resilience [ALH16, VL11, BS19b, KPG19, RW14a, SY15, WB19]. Resin [Cab18]. Resistance [WPA17, ACMK12, ABA11, CDO17, CL17a, Che16, CGW18, EJ16, FM10, FM14a, GSYS10, GY19, Goe15, GA14, HNP18, HT18, HPvdB17, HLTR19, JSF^{+}11, Kle14, LWRE14, LvdBP12, MPM10, MB18a, NAK^{+}11, SBM^{+}16, Sch17, Sch23, SS11, TB15, VL11, WSVT14, XSMF12]. resistant [Car17, CL18, CBGS18, LP19, PDW10, TDKJ15, WMN18]. resistivity [Tak17a]. resistivity/conductivity [Tak17a]. resolution [ASL^{+}18, CZPC^{+}18, NGND12, ZSW11, vdBD10]. resolve [JEA18, KMMS17, MRPN12, Tjø10, WMD^{+}11]. resolved [Ste13]. resolving [SL17]. resonance [BDLR12, CE17, GL12b, IDI11, LL13b]. resonant [GP11b]. resorption [FGAM12]. resorptive [CLA^{+}16]. Resource [DBB14a, GSRR17, AR11, AdVG15, GCZ^{+}12, GW14, HMTA17, IL13, KB15, MG14b, MLT10, RC11, Rev15, RB13, SCH^{+}19a, SSD11b, SSD12, SB17, SY12, SA14, TSL12, TLW18, TRJD19, WD11, YSST13, tBdR18]. resource-based [Rev15]. resource-consumer [SSD12]. resources [BV13, GW14, HWGW18, KRSD11, LJI17, RC11, SMG18, SJK18, VK10]. respect [BAM^{+}15, SMG18]. respectively [SGD^{+}16]. respiratory [BTK14, CXT19, Don16, EAN14, JB19b, LCJ16, Str13]. responds [BBS18b, Lai18]. responds-promote [Lai18]. Response [JAHKH12, KSK^{+}11, SNC12, ARG14, AJD11, ABV19, ARD14, ABN15, AHAFA14, BBJ^{+}10, BS13, BPPC15, BPLM12, BPGS12a, BPGS12b, BKPV15, BAM^{+}11, CKB18, CNM^{+}13, CMJD11, DFS11, DLSD15, DHT19, Eli17, FMLM12, FM15b, FHS^{+}14, FGJ10, GVCG12, GVGC15, GG12, GG14, Har15, HBVP15, JB18a, KHH10, KHS13, KB19, KMN16, Kor18, KP16, LDHD14, LMYL15, LNH13, LvdBP12, LG12, Lüt16, MCL19, Mal17, MG10b, Mor10b, NKB16b, NBW11, OM10, OFT15, OHWS18, PDG17, PGHC12, PMKS^{+}15, RZRSC19, RGA^{+}10, RTFP^{+}17, RG10, RPD14, RTEKG15, RTEKG19b, RZIR16, RW12, RDTHSM13, Sat10, SCLA10, SBSR13, SR12, SMA11, SVA18, SH16, SKD^{+}10, SW19, SSKK17, Str15, TCH14, TALC16, VAG16, VBVD17, WPH^{+}12, YCH^{+}17a, YCH^{+}17b, YFB^{+}12, ZCSRG12, ZZL^{+}11, ZLM12, ZSCL16a, ZSCL16b, vLBJ^{+}13]. response/signaling [YFB^{+}12]. responses [AF10, Abr14, ABLH18, CML11, CB10b, CTL^{+}15, DS13, GGQ^{+}12, GS13a, GS14, JDPK15, Jen10, JSDEK14, KC14, LLTP19, LCMC14, Lüt16, ME15, MBBV14, PP17b, Rev15, RBB^{+}12, RR12, WDH^{+}11]. responsiveness [PDT^{+}10, SN17]. rest [TKTB18]. rest-inserted [TKTB18]. restitution [WMT15]. Restoration [GDB^{+}19, MI11b]. restore [KP16]. restored [MH17]. restorer [YYB^{+}19]. restoring [RRL^{+}13]. Restricted [MR11a, SHN12, BYY^{+}11, HRCA19, MBCM12]. Restriction [ESG10, KDL16, RTR^{+}13, SGS15]. Restriction-modification [ESG10]. restuans [YMZ18]. result [KK18, Wax11a, WWB14]. resulting [Alv18, IU18, KKS^{+}14, MPLK14]. Results [CCB11a, LHD^{+}17, Smi11, ACM16, IMH15, MS14b, NSK13, PM10a, RGA^{+}10, RG10, SW17, VPFV13]. resurgence [ZLCH18]. Retaliation [Wol12]. retardation [CKZ^{+}17]. rete [OB15]. retention [ADC19]. retentive [FMS^{+}10]. Rethinking [GAGP^{+}14, RD17]. reticular [DL12b, DL16]. reticulum [SGGHS19]. retinal [DFG^{+}18b]. retinitis [CW13, CPW16, RGL^{+}17]. Retinogenesis [BF15]. retinol [NHTS14]. RETRACTED [BBM^{+}97, BBM^{+}13c, GZ14a]. Retraction [BBM^{+}13c, GZ14b]. Retroactivity [PHÁBAI^{+}16]. retrospective [LCJ16]. retrotransposons [BWP10]. retroviral [VDD12]. retrovirus [WHZ^{+}17]. retroviruses [LHL^{+}14]. returns [FOT^{+}15, WSW10]. Retz [RAA^{+}16]. reveal [APW10, GRB^{+}13, LPS^{+}14, MSN14, SHN12, SS15a, SBMH10, TWR^{+}18]. Revealed [PS13, SYY17a, SYY17b]. Revealing [HII^{+}19, KSP^{+}14, Sha14]. reveals [BPC15, BGCB12, DLEMP19, EKvdKvFK13, FB12, GLZ^{+}11, GSS19, HPM^{+}17, HS14a, KBF18, LMM^{+}13, Luo14, SW19, WLL^{+}14]. Revenue [TLW18]. Revenue-sharing [TLW18]. reversal [EA19, Gal11a, gSxFH^{+}12, SBB18]. Reverse [RGBR17, TP10c, TVA18, KSP18, LMM^{+}13, NdlPZA10, SWTO15]. Reverse-engineering [TP10c]. Reversibility [TP10b]. Reversible [SMD^{+}16, CWF^{+}16, CK15, UBP12, YS14]. reversion [CST16]. Review [AS12, WH14, BCPM^{+}16, Bur19, GBJ18, PBSM19, SBP13]. revised [JLH15a, KFG^{+}14]. Revising [MMB14]. revisited [ASGE14, AN15, AR17, CNG^{+}12, CB13b, DGJ14, FDM^{+}11, Goe15, GGR11, HMM17, IRS11, JLH15a, KV15, LDB^{+}14, MMB11, SDJ^{+}15, WWY12, Wol11, XF13, XG12]. Revisiting [BA19, BK19, BCK19, FS11b, GGG19, GHBC14, Gre15, Sat17, TP10a]. revival [HAB^{+}17]. reward [Hau10, RM17a, SU11]. Rewarding [WSP14, CGD13]. rewards [IOY15]. Reynolds [HM11a, LS10b, SV14]. RFCRYS [JM12]. rhamnosus [RZIR16]. rhesus [WVO^{+}15, HII^{+}19]. Rheumatoid [MF19]. Rhinonicteris [AK11]. rhinovirus [AK13]. rhythm [HBVP15, LW12]. rhythms [GI15, KA16, SCLA10]. ribosomal [KMLT14, KPLT14, RBRB16, Xie13a, ZLZ^{+}16, ZLB13]. Ribosome [SC11a, BBRT12, RBRB15, RBRB16, SV14, Xie13b, Xie15]. Ribosomes [RBRB16]. riboswitch [GWW^{+}18, XYZ15, dS15a]. riboswitch-binding [dS15a]. rice [DXW^{+}16, JSF^{+}11]. rich [CT14b, KB11, MBCM12]. richardii [SH17]. Richards [TT10a, WWY12]. Richards-model [TT10a]. Richardson [NBW11]. richness [DSA^{+}16, MB15]. ridge [KC13, WMK14]. rifampin [GBJ^{+}11]. Rift [XSCS12]. right [Arc11, Mar11, Pan11, SLL18, SWSLMJ19]. rigidity [POP^{+}19, THBM10]. rigorous [DHK13, Shu13]. RIIIb [OBK^{+}11]. Riluzole [BGCB12]. ring [ACvKA10, CMB^{+}12, Hor11b, KHB^{+}18, Opr10, Pag19, SYSM11, Sht17]. rings [DS19b]. riparian [TCR13]. rise [Laz13, SAB17]. rises [JV14]. Risk [HYSG19, cSGFB17, Ada16, BZ10, EMM10, FLR14, HYN19, HN19, HZL^{+}11, HvLPdB15, HRG11, IMW15, JPBM17, KOS13, KKUM10, Lar10, MDE11, MSBB13, MPLK14, PBB10, PGF11, RT19, SI11a, SP11, TDZ^{+}18, TCB13, UI13]. risk-averse [MSBB13]. risks [TLS^{+}13, ZZX^{+}16]. river [MBRRI19]. riverine [JB18a]. rivers [LS11b]. RNA [AH18, AEZR^{+}16, BS19a, BB13, CLC11, CDC18, CzST^{+}18, CKS15, Das18, DS19b, Deu18, DHBS19, Di 12, DMCP14, Dun11, Fan11, Fra16, FYZ15, JSZ12, JWS^{+}10, Koe15, KIT^{+}16, LL12, LTP19, LZMM15, MMY^{+}12, MWSM13, Mas18, Mau15, MMCZM12, Nah14, Nak12, NKB16a, PDM17, RBR18, Rou14, SIK^{+}18, SE10, Sel12a, Sel16, Sos18, SM11a, TTC19, TXCW15, TI12b, TH17, WYL^{+}19, WSK^{+}19, dS15a]. RNA-assisted [BB13]. RNA-binding [JSZ12]. RNA-guided [TXCW15]. RNA-protein [AEZR^{+}16, MMY^{+}12, WYL^{+}19]. RNA-RNA [MMCZM12]. RNA-world [CKS15]. RNA/protein [dS15a]. RNAs [CM17, DMCP14, FZL18, GMZM15, Sel13, VHS17]. RNase [ACdlRMR^{+}11]. RNN [MP12]. Ro [AP13a, BFGS10]. Roadmap [GM17]. Robert [HS16, CB15]. Robust [Fra19b, GCO^{+}11, LWW14, TW13, BWM^{+}19, Fra19a, GSN^{+}11, Hir16, iHIM12, pNMK^{+}10, PJ13, RG12, SCÁBM10, SCÁBM11, TF15, YM16]. robustly [TM16]. Robustness [BV13, ITO16, AUE15, BBJDS11, CZ14, GJS^{+}10, GSS19, HH14, IMA14, KSvdH18, LKO^{+}17, MK11, MH17, PCC16, PPM12, PMP13, SGD12, SI11b, SJK18, UPWK15, WB10]. rock [BMdA17, Mob10, NIiT18b, NiTI18, NIiT19]. rock-paper-scissors [BMdA17, Mob10, NIiT18b, NIiT19]. rock-scissors-paper [NiTI18]. rodent [AGPK13, SWG^{+}15]. Rogers [KMMS17]. rogue [SAG19]. Role [ABR19, ARG17, Ber14, GCYL18, Jam16a, KSSM11, KSP12, KB10b, LHV16, MK18b, PR17, RD12, SCH^{+}19a, SSV^{+}15, YX15, dG10b, AZOLVH18, AEE19, APS^{+}13, ABKS11, AdlPLMZ13, AGC18, AN13, AK11, AIB^{+}19, BA10, Bar19, BZ18, BCKS13, BRN15, BLNR15, BN12, BPR15, BPLM12, BLS^{+}17b, CRGS18, Cam11, CT14a, CFF12, CST16, CW11c, CFGRB17, CSB15, Cle18a, CV19, Cor18, Di 17a, DBD16, DFC^{+}12, EdKM11, EKvdKvFK13, GWNW15, GML10, Gru18, dOGL10, dOGL13, HJLNZ11, HLI15, Höh15, KCZ^{+}19, KSP^{+}14, KAZ11, KPD18, KPS17, KC11a, KGM15b, KST17, LS16a, LCCC10, LCC15, LPGTBMA19, LLD^{+}17, LBF13, LVLC18, MCL19, Mal10, MM19, MGCL13, MKMT14, MPZK16, MGB17, MK14d, MBKB13, MH14, NCS^{+}13, NG17, PR18, PKZ12, PA15, PZS^{+}10, PBKR13, PRM12, PBA12, RTRRS^{+}17, RFS^{+}15, RPGG^{+}19]. role [RKK^{+}19, RBKW19, SDD15, SPLP12, SAB^{+}19, SV14, SAGAGB17, SW13b, SW19, SM11b, Tox17, Unc11, WKB13, WMT15, Wol12, WSM12, ZHAK14, ZGS^{+}10, ZFWK17, dSKBS10, dBE16, vNÁBG12]. Roles [SZ18, ZZS19, BN15, CDGV10, GH11, GK10, HSiI^{+}19, HKS15, MNSZ16, MHMM11, PM11a, RAMS11, TC11, TWB16, WLL^{+}14]. root [BBS18b, Che16, DKP^{+}18, Di 19b, DV12, FM14b, MBK^{+}11, MBKB13, MLL^{+}16, PMKM10, TP14, TLR14]. rooted [Kae17, Mul11, Mul12]. roots [FM13, SSF15, VW10]. rorqual [PGS10]. ROS [SAB^{+}19, XCW^{+}18]. Rosen [HS16, CB15, PWG16]. Rössler [BA19]. rostrum [IST11]. rotating [CC10, KNT10, VDH^{+}15]. rotation [CJ12b, Cle18a]. Rotational [MED12]. rough [CZW^{+}11, CKNB19]. roughness [PIT12, QA15]. route [Lio18, Lio24, Sil16]. routes [HB12b, MSB16, SPRF13]. routine [Hou16]. rpiCOOL [AEZR^{+}16]. rRNA [WSTL16, WZ17]. rRNAs [RBRB16].

RRSM [HW11, HW13]. RSV [GPD18]. rule [AW12, Fer11, HCS^{+}19, HKR^{+}19, IDI11, Mar11, MS14d, NMZ19, Ros10a, SFD12, SSD12, TMHM12, Voe18, ZZC10, ZLY^{+}17, vV11b, vVAH^{+}17, Tay16b]. ruler [Kee10]. rules [CC10, GG16, JD13, Lug15, MKJS13, MB18c, Pan11, PMW12, SHM^{+}18, Tak16, US10, WM17, WMT10, ZZRZ11]. rumen [WJL^{+}16, vLFM^{+}19]. ruminant [HFD17]. Run [AA12, Ran12]. Run-and-tumble [AA12]. runaway [NPB13]. running [ARB13, BBJDS11, Fue14, Fue15, Fue16, LGR^{+}12, SS18]. rural [LG18]. Russia [AL17]. rust [LK15]. ryanodine [Gri11].

S [BD10b, JAM18, JZ17, JW18, Pan19, HKR^{+}19, KGM15b, WZM19, ZNA^{+}16]. S-FLN [JAM18]. S-glutathionylation [ZNA^{+}16]. S-loop [KGM15b]. S-prenylation [HKR^{+}19]. S-sulfenylation [JZ17, JW18, WZM19]. S-SulfPred [JZ17]. s.s [ÅCHS19]. S2 [KGM15b]. SAAC [AH18, AHJ18b, HK12]. Saccharomyces [AM12, BLP10, HZW^{+}14, LLCC12, LXC15, OB10, SSJK18, TK10b]. saddle [SSN^{+}14]. saddle-point [SSN^{+}14]. safety [FGLS10]. Saharan [MS16]. SAIR [Gru18]. Sakishima [SFT14]. saliva [HR12, PSS^{+}10, PSP^{+}12, PSS^{+}13, dLMD^{+}11]. salmon [GDJC11, HIT18, SGCD14, SWD16]. Salmonella [LPS^{+}14, SPSM15, SBSR13, ZVMB10]. salt [FM13, FM15b]. SAM [XYZ15]. SAM-II [XYZ15]. same [BHBR12, CRV16, CB15, KBK16]. sample [Bar16, WSW10]. samples [GZT15, WCC14]. Sampling [BGLR^{+}19, CC11b, Sta10, JHE15a, JHE15b, LAS14, MSP19, MM18, NTFK11, SSP15, SS15d, WHL13]. Sampling-through-time [Sta10]. sanctioning [SHBT14]. Sanctions [SHBT14, NSKO18]. sapiens [KHH^{+}17]. sapwood [WCF12]. sarcomas [lGLLL14]. sarcomere [Giv10, TSB10, ZPM^{+}15]. sarcomerogenesis [GAPK10]. sarcoplasmic [SGGHS19]. sarcoplasmic-reticulum [SGGHS19]. sarcoptic [BCF19]. SAT [MKF^{+}14]. satellite [GMOP12]. satellites [RGP13]. satiety [DZJR10]. saturation [CMS^{+}19]. Savanna [DAVS11, ADV^{+}10, Urs14, VLCM10]. savannas [DDS13, MGCL13, STLJ18]. saving [TTW13]. savings [MPJH13]. scaffold [BBDB13, HHA17]. scaffolding [KFL12, Smi14]. scalable [MC19]. Scalar [OABB18, RWH16, QA15]. Scale [BAM^{+}11, BP15, BVJE17, BC15, CAGM^{+}17, CL17b, CDC18, CMJD11, Cox10, EBP15, FMS^{+}12, FGGT15, Giv10, GJS^{+}10, HXL16, HY10, ITO16, Ing16, KCZ^{+}19, KPD18, KSKS13, LKL13, LDB^{+}14, Mac11, MAM16, Mei14, Mic14, Mic15a, MP13c, MLL^{+}16, Ots11, PAK11, PBD13, PMSY17, RA10, Rev15, SH15, TvMG16, TB16, WZ17, WGC13, XLSF19, dlCGSA16]. scale-free [BC15, GJS^{+}10, KSKS13, LKL13, XLSF19]. scales [BBWS^{+}13, DLSD15, GABM12, GHS18, GHS20, MSIR10, NGND12, NP15, TZ11]. Scaling [CNM^{+}13, LLTP19, PGS10, RV16, RHS14a, ZG10a, ZG10b, BZN17, BS13, BDE^{+}14, DSF11, DCR18, GMMGVD12, Gla13, LLLV11, Mac11, Rei12, TKKE19]. scallop [WL13]. scan [BR16]. scapularis [DB11, WDL^{+}13, DDSDW13]. scar [PD16, YWP13]. scarce [BC15]. scattering [SHB^{+}17]. scenario [BP16]. scenarios [BHH^{+}14, PGGvL^{+}19, QF10, RMF^{+}18, SS15d]. scenes [BZ18]. SCFG [MWSM13]. SCFG-modeled [MWSM13]. scFv [MNSZ16]. schedule [MB12a, OT13]. schedules [BI10, FM10]. scheduling [BH13c, GSSBF18, LF19, MSBB13, NH19]. scheme [DZJR10, FP18, Mil11]. schistosomiasis [MCC^{+}17]. schizophrenia [SST19]. Schlögl [Fra16]. school [NTY16]. SCID [CS14b]. science [MPCTGJ^{+}15, Ros10a]. scissors [BMdA17, Mob10, NIiT18b, NiTI18, NIiT19]. Sclerosis [SHG16, BGCB12, KKS^{+}14]. scope [HL10, KCE^{+}11]. scores [DTY12, ZLW^{+}19]. scoring [DLL^{+}17, FP18, KHK^{+}17, MIH16, QLC^{+}18, SLDP13, WYX^{+}17]. Scorpion [DS10b]. Scoters [Luk14]. Scots [CRLH^{+}19]. scrambled [BKC^{+}16]. scratch [JSP^{+}16, JRB^{+}16]. screening [BYY^{+}11, GBC^{+}16, KJM17, Kha10, KMLT14, KPLT14, NM11, SBR16, TVA18]. screenings [LCT11]. sea [LLLV11, NSO15, SCS14, vHHKB14]. seabird [PFB10]. seagrass [Alv19]. seahorses [RVA^{+}11]. seal [HH10, SH17]. sealing [KHB^{+}18, KFS^{+}13, XWW^{+}19]. search [BH13a, CCAdS13, HG18a, LFB^{+}16, LDWW14, RSC14, RW15b, RPCW18, SM11a, VM12, WSW10]. searching [BC15, HS14c, TBQG14, XJ19]. seashells [MGC12]. season [HHE19]. Seasonal [CSLE11, TWS15, BI10, BI12, BG12, BM10, BTW11, DFM^{+}14, ECP^{+}16, GCG14, LKK13, OS13, OVKL14, PLF19, TA13, VR17]. Seasonality [MGT17, TF16, UD10, ABKS11, CB13a, CH18, HAK^{+}19, LHV16, OA11]. seasonally [GA15, LC17, TGZF11, US12]. seaweed [MJV16]. seaweeds [MR10]. second [ARM18, CW14, HK16, KK15, PVCEC18, RLM^{+}14, Zhe16]. second-hand [PVCEC18]. second-order [CW14]. Secondary [Alv19, ABKS11, CM17, FYZ15, HWHL15, IIKH16, KBK16, KZL14, KF15, LRL^{+}16, LJ10, LZMM15, PM14, STN19a, SAGC12, WDH^{+}16, ZKHL16]. secreted [LZ18, YGL^{+}10]. secretion [GTB10, NF16, PSS^{+}10, PSP^{+}12, PSS^{+}13, SW13b, SW19, WTC10]. secretory [Hua12]. SecretP [YGL^{+}10]. section [WCF12]. sectional [KKOM18]. sections [KCJ^{+}11]. sedentary [YSYI13]. sediment [JCLS^{+}11, JCLS^{+}21]. seed [BTG^{+}15, BS19b, GCvWE^{+}14, KKWA18, LHD^{+}17, MK14c, ST17a, TI13]. seed-bank [GCvWE^{+}14]. seeding [TI13]. seeds [Alv19, SIHO15]. seek [ZLT^{+}19]. segment [vE11]. segmentally [Sat19]. segmentation [CHL^{+}11]. segmented [HME12, Hua16]. segments [dSMP^{+}11]. segregation [LCTG15, RA13, Rad16]. SEIAR [MZAI19]. seismic [BJB18]. seizure [DSLR18]. seizures [HZG^{+}17, MGB17]. select [BWB11, Ghu18, LGPS17]. selectable [Mor11]. selected [DKP^{+}18, LT14, PCB14, PBKR13]. Selection [FGLS10, Gra18, LE19, Miy16, Miy17, OKVN18, Rey13, YA14, AH19, ACH18, AHP^{+}19, Ano19-50, AM11, BWB11, Bon12, BWL12, CL18, CB17a, CNDK15, CS15b, Cro18, Cro17, CB10c, Del12, DEG^{+}14, GCM14, Gar11, GL18b, Gri19, lGLLL14, GGCJ16, HTN14, HPvdB17, HG18b, HCS^{+}19, HT10, HT13, JSZ12, JZ17, JD16b, JSP^{+}16, JJ15, JW18, KKD18, KPK^{+}13, KHX^{+}19, KS14, KF11, LHD15, LZA^{+}19, LvBJ16, LTS^{+}16, Lio18, Lio24, LPF11, LDH^{+}14, Luo14, MS14a, MMH10, MSN14, MO12, MK16a, MFZ18, MB16, MB19, MM18, Mun12, NKM^{+}12, NPS10, NI19, NPB13, NPS17, NPS18, Orr19, OGO19, PD13, PGG^{+}15, RM17b, Rei12, RRÚJ19, RG12, SC10a, Sal10, SL10, SSM15, SFD12, Sim14, SD16, Soz13, Tak17b, TGL15, Tos11, TMHM12, Uit11, VDTF15, jWlGQ19, WYL^{+}19, WWB14, Woo10, Yam16a, YHZ14]. selection [ZFWK17, ZW16, vVGSE12, vVLS14, vV18, vTMW15]. selective [KGB13, MPLK14, SG15b, TA16]. selectively [HA15b]. selenol [KIH16]. Self [EdKM11, Kue16, LPB17, LS15b, LYM^{+}19, MDS16, PD13, QLZQ11, Rei12, RA18, VY12, ABD^{+}15, AY15, AIY16, CJ12b, CDD12, DPRSS11, Fan11, FWSG17, GMK18, IU18, JOM16, KHB^{+}18, KKG^{+}14, KFS^{+}13, Lar10, MMJB17, Mau15, MBBD13, Mic15b, OF10, Pai19, PCBMM12, RL15, Sel16, SHM^{+}18, SLS17, SG18, TDT16, VSLVB15, Vel17, WD12, ZCAB17, Coh15b, LPB18]. self- [OF10]. self-assembly [FWSG17]. Self-binding [LYM^{+}19]. self-complementary [Mic15b]. self-directed [ABD^{+}15]. Self-extinction [PD13]. self-hybridization [Sel16]. self-incompatibility [VSLVB15]. self-interactions [ZCAB17]. self-maintenance [PCBMM12]. self-matching [MBBD13]. Self-mechanical [Kue16]. self-organisation [KKG^{+}14, MMJB17, Pai19]. Self-organization [LS15b, MDS16, DPRSS11, SHM^{+}18]. Self-organized [VY12, CDD12, IU18]. self-organizing [GMK18]. self-paired [Fan11]. self-propel [RL15]. self-regulated [JOM16]. Self-regulation [EdKM11, WD12]. self-regulatory [Vel17]. self-replicating [AIY16, Mau15, SG18]. self-replicator [AY15]. self-repressing [SLS17]. self-sealing [KHB^{+}18, KFS^{+}13]. Self-similarity [QLZQ11, Rei12]. self-support [TDT16]. Self-sustaining [RA18]. self-thinning [Lar10]. Self-tolerance [LPB17, LPB18]. selfing [VDTF15, YYB^{+}19]. Selfish [CW14, ASS19, GCM17b, OO17a]. Selfishness [SS12b, ZSS10, SC12]. Semantic [ZHC10, ZT16]. semi [BS19b, SAKG13, VG13, ZNCM15]. semi-deserts [BS19b]. semi-mechanistic [SAKG13]. semi-stochastic [VG13]. semi-supervised [ZNCM15]. Seminal [KGM15b, WZ18]. senescence [DKS15, DBB14a, EBX17, WWB14]. sense [LLSO13, STH18]. sensing [BSL^{+}17, Chu17, FBK11, GSS19, KAN11, LDHD14, MKPVH16, NBW10, NBW11, ONO15, PLHL19, RAF^{+}14, RRC^{+}11]. sensitisation [HLTR19]. sensitive [CW11c, JZ17, LSP13, TSMB14, YI18a]. sensitivities [BV13, YIS^{+}17]. Sensitivity [GZFX14, MF15, WST15, AKNP18, CCS^{+}16, CNG^{+}12, CMN^{+}11, DJD10, DDSDW13, FDM^{+}11, FR17, KMM18, KLHS17, Lug15, MK11, MBBV14]. sensor [GT15]. sensors [KTN^{+}10]. Sensory [CV15, BA10]. sentences [FIS11, FIS16a]. sentimental [ZSW^{+}14]. separate [XY14]. separated [Di 12, PP14]. separation [PAK11, SNY^{+}13, YS14]. Sepsis [CA17]. septa [HWM^{+}16]. septal [HWMT17]. September [Ano10-29, Ano10-36, Ano11-34, Ano11-38, Ano12-29, Ano12-39, Ano13-39, Ano13-36, Ano14-37, Ano14-41, Ano15-45, Ano16-46, Ano17-31, Ano17-38, Ano18-47, Ano18-31, Ano19-48, Ano19-34]. Sequence [AP13b, CJ12a, DGL12, DSPM14, JM12, LLL^{+}17, LLCC12, MS14a, PDM17, AH12b, BMSEE14, CLL18, DLYZ11, DYS^{+}13, DZW10, ESW13, EJ16, Fer12, GKPB13, GGX^{+}10, dOGL10, dOGL13, GAB14, HA15a, HK11a, HKR^{+}19, JAM18, KCM^{+}11, KG15, Kin18, LGK^{+}09, LGK^{+}12, LL13a, LWL^{+}11, LlLYB13, MGO^{+}15, Miy17, MP14b, NGND12, NBL10, PMYHR12, RBMS17, SIK^{+}18, She11b, SSP15, SLT^{+}18, Tri10, WSH^{+}10, WSK^{+}19, WZY14, WS10, WGDZ18, XLL^{+}10, XSLZ16, YW13b, YCY14, YC15, YLF^{+}17, ZPHS11, ZLZ^{+}15b, ZLZ^{+}16, pZWZ^{+}16, ZNA^{+}16, dBJ11]. Sequence-based [DSPM14, JM12, MS14a, PDM17, HKR^{+}19, JAM18, MP14b, RBMS17, YC15]. Sequence-dependent [LLCC12]. sequence-derived [KCM^{+}11, ZNA^{+}16]. sequence-protein [dBJ11]. sequence-specific [Fer12]. Sequences [SBT^{+}18, AH18, Boz15, Bry13, CBL14, CK16, CL17b, DDLW10, DLYW13, FHW^{+}10, FS15a, FS16a, GKNT10, aHLZ^{+}12, HLHY17, HYZ^{+}15, HZWH10, HZLX11, HYW11, HY16, IA17, LLL^{+}17, LTP19, LHH12, MPS^{+}11, MP12, NMZ19, PM11b, QQL10, SFV16, SSCJ16, SQS^{+}12, SCS10, SCLC13, SM11b, Tak17a, TTP17, WTQL10, WRZ^{+}12, WZ17, WXC10, WZWY11, WRWX19, XM11, XWW^{+}14, YZW10, YZZ12, YW13a, YLC16, YYH^{+}14, Yin17, YC13, YCR^{+}15, ZMS11, ZZK14, ZLZ^{+}16, ZLB13, dF13]. Sequential [Lai18, CHL^{+}11, VBV10, WCS19, WSW10]. sequestration [Ber14, HSR12, PR18]. Serial [Lan17, SR14]. series [AIY16, KHRS10, LJM15, MR19, NMI18, SCLC13, Woh15, XSKA17, ZHAK14]. serine [IHNS16, Sel12b]. serotype [KM15]. Serum [VM16, SMM^{+}13]. server [GDPPSS^{+}11, SLW^{+}18a, WXC10]. service [EVRE14]. sesquiterpene [LCQ^{+}18]. set [CZW^{+}11, MP13b, NBL10, NLB14, SM18]. set-based [CZW^{+}11]. sets [GKNT10, MFKS13, SVB^{+}10]. setting [GAY^{+}11, SSG^{+}19]. settings [MHMM18, PT10, RK19, SPRF13]. settlement [BPG^{+}18, FMS^{+}10]. settling [MN12]. setup [FE11]. Seven [MB18c]. Several [DZW10, JRR^{+}12, MKBE17a, MKBE17b, RPD14, TD18]. severe [DD18, Sat18, SGDL12]. severity [ZZL19]. Sex [BS11, Del12, EHSR17, Gar10, NG11, AM10, Arg12, Chu10, GGM12, Jam11, Jam12, Jam13, JV14, Jam16b, KMHdlP10, LL18, MK16a, MYK17, MB10, MM15a, MFP^{+}14, MLBA12, MLBA13, NT14, PKH11, Rad16, Ram10, SYI17, Sek12, gSxFH^{+}12, TEY10, TEW12, TEY16, VMZK^{+}19, WKH16, YSST13, YYST13, YHI14, Yam16a, YI18a, YI18b, YTK10, YK11, YT12, ZP11]. Sex-biased [Gar10]. sex-determining [PKH11]. sex-specific [Sek12]. sex-structured [LL18]. sexes [XY14]. Sexual [LN19, YYST13, BHBR12, CHK16, Cro17, EHSR17, EBSW17, Fis19, Fre10, Gra14, KK12a, Mor13, MM18, NI19, SLL18, SSD11b, TCB13, WSRG18, ZRZ15]. sexually [HE16, HHA^{+}13, MK16a, MBC^{+}12a, TB18, YHI14, YZMY18, ZAL^{+}19].

SFM [YC15]. SGB [XXD^{+}17]. SGK1 [KSP^{+}14]. SH3 [HWW^{+}14]. Shade [DHK13, Veg10]. shadow [LJM^{+}17]. shall [GCM14]. Shannon [CDM^{+}14, Nem17, Pia19, RFME^{+}12, SSP15]. shape [BTR18, CdPD11, DMM^{+}14, KTO^{+}14, LKSM14, LSDD13, MBM11, PRSC11, PP17b, Spe15, WGO^{+}15, LZMM15]. shaped [BORA10, DNB12, KP15, UI13, Yam16a]. shapes [FZ14, GVSLG16, HMH14, IS17, MOSS15, TDKJ15]. shaping [BPLM12]. Shapley [WF17]. Shared [CGD13, IEN15, KB15, WL19, ZZCZ17]. sharing [BORA10, PEW18, TLW18]. sharks [IPMH12]. sharp [PJ13]. sharpness [ZKMB19]. Shear [FPG11, BB15, DPCM16, She10, THBM10, WYL^{+}18]. shedding [CMN^{+}11, HR12]. sheet [Mau15, ZZRZ11]. sheets [Han10]. shell [BB15, KBS^{+}19, KBS^{+}20, SCsMzX18, She10, SK11a, VH11a]. shells [LCG12, Nos14, NSS16]. shielding [Kar15]. shift [AN12, BRG^{+}12, BH13b, FL12, FL13, RG18, SI11b, SSH^{+}19, SFS13, SY15]. Shifted [MPL16]. Shifts [EVRE14, CVO^{+}15, HBT13, JAB18, KF15, LG13, PSV17, SI10b, Tou14, WS16, tBdR18]. Shigesada [YLH12]. shikimate [AEE19]. Shine [SGD^{+}16]. shock [MG10b, RZIR16]. shoot [TP14]. Short [FS15a, FS16a, WGS15, ABLH18, BB11b, JP19, MS12a, SG15b, TSB10, WMT15, WGC13, YW13a, Zol14, ZPM^{+}15]. short- [ABLH18, JP19]. Short-term [WGS15, BB11b]. shortages [Che18]. shortening [GSV11, RBKW19, SKK^{+}12]. shortest [BMV12, YZZ12, vTMW15]. shot [AS12, Lai18]. should [CT14b, PW18a]. show [Laz13]. showing [AGRRTR11]. shrimp [NR14]. shrinkage [Yat14]. Shrinking [IDM15]. shuffled [PPF17]. shunt [NAK^{+}11]. Shutters [DHK13]. shuttle [Mar12]. shuttling [SWR11]. sib [GZ19]. sib-pair [GZ19]. sibs [Jam16b]. SicA [LPS^{+}14]. sickness [For14]. side [Gra18, LE19, dM10b, dMP11]. sidearm [Sel14a]. sided [JZ17]. Siderophore [LvBJ16, FE11]. sighting [TLS^{+}13]. sightings [KDS18, NdMLLBLC13]. sigma [dEG11]. sigma-factor [dEG11]. sign [PTNKT11, VDTF15]. Signal [HM10b, JC10, VLPH17, ANK10, ANY14, CL13a, DFM^{+}19, GdGL17, HYSG19, HMM16, JAB18, JRMS12, JR17b, KFG15, KP16, LDHD14, LCGMH12, Lug15, MÇ15a, MÇ16b, NF16, OS11, PPM12, RW10, SAA10, Tsu19]. signal-integration [PPM12]. signaling [Akç12, BB11a, BJJR10, Chu10, FLCS^{+}15, FTS15, GM13, Gri11, HSiI^{+}19, HWD^{+}14, Ken19, KB19, LSMC18, Lee16b, LXZQ14, NS18, NSS^{+}11, PR18, PC10b, RSV10, SPS11, UI13, YFB^{+}12, Zhu11, ZXC^{+}10, ZG10c]. Signalling [DTGD19, AP17, FD11a, MBM11, MBKB13, MLL^{+}16, OTGT10, Par19, PSD^{+}18, SNC12, STX^{+}11, TC11, WPH^{+}12]. Signalling-modulated [DTGD19]. signals [BR13, HWGT17, MOSS15, Nak16b, SM14, SHBT14, SVSP15, SKK^{+}11, VH11a]. signature [CVO^{+}15, DCS14, TLY^{+}19]. Signatures [SHM^{+}18, TST^{+}13, Bar18, NS16, NKM^{+}12, NBBCC18, PPBD10, SSCJ16, WSK^{+}19, WDN^{+}19]. signed [cLCJ^{+}10]. Significance [MAH12, CB13b, GK17, iHM17, SWPC^{+}16]. significant [SDRA^{+}15, SPA18]. silencers [CWW18]. Silico [AEZR^{+}16, ACM16, AMFR^{+}16, ASRM15, ALM^{+}19, ACvKA10, BGCB12, CRH^{+}15, CAGM^{+}17, CGRRGM^{+}16, DBD15, EME^{+}16, GHJ14, HH14, Hal17, HCHI18, HNP18, HMMSRSD15, JSC^{+}16, KJM17, KRDJ15, LZL^{+}17, LKAJ18, LMM18, LZL^{+}19, MÇ15a, MÇ16b, MPZK16, MG15, NTC^{+}11, NRS^{+}16, NGJ^{+}14, PSG^{+}17, PMCS16, PEZ^{+}13, RRG^{+}10, RRG^{+}12, RAA^{+}16, RM19, SBR16, SDRA^{+}15, SGD^{+}16, SSRA16, SB15, TKTB18, WB15, ZNCM15, CMHM19]. Silk [RGCML10, FMS^{+}10, YLY15]. Silver [MHD18]. Simian [SBCR10, HII^{+}19]. simian/human [HII^{+}19]. similar [DG16]. similarities [CLL18, MJ11, Pav19]. Similarity [SCS10, YZZ12, CBP12, lDsWM^{+}15, GZT12, Hou13, HZWH10, IK15, LSDD13, Ots11, QLZQ11, QLWC19, Rei12, WTQL10, WSK^{+}19, WGDZ18, YZW10, YW13a, YCY14, ZZK14, ZT16, dASdC^{+}15]. similarity-essential [dASdC^{+}15]. Simple [BJPP12, Mar12, PDL^{+}17, SSME10, Tak16, ACdlRMR^{+}11, BRB18, BSR^{+}11, BLS^{+}12, DDLW10, DLYW13, DDSDW13, FBFM12, GI15, HS12, HZWH10, KF15, MGML10, NSO15, OKVN18, OSF11, PAA11, PCBMM12, PML18, QQJW13, RG18, RHT18, RAMS11, SA13, ST17b, TA10, Tal12, VK10, VLCM10, ZBA14, ZLY^{+}17, vVLS14]. simplest [JMZ13]. Simplex [KB10a, BYY^{+}11]. Simplex- [KB10a]. simplicial [ER18]. Simplification [Jia16, UBP12]. Simplified [Pat16, HM11a, HWHL15, LS10b, RLSM17]. simulate [ASL^{+}18, CMJD11]. simulated [BWL12, Don17]. Simulating [HNV^{+}16, HWGT15b, LLD^{+}17, LC16, NR14, POK^{+}12, HGGCR13, LKL13, LHPF18]. Simulation [DLL^{+}18, JL11, MB12a, SNOV11, WGH^{+}14, dlCGSA16, ADS^{+}19, BUC14, Boc12, BTK15, BMF^{+}18, Cox10, CT16, DHB15, DV12, FKMG15, FWSG17, GJ11, GHBI16, GKG^{+}18, HAuR^{+}18, ICG16, KMLT14, KPLT14, LSP13, LCA^{+}15, LHD^{+}17, NI19, OA15, PAOMV17, PRMH14, STNT17, SK15, SKD^{+}10, ST17b, TD17, TMS13, VF12, VG13, WWIG19, WMD16, WC10, XYZ15, ZLW14, ZZL^{+}11]. Simulations [ARB13, Laz13, SBB14, APW10, CK17, For10, KGP^{+}15, KA15, KTJ19, Lai18, LSSG10, MPP^{+}16a, MSN14, MAA^{+}17, NZZ19, SMS19, SMJS14, SRH19, Zha10a, ZZ18]. simulator [MZ17]. Simultaneous [Eft13, VBV10, AB16, BHBR12, KMHdlP10, YvBS18]. since [SS19]. sinensis [LCQ^{+}18, PWH^{+}13]. Single [AAAV14, Dol16, BFR14, BJOS13, BF16, BGD^{+}13, BAR14, EUM^{+}16, ESGA15, FLA^{+}19, HB12a, HY13, III17, KB10b, Lai18, LM15, Lei09, Lei10, LCG^{+}15, LW10a, LW10b, LAJC19, MV10, MB10, MG15, NS18, RB13, RHS14a, SKH17, SC11a, SSKS18, SSKS20, SLC12, SL19, TLY^{+}19, TST^{+}13, WL10, XWC11, ZZ10]. Single-cell [AAAV14]. single-dependent [MB10]. single-linkage [SKH17]. single-molecule [LAJC19]. single-peak [FLA^{+}19, LCG^{+}15]. single-shot [Lai18]. single-specie [LW10a, LW10b]. single-species [RB13]. singular [Lan17]. Sink [Gro17, SSI^{+}16, SSZB15]. sinking [DD10]. sinus [ASC16, SWPC^{+}16]. sinusoidal [WSLC14]. sinusoids [DLL^{+}18]. SIQS [LCT11]. SIR [BBR16, LY11, TGZF11, YZMY18, YS11, ZYJL18, ZTT18, dM10b]. SIR-SVS [LY11]. siRNA [KP16]. siRNA-based [KP16]. SIRWS [LHFM16]. SIS [HT19, KKYV18, LBZ18, ZLL17]. Sister [Fer12].

Sister-Chromatid [Fer12]. Sit [RW15b]. Sit-and-wait [RW15b]. site [AR17, CCY^{+}19, CK15, Cro18, DK13c, Gal10, HNL^{+}11, HG18a, JMH^{+}19, Kar12, NTC^{+}11, NPS10, NPS17, NPS18, QZY17, SMJS14]. site-selection [NPS17]. site-specific [CK15]. Sites [WZM19, AH18, AR17, BGCB12, CST16, CK15, CGRRGM^{+}16, DSPM14, FGAM12, GYWJ10, GRRG16, HG18a, HY13, HKR^{+}19, JRR^{+}12, JWS^{+}10, JLX^{+}16, JZ17, JCG15, JW18, KJH^{+}19, KB10a, KBK16, KIT^{+}16, LTSTD15, LL12, MSND12, MSRW16, MSSM18, NMZ19, QW12, SIK^{+}18, TTC19, Tun13, UT17, XWC11, XWD^{+}10, XWW^{+}14, ZZC^{+}17, ZNCM15, ZNC^{+}15, ZNA^{+}16, ZZN^{+}19]. situ [GCSP17, KFG^{+}14, METC12, NWB^{+}10, SGGY10]. SIV [Lev14]. six [CCY^{+}19, RBMS17, TMD14]. Sizable [TP10a]. Size [PCBMM12, RU16, SA17, Spe15, AMM16, AY15, AC12, BB15, Bar11, Bar16, BR15, CAG13, Cut15, DRPM17, EzMMH18, Fil10, Fil15, GSL13, GTB10, HAB11, HTT19, JL11, KSP12, KFG15, KG15, KDK14a, KDK14b, Lar10, Lee16a, LSY^{+}10, MJS11, MB18a, MP13a, Mor13, MI11c, NYSM12, OT13, OGA16, PIPB10, PN16, PN18, RBB16a, RGCML10, Sch14, SN18, SAB10, SJSG13, SDPC11, YK11, YT12, ZADB15, ZTT18, ZL18b, Zhu11, ZEJA11]. Size-dependent [RU16, SA17]. size-spectrum [MP13a]. size-structured [HAB11, RBB16a, ZADB15]. sizes [YHI14]. Skeletal [NF16, BRN15, BSS11, GHS18, GHS20, HGGCR13, KCZ^{+}19, LO15, LCGMH12, LHPF18, RGA^{+}10, SAD17, TSB10, WC10, XYN15]. skeleton [CS11b, GGG19, GBN14]. skeptics [SGA^{+}12]. skew [Non10, SR10, SRV11]. skill [Nak19]. skin [GBJ18, HPM^{+}17, KC13, PMD^{+}18, ZTK12]. skinned [Smi14]. skip [ZLCH18]. skip-and-resurgence [ZLCH18]. skull [CW11a, SCF^{+}12, TF15]. skunk [BLS^{+}12]. skyline [DM15]. slab [MKBE17b]. sleep [ARR14, FPR10, KAPR12, Pat16, PR17, PRK13, PFPR11, RTR^{+}13, RRL^{+}13, RLW^{+}14, ZKR15, ZLL^{+}12]. sleep-active [Pat16]. sleep-wake [PR17]. slender [Che16]. slender-body [Che16]. sliding [MZAI19]. slime [AABS16, Tre19]. slopes [GW12]. SLOSS [Tjø10]. Slow [SL19, WTZ^{+}11, ZKR15, BRB18, DOC17, Hor11b, ID14, LM15, LG13, MB18b, PB16, SZ15b, VWF19]. slow-diffusing [PB16]. Slow-wave [ZKR15]. slowdown [ATG12]. slows [Van17a, YRD17]. Sm [ZJS11]. Small [RL15, YSST13, Ano19-50, EzMMH18, Fue16, GY19, HB16b, HSR12, HM10b, ISM^{+}11, KFA12, Laz13, LPTR14, MZM^{+}16, TLL^{+}12, TDZ^{+}18]. Small-amplitude [RL15]. small-world [HM10b]. smallest [JLH^{+}15b]. Smallness [YI16]. smoking [CJL15]. smoking-tuberculosis [CJL15]. smooth [AJD11, AJC12, Don13, Kro11, MAH12, MH14, SS14, SKH11, SH16]. smoothed [IK16]. smoothing [SRD^{+}15]. SMOTE [ZZG^{+}16]. snail [XWW^{+}19]. snout [RVA^{+}11]. Snowdrift [Lai18, FNH10, IOY15, QMJW15, SPSP12]. SNP [HZL^{+}11]. SNPs [Kha10, LZL^{+}17, MPP^{+}16a]. soaring [vLSBBD11]. Social [AB11, MK14b, MPP^{+}18, RFME^{+}12, TSL12, AMFL10, AS12, CBBO16, CW14, DBD12, Déb17, HNR14, HS14b, HBL10, Kea16, KMMS17, LX10, MJI14, MR12, NO14a, NO14b, NBS^{+}13, NR16, NC15, PN18, PB18, PS12, RANO10, RMB15, RDD14, RDSD^{+}12, Ros10a, Sal10, SAB17, SSC13, TC12, TF14, TT10b, UI13, WGS10, WMD^{+}11, WB19, WLW19, WLW20, WL15, WL12c, ZGWL14, ZLB^{+}18]. sociality [PBA12, vVGA10]. societies [GSF13, KP12, LF13, NC15, Sal15, SC11b]. society [HCW18, TT19]. socio [DCR18, GAY^{+}11]. socio-demographic [DCR18]. socio-demography [GAY^{+}11]. sociological [SY11]. sockeye [SGCD14, SWD16]. sodium [BGCB12, PDG17, PBS^{+}12]. soft [BZA16, DS16, Kro10, WHWZ18]. soil [CMB^{+}12, CG11, DAVS11, DDS13, FWMC14, MCWF14]. soil-borne [CG11]. Solar [KTO^{+}14]. Solenopsis [AC12]. solid [BAM^{+}11, DS16, MCL19, MH13, PGHC12, SKD^{+}10, XLLF17, YLLL12]. solitary [OSN18]. solubility [XNJ^{+}13]. Soluble [LSMC18]. solute [SSvdM10]. solution [DZJR10, Mil11, OYA14, OSN18, VLCM10, ZS10]. solutions [CN12a, JBSFB12, LHFM16, MA11a, MRS14, Ou19, Pie10, ZM10]. solvation [KF10]. Somatic [PPW16]. Some [BAM^{+}11, Cho11, Di 17b, DMP15, MP16, PG18b, PBRW11, Ser16, Gia13, LMM^{+}13, LT14, NNK^{+}15, dS15a]. song [BORA10, GZT12]. Sonication [Ste12]. Sonication-induced [Ste12]. sorties [KTO^{+}14]. sorting [CSBK15, bCRS17, CVO^{+}15, Eft13, MPK^{+}12, PWE15, TF17, Tre19, VBD10]. SOS [BCKS13, BKPV15]. SOS-induced [BCKS13]. SOS-response [BKPV15]. Sottas [AD15]. sound [DWG^{+}19]. source [ACM16, BL15b, MGGM10a, SSI^{+}16, SSZB15, SC13a, SC13b, YZ19]. Sources [JB19b, CB16, DNB12, PCC15, SZ10, TD17]. South [PDW11, SKPK17, WCJ11, DB18, Kur14, LJ15]. southern [MFP^{+}14]. space [AH12b, AH16, AWRD15, ATB14, AP13b, Bau18, BP15, BLP10, BLS17a, CCAdS13, Cut15, DAVS11, Gau11, GD16, GNB^{+}13, HT19, HTK14, HMW16, Kin18, LWLM18, LAPK14, Mal15, MDS16, MWSM13, MBCM12, MS14c, MSIR10, PBR17b, PD16, RD14, RBB16b, Spe15, TAI^{+}18, WO11, WPA17, Wil11, YDC^{+}13]. space-fractional-order [PD16]. space-time [MDS16, MS14c]. spaced [JCG16, MMH^{+}12, Tun13]. spaces [BLS17a, Cre10, GKNT10, HTK14, IMD16]. spacing [RG18, Smi14]. spacings [Smi11]. Spain [VPGDP^{+}14]. spark [SVCS10]. Sparse [WZY14, PBEI12]. Spatial [ACMK12, Ayd18, CLPZ14, CNL14, CVO^{+}15, CPPLAB18, GB12, GABM12, JAB18, LLLV11, MMC19a, MKKS14, MGCL13, NvHM13, POW18, RSD12, SS15a, ŚKSRW16, VDD^{+}17, YvBS18, YBT^{+}17, BBT^{+}15, BZ18, BH13a, BCPG18, BTK15, CN10, CLA17, CSBK15, CF11, DE17, DCN13, Eft13, EPP19, FKV19, GDS11, HT19, HM11b, HASM17, HB13b, IGL^{+}12, IRS11, JH11, JC16, KDST15, Kar12, LJ15, LFB^{+}16, LLH^{+}13, Lio16, LXJ15, LVLC18, MCL19, MMAS13, MBS19, MKMG^{+}14, MCWF14, MS16, MB19, MAL^{+}11, POW17, PTC13, PRM12, RDP16, RAMS11, RM17d, SDJ^{+}15, SCS15, SLL18, SJSK18, SLJ^{+}10, SS11, SY12, SS12b, SSC13, TB16, TGL15, VY12, WH11, WKB13, YISG14, YSI18]. Spatial-Stochastic [MMC19a]. spatial-temporal [LJ15]. Spatially [TKB^{+}18, AdVG15, AGGME18, BI19, BLS^{+}17b, CCS15, GA16, HDNB10, LDHD14, NdORC11, PNL15, PDT^{+}10, SRS^{+}15, SMdA16, SSS15, SB17, SS11, TM15b, VZ19, ZCS^{+}15]. spatially-averaged [ZCS^{+}15]. spatially-explicit [TM15b]. spatially-extended [LDHD14]. spatially-structured [VZ19]. Spatio [HLTR19, MS10a, QJR^{+}16, STX^{+}11, TC11, ABB^{+}19, AGPK13, CH16, MTdS^{+}16, PM10a, TP10c, ZVMB10]. Spatio-Genetic [HLTR19]. Spatio-temporal [MS10a, QJR^{+}16, STX^{+}11, TC11, ABB^{+}19, AGPK13, CH16, MTdS^{+}16, PM10a, TP10c, ZVMB10]. Spatiotemporal [ARD14, DHRA10, PLHL19, Ros10b, d’O12, BLS^{+}12, HH16, HHD^{+}16, Sht17]. Spatiotemporally [HRDL14, HRD14, Zhu11, IH17]. spatula [FPG11]. spatula-like [FPG11]. spawning [HCHI18, MDB12b]. speakers [SN16]. special [Cle16, GKK13, TK19]. specialist [RT19]. specialization [NP11, NP13, RD17, WF18]. Specialized [KMMS17, LIPD12, IEN15, SKH17]. speciation [AWRD15, ACvKA10, BdABY13, KG15, Mar12, MYN^{+}15, SIVH17, SS12a, SGW^{+}18, WCPF15, WK18b, YI16, YI17]. specie [AT10, LW10a, LW10b]. Species [CC11a, CCB11a, DMSW10, HSM17, LLH^{+}13, LZT12, LZT13, PC13, RM17c, Sel14b, SCF^{+}18, TLM14, YLW^{+}14, dBJ11, AP19, ADR11a, AHP^{+}19, ACvKA10, AMRRI10, Bac15, BJOS13, BWY^{+}17, BCF16b, BJPP12, BP16, CBL14, CHD^{+}10, CCA17, CMCS18, CW11a, bCR19, CH11, CK15, Cre10, DBD15, DSA^{+}16, DZE11, FB12, FVC15, GDF18, GK16, GZT12, GABM12, GDJC11, Han12, HF17, HMM17, IC11, JYZ18, KDS18, LSDD13, cLCJ^{+}10, LRGA13, MG14b, Mal17, MPBS17, MYN^{+}15, MB15, MBLC17, MI11b, MI11c, NBS^{+}13, NvHM13, NdlPZA10, PC10a, PB16, PPTC19, RB13, RR12, Shi19, SLL18, SLHS13, Su16, SKS^{+}19, SBM^{+}12, SA14, TIS10, TGLK19, TCR13, TD18, TSS16, Tjø10, VDTF15, YSYI13, ZGG^{+}19, ZNA^{+}16, ZSH^{+}16, ZEJA11, ZYD15, dS15a, vLBJ^{+}13]. species-abundance [FVC15]. species-area [GABM12, ZGG^{+}19]. species-diversity [Tjø10]. species-specific [CBL14, CHD^{+}10, JYZ18, ZNA^{+}16]. specific [CBL14, CHD^{+}10, CK15, CDK11, CTL^{+}15, DL15a, DLL^{+}17, DBBM11, Edu16, Fer12, FMLM12, GMB11, GKG^{+}18, HZG^{+}17, HM13, IK15, JRR^{+}12, JYZ18, JD16a, JD17, JAHKH12, JP13, JK12, KHK^{+}17, LCA^{+}15, MDE11, MDMG14, NiTI18, NK18, OF10, PBYB18, QLC^{+}18, RMSTG13, RW12, SFMS16, Sek12, SLDP13, SSKK17, TDHC^{+}18, WYX^{+}17, WCC14, XDD^{+}15, XJ19, ZNA^{+}16, dEG11, dBJ11]. specification [ANMH11, BKL14, SCÁBM10, SCÁBM11, ZZX^{+}16]. specificity [ZZ10]. specifying [MBKB13, MLL^{+}16]. spectra [NSK13, SM17b]. Spectral [HRD14, BWS10, BTK15, CJ12b, SPMGR10, TP17, WGDZ18]. spectroscopy [SKF16]. Spectrum [CMM19, GSL13, HYZ^{+}15, LAPK14, MP13a, PZLF19, VHM^{+}17, WY13, ZDY11]. speculating [Mal10]. Speed [FOT^{+}15, GFM^{+}19, WHHS15]. speeding [SSF15]. speeds [BSL^{+}17, Fue16, GGBW14, GWBG14, MMG^{+}16, NP15, dRFI15]. spent [PTC13]. Sperm [PIPB10, CKBCG12, Ish16, ICG16, IGG^{+}18, OSF11, SOCF14]. spermathecal [FKMG15]. spermatozoa [HB16a, KG12, LXZQ14]. spermatozoa-oocyte [HB16a]. spermatozoan [IG14]. sphere [KKGN16, SGKCP19]. spheres [ICG16]. spherical [KMM18, SCsMzX18]. spheroid [AP13a, BPF^{+}19]. spheroids [BFGS10, DFG^{+}18a, HHA17, MFL^{+}18, TDSM12]. sphingomyelin [TWH^{+}13]. spider [DEG^{+}14, Mac10, QJR^{+}16, RGCML10]. spike [MKBE17b, Opr10, WCC13]. spiking [BPF15a, BPF15b, Fer11, Hor11b]. spillover [ML11a]. spillovers [VABS18]. spinal [ISZ18]. spindle [ARR14, BZL17]. spine [LSP13]. spiny [RKMG12]. spiral [RLSM17]. spiraled [FKMG15]. spiriferide [SK11a]. spite [KM12]. spiteful [GCM14]. splice [MSRW16]. spliced [BLZ^{+}19, ZLB13]. splicing [QYO10]. split [CS15a, Di 12, Di 14b]. Spodoptera [KSK^{+}17]. sponges [DHBS19]. Spontaneous [CBP12, Kon11b, CP14, CKB19, KB10b, MTdS^{+}16, OS11, SN12, TK10a]. spore [UI13]. spot [HTM15]. spots [CZW^{+}11, ZK18]. Spotted [PLF18]. spp [ADCG14]. spread [AZOLVH18, AL17, AP19, BJPP12, CDO17, CL17a, Cha18, CSLE11, CMN^{+}11, DK13a, HLTW14, HHT^{+}19, KMCJ17, Kle14, KF11, LS10a, LCCK12, LCJ18, LK15, LL18, LCBO^{+}12, MPS16, MGC13, MS12b, MS12c, MPLK14, MLBA12, MLBA13, PSL^{+}10, SHH15, SZ15b, VMZK^{+}19, WHT17, WDH^{+}11, XLSF19, ZJY19, ZVMB10]. Spreading [GGBW14, GWBG14, BHKR14, CGW18, FS16c, HPM^{+}17, KPD18, KPD19, LLJ18, MSM^{+}14, MBS19, PTC13, RPP16, SHvHB16, WMCL18, ZCXF13]. SPrenylC [HKR^{+}19]. SPrenylC-PseAAC [HKR^{+}19]. spring [WGH^{+}14]. squamous [MGM13b]. Src [HWW^{+}14]. SSBP [KGP^{+}15]. SSR [NNK^{+}15]. SST [CH16]. Stability [CL11, Gou16, LR16, MK14d, NPS10, SX12, TF17, VL11, WM17, ALH16, BS11, BBJDS11, BR13, Cha17, CLP11, Cre10, DKR16, DBM^{+}18, DHB15, Ezo12, FB18a, FS11b, FR13a, GAL11b, GCM17a, GCZ^{+}12, HGHL10, HK10, HYQ^{+}16, IG14, IvdSW16, Jam12, Kři10, LS14a, LS19, Lun13, MSZW19, Mil11, Miy16, Miy17, MD16, MKPVH16, NWZ15, Opr10, PJ18, Pav14, PMR11, PLN14, PDG12, PJL14, PSG^{+}17, PKH11, PML18, RW16, RCH14, RS10, RBMPP^{+}13, RM13, SW18, SS15c, SS18, SBS17, THM10, TRG16, TRJD19, WL12a, WL12b, Wal16, WR14, XDC^{+}11, XS13, ZCT18, Zha10a, Zha11, ZW16, ZWG10]. Stability-based [TF17]. stability-complexity [THM10]. stability/foldability [Miy17]. stabilized [LIPD12, NiTI18, PC13]. stabilizes [AS13, Smi14]. Stabilizing [SSD13, APS^{+}13, HMMSRSD15, RAMS11, TMD14]. stabilty [Ayd18]. Stable [HWGT17, Kar12, MN14, OO17b, RRÚJ19, BT17, BI12, Bro13, BHBR12, Cor16, CMGN17, DWL^{+}14, DI10a, Kis16, KM12, LFZN11, ML09, ML12a, Mit17, Rem15, SMG18, SKS^{+}19, TM14, VA10, XKCG15, ZLY^{+}17]. stably [ATRR10]. stack [FIS11, FIS16a]. stacking [Tri10, Zol14]. staff [HLTW14]. stage [BWY^{+}17, CKZ^{+}17, CT18a, CEP14, GWBG14, HYA14, JSWY19, JMK^{+}17, KK12a, Lab16, LP12b, LZTD18, MBP16, NTED^{+}19, RC16, SvK12, SMA11, TRJD19, WcW14, WTC16, YD14, ZLT^{+}19, ZK10]. stage-dependent [RC16]. stage-structure [BWY^{+}17, LZTD18]. stage-structured [CT18a, JMK^{+}17, LP12b, MBP16, RC16, SvK12, TRJD19]. stages [BAM^{+}11, FD17, HVM13, MM12, SBK16, WRC^{+}19]. staging [KKOM18]. stakes [CT14b]. stalk [UI13]. stalk/spore [UI13]. stalking [BR12a]. Stand [AdlPLMZ13, Lar10]. standard [BWM^{+}19, LAPK14, LPH19, Mas16, OUMA10, SSS13, ZGW16]. standing [SNCM12]. Staphylococcus [PDW10]. star [FBAPMD13, FLCS^{+}15, LDWW14, MYOS14, NIiT18a]. Start [RGP13, KK12b, Tox18]. Start/stop [RGP13]. starting [TSB10]. startle [Bac15]. starvation [AK15, HMH14, MBC^{+}12b]. starvation-predation [HMH14]. STAT [RG12]. state [AOM19, BTF19, BKSW16, BCPG18, CN12a, CMGN17, Cut15, DEK15, Dim17, Ell15, FTPN10, FT12, GGX^{+}10, JRG14a, JRG14b, JRG14c, KTI18, KL11, Kla10, LRA^{+}13, Lüt16, MBLV10, MBCM12, MS14d, MCCC^{+}10, PRK13, RFS^{+}15, RR16, RCL^{+}10, RCD16, SBCR10, SKD^{+}10, ZTT18, ZL18b]. state-dependent [ZTT18]. state-discrete [SKD^{+}10]. states [BVK10, BDR10, Bro13, IvdSW16, Jia16, LSMP14, MS18, MS10b, NWWL17, OABI12, RNN15, Sch19c, SKS^{+}19, TM14, UBP12, VA10, XGZ17, XNW17]. Static [MMRCCC10, FD17, PBEI12]. Stationary [Sch18, Sch21, ALH10, Kae17, Kär11, LS13a, SC11b, SVCS10, ZWG10]. Statistical [AM14, Cro17, GLZ^{+}13, Ing16, LCG^{+}15, MSS10, NdMLLBLC13, SS19, ABD^{+}15, AD16, AAJGCD15, BK19, BORA10, BRK19, CLW12, GGR11, KJH^{+}19, KCD11a, KCD11b, LKSM14, Mas18, NGND12, OGE10, PDW11, RSR11, SSKS18, SSKS20, Wal18, WH14, YW13b, ZMC^{+}18]. statistical-reaction-diffusion [RSR11]. statistically [SSS13, SDRA^{+}15]. statistics [LWL^{+}11, ZRGW19, vVGSE12]. status [DBD12, JS15]. staying [GN14, ON14, TTN13]. Steady [CN12a, IvdSW16, KTI18, BKSW16, DEK15, Dim17, FTPN10, FT12, JRG14a, JRG14b, JRG14c, Lüt16, MS10b, RR16, SBCR10]. Steady-state [CN12a, KTI18, Dim17, Lüt16, RR16]. stearoyl [PWH^{+}13]. Stegomyia [ASF^{+}15]. stellate [GWW^{+}10, TWTA^{+}18]. Steller [LLLV11]. Stem [BGL^{+}19, Hor13, HMJB15, BZL17, CDC18, CMS16, DBBW09, DBBW11, DBM^{+}18, EUM^{+}16, JMC^{+}10, Kar15, Mal18, RJSC18, ST16, TRM^{+}14, TK10c, VKKA12, VW10, Wal12, YAK17, ZZL^{+}11, dSdS13]. stenosis [Fok12, MP19]. stent [MAR^{+}17c]. step [Das18, NMZ19]. stepping [AP19, EKFV18, SV17]. stepping-stone [EKFV18, SV17]. Steps [LCCB11, HKR^{+}19]. stepwise [CJKR10]. stereochemical [RMRC^{+}16]. sterile [IH17, LBGW13]. sterility [YYB^{+}19]. sterilizing [JB18b]. steroid [NCS^{+}13, ZP11]. steroidogenesis [HH14, HCHI18]. steroids [WZ12]. sterol [BSM^{+}14]. STI [MLBA13, Cor18, MLBA12]. stick [MR12]. stiffness [AJC12, BSP18, DAA17, FS11a, SS15b, SS18, SHW16]. stiffnesses [SS15c]. stigmas [RSH19]. still [Zhe15]. STIM [SAB^{+}19]. STIM-ORAI [SAB^{+}19]. stimulation [FW15, FR13b, FR14, Kar16, WPH^{+}12]. stimulations [XNW17]. stimulator [AANF16]. stimuli [DHT19, VBVD17]. Stimulus [MAS15b, Str15, CPS19, Cox10, JP13, TKTB18, WSLC14].

Stimulus-dependent [MAS15b]. Stimulus-response [Str15]. stinginess [FZW17]. Stochastic [BM10, BDLR12, CCA17, CK16, DLEMP19, FMLM12, GMOP12, GL18a, GN14, GSV11, GL12b, IDI11, JP15, KKYV18, Kun16, LTZ19, LL13b, LPH19, LSSG10, MMC19a, MH11, MRPN12, MG10b, Oht10, PAK11, PAOMV17, RAD14, STA15, SC11a, SR14, SZ18, SLvdBMP10, VABS18, WR14, ZZS19, dlCGSA16, vE11, AGRRTR11, ATG12, BFR14, BJJR10, BR14, BCFR10, CCF^{+}14, CdMB10, CH16, Chu17, Fil10, Gal11a, GCG14, GMBK14, GAV15, HMM16, HK16, IOK19, IGL^{+}12, Jab10, KMA10, KTN^{+}10, KAKK19, LDF^{+}11a, LdLK11, LW10a, LW10b, LLL13, LPTR14, LJW^{+}16, LDM16, MCL19, MK19, MBCM12, NBC16a, NBC16b, OT13, OOY16, OAJK10, OBHS19, PR18, PCHM10, PMW12, PGF11, RBKW19, Ros10b, RT15, RM13, SW16, SI12, SW17, SSLB15, SB19, SLS17, TCR13, TD18, UK17, Veg10, VG13, WMPF^{+}15, WL10, XXD^{+}17]. stochastic [YD15, ZS10, ZZ10, ZWG10, ZMC^{+}18, ZSS10, vdBD10]. stochastic-deterministic [IGL^{+}12, RBKW19]. stochastically [EB15, GW15, Mul11, Mul12, MHF^{+}13]. Stochasticity [BF15, Lei09, Lei10, TDW16, ACH18, BGM19, BLNR15, BFW10, DSA^{+}16, Kam11, LP14, MPM10, MGCL13, Oku15, PDSP13, PM11a, PCBMM12, PR13, STN^{+}19b]. Stocking [MNH^{+}12]. stockpile [HSLW16]. stocks [DHP11, NBA^{+}18, TM14]. stoichiometric [FTPN10, PPW16]. stomach [KP12, LPGTBMA19, SK16a]. stomatal [KC14, LDM16]. stone [EKFV18, SV17]. stones [AP19]. stony [KRSD11]. Stop [MSBB13, KK12b, RGP13]. storage [GS13a, GS14, GW14, Gro17, KF10, MK14a, MGT17]. store [DHT19]. store/recall [DHT19]. stores [TvMG16]. STR [AEM14, AEM13]. Strain [TKTB18, ABKS11, ABN15, BDR10, CDK11, CP11, Cox10, DLMK12, DGD^{+}11, GVSLG16, GKG^{+}18, JCW13, PCC15, SA17, SCC11, SR12, XYN15, ZN18]. strain- [GKG^{+}18]. strain-specific [CDK11]. strains [BC19, BTW11, MSE^{+}14, OAS10, OS13, PB18, PDW10, SKH11]. strand [BKL^{+}15, Kae17, KDL16, SNS17, ZZRZ11, Zha15]. strand-symmetric [Kae17]. stranded [Kue16]. strands [CJ12a, ZDG^{+}10]. strangling [Oka15]. strategic [BBR12, BC13, HMH14]. Strategies [MI19, VKKA12, AK13, AKS^{+}19, AMFL10, BFH^{+}15, BC15, BB10, Bos12, BI19, CLA17, Cha18, CZ14, CSS^{+}18, CFF11, CG11, CBGS18, DG16, DI10a, EL12, GT11b, GCG14, GBRS19, GW19, HS15, HWTN15, HHE19, HHS^{+}10, HS14c, IM18, KLJ17, KdlRJ18, KKYV18, LCC15, LC17, LY11, LA15, LPvSP11, LPvSP12, MvdDW13, MFMPS10, MGT17, MG10a, ML10, MPLK14, NW10, OGA16, PW19, PBSM19, RK19, SSM^{+}18, SMC^{+}13, SNCM12, UI10, Ush16, WCJ11, WSW10, WHYMG17, ZLW18]. Strategy [GT11b, LXS15, Woo10, Ber14, CW15a, CJL15, DWL^{+}14, GVCG12, Has14, HRHAAA15, ITO16, KOS13, KJ19, LS15a, LL18, LFW^{+}15, LDJW16, LHD^{+}17, MW14, MR19, NPS17, QMJW15, SI11a, ST17a, SC10b, SSRA16, VCF^{+}19, ZWC^{+}13, ZWG10]. stratification [SM17b, cSGFB17]. stratified [MYLK11]. stratigraphic [SGW^{+}18]. stratum [SSvdM10]. streamline [KSM^{+}15]. streams [MN12]. Strehler [TMF17]. strength [JZZK11, KL13, RR12, VL11]. Strengthening [DGC11]. Stress [BIMC17, ASK17, AAJ15, BSP18, BYM^{+}18, CPF^{+}13, CFGRB17, CB10b, CDGV10, DPCM16, DBJ12, FM15b, GRCdL16, GTSP16, GA14, KHH10, KMN16, LLTP19, Mal15, MR15, MBPS17, NF14a, REC10, SGAM12, SWSLMJ19, SAI14, XYN15]. stress-driven [CFGRB17]. Stress-mediated [BIMC17]. stress-strain [XYN15]. stressed [BHBR12, DLMK12, OBAI11]. stresses [Ren13, ZR16]. stressors [DL15a]. stretch [MSA^{+}16]. stretch-induced [MSA^{+}16]. stretches [TSB10]. stretching [KSK^{+}11, KHH10]. striatus [SKSO12]. stroma [GCSP17, Kar16]. stromal [MGGM10a, MGGM10b, NJP18]. Strong [FS12, GCvWE^{+}14, Kur17, DGC11, JB19a]. stronger [HNP18]. strongly [Ala15, Kon11b, PP17b]. Structural [ÁTC^{+}14, CNG^{+}12, FDM^{+}11, KJSS10, KF12, MPP^{+}16a, PBKR13, SPG^{+}18, UD10, ZKHL16, AKNP18, ASRM15, AA16, CT18b, CzST^{+}18, CM17, DPRSS11, DYQ^{+}14, Don17, EBSW17, FS11a, HH14, IK15, KZL14, LJ10, LCQ^{+}18, LKAJ18, MPCTGJ^{+}15, MF15, OYY16, PDB^{+}15, PPC^{+}17, RSD^{+}16, SBR16, SSSD17, SG15a, Smi14, SLT^{+}18, USF^{+}18, WM17, YX15, Zha10a, Zha11, ZLY14, ZZK14, ZLT^{+}19, dBJ11]. structuration [Nak19]. Structure [KKM12, KI15, dBB14b, AAGO^{+}17, ACMK12, AP13b, BBT^{+}15, BKR14, BZ18, BSP18, BJO^{+}16, BWY^{+}17, BHKR14, BORA10, CTA15, CLC11, CVPCV^{+}15, CV19, DHB15, ESW13, EVRE14, FW15, FLM18, FGPR10, FYZ15, GGQ^{+}12, GDS11, GHBI16, GB12, GGM12, GCZ^{+}12, GHS18, GHS20, HTK14, HWHL15, IIKH16, KDST15, KKGN16, KGP^{+}15, KBK16, KZL14, LS15a, LLS^{+}15, LL12, LL13a, LGBT15, LMJ^{+}16, cLCJ^{+}10, LZMM15, LZTD18, MS10b, MGM13b, MHH13, MK14d, NBL14, NGS^{+}16, OAJK10, PM14, PBEI12, QQL10, RW15a, RK18, RBHK14, RKJ^{+}11, SAGC12, SJSK18, SDPC11, TA10, WSH^{+}10, WSLC14, WCPF15, WDH^{+}16, WKB13, WZLW15, WGO^{+}15, XLSF19, YISG14, YSH^{+}14, ZGZ^{+}10, ZZRZ11, ZKHL16, ZL18b, dBJ11, vV11a]. structure-aided [WDH^{+}16]. structure-antioxidant [LL13a]. Structure-function [dBB14b]. structure-sequence [WSH^{+}10]. structured [AdlPLMZ13, BV12, CT18a, Déb17, FBM19, GvVT14, GWBG14, Gra15, GMR16, HAB11, HI12, JMK^{+}17, KCS16, KFB18, KSPA17, LM11, LS16b, LW15, LL18, Lio18, Lio24, LP12b, LCDH15, Mac14, MBP16, NTED^{+}19, NP11, NPS18, OT13, PNL15, PTB12, PP12, RBB16a, RC16, RRR15, SPSP12, SvK12, SMdA16, SKS15, Tak16, TLCZ12, TRJD19, VZ19, WL14, WTC16, WL19, YLW^{+}14, ZADB15, ZLW18, dTKDV15]. structures [CAA11, CM17, EG12, FWSG17, FYZ15, GZY12, HY10, HBW^{+}11, Kee10, KCJ^{+}11, KF15, MDVT10, Mas18, NHTS14, SSI^{+}16, Str13, ZSW11, ZZ14a]. structuring [LG10a]. struggle [LM14]. STTNC [AH18]. stuck [Cza14]. studied [BHBH11]. Studies [Zha10a, ZZRZ11, BWS10, CW11c, KJM17, KSK^{+}17, LSMP14, MTE15, MBF^{+}15, PS13, PEZ^{+}13, RMM^{+}16b, TWR^{+}18, TAR16, TMS13, YNY^{+}10, Zha11, ZZ14a, ZK10]. studiosus [QJR^{+}16].

Study [DYS^{+}13, RNVP10, SM10a, ASK17, AHD^{+}18, AKR^{+}18, AABS16, ALH10, ACvKA10, BRA15, BC19, BB15, BZ18, BLV18, BD10a, Boc12, BVK10, BR12b, CFF12, CPS10, Che19, CH16, Cle16, CDGV10, Cro19, DRW14, DCL18, DK13c, DGD^{+}11, EJ17, EJ19, GKMC10, GKNT10, GHJ14, GDPPSS^{+}11, GXF13, HHJR11, HI19, HTM15, HWW^{+}14, HME12, HM11a, Hir16, HWGT17, HM15, iITIM15, IG14, ICG16, JL11, JRG14a, JRG14b, JRG14c, KKD18, KBL15, KRR14, KCJ^{+}11, KTJ19, LBMG16, LWLM18, LHW14, LS10b, LCQ^{+}18, LS10c, LCGMH12, MFM16, MRPH17, MS18, MPS^{+}11, MBC^{+}12b, NTC^{+}11, NZZ19, NI19, OBK^{+}11, PCL^{+}15, PRV^{+}14, PVGV19, PMKS^{+}15, RCH14, RCL^{+}10, RP18, RSV10, SPSM15, STNT17, Sch17, Sch23, SSG^{+}19, SG15a, SGDL12, SRS18, SKD^{+}10, SCS10, SAI14, SIHO15, TDZ^{+}18, TA16, VG13, WWIG19]. study [WC10, Wu14, XLG^{+}15, YSST13, ZSH^{+}16, ZZL^{+}11]. Studying [ARG14, BPM^{+}12, DHK13, KFG^{+}14, Kon17, Shu13]. Stylophora [KRSD11]. Sub [USF^{+}18, WN19, LLW^{+}18, MS16, ZLW^{+}19]. Sub- [WN19]. sub-Golgi [ZLW^{+}19]. sub-network [LLW^{+}18]. sub-Saharan [MS16]. Subcellular [DTY12, SRD^{+}15, AAJGCD15, CXC18, DHS^{+}15, FL12, HLS^{+}12, LLW^{+}18, LTL^{+}15, Mei12b, Mei13b, SC10b, STG19, TKK14, USF^{+}18, VLCT19, WMK13, WMK14, WMK15, XWC11, YLL^{+}14, ZL18a, ZD18]. subchloroplast [HY13, LWW14]. subcutaneous [MYK^{+}11, SAI14]. subdiffusion [SF15]. subdivided [PBR17a]. subgraph [LCL14]. subject [LPTR14, O’F11, SVB^{+}10, SLHS13, ZMW10]. subjected [GRCdL16, ZG19]. subjective [FPR10]. subjects [BRK16]. sublocalization [PWHW16]. submerged [STI13]. submitochondria [Mei12a, Mei13a, ZMS11]. submitochondrial [JD17, QLC^{+}18]. submitted [LSMP14]. suboptimal [FYZ15]. suboptimality [Jam15]. subordinate [NBS^{+}13, SYI17]. subpopulation [DS19a, NiTI18, SSZB15]. subpopulation-specific [NiTI18]. subpopulations [DGJ15, MK19]. subsample [MSK10]. subsequent [WK17]. subsidizing [KTJ19]. subsidy [EKFV18, LHV16, SV17]. subsp [AMSSG16, AHMA^{+}19, CRH^{+}15]. subspace [RAF^{+}19]. Substitution [AKR^{+}18, AD15, ADS^{+}19, BM11, CK15, DKP^{+}18, HS11, Orr19, ZPdFJ19]. substitutions [CzST^{+}18]. Substrate [BSP18, KPS17, KHH^{+}17, Aya12, KGM^{+}15a, Sar10, Str15, VACGF17, WL12c]. substrate-dependent [Aya12]. substrates [Sar10]. subsystems [MDB12a]. subterranean [JL11]. subtilis [BLP10, ZCSRG12]. Subtle [VK10]. subtree [BLS17a]. subtype [HZL^{+}13]. subunit [ZJS11]. Suc [DLL^{+}17]. success [ASGG17, BM14, BPG^{+}18, BAR14, EA17, FS12, GZT12, HH10, LYF^{+}15, MvdDW13]. successful [MFN^{+}18, OS13]. succession [Cam11]. Successional [Kla10]. successive [SSvdM10]. succinylation [DLL^{+}17, JLX^{+}16, ZNCM15]. succulent [KFS^{+}13]. Suction [VDH^{+}15, RVA^{+}11]. Sudan [OL19]. sudden [BIMC17]. sufficiency [vV18]. Sufficient [TL15, CHL^{+}11, FT12, GC16, Tox18]. sufficiently [XJ19]. sugar [JBSFB12]. sugarcane [EJ16]. sugars [OHWS18]. suggest [PEZ^{+}13, SBB14]. Suggested [dF13, Di 16]. suggests [DAA17, Kun16, MPP^{+}15, RRR15, SBB18, SH17, SM11b]. SUI [Yan10a]. suicide [JDSPK15, NPB13]. suitable [YW13a]. sulcus [SK11a]. sulfate [CZT^{+}16, SABB15]. sulfenylation [JZ17, JW18, WZM19]. SulfPred [JZ17, WZM19]. super [FWSG17, KF15]. super-hydrophobic [FWSG17]. super-secondary [KF15]. supercritical [NYSM12, WN19]. superficial [BMGC11a, BMGC11b]. Superinfection [PGPD19]. superior [Mas16]. Superiority [ZZ10]. superovulation [YD14]. superoxide [DWM15, Kun16]. superspreading [LCJ16]. superstars [JLH15a]. superstition [HLI15, TI12a]. superstructure [RFdL15]. supertree [MP16]. supervised [GMK18, NNK^{+}15, ZNCM15]. supplemental [JMH^{+}19]. supplements [SVA18]. supply [ASGE14, GLOC10, KFA12, KK18, LAPK14, TP14]. supply-demand [LAPK14]. support [Di 13b, GYWJ10, HZL^{+}11, Jam11, Jan15, JW18, KHX^{+}19, KSKK15, LYF^{+}15, MBE11, NHSX14, TDT16, XWD^{+}10, ZDG^{+}10, pZWZ^{+}16, ZZC^{+}17]. supports [BSB^{+}13]. suppress [AANF16, LSDI15, PPBT11]. suppressed [AR17]. suppression [CdMB10, CDK11, GC18, HHSA15, HA15b, HLH^{+}18, MYK17, SHF16, SSN^{+}14, YZY^{+}16]. suppressive [STI13, WR14, WML^{+}17]. suppressor [HTN14, SSM^{+}18]. suprachiasmatic [KTH16, KB10b, VH11a]. Surf [Luk14]. surface [AdSG14, AAJ15, CEKM^{+}19, HHMT19, HGB^{+}13, IMH15, IG14, SKGM19, TAI^{+}18, TSP15, WBB^{+}18, ZDF^{+}12]. surface-associated [ZDF^{+}12]. surface-dependent [TSP15]. surfaces [CKS15, MPY14, PVGA12, XLG^{+}15]. surfacing [MDB12b]. surgical [JMH^{+}19]. surname [MSK10]. Surrogate [SKK^{+}11]. surrounded [SA17]. surveillance [MOEdP16, MW13, PGF11, ZZL19]. survey [QZ14, WPH^{+}12]. surveys [TKB^{+}18]. Survival [GCM17b, AP13a, HHS^{+}10, KCD11a, KCD11b, LS10a, MSB16, NR14, TSL12, ZZC10]. surviving [ML09, ML12a]. susceptibility [HR14, HZL^{+}11, LLC15, SY11, SKS^{+}19]. susceptible [Gru18, ITR^{+}18, TDKJ15, Zuc14]. suspicious [PCN17]. sustain [MKJS13]. Sustainability [KG13, RDMP11, TLW18]. sustainable [BR14, GL12a, GK13, TM15b]. Sustained [BF13, NCLB16, SH16]. sustaining [MSN14, RA18]. sustains [PMSY17]. sutures [BFGAGA16, YKO^{+}16]. SVM [BMSEE14, HYA14, HZC^{+}10, JCG16, LT14, Mei14, ZZC^{+}17]. SVS [LY11]. swallowing [dLMD^{+}11]. swarm [FCS18, MMAS13, PP18, PP19, YYC19, ZZN^{+}19]. Swarming [LX10, RA17]. swarms [Gal10, LTSTD15, RLK10]. Sweden [FGH^{+}14]. swelling [LSP^{+}17, XLLF17]. Swift [MROS15]. swim [Gro17]. swimmer [HFT15]. swimmers [LCBO^{+}12, RL15]. Swimming [SKSO12, BPG^{+}18, CT16, HN10, IG14, Ish16, IGG^{+}18, Luk14, TTW13, WGH^{+}14]. Swinger [Sel16, Sel15b]. Switch [SSKK17, CSKZ19, DD13, GC16, HL14, Kar16, OM10, PRH^{+}11, Sos18, SBJ^{+}18, VDRL14, Voh17, WR14, WKS14, vLBJ^{+}13, vdSS12]. Switch-like [SSKK17]. Switches [CKF17, ANY14, Gol10b, SX12, UJLG14]. Switching [LKP^{+}12, KAN11, LW10a, LW10b, LZT12, LZT13, MHF^{+}13, NS18, Pat16, PLN14, PHPM18, RFS^{+}15, WFZW13, ZJY19]. syllable [GZT12]. sylvatic [KZ14, PKZ12]. sylvestre [RAA^{+}16]. Symbiodinium [CMGN17]. Symbiogenesis [AE17]. symbioses [CMGN17]. Symbiosis [LGEM17, APBS15, DMS^{+}16, PMK15]. Symbiotic [BLW11]. Symbol [Fer11]. Symbolic [PM17]. symmetric [DGJ14, HMJB15, Kae17, MMJB17, MG10a, Mig16, MS14d, RPD14, ST16, Wu14]. symmetries [TS17b]. Symmetry [RVP19, WN19, ABG^{+}13, Len14, LJW^{+}16, yTRSC13, She11b, SWSLMJ19, SG18, TM15a, Zha15]. symmetry-breaking [SWSLMJ19]. sympathetic [BPPC15]. sympatric [SA14]. symptoms [PZLF19, Sat18]. synapse [JSZ10, MYC12, TTG19]. synapses [HKS15, KLHS17]. Synaptic [ML11a, VGZS18, BB11b, FIS11, FIS16a, FR14, GL12b, HKS16, Hor11b, Rob11]. Synaptotagmins [GU13]. synchronisation [BTO15]. synchronization [ABJ12, BTR18, CTJ10, CFF12, DS10a, HTK^{+}18, SN12, ZLL^{+}12]. Synchronized [CH11, AANF16, OKVN18, PJ13, TIS10]. synchronizes [KDL16]. synchronizing [GH11]. Synchronous [XNW17, AM10]. synchrony [DBD12, JP13, PJ13]. syncytial [KKG^{+}14]. syndrome [AD15, EAN14, LCJ16, MSB16, Veg10, YXH^{+}14]. syndromes [JHE15a, JHE15b]. synergism [Wal16]. Synergistic [CS15b, YNY^{+}10, BZJP18, LW17, LW15, PBA12, Tay13, Tay16b]. synergists [Bon10]. Synergy [DL15a, MV18, Oht12, SOF16]. synonymous [LY14, MV10]. syntax [WSH^{+}10]. synthase [EJ16, KHH^{+}17, LCQ^{+}18, TWH^{+}13]. Synthesis [Sak10, BFJ^{+}18, EHBC10, HL11, HMTA17, JS17, JS18, NCS^{+}13, SC11a, SBB18]. synthetases [Di 17a, Rog19]. synthetic [ASRM15, ABIM10, BPP^{+}16, DGW^{+}18, FWR19, HMTA17, HAuR^{+}18, JZ17, MMC19a, PFJS15, PPF17]. syntrophic [SK16b]. System [GCB17, LY14, ASK17, ANMH11, ATI12, AA16, ARZ15, BCKS13, BBS18b, BAM^{+}11, CdMB10, CST19, CSBK15, CTW10, CKF17, CXT19, CKN^{+}12, DKP^{+}18, DFMR19, EBSW17, FRP14, FJR19, FMLM12, FK13, FIJJ11, GB13, GPR^{+}16, GBM18, GGG12, HJLNZ11, HGM15, HI19, HAB^{+}17, HWGT15b, Kär11, KNA^{+}18, LW17, LGR^{+}12, LYW11, LLL13, MI11a, OO18, OHWS18, PWG16, PR17, PHPM18, PZS^{+}10, PBEI12, RB13, SHG16, SLR17, Sat19, SHH15, SKPK17, SR12, SX12, SKK18, SY12, TWW19, TWS15, Thu15, TAORS10, VWF19, VAG16, VC10, Voh17, WD11, Wan19, YiTS19, ZCSRG12, ZZ14b, ZZN^{+}19, ZMW10, d’O12]. Systematic [GBRS19, HWD^{+}14, Oka12, Sel13, Sel15b, Sel15a, VM12, WPH^{+}12]. systematically [Sel14b, Sel16, Sel17]. systemic [CSDR15, DL15a, NAK^{+}11, SCA13, TvMG16]. Systems [CKS15, MTE15, MMLK11, Abb10, AF10, ASK17, AENK12, BWB11, BA19, Ben14, BSL^{+}17, BJ17, BRP^{+}18, BE14, CR14, CFZ14, CLK18, CGHF14, DS13, DGJ14, DG10a, DV12, EA17, ESG10, Ezo19, FLA^{+}19, FKV19, GZFX14, GVSLG16, GMM^{+}13, GP11b, GP12, GA15, GT15, GW19, HJWC11, HAK^{+}19, IMA14, JZL13, JCW13, KSP12, LBMG16, LCSH14, LS16a, LDH^{+}12, LL13a, LLZ13, LMdL11, MH12, Mit14, MH13, MAF^{+}19, NH12, NBC16a, NBC16b, OA15, ONO15, OUMA10, Pag19, PJL14, PSL^{+}10, RV16, RR16, RDMP11, RGBR17, RAD14, RA18, SYSM11, SBSE14, SVA18, Sue12, Sza15b, TLW18, VR17, Wal18, WZ18, WW19, WSM12, WFM^{+}13, WFC^{+}14, YYST13, dlEBRM15]. systole [ASC15].

T [DHK13, DL16, KL11, SI11a, SRAL12, AHAFA14, BCPL10, BYY^{+}11, CSR^{+}05, CTL^{+}15, DP13a, DKR16, DL12b, EH17, EME^{+}16, FMLM12, GML10, GBIR13, JDPK15, JB19a, Lab16, LAH^{+}16, LFB^{+}16, LG12, MEKK11, MYC12, OTGT10, RAF^{+}14, RW12, SI10a, SDD15, SDT17, STN19a, SLvdBMP10, SSRA16, TXWW12, XJ19]. T-Cell [STN19a, BYY^{+}11, DKR16, DL12b, GBIR13, JDPK15, SDT17, SSRA16, XJ19]. T. [Mac11]. T4 [ARZ15]. T4/Escherichia [ARZ15]. T7 [dRFI15]. table [Nah12, TLR14]. Tag [GvVT14, Lai11, MBBD13, MNSZ16, WFZW13]. tail [JZZK11, OYIM10]. tailed [LK19]. tails [MDM16, SSZR17]. take [MS14b]. taken [MPNP12]. takes [vVGA10]. Taking [MA10, FD18]. tale [VHS17]. talk [BJJR10, FZL18]. talking [CDM^{+}14, YST14]. tall [Gou15]. talus [PRSC11]. Tangled [CTA15]. Tank [HKH18]. Target [Fil10, BCPM^{+}16, BJ18, HII^{+}19, HG18a, HW11, HW13, IH17, ITR^{+}18, KS10, LN13, RAF^{+}19, TAR16, XJ19]. Targeted [SGW17, CFF11, DK13a, GGQ^{+}12, HS15, MKMG^{+}14, ZZG^{+}16]. targeting [MSB16, SSM^{+}18, SH15, WHH17]. targets [HRCA19, HZC^{+}10, MG14a, PYG^{+}19, RAA^{+}16, YWW^{+}14]. Task [GSF13, GXFF13, JAHKH12, KP12]. Task-partitioning [GSF13]. task-specific [JAHKH12]. tat [ARG19, BYJ17, YBC17]. taxa [SAG19]. taxon [SSS13, SS15d]. taxonomically [SKH17]. taxonomy [BBDB13]. Taylor [KHRS10, KMM16]. TB [GSV16, PDW11]. TCF [SWR11]. TCP [KBV11]. TCR [LCHMP16]. TD [SAKG13]. teacher [CB17a]. teaching [CT14a, CCNT19]. TEAD4 [NZZ19]. technique [BGF^{+}14, CL17b, JZ17, LLC15, SLDP13, ZZG^{+}16, ZNCM15]. techniques [AENK12, LBGW13, SSJK18, TTW13, WP17]. teeth [LBB^{+}13]. telencephalon [BF13]. teleological [GJ15a]. Teleonomy [TP14]. teleost [Yam16a]. tell [LWLM18]. telomerase [PVGV19]. telomere [EBX17, GSV11, HWPL12, OB10, RBKW19, SRH19]. temozolomide [HL18]. temperate [ECP^{+}16]. Temperature [CMM19, KH19, RL17, YI18a, YMZ18, APS^{+}13, AC12, CFL^{+}15, DLSD15, ECP^{+}16, GMZM15, NSO15, RBB^{+}12, VBV10, WTC16, WDL^{+}13]. Temperature-dependent [KH19, YI18a]. Temperature-driven [YMZ18, WDL^{+}13]. temperature-mediated [AC12]. temperatures [SSV^{+}15]. template [BRB18, HS12, IDM15]. template-based [HS12]. templating [Dun11]. Temporal [CzST^{+}18, ABB^{+}19, AGPK13, CNL14, CH16, KMM16, LJ15, LS11b, MTdS^{+}16, MS10a, OABB18, PM11a, PM10a, QJR^{+}16, STX^{+}11, TP10c, TBA14, TC11, ZVMB10]. temporally [AdVG15, Ros10b]. temporary [ESS19, WW12a]. tenase [KPS17]. tendency [FB12, NPB13]. tendon [FS11a, RS12, WYL^{+}18]. tendons [RS14a]. Tenofovir [RVZ14]. tensegrity [AA16]. tensile [WYL^{+}18]. tension [GGC14, MHMM11, TSB10, WHWZ18]. Tensor [PH13]. tensors [SC10c]. Teramoto [YLH12]. term [ABLH18, BB11b, FHM18, GA14, HKS15, HKS16, MPP^{+}16a, MBP16, MI11b, MSL^{+}16, SBB14, VZB^{+}15, WMK13, WMT15, WGS15]. term-frequency [WMK13]. terminal [BCS^{+}16, GSGV11, RSBB10, SSZR17]. terminals [GU13]. termination [TH17, dSKBS10]. termites [JL11]. terms [Hua12, ZT16]. ternary [TKK14]. terrain [ASL^{+}18]. terrestrial [LYK12, Pon12]. territorial [BR12b, GSRR17, MGS16]. territoriality [HK10, TIY^{+}11]. territories [GSRR17]. territory [JL11, MN14]. tertiary [Mas18, NHTS14]. test [All11, ARZ15, GZ19, HPB^{+}14, LSSG10, NBW10, NBW11, RRG^{+}12, RRC^{+}11, ZPdFJ19]. test-based [LSSG10]. testicular [SFMS16]. Testing [NBW10, NBW11, RRC^{+}11, RCD19, SAG19, CS14a, CS11b, JZZK11, NM11]. testosterone [GS17b, HS14a, Jam14, ZP11]. tests [GK16, SD16, TW13]. tethered [CKBCG12]. tetra [Sel15a]. tetra- [Sel15a]. tetrahedron [Den19]. Tetrahymena [ZGZ^{+}10]. Texas [BLS^{+}12]. texture [CLL18, KSM^{+}15, NSBL10]. textures [DDF^{+}14]. TGF [CWP^{+}18]. TGF- [CWP^{+}18]. Th1 [EBE10, GMB11]. Th1-Th2-Treg [GMB11]. TH17 [SGDL12]. Th2 [GMB11, EBE10]. thalamus [HZG^{+}17]. thaliana [BHH^{+}14, HNA15, MBK^{+}11, MBKB13, MLL^{+}16, OFT15, SCÁBM10, SCÁBM11]. their [AP17, AHJ18b, BFS18, CFF12, CT18b, DLYZ11, EME^{+}16, FD18, FS15b, GL14, GXF13, HPB^{+}14, HM15, HZL^{+}13, JLX^{+}15, JCG15, KCMF11, KAN11, Kur10, LLLV11, LL13a, LD11, LCBO^{+}12, MYK^{+}11, MAW12, MP12, Pai19, PDC^{+}17, PMCS16, QW11, RMSTG13, RT19, RBRB16, SMD^{+}16, SS17b, Wal18, WSRG18, WXC10, XM11, ZADB15]. them [DBBW09, DBBW11, Di 12]. Theor [AEM14, BPF15b, BBM^{+}13a, BBM^{+}13b, BPGS12a, BMGC11a, CLG^{+}11, CNG^{+}12, DBBW11, EPJ^{+}11, FIS16a, FS16a, GZ14b, GS14, dOGL13, Gün13, JN14, JRG14a, JRG14b, JCLS^{+}21, KDK14a, KMLT14, LBS^{+}11, LGK^{+}12, Lei10, Lio24, LZT13, LHPF18, LPvSP12, ML12a, Mei13a, Mei13b, MLBA13, Mul12, NO14a, PSS^{+}13, SCÁBM11, Sch21, Sch23, SZ16, SSKS20, SC13a, SRAL12, TSM13a, Van16b, WL12a, ZG10a, ZSCL16a, ZZR13]. Theorem [Gra18, LE19, WSM12, vV18]. Theoret [PL14a]. Theoretic [GDPPSS^{+}11, Bar18, BBR12, DH16, DMSW10, MWSM13, MGS16, TWB16, WHYMG17, ZMAM19]. Theoretic-experimental [GDPPSS^{+}11]. Theoretical [AK15, BB15, CMD^{+}10, DS19b, DHK13, DHT16, EJ19, FKMP10, Fer12, Fue16, GHS20, GDB^{+}19, JS18, LPB18, LBW^{+}13, MBF^{+}15, MdSPBL16, Mol18, PCT19, Pan19, PPC^{+}17, PZS^{+}10, PP19, QF10, RG10, RTEKG19b, RTEKG19a, RZIR16, RS14b, SK11a, SPA18, SYY17a, WLW20, XWW^{+}19, XLG^{+}15, Ano19-27, BD10a, BCKS13, BWM^{+}19, BBDB13, CFF12, CPS10, CW11c, Che18, Che19, Cro19, CG11, DOC17, GGL^{+}16, Gri15a, HDZ^{+}19, HHS^{+}10, IKHL16, KFG15, KCE^{+}11, LPvSP11, LPvSP12, MR10, MYK^{+}11, Mal10, MFMB12, Mor11, RDP16, SMM15, SMHB10, SMS17, Sat17, SYI17, Ser16, SPMGR10, SGGY10, SNCM12, Sza15b, Voe18, YSST13, YLH12, ZPdFJ19, GGG12]. theoretically [Rao12]. theories [Di 16, Di 17b, HL11, Mor16, RB15, RBF15, Sak10, TGL15, YYST13, YMW12]. theorists [Ran12]. Theory [MMB18, MBS19, OT13, PWL^{+}11, YYB^{+}19, AE17, ASH15, AAJ15, Arc11, AS12, ABLH18, BCBD19, BCPM^{+}16, BWP10, BORA10, BM16, BAM^{+}11, Che12, Che16, CDM^{+}14, Cle16, DMO^{+}17, Di 10, Di 16, Di 17a, Di 17b, FHSG15, Fra19a, FHR13, FR13b, FR14, GBC^{+}16, Gre15, HRDL14, HRD14, HKB10, HZWH10, IST11, KFG^{+}14, KDMK16, KP16, Kue16, LP17, MP14a, MvAKR17, Nah12, NAK^{+}11, NTFK11, RSH19, RC11, RH19, Rob11, Ros10a, RWH16, SA17, SGGM11, SSFG15, SC11a, SR10, SSD12, TR17, THBM10, Wal16, WRC^{+}19, XNJ^{+}13, XLLF17, vTMW15]. theory-the [Nah12]. therapeutic [CBGS18, DL12a, HAuR^{+}18, HB18, LN13, MB12a, SPSM15, SSM^{+}18, SH15, WB15, WFC^{+}14]. therapeutics [GGQ^{+}12, PYG^{+}19]. therapies [Ala15, DH18, HL19, PDF18, RRTR12, SOF16]. therapy [AGRRTR11, AGRR^{+}15, BAGG14, BGCB12, CFCM13, CLA^{+}16, DDTL19, FKDW15, FM10, FM14a, GBJ^{+}11, HBA10, HB18, HSLW16, KK18, LMM18, LNRK11, Mal10, MKMG^{+}14, MSL^{+}16, NH19, NSH^{+}10, PMCS16, RRRRSPTR16, RAR19, SGAM12, SSN^{+}14, TXWW12, TXCW15, VZB^{+}15, WR14, WML^{+}17, YCL^{+}17, ZCT18]. therapy-induced [KK18]. there [Di 18, Ros10a]. Thermal [MK18b, BUL12a, JMS12, MK17, Pav14, RTRRS^{+}17, RCL^{+}10, ZG19]. thermo [MDD13]. Thermodynamic [Cha17, ABGM11, FTPN10, Ghu18, GGGM16, LGPS17, LD18, LWvB^{+}19, Mar12, MH13, NN18, OABI12, PGT^{+}19, WdLS17]. thermodynamically [vLFM^{+}19]. thermodynamics [AH12a, FJ10, KK15, Tsu19]. thermogenic [RBB^{+}12]. Thermomyces [KGP^{+}15]. thermophilic [FLW16, LHH12]. thermoregulation [STKE12]. Thermostable [KGP^{+}15]. theta [FR14, KA16]. thick [Don13]. thickening [Fok12]. thickness [BMM^{+}14, IST11]. Thickveins [Che19]. thing [OAP14]. thinning [Lar10]. Third [JR17b, KDST15, KB10a]. third-order [KDST15]. Third-party [JR17b]. Thomas [BA19, TK19]. Thomas-Rössler [BA19]. thoughts [PBRW11, SB15]. threading [Bry13]. threaten [RDMP11]. threats [Dra19]. Three [ANY14, FJC^{+}10, Kur19, XM11, BTH^{+}13, BMN16, BCF16b, CRLH^{+}19, DKF17, GDJC11, HVM13, HY13, KK12b, Kro10, MPY14, Mal15, MNSZ16, NO14a, NO14b, NTED^{+}19, NOS17, SNY^{+}13, SMA11, SMZ^{+}17, VHS17, VBVD17, WPPD16, WCC13, ZR16]. Three-dimensional [FJC^{+}10, BTH^{+}13, MPY14, SMZ^{+}17, VBVD17, ZR16]. three-generation [CRLH^{+}19]. three-person [NOS17]. Three-player [Kur19]. three-species [GDJC11]. three-stage [NTED^{+}19, SMA11]. three-tiered [WPPD16]. three-trophic [DKF17]. threonine [HKS15]. Threshold [GP12, BV12, BGL10, BN15, GDS11, HBSF11, HW13, HK16, KEKB18, LCG^{+}15, MAS15a, PR13, SE10, SHF16, SS11, TL15, TMM10, ZTT18]. Thresholds [RW10, WB19, BN12, GSCS11, GGL^{+}16, GRBL13, HBT13, MCRG18, MGC13, OF10, RMF^{+}18, SSZB15, Str15, TH17, ZXH19]. thrombin [SPA18]. thrombocytopenia [ZML19]. thrombopoiesis [SGL10, WVO^{+}15]. Thrombus [MP19, MTdS^{+}16, TAB^{+}13]. thumb [Voe18]. thuringiensis [RBSD10]. thwart [Cro19]. thy [WMD^{+}11]. thylakoid [LL17b]. thymocyte [MZMM^{+}14]. thymus [MZMM^{+}14]. Thyroid [MM15a, For14]. tibia [Cle18a]. Tibor [HS16, CB15, Sza15b]. TIBS [YXH^{+}14]. tick [DB11, DDSDW13, WDL^{+}13]. tick-borne [DDSDW13]. tidal [SM16]. tier [SKK18]. tiered [WPPD16, XDC^{+}11]. tight [KRR14]. Time [BJOS13, NKB16b, PW19, Yam16a, AIY16, BGM19, BP15, BSKV18, BUC14, BH11, BCK19, BD12, Cam15, CK15, CV19, Cor16, CMJD11, DLSD15, DD10, DHP11, EAN14, FGGT15, FS16c, GLR^{+}17, GCM17a, GFM^{+}19, Gol16, Gre15, GBM18, Hal12, HY10, Höh15, JSK16, Kar16, KHRS10, KAKK19, KII19, LJM15, LS16a, LSM17, LBZ18, LP12b, LBS^{+}14, MDS16, MTNM12, MMH^{+}12, MYN^{+}13, MYN^{+}15, MS14c, MBP16, MSIR10, NIiT19, NS11, NMI18, NNvdB17, Opr10, PAK11, PDSP13, PTC13, PBR17b, PR13, PKH11, Rev15, STH18, Sim14, SVSP15, SS15d, Sta10, SMC^{+}13, SBM^{+}12, Voe18, WYKT17, Wax11a, Wei12, Wil11, WGC13, XSKA17, XKCG15, vDRT14, vE11]. Time-delayed [NKB16b]. time-dependent [Höh15, Wax11a]. time-linked [BUC14]. time-reversible [CK15]. time-scale [FGGT15, PAK11, WGC13]. time-series [NMI18]. time-space [PBR17b]. time-varying [DHP11, NIiT19, NNvdB17, SMC^{+}13]. timed [ANY14]. times [CW15c, CK19, DBBW09, DBBW11, FS12, KC17, Lab16, LCC15, LPKH15b, Par15, Sch18, Sch21, SM11a]. Timescale [RWW^{+}15, PP14]. timescales [ESS19, WD12]. timid [LG19a]. timidity [GG14]. Timing [OS13, BZA16, BI12, BHLH12, HJWC11, OABB18, TTC10, WSRG18]. TIMP [SMS17]. Tinker [Ou19]. TiO [ALM^{+}19]. tip [EdKM11, LLK14]. tip-growth [EdKM11]. Tipping [CNDK15, KMN16, YI17]. tire [PVCEC18]. Tissue [GSSBF18, GL18b, RW14a, BSP18, BJO^{+}16, BJWL15, CF11, CLCH10, CGRRGM^{+}16, DBBW09, DBBW11, HPP10, Hir16, LMC^{+}13, LW18, MYK^{+}11, MBBV14, MA11a, MAMEA15, MH13, MHMM11, NHE^{+}16, NTOI16, POP12, PDL^{+}17, RTRRS^{+}17, Rem15, RJSC18, RC13, SBSR13, SGS15, VH11b, WMPF^{+}15, Wod18, Yat14, vDVFH^{+}18]. tissue-diet [Rem15]. tissues [BTF19, DS16, EHBC10, FAFA17, FM14b, Kro10, KBF18, LLB^{+}18, MYK^{+}11, RHS14b, RHS14a, Wal16, ZGS^{+}10]. tit [BYJ17, YBC17]. tit-for-tat [BYJ17, YBC17]. titration [Mar17b]. Tiwanaku [FBU11]. TK [SAKG13]. TK/TD [SAKG13]. Tkv [Che19]. TLR4 [Gün13, Gün12]. TNF [MMFK10, RAR19, WPH^{+}12]. TNF- [RAR19]. toe [SHW16]. toes [LCBO^{+}12]. together [GN14, Mor10b, ON14, SSP15, TTN13]. tolerance [LPB17, Moo14, VB19b, LPB18]. tolerated [BH13c]. Toll [AES^{+}13]. Toll-like [AES^{+}13]. tomato [AKdV^{+}14]. Tomlinson [BR14]. tomography [CZPC^{+}18, LDB^{+}14]. tongue [TF16, WHWZ18]. too [OAP14, VWF19]. tool [AEZR^{+}16, BVJE17, HYSG19, MdFDM10, PCL^{+}15, SVA18, dEG11]. Tooth [Fre10, TIMI^{+}15, AKR^{+}18, CGRRGM^{+}16, IvLP12]. Top [ZYD15, MPS^{+}11, SC10b]. top-down [MPS^{+}11, SC10b]. topographic [SCT12]. Topological [HNO18, JC16, MEJ18, ACdlRMR^{+}11, FBAPMD13, HZL^{+}11, cLCJ^{+}10, MBE10, YWW^{+}14, YWZ^{+}16]. topologies [ADR11a, BKL14, ZSO16]. topology [BBS18b, CL13a, CK15, GJS^{+}10, IMA14, LS14b, LCL14, MvdDW13, MD16, PJL14]. toppling [Lar10]. torque [GY18]. torsional [APW10]. Tortuosity [LXJ15, PA15]. tortuous [Che14]. torus [SGKCP19]. Total [Gar14, Gra14, JRG14a, JRG14b, JRG14c, RTR^{+}13, RLW^{+}14, RHT18, TKB^{+}18, TW12]. touch [ARB13]. touch-down [ARB13]. tourists [ZZL19]. toxicant [HAK^{+}19, LLL13, PPW16]. toxicants [NNK10]. toxicities [QLWC19]. toxicity [FSA15, RWW^{+}15, RGL^{+}17, TTB^{+}18]. toxicodynamic [SAKG13]. toxicokinetic [SAKG13]. toxicokinetic-toxicodynamic [SAKG13]. Toxin [DHV19, BSL^{+}17, HSR12, YWW^{+}14]. Toxin-mediated [DHV19]. toxins [HWL15, PWZ^{+}19]. Toxoplasma [JSSZ12]. trabecular [ASH15, SAB10]. tracer [BGF^{+}14, BHSB11, BvLH14]. tracer-transport [BHSB11]. tracers [BSMK11]. Tracing [CW13, BGW15, CNL14]. tractable [SBJ13]. Trade [MLT10, PBB10, Chu18, FtL18, HMH14, JB18b, KZ14, MLMG^{+}15, PVCEC18, PBD11, WK15]. Trade-off [PBB10, HMH14, JB18b, KZ14, WK15]. Trade-offs [MLT10, Chu18, FtL18, PBD11]. tradeoff [Lug15, PHPM18]. Tradeoffs [KL13, DBB14a, Fra19a, LSM17]. Trading [AS13, DMS^{+}16, ESE15]. traditionally [AM14]. traditions [KLM14]. traffic [MSA^{+}16]. trafficking [CN10]. tragedy [MB18c, VSP11]. trail [Amo15, RRBM12]. trail-laying [RRBM12]. trained [LT14]. training [CXC18, SG15b]. trait [BCF^{+}16a, BBS18b, CVO^{+}15, Cre10, GCZ^{+}12, Gri15b, HD11, HT13, Jab10, MH18, Mog15, RK18, RBB16b, SL10, SLL18, SA13, SYY17a, SYY17b, TGLK19, Van16a]. trait-based [Jab10, SL10]. trait-dependent [TGLK19]. trait-mediated [MH18, RK18]. traitorousness [Lai11]. traits [BBS18b, BGLR^{+}19, CRLH^{+}19, CL18, DXW^{+}16, JPDP10, LIPD12, POW18, RCD19, SB12, SMG18, SQZZ17, ZMW10, ZMT11]. trajectories [CHL^{+}19, LWLM18, LAJC19, PRSC11]. tramadol [ACR^{+}17]. transactions [MLMG^{+}15]. transboundary [NBA^{+}18]. Transcellular [Sil16]. transcranial [FR13b]. transcript [SC11a]. Transcription [HG18a, KEHK17, MSND12, Bar19, Das18, FLM18, GY18, KHS13, LLY15, MMH^{+}12, Mic11, MHD18, NZZ19, PHÁBAI^{+}16, Sel15b, Sel16, Sel17, YST14]. Transcription-associated [KEHK17]. Transcriptional [Pol12, ARG17, BPP^{+}16, DM19, Hal17, HCHI18, NZZ19, Sel15a, TK10b]. transcriptome [SLW^{+}18a]. transcriptomic [CTSL11]. transduction [ANK10, ANY14, DFM^{+}19, JC10, RW10, RHJ11, SAA10, Tsu19, VLPH17]. transfer [Di 12, FXML18, LWW14, LWL^{+}11, MTE15, Mau15, Mei12a, Mei12b, Mei13a, Mei13b, Mei14, NNK10, PMSY17, RS19, SSS13, SLHS13, ZDF^{+}12]. transferrin [GHBI16, YCL^{+}17]. transferrin-CRM107 [YCL^{+}17]. transfers [BCBD19, bCRS17, CL13b, DLM^{+}19, GSF13]. Transform [YCY14, ATRR10, AAJGCD15, CL17b, GT14, YZW10, YY15, ZLDZ13]. transformation [CMHM19, FH14, Gla13]. transformations [Bal13, FGGS15, MH13, RKMG12]. transformed [KNA^{+}18, SCsMzX18]. Transforming [HMW16, MHMM11]. transforms [JLX^{+}15, SVCS10]. transfusion [HSLW16]. transgenic [BAR14, LBGW13]. Transient [BB11a, KK18, BF16, FGH^{+}14, Hor11b, LLTP19, Lüt16, PW19, RW10]. transients [MBP16, SHF16]. transit [KS15, SSZB15, Wei12]. transition [BBR10, CCdS15, DB10, KL11, Kin18, MS14d, OTTF11, TWB16, TK10c, ZLW16]. transitional [HSiI^{+}19]. Transitions [MVGGB18, WD11, DS10a, GJ15a, MB17, RA17, SCA13, TBM^{+}13, WK18a, XNW17]. transitive [DN18]. translation [BBRT12, DCL18, GYWJ10, GRRG16, KSvdH18, SGD^{+}16, Van17b, Xie13b, ZTWL12, dSKBS10]. Translational [Tak17a, BPP^{+}16, EBSW17, KSvdH18, XG12]. translocation [ACH15]. transmembrane [CZT^{+}16, EG12, IHNS16]. transmigration [Kun16]. transmissibilities [WN19]. Transmission [EN15, FRP14, KR15, MCCC^{+}10, XSTW14, ABB^{+}19, BBR16, BDR10, BHO^{+}18, CDO17, CT18a, CLK18, CBC^{+}18, CG14, DFM^{+}14, DB11, Déb17, DBBM11, ERT13, GSYS10, GWCA14, GC18, GL12b, GAY^{+}11, HAP^{+}16, HHE19, HBL10, HLTW14, JSSZ12, KGK10, KLJ17, Kle14, KZ14, LLN^{+}19, Lan16, LKK13, LCC15, LJ15, LKO^{+}17, LP19, LN19, LMCW18, MHMM18, MHX^{+}14, MCC^{+}17, MG10a, MSB16, yTRSC13, MGC13, MMRCCC10, NKL10, NT14, NDZMA14, NTED^{+}19, NYSM12, OVKL14, OL19, PSJ15, PNVN18, PaCZ10, PT10, PKZ12, PSL^{+}10, PEZ^{+}13, RC16, RDSD^{+}12, Ros10b, RCL16, SW16, ST17a, SN11, SSST19, SY11, TC12, TDM14, TCB13, VGZS18, WW12a, WZJ^{+}13, WL15, WCJ11, YTK10, ZZL19, ZN18, ZXH19, ZZR10, ZZR13, ZRZ15]. transmissions [KK12a]. transmissivity [Cam11]. transmitted [BPFR16, BCF19, CCW16, HE16, HHA^{+}13, MPM10, MBC^{+}12a, TB18, YZMY18]. transmural [DLMK12]. transplant [WLL^{+}14]. transplantation [EUM^{+}16, MI11b]. transplanting [MI11c]. transplants [JEA18]. Transport [ISM^{+}11, SF15, AMM16, BD10a, BHSB11, BvLH14, CN10, DLEMP19, DEK15, DS16, DFG^{+}18b, FZ14, FM13, FM14b, Fue14, Fue15, GDF18, GSSBF18, GZFX14, GSGV11, HPP10, JSK16, JG14, KBS^{+}19, KBS^{+}20, KK17b, KH19, LM15, LS15b, LCT11, LFD11, LBZ18, MAFK12, MN12, MSC10, MK17, ML12b, Mor10a, MLL^{+}16, OFB10, ØW16, PCT11, PCT19, PWC^{+}15, Pan19, Pel18, PA15, QA15, RHS14a, SCH^{+}19a, SSFG15, SL19, SB15, SPL14, TLL^{+}12, TC11, WFC^{+}14, ZGS^{+}10]. transport-related [LCT11, LBZ18]. transportation [GSN^{+}11, KHP^{+}12, ZG10a, ZG10b]. transporter [Dim17]. transporters [PGT^{+}19]. transsulfuration [VHM^{+}17]. transversally [SGB12]. traps [CGD13, LIPD12]. travel [DB18, MSBB13]. Traveler [MRPN12]. Traveling [LDHD14, KO18, LW18]. treat [PPC^{+}17]. treated [DDTL19, HS14b, NMAZP16, OOY16]. treating [BFJ^{+}18]. Treatment [HSLW16, AKS^{+}19, BS15a, BCF19, BH13c, BBI14, BCPG18, CWM11, CB13a, DD18, DGD^{+}11, GSSBF18, GPD18, GBRS19, HLTR19, HGM^{+}16, IBB^{+}15, KBV11, KCSB14, KST17, LCTG15, LF19, LCCC10, LGMV14, LZL^{+}19, MS13, MVH11, NTED^{+}19, PCL^{+}15, RCH14, SOF16, TDKJ15, TGZF11, VZB^{+}15, WB15, WMN18, WCJ11, WCM15, ZZL^{+}11]. Treatment-donation-stockpile [HSLW16]. treatments [LXLF14, MBPS17, TDHC^{+}18, TALC16, TWB16]. Tree [ADV^{+}10, DMSW10, ADR11a, AdlPLMZ13, CTA15, bCR19, CK15, CS15a, CMJD11, DGL12, DAVS11, Di 19b, Di 19c, FM18, GK16, GJ15b, GRCdL16, Hay16, HDNB10, HMW16, LWH^{+}11, LMJ^{+}16, Mig16, QQL10, RM17c, SM17a, SMD^{+}16, SLHS13, TGLK19, VW10, vE11]. tree-based [Hay16]. tree-child [FM18]. Tree-grass [ADV^{+}10]. tree-grass-soil [DAVS11]. trees [AM10, AHR10, BNG16, DDS13, GD16, Gou15, GRR^{+}14, HS11, Kär11, Lar10, LG18, MS14d, Mul11, Mul12, MC15b, NP15, Oka15, SW16, SSS13, SWG^{+}15, SMC17, Sta10, SSKS18, SSKS20, Ste13, TI13, YLH12, vIK11]. Treg [GMB11]. TREX1 [GSS19]. tri [KPG19, Kři14, LTL^{+}15]. tri-gram [LTL^{+}15]. tri-trophic [KPG19, Kři14]. trials [DH18]. triangulation [QLWC19]. Triatoma [MMG^{+}16]. trichome [YLY15]. trifallax [BKC^{+}16]. trigger [ZLY^{+}13]. triggered [HL14, WRC^{+}19]. triggerfish [YSST13]. triggering [RRG^{+}10]. triggers [HBSF11]. trilobite [SKSO12]. trimorphic [Kar12]. trimorphisms [RQB17]. Trinary [TSM13a, TSM13b]. trinucleotide [FS15b, MP13b, Mic15b, MPP16b, ZLDZ13]. trinucleotides [RGP13]. trios [MHKA16]. Tripathi [Pan19]. triple [NJP18]. triple-negative [NJP18]. triplet [DMCP14, FW15, HM11b, dSMP^{+}11]. triplet-structure [FW15]. triplet-wise [dSMP^{+}11]. trisporic [EBSW17]. trisporoid [ESW13]. Trivers [BD10b, Jam13]. trivial [HA15b]. tRNA [BB13, Di 12, Di 17a, Di 19a, Rog19, Sel14a, VHS17, Xie13b, YLC16]. tRNAfeature [YLC16]. tRNAs [Di 13b, Sel11, Sel12b, Sel14a, dF13]. trophic [All11, DKF17, KPG19, Kři14, PMSY17, PML18, RM13, SL10, SRS18]. trophically [BPFR16]. trophy [HF17]. trophy-hunted [HF17]. tropical [HDNB10, SS15a]. tropism [OAP14]. troponin [RG18, VM16]. trout [KSKS13, MNH^{+}12]. trp [HVSZ10, ZCSRG12]. TRPV1 [MK18b]. truncated [Den19, JZZK11]. Truncation [MB16, MB19, VDTF15]. trustful [PCN17]. Trypanosoma [KZ14, PKZ12]. tsutsugamushi [KGK10]. Tuberculosis [MEKK11, BCPG18, CJL15, Ger16, GBJ^{+}11, GDB^{+}19, GAY^{+}11, GAY^{+}15, HYN19, HN19, KdlRJ18, LBMG16, LGMV14, LKAJ18, MM12, MMFK10, MG14a, OO11, PDRC16, PL14a, PL14b, PCL^{+}15, RK19, TDM14, WDL^{+}17, WCJ11, ZLZ15a, KS10, PCC15, SCC11]. tuberculous [EKvdKvFK13]. tubes [CKD11]. tubule [Hir16]. tubules [She11a]. tubulin [BJ18, DHHP14, KGM15b, SSZR17]. tubulin-driven [DHHP14]. tubulogenesis [HHA17]. tufted [VB19a]. Tug [SP14, Non10, SR10]. Tug-of-war [SP14, Non10, SR10]. tumble [AA12]. Tumor [CL14, CL19, GCSP17, KBV12, AGRR^{+}15, ABV19, BPC15, BSV^{+}10, CFCM13, CL18, CMHM19, CCAdS13, DFG^{+}18a, DS16, EBE10, EKvdKvFK13, FJC^{+}10, GLR^{+}18, GL18a, HM10a, KHX^{+}19, Kar16, KRDJ15, KRR14, KL18, KC11a, KK18, LL14, LBF13, LS16c, MOEdP16, MRF19, MKKS14, MFL^{+}18, MGM13a, NF17, NJP18, OAP14, PA15, RTEKG12, RTEKG15, RTEKG19b, RTEKG19a, RRTR12, RRRRSPTR16, SGB19, SFMS16, SSM^{+}18, SZ15b, SKD^{+}10, TTB^{+}18, VDD^{+}17, Wod18, WFM^{+}13, WFC^{+}14, YZY^{+}16, YBT^{+}17, ZZL^{+}11, d’O12]. tumor-immune [GLR^{+}18, MOEdP16, MRF19, MGM13a, RTEKG12, RTEKG19a]. tumor-induced [Kar16, SSM^{+}18]. tumor-tropism [OAP14]. tumoral [LDF^{+}11a, PDF18]. tumorigenesis [DM11]. tumorous [ZZS19]. tumors [AGRRTR11, AK16, BAM^{+}11, LS15c, PMK15, TW12, XLLF17, YLLL12, ZSSM14]. Tumour [CdMB10, MGGM10b, SGGM11, AKS^{+}19, AHW13, BBJ^{+}10, BFGS10, BJO^{+}16, BB19, CXWL11, DBBW09, DBBW11, DTGC14, HNP18, HB13b, IBB^{+}15, JYY^{+}18, JMC^{+}10, KHP^{+}12, PGHC12, PM13, SLD^{+}17, TDSM12]. tumour-associated [BB19]. tumour-induced [SLD^{+}17]. Tumour-stromal [MGGM10b]. tumours [BSB^{+}13, Car17, JKF19, MCL19, MGGM10a, SHvHB16, dG10b]. tuna [RDD14, TTW13]. tunability [MK18a]. tunable [BN12]. tune [SSZR17]. Tuning [JRMS12, LXS15, OAJK10]. tuple [LFW^{+}15]. turbid [MMB18]. Turbulent [MN12, WST15]. Turgidity [GRCdL16]. Turgidity-dependent [GRCdL16]. Turing [FK13, Kon17, KGM^{+}15a, SGKCP19, ZSO16]. Turk [Ran12]. turn [SHLL11, Toz15]. turnover [DP13a, Gin10, LMC^{+}13, PMK15, SHB^{+}17]. turtle [NSO15]. twice [Gar14]. TWIG [CMJD11]. Twist [Zol14]. Two [BC15, BKL14, Coh15b, DGJ14, GW14, HZW^{+}14, KC17, MRPLAS15, Mit14, Mog15, O’F11, Pan11, RKJ^{+}11, ATB14, AC15, AHJ18b, BLV18, BPFR16, BBS18a, Boc12, BP16, CS14a, Chu10, Cre10, DEG^{+}14, DGJ15, Di 12, Di 19a, FZ14, FK13, GMZM15, GI10, GGM12, GPR^{+}16, GT15, HYA14, Han12, Has14, HTK^{+}18, HB13b, JV14, JC10, JSWY19, JCG15, KA11, Kar12, KB10a, KMCJ17, KCSB14, KN11, Lai11, LCC15, LBZ18, Mal15, MMFK10, MB10, MBLC17, MSS10, MPLK14, MHMM11, NBS^{+}13, New11, NTOI16, OUMA10, PBB10, PBvdG10, PC10a, Pła10, RMM^{+}16a, Rei12, RBB16b, SC10a, SL10, SvK12, SBR10, She11b, TWW19, TBB14, TT10a, VCF^{+}19, WL12a, WL12b, WMK16, WcW14, Wan19, WBMM18, WF17, XDC^{+}11, XGZ17, ZLL17, ZN18, ZXH19, ZK10, vIK11]. two-allele [WL12a, WL12b]. two-compartment [MHMM11, PBvdG10]. two-compartmental [MMFK10]. two-component [OUMA10]. two-delay [KCSB14]. two-dimensional [ATB14, FZ14]. two-fold [KB10a, She11b]. Two-intermediate [RKJ^{+}11]. two-layer [AC15, FK13, WMK16]. two-level [AHJ18b, JCG15]. Two-part [Mog15]. Two-patch [DGJ14, BLV18, KA11, LCC15, LBZ18, TWW19, Wan19]. two-person [Pła10]. two-phase [WBMM18]. two-phenotype [BBS18a]. Two-player [KC17]. Two-scale [BC15]. two-sensor [GT15]. two-sex [Chu10, GGM12, MB10]. two-species [Cre10, Han12, PC10a]. two-stage [HYA14, JSWY19, ZK10]. two-states [XGZ17]. two-strain [ZN18]. two-tag [Lai11]. two-thresholds [ZXH19]. two-tiered [XDC^{+}11]. two-trophic-level [SL10]. Type [JZL13, AF10, AD15, BSR14, BPF15a, BPF15b, CNM^{+}13, DS13, EMM10, GKW16, HHJR11, HA15b, JDPK15, JDSPK15, KSU^{+}18, KB15, KDL16, KF10, LJ18, Mor10b, MMB14, NS11, NKOS11, PSD^{+}18, PRN10, RFdL15, RAD14, SDKF12, SAD17, SVA18, Wei17, WHS^{+}13, Zha10a, ZXC^{+}10, DFMR19]. type-2 [AF10]. type-dependent [PRN10]. types [AH15, AHJ18b, BK19, CL13a, CCY^{+}19, Di 19c, HYA14, HK11a, HK12, JD16a, JD16b, KDG13, Kar12, KHKI17, LZ18, LD11, MJ11, NO14a, NO14b, SM17a, SM18, SN16, SBR10, SVB^{+}10, SHLL11, TP17, WMK16, WXmH10, WK18a, ZADB15, ZZG^{+}16]. Typhimurium [SPSM15]. typical [TTP17]. Tyr422Ala [NZZ19]. tyrosine [CFR^{+}14, CMR^{+}18].

UAR [Sel12b]. Ubiquitin [PRH^{+}11]. ubiquitous [MIH16, YY18]. Uhlenbeck [BGKL17]. ulcers [FKDW15]. Ultimate [BC15]. Ultimatum [IRS11, NOS17]. ultra [FS11a]. ultra-structural [FS11a]. ultradian [HBA17, PRK13]. ultradiscrete [GI15]. ultrafiltration [dlRVFK^{+}16]. ultrametric [GD16]. ultrasensitive [MBM11]. ultrasensitivity [JRG14a, JRG14b, JRG14c, Str15, XG12]. umbrella [MSA^{+}16]. un-biased [KHKI17]. unassisted [Mar12]. Unb [KHKI17]. Unb-DPC [KHKI17]. unbalanced [MR15]. unbiased [NPH^{+}12]. uncertain [Ala15, KM18, KDS18, MZAI19]. uncertainties [KLHS17]. Uncertainty [YD15, CT14b, Che14, MWSM13, SJSK18]. uncompetitive [VACGF17]. uncoupling [Mar12]. Uncovering [CML10, CzST^{+}18]. Undaria [MJV16]. underdominance [ATRR10, EA17, EA19, KEKB18]. underdominant [LR16]. undergoing [BBR10, CKBCG12, DB10, RM10]. underlying [AP13a, HK14, KAN11, MvAKR17, RDD14, SN12]. undermine [ZCW13]. underreporting [MCCC^{+}10]. undersampling [JZ17]. undersampling-synthetic [JZ17]. understand [ÁTC^{+}14, HKS16, Mal18, MLBA12, MLBA13, RDP16, SCC11, YG18]. Understanding [AIB^{+}19, BN15, CS14b, DG12, DBG18, GN10, JDPK15, Leh18, NS11, PDRC16, Rao12, SV14, SBJ^{+}18, ZXC^{+}10, vLSBBD11, ANK10, CR14, CB10b, Di 19a, FM15b, HGM15, LCCB11, MRPLAS15, MMLK11, NT14, dLMD^{+}11]. undetected [Hou15]. unexpected [Eft13]. Unfolding [DG16, Ste12]. uni [WD11]. uni-directional [WD11]. unidirectional [HHJR11, Hor11b]. Unified [MvAKR17, CL17a, KFG^{+}14, LN13, NHM10, RTR^{+}13, SR10, TT10a, Xie15]. Uniform [DLMK12, CR19, KBKK14]. uniformities [Giv10]. uniformity [Yin17]. Unifying [FB18a, Luo14, PBB^{+}15a, Ros10a]. unihemispheric [KAPR12]. unimodal [FH13b]. uninfected [HII^{+}19]. Unique [MI11a, CRH^{+}15]. unit [BCK19, CS11a, WBM^{+}15, WK12]. unitary [XY14]. units [MS10a]. Universal [JSC^{+}16, MSK10, Bar19, CCdS15, Di 19c, Hay16, JBSFB12, OM10, Rou14, WTQL10]. universalities [BCGD10]. universe [YDC^{+}13]. unknown [BJOS13]. unloading [MYK^{+}11]. unmineralized [KF10]. unordered [BNG16, Cut15]. unpredictability [IS10]. unpredictable [Bon12]. Unraveling [DKS15, JDSPK15, NBBCC18]. Unravelling [MGO^{+}15]. unreliable [GL12b]. Unresolved [RBF15, Coh15b]. unrooted [CK15]. unspecific [HG18a]. unstable [SSN^{+}14]. unsteady [GI10, Whi11]. unstirred [DK13b]. unstressed [BHBR12]. unstructured [MKKS14]. unsupervised [NNK^{+}15]. unsynchronized [MA10]. untie [NvHM13]. until [Gre15]. Unusual [Mou12a, CEdLSG^{+}16]. unveils [DS15b]. unwinding [Xie15, Xie18]. update [CV15, Jam14, WMT10]. updated [Kon17]. updating [QMJW15]. upon [LYM^{+}19, PMP13, RAD14]. upper [CLP11, RRB10]. upper-body [CLP11]. upregulate [MHD18]. ups [vdBD10]. Upstream [IM10, PPBT11]. uptake [Ber14, Che16, CFR^{+}14, CMR^{+}18, FM13, FM14b, JSWY19, KF12, LLB^{+}18, PWL^{+}11, SSB^{+}15, TW12, WST15]. urchin [SCS14, vHHKB14]. ureteral [VF12]. ureteric [LF16]. uroguanylin [MPP^{+}16a]. ursinus [BCF19]. usage [EBS11, LY14, SS19]. Use [CLL18, WHL13, ZLT^{+}19, ADB^{+}13, BW13, Eft13, Gau11, GWE11, GRCdL16, HSR12, HMH14, JMH^{+}19, KGJ12, LWLM18, MPT19, PH13, PWH^{+}13, SKK^{+}12, SS15c, SD16, TT10a, TDE12, WB15, Wil11, ZKMB19]. used [GvVT14, SOIO10]. useful [Gla13]. users [DK13a, MVH11, RDSD^{+}12]. uses [EHSR17]. Using [ATRR10, BGKL17, BTF19, CZPC^{+}18, DBBW09, DBBW11, ESG^{+}16, EMM10, FKB^{+}12, GMZM15, JSF^{+}11, KK15, Kri18, LlLYB13, LXC15, MFG^{+}13, NSH^{+}10, XNJ^{+}13, ZZG^{+}16, AFS19, AGB^{+}14, AH12b, AEZR^{+}16, AJC12, Ala15, AH15, AH16, AHMA^{+}19, ABIM10, AAJ15, BS15a, BMI19, BJJR10, BMN16, BGF^{+}14, BSKV18, BF16, BJO^{+}16, BCFR10, BR12b, BMB^{+}18b, BPGS12a, BPGS12b, BBP13, CMPS17, CZW^{+}11, Che12, CWP^{+}18, CLK18, CJL15, DGM15, DKP^{+}18, DLL^{+}17, DROC11, DBG18, DZJR10, EAN14, FHG15, FLCS^{+}15, FKDW15, FWLW11, GYWJ10, GXFF13, GKNT10, GMM^{+}13, lGLLL14, HBK12, HMMSRSD15, HAB^{+}17, HTK14, HM11b, HYZ^{+}15, HZL^{+}13, HY16, IMW15, JD16a, JD16b, JW18, KKD18, KHX^{+}19, KCZ^{+}19, KSP^{+}14, KKGN16, KMA10, KHK15, KB19, KJ19, KCSB14, Kit10, KFW12, KDS18, Kon12, KAKK19, LCMC14, LWH^{+}11, LCJ18, LL12, LL13a, LL14, LYF^{+}15]. using [LTP19, LT14, LJ10, LLB^{+}18, wLW10, LBS^{+}14, LPD^{+}16, MS14a, MPJH13, MSND12, MFZ18, MALAN17, MIH16, MBE11, MAR^{+}17c, Mou12b, NH19, NBS^{+}13, NSBL10, NNK^{+}15, NK18, NMZ19, NI11, NHSX14, OSN18, PAK11, PDM17, PS13, PBEI12, PPF17, PKH11, PP18, PP19, PSAA13, PPT^{+}16, QW12, RSS^{+}18, RSD^{+}16, RMM^{+}16a, RA10, RNVP10, RP18, RRTR12, RBF15, RM17c, SIK^{+}18, SS17a, SRD^{+}15, SM14, ST17a, SM17a, SMG18, SGW17, SLDP13, SBT^{+}18, SSSD17, SMS19, SHLL11, SCF^{+}12, SSS15, SGDL12, SLT^{+}18, SWPC^{+}16, SKK18, SCLC13, SSRA16, SVCS10, TBQG14, TAI^{+}18, Tay16a, TLT^{+}15, TBMM19, TF18, TWC^{+}19, THBM10, TGZF11, Tun13, USF^{+}18, UPWK15, VF12, VZB^{+}15, Wal12, WMK14, WH14, WDN^{+}19, WHS^{+}13, WGH^{+}14, WZ15, XWD^{+}10, XXD^{+}17, YMW12, Yin17, YLH12, ZDG^{+}10, ZLY^{+}13, ZSZM14, ZKHL16, ZRGW19, ZNCM15, ZZN^{+}19, vdSS12, vTMW15]. usually [KB10a]. utility [Bry13, RWH16]. Utilization [KK17b, MHC16]. utilizing [QJR^{+}16, WYX^{+}17]. UTRs [SGD^{+}16]. UV [BKPV15, DM15, PHTP^{+}12]. UV-induced [BKPV15, PHTP^{+}12]. Uygur [WZJ^{+}13].

V [BTF19, BZJP18]. VABP [KKS^{+}14]. vacant [GSN^{+}11]. vacant-particle [GSN^{+}11]. vaccinate [KTJ19]. vaccinating [BB10]. Vaccination [WLW19, WLW20, ZTT18, CR14, Cha18, CT14c, CFF11, DB19, FWW18, FHM18, HB16b, Hou16, KM15, KTJ19, LY11, LZY^{+}15, MvdDW13, MFN^{+}18, PaCZ10, PRM14, RK19, SBCR10]. vaccination-subsidizing [KTJ19]. vaccinations [Che18]. vaccinators [CT18a]. Vaccine [JSC^{+}16, dM10b, BBT^{+}15, CRH^{+}15, Che18, GVSLG16, HHE19, KRDJ15, KR15, KJ19, NGJ^{+}14, SGA^{+}12, SSRA16, TWR^{+}18, dMP11]. vaccine-derived [KR15]. vaccines [MIJ16, MDMG14, RKK^{+}19]. vaginal [DBBM11]. valence [BTF19]. validated [LSMP14]. Validating [Tay16a]. Validation [GCB17, GMOP12, WWY12, AENK12, ALD^{+}11, DDF^{+}14, JAHKH12, PFJS15]. Valley [XSCS12]. valuations [CB10c]. Value [RMB15, WF17, BH13a, DBB^{+}17, Gar14, Gri15a, KMMS17, LV17, Mac14, MPC12]. valued [HMWB13]. Values [CRLH^{+}19, AIY16, Gra14, SS15b, vTMW15]. valve [ASC15, LCA^{+}15]. variabilities [Gol16]. Variability [BI12, MSL^{+}16, PN16, APS^{+}13, ALH16, BHKR14, CH16, DPRSS11, DTGC14, Fra19b, JB19b, KEHK17, KAKK19, LS11b, MKMT14, PW18b, PMSY17, RRRRSPTR16, STLJ18, WMPF^{+}15]. Variable [CB10c, AdVG15, DTP19, Gou16, Hal12, KL13, KK17a, KM19, KK13, LRL^{+}16, Laz13, Lug15, MGM13b, NP15, WTZ^{+}11]. variance [FS12, GJ15b, WCS19]. variant [MP13b, ZGW16]. variants [GZ19]. Variation [CBL14, BBS18b, CC10, CK15, DMM^{+}14, GGBW14, Gri15b, GAV15, Jam12, Jam13, LLH^{+}13, NNK10, PCB14, RMM^{+}16b, RR12, TA13, TEY16, WS10]. variational [CHH10, FMS^{+}12]. variations [FBU11, LVLC18, Mal15, Oka12]. varicella [GMR16]. Varicelloviruses [KB10a]. varied [SRS^{+}15]. varies [MG14b]. various [BI19, Kri18, MOSS15, OFT15, PWZ^{+}19, QLWC19, SM17a, SN16, TWC^{+}19, YW13a, ZMS11]. vary [NG11]. varying [BV13, DHP11, HRDL14, HRD14, LC17, NIiT19, NNvdB17, SMC^{+}13, YSST13]. Vascular [GCB17, HGM^{+}16, BJO^{+}16, BPPC15, FRTP13, Hal16, HTM15, HB13b, LSMC18, LC10, NF17, PHA^{+}17, SWG^{+}15, SBP13, SLD^{+}17, dlRVFK^{+}16, WFM^{+}13, WFC^{+}14, dG10b]. vascularised [dG10b]. Vascularized [RHS14b, LS15c]. vasculature [NJP18]. Vasomotion [POP12, HPP10]. vault [BFGAGA16]. Vector [Hua16, BZ10, CBC^{+}18, DDLW10, GYWJ10, GWCA14, HLHY17, HB16b, HTB^{+}13, HZL^{+}11, JWTW17, JW18, KHX^{+}19, KEKB18, KMCJ17, KSKK15, LYF^{+}15, MRPH17, MS16, MMG^{+}16, MBE11, MYOS14, NTED^{+}19, NHSX14, OL19, PRSC11, PKZ12, SHLL11, TXTW16, TEY10, WDL^{+}13, XWD^{+}10, ZDG^{+}10, pZWZ^{+}16, ZZC^{+}17, ZN18]. vector-bias [KMCJ17, OL19]. vector-borne [BZ10, CBC^{+}18, MRPH17, MYOS14, PKZ12]. vectored [WW12b]. vectors [DROC11, OA12, RC16]. Veelen [Mar11]. Vegetation [BFM10, WZ18, BS19b, SYSM11, TCR13]. vegetation-ring [SYSM11]. vegetative [RAAS15]. VEGF [LF19, LDF^{+}11a, LRA^{+}13, WHL13]. VEGF-A [LDF^{+}11a]. VEGFR1 [LSMC18]. VEGFR2 [BZJP18]. vein [GB13, LFM14]. velocimetry [MAR^{+}17c]. velocities [SL19]. Velocity [TCH14, GMMN18, LM15, WMT15]. vena [NPMM15]. venom [DS10b, WB15]. venoms [DS10b]. venous [FKDW15]. ventilation [Don16, SEK^{+}10, STKE12, SCT12]. ventilation-perfusion [SEK^{+}10]. Ventilatory [HLVR16]. ventral [SSKK17]. ventricle [KCS^{+}15]. ventricular [STNT17]. Venus [Leh18]. veridical [MMH10]. verification [ARR14, BFS18]. vermifuge [KKUM10]. vernalization [Sat10]. versatility [Cor18]. version [PMW12]. versus [ACT12, BH13c, BID15, CL17b, CFR13, GCM17b, HvLPdB15, LM15, LYK12, LYM^{+}19, ME15, RW15b, SL19, TKTB18, VBV10, ZSS10, Zol14, vDRT14]. vertebrate [Kri16, SWSLMJ19]. vertebrates [MPL16]. vertex [HNO18, MFKS13]. Vertical [PSL^{+}10, RRB10, Cle18a, Cle18b, HCS^{+}19, JLH^{+}15b, MYLK11, MA10, MK16b, NNK10, Gro17]. very [Ano19-50, HS14c, UI10, VWF19]. vesicle [NHM10, Pel18]. vesicles [PWC^{+}15, SWSLMJ19]. Vessel [Fok12, CPS19, FWR19, LSC^{+}18, dG10b]. vessels [CCG^{+}18, HPP10, LMM^{+}13, MGGM10a, POP12]. vestibular [LSB10, LB11, LB12]. Vestigial [Zhu11]. VI [NHTS14]. via [AFS19, AD16, BdABY13, CW13, CPW16, CCS18, DGL12, DGMY18, Don13, DvDBD15, FW15, FD11a, FE10, GTB10, HPB^{+}14, HY13, HKR^{+}19, JYZ18, KSvdH18, KDK14a, KDK14b, LBGW13, LHW14, LZ18, MMY^{+}12, MGT17, MKRE18, Mas14, NMZ19, NS13, STG19, SKK18, UI13, VSP11, WZM19, ZDY11]. Viability [Sue12, Pan11, TT17]. Vibration [AAJ15, THBM10, BB15]. Vibrio [FBK11, HK14, SDK11]. Vicodin [CFS^{+}19]. victory [NiTI18]. view [GJ15a, SLM17]. viewpoint [AH12b, ZZ10]. vigilance [BR16, RLW^{+}14]. vigilant [SP11]. village [WW12a]. villages [MYOS14]. villages-and-city [MYOS14]. villi [GSGV11, SSB^{+}15]. villosus [EBS11]. villous [LMJ^{+}16]. Viral [FLA^{+}19, Ric17, YHY14, BL14, CST16, FHS^{+}14, GC18, GN10, GW19, JYY^{+}18, KEHK17, Kor18, LM14, LG12, OS13, OMO13, PVB^{+}10, PGPD19, RM10, RW14a, STA15, SSCJ16, SH15, WR14, WML^{+}17, WBMH12, ZXC^{+}10]. viral-tumour [JYY^{+}18]. viremia [CLPZ14]. virion [Nak16a]. virions [MR11a]. virophage [TCW14]. virotherapies [ME15]. virotherapy [ARM18, JKF19, OAP14]. virtual [BYY^{+}11, DXW^{+}16, GBC^{+}16, KJM17, KMLT14, KPLT14, PWC^{+}12, SBR16]. virulence [AHMA^{+}19, FIJJ11, KGB13, LPS^{+}14, MA11b, MB12b, SB17, TDKJ15, WK18a, XSMF12]. virulent [ARZ15, PB18]. Virus [DK13a, XWC11, ZXH19, dTKDV15, AWK15, ARM18, BYY^{+}11, BMB^{+}18a, BB19, CM12, CS14b, FCC^{+}10, FHS^{+}14, GM16a, GC18, GRBL13, GW19, HAP^{+}16, HII^{+}19, HR12, IGL^{+}12, KMN^{+}15, KNA^{+}18, KW10, LLW15, LCJ18, LN19, LH14, LY14, LMYL15, MRPH17, MLT10, MG15, MSL^{+}16, NS11, Nak12, Nak16a, NKB16a, OA11, PaCZ10, PBB^{+}15b, PBSM19, PMKS^{+}15, RC16, RCL16, SBCR10, SZ15b, TXWW12, TXCW15, TCW14, UD10, ZZR10, ZZR13, ZRZ15, dRFI15]. virus-host [CS14b]. virus-immune [TXCW15]. virus-microbe [GW19]. Viruses [VW10, Gru18, Hua16, KR15, Mic15b, yTRSC13]. visco [JH12, PAV10]. visco-elastic [JH12]. Viscoelastic [DFG^{+}18a, WYL^{+}18, RL15, SDKF12]. viscoelasticity [CMJD11, KSSM11]. viscosities [PCT11, PCT19, Pan19]. viscosity [ASH15, IGG^{+}18, KH19, TM15a, dLMD^{+}11]. viscous [YG12]. visitor [DS19a]. visitors [HLTW14]. visual [Esc11, HRDL14, HRD14, Möl12, Yam16b]. visualize [KCJ^{+}11]. vitamin [SB15, WLD^{+}11]. vitro [AH12a, BE14, Car17, CTSL11, CCAdS13, DBBW09, DBBW11, GMOP12, HHA17, JRB^{+}16, KMN^{+}15, NTC^{+}11, Por13, PMKM10, POLT12, PEZ^{+}13, THC^{+}11, TSMB14, VG13, YD14, YD15]. vitulina [SH17]. vivax [CB13a, EN15, NKL10]. viverrini [BHO^{+}18]. vivid [RAR19]. vivo [ADR^{+}11b, BCPL10, BSC^{+}18, DBBW09, DBBW11, GM16a, LLB^{+}18, NAK^{+}11, POLT12, PEZ^{+}13, RGA^{+}10, WC10]. vocal [KFG15]. Voigt [SAD17]. volatile [vLFM^{+}19]. volition [HFT^{+}18]. voltage [LSP13, LSMP14, PDG17, ZCA^{+}14]. Volterra [DG11, DZJR10, Mil11, THM10]. volume [BHBH11, CCS18, CN12a, HGB^{+}13, KKO^{+}18, KFG^{+}14, MFM16, ØW16]. volumetric [KFG^{+}14]. voluntarism [CB10c]. Voluntary [Che18, FWW18, DDRP13]. volunteer [CGD13]. volunteering [XWZ10]. volunteers [WXWL12]. Vombatus [BCF19]. Voronoi [Arc16, MMY^{+}12, OO17a]. vortex [ASC16]. vortices [SWPC^{+}16]. voter [HH16]. Voting [AH15, AH19, YZZ13]. vowel [SFS13]. voxels [BDTR15]. Vpu [SPG^{+}18]. VR [GJ15b]. VR-BFDT [GJ15b]. vRNA [dTKDV15]. vs [BF13, EPP19, FNH10, Gol10b, JLH^{+}15b, RS12, SFV16, YBT^{+}17, vVGSE12].

Wagner [LP12a, TMHM12]. wait [RW15b]. Wake [Whi11, BVK10, Pat16, PR17, ZKR15]. wake-active [Pat16]. Wake-based [Whi11]. walk [CC11b, Che10, MSIR10, Rey13, RSC14]. walkers [NIiT18a]. walking [BRB18, CLP11, LGR^{+}12, SHW16, WHHS15, ZHAK14]. walks [BP15, Est10, HS14c, MAA18, PNK16, Rey13]. walks-based [Est10]. Wall [SWSLMJ19, AMM16, DQY^{+}15, DPCM16, DBJ12, EdKM11, EJK16, JSK16, Js12, JG14]. walls [Ren13, ZR16]. waning [FGH^{+}14]. war [IK16, Non10, SR10, SP14, vIK11]. Warburg [Arc14]. ward [CL17a]. Warfare [MMC19b]. warming [KTO^{+}14]. warning [JAB18]. warping [LBS^{+}14, SVSP15]. Wars [JV14]. was [AOR17, Sos18]. wash [MW13]. wasp [KP12, NBS^{+}13]. wasting [OVKL14]. Water [MDVT10, AMM16, BKR14, BBS18b, Che16, DAVS11, DDS13, GRCdL16, KFG^{+}14, KH19, KF10, MCRG18, MAM16, MYLK11, MS12b, SV14, SYSM11, TLR14, XWW^{+}19]. water-borne [MCRG18]. water-limited [SYSM11]. water-related [MCRG18]. water-uptake [Che16]. waterborne [CG14, ERT13]. watering [Gou15]. Watts [MZM^{+}16]. Wave [KM19, BJB18, CNL14, DOC17, Ghu18, LGPS17, LW18, OTTF11, PRV^{+}14, RLSM17, SM14, WML16, ZKR15]. waveforms [KTT^{+}19, SMM15]. wavefront [MMB11]. Wavelet [GZT15, AAJGCD15, JLX^{+}15, SVCS10, TWC^{+}19]. wavelets [ZLDZ13]. Waves [KKWA18, AGPK13, EG10, Fra16, IPS14, LDHD14, MDEKH13, PSP^{+}12, SMZ^{+}17]. Waxman [RRC^{+}11]. way [CRV16]. WB [MAFB18]. WB-PBPK [MAFB18]. WD40 [WDH^{+}16]. weak [GNB^{+}13, HM10b, RG10]. weaken [SJSK18]. weakening [RM13]. weakly [DHV19]. weapons [ACMK12]. weather [vNÁBG12]. web [AD13, AWRD15, DKF17, GDPPSS^{+}11, HAB11, HY10, Kři14, Mac10, MD16, OSV^{+}16, PBRW11, PDG12, QJR^{+}16, RPD14, RBB16b, SLW^{+}18a, WPPD16, WXC10, YXH^{+}14]. web-server [WXC10]. webs [BV13, BEK10, CAG13, KDG13, cLCJ^{+}10, MB15, MD16, MEJ18, RBB16a, RM13, TBM^{+}13, TD18, VL11, ZG10a, ZG10b, ZEJA11]. Weddell [HH10]. weed [RDMP11]. weekday [TC12]. weekly [TXTW16]. weevil [IST11]. weight [DMM^{+}14, iHM17, Hor11a, MFM16]. Weighted [YLF^{+}17, COWA11, LWW14, LXC15, MKF^{+}14, QLZQ11, YZ19, ZG10a, ZG10b]. Weighting [MLBA12, MLBA13]. weights [ZCXF13]. well [GF11, NWWL17, Pag19, PMW12, WL19, ZDF^{+}12, ZLW18]. well-known [PMW12]. well-mixed [NWWL17, Pag19, WL19, ZDF^{+}12, ZLW18]. wenxiang [Zho11]. west [DB18, DE17, MRPH17, RC16, ZXH19]. whales [FM15a, PGS10, WF18]. wheats [ASRM15]. Wheeler [YZW10]. Where [cSGFB17, MG14b, MDVT10, SI19]. whether [NRKE18, SN17]. Which [Tos11, VC10, CV15, Cle16, Di 16, HC17, HPvdB17, HBSF11, Lai18, LPH19]. while [GT11a]. white [LK15, MSB16, YLY15]. white-nose [MSB16]. whitefly [CEdLSG^{+}16]. whitetail [EEHMH18]. Who [ESG10, MFP^{+}14, PBSM19, RCL16]. Whole [BBS18b, MS14c, MAFB18, YY15]. whole-organ [MS14c]. whom [MFP^{+}14, Sak10]. wide [HHJR11, MKBE17a, MKBE17b, MKRE18, TDZ^{+}18]. Wild [CR14, Hou15, HT13, Hur12, Zha10a]. wild-domestic [HT13]. wild-type [Zha10a]. Will [ZL18b, HC17, RW14b]. Willard [BD10b, Jam13]. Williamson [ZG19]. Wilms [KRDJ15]. Wilson [CV19]. Wind [Sac13, TSdL14, ASL^{+}18, TTG19]. wind-up [TTG19]. Windkessel [AV19]. window [MM15a]. Wing [FH14, Che19, IS17, NC16, Zhu11]. Wingless [Zhu11, iHIM12]. wings [DS12, LS18]. winnerless [HGM15]. wins [ESG10]. winter [HHE19, SI12]. wires [AFD^{+}17]. wise [KI17, NWZ15, dSMP^{+}11]. wisely [FtL18]. With-in [RMF^{+}18]. withhold [GvVT14]. Within [Hur12, LMCW18, VDD12, AHKN10, BHBR12, CFZ14, CNL14, CMN^{+}11, DMM^{+}14, DRPM17, GDF18, HWW^{+}14, KLN^{+}12, KDFM14, Kor18, MFG^{+}13, NS11, NBBCC18, PJ13, PVB^{+}10, PBB^{+}15b, PML18, ST17a, TEY10, VW10, WBMH12, ZVMB10, dlCGSA16]. Within-host [Hur12, LMCW18, VDD12, CFZ14, DRPM17, Kor18, MFG^{+}13, NBBCC18, PVB^{+}10, PBB^{+}15b, WBMH12]. within-vector [TEY10]. without [AS12, IvdSW16, MBBD13]. WLFs [Pie12]. Wnt11 [SSKK17]. Wolbachia [HHT^{+}19, HLH^{+}18, LL18, SKK^{+}12, TGC^{+}17, ZLT^{+}19]. Wolbachia-infected [ZLT^{+}19]. wombats [BCF19]. women [KMH19]. Womersley [AV19]. word [ABG^{+}13, DLYZ11, DYS^{+}13, DLYW13, SS19, SD16, YZZ12, YZZ13, YW13a, YW13b, YZFY16, ZLY14]. words [GAB14, HRC^{+}12, SFV16]. worker [OAV^{+}16]. workers [LWY^{+}16]. Working [TIMI^{+}15, SMM^{+}13]. World [JV14, CKS15, Fra16, HM10b, Koe15, MZM^{+}16, Mau15, Sos18, dS15a]. worm [CHH10]. worm-like [CHH10]. worms [HASM17]. would [Di 18]. Wound [FBFM12, CSDR15, GBJ18, MHMM11, NCLB16, PC10b, VG13, VAG16, YWP13]. wounds [BPS15, BMM^{+}14, TBMM19]. Wright [CS18, LPH19, SW17, TT10b, jWlGQ19, Wax11b, ZLW14, ZGW16]. wrong’ [Mar11]. wrongs [Pan11]. Wu [JN14]. Wyman [Oko16].

X [Gar14, AKR^{+}18, CZPC^{+}18, Kon12, KPS17, Mic15b, PWG16, Sek12, Unc11, Xie11, ZZ14a]. X-linkage [Gar14]. X-linked [AKR^{+}18, Sek12]. X-machine [PWG16]. X-ray [CZPC^{+}18, ZZ14a]. Xanthomonas [JSF^{+}11]. XARNES [Kon12]. xenobiotic [SBMH10]. xenograft [TTB^{+}18]. xenografted [LBW^{+}13]. Xenopus [YM14, SSKK17]. Xi’an [TXY^{+}12]. XII [TSP15]. Xinjiang [WZJ^{+}13]. Xor [Mou12b]. Xylella [AHD^{+}18]. xylem [DDB10, LMM^{+}13, LSC^{+}18].

Y-linked [AGM11, GGM12]. Y-STR [AEM14, AEM13]. year [Dol16]. years [Kir11, Mar17a]. Yeast [TK10b, ADC19, JL16, LSY^{+}10, MMK16, PRM11, TGB^{+}18, WKS14, WZWY11]. Yemen [HWGW18]. Yield [DBJ12, BR14, BKL14, GL12a, GK13, KPG19]. yields [ARZ15, JAK19, KSA16, PMKM10, TM15b, ZEJA11]. yjdF [GWW^{+}18]. yo [IPMH12]. yo-yo [IPMH12]. young [SRV11]. Yule [MS14d, MC15b].

Zanamivir [TLT^{+}15]. Zealand [MR11b]. zebrafish [Hal17, YSI17, YSI18, ZLB^{+}18, ZBA18]. Zero [IM18, JRG14a, JRG14b, JRG14c, XG12, BR13, CZ14, HWTN15, Mog15]. zero-cost [BR13]. Zero-determinant [IM18, CZ14, HWTN15]. zero-inflated [Mog15]. Zero-order [JRG14a, JRG14b, JRG14c, XG12]. Zika [CBC^{+}18]. zinc [AHD^{+}18, HMMSRSD15]. Ziv [YHY14, ZLY14]. zona [KG12]. zonation [SBMH10]. zones [ACMK12]. zooplankton [BSC19, Mor10b, MA10, MK16b]. zygomycete [ESW13].

References

Allena:2012:RTL

[AA12] Rachele Allena and Denis Aubry. ‘run-and-tumble’ or ‘look-and-run’? A mechanical model to explore the behavior of a migrating amoeboid cell. Journal of Theoretical Biology, 306(??):15–31, August 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001737.

Aranda-Anzaldo:2016:IMC

[AA16] Armando Aranda-Anzaldo. The interphase mammalian chromosome as a structural system based on tensegrity. Journal of Theoretical Biology, 393(??):51–59, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000187.

Abner:2014:SCM

[AAAV14] Kristo Abner, Tõnis Aaviksaar, Kaarel Adamberg, and Raivo Vilu. Single-cell model of prokaryotic cell cycle. Journal of Theoretical Biology, 341(??):78–87, January 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004815.

Ali:2016:BGF

[AABS16] N. Ali, Z. Asghar, O. Anwar Bég, and M. Sajid. Bacterial gliding fluid dynamics on a layer of non-Newtonian slime: Perturbation and numerical study. Journal of Theoretical Biology, 397(??):22–32, May 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001016.

Al-Anzi:2017:MAM

[AAGO^{+}17] Bader Al-Anzi, Sherif Gerges, Noah Olsman, Christopher Ormerod, Georgios Piliouras, John Ormerod, and Kai Zinn. Modeling and analysis of modular structure in diverse biological networks. Journal of Theoretical Biology, 422(??):18–30, June 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301534.

Arani:2015:VBF

[AAJ15] A. Ghorbanpour Arani, M. Abdollahian, and M. H. Jalaei. Vibration of bioliquid-filled microtubules embedded in cytoplasm including surface effects using modified couple stress theory. Journal of Theoretical Biology, 367(??):29–38, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006717.

Arango-Argoty:2015:FES

[AAJGCD15] G. A. Arango-Argoty, J. A. Jaramillo-Garzón, and G. Castellanos-Domínguez. Feature extraction by statistical contact potentials and wavelet transform for predicting subcellular localizations in gram negative bacterial proteins. Journal of Theoretical Biology, 364(??):121–130, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005384.

Amiri:2016:NBI

[AANF16] Masoud Amiri, Mahmood Amiri, Soheila Nazari, and Karim Faez. A new bio-inspired stimulator to suppress hypersynchronized neural firing in a cortical network. Journal of Theoretical Biology, 410(??):107–118, December 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302958.

Andre:2011:SOE

[AB11] Jean-Baptiste André and Nicolas Baumard. Social opportunities and the evolution of fairness. Journal of Theoretical Biology, 289(??):128–135, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004012.

Andrews:2016:ISD

[AB16] Michael A. Andrews and Chris T. Bauch. The impacts of simultaneous disease intervention decisions on epidemic outcomes. Journal of Theoretical Biology, 395(??):1–10, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000667.

Alphey:2011:MRG

[ABA11] Nina Alphey, Michael B. Bonsall, and Luke Alphey. Modeling resistance to genetic control of insects. Journal of Theoretical Biology, 270(1):42–55, February 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006016.

Abbott:2010:BEC

[Abb10] Kevin R. Abbott. Background evolution in camouflage systems: a predator–prey/pollinator-flower game. Journal of Theoretical Biology, 262(4):662–678, February 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004214.

Abdelheq:2019:PST

[ABB^{+}19] Mezouaghi Abdelheq, Omar Belhamiti, Leila Bouzid, Deccy Y. Trejos, and Jose C. Valverde. A predictive spatio-temporal model for bovine Babesiosis epidemic transmission. Journal of Theoretical Biology, 480(??):192–204, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303005.

Ascenzi:2011:ELL

[ABD^{+}11] Maria-Grazia Ascenzi, Christian Blanco, Ian Drayer, Hannah Kim, Ryan Wilson, Kelsey N. Retting, Karen M. Lyons, and George Mohler. Effect of localization, length and orientation of chondrocytic primary cilium on murine growth plate organization. Journal of Theoretical Biology, 285(1):147–155, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003158.

Agliari:2015:ASD

[ABD^{+}15] Elena Agliari, Adriano Barra, Gino Del Ferraro, Francesco Guerra, and Daniele Tantari. Anergy in self-directed B lymphocytes: a statistical mechanics perspective. Journal of Theoretical Biology, 375(??):21–31, June 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400277X.

Afreixo:2013:BWS

[ABG^{+}13] Vera Afreixo, Carlos A. C. Bastos, Sara P. Garcia, João M. O. S. Rodrigues, Armando J. Pinho, and Paulo J. S. G. Ferreira. The breakdown of the word symmetry in the human genome. Journal of Theoretical Biology, 335(??):153–159, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003044.

Agliari:2011:TPI

[ABGM11] Elena Agliari, Adriano Barra, Francesco Guerra, and Francesco Moauro. A thermodynamic perspective of immune capabilities. Journal of Theoretical Biology, 287(??):48–63, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003778.

Almeida:2011:MMD

[ABH^{+}11] Luís Almeida, Patrizia Bagnerini, Abderrahmane Habbal, Stéphane Noselli, and Fanny Serman. A mathematical model for dorsal closure. Journal of Theoretical Biology, 268(1):105–119, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005035.

Ang:2010:CUI

[ABIM10] Jordan Ang, Sangram Bagh, Brian P. Ingalls, and David R. McMillen. Considerations for using integral feedback control to construct a perfectly adapting synthetic gene network. Journal of Theoretical Biology, 266(4):723–738, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003917.

Amiri:2012:FCA

[ABJ12] Mahmood Amiri, Fariba Bahrami, and Mahyar Janahmadi. Functional contributions of astrocytes in synchronization of a neuronal network model. Journal of Theoretical Biology, 292(??):60–70, January 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004693.

Asano:2012:QLM

[ABK^{+}12] Masanari Asano, Irina Basieva, Andrei Khrennikov, Masanori Ohya, Yoshiharu Tanaka, and Ichiro Yamato. Quantumlike model of diauxie in Escherichia coli: Operational description of precultivation effect. Journal of Theoretical Biology, 314(??):130–137, December 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004444.

Aguiar:2011:RSI

[ABKS11] Maíra Aguiar, Sebastien Ballesteros, Bob W. Kooi, and Nico Stollenwerk. The role of seasonality and import in a minimalistic multi-strain dengue model capturing differences between primary and secondary infections: Complex dynamics and its implications for data analysis. Journal of Theoretical Biology, 289(??):181–196, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004462.

Arnoldi:2018:HER

[ABLH18] J.-F. Arnoldi, A. Bideault, M. Loreau, and B. Haegeman. How ecosystems recover from pulse perturbations: a theory of short- to long-term responses. Journal of Theoretical Biology, 436(??):79–92, January 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304587.

Aquino:2015:CEE

[ABN15] Tomás Aquino, Diogo Bolster, and Ana Nunes. Characterization of the endemic equilibrium and response to mutant injection in a multi-strain disease model. Journal of Theoretical Biology, 368(??):27–36, March 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006900.

Afreixo:2011:GAD

[ABP^{+}11] Vera Afreixo, Carlos A. C. Bastos, Armando J. Pinho, Sara P. Garcia, and Paulo J. S. G. Ferreira. Genome analysis with distance to the nearest dissimilar nucleotide. Journal of Theoretical Biology, 275(1):52–58, April 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000646.

Allen:2011:MLR

[ABR11] R. J. Allen, I. D. L. Bogle, and A. J. Ridley. A model of localised Rac1 activation in endothelial cells due to fluid flow. Journal of Theoretical Biology, 280(1):34–42, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001676.

Abrams:2014:EBM

[Abr14] Peter A. Abrams. The evolutionary and behavioral modification of consumer responses to environmental change. Journal of Theoretical Biology, 343(??):162–173, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005171.

AlBasir:2019:RFA

[ABR19] Fahad Al Basir, Arnab Banerjee, and Santanu Ray. Role of farming awareness in crop pest management — a mathematical model. Journal of Theoretical Biology, 461(??):59–67, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305241.

Alvarez:2019:NMM

[ABV19] Robinson F. Alvarez, José A. M. Barbuto, and Roberto Venegeroles. A nonlinear mathematical model of cell-mediated immune response for tumor phenotypic heterogeneity. Journal of Theoretical Biology, 471(??):42–50, June 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301328.

Asano:2012:MTM

[AC12] Erika Asano and Deby L. Cassill. Modeling temperaturemediated fluctuation in colony size in the fire ant, solenopsis invicta. Journal of Theoretical Biology, 305(??):70–77, July 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001282.

Aram:2015:TLC

[AC15] Reza Zohouri Aram and Nasrollah Moghadam Charkari. A two-layer classification framework for protein fold recognition. Journal of Theoretical Biology, 365(??):32–39, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314005712.

Arenas:2011:JEC

[ACCR11] Alex Arenas, Juan Camacho, José A. Cuesta, and Rubén J. Requejo. The joker effect: Cooperation driven by destructive agents. Journal of Theoretical Biology, 279(1):113–119, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001639.

Aguero-Chapin:2011:NLM

[ACdlRMR^{+}11] Guillermin Agüero-Chapin, Gustavo A. de la Riva, Reinaldo Molina-Ruiz, Aminael Sánchez-Rodríguez, Gisselle Pérez-Machado, Vítor Vasconcelos, and Agostinho Antunes. Nonlinear models based on simple topological indices to identify RNase III protein members. Journal of Theoretical Biology, 273(1):167–178, March 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310006740.

Alshehri:2015:EBD

[ACH15] Mansoor H. Alshehri, Barry J. Cox, and James M. Hill. Energy behaviour for DNA translocation through graphene nanopores. Journal of Theoretical Biology, 387(??):68–75, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004634.

Akman:2018:MSI

[ACH18] Olcay Akman, Timothy D. Comar, and Daniel Hrozencik. Model selection for integrated pest management with stochasticity. Journal of Theoretical Biology, 442(??):110–122, April 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317305386.

Aastrand:2019:PLB

[ÅCHS19] Mia Åstrand, Julia Cuellar, Jukka Hytönen, and Tiina A. Salminen. Predicting the ligand-binding properties of Borrelia burgdorferi s.s. Bmp proteins in light of the conserved features of related Borrelia proteins. Journal of Theoretical Biology, 462(??):97–108, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305514.

Aftabi:2016:SAI

[ACM16] Younes Aftabi, Abasalt Hosseinzadeh Colagar, and Faramarz Mehrnejad. An in silico approach to investigate the source of the controversial interpretations about the phenotypic results of the human AhR-gene g1661a polymorphism. Journal of Theoretical Biology, 393(??):1–15, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000023.

Allstadt:2012:ICI

[ACMK12] Andrew Allstadt, Thomas Caraco, F. Molnár, and G. Korniss. Interference competition and invasion: Spatial structure, novel weapons and resistance zones. Journal of Theoretical Biology, 306(??):46–60, August 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001981.

Abbiati:2017:PBM

[ACR^{+}17] Roberto Andrea Abbiati, Petra Cagnardi, Giuliano Ravasio, Roberto Villa, and Davide Manca. A physiologically based model for tramadol pharmacokinetics in horses. Journal of Theoretical Biology, 429(??):46–51, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303053.

Agarwala:2012:PFV

[ACT12] Edward K. Agarwala, Hillel J. Chiel, and Peter J. Thomas. Pursuit of food versus pursuit of information in a Markovian perception-action loop model of foraging. Journal of Theoretical Biology, 304(??):235–272, July 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000896.

Ashlock:2010:EIR

[ACvKA10] Daniel Ashlock, Elizabeth L. Clare, Taika E. von Königslöw, and Wendy Ashlock. Evolution and instability in ring species complexes: An in silico approach to the study of speciation. Journal of Theoretical Biology, 264(4):1202–1213, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001402.

Allhoff:2013:WDE

[AD13] Korinna T. Allhoff and Barbara Drossel. When do evolutionary food web models generate complex networks? Journal of Theoretical Biology, 334(??):122–129, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002804.

Agrahari:2015:III

[AD15] Ashish Agrahari and C. George Priya Doss. Impact of i30t and i30m substitution in MPZ gene associated with Dejerine– Sottas syndrome type B (DSSB): a molecular modeling and dynamics. Journal of Theoretical Biology, 382(??):23–33, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002982.

Amiri:2016:CGD

[AD16] Saeid Amiri and Ivo D. Dinov. Comparison of genomic data via statistical distribution. Journal of Theoretical Biology, 407(??):318–327, October 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302168.

Adams:2016:HDD

[Ada16] Ben Adams. Household demographic determinants of Ebola epidemic risk. Journal of Theoretical Biology, 392(??):99–106, March 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005809.

Agusto:2013:IBN

[ADB^{+}13] Folashade B. Agusto, Sara Y. Del Valle, Kbenesh W. Blayneh, Calistus N. Ngonghala, Maria J. Goncalves, Nianpeng Li, Ruijun Zhao, and Hongfei Gong. The impact of bed-net use on malaria prevalence. Journal of Theoretical Biology, 320(??):58–65, March 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006315.

Ali:2019:ADR

[ADC19] Qasim Ali, Riccardo Dainese, and Marija Cvijovic. Adaptive damage retention mechanism enables healthier yeast population. Journal of Theoretical Biology, 473(??):52–66, July 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301468.

Ali:2014:BAD

[ADCG14] Shimaila Ali, Jin Duan, Trevor C. Charles, and Bernard R. Glick. A bioinformatics approach to the determination of genes involved in endophytic behavior in Burkholderia spp. Journal of Theoretical Biology, 343(??):193–198, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313004992.

Apri:2012:CRP

[AdGM12] Mochamad Apri, Maarten de Gee, and Jaap Molenaar. Complexity reduction preserving dynamical behavior of biochemical networks. Journal of Theoretical Biology, 304(??):16–26, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001439.

Angulo:2013:SDT

[AdlPLMZ13] Óscar Angulo, Rafael Bravo de la Parra, Juan C. López-Marcos, and Miguel A. Zavala. Stand dynamics and tree coexistence in an analytical structured model: The role of recruitment. Journal of Theoretical Biology, 333(??):91–101, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002415.

Armiento:2016:EMM

[ADMR16] Aurora Armiento, Marie Doumic, Philippe Moireau, and H. Rezaei. Estimation from moments measurements for amyloid depolymerisation. Journal of Theoretical Biology, 397(??):68–88, May 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001399.

Allman:2011:DST

[ADR11a] Elizabeth S. Allman, James H. Degnan, and John A. Rhodes. Determining species tree topologies from clade probabilities under the coalescent. Journal of Theoretical Biology, 289(??):96–106, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003924.

Aston:2011:MAP

[ADR^{+}11b] Philip J. Aston, Gianne Derks, Adewale Raji, Balaji M. Agoram, and Piet H. van der Graaf. Mathematical analysis of the pharmacokinetic-pharmacodynamic (PKPD) behaviour of monoclonal antibodies: Predicting in vivo potency. Journal of Theoretical Biology, 281(1):113–121, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002311.

Agrahari:2019:MIG

[ADS^{+}19] Ashish Kumar Agrahari, George Priya C. Doss, R. Siva, R. Magesh, and Hatem Zayed. Molecular insights of the g2019s substitution in LRRK2 kinase domain associated with Parkinson’s disease: a molecular dynamics simulation approach. Journal of Theoretical Biology, 469(??):163–171, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300967.

Amarante:2014:EAS

[AdSG14] Tauanne D. Amarante, Jafferson K. L. da Silva, and Guilherme J. M. Garcia. Effects of airway surface liquid height on the kinetics of extracellular nucleotides in airway epithelia. Journal of Theoretical Biology, 363(??):427–435, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004913.

Accatino:2010:TGC

[ADV^{+}10] Francesco Accatino, Carlo De Michele, Renata Vezzoli, Davide Donzelli, and Robert J. Scholes. Tree-grass co-existence in savanna: Interactions of rain and fire. Journal of Theoretical Biology, 267(2):235–242, November 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004212.

Aguilee:2015:DER

[AdVG15] Robin Aguilée, Pierre de Villemereuil, and Jean-Michel Guillon. Dispersal evolution and resource matching in a spatially and temporally variable environment. Journal of Theoretical Biology, 370(??):184–196, April 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000272.

Aanen:2017:SBE

[AE17] Duur K. Aanen and Paul Eggleton. Symbiogenesis: Beyond the endosymbiosis theory? Journal of Theoretical Biology, 434(??):99–103, December 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303612.

Adams:2019:RRS

[AEE19] Z. P. Adams, J. Ehlting, and R. Edwards. The regulatory role of shikimate in plant phenylalanine metabolism. Journal of Theoretical Biology, 462(??):158–170, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305526.

Andersen:2013:DLE

[AEM13] Mikkel Meyer Andersen, Poul Svante Eriksen, and Niels Morling. The discrete Laplace exponential family and estimation of y-STR haplotype frequencies. Journal of Theoretical Biology, 329(??):39–51, July 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519313001136. See corrigendum [AEM14].

Andersen:2014:CDL

[AEM14] Mikkel Meyer Andersen, Poul Svante Eriksen, and Niels Morling. Corrigendum to “The discrete Laplace exponential family and estimation of Y-STR haplotype frequencies” [J. Theor. Biol. 329 (2013) 39–51]. Journal of Theoretical Biology, 350(??):110, June 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001362. See [AEM13].

Ackermann:2012:RTN

[AENK12] J. Ackermann, J. Einloft, J. Nöthen, and I. Koch. Reduction techniques for network validation in systems biology. Journal of Theoretical Biology, 315(??):71–80, December 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200464X.

Arciero:2013:MIB

[AES^{+}13] Julia Arciero, G. Bard Ermentrout, Richard Siggers, Amin Afrazi, David Hackam, Yoram Vodovotz, and Jonathan Rubin. Modeling the interactions of bacteria and toll-like receptormediated inflammation in necrotizing enterocolitis. Journal of Theoretical Biology, 321(??):83–99, March 21, 2013.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006261.

Akbaripour-Elahabad:2016:RTS

[AEZR^{+}16] Mohammad Akbaripour-Elahabad, Javad Zahiri, Reza Rafeh, Morteza Eslami, and Mahboobeh Azari. rpiCOOL: a tool for in silico RNA-protein interaction detection using random forest. Journal of Theoretical Biology, 402(??):1–8, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300534.

Abrams:2010:PPA

[AF10] Peter A. Abrams and Simon R. Fung. Prey persistence and abundance in systems with intraguild predation and type-2 functional responses. Journal of Theoretical Biology, 264(3):1033–1042, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001207.

Aydogdu:2017:MBW

[AFD^{+}17] A. Aydoğdu, P. Frasca, C. D’Apice, R. Manzo, J. M. Thornton, B. Gachomo, T. Wilson, B. Cheung, U. Tariq, W. Saidel, and B. Piccoli. Modeling birds on wires. Journal of Theoretical Biology, 415(??):102–112, February 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304064.

Ahmed:2010:EMF

[AFM10] Ahmed Ahmed, Gabriel Frey, and Christian J. Michel. Essential molecular functions associated with the circular code evolution. Journal of Theoretical Biology, 264(2):613–622, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000731.

Adilina:2019:EDB

[AFS19] Sheikh Adilina, Dewan Md Farid, and Swakkhar Shatabda. Effective DNA binding protein prediction by using key features via Chou’s general PseAAC. Journal of Theoretical Biology, 460(??):64–78, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305034.

Agnew:2014:DBM

[AGB^{+}14] D. J. G. Agnew, J. E. F. Green, T. M. Brown, M. J. Simpson, and B. J. Binder. Distinguishing between mechanisms of cell aggregation using pair-correlation functions. Journal of Theoretical Biology, 352(??):16–23, July 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001088.

Angulo:2018:IRE

[AGC18] O. Angulo, O. Gandrillon, and F. Crauste. Investigating the role of the experimental protocol in phenylhydrazineinduced anemia on mice recovery. Journal of Theoretical Biology, 437(??):286–298, January 21, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305003.

Arias:2018:EPM

[AGGME18] Juddy H. Arias, Jesus Gómez-Gardeñes, Sandro Meloni, and Ernesto Estrada. Epidemics on plants: Modeling long-range dispersal on spatially embedded networks. Journal of Theoretical Biology, 453(??):1–13, September 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302303.

Alsmeyer:2011:LGF

[AGM11] G. Alsmeyer, C. Gutiérrez, and R. Martínez. Limiting genotype frequencies of y-linked genes through bisexual branching processes with blind choice. Journal of Theoretical Biology, 275(1):42–51, April 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000361.

Abramson:2013:QOD

[AGPK13] Guillermo Abramson, Luca Giuggioli, Robert R. Parmenter, and V. M. Kenkre. Quasi-one-dimensional waves in rodent populations in heterogeneous habitats: a consequence of elevational gradients on spatio-temporal dynamics. Journal of Theoretical Biology, 319(??):96–101, February 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006157.

Albano:2015:EDE

[AGRR^{+}15] Giuseppina Albano, Virginia Giorno, Patricia Román-Román, Sergio Román-Román, and Francisco Torres-Ruiz. Estimating and determining the effect of a therapy on tumor dynamics by means of a modified Gompertz diffusion process. Journal of Theoretical Biology, 364(??):206–219, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314005530.

Albano:2011:IET

[AGRRTR11] Giuseppina Albano, Virginia Giorno, Patricia Román-Román, and Francisco Torres-Ruiz. Inferring the effect of therapy on tumors showing stochastic Gompertzian growth. Journal of Theoretical Biology, 276(1):67–77, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100066X.

Antoneli:2018:HFF

[AGS18] Fernando Antoneli, Martin Golubitsky, and Ian Stewart. Homeostasis in a feed forward loop gene regulatory motif. Journal of Theoretical Biology, 445(??):103–109, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300869.

Aita:2012:BCB

[AH12a] Takuyo Aita and Yuzuru Husimi. Biophysical connection between evolutionary dynamics and thermodynamics in in vitro evolution. Journal of Theoretical Biology, 294(??):122–129, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005595.

Aita:2012:TEC

[AH12b] Takuyo Aita and Yuzuru Husimi. Toward an evolutionary containment of evolving pathogen-receptors by using an ensemble of multiple mutant ligands: From the viewpoint of fitness landscape in sequence space. Journal of Theoretical Biology, 296(??):49–55, March 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100600X.

Ali:2015:CMP

[AH15] Farman Ali and Maqsood Hayat. Classification of membrane protein types using voting feature interval in combination with Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 384(??):78–83, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003926.

Ali:2016:MLA

[AH16] Farman Ali and Maqsood Hayat. Machine learning approaches for discrimination of extracellular matrix proteins using hybrid feature space. Journal of Theoretical Biology, 403(??):30–37, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300893.

Alves:2017:HCA

[AH17] Mickaël Teixeira Alves and Frank M. Hilker. Hunting cooperation and Allee effects in predators. Journal of Theoretical Biology, 419(??):13–22, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300577.

Akbar:2018:ISI

[AH18] Shahid Akbar and Maqsood Hayat. iMethyl-STTNC: Identification of N^{6}-methyladenosine sites by extending the idea of SAAC into Chou’s PseAAC to formulate RNA sequences. Journal of Theoretical Biology, 455(??):205–211, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303436.

Ahmad:2019:MMV

[AH19] Jamal Ahmad and Maqsood Hayat. MFSC: Multi-voting based feature selection for classification of Golgi proteins by adopting the general form of Chou’s PseAAC components. Journal of Theoretical Biology, 463(??):99–109, February 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318306076.

Arias:2014:MMC

[AHAFA14] Clemente F. Arias, Miguel A. Herrero, Francisco J. Acosta, and Cristina Fernandez-Arias. A mathematical model for a T cell fate decision algorithm during immune response. Journal of Theoretical Biology, 349(??):109–120, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000587.

Aggarwal:2018:FMA

[AHD^{+}18] Manu Aggarwal, M. Y. Hussaini, Leonardo De La Fuente, Fernando Navarrete, and N. G. Cogan. A framework for model analysis across multiple experiment regimes: Investigating effects of zinc on Xylella fastidiosa as a case study. Journal of Theoretical Biology, 457(??):88–100, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304077.

Ahn:2018:INE

[AHJ^{+}18a] Inkyung Ahn, Seongman Heo, Seunghyun Ji, Kyung Hyun Kim, Taehwan Kim, Eun Joo Lee, Jooyoung Park, and Keehoon Sung. Investigation of nonlinear epidemiological models for analyzing and controlling the MERS outbreak in Korea. Journal of Theoretical Biology, 437(??):17–28, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304666.

Arif:2018:ITL

[AHJ18b] Muhammad Arif, Maqsood Hayat, and Zahoor Jan. iMem-2LSAAC: a two-level model for discrimination of membrane proteins and their types by extending the notion of SAAC into Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 442(??):11–21, April 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830016X.

Ait-Haddou:2010:CBD

[AHKN10] Rachid Ait-Haddou, Yoshihisa Kurachi, and Taishin Nomura. On calcium-buffer dynamics within the excess buffer regime. Journal of Theoretical Biology, 264(1):55–65, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000299.

He:2012:GRP

[aHLZ^{+}12] Ping an He, Dan Li, Yanping Zhang, Xin Wang, and Yuhua Yao. A 3d graphical representation of protein sequences based on the Gray code. Journal of Theoretical Biology, 304(??):81– 87, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001476.

Alonso-Hearn:2019:DVM

[AHMA^{+}19] Marta Alonso-Hearn, Gesham Magombedze, Naiara Abendaño, Mariana Landin, and Ramon A. Juste. Deciphering the virulence of mycobacterium avium subsp. paratuberculosis isolates in animal macrophages using mathematical models. Journal of Theoretical Biology, 468(??):82–91, May 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300347.

Anceschi:2019:NNF

[AHP^{+}19] Niccolo Anceschi, Jorge Hidalgo, Carlos A. Plata, Tommaso Bellini, Amos Maritan, and Samir Suweis. Neutral and niche forces as drivers of species selection. Journal of Theoretical Biology, 483(??):Article 109969, December 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303066.

Allman:2010:ETF

[AHR10] Elizabeth S. Allman, Mark T. Holder, and John A. Rhodes. Estimating trees from filtered data: Identifiability of models for morphological phylogenetics. Journal of Theoretical Biology, 263(1):108–119, March 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005682.

Al-Husari:2013:RTI

[AHW13] Maymona Al-Husari and Steven D. Webb. Regulation of tumour intracellular pH: a mathematical model examining the interplay between H^{+} and lactate. Journal of Theoretical Biology, 322(??):58–71, April 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300026X.

Ashby:2019:URE

[AIB^{+}19] Ben Ashby, Ryosuke Iritani, Alex Best, Andy White, and Mike Boots. Understanding the role of eco-evolutionary feedbacks in host-parasite coevolution. Journal of Theoretical Biology, 464(??):115–125, March 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306313.

Aita:2016:IFV

[AIY16] Takuyo Aita, Norikazu Ichihashi, and Tetsuya Yomo. Inference of fitness values and putative appearance time points for evolvable self-replicating molecules from time series of occurrence frequencies in an evolution reactor. Journal of Theoretical Biology, 401(??):38–42, July 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630042X.

Al-Jumaily:2012:EAS

[AJC12] Ahmed Al-Jumaily and Leizhi Chen. Estimation of airway smooth muscle stiffness changes due to length oscillation using artificial neural network. Journal of Theoretical Biology, 310(??):115–118, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002974.

Al-Jumaily:2011:FMM

[AJD11] A. M. Al-Jumaily and Y. Du. Fading memory model for airway smooth muscle dynamic response. Journal of Theoretical Biology, 283(1):10–13, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002669.

Abou-Jaoude:2019:LBD

[AJM19] Wassim Abou-Jaoudé and Pedro T. Monteiro. On logical bifurcation diagrams. Journal of Theoretical Biology, 466 (??):39–63, April 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519319300086.

Armstrong:2011:MPD

[AK11] Kyle N. Armstrong and Leonard J. Kerry. Modelling the prey detection performance of Rhinonicteris aurantia (Chiroptera: Hipposideridae) in different atmospheric conditions discounts the notional role of relative humidity in adaptive evolution. Journal of Theoretical Biology, 278(1):44–54, June 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001287.

Adler:2013:MCS

[AK13] Frederick R. Adler and Peter S. Kim. Models of contrasting strategies of rhinovirus immune manipulation. Journal of Theoretical Biology, 327(??):1–10, June 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000787.

Akita:2015:TLI

[AK15] Tetsuya Akita and Masashi Kamo. Theoretical lessons for increasing algal biofuel: Evolution of oil accumulation to avert carbon starvation in microalgae. Journal of Theoretical Biology, 380(??):183–191, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002623.

Asatryan:2016:EGI

[AK16] Ani D. Asatryan and Natalia L. Komarova. Evolution of genetic instability in heterogeneous tumors. Journal of Theoretical Biology, 396(??):1–12, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005937.

Akcay:2012:IFI

[Akç12] Erol Akçay. Incentives in the family II: Behavioral dynamics and the evolution of non-costly signaling. Journal of Theoretical Biology, 294(??):9–18, February 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005546.

Apri:2014:MCD

[AKdV^{+}14] Mochamad Apri, Johannes Kromdijk, Pieter H. B. de Visser, Maarten de Gee, and Jaap Molenaar. Modelling cell division and endoreduplication in tomato fruit pericarp. Journal of Theoretical Biology, 349(??):32–43, May 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000484.

Aldebert:2018:SSP

[AKNP18] Clement Aldebert, Bob W. Kooi, David Nerini, and Jean-Christophe Poggiale. Is structural sensitivity a problem of oversimplified biological models? Insights from nested Dynamic Energy Budget models. Journal of Theoretical Biology, 448(??):1–8, July 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301334.

Agrahari:2018:SIH

[AKR^{+}18] Ashish Kumar Agrahari, Amit Kumar, Siva R., Hatem Zayed, and George Priya Doss C. Substitution impact of highly conserved arginine residue at position 75 in GJB1 gene in association with X-linked Charcot–Marie-tooth disease: a computational study. Journal of Theoretical Biology, 437(??):305–317, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304976.

Akhmetzhanov:2019:MBT

[AKS^{+}19] Andrei R. Akhmetzhanov, Jong Wook Kim, Ryan Sullivan, Robert A. Beckman, Pablo Tamayo, and Chen-Hsiang Yeang. Modelling bistable tumour population dynamics to design effective treatment strategies. Journal of Theoretical Biology, 474(??):88–102, August 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301900.

Ajelli:2017:ECP

[AL17] Marco Ajelli and Maria Litvinova. Estimating contact patterns relevant to the spread of infectious diseases in Russia. Journal of Theoretical Biology, 419(??):1–7, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300504.

Alamir:2015:PCC

[Ala15] Mazen Alamir. On probabilistic certification of combined cancer therapies using strongly uncertain models. Journal of Theoretical Biology, 384(??):59–69, November 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004002.

Aylaj:2011:MPD

[ALD^{+}11] Bouchra Aylaj, Flavie Luciani, Veronique Delmas, Lionel Larue, and Florian De Vuyst. Melanoblast proliferation dynamics during mouse embryonic development. Modeling and validation. Journal of Theoretical Biology, 276(1):86–98, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000671.

Artalejo:2010:QSR

[ALH10] J. R. Artalejo and M. J. Lopez-Herrero. Quasi-stationary and ratio of expectations distributions: a comparative study. Journal of Theoretical Biology, 266(2):264–274, September 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310003206.

Arnoldi:2016:RRV

[ALH16] J-F. Arnoldi, M. Loreau, and B. Haegeman. Resilience, reactivity and variability: a mathematical comparison of ecological stability measures. Journal of Theoretical Biology, 389(??):47–59, January 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005056.

Allesina:2011:PTR

[All11] Stefano Allesina. Predicting trophic relations in ecological networks: a test of the allometric diet breadth model. Journal of Theoretical Biology, 279(1):161–168, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003371.

Allen:2017:CHG

[All17] John F. Allen. The CoRR hypothesis for genes in organelles. Journal of Theoretical Biology, 434(??):50–57, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301571.

Anupama:2019:IIK

[ALM^{+}19] Rani Anupama, Sajitha Lulu, Rout Madhusmita, Sundararajan Vino, Amitava Mukherjee, and Subramanian Babu. Insights into the interaction of key biofilm proteins in Pseudomonas aeruginosa PAO1 with TiO 2 nanoparticle: An in silico analysis. Journal of Theoretical Biology, 462(??):12–25, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305393.

Alvarez:2018:MOL

[Alv18] A. Alvarez. Modeling the odor-landscape resulting from the pumping behavior of bivalve clams in the presence of predators. Journal of Theoretical Biology, 453(??):40–47, September 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302595.

Alvarez:2019:SDS

[Alv19] A. Alvarez. Secondary dispersal of seagrass seeds in complex microtopographies. Journal of Theoretical Biology, 473 (??):28–37, July 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519319301717.

Akita:2010:WDS

[AM10] Tetsuya Akita and Hiroyuki Matsuda. Why do sex ratio dimorphisms exist in Quercus masting? Evolution of imperfect synchronous reproduction in monoecious trees. Journal of Theoretical Biology, 264(2):223–236, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000500.

Assaf:2011:FDA

[AM11] Michael Assaf and Mauro Mobilia. Fixation of a deleterious allele under mutation pressure and finite selection intensity. Journal of Theoretical Biology, 275(1):93–103, April 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000373.

Apostu:2012:MMG

[AM12] Raluca Apostu and Michael C. Mackey. Mathematical model of GAL regulon dynamics in Saccharomyces cerevisiae. Journal of Theoretical Biology, 293(??):219–235, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005273.

Artoisenet:2014:SGT

[AM14] Pierre Artoisenet and Laure-Anne Minsart. Statistical genetics in traditionally cultivated crops. Journal of Theoretical Biology, 360(??):208–221, November 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003713.

Al-Mamun:2013:HCM

[AMBH^{+}13] M. A. Al-Mamun, L. J. Brown, M. A. Hossain, C. Fall, L. Wagstaff, and R. Bass. A hybrid computational model for the effects of maspin on cancer cell dynamics. Journal of Theoretical Biology, 337(??):150–160, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003901.

Arbilly:2010:CEL

[AMFL10] Michał Arbilly, Uzi Motro, Marcus W. Feldman, and Arnon Lotem. Co-evolution of learning complexity and social foraging strategies. Journal of Theoretical Biology, 267(4):573–581, December 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000500X.

Al-Mamun:2016:SMD

[AMFR^{+}16] M. A. Al-Mamun, D. M. Farid, L. Ravenhil, M. A. Hossain, C. Fall, and R. Bass. An in silico model to demonstrate the effects of maspin on cancer cell dynamics. Journal of Theoretical Biology, 388(??):37–49, January 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005007.

Abdel-Mageed:2016:CMC

[AMM16] Samir M. Abdel-Mageed and Ehab I. Mohamed. Comparative modeling of combined transport of water and gradedsize molecules across the glomerular capillary wall. Journal of Theoretical Biology, 394(??):109–116, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000527.

Amorim:2015:MAF

[Amo15] Paulo Amorim. Modeling ant foraging: a chemotaxis approach with pheromones and trail formation. Journal of Theoretical Biology, 385(??):160–173, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004270.

Adams:2012:ARE

[AMP12] M. P. Adams, D. G. Mallet, and G. J. Pettet. Active regulation of the epidermal calcium profile. Journal of Theoretical Biology, 301(??):112–121, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000902.

Azaele:2010:IPE

[AMRRI10] S. Azaele, R. Muneepeerakul, A. Rinaldo, and I. Rodriguez-Iturbe. Inferring plant ecosystem organization from species occurrences. Journal of Theoretical Biology, 262(2):323–329, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004561.

Al-Mamun:2016:MMA

[AMSSG16] Mohammad A. Al-Mamun, Rebecca L. Smith, Ynte H. Schukken, and Yrjö T. Gröhn. Modeling of mycobacterium avium subsp. paratuberculosis dynamics in a dairy herd: An individual based approach. Journal of Theoretical Biology, 408(??):105–117, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302478.

Arazi:2010:MIC

[AN10] A. Arazi and A. U. Neumann. Modeling immune complexmediated autoimmune inflammation. Journal of Theoretical Biology, 267(3):426–436, December 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000456X.

Allen:2012:ESD

[AN12] Benjamin Allen and Martin A. Nowak. Evolutionary shift dynamics on a cycle. Journal of Theoretical Biology, 311(??):28–39, October 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003402.

Arazi:2013:RPF

[AN13] A. Arazi and A. U. Neumann. The role of positive feedback loops involving anti-dsDNA and anti-anti-dsDNA antibodies in autoimmune glomerulonephritis. Journal of Theoretical Biology, 319(??):8–22, February 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200495X.

Allen:2015:GAR

[AN15] Benjamin Allen and Martin A. Nowak. Games among relatives revisited. Journal of Theoretical Biology, 378(??):103–116, August 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002131.

Arthur:2017:EMG

[AN17] R. Arthur and A. Nicholson. An entropic model of Gaia. Journal of Theoretical Biology, 430(??):177–184, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303302.

Anastasio:2011:DDM

[Ana11] Thomas J. Anastasio. Data-driven modeling of Alzheimer Disease pathogenesis. Journal of Theoretical Biology, 290 (??):60–72, December 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519311004413.

Al-Nuaimi:2012:PMM

[ANGPB12] Yusur Al-Nuaimi, Marc Goodfellow, Ralf Paus, and Gerold Baier. A prototypic mathematical model of the human hair cycle. Journal of Theoretical Biology, 310(??):143–159, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200269X.

Alam-Nazki:2010:MMF

[ANK10] Aiman Alam-Nazki and J. Krishnan. A mathematical modelling framework for understanding chemorepulsive signal transduction in Dictyostelium. Journal of Theoretical Biology, 266(1):140–153, September 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002523.

Akhtar:2011:DES

[ANMH11] Noreen Akhtar, Muaz Niazi, Farah Mustafa, and Amir Hussain. A discrete event system specification (DEVS)- based model of consanguinity. Journal of Theoretical Biology, 285(1):103–112, September 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003031.

Anonymous:2010:EBa

[Ano10a] Anonymous. Editorial Board. Journal of Theoretical Biology, 262(1):ifc, January 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005086.

Anonymous:2010:EBb

[Ano10b] Anonymous. Editorial Board. Journal of Theoretical Biology, 262(2):ifc, January 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005384.

Anonymous:2010:EBc

[Ano10c] Anonymous. Editorial Board. Journal of Theoretical Biology, 262(3):ifc, February 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005463.

Anonymous:2010:EBd

[Ano10d] Anonymous. Editorial Board. Journal of Theoretical Biology, 262(4):ifc, February 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005876.

Anonymous:2010:EBe

[Ano10e] Anonymous. Editorial Board. Journal of Theoretical Biology, 263(1):ifc, March 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000214.

Anonymous:2010:EBf

[Ano10f] Anonymous. Editorial Board. Journal of Theoretical Biology, 263(2):ifc, March 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000408.

Anonymous:2010:EBg

[Ano10g] Anonymous. Editorial Board. Journal of Theoretical Biology, 263(3):ifc, April 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000524.

Anonymous:2010:EBh

[Ano10h] Anonymous. Editorial Board. Journal of Theoretical Biology, 263(4):ifc, April 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000652.

Anonymous:2010:EBi

[Ano10i] Anonymous. Editorial Board. Journal of Theoretical Biology, 264(1):ifc, May 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000834.

Anonymous:2010:EBj

[Ano10j] Anonymous. Editorial Board. Journal of Theoretical Biology, 264(2):ifc, May 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001499.

Anonymous:2010:EBk

[Ano10k] Anonymous. Editorial Board. Journal of Theoretical Biology, 264(3):ifc, June 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002079.

Anonymous:2010:EBl

[Ano10l] Anonymous. Editorial Board. Journal of Theoretical Biology, 264(4):ifc, June 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002201.

Anonymous:2010:EBm

[Ano10m] Anonymous. Editorial Board. Journal of Theoretical Biology, 265(1):ifc, July 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002456.

Anonymous:2010:EBn

[Ano10n] Anonymous. Editorial Board. Journal of Theoretical Biology, 265(2):ifc, July 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002602.

Anonymous:2010:EBo

[Ano10o] Anonymous. Editorial Board. Journal of Theoretical Biology, 265(3):ifc, August 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003012.

Anonymous:2010:EBp

[Ano10p] Anonymous. Editorial Board. Journal of Theoretical Biology, 265(4):ifc, August 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003279.

Anonymous:2010:EBq

[Ano10q] Anonymous. Editorial Board. Journal of Theoretical Biology, 266(1):ifc, September 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003516.

Anonymous:2010:EBr

[Ano10r] Anonymous. Editorial Board. Journal of Theoretical Biology, 266(2):ifc, September 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000367X.

Anonymous:2010:EBs

[Ano10s] Anonymous. Editorial Board. Journal of Theoretical Biology, 266(3):ifc, October 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004078.

Anonymous:2010:EBt

[Ano10t] Anonymous. Editorial Board. Journal of Theoretical Biology, 266(4):ifc, October 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000439X.

Anonymous:2010:EBu

[Ano10u] Anonymous. Editorial Board. Journal of Theoretical Biology, 267(1):ifc, November 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004613.

Anonymous:2010:EBv

[Ano10v] Anonymous. Editorial Board. Journal of Theoretical Biology, 267(2):ifc, November 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004819.

Anonymous:2010:EBw

[Ano10w] Anonymous. Editorial Board. Journal of Theoretical Biology, 267(3):ifc, December 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005187.

Anonymous:2010:EBx

[Ano10x] Anonymous. Editorial Board. Journal of Theoretical Biology, 267(4):ifc, December 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005473.

Anonymous:2010:PNa

[Ano10y] Anonymous. Pages 1–128 (7 November 2010). Journal of Theoretical Biology, 267(1):??, November 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PMa

[Ano10z] Anonymous. Pages 1–160 (7 March 2010). Journal of Theoretical Biology, 263(1):??, March 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PMc

[Ano10-27] Anonymous. Pages 1–160 (7 May 2010). Journal of Theoretical Biology, 264(1):??, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PJa

[Ano10-28] Anonymous. Pages 1–196 (7 January 2010). Journal of Theoretical Biology, 262(1):??, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PSa

[Ano10-29] Anonymous. Pages 1–210 (7 September 2010). Journal of Theoretical Biology, 266(1):??, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PJe

[Ano10-30] Anonymous. Pages 1–94 (7 July 2010). Journal of Theoretical Biology, 265(1):??, July 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PJd

[Ano10-31] Anonymous. Pages 1089–1298 (21 June 2010). Journal of Theoretical Biology, 264(4):??, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PNb

[Ano10-32] Anonymous. Pages 129–264 (21 November 2010). Journal of Theoretical Biology, 267(2):??, November 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PMb

[Ano10-33] Anonymous. Pages 161–268 (21 March 2010). Journal of Theoretical Biology, 263(2):??, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PMd

[Ano10-34] Anonymous. Pages 161–644 (21 May 2010). Journal of Theoretical Biology, 264(2):??, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PJb

[Ano10-35] Anonymous. Pages 197–382 (21 January 2010). Journal of Theoretical Biology, 262(2):??, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PSb

[Ano10-36] Anonymous. Pages 211–342 (21 September 2010). Journal of Theoretical Biology, 266(2):??, September 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PAc

[Ano10-37] Anonymous. Pages 218–492 (7 August 2010). Journal of Theoretical Biology, 265(3):??, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PDa

[Ano10-38] Anonymous. Pages 265–470 (7 December 2010). Journal of Theoretical Biology, 267(3):??, December 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PAa

[Ano10-39] Anonymous. Pages 269–392 (7 April 2010). Journal of Theoretical Biology, 263(3):??, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:POa

[Ano10-40] Anonymous. Pages 343–484 (7 October 2010). Journal of Theoretical Biology, 266(3):??, October 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PFa

[Ano10-41] Anonymous. Pages 383–566 (7 February 2010). Journal of Theoretical Biology, 262(3):??, February 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PAb

[Ano10-42] Anonymous. Pages 393–590 (21 April 2010). Journal of Theoretical Biology, 263(4):??, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PDb

[Ano10-43] Anonymous. Pages 471–684 (21 December 2010). Journal of Theoretical Biology, 267(4):??, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:POb

[Ano10-44] Anonymous. Pages 485–742 (21 October 2010). Journal of Theoretical Biology, 266(4):??, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PAd

[Ano10-45] Anonymous. Pages 493–734 (21 August 2010). Journal of Theoretical Biology, 265(4):??, August 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PFb

[Ano10-46] Anonymous. Pages 567–764 (21 February 2010). Journal of Theoretical Biology, 262(4):??, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PJc

[Ano10-47] Anonymous. Pages 645–1088 (7 June 2010). Journal of Theoretical Biology, 264(3):??, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2010:PJf

[Ano10-48] Anonymous. Pages 95–218 (21 July 2010). Journal of Theoretical Biology, 265(2):??, July 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. Journal of Theoretical Biology, 268(1):ifc, January 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005679.

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. Journal of Theoretical Biology, 269(1):ifc, January 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310006120.

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board. Journal of Theoretical Biology, 270(1):ifc, February 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310006491.

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Journal of Theoretical Biology, 271(1):ifc, February 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310006922.

Anonymous:2011:EBe

[Ano11e] Anonymous. Editorial Board. Journal of Theoretical Biology, 272(1):ifc, March 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000129.

Anonymous:2011:EBf

[Ano11f] Anonymous. Editorial Board. Journal of Theoretical Biology, 273(1):ifc, March 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000476.

Anonymous:2011:EBg

[Ano11g] Anonymous. Editorial Board. Journal of Theoretical Biology, 274(1):ifc, April 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000853.

Anonymous:2011:EBh

[Ano11h] Anonymous. Editorial Board. Journal of Theoretical Biology, 275(1):ifc, April 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001111.

Anonymous:2011:EBi

[Ano11i] Anonymous. Editorial Board. Journal of Theoretical Biology, 276(1):ifc, May 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931100141X.

Anonymous:2011:EBj

[Ano11j] Anonymous. Editorial Board. Journal of Theoretical Biology, 277(1):ifc, May 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001585.

Anonymous:2011:EBk

[Ano11k] Anonymous. Editorial Board. Journal of Theoretical Biology, 278(1):ifc, June 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001834.

Anonymous:2011:EBl

[Ano11l] Anonymous. Editorial Board. Journal of Theoretical Biology, 279(1):ifc, June 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002050.

Anonymous:2011:EBm

[Ano11m] Anonymous. Editorial Board. Journal of Theoretical Biology, 280(1):ifc, July 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002414.

Anonymous:2011:EBn

[Ano11n] Anonymous. Editorial Board. Journal of Theoretical Biology, 281(1):ifc, July 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002712.

Anonymous:2011:EBo

[Ano11o] Anonymous. Editorial Board. Journal of Theoretical Biology, 282(1):ifc, August 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002918.

Anonymous:2011:EBp

[Ano11p] Anonymous. Editorial Board. Journal of Theoretical Biology, 283(1):ifc, August 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003274.

Anonymous:2011:EBq

[Ano11q] Anonymous. Editorial Board. Journal of Theoretical Biology, 284(1):ifc, September 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003535.

Anonymous:2011:EBr

[Ano11r] Anonymous. Editorial Board. Journal of Theoretical Biology, 285(1):ifc, September 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003808.

Anonymous:2011:EBs

[Ano11s] Anonymous. Editorial Board. Journal of Theoretical Biology, 286(??):ifc, October 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003948.

Anonymous:2011:EBt

[Ano11t] Anonymous. Editorial Board. Journal of Theoretical Biology, 287(??):ifc, October 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004164.

Anonymous:2011:EBu

[Ano11u] Anonymous. Editorial Board. Journal of Theoretical Biology, 288(??):ifc, November 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004504.

Anonymous:2011:EBv

[Ano11v] Anonymous. Editorial Board. Journal of Theoretical Biology, 289(??):ifc, November 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004784.

Anonymous:2011:EBw

[Ano11w] Anonymous. Editorial Board. Journal of Theoretical Biology, 290(??):ifc, December 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004954.

Anonymous:2011:EBx

[Ano11x] Anonymous. Editorial Board. Journal of Theoretical Biology, 291(??):ifc, December 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005170.

Anonymous:2011:PAc

[Ano11y] Anonymous. Pages 1–100 (7 August 2011). Journal of Theoretical Biology, 282(1):??, August 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PAb

[Ano11z] Anonymous. Pages 1–104 (21 April 2011). Journal of Theoretical Biology, 275(1):??, April 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PDa

[Ano11-27] Anonymous. Pages 1–104 (7 December 2011). Journal of Theoretical Biology, 290(??):1–104, December 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PNa

[Ano11-28] Anonymous. Pages 1–104 (7 November 2011). Journal of Theoretical Biology, 288(??):1–104, November 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:POa

[Ano11-29] Anonymous. Pages 1–114 (7 October 2011). Journal of Theoretical Biology, 286(??):1–114, October 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PMd

[Ano11-30] Anonymous. Pages 1–122 (21 May 2011). Journal of Theoretical Biology, 277(1):??, May 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PDb

[Ano11-31] Anonymous. Pages 1–128 (21 December 2011). Journal of Theoretical Biology, 291(??):1–128, December 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PJc

[Ano11-32] Anonymous. Pages 1–134 (7 June 2011). Journal of Theoretical Biology, 278(1):??, June 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PJf

[Ano11-33] Anonymous. Pages 1–140 (21 July 2011). Journal of Theoretical Biology, 281(1):??, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PSa

[Ano11-34] Anonymous. Pages 1–158 (7 September 2011). Journal of Theoretical Biology, 284(1):??, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PJa

[Ano11-35] Anonymous. Pages 1–160 (7 January 2011). Journal of Theoretical Biology, 268(1):??, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:POb

[Ano11-36] Anonymous. Pages 1–170 (21 October 2011). Journal of Theoretical Biology, 287(??):1–170, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PJd

[Ano11-37] Anonymous. Pages 1–176 (21 June 2011). Journal of Theoretical Biology, 279(1):??, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PSb

[Ano11-38] Anonymous. Pages 1–182 (21 September 2011). Journal of Theoretical Biology, 285(1):??, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PFa

[Ano11-39] Anonymous. Pages 1–196 (7 February 2011). Journal of Theoretical Biology, 270(1):??, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PJe

[Ano11-40] Anonymous. Pages 1–198 (7 July 2011). Journal of Theoretical Biology, 280(1):??, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PMa

[Ano11-41] Anonymous. Pages 1–200 (7 March 2011). Journal of Theoretical Biology, 272(1):??, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PAa

[Ano11-42] Anonymous. Pages 1–204 (7 April 2011). Journal of Theoretical Biology, 274(1):??, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PFb

[Ano11-43] Anonymous. Pages 1–212 (21 February 2011). Journal of Theoretical Biology, 271(1):??, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PAd

[Ano11-44] Anonymous. Pages 1–238 (21 August 2011). Journal of Theoretical Biology, 283(1):??, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PMb

[Ano11-45] Anonymous. Pages 1–248 (21 March 2011). Journal of Theoretical Biology, 273(1):??, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PNb

[Ano11-46] Anonymous. Pages 1–270 (21 November 2011). Journal of Theoretical Biology, 289(??):1–270, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PMc

[Ano11-47] Anonymous. Pages 1–276 (7 May 2011). Journal of Theoretical Biology, 276(1):??, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2011:PJb

[Ano11-48] Anonymous. Pages 1–366 (21 January 2011). Journal of Theoretical Biology, 269(1):??, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Journal of Theoretical Biology, 292(??):ifc, January 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005443.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. Journal of Theoretical Biology, 293(??):ifc, January 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005765.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. Journal of Theoretical Biology, 294(??):ifc, February 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006060.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Journal of Theoretical Biology, 295(??):ifc, February 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006242.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board. Journal of Theoretical Biology, 296(??):ifc, March 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006370.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board. Journal of Theoretical Biology, 297(??):ifc, March 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000185.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board. Journal of Theoretical Biology, 298(??):ifc, April 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000471.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board. Journal of Theoretical Biology, 299(??):ifc, April 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000690.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Journal of Theoretical Biology, 300(??):ifc, May 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000987.

Anonymous:2012:EBj

[Ano12j] Anonymous. Editorial Board. Journal of Theoretical Biology, 301(??):ifc, May 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001361.

Anonymous:2012:EBk

[Ano12k] Anonymous. Editorial Board. Journal of Theoretical Biology, 302(??):ifc, June 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001518.

Anonymous:2012:EBl

[Ano12l] Anonymous. Editorial Board. Journal of Theoretical Biology, 303(??):ifc, June 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001841.

Anonymous:2012:EBm

[Ano12m] Anonymous. Editorial Board. Journal of Theoretical Biology, 304(??):ifc, July 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002172.

Anonymous:2012:EBn

[Ano12n] Anonymous. Editorial Board. Journal of Theoretical Biology, 305(??):ifc, July 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931200238X.

Anonymous:2012:EBo

[Ano12o] Anonymous. Editorial Board. Journal of Theoretical Biology, 306(??):ifc, August 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002494.

Anonymous:2012:EBp

[Ano12p] Anonymous. Editorial Board. Journal of Theoretical Biology, 307(??):ifc, August 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002883.

Anonymous:2012:EBq

[Ano12q] Anonymous. Editorial Board. Journal of Theoretical Biology, 308(??):ifc, September 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003189.

Anonymous:2012:EBr

[Ano12r] Anonymous. Editorial Board. Journal of Theoretical Biology, 309(??):ifc, September 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003505.

Anonymous:2012:EBs

[Ano12s] Anonymous. Editorial Board. Journal of Theoretical Biology, 310(??):ifc, October 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003657.

Anonymous:2012:EBt

[Ano12t] Anonymous. Editorial Board. Journal of Theoretical Biology, 311(??):ifc, October 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004201.

Anonymous:2012:EBu

[Ano12u] Anonymous. Editorial Board. Journal of Theoretical Biology, 312(??):ifc, November 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004353.

Anonymous:2012:EBv

[Ano12v] Anonymous. Editorial Board. Journal of Theoretical Biology, 313(??):ifc, November 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004778.

Anonymous:2012:EBw

[Ano12w] Anonymous. Editorial Board. Journal of Theoretical Biology, 314(??):ifc, December 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931200519X.

Anonymous:2012:EBx

[Ano12x] Anonymous. Editorial Board. Journal of Theoretical Biology, 315(??):ifc, December 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005413.

Anonymous:2012:PMa

[Ano12y] Anonymous. Pages 1–102 (7 March 2012). Journal of Theoretical Biology, 296(??):1–102, March 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PJa

[Ano12z] Anonymous. Pages 1–128 (7 January 2012). Journal of Theoretical Biology, 292(??):1–128, January 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PJf

[Ano12-27] Anonymous. Pages 1–144 (21 July 2012). Journal of Theoretical Biology, 305(??):1–144, July 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PAb

[Ano12-28] Anonymous. Pages 1–146 (7 August 2012). Journal of Theoretical Biology, 306(??):1–146, August 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PSa

[Ano12-29] Anonymous. Pages 1–146 (7 September 2012). Journal of Theoretical Biology, 308(??):1–146, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PDb

[Ano12-30] Anonymous. Pages 1–150 (21 December 2012). Journal of Theoretical Biology, 315(??):1–150, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:POb

[Ano12-31] Anonymous. Pages 1–152 (21 October 2012). Journal of Theoretical Biology, 311(??):1–152, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PJd

[Ano12-32] Anonymous. Pages 1–154 (21 June 2012). Journal of Theoretical Biology, 303(??):1–154, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PAa

[Ano12-33] Anonymous. Pages 1–156 (7 April 2012). Journal of Theoretical Biology, 298(??):1–156, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PNa

[Ano12-34] Anonymous. Pages 1–158 (7 November 2012). Journal of Theoretical Biology, 312(??):1–158, November 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PMd

[Ano12-35] Anonymous. Pages 1–174 (21 May 2012). Journal of Theoretical Biology, 301(??):1–174, May 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PFa

[Ano12-36] Anonymous. Pages 1–186 (7 February 2012). Journal of Theoretical Biology, 294(??):1–186, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PMb

[Ano12-37] Anonymous. Pages 1–188 (21 March 2012). Journal of Theoretical Biology, 297(??):1–188, March 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PFb

[Ano12-38] Anonymous. Pages 1–204 (21 February 2012). Journal of Theoretical Biology, 295(??):1–204, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PSb

[Ano12-39] Anonymous. Pages 1–204 (21 September 2012). Journal of Theoretical Biology, 309(??):1–204, September 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PAc

[Ano12-40] Anonymous. Pages 1–212 (21 August 2012). Journal of Theoretical Biology, 307(??):1–212, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PNb

[Ano12-41] Anonymous. Pages 1–218 (21 November 2012). Journal of Theoretical Biology, 313(??):1–218, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PDa

[Ano12-42] Anonymous. Pages 1–218 (7 December 2012). Journal of Theoretical Biology, 314(??):1–218, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PJb

[Ano12-43] Anonymous. Pages 1–236 (21 January 2012). Journal of Theoretical Biology, 293(??):1–236, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:POa

[Ano12-44] Anonymous. Pages 1–256 (7 October 2012). Journal of Theoretical Biology, 310(??):1–256, October 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PJe

[Ano12-45] Anonymous. Pages 1–308 (7 July 2012). Journal of Theoretical Biology, 304(??):1–308, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PMc

[Ano12-46] Anonymous. Pages 1–376 (7 May 2012). Journal of Theoretical Biology, 300(??):1–376, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2012:PJc

[Ano12-47] Anonymous. Pages 1–96 (7 June 2012). Journal of Theoretical Biology, 302(??):1–96, June 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Journal of Theoretical Biology, 316(??):ifc, January 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005541.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board. Journal of Theoretical Biology, 317(??):ifc, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005991.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board. Journal of Theoretical Biology, 318(??):ifc, February 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312006443.

Anonymous:2013:EBd

[Ano13d] Anonymous. Editorial Board. Journal of Theoretical Biology, 319(??):ifc, February 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000118.

Anonymous:2013:EBe

[Ano13e] Anonymous. Editorial Board. Journal of Theoretical Biology, 320(??):ifc, March 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000301.

Anonymous:2013:EBf

[Ano13f] Anonymous. Editorial Board. Journal of Theoretical Biology, 321(??):ifc, March 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000544.

Anonymous:2013:EBg

[Ano13g] Anonymous. Editorial Board. Journal of Theoretical Biology, 322(??):ifc, April 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000672.

Anonymous:2013:EBh

[Ano13h] Anonymous. Editorial Board. Journal of Theoretical Biology, 323(??):ifc, April 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000854.

Anonymous:2013:EBi

[Ano13i] Anonymous. Editorial Board. Journal of Theoretical Biology, 324(??):ifc, May 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300101X.

Anonymous:2013:EBj

[Ano13j] Anonymous. Editorial Board. Journal of Theoretical Biology, 325(??):ifc, May 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001264.

Anonymous:2013:EBk

[Ano13k] Anonymous. Editorial Board. Journal of Theoretical Biology, 326(??):ifc, June 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300132X.

Anonymous:2013:EBl

[Ano13l] Anonymous. Editorial Board. Journal of Theoretical Biology, 327(??):ifc, June 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001392.

Anonymous:2013:EBm

[Ano13m] Anonymous. Editorial Board. Journal of Theoretical Biology, 328(??):ifc, July 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001732.

Anonymous:2013:EBn

[Ano13n] Anonymous. Editorial Board. Journal of Theoretical Biology, 329(??):ifc, July 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002063.

Anonymous:2013:EBo

[Ano13o] Anonymous. Editorial Board. Journal of Theoretical Biology, 330(??):ifc, August 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002282.

Anonymous:2013:EBp

[Ano13p] Anonymous. Editorial Board. Journal of Theoretical Biology, 331(??):ifc, August 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002361.

Anonymous:2013:EBq

[Ano13q] Anonymous. Editorial Board. Journal of Theoretical Biology, 332(??):ifc, September 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002683.

Anonymous:2013:EBr

[Ano13r] Anonymous. Editorial Board. Journal of Theoretical Biology, 333(??):ifc, September 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313003068.

Anonymous:2013:EBs

[Ano13s] Anonymous. Editorial Board. Journal of Theoretical Biology, 334(??):ifc, October 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313003524.

Anonymous:2013:EBt

[Ano13t] Anonymous. Editorial Board. Journal of Theoretical Biology, 335(??):ifc, October 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313003627.

Anonymous:2013:EBu

[Ano13u] Anonymous. Editorial Board. Journal of Theoretical Biology, 336(??):ifc, November 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313004104.

Anonymous:2013:EBv

[Ano13v] Anonymous. Editorial Board. Journal of Theoretical Biology, 337(??):ifc, November 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313004608.

Anonymous:2013:EBw

[Ano13w] Anonymous. Editorial Board. Journal of Theoretical Biology, 338(??):ifc, December 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313004840.

Anonymous:2013:EBx

[Ano13x] Anonymous. Editorial Board. Journal of Theoretical Biology, 339(??):ifc, December 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313004682.

Anonymous:2013:PJf

[Ano13y] Anonymous. Pages 1–100 (21 July 2013). Journal of Theoretical Biology, 329(??):1–100, July 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PMb

[Ano13z] Anonymous. Pages 1–100 (21 March 2013). Journal of Theoretical Biology, 321(??):1–100, March 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PMd

[Ano13-27] Anonymous. Pages 1–102 (21 May 2013). Journal of Theoretical Biology, 325(??):1–102, May 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PJe

[Ano13-28] Anonymous. Pages 1–102 (7 July 2013). Journal of Theoretical Biology, 328(??):1–102, July 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PJa

[Ano13-29] Anonymous. Pages 1–106 (7 January 2013). Journal of Theoretical Biology, 316(??):1–106, January 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PMc

[Ano13-30] Anonymous. Pages 1–108 (7 May 2013). Journal of Theoretical Biology, 324(??):1–108, May 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PD

[Ano13-31] Anonymous. Pages 1–112 (7 December 2013). Journal of Theoretical Biology, 338(??):1–112, December 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PFb

[Ano13-32] Anonymous. Pages 1–148 (21 February 2013). Journal of Theoretical Biology, 319(??):1–148, February 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PMa

[Ano13-33] Anonymous. Pages 1–170 (7 March 2013). Journal of Theoretical Biology, 320(??):1–170, March 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PNb

[Ano13-34] Anonymous. Pages 1–190 (21 November 2013). Journal of Theoretical Biology, 337(??):1–190, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:POa

[Ano13-35] Anonymous. Pages 1–204 (7 October 2013). Journal of Theoretical Biology, 334(??):1–204, October 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PSb

[Ano13-36] Anonymous. Pages 1–210 (21 September 2013). Journal of Theoretical Biology, 333(??):1–210, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PFa

[Ano13-37] Anonymous. Pages 1–226 (7 February 2013). Journal of Theoretical Biology, 318(??):1–226, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PNa

[Ano13-38] Anonymous. Pages 1–258 (7 November 2013). Journal of Theoretical Biology, 336(??):1–258, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PSa

[Ano13-39] Anonymous. Pages 1–290 (7 September 2013). Journal of Theoretical Biology, 332(??):1–290, September 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:POb

[Ano13-40] Anonymous. Pages 1–298 (21 October 2013). Journal of Theoretical Biology, 335(??):1–298, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PJb

[Ano13-41] Anonymous. Pages 1–430 (21 January 2013). Journal of Theoretical Biology, 317(??):1–430, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PJc

[Ano13-42] Anonymous. Pages 1–82 (7 June 2013). Journal of Theoretical Biology, 326(??):1–82, June 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PAc

[Ano13-43] Anonymous. Pages 1–88 (7 August 2013). Journal of Theoretical Biology, 330(??):1–88, August 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PAb

[Ano13-44] Anonymous. Pages 1–90 (21 April 2013). Journal of Theoretical Biology, 323(??):1–90, April 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PAd

[Ano13-45] Anonymous. Pages 1–90 (21 August 2013). Journal of Theoretical Biology, 331(??):1–90, August 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PJd

[Ano13-46] Anonymous. Pages 1–90 (21 June 2013). Journal of Theoretical Biology, 327(??):1–90, June 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2013:PAa

[Ano13-47] Anonymous. Pages 1–94 (7 April 2013). Journal of Theoretical Biology, 322(??):1–94, April 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. Journal of Theoretical Biology, 340(??):ifc, January 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005031.

Anonymous:2014:EBb

[Ano14b] Anonymous. Editorial Board. Journal of Theoretical Biology, 341(??):ifc, January 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005213.

Anonymous:2014:EBc

[Ano14c] Anonymous. Editorial Board. Journal of Theoretical Biology, 342(??):ifc, February 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000058.

Anonymous:2014:EBd

[Ano14d] Anonymous. Editorial Board. Journal of Theoretical Biology, 343(??):ifc, February 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005766.

Anonymous:2014:EBe

[Ano14e] Anonymous. Editorial Board. Journal of Theoretical Biology, 344(??):ifc, March 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005845.

Anonymous:2014:EBf

[Ano14f] Anonymous. Editorial Board. Journal of Theoretical Biology, 345(??):ifc, March 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000265.

Anonymous:2014:EBg

[Ano14g] Anonymous. Editorial Board. Journal of Theoretical Biology, 346(??):ifc, April 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000691.

Anonymous:2014:EBh

[Ano14h] Anonymous. Editorial Board. Journal of Theoretical Biology, 347(??):ifc, April 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000770.

Anonymous:2014:EBi

[Ano14i] Anonymous. Editorial Board. Journal of Theoretical Biology, 348(??):ifc, May 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001131.

Anonymous:2014:EBj

[Ano14j] Anonymous. Editorial Board. Journal of Theoretical Biology, 349(??):ifc, May 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001271.

Anonymous:2014:EBk

[Ano14k] Anonymous. Editorial Board. Journal of Theoretical Biology, 350(??):ifc, June 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001726.

Anonymous:2014:EBl

[Ano14l] Anonymous. Editorial Board. Journal of Theoretical Biology, 351(??):ifc, June 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001994.

Anonymous:2014:EBm

[Ano14m] Anonymous. Editorial Board. Journal of Theoretical Biology, 352(??):ifc, July 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002057.

Anonymous:2014:EBn

[Ano14n] Anonymous. Editorial Board. Journal of Theoretical Biology, 353(??):ifc, July 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002240.

Anonymous:2014:EBo

[Ano14o] Anonymous. Editorial Board. Journal of Theoretical Biology, 354(??):ifc, August 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002495.

Anonymous:2014:EBp

[Ano14p] Anonymous. Editorial Board. Journal of Theoretical Biology, 355(??):ifc, August 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003087.

Anonymous:2014:EBq

[Ano14q] Anonymous. Editorial Board. Journal of Theoretical Biology, 356(??):ifc, September 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003269.

Anonymous:2014:EBr

[Ano14r] Anonymous. Editorial Board. Journal of Theoretical Biology, 357(??):ifc, September 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003488.

Anonymous:2014:EBs

[Ano14s] Anonymous. Editorial Board. Journal of Theoretical Biology, 358(??):ifc, October 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003853.

Anonymous:2014:EBt

[Ano14t] Anonymous. Editorial Board. Journal of Theoretical Biology, 359(??):ifc, October 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004251.

Anonymous:2014:EBu

[Ano14u] Anonymous. Editorial Board. Journal of Theoretical Biology, 360(??):ifc, November 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005220.

Anonymous:2014:EBv

[Ano14v] Anonymous. Editorial Board. Journal of Theoretical Biology, 361(??):ifc, November 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931400513X.

Anonymous:2014:EBw

[Ano14w] Anonymous. Editorial Board. Journal of Theoretical Biology, 362(??):ifc, December 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004640.

Anonymous:2014:EBx

[Ano14x] Anonymous. Editorial Board. Journal of Theoretical Biology, 363(??):ifc, December 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006389.

Anonymous:2014:PFa

[Ano14y] Anonymous. Pages 1–106 (7 February 2014). Journal of Theoretical Biology, 342(??):1–106, February 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PAa

[Ano14z] Anonymous. Pages 1–108 (7 April 2014). Journal of Theoretical Biology, 346(??):1–108, April 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PJe

[Ano14-27] Anonymous. Pages 1–108 (7 July 2014). Journal of Theoretical Biology, 352(??):1–108, July 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PMb

[Ano14-28] Anonymous. Pages 1–110 (21 March 2014). Journal of Theoretical Biology, 345(??):1–110, March 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PJc

[Ano14-29] Anonymous. Pages 1–112 (7 June 2014). Journal of Theoretical Biology, 350(??):1–112, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PMc

[Ano14-30] Anonymous. Pages 1–122 (7 May 2014). Journal of Theoretical Biology, 348(??):1–122, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PJb

[Ano14-31] Anonymous. Pages 1–142 (21 January 2014). Journal of Theoretical Biology, 341(??):1–142, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PAc

[Ano14-32] Anonymous. Pages 1–146 (7 August 2014). Journal of Theoretical Biology, 354(??):1–146, August 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PMd

[Ano14-33] Anonymous. Pages 1–172 (21 May 2014). Journal of Theoretical Biology, 349(??):1–172, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PJf

[Ano14-34] Anonymous. Pages 1–178 (21 July 2014). Journal of Theoretical Biology, 353(??):1–178, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PAb

[Ano14-35] Anonymous. Pages 1–190 (21 April 2014). Journal of Theoretical Biology, 347(??):1–190, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PFb

[Ano14-36] Anonymous. Pages 1–210 (21 February 2014). Journal of Theoretical Biology, 343(??):1–210, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PSa

[Ano14-37] Anonymous. Pages 1–222 (7 September 2014). Journal of Theoretical Biology, 356(??):1–222, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:POb

[Ano14-38] Anonymous. Pages 1–238 (21 October 2014). Journal of Theoretical Biology, 359(??):1–238, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PAd

[Ano14-39] Anonymous. Pages 1–242 (21 August 2014). Journal of Theoretical Biology, 355(??):1–242, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:POa

[Ano14-40] Anonymous. Pages 1–246 (7 October 2014). Journal of Theoretical Biology, 358(??):1–246, October 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PSb

[Ano14-41] Anonymous. Pages 1–272 (21 September 2014). Journal of Theoretical Biology, 357(??):1–272, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PJa

[Ano14-42] Anonymous. Pages 1–294 (7 January 2014). Journal of Theoretical Biology, 340(??):1–294, January 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PN

[Ano14-43] Anonymous. Pages 1–318 (7 November 2014). Journal of Theoretical Biology, 360(??):1–318, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PD

[Ano14-44] Anonymous. Pages 1–454 (21 December 2014). Journal of Theoretical Biology, 363(??):1–454, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PMa

[Ano14-45] Anonymous. Pages 1–88 (7 March 2014). Journal of Theoretical Biology, 344(??):1–88, March 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PJd

[Ano14-46] Anonymous. Pages 1–96 (21 June 2014). Journal of Theoretical Biology, 351(??):1–96, June 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2014:PRR

[Ano14-47] Anonymous. Pages R1—R2, 1–206 (21 November 2014). Journal of Theoretical Biology, 361(??):R1–, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board. Journal of Theoretical Biology, 364(??):ifc, January 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006572.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. Journal of Theoretical Biology, 365(??):ifc, January 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931400681X.

Anonymous:2015:EBc

[Ano15c] Anonymous. Editorial Board. Journal of Theoretical Biology, 366(??):ifc, February 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006985.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial Board. Journal of Theoretical Biology, 367(??):ifc, February 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000107.

Anonymous:2015:EBe

[Ano15e] Anonymous. Editorial Board. Journal of Theoretical Biology, 368(??):ifc, March 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000429.

Anonymous:2015:EBf

[Ano15f] Anonymous. Editorial Board. Journal of Theoretical Biology, 369(??):ifc, March 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000818.

Anonymous:2015:EBg

[Ano15g] Anonymous. Editorial Board. Journal of Theoretical Biology, 370(??):ifc, April 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000879.

Anonymous:2015:EBh

[Ano15h] Anonymous. Editorial Board. Journal of Theoretical Biology, 371(??):ifc, April 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001162.

Anonymous:2015:EBi

[Ano15i] Anonymous. Editorial Board. Journal of Theoretical Biology, 372(??):ifc, May 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001332.

Anonymous:2015:EBj

[Ano15j] Anonymous. Editorial Board. Journal of Theoretical Biology, 373(??):ifc, May 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001617.

Anonymous:2015:EBk

[Ano15k] Anonymous. Editorial Board. Journal of Theoretical Biology, 374(??):ifc, June 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001885.

Anonymous:2015:EBl

[Ano15l] Anonymous. Editorial Board. Journal of Theoretical Biology, 375(??):ifc, June 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001824.

Anonymous:2015:EBm

[Ano15m] Anonymous. Editorial Board. Journal of Theoretical Biology, 376(??):ifc, July 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002052.

Anonymous:2015:EBn

[Ano15n] Anonymous. Editorial Board. Journal of Theoretical Biology, 377(??):ifc, July 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002258.

Anonymous:2015:EBo

[Ano15o] Anonymous. Editorial Board. Journal of Theoretical Biology, 378(??):ifc, August 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002519.

Anonymous:2015:EBp

[Ano15p] Anonymous. Editorial Board. Journal of Theoretical Biology, 379(??):ifc, August 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002738.

Anonymous:2015:EBq

[Ano15q] Anonymous. Editorial Board. Journal of Theoretical Biology, 380(??):ifc, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003422.

Anonymous:2015:EBr

[Ano15r] Anonymous. Editorial Board. Journal of Theoretical Biology, 381(??):ifc, September 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003604.

Anonymous:2015:EBs

[Ano15s] Anonymous. Editorial Board. Journal of Theoretical Biology, 382(??):ifc, October 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003847.

Anonymous:2015:EBt

[Ano15t] Anonymous. Editorial Board. Journal of Theoretical Biology, 383(??):ifc, October 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004166.

Anonymous:2015:EBu

[Ano15u] Anonymous. Editorial Board. Journal of Theoretical Biology, 384(??):ifc, November 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004440.

Anonymous:2015:EBv

[Ano15v] Anonymous. Editorial Board. Journal of Theoretical Biology, 385(??):ifc, November 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004701.

Anonymous:2015:EBw

[Ano15w] Anonymous. Editorial Board. Journal of Theoretical Biology, 386(??):ifc, December 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004919.

Anonymous:2015:EBx

[Ano15x] Anonymous. Editorial Board. Journal of Theoretical Biology, 387(??):ifc, December 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005093.

Anonymous:2015:PMb

[Ano15y] Anonymous. Pages 1–110 (21 March 2015). Journal of Theoretical Biology, 369(??):1–110, March 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PAc

[Ano15z] Anonymous. Pages 1–116 (7 August 2015). Journal of Theoretical Biology, 378(??):1–116, August 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PJe

[Ano15-27] Anonymous. Pages 1–118 (21 July 2015). Journal of Theoretical Biology, 377(??):1–118, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PFa

[Ano15-28] Anonymous. Pages 1–130 (7 February 2015). Journal of Theoretical Biology, 366(??):1–130, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PMd

[Ano15-29] Anonymous. Pages 1–132 (21 May 2015). Journal of Theoretical Biology, 373(??):1–132, May 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PJd

[Ano15-30] Anonymous. Pages 1–134 (7 July 2015). Journal of Theoretical Biology, 376(??):1–134, July 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PMa

[Ano15-31] Anonymous. Pages 1–144 (7 March 2015). Journal of Theoretical Biology, 368(??):1–144, March 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PNa

[Ano15-32] Anonymous. Pages 1–152 (7 November 2015). Journal of Theoretical Biology, 384(??):1–152, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:POb

[Ano15-33] Anonymous. Pages 1–156 (21 October 2015). Journal of Theoretical Biology, 383(??):1–156, October 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PAb

[Ano15-34] Anonymous. Pages 1–166 (21 April 2015). Journal of Theoretical Biology, 371(??):1–166, April 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PNb

[Ano15-35] Anonymous. Pages 1–174 (21 November 2015). Journal of Theoretical Biology, 385(??):1–174, November 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PJc

[Ano15-36] Anonymous. Pages 1–182 (7 June 2015). Journal of Theoretical Biology, 374(??):1–182, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PDa

[Ano15-37] Anonymous. Pages 1–190 (7 December 2015). Journal of Theoretical Biology, 386(??):1–190, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PAa

[Ano15-38] Anonymous. Pages 1–210 (7 April 2015). Journal of Theoretical Biology, 370(??):1–210, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PMc

[Ano15-39] Anonymous. Pages 1–222 (7 May 2015). Journal of Theoretical Biology, 372(??):1–222, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PDb

[Ano15-40] Anonymous. Pages 1–262 (21 December 2015). Journal of Theoretical Biology, 387(??):1–262, December 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PFb

[Ano15-41] Anonymous. Pages 1–296 (21 February 2015). Journal of Theoretical Biology, 367(??):1–296, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:POa

[Ano15-42] Anonymous. Pages 1–404 (7 October 2015). Journal of Theoretical Biology, 382(??):1–404, October 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PJa

[Ano15-43] Anonymous. Pages 1–438 (7 January 2015). Journal of Theoretical Biology, 364(??):1–438, January 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PJb

[Ano15-44] Anonymous. Pages 1–496 (21 January 2015). Journal of Theoretical Biology, 365(??):1–496, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PS

[Ano15-45] Anonymous. Pages 1–590 (7 September 2015). Journal of Theoretical Biology, 380(??):1–590, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2015:PAd

[Ano15-46] Anonymous. Pages 1–98 (21 August 2015). Journal of Theoretical Biology, 379(??):1–98, August 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial Board. Journal of Theoretical Biology, 388(??):ifc, January 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005421.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial Board. Journal of Theoretical Biology, 389(??):ifc, January 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005536.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. Journal of Theoretical Biology, 390(??):ifc, February 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005822.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial Board. Journal of Theoretical Biology, 391(??):ifc, February 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000060.

Anonymous:2016:EBe

[Ano16e] Anonymous. Editorial Board. Journal of Theoretical Biology, 392(??):ifc, March 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000618.

Anonymous:2016:EBf

[Ano16f] Anonymous. Editorial Board. Journal of Theoretical Biology, 393(??):ifc, March 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000862.

Anonymous:2016:EBg

[Ano16g] Anonymous. Editorial Board. Journal of Theoretical Biology, 394(??):ifc, April 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001077.

Anonymous:2016:EBh

[Ano16h] Anonymous. Editorial Board. Journal of Theoretical Biology, 395(??):ifc, April 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001326.

Anonymous:2016:EBi

[Ano16i] Anonymous. Editorial Board. Journal of Theoretical Biology, 396(??):ifc, May 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001685.

Anonymous:2016:EBj

[Ano16j] Anonymous. Editorial Board. Journal of Theoretical Biology, 397(??):ifc, May 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300042.

Anonymous:2016:EBk

[Ano16k] Anonymous. Editorial Board. Journal of Theoretical Biology, 398(??):ifc, June 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300339.

Anonymous:2016:EBl

[Ano16l] Anonymous. Editorial Board. Journal of Theoretical Biology, 399(??):ifc, June 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300625.

Anonymous:2016:EBm

[Ano16m] Anonymous. Editorial Board. Journal of Theoretical Biology, 400(??):ifc, July 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300807.

Anonymous:2016:EBn

[Ano16n] Anonymous. Editorial Board. Journal of Theoretical Biology, 401(??):ifc, July 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300996.

Anonymous:2016:EBo

[Ano16o] Anonymous. Editorial Board. Journal of Theoretical Biology, 402(??):ifc, August 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301096.

Anonymous:2016:EBp

[Ano16p] Anonymous. Editorial Board. Journal of Theoretical Biology, 403(??):ifc, August 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301345.

Anonymous:2016:EBq

[Ano16q] Anonymous. Editorial Board. Journal of Theoretical Biology, 404(??):ifc, September 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931630176X.

Anonymous:2016:EBr

[Ano16r] Anonymous. Editorial Board. Journal of Theoretical Biology, 405(??):ifc, September 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301862.

Anonymous:2016:EBs

[Ano16s] Anonymous. Editorial Board. Journal of Theoretical Biology, 406(??):ifc, October 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302296.

Anonymous:2016:EBt

[Ano16t] Anonymous. Editorial Board. Journal of Theoretical Biology, 407(??):ifc, October 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302636.

Anonymous:2016:EBu

[Ano16u] Anonymous. Editorial Board. Journal of Theoretical Biology, 408(??):ifc, November 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302892.

Anonymous:2016:EBv

[Ano16v] Anonymous. Editorial Board. Journal of Theoretical Biology, 409(??):ifc, November 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303071.

Anonymous:2016:EBw

[Ano16w] Anonymous. Editorial Board. Journal of Theoretical Biology, 410(??):ifc, December 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303265.

Anonymous:2016:EBx

[Ano16x] Anonymous. Editorial Board. Journal of Theoretical Biology, 411(??):ifc, December 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303411.

Anonymous:2016:PDb

[Ano16y] Anonymous. Pages 1–102 (21 December 2016). Journal of Theoretical Biology, 411(??):1–102, December 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PFb

[Ano16z] Anonymous. Pages 1–112 (21 February 2016). Journal of Theoretical Biology, 391(??):1–112, February 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PJf

[Ano16-27] Anonymous. Pages 1–122 (21 July 2016). Journal of Theoretical Biology, 401(??):1–122, July 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PMa

[Ano16-28] Anonymous. Pages 1–126 (7 March 2016). Journal of Theoretical Biology, 392(??):1–126, March 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PJa

[Ano16-29] Anonymous. Pages 1–130 (7 January 2016). Journal of Theoretical Biology, 388(??):1–130, January 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PJe

[Ano16-30] Anonymous. Pages 1–158 (7 July 2016). Journal of Theoretical Biology, 400(??):1–158, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PFa

[Ano16-31] Anonymous. Pages 1–164 (7 February 2016). Journal of Theoretical Biology, 390(??):1–164, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PAc

[Ano16-32] Anonymous. Pages 1–170 (7 August 2016). Journal of Theoretical Biology, 402(??):1–170, August 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PNb

[Ano16-33] Anonymous. Pages 1–172 (21 November 2016). Journal of Theoretical Biology, 409(??):1–172, November 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PJd

[Ano16-34] Anonymous. Pages 1–174 (21 June 2016). Journal of Theoretical Biology, 399(??):1–174, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PDa

[Ano16-35] Anonymous. Pages 1–180 (7 December 2016). Journal of Theoretical Biology, 410(??):1–180, December 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PJc

[Ano16-36] Anonymous. Pages 1–180 (7 June 2016). Journal of Theoretical Biology, 398(??):1–180, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:POa

[Ano16-37] Anonymous. Pages 1–192 (7 October 2016). Journal of Theoretical Biology, 406(??):1–192, October 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PMd

[Ano16-38] Anonymous. Pages 1–206 (21 May 2016). Journal of Theoretical Biology, 397(??):1–206, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PMc

[Ano16-39] Anonymous. Pages 1–210 (7 May 2016). Journal of Theoretical Biology, 396(??):1–210, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PAd

[Ano16-40] Anonymous. Pages 1–218 (21 August 2016). Journal of Theoretical Biology, 403(??):1–218, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PMb

[Ano16-41] Anonymous. Pages 1–228 (21 March 2016). Journal of Theoretical Biology, 393(??):1–228, March 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PAb

[Ano16-42] Anonymous. Pages 1–246 (21 April 2016). Journal of Theoretical Biology, 395(??):1–246, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PAa

[Ano16-43] Anonymous. Pages 1–254 (7 April 2016). Journal of Theoretical Biology, 394(??):1–254, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PNa

[Ano16-44] Anonymous. Pages 1–290 (7 November 2016). Journal of Theoretical Biology, 408(??):1–290, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PJb

[Ano16-45] Anonymous. Pages 1–310 (21 January 2016). Journal of Theoretical Biology, 389(??):1–310, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:PS

[Ano16-46] Anonymous. Pages 1–398 (7 September 2016). Journal of Theoretical Biology, 404(??):1–398, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2016:POb

[Ano16-47] Anonymous. Pages 1–428 (21 October 2016). Journal of Theoretical Biology, 407(??):1–428, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board. Journal of Theoretical Biology, 412(??):ifc, January 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303691.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board. Journal of Theoretical Biology, 413(??):ifc, January 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303976.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial Board. Journal of Theoretical Biology, 414(??):ifc, February 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931730084X.

Anonymous:2017:EBd

[Ano17d] Anonymous. Editorial Board. Journal of Theoretical Biology, 415(??):ifc, February 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316304398.

Anonymous:2017:EBe

[Ano17e] Anonymous. Editorial Board. Journal of Theoretical Biology, 416(??):ifc, March 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300413.

Anonymous:2017:EBf

[Ano17f] Anonymous. Editorial Board. Journal of Theoretical Biology, 417(??):ifc, March 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301017.

Anonymous:2017:EBg

[Ano17g] Anonymous. Editorial Board. Journal of Theoretical Biology, 418(??):ifc, April 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301078.

Anonymous:2017:EBh

[Ano17h] Anonymous. Editorial Board. Journal of Theoretical Biology, 419(??):ifc, April 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301297.

Anonymous:2017:EBi

[Ano17i] Anonymous. Editorial Board. Journal of Theoretical Biology, 420(??):ifc, May 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301595.

Anonymous:2017:EBj

[Ano17j] Anonymous. Editorial Board. Journal of Theoretical Biology, 421(??):ifc, May 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301741.

Anonymous:2017:EBk

[Ano17k] Anonymous. Editorial Board. Journal of Theoretical Biology, 422(??):ifc, June 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301893.

Anonymous:2017:EBl

[Ano17l] Anonymous. Editorial Board. Journal of Theoretical Biology, 423(??):ifc, June 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302096.

Anonymous:2017:EBm

[Ano17m] Anonymous. Editorial Board. Journal of Theoretical Biology, 424(??):ifc, July 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302321.

Anonymous:2017:EBn

[Ano17n] Anonymous. Editorial Board. Journal of Theoretical Biology, 425(??):ifc, July 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302400.

Anonymous:2017:EBo

[Ano17o] Anonymous. Editorial Board. Journal of Theoretical Biology, 426(??):ifc, August 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302667.

Anonymous:2017:EBp

[Ano17p] Anonymous. Editorial Board. Journal of Theoretical Biology, 427(??):ifc, August 1, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302941.

Anonymous:2017:EBq

[Ano17q] Anonymous. Editorial Board. Journal of Theoretical Biology, 428(??):ifc, September 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303132.

Anonymous:2017:EBr

[Ano17r] Anonymous. Editorial Board. Journal of Theoretical Biology, 429(??):ifc, September 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303351.

Anonymous:2017:EBs

[Ano17s] Anonymous. Editorial Board. Journal of Theoretical Biology, 430(??):ifc, October 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303685.

Anonymous:2017:EBt

[Ano17t] Anonymous. Editorial Board. Journal of Theoretical Biology, 431(??):ifc, October 27, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303843.

Anonymous:2017:EBu

[Ano17u] Anonymous. Editorial Board. Journal of Theoretical Biology, 432(??):ifc, November 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304046.

Anonymous:2017:EBv

[Ano17v] Anonymous. Editorial Board. Journal of Theoretical Biology, 433(??):ifc, November 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304216.

Anonymous:2017:EBw

[Ano17w] Anonymous. Editorial Board. Journal of Theoretical Biology, 434(??):ifc, December 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304502.

Anonymous:2017:EBx

[Ano17x] Anonymous. Editorial Board. Journal of Theoretical Biology, 435(??):ifc, December 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304605.

Anonymous:2017:PNb

[Ano17y] Anonymous. Pages 1–106 (21 November 2017). Journal of Theoretical Biology, 433(??):1–106, November 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PJe

[Ano17z] Anonymous. Pages 1–110 (7 July 2017). Journal of Theoretical Biology, 424(??):1–110, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PJf

[Ano17-27] Anonymous. Pages 1–112 (21 July 2017). Journal of Theoretical Biology, 425(??):1–112, July 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PAa

[Ano17-28] Anonymous. Pages 1–138 (7 April 2017). Journal of Theoretical Biology, 418(??):1–138, April 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PMb

[Ano17-29] Anonymous. Pages 1–142 (21 March 2017). Journal of Theoretical Biology, 417(??):1–142, March 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PNa

[Ano17-30] Anonymous. Pages 1–168 (7 November 2017). Journal of Theoretical Biology, 432(??):1–168, November 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PSa

[Ano17-31] Anonymous. Pages 1–168 (7 September 2017). Journal of Theoretical Biology, 428(??):1–168, September 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PFb

[Ano17-32] Anonymous. Pages 1–170 (21 February 2017). Journal of Theoretical Biology, 415(??):1–170, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PAc

[Ano17-33] Anonymous. Pages 1–180 (7 August 2017). Journal of Theoretical Biology, 426(??):1–180, August 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PMd

[Ano17-34] Anonymous. Pages 1–202 (21 May 2017). Journal of Theoretical Biology, 421(??):1–202, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PJa

[Ano17-35] Anonymous. Pages 1–204 (7 January 2017). Journal of Theoretical Biology, 412(??):1–204, January 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PMa

[Ano17-36] Anonymous. Pages 1–208 (7 March 2017). Journal of Theoretical Biology, 416(??):1–208, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PD

[Ano17-37] Anonymous. Pages 1–264 (21 December 2017). Journal of Theoretical Biology, 435(??):1–264, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PSb

[Ano17-38] Anonymous. Pages 1–266 (21 September 2017). Journal of Theoretical Biology, 429(??):1–266, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PFa

[Ano17-39] Anonymous. Pages 1–268 (7 February 2017). Journal of Theoretical Biology, 414(??):1–268, February 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:POa

[Ano17-40] Anonymous. Pages 1–282 (7 October 2017). Journal of Theoretical Biology, 430(??):1–282, October 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PMc

[Ano17-41] Anonymous. Pages 1–324 (7 May 2017). Journal of Theoretical Biology, 420(??):1–324, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PAb

[Ano17-42] Anonymous. Pages 1–394 (21 April 2017). Journal of Theoretical Biology, 419(??):1–394, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PJd

[Ano17-43] Anonymous. Pages 1–70 (21 June 2017). Journal of Theoretical Biology, 423(??):1–70, June 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PJb

[Ano17-44] Anonymous. Pages 1–86 (21 January 2017). Journal of Theoretical Biology, 413(??):1–86, January 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PAd

[Ano17-45] Anonymous. Pages 1–90 (1 August 2017). Journal of Theoretical Biology, 427(??):1–90, August 1, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:PJc

[Ano17-46] Anonymous. Pages 1–90 (7 June 2017). Journal of Theoretical Biology, 422(??):1–90, June 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2017:POb

[Ano17-47] Anonymous. Pages 1–96 (27 October 2017). Journal of Theoretical Biology, 431(??):1–96, October 27, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Journal of Theoretical Biology, 436(??):ifc, January 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304903.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Journal of Theoretical Biology, 437(??):ifc, January 21, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305106.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Journal of Theoretical Biology, 438(??):ii, February 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305441.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Journal of Theoretical Biology, 439(??):ii, February 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305672.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Journal of Theoretical Biology, 440(??):ii, March 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300109.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. Journal of Theoretical Biology, 441(??):ii, March 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300274.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial Board. Journal of Theoretical Biology, 442(??):ii, April 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830050X.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board. Journal of Theoretical Biology, 443(??):ii, April 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300705.

Anonymous:2018:EBi

[Ano18i] Anonymous. Editorial Board. Journal of Theoretical Biology, 444(??):ii, May 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301036.

Anonymous:2018:EBj

[Ano18j] Anonymous. Editorial Board. Journal of Theoretical Biology, 445(??):ii, May 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301152.

Anonymous:2018:EBk

[Ano18k] Anonymous. Editorial Board. Journal of Theoretical Biology, 446(??):ii, June 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301553.

Anonymous:2018:EBl

[Ano18l] Anonymous. Editorial Board. Journal of Theoretical Biology, 447(??):ii, June 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830184X.

Anonymous:2018:EBm

[Ano18m] Anonymous. Editorial Board. Journal of Theoretical Biology, 448(??):ii, July 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301991.

Anonymous:2018:EBn

[Ano18n] Anonymous. Editorial Board. Journal of Theoretical Biology, 449(??):ii, July 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302145.

Anonymous:2018:EBo

[Ano18o] Anonymous. Editorial Board. Journal of Theoretical Biology, 450(??):ii, August 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302406.

Anonymous:2018:EBp

[Ano18p] Anonymous. Editorial Board. Journal of Theoretical Biology, 451(??):ii, August 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302522.

Anonymous:2018:EBq

[Ano18q] Anonymous. Editorial Board. Journal of Theoretical Biology, 452(??):ii, September 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302698.

Anonymous:2018:EBr

[Ano18r] Anonymous. Editorial Board. Journal of Theoretical Biology, 453(??):ii, September 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830290X.

Anonymous:2018:EBs

[Ano18s] Anonymous. Editorial Board. Journal of Theoretical Biology, 454(??):ii, October 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830362X.

Anonymous:2018:EBt

[Ano18t] Anonymous. Editorial Board. Journal of Theoretical Biology, 455(??):ii, October 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303904.

Anonymous:2018:EBu

[Ano18u] Anonymous. Editorial Board. Journal of Theoretical Biology, 456(??):ii, November 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830420X.

Anonymous:2018:EBv

[Ano18v] Anonymous. Editorial Board. Journal of Theoretical Biology, 457(??):ii, November 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304508.

Anonymous:2018:EBw

[Ano18w] Anonymous. Editorial Board. Journal of Theoretical Biology, 458(??):ii, December 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304934.

Anonymous:2018:EBx

[Ano18x] Anonymous. Editorial Board. Journal of Theoretical Biology, 459(??):ii, December 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305150.

Anonymous:2018:PAc

[Ano18y] Anonymous. Pages 1–104 (7 August 2018). Journal of Theoretical Biology, 450(??):1–104, August 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PAd

[Ano18z] Anonymous. Pages 1–122 (14 August 2018). Journal of Theoretical Biology, 451(??):1–122, August 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PJf

[Ano18-27] Anonymous. Pages 1–124 (14 July 2018). Journal of Theoretical Biology, 449(??):1–124, July 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PMc

[Ano18-28] Anonymous. Pages 1–124 (7 May 2018). Journal of Theoretical Biology, 444(??):1–124, May 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PMa

[Ano18-29] Anonymous. Pages 1–132 (7 March 2018). Journal of Theoretical Biology, 440(??):1–132, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PJa

[Ano18-30] Anonymous. Pages 1–134 (7 January 2018). Journal of Theoretical Biology, 436(??):1–134, January 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PSb

[Ano18-31] Anonymous. Pages 1–156 (14 September 2018). Journal of Theoretical Biology, 453(??):1–156, September 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PJe

[Ano18-32] Anonymous. Pages 1–156 (7 July 2018). Journal of Theoretical Biology, 448(??):1–156, July 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PAa

[Ano18-33] Anonymous. Pages 1–158 (7 April 2018). Journal of Theoretical Biology, 442(??):1–158, April 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PDb

[Ano18-34] Anonymous. Pages 1–170 (14 December 2018). Journal of Theoretical Biology, 459(??):1–170, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PFa

[Ano18-35] Anonymous. Pages 1–174 (7 February 2018). Journal of Theoretical Biology, 438(??):1–174, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PAb

[Ano18-36] Anonymous. Pages 1–176 (14 April 2018). Journal of Theoretical Biology, 443(??):1–176, April 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PMd

[Ano18-37] Anonymous. Pages 1–186 (14 May 2018). Journal of Theoretical Biology, 445(??):1–186, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PJd

[Ano18-38] Anonymous. Pages 1–200 (14 June 2018). Journal of Theoretical Biology, 447(??):1–200, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PDa

[Ano18-39] Anonymous. Pages 1–206 (7 December 2018). Journal of Theoretical Biology, 458(??):1–206, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PFb

[Ano18-40] Anonymous. Pages 1–232 (14 February 2018). Journal of Theoretical Biology, 439(??):1–232, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PJc

[Ano18-41] Anonymous. Pages 1–260 (7 June 2018). Journal of Theoretical Biology, 446(??):1–260, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PNb

[Ano18-42] Anonymous. Pages 1–270 (14 November 2018). Journal of Theoretical Biology, 457(??):1–270, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PNa

[Ano18-43] Anonymous. Pages 1–314 (7 November 2018). Journal of Theoretical Biology, 456(??):1–314, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PJb

[Ano18-44] Anonymous. Pages 1–318 (21 January 2018). Journal of Theoretical Biology, 437(??):1–318, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:POb

[Ano18-45] Anonymous. Pages 1–370 (14 October 2018). Journal of Theoretical Biology, 455(??):1–370, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:POa

[Ano18-46] Anonymous. Pages 1–424 (7 October 2018). Journal of Theoretical Biology, 454(??):1–424, October 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PSa

[Ano18-47] Anonymous. Pages 1–84 (7 September 2018). Journal of Theoretical Biology, 452(??):1–84, September 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2018:PMb

[Ano18-48] Anonymous. Pages 1–92 (14 March 2018). Journal of Theoretical Biology, 441(??):1–92, March 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:Db

[Ano19a] Anonymous. 21 December 2019. Journal of Theoretical Biology, 483(??):??, December 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:Da

[Ano19b] Anonymous. 7 December 2019. Journal of Theoretical Biology, 482(??):??, December 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:EBa

[Ano19c] Anonymous. Editorial Board. Journal of Theoretical Biology, 460(??):ii, January 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305435.

Anonymous:2019:EBb

[Ano19d] Anonymous. Editorial Board. Journal of Theoretical Biology, 461(??):ii, January 14, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305599.

Anonymous:2019:EBc

[Ano19e] Anonymous. Editorial Board. Journal of Theoretical Biology, 462(??):ii, February 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318306118.

Anonymous:2019:EBd

[Ano19f] Anonymous. Editorial Board. Journal of Theoretical Biology, 463(??):ii, February 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318306404.

Anonymous:2019:EBe

[Ano19g] Anonymous. Editorial Board. Journal of Theoretical Biology, 464(??):ii, March 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300244.

Anonymous:2019:EBf

[Ano19h] Anonymous. Editorial Board. Journal of Theoretical Biology, 465(??):ii, March 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300414.

Anonymous:2019:EBg

[Ano19i] Anonymous. Editorial Board. Journal of Theoretical Biology, 466(??):ii, April 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300657.

Anonymous:2019:EBh

[Ano19j] Anonymous. Editorial Board. Journal of Theoretical Biology, 467(??):ii, April 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300797.

Anonymous:2019:EBi

[Ano19k] Anonymous. Editorial Board. Journal of Theoretical Biology, 468(??):ii, May 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931930102X.

Anonymous:2019:EBj

[Ano19l] Anonymous. Editorial Board. Journal of Theoretical Biology, 469(??):ii, May 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301183.

Anonymous:2019:EBk

[Ano19m] Anonymous. Editorial Board. Journal of Theoretical Biology, 470(??):ii, June 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301389.

Anonymous:2019:EBl

[Ano19n] Anonymous. Editorial Board. Journal of Theoretical Biology, 471(??):ii, June 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931930164X.

Anonymous:2019:EBm

[Ano19o] Anonymous. Editorial Board. Journal of Theoretical Biology, 472(??):ii, July 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301808.

Anonymous:2019:EBn

[Ano19p] Anonymous. Editorial Board. Journal of Theoretical Biology, 473(??):ii, July 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301973.

Anonymous:2019:EBo

[Ano19q] Anonymous. Editorial Board. Journal of Theoretical Biology, 474(??):ii, August 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302061.

Anonymous:2019:EBp

[Ano19r] Anonymous. Editorial Board. Journal of Theoretical Biology, 475(??):ii, August 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931930222X.

Anonymous:2019:EBq

[Ano19s] Anonymous. Editorial Board. Journal of Theoretical Biology, 476(??):ii, September 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302450.

Anonymous:2019:EBr

[Ano19t] Anonymous. Editorial Board. Journal of Theoretical Biology, 477(??):ii, September 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302632.

Anonymous:2019:EBs

[Ano19u] Anonymous. Editorial Board. Journal of Theoretical Biology, 478(??):ii, October 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302930.

Anonymous:2019:EBt

[Ano19v] Anonymous. Editorial Board. Journal of Theoretical Biology, 479(??):ii, October 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319303133.

Anonymous:2019:EBu

[Ano19w] Anonymous. Editorial Board. Journal of Theoretical Biology, 480(??):ii, November 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319303352.

Anonymous:2019:EBv

[Ano19x] Anonymous. Editorial Board. Journal of Theoretical Biology, 481(??):ii, November 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319303649.

Anonymous:2019:EBw

[Ano19y] Anonymous. Editorial Board. Journal of Theoretical Biology, 482(??):Article 110010, December 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303807.

Anonymous:2019:EBx

[Ano19z] Anonymous. Editorial Board. Journal of Theoretical Biology, 483(??):Article 110023, December 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303935.

Anonymous:2019:WCE

[Ano19-27] Anonymous. Have we come to the end of theoretical and mathematical biology? Journal of Theoretical Biology, 481(??):1–2, November 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319303571.

Anonymous:2019:O

[Ano19-28] Anonymous. Overview. Journal of Theoretical Biology, 481(??):3–5, November 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319303583.

Anonymous:2019:PAc

[Ano19-29] Anonymous. Pages 1–102 (7 August 2019). Journal of Theoretical Biology, 474(??):1–102, August 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:POb

[Ano19-30] Anonymous. Pages 1–106 (21 October 2019). Journal of Theoretical Biology, 479(??):1–106, October 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PJc

[Ano19-31] Anonymous. Pages 1–120 (7 June 2019). Journal of Theoretical Biology, 470(??):1–120, June 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PJd

[Ano19-32] Anonymous. Pages 1–124 (21 June 2019). Journal of Theoretical Biology, 471(??):1–124, June 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PJe

[Ano19-33] Anonymous. Pages 1–124 (7 July 2019). Journal of Theoretical Biology, 472(??):1–124, July 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PSb

[Ano19-34] Anonymous. Pages 1–126 (21 September 2019). Journal of Theoretical Biology, 477(??):1–126, September 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PMb

[Ano19-35] Anonymous. Pages 1–134 (21 March 2019). Journal of Theoretical Biology, 465(??):1–134, March 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PAa

[Ano19-36] Anonymous. Pages 1–144 (7 April 2019). Journal of Theoretical Biology, 466(??):1–144, April 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PFb

[Ano19-37] Anonymous. Pages 1–166 (21 February 2019). Journal of Theoretical Biology, 463(??):1–166, February 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:POa

[Ano19-38] Anonymous. Pages 1–168 (7 October 2019). Journal of Theoretical Biology, 478(??):1–168, October 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PAb

[Ano19-39] Anonymous. Pages 1–174 (21 April 2019). Journal of Theoretical Biology, 467(??):1–174, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PMa

[Ano19-40] Anonymous. Pages 1–182 (7 March 2019). Journal of Theoretical Biology, 464(??):1–182, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PMd

[Ano19-41] Anonymous. Pages 1–202 (21 May 2019). Journal of Theoretical Biology, 469(??):1–202, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PJa

[Ano19-42] Anonymous. Pages 1–274 (7 January 2019). Journal of Theoretical Biology, 460(??):1–274, January 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PNa

[Ano19-43] Anonymous. Pages 1–284 (7 November 2019). Journal of Theoretical Biology, 480(??):1–284, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PJb

[Ano19-44] Anonymous. Pages 1–290 (14 January 2019). Journal of Theoretical Biology, 461(??):1–290, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PFa

[Ano19-45] Anonymous. Pages 1–582 (7 February 2019). Journal of Theoretical Biology, 462(??):1–582, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PMc

[Ano19-46] Anonymous. Pages 1–92 (7 May 2019). Journal of Theoretical Biology, 468(??):1–92, May 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PJf

[Ano19-47] Anonymous. Pages 1–94 (21 July 2019). Journal of Theoretical Biology, 473(??):1–94, July 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PSa

[Ano19-48] Anonymous. Pages 1-108 (7 September 2019). Journal of Theoretical Biology, 476(??):1–108, September 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PA

[Ano19-49] Anonymous. Pages 1-72 (21 August 2019). Journal of Theoretical Biology, 475(??):1–72, August 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:2019:PPK

[Ano19-50] Anonymous. Professor Philip K. Maini FRS: a very small selection of his research contributions. Journal of Theoretical Biology, 481(??):6–9, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931930356X.

Anonymous:2019:PN

[Ano19-51] Anonymous. Publisher’s Note. Journal of Theoretical Biology, 482(??):Article 109988, December 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303315.

Antrup:2013:CEI

[Ant13] A. H. Antrup. Co-evolution of institutions and preferences: The case of the (human) mating market. Journal of Theoretical Biology, 332(??):9–19, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001628.

Armbruster:2014:TLS

[ANY14] Dieter Armbruster, John Nagy, and Jon Young. Three level signal transduction cascades lead to reliably timed switches. Journal of Theoretical Biology, 361(??):69–80, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004056.

Afenya:2019:MMB

[AOM19] Evans K. Afenya, Rachid Ouifki, and Suneel D. Mundle. Mathematical modeling of bone marrow — peripheral blood dynamics in the disease state based on current emerging paradigms, part II. Journal of Theoretical Biology, 460(??):37–55, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304727.

Alonso:2017:OMW

[AOR17] Diego López Alonso and Isabel M. Ortiz-Rodríguez. Offspring mortality was a determinant factor in the evolution of paternal investment in humans: An evolutionary game approach. Journal of Theoretical Biology, 419(??):44–51, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300528.

Asano:2011:QLM

[AOT^{+}11] Masanari Asano, Masanori Ohya, Yoshiharu Tanaka, Irina Basieva, and Andrei Khrennikov. Quantum-like model of brain’s functioning: Decision making from decoherence. Journal of Theoretical Biology, 281(1):56–64, July 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002220.

Angus:2013:NME

[AP13a] Simon D. Angus and Monika Joanna Piotrowska. A numerical model of EMT6/Ro spheroid dynamics under irradiation: Calibration and estimation of the underlying irradiation-induced cell survival probability. Journal of Theoretical Biology, 320(??):23–32, March 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006248.

Arodz:2013:SSS

[AP13b] Tomasz Arod’z and Przemyslaw M. Plonka. Sequence and structure space model of protein divergence driven by point mutations. Journal of Theoretical Biology, 330(??):1–8, August 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001197.

Allen:2017:MMA

[AP17] Henry R. Allen and Mariya Ptashnyk. Mathematical modelling and analysis of the brassinosteroid and gibberellin signalling pathways and their interactions. Journal of Theoretical Biology, 432(??):109–131, November 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730379X.

Alharbi:2019:ECL

[AP19] Weam Alharbi and Sergei Petrovskii. Effect of complex landscape geometry on the invasive species spread: Invasion with stepping stones. Journal of Theoretical Biology, 464(??):85– 97, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318306143.

Addison:2015:IMI

[APBS15] Audrey Addison, James A. Powell, Barbara J. Bentz, and Diana L. Six. Integrating models to investigate critical phenological overlaps in complex ecological interactions: The mountain pine beetle-fungus symbiosis. Journal of Theoretical Biology, 368(??):55–66, March 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314007085.

Addison:2013:RTV

[APS^{+}13] A. L. Addison, J. A. Powell, D. L. Six, M. Moore, and B. J. Bentz. The role of temperature variability in stabilizing the mountain pine beetle-fungus mutualism. Journal of Theoretical Biology, 335(??):40–50, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002841.

Anjukandi:2010:LDS

[APW10] Padmesh Anjukandi, Gerald G. Pereira, and Martin A. K. Williams. Langevin dynamics simulations reveal biologically relevant folds arising from the incorporation of a torsional potential. Journal of Theoretical Biology, 265(3):245–249, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002407.

Abrams:2011:HDA

[AR11] Peter A. Abrams and Lasse Ruokolainen. How does adaptive consumer movement affect population dynamics in consumerresource metacommunities with homogeneous patches? Journal of Theoretical Biology, 277(1):99–110, May 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001214.

Argasinski:2017:NSL

[AR17] Krzysztof Argasinski and Ryszard Rudnicki. Nest site lottery revisited: towards a mechanistic model of population growth suppressed by the availability of nest sites. Journal of Theoretical Biology, 420(??):279–289, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301157.

Andrada:2013:GRQ

[ARB13] Emanuel Andrada, Christian Rode, and Reinhard Blickhan. Grounded running in quails: Simulations indicate benefits of observed fixed aperture angle between legs before touchdown. Journal of Theoretical Biology, 335(??):97–107, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003032.

Archetti:2011:CTE

[Arc11] Marco Archetti. Contract theory for the evolution of cooperation: The right incentives attract the right partners. Journal of Theoretical Biology, 269(1):201–207, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931000559X.

Archetti:2013:EPM

[Arc13] Marco Archetti. Evolution of polygamous marriage by maximization of inclusive fitness. Journal of Theoretical Biology, 319(??):134–143, February 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005929.

Archetti:2014:EDW

[Arc14] Marco Archetti. Evolutionary dynamics of the Warburg effect: Glycolysis as a collective action problem among cancer cells. Journal of Theoretical Biology, 341(??):1–8, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004451.

Archetti:2016:CAC

[Arc16] Marco Archetti. Cooperation among cancer cells as public goods games on Voronoi networks. Journal of Theoretical Biology, 396(??):191–203, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001235.

Andersen:2012:LE

[ARD12] Michael S. Andersen, John Rasmussen, and Michael Damsgaard. Letter to the Editor. Journal of Theoretical Biology, 298(??):154–155, April 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000161.

Aquino:2014:SHR

[ARD14] K. M. Aquino, P. A. Robinson, and P. M. Drysdale. Spatiotemporal hemodynamic response functions derived from physiology. Journal of Theoretical Biology, 347(??):118–136, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005924.

Argasinski:2012:DSR

[Arg12] Krzysztof Argasinski. The dynamics of sex ratio evolution dynamics of global population parameters. Journal of Theoretical Biology, 309(??):134–146, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002676.

Aguilera:2014:SHL

[ARG14] Luis U. Aguilera and Jesús Rodríguez-González. Studying HIV latency by modeling the interaction between HIV proteins and the innate immune response. Journal of Theoretical Biology, 360(??):67–77, November 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003683.

Almeira:2017:RTB

[ARG17] N. Almeira and S. Risau-Gusman. Role of transcriptional bursts in cellular oscillations. Journal of Theoretical Biology, 426(??):49–56, August 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302515.

Aguilera:2019:MET

[ARG19] Luis U. Aguilera and Jesús Rodríguez-González. Modeling the effect of tat inhibitors on HIV latency. Journal of Theoretical Biology, 473(??):20–27, July 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301596.

Ashyani:2018:MAE

[ARM18] A. Ashyani, O. RabieiMotlagh, and H. M. Mohammadinejad. A mathematical approach to effects of CTLs on cancer virotherapy in the second injection of virus. Journal of Theoretical Biology, 453(??):78–87, September 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302583.

Abeysuriya:2014:PVN

[ARR14] R. G. Abeysuriya, C. J. Rennie, and P. A. Robinson. Prediction and verification of nonlinear sleep spindle harmonic oscillations. Journal of Theoretical Biology, 344(??):70–77, March 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005407.

Aviram:2015:MYV

[ARZ15] Ira Aviram, Avinoam Rabinovitch, and Arieh Zaritsky. Maximizing yields of virulent phage: The T4/Escherichia coli system as a test case. Journal of Theoretical Biology, 364(??):428–432, January 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005591.

Archetti:2012:RGT

[AS12] Marco Archetti and István Scheuring. Review: Game theory of public goods in one-shot social dilemmas without assortment. Journal of Theoretical Biology, 299(??):9–20, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311003171.

Archetti:2013:TPG

[AS13] Marco Archetti and István Scheuring. Trading public goods stabilizes interspecific mutualism. Journal of Theoretical Biology, 318(??):58–67, February 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005619.

Archetti:2016:EOH

[AS16] Marco Archetti and István Scheuring. Evolution of optimal Hill coefficients in nonlinear public goods games. Journal of Theoretical Biology, 406(??):73–82, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301643.

Aboelkassem:2015:MMA

[ASC15] Yasser Aboelkassem, Dragana Savic, and Stuart G. Campbell. Mathematical modeling of aortic valve dynamics during systole. Journal of Theoretical Biology, 365(??):280–288, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006171.

Aboelkassem:2016:MPC

[ASC16] Yasser Aboelkassem, Dragana Savic, and Stuart G. Campbell. Modeling predicts a connection between sinus vortex effects and aortic compliance. Journal of Theoretical Biology, 389(??):306–309, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500538X.

Aznar:2015:MDA

[ASF^{+}15] Victoria Romeo Aznar, María Sol De Majo, Sylvia Fischer, Diego Francisco, Mario A. Natiello, and Hernán G. Solari. A model for the development of Aedes (Stegomyia) aegypti as a function of the available food. Journal of Theoretical Biology, 365(??):311–324, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006055.

Al-Shammari:2014:MCO

[ASGE14] Abdullah A. Al-Shammari, Eamonn A. Gaffney, and Stuart Egginton. Modelling capillary oxygen supply capacity in mixed muscles: Capillary domains revisited. Journal of Theoretical Biology, 356(??):47–61, September 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002367.

Ashcroft:2017:EPG

[ASGG17] Peter Ashcroft, Cassandra E. R. Smith, Matthew Garrod, and Tobias Galla. Effects of population growth on the success of invading mutants. Journal of Theoretical Biology, 420(??):232– 240, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301261.

Abdalrahman:2015:TBP

[ASH15] T. Abdalrahman, S. Scheiner, and C. Hellmich. Is trabecular bone permeability governed by molecular ordering-induced fluid viscosity gain? Arguments from re-evaluation of experimental data in the framework of homogenization theory. Journal of Theoretical Biology, 365(??):433–444, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006006.

Abroudi:2017:CCS

[ASK17] Ali Abroudi, Sandhya Samarasinghe, and Don Kulasiri. A comprehensive complex systems approach to the study and analysis of mammalian cell cycle control system in the presence of DNA damage stress. Journal of Theoretical Biology, 429(??):204–228, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302898.

Aurbach:2018:CBC

[ASL^{+}18] Annika Aurbach, Baptiste Schmid, Felix Liechti, Ndaona Chokani, and Reza Abhari. Complex behaviour in complex terrain — modelling bird migration in a high resolution wind field across mountainous terrain to simulate observed patterns. Journal of Theoretical Biology, 454(??):126–138, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302868.

Arnold:2013:CCE

[ASP13] Christian Arnold, Peter F. Stadler, and Sonja J. Prohaska. Chromatin computation: Epigenetic inheritance as a pattern reconstruction problem. Journal of Theoretical Biology, 336(??):61–74, November 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003354.

Ali:2015:MCP

[ASRM15] Iftikhar Ali, Zainab Sardar, Awais Rasheed, and Tariq Mahmood. Molecular characterization of the puroindolinea and b alleles in synthetic hexaploid wheats and in silico functional and structural insights into Pina-D1. Journal of Theoretical Biology, 376(??):1–7, July 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001551.

Algar:2019:ASH

[ASS19] Shannon Dee Algar, Thomas Stemler, and Michael Small. The active selfish herd. Journal of Theoretical Biology, 471(??):82– 90, June 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301286.

Atkins:2010:LAE

[AT10] K. E. Atkins and J. M. J. Travis. Local adaptation and the evolution of species’ ranges under climate change. Journal of Theoretical Biology, 266(3):449–457, October 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003632.

Antonioni:2014:RDC

[ATB14] Alberto Antonioni, Marco Tomassini, and Pierre Buesser. Random diffusion and cooperation in continuous twodimensional space. Journal of Theoretical Biology, 344(??):40– 48, March 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005456.

Avila:2014:SIG

[ÁTC^{+}14] Marco Fidel Ávila, Daniel Torrente, Ricardo Cabezas, Ludis Morales, Luis Miguel García-Segura, Janneth Gonzalez, and George E. Barreto. Structural insights from GRP78-NF-κB binding interactions: a computational approach to understand a possible neuroprotective pathway in brain injuries. Journal of Theoretical Biology, 345(??):43–51, March 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005614.

Altrock:2012:MSS

[ATG12] Philipp M. Altrock, Arne Traulsen, and Tobias Galla. The mechanics of stochastic slowdown in evolutionary games. Journal of Theoretical Biology, 311(??):94–106, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003372.

Akita:2012:MEG

[ATI12] Tetsuya Akita, Shohei Takuno, and Hideki Innan. Modeling evolutionary growth of a microRNA-mediated regulation system. Journal of Theoretical Biology, 311(??):54–65, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003451.

Akiyama:2010:MMC

[ATK10] Masakazu Akiyama, Atsushi Tero, and Ryo Kobayashi. A mathematical model of cleavage. Journal of Theoretical Biology, 264(1):84–94, May 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005839.

Altrock:2010:UUB

[ATRR10] Philipp M. Altrock, Arne Traulsen, R. Guy Reeves, and Floyd A. Reed. Using underdominance to bi-stably transform local populations. Journal of Theoretical Biology, 267(1):62–75, November 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004030.

Allen:2012:HMA

[ATTN12] Benjamin Allen, Arne Traulsen, Corina E. Tarnita, and Martin A. Nowak. How mutation affects evolutionary games on graphs. Journal of Theoretical Biology, 299(??):97–105, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001809.

Ahmed:2015:RPR

[AUE15] Hafiz Ahmed, Rosane Ushirobira, and Denis Efimov. On robustness of phase resetting to cell division under entrainment. Journal of Theoretical Biology, 387(??):206–213, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931500483X.

Aboelkassem:2019:HWW

[AV19] Yasser Aboelkassem and Zdravko Virag. A hybrid Windkessel– Womersley model for blood flow in arteries. Journal of Theoretical Biology, 462(??):499–513, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305952.

Alger:2012:GHR

[AW12] Ingela Alger and Jörgen W. Weibull. A generalization of Hamilton’s rule — love others how much? Journal of Theoretical Biology, 299(??):42–54, April 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002505.

Asatryan:2015:NVD

[AWK15] Ani Asatryan, Dominik Wodarz, and Natalia L. Komarova. New virus dynamics in the presence of multiple infection. Journal of Theoretical Biology, 377(??):98–109, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001769.

Allhoff:2015:ISD

[AWRD15] Korinna T. Allhoff, Eva Marie Weiel, Tobias Rogge, and Barbara Drossel. On the interplay of speciation and dispersal: An evolutionary food web model in space. Journal of Theoretical Biology, 366(??):46–56, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006493.

Axelrod:2012:LEC

[Axe12] Robert Axelrod. Launching “the evolution of cooperation”. Journal of Theoretical Biology, 299(??):21–24, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002153.

Aita:2015:EDP

[AY15] Takuyo Aita and Tetsuya Yomo. Evolutionary dynamics of a polymorphic self-replicator population with a finite population size and hyper mutation rate. Journal of Theoretical Biology, 382(??):298–308, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003367.

Ayati:2012:MDB

[Aya12] Bruce P. Ayati. Microbial dormancy in batch cultures as a function of substrate-dependent mortality. Journal of Theoretical Biology, 293(??):34–40, January 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005108.

Aydogmus:2018:DEN

[Ayd18] Ozgur Aydogmus. Discovering the effect of nonlocal payoff calculation on the stabilty of ESS: Spatial patterns of hawk–dove game in metapopulations. Journal of Theoretical Biology, 442(??):87–97, April 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300304.

Acuna-Zegarra:2018:RAG

[AZOLVH18] Manuel Adrian Acuña-Zegarra, Daniel Olmos-Liceaga, and Jorge X. Velasco-Hernández. The role of animal grazing in the spread of Chagas disease. Journal of Theoretical Biology, 457(??):19–28, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304041.

Balakrishnan:2010:RHB

[BA10] J. Balakrishnan and B. Ashok. The role of Hopf bifurcation dynamics in sensory processes. Journal of Theoretical Biology, 265(2):126–135, July 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000189X.

Basios:2019:HLC

[BA19] Vasileios Basios and Chris G. Antonopoulos. Hyperchaos and labyrinth chaos: Revisiting Thomas-Rössler systems. Journal of Theoretical Biology, 460(??):153–159, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305010.

Bach:2015:CMB

[Bac15] Dominik R. Bach. A cost minimisation and Bayesian inference model predicts startle reflex modulation across species. Journal of Theoretical Biology, 370(??):53–60, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000405.

Barua:2014:RBF

[BAGG14] Amlan Barua, Jhankar Acharya, Saroj Ghaskadbi, and Pranay Goel. The relationship between fasting plasma glucose and HbA 1c during intensive periods of glucose control in antidiabetic therapy. Journal of Theoretical Biology, 363(??):158–163, December 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004755.

Balasuriya:2010:IDD

[Bal10] Sanjeeva Balasuriya. Invasions with density-dependent ecological parameters. Journal of Theoretical Biology, 266(4):

657–666, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519310003929.

Ballantyne:2013:EMF

[Bal13] Ford Ballantyne IV. Evaluating model fit to determine if logarithmic transformations are necessary in allometry: a comment on the exchange between Packard (2009) and Kerkhoff and Enquist (2009). Journal of Theoretical Biology, 317(??):418–421, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005139.

Buzea:2011:SIS

[BAM^{+}11] C. Gh. Buzea, M. Agop, Evelina Moraru, Bogdan A. Stana, Manuela Gîrţu, and D. Iancu. Some implications of scale relativity theory in avascular stages of growth of solid tumors in the presence of an immune system response. Journal of Theoretical Biology, 282(1):52–64, August 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002359.

Blickhan:2015:PHR

[BAM^{+}15] Reinhard Blickhan, Emanuel Andrada, Roy Müller, Christian Rode, and Naomichi Ogihara. Positioning the hip with respect to the COM: Consequences for leg operation. Journal of Theoretical Biology, 382(??):187–197, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500315X.

Barclay:2011:CHI

[Bar11] Pat Barclay. Competitive helping increases with the size of biological markets and invades defection. Journal of Theoretical Biology, 281(1):47–55, July 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002232.

Boete:2014:IMB

[BAR14] C. Boëte, F. B. Agusto, and R. G. Reeves. Impact of mating behaviour on the success of malaria control through a single inundative release of transgenic mosquitoes. Journal of Theoretical Biology, 347(??):33–43, April 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000198.

Bartoszek:2016:PES

[Bar16] Krzysztof Bartoszek. Phylogenetic effective sample size. Journal of Theoretical Biology, 407(??):371–386, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630159X.

Barbosa:2018:ITS

[Bar18] Valmir C. Barbosa. Information-theoretic signatures of biodiversity in the barcoding gene. Journal of Theoretical Biology, 451(??):111–116, August 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302340.

Barik:2019:HUT

[Bar19] Sailen Barik. How universal is the transcription regulatory moonlighting role of fructose 1,6-bisphosphate aldolase? Journal of Theoretical Biology, 464(??):112–114, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830643X.

Baum:2018:OEE

[Bau18] David A. Baum. The origin and early evolution of life in chemical composition space. Journal of Theoretical Biology, 456(??):295–304, November 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303953.

Bhattacharyya:2010:GDM

[BB10] Samit Bhattacharyya and C. T. Bauch. A game dynamic model for delayer strategies in vaccinating behaviour for pediatric infectious diseases. Journal of Theoretical Biology, 267(3):276–282, December 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004716.

Banerjee:2011:TPF

[BB11a] Subhasis Banerjee and Indrani Bose. Transient pulse formation in jasmonate signaling pathway. Journal of Theoretical Biology, 273(1):188–196, March 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310007009.

Bose:2011:CEA

[BB11b] Amitabha Bose and Victoria Booth. Co-existent activity patterns in inhibitory neuronal networks with short-term synaptic depression. Journal of Theoretical Biology, 272(1):42–54, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006430.

Biro:2013:CRA

[BB13] Jan C. Biro and Josephine M. K. Biro. The concept of RNA-assisted protein folding: Representation of amino acid kinetics at the tRNA level. Journal of Theoretical Biology, 317(??):168–174, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005103.

Baninajjaryan:2015:TSE

[BB15] Alireza Baninajjaryan and Yaghoub Tadi Beni. Theoretical study of the effect of shear deformable shell model, elastic foundation and size dependency on the vibration of protein microtubule. Journal of Theoretical Biology, 382(??):111–121, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003173.

Boemo:2019:MMH

[BB19] Michael A. Boemo and Helen M. Byrne. Mathematical modelling of a hypoxia-regulated oncolytic virus delivered by tumour-associated macrophages. Journal of Theoretical Biology, 461(??):102–116, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305265.

Berec:2018:EMF

[BBB18] Ludek Berec, Veronika Bernhauerová, and Barbara Boldin. Evolution of mate-finding Allee effect in prey. Journal of Theoretical Biology, 441(??):9–18, March 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305635.

Brush:2018:IRN

[BBD18] Eleanor Brush, Åke Brännström, and Ulf Dieckmann. Indirect reciprocity with negative assortment and limited information can promote cooperation. Journal of Theoretical Biology, 443(??):56–65, April 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300055.

Boutwell:2013:EOL

[BBDB13] Brian B. Boutwell, J. C. Barnes, Raelynn Deaton, and Kevin M. Beaver. On the evolutionary origins of lifecourse persistent offending: a theoretical scaffold for Moffitt’s developmental taxonomy. Journal of Theoretical Biology, 322(??):72–80, April 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000246.

Berezansky:2014:ETG

[BBI14] Leonid Berezansky, Elena Braverman, and Lev Idels. Effect of treatment on the global dynamics of delayed pathological angiogenesis models. Journal of Theoretical Biology, 363(??):13–21, December 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004603.

Barazzuol:2010:MMB

[BBJ^{+}10] Lara Barazzuol, Neil G. Burnet, Raj Jena, Bleddyn Jones, Sarah J. Jefferies, and Norman F. Kirkby. A mathematical model of brain tumour response to radiotherapy and chemotherapy considering radiobiological aspects. Journal of Theoretical Biology, 262(3):553–565, February 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005025.

Blum:2011:DCL

[BBJDS11] Yvonne Blum, Aleksandra Birn-Jeffery, Monica A. Daley, and Andre Seyfarth. Does a crouched leg posture enhance running stability and robustness? Journal of Theoretical Biology, 281(1):97–106, July 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100230X.

Boldrini:1997:RNL

[BBM^{+}97] J. L. Boldrini, R. C. Bassanezi, A. C. Moretti, F. J. Von Zuben, W. A. C. Godoy, C. J. Von Zuben, and S. F. dos Reis. RETRACTED: Non-local interactions and the dynamics of dispersal in immature insects. Journal of Theoretical Biology, 185(4):523–531, April 21, 1997. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519396903203. See [BBM^{+}13c, BBM^{+}13a, BBM^{+}13b].

Boldrini:2013:CNLa

[BBM^{+}13a] J. L. Boldrini, R. C. Bassanezi, A. C. Moretti, F. J. Von Zuben, W. A. C. Godoy, C. J. Von Zuben, and S. F. dos Reis. Corrigendum to “Non-local interactions and the dynamics of dispersal in immature insects” [J. Theor. Biol. 185 (4) (1997) 523–531]. Journal of Theoretical Biology, 336(??):251, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003445. See [BBM^{+}97,

BBM^{+}13c].

Boldrini:2013:CNLb

[BBM^{+}13b] J. L. Boldrini, R. C. Bassanezi, A. C. Moretti, F. J. Von Zuben, W. A. C. Godoy, C. J. Von Zuben, and S. F. dos Reis. Corrigendum to “Non-local interactions and the dynamics of dispersal in immature insects” [J. Theor. Biol. 185 (4) (1997) 523–531]. Journal of Theoretical Biology, 336(??):252–258, November 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003433. See [BBM^{+}97].

Boldrini:2013:RRN

[BBM^{+}13c] J. L. Boldrini, R. C. Bassanezi, A. C. Moretti, F. J. Von Zuben, W. A. C. Godoy, C. J. Von Zuben, and S. F. dos Reis. RETRACTED: Retraction notice to “non-local interactions and the dynamics of dispersal in immature insects”. Journal of Theoretical Biology, 336(??):250, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000441. See [BBM^{+}97].

Butail:2013:ACC

[BBP13] Sachit Butail, Erik M. Bollt, and Maurizio Porfiri. Analysis and classification of collective behavior using generative modeling and nonlinear manifold learning. Journal of Theoretical Biology, 336(??):185–199, November 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003706.

Bazil:2010:BMM

[BBR10] Jason N. Bazil, Gregery T. Buzzard, and Ann E. Rundell. A bioenergetic model of the mitochondrial population undergoing permeability transition. Journal of Theoretical Biology, 265(4):672–690, August 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002821.

Barker:2012:GTM

[BBR12] Heather A. Barker, Mark Broom, and Jan Rychtár. A game theoretic model of kleptoparasitism with strategic arrivals and departures of beetles at dung pats. Journal of Theoretical Biology, 300(??):292–298, May 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000586.

Ballard:2016:PEF

[BBR16] P. G. Ballard, N. G. Bean, and J. V. Ross. The probability of epidemic fade-out is non-monotonic in transmission rate for the Markovian SIR model with demography. Journal of Theoretical Biology, 393(??):170–178, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931600045X.

Brackley:2012:MPM

[BBRT12] Chris A. Brackley, David S. Broomhead, M. Carmen Romano, and Marco Thiel. A max-plus model of ribosome dynamics during mRNA translation. Journal of Theoretical Biology, 303(??):128–140, June 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001245.

Bougaran:2010:MCC

[BBS10] Gaël Bougaran, Olivier Bernard, and Antoine Sciandra. Modeling continuous cultures of microalgae colimited by nitrogen and phosphorus. Journal of Theoretical Biology, 265(3):443–454, August 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002018.

Bayer:2018:TPM

[BBS18a] Péter Bayer, Joel S. Brown, and Katerina Stanková. A twophenotype model of immune evasion by cancer cells. Journal of Theoretical Biology, 455(??):191–204, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303382.

Bouda:2018:WRS

[BBS18b] Martin Bouda, Craig Brodersen, and James Saiers. Whole root system water conductance responds to both axial and radial traits and network topology over natural range of trait variation. Journal of Theoretical Biology, 456(??):49–61, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303576.

Bafghi:2015:BAC

[BBT^{+}15] M. Ghaemi Bafghi, M. R. Bassami, Gh. R. Hashemi Tabar, M. R. Saberi, A. R. Haghparast, and H. Dehghani. A bioinformatic approach to check the spatial epitope structure of an immunogenic protein coded by DNA vaccine plasmids. Journal of Theoretical Biology, 380(??):315–320, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931500212X.

Belmonte-Beitia:2013:MBI

[BBWS^{+}13] J. Belmonte-Beitia, T. E. Woolley, J. G. Scott, P. K. Maini, and E. A. Gaffney. Modelling biological invasions: Individual to population scales at interfaces. Journal of Theoretical Biology, 334(??):1–12, October 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002646.

Broom:2013:DNP

[BC13] Mark Broom and Chris Cannings. A dynamic network population model with strategic link formation governed by individual preferences. Journal of Theoretical Biology, 335(??):160–168, October 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002968.

Benhamou:2015:UFL

[BC15] Simon Benhamou and Julien Collet. Ultimate failure of the Lévy foraging hypothesis: Two-scale searching strategies outperform scale-free ones even when prey are scarce and cryptic. Journal of Theoretical Biology, 387(??):221–227, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004841.

Bai:2019:RDC

[BC19] Xiaolu Bai and Xiaolin Chen. Rational design, conformational analysis and membrane-penetrating dynamics study of bac2a-derived antimicrobial peptides against grampositive clinical strains isolated from pyemia. Journal of Theoretical Biology, 473(??):44–51, July 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301250.

Bisson:2019:MFT

[BCBD19] Anne Bisson, Céline Casenave, Simon Boudsocq, and Tanguy Daufresne. Maximization of fertility transfers from rangeland to cropland: The contribution of control theory. Journal of Theoretical Biology, 469(??):187–200, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300323.

Billiard:2016:ECH

[BCF^{+}16a] Sylvain Billiard, Pierre Collet, Régis Ferrière, Sylvie Méléard, and Viet Chi Tran. The effect of competition and horizontal trait inheritance on invasion, fixation, and polymorphism. Journal of Theoretical Biology, 411(??):48–58, December 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303344.

Bologna:2016:NLM

[BCF16b] Mauro Bologna, Kristopher J. Chandía, and J. C. Flores. A non-linear mathematical model for a three species ecosystem: Hippos in Lake Edward. Journal of Theoretical Biology, 389(??):83–87, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005251.

Beeton:2019:MTE

[BCF19] Nicholas J. Beeton, Scott Carver, and Lawrence K. Forbes. A model for the treatment of environmentally transmitted sarcoptic mange in bare-nosed wombats (Vombatus ursinus). Journal of Theoretical Biology, 462(??):466–474, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305861.

Broom:2010:SMK

[BCFR10] Mark Broom, Mary L. Crowe, Meghan R. Fitzgerald, and Jan Rychtár. The stochastic modelling of kleptoparasitism using a Markov process. Journal of Theoretical Biology, 264(2):266– 272, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000160.

Barberis:2010:CGP

[BCGD10] L. Barberis, C. A. Condat, A. S. Gliozzi, and P. P. Delsanto. Concurrent growth of phenotypic features: a phenomenological universalities approach. Journal of Theoretical Biology, 264(1):123–129, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005979.

Broom:2019:RFA

[BCK19] Mark Broom, Ross Cressman, and Vlastimil Krivan. Revisiting the “fallacy of averages” in ecology: Expected gain per unit time equals expected gain divided by expected time. Journal of Theoretical Biology, 483(??):Article 109993, December 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303418.

Belov:2013:RBM

[BCKS13] Oleg V. Belov, Ochbadrakh Chuluunbaatar, Mikhail I. Kapralov, and Nasser H. Sweilam. The role of the bacterial mismatch repair system in SOS-induced mutagenesis: a theoretical background. Journal of Theoretical Biology, 332(??):30–41, September 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300194X.

Bowness:2018:MEB

[BCPG18] Ruth Bowness, Mark A. J. Chaplain, Gibin G. Powathil, and Stephen H. Gillespie. Modelling the effects of bacterial cell state and spatial location on tuberculosis treatment: Insights from a hybrid multiscale cellular automaton model. Journal of Theoretical Biology, 446(??):87–100, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301206.

Baltcheva:2010:LDH

[BCPL10] Irina Baltcheva, Laura Codarri, Giuseppe Pantaleo, and Jean-Yves Le Boudec. Lifelong dynamics of human CD4 + CD25 + regulatory T cells: Insights from in vivo data and mathematical modeling. Journal of Theoretical Biology, 266(2):307–322, September 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003140.

Bodgi:2016:MMR

[BCPM^{+}16] Larry Bodgi, Aurélien Canet, Laurent Pujo-Menjouet, Annick Lesne, Jean-Marc Victor, and Nicolas Foray. Mathematical models of radiation action on living cells: From the target theory to the modern approaches. A historical and critical review. Journal of Theoretical Biology, 394(??):93–101, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316000515.

Chan:2019:RGN

[bCR19] Yao ban Chan and Charles Robin. Reconciliation of a gene network and species tree. Journal of Theoretical Biology, 472(??):54–66, July 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301341.

Chan:2017:IIL

[bCRS17] Yao ban Chan, Vincent Ranwez, and Céline Scornavacca. Inferring incomplete lineage sorting, duplications, transfers and losses with reconciliations. Journal of Theoretical Biology, 432(??):1–13, November 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303740.

Butner:2016:HAB

[BCS^{+}16] Joseph D. Butner, Yao-Li Chuang, Eman Simbawa, A. S. Al-Fhaid, S. R. Mahmoud, Vittorio Cristini, and Zhihui Wang. A hybrid agent-based model of the developing mammary terminal end bud. Journal of Theoretical Biology, 407(??):259–270, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302235.

Barta:2010:GTM

[BD10a] Efrath Barta and Arieh Drugan. Glucose transport from mother to fetus — a theoretical study. Journal of Theoretical Biology, 263(3):295–302, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005773.

Blanchard:2010:ECT

[BD10b] Pierrick Blanchard and Sébastien Devillard. On the evolutionary consequences of a Trivers–Willard effect in humans: Further analysis of Dr S. Kanazawa’s paper on ‘beautiful parents’. Journal of Theoretical Biology, 264(4):1296–1298, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001797.

Bulelzai:2012:LTE

[BD12] M. A. K. Bulelzai and Johan L. A. Dubbeldam. Long time evolution of atherosclerotic plaques. Journal of Theoretical Biology, 297(??):1–10, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005935.

Baptestini:2013:CNS

[BdABY13] Elizabeth M. Baptestini, Marcus A. M. de Aguiar, and Yaneer Bar-Yam. Conditions for neutral speciation via isolation by distance. Journal of Theoretical Biology, 335(??):51–56, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931300283X.

Boukal:2014:LHI

[BDE^{+}14] David S. Boukal, Ulf Dieckmann, Katja Enberg, Mikko Heino, and Christian Jørgensen. Life-history implications of the allometric scaling of growth. Journal of Theoretical Biology, 359(??):199–207, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002951.

Berthaume:2012:PFE

[BDID^{+}12] Michael A. Berthaume, Paul C. Dechow, Jose Iriarte-Diaz, Callum F. Ross, David S. Strait, Qian Wang, and Ian R. Grosse. Probabilistic finite element analysis of a craniofacial finite element model. Journal of Theoretical Biology, 300(??):242–253, May 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000410.

Borgogno:2012:SRC

[BDLR12] Fabio Borgogno, Paolo D’Odorico, Francesco Laio, and Luca Ridolfi. Stochastic resonance and coherence resonance in groundwater-dependent plant ecosystems. Journal of Theoretical Biology, 293(??):65–73, January 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004711.

Breban:2010:GMS

[BDR10] Romulus Breban, John M. Drake, and Pejman Rohani. A general multi-strain model with environmental transmission: Invasion conditions for the disease-free and endemic states. Journal of Theoretical Biology, 264(3):729–736, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001281.

Barbosa:2012:QDN

[BDS12] Valmir C. Barbosa, Raul Donangelo, and Sergio R. Souza. Quasispecies dynamics with network constraints. Journal of Theoretical Biology, 312(??):114–119, November 7, 2012.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003852.

Baker:2015:MFM

[BDTR15] C. Baker, N. Dowson, P. Thomas, and S. Rose. Modelling of FDG metabolism in liver voxels. Journal of Theoretical Biology, 365(??):390–402, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006183.

Buchelnikov:2014:QCV

[BE14] Anatoly S. Buchelnikov and Maxim P. Evstigneev. Quantitative correlation of the in vitro biological effect with parameters of molecular complexation in mutagen-interceptor systems. Journal of Theoretical Biology, 357(??):268–271, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314003592.

Becchetti:2014:EFG

[Bec14] Andrea Becchetti. Empirically founded genotype-phenotype maps from mammalian cyclic nucleotide-gated ion channels. Journal of Theoretical Biology, 363(??):205–215, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004998.

Berec:2010:AFD

[BEK10] Ludek Berec, Jan Eisner, and Vlastimil Křivan. Adaptive foraging does not always lead to more complex food webs. Journal of Theoretical Biology, 266(2):211–218, September 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310003255.

Benhamou:2014:PIC

[Ben14] Simon Benhamou. Path integration and coordinate systems. Journal of Theoretical Biology, 349(??):163–166, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000873.

Bertrand:2014:LPA

[Ber14] Robert L. Bertrand. Lag phase-associated iron accumulation is likely a microbial counter-strategy to host iron sequestration: Role of the ferric uptake regulator (fur). Journal of Theoretical Biology, 359(??):72–79, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003221.

Barton:2013:SVO

[BF13] A. Barton and A. J. Fendrik. Sustained vs. oscillating expressions of Ngn2, Dll1 and Hes1: a model of neural differentiation of embryonic telencephalon. Journal of Theoretical Biology, 328(??):1–8, July 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000969.

Barton:2015:RSC

[BF15] A. Barton and A. J. Fendrik. Retinogenesis: Stochasticity and the competency model. Journal of Theoretical Biology, 373(??):73–81, May 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001241.

Bernstein:2016:ASP

[BF16] Jason Bernstein and John Fricks. Analysis of single particle diffusion with transient binding using particle filtering. Journal of Theoretical Biology, 401(??):109–121, July 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300418.

Burgos-Florez:2016:FBS

[BFGAGA16] F. J. Burgos-Flórez, M. E. Gavilán-Alfonso, and D. A. Garzón-Alvarado. Flat bones and sutures formation in the human cranial vault during prenatal development and infancy: a computational model. Journal of Theoretical Biology, 393(??):127–144, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000217.

Bertuzzi:2010:NCE

[BFGS10] Alessandro Bertuzzi, Antonio Fasano, Alberto Gandolfi, and Carmela Sinisgalli. Necrotic core in EMT6/Ro tumour spheroids: Is it caused by an ATP deficit? Journal of Theoretical Biology, 262(1):142–150, January 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004548.

Banks:2015:IMS

[BFH^{+}15] H. T. Banks, Kevin B. Flores, Shuhua Hu, Eric Rosenberg, Maria Buzon, Xu Yu, and Matthias Lichterfeld. Immunomodulatory strategies for reduction of HIV reservoir cells. Journal of Theoretical Biology, 372(??):146–158, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500065X.

Bara:2018:TMF

[BFJ^{+}18] Ouassim Bara, Michel Fliess, Cédric Join, Judy Day, and Seddik M. Djouadi. Toward a model-free feedback control synthesis for treating acute inflammation. Journal of Theoretical Biology, 448(??):26–37, July 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301632.

Bryant:2017:QMP

[BFLS17] Christopher Bryant, Mareike Fischer, Simone Linz, and Charles Semple. On the quirks of maximum parsimony and likelihood on phylogenetic networks. Journal of Theoretical Biology, 417(??):100–108, March 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300115.

Borthagaray:2010:VPF

[BFM10] Ana I. Borthagaray, Miguel A. Fuentes, and Pablo A. Marquet. Vegetation pattern formation in a fog-dependent ecosystem. Journal of Theoretical Biology, 265(1):18–26, July 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310002031.

Barton:2014:SMN

[BFR14] A. Barton, A. J. Fendrik, and E. Rotondo. A stochastic model of neurogenesis controlled by a single factor. Journal of Theoretical Biology, 355(??):77–82, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001945.

Baudier:2018:GRM

[BFS18] Adrien Baudier, François Fages, and Sylvain Soliman. Graphical requirements for multistationarity in reaction networks and their verification in BioModels. Journal of Theoretical Biology, 459(??):79–89, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304600.

Bode:2010:MNE

[BFW10] Nikolai W. F. Bode, Daniel W. Franks, and A. Jamie Wood. Making noise: Emergent stochasticity in collective motion. Journal of Theoretical Biology, 267(3):292–299, December 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310004571.

Biton:2012:SEO

[BG12] Eli Biton and Hezi Gildor. The seasonal effect in onedimensional daisyworld. Journal of Theoretical Biology, 314(??):145–156, December 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004651.

Bello:2012:SDR

[BGCB12] Omar Sierra Bello, Janneth Gonzalez, Francisco Capani, and George E. Barreto. In silico docking reveals possible riluzole binding sites on nav1.6 sodium channel: Implications for amyotrophic lateral sclerosis therapy. Journal of Theoretical Biology, 315(??):53–63, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004729.

Bodgi:2013:SFD

[BGD^{+}13] Larry Bodgi, Adeline Granzotto, Clément Devic, Guillaume Vogin, Annick Lesne, Jean-François Bottollier-Depois, Jean-Marc Victor, Mira Maalouf, Georges Fares, and Nicolas Foray. A single formula to describe radiation-induced protein relocalization: Towards a mathematical definition of individual radiosensitivity. Journal of Theoretical Biology, 333(??):135–145, September 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300249X.

Berends:2014:EST

[BGF^{+}14] H. Berends, W. J. J. Gerrits, J. France, J. L. Ellis, S. M. van Zijderveld, and J. Dijkstra. Evaluation of the SF 6 tracer technique for estimating methane emission rates with reference to dairy cows using a mechanistic model. Journal of Theoretical Biology, 353(??):1–8, July 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001246.

Bartoszek:2017:UOU

[BGKL17] Krzysztof Bartoszek, Sylvain Glémin, Ingemar Kaj, and Martin Lascoux. Using the Ornstein–Uhlenbeck process to model the evolution of interacting populations. Journal of Theoretical Biology, 429(??):35–45, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302801.

Barthelemy:2010:FEM

[BGL10] Marc Barthélemy, Claude Godrèche, and Jean-Marc Luck. Fluctuation effects in metapopulation models: Percolation and pandemic threshold. Journal of Theoretical Biology, 267(4):554–564, December 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004893.

Becker:2019:SCH

[BGL^{+}19] Nils B. Becker, Matthias Günther, Congxin Li, Adrien Jolly, and Thomas Höfer. Stem cell homeostasis by integral feedback through the niche. Journal of Theoretical Biology, 481(??):100–109, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306301.

Bourhis:2019:SDA

[BGLR^{+}19] Yoann Bourhis, Timothy R. Gottwald, Francisco J. Lopez-Ruiz, Sujin Patarapuwadol, and Frank van den Bosch. Sampling for disease absence-deriving informed monitoring from epidemic traits. Journal of Theoretical Biology, 461(??):8–16, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305204.

Bajeux:2019:ICP

[BGM19] Nicolas Bajeux, Frédéric Grognard, and Ludovic Mailleret. Influence of the components of propagule pressure, Allee effects, and stochasticity on the time to establish introduced populations. Journal of Theoretical Biology, 471(??):91–107, June 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301134.

Browne:2015:MCT

[BGW15] Cameron Browne, Hayriye Gulbudak, and Glenn Webb. Modeling contact tracing in outbreaks with application to Ebola. Journal of Theoretical Biology, 384(??):33–49, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315003823.

Blackwood:2011:ETD

[BH11] Julie C. Blackwood and Alan Hastings. The effect of time delays on Caribbean coral-algal interactions. Journal of Theoretical Biology, 273(1):37–43, March 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006776.

Bate:2012:RPB

[BH12] Andrew M. Bate and Frank M. Hilker. Rabbits protecting birds: Hypopredation and limitations of hyperpredation.

Journal of Theoretical Biology, 297(??):103–115, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931100614X.

Barton:2013:PDV

[BH13a] Kamil A. Bartoń and Thomas Hovestadt. Prey density, value, and spatial distribution affect the efficiency of area-concentrated search. Journal of Theoretical Biology, 316(??):61–69, January 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004705.

Bate:2013:PPO

[BH13b] Andrew M. Bate and Frank M. Hilker. predator–prey oscillations can shift when diseases become endemic. Journal of Theoretical Biology, 316(??):1–8, January 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004912.

Benzekry:2013:MTD

[BH13c] Sébastien Benzekry and Philip Hahnfeldt. Maximum tolerated dose versus metronomic scheduling in the treatment of metastatic cancers. Journal of Theoretical Biology, 335(??):235–244, October 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313003196.

Berndt:2011:ICC

[BHBH11] Nikolaus Berndt, Sabrina Hoffmann, Jan Benda, and Hermann-Georg Holzhütter. The influence of the chloride currents on action potential firing and volume regulation of excitable cells studied by a kinetic model. Journal of Theoretical Biology, 276(1):42–49, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000348.

Broom:2012:ESS

[BHBR12] M. Broom, R. N. Hughes, M. T. Burrows, and G. D. Ruxton. Evolutionarily stable sexual allocation by both stressed and unstressed potentially simultaneous hermaphrodites within the same population. Journal of Theoretical Biology, 309(??):96–102, September 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002810.

Biermans:2014:DDM

[BHH^{+}14] Geert Biermans, Nele Horemans, Niel Hens, Jordi Vives i Batlle, Hildegarde Vandenhove, and Ann Cuypers. A dynamic dosimetry model for radioactive exposure scenarios in Arabidopsis thaliana. Journal of Theoretical Biology, 347(??):54– 62, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000216.

Black:2014:ECP

[BHKR14] Andrew J. Black, Thomas House, Matt J. Keeling, and Joshua V. Ross. The effect of clumped population structure on the variability of spreading dynamics. Journal of Theoretical Biology, 359(??):45–53, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003312.

Bromage:2012:PEE

[BHLH12] Timothy G. Bromage, Russell T. Hogg, Rodrigo S. Lacruz, and Chen Hou. Primate enamel evinces long period biological timing and regulation of life history. Journal of Theoretical Biology, 305(??):131–144, July 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001804.

Burli:2018:MAT

[BHO^{+}18] Christine Bürli, Helmut Harbrecht, Peter Odermatt, Somphou Sayasone, and Nakul Chitnis. Mathematical analysis of the transmission dynamics of the liver fluke, Opisthorchis viverrini. Journal of Theoretical Biology, 439(??):181–194, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305271.

Bica:2017:ABP

[BHP17] Ion Bica, Thomas Hillen, and Kevin J. Painter. Aggregation of biological particles under radial directional guidance. Journal of Theoretical Biology, 427(??):77–89, August 1, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317302606.

Broom:2010:OIA

[BHR10] Mark Broom, Andrew D. Higginson, and Graeme D. Ruxton. Optimal investment across different aspects of anti-predator defences. Journal of Theoretical Biology, 263(4):579–586, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000068.

Buhler:2011:AML

[BHSB11] Jonas Bühler, Gregor Huber, Friederike Schmid, and Peter Blümler. Analytical model for long-distance tracer-transport in plants. Journal of Theoretical Biology, 270(1):70–79, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005904.

Bessho:2010:OSS

[BI10] Kazuhiro Bessho and Yoh Iwasa. Optimal seasonal schedules and the relative dominance of heteromorphic and isomorphic life cycles in macroalgae. Journal of Theoretical Biology, 267(2):201–212, November 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000425X.

Bessho:2012:VES

[BI12] Kazuhiro Bessho and Yoh Iwasa. Variability in the evolutionarily stable seasonal timing of germination and maturation of annuals and the mode of competition. Journal of Theoretical Biology, 304(??):66–80, July 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001440.

Braverman:2019:IHV

[BI19] Elena Braverman and Ilia Ilmer. On the interplay of harvesting and various diffusion strategies for spatially heterogeneous populations. Journal of Theoretical Biology, 466(??):106–118, April 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300360.

Bessho:2015:EAH

[BID15] Kazuhiro Bessho, Yoh Iwasa, and Troy Day. The evolutionary advantage of haploid versus diploid microbes in nutrient-poor environments. Journal of Theoretical Biology, 383(??):116–129, October 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003744.

Booton:2017:SMA

[BIMC17] Ross D. Booton, Yoh Iwasa, James A. R. Marshall, and Dylan Z. Childs. Stress-mediated Allee effects can cause the sudden collapse of honey bee colonies. Journal of Theoretical Biology, 420(??):213–219, May 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301212.

Brinck:2017:EEA

[BJ17] Katharina Brinck and Henrik Jeldtoft Jensen. The evolution of ecosystem ascendency in a complex systems based model. Journal of Theoretical Biology, 428(??):18–25, September 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302825.

Byun:2018:MMA

[BJ18] Jong Hyuk Byun and Il Hyo Jung. Mathematical modeling of antibody drug conjugates with the target and tubulin dynamics to predict AUC. Journal of Theoretical Biology, 443(??):113–124, April 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300456.

Blanco:2018:SWM

[BJB18] R. Ernesto Blanco, Washington W. Jones, and Nicolás Benech. The seismic wave motion camouflage of large carnivorous dinosaurs. Journal of Theoretical Biology, 459(??):154–161, December 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304806.

Bauer:2010:RCT

[BJJR10] Amy L. Bauer, Trachette L. Jackson, Yi Jiang, and Thimo Rohlf. Receptor cross-talk in angiogenesis: Mapping environmental cues to cell phenotype using a stochastic, Boolean signaling network model. Journal of Theoretical Biology, 264(3):838–846, June 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000158X.

Besenhard:2016:MTC

[BJO^{+}16] Maximilian O. Besenhard, Monika Jarzabek, Alice C. O’Farrell, John J. Callanan, Jochen H. M. Prehn, Annette T. Byrne, and Heinrich J. Huber. Modelling tumour cell proliferation from vascular structure using tissue decomposition into avascular elements. Journal of Theoretical Biology, 402(??):129–143, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931630056X.

Bartoszek:2013:TSH

[BJOS13] Krzysztof Bartoszek, Graham Jones, Bengt Oxelman, and Serik Sagitov. Time to a single hybridization event in a group of species with unknown ancestral history. Journal of Theoretical Biology, 322(??):1–6, April 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000209.

Brown:2012:HFS

[BJPP12] R. G. Brown, A. F. James, J. W. Pitchford, and M. J. Plank. Habitat fragmentation: Simple models for local persistence and the spread of invasive species. Journal of Theoretical Biology, 310(??):231–238, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003268.

Bokka:2015:MIP

[BJWL15] Kishore K. Bokka, Edwin C. Jesudason, David Warburton, and Sharon R. Lubkin. Morphogenetic implications of peristaltic fluid-tissue dynamics in the embryonic lung. Journal of Theoretical Biology, 382(??):378–385, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500301X.

Bornholdt:2019:EDC

[BK19] Stefan Bornholdt and Stuart Kauffman. Ensembles, dynamics, and cell types: Revisiting the statistical mechanics perspective on cellular regulation. Journal of Theoretical Biology, 467(??):15–22, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300530.

Burns:2016:RPA

[BKC^{+}16] Jonathan Burns, Denys Kukushkin, Xiao Chen, Laura F. Landweber, Masahico Saito, and Nataša Jonoska. Recurring patterns among scrambled genes in the encrypted genome of the ciliate Oxytricha trifallax. Journal of Theoretical Biology, 410(??):171–180, December 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302806.

Brown:2014:TDN

[BKL14] L. E. Brown, J. R. King, and M. Loose. Two different network topologies yield bistability in models of mesoderm and anterior mesendoderm specification in amphibians. Journal of Theoretical Biology, 353(??):67–77, July 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001507.

Belov:2015:QMM

[BKL^{+}15] Oleg V. Belov, Eugene A. Krasavin, Marina S. Lyashko, Munkhbaatar Batmunkh, and Nasser H. Sweilam. A quantitative model of the major pathways for radiation-induced DNA double-strand break repair. Journal of Theoretical Biology, 366(??):115–130, February 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005633.

Bugay:2015:MNE

[BKPV15] Aleksandr N. Bugay, Evgeny A. Krasavin, Aleksandr Yu. Parkhomenko, and Maria A. Vasilyeva. Modeling nucleotide excision repair and its impact on UV-induced mutagenesis during SOS-response in bacterial cells. Journal of Theoretical Biology, 364(??):7–20, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005025.

Banach:2014:FOD

[BKR14] Mateusz Banach, Leszek Konieczny, and Irena Roterman. The fuzzy oil drop model, based on hydrophobicity density distribution, generalizes the influence of water environment on protein structure and function. Journal of Theoretical Biology, 359(??):6–17, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002781.

Boie:2016:EQS

[BKSW16] Sebastian Boie, Vivien Kirk, James Sneyd, and Martin Wechselberger. Effects of quasi-steady-state reduction on biophysical models with oscillations. Journal of Theoretical Biology, 393(??):16–31, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315006025.

Bonachela:2014:ECB

[BL14] Juan A. Bonachela and Simon A. Levin. Evolutionary comparison between viral lysis rate and latent period. Journal of Theoretical Biology, 345(??):32–42, March 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005572.

Barroso:2015:LCL

[BL15a] Gustavo V. Barroso and David R. Luz. On the limits of complexity in living forms. Journal of Theoretical Biology, 379(??):89–90, August 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002143.

Behar:2015:CPS

[BL15b] Hilla Behar and Yoram Louzoun. Commensal pathogens as a source of a coexistence mechanism. Journal of Theoretical Biology, 370(??):45–52, April 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000399.

Binder:2015:IPR

[BLNR15] Benjamin J. Binder, Kerry A. Landman, Donald F. Newgreen, and Joshua V. Ross. Incomplete penetrance: The role of stochasticity in developmental cell colonization. Journal of Theoretical Biology, 380(??):309–314, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002635.

Boghigian:2010:CAP

[BLP10] Brett A. Boghigian, Kyongbum Lee, and Blaine A. Pfeifer. Computational analysis of phenotypic space in heterologous polyketide biosynthesis — applications to Escherichia coli, Bacillus subtilis, and Saccharomyces cerevisiae. Journal of Theoretical Biology, 262(2):197–207, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004810.

Borchering:2012:SSR

[BLS^{+}12] Rebecca K. Borchering, Hao Liu, Mara C. Steinhaus, Carl L. Gardner, and Yang Kuang. A simple spatiotemporal rabies model for skunk and bat interaction in northeast Texas. Journal of Theoretical Biology, 314(??):16–22, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004559.

Bordewich:2017:LSG

[BLS17a] Magnus Bordewich, Simone Linz, and Charles Semple. Lost in space? Generalising subtree prune and regraft to spaces of phylogenetic networks. Journal of Theoretical Biology, 423(??):1–12, June 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301674.

Burgess:2017:ERI

[BLS^{+}17b] Andrew E. F. Burgess, Tommaso Lorenzi, Pietà G. Schofield, Stephen F. Hubbard, and Mark A. J. Chaplain. Examining the role of individual movement in promoting coexistence in a spatially explicit prisoner’s dilemma. Journal of Theoretical Biology, 419(??):323–332, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300917.

Barrios:2018:AED

[BLV18] Edwin Barrios, Sunmi Lee, and Olga Vasilieva. Assessing the effects of daily commuting in two-patch dengue dynamics: a case study of Cali, Colombia. Journal of Theoretical Biology, 453(??):14–39, September 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302431.

Boyle:2011:SPP

[BLW11] Richard A. Boyle, Timothy M. Lenton, and Andrew J. Watson. Symbiotic physiology promotes homeostasis in daisyworld. Journal of Theoretical Biology, 274(1):170–182, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931100004X.

Bo:2019:PRB

[BLZ^{+}19] Suling Bo, Hong Li, Qiang Zhang, Zhanyuan Lu, Tonglaga Bao, and Xiaoqing Zhao. Potential relations between postspliced introns and mature mRNAs in the Caenorhabditis elegans genome. Journal of Theoretical Biology, 467(??):7–14, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300487.

Black:2010:SAE

[BM10] Andrew J. Black and Alan J. McKane. Stochastic amplification in an epidemic model with seasonal forcing. Journal of Theoretical Biology, 267(1):85–94, November 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004236.

Benard:2011:GSE

[BM11] Emmanuel Benard and Christian J. Michel. A generalization of substitution evolution models of nucleotides to genetic motifs. Journal of Theoretical Biology, 288(??):73–83, November 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003596.

Bansal:2012:IPE

[BM12] Shweta Bansal and Lauren Ancel Meyers. The impact of past epidemics on future disease dynamics. Journal of Theoretical Biology, 309(??):176–184, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002998.

Berec:2014:WPP

[BM14] L. Berec and D. Maxin. Why have parasites promoting mating success been observed so rarely? Journal of Theoretical Biology, 342(??):47–61, February 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005109.

Burger:2016:ETA

[BM16] Oskar Burger and Trifon I. Missov. Evolutionary theory of ageing and the problem of correlated Gompertz parameters. Journal of Theoretical Biology, 408(??):34–41, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302363.

Bocharov:2018:IBR

[BMB^{+}18a] G. Bocharov, A. Meyerhans, N. Bessonov, S. Trofimchuk, and V. Volpert. Interplay between reaction and diffusion processes in governing the dynamics of virus infections. Journal of Theoretical Biology, 457(??):221–236, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304168.

Browning:2018:IPL

[BMB^{+}18b] Alexander P. Browning, Scott W. McCue, Rachelle N. Binny, Michael J. Plank, Esha T. Shah, and Matthew J. Simpson. Inferring parameters for a lattice-free model of cell migration and proliferation using experimental data. Journal of Theoretical Biology, 437(??):251–260, January 21, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305015.

Barreto:2017:GAR

[BMdA17] Wendell P. Barreto, Flavia M. D. Marquitti, and Marcus A. M. de Aguiar. A genetic approach to the rock-paper-scissors game. Journal of Theoretical Biology, 421(??):146–152, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301510.

Bridge:2018:MSB

[BMF^{+}18] L. J. Bridge, J. Mead, E. Frattini, I. Winfield, and G. Ladds. Modelling and simulation of biased agonism dynamics at a g protein-coupled receptor. Journal of Theoretical Biology, 442(??):44–65, April 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300183.

Bunimovich-Mendrazitsky:2011:CMM

[BMGC11a] Svetlana Bunimovich-Mendrazitsky, Jean Claude Gluckman, and Joel Chaskalovic. Corrigendum to “A mathematical model of combined bacillus Calmette–Guerin (BCG) and interleukin(IL)-2 immunotherapy of superficial bladder cancer” [J. Theor. Biol. 277(1) (2011) 27–40]. Journal of Theoretical Biology, 281(1):140, July 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001962. See [BMGC11b].

Bunimovich-Mendrazitsky:2011:MMC

[BMGC11b] Svetlana Bunimovich-Mendrazitsky, Jean Claude Gluckman, and Joel Chaskalovic. A mathematical model of combined bacillus Calmette–Guerin (BCG) and interleukin (IL)-2 immunotherapy of superficial bladder cancer. Journal of Theoretical Biology, 277(1):27–40, May 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519311001020. See corrigendum [BMGC11a].

Barac:2019:GOU

[BMI19] Diana Barac, Michael D. Multerer, and Dagmar Iber. Global optimization using Gaussian processes to estimate biological parameters from image data. Journal of Theoretical Biology, 481(??):233–248, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305927.

Bowden:2014:ODE

[BMM^{+}14] L. G. Bowden, P. K. Maini, D. E. Moulton, J. B. Tang, X. T. Wang, P. Y. Liu, and H. M. Byrne. An ordinary differential equation model for full thickness wounds and the effects of diabetes. Journal of Theoretical Biology, 361(??):87–100, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003944.

Behbahani:2016:ACL

[BMN16] Mandana Behbahani, Hassan Mohabatkar, and Mokhtar Nosrati. Analysis and comparison of lignin peroxidases between fungi and bacteria using three different modes of Chou’s general pseudo amino acid composition. Journal of Theoretical Biology, 411(??):1–5, December 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302843.

Bakhtiarizadeh:2014:NNS

[BMSEE14] Mohammad Reza Bakhtiarizadeh, Mohammad Moradi-Shahrbabak, Mansour Ebrahimi, and Esmaeil Ebrahimie. Neural network and SVM classifiers accurately predict lipid binding proteins, irrespective of sequence homology. Journal of Theoretical Biology, 356(??):213–222, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002768.

Bonifaci:2012:PCC

[BMV12] Vincenzo Bonifaci, Kurt Mehlhorn, and Girish Varma. Physarum can compute shortest paths. Journal of Theoretical Biology, 309(??):121–133, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003049.

Blyuss:2012:RTA

[BN12] K. B. Blyuss and L. B. Nicholson. The role of tunable activation thresholds in the dynamics of autoimmunity. Journal of Theoretical Biology, 308(??):45–55, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002615.

Blyuss:2015:URA

[BN15] K. B. Blyuss and L. B. Nicholson. Understanding the roles of activation threshold and infections in the dynamics of autoimmune disease. Journal of Theoretical Biology, 375(??):13–20, June 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004743.

Ben-Naoum:2016:AHC

[BNG16] Farah Ben-Naoum and Christophe Godin. Algorithmic height compression of unordered trees. Journal of Theoretical Biology, 389(??):237–252, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005299.

Bocianowski:2012:CTM

[Boc12] Jan Bocianowski. A comparison of two methods to estimate additive-by-additive interaction of QTL effects by a simulation study. Journal of Theoretical Biology, 308(??):20–24, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002603.

Bonsall:2010:INL

[Bon10] Michael B. Bonsall. The impact of non-lethal synergists on the population and evolutionary dynamics of host-pathogen interactions. Journal of Theoretical Biology, 262(4):567–575, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005244.

Bonneuil:2012:MPM

[Bon12] Noël Bonneuil. Multiallelic polymorphism maintained under unpredictable migration and selection. Journal of Theoretical Biology, 293(??):189–196, January 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004917.

Briefer:2010:BSC

[BORA10] Elodie Briefer, Tomasz S. Osiejuk, Fanny Rybak, and Thierry Aubin. Are bird song complexity and song sharing shaped by habitat structure? An information theory and statistical approach. Journal of Theoretical Biology, 262(1):151–164, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004408.

Boswell:2012:MCS

[Bos12] Graeme P. Boswell. Modelling combat strategies in fungal mycelia. Journal of Theoretical Biology, 304(??):226–234, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312001683.

Bozorgmehr:2015:QPS

[Boz15] Joseph E. Hannon Bozorgmehr. Quantifying protein sequences with reference to the genetic code. Journal of Theoretical Biology, 372(??):39–46, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000764.

Bearup:2015:TSI

[BP15] Daniel Bearup and Sergei Petrovskii. On time scale invariance of random walks in confined space. Journal of Theoretical Biology, 367(??):230–245, February 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006791.

Burgos:2016:CAI

[BP16] Andrés C. Burgos and Daniel Polani. Cooperation and antagonism in information exchange in a growth scenario with two species. Journal of Theoretical Biology, 399(??):117–133, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300261.

Band:2018:PIM

[BP18] Leah R. Band and Simon P. Preston. Parameter inference to motivate asymptotic model reduction: An analysis of the gibberellin biosynthesis pathway. Journal of Theoretical Biology, 457(??):66–78, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302753.

Barberis:2015:JFR

[BPC15] L. Barberis, M. A. Pasquale, and C. A. Condat. Joint fitting reveals hidden interactions in tumor growth. Journal of Theoretical Biology, 365(??):420–432, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006286.

Bodova:2015:CSN

[BPF15a] Katarína Bod’ová, David Paydarfar, and Daniel B. Forger. Characterizing spiking in noisy type II neurons. Journal of Theoretical Biology, 365(??):40–54, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005888. See erratum [BPF15b].

Bodova:2015:ECS

[BPF15b] Katarína Bod’ová, David Paydarfar, and Daniel B. Forger. Erratum to: “Characterizing spiking in noisy type II neurons” [J. Theor. Biol. 365 (2015) 40–54]. Journal of Theoretical Biology, 373(??):132, May 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001228. See [BPF15a].

Bubba:2019:CBE

[BPF^{+}19] Federica Bubba, Camille Pouchol, Nathalie Ferrand, Guillaume Vidal, Luis Almeida, Benoît Perthame, and Michèle Sabbah. A chemotaxis-based explanation of spheroid formation in 3D cultures of breast cancer cells. Journal of Theoretical Biology, 479(??):73–80, October 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302796.

Baudrot:2016:CHC

[BPFR16] Virgile Baudrot, Antoine Perasso, Clémentine Fritsch, and Francis Raoul. Competence of hosts and complex foraging behavior are two cornerstones in the dynamics of trophically transmitted parasites. Journal of Theoretical Biology, 397(??):158–168, May 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931600151X.

Berenshtein:2018:ACD

[BPG^{+}18] Igal Berenshtein, Claire B. Paris, Hezi Gildor, Erick Fredj, Yael Amitai, Omri Lapidot, and Moshe Kiflawi. Autocorrelated directional swimming can enhance settlement success and connectivity in fish larvae. Journal of Theoretical Biology, 439(??):76–85, February 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305167.

Buenzli:2012:CMA

[BPGS12a] Pascal R. Buenzli, Peter Pivonka, Bruce S. Gardiner, and David W. Smith. Corrigendum to “Modelling the anabolic response of bone using a cell population model” [J. Theor. Biol. 307 (2012) 42–52]. Journal of Theoretical Biology, 309(??):204, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200327X. See [BPGS12b].

Buenzli:2012:MAR

[BPGS12b] Pascal R. Buenzli, Peter Pivonka, Bruce S. Gardiner, and David W. Smith. Modelling the anabolic response of bone using a cell population model. Journal of Theoretical Biology, 307(??):42–52, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519312002007. See corrigendum [BPGS12a].

Brooks:2012:NDA

[BPLM12] Grace Brooks, Gabriel Provencher, Jinzhi Lei, and Michael C. Mackey. Neutrophil dynamics after chemotherapy and g-CSF: The role of pharmacokinetics in shaping the response. Journal of Theoretical Biology, 315(??):97–109, December 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200450X.

Bartoszek:2012:PCM

[BPM^{+}12] Krzysztof Bartoszek, Jason Pienaar, Petter Mostad, Staffan Andersson, and Thomas F. Hansen. A phylogenetic comparative method for studying multivariate adaptation. Journal of Theoretical Biology, 314(??):204–215, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003918.

Bandiera:2016:EMM

[BPP^{+}16] Lucia Bandiera, Alice Pasini, Lorenzo Pasotti, Susanna Zucca, Giuliano Mazzini, Paolo Magni, Emanuele Giordano, and Simone Furini. Experimental measurements and mathematical modeling of biological noise arising from transcriptional and translational regulation of basic synthetic gene circuits. Journal of Theoretical Biology, 395(??):153–160, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000941.

Briant:2015:MVR

[BPPC15] Linford J. B. Briant, Julian F. R. Paton, Anthony E. Pickering, and Alan R. Champneys. Modelling the vascular response to sympathetic postganglionic nerve activity. Journal of Theoretical Biology, 371(??):102–116, April 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000545.

Bolzoni:2015:RHI

[BPR15] L. Bolzoni, A. Pugliese, and R. Rosà. The role of heterogeneity on the invasion probability of mosquito-borne diseases in multi-host models. Journal of Theoretical Biology, 377(??):25– 35, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001423.

Bianchi:2015:MML

[BPS15] Arianna Bianchi, Kevin J. Painter, and Jonathan A. Sherratt. A mathematical model for lymphangiogenesis in normal and diabetic wounds. Journal of Theoretical Biology, 383(??):61–86, October 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003689.

Baker:2019:FBM

[BPS19] Ruth E. Baker, Andrew Parker, and Matthew J. Simpson. A free boundary model of epithelial dynamics. Journal of Theoretical Biology, 481(??):61–74, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306246.

Broom:2012:PAP

[BR12a] Mark Broom and Graeme D. Ruxton. Perceptual advertisement by the prey of stalking or ambushing predators. Journal of Theoretical Biology, 315(??):9–16, December 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004481.

Broom:2012:GFA

[BR12b] Mark Broom and Jan Rychtár. A general framework for analysing multiplayer games in networks using territorial interactions as a case study. Journal of Theoretical Biology, 302(??):70–80, June 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001075.

Broom:2013:ESZ

[BR13] M. Broom and G. D. Ruxton. On the evolutionary stability of zero-cost pooled-equilibrium signals. Journal of Theoretical Biology, 323(??):69–75, April 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000428.

Bordet:2014:SPT

[BR14] Charles Bordet and Louis-Paul Rivest. A stochastic Pella Tomlinson model and its maximum sustainable yield. Journal of Theoretical Biology, 360(??):46–53, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003555.

Black:2015:CEF

[BR15] Andrew J. Black and J. V. Ross. Computation of epidemic final size distributions. Journal of Theoretical Biology, 367(??):159–165, February 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006882.

Beauchamp:2016:MSI

[BR16] G. Beauchamp and G. D. Ruxton. Modeling scan and interscan durations in antipredator vigilance. Journal of Theoretical Biology, 390(??):86–96, February 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005652.

Bag:2015:FCE

[BRA15] Susmita Bag, Sudha Ramaiah, and Anand Anbarasu. fabp4 is central to eight obesity associated genes: a functional gene network-based polymorphic study. Journal of Theoretical Biology, 364(??):344–354, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005736.

Biswas:2018:SEI

[BRB18] Tirthabir Biswas, Suhas Rao, and Vikas Bhandawat. A simple extension of inverted pendulum template to explain features of slow walking. Journal of Theoretical Biology, 457(??):112–123, November 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304065.

Baldazzi:2012:IMC

[BRG^{+}12] Valentina Baldazzi, Delphine Ropers, Johannes Geiselmann, Daniel Kahn, and Hidde de Jong. Importance of metabolic coupling for the dynamics of gene expression following a diauxic shift in Escherichia coli. Journal of Theoretical Biology, 295(??):100–115, February 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005728.

Bachar:2016:IMM

[BRK16] Mostafa Bachar, Jochen G. Raimann, and Peter Kotanko. Impulsive mathematical modeling of ascorbic acid metabolism in healthy subjects. Journal of Theoretical Biology, 392(??):35– 47, March 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005950.

Butt:2019:PAP

[BRK19] Ahmad Hassan Butt, Nouman Rasool, and Yaser Daanial Khan. Prediction of antioxidant proteins by incorporating statistical moments based features into Chou’s PseAAC. Journal of Theoretical Biology, 473(??):1–8, July 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301602.

Bilinsky:2015:RSM

[BRN15] L. M. Bilinsky, M. C. Reed, and H. F. Nijhout. The role of skeletal muscle in liver glutathione metabolism during acetaminophen overdose. Journal of Theoretical Biology, 376(??):118–133, July 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500168X.

Brogioli:2013:LCR

[Bro13] Doriano Brogioli. Looking for chemical reaction networks exhibiting a drift along a manifold of marginally stable states. Journal of Theoretical Biology, 318(??):110–123, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005851.

Bruineberg:2018:FEM

[BRP^{+}18] Jelle Bruineberg, Erik Rietveld, Thomas Parr, Leendert van Maanen, and Karl J. Friston. Free-energy minimization in joint agent-environment systems: a niche construction perspective. Journal of Theoretical Biology, 455(??):161–178, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303151.

Brylinski:2013:UAE

[Bry13] Michał Brylinski. The utility of artificially evolved sequences in protein threading and fold recognition. Journal of Theoretical Biology, 328(??):77–88, July 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001227.

Binmore:2011:SES

[BS11] Ken Binmore and Larry Samuelson. Sex and evolutionary stability. Journal of Theoretical Biology, 278(1):63–73, June 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001263.

Bishop:2013:IER

[BS13] Charles M. Bishop and Robin J. Spivey. Integration of exercise response and allometric scaling in endotherms. Journal of Theoretical Biology, 323(??):11–19, April 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000210.

Babaei:2015:PDA

[BS15a] Naser Babaei and Metin U. Salamci. Personalized drug administration for cancer treatment using model reference adaptive control. Journal of Theoretical Biology, 371(??):24–44, April 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315000557.

Bartoszek:2015:CEE

[BS15b] Krzysztof Bartoszek and Serik Sagitov. A consistent estimator of the evolutionary rate. Journal of Theoretical Biology, 371(??):69–78, April 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000284.

Belitsky:2019:RPI

[BS19a] V. Belitsky and G. M. Schütz. RNA polymerase interactions and elongation rate. Journal of Theoretical Biology, 462(??):370–380, February 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305794.

Bennett:2019:LDS

[BS19b] Jamie J. R. Bennett and Jonathan A. Sherratt. Longdistance seed dispersal affects the resilience of banded vegetation patterns in semi-deserts. Journal of Theoretical Biology, 481(??):151–161, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304740.

Bonello:2013:BIS

[BSB^{+}13] Nicolas Bonello, James Sampson, John Burn, Ian J. Wilson, Gail McGrown, Geoff P. Margison, Mary Thorncroft, Philip Crossbie, Andrew C. Povey, Mauro Santibanez-Koref, and Kevin Walters. Bayesian inference supports a location and neighbour-dependent model of DNA methylation propagation at the MGMT gene promoter in lung tumours. Journal of Theoretical Biology, 336(??):87–95, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003421.

Baker:2014:GMC

[BSB14] C. J. Baker, M. Sterling, and P. Berry. A generalised model of crop lodging. Journal of Theoretical Biology, 363(??):1–12, December 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004445.

Bunte:2018:LPM

[BSC^{+}18] Kerstin Bunte, David J. Smith, Michael J. Chappell, Zaki K. Hassan-Smith, Jeremy W. Tomlinson, Wiebke Arlt, and Peter Tiňo. Learning pharmacokinetic models for in vivo glucocorticoid activation. Journal of Theoretical Biology, 455(??):222–231, October 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303497.

Banerjee:2019:ECC

[BSC19] Swarnendu Banerjee, Ram Rup Sarkar, and Joydev Chattopadhyay. Effect of copper contamination on zooplankton epidemics. Journal of Theoretical Biology, 469(??):61–74, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300852.

Beckmann:2019:PAE

[BSH19] Aike Beckmann, C-Elisa Schaum, and Inga Hense. Phytoplankton adaptation in ecosystem models. Journal of Theoretical Biology, 468(??):60–71, May 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300335.

Bergman:2018:IPG

[BSKV18] Juraj Bergman, Dominik Schrempf, Carolin Kosiol, and Claus Vogl. Inference in population genetics using forward and backward, discrete and continuous time processes. Journal of Theoretical Biology, 439(??):166–180, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305477.

Bewick:2017:ISM

[BSL^{+}17] Sharon Bewick, Phillip P. A. Staniczenko, Bingtuan Li, David K. Karig, and William F. Fagan. Invasion speeds in microbial systems with toxin production and quorum sensing. Journal of Theoretical Biology, 420(??):290–303, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300358.

Bhattacharya:2014:MMS

[BSM^{+}14] Bonhi S. Bhattacharya, Peter K. Sweby, Anne-Marie Minihane, Kim G. Jackson, and Marcus J. Tindall. A mathematical model of the sterol regulatory element binding protein 2 cholesterol biosynthesis pathway. Journal of Theoretical Biology, 349(??):150–162, May 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000228.

Bass:2011:ATL

[BSMK11] Ludvik Bass, Michael Sørensen, Ole Lajord Munk, and Susanne Keiding. Analogue tracers and lumped constant in capillary beds. Journal of Theoretical Biology, 285(1):177–181, September 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003390.

Basu:2018:SSM

[BSP18] S. Basu, S. Sutradhar, and R. Paul. Substrate stiffness and mechanical stress due to intercellular cooperativity guides tissue structure. Journal of Theoretical Biology, 457(??):124–136, November 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303710.

Boer:2011:SMM

[BSR^{+}11] H. M. T. Boer, C. Stötzel, S. Röblitz, P. Deuflhard, R. F. Veerkamp, and H. Woelders. A simple mathematical model of the bovine estrous cycle: Follicle development and endocrine interactions. Journal of Theoretical Biology, 278(1):20–31, June 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001238.

Balakrishnan:2014:PMM

[BSR14] Naviyn Prabhu Balakrishnan, Lakshminarayanan Samavedham, and Gade Pandu Rangaiah. Personalized mechanistic models for exercise, meal and insulin interventions in children and adolescents with type 1 diabetes. Journal of Theoretical Biology, 357(??):62–73, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002720.

Bol:2011:PMF

[BSS11] Markus Böl, Heiko Stark, and Nadja Schilling. On a phenomenological model for fatigue effects in skeletal muscles. Journal of Theoretical Biology, 281(1):122–132, July 21, 2011.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000127X.

Bauer:2014:CIE

[BST14] Benedikt Bauer, Reiner Siebert, and Arne Traulsen. Cancer initiation with epistatic interactions between driver and passenger mutations. Journal of Theoretical Biology, 358(??):52–60, October 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002896.

Bastogne:2010:PMT

[BSV^{+}10] T. Bastogne, A. Samson, P. Vallois, S. Wantz-Mézières, S. Pinel, D. Bechet, and M. Barberi-Heyob. Phenomenological modeling of tumor diameter growth based on a mixed effects model. Journal of Theoretical Biology, 262(3):544–552, February 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004834.

Banerjee:2017:MDP

[BT17] Malay Banerjee and Yasuhiro Takeuchi. Maturation delay for the predators can enhance stable coexistence for a class of prey-predator models. Journal of Theoretical Biology, 412(??):154–171, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303472.

Bilinsky:2019:UMM

[BTF19] Lydia M. Bilinsky, David J. Thomas, and Jeffrey W. Fisher. Using mathematical modeling to infer the valence state of arsenicals in tissues: a PBPK model for dimethylarsinic acid (DMA V) and dimethylarsinous acid (DMA III) in mice. Journal of Theoretical Biology, 461(??):215–229, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305332.

Barbillon:2015:NIP

[BTG^{+}15] Pierre Barbillon, Mathieu Thomas, Isabelle Goldringer, Frédéric Hospital, and Stéphane Robin. Network impact on persistence in a finite population dynamic diffusion model: Application to an emergent seed exchange network. Journal of Theoretical Biology, 365(??):365–376, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006225.

Barber:2013:TDM

[BTH^{+}13] Jared Barber, Mark Tronzo, C. Harold Horvat, Gilles Clermont, Jeffrey Upperman, Yoram Vodovotz, and Ivan Yotov. A three-dimensional mathematical and computational model of necrotizing enterocolitis. Journal of Theoretical Biology, 322(??):17–32, April 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005930.

Borzan:2014:HPR

[BTK14] Vladimir Borzan, Boris Tomasević, and Sven Kurbel. Hypothesis: Possible respiratory advantages for heterozygote carriers of cystic fibrosis linked mutations during dusty climate of last glaciation. Journal of Theoretical Biology, 363(??):164–168, December 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931400469X.

Brand:2015:RSS

[BTK15] Samuel P. C. Brand, Michael J. Tildesley, and Matthew J. Keeling. Rapid simulation of spatial epidemics: a spectral method. Journal of Theoretical Biology, 370(??):121–134, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315000363.

Barrack:2014:MCB

[BTO14] Duncan S. Barrack, Rüdiger Thul, and Markus R. Owen. Modelling the coupling between intracellular calcium release and the cell cycle during cortical brain development. Journal of Theoretical Biology, 347(??):17–32, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000137.

Barrack:2015:MCC

[BTO15] Duncan S. Barrack, Rüdiger Thul, and Markus R. Owen. Modelling cell cycle synchronisation in networks of coupled radial glial cells. Journal of Theoretical Biology, 377(??):85–97, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001757.

Baghdadi:2018:MMM

[BTR18] Golnaz Baghdadi, Farzad Towhidkhah, and Reza Rostami. A mathematical model to mimic the shape of event related desynchronization/synchronization. Journal of Theoretical Biology, 453(??):117–124, September 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302650.

Bourouiba:2011:HPA

[BTW11] L. Bourouiba, A. Teslya, and J. Wu. Highly pathogenic avian influenza outbreak mitigated by seasonal low pathogenic strains: Insights from dynamic modeling. Journal of Theoretical Biology, 271(1):181–201, February 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005989.

Birch:2014:IFO

[BUC14] Elsa W. Birch, Madeleine Udell, and Markus W. Covert. Incorporation of flexible objectives and time-linked simulation with flux balance analysis. Journal of Theoretical Biology, 345(??):12–21, March 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005602.

Buenzli:2015:ORB

[Bue15] Pascal R. Buenzli. Osteocytes as a record of bone formation dynamics: a mathematical model of osteocyte generation in bone matrix. Journal of Theoretical Biology, 364(??):418–427, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005670.

Bjorn:2012:EIT

[BUL12a] Lars Olof Björn, Per Uvdal, and Shaoshan Li. Ecological importance of the thermal emissivity of avian eggshells. Journal of Theoretical Biology, 301(??):62–66, May 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000847.

Bullock:2012:EAE

[Bul12b] Seth Bullock. An evolutionary advantage for extravagant honesty. Journal of Theoretical Biology, 292(??):30–38, January 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004887.

Burton:2010:EDC

[Bur10] Richard F. Burton. Evolutionary determinants of the concentrations of inorganic ions in body fluids. Journal of Theoretical Biology, 264(2):639–640, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000470.

Burke:2019:MMC

[Bur19] Meghan A. Burke. Maini’s many contributions to mathematical enzyme kinetics: a review. Journal of Theoretical Biology, 481(??):24–27, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305939.

Balcan:2012:ITS

[BV12] Duygu Balcan and Alessandro Vespignani. Invasion threshold in structured populations with recurrent mobility patterns. Journal of Theoretical Biology, 293(??):87–100, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100525X.

Bellingeri:2013:REF

[BV13] Michele Bellingeri and Simone Vincenzi. Robustness of empirical food webs with varying consumer’s sensitivities to loss of resources. Journal of Theoretical Biology, 333(??):18–26, September 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002014.

Beaunee:2017:CBP

[BVJE17] Gaël Beaunée, Elisabeta Vergu, Alain Joly, and Pauline Ezanno. Controlling bovine paratuberculosis at a regional scale: Towards a decision modelling tool. Journal of Theoretical Biology, 435(??):157–183, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304290.

Bojic:2010:MEF

[BVK10] Tijana Bojić, Aleksandra Vuckovic, and Aleksandar Kalauzi. Modeling EEG fractal dimension changes in wake and drowsy states in humans — a preliminary study. Journal of Theoretical Biology, 262(2):214–222, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004767.

Buhler:2014:CCM

[BvLH14] Jonas Bühler, Eric von Lieres, and Gregor Huber. A class of compartmental models for long-distance tracer transport in plants. Journal of Theoretical Biology, 341(??):131–142, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004530.

Boyles:2013:NFP

[BW13] Justin G. Boyles and Robin W. Warne. A novel framework for predicting the use of facultative heterothermy by endotherms.

Journal of Theoretical Biology, 336(??):242–245, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003846.

Bacelar:2011:LHM

[BWB11] Flora S. Bacelar, Andrew White, and Mike Boots. Life history and mating systems select for male biased parasitism mediated through natural selection and ecological feedbacks. Journal of Theoretical Biology, 269(1):131–137, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310005321.

Boyle:2012:NSC

[BWL12] Richard A. Boyle, Hywel T. P. Williams, and Timothy M. Lenton. Natural selection for costly nutrient recycling in simulated microbial metacommunities. Journal of Theoretical Biology, 312(??):1–12, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003621.

Blazej:2019:MAT

[BWM^{+}19] Paweł Blazej, Małgorzata Wnetrzak, Dorota Mackiewicz, Przemyslaw Gagat, and Paweł Mackiewicz. Many alternative and theoretical genetic codes are more robust to amino acid replacements than the standard genetic code. Journal of Theoretical Biology, 464(??):21–32, March 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306295.

Bousios:2010:ACF

[BWP10] Alexandros Bousios, David Waxman, and Stephen R. Pearce. Absence of close-facing retrotransposons: a comparison of molecular data and theory. Journal of Theoretical Biology, 264(2):205–210, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005566.

Bielinska-Waz:2010:CSB

[BWS10] Dorota Bielińska-Waz̧ and Shankar Subramaniam. Classification studies based on a spectral representation of DNA. Journal of Theoretical Biology, 266(4):667–674, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310003954.

Bewick:2017:IDC

[BWY^{+}17] Sharon Bewick, Guoqing Wang, Hannah Younes, Bingtuan Li, and William F. Fagan. Invasion dynamics of competing species with stage-structure. Journal of Theoretical Biology, 435(??):12–21, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303624.

Baek:2017:DBC

[BYJ17] Seung Ki Baek, Su Do Yi, and Hyeong-Chai Jeong. Duality between cooperation and defection in the presence of tit-for-tat in replicator dynamics. Journal of Theoretical Biology, 430(??):215–220, October 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303570.

Booton:2018:IBI

[BYM^{+}18] Ross D. Booton, Ryo Yamaguchi, James A. R. Marshall, Dylan Z. Childs, and Yoh Iwasa. Interactions between immunotoxicants and parasite stress: Implications for host health. Journal of Theoretical Biology, 445(??):120–127, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931830078X.

Bi:2011:KBV

[BYY^{+}11] Jianjun Bi, Huilan Yang, Huacheng Yan, Rengang Song, and Jianyong Fan. Knowledge-based virtual screening of HLA-a*0201-restricted CD8 + T-cell epitope peptides from herpes simplex virus genome. Journal of Theoretical Biology, 281(1):133–139, July 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002189.

Brooks:2010:NMC

[BZ10] Christopher P. Brooks and Haimeng Zhang. A null model of community disassembly effects on vector-borne disease risk. Journal of Theoretical Biology, 264(3):866–873, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001396.

Barranca:2018:CSR

[BZ18] Victor J. Barranca and Xiuqi George Zhu. A computational study of the role of spatial receptive field structure in processing natural and non-natural scenes. Journal of Theoretical Biology, 454(??):268–277, October 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303059.

Bertacchi:2016:TLH

[BZA16] Daniela Bertacchi, Fabio Zucca, and Roberto Ambrosini. The timing of life history events in the presence of soft disturbances. Journal of Theoretical Biology, 389(??):287–303, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005275.

Ben-Zion:2010:MED

[BZCS10] Yossi Ben-Zion, Yahel Cohen, and Nadav M. Shnerb. Modeling epidemics dynamics on heterogeneous networks. Journal of Theoretical Biology, 264(2):197–204, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000494.

Bazzazi:2018:CMS

[BZJP18] Hojjat Bazzazi, Yu Zhang, Mohammad Jafarnejad, and Aleksander S. Popel. Computational modeling of synergistic interaction between αvβ3 integrin and VEGFR2 in endothelial cells: Implications for the mechanism of action of angiogenesis-modulating integrin-binding peptides. Journal of Theoretical Biology, 455(??):212–221, October 14, 2018.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303199.

Bertalan:2017:MMC

[BZL17] Zsolt Bertalan, Stefano Zapperi, and Caterina A. M. La Porta. Modeling mechanical control of spindle orientation of intestinal crypt stem cells. Journal of Theoretical Biology, 430(??):103–108, October 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303429.

Bassil:2017:ASE

[BZN17] Guillaume Bassil, Manuel Zarzoso, and Sami F. Noujaim. Allometric scaling of electrical excitation and propagation in the mammalian heart. Journal of Theoretical Biology, 419(??):238–242, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303216.

Cockrell:2017:SRI

[CA17] Chase Cockrell and Gary An. Sepsis reconsidered: Identifying novel metrics for behavioral landscape characterization with a high-performance computing implementation of an agentbased model. Journal of Theoretical Biology, 430(??):157–168, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303491.

Colombo:2018:NPD

[CA18] E. H. Colombo and C. Anteneodo. Nonlinear population dynamics in a bounded habitat. Journal of Theoretical Biology, 446(??):11–18, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300912.

Cortes:2011:MME

[CAA11] L. M. Peinado Cortés, J. C. Vanegas Acosta, and D. A. Garzón Alvarado. A mechanobiological model of epiphysis structures formation. Journal of Theoretical Biology, 287(??):13–25, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003614.

Cabrita:2018:RFC

[Cab18] Paulo Cabrita. Resin flow in conifers. Journal of Theoretical Biology, 453(??):48–57, September 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302571.

Capitan:2013:DIF

[CAG13] José A. Capitán, Alex Arenas, and Roger Guimerà. Degree of intervality of food webs: From body-size data to models. Journal of Theoretical Biology, 334(??):35–44, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002762.

Castro-Abril:2017:CSM

[CAGM^{+}17] H. A. Castro-Abril, J. M. Guevara, M. A. Moncayo, S. J. Shefelbine, L. A. Barrera, and D. A. Garzón-Alvarado. Cellular scale model of growth plate: An in silico model of chondrocyte hypertrophy. Journal of Theoretical Biology, 428(??):87–97, September 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302217.

Cammarano:2011:CDS

[Cam11] Mario Cammarano. Co-dominance and succession in forest dynamics: The role of interspecific differences in crown transmissivity. Journal of Theoretical Biology, 285(1):46–57, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003365.

Campbell:2015:EIC

[Cam15] R. B. Campbell. The effect of inbreeding constraints and offspring distribution on time to the most recent common ancestor. Journal of Theoretical Biology, 382(??):74–80, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003161.

Carrere:2017:OVC

[Car17] Cécile Carrère. Optimization of an in vitro chemotherapy to avoid resistant tumours. Journal of Theoretical Biology, 413(??):24–33, January 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303666.

Cooney:2016:AEC

[CAV16] Daniel Cooney, Benjamin Allen, and Carl Veller. Assortment and the evolution of cooperation in a Moran process with exponential fitness. Journal of Theoretical Biology, 409(??):38–46, November 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302600.

Chu:2010:MFM

[CB10a] Dominique Chu and David J. Barnes. Modeling fimbriae mediated parasite-host interactions. Journal of Theoretical Biology, 264(4):1169–1176, June 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001724.

Couee:2010:EIM

[CB10b] Ivan Couée and Françoise Bringel. Expanding importance of mRNA expression in understanding stress and stress responses. Journal of Theoretical Biology, 266(3):479–482, October 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003607.

Crowley:2010:VVV

[CB10c] Philip H. Crowley and Kyung Hwan Baik. Variable valuations and voluntarism under group selection: An evolutionary public goods game. Journal of Theoretical Biology, 265(3):238–244, August 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002341.

Chamchod:2013:MPV

[CB13a] Farida Chamchod and John C. Beier. Modeling plasmodium vivax: Relapses, treatment, seasonality, and G6PD deficiency. Journal of Theoretical Biology, 316(??):25–34, January 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004468.

Cornish-Bowden:2013:PSN

[CB13b] Athel Cornish-Bowden. The physiological significance of negative cooperativity revisited. Journal of Theoretical Biology, 319(??):144–147, February 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312006388.

Cornish-Bowden:2015:TGR

[CB15] Athel Cornish-Bowden. Tibor Gánti and Robert Rosen: Contrasting approaches to the same problem. Journal of Theoretical Biology, 381(??):6–10, September 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500243X. See comment [HS16].

Codling:2016:BDI

[CB16] Edward A. Codling and Nikolai W. F. Bode. Balancing direct and indirect sources of navigational information in a leaderless model of collective animal movement. Journal of Theoretical Biology, 394(??):32–42, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000278.

Cimentepe:2017:POT

[CB17a] Ibrahim Cimentepe and Haluk O. Bingol. Parent oriented teacher selection causes language diversity. Journal of Theoretical Biology, 429(??):142–148, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303089.

Cornish-Bowden:2017:LMO

[CB17b] Athel Cornish-Bowden. Lynn Margulis and the origin of the eukaryotes. Journal of Theoretical Biology, 434(??):1, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304459.

Caiado:2016:FLA

[CBBO16] Camila C. S. Caiado, William A. Brock, R. Alexander Bentley, and Michael J. O’Brien. Fitness landscapes among many options under social influence. Journal of Theoretical Biology, 405(??):5–16, September 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931600014X.

Cornish-Bowden:2017:LBL

[CBC17] Athel Cornish-Bowden and María Luz Cárdenas. Life before LUCA. Journal of Theoretical Biology, 434(??):68–74, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302370.

Chen:2018:MIL

[CBC^{+}18] Jing Chen, John C. Beier, Robert Stephen Cantrell, Chris Cosner, Douglas O. Fuller, Yongtao Guan, Guoyan Zhang, and Shigui Ruan. Modeling the importation and local transmission of vector-borne diseases in Florida: The case of Zika outbreak in 2016. Journal of Theoretical Biology, 455(??):342–356, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303503.

Cunningham:2018:OCD

[CBGS18] Jessica J. Cunningham, Joel S. Brown, Robert A. Gatenby, and Katerina Stanková. Optimal control to develop therapeutic strategies for metastatic castrate resistant prostate cancer. Journal of Theoretical Biology, 459(??):67–78, December 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304582.

Calistri:2014:VCS

[CBL14] Elisa Calistri, Marcello Buiatti, and Roberto Livi. Variation and constraints in species-specific promoter sequences. Journal of Theoretical Biology, 363(??):357–366, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004548.

Colman:2012:SSD

[CBP12] Andrew M. Colman, Lindsay Browning, and Briony D. Pulford. Spontaneous similarity discrimination in the evolution of cooperation. Journal of Theoretical Biology, 299(??):162–171, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311002645.

Castro-Chavez:2010:RVA

[CC10] Fernando Castro-Chavez. The rules of variation: Amino acid exchange according to the rotating circular genetic code. Journal of Theoretical Biology, 264(3):711–721, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001815.

Capitan:2011:SAMa

[CC11a] José A. Capitán and José A. Cuesta. Species assembly in model ecosystems, i: Analysis of the population model and the invasion dynamics. Journal of Theoretical Biology, 269(1):330–343, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005096.

Castellano:2011:SAA

[CC11b] Sergio Castellano and Paolo Cermelli. Sampling and assessment accuracy in mate choice: a random-walk model of information processing in mating decision. Journal of Theoretical Biology, 274(1):161–169, April 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000026.

Capitan:2017:SCE

[CCA17] José A. Capitán, Sara Cuenda, and David Alonso. Stochastic competitive exclusion leads to a cascade of species extinctions. Journal of Theoretical Biology, 419(??):137–151, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300632.

Costa:2013:BIM

[CCAdS13] F. H. S. Costa, M. Campos, O. E. Aiéllo, and M. A. A. da Silva. Basic ingredients for mathematical modeling of tumor growth in vitro: Cooperative effects and search for space. Journal of Theoretical Biology, 337(??):24–29, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003731.

Capitan:2011:SAMb

[CCB11a] José A. Capitán, José A. Cuesta, and Jordi Bascompte. Species assembly in model ecosystems, II: Results of the assembly process. Journal of Theoretical Biology, 269(1): 344–355, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519310005655.

Chatelain:2011:MCE

[CCB11b] Clément Chatelain, Pasquale Ciarletta, and Martine Ben Amar. Morphological changes in early melanoma development: Influence of nutrients, growth inhibitors and cell-adhesion mechanisms. Journal of Theoretical Biology, 290(??):46–59, December 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004322.

Costa:2015:UGR

[CCdS15] F. H. S. Costa, M. Campos, and M. A. A. da Silva. The universal growth rate behavior and regime transition in adherent cell colonies. Journal of Theoretical Biology, 387(??):181–188, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004853.

Cantini:2014:LNA

[CCF^{+}14] Laura Cantini, Claudia Cianci, Duccio Fanelli, Emma Massi, and Luigi Barletti. Linear noise approximation for stochastic oscillations of intracellular calcium. Journal of Theoretical Biology, 349(??):92–99, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931400054X.

Casha:2018:MMP

[CCG^{+}18] Aaron R. Casha, Liberato Camilleri, Marilyn Gauci, Ruben Gatt, David Sladden, Stanley Chetcuti, and Joseph N. Grima. A mathematical model for pressure-based organs behaving as biological pressure vessels. Journal of Theoretical Biology, 450(??):37–42, August 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302170.

Castro:2019:ATD

[CCNT19] Laureano Castro, Miguel Ángel Castro-Nogueira, and Miguel Ángel Toro. Assessor teaching and the development of the capacity to innovate and to imitate. Journal of Theoretical Biology, 472(??):88–94, July 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301456.

Calvetti:2015:SDC

[CCS15] Daniela Calvetti, Yougan Cheng, and Erkki Somersalo. A spatially distributed computational model of brain cellular metabolism. Journal of Theoretical Biology, 376(??):48–65, July 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001526.

Chen:2016:RBC

[CCS^{+}16] Xianli Chen, Jia Chen, Bin Shao, Linjie Zhao, Haicen Yue, and Qi Ouyang. Relationship between cancer mutations and parameter sensitivity in Rb pathway. Journal of Theoretical Biology, 404(??):120–125, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300881.

Carrillo:2018:AVC

[CCS18] José Antonio Carrillo, Annachiara Colombi, and Marco Scianna. Adhesion and volume constraints via nonlocal interactions determine cell organisation and migration profiles. Journal of Theoretical Biology, 445(??):75–91, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300821.

Caraco:2016:ETP

[CCW16] Thomas Caraco, Carrie A. Cizauskas, and Ing-Nang Wang. Environmentally transmitted parasites: Host-jumping in a heterogeneous environment. Journal of Theoretical Biology, 397(??):33–42, May 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001211.

Chen:2019:PFA

[CCY^{+}19] Guodong Chen, Man Cao, Jialin Yu, Xinyun Guo, and Shaoping Shi. Prediction and functional analysis of prokaryote lysine acetylation site by incorporating six types of features into Chou’s general PseAAC. Journal of Theoretical Biology, 461(??):92–101, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305290.

Cushing:2016:MGD

[CD16] J. M. Cushing and Odo Diekmann. The many guises of R_{0} (a didactic note). Journal of Theoretical Biology, 404(??):295–302, September 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301515.

Chiu:2018:LSF

[CDC18] Jimmy Ka Ho Chiu, Tharam S. Dillon, and Yi-Ping Phoebe Chen. Large-scale frequent stem pattern mining in RNA families. Journal of Theoretical Biology, 455(??):131–139, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303394.

Collignon:2012:LBS

[CDD12] Bertrand Collignon, Jean Louis Deneubourg, and Claire Detrain. Leader-based and self-organized communication: Modelling group-mass recruitment in ants. Journal of Theoretical Biology, 313(??):79–86, November 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003773.

Crauste:2010:MSF

[CDGV10] Fabien Crauste, Ivan Demin, Olivier Gandrillon, and Vitaly Volpert. Mathematical study of feedback control roles and relevance in stress erythropoiesis. Journal of Theoretical Biology, 263(3):303–316, April 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005992.

Ciupe:2011:PSP

[CDK11] Stanca M. Ciupe, Patrick De Leenheer, and Thomas B. Kepler. Paradoxical suppression of poly-specific broadly neutralizing antibodies in the presence of strain-specific neutralizing antibodies following HIV infection. Journal of Theoretical Biology, 277(1):55–66, May 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000762.

Ciulla:2014:ACT

[CDM^{+}14] Michele M. Ciulla, Federico De Marco, Elisa Montelatici, Lorenza Lazzari, Gianluca L. Perrucci, and Fabio Magrini. Assessing cytokines’ talking patterns following experimental myocardial damage by applying Shannon’s information theory. Journal of Theoretical Biology, 343(??):25–31, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005183.

Caravagna:2010:TSI

[CdMB10] Giulio Caravagna, Alberto d’Onofrio, Paolo Milazzo, and Roberto Barbuti. Tumour suppression by immune system through stochastic oscillations. Journal of Theoretical Biology, 265(3):336–345, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002420.

Canon:2017:MMS

[CDO17] Mario Cañón, Hernando Diaz, and Andrés Olarte. Mathematical model for the spread of drug resistance in Plasmodium falciparum parasite considering transmission conditions. Journal of Theoretical Biology, 435(??):1–11, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304101.

Couturier:2011:FLG

[CdPD11] Etienne Couturier, Sylvain Courrech du Pont, and Stéphane Douady. The filling law: a general framework for leaf folding and its consequences on leaf shape diversity. Journal of Theoretical Biology, 289(??):47–64, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004231.

Chen:2017:RBC

[CE17] Shiyang Chen and Bogdan Epureanu. Regular biennial cycles in epidemics caused by parametric resonance. Journal of Theoretical Biology, 415(??):137–144, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304246.

Calle-Espinosa:2016:NLW

[CEdLSG^{+}16] Jorge Calle-Espinosa, Miguel Ponce de Leon, Diego Santos-Garcia, Francisco J. Silva, Francisco Montero, and Juli Peretó. Nature lessons: The whitefly bacterial endosymbiont is a minimal amino acid factory with unusual energetics. Journal of Theoretical Biology, 407(??):303–317, October 21, 2016.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302077.

Cusseddu:2019:CBS

[CEKM^{+}19] D. Cusseddu, L. Edelstein-Keshet, J. A. Mackenzie, S. Portet, and A. Madzvamuse. A coupled bulk-surface model for cell polarisation. Journal of Theoretical Biology, 481(??):119– 135, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519318304387.

Crouzy:2014:BPA

[CEP14] Benoit Crouzy, Katharina Edmaier, and Paolo Perona. Biomechanics of plant anchorage at early development stage. Journal of Theoretical Biology, 363(??):22–29, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004469.

Cherry:2011:EBG

[CF11] Elizabeth M. Cherry and Flavio H. Fenton. Effects of boundaries and geometry on the spatial distribution of action potential duration in cardiac tissue. Journal of Theoretical Biology, 285(1):164–176, September 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003444.

Calzada:2013:OCO

[CFCM13] M. Carmen Calzada, Enrique Fernández-Cara, and Mercedes Marín. Optimal control oriented to therapy for a free-boundary tumor growth model. Journal of Theoretical Biology, 325(??):1–11, May 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000647.

Chow:2011:DME

[CFF11] Lamwah Chow, Meng Fan, and Zhilan Feng. Dynamics of a multigroup epidemiological model with grouptargeted vaccination strategies. Journal of Theoretical Biology, 291(??):56–64, December 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004760.

Catacuzzeno:2012:TSR

[CFF12] Luigi Catacuzzeno, Bernard Fioretti, and Fabio Franciolini. A theoretical study on the role of Ca^{2+}-activated K^{+} channels in the regulation of hormone-induced Ca^{2+} oscillations and their synchronization in adjacent cells. Journal of Theoretical Biology, 309(??):103–112, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002354.

Cherubini:2017:NSD

[CFGRB17] Christian Cherubini, Simonetta Filippi, Alessio Gizzi, and Ricardo Ruiz-Baier. A note on stress-driven anisotropic diffusion and its role in active deformable media. Journal of Theoretical Biology, 430(??):221–228, October 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303417.

Chen:2015:DTL

[CFL^{+}15] Ming Chen, Meng Fan, Rui Liu, Xiaoyu Wang, Xing Yuan, and Huaiping Zhu. The dynamics of temperature and light on the growth of phytoplankton. Journal of Theoretical Biology, 385(??):8–19, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004014.

Cook:2013:EVE

[CFR13] Zoe Cook, Daniel W. Franks, and Elva J. H. Robinson. Exploration versus exploitation in polydomous ant colonies. Journal of Theoretical Biology, 323(??):49–56, April 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000477.

Crompton:2014:IDM

[CFR^{+}14] L. A. Crompton, J. France, C. K. Reynolds, J. A. N. Mills, M. D. Hanigan, J. L. Ellis, A. Bannink, B. J. Bequette, and J. Dijkstra. An isotope dilution model for partitioning phenylalanine and tyrosine uptake by the mammary gland of lactating dairy cows. Journal of Theoretical Biology, 359(??):54–60, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002835.

Caldwell:2019:VAP

[CFS^{+}19] Wendy K. Caldwell, Benjamin Freedman, Luke Settles, Michael M. Thomas, Erika T. Camacho, and Stephen Wirkus. The Vicodin abuse problem: a mathematical approach. Journal of Theoretical Biology, 483(??):Article 110003, December 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930373X.

Cencini:2016:CFG

[CFSB16] M. Cencini, M. Franchino, F. Santamaria, and G. Boffetta. Centripetal focusing of gyrotactic phytoplankton. Journal of Theoretical Biology, 399(??):62–70, June 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300133.

Centler:2011:MPP

[CFT11] Florian Centler, Ingo Fetzer, and Martin Thullner. Modeling population patterns of chemotactic bacteria in homogeneous porous media. Journal of Theoretical Biology, 287(??):82–91, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003742.

Cen:2014:EDD

[CFZ14] Xiuli Cen, Zhilan Feng, and Yulin Zhao. Emerging disease dynamics in a model coupling within-host and betweenhost systems. Journal of Theoretical Biology, 361(??):141– 151, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519314004421.

Crane:2010:CMA

[CG10] Kenneth W. Crane and James P. Grover. Coexistence of mixotrophs, autotrophs, and heterotrophs in planktonic microbial communities. Journal of Theoretical Biology, 262(3):517–527, February 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005165.

Cunniffe:2011:TFB

[CG11] Nik J. Cunniffe and Christopher A. Gilligan. A theoretical framework for biological control of soil-borne plant pathogens: Identifying effective strategies. Journal of Theoretical Biology, 278(1):32–43, June 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001251.

Collins:2014:IHT

[CG14] O. C. Collins and K. S. Govinder. Incorporating heterogeneity into the transmission dynamics of a waterborne disease model. Journal of Theoretical Biology, 356(??):133–143, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002422.

Crona:2013:PGF

[CGB13] Kristina Crona, Devin Greene, and Miriam Barlow. The peaks and geometry of fitness landscapes. Journal of Theoretical Biology, 317(??):1–10, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005061.

Chen:2013:SRO

[CGD13] Xiaojie Chen, Thilo Gross, and Ulf Dieckmann. Shared rewarding overcomes defection traps in generalized volunteer’s dilemmas. Journal of Theoretical Biology, 335(??):13–21, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002865.

Chivers:2014:PPS

[CGHF14] W. J. Chivers, W. Gladstone, R. D. Herbert, and M. M. Fuller. predator–prey systems depend on a prey refuge. Journal of Theoretical Biology, 360(??):271–278, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004135.

Corredor-Gomez:2016:IOM

[CGRRGM^{+}16] Jennifer Paola Corredor-Gómez, Andrés Mauricio Rueda-Ramírez, Miguel Alejandro Gamboa-Márquez, Carolina Torres-Rodríguez, and Carlos Julio Cortés-Rodríguez. An intramembranous ossification model for the in silico analysis of bone tissue formation in tooth extraction sites. Journal of Theoretical Biology, 401(??):64–77, July 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300510.

Carlier:2015:OCD

[CGvG^{+}15] Aurélie Carlier, Liesbet Geris, Nick van Gastel, Geert Carmeliet, and Hans Van Oosterwyck. Oxygen as a critical determinant of bone fracture healing — a multiscale model. Journal of Theoretical Biology, 365(??):247–264, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006018. See corri-

gendum [Van16b].

Cui:2018:ESD

[CGW18] Peng-Bi Cui, Lei Gao, and Wei Wang. Epidemic spreading dynamics with drug resistance and heterogeneous contacts. Journal of Theoretical Biology, 441(??):19–27, March 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300031.

Chen:2011:SRP

[CH11] Yu-Yun Chen and Sze-Bi Hsu. Synchronized reproduction promotes species coexistence through reproductive facilitation. Journal of Theoretical Biology, 274(1):136–144, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000257.

Cheng:2016:SST

[CH16] Zhang Cheng and Alexander Hoffmann. A stochastic spatiotemporal (SST) model to study cell-to-cell variability in HIV-1 infection. Journal of Theoretical Biology, 395(??):87–96, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931600076X.

Cooper:2018:ISD

[CH18] A. J. Cooper and T. Déirdre Hollingsworth. The impact of seasonality on the dynamics and control of Ascaris lumbricoides infections. Journal of Theoretical Biology, 453(??):96–107, September 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302649.

Chakravarty:2017:TSB

[Cha17] Ashim K. Chakravarty, FNASc. Thermodynamic stability of biomolecules and evolution. Journal of Theoretical Biology, 427(??):8–9, August 1, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302564.

Chamchod:2018:MSC

[Cha18] Farida Chamchod. Modeling the spread of capripoxvirus among livestock and optimal vaccination strategies. Journal of Theoretical Biology, 437(??):179–186, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304721.

Callender:2010:MSS

[CHD^{+}10] Hannah L. Callender, Mary Ann Horn, Dianne L. De-Camp, Paul C. Sternweis, and H. Alex Brown. Modeling species-specific diacylglycerol dynamics in the RAW 264.7 macrophage. Journal of Theoretical Biology, 262(4):679–690, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005190.

Cheung:2010:FMR

[Che10] Allen Cheung. The fourth moment of the radial displacement of a discrete correlated/persistent random walk. Journal of Theoretical Biology, 264(2):641–644, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000113X.

Chen:2012:MAP

[Che12] Frederick Chen. A mathematical analysis of public avoidance behavior during epidemics using game theory. Journal of Theoretical Biology, 302(??):18–28, June 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001191.

Cheung:2014:API

[Che14] Allen Cheung. Animal path integration: a model of positional uncertainty along tortuous paths. Journal of Theoretical Biology, 341(??):17–33, January 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004773.

Chen:2016:HRP

[Che16] Kang Ping Chen. Hydraulic resistance of a plant root to water-uptake: a slender-body theory. Journal of Theoretical Biology, 396(??):63–74, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931600120X.

Chen:2018:VVV

[Che18] Frederick Chen. Voluntary vaccinations and vaccine shortages: a theoretical analysis. Journal of Theoretical Biology, 446(??):19–32, June 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300924.

Chen:2019:FTT

[Che19] Zhan Chen. The formation of the thickveins (Tkv) gradient in Drosophila wing discs: a theoretical study. Journal of Theoretical Biology, 474(??):25–41, August 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301560.

Chan:2010:FEF

[CHH10] Yue Chan, Richard G. Haverkamp, and James M. Hill. Force-extension formula for the worm-like chain model from a variational principle. Journal of Theoretical Biology, 262(3):498–504, February 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004846.

Chan:2016:ELA

[CHK16] Matthew H. Chan, Kristen Hawkes, and Peter S. Kim. Evolution of longevity, age at last birth and sexual conflict with grandmothering. Journal of Theoretical Biology, 393(??):145–157, March 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000151.

Chisholm:2011:WCO

[CHL^{+}11] Rebecca H. Chisholm, Barry D. Hughes, Kerry A. Landman, Georg Mayer, and Paul M. Whitington. When are cellular oscillators sufficient for sequential segmentation? Journal of Theoretical Biology, 279(1):150–160, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100124X.

Cai:2019:ELN

[CHL^{+}19] Rui Cai, Ziying He, Yancai Liu, Jinqiao Duan, Jürgen Kurths, and Xiaofan Li. Effects of Lévy noise on the Fitzhugh– Nagumo model: a perspective on the maximal likely trajectories. Journal of Theoretical Biology, 480(??):166–174, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303224.

Craig:2015:NDD

[CHN^{+}15] Morgan Craig, Antony R. Humphries, Fahima Nekka, Jacques Bélair, Jun Li, and Michael C. Mackey. Neutrophil dynamics during concurrent chemotherapy and g-CSF administration: Mathematical modelling guides dose optimisation to minimise neutropenia. Journal of Theoretical Biology, 385(??):77–89, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004087.

Chou:2011:SRP

[Cho11] Kuo-Chen Chou. Some remarks on protein attribute prediction and pseudo amino acid composition. Journal of Theoretical Biology, 273(1):236–247, March 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000679X.

Chu:2010:TSL

[Chu10] C. Y. Cyrus Chu. A two-sex life history model of handicap signaling. Journal of Theoretical Biology, 263(2):219–226, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900561X.

Chu:2017:LSM

[Chu17] Dominique Chu. Limited by sensing — a minimal stochastic model of the lag-phase during diauxic growth. Journal of Theoretical Biology, 414(??):137–146, February 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303551.

Chu:2018:PLT

[Chu18] Dominique Chu. Performance limits and trade-offs in entropy-driven biochemical computers. Journal of Theoretical Biology, 443(??):1–9, April 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300389.

Caudron:2012:SCA

[CJ12a] B. Caudron and J. L. Jestin. Sequence criteria for the antiparallel character of protein beta-strands. Journal of Theoretical Biology, 315(??):146–149, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004894.

Chen:2012:NSA

[CJ12b] Bo Chen and Ping Ji. Numericalization of the self adaptive spectral rotation method for coding region prediction. Journal of Theoretical Biology, 296(??):95–102, March 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006023.

Caliebe:2010:MCD

[CJKR10] Amke Caliebe, Arne Jochens, Michael Krawczak, and Uwe Rösler. A Markov chain description of the stepwise mutation model: Local and global behaviour of the allele process. Journal of Theoretical Biology, 266(2):336–342, September 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003243.

Choi:2015:OIS

[CJL15] Sunhwa Choi, Eunok Jung, and Seok-Min Lee. Optimal intervention strategy for prevention tuberculosis using a smoking-tuberculosis model. Journal of Theoretical Biology, 380(??):256–270, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500257X.

Chen:2011:PAE

[CJRR11] Frederick Chen, Miaohua Jiang, Scott Rabidoux, and Stephen Robinson. Public avoidance and epidemics: Insights from an economic model. Journal of Theoretical Biology, 278(1):107– 119, June 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001366.

Chifman:2015:IUS

[CK15] Julia Chifman and Laura Kubatko. Identifiability of the unrooted species tree topology under the coalescent model with time-reversible substitution processes, site-specific rate variation, and invariable sites. Journal of Theoretical Biology, 374(??):35–47, June 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001095.

Chaley:2016:SMH

[CK16] Maria Chaley and Vladimir Kutyrkin. Stochastic model of homogeneous coding and latent periodicity in DNA sequences. Journal of Theoretical Biology, 390(??):106–116, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315005640.

Clegg:2017:RDB

[CK17] Robert J. Clegg and Jan-Ulrich Kreft. Reducing discrepancies between 3D and 2D simulations due to cell packing density. Journal of Theoretical Biology, 423(??):26–30, June 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301807.

Choi:2018:PPE

[CK18] Jieun Choi and Yong-Jung Kim. predator–prey equations with constant harvesting and planting. Journal of Theoretical Biology, 458(??):47–57, December 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304235.

Cressman:2019:BGI

[CK19] Ross Cressman and Vlastimil Křivan. Bimatrix games that include interaction times alter the evolutionary outcome: The Owner–Intruder game. Journal of Theoretical Biology, 460(??):262–273, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305095.

Chenar:2018:MMI

[CKB18] F. Fatehi Chenar, Y. N. Kyrychko, and K. B. Blyuss. Mathematical model of immune response to hepatitis B. Journal of Theoretical Biology, 447(??):98–110, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301395.

Cortes:2019:OSP

[CKB19] Michael G. Cortes, Jonathan Krog, and Gábor Balázsi. Optimality of the spontaneous prophage induction rate. Journal of Theoretical Biology, 483(??):Article 110005, December 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303753.

Curtis:2012:MTM

[CKBCG12] M. P. Curtis, J. C. Kirkman-Brown, T. J. Connolly, and E. A. Gaffney. Modelling a tethered mammalian sperm cell undergoing hyperactivation. Journal of Theoretical Biology, 309(??):1–10, September 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002779.

Chavarria-Krauser:2011:MPM

[CKD11] Andrés Chavarría-Krauser and Du Yejie. A model of plasma membrane flow and cytosis regulation in growing pollen tubes. Journal of Theoretical Biology, 285(1):10–24, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311003079.

Chen:2017:SGR

[CKF17] Xi Chen, Yan-Mei Kang, and Yu-Xuan Fu. Switches in a genetic regulatory system under multiplicative non-Gaussian noise. Journal of Theoretical Biology, 435(??):134–144, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304265.

Chifman:2012:CCS

[CKN^{+}12] J. Chifman, A. Kniss, P. Neupane, I. Williams, B. Leung, Z. Deng, P. Mendes, V. Hower, F. M. Torti, S. A. Akman, S. V. Torti, and R. Laubenbacher. The core control system of intracellular iron homeostasis: a mathematical model. Journal of Theoretical Biology, 300(??):91–99, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000343.

Chu:2019:ERE

[CKNB19] Sherry Chu, Mehran Kardar, David R. Nelson, and Daniel A. Beller. Evolution in range expansions with competition at rough boundaries. Journal of Theoretical Biology, 478(??):153–160, October 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931930253X.

Czaran:2015:MCR

[CKS15] Tamás Czárán, Balázs Könnyű, and Eörs Szathmáry. Metabolically coupled replicator systems: Overview of an RNA-world model concept of prebiotic evolution on mineral surfaces. Journal of Theoretical Biology, 381(??):39– 54, September 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519315002817.

Cao:2017:MMQ

[CKZ^{+}17] Pengxing Cao, Nectarios Klonis, Sophie Zaloumis, David S. Khoury, Deborah Cromer, Miles P. Davenport, Leann Tilley, Julie A. Simpson, and James M. McCaw. A mechanistic model quantifies artemisinin-induced parasite growth retardation in blood-stage Plasmodium falciparum infection. Journal of Theoretical Biology, 430(??):117–127, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730348X.

Cibert:2010:CFI

[CL10] Christian Cibert and Andrei Ludu. Is the curvature of the flagellum involved in the apparent cooperativity of the dynein arms along the “9+2” axoneme? Journal of Theoretical Biology, 265(2):95–103, July 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001876.

Chechetkin:2011:SGC

[CL11] V. R. Chechetkin and V. V. Lobzin. Stability of the genetic code and optimal parameters of amino acids. Journal of Theoretical Biology, 269(1):57–63, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005436.

Chen:2013:PMP

[CL13a] Yen-Kuang Chen and Kuo-Bin Li. Predicting membrane protein types by incorporating protein topology, domains, signal peptides, and physicochemical properties into the general form of Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 318(??):1–12, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005735.

Cyrus:2013:EID

[CL13b] Chu C. Y. Cyrus and Ronald D. Lee. On the evolution of intergenerational division of labor, menopause and transfers among adults and offspring. Journal of Theoretical Biology, 332(??):171–180, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001999.

Chen:2014:TGC

[CL14] Ying Chen and John S. Lowengrub. Tumor growth in complex, evolving microenvironmental geometries: a diffuse domain approach. Journal of Theoretical Biology, 361(??):14–30, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003671.

Caudill:2017:UIH

[CL17a] Lester Caudill and Barry Lawson. A unified inter-host and in-host model of antibiotic resistance and infection spread in a hospital ward. Journal of Theoretical Biology, 421(??):112– 126, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301443.

Chechetkin:2017:LSC

[CL17b] V. R. Chechetkin and V. V. Lobzin. Large-scale chromosome folding versus genomic DNA sequences: a discrete double Fourier transform technique. Journal of Theoretical Biology, 426(??):162–179, August 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302539.

Cho:2018:MCI

[CL18] H. Cho and D. Levy. Modeling the chemotherapy-induced selection of drug-resistant traits during tumor growth. Journal of Theoretical Biology, 436(??):120–134, January 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304654.

Chen:2019:TGC

[CL19] Ying Chen and John S. Lowengrub. Tumor growth and calcification in evolving microenvironmental geometries. Journal of Theoretical Biology, 463(??):138–154, February 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305964.

Coelho:2016:DMB

[CLA^{+}16] Rui Moura Coelho, João Miranda Lemos, Irina Alho, Duarte Valério, Arlindo R. Ferreira, Luís Costa, and Susana Vinga. Dynamic modeling of bone metastasis, microenvironment and therapy: Integrating parathyroid hormone (PTH) effect, anti-resorptive and anti-cancer therapy. Journal of Theoretical Biology, 391(??):1–12, February 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005755.

Cazaubiel:2017:CSC

[CLA17] Annette Cazaubiel, Alessandra F. Lütz, and Jeferson J. Arenzon. Collective strategies and cyclic dominance in asymmetric predator–prey spatial games. Journal of Theoretical Biology, 430(??):45–52, October 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303272.

Chen:2011:IBD

[CLC11] Qingfeng Chen, Gang Li, and Yi-Ping Phoebe Chen. Intervalbased distance function for identifying RNA structure candidates. Journal of Theoretical Biology, 269(1):280–286, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005874.

Chung:2010:HCA

[CLCH10] C. A. Chung, Tze-Hung Lin, Shih-Di Chen, and Hsing-I Huang. Hybrid cellular automaton modeling of nutrient modulated cell growth in tissue engineering constructs. Journal of Theoretical Biology, 262(2):267–278, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900469X.

Liu:2010:QIS

[cLCJ^{+}10] Wei chung Liu, Hsuan-Wien Chen, Ferenc Jordán, Wen-Hsieh Lin, and Chester Wai-Jen Liu. Quantifying the interaction structure and the topological importance of species in food webs: a signed digraph approach. Journal of Theoretical Biology, 267(3):355–362, December 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004583.

Cleaves:2010:OBC

[Cle10] H. James Cleaves II. The origin of the biologically coded amino acids. Journal of Theoretical Biology, 263(4):490–498, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005815.

Clerc:2016:NEE

[Cle16] Daryl G. Clerc. Nonlinear effects in evolution — an ab initio study: a model in which the classical theory of evolution occurs as a special case. Journal of Theoretical Biology, 401(??):94–108, July 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001739.

Cleather:2018:IRB

[Cle18a] Daniel J. Cleather. An important role of the biarticular hamstrings is to exert internal/external rotation moments on the tibia during vertical jumping. Journal of Theoretical Biology, 455(??):101–108, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303369.

Cleather:2018:PMD

[Cle18b] Daniel John Cleather. The patella: a mechanical determinant of coordination during vertical jumping. Journal of Theoretical Biology, 446(??):205–211, June 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301279.

Cardenas:2010:CEC

[CLG^{+}10] María Luz Cárdenas, Juan-Carlos Letelier, Claudio Gutierrez, Athel Cornish-Bowden, and Jorge Soto-Andrade. Closure to efficient causation, computability and artificial life. Journal of Theoretical Biology, 263(1):79–92, March 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519309005360. See corrigendum [CLG^{+}11].

Cardenas:2011:CCE

[CLG^{+}11] María Luz Cárdenas, Juan-Carlos Letelier, Claudio Gutierrez, Athel Cornish-Bowden, and Jorge Soto-Andrade. Corrigendum to “Closure to efficient causation, computability and artificial life” [J. Theor. Biol. 263 (2010) 79–92]. Journal of Theoretical Biology, 274(1):186, April 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000427. See [CLG^{+}10].

Chauhan:2011:RMC

[CLHB11] Anuradha Chauhan, Stephan Lorenzen, Hanspeter Herzel, and Samuel Bernard. Regulation of mammalian cell cycle progression in the regenerating liver. Journal of Theoretical Biology, 283(1):103–112, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002682.

Chen:2018:MEC

[CLK18] Can Chen, Chentong Li, and Yanmei Kang. Modelling the effects of cutting off infected branches and replanting on fireblight transmission using Filippov systems. Journal of Theoretical Biology, 439(??):127–140, February 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305246.

Chen:2018:UIT

[CLL18] Weiyang Chen, Bo Liao, and Weiwei Li. Use of image texture analysis to find DNA sequence similarities. Journal of Theoretical Biology, 455(??):1–6, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830328X.

Chyou:2011:UBC

[CLP11] T. Chyou, G. F. Liddell, and M. G. Paulin. An upper-body can improve the stability and efficiency of passive dynamic walking. Journal of Theoretical Biology, 285(1):126–135, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003377.

Cardozo:2014:SMH

[CLPZ14] E. Fabian Cardozo, Rutao Luo, Michael J. Piovoso, and Ryan Zurakowski. Spatial modeling of HIV cryptic viremia and 2-LTR formation during raltegravir intensification. Journal of Theoretical Biology, 345(??):61–69, March 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005717.

Cohen:2012:BSA

[CLW12] Andrew L. Cohen, Tanya L. Leise, and David K. Welsh. Bayesian statistical analysis of circadian oscillations in fibroblasts. Journal of Theoretical Biology, 314(??):182–191, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004602.

Chakrabarty:2012:MHC

[CM12] Siddhartha P. Chakrabarty and John M. Murray. Modelling hepatitis C virus infection and the development of hepatocellular carcinoma. Journal of Theoretical Biology, 305(??):24–29, July 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001622.

Constable:2014:PGI

[CM14] George W. A. Constable and Alan J. McKane. Population genetics on islands connected by an arbitrary network: An analytic approach. Journal of Theoretical Biology, 358(??):149–165, October 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003063.

Cuesta:2017:ESS

[CM17] Jose A. Cuesta and Susanna Manrubia. Enumerating secondary structures and structural moieties for circular RNAs. Journal of Theoretical Biology, 419(??):375–382, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300796.

Carteni:2012:NPS

[CMB^{+}12] Fabrizio Cartenì, Addolorata Marasco, Giuliano Bonanomi, Stefano Mazzoleni, Max Rietkerk, and Francesco Giannino. Negative plant soil feedback explaining ring formation in clonal plants. Journal of Theoretical Biology, 313(??):153–161, November 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003943.

Chaput:2019:HBC

[CMBP19] Romain Chaput, John E. Majoris, Peter M. Buston, and Claire B. Paris. Hydrodynamic and biological constraints on group cohesion in plankton. Journal of Theoretical Biology, 482(??):Article 109987, December 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303303.

Cenci:2018:EER

[CMCS18] Simone Cenci, Ana Montero-Castaño, and Serguei Saavedra. Estimating the effect of the reorganization of interactions on the adaptability of species to changing environments. Journal of Theoretical Biology, 437(??):115–125, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304794.

Chiyaka:2010:TAM

[CMD^{+}10] Christinah Chiyaka, Zindoga Mukandavire, Prasenjit Das, Farai Nyabadza, Senelani D. Hove-Musekwa, and Henry Mwambi. Theoretical analysis of mixed Plasmodium malariae and Plasmodium falciparum infections with partial crossimmunity. Journal of Theoretical Biology, 263(2):169–178, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005219.

Cunning:2017:DBM

[CMGN17] Ross Cunning, Erik B. Muller, Ruth D. Gates, and Roger M. Nisbet. A dynamic bioenergetic model for coral-Symbiodinium symbioses and coral bleaching as an alternate stable state. Journal of Theoretical Biology, 431(??):49–62, October 27, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303636.

Chowkwale:2019:MSM

[CMHM19] M. Chowkwale, G. J. Mahler, P. Huang, and B. T. Murray. A multiscale in silico model of endothelial to mesenchymal transformation in a tumor microenvironment. Journal of Theoretical Biology, 480(??):229–240, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303248.

Coutand:2011:TMS

[CMJD11] Catherine Coutand, Jean-Denis Mathias, Georges Jeronimidis, and Jean-François Destrebecq. TWIG: a model to simulate the gravitropic response of a tree axis in the frame of elasticity and viscoelasticity, at intra-annual time scale. Journal of Theoretical Biology, 273(1):115–129, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931000682X.

Campos:2010:PRM

[CML10] Daniel Campos, Vicenç Méndez, and Isaac Llopis. Persistent random motion: Uncovering cell migration dynamics. Journal of Theoretical Biology, 267(4):526–534, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310004960.

Comisar:2011:IOL

[CML11] W. A. Comisar, D. J. Mooney, and J. J. Linderman. Integrin organization: Linking adhesion ligand nanopatterns with altered cell responses. Journal of Theoretical Biology, 274(1):120–130, April 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000191.

Corkrey:2019:BST

[CMM19] Ross Corkrey, Cameron Macdonald, and Tom McMeekin. The biokinetic spectrum for temperature and optimal Darwinian fitness. Journal of Theoretical Biology, 462(??):171–183, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305356.

Corbacho:2013:BDB

[CMMR13] Casimiro Corbacho, Manuel Molina, Manuel Mota, and Alfonso Ramos. Birth–death branching models. Application to African elephant populations. Journal of Theoretical Biology, 332(??):108–116, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001987.

Courcoul:2011:MEH

[CMN^{+}11] Aurélie Courcoul, Hervé Monod, Mirjam Nielen, Don Klinkenberg, Lenny Hogerwerf, François Beaudeau, and Elisabeta Vergu. Modelling the effect of heterogeneity of shedding on the within herd Coxiella burnetii spread and identification of key parameters by sensitivity analysis. Journal of Theoretical Biology, 284(1):130–141, September 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100316X.

Chaumont:2017:RPU

[CMPS17] Loïc Chaumont, Valéry Malécot, Richard Pymar, and Chaker Sbai. Reconstructing pedigrees using probabilistic analysis of ISSR amplification. Journal of Theoretical Biology, 412(??):8–16, January 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303022.

Crompton:2018:IDM

[CMR^{+}18] L. A. Crompton, L. L. McKnight, C. K. Reynolds, J. A. N. Mills, J. L. Ellis, M. D. Hanigan, J. Dijkstra, B. J. Bequette, A. Bannink, and J. France. An isotope dilution model for partitioning of phenylalanine and tyrosine uptake by the liver of lactating dairy cows. Journal of Theoretical Biology, 444(??):100–107, May 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305556.

Crowell:2016:FMC

[CMS16] Helena L. Crowell, Adam L. MacLean, and Michael P. H. Stumpf. Feedback mechanisms control coexistence in a stem cell model of acute myeloid leukaemia. Journal of Theoretical Biology, 401(??):43–53, July 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300212.

Carrillo:2019:PDM

[CMS^{+}19] Jose A. Carrillo, Hideki Murakawa, Makoto Sato, Hideru Togashi, and Olena Trush. A population dynamics model of cell–cell adhesion incorporating population pressure and density saturation. Journal of Theoretical Biology, 474(??):14–24, August 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301730.

Cangiani:2010:SMC

[CN10] A. Cangiani and R. Natalini. A spatial model of cellular molecular trafficking including active transport along microtubules. Journal of Theoretical Biology, 267(4):614–625, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004261.

Cha:2012:SSS

[CN12a] Chae Young Cha and Akinori Noma. Steady-state solutions of cell volume in a cardiac myocyte model elaborated for membrane excitation, ion homeostasis and Ca^{2+} dynamics. Journal of Theoretical Biology, 307(??):70–81, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312002068.

Cropp:2012:CE

[CN12b] Roger Cropp and John Norbury. Constructing ecologies. Journal of Theoretical Biology, 294(??):1–8, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005510.

Cooper:2015:TMS

[CNDK15] Jacob D. Cooper, Claudia Neuhauser, Antony M. Dean, and Benjamin Kerr. Tipping the mutation-selection balance: Limited migration increases the frequency of deleterious mutants.

Journal of Theoretical Biology, 380(??):123–133, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002313.

Cordoleani:2012:CSS

[CNG^{+}12] Flora Cordoleani, David Nerini, Mathias Gauduchon, Andrew Morozov, and Jean-Christophe Poggiale. Corrigendum to “Structural sensitivity of biological models revisited” [J. Theor. Biol. 283 (2011) 82–91]. Journal of Theoretical Biology, 300(??):376, May 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006497. See [FDM^{+}11].

Cheeseman:2014:STD

[CNL14] Bevan L. Cheeseman, Donald F. Newgreen, and Kerry A. Landman. Spatial and temporal dynamics of cell generations within an invasion wave: a link to cell lineage tracing. Journal of Theoretical Biology, 363(??):344–356, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004706.

Cordoleani:2013:SPF

[CNM^{+}13] Flora Cordoleani, David Nerini, Andrey Morozov, Mathias Gauduchon, and Jean-Christophe Poggiale. Scaling up the predator functional response in heterogeneous environment: When Holling type III can emerge? Journal of Theoretical Biology, 336(??):200–208, November 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003342.

Chaves:2019:AGM

[COG19] Madalena Chaves, Diego A. Oyarzún, and Jean-Luc Gouzé. Analysis of a genetic-metabolic oscillator with piecewise linear models. Journal of Theoretical Biology, 462(??):259–269, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305022.

Cohn:2015:ARP

[Coh15a] Melvin Cohn. Autoimmunity: Rationalizing possible pathways from initiation to disease. Journal of Theoretical Biology, 375(??):40–51, June 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003038.

Cohn:2015:TUP

[Coh15b] Melvin Cohn. Two unresolved problems facing models of the Self–Nonself discrimination. Journal of Theoretical Biology, 387(??):31–38, December 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004646.

Cooper:2016:FRE

[Coo16] William E. Cooper. Fleeing to refuge: Escape decisions in the race for life. Journal of Theoretical Biology, 406(??):129–136, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301576.

Cortez:2016:HED

[Cor16] Michael H. Cortez. Hydra effects in discrete-time models of stable communities. Journal of Theoretical Biology, 411(??):59–67, December 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303198.

Cortes:2018:HRV

[Cor18] Andrés J. Cortés. On how role versatility boosts an STI. Journal of Theoretical Biology, 440(??):66–69, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305568.

Cao:2011:ECA

[COWA11] Lang Cao, Hisashi Ohtsuki, Bing Wang, and Kazuyuki Aihara. Evolution of cooperation on adaptively weighted networks. Journal of Theoretical Biology, 272(1):8–15, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006636.

Cox:2010:MSD

[Cox10] Brian Cox. A multi-scale, discrete-cell simulation of organogenesis: Application to the effects of strain stimulus on collective cell behavior during ameloblast migration. Journal of Theoretical Biology, 262(1):58–72, January 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004305.

Cobey:2011:CHH

[CP11] Sarah Cobey and Mercedes Pascual. Consequences of host heterogeneity, epitope immunodominance, and immune breadth for strain competition. Journal of Theoretical Biology, 270(1):80–87, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005941.

Caceres:2014:BBE

[CP14] María J. Cáceres and Benoît Perthame. Beyond blowup in excitatory integrate and fire neuronal networks: Refractory period and spontaneous activity. Journal of Theoretical Biology, 350(??):81–89, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000666.

Carloni:2013:HDM

[CPF^{+}13] A. Carloni, V. Poletti, L. Fermo, N. Bellomo, and M. Chilosi. Heterogeneous distribution of mechanical stress in human lung: a mathematical approach to evaluate abnormal remodeling in IPF. Journal of Theoretical Biology, 332(??):136–140, September 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002130.

Carbonell-Pascual:2016:CBH

[CPGF^{+}16] Beatriz Carbonell-Pascual, Eduardo Godoy, Ana Ferrer, Lucia Romero, and Jose M. Ferrero. Comparison between Hodgkin– Huxley and Markov formulations of cardiac ion channels. Journal of Theoretical Biology, 399(??):92–102, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300157.

Chen:2014:EHA

[CPH^{+}14] Xiao Chen, Qinke Peng, Libin Han, Tao Zhong, and Tao Xu. An effective haplotype assembly algorithm based on hypergraph partitioning. Journal of Theoretical Biology, 358(??):85–92, October 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003178.

Cortes-Poza:2018:SDF

[CPPLAB18] Yuriria Cortes-Poza, Pablo Padilla-Longoria, and Elena Alvarez-Buylla. Spatial dynamics of floral organ formation. Journal of Theoretical Biology, 454(??):30–40, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318302790.

Chara:2010:NFE

[CPS10] Osvaldo Chara, Diego E. Pafundo, and Pablo J. Schwarzbaum. Negative feedback of extracellular ADP on ATP release in goldfish hepatocytes: a theoretical study. Journal of Theoretical Biology, 264(4):1147–1158, June 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001451.

Calvetti:2019:EHS

[CPS19] D. Calvetti, J. Prezioso, and E. Somersalo. Estimating hemodynamic stimulus and blood vessel compliance from cerebral blood flow data. Journal of Theoretical Biology, 460(??):243–261, January 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304818.

Chavez:2016:MIP

[CPV16] Vasthi Alonso Chavez, Stephen Parnell, and Frank Van den Bosch. Monitoring invasive pathogens in plant nurseries for early-detection and to minimise the probability of escape. Journal of Theoretical Biology, 407(??):290–302, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302247.

Camacho:2016:QMC

[CPW16] Erika T. Camacho, Claudio Punzo, and Stephen A. Wirkus. Quantifying the metabolic contribution to photoreceptor death in retinitis pigmentosa via a mathematical model. Journal of Theoretical Biology, 408(??):75–87, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302272.

Calenge:2014:BMH

[CR14] Clément Calenge and Sophie Rossi. Bayesian modelling of hunting data may improve the understanding of host-parasite systems: Wild boar diseases and vaccination as an example. Journal of Theoretical Biology, 343(??):32–43, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005389.

Coons:2016:NPI

[CR16] Jane Ivy Coons and Joseph Rusinko. A note on the path interval distance. Journal of Theoretical Biology, 398(??):145–149, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001740.

Chang:2019:MNU

[CR19] Shyr-Shea Chang and Marcus Roper. Microvascular networks with uniform flow. Journal of Theoretical Biology, 462(??):48–64, February 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305319.

Caiazzo:2015:MMP

[CRC15] Alfonso Caiazzo and Ignacio Ramis-Conde. Multiscale modelling of palisade formation in gliobastoma multiforme. Journal of Theoretical Biology, 383(??):145–156, October 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003586.

Cressman:2010:CND

[Cre10] Ross Cressman. CSS, NIS and dynamic stability for twospecies behavioral models with continuous trait spaces. Journal of Theoretical Biology, 262(1):80–89, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004391.

Calvetti:2018:CMI

[CRGS18] D. Calvetti, G. Capo Rangel, L. Gerardo Giorda, and E. Somersalo. A computational model integrating brain electrophysiology and metabolism highlights the key role of extracellular potassium and oxygen. Journal of Theoretical Biology, 446(??):238–258, June 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300900.

Carlos:2015:SEA

[CRH^{+}15] Perla Carlos, Virginie Roupie, Sébastien Holbert, Felipe Ascencio, Kris Huygen, Gracia Gomez-Anduro, Maxime Branger, Martha Reyes-Becerril, and Carlos Angulo. In silico epitope analysis of unique and membrane associated proteins from mycobacterium avium subsp. paratuberculosis for immunogenicity and vaccine evaluation. Journal of Theoretical Biology, 384(??):1–9, November 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003811.

Calleja-Rodriguez:2019:APE

[CRLH^{+}19] Ainhoa Calleja-Rodriguez, Zitong Li, Henrik R. Hallingbäck, Mikko J. Sillanpää, Harry X. Wu, Sara Abrahamsson, and Maria Rosario García-Gil. Analysis of phenotypic- and Estimated Breeding Values (EBV) to dissect the genetic architecture of complex traits in a Scots pine three-generation pedigree design. Journal of Theoretical Biology, 462(??):283–292, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931830554X.

Crouch:2017:SAE

[Cro17] Daniel J. M. Crouch. Statistical aspects of evolution under natural selection, with implications for the advantage of sexual reproduction. Journal of Theoretical Biology, 431(??):79–86, October 27, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303521.

Cronin:2018:ABM

[Cro18] Adam L. Cronin. An agent-based model of nest-site selection in a mass-recruiting ant. Journal of Theoretical Biology, 455(??):54–63, October 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303175.

Cronin:2019:FGT

[Cro19] Adam L. Cronin. Fussy groups thwart the collective burden of choice: a theoretical study of house-hunting ants. Journal of Theoretical Biology, 483(??):Article 110000, December 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303698.

Comper:2016:FPP

[CRV16] W. D. Comper, L. M. Russo, and J. Vuchkova. Are filtered plasma proteins processed in the same way by the kidney? Journal of Theoretical Biology, 410(??):18–24, December 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303010.

Calvetti:2011:DAM

[CS11a] Daniela Calvetti and Erkki Somersalo. Dynamic activation model for a glutamatergic neurovascular unit. Journal of Theoretical Biology, 274(1):12–29, April 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006624.

Chakra:2011:HCM

[CS11b] Maria Abou Chakra and Jonathon Richard Stone. Holotestoid: a computational model for testing hypotheses about echinoid skeleton form and growth. Journal of Theoretical Biology, 285(1):113–125, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003183.

Campenni:2014:PCP

[CS14a] Marco Campennì and Gabriele Schino. Partner choice promotes cooperation: The two faces of testing with agent-based models. Journal of Theoretical Biology, 344(??):49–55, March 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005468.

Ciupe:2014:UVH

[CS14b] Stanca M. Ciupe and Elissa J. Schwartz. Understanding virushost dynamics following EIAV infection in SCID horses. Journal of Theoretical Biology, 343(??):1–8, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005286.

Chor:2015:DTS

[CS15a] Benny Chor and Mike Steel. Do tree split probabilities determine the branch lengths? Journal of Theoretical Biology, 374(??):54–59, June 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001393.

Corning:2015:SSD

[CS15b] Peter A. Corning and Eörs Szathmáry. “synergistic selection”: a Darwinian frame for the evolution of complexity. Journal of Theoretical Biology, 371(??):45–58, April 21, 2015.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000612.

Chalub:2018:FPQ

[CS18] Fabio A. C. C. Chalub and Max O. Souza. Fitness potentials and qualitative properties of the Wright–Fisher dynamics. Journal of Theoretical Biology, 457(??):57–65, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304028.

Cleather:2015:RBH

[CSB15] Daniel J. Cleather, Dominic F. L. Southgate, and Anthony M. J. Bull. The role of the biarticular hamstrings and gastrocnemius muscles in closed chain lower limb extension. Journal of Theoretical Biology, 365(??):217–225, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314006092.

Chan:2015:MMS

[CSBK15] Matthew H. Chan, Richard Shine, Gregory P. Brown, and Peter S. Kim. Mathematical modelling of spatial sorting and evolution in a host-parasite system. Journal of Theoretical Biology, 380(??):530–541, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003069.

Cooper:2015:MES

[CSDR15] Racheal L. Cooper, Rebecca A. Segal, Robert F. Diegelmann, and Angela M. Reynolds. Modeling the effects of systemic mediators on the inflammatory phase of wound healing. Journal of Theoretical Biology, 367(??):86–99, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006511.

Suen:2017:RSC

[cSGFB17] Sze chuan Suen, Jeremy D. Goldhaber-Fiebert, and Margaret L. Brandeau. Risk stratification in compartmental epidemic models: Where to draw the line? Journal of Theoretical Biology, 428(??):1–17, September 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730276X.

Coskun:2010:MMP

[CSKF10] Huseyin Coskun, Taryn L. S. Summerfield, Douglas A. Kniss, and Avner Friedman. Mathematical modeling of preadipocyte fate determination. Journal of Theoretical Biology, 265(1):87– 94, July 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001827.

Chong:2019:MMC

[CSKZ19] Ket Hing Chong, Sandhya Samarasinghe, Don Kulasiri, and Jie Zheng. Mathematical modelling of core regulatory mechanism in p53 protein that activates apoptotic switch. Journal of Theoretical Biology, 462(??):134–147, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305551.

Charron:2011:SSC

[CSLE11] Maud V. P. Charron, Henri Seegers, Michel Langlais, and Pauline Ezanno. Seasonal spread and control of bluetongue in cattle. Journal of Theoretical Biology, 291(??):1–9, December 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004449.

Christian:2014:MBM

[CSM^{+}14] Nils Christian, Alexander Skupin, Silvia Morante, Karl Jansen, Giancarlo Rossi, and Oliver Ebenhöh. Mesoscopic behavior from microscopic Markov dynamics and its application to calcium release channels. Journal of Theoretical Biology, 343(??):102–112, February 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005377.

Casal:2005:ABM

[CSR^{+}05] Arancha Casal, Cenk Sumen, Timothy E. Reddy, Mark S. Alber, and Peter P. Lee. Agent-based modeling of the context dependency in T cell recognition. Journal of Theoretical Biology, 236(4):376–391, October 21, 2005. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519305001281. See corrigendum [SRAL12].

Chitnis:2018:MAP

[CSS^{+}18] Nakul Chitnis, Allan Schapira, Christian Schindler, Melissa A. Penny, and Thomas A. Smith. Mathematical analysis to prioritise strategies for malaria elimination. Journal of Theoretical Biology, 455(??):118–130, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830331X.

Cavaliere:2012:PAI

[CST^{+}12] Matteo Cavaliere, Sean Sedwards, Corina E. Tarnita, Martin A. Nowak, and Attila Csikász-Nagy. Prosperity is associated with instability in dynamical networks. Journal of Theoretical Biology, 299(??):126–138, April 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004619.

Chan:2016:MRI

[CST16] Carmen H. S. Chan, Lloyd P. Sanders, and Mark M. Tanaka. Modelling the role of immunity in reversion of viral antigenic sites. Journal of Theoretical Biology, 392(??):23–34, March 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315005998.

Cervantes:2019:HFM

[CST19] Ivan Colorado Cervantes, Vittorio Sansalone, and Luciano Teresi. The heart function as a motor-brake system. Journal of Theoretical Biology, 467(??):23–30, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300517.

Cressman:2012:CED

[CSZT12] Ross Cressman, Jie-Wen Song, Bo-Yu Zhang, and Yi Tao. Cooperation and evolutionary dynamics in the public goods game with institutional incentives. Journal of Theoretical Biology, 299(??):144–151, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003869.

Castro:2014:CCE

[CT14a] Laureano Castro and Miguel A. Toro. Cumulative cultural evolution: The role of teaching. Journal of Theoretical Biology, 347(??):74–83, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000150.

Chakra:2014:UHS

[CT14b] Maria Abou Chakra and Arne Traulsen. Under high stakes and uncertainty the rich should lend the poor a helping hand. Journal of Theoretical Biology, 341(??):123–130, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313004967.

Chen:2014:EV

[CT14c] Frederick Chen and Flavio Toxvaerd. The economics of vaccination. Journal of Theoretical Biology, 363(??):105–117, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004512.

Curatolo:2016:MSF

[CT16] Michele Curatolo and Luciano Teresi. Modeling and simulation of fish swimming with active muscles. Journal of Theoretical Biology, 409(??):18–26, November 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302594.

Caraco:2018:PTS

[CT18a] Thomas Caraco and Wendy C. Turner. Pathogen transmission at stage-structured infectious patches: Killers and vaccinators. Journal of Theoretical Biology, 436(??):51–63, January 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304472.

Contreras-Torres:2018:PSC

[CT18b] Ernesto Contreras-Torres. Predicting structural classes of proteins by incorporating their global and local physicochemical and conformational properties into general Chou’s PseAAC. Journal of Theoretical Biology, 454(??):139–145, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318302807.

Canko:2015:PTC

[CTA15] Osman Canko, Ferhat Taskin, and Kamil Argin. Phylogenetic tree and community structure from a tangled nature model. Journal of Theoretical Biology, 382(??):216–222, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003343.

Choudhury:2018:MMN

[CTB18] M. Jabed A. Choudhury, Philip M. J. Trevelyan, and Graeme P. Boswell. A mathematical model of nutrient influence on fungal competition. Journal of Theoretical Biology, 438(??):9–20, February 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305076.

Caplan:2010:EMG

[CTJ10] S. Roy Caplan, Gloria S. Tannenbaum, and Rose M. Johnstone. An enzymatic model of the growth hormone-releasing hormone oscillator incorporating neuronal synchronization. Journal of Theoretical Biology, 264(3):984–989, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001177.

Crauste:2015:PPS

[CTL^{+}15] F. Crauste, E. Terry, I. Le Mercier, J. Mafille, S. Djebali, T. Andrieu, B. Mercier, G. Kaneko, C. Arpin, J. Marvel, and O. Gandrillon. Predicting pathogen-specific CD8 T cell immune responses from a modeling approach. Journal of Theoretical Biology, 374(??):66–82, June 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001484.

Cibert:2010:BAA

[CTPB10] Christian Cibert, Jérémy Toscano, Vincent Pensée, and Guy Bonnet. Bending of the “9+2” axoneme analyzed by the finite element method. Journal of Theoretical Biology, 264(4):1089–1101, June 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000175X.

Cheng:2011:VTP

[CTSL11] Feng Cheng, Dan Theodorescu, Ira G. Schulman, and Jae K. Lee. In vitro transcriptomic prediction of hepatotoxicity for early drug discovery. Journal of Theoretical Biology, 290(??):27–36, December 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004048.

Chen:2010:MPG

[CTW10] Cheng-Liang Chen, Hong-Wen Tsai, and Sio-Si Wong. Modeling the physiological glucose-insulin dynamic system on diabetics. Journal of Theoretical Biology, 265(3):314–322, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002316.

Cuthill:2015:SCS

[Cut15] Jennifer Hoyal Cuthill. The size of the character state space affects the occurrence and detection of homoplasy: Modelling the probability of incompatibility for unordered phylogenetic characters. Journal of Theoretical Biology, 366(??):24–32, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006237.

Cheung:2015:SUE

[CV15] Allen Cheung and Robert Vickerstaff. Sensory and update errors which can affect path integration. Journal of Theoretical Biology, 372(??):217–221, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000508.

Conti:2019:RNS

[CV19] Federica Conti and Robert A. Van Gorder. The role of network structure and time delay in a metapopulation Wilson–Cowan model. Journal of Theoretical Biology, 477(??):1–13, September 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302024.

Curtis:2018:PPM

[CvBF18] Louis T. Curtis, Victor H. van Berkel, and Hermann B. Frieboes. Pharmacokinetic/pharmacodynamic modeling of combination-chemotherapy for lung cancer. Journal of Theoretical Biology, 448(??):38–52, July 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301498.

Camacho:2010:MMP

[CVH^{+}10] Erika T. Camacho, Miguel A. Colón Vélez, Daniel J. Hernández, Ubaldo Rodríguez Bernier, Jon Van Laarhoven, and Stephen Wirkus. A mathematical model for photoreceptor interactions. Journal of Theoretical Biology, 267(4): 638–646, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519310004728.

Cobben:2015:SSR

[CVO^{+}15] M. M. P. Cobben, J. Verboom, P. F. M. Opdam, R. F. Hoekstra, R. Jochem, and M. J. M. Smulders. Spatial sorting and range shifts: Consequences for evolutionary potential and genetic signature of a dispersal trait. Journal of Theoretical Biology, 373(??):92–99, May 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001289.

Chimal-Vega:2015:ESC

[CVPCV^{+}15] Brenda Chimal-Vega, Norma Paniagua-Castro, Jonathan Carrillo Vazquez, Jorge L. Rosas-Trigueros, Absalom Zamorano-Carrillo, and Claudia G. Benítez-Cardoza. Exploring the structure and conformational landscape of human leptin. A molecular dynamics approach. Journal of Theoretical Biology, 385(??):90–101, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004075.

Chamoli:2011:ADM

[CW11a] Uphar Chamoli and Stephen Wroe. Allometry in the distribution of material properties and geometry of the felid skull: Why larger species may need to change and how they may achieve it. Journal of Theoretical Biology, 283(1):217–226, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311002621.

Chang:2011:GAC

[CW11b] Guisong Chang and Tianming Wang. Genome analysis with the conditional multinomial distribution profile. Journal of Theoretical Biology, 271(1):44–50, February 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006296.

Chen:2011:FRA

[CW11c] Bing-Shuo Chen and Sheng-Nan Wu. Functional role of the activity of ATP-sensitive potassium channels in electrical behavior of hippocampal neurons: Experimental and theoretical studies. Journal of Theoretical Biology, 272(1):16–25, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006454.

Camacho:2013:TPR

[CW13] Erika T. Camacho and Stephen Wirkus. Tracing the progression of retinitis pigmentosa via photoreceptor interactions. Journal of Theoretical Biology, 317(??):105–118, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005127.

Cui:2014:SPA

[CW14] Pengbi Cui and Zhi-Xi Wu. Selfish punishment with avoiding mechanism can alleviate both first-order and second-order social dilemma. Journal of Theoretical Biology, 361(??):111–123, November 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004226.

Chen:2015:IAI

[CW15a] Li Chen and Dongyi Wang. An improved acquaintance immunization strategy for complex network. Journal of Theoretical Biology, 385(??):58–65, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003987.

Chen:2015:GDA

[CW15b] Yongxue Chen and Yongxian Wen. Global dynamic analysis of a H7N9 avian–human influenza model in an outbreak region. Journal of Theoretical Biology, 367(??):180–188, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006912.

Chou:2015:FTD

[CW15c] Tom Chou and Yu Wang. Fixation times in differentiation and evolution in the presence of bottlenecks, deserts, and oases. Journal of Theoretical Biology, 372(??):65–73, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000922.

Chapman:2017:MMC

[CWB^{+}17] Lloyd A. C. Chapman, Jonathan P. Whiteley, Helen M. Byrne, Sarah L. Waters, and Rebecca J. Shipley. Mathematical modelling of cell layer growth in a hollow fibre bioreactor. Journal of Theoretical Biology, 418(??):36–56, April 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300164.

Chen:2016:OPE

[CWF^{+}16] Xiufang Chen, Yue Wang, Tianquan Feng, Ming Yi, Xingan Zhang, and Da Zhou. The overshoot and phenotypic equilibrium in characterizing cancer dynamics of reversible phenotypic plasticity. Journal of Theoretical Biology, 390(??):40–49, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005512.

Camacho:2011:MMF

[CWM11] Erika T. Camacho, Stephen Wirkus, and Pamela A. Marshall. Mathematical modeling of fungal infection in immune compromised individuals: Implications for drug treatment. Journal of Theoretical Biology, 281(1):9–17, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001974.

Chen:2018:ICM

[CWP^{+}18] Michael J. Chen, Jonathan P. Whiteley, Colin P. Please, Andrea Schwab, Franziska Ehlicke, Sarah L. Waters, and Helen M. Byrne. Inducing chondrogenesis in MSC/chondrocyte co-cultures using exogenous TGF-β: a mathematical model. Journal of Theoretical Biology, 439(??):1–13, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317305313.

Chen:2018:DSE

[CWW18] Tong Chen, Zhenghong Wu, and Le Wang. Disseminators or silencers: The effect of information diffusion intensity on cooperation in public goods game. Journal of Theoretical Biology, 452(??):47–55, September 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301930.

Cheng:2018:PMP

[CXC18] Xiang Cheng, Xuan Xiao, and Kuo-Chen Chou. pLoc balmGneg: Predict subcellular localization of Gram-negative bacterial proteins by quasi-balancing training dataset and general PseAAC. Journal of Theoretical Biology, 458(??):92–102, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304351.

Chen:2019:AQI

[CXT19] Daipeng Chen, Yanni Xiao, and Sanyi Tang. Air quality index induced nonsmooth system for respiratory infection. Journal of Theoretical Biology, 460(??):160–169, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304879.

Cai:2011:CMT

[CXWL11] Yan Cai, Shixiong Xu, Jie Wu, and Quan Long. Coupled modelling of tumour angiogenesis, tumour growth and blood perfusion. Journal of Theoretical Biology, 279(1):90–101, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001196.

Chen:2014:RZD

[CZ14] Jing Chen and Aleksey Zinger. The robustness of zerodeterminant strategies in iterated prisoner’s dilemma games. Journal of Theoretical Biology, 357(??):46–54, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002744.

Czaczkes:2014:HGS

[Cza14] Tomer J. Czaczkes. How to not get stuck — negative feedback due to crowding maintains flexibility in ant foraging. Journal of Theoretical Biology, 360(??):172–180, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314003981.

Cheng:2014:EDR

[CZJQ14] Xiao-Sheng Cheng, Heping Zhang, Xian’an Jin, and Wen-Yuan Qiu. Ear decomposition of 3-regular polyhedral links with applications. Journal of Theoretical Biology, 359(??):146–154, October 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003464.

Chatterjee:2012:EGD

[CZN12] Krishnendu Chatterjee, Damien Zufferey, and Martin A. Nowak. Evolutionary game dynamics in populations with different learners. Journal of Theoretical Biology, 301(??):161– 173, May 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200094X.

Cooper:2018:UHR

[CZPC^{+}18] L. J. Cooper, B. Zeller-Plumhoff, G. F. Clough, B. Ganapathisubramani, and T. Roose. Using high resolution X-ray computed tomography to create an image based model of a lymph node. Journal of Theoretical Biology, 449(??):73–82, July 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301899.

Costa:2018:TOS

[CzST^{+}18] Maria Beatriz Walter Costa, Christian Höner zu Siederdissen, Dan Tulpan, Peter F. Stadler, and Katja Nowick. Temporal ordering of substitutions in RNA evolution: Uncovering the structural evolution of the human accelerated region 1. Journal of Theoretical Biology, 438(??):143–150, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305222.

Chernyshova:2016:IMS

[CZT^{+}16] Ekaterina S. Chernyshova, Yulia S. Zaikina, Galina A. Tsvetovskaya, Dmitry I. Strokotov, Maxim A. Yurkin, Elena S. Serebrennikova, Leonid Volkov, Valeri P. Maltsev, and Andrei V. Chernyshev. Influence of magnesium sulfate on HCO_{3} /Cl transmembrane exchange rate in human erythrocytes. Journal of Theoretical Biology, 393(??):194–202, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000291.

Chen:2011:PPI

[CZW^{+}11] Ruoying Chen, Zhiwang Zhang, Di Wu, Peng Zhang, Xinyang Zhang, Yong Wang, and Yong Shi. Prediction of protein interaction hot spots using rough set-based multiple criteria linear programming. Journal of Theoretical Biology, 269(1):174–180, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005552.

Deveraux:2017:NMS

[DAA17] Solenne Deveraux, Rachele Allena, and Denis Aubry. A numerical model suggests the interplay between nuclear plasticity and stiffness during a perfusion assay. Journal of Theoretical Biology, 435(??):62–77, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304198.

Dai:2012:NMF

[Dai12] Jianghong Dai. Novel molecular fossils of bacteria: Insights into hydrothermal origin of life. Journal of Theoretical Biology, 310(??):249–256, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003347.

Dalmau:2017:CMM

[Dal17] Joseba Dalmau. Convergence of a Moran model to Eigen’s quasispecies model. Journal of Theoretical Biology, 420 (??):36–40, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251931730098X.

Dasgupta:2012:DME

[Das12] Partha Dasgupta. Dark matters: Exploitation as cooperation. Journal of Theoretical Biology, 299(??):180–187, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002207.

Dasmahapatra:2013:MHP

[Das13] Srinandan Dasmahapatra. Model of haplotype and phenotype in the evolution of a duplicated autoregulatory activator. Journal of Theoretical Biology, 325(??):83–102, May 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000507.

Dasmeh:2018:MSR

[Das18] Pouria Dasmeh. Multi-step regulation of transcription kinetics explains the non-linear relation between RNA polymerase II density and mRNA expression in dosage compensation. Journal of Theoretical Biology, 438(??):92–95, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305052.

deAguiar:2015:ECP

[dASdC^{+}15] Marcus A. M. de Aguiar, David M. Schneider, Eduardo do Carmo, Paulo R. A. Campos, and Ayana B. Martins. Error catastrophe in populations under similarity-essential recombination. Journal of Theoretical Biology, 374(??):48–53, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001435.

DeMichele:2011:SDH

[DAVS11] C. De Michele, F. Accatino, R. Vezzoli, and R. J. Scholes. Savanna domain in the herbivores-fire parameter space exploiting a tree-grass-soil water dynamic model. Journal of Theoretical Biology, 289(??):74–82, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004097.

Dang:2010:PBA

[DB10] Utkarsh Dang and Chris T. Bauch. A population biological approach to the collective dynamics of countries undergoing demographic transition. Journal of Theoretical Biology, 265(2):167–176, July 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001918.

Davis:2011:LAP

[DB11] Stephen Davis and Stephen J. Bent. Loop analysis for pathogens: niche partitioning in the transmission graph for pathogens of the North American tick ixodes scapularis. Journal of Theoretical Biology, 269(1):96–103, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005394.

Diekman:2018:RCP

[DB18] Casey O. Diekman and Amitabha Bose. Reentrainment of the circadian pacemaker during jet lag: East-west asymmetry and the effects of north-south travel. Journal of Theoretical Biology, 437(??):261–285, January 21, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304575.

Deka:2019:GDM

[DB19] Aniruddha Deka and Samit Bhattacharyya. Game dynamic model of optimal budget allocation under individual vaccination choice. Journal of Theoretical Biology, 470(??):108–118, June 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301213.

Davison:2014:RAD

[DBB14a] Raziel Davison, Carol L. Boggs, and Annette Baudisch. Resource allocation as a driver of senescence: Life history tradeoffs produce age patterns of mortality. Journal of Theoretical Biology, 360(??):251–262, November 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004123.

deBorst:2014:SFR

[dBB14b] K. de Borst and T. K. Bader. Structure-function relationships in hardwood — insight from micromechanical modelling. Journal of Theoretical Biology, 345(??):78–91, March 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931300564X.

Dragicevic:2017:NCV

[DBB^{+}17] Arnaud Dragicevic, Vincent Boulanger, Max Bruciamacchie, Sandrine Chauchard, Jean-Luc Dupouey, and Anne Stenger. Network connectivity value. Journal of Theoretical Biology, 419(??):310–322, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300814.

Dimitrov:2011:MGS

[DBBM11] Dobromir T. Dimitrov, Marie-Claude Boily, Rebecca F. Baggaley, and Benoit Masse. Modeling the gender-specific impact of vaginal microbicides on HIV transmission. Journal of Theoretical Biology, 288(??):9–20, November 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003870.

Daukste:2009:USC

[DBBW09] Liene Daukste, Britta Basse, Bruce C. Baguley, and David J. N. Wall. Using a stem cell and progeny model to illustrate the relationship between cell cycle times of in vivo human tumour cell tissue populations, in vitro primary cultures and the cell lines derived from them. Journal of Theoretical Biology, 260(4):563–571, October 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309002963. See corrigen-

dum [DBBW11].

Daukste:2011:CUS

[DBBW11] Liene Daukste, Britta Basse, Bruce C. Baguley, and David J. N. Wall. Corrigendum to “Using a stem cell and progeny model to illustrate the relationship between cell cycle times of in vivo human tumour cell tissue populations, in vitro primary cultures and the cell lines derived from them” [J. Theor. Biol. 260 (2009) 563–571]. Journal of Theoretical Biology, 289(??):269, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004577. See [DBBW09].

David-Barrett:2012:CBS

[DBD12] Tamás Dávid-Barrett and R. I. M. Dunbar. Cooperation, behavioural synchrony and status in social networks. Journal of Theoretical Biology, 308(??):88–95, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002330.

Daly:2015:IIE

[DBD15] Aisling J. Daly, Jan M. Baetens, and Bernard De Baets. The impact of initial evenness on biodiversity maintenance for a four-species in silico bacterial community. Journal of Theoretical Biology, 387(??):189–205, December 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004865.

Dormanns:2016:RNO

[DBD16] K. Dormanns, R. G. Brown, and T. David. The role of nitric oxide in neurovascular coupling. Journal of Theoretical Biology, 394(??):1–17, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000308.

David-Barrett:2017:FKE

[DBD17] Tamas David-Barrett and Robin I. M. Dunbar. Fertility, kinship and the evolution of mass ideologies. Journal of Theoretical Biology, 417(??):20–27, March 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300152.

denBreems:2016:RPM

[dBE16] Nicoline Y. den Breems and Raluca Eftimie. The repolarisation of M2 and M1 macrophages and its role on cancer outcomes. Journal of Theoretical Biology, 390(??):23–39, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005330.

Dini:2018:UIB

[DBG18] S. Dini, B. J. Binder, and J. E. F. Green. Understanding interactions between populations: Individual based modelling and quantification using pair correlation functions. Journal of Theoretical Biology, 439(??):50–64, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305210.

deBrevern:2011:SSA

[dBJ11] Alexandre G. de Brevern and Agnel Praveen Joseph. Species specific amino acid sequence-protein local structure relationships: An analysis in the light of a structural alphabet. Journal of Theoretical Biology, 276(1):209–217, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000737.

Dyson:2012:MCK

[DBJ12] R. J. Dyson, L. R. Band, and O. E. Jensen. A model of crosslink kinetics in the expanding plant cell wall: Yield stress and enzyme action. Journal of Theoretical Biology, 307(??):125–136, August 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002251.

Djema:2018:CDE

[DBM^{+}18] Walid Djema, Catherine Bonnet, Frédéric Mazenc, Jean Clairambault, Emilia Fridman, Pierre Hirsch, and François Delhommeau. Control in dormancy or eradication of cancer stem cells: Mathematical modeling and stability issues. Journal of Theoretical Biology, 449(??):103–123, July 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301607.

Devi:2018:MMS

[DCL18] Martina Devi, Esther Chingbiaknem, and R. H. Duncan Lyngdoh. A molecular mechanics study on GA codon box translation. Journal of Theoretical Biology, 441(??):28–43, March 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931830002X.

Dimitrio:2013:SPM

[DCN13] Luna Dimitrio, Jean Clairambault, and Roberto Natalini. A spatial physiological model for p53 intracellular dynamics. Journal of Theoretical Biology, 316(??):9–24, January 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004572.

Dinh:2018:GSE

[DCR18] L. Dinh, G. Chowell, and R. Rothenberg. Growth scaling for the early dynamics of HIV/AIDS epidemics in Brazil and the influence of socio-demographic factors. Journal of Theoretical Biology, 442(??):79–86, April 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305775.

Ding:2014:ICO

[DCS14] Xiao Ding, Chang-Chang Cao, and Xiao Sun. Intrinsic correlation of oligonucleotides: a novel genomic signature for metagenome analysis. Journal of Theoretical Biology, 353(??):9–18, July 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001234.

Delhez:2010:RTE

[DD10] Éric J. M. Delhez and Eric Deleersnijder. Residence time and exposure time of sinking phytoplankton in the euphotic layer. Journal of Theoretical Biology, 262(3):505–516, February 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004792.

DeCaluwe:2013:PTA

[DD13] Joëlle De Caluwé and Geneviève Dupont. The progression towards Alzheimer’s disease described as a bistable switch arising from the positive loop between amyloids and Ca^{2+}. Journal of Theoretical Biology, 331(??):12–18, August 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313001835.

Deecke:2018:ITS

[DD18] Lucas Deecke and Hana M. Dobrovolny. Intermittent treatment of severe influenza. Journal of Theoretical Biology, 442(??):129–138, April 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300201.

Drew:2010:CPB

[DDB10] David M. Drew, Geoffrey M. Downes, and Michael Battaglia. CAMBIUM, a process-based model of daily xylem development in eucalyptus. Journal of Theoretical Biology, 264(2):395–406, May 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000809.

Doyennette:2014:MIP

[DDF^{+}14] M. Doyennette, I. Déléris, G. Féron, E. Guichard, I. Souchon, and I. C. Trelea. Main individual and product characteristics influencing in-mouth flavour release during eating masticated food products with different textures: Mechanistic modelling and experimental validation. Journal of Theoretical Biology, 340(??):209–221, January 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004268.

Danesh:2012:BPM

[DDHM12] Kaveh Danesh, Rick Durrett, Laura J. Havrilesky, and Evan Myers. A branching process model of ovarian cancer. Journal of Theoretical Biology, 314(??):10–15, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200447X.

Ding:2010:SFR

[DDLW10] Shuyan Ding, Qi Dai, Hongmei Liu, and Tianming Wang. A simple feature representation vector for phylogenetic analysis of DNA sequences. Journal of Theoretical Biology, 265(4):618–623, August 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002705.

Das:2012:GNF

[DDP12] Dipjyoti Das, Dibyendu Das, and Ashok Prasad. Giant number fluctuations in microbial ecologies. Journal of Theoretical Biology, 308(??):96–104, September 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200272X.

Dercole:2013:ONF

[DDRP13] Fabio Dercole, Marco De Carli, Fabio Della Rossa, and Alessandro V. Papadopoulos. Overpunishing is not necessary to fix cooperation in voluntary public goods games. Journal of Theoretical Biology, 326(??):70–81, June 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006236.

Donzelli:2013:CBT

[DDS13] D. Donzelli, C. De Michele, and R. J. Scholes. Competition between trees and grasses for both soil water and mineral nitrogen in dry savannas. Journal of Theoretical Biology, 332(??):181–190, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001562.

Dunn:2013:SME

[DDSDW13] J. M. Dunn, S. Davis, A. Stacey, and M. A. Diuk-Wasser. A simple model for the establishment of tick-borne pathogens of Ixodes scapularis: a global sensitivity analysis of R_{0}. Journal of Theoretical Biology, 335(??):213–221, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003184.

Draghi:2019:PIM

[DDTL19] Clément Draghi, Fabrice Denis, Alain Tolédano, and Christophe Letellier. Parameter identification of a model for prostate cancer treated by intermittent therapy. Journal of Theoretical Biology, 461(??):117–132, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304764.

DeJaegher:2017:PME

[De 17] Kris De Jaegher. By-product mutualism with evolving common enemies. Journal of Theoretical Biology, 420(??):158– 173, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300929.

DeJaegher:2019:HEM

[De 19] Kris De Jaegher. Harsh environments: Multi-player cooperation with excludability and congestion. Journal of Theoretical Biology, 460(??):18–36, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304788.

DSilva:2017:MSI

[DE17] Jeremy P. D’Silva and Marisa C. Eisenberg. Modeling spatial invasion of Ebola in West Africa. Journal of Theoretical Biology, 428(??):65–75, September 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302576.

Debarre:2017:FPO

[Déb17] F. Débarre. Fidelity of parent-offspring transmission and the evolution of social behavior in structured populations. Journal of Theoretical Biology, 420(??):26–35, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300826.

deAvilaeSilva:2011:BBP

[dEG11] Scheila de Avila e Silva, Sergio Echeverrigaray, and Günther J. L. Gerhardt. BacPP: Bacterial promoter prediction — a tool for accurate sigma-factor specific assignment in enterobacteria. Journal of Theoretical Biology, 287(??):92–99, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003675.

Deng:2014:MSO

[DEG^{+}14] Bo Deng, Alex Estes, Brett Grieb, Douglas Richard, Brittney Hinds, and Eileen Hebets. A male spider’s ornamentation polymorphism maintained by opposing selection with two niches. Journal of Theoretical Biology, 357(??):103–111, September 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002689.

Dauvergne:2015:AQS

[DEK15] Duncan Dauvergne and Leah Edelstein-Keshet. Application of quasi-steady state methods to molecular motor transport on microtubules in fungal hyphae. Journal of Theoretical Biology, 379(??):47–58, August 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002155.

DelGiudice:2012:SRD

[Del12] Marco Del Giudice. Sex ratio dynamics and fluctuating selection on personality. Journal of Theoretical Biology, 297(??):48–60, March 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006138.

Demongeot:2011:IIC

[DEN^{+}11] Jacques Demongeot, Adrien Elena, Mathilde Noual, Sylvain Sené, and Florence Thuderoz. “Immunetworks”, intersecting circuits and dynamics. Journal of Theoretical Biology, 280(1):19–33, July 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931100169X.

Deng:2019:CDP

[Den19] Tao Deng. Configuration of DNA polyhedra of truncated tetrahedron, cuboctahedron, truncated octahedron. Journal of Theoretical Biology, 472(??):4–10, July 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301274.

Deutsch:2018:CMG

[Deu18] J. M. Deutsch. Computational mechanisms in genetic regulation by RNA. Journal of Theoretical Biology, 458(??):156–168, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304466.

deFarias:2013:SPT

[dF13] Sávio Torres de Farias. Suggested phylogeny of tRNAs based on the construction of ancestral sequences. Journal of Theoretical Biology, 335(??):245–248, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003159.

Dunn:2012:MRB

[DFC^{+}12] Sara-Jane Dunn, Alexander G. Fletcher, S. Jonathan Chapman, David J. Gavaghan, and James M. Osborne. Modelling the role of the basement membrane beneath a growing epithelial monolayer. Journal of Theoretical Biology, 298(??):82–91, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311006229.

Dechriste:2018:VMF

[DFG^{+}18a] Guillaume Dechristé, Jérôme Fehrenbach, Elena Griseti, Valérie Lobjois, and Clair Poignard. Viscoelastic modeling of the fusion of multicellular tumor spheroids in growth phase. Journal of Theoretical Biology, 454(??):102–109, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302315.

Dvoriashyna:2018:OEF

[DFG^{+}18b] Mariia Dvoriashyna, Alexander J. E. Foss, Eamonn A. Gaffney, Oliver E. Jensen, and Rodolfo Repetto. Osmotic and electroosmotic fluid transport across the retinal pigment epithelium: a mathematical model. Journal of Theoretical Biology, 456(??):233–248, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303825.

Dafilis:2014:DCS

[DFM^{+}14] Mathew P. Dafilis, Federico Frascoli, Jodie McVernon, Jane M. Heffernan, and James M. McCaw. The dynamical consequences of seasonal forcing, immune boosting and demographic change in a model of disease transmission. Journal of Theoretical Biology, 361(??):124–132, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004408.

Dorvash:2019:DMS

[DFM^{+}19] Mohammadreza Dorvash, Mohammad Farahmandnia, Pouria Mosaddeghi, Mitra Farahmandnejad, Hosein Saber, Mohammadhossein Khorraminejad-Shirazi, Amir Azadi, and Iman Tavassoly. Dynamic modeling of signal transduction by mTOR complexes in cancer. Journal of Theoretical Biology, 483(??):Article 109992, December 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303431.

Doumic:2019:BMN

[DFMR19] Marie Doumic, Klemens Fellner, Mathieu Mezache, and Human Rezaei. A bi-monomeric, nonlinear Becker–Döring-type system to capture oscillatory aggregation kinetics in prion dynamics. Journal of Theoretical Biology, 480(??):241–261, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303194.

Day:2011:MHR

[DFS11] Judy Day, Avner Friedman, and Larry S. Schlesinger. Modeling the host response to inhalation anthrax. Journal of Theoretical Biology, 276(1):199–208, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000804.

DeCaluwe:2017:MPE

[DFT^{+}17] Joëlle De Caluwé, José Romário Fernandes de Melo, Alen Tosenberger, Christian Hermans, Nathalie Verbruggen, Jean-Christophe Leloup, and Didier Gonze. Modeling the photoperiodic entrainment of the plant circadian clock. Journal of Theoretical Biology, 420(??):220–231, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301170.

Dharmadi:2010:ENR

[DG10a] Yandi Dharmadi and Ramon Gonzalez. Elementary network reconstruction: a framework for the analysis of regulatory networks in biological systems. Journal of Theoretical Biology, 263(4):499–509, April 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005748.

dOnofrio:2010:CVT

[dG10b] Alberto d’Onofrio and Alberto Gandolfi. Chemotherapy of vascularised tumours: Role of vessel density and the effect of vascular “pruning”. Journal of Theoretical Biology, 264(2):253–265, May 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000354.

Das:2011:MMF

[DG11] S. Das and P. K. Gupta. A mathematical model on fractional Lotka–Volterra equations. Journal of Theoretical Biology, 277(1):1–6, May 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000609.

Damore:2012:UMC

[DG12] James A. Damore and Jeff Gore. Understanding microbial cooperation. Journal of Theoretical Biology, 299(??):31–41, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001482.

Dercole:2016:URI

[DG16] Fabio Dercole and Stefan A. H. Geritz. Unfolding the resident-invader dynamics of similar strategies. Journal of Theoretical Biology, 394(??):231–254, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005974.

Deng:2011:SSR

[DGC11] Kuiying Deng, Herbert Gintis, and Tianguang Chu. Strengthening strong reciprocity. Journal of Theoretical Biology, 268(1):141–145, January 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000514X.

Dobrovolny:2011:NIT

[DGD^{+}11] Hana M. Dobrovolny, Ronald Gieschke, Brian E. Davies, Nelson L. Jumbe, and Catherine A. A. Beauchemin. Neuraminidase inhibitors for treatment of human and avian strain influenza: a comparative modeling study. Journal of Theoretical Biology, 269(1):234–244, January 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000545X.

Dey:2014:ESA

[DGJ14] Snigdhadip Dey, Bedartha Goswami, and Amitabh Joshi. Effects of symmetric and asymmetric dispersal on the dynamics of heterogeneous metapopulations: Two-patch systems revisited. Journal of Theoretical Biology, 345(??):52–60, March 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005560.

Dey:2015:PMA

[DGJ15] Snigdhadip Dey, Bedartha Goswami, and Amitabh Joshi. A possible mechanism for the attainment of out-of-phase periodic dynamics in two chaotic subpopulations coupled at low dispersal rate. Journal of Theoretical Biology, 367(??):100–110, February 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006870.

Dai:2012:SCP

[DGL12] Qi Dai, Xiaodong Guo, and Lihua Li. Sequence comparison via polar coordinates representation and curve tree. Journal of Theoretical Biology, 292(??):78–85, January 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005017.

Dawson:2015:DDO

[DGM15] Peter Dawson, Ralph Gailis, and Alaster Meehan. Detecting disease outbreaks using a combined Bayesian network and particle filter approach. Journal of Theoretical Biology, 370(??):171–183, April 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000326.

Damti:2018:CED

[DGMY18] Yanir Damti, Ilan Gronau, Shlomo Moran, and Irad Yavneh. Comparing evolutionary distances via adaptive distance functions. Journal of Theoretical Biology, 440(??):88–99, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317305611.

DeVita:2017:MMI

[DGNT17] R. De Vita, R. Grange, P. Nardinocchi, and L. Teresi. Mathematical model for isometric and isotonic muscle contractions. Journal of Theoretical Biology, 425(??):1–10, July 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317302151.

Dobay:2013:RME

[DGRB13] Akos Dobay, Gabriella E. C. Gall, Daniel J. Rankin, and Homayoun C. Bagheri. Renaissance model of an epidemic with quarantine. Journal of Theoretical Biology, 317(??):348–358, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005279.

Dalwadi:2018:AAM

[DGW^{+}18] Mohit P. Dalwadi, Marco Garavaglia, Joseph P. Webb, John R. King, and Nigel P. Minton. Applying asymptotic methods to synthetic biology: Modelling the reaction kinetics of the mevalonate pathway. Journal of Theoretical Biology, 439(??):39–49, February 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305295.

DeJaegher:2016:PMA

[DH16] Kris De Jaegher and Britta Hoyer. By-product mutualism and the ambiguous effects of harsher environments — a game-theoretic model. Journal of Theoretical Biology, 393(??):82–97, March 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000424.

Davies:2018:WAT

[DH18] Adrian M. Davies and Alan G. Holt. Why antioxidant therapies have failed in clinical trials. Journal of Theoretical Biology, 457(??):1–5, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830393X.

Doghaei:2015:MCE

[DHB15] A. Verdian Doghaei, M. R. Housaindokht, and M. R. Bozorgmehr. Molecular crowding effects on conformation and stability of G-quadruplex DNA structure: Insights from molecular dynamics simulation. Journal of Theoretical Biology, 364(??):103–112, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005542.

Dhawan:2019:EMS

[DHBS19] Andrew Dhawan, Adrian L. Harris, Francesca M. Buffa, and Jacob G. Scott. Endogenous miRNA sponges mediate the generation of oscillatory dynamics for a non-coding RNA network. Journal of Theoretical Biology, 481(??):54–60, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931830537X.

Diehl:2014:ODM

[DHHP14] S. Diehl, E. Henningsson, A. Heyden, and S. Perna. A onedimensional moving-boundary model for tubulin-driven axonal growth. Journal of Theoretical Biology, 358(??):194–207, October 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003622.

Doebeli:2013:CTR

[DHK13] Michael Doebeli, Christoph Hauert, and Timothy Killingback. A comment on “Towards a rigorous framework for studying 2-player continuous games” by Shade T. Shutters, Journal of Theoretical Biology 321, 40–43, 2013. Journal of Theoretical Biology, 336(??):240–241, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300266X.

Duncan:2011:OHF

[DHP11] Stephen Duncan, Cameron Hepburn, and Antonis Papachristodoulou. Optimal harvesting of fish stocks under a time-varying discount rate. Journal of Theoretical Biology, 269(1):166–173, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005308.

Drysdale:2010:SBD

[DHRA10] P. M. Drysdale, J. P. Huber, P. A. Robinson, and K. M. Aquino. Spatiotemporal BOLD dynamics from a poroelastic hemodynamic model. Journal of Theoretical Biology, 265(4):524–534, August 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002699.

Dehzangi:2015:GPG

[DHS^{+}15] Abdollah Dehzangi, Rhys Heffernan, Alok Sharma, James Lyons, Kuldip Paliwal, and Abdul Sattar. Gram-positive and Gram-negative protein subcellular localization by incorporating evolutionary-based descriptors into Chou’s general PseAAC. Journal of Theoretical Biology, 364(??):284–294, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005682.

Dumont:2016:TCB

[DHT16] Grégory Dumont, Jacques Henry, and Carmen Oana Tarniceriu. Theoretical connections between mathematical neuronal models corresponding to different expressions of noise. Journal of Theoretical Biology, 406(??):31–41, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301564.

Demongeot:2019:MPB

[DHT19] Jacques Demongeot, Hana Hasgui, and Michel Thellier. Memory in plants: Boolean modeling of the learning and store/recall memory functions in response to environmental stimuli. Journal of Theoretical Biology, 467(??):123–133, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300190.

Dean:2019:TMC

[DHV19] Andrew D. Dean, Malcolm J. Horsburgh, and Bakhti Vasiev. Toxin-mediated competition in weakly motile bacteria. Journal of Theoretical Biology, 480(??):205–217, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930298X.

DiGiulio:2010:BEA

[Di 10] Massimo Di Giulio. Biological evidence against the panspermia theory. Journal of Theoretical Biology, 266(4):569–572, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003747.

DiGiulio:2012:RST

[Di 12] Massimo Di Giulio. The ‘recently’ split transfer RNA genes may be close to merging the two halves of the tRNA rather than having just separated them. Journal of Theoretical Biology, 310(??):1–2, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003098.

DiGiulio:2013:OGC

[Di 13a] Massimo Di Giulio. The origin of the genetic code in the ocean abysses: New comparisons confirm old observations. Journal of Theoretical Biology, 333(??):109–116, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002488.

DiGiulio:2013:PMO

[Di 13b] Massimo Di Giulio. A polyphyletic model for the origin of tRNAs has more support than a monophyletic model. Journal of Theoretical Biology, 318(??):124–128, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005875.

DiGiulio:2014:GCD

[Di 14a] Massimo Di Giulio. The genetic code did not originate from an mRNA codifying polyglycine because the proto-mRNAs already codified for an amino acid number greater than one. Journal of Theoretical Biology, 361(??):204–205, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400544X.

DiGiulio:2014:SGN

[Di 14b] Massimo Di Giulio. The split genes of Nanoarchaeum equitans have not originated in its lineage and have been merged in another Nanoarchaeota: a reply to Podar et al. Journal of Theoretical Biology, 349(??):167–169, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000861.

DiGiulio:2016:LFM

[Di 16] Massimo Di Giulio. The lack of foundation in the mechanism on which are based the physico-chemical theories for the origin of the genetic code is counterposed to the credible and natural mechanism suggested by the coevolution theory. Journal of Theoretical Biology, 399(??):134–140, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300248.

DiGiulio:2017:ATS

[Di 17a] Massimo Di Giulio. The aminoacyl-tRNA synthetases had only a marginal role in the origin of the organization of the genetic code: Evidence in favor of the coevolution theory. Journal of Theoretical Biology, 432(??):14–24, November 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730365X.

DiGiulio:2017:SPA

[Di 17b] Massimo Di Giulio. Some pungent arguments against the physico-chemical theories of the origin of the genetic code and corroborating the coevolution theory. Journal of Theoretical Biology, 414(??):1–4, February 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303794.

DiGiulio:2018:ETW

[Di 18] Massimo Di Giulio. On Earth, there would be a number of fundamental kinds of primary cells — cellular domains — greater than or equal to four. Journal of Theoretical Biology, 443(??):10–17, April 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300419.

DiGiulio:2019:CBT

[Di 19a] Massimo Di Giulio. A comparison between two models for understanding the origin of the tRNA molecule. Journal of Theoretical Biology, 480(??):99–103, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303054.

DiGiulio:2019:QCI

[Di 19b] Massimo Di Giulio. A qualitative criterion for identifying the root of the tree of life. Journal of Theoretical Biology, 464(??):126–131, March 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306453.

DiGiulio:2019:UAD

[Di 19c] Massimo Di Giulio. The universal ancestor, the deeper nodes of the tree of life, and the fundamental types of primary cells (cellular domains). Journal of Theoretical Biology, 460(??):142–143, January 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830496X.

Doebeli:2010:CSS

[DI10a] Michael Doebeli and Iaroslav Ispolatov. Continuously stable strategies as evolutionary branching points. Journal of Theoretical Biology, 266(4):529–535, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003334.

Doebeli:2010:MED

[DI10b] Michael Doebeli and Iaroslav Ispolatov. A model for the evolutionary diversification of religions. Journal of Theoretical Biology, 267(4):676–684, December 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004790.

Dawes:2013:CGM

[DI13] Adriana T. Dawes and David Iron. Cortical geometry may influence placement of interface between Par protein domains in early Caenorhabditis elegans embryos. Journal of Theoretical Biology, 333(??):27–37, September 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001926.

Dilao:2014:RGE

[Dil14] Rui Dilão. The regulation of gene expression in eukaryotes: Bistability and oscillations in repressilator models. Journal of Theoretical Biology, 340(??):199–208, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004384.

Dimonte:2010:CKM

[Dim10] Guy Dimonte. A cell kinetics model for prostate cancer and its application to clinical data and individual patients. Journal of Theoretical Biology, 264(2):420–442, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000986.

Dimitrov:2017:AED

[Dim17] A. G. Dimitrov. An approach to expand description of the pump and co-transporter steady-state current. Journal of Theoretical Biology, 412(??):94–99, January 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630337X.

Dancik:2010:PES

[DJD10] Garrett M. Dancik, Douglas E. Jones, and Karin S. Dorman. Parameter estimation and sensitivity analysis in an agent-based model of Leishmania major infection. Journal of Theoretical Biology, 262(3):398–412, February 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004822.

DeVos:2013:ETI

[DK13a] Anneke S. De Vos and Mirjam E. E. Kretzschmar. The efficiency of targeted intervention in limiting the spread of HIV and hepatitis C virus among injecting drug users. Journal of Theoretical Biology, 333(??):126–134, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002464.

Deonikar:2013:CMP

[DK13b] Prabhakar Deonikar and Mahendra Kavdia. Contribution of membrane permeability and unstirred layer diffusion to nitric oxide-red blood cell interaction. Journal of Theoretical Biology, 317(??):321–330, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005644.

Devipriya:2013:CDD

[DK13c] B. Devipriya and P. Kumaradhas. Charge density distribution and the electrostatic moments of CTPB in the active site of p300 enzyme: a DFT and charge density study. Journal of Theoretical Biology, 335(??):119–129, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002579.

Dutta:2017:IPO

[DKF17] Partha Sharathi Dutta, Bob W. Kooi, and Ulrike Feudel. The impact of a predator on the outcome of competition in the three-trophic food web. Journal of Theoretical Biology, 417(??):28–42, March 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300206.

DeLaFuenteCanto:2018:MGC

[DKP^{+}18] C. De La Fuente Canto, D. I. Kalogiros, M. Ptashnyk, T. S. George, R. Waugh, A. G. Bengough, J. Russell, and L. X. Dupuy. Morphological and genetic characterisation of the root system architecture of selected barley recombinant chromosome substitution lines using an integrated phenotyping approach. Journal of Theoretical Biology, 447(??):84–97, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301346.

Diebner:2016:ESP

[DKR16] Hans H. Diebner, Jörg Kirberg, and Ingo Roeder. An evolutionary stability perspective on oncogenesis control in mature T-cell populations. Journal of Theoretical Biology, 389(??):88–100, January 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005329.

Danko:2015:UNS

[DKS15] Maciej J. Dańko, Jan Kozlowski, and Ralf Schaible. Unraveling the non-senescence phenomenon in hydra. Journal of Theoretical Biology, 382(??):137–149, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003227.

Delitala:2012:MMD

[DL12a] Marcello Delitala and Tommaso Lorenzi. A mathematical model for the dynamics of cancer hepatocytes under therapeutic actions. Journal of Theoretical Biology, 297(??):88–102, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005923.

Donovan:2012:CMR

[DL12b] Graham M. Donovan and Grant Lythe. T-cell movement on the reticular network. Journal of Theoretical Biology, 295(??):59–67, February 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005637.

Day:2015:SLR

[DL15a] J. D. Day and E. K. LeGrand. Synergy of local, regional, and systemic non-specific stressors for host defense against pathogens. Journal of Theoretical Biology, 367(??):39–48, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006651.

Deutsch:2015:MFI

[DL15b] J. M. Deutsch and Ian P. Lewis. Motor function in interpolar microtubules during metaphase. Journal of Theoretical Biology, 370(??):1–10, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500020X.

Donovan:2016:CRN

[DL16] Graham M. Donovan and Grant Lythe. T cell and reticular network co-dependence in HIV infection. Journal of Theoretical Biology, 395(??):211–220, April 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000813.

delaCruz:2016:SMS

[dlCGSA16] Roberto de la Cruz, Pilar Guerrero, Fabian Spill, and Tomás Alarcón. Stochastic multi-scale models of competition within heterogeneous cellular populations: Simulation methods and mean-field analysis. Journal of Theoretical Biology, 407(??):161–183, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302107.

delaEscosura:2015:SPC

[dlEBRM15] Andrés de la Escosura, Carlos Briones, and Kepa Ruiz-Mirazo. The systems perspective at the crossroads between chemistry and biology. Journal of Theoretical Biology, 381(??):11–22, September 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002180.

Dallon:2019:SMR

[DLEMP19] J. C. Dallon, Cécile Leduc, Sandrine Etienne-Manneville, and Stéphanie Portet. Stochastic modeling reveals how motor protein and filament properties affect intermediate filament transport. Journal of Theoretical Biology, 464(??):132–148, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318306192.

Duan:2011:MMA

[DLHS11] Wen Duan, Kiho Lee, Allan E. Herbison, and James Sneyd. A mathematical model of adult GnRH neurons in mouse brain and its bifurcation analysis. Journal of Theoretical Biology, 276(1):22–34, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000610.

Dehzangi:2017:PSA

[DLL^{+}17] Abdollah Dehzangi, Yosvany López, Sunil Pranit Lal, Ghazaleh Taherzadeh, Jacob Michaelson, Abdul Sattar, Tatsuhiko Tsunoda, and Alok Sharma. PSSM-suc: Accurately predicting succinylation using position specific scoring matrix into bigram for feature extraction. Journal of Theoretical Biology, 425(??):97–102, July 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302072.

Ding:2018:SBO

[DLL^{+}18] Weiping Ding, Sen Liu, Shibo Li, Duobiao Ge, Fenfen Li, and Dayong Gao. Simulation of blood and oxygen distributions in a hepatic lobule with sinusoids obstructed by cancer cells. Journal of Theoretical Biology, 446(??):229–237, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301309.

Dyson:2019:DIP

[DLM^{+}19] Janet Dyson, Frank Le Foll, Pierre Magal, Ahmed Noussair, and Jennifer Pasquier. Direct and indirect P-glycoprotein transfers in MCF7 breast cancer cells. Journal of Theoretical Biology, 461(??):239–253, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305228.

deLoubens:2011:BMS

[dLMD^{+}11] Clément de Loubens, Albert Magnin, Marion Doyennette, Ioan Cristian Tréléa, and Isabelle Souchon. A biomechanical model of swallowing for understanding the influence of saliva and food bolus viscosity on flavor release. Journal of Theoretical Biology, 280(1):180–188, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002165.

Destrade:2012:UTS

[DLMK12] Michel Destrade, Yi Liu, Jeremiah G. Murphy, and Ghassan S. Kassab. Uniform transmural strain in pre-stressed arteries occurs at physiological pressure. Journal of Theoretical Biology, 303(??):93–97, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001270.

deLoubens:2010:LAP

[dLMV^{+}10] Clément de Loubens, Albert Magnin, Eric Verin, Marion Doyennette, Ioan Cristian Tréléa, and Isabelle Souchon. A lubrication analysis of pharyngeal peristalsis: Application to flavour release. Journal of Theoretical Biology, 267(3):300–311, December 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004698.

V:2016:PBM

[dlRVFK^{+}16] Aurelio A. de los Reyes V., Doris H. Fuertinger, Franz Kappel, Anna Meyring-Wösten, Stephan Thijssen, and Peter Kotanko. A physiologically based model of vascular refilling during ultrafiltration in hemodialysis. Journal of Theoretical Biology, 390(??):146–155, February 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005627.

DeBlasio:2015:MGD

[DLSD15] Fabio Vittorio De Blasio, Lee Hsiang Liow, Tore Schweder, and Birgitte Freiesleben De Blasio. A model for global diversity in response to temperature change over geological time scales, with reference to planktic organisms. Journal of Theoretical Biology, 365(??):445–456, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006213.

Ding:2013:SWI

[DLYW13] Shuyan Ding, Yang Li, Xiwu Yang, and Tianming Wang. A simple k-word interval method for phylogenetic analysis of DNA sequences. Journal of Theoretical Biology, 317(??):192–199, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005358.

Dai:2011:NCW

[DLYZ11] Qi Dai, Xiaoqing Liu, Yuhua Yao, and Fukun Zhao. Numerical characteristics of word frequencies and their application to dissimilarity measure for sequence comparison. Journal of Theoretical Biology, 276(1):174–180, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000993.

Dilao:2010:MDE

[DM10a] Rui Dilão and Daniele Muraro. mRNA diffusion explains protein gradients in Drosophila early development. Journal of Theoretical Biology, 264(3):847–853, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001359.

dOnofrio:2010:VDD

[dM10b] Alberto d’Onofrio and Piero Manfredi. Vaccine demand driven by vaccine side effects: Dynamic implications for SIR diseases. Journal of Theoretical Biology, 264(2):237–252, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310000743.

DeVargasRoditi:2011:EDB

[DM11] Laura De Vargas Roditi and Franziska Michor. Evolutionary dynamics of BRCA1 alterations in breast tumorigenesis. Journal of Theoretical Biology, 273(1):207–215, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006880.

Differt:2015:IMI

[DM15] Dario Differt and Ralf Möller. Insect models of illuminationinvariant skyline extraction from UV and green channels. Journal of Theoretical Biology, 380(??):444–462, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002994.

Dilao:2019:TRP

[DM19] Rui Dilão and Bárbara Mota. The transcriptional regulation of PER protein in Drosophila. Journal of Theoretical Biology, 469(??):12–17, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300918.

DiGiulio:2014:REM

[DMCP14] Massimo Di Giulio, Marco Moracci, and Beatrice Cobucci-Ponzano. RNA editing and modifications of RNAs might have favoured the evolution of the triplet genetic code from an ennuplet code. Journal of Theoretical Biology, 359(??):1–5, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003208.

Davis:2014:PVH

[DMM^{+}14] Gregg S. Davis, Simeone Marino, Carl F. Marrs, Janet R. Gilsdorf, Suzanne Dawid, and Denise E. Kirschner. Phase variation and host immunity against high molecular weight (HMW) adhesins shape population dynamics of nontypeable haemophilus influenzae within human hosts. Journal of Theoretical Biology, 355(??):208–218, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002197.

DeLeenheer:2017:PPM

[DMO^{+}17] Patrick De Leenheer, Anushaya Mohapatra, Haley A. Ohms, David A. Lytle, and J. M. Cushing. The puzzle of partial migration: Adaptive dynamics and evolutionary game theory perspectives. Journal of Theoretical Biology, 412(??):172–185, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303496.

dOnofrio:2011:IVS

[dMP11] Alberto d’Onofrio, Piero Manfredi, and Piero Poletti. The impact of vaccine side effects on the natural history of immunization programmes: An imitation-game approach. Journal of Theoretical Biology, 273(1):63–71, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006843.

Doherty:2015:SMM

[DMP15] Kevin Doherty, Martin Meere, and Petri T. Piiroinen. Some mathematical models of intermolecular autophosphorylation. Journal of Theoretical Biology, 370(??):27–38, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000247.

Dean:2016:HCN

[DMS^{+}16] Andrew D. Dean, Ewan J. A. Minter, Megan E. S. Sørensen, Christopher D. Lowe, Duncan D. Cameron, Michael A. Brockhurst, and A. Jamie Wood. Host control and nutrient trading in a photosynthetic symbiosis. Journal of Theoretical Biology, 405(??):82–93, September 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931600117X.

Dress:2010:SCT

[DMSW10] Andreas Dress, Vincent Moulton, Mike Steel, and Taoyang Wu. Species, clusters and the ‘tree of life’: a graph-theoretic perspective. Journal of Theoretical Biology, 265(4):535–542, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002742.

Doi:2018:CTI

[DN18] Kazuto Doi and Mayuko Nakamaru. The coevolution of transitive inference and memory capacity in the hawk-dove game. Journal of Theoretical Biology, 456(??):91–107, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303722.

Dalessi:2012:MMG

[DNB12] S. Dalessi, A. Neves, and S. Bergmann. Modeling morphogen gradient formation from arbitrary realistically shaped sources. Journal of Theoretical Biology, 294(??):130–138, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311005297.

dOnofrio:2012:SEP

[d’O12] Alberto d’Onofrio. Spatiotemporal effects of a possible chemorepulsion of tumor cells by immune system effectors. Journal of Theoretical Biology, 296(??):41–48, March 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005832.

Du:2017:TAA

[DOC17] Peng Du, Gregory O’Grady, and Leo K. Cheng. A theoretical analysis of anatomical and functional intestinal slow wave re-entry. Journal of Theoretical Biology, 425(??):72–79, July 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301856.

Guerra:2010:RPE

[dOGL10] João Carlos de Oliveira Guerra and Pedro Licinio. The role played by exons in genomic DNA sequence correlations. Journal of Theoretical Biology, 264(3):830–837, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000130X. See erratum [dOGL13].

Guerra:2013:ERP

[dOGL13] João Carlos de Oliveira Guerra and Pedro Licinio. Erratum to “The role played by exons in genomic DNA sequence correlations” [J. Theor. Biol. 264 (2010) 830–837]. Journal of Theoretical Biology, 328(??):99, July 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000829. See [dOGL10].

Dolejs:2016:SPI

[Dol16] Josef Dolejs. Single parameter of inverse proportion between mortality and age could determine all mortality indicators in the first year of life. Journal of Theoretical Biology, 397(??):193–198, May 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931600148X.

Doluca:2019:GGQ

[Dol19] Osman Doluca. G4Catchall: a G-quadruplex prediction approach considering atypical features. Journal of Theoretical Biology, 463(??):92–98, February 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305976.

Donovan:2013:MAS

[Don13] Graham M. Donovan. Modelling airway smooth muscle passive length adaptation via thick filament length distributions. Journal of Theoretical Biology, 333(??):102–108, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002427.

Donovan:2016:CVD

[Don16] Graham M. Donovan. Clustered ventilation defects and bilinear respiratory reactance in asthma. Journal of Theoretical Biology, 406(??):166–175, October 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301692.

Donovan:2017:IAS

[Don17] G. M. Donovan. Inter-airway structural heterogeneity interacts with dynamic heterogeneity to determine lung function and flow patterns in both asthmatic and control simulated lungs. Journal of Theoretical Biology, 435(??):98–105, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730396X.

Doolittle:2017:DG

[Doo17] W. Ford Doolittle. Darwinizing Gaia. Journal of Theoretical Biology, 434(??):11–19, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300723.

DeBoer:2013:QLT

[DP13a] Rob J. De Boer and Alan S. Perelson. Quantifying T lymphocyte turnover. Journal of Theoretical Biology, 327(??):45–87, June 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313000076.

Droz:2013:MAP

[DP13b] Michel Droz and Andrzej Pekalski. Model of annual plants dynamics with facilitation and competition. Journal of Theoretical Biology, 335(??):1–12, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002828.

Dobreva:2015:MMA

[DPC15] Atanaska Dobreva, Ralf Paus, and N. G. Cogan. Mathematical model for alopecia areata. Journal of Theoretical Biology, 380(??):332–345, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002684.

Doutel:2016:LBD

[DPCM16] E. Doutel, S. I. S. Pinto, J. B. L. M. Campos, and J. M. Miranda. Link between deviations from Murray’s law and occurrence of low wall shear stress regions in the left coronary artery. Journal of Theoretical Biology, 402(??):89–99, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300728.

Damos:2011:ELI

[DPRSS11] Petros T. Damos, Nikos T. Papadopoulos, Alexandros Rigas, and Matilda Savopoulou-Soultani. Energetic loads and informational entropy during insect metamorphosis: Measuring structural variability and self-organization. Journal of Theoretical Biology, 286(??):1–12, October 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003341.

Denis:2012:PAC

[DPvBvA12] Damien Denis, Jean-Sébastien Pierre, Joan van Baaren, and Jacques J. M. van Alphen. Physiological adaptations to climate change in pro-ovigenic parasitoids. Journal of Theoretical Biology, 309(??):67–77, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002822.

Deng:2015:NBI

[DQS^{+}15] Shiguo Deng, Jingchao Qi, Mutua Stephen, Lu Qiu, and Huijie Yang. Network-based identification of reliable bio-markers for cancers. Journal of Theoretical Biology, 383(??):20–27, October 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003719.

Dai:2015:CAA

[DQY^{+}15] Jianghong Dai, Hong Qu, Zhisheng Yu, Jiangke Yang, and Hongxun Zhang. Computational analysis of AnmK-like kinase: New insights into the cell wall metabolism of fungi. Journal of Theoretical Biology, 379(??):59–65, August 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002325.

Dragicevic:2019:CFG

[Dra19] Arnaud Z. Dragicevic. Comparing forest governance models against invasive biological threats. Journal of Theoretical Biology, 462(??):270–282, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305678.

Didier:2011:MMB

[DRC11] Gilles Didier, Elisabeth Remy, and Claudine Chaouiya. Mapping multivalued onto Boolean dynamics. Journal of Theoretical Biology, 270(1):177–184, February 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004911.

Didier:2012:REP

[DRCL12] Gilles Didier, Manuela Royer-Carenzi, and Michel Laurin. The reconstructed evolutionary process with the fossil record. Journal of Theoretical Biology, 315(??):26–37, December 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004687.

deRioja:2015:FPS

[dRFI15] V. L. de Rioja, J. Fort, and N. Isern. Front propagation speeds of t7 virus mutants. Journal of Theoretical Biology, 385(??):112–118, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003938.

Diaz:2011:MCM

[DROC11] H. Diaz, A. A. Ramirez, A. Olarte, and C. Clavijo. A model for the control of malaria using genetically modified vectors. Journal of Theoretical Biology, 276(1):57–66, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000798.

Dybowski:2017:IWH

[DRPM17] R. Dybowski, O. Restif, D. J. Price, and P. Mastroeni. Inferring within-host bottleneck size: a Bayesian approach. Journal of Theoretical Biology, 435(??):218–228, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304289.

DeJaegher:2014:EPC

[DRW14] Kris De Jaegher, Stephanie Rosenkranz, and Utz Weitzel. Economic principles in communication: An experimental study.

Journal of Theoretical Biology, 363(??):62–73, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004470.

Dostalkova:2010:WGC

[DS10a] Iva Dostálková and Marek Spinka. When to go with the crowd: Modelling synchronization of all-or-nothing activity transitions in grouped animals. Journal of Theoretical Biology, 263(4):437–448, April 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000032.

DSuze:2010:SVC

[DS10b] Gina D’Suze and Carlos Sevcik. Scorpion venom complexity fractal analysis. Its relevance for comparing venoms. Journal of Theoretical Biology, 267(3):405–416, December 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004753.

Du:2012:AEC

[DS12] Gang Du and Mao Sun. Aerodynamic effects of corrugation and deformation in flapping wings of hovering hoverflies. Journal of Theoretical Biology, 300(??):19–28, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000112.

Dawes:2013:DHT

[DS13] J. H. P. Dawes and M. O. Souza. A derivation of Holling’s type i, II and III functional responses in predator–prey systems. Journal of Theoretical Biology, 327(??):11–22, June 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931300091X.

daSilva:2015:RWR

[dS15a] José A. L. da Silva. From the RNA world to the RNA/protein world: Contribution of some riboswitch-binding species? Journal of Theoretical Biology, 370(??):197–201, April 7, 2015.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314007103.

Dahlstrom:2015:MCI

[DS15b] Käthe M. Dahlström and Tiina A. Salminen. 3D model for cancerous inhibitor of protein phosphatase 2A armadillo domain unveils highly conserved protein–protein interaction characteristics. Journal of Theoretical Biology, 386(??):78–88, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004531.

Dey:2016:HCE

[DS16] Bibaswan Dey and G. P. Raja Sekhar. Hydrodynamics and convection enhanced macromolecular fluid transport in soft biological tissues: Application to solid tumor. Journal of Theoretical Biology, 395(??):62–86, April 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000709.

Danino:2018:FAF

[DS18] Matan Danino and Nadav M. Shnerb. Fixation and absorption in a fluctuating environment. Journal of Theoretical Biology, 441(??):84–92, March 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300043.

Dansu:2019:MED

[DS19a] Emmanuel J. Dansu and Hiromi Seno. A model for epidemic dynamics in a community with visitor subpopulation. Journal of Theoretical Biology, 478(??):115–127, October 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319302553.

Demongeot:2019:TMR

[DS19b] Jacques Demongeot and Hervé Seligmann. Theoretical minimal RNA rings recapitulate the order of the genetic code’s codon-amino acid assignments. Journal of Theoretical Biology, 471(??):108–116, June 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301316.

Danino:2016:EES

[DSA^{+}16] Matan Danino, Nadav M. Shnerb, Sandro Azaele, William E. Kunin, and David A. Kessler. The effect of environmental stochasticity on species richness in neutral communities. Journal of Theoretical Biology, 409(??):155–164, November 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302715.

dosSantos:2013:NKC

[dSdS13] Renato Vieira dos Santos and Linaena Méricy da Silva. The noise and the KISS in the cancer stem cells niche. Journal of Theoretical Biology, 335(??):79–87, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300297X.

Deeds:2011:CMS

[DSF11] Eric J. Deeds, Van Savage, and Walter Fontana. Curvature in metabolic scaling: a reply to MacKay. Journal of Theoretical Biology, 280(1):197–198, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001883.

deSilva:2010:MMF

[dSKBS10] Eric de Silva, J. Krishnan, Russell Betney, and Ian Stansfield. A mathematical modelling framework for elucidating the role of feedback control in translation termination. Journal of Theoretical Biology, 264(3):808–821, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000196.

Deeba:2018:DAS

[DSLR18] F. Deeba, Paula Sanz-Leon, and P. A. Robinson. Dependence of absence seizure dynamics on physiological parameter evolution. Journal of Theoretical Biology, 454(??):11–21, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302765.

deSousa:2011:AAP

[dSMP^{+}11] Miguel M. de Sousa, Cristian R. Munteanu, Alejandro Pazos, Nuno A. Fonseca, Rui Camacho, and A. L. Magalhães. Amino acid pair- and triplet-wise groupings in the interior of α-helical segments in proteins. Journal of Theoretical Biology, 271(1):136–144, February 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006235.

Dhole:2014:SBP

[DSPM14] Kaustubh Dhole, Gurdeep Singh, Priyadarshini P. Pai, and Sukanta Mondal. Sequence-based prediction of protein– protein interaction sites with l1-logreg classifier. Journal of Theoretical Biology, 348(??):47–54, May 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000459.

dosSantos:2015:MAE

[dSRM15] Renato Vieira dos Santos, Fabiano L. Ribeiro, and Alexandre Souto Martinez. Models for Allee effect based on physical principles. Journal of Theoretical Biology, 385(??):143– 152, November 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519315004117.

Dessavre:2019:PDW

[DSTD19] Adjani Gama Dessavre, Emma Southall, Michael J. Tildesley, and Louise Dyson. The problem of detrending when analysing potential indicators of disease elimination. Journal of Theoretical Biology, 481(??):183–193, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301523.

Domschke:2014:MMC

[DTGC14] Pia Domschke, Dumitru Trucu, Alf Gerisch, and Mark A. J. Chaplain. Mathematical modelling of cancer invasion: Implications of cell adhesion variability for tumour infiltrative growth patterns. Journal of Theoretical Biology, 361(??):41–60, November 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004032.

DeCaluwe:2019:SMG

[DTGD19] Joëlle De Caluwé, Alen Tosenberger, Didier Gonze, and Geneviève Dupont. Signalling-modulated gene regulatory networks in early mammalian development. Journal of Theoretical Biology, 463(??):56–66, February 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305988.

deTrixhe:2015:VSP

[dTKDV15] X. Woot de Trixhe, W. Krzyzanski, F. De Ridder, and A. Vermeulen. vRNA structured population model for hepatitis C virus dynamics. Journal of Theoretical Biology, 378(??):1–11, August 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001800.

Du:2019:VPM

[DTP19] Huan Du, Thi-Bich-Thuy Tran, and Patrick Perré. A 3-variable PDE model for predicting fungal growth derived from microscopic mechanisms. Journal of Theoretical Biology, 470(??):90–100, June 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301225.

Du:2012:SLP

[DTY12] Pufeng Du, Yang Tian, and Yan Yan. Subcellular localization prediction for human internal and organelle membrane proteins with projected gene ontology scores. Journal of Theoretical Biology, 313(??):61–67, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004304.

Dunn:2011:RTM

[Dun11] Ian S. Dunn. RNA templating of molecular assembly and covalent modification patterning in early molecular evolution and modern biosystems. Journal of Theoretical Biology, 284(1):32–41, September 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003080.

Dupuy:2012:ASG

[DV12] Lionel Xavier Dupuy and Matthieu Vignes. An algorithm for the simulation of the growth of root systems on deformable domains. Journal of Theoretical Biology, 310(??):164–174, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003128.

Dormanns:2015:NCI

[DvDBD15] K. Dormanns, E. M. J. van Disseldorp, R. G. Brown, and T. David. Neurovascular coupling and the influence of luminal agonists via the endothelium. Journal of Theoretical Biology, 364(??):49–70, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004901.

Dudziuk:2019:BSF

[DWG^{+}19] Grzegorz Dudziuk, Weronika Wronowska, Anna Gambin, Zuzanna Szymańska, and Mikołaj Rybiński. Biologically sound formal model of Hsp70 heat induction. Journal of Theoretical Biology, 478(??):74–101, October 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301936.

Deng:2014:BBE

[DWL^{+}14] Xinyang Deng, Zhen Wang, Qi Liu, Yong Deng, and Sankaran Mahadevan. A belief-based evolutionarily stable strategy. Journal of Theoretical Biology, 361(??):81–86, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004020.

Ding:2015:CIS

[DWM15] Yanrui Ding, Xueqin Wang, and Zhaolin Mou. Communities in the iron superoxide dismutase amino acid network. Journal of Theoretical Biology, 367(??):278–285, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314006894.

deWeerd:2011:EAP

[dWV11] Harmen de Weerd and Rineke Verbrugge. Evolution of altruistic punishment in heterogeneous populations. Journal of Theoretical Biology, 290(??):88–103, December 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004371.

Ding:2016:GAB

[DXW^{+}16] Weilong Ding, Lifeng Xu, Yang Wei, Fuli Wu, Defeng Zhu, Yuping Zhang, and Nelson Max. Genetic algorithm based approach to optimize phenotypical traits of virtual rice. Journal of Theoretical Biology, 403(??):59–67, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300765.

Ding:2014:PSC

[DYQ^{+}14] Shuyan Ding, Shoujiang Yan, Shuhua Qi, Yan Li, and Yuhua Yao. A protein structural classes prediction method based on PSI-BLAST profile. Journal of Theoretical Biology, 353(??):19–23, July 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931400109X.

Dai:2013:SLW

[DYS^{+}13] Qi Dai, Zhaofang Yan, Zhuoxing Shi, Xiaoqing Liu, Yuhua Yao, and Pingan He. Study of LZ-word distribution and its application for sequence comparison. Journal of Theoretical Biology, 336(??):52–60, November 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003299.

Du:2011:NDD

[DZE11] Xiaoguang Du, Shurong Zhou, and Rampal S. Etienne. Negative density dependence can offset the effect of species competitive asymmetry: a niche-based mechanism for neutrallike patterns. Journal of Theoretical Biology, 278(1):127–134, June 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001329.

Dubey:2010:SAP

[DZJR10] B. Dubey, T. G. Zhao, M. Jonsson, and H. Rahmanov. A solution to the accelerated-predator-satiety Lotka–Volterra predator–prey problem using Boubaker polynomial expansion scheme. Journal of Theoretical Biology, 264(1):154–160, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931000038X.

Dou:2010:SAB

[DZW10] Yongchao Dou, Xiaoqi Zheng, and Jun Wang. Several appropriate background distributions for entropy-based protein sequence conservation measures. Journal of Theoretical Biology, 262(2):317–322, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004688.

Edgington:2017:CSE

[EA17] Matthew P. Edgington and Luke S. Alphey. Conditions for success of engineered underdominance gene drive systems. Journal of Theoretical Biology, 430(??):128–140, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303466.

Edgington:2019:MMR

[EA19] Matthew P. Edgington and Luke S. Alphey. Modeling the mutation and reversal of engineered underdominance gene drives. Journal of Theoretical Biology, 479(??):14–21, October 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302590.

Ejima:2014:PDD

[EAN14] Keisuke Ejima, Kazuyuki Aihara, and Hiroshi Nishiura. Probabilistic differential diagnosis of middle East respiratory syndrome (MERS) using the time from immigration to illness onset among imported cases. Journal of Theoretical Biology, 346(??):47–53, April 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005894.

Enns:2015:LRC

[EB15] Eva A. Enns and Margaret L. Brandeau. Link removal for the control of stochastically evolving epidemics over networks: a comparison of approaches. Journal of Theoretical Biology, 371(??):154–165, April 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000648.

Eftimie:2010:MAT

[EBE10] Raluca Eftimie, Jonathan L. Bramson, and David J. D. Earn. Modeling anti-tumor Th1 and Th2 immunity in the rejection of melanoma. Journal of Theoretical Biology, 265(3):467–480, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002298.

ElCheikh:2014:MCC

[EBE14] R. El Cheikh, S. Bernard, and N. El Khatib. Modeling circadian clock-cell cycle interaction effects on cell population growth rates. Journal of Theoretical Biology, 363(??):318–331, December 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004561.

ElCheikh:2017:MMA

[EBE17] Raouf El Cheikh, Samuel Bernard, and Nader El Khatib. A multiscale modelling approach for the regulation of the cell cycle by the circadian clock. Journal of Theoretical Biology, 426(??):117–125, August 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302291.

El-Bouri:2015:MSH

[EBP15] Wahbi K. El-Bouri and Stephen J. Payne. Multi-scale homogenization of blood flow in 3-dimensional human cerebral microvascular networks. Journal of Theoretical Biology, 380(??):40–47, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002398.

Einarsson:2011:DEB

[EBS11] Baldvin Einarsson, Björn Birnir, and Sven Sigurdsson. A dynamic energy budget (DEB) model for the energy usage and reproduction of the Icelandic capelin (Mallotus villosus). Journal of Theoretical Biology, 281(1):1–8, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001779.

Ellenberger:2017:PTR

[EBSW17] Sabrina Ellenberger, Anke Burmester, Stefan Schuster, and Johannes Wöstemeyer. Post-translational regulation by structural changes of 4-dihydromethyltrisporate dehydrogenase, a key enzyme in sexual and parasitic communication mediated by the trisporic acid pheromone system, of the fungal fusion parasite Parasitella parasitica. Journal of Theoretical Biology, 413(??):50–57, January 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303642.

Eugene:2017:EIT

[EBX17] Sarah Eugène, Thibault Bourgeron, and Zhou Xu. Effects of initial telomere length distribution on senescence onset and heterogeneity. Journal of Theoretical Biology, 413(??):58–65, January 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303605.

Ewing:2016:MET

[ECP^{+}16] D. A. Ewing, C. A. Cobbold, B. V. Purse, M. A. Nunn, and S. M. White. Modelling the effect of temperature on the seasonal population dynamics of temperate mosquitoes. Journal of Theoretical Biology, 400(??):65–79, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300285.

Edgar:2019:AAE

[Edg19] John Alexander Edgar. L-ascorbic acid and the evolution of multicellular eukaryotes. Journal of Theoretical Biology, 476(??):62–73, September 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302280.

Eggen:2011:SRT

[EdKM11] Eelco Eggen, M. Niels de Keijzer, and Bela M. Mulder. Self-regulation in tip-growth: The role of cell wall ageing. Journal of Theoretical Biology, 283(1):113–121, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002864.

Eduardo:2016:MDB

[Edu16] Anderson A. Eduardo. Multiple dimensions of biodiversity and ecosystem processes: Exploring the joint influence of intraspecific, specific and interspecific diversity. Journal of Theoretical Biology, 404(??):215–221, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301321.

Ekanayake:2018:MRW

[EEHMH18] Dinesh B. Ekanayake, Amy J. Ekanayake, Jason Hunt, and Catherine L. Miller-Hunt. Modeling reproduction of whitetail deer and its applications. Journal of Theoretical Biology, 459(??):90–102, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304685.

Eftimie:2013:SUD

[Eft13] Raluca Eftimie. Simultaneous use of different communication mechanisms leads to spatial sorting and unexpected collective behaviours in animal groups. Journal of Theoretical Biology, 337(??):42–53, November 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003718.

Edwards:2010:MCW

[EG10] James R. Edwards and William G. Gibson. A model for Ca^{2+} waves in networks of glial cells incorporating both intercellular and extracellular communication pathways. Journal of Theoretical Biology, 263(1):45–58, March 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005694.

Emerson:2012:RCA

[EG12] I. Arnold Emerson and K. M. Gothandam. Residue centrality in alpha helical polytopic transmembrane protein structures. Journal of Theoretical Biology, 309(??):78–87, September 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002792.

Eftimie:2017:MIC

[EH17] Raluca Eftimie and Haneen Hamam. Modelling and investigation of the CD4+ T cells — macrophages paradox in melanoma immunotherapies. Journal of Theoretical Biology, 420(??):82–104, May 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300772.

El-Haroun:2010:MMN

[EHBC10] Ehab R. El-Haroun, Dominique P. Bureau, and John P. Cant. A mechanistic model of nutritional control of protein synthesis in animal tissues. Journal of Theoretical Biology, 262(2):361–369, January 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900472X.

Edhan:2017:SNR

[EHSR17] Omer Edhan, Ziv Hellman, and Dana Sherill-Rofe. Sex with no regrets: How sexual reproduction uses a no regret learning algorithm for evolutionary advantage. Journal of Theoretical Biology, 426(??):67–81, August 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302266.

Ezoe:2013:CMN

[EI13] Hideo Ezoe and Yusuke Ikegawa. Coexistence of mutualists and non-mutualists in a dual-lattice model. Journal of Theoretical Biology, 332(??):1–8, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001847.

Evans:2016:EMA

[EJ16] Dyfed Lloyd Evans and Shailesh Vinay Joshi. Elucidating modes of activation and herbicide resistance by sequence assembly and molecular modelling of the acetolactate synthase complex in sugarcane. Journal of Theoretical Biology, 407(??):184–197, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302089.

Eisenberg:2017:CBE

[EJ17] Marisa C. Eisenberg and Harsh V. Jain. A confidence building exercise in data and identifiability: Modeling cancer chemotherapy as a case study. Journal of Theoretical Biology, 431(??):63–78, October 27, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303454.

Eisenberg:2019:CCB

[EJ19] Marisa C. Eisenberg and Harsh V. Jain. Corrigendum to “A confidence building exercise in data and identifiability: Modeling cancer chemotherapy as a case study” [Journal of Theoretical Biology 431 (2017) 63–78]. Journal of Theoretical Biology, 464(??):179, March 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306489.

Eskandari:2016:EDA

[EJK16] Mona Eskandari, Ali Javili, and Ellen Kuhl. Elastosis during airway wall remodeling explains multiple co-existing instability patterns. Journal of Theoretical Biology, 403(??):209–218, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301060.

Eide:2018:PPS

[EKFV18] Ragna M. Eide, Andrew L. Krause, Nabil T. Fadai, and Robert A. Van Gorder. predator–prey-subsidy population dynamics on stepping-stone domains with dispersal delays. Journal of Theoretical Biology, 451(??):19–34, August 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318302066.

El-Kebir:2013:CMT

[EKvdKvFK13] M. El-Kebir, M. van der Kuip, A. M. van Furth, and D. E. Kirschner. Computational modeling of tuberculous meningitis reveals an important role for tumor necrosis factorα. Journal of Theoretical Biology, 328(??):43–53, July 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313001124.

Ehn:2012:ASC

[EL12] Micael Ehn and Kevin Laland. Adaptive strategies for cumulative cultural learning. Journal of Theoretical Biology, 301(??):103–111, May 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000756.

Elias:2017:PEM

[Eli17] Ján Elias. Positive effect of Mdm2 on p53 expression explains excitability of p53 in response to DNA damage. Journal of Theoretical Biology, 418(??):94–104, April 7, 2017.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300383.

Elliot:2015:IIA

[Ell15] Michael G. Elliot. Identical inferences about correlated evolution arise from ancestral state reconstruction and independent contrasts. Journal of Theoretical Biology, 364(??):321–325, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005724.

Edwards:2011:IDH

[EM11] James R. Edwards and Mary R. Myerscough. Intelligent decisions from the hive mind: Foragers and nectar receivers of Apis mellifera collaborate to optimise active forager numbers. Journal of Theoretical Biology, 271(1):64–77, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006223.

ElSoufi:2016:CCM

[EM16] Karim El Soufi and Christian J. Michel. Circular code motifs in genomes of eukaryotes. Journal of Theoretical Biology, 408(??):198–212, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302053.

Ehling:2016:HCS

[EME^{+}16] Petra Ehling, Patrick Meuth, Paul Eichinger, Alexander M. Herrmann, Stefan Bittner, Matthias Pawlowski, Susann Pankratz, Michael Herty, Thomas Budde, and Sven G. Meuth. Human T cells in silico: Modelling their electrophysiological behaviour in health and disease. Journal of Theoretical Biology, 404(??):236–250, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301291.

Esmaeili:2010:UCC

[EMM10] Maryam Esmaeili, Hassan Mohabatkar, and Sasan Mohsenzadeh. Using the concept of Chou’s pseudo amino acid composition for risk type prediction of human papillomaviruses. Journal of Theoretical Biology, 263(2):203–209, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930900558X.

Endo:2015:TDV

[EN15] Akira Endo and Hiroshi Nishiura. Transmission dynamics of vivax malaria in the Republic of Korea: Effectiveness of anti-malarial mass chemoprophylaxis. Journal of Theoretical Biology, 380(??):499–505, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003033.

Evans:2009:MMD

[EPJ^{+}09] C. J. Evans, R. M. Phillips, P. F. Jones, P. M. Loadman, B. D. Sleeman, C. J. Twelves, and S. W. Smye. A mathematical model of doxorubicin penetration through multicellular layers. Journal of Theoretical Biology, 257(4):598–608, April 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519308006048. See corri-

gendum [EPJ^{+}11].

Evans:2011:CMM

[EPJ^{+}11] C. J. Evans, R. M. Phillips, P. F. Jones, P. M. Loadman, B. D. Sleeman, C. J. Twelves, and S. W. Smye. Corrigendum to “A mathematical model of doxorubicin penetration through multicellular layers” [J. Theor. Biol. 257 (2009) 598–608]. Journal of Theoretical Biology, 273(1):235, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006855. See [EPJ^{+}09].

Ellis:2019:EDD

[EPP19] John Ellis, Natalia Petrovskaya, and Sergei Petrovskii. Effect of density-dependent individual movement on emerging spatial population distribution: Brownian motion vs Levy flights. Journal of Theoretical Biology, 464(??):159–178, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318306064.

Ellingsen:2012:EPI

[ER12] Tore Ellingsen and Jack Robles. The evolution of parental investment: Re-examining the anisogamy argument. Journal of Theoretical Biology, 299(??):113–119, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005029.

Estrada:2018:CSC

[ER18] Ernesto Estrada and Grant J. Ross. Centralities in simplicial complexes. Applications to protein interaction networks. Journal of Theoretical Biology, 438(??):46–60, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317305040.

Eisenberg:2013:IEM

[ERT13] Marisa C. Eisenberg, Suzanne L. Robertson, and Joseph H. Tien. Identifiability and estimation of multiple transmission pathways in cholera and waterborne disease. Journal of Theoretical Biology, 324(??):84–102, May 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000039.

Escobar:2011:QFN

[Esc11] Alex Escobar. Qualia as the fundamental nature of visual awareness. Journal of Theoretical Biology, 279(1):172–176, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001743.

Enyeart:2015:MME

[ESE15] Peter J. Enyeart, Zachary B. Simpson, and Andrew D. Ellington. A microbial model of economic trading and comparative advantage. Journal of Theoretical Biology, 364(??):326–343, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005694.

Enikeeva:2010:RMS

[ESG10] Farida N. Enikeeva, Konstantin V. Severinov, and Mikhail S. Gelfand. Restriction-modification systems and bacteriophage invasion: Who wins? Journal of Theoretical Biology, 266(4):550–559, October 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003498.

Eastman:2016:UDD

[ESG^{+}16] Justin Eastman, Julian Sass, Johnny M. Gomes, Rodrigo Weber dos Santos, and Elizabeth M. Cherry. Using delay differential equations to induce alternans in a model of cardiac electrophysiology. Journal of Theoretical Biology, 404(??):262–272, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301436.

Escribano:2015:MFC

[ESGA15] J. Escribano, M. T. Sánchez, and J. M. García-Aznar. Modeling the formation of cell-matrix adhesions on a single 3D matrix fiber. Journal of Theoretical Biology, 384(??):84–94, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003525.

Eilertsen:2019:CCM

[ESS19] Justin Eilertsen, Wylie Stroberg, and Santiago Schnell. Characteristic, completion or matching timescales? An analysis of temporary boundaries in enzyme kinetics. Journal of Theoretical Biology, 481(??):28–43, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300050.

Estrada:2010:GWB

[Est10] Ernesto Estrada. Generalized walks-based centrality measures for complex biological networks. Journal of Theoretical Biology, 263(4):556–565, April 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000184.

Ellenberger:2013:CBS

[ESW13] Sabrina Ellenberger, Stefan Schuster, and Johannes Wöstemeyer. Correlation between sequence, structure and function for trisporoid processing proteins in the model zygomycete Mucor mucedo. Journal of Theoretical Biology, 320(??):66–75, March 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312006339.

Ema:2016:RDS

[EUM^{+}16] Hideo Ema, Kouki Uchinomiya, Yohei Morita, Toshio Suda, and Yoh Iwasa. Repopulation dynamics of single haematopoietic stem cells in mouse transplantation experiments: Importance of stem cell composition in competitor cells. Journal of Theoretical Biology, 394(??):57–67, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000382.

Encinas-Viso:2014:SPP

[EVRE14] Francisco Encinas-Viso, Tomás A. Revilla, and Rampal S. Etienne. Shifts in pollinator population structure may jeopardize pollination service. Journal of Theoretical Biology, 352(??):24– 30, July 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001052.

Ez-zizi:2018:OGS

[EzMMH18] Adnane Ez-zizi, John M. McNamara, Gaurav Malhotra, and Alasdair I. Houston. Optimal gut size of small birds and its dependence on environmental and physiological parameters. Journal of Theoretical Biology, 454(??):357–366, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302352.

Ezoe:2012:ESO

[Ezo12] Hideo Ezoe. Evolutionary stability of one-to-many mutualisms. Journal of Theoretical Biology, 314(??):138–144, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004596.

Ezoe:2019:APR

[Ezo19] Hideo Ezoe. Adaptive partner recruitment can help maintain an intra-guild diversity in mutualistic systems. Journal of Theoretical Biology, 478(??):40–47, October 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302528.

Fallah:2017:MMR

[FAFA17] A. Fallah, M. T. Ahmadian, K. Firozbakhsh, and M. M. Aghdam. Micromechanical modeling of rate-dependent behavior of connective tissues. Journal of Theoretical Biology, 416(??):119–128, March 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300127.

Fang:2011:EED

[Fan11] Li Tai Fang. The end-to-end distance of RNA as a randomly self-paired polymer. Journal of Theoretical Biology, 280(1):101–107, July 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002001.

Fann:2012:ECR

[FB12] Sarah L. Fann and Stuart R. Borrett. Environ centrality reveals the tendency of indirect effects to homogenize the functional importance of species in ecosystems. Journal of Theoretical Biology, 294(??):74–86, February 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005534.

Feng:2018:URB

[FB18a] Wenfeng Feng and Richard M. Bailey. Unifying relationships between complexity and stability in mutualistic ecological communities. Journal of Theoretical Biology, 439(??):100–126, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305337.

Ferris:2018:EHD

[FB18b] Charlotte Ferris and Alex Best. The evolution of host defence to parasitism in fluctuating environments. Journal of Theoretical Biology, 440(??):58–65, March 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305398.

Fernandez-Blanco:2013:RFC

[FBAPMD13] Enrique Fernández-Blanco, Vanessa Aguiar-Pulido, Cristian Robert Munteanu, and Julian Dorado. Random forest classification based on star graph topological indices for antioxidant proteins. Journal of Theoretical Biology, 317(??):331–337, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005310.

Fletcher:2012:MMM

[FBC12] Alexander G. Fletcher, Christopher J. W. Breward, and S. Jonathan Chapman. Mathematical modeling of monoclonal conversion in the colonic crypt. Journal of Theoretical Biology, 300(??):118–133, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000318.

Flegg:2012:WHA

[FBFM12] Jennifer A. Flegg, Helen M. Byrne, Mark B. Flegg, and D. L. Sean McElwain. Wound healing angiogenesis: The clinical implications of a simple mathematical model. Journal of Theoretical Biology, 300(??):309–316, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200063X.

Fenley:2011:CMD

[FBK11] Andrew T. Fenley, Suman K. Banik, and Rahul V. Kulkarni. Computational modeling of differences in the quorum sensing induced luminescence phenotypes of Vibrio harveyi and Vibrio cholerae. Journal of Theoretical Biology, 274(1):145–153, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000208.

Ford:2019:LSM

[FBM19] Hugh Z. Ford, Helen M. Byrne, and Mary R. Myerscough. A lipid-structured model for macrophage populations in atherosclerotic plaques. Journal of Theoretical Biology, 479(??):48–63, October 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302802.

Flores:2011:MMA

[FBU11] J. C. Flores, Mauro Bologna, and Deterlino Urzagasti. A mathematical model for the Andean Tiwanaku civilization collapse: Climate variations. Journal of Theoretical Biology, 291(??):29–32, December 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004747.

Fabre:2010:AIP

[FCC^{+}10] Frédéric Fabre, Joël Chadoeuf, Caroline Costa, Hervé Lecoq, and Cécile Desbiez. Asymmetrical over-infection as a process of plant virus emergence. Journal of Theoretical Biology, 265(3):377–388, August 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002183.

Fu:2018:RDP

[FCS18] Yi Fu, Zhiguo Chen, and Jun Sun. Random drift particle swarm optimisation algorithm for highly flexible protein-ligand docking. Journal of Theoretical Biology, 457(??):180–189, November 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304144.

Farr:2011:MNC

[FD11a] Hannah Farr and Tim David. Models of neurovascular coupling via potassium and EET signalling. Journal of Theoretical Biology, 286(??):13–23, October 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100350X.

Fernandez:2011:FBM

[FD11b] Aina Astudillo Fernandez and Jean Louis Deneubourg. On following behaviour as a mechanism for collective movement. Journal of Theoretical Biology, 284(1):7–15, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002992.

Fiorentino:2017:ICM

[FD17] Jonathan Fiorentino and Andrea De Martino. Independent channels for miRNA biosynthesis ensure efficient static and dynamic control in the regulation of the early stages of myogenesis. Journal of Theoretical Biology, 430(??):53–63, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303247.

Facchiano:2018:GCO

[FD18] Angelo Facchiano and Massimo Di Giulio. The genetic code is not an optimal code in a model taking into account both the biosynthetic relationships between amino acids and their physicochemical properties. Journal of Theoretical Biology, 459(??):45–51, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304570.

Flora:2011:SSB

[FDM^{+}11] Cordoleani Flora, Nerini David, Gauduchon Mathias, Morozov Andrew, and Poggiale Jean-Christophe. Structural sensitivity of biological models revisited. Journal of Theoretical Biology, 283(1):82–91, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519311002633. See corrigendum [CNG^{+}12].

Fimmel:2013:DCB

[FDS13] Elena Fimmel, Alberto Danielli, and Lutz Strüngmann. On dichotomic classes and bijections of the genetic code. Journal of Theoretical Biology, 336(??):221–230, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003500.

Fgaier:2010:CMB

[FE10] Hedia Fgaier and Hermann J. Eberl. A competition model between Pseudomonas fluorescens and pathogens via iron chelation. Journal of Theoretical Biology, 263(4):566–578, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309005700.

Fgaier:2011:ACM

[FE11] Hedia Fgaier and Hermann J. Eberl. Antagonistic control of microbial pathogens under iron limitations by siderophore producing bacteria in a chemostat setup. Journal of Theoretical Biology, 273(1):103–114, March 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006892.

Fernando:2011:SMR

[Fer11] Chrisantha Fernando. Symbol manipulation and rule learning in spiking neuronal networks. Journal of Theoretical Biology, 275(1):29–41, April 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931100021X.

Fernandez:2012:TAD

[Fer12] José Luis Fernández. Theoretical approaches for detection of sequence-specific sister-chromatid exchanges in interphase cells. Journal of Theoretical Biology, 310(??):160–163, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003165.

Fernandez:2012:PCP

[FGAM12] J. R. Fernández, J. M. García-Aznar, and R. Martínez. Piezoelectricity could predict sites of formation/resorption in bone remodelling and modelling. Journal of Theoretical Biology, 292(??):86–92, January 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005030.

Fimmel:2015:DCC

[FGGS15] Elena Fimmel, Simone Giannerini, Diego Luis Gonzalez, and Lutz Strüngmann. Dinucleotide circular codes and bijective transformations. Journal of Theoretical Biology, 386(??):159–165, December 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004397.

Fortelius:2015:ADE

[FGGT15] Mikael Fortelius, Stefan Geritz, Mats Gyllenberg, and Jaakko Toivonen. Adaptive dynamics on an environmental gradient that changes over a geological time-scale. Journal of Theoretical Biology, 376(??):91–104, July 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001514.

Feng:2014:MRI

[FGH^{+}14] Zhilan Feng, John W. Glasser, Andrew N. Hill, Mikael A. Franko, Rose-Marie Carlsson, Hans Hallander, Peet Tüll, and Patrick Olin. Modeling rates of infection with transient maternal antibodies and waning active immunity: Application to Bordetella pertussis in Sweden. Journal of Theoretical Biology, 356(??):123–132, September 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002409.

Funk:2010:EDA

[FGJ10] S. Funk, E. Gilad, and V. A. A. Jansen. Endemic disease, awareness, and local behavioural response. Journal of Theoretical Biology, 264(2):501–509, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001074.

Flynn:2010:SFI

[FGLS10] Kevin J. Flynn, H. Christopher Greenwell, Robert W. Lovitt, and Robin J. Shields. Selection for fitness at the individual or population levels: Modelling effects of genetic modifications in microalgae on productivity and environmental safety. Journal of Theoretical Biology, 263(3):269–280, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309005943.

Filotas:2010:PIE

[FGPR10] Elise Filotas, Martin Grant, Lael Parrott, and Per Arne Rikvold. Positive interactions and the emergence of community structure in metacommunities. Journal of Theoretical Biology, 266(3):419–429, October 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003462.

Faruque:2010:DIFa

[FH10a] Imraan Faruque and J. Sean Humbert. Dipteran insect flight dynamics. Part 1. Longitudinal motion about hover. Journal of Theoretical Biology, 264(2):538–552, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000937.

Faruque:2010:DIFb

[FH10b] Imraan Faruque and J. Sean Humbert. Dipteran insect flight dynamics. Part 2: Lateral-directional motion about hover. Journal of Theoretical Biology, 265(3):306–313, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310002328.

Fishman:2013:PMR

[FH13a] Michael A. Fishman and Lilach Hadany. Pollinators’ mating rendezvous and the evolution of floral advertisement. Journal of Theoretical Biology, 316(??):99–106, January 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004742.

Franco:2013:ALC

[FH13b] Daniel Franco and Frank M. Hilker. Adaptive limiter control of unimodal population maps. Journal of Theoretical Biology, 337(??):161–173, November 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003925.

Faruque:2014:WMT

[FH14] Imraan A. Faruque and J. Sean Humbert. Wing motion transformation to evaluate aerodynamic coupling in flapping wing flight. Journal of Theoretical Biology, 363(??):198–204, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400438X.

Fatemi:2015:QPH

[FHG15] Mohammad H. Fatemi, Afsane Heidari, and Sajjad Gharaghani. QSAR prediction of HIV-1 protease inhibitory activities using docking derived molecular descriptors. Journal of Theoretical Biology, 369(??):13–22, March 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500017X.

Fasano:2010:DGP

[FHLM10] A. Fasano, M. A. Herrero, J. M. López, and E. Medina. On the dynamics of the growth plate in primary ossification. Journal of Theoretical Biology, 265(4):543–553, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002730.

Flaig:2018:CMA

[FHM18] Julien Flaig, Nicolas Houy, and Philippe Michel. Canonical modeling of anticipatory vaccination behavior and long term epidemic recurrence. Journal of Theoretical Biology, 436(??):26–38, January 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931730440X.

Fung:2013:NFTa

[FHR13] P. K. Fung, A. L. Haber, and P. A. Robinson. Neural field theory of plasticity in the cerebral cortex. Journal of Theoretical Biology, 318(??):44–57, February 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005085.

Fribourg:2014:MIV

[FHS^{+}14] M. Fribourg, B. Hartmann, M. Schmolke, N. Marjanovic, R. A. Albrecht, A. García-Sastre, S. C. Sealfon, C. Jayaprakash, and F. Hayot. Model of influenza A virus infection: Dynamics of viral antagonism and innate immune response. Journal of Theoretical Biology, 351(??):47–57, June 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001040.

Feng:2015:ETA

[FHSG15] Zhilan Feng, Andrew N. Hill, Philip J. Smith, and John W. Glasser. An elaboration of theory about preventing outbreaks in homogeneous populations to include heterogeneity or preferential mixing. Journal of Theoretical Biology, 386(??):177–187, December 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004415.

Feng:2010:NMC

[FHW^{+}10] Jie Feng, Yong Hu, Ping Wan, Aibing Zhang, and Weizhong Zhao. New method for comparing DNA primary sequences based on a discrimination measure. Journal of Theoretical Biology, 266(4):703–707, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003978.

Fumanelli:2011:MMB

[FIJJ11] Laura Fumanelli, Mimmo Iannelli, Hussnain Ahmed Janjua, and Olivier Jousson. Mathematical modeling of bacterial virulence and host-pathogen interactions in the Dictyostelium/Pseudomonas system. Journal of Theoretical Biology, 270(1):19–24, February 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000603X.

Filin:2010:TSO

[Fil10] Ido Filin. Target size and optimal life history when individual growth and energy budget are stochastic. Journal of Theoretical Biology, 264(2):510–516, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001062.

Filin:2015:RBM

[Fil15] Ido Filin. The relation between maternal phenotype and offspring size, explained by overhead material costs of reproduction. Journal of Theoretical Biology, 364(??):168–178, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005451.

Fedor:2011:PRS

[FIS11] Anna Fedor, Péter Ittzés, and Eörs Szathmáry. Parsing recursive sentences with a connectionist model including a neural stack and synaptic gating. Journal of Theoretical Biology, 271(1):100–105, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519310006211. See corrigendum [FIS16a].

Fedor:2016:CPR

[FIS16a] Anna Fedor, Péter Ittzés, and Eörs Szathmáry. Corrigendum to: “Parsing recursive sentences with a connectionist model including a neural stack and synaptic gating” [J. Theor. Biol. 271 (2011) 100–105]. Journal of Theoretical Biology, 395(??):245, April 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000758. See [FIS11].

Fishman:2016:PEG

[Fis16b] Michael A. Fishman. Polymorphic evolutionary games. Journal of Theoretical Biology, 398(??):130–135, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001533.

Fishman:2019:ACD

[Fis19] Michael A. Fishman. Animal conflicts in diploid populations with sexual reproduction. Journal of Theoretical Biology, 462(??):475–478, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305940.

Fang:2010:GTP

[FJ10] Yi Fang and Junmei Jing. Geometry, thermodynamics, and protein. Journal of Theoretical Biology, 262(3):383–390, February 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004330.

Frieboes:2010:TDM

[FJC^{+}10] Hermann B. Frieboes, Fang Jin, Yao-Li Chuang, Steven M. Wise, John S. Lowengrub, and Vittorio Cristini. Threedimensional multispecies nonlinear tumor growth-II: Tumor invasion and angiogenesis. Journal of Theoretical Biology, 264(4):1254–1278, June 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001116.

Ferdousi:2019:NHG

[FJR19] M. Ferdousi, T. Babaie Janvier, and P. A. Robinson. Nonlinear harmonic generation in the corticothalamic system. Journal of Theoretical Biology, 460(??):184–194, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304831.

Fan:2019:NEM

[FJS^{+}19] Dongmei Fan, Guo-Ping Jiang, Yu-Rong Song, Yin-Wei Li, and Guanrong Chen. Novel epidemic models on PSObased networks. Journal of Theoretical Biology, 477(??): 36–43, September 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519319302334.

Fujita:2013:PFT

[FK13] Hironori Fujita and Masayoshi Kawaguchi. Pattern formation by two-layer Turing system with complementarysynthesis. Journal of Theoretical Biology, 322(??):33–45, April 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000271.

Fallqvist:2016:CMC

[FK16] B. Fallqvist and M. Kroon. Constitutive modelling of composite biopolymer networks. Journal of Theoretical Biology, 395(??):51–61, April 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000734.

Faure:2018:CPM

[FK18] Adrien Fauré and Shizuo Kaji. A circuit-preserving mapping from multilevel to Boolean dynamics. Journal of Theoretical Biology, 440(??):71–79, March 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305532.

Fischer:2012:MEI

[FKB^{+}12] S. Fischer, P. Kurbatova, N. Bessonov, O. Gandrillon, V. Volpert, and F. Crauste. Modeling erythroblastic islands: Using a hybrid model to assess the function of central macrophage. Journal of Theoretical Biology, 298(??):92–106, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000033.

Fozard:2013:MAE

[FKB13] J. A. Fozard, J. R. King, and M. J. Bennett. Modelling auxin efflux carrier phosphorylation and localization. Journal of Theoretical Biology, 319(??):34–49, February 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005863.

Flegg:2015:HVU

[FKDW15] Jennifer A. Flegg, Jessica Kasza, Ian Darby, and Carolina D. Weller. Healing of venous ulcers using compression therapy: Predictions of a mathematical model. Journal of Theoretical Biology, 379(??):1–9, August 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002106.

Fallqvist:2014:MCL

[FKK14] B. Fallqvist, A. Kulachenko, and M. Kroon. Modelling of cross-linked actin networks — influence of geometrical parameters and cross-link compliance. Journal of Theoretical Biology, 350(??):57–69, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000496.

Filippov:2015:MPP

[FKMG15] Alexander Filippov, Alexander Kovalev, Yoko Matsumura, and Stanislav N. Gorb. Male penile propulsion into spiraled spermathecal ducts of female chrysomelid beetles: a numerical simulation approach. Journal of Theoretical Biology, 384(??):140–146, November 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500380X.

Fedorov:2010:TEI

[FKMP10] Alexey A. Fedorov, Vladimir E. Kurochkin, Alexandr I. Martynov, and Ram V. Petrov. Theoretical and experimental investigation of immunoprecipitation pattern formation in gel medium. Journal of Theoretical Biology, 264(1):37–44, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310000044.

Fussell:2019:HAM

[FKV19] Elizabeth F. Fussell, Andrew L. Krause, and Robert A. Van Gorder. Hybrid approach to modeling spatial dynamics of systems with generalist predators. Journal of Theoretical Biology, 462(??):26–47, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305368.

Fan:2012:PMP

[FL12] Guo-Liang Fan and Qian-Zhong Li. Predict mycobacterial proteins subcellular locations by incorporating pseudo-average chemical shift into the general form of Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 304(??):88– 95, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001348.

Fan:2013:DBP

[FL13] Guo-Liang Fan and Qian-Zhong Li. Discriminating bioluminescent proteins by incorporating average chemical shift and evolutionary information into the general form of Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 334(??):45–51, October 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002750.

Fornes:2019:VRM

[FLA^{+}19] Joan Fornés, J. Tomás Lázaro, Tomás Alarcón, Santiago F. Elena, and Josep Sardanyés. Viral replication modes in single-peak fitness landscapes: a dynamical systems analysis. Journal of Theoretical Biology, 460(??):170–183, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304776.

Fernandez-Lozano:2015:CSP

[FLCS^{+}15] Carlos Fernandez-Lozano, Rubén F. Cuiñas, José A. Seoane, Enrique Fernández-Blanco, Julian Dorado, and Cristian R. Munteanu. Classification of signaling proteins based on molecular star graph descriptors using machine learning models. Journal of Theoretical Biology, 384(??):50–58, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315003999.

Fernandez-Lozano:2014:MMP

[FLGGD^{+}14] Carlos Fernandez-Lozano, Marcos Gestal, Humberto González-Díaz, Julián Dorado, Alejandro Pazos, and Cristian R. Munteanu. Markov mean properties for cell death-related protein classification. Journal of Theoretical Biology, 349(??):12– 21, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000526.

Feng:2018:RLR

[FLM18] Zhen-Xing Feng, Qian-Zhong Li, and Jian-Jun Meng. Recognition of the long range enhancer-promoter interactions by further adding DNA structure properties and transcription factor binding motifs in human cell lines. Journal of Theoretical Biology, 445(??):136–150, May 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300833.

Flores:2011:DCM

[Flo11] J. C. Flores. Diffusion coefficient of modern humans outcompeting Neanderthals. Journal of Theoretical Biology, 280(1):189–190, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519311001986.

Foo:2014:MRR

[FLR14] Jasmine Foo, Kevin Leder, and Marc D. Ryser. Multifocality and recurrence risk: a quantitative model of field cancerization. Journal of Theoretical Biology, 355(??):170–184, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002203.

Fan:2016:ITP

[FLW16] Guo-Liang Fan, Yan-Ling Liu, and Hui Wang. Identification of thermophilic proteins by incorporating evolutionary and acid dissociation information into Chou’s general pseudo amino acid composition. Journal of Theoretical Biology, 407(??):138–142, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630193X.

Foo:2010:ERA

[FM10] Jasmine Foo and Franziska Michor. Evolution of resistance to anti-cancer therapy during general dosing schedules. Journal of Theoretical Biology, 263(2):179–188, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005645.

Foster:2013:MMU

[FM13] Kylie J. Foster and Stanley J. Miklavcic. Mathematical modelling of the uptake and transport of salt in plant roots. Journal of Theoretical Biology, 336(??):132–143, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003482.

Foo:2014:EAR

[FM14a] Jasmine Foo and Franziska Michor. Evolution of acquired resistance to anti-cancer therapy. Journal of Theoretical Biology, 355(??):10–20, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001003.

Foster:2014:CUT

[FM14b] Kylie J. Foster and Stanley J. Miklavcic. On the competitive uptake and transport of ions through differentiated root tissues. Journal of Theoretical Biology, 340(??):1–10, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004256.

Filatova:2015:ABM

[FM15a] Olga A. Filatova and Patrick J. O. Miller. An agentbased model of dialect evolution in killer whales. Journal of Theoretical Biology, 373(??):82–91, May 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001290.

Foster:2015:TBU

[FM15b] Kylie J. Foster and Stanley J. Miklavcic. Toward a biophysical understanding of the salt stress response of individual plant cells. Journal of Theoretical Biology, 385(??):130–142, November 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004257.

Francis:2018:ITC

[FM18] Andrew Francis and Vincent Moulton. Identifiability of tree-child phylogenetic networks under a probabilistic recombination-mutation model of evolution. Journal of Theoretical Biology, 446(??):160–167, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301255.

Figueroa-Morales:2012:SAC

[FMLM12] Nuris Figueroa-Morales, Kalet León, and Roberto Mulet. Stochastic approximation to the T cell mediated specific response of the immune system. Journal of Theoretical Biology, 295(??):37–46, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005650.

Fernandez:2010:ERM

[FMS^{+}10] Aina Astudillo Fernandez, Anne-Catherine Mailleux, Fabrice Saffre, Bertrand Krafft, and Jean Louis Deneubourg. Effect of retentive markers on the dynamics of settlement: The case of arthropod silk. Journal of Theoretical Biology, 265(3):411–421, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002432.

Fleming:2012:VPC

[FMS^{+}12] R. M. T. Fleming, C. M. Maes, M. A. Saunders, Y. Ye, and B. Ø. Palsson. A variational principle for computing nonequilibrium fluxes and potentials in genome-scale biochemical networks. Journal of Theoretical Biology, 292(??):71–77, January 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004930.

Fu:2010:IEC

[FNH10] Feng Fu, Martin A. Nowak, and Christoph Hauert. Invasion and expansion of cooperators in lattice populations: Prisoner’s dilemma vs. snowdrift games. Journal of Theoretical Biology, 266(3):358–366, October 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003395.

Fok:2012:MMI

[Fok12] Pak-Wing Fok. Mathematical model of intimal thickening in atherosclerosis: Vessel stenosis as a free boundary problem. Journal of Theoretical Biology, 314(??):23–33, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003827.

Fortier:2010:DER

[For10] Pierre A. Fortier. Detecting and estimating rectification of gap junction conductance based on simulations of dual-cell recordings from a pair and a network of coupled cells. Journal of Theoretical Biology, 265(2):104–114, July 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001839.

Forbes:2014:PSP

[For14] Scott Forbes. Pregnancy sickness and parent-offspring conflict over thyroid function. Journal of Theoretical Biology, 355(??):61–67, August 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001970.

Fumagalli:2015:SEL

[FOT^{+}15] Maria R. Fumagalli, Matteo Osella, Philippe Thomen, Francois Heslot, and Marco Cosentino Lagomarsino. Speed of evolution in large asexual populations with diminishing returns. Journal of Theoretical Biology, 365(??):23–31, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400589X.

Farina:2018:FBI

[FP18] Lorenzo Farina and Paola Paci. A feature-based integrated scoring scheme for cell cycle-regulated genes prioritization. Journal of Theoretical Biology, 459(??):130–141, December 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304612.

Fronhofer:2015:EDD

[FPD15] Emanuel A. Fronhofer, Hans Joachim Poethke, and Ulf Dieckmann. Evolution of dispersal distance: Maternal investment leads to bimodal dispersal kernels. Journal of Theoretical Biology, 365(??):270–279, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006146.

Flores:2010:PEH

[FPdRdH10] J. Flores, E. Corvera Poiré, J. A. del Río, and M. López de Haro. A plausible explanation for heart rates in mammals. Journal of Theoretical Biology, 265(4):599–603, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002869.

Filippov:2011:SIA

[FPG11] Alexander Filippov, Valentin L. Popov, and Stanislav N. Gorb. Shear induced adhesion: Contact mechanics of biological spatula-like attachment devices. Journal of Theoretical Biology, 276(1):126–131, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000750.

Fulcher:2010:QPB

[FPR10] B. D. Fulcher, A. J. K. Phillips, and P. A. Robinson. Quantitative physiologically based modeling of subjective fatigue during sleep deprivation. Journal of Theoretical Biology, 264(2):407–419, May 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001037.

Fowler:2013:CCC

[FR13a] Mike S. Fowler and Lasse Ruokolainen. Colonization, covariance and colour: Environmental and ecological drivers of diversity-stability relationships. Journal of Theoretical Biology, 324(??):32–41, May 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000416.

Fung:2013:NFTb

[FR13b] P. K. Fung and P. A. Robinson. Neural field theory of calcium dependent plasticity with applications to transcranial magnetic stimulation. Journal of Theoretical Biology, 324(??):72– 83, May 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313000386.

Fung:2014:NFT

[FR14] P. K. Fung and P. A. Robinson. Neural field theory of synaptic metaplasticity with applications to theta burst stimulation. Journal of Theoretical Biology, 340(??):164–176, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004499.

Forment:2017:MNC

[FR17] Marzo Forment and Guillermo Rodrigo. Molecular noise can minimize the collective sensitivity of a clonal heterogeneous cell population. Journal of Theoretical Biology, 416(??):38–44, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316304337.

Frank:2016:FWE

[Fra16] T. D. Frank. Front waves in the early RNA world: The Schlögl model and the logistic growth model. Journal of Theoretical Biology, 392(??):62–68, March 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005913.

Frank:2019:EDRa

[Fra19a] Steven A. Frank. Evolutionary design of regulatory control. I. A robust control theory analysis of tradeoffs. Journal of Theoretical Biology, 463(??):121–137, February 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306180.

Frank:2019:EDRb

[Fra19b] Steven A. Frank. Evolutionary design of regulatory control. II. robust error-correcting feedback increases genetic and phenotypic variability. Journal of Theoretical Biology, 468(??):72– 81, May 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931930075X.

French:2010:RTH

[Fre10] Robert M. French. The red tooth hypothesis: a computational model of predator–prey relations, protean escape behavior and sexual reproduction. Journal of Theoretical Biology, 262(1):165–176, January 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004366.

Foster:2013:DLM

[FRG^{+}13] David V. Foster, Mary M. Rorick, Tanja Gesell, Laura M. Feeney, and Jacob G. Foster. Dynamic landscapes: a model of context and contingency in evolution. Journal of Theoretical Biology, 334(??):162–172, October 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002610.

Franco:2015:CCI

[FRH15] Daniel Franco and Alfonso Ruiz-Herrera. To connect or not to connect isolated patches. Journal of Theoretical Biology, 370(??):72–80, April 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000387.

Faria:2014:TIC

[FRP14] L. C. B. Faria, A. S. L. Rocha, and R. Palazzo. Transmission of intra-cellular genetic information: a system proposal. Journal of Theoretical Biology, 358(??):208–231, October 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003233.

Flores:2013:FAV

[FRTP13] Joaquín Flores, Alejandro Meza Romero, Rui D. M. Travasso, and Eugenia Corvera Poiré. Flow and anastomosis in vascular networks. Journal of Theoretical Biology, 317(??):257–270, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005309.

Fessel:2011:ESA

[FS11a] Gion Fessel and Jess G. Snedeker. Equivalent stiffness after glycosaminoglycan depletion in tendon — an ultra-structural finite element model and corresponding experiments. Journal of Theoretical Biology, 268(1):77–83, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005357.

Field:2011:MPS

[FS11b] C. M. Field and D. K. Summers. Multicopy plasmid stability: Revisiting the dimer catastrophe. Journal of Theoretical Biology, 291(??):119–127, December 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004620.

Favre:2012:SGD

[FS12] Maroussia Favre and Didier Sornette. Strong gender differences in reproductive success variance, and the times to the most recent common ancestors. Journal of Theoretical Biology, 310(??):43–54, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200313X.

Filatov:2015:SNS

[FS15a] Felix Filatov and Alexander Shargunov. Short nucleotide sequences in herpesviral genomes identical to the human DNA. Journal of Theoretical Biology, 372(??):12–21, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000788. See corrigendum [FS16a].

Fimmel:2015:HTC

[FS15b] Elena Fimmel and Lutz Strüngmann. On the hierarchy of trinucleotide n-circular codes and their corresponding amino acids. Journal of Theoretical Biology, 364(??):113–120, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005505.

Filatov:2016:CSN

[FS16a] Felix P. Filatov and Alexander Shargunov. Corrigendum to “Short nucleotide sequences in herpesviral genomes identical to the human DNA” [J. Theor. Biol. 372 (2015) 12– 21]. Journal of Theoretical Biology, 392(??):125, March 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315005500. See [FS15a].

Fimmel:2016:MDC

[FS16b] Elena Fimmel and Lutz Strüngmann. Maximal dinucleotide comma-free codes. Journal of Theoretical Biology, 389(??):206–213, January 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005214.

Frasca:2016:DTM

[FS16c] Mattia Frasca and Kieran J. Sharkey. Discrete-time moment closure models for epidemic spreading in populations of interacting individuals. Journal of Theoretical Biology, 399(??):13– 21, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931600165X.

Fatemi:2015:DQI

[FSA15] M. H. Fatemi, E. Mousa Shahroudi, and Z. Amini. Development of quantitative interspecies toxicity relationship modeling of chemicals to fish. Journal of Theoretical Biology, 380(??):16–23, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002453.

Ferretti:2016:MEF

[FSW^{+}16] Luca Ferretti, Benjamin Schmiegelt, Daniel Weinreich, Atsushi Yamauchi, Yutaka Kobayashi, Fumio Tajima, and Guillaume Achaz. Measuring epistasis in fitness landscapes: The correlation of fitness effects of mutations. Journal of Theoretical Biology, 396(??):132–143, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000771.

Fleming:2012:MCE

[FT12] R. M. T. Fleming and I. Thiele. Mass conserved elementary kinetics is sufficient for the existence of a nonequilibrium steady state concentration. Journal of Theoretical Biology, 314(??):173–181, December 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004432.

Fleischer:2018:PYT

[FtL18] Samuel R. Fleischer, Casey P. terHorst, and Jing Li. Pick your trade-offs wisely: predator–prey eco-evo dynamics are qualitatively different under different trade-offs. Journal of Theoretical Biology, 456(??):201–212, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303862.

Fleming:2010:IST

[FTPN10] R. M. T. Fleming, I. Thiele, G. Provan, and H. P. Nasheuer. Integrated stoichiometric, thermodynamic and kinetic modelling of steady state metabolism. Journal of Theoretical Biology, 264(3):683–692, June 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001190.

Fox:2015:DAS

[FTS15] Michael J. Fox, Behrouz Touri, and Jeff S. Shamma. Dynamics in atomic signaling games. Journal of Theoretical Biology, 376(??):82–90, July 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001538.

Fuentes:2014:MCT

[Fue14] Mauricio A. Fuentes. The mechanical cost of transport of fast running animals. Journal of Theoretical Biology, 345(??):22–31, March 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005535.

Fuentes:2015:CMC

[Fue15] Mauricio A. Fuentes. Corrigendum for “the mechanical cost of transport of fast running animals”. Journal of Theoretical Biology, 370(??):209, April 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000235.

Fuentes:2016:TCM

[Fue16] Mauricio A. Fuentes. Theoretical considerations on maximum running speeds for large and small animals. Journal of Theoretical Biology, 390(??):127–135, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005603.

Fuentes:2018:BNA

[Fue18] Marcelino Fuentes. Biological novelty in the anthropocene. Journal of Theoretical Biology, 437(??):137–140, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304964.

Fung:2015:AFC

[FVC15] Tak Fung, Laura Villain, and Ryan A. Chisholm. Analytical formulae for computing dominance from species-abundance distributions. Journal of Theoretical Biology, 386(??):147–158, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004543.

Fleming:2016:CDB

[FVTS16] Ronan M. T. Fleming, Nikos Vlassis, Ines Thiele, and Michael A. Saunders. Conditions for duality between fluxes and concentrations in biochemical networks. Journal of Theoretical Biology, 409(??):1–10, November 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301679.

Fan:2015:IDP

[FW15] Denggui Fan and Qingyun Wang. Improving desynchronization of Parkinsonian neuronal network via tripletstructure coordinated reset stimulation. Journal of Theoretical Biology, 370(??):157–170, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000570.

Fagerlind:2012:DMC

[FWB^{+}12] Magnus G. Fagerlind, Jeremy S. Webb, Nicolas Barraud, Diane McDougald, Andreas Jansson, Patric Nilsson, Mikael Harlén, Staffan Kjelleberg, and Scott A. Rice. Dynamic modelling of cell death during biofilm development. Journal of Theoretical Biology, 295(??):23–36, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005145.

Fu:2011:MFF

[FWLW11] Guifang Fu, Zhong Wang, Jiahan Li, and Rongling Wu. A mathematical framework for functional mapping of complex phenotypes using delay differential equations. Journal of Theoretical Biology, 289(??):206–216, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311003882.

Fowler:2014:OSB

[FWMC14] A. C. Fowler, H. F. Winstanley, M. J. McGuinness, and L. B. Cribbin. Oscillations in soil bacterial redox reactions. Journal of Theoretical Biology, 342(??):33–38, February 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005080.

Fredrich:2019:DVA

[FWR19] Thierry Fredrich, Michael Welter, and Heiko Rieger. Dynamic vessel adaptation in synthetic arteriovenous networks. Journal of Theoretical Biology, 483(??):Article 109989, December 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303406.

Filippov:2017:NSC

[FWSG17] Alexander É. Filippov, Jonas O. Wolff, Michael Seiter, and Stanislav N. Gorb. Numerical simulation of colloidal selfassembly of super-hydrophobic arachnid cerotegument structures. Journal of Theoretical Biology, 430(??):1–8, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303260.

Feng:2018:VVD

[FWW18] Xue Feng, Bin Wu, and Long Wang. Voluntary vaccination dilemma with evolving psychological perceptions. Journal of Theoretical Biology, 439(??):65–75, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305180.

Fan:2018:HGT

[FXML18] Yuhang Fan, Yandong Xiao, Babak Momeni, and Yang-Yu Liu. Horizontal gene transfer can help maintain the equilibrium of microbial communities. Journal of Theoretical Biology, 454(??):53–59, October 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302832.

Fu:2015:RSS

[FYZ15] Haoyue Fu, Lianping Yang, and Xiangde Zhang. An RNA secondary structure prediction method based on minimum and suboptimal free energy structures. Journal of Theoretical Biology, 380(??):473–479, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002350.

Feng:2014:BMT

[FZ14] Shiliang Feng and Weiping Zhu. Bidirectional molecular transport shapes cell polarization in a two-dimensional model of eukaryotic chemotaxis. Journal of Theoretical Biology, 363(??):235–246, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004949.

Fumagalli:2018:ICT

[FZL18] Maria Rita Fumagalli, Stefano Zapperi, and Caterina A. M. La Porta. Impact of the cross-talk between circular and messenger RNAs on cell regulation. Journal of Theoretical Biology, 454(??):386–395, October 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303230.

Feng:2017:ESG

[FZW17] Xue Feng, Yanling Zhang, and Long Wang. Evolution of stinginess and generosity in finite populations. Journal of Theoretical Biology, 421(??):71–80, May 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301388.

Garzon-Alvarado:2013:HFP

[GA13] Diego A. Garzón-Alvarado. A hypothesis on the formation of the primary ossification centers in the membranous neurocranium: a mathematical and computational model. Journal of Theoretical Biology, 317(??):366–376, January 21, 2013.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004936.

Graham:2014:LTM

[GA14] Erica J. Graham and Frederick R. Adler. Long-term models of oxidative stress and mitochondrial damage in insulin resistance progression. Journal of Theoretical Biology, 340(??):238–250, January 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300458X.

Greenman:2015:EPS

[GA15] J. V. Greenman and B. Adams. The exclusion problem in seasonally forced epidemiological systems. Journal of Theoretical Biology, 367(??):49–60, February 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006547.

Gaythorpe:2016:DDO

[GA16] Katy Gaythorpe and Ben Adams. Disease and disaster: Optimal deployment of epidemic control facilities in a spatially heterogeneous population with changing behaviour. Journal of Theoretical Biology, 397(??):169–178, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001478.

Gunasinghe:2014:EHQ

[GAB14] Upuli Gunasinghe, Damminda Alahakoon, and Susan Bedingfield. Extraction of high quality k-words for alignmentfree sequence comparison. Journal of Theoretical Biology, 358(??):31–51, October 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002872.

Grilli:2012:SAS

[GABM12] Jacopo Grilli, Sandro Azaele, Jayanth R. Banavar, and Amos Maritan. Spatial aggregation and the species-area relationship across scales. Journal of Theoretical Biology, 313(??):87–97, November 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003839.

Garcia-Algarra:2014:RLA

[GAGP^{+}14] Javier García-Algarra, Javier Galeano, Juan Manuel Pastor, José María Iriondo, and José J. Ramasco. Rethinking the logistic approach for population dynamics of mutualistic interactions. Journal of Theoretical Biology, 363(??):332–343, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005001.

Galla:2010:IIN

[Gal10] Tobias Galla. Independence and interdependence in the nest-site choice by honeybee swarms: Agent-based models, analytical approaches and pattern formation. Journal of Theoretical Biology, 262(1):186–196, January 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004275.

Galla:2011:IIA

[Gal11a] Tobias Galla. Imitation, internal absorption and the reversal of local drift in stochastic evolutionary games. Journal of Theoretical Biology, 269(1):46–56, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005126.

Gao:2011:PAF

[GAL11b] Na Gao, Hikaru Aono, and Hao Liu. Perturbation analysis of 6DoF flight dynamics and passive dynamic stability of hovering fruit fly Drosophila melanogaster. Journal of Theoretical Biology, 270(1):98–111, February 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006077.

Goktepe:2010:MME

[GAPK10] Serdar Göktepe, Oscar John Abilez, Kevin Kit Parker, and Ellen Kuhl. A multiscale model for eccentric and concentric cardiac growth through sarcomerogenesis. Journal of Theoretical Biology, 265(3):433–442, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002146.

Gardner:2010:SBD

[Gar10] Andy Gardner. Sex-biased dispersal of adults mediates the evolution of altruism among juveniles. Journal of Theoretical Biology, 262(2):339–345, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004585.

Gardner:2011:KSU

[Gar11] Andy Gardner. Kin selection under blending inheritance. Journal of Theoretical Biology, 284(1):125–129, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931100333X.

Gardner:2014:TRV

[Gar14] Andy Gardner. Total reproductive value of juvenile females is twice that of juvenile males under x-linkage and haplodiploidy. Journal of Theoretical Biology, 359(??):236–237, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314003798.

Gautestad:2011:MMI

[Gau11] Arild O. Gautestad. Memory matters: Influence from a cognitive map on animal space use. Journal of Theoretical Biology, 287(??):26–36, October 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003602.

Guerberoff:2015:SMM

[GAV15] Gustavo Guerberoff and Fernando Alvarez-Valin. A stochastic microscopic model for the dynamics of antigenic variation.

Journal of Theoretical Biology, 380(??):489–498, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315003045.

Guzzetta:2011:MSD

[GAY^{+}11] Giorgio Guzzetta, Marco Ajelli, Zhenhua Yang, Stefano Merler, Cesare Furlanello, and Denise Kirschner. Modeling socio-demography to capture tuberculosis transmission dynamics in a low burden setting. Journal of Theoretical Biology, 289(??):197–205, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004358.

Guzzetta:2015:ECI

[GAY^{+}15] Giorgio Guzzetta, Marco Ajelli, Zhenhua Yang, Leonard N. Mukasa, Naveen Patil, Joseph H. Bates, Denise E. Kirschner, and Stefano Merler. Effectiveness of contact investigations for tuberculosis control in Arkansas. Journal of Theoretical Biology, 380(??):238–246, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002660.

Gilarranz:2012:SNS

[GB12] Luis J. Gilarranz and Jordi Bascompte. Spatial network structure and metapopulation persistence. Journal of Theoretical Biology, 297(??):11–16, March 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005972.

Garbey:2013:DSD

[GB13] Marc Garbey and Scott A. Berceli. A dynamical system that describes vein graft adaptation and failure. Journal of Theoretical Biology, 336(??):209–220, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003275.

Gogoi:2016:PBV

[GBC^{+}16] Dhrubajyoti Gogoi, Vishwa Jyoti Baruah, Amrita Kashyap Chaliha, Bibhuti Bhushan Kakoti, Diganta Sarma, and Alak Kumar Buragohain. 3D pharmacophore-based virtual screening, docking and density functional theory approach towards the discovery of novel human epidermal growth factor receptor-2 (HER2) inhibitors. Journal of Theoretical Biology, 411(??):68–80, December 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303058.

Greene:2013:BMP

[GBIR13] Joshua Greene, Marc R. Birtwistle, Leszek Ignatowicz, and Grzegorz A. Rempala. Bayesian multivariate Poisson abundance models for T-cell receptor data. Journal of Theoretical Biology, 326(??):1–10, June 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000775.

Goutelle:2011:MMP

[GBJ^{+}11] Sylvain Goutelle, Laurent Bourguignon, Roger W. Jelliffe, John E. Conte, and Pascal Maire. Mathematical modeling of pulmonary tuberculosis therapy: Insights from a prototype model with rifampin. Journal of Theoretical Biology, 282(1):80–92, August 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002554.

Guerra:2018:MSW

[GBJ18] Ana Guerra, Jorge Belinha, and Renato Natal Jorge. Modelling skin wound healing angiogenesis: a review. Journal of Theoretical Biology, 459(??):1–17, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304569.

Greenspoon:2018:ISH

[GBM18] Philip B. Greenspoon, Sydney Banton, and Nicole Mideo. Immune system handling time may alter the outcome of competition between pathogens and the immune system. Journal of Theoretical Biology, 447(??):25–31, June 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301243.

Glimm:2014:MMG

[GBN14] T. Glimm, R. Bhat, and S. A. Newman. Modeling the morphodynamic galectin patterning network of the developing avian limb skeleton. Journal of Theoretical Biology, 346(??):86–108, April 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005559.

Gromov:2019:SEP

[GBRS19] Dmitry Gromov, Ingo Bulla, and Ethan O. Romero-Severson. Systematic evaluation of the population-level effects of alternative treatment strategies on the basic reproduction number. Journal of Theoretical Biology, 462(??):381–390, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305824.

Grilo:2011:EAE

[GC11] Carlos Grilo and Luís Correia. Effects of asynchronism on evolutionary games. Journal of Theoretical Biology, 269(1):109–122, January 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005564.

Gray:2016:ASM

[GC16] Catheryn W. Gray and Adelle C. F. Coster. The Akt switch model: Is location sufficient? Journal of Theoretical Biology, 398(??):103–111, June 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001466.

Goyal:2018:DIV

[GC18] Ashish Goyal and Ranjit Chauhan. The dynamics of integration, viral suppression and cell–cell transmission in the development of occult Hepatitis B virus infection. Journal of Theoretical Biology, 455(??):269–280, October 14, 2018.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303187.

Garbey:2017:VAP

[GCB17] Marc Garbey, Stefano Casarin, and Scott A. Berceli. Vascular adaptation: Pattern formation and cross validation between an agent based model and a dynamical system. Journal of Theoretical Biology, 429(??):149–163, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317302849.

Gomez-Corral:2014:CSS

[GCG14] A. Gómez-Corral and M. López García. Control strategies for a stochastic model of host-parasite interaction in a seasonal environment. Journal of Theoretical Biology, 354(??):1–11, August 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400160X.

Garay:2014:UMS

[GCM14] József Garay, Villő Csiszár, and Tamás F. Móri. Under multilevel selection: “When shall you be neither spiteful nor envious?”. Journal of Theoretical Biology, 340(??):73–84, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004050.

Garay:2017:ESM

[GCM17a] József Garay, Villő Csiszár, and Tamás F. Móri. Evolutionary stability for matrix games under time constraints. Journal of Theoretical Biology, 415(??):1–12, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931630409X.

Garay:2017:SPS

[GCM17b] József Garay, Villő Csiszár, and Tamás F. Móri. Survival phenotype, selfish individual versus Darwinian phenotype. Journal of Theoretical Biology, 430(??):86–91, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303090.

Gobel:2011:RMA

[GCO^{+}11] B. Göbel, M. Chung, K. M. Oltmanns, A. Peters, and D. Langemann. Robust modeling of appetite regulation. Journal of Theoretical Biology, 291(??):65–75, December 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004681.

Grujic:2012:CCD

[GCS12] Jelena Grujić, José A. Cuesta, and Angel Sánchez. On the coexistence of cooperators, defectors and conditional cooperators in the multiplayer iterated Prisoner’s dilemma. Journal of Theoretical Biology, 300(??):299–308, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000677.

Garcia-Camacho:2012:MMP

[GCSMMG^{+}12] F. García-Camacho, A. Sánchez-Mirón, E. Molina-Grima, F. Camacho-Rubio, and J. C. Merchuck. A mechanistic model of photosynthesis in microalgae including photoacclimation dynamics. Journal of Theoretical Biology, 304(??):1–15, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312001452.

Gallinato:2017:TGM

[GCSP17] Olivier Gallinato, Thierry Colin, Olivier Saut, and Clair Poignard. Tumor growth model of ductal carcinoma: from in situ phase to stroma invasion. Journal of Theoretical Biology, 429(??):253–266, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302990.

Gonzalez-Casanova:2014:SSB

[GCvWE^{+}14] Adrián González-Casanova, Eneas Aguirre von Wobeser, Guadalupe Espín, Luis Servín-González, Noemi Kurt, Dario Spanò, Jochen Blath, and Gloria Soberón-Chávez. Strong seed-bank effects in bacterial evolution. Journal of Theoretical Biology, 356(??):62–70, September 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002185.

Guo:2018:MMA

[GCYL18] Muyi Guo, Yan Cai, Xinke Yao, and Zhiyong Li. Mathematical modeling of atherosclerotic plaque destabilization: Role of neovascularization and intraplaque hemorrhage. Journal of Theoretical Biology, 450(??):53–65, August 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830208X.

Gove:2012:ECN

[GCZ^{+}12] Robert P. Gove, William Chen, Nicholas B. Zweber, Rebecca Erwin, Jan Rychtář, and David L. Remington. Effects of causal networks on the structure and stability of resource allocation trait correlations. Journal of Theoretical Biology, 293(??):1–14, January 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005054.

Gavryushkin:2016:SUP

[GD16] Alex Gavryushkin and Alexei J. Drummond. The space of ultrametric phylogenetic trees. Journal of Theoretical Biology, 403(??):197–208, August 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630073X.

Gupta:2019:RCC

[GDB^{+}19] Ankur Gupta, Phonindra Nath Das, Rania Bouzeyen, Sonali Porey Karmakar, Ramandeep Singh, Nandadulal Bairagi, and Samrat Chatterjee. Restoration of cytosolic calcium inhibits mycobacterium tuberculosis intracellular growth: Theoretical evidence and experimental observation. Journal of Theoretical Biology, 472(??):110–123, July 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301584.

Gabriel:2017:QIO

[GDF17] Sargon A. Gabriel, Yan Ding, and Yuqing Feng. Quantifying the influence of oscillatory flow disturbances on blood flow. Journal of Theoretical Biology, 430(??):195–206, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303338.

Gabriel:2018:MPA

[GDF18] Sargon A. Gabriel, Yan Ding, and Yuqing Feng. Modelling the period-average transport of species within pulsatile blood flow. Journal of Theoretical Biology, 457(??):258–269, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303308.

Goede:2017:PGI

[GdGL17] Simon L. Goede, Bastiaan E. de Galan, and Melvin Khee Shing Leow. Personalized glucose-insulin model based on signal analysis. Journal of Theoretical Biology, 419(??):333–342, April 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316304283.

Guill:2011:TSM

[GDJC11] Christian Guill, Barbara Drossel, Wolfram Just, and Eddy Carmack. A three-species model explaining cyclic dominance of Pacific salmon. Journal of Theoretical Biology, 276(1):16– 21, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000622.

Gonzalez-Diaz:2011:NMB

[GDPPSS^{+}11] Humberto González-Díaz, Francisco Prado-Prado, Eduardo Sobarzo-Sánchez, Mohamed Haddad, Séverine Maurel Chevalley, Alexis Valentin, Joëlle Quetin-Leclercq, María A. Dea-Ayuela, María Teresa Gomez-Muños, Cristian R. Munteanu, Juan José Torres-Labandeira, Xerardo García-Mera, Ricardo A. Tapia, and Florencio M. Ubeira. NL MIND-BEST: a web server for ligands and proteins discovery — theoreticexperimental study of proteins of Giardia lamblia and new compounds active against Plasmodium falciparum. Journal of Theoretical Biology, 276(1):229–249, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000221.

Garcia-Domingo:2011:ETS

[GDS11] Josep L. Garcia-Domingo and Joan Saldaña. Extinction threshold for spatial forest dynamics with height structure. Journal of Theoretical Biology, 276(1):138–149, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100083X.

Glass:2018:HMG

[GE18] Leon Glass and Roderick Edwards. Hybrid models of genetic networks: Mathematical challenges and biological relevance. Journal of Theoretical Biology, 458(??):111–118, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304429.

Gerberry:2016:PAB

[Ger16] David J. Gerberry. Practical aspects of backward bifurcation in a mathematical model for tuberculosis. Journal of Theoretical Biology, 388(??):15–36, January 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004968.

Ghosal:2011:DBI

[GF11] Sandip Ghosal and Yoshio Fukui. Does buckling instability of the pseudopodium limit how well an amoeba can climb? Journal of Theoretical Biology, 271(1):202–204, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006314.

Gatti:2018:NEA

[GFH^{+}18] Roberto Cazzolla Gatti, Brian Fath, Wim Hordijk, Stuart Kauffman, and Robert Ulanowicz. Niche emergence as an autocatalytic process in the evolution of ecosystems. Journal of Theoretical Biology, 454(??):110–117, October 7, 2018.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302856.

Gavagnin:2019:ISC

[GFM^{+}19] Enrico Gavagnin, Matthew J. Ford, Richard L. Mort, Tim Rogers, and Christian A. Yates. The invasion speed of cell migration models with realistic cell cycle time distributions. Journal of Theoretical Biology, 481(??):91–99, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304417.

Guillamon:2015:BAO

[GFS15] Antoni Guillamon, Ernest Fontich, and Josep Sardanyés. Bifurcations analysis of oscillating hypercycles. Journal of Theoretical Biology, 387(??):23–30, December 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004610.

Geritz:2012:MDD

[GG12] Stefan Geritz and Mats Gyllenberg. A mechanistic derivation of the DeAngelis–Beddington functional response. Journal of Theoretical Biology, 314(??):106–108, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004523.

Geritz:2014:DBF

[GG14] Stefan A. H. Geritz and Mats Gyllenberg. The DeAngelis– Beddington functional response and the evolution of timidity of the prey. Journal of Theoretical Biology, 359(??):37–44, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002860.

Gaunt:2016:PRA

[GG16] Stephen J. Gaunt and Alexander L. Gaunt. Possible rules for the ancestral origin of Hox gene collinearity. Journal of Theoretical Biology, 410(??):1–8, December 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302971.

Gilbert:2014:SSP

[GGBW14] Mark A. Gilbert, Eamonn A. Gaffney, James M. Bullock, and Steven M. White. Spreading speeds for plant populations in landscapes with low environmental variation. Journal of Theoretical Biology, 363(??):436–452, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004779.

Gao:2014:ITH

[GGC14] Zhipeng Gao, Hongmei Guo, and Weiyi Chen. Initial tension of the human extraocular muscles in the primary eye position. Journal of Theoretical Biology, 353(??):78–83, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314001532.

Guo:2016:CBG

[GGCJ16] Shun Guo, Donghui Guo, Lifei Chen, and Qingshan Jiang. A centroid-based gene selection method for microarray data classification. Journal of Theoretical Biology, 400(??):32–41, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300108.

Guria:2012:IFR

[GGG12] Konstantin G. Guria, Anita R. Gagarina, and Georgy Th. Guria. Instabilities in fibrinolytic regulatory system. Theoretical analysis of blow-up phenomena. Journal of Theoretical Biology, 304(??):27–38, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001634.

Gerard:2019:RSM

[GGG19] Claude Gérard, Didier Gonze, and Albert Goldbeter. Revisiting a skeleton model for the mammalian cell cycle: From bistability to Cdk oscillations and cellular heterogeneity. Journal of Theoretical Biology, 461(??):276–290, January 14, 2019.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305253.

Granda:2016:IPI

[GGGM16] Pablo Araujo Granda, Anna Gras, Marta Ginovart, and Vincent Moulton. INDISIM-Paracoccus, an individual-based and thermodynamic model for a denitrifying bacterium. Journal of Theoretical Biology, 403(??):45–58, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300959.

Gomes:2016:TFI

[GGL^{+}16] M. Gabriela M. Gomes, Erida Gjini, Joao S. Lopes, Caetano Souto-Maior, and Carlota Rebelo. A theoretical framework to identify invariant thresholds in infectious disease epidemiology. Journal of Theoretical Biology, 395(??):97–102, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000680.

Gonzalez:2012:ECL

[GGM12] M. González, C. Gutiérrez, and R. Martínez. Extinction conditions for y-linked mutant-alleles through two-sex branching processes with blind-mating structure. Journal of Theoretical Biology, 307(??):104–116, August 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002305.

Greenhalgh:2015:DER

[GGM15] Scott Greenhalgh, Alison P. Galvani, and Jan Medlock. Disease elimination and re-emergence in differential-equation models. Journal of Theoretical Biology, 387(??):174–180, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004877.

Gabriel:2012:CAS

[GGQ^{+}12] Pierre Gabriel, Shawn P. Garbett, Vito Quaranta, Darren R. Tyson, and Glenn F. Webb. The contribution of age structure to cell population responses to targeted therapeutics.

Journal of Theoretical Biology, 311(??):19–27, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312003359.

Gonzalez:2011:CCR

[GGR11] D. L. Gonzalez, S. Giannerini, and R. Rosa. Circular codes revisited: a statistical approach. Journal of Theoretical Biology, 275(1):21–28, April 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000403.

Guang:2010:PSC

[GGX^{+}10] Xuanmin Guang, Yanzhi Guo, Jiamin Xiao, Xia Wang, Jing Sun, Wenjia Xiong, and Menglong Li. Predicting the state of cysteines based on sequence information. Journal of Theoretical Biology, 267(3):312–318, December 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004686.

Goldwyn:2011:RME

[GH11] Eli E. Goldwyn and Alan Hastings. The roles of the Moran effect and dispersal in synchronizing oscillating populations. Journal of Theoretical Biology, 289(??):237–246, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931100436X.

Ghachem:2016:IBI

[Gha16] Montasser Ghachem. The institution as a blunt instrument: Cooperation through imperfect observability. Journal of Theoretical Biology, 396(??):182–190, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931600103X.

Gjini:2014:RDA

[GHBC14] Erida Gjini, Daniel T. Haydon, J. David Barry, and Christina A. Cobbold. Revisiting the diffusion approximation to estimate evolutionary rates of gene family diversification. Journal of Theoretical Biology, 341(??):111–122, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004931.

Ghanbari:2016:EGT

[GHBI16] Z. Ghanbari, M. R. Housaindokht, M. R. Bozorgmehr, and M. Izadyar. The effect of glycosylation on the transferrin structure: a molecular dynamic simulation analysis. Journal of Theoretical Biology, 404(??):73–81, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630114X.

Ghorbaniaghdam:2014:SSR

[GHJ14] Atefeh Ghorbaniaghdam, Olivier Henry, and Mario Jolicoeur. An in-silico study of the regulation of CHO cells glycolysis. Journal of Theoretical Biology, 357(??):112–122, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002653.

Gunther:2018:BMS

[GHS18] Michael Günther, Daniel F. B. Haeufle, and Syn Schmitt. The basic mechanical structure of the skeletal muscle machinery: One model for linking microscopic and macroscopic scales. Journal of Theoretical Biology, 456(??):137–167, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303473. See [GHS20].

Gunther:2020:CBM

[GHS20] Michael Günther, Daniel F. B. Haeufle, and Syn Schmitt. Corrigendum to “The basic mechanical structure of the skeletal muscle machinery: One model for linking microscopic and macroscopic scales” [Journal of Theoretical Biology 456 (2018) 137–167]. Journal of Theoretical Biology, 488(??):Article 110143, March 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319305132. See [GHS18].

Ghuchani:2018:CET

[Ghu18] M. S. Ghuchani. Comment on: An engineering thermodynamic approach to select the electromagnetic wave effective on cell growth. Journal of Theoretical Biology, 440(??):70, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730557X.

Giacche:2010:HIT

[GI10] Davide Giacché and Takuji Ishikawa. Hydrodynamic interaction of two unsteady model microorganisms. Journal of Theoretical Biology, 267(2):252–263, November 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004029.

Gibo:2015:DUM

[GI15] Shingo Gibo and Hiroshi Ito. Discrete and ultradiscrete models for biological rhythms comprising a simple negative feedback loop. Journal of Theoretical Biology, 378(??):89–95, August 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002003.

Giaimo:2013:SCM

[Gia13] Stefano Giaimo. Are some conventional measures of the rate of ageing consistent with antagonistic pleiotropy? Journal of Theoretical Biology, 323(??):20–24, April 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000027.

Ginn:2010:IGH

[Gin10] Brian R. Ginn. The implications of gene heterozygosity for protein folding and protein turnover. Journal of Theoretical Biology, 265(4):554–564, August 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002584.

Givli:2010:TMS

[Giv10] Sefi Givli. Towards multi-scale modeling of muscle fibers with sarcomere non-uniformities. Journal of Theoretical Biology, 264(3):882–892, June 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001244.

Ginsburg:2010:EAL

[GJ10] Simona Ginsburg and Eva Jablonka. The evolution of associative learning: a factor in the Cambrian explosion. Journal of Theoretical Biology, 266(1):11–20, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003073.

Gardner:2011:EMS

[GJ11] Carl L. Gardner and Jeremiah R. Jones. Electrodiffusion model simulation of the potassium channel. Journal of Theoretical Biology, 291(??):10–13, December 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004668.

Gaucherel:2012:OEN

[GJ12] Cédric Gaucherel and Henrik Jeldtoft Jensen. Origins of evolution: Non-acquired characters dominates over acquired characters in changing environment. Journal of Theoretical Biology, 304(??):111–120, July 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001130.

Ginsburg:2015:TTE

[GJ15a] Simona Ginsburg and Eva Jablonka. The teleological transitions in evolution: a Gántian view. Journal of Theoretical Biology, 381(??):55–60, September 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001691.

Golzari:2015:VBV

[GJ15b] Fahimeh Golzari and Saeed Jalili. VR-BFDT: a variance reduction based binary fuzzy decision tree induction method for protein function prediction. Journal of Theoretical Biology, 377(??):10–24, July 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001381.

Greenbury:2010:ESF

[GJS^{+}10] Sam F. Greenbury, Iain G. Johnston, Matthew A. Smith, Jonathan P. K. Doye, and Ard A. Louis. The effect of scale-free topology on the robustness and evolvability of genetic regulatory networks. Journal of Theoretical Biology, 267(1):48–61, November 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004054.

Golubski:2010:CMC

[GK10] Antonio J. Golubski and Christopher A. Klausmeier. Control in mutualisms: Combined implications of partner choice and bargaining roles. Journal of Theoretical Biology, 267(4):535–545, December 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004972.

Ghosh:2013:PEI

[GK13] Bapan Ghosh and T. K. Kar. Possible ecosystem impacts of applying maximum sustainable yield policy in food chain models. Journal of Theoretical Biology, 329(??):6–14, July 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313001185.

Gaither:2016:HTP

[GK16] Jeff Gaither and Laura Kubatko. Hypothesis tests for phylogenetic quartets, with applications to coalescentbased species tree inference. Journal of Theoretical Biology, 408(??):179–186, November 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302508.

Garnier:2017:ASP

[GK17] Simon Garnier and Daniel J. C. Kronauer. The adaptive significance of phasic colony cycles in army ants. Journal of Theoretical Biology, 428(??):43–47, September 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730187X.

Gussmann:2018:SCH

[GKG^{+}18] Maya Gussmann, Carsten Kirkeby, Kaare Græsbøll, Michael Farre, and Tariq Halasa. A strain-, cow-, and herd-specific bio-economic simulation model of intramammary infections in dairy cattle herds. Journal of Theoretical Biology, 449(??):83– 93, July 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301905.

Geritz:2013:ISI

[GKK13] Stefan A. H. Geritz, Éva Kisdi, and Christopher A. Klausmeier. Introduction to special issue. Journal of Theoretical Biology, 339(??):1–2, December 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004293.

Gaeta:2010:FCI

[GKMC10] Stephen A. Gaeta, Trine Krogh-Madsen, and David J. Christini. Feedback-control induced pattern formation in cardiac myocytes: a mathematical modeling study. Journal of Theoretical Biology, 266(3):408–418, October 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003383.

Georgiou:2010:SEC

[GKNT10] D. N. Georgiou, T. E. Karakasidis, Juan J. Nieto, and A. Torres. A study of entropy/clarity of genetic sequences using metric spaces and fuzzy sets. Journal of Theoretical Biology, 267(1):95–105, November 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004194.

Giorgi:2013:MSE

[GKPB13] Elena E. Giorgi, Bette T. Korber, Alan S. Perelson, and Tanmoy Bhattacharya. Modeling sequence evolution in HIV-1 infection with recombination. Journal of Theoretical Biology, 329(??):82–93, July 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001525.

Gyllenberg:2016:EPT

[GKW16] Mats Gyllenberg, Éva Kisdi, and Helene C. Weigang. On the evolution of patch-type dependent immigration. Journal of Theoretical Biology, 395(??):115–125, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000837.

Graham:2017:OCA

[GKWG17] Jason M. Graham, Albert B. Kao, Dylana A. Wilhelm, and Simon Garnier. Optimal construction of army ant living bridges. Journal of Theoretical Biology, 435(??):184–198, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304356.

Gecek:2012:IMS

[GL12a] Suncana Gecek and Tarzan Legović. Impact of maximum sustainable yield on competitive community. Journal of Theoretical Biology, 307(??):96–103, August 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002081.

Guo:2012:SRH

[GL12b] Daqing Guo and Chunguang Li. Stochastic resonance in Hodgkin–Huxley neuron induced by unreliable synaptic transmission. Journal of Theoretical Biology, 308(??):105–114, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002767.

Georgelin:2014:DCM

[GL14] E. Georgelin and N. Loeuille. Dynamics of coupled mutualistic and antagonistic interactions, and their implications for ecosystem management. Journal of Theoretical Biology, 346(??):67–74, April 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005638.

George:2018:SMT

[GL18a] Jason T. George and Herbert Levine. Stochastic modeling of tumor progression and immune evasion. Journal of Theoretical Biology, 458(??):148–155, December 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304430.

George:2018:TGM

[GL18b] Uduak Z. George and Sharon R. Lubkin. Tissue geometry may govern lung branching mode selection. Journal of Theoretical Biology, 442(??):22–30, April 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305787.

Glazier:2013:LTU

[Gla13] Douglas S. Glazier. Log-transformation is useful for examining proportional relationships in allometric scaling. Journal of Theoretical Biology, 334(??):200–203, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002890.

Greene:2015:MIH

[GLF^{+}15] James M. Greene, Doron Levy, King Leung Fung, Paloma S. Souza, Michael M. Gottesman, and Orit Lavi. Modeling intrinsic heterogeneity and growth of cancer cells. Journal of Theoretical Biology, 367(??):262–277, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006699.

Gobel:2010:CEM

[GLOC10] Britta Göbel, Dirk Langemann, Kerstin M. Oltmanns, and Matthias Chung. Compact energy metabolism model: Brain controlled energy supply. Journal of Theoretical Biology, 264(4):1214–1224, June 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001086.

Gamez:2017:EMD

[GLR^{+}17] M. Gámez, I. López, C. Rodríguez, Z. Varga, and J. Garay. Ecological monitoring in a discrete-time prey-predator model. Journal of Theoretical Biology, 429(??):52–60, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303028.

Gallaher:2018:MDK

[GLR^{+}18] Jill Gallaher, Kamila Larripa, Marissa Renardy, Blerta Shtylla, Nessy Tania, Diana White, Karen Wood, Li Zhu, Chaitali Passey, Michael Robbins, Natalie Bezman, Suresh Shelat, Hearn Jay Cho, and Helen Moore. Methods for determining key components in a mathematical model for tumor-immune dynamics in multiple myeloma. Journal of Theoretical Biology, 458(??):31–46, December 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304259.

Gong:2011:DED

[GLZ^{+}11] Binsheng Gong, Tao Liu, Xiaoyu Zhang, Xi Chen, Jiang Li, Hongchao Lv, Yi Zou, Xia Li, and Shaoqi Rao. Disease embryo development network reveals the relationship between disease genes and embryo development genes. Journal of Theoretical Biology, 287(??):100–108, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003687.

Gao:2013:SMJ

[GLZ^{+}13] Huijiang Gao, Yongxin Liu, Tingting Zhang, Runqing Yang, and Huanmin Yang. Statistical models for jointly analyzing multiple allometries. Journal of Theoretical Biology, 318(??):205–209, February 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005760.

Gonzalez-Miranda:2013:ECO

[GM13] J. M. González-Miranda. On the effect of circadian oscillations on biochemical cell signaling by NF-κB. Journal of Theoretical Biology, 335(??):283–294, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002993.

Goyal:2016:DVH

[GM16a] Ashish Goyal and John M. Murray. Dynamics of in vivo hepatitis D virus infection. Journal of Theoretical Biology, 398(??):9–19, June 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001594.

Grbatinic:2016:CAH

[GM16b] Ivan Grbatinić and Nebojsa Milosević. Classification of adult human dentate nucleus border neurons: Artificial neural networks and multidimensional approach. Journal of Theoretical Biology, 404(??):273–284, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630145X.

Goyal:2017:RCH

[GM17] Ashish Goyal and John M. Murray. Roadmap to control HBV and HDV epidemics in China. Journal of Theoretical Biology, 423(??):41–52, June 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301558.

Garira:2019:CMM

[GM19] Winston Garira and Dephney Mathebula. A coupled multiscale model to guide malaria control and elimination. Journal of Theoretical Biology, 475(??):34–59, August 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302103.

Gross:2011:MMA

[GMB11] Fridolin Groß, Gerhard Metzner, and Ulrich Behn. Mathematical modeling of allergy and specific immunotherapy: Th1-Th2-Treg interactions. Journal of Theoretical Biology, 269(1):70–78, January 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005412.

Gomez:2014:MSG

[GMBK14] David Gomez, Rahul Marathe, Veronika Bierbaum, and Stefan Klumpp. Modeling stochastic gene expression in growing cells. Journal of Theoretical Biology, 348(??):1–11, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000344.

Grbatinic:2018:NCP

[GMK18] Ivan Grbatinić, Nebojsa Milosević, and Bojana Krstonosić. The neuromorphological caudate-putaminal clustering of neostriate interneurons: Kohonen self-organizing maps and supervised artificial neural networks with multivariate analysis. Journal of Theoretical Biology, 438(??):96–115, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317305209.

Garcia-Martinez:2010:MRI

[GML10] Karina García-Martínez and Kalet León. Modeling the role of IL-2 in the interplay between CD4+ helper and regulatory T cells: Assessing general dynamical properties. Journal of Theoretical Biology, 262(4):720–732, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005062.

Gong:2013:PLN

[GMM^{+}13] Chang Gong, Joshua T. Mattila, Mark Miller, JoAnne L. Flynn, Jennifer J. Linderman, and D. Kirschner. Predicting lymph node output efficiency using systems biology. Journal of Theoretical Biology, 335(??):169–184, October 21, 2013.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002889.

Grbatinic:2015:NAH

[GMM15] Ivan Grbatinić, Dusica L. Marić, and Nebojsa T. Milosević. Neurons from the adult human dentate nucleus: Neural networks in the neuron classification. Journal of Theoretical Biology, 370(??):11–20, April 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000338.

Garcia-Manso:2012:LSL

[GMMGVD12] J. M. García-Manso, J. M. Martín-González, D. Vaamonde, and M. E. Da Silva-Grigoletto. The limitations of scaling laws in the prediction of performance in endurance events. Journal of Theoretical Biology, 300(??):324–329, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000380.

Grec:2018:MLA

[GMMN18] Bérénice Grec, Bertrand Maury, Nicolas Meunier, and Laurent Navoret. A 1D model of leukocyte adhesion coupling bond dynamics with blood velocity. Journal of Theoretical Biology, 452(??):35–46, September 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830081X.

Gong:2014:NID

[GMNY14] Xue Gong, Gregory Moses, Alexander B. Neiman, and Todd Young. Noise-induced dispersion and breakup of clusters in cell cycle dynamics. Journal of Theoretical Biology, 355(??):160–169, August 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001817.

Garijo:2012:SCA

[GMOP12] N. Garijo, R. Manzano, R. Osta, and M. A. Perez. Stochastic cellular automata model of cell migration, proliferation and differentiation: Validation with in vitro cultures of muscle satellite cells. Journal of Theoretical Biology, 314(??):1–9, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003906.

Gwiazda:2016:BIA

[GMR16] Piotr Gwiazda, Blazej Miasojedow, and Magdalena Rosińska. Bayesian inference for age-structured population model of infectious disease with application to varicella in Poland. Journal of Theoretical Biology, 407(??):38–50, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301916.

Gutting:2013:MLD

[GMS^{+}13] Bradford W. Gutting, David Marchette, Robert Sherwood, George A. Andrews, Alison Director-Myska, Stephen R. Channel, Daniel Wolfe, Alan E. Berger, Ryan S. Mackie, Brent J. Watson, and Andrey Rukhin. Modeling low-dose mortality and disease incubation period of inhalational anthrax in the rabbit. Journal of Theoretical Biology, 329(??):20–31, July 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313001240.

Ganjtabesh:2015:UTE

[GMZM15] Mohammad Ganjtabesh, Soheila Montaseri, and Fatemeh Zare-Mirakabad. Using temperature effects to predict the interactions between two RNAs. Journal of Theoretical Biology, 364(??):98–102, January 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005396.

Guedj:2010:UHC

[GN10] J. Guedj and A. U. Neumann. Understanding hepatitis C viral dynamics with direct-acting antiviral agents due to the interplay between intracellular replication and cellular infection dynamics. Journal of Theoretical Biology, 267(3):330–340, December 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004595.

Ghang:2014:SES

[GN14] Whan Ghang and Martin A. Nowak. Stochastic evolution of staying together. Journal of Theoretical Biology, 360(??):129–136, November 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003695.

Ghang:2015:IRO

[GN15] Whan Ghang and Martin A. Nowak. Indirect reciprocity with optional interactions. Journal of Theoretical Biology, 365(??):1–11, January 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005839.

Gullotto:2013:PPS

[GNB^{+}13] Danilo Gullotto, Mario Salvatore Nolassi, Andrea Bernini, Ottavia Spiga, and Neri Niccolai. Probing the protein space for extending the detection of weak homology folds. Journal of Theoretical Biology, 320(??):152–158, March 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006297.

Goel:2015:IRH

[Goe15] Pranay Goel. Insulin resistance or hypersecretion? The βIG picture revisited. Journal of Theoretical Biology, 384 (??):131–139, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519315003914.

Golubev:2010:EMG

[Gol10a] A. Golubev. Exponentially modified Gaussian (EMG) relevance to distributions related to cell proliferation and differentiation. Journal of Theoretical Biology, 262(2):257–266, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004809.

Golubev:2010:RDC

[Gol10b] A. Golubev. Random discrete competing events vs. dynamic bistable switches in cell proliferation in differentiation. Journal of Theoretical Biology, 267(3):341–354, December 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004558.

Golubev:2016:AIE

[Gol16] A. Golubev. Applications and implications of the exponentially modified gamma distribution as a model for time variabilities related to cell proliferation and gene expression. Journal of Theoretical Biology, 393(??):203–217, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000345.

Gonze:2013:MEC

[Gon13] Didier Gonze. Modeling the effect of cell division on genetic oscillators. Journal of Theoretical Biology, 325(??):22–33, May 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000593.

Gouin:2015:WTT

[Gou15] Henri Gouin. The watering of tall trees — embolization and recovery. Journal of Theoretical Biology, 369(??):42–50, March 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000181.

Goufo:2016:SCA

[Gou16] Emile Franc Doungmo Goufo. Stability and convergence analysis of a variable order replicator-mutator process in a moving medium. Journal of Theoretical Biology, 403(??):178–187, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300856.

Gatesy:2011:AOA

[GP11a] Stephen M. Gatesy and Nancy S. Pollard. Apples, oranges, and angles: Comparative kinematic analysis of disparate limbs. Journal of Theoretical Biology, 282(1):7–13, August 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002517.

Greenman:2011:PCR

[GP11b] J. V. Greenman and V. B. Pasour. Phase control of resonant systems: Interference, chaos and high periodicity. Journal of Theoretical Biology, 278(1):74–86, June 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001317.

Greenman:2012:TDP

[GP12] J. V. Greenman and V. B. Pasour. Threshold dynamics for periodically forced ecological systems: The control of population invasion and exclusion. Journal of Theoretical Biology, 295(??):154–167, February 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005856.

Gonzalez-Parra:2018:MFI

[GPD18] Gilberto González-Parra and Hana M. Dobrovolny. Modeling of fusion inhibitor treatment of RSV in African green monkeys. Journal of Theoretical Biology, 456(??):62–73, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303539.

Gramlich:2016:IDP

[GPR^{+}16] P. Gramlich, S. J. Plitzko, L. Rudolf, B. Drossel, and T. Gross. The influence of dispersal on a predator–prey system with two habitats. Journal of Theoretical Biology, 398(??):150–161, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001569.

Guiver:2017:NCD

[GPT17] Chris Guiver, David Packman, and Stuart Townley. A necessary condition for dispersal driven growth of populations with discrete patch dynamics. Journal of Theoretical Biology, 424(??):11–25, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931730156X.

Grafen:2014:TRV

[Gra14] Alan Grafen. Total reproductive values for females and for males in sexual diploids are not equal. Journal of Theoretical Biology, 359(??):233–235, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002926.

Grafen:2015:BFP

[Gra15] Alan Grafen. Biological fitness and the price equation in class-structured populations. Journal of Theoretical Biology, 373(??):62–72, May 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000739.

Grafen:2018:LHS

[Gra18] Alan Grafen. The left hand side of the fundamental theorem of natural selection. Journal of Theoretical Biology, 456(??):175–189, November 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303485.

Guedon:2013:PIC

[GRB^{+}13] Yann Guédon, Yassin Refahi, Fabrice Besnard, Etienne Farcot, Christophe Godin, and Teva Vernoux. Pattern identification and characterization reveal permutations of organs as a key genetically controlled property of post-meristematic phyllotaxis. Journal of Theoretical Biology, 338(??):94–110, December 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003494.

Goyens:2013:DTM

[GRBL13] J. Goyens, J. Reijniers, B. Borremans, and H. Leirs. Density thresholds for Mopeia virus invasion and persistence in its host Mastomys natalensis. Journal of Theoretical Biology, 317(??):55–61, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005176.

Gonzalez-Rodriguez:2016:TDP

[GRCdL16] David Gonzalez-Rodriguez, Paul-Henry Cournède, and Emmanuel de Langre. Turgidity-dependent petiole flexibility enables efficient water use by a tree subjected to water stress. Journal of Theoretical Biology, 398(??):20–31, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001521.

Greenbaum:2015:RTU

[Gre15] Gili Greenbaum. Revisiting the time until fixation of a neutral mutant in a finite population — a coalescent theory approach. Journal of Theoretical Biology, 380(??):98–102, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002477.

Gregorius:2016:ENP

[Gre16] Hans-Rolf Gregorius. Effective numbers in the partitioning of biological diversity. Journal of Theoretical Biology, 409(??):133–147, November 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931630279X.

Griswold:2011:MPB

[Gri11] Cortland K. Griswold. A model of the physiological basis of a multivariate phenotype that is mediated by Ca^{2+} signaling and controlled by ryanodine receptor composition. Journal of Theoretical Biology, 282(1):14–22, August 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002372.

Griesemer:2015:EVG

[Gri15a] James Griesemer. The enduring value of Gánti’s chemoton model and life criteria: Heuristic pursuit of exact theoretical biology. Journal of Theoretical Biology, 381(??):23–28, September 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002441.

Griswold:2015:AGV

[Gri15b] Cortland K. Griswold. Additive genetic variation and evolvability of a multivariate trait can be increased by epistatic gene action. Journal of Theoretical Biology, 387(??):241–257, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931500466X.

Griswold:2019:APS

[Gri19] Cortland K. Griswold. An ancestral process with selection in an ecological community. Journal of Theoretical Biology, 466(??):128–144, April 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318306325.

Grover:2017:SSV

[Gro17] James P. Grover. Sink or swim? Vertical movement and nutrient storage in phytoplankton. Journal of Theoretical Biology, 432(??):38–48, November 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303788.

Gutierrez:2014:DTA

[GRR^{+}14] Sonia L. Mestizo Gutiérrez, Marisol Herrera Rivero, Nicandro Cruz Ramírez, Elena Hernández, and Gonzalo E. Aranda-Abreu. Decision trees for the analysis of genes involved in Alzheimer’s disease pathology. Journal of Theoretical Biology, 357(??):21–25, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002690.

Gupta:2016:CAC

[GRRG16] Paras Gupta, Latha Rangan, T. Venkata Ramesh, and Mudit Gupta. Comparative analysis of contextual bias around the translation initiation sites in plant genomes. Journal of Theoretical Biology, 404(??):303–311, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301497.

Grunnill:2018:ERA

[Gru18] Martin Grunnill. An exploration of the role of asymptomatic infections in the epidemiology of dengue viruses through susceptible, asymptomatic, infected and recovered (SAIR) models. Journal of Theoretical Biology, 439(??):195–204, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305489.

Gunther:2010:MAD

[GS10] Michael Günther and Syn Schmitt. A macroscopic ansatz to deduce the Hill relation. Journal of Theoretical Biology, 263(4):407–418, April 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309006006.

Giacomini:2013:ARE

[GS13a] Henrique C. Giacomini and Brian J. Shuter. Adaptive responses of energy storage and fish life histories to climatic gradients. Journal of Theoretical Biology, 339(??):100–111, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003949. See corri-

gendum [GS14].

Gifford:2013:MCP

[GS13b] Danna R. Gifford and Sijmen E. Schoustra. Modelling colony population growth in the filamentous fungus Aspergillus nidulans. Journal of Theoretical Biology, 320(??):124–130, March 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312006273.

Giacomini:2014:CAR

[GS14] Henrique C. Giacomini and Brian J. Shuter. Corrigendum to “Adaptive responses of energy storage and fish life histories to climatic gradients” [J. Theor. Biol. 339 (2013) 100– 111]. Journal of Theoretical Biology, 363(??):454, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314005517. See [GS13a].

Gabora:2017:ANC

[GS17a] Liane Gabora and Mike Steel. Autocatalytic networks in cognition and the origin of culture. Journal of Theoretical Biology, 431(??):87–95, October 27, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303533.

Graham:2017:MOR

[GS17b] Erica J. Graham and James F. Selgrade. A model of ovulatory regulation examining the effects of insulin-mediated testosterone production on ovulatory function. Journal of Theoretical Biology, 416(??):149–160, March 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300073.

Givan:2011:PET

[GSCS11] Or Givan, Nehemia Schwartz, Assaf Cygelberg, and Lewi Stone. Predicting epidemic thresholds on complex networks: Limitations of mean-field approaches. Journal of Theoretical Biology, 288(??):21–28, November 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003651.

Gruter:2013:TPI

[GSF13] Christoph Grüter, Roger Schürch, and Walter M. Farina. Task-partitioning in insect societies: Non-random direct material transfers affect both colony efficiency and information flow. Journal of Theoretical Biology, 327(??):23–33, June 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000817.

Gill:2011:MOT

[GSGV11] J. S. Gill, C. M. Salafia, D. Grebenkov, and D. D. Vvedensky. Modeling oxygen transport in human placental terminal villi. Journal of Theoretical Biology, 291(??):33–41, December 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004644.

Giacomini:2013:PBP

[GSL13] Henrique C. Giacomini, Brian J. Shuter, and Nigel P. Lester. Predator bioenergetics and the prey size spectrum: Do foraging costs determine fish production? Journal of Theoretical Biology, 332(??):249–260, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002191.

Gunji:2011:ARB

[GSN^{+}11] Yukio-Pegio Gunji, Tomohiro Shirakawa, Takayuki Niizato, Masaki Yamachiyo, and Iori Tani. An adaptive and robust biological network based on the vacant-particle transportation model. Journal of Theoretical Biology, 272(1):187–200, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006685.

Galanter:2017:RCA

[GSRR17] Nina Galanter, Dennis Silva, Jonathan T. Rowell, and Jan Rychtár. Resource competition amid overlapping territories: The territorial raider model applied to multi-group interactions. Journal of Theoretical Biology, 412(??):100–106, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303393.

Gregg:2019:MMC

[GSS19] Robert W. Gregg, Saumendra N. Sarkar, and Jason E. Shoemaker. Mathematical modeling of the cGAS pathway reveals robustness of DNA sensing to TREX1 feedback. Journal of Theoretical Biology, 462(??):148–157, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305484.

Ganz:2018:TTA

[GSSBF18] Dan E. Ganz, Briana Sexton-Stallone, Emily L. Brackett, and Neil S. Forbes. Tissue transport affects how treatment scheduling increases the efficacy of chemotherapeutic drugs. Journal of Theoretical Biology, 438(??):21–33, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304861.

Grasman:2011:SML

[GSV11] J. Grasman, H. M. Salomons, and S. Verhulst. Stochastic modeling of length-dependent telomere shortening in Corvus monedula. Journal of Theoretical Biology, 282(1):1–6, August 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311002268.

Grobler:2016:MND

[GSV16] A. Grobler, Z. Perez Sierra, and H. J. Viljoen. Modeling nanoparticle delivery of TB drugs to granulomas. Journal of Theoretical Biology, 388(??):85–95, January 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500497X.

Shen:2012:LEA

[gSxFH^{+}12] Zhi gang Shen, Qi xue Fan, Margaret Anne Hurley, Cong xin Xie, Wei Yang, and Yun long Zhang. A letter to the editor about the article “a model for environmental sex reversal in fish”. Journal of Theoretical Biology, 294(??):185, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005522.

Gehring:2010:NBA

[GSYS10] Ronette Gehring, Phillip Schumm, Mina Youssef, and Caterina Scoglio. A network-based approach for resistance transmission in bacterial populations. Journal of Theoretical Biology, 262(1):97–106, January 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004226.

Garcia:2011:MMA

[GT11a] Jacqueline García and Néstor Torres. Mathematical modelling and assessment of the pH homeostasis mechanisms in Aspergillus niger while in citric acid producing conditions. Journal of Theoretical Biology, 282(1):23–35, August 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002281.

Gokhale:2011:SAE

[GT11b] Chaitanya S. Gokhale and Arne Traulsen. Strategy abundance in evolutionary many-player games with multiple strategies. Journal of Theoretical Biology, 283(1):180–191, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002839.

Garcia:2012:LLA

[GT12] Julián García and Arne Traulsen. Leaving the loners alone: Evolution of cooperation in the presence of antisocial punishment. Journal of Theoretical Biology, 307(??):168–173, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002457.

Guo:2014:FTD

[GT14] Ya Guo and Jinglu Tan. Fourier transform of delayed fluorescence as an indicator of herbicide concentration. Journal of Theoretical Biology, 363(??):271–276, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004780.

Grieshaber:2015:CMI

[GT15] Beverley J. Grieshaber and John S. Terblanche. A computational model of insect discontinuous gas exchange: a two-sensor, control systems approach. Journal of Theoretical Biology, 374(??):138–151, June 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001459.

Gin:2010:MCL

[GTB10] Elan Gin, Elly M. Tanaka, and Lutz Brusch. A model for cyst lumen expansion and size regulation via fluid secretion. Journal of Theoretical Biology, 264(3):1077–1088, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931000144X.

Ghosh:2019:EAC

[GTC19] Indrajit Ghosh, Pankaj Kumar Tiwari, and Joydev Chattopadhyay. Effect of active case finding on dengue control: Implications from a mathematical model. Journal of Theoretical Biology, 464(??):50–62, March 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306271.

Ganguly:2015:HMF

[GTS15] Bhaskar Ganguly, Kamal Tewari, and Rashmi Singh. Homology modeling, functional annotation and comparative genomics of outer membrane protein H of Pasteurella multocida. Journal of Theoretical Biology, 386(??):18–24, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004294.

Gorban:2016:EAM

[GTSP16] Alexander N. Gorban, Tatiana A. Tyukina, Elena V. Smirnova, and Lyudmila I. Pokidysheva. Evolution of adaptation mechanisms: Adaptation energy, stress, and oscillating death. Journal of Theoretical Biology, 405(??):127– 139, September 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316000205.

Gundersen:2013:SMR

[GU13] Cameron B. Gundersen and Joy A. Umbach. Synaptotagmins 1 and 2 as mediators of rapid exocytosis at nerve terminals: The dyad hypothesis. Journal of Theoretical Biology, 332(??):149–160, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001975.

Gunel:2012:MIB

[Gün12] Aylin Günel. Modelling the interactions between TLR4 and IFNβ pathways. Journal of Theoretical Biology, 307 (??):137–148, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251931200207X.

See erratum [Gün13].

Gunel:2013:EMI

[Gün13] Aylin Günel. Erratum to “Modelling the interactions between TLR4 and IFNβ pathways” [J. Theor. Biol. 307 (2012) 137–148]. Journal of Theoretical Biology, 322(??):94, April 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312006510. See [Gün12].

Garay:2012:ONF

[GVCG12] József Garay, Zoltán Varga, Tomás Cabello, and Manuel Gámez. Optimal nutrient foraging strategy of an omnivore: Liebig’s law determining numerical response. Journal of Theoretical Biology, 310(??):31–42, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003086.

Garay:2015:FRP

[GVGC15] József Garay, Zoltán Varga, Manuel Gámez, and Tomás Cabello. Functional response and population dynamics for fighting predator, based on activity distribution. Journal of Theoretical Biology, 368(??):74–82, March 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314007097.

Gjini:2016:HDC

[GVSLG16] Erida Gjini, Carina Valente, Raquel Sá-Leão, and M. Gabriela M. Gomes. How direct competition shapes coexistence and vaccine effects in multi-strain pathogen systems. Journal of Theoretical Biology, 388(??):50–60, January 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004816.

Garcia:2014:EGB

[GvVT14] Julián García, Matthijs van Veelen, and Arne Traulsen. Evil green beards: Tag recognition can also be used to withhold cooperation in structured populations. Journal of Theoretical Biology, 360(??):181–186, November 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003956.

Gunther:2012:CHP

[GW12] Michael Günther and Tom Weihmann. Climbing in hexapods: a plain model for heavy slopes. Journal of Theoretical Biology, 293(??):82–86, January 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005261.

Gibson:2013:IBC

[GW13] William T. Gibson and William G. Wilson. Individual-based chaos: Extensions of the discrete logistic model. Journal of Theoretical Biology, 339(??):84–92, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004220.

Grover:2014:CAR

[GW14] James P. Grover and Feng-Bin Wang. Competition and allelopathy with resource storage: Two resources. Journal of Theoretical Biology, 351(??):9–24, June 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000885.

Gan:2015:HHS

[GW15] Xinjun Gan and David Waxman. Hard harvesting of a stochastically changing population. Journal of Theoretical Biology, 380(??):559–568, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002945.

Gulbudak:2019:HVS

[GW19] Hayriye Gulbudak and Joshua S. Weitz. Heterogeneous viral strategies promote coexistence in virus-microbe systems. Journal of Theoretical Biology, 462(??):65–84, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305381.

Gilbert:2014:SSS

[GWBG14] Mark A. Gilbert, Steven M. White, James M. Bullock, and Eamonn A. Gaffney. Spreading speeds for stage structured plant populations in fragmented landscapes. Journal of Theoretical Biology, 349(??):135–149, May 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000411.

Galante:2012:MLI

[GWBL12] Amanda Galante, Susanne Wisen, Devaki Bhaya, and Doron Levy. Modeling local interactions during the motion of cyanobacteria. Journal of Theoretical Biology, 309(??):147–158, September 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003001.

Ghosh:2014:CAA

[GWCA14] Suma Ghosh, Jessica L. Waite, Dale H. Clayton, and Frederick R. Adler. Can antibodies against flies alter malaria transmission in birds by changing vector behavior? Journal of Theoretical Biology, 358(??):93–101, October 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002914.

Gompert:2011:HPM

[GWE11] Zachariah Gompert, Keith Willmott, and Marianne Elias. Heterogeneity in predator micro-habitat use and the maintenance of Müllerian mimetic diversity. Journal of Theoretical Biology, 281(1):39–46, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002244.

Gao:2015:PHR

[GWNW15] Shiping Gao, Te Wu, Suli Nie, and Long Wang. Promote or hinder? The role of punishment in the emergence of cooperation. Journal of Theoretical Biology, 386(??):69–77, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931500452X.

Green:2010:NLM

[GWW^{+}10] J. E. F. Green, S. L. Waters, J. P. Whiteley, L. Edelstein-Keshet, K. M. Shakesheff, and H. M. Byrne. Non-local models for the formation of hepatocyte-stellate cell aggregates. Journal of Theoretical Biology, 267(1):106–120, November 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310004224.

Gong:2018:GRM

[GWW^{+}18] Sha Gong, Yanli Wang, Zhen Wang, Yujie Wang, and Wenbing Zhang. Genetic regulation mechanism of the yjdF riboswitch. Journal of Theoretical Biology, 439(??):152–159, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305404.

Guo:2017:MPM

[GWX17] Si-Kao Guo, Peng-Ye Wang, and Ping Xie. A model of processive movement of dimeric kinesin. Journal of Theoretical Biology, 414(??):62–75, February 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304027.

Govan:2013:CSN

[GXF13] G. Govan, A. Xenos, and P. Frisco. A critical study of network models for neural networks and their dynamics. Journal of Theoretical Biology, 336(??):1–10, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003263.

Gao:2013:PCC

[GXFF13] Shan Gao, Shuo Xu, Yaping Fang, and Jianwen Fang. Prediction of core cancer genes using multi-task classification framework. Journal of Theoretical Biology, 317(??):62–70, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200505X.

Grinevich:2018:IDT

[GY18] Andrey A. Grinevich and Ludmila V. Yakushevich. The influence of the DNA torque on the dynamics of transcription bubbles in plasmid PTTQ18. Journal of Theoretical Biology, 453(??):68–77, September 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302194.

Ghosh:2019:HBA

[GY19] Avirup Ghosh and Hong Yan. Hydrogen bond analysis of the EGFR-ErbB3 heterodimer related to non-small cell lung cancer and drug resistance. Journal of Theoretical Biology, 464(??):63–71, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318306350.

Gao:2010:ITI

[GYWJ10] Tingting Gao, Zhixia Yang, Yong Wang, and Ling Jing. Identifying translation initiation sites in prokaryotes using support vector machine. Journal of Theoretical Biology, 262(4):644–649, February 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005049.

Ge:2014:RIH

[GZ14a] Huihua Ge and Guangya Zhang. RETRACTED: Identifying halophilic proteins based on random forests with preprocessing of the pseudo-amino acid composition. Journal of Theoretical Biology, 361(??):175, November 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004147.

Ge:2014:RNI

[GZ14b] Huihua Ge and Guangya Zhang. Retraction notice to “Identifying halophilic proteins based on random forests with preprocessing of the pseudo-amino acid composition”

[J. Theor. Biol. 361 (2014) 175–181]. Journal of Theoretical Biology, 361(??):R1, November 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006109.

Guo:2019:MAT

[GZ19] Yixing Guo and Ying Zhou. A modified association test for rare and common variants based on affected sib-pair design. Journal of Theoretical Biology, 467(??):1–6, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300141.

Gao:2014:SAP

[GZFX14] Kuikui Gao, Xu Zhang, Enmin Feng, and Zhilong Xiu. Sensitivity analysis and parameter identification of nonlinear hybrid systems for glycerol transport mechanisms in continuous culture. Journal of Theoretical Biology, 347(??):137–143, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005900.

Garamszegi:2012:RBS

[GZT12] László Zsolt Garamszegi, Sándor Zsebők, and János Török. The relationship between syllable repertoire similarity and pairing success in a passerine bird species with complex song. Journal of Theoretical Biology, 295(??):68–76, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931100573X.

Guo:2015:WAP

[GZT15] Ya Guo, Yesen Zhou, and Jinglu Tan. Wavelet analysis of pulse-amplitude-modulated chlorophyll fluorescence for differentiation of plant samples. Journal of Theoretical Biology, 370(??):116–120, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000582.

Guan:2012:RBG

[GZY12] Pearl P. Guan, Weiqiang Zhou, and Hong Yan. The relationship between geometric patterns of hydrogen bonds and periodic dinucleotides in nucleosome structures. Journal of Theoretical Biology, 313(??):136–141, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200433X.

Hashim:2015:RMD

[HA15a] Ezzeddin Kamil Mohamed Hashim and Rosni Abdullah. Rare k-mer DNA: Identification of sequence motifs and prediction of CpG island and promoter. Journal of Theoretical Biology, 387(??):88–100, December 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004579.

Hirata:2015:AIA

[HA15b] Yoshito Hirata and Kazuyuki Aihara. Ability of intermittent androgen suppression to selectively create a non-trivial periodic orbit for a type of prostate cancer patients. Journal of Theoretical Biology, 384(??):147–152, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004038.

Hartvig:2011:FWF

[HAB11] Martin Hartvig, Ken H. Andersen, and Jan E. Beyer. Food web framework for size-structured populations. Journal of Theoretical Biology, 272(1):113–122, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006612.

Haseeb:2017:PRU

[HAB^{+}17] Muhammad Haseeb, Shumaila Azam, A. I. Bhatti, Rizwan Azam, Mukhtar Ullah, and Sahar Fazal. On p53 revival using system oriented drug dosage design. Journal of Theoretical Biology, 415(??):53–57, February 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304167.

Hendrickx:2019:IBI

[HAH19] Diana M. Hendrickx, Steven Abrams, and Niel Hens. The impact of behavioral interventions on co-infection dynamics: An exploration of the effects of home isolation. Journal of Theoretical Biology, 476(??):5–18, September 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302152.

Hassan:2019:ESP

[HAK^{+}19] Md. Nazmul Hassan, Lale Asik, Jackson Kulik, K. R. Long, and Angela Peace. Environmental seasonality on predatorprey systems under nutrient and toxicant constraints. Journal of Theoretical Biology, 480(??):71–80, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303078.

Hallinan:2012:GTV

[Hal12] Nathaniel Hallinan. The generalized time variable reconstructed birth-death process. Journal of Theoretical Biology, 300(??):265–276, May 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000616.

Hald:2016:GMA

[Hal16] Bjørn Olav Hald. A generative modeling approach to connectivity — electrical conduction in vascular networks. Journal of Theoretical Biology, 399(??):1–12, June 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300029.

Hala:2017:SPR

[Hal17] D. Hala. In silico predicted reproductive endocrine transcriptional regulatory networks during zebrafish (Danio rerio) development. Journal of Theoretical Biology, 417(??):51–60, March 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300279.

Hansma:2010:POL

[Han10] Helen Greenwood Hansma. Possible origin of life between mica sheets. Journal of Theoretical Biology, 266(1):175–188, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003061.

Hanert:2012:FDT

[Han12] Emmanuel Hanert. Front dynamics in a two-species competition model driven by Lévy flights. Journal of Theoretical Biology, 300(??):134–142, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931200032X.

Hansen:2017:DMF

[Han17] Thomas F. Hansen. On the definition and measurement of fitness in finite populations. Journal of Theoretical Biology, 419(??):36–43, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316304350.

Hamelin:2016:EPV

[HAP^{+}16] Frédéric M. Hamelin, Linda J. S. Allen, Holly R. Prendeville, M. Reza Hajimorad, and Michael J. Jeger. The evolution of plant virus transmission pathways. Journal of Theoretical Biology, 396(??):75–89, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001132.

Hartung:2015:CMF

[Har15] John Hartung. Cancer might be a failed response to renegade mitochondria. Journal of Theoretical Biology, 379(??):91–93, August 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002234.

Hashimoto:2014:MEF

[Has14] Koh Hashimoto. Multigame effect in finite populations induces strategy linkage between two games. Journal of Theoretical Biology, 345(??):70–77, March 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005584.

Hiebeler:2017:EMI

[HASM17] David E. Hiebeler, Andrew Audibert, Emma Strubell, and Isaac J. Michaud. An epidemiological model of Internet worms with hierarchical dispersal and spatial clustering of hosts. Journal of Theoretical Biology, 418(??):8–15, April 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300334.

Hauert:2010:RDR

[Hau10] Christoph Hauert. Replicator dynamics of reward and reputation in public goods games. Journal of Theoretical Biology, 267(1):22–28, November 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004182.

Hassan:2018:ESM

[HAuR^{+}18] Mubashir Hassan, Muhammad Athar Abbasi, Aziz ur Rehman, Sabahat Zahra Siddiqui, Ghulam Hussain, Syed Adnan Ali Shah, Muhammad Shahid, and Sung-Yum Seo. Exploration of synthetic multifunctional amides as new therapeutic agents for Alzheimer’s disease through enzyme inhibition, chemoinformatic properties, molecular docking and dynamic simulation insights. Journal of Theoretical Biology, 458(??):169–183, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304545.

Hayamizu:2016:EIM

[Hay16] Momoko Hayamizu. On the existence of infinitely many universal tree-based networks. Journal of Theoretical Biology, 396(??):204–206, May 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001193.

Hernandez-Bermejo:2010:RGA

[HB10] Benito Hernández-Bermejo. Renormalization group approach to power-law modeling of complex metabolic networks. Journal of Theoretical Biology, 265(3):422–432, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002158.

Hambli:2012:PBF

[HB12a] Ridha Hambli and Abdelwahed Barkaoui. Physically based 3d finite element model of a single mineralized collagen microfibril. Journal of Theoretical Biology, 301(??):28–41, May 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000781.

Hordijk:2012:DRR

[HB12b] Wim Hordijk and Olivier Broennimann. Dispersal routes reconstruction and the minimum cost arborescence problem. Journal of Theoretical Biology, 308(??):115–122, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312002846.

Holt:2013:DPP

[HB13a] Robert D. Holt and Michael Barfield. Direct plant-predator interactions as determinants of food chain dynamics. Journal of Theoretical Biology, 339(??):47–57, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002026.

Hubbard:2013:MMV

[HB13b] M. E. Hubbard and H. M. Byrne. Multiphase modelling of vascular tumour growth in two spatial dimensions. Journal of Theoretical Biology, 316(??):70–89, January 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005097.

Hedrih:2016:EFI

[HB16a] Andjelka Hedrih and Milan Banić. The effect of friction and impact angle on the spermatozoa-oocyte local contact dynamics. Journal of Theoretical Biology, 393(??):32–42, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000394.

Hendron:2016:IVV

[HB16b] Ross-William S. Hendron and Michael B. Bonsall. The interplay of vaccination and vector control on small dengue networks. Journal of Theoretical Biology, 407(??):349–361, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302181.

Horvath:2018:NCF

[HB18] D. Horvath and B. Brutovsky. A new conceptual framework for the therapy by optimized multidimensional pulses of therapeutic activity. The case of multiple myeloma model. Journal of Theoretical Biology, 454(??):292–309, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303096.

Hirata:2010:DMM

[HBA10] Yoshito Hirata, Nicholas Bruchovsky, and Kazuyuki Aihara. Development of a mathematical model that predicts the outcome of hormone therapy for prostate cancer. Journal of Theoretical Biology, 264(2):517–527, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001025.

Huard:2017:MID

[HBA17] B. Huard, A. Bridgewater, and M. Angelova. Mathematical investigation of diabetically impaired ultradian oscillations in the glucose-insulin regulation. Journal of Theoretical Biology, 418(??):66–76, April 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300395.

Harding:2012:IMH

[HBEW12] K. C. Harding, M. Begon, A. Eriksson, and B. Wennberg. Increased migration in host-pathogen metapopulations can cause host extinction. Journal of Theoretical Biology, 298(??): 1–7, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311006187.

Hadjichrysanthou:2012:AED

[HBK12] Christoforos Hadjichrysanthou, Mark Broom, and Istvan Z. Kiss. Approximating evolutionary dynamics on networks using a neighbourhood configuration model. Journal of Theoretical Biology, 312(??):13–21, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003591.

Hoppitt:2010:DST

[HBL10] William Hoppitt, Neeltje J. Boogert, and Kevin N. Laland. Detecting social transmission in networks. Journal of Theoretical Biology, 263(4):544–555, April 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000081.

Howells:2011:BAD

[HBSF11] Christopher C. Howells, William T. Baumann, David C. Samuels, and Carla V. Finkielstein. The bcl-2-associated death promoter (BAD) lowers the threshold at which the bcl-2-interacting domain death agonist (BID) triggers mitochondria disintegration. Journal of Theoretical Biology, 271(1):114–123, February 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310006351.

Henry:2013:EED

[HBT13] Roslyn C. Henry, Greta Bocedi, and Justin M. J. Travis. Ecoevolutionary dynamics of range shifts: Elastic margins and critical thresholds. Journal of Theoretical Biology, 321(??):1– 7, March 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312006285.

Hartzell:2015:NDC

[HBVP15] Samantha Hartzell, Mark S. Bartlett, Lawrence Virgin, and Amilcare Porporato. Nonlinear dynamics of the CAM circadian rhythm in response to environmental forcing. Journal of Theoretical Biology, 368(??):83–94, March 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314007073.

Hu:2011:IDA

[HBW^{+}11] S. Hu, T. Biben, X. Wang, P. Jurdic, and J.-C. Géminard. Internal dynamics of actin structures involved in the cell motility and adhesion: Modeling of the podosomes at the molecular level. Journal of Theoretical Biology, 270(1):25–30, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005990.

Handelman:2017:CWA

[HC17] Tomer Handelman and Benny Chor. Cases in which ancestral maximum likelihood will be confusingly misleading. Journal of Theoretical Biology, 420(??):318–323, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301133.

Hala:2018:SPT

[HCHI18] D. Hala, J. A. Cullen, B. Hernout, and I. Ivanov. In silico predicted transcriptional regulatory control of steroidogenesis in spawning female fathead minnows (Pimephales promelas). Journal of Theoretical Biology, 455(??):179–190, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303448.

Havelka:2011:HFE

[HCK^{+}11] D. Havelka, M. Cifra, O. Kucera, J. Pokorný, and J. Vrba. High-frequency electric field and radiation characteristics of cellular microtubule network. Journal of Theoretical Biology, 286(??):31–40, October 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003511.

Hrycik:2019:MRS

[HCS^{+}19] Allison R. Hrycik, Paris D. Collingsworth, Timothy M. Sesterhenn, Daisuke Goto, and Tomas O. Höök. Movement rule selection through eco-genetic modeling: Application to diurnal vertical movement. Journal of Theoretical Biology, 478(??):128–138, October 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302541.

Huang:2018:ECH

[HCW18] Feng Huang, Xiaojie Chen, and Long Wang. Evolution of cooperation in a hierarchical society with corruption control. Journal of Theoretical Biology, 449(??):60–72, July 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301784.

Herron:2011:ADP

[HD11] Matthew D. Herron and Michael Doebeli. Adaptive diversification of a plastic trait in a predictably fluctuating environment. Journal of Theoretical Biology, 285(1):58–68, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003067.

Hamm:2017:HHH

[HD17] Michaela Hamm and Barbara Drossel. Habitat heterogeneity hypothesis and edge effects in model metacommunities. Journal of Theoretical Biology, 426(??):40–48, August 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302308.

Horvat:2010:SEM

[HDNB10] Sz. Horvát, A. Derzsi, Z. Néda, and A. Balog. A spatially explicit model for tropical tree diversity patterns. Journal of Theoretical Biology, 265(4):517–523, August 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002754.

Hu:2019:OMP

[HDZ^{+}19] Bing Hu, Xiyezi Diao, Liqing Zong, Jingsong Zhang, Liang Yu, Hao Dai, and Luonan Chen. The onset mechanism of Parkinson’s beta oscillations: a theoretical analysis. Journal of Theoretical Biology, 470(??):1–16, June 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301092.

Hayashi:2016:EAP

[HE16] Michael A. L. Hayashi and Marisa C. Eisenberg. Effects of adaptive protective behavior on the dynamics of sexually transmitted infections. Journal of Theoretical Biology, 388(??):119–130, January 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004233.

Hochard:2017:CJM

[HF17] Jacob Hochard and David Finnoff. Cross-jurisdictional management of a trophy-hunted species. Journal of Theoretical Biology, 420(??):41–52, May 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300565.

Hook:2017:FAP

[HFD17] Sarah E. Hook, James France, and Jan Dijkstra. Further assessment of the protozoal contribution to the nutrition of the ruminant animal. Journal of Theoretical Biology, 416(??):8– 15, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316304192.

Hamlet:2015:EIM

[HFT15] Christina Hamlet, Lisa J. Fauci, and Eric D. Tytell. The effect of intrinsic muscular nonlinearities on the energetics of locomotion in a computational model of an anguilliform swimmer. Journal of Theoretical Biology, 385(??):119–129, November 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004245.

He:2018:AEG

[HFT^{+}18] Qiao-Qiao He, Tian-Jiao Feng, Yi Tao, Boyu Zhang, and Ting Ji. Asymmetric evolutionary game dynamics based on individuals’ own volition. Journal of Theoretical Biology, 454(??):118–125, October 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302479.

Hettich:2018:TFT

[HG18a] J. Hettich and J. C. M. Gebhardt. Transcription factor target site search and gene regulation in a background of unspecific binding sites. Journal of Theoretical Biology, 454(??):91–101, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302844.

Hickey:2018:ARW

[HG18b] Donal A. Hickey and G. Brian Golding. The advantage of recombination when selection is acting at many genetic loci. Journal of Theoretical Biology, 442(??):123–128, April 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300341.

Herschlag:2013:MMA

[HGB^{+}13] Gregory Herschlag, Guilherme J. M. Garcia, Brian Button, Robert Tarran, Brandon Lindley, Benjamin Reinhardt, Timothy C. Elston, and M. Gregory Forest. A mechanochemical model for auto-regulation of lung airway surface layer volume. Journal of Theoretical Biology, 325(??):42–51, May 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000489.

Hernandez-Gascon:2013:EMC

[HGGCR13] B. Hernández-Gascón, J. Grasa, B. Calvo, and J. F. Rodríguez. A 3D electro-mechanical continuum model for simulating skeletal muscle contraction. Journal of Theoretical Biology, 335(??):108–118, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519313003019. See corrigendum [LHPF18].

Hauzy:2010:DDD

[HGHL10] Céline Hauzy, Mathias Gauduchon, Florence D. Hulot, and Michel Loreau. Density-dependent dispersal and relative dispersal affect the stability of predator–prey metacommunities. Journal of Theoretical Biology, 266(3):458–469, October 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003577.

Hua:2010:CGE

[HgLL^{+}10] Lin Hua, Dong guo Li, Hui Lin, Lin Li, Xia Li, and Zhi cheng Liu. The correlation of gene expression and co-regulated gene patterns in characteristic KEGG pathways. Journal of Theoretical Biology, 266(2):242–249, September 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000319X.

Hadaeghi:2015:TCS

[HGM15] Fatemeh Hadaeghi, Mohammad Reza Hashemi Golpayegani, and Greg Murray. Towards a complex system understanding of bipolar disorder: a map based model of a complex winnerless competition. Journal of Theoretical Biology, 376(??):74– 81, July 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931500079X.

Hutchinson:2016:VPI

[HGM^{+}16] L. G. Hutchinson, E. A. Gaffney, P. K. Maini, J. Wagg, A. Phipps, and H. M. Byrne. Vascular phenotype identification and anti-angiogenic treatment recommendation: a pseudo-multiscale mathematical model of angiogenesis. Journal of Theoretical Biology, 398(??):162–180, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001429.

Hoffman:2018:MMA

[HGO^{+}18] F. Hoffman, D. Gavaghan, J. Osborne, I. P. Barrett, T. You, H. Ghadially, R. Sainson, R. W. Wilkinson, and H. M. Byrne. A mathematical model of antibodydependent cellular cytotoxicity (ADCC). Journal of Theoretical Biology, 436(??):39–50, January 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304551.

Hindle:2010:EBH

[HH10] Allyson G. Hindle and Markus Horning. Energetics of breath-hold hunting: Modeling the effects of aging on foraging success in the Weddell seal. Journal of Theoretical Biology, 264(3):673–682, June 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001803.

Huang:2013:IRI

[HH13] Shanran Huang and Jitao T. Huang. Inter-residue interaction is a determinant of protein folding kinetics. Journal of Theoretical Biology, 317(??):224–228, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005280.

Hala:2014:SPS

[HH14] D. Hala and D. B. Huggett. In silico predicted structural and functional robustness of piscine steroidogenesis. Journal of Theoretical Biology, 345(??):99–108, March 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005547.

Hiebeler:2016:LDPa

[HH16] David E. Hiebeler and Jack L. Hill. Locally dispersing populations in heterogeneous dynamic landscapes with spatiotemporal correlations. II. habitat driven by voter dynamics. Journal of Theoretical Biology, 407(??):81–89, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302144.

Heijne:2013:CPR

[HHA^{+}13] Janneke C. M. Heijne, Sereina A. Herzog, Christian L. Althaus, Nicola Low, and Mirjam Kretzschmar. Case and partnership reproduction numbers for a curable sexually transmitted infection. Journal of Theoretical Biology, 331(??):38–47, August 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931300163X.

Hirashima:2017:VTM

[HHA17] Tsuyoshi Hirashima, Miho Hoshuyama, and Taiji Adachi. In vitro tubulogenesis of Madin–Darby canine kidney (MDCK) spheroids occurs depending on constituent cell number and scaffold gel concentration. Journal of Theoretical Biology, 435(??):110–115, December 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304277.

Hiebeler:2016:LDPb

[HHD^{+}16] David E. Hiebeler, Jennifer Houle, Frank Drummond, Peter Bilodeau, and Jeffery Merckens. Locally dispersing populations in heterogeneous dynamic landscapes with spatiotemporal correlations. I. Block disturbance. Journal of Theoretical Biology, 407(??):212–224, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302156.

Ho:2019:MMT

[HHE19] Shing Hei Ho, Daihai He, and Raluca Eftimie. Mathematical models of transmission dynamics and vaccine strategies in Hong Kong during the 2017–2018 winter influenza season. Journal of Theoretical Biology, 476(??):74–94, September 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302127.

Hughes:2011:MCA

[HHF11] John Hughes, William O. Hancock, and John Fricks. A matrix computational approach to kinesin neck linker extension. Journal of Theoretical Biology, 269(1):181–194, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005333.

Harvey:2014:QED

[HHG14] Emily Harvey, Jeffrey Heys, and Tomás Gedeon. Quantifying the effects of the division of labor in metabolic pathways. Journal of Theoretical Biology, 360(??):222–242, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004044.

Hakkinen:2011:GWS

[HHJR11] Antti Häkkinen, Shannon Healy, Howard T. Jacobs, and Andre S. Ribeiro. Genome wide study of NF-y type CCAAT boxes in unidirectional and bidirectional promoters in human and mouse. Journal of Theoretical Biology, 281(1):74– 83, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100227X.

Haber:2019:WDM

[HHMT19] Shimon Haber, Zeev Holtzman, Steven J. Mentzer, and Akira Tsuda. Why do myofibroblasts preferentially accumulate on the convex surface of the remodeling lung after pneumonectomy? Journal of Theoretical Biology, 479(??):90–96, October 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931930267X.

Humplik:2014:EDI

[HHN14] Jan Humplik, Alison L. Hill, and Martin A. Nowak. Evolutionary dynamics of infectious diseases in finite populations. Journal of Theoretical Biology, 360(??):149–162, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314003907.

Hummert:2010:GTM

[HHS^{+}10] S. Hummert, C. Hummert, A. Schröter, B. Hube, and S. Schuster. Game theoretical modelling of survival strategies of Candida albicans inside macrophages. Journal of Theoretical Biology, 264(2):312–318, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000342.

Hatano:2015:CBM

[HHSA15] Takuma Hatano, Yoshito Hirata, Hideyuki Suzuki, and Kazuyuki Aihara. Comparison between mathematical models of intermittent androgen suppression for prostate cancer. Journal of Theoretical Biology, 366(??):33–45, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314006249.

Hu:2019:WSD

[HHT^{+}19] Linchao Hu, Mugen Huang, Moxun Tang, Jianshe Yu, and Bo Zheng. Wolbachia spread dynamics in multiregimes of environmental conditions. Journal of Theoretical Biology, 462(??):247–258, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305629.

Hauert:2012:EGD

[HI12] Christoph Hauert and Lorens A. Imhof. Evolutionary games in deme structured, finite populations. Journal of Theoretical Biology, 299(??):106–112, April 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003092.

Hara:2017:WAI

[HI17] Akane Hara and Yoh Iwasa. When is allergen immunotherapy effective? Journal of Theoretical Biology, 425(??):23–42, July 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730200X.

Hara:2019:CDI

[HI19] Akane Hara and Yoh Iwasa. Coupled dynamics of intestinal microbiome and immune system — a mathematical study. Journal of Theoretical Biology, 464(??):9–20, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306179.

Hara:2019:RUI

[HII^{+}19] Akane Hara, Shoya Iwanami, Yusuke Ito, Tomoyuki Miura, Shinji Nakaoka, and Shingo Iwami. Revealing uninfected and infected target cell dynamics from peripheral blood data in highly and less pathogenic simian/human immunodeficiency virus infected Rhesus macaque. Journal of Theoretical Biology, 479(??):29–36, October 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302826.

Hirashima:2016:MSR

[Hir16] Tsuyoshi Hirashima. Mathematical study on robust tissue pattern formation in growing epididymal tubule. Journal of Theoretical Biology, 407(??):71–80, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301825.

Horita:2018:EBL

[HIT18] Junnosuke Horita, Yoh Iwasa, and Yuuya Tachiki. Evolutionary bistability of life history decision in male masu salmon. Journal of Theoretical Biology, 448(??):104–111, July 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301681.

Hackett-Jones:2011:RDA

[HJLNZ11] Emily J. Hackett-Jones, Kerry A. Landman, Donald F. Newgreen, and Dongcheng Zhang. On the role of differential adhesion in gangliogenesis in the enteric nervous system. Journal of Theoretical Biology, 287(??):148–159, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003638.

Hou:2012:DRE

[HJR12] Qiang Hou, Zhen Jin, and Shigui Ruan. Dynamics of rabies epidemics and the impact of control efforts in Guangdong Province, China. Journal of Theoretical Biology, 300(??):39– 47, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000070.

Hackett-Jones:2011:EDT

[HJWC11] Emily Hackett-Jones, Andrew White, and Christina A. Cobbold. The evolution of developmental timing in natural enemy systems. Journal of Theoretical Biology, 275(1):1–11, April 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310007034.

Hinsch:2010:DIE

[HK10] Martin Hinsch and Jan Komdeur. Defence, intrusion and the evolutionary stability of territoriality. Journal of Theoretical Biology, 266(4):606–613, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003905.

Hayat:2011:PMP

[HK11a] Maqsood Hayat and Asifullah Khan. Predicting membrane protein types by fusing composite protein sequence features into pseudo amino acid composition. Journal of Theoretical Biology, 271(1):10–17, February 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006028.

House:2011:EPC

[HK11b] Thomas House and Matt J. Keeling. Epidemic prediction and control in clustered populations. Journal of Theoretical Biology, 272(1):1–7, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006648.

Hayat:2012:MPM

[HK12] Maqsood Hayat and Asifullah Khan. MemHyb: Predicting membrane protein types by hybridizing SAAC and PSSM. Journal of Theoretical Biology, 292(??):93–102, January 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004905.

Hunter:2014:MUA

[HK14] Geoffrey A. M. Hunter and James P. Keener. Mechanisms underlying the additive and redundant Qrr phenotypes in Vibrio harveyi and Vibrio cholerae. Journal of Theoretical Biology, 340(??):38–49, January 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004086.

Hubai:2016:MGN

[HK16] András G. Hubai and Ádám Kun. Maximal gene number maintainable by stochastic correction — the second error threshold. Journal of Theoretical Biology, 405(??):29– 35, September 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316000977.

Harish:2017:MCB

[HK17] Ajith Harish and Charles G. Kurland. Mitochondria are not captive bacteria. Journal of Theoretical Biology, 434 (??):88–98, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519317303442.

Hooper:2010:TLH

[HKB10] Paul L. Hooper, Hillard S. Kaplan, and James L. Boone. A theory of leadership in human cooperative groups. Journal of Theoretical Biology, 265(4):633–646, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002778.

Husain:2018:PIC

[HKH18] Shahrukh Husain, Vijay Kumar, and Md. Imtaiyaz Hassan. Phosphorylation-induced changes in the energetic frustration in human tank binding kinase 1. Journal of Theoretical Biology, 449(??):14–22, July 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301772.

Horvath:2012:LMC

[HKM12] Gergely Horváth, Jaromír Kovárík, and Friederike Mengel. Limited memory can be beneficial for the evolution of cooperation. Journal of Theoretical Biology, 300(??):193–205, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000446.

Hussain:2019:SPS

[HKR^{+}19] Waqar Hussain, Yaser Daanial Khan, Nouman Rasool, Sher Afzal Khan, and Kuo-Chen Chou. SPrenylC-PseAAC: a sequence-based model developed via Chou’s 5-steps rule and general PseAAC for identifying S-prenylation sites in proteins. Journal of Theoretical Biology, 468(??):1–11, May 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300700.

He:2015:MDC

[HKS15] Y. He, D. Kulasiri, and S. Samarasinghe. Modelling the dynamics of CaMKII-NMDAR complex related to memory formation in synapses: The possible roles of threonine 286 autophosphorylation of CaMKII in long term potentiation. Journal of Theoretical Biology, 365(??):403–419, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400633X.

He:2016:MBM

[HKS16] Yao He, Don Kulasiri, and Sandhya Samarasinghe. Modelling bidirectional modulations in synaptic plasticity: a biochemical pathway model to understand the emergence of long term potentiation (LTP) and long term depression (LTD). Journal of Theoretical Biology, 403(??):159–177, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300935.

Hallen:2010:ESQ

[HL10] Mark A. Hallen and Anita T. Layton. Expanding the scope of quantitative FRAP analysis. Journal of Theoretical Biology, 262(2):295–305, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004950.

Haegeman:2011:MSN

[HL11] Bart Haegeman and Michel Loreau. A mathematical synthesis of niche and neutral theories in community ecology. Journal of Theoretical Biology, 269(1):150–165, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310005345.

He:2014:IOA

[HL14] Qinbin He and Zengrong Liu. Investigation of oscillation accumulation triggered genetic switch in gene regulatory networks. Journal of Theoretical Biology, 353(??):61–66, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314001520.

Houy:2018:ATC

[HL18] Nicolas Houy and François Le Grand. Administration of temozolomide: Comparison of conventional and metronomic chemotherapy regimens. Journal of Theoretical Biology, 446(??):71–78, June 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301085.

Houy:2019:OIC

[HL19] Nicolas Houy and François Le Grand. Optimizing immune cell therapies with artificial intelligence. Journal of Theoretical Biology, 461(??):34–40, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304375.

Huang:2018:AEW

[HLH^{+}18] Mugen Huang, Jiaowan Luo, Linchao Hu, Bo Zheng, and Jianshe Yu. Assessing the efficiency of Wolbachia driven Aedes mosquito suppression by delay differential equations. Journal of Theoretical Biology, 440(??):1–11, March 7, 2018.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305519.

He:2017:NAF

[HLHY17] Lily He, Yongkun Li, Rong Lucy He, and Stephen S.-T. Yau. A novel alignment-free vector method to cluster protein sequences. Journal of Theoretical Biology, 427(??):41–52, August 1, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302631.

Hara:2015:CEH

[HLI15] Makoto Hara, Joung-Hun Lee, and Yoh Iwasa. Cultural evolution of hinoeuma superstition controlling human mate choice: The role of half-believer. Journal of Theoretical Biology, 385(??):40–49, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004099.

Hu:2012:PGP

[HLS^{+}12] Yinxia Hu, Tonghua Li, Jiangming Sun, Shengnan Tang, Wenwei Xiong, Dapeng Li, Guanyan Chen, and Peisheng Cong. Predicting Gram-positive bacterial protein subcellular localization based on localization motifs. Journal of Theoretical Biology, 308(??):135–140, September 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002731.

Hodgkinson:2019:SGP

[HLTR19] Arran Hodgkinson, Laurent Le Cam, Dumitru Trucu, and Ovidiu Rădulescu. Spatio-genetic and phenotypic modelling elucidates resistance and re-sensitisation to treatment in heterogeneous melanoma. Journal of Theoretical Biology, 466(??):84–105, April 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305903.

Hsieh:2014:IVH

[HLTW14] Ying-Hen Hsieh, Junli Liu, Yun-Huei Tzeng, and Jianhong Wu. Impact of visitors and hospital staff on nosocomial transmission and spread to community. Journal of Theoretical Biology, 356(??):20–29, September 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002124.

Hermand:2016:VOE

[HLVR16] Eric Hermand, François J. Lhuissier, Nicolas Voituron, and Jean-Paul Richalet. Ventilatory oscillations at exercise in hypoxia: a mathematical model. Journal of Theoretical Biology, 411(??):92–101, December 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303241.

Haeno:2010:ETM

[HM10a] Hiroshi Haeno and Franziska Michor. The evolution of tumor metastases during clonal expansion. Journal of Theoretical Biology, 263(1):30–44, March 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005311.

Hong:2010:SPS

[HM10b] Dawei Hong and Shushuang Man. Signal propagation in small-world biological networks with weak noise. Journal of Theoretical Biology, 262(2):370–380, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004949.

Herschlag:2011:RNL

[HM11a] Gregory Herschlag and Laura Miller. Reynolds number limits for jet propulsion: a numerical study of simplified jellyfish. Journal of Theoretical Biology, 285(1):84–95, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311002876.

Hiebeler:2011:PTA

[HM11b] David E. Hiebeler and Nicholas E. Millett. Pair and triplet approximation of a spatial lattice population model with multiscale dispersal using Markov chains for estimating spatial autocorrelation. Journal of Theoretical Biology, 279(1):74–82, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001731.

Hyun:2013:IPS

[HM13] Alice Z. Hyun and Paul Macklin. Improved patient-specific calibration for agent-based cancer modeling. Journal of Theoretical Biology, 317(??):422–424, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005504.

Hoover:2015:NSB

[HM15] Alexander Hoover and Laura Miller. A numerical study of the benefits of driving jellyfish bells at their natural frequency. Journal of Theoretical Biology, 374(??):13–25, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315001253.

Hemmateenejad:2012:SPC

[HME12] Bahram Hemmateenejad, Ramin Miri, and Maryam Elyasi. A segmented principal component analysis-regression approach to QSAR study of peptides. Journal of Theoretical Biology, 305(??):37–44, July 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001609.

Higginson:2014:SPT

[HMH14] Andrew D. Higginson, John M. McNamara, and Alasdair I. Houston. The starvation-predation trade-off shapes the strategic use of protein for energy during fasting. Journal of Theoretical Biology, 359(??):208–219, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003580.

Hoyem:2015:SCR

[HMJB15] Marte Rørvik Høyem, Frode Måløy, Per Jakobsen, and Bjørn Olav Brandsdal. Stem cell regulation: Implications when differentiated cells regulate symmetric stem cell division. Journal of Theoretical Biology, 380(??):203–219, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002374.

Huang:2011:GEB

[HMM11] Jianxiong Huang, Xiangcheng Mi, and Keping Ma. A genome evolution-based framework for measures of originality for clades. Journal of Theoretical Biology, 276(1):99–105, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000634.

Hong:2016:SMS

[HMM16] Dawei Hong, Shushuang Man, and Joseph V. Martin. A stochastic mechanism for signal propagation in the brain: Force of rapid random fluctuations in membrane potentials of individual neurons. Journal of Theoretical Biology, 389(??):225–236, January 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005342.

Holden:2017:HPR

[HMM17] Matthew H. Holden and Eve McDonald-Madden. High prices for rare species can drive large populations extinct: the anthropogenic Allee effect revisited. Journal of Theoretical Biology, 429(??):170–180, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302916.

Hamzeh-Mivehroud:2015:IKI

[HMMSRSD15] Maryam Hamzeh-Mivehroud, Hakimeh Moghaddas-Sani, Mahdieh Rahbar-Shahrouziasl, and Siavoush Dastmalchi. Identifying key interactions stabilizing DOF zinc finger-DNA complexes using in silico approaches. Journal of Theoretical Biology, 382(??):150–159, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002921.

Halter:2017:RDP

[HMTA17] Wolfgang Halter, Jan Maximilian Montenbruck, Zoltan A. Tuza, and Frank Allgöwer. A resource dependent protein synthesis model for evaluating synthetic circuits. Journal of Theoretical Biology, 420(??):267–278, May 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301169.

Huber:2016:TPN

[HMW16] Katharina T. Huber, Vincent Moulton, and Taoyang Wu. Transforming phylogenetic networks: Moving beyond tree space. Journal of Theoretical Biology, 404(??):30–39, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301199.

Hiebeler:2013:HAP

[HMWB13] David E. Hiebeler, Isaac J. Michaud, Ben A. Wasserman, and Timothy D. Buchak. Habitat association in populations on landscapes with continuous-valued heterogeneous habitat quality. Journal of Theoretical Biology, 317(??):47–54, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004869.

Henderson:2010:PBE

[HN10] Donald M. Henderson and Darren Naish. Predicting the buoyancy, equilibrium and potential swimming ability of giraffes by computational analysis. Journal of Theoretical Biology, 265(2):151–159, July 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001906.

Hamaguchi:2019:EARb

[HN19] Yuko Hamaguchi and Hiroshi Nishiura. Estimate of the annual risk of tuberculosis infection in a general population of Japan. Journal of Theoretical Biology, 472(??):1–3, July 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301444.

Hirao:2015:NPC

[HNA15] Koudai Hirao, Atsushi J. Nagano, and Akinori Awazu. Noise-plasticity correlations of gene expression in the multicellular organism Arabidopsis thaliana. Journal of Theoretical Biology, 387(??):13–22, December 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004609.

Harrington:2011:AIE

[HNL^{+}11] Timothy D. Harrington, Nariman Naber, Adam G. Larson, Roger Cooke, Sarah E. Rice, and Edward Pate. Analysis of the interaction of the Eg5 Loop5 with the nucleotide site. Journal of Theoretical Biology, 289(??):107–115, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004127.

Hashimoto:2018:TGD

[HNO18] Atsushi Hashimoto, Atsuki Nagao, and Satoru Okuda. Topological graph description of multicellular dynamics based on vertex model. Journal of Theoretical Biology, 437(??):187–201, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304873.

Hamis:2018:WDK

[HNP18] Sara Hamis, Perumal Nithiarasu, and Gibin G. Powathil. What does not kill a tumour may make it stronger: In silico insights into chemotherapeutic drug resistance. Journal of Theoretical Biology, 454(??):253–267, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303084.

Hauser:2014:PDP

[HNR14] Oliver P. Hauser, Martin A. Nowak, and David G. Rand. Punishment does not promote cooperation under exploration dynamics when anti-social punishment is possible. Journal of Theoretical Biology, 360(??):163–171, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003920.

Haarsma:2016:SEP

[HNV^{+}16] Loren Haarsma, Serita Nelesen, Ethan VanAndel, James Lamine, and Peter VandeHaar. Simulating evolution of protein complexes through gene duplication and co-option. Journal of Theoretical Biology, 399(??):22–32, June 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001752.

Hohna:2015:TDR

[Höh15] Sebastian Höhna. The time-dependent reconstructed evolutionary process with a key-role for mass-extinction events. Journal of Theoretical Biology, 380(??):321–331, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002842.

Horgan:2011:BNM

[Hor11a] G. W. Horgan. The behaviour of a neutral model of weight regulated only by body mass. Journal of Theoretical Biology, 270(1):1–6, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005862.

Horikawa:2011:ETP

[Hor11b] Yo Horikawa. Exponential transient propagating oscillations in a ring of spiking neurons with unidirectional slow inhibitory synaptic coupling. Journal of Theoretical Biology, 289(??):151–159, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004280.

Hormoz:2013:SCP

[Hor13] Sahand Hormoz. Stem cell population asymmetry can reduce rate of replicative aging. Journal of Theoretical Biology, 331(??):19–27, August 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001860.

Hordijk:2017:ACC

[Hor17] Wim Hordijk. Autocatalytic confusion clarified. Journal of Theoretical Biology, 435(??):22–28, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304150.

Houy:2013:ECS

[Hou13] Nicolas Houy. Evolution of cooperation with similarity to an archetype. Journal of Theoretical Biology, 332(??):78–88, September 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001549.

Houy:2015:PUW

[Hou15] Nicolas Houy. The probability of undetected wild poliovirus circulation: Can we do better? Journal of Theoretical Biology, 382(??):272–278, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003252.

Houy:2016:CPO

[Hou16] Nicolas Houy. The case for periodic OPV routine vaccination campaigns. Journal of Theoretical Biology, 389(??):20–27, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004828.

Handy:2012:EIR

[HP12] Gregory A. Handy and Bradford E. Peercy. Extending the IP 3 receptor model to include competition with partial agonists. Journal of Theoretical Biology, 310(??):97–104, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002986.

Hanin:2016:QIM

[HP16] Leonid Hanin and Lyudmila Pavlova. A quantitative insight into metastatic relapse of breast cancer. Journal of Theoretical Biology, 394(??):172–181, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000473.

Hajisharifi:2014:PAP

[HPB^{+}14] Zohre Hajisharifi, Moien Piryaiee, Majid Mohammad Beigi, Mandana Behbahani, and Hassan Mohabatkar. Predicting anticancer peptides with Chou’s pseudo amino acid composition and investigating their mutagenicity via Ames test. Journal of Theoretical Biology, 341(??):34–40, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313004190.

Haridas:2017:QRC

[HPM^{+}17] Parvathi Haridas, Catherine J. Penington, Jacqui A. McGovern, D. L. Sean McElwain, and Matthew J. Simpson. Quantifying rates of cell migration and cell proliferation in coculture barrier assays reveals how skin and melanoma cells interact during melanoma spreading and invasion. Journal of Theoretical Biology, 423(??):13–25, June 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301819.

Hoare:2018:KOM

[HPML18] Samuel R. J. Hoare, Nicolas Pierre, Arturo Gonzalez Moya, and Brad Larson. Kinetic operational models of agonism for g-protein-coupled receptors. Journal of Theoretical Biology, 446(??):168–204, June 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300663.

Hapuarachchi:2010:QEV

[HPP10] Tharindi Hapuarachchi, Chang Sub Park, and Stephen Payne. Quantification of the effects of vasomotion on mass transport to tissue from axisymmetric blood vessels. Journal of Theoretical Biology, 264(2):553–559, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001256.

Helps:2017:ICU

[HPvdB17] J. C. Helps, N. D. Paveley, and F. van den Bosch. Identifying circumstances under which high insecticide dose increases or decreases resistance selection. Journal of Theoretical Biology, 428(??):153–167, September 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302783.

Huang:2013:MIC

[HPW^{+}13] Qihua Huang, Laura Parshotam, Hao Wang, Caroline Bampfylde, and Mark A. Lewis. A model for the impact of contaminants on fish population dynamics. Journal of Theoretical Biology, 334(??):71–79, October 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002476.

Huynh:2012:MDV

[HR12] Giao T. Huynh and Libin Rong. Modeling the dynamics of virus shedding into the saliva of Epstein–Barr virus positive individuals. Journal of Theoretical Biology, 310(??):105–114, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002743.

Hickson:2014:HPH

[HR14] R. I. Hickson and M. G. Roberts. How population heterogeneity in susceptibility and infectivity influences epidemic dynamics. Journal of Theoretical Biology, 350(??):70–80, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400023X.

Hackenberg:2012:CDW

[HRC^{+}12] Michael Hackenberg, Antonio Rueda, Pedro Carpena, Pedro Bernaola-Galván, Guillermo Barturen, and José L. Oliver. Clustering of DNA words and biological function: a proof of principle. Journal of Theoretical Biology, 297(??):127–136, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311006485.

Hou:2019:NDN

[HRCA19] Wenpin Hou, Peiying Ruan, Wai-Ki Ching, and Tatsuya Akutsu. On the number of driver nodes for controlling a Boolean network when the targets are restricted to attractors. Journal of Theoretical Biology, 463(??):1–11, February 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318306027.

Henke:2014:SVVb

[HRD14] H. Henke, P. A. Robinson, and P. M. Drysdale. Spatiotemporally varying visual hallucinations: II. spectral classification and comparison with theory. Journal of Theoretical Biology, 357(??):210–219, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002975.

Henke:2014:SVVa

[HRDL14] H. Henke, P. A. Robinson, P. M. Drysdale, and P. N. Loxley. Spatiotemporally varying visual hallucinations: I. corticothalamic theory. Journal of Theoretical Biology, 357(??):200–209, September 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002987.

Hugon:2011:NBD

[HRG11] J. Hugon, J.-C. Rostain, and B. Gardette. A new biophysical decompression model for estimating the risk of articular bends during and after decompression. Journal of Theoretical Biology, 283(1):168–179, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002360.

Herrera-Rivero:2015:SFM

[HRHAAA15] Marisol Herrera-Rivero, María Elena Hernández-Aguilar, and Gonzalo Emiliano Aranda-Abreu. A strategy focused on MAPT, APP, NCSTN and BACE1 to build blood classifiers for Alzheimer’s disease. Journal of Theoretical Biology, 376(??):32–38, July 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500154X.

Holder:2011:EPT

[HS11] Mark T. Holder and Mike Steel. Estimating phylogenetic trees from pairwise likelihoods and posterior probabilities of substitution counts. Journal of Theoretical Biology, 280(1):159–166, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001950.

Hordijk:2012:PTB

[HS12] Wim Hordijk and Mike Steel. Predicting template-based catalysis rates in a simple catalytic reaction model. Journal of Theoretical Biology, 295(??):132–138, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005947.

Hendrix:2014:BAM

[HS14a] Angelean O. Hendrix and James F. Selgrade. Bifurcation analysis of a menstrual cycle model reveals multiple mechanisms linking testosterone and classical PCOS. Journal of Theoretical Biology, 361(??):31–40, November 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004214.

Honjo:2014:PMN

[HS14b] Keita Honjo and Akiko Satake. N-player mosquito net game: Individual and social rationality in the misuse of insecticidetreated nets. Journal of Theoretical Biology, 342(??):39–46, February 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005262.

Humphries:2014:OFS

[HS14c] Nicolas E. Humphries and David W. Sims. Optimal foraging strategies: Lévy walks balance searching and patch exploitation under a very broad range of conditions. Journal of Theoretical Biology, 358(??):179–193, October 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003051.

Hadjichrysanthou:2015:ECA

[HS15] Christoforos Hadjichrysanthou and Kieran J. Sharkey. Epidemic control analysis: Designing targeted intervention strategies against epidemics propagated on contact networks. Journal of Theoretical Biology, 365(??):84–95, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005955.

Hordijk:2016:CTG

[HS16] Wim Hordijk and Mike Steel. Comment on “Tibor Gánti and Robert Rosen” by Athel Cornish-Bowden. Journal of Theoretical Biology, 392(??):122–123, March 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005718. See [CB15].

Hanert:2011:FDF

[HSD11] Emmanuel Hanert, Eva Schumacher, and Eric Deleersnijder. Front dynamics in fractional-order epidemic models. Journal of Theoretical Biology, 279(1):9–16, June 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001524.

Hu:2018:OBC

[HSG^{+}18] Bing Hu, Qianqian Shi, Yu Guo, Xiyezi Diao, Heng Guo, Jinsong Zhang, Liang Yu, Hao Dai, and Luonan Chen. The oscillatory boundary conditions of different frequency bands in Parkinson’s disease. Journal of Theoretical Biology, 451(??):67–79, August 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302248.

Hatanaka:2019:CRR

[HSiI^{+}19] Naoya Hatanaka, Takao Seki, Jun ichiro Inoue, Atsushi Tero, and Takashi Suzuki. Critical roles of IκBα and RelA phosphorylation in transitional oscillation in NF-κB signaling module. Journal of Theoretical Biology, 462(??):479–489, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305782.

Huo:2016:TDS

[HSLW16] Xi Huo, Xiaodan Sun, Kunquan Lan, and Jianhong Wu. Treatment-donation-stockpile dynamics in Ebola convalescent blood transfusion therapy. Journal of Theoretical Biology, 392(??):53–61, March 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005962.

Hidalgo:2017:SCN

[HSM17] Jorge Hidalgo, Samir Suweis, and Amos Maritan. Species coexistence in a neutral dynamics with environmental noise. Journal of Theoretical Biology, 413(??):1–10, January 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316303617.

Hauert:2019:AEG

[HSM19] Christoph Hauert, Camille Saade, and Alex McAvoy. Asymmetric evolutionary games with environmental feedback. Journal of Theoretical Biology, 462(??):347–360, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305721.

Hernandez-Suarez:2019:BMP

[HSMCCR19] Carlos Hernandez-Suarez, Paula Medone, Carlos Castillo-Chavez, and Jorge Rabinovich. Building matrix population models when individuals are non-identifiable. Journal of Theoretical Biology, 460(??):13–17, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304843.

Higginson:2012:EAP

[HSR12] Andrew D. Higginson, Michael P. Speed, and Graeme D. Ruxton. Effects of anti-predator defence through toxin sequestration on use of alternative food microhabitats by small herbivores. Journal of Theoretical Biology, 300(??):368–375, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000288.

Huang:2010:FPR

[HT10] Weini Huang and Arne Traulsen. Fixation probabilities of random mutants under frequency dependent selection. Journal of Theoretical Biology, 263(2):262–268, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005670.

Huisman:2013:MWD

[HT13] Jisca Huisman and Jarle Tufto. Modelling wild-domestic interbreeding: How selection on a quantitative trait affects gene flow at a neutral locus. Journal of Theoretical Biology, 332(??):42–51, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001902.

Hossjer:2016:MDM

[HT16] Ola Hössjer and Peder A. Tyvand. A monoecious and diploid Moran model of random mating. Journal of Theoretical Biology, 394(??):182–196, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000357.

Haridas:2018:MEE

[HT18] C. V. Haridas and Brigitte Tenhumberg. Modeling effects of ecological factors on evolution of polygenic pesticide resistance. Journal of Theoretical Biology, 456(??):224–232, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303588.

Hamada:2019:EPS

[HT19] Miki Hamada and Fugo Takasu. Equilibrium properties of the spatial SIS model as a point pattern dynamics — how is infection distributed over space? Journal of Theoretical Biology, 468(??):12–26, May 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300608.

Hu:2013:EAD

[HTB^{+}13] K. Hu, C. Thoens, S. Bianco, S. Edlund, M. Davis, J. Douglas, and J. H. Kaufman. The effect of antibodydependent enhancement, cross immunity, and vector population on the dynamics of dengue fever. Journal of Theoretical Biology, 319(??):62–74, February 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005966.

Huang:2012:HEP

[HTK12] Gang Huang, Yasuhiro Takeuchi, and Andrei Korobeinikov. HIV evolution and progression of the infection to AIDS. Journal of Theoretical Biology, 307(??):149–159, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312002470.

Hayat:2014:PPS

[HTK14] Maqsood Hayat, Muhammad Tahir, and Sher Afzal Khan. Prediction of protein structure classes using hybrid space of multi-profile Bayes and bi-gram probability feature spaces. Journal of Theoretical Biology, 346(??):8–15, April 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005663.

Hayashi:2018:IFM

[HTK^{+}18] Tatsuya Hayashi, Tetsuji Tokihiro, Hiroki Kurihara, Fumimasa Nomura, and Kenji Yasuda. Integrate and fire model with refractory period for synchronization of two cardiomyocytes. Journal of Theoretical Biology, 437(??):141–148, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730471X.

Hayakawa:2015:MSM

[HTM15] Yoshinori Hayakawa, Masashi Tachikawa, and Atsushi Mochizuki. Mathematical study for the mechanism of vascular and spot patterns by auxin and pin dynamics in plant development. Journal of Theoretical Biology, 365(??):12–22, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005864.

Hayakawa:2016:FLF

[HTM16] Yoshinori Hayakawa, Masashi Tachikawa, and Atsushi Mochizuki. Flat leaf formation realized by cell-division control and mutual recessive gene regulation. Journal of Theoretical Biology, 404(??):206–214, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301394.

Hauser:2014:HBF

[HTN14] Oliver P. Hauser, Arne Traulsen, and Martin A. Nowak. Heterogeneity in background fitness acts as a suppressor of selection. Journal of Theoretical Biology, 343(??):178–185, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005110.

He:2019:FAI

[HTT19] Yang He, Suyan Tian, and Pu Tian. Fundamental asymmetry of insertions and deletions in genomes size evolution. Journal of Theoretical Biology, 482(??):Article 109983, December 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303261.

Huang:2012:RGO

[Hua12] Wen-Lin Huang. Ranking gene ontology terms for predicting non-classical secretory proteins in eukaryotes and prokaryotes. Journal of Theoretical Biology, 312(??):105–113, November 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003797.

Huang:2016:EDM

[Hua16] Hsin-Hsiung Huang. An ensemble distance measure of k-mer and natural vector for the phylogenetic analysis of multiple-segmented viruses. Journal of Theoretical Biology, 398(??):136–144, June 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001442.

Hurtado:2012:WHD

[Hur12] Paul J. Hurtado. Within-host dynamics of mycoplasma infections: Conjunctivitis in wild passerine birds. Journal of Theoretical Biology, 306(??):73–92, August 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001993.

Herrera-Valdez:2011:RMP

[HVL11] Marco Arieli Herrera-Valdez and Joceline Lega. Reduced models for the pacemaker dynamics of cardiac cells. Journal of Theoretical Biology, 270(1):164–176, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005254.

Huang:2015:DVF

[HvLPdB15] Zheng Y. X. Huang, Frank van Langevelde, Herbert H. T. Prins, and Willem F. de Boer. Dilution versus facilitation: Impact of connectivity on disease risk in metapopulations. Journal of Theoretical Biology, 376(??):66–73, July 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315001599.

Hernandez-Vargas:2013:MTS

[HVM13] Esteban A. Hernandez-Vargas and Richard H. Middleton. Modeling the three stages in HIV infection. Journal of Theoretical Biology, 320(??):33–40, March 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006170.

Hernandez-Valdez:2010:CEC

[HVSZ10] Areli Hernández-Valdez, Moisés Santillán, and Eduardo S. Zeron. Cycling expression and cooperative operator interaction in the trp operon of Escherichia coli. Journal of Theoretical Biology, 263(3):340–352, April 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005724.

Hsieh:2011:HMT

[HW11] Wan J. Hsieh and Hsiuying Wang. Human microRNA target identification by RRSM. Journal of Theoretical Biology, 286(??):79–84, October 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003213.

Hsieh:2013:RDD

[HW13] Wan J. Hsieh and Hsiuying Wang. RRSM with a datadependent threshold for miRNA target prediction. Journal of Theoretical Biology, 337(??):54–60, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300372X.

Huang:2014:SMI

[HWD^{+}14] Can Huang, Ming Wu, Jun Du, Di Liu, and Christina Chan. Systematic modeling for the insulin signaling network mediated by IRS 1 and IRS 2. Journal of Theoretical Biology, 355(??):40–52, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001775.

Holtz:2015:NCM

[HWGT15a] Lena-Maria Holtz, Dieter Wolf-Gladrow, and Silke Thoms. Numerical cell model investigating cellular carbon fluxes in Emiliania huxleyi. Journal of Theoretical Biology, 364(??): 305–315, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519314005013.

Holtz:2015:SEL

[HWGT15b] Lena-Maria Holtz, Dieter Wolf-Gladrow, and Silke Thoms. Simulating the effects of light intensity and carbonate system composition on particulate organic and inorganic carbon production in Emiliania huxleyi. Journal of Theoretical Biology, 372(??):192–204, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315000958.

Holtz:2017:SCI

[HWGT17] Lena-Maria Holtz, Dieter Wolf-Gladrow, and Silke Thoms. Stable carbon isotope signals in particulate organic and inorganic carbon of coccolithophores — a numerical model study for Emiliania huxleyi. Journal of Theoretical Biology, 420(??):117–127, May 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300309.

He:2018:MYC

[HWGW18] Daihai He, Xueying Wang, Daozhou Gao, and Jin Wang. Modeling the 2016–2017 Yemen cholera outbreak with the impact of limited medical resources. Journal of Theoretical Biology, 451(??):80–85, August 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830225X.

Huang:2015:PPF

[HWHL15] Jitao T. Huang, Titi Wang, Shanran R. Huang, and Xin Li. Prediction of protein folding rates from simplified secondary structure alphabet. Journal of Theoretical Biology, 383(??):1–6, October 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003690.

Huang:2015:IET

[HWL15] Qihua Huang, Hao Wang, and Mark A. Lewis. The impact of environmental toxins on predator–prey dynamics. Journal of Theoretical Biology, 378(??):12–30, August 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001952.

Haber:2016:IFF

[HWM^{+}16] Shimon Haber, Michał Weisbord, Michiaki Mishima, Steve J. Mentzer, and Akira Tsuda. Interstitial fluid flow of alveolar primary septa after pneumonectomy. Journal of Theoretical Biology, 400(??):118–128, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300091.

Haber:2017:ASP

[HWMT17] Shimon Haber, Michał Weisbord, Steven J. Mentzer, and Akira Tsuda. Alveolar septal patterning during compensatory lung growth: Part II the effect of parenchymal pressure gradients. Journal of Theoretical Biology, 421(??):168–178, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301406.

Hirt:2012:ETD

[HWPL12] Bartholomäus V. Hirt, Jonathan A. D. Wattis, Simon P. Preston, and Charles A. Laughton. The effects of a telomere destabilizing agent on cancer cell-cycle dynamics — integrated modelling and experiments. Journal of Theoretical Biology, 295(??):9–22, February 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005613.

Hilbe:2015:EPZ

[HWTN15] Christian Hilbe, Bin Wu, Arne Traulsen, and Martin A. Nowak. Evolutionary performance of zero-determinant strategies in multiplayer games. Journal of Theoretical Biology, 374(??):115–124, June 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001472.

He:2014:WLC

[HWW^{+}14] Ping He, Wei Wu, Hai-Dong Wang, Ke-Long Liao, Wei Zhang, Feng-Lin Lv, and Kang Yang. Why ligand crossreactivity is high within peptide recognition domain families? A case study on human c-Src SH3 domain. Journal of Theoretical Biology, 340(??):30–37, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004001.

He:2016:EAA

[HXL16] Qinbin He, Zhile Xia, and Bin Lin. An efficient approach of attractor calculation for large-scale Boolean gene regulatory networks. Journal of Theoretical Biology, 408(??):137–144, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302399.

He:2018:PAA

[HXL18a] Qinbin He, Zhile Xia, and Bin Lin. P UNSAT approach of attractor calculation for Boolean gene regulatory networks. Journal of Theoretical Biology, 447(??):171–177, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301516.

Hu:2018:EEB

[HXL18b] Di Hu, Cai-Hua Xiong, and Ming-Jin Liu. Exploring the existence of better hands for manipulation than the human hand based on hand proportions. Journal of Theoretical Biology, 440(??):100–111, March 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305738.

Hironaga:2010:EEF

[HY10] Ryo Hironaga and Norio Yamamura. Effects of extinction on food web structures on an evolutionary time scale. Journal of Theoretical Biology, 263(2):161–168, March 21, 2010.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005608.

Huang:2013:PPS

[HY13] Chao Huang and Jing-Qi Yuan. Predicting protein subchloroplast locations with both single and multiple sites via three different modes of Chou’s pseudo amino acid compositions. Journal of Theoretical Biology, 335(??):205–212, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003160.

Huang:2016:CDS

[HY16] Hsin-Hsiung Huang and Chenglong Yu. Clustering DNA sequences using the out-of-place measure with reduced ngrams. Journal of Theoretical Biology, 406(??):61–72, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301631.

Han:2014:TSS

[HYA14] Guo-Sheng Han, Zu-Guo Yu, and Vo Anh. A two-stage SVM method to predict membrane protein types by incorporating amino acid classifications and physicochemical properties into a general form of Chou’s PseAAC. Journal of Theoretical Biology, 344(??):31–39, March 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005444.

Hamaguchi:2019:EARa

[HYN19] Yuko Hamaguchi, Takayuki Yamaguchi, and Hiroshi Nishiura. Estimating the annual risk of tuberculosis infection in Japan from interferon-gamma release assay data. Journal of Theoretical Biology, 460(??):125–133, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304867.

Huang:2016:ESH

[HYQ^{+}16] Yunxin Huang, Nianwan Yang, Yun Qin, Feng An, Zhaohua Li, and Fanghao Wan. Enhanced stability in hostparasitoid interactions with autoparasitism and parasitoid migration. Journal of Theoretical Biology, 393(??):43–50, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000400.

Hassani:2019:RMS

[HYSG19] Hossein Hassani, Mohammad Reza Yeganegi, Emmanuel Sirimal Silva, and Fatemeh Ghodsi. Risk management, signal processing and econometrics: a new tool for forecasting the risk of disease outbreaks. Journal of Theoretical Biology, 467(??):57– 62, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300499.

Huang:2011:PAD

[HYW11] Yujuan Huang, Lianping Yang, and Tianming Wang. Phylogenetic analysis of DNA sequences based on the generalized pseudo-amino acid composition. Journal of Theoretical Biology, 269(1):217–223, January 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005618.

Hoang:2015:NMC

[HYZ^{+}15] Tung Hoang, Changchuan Yin, Hui Zheng, Chenglong Yu, Rong Lucy He, and Stephen S.-T. Yau. A new method to cluster DNA sequences using Fourier power spectrum. Journal of Theoretical Biology, 372(??):135–145, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000971.

Hu:2014:RIN

[HYZS14] Guang Hu, Wenying Yan, Jianhong Zhou, and Bairong Shen. Residue interaction network analysis of Dronpa and a DNA clamp. Journal of Theoretical Biology, 348(??):55–64, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400040X.

Huang:2010:PPD

[HZC^{+}10] Chen Huang, Ruijie Zhang, Zhiqiang Chen, Yongshuai Jiang, Zhenwei Shang, Peng Sun, Xuehong Zhang, and Xia Li. Predict potential drug targets from the ion channel proteins based on SVM. Journal of Theoretical Biology, 262(4):750–756, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005281.

Hu:2017:CAE

[HZG^{+}17] Bing Hu, Xiaoqiang Zou, Yu Guo, Zhejia Yang, Feng Shi, and Wangyuan Dong. Control of absence epilepsy seizures in specific relay nuclei of thalamus. Journal of Theoretical Biology, 435(??):50–61, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304186.

Hua:2011:MSG

[HZL^{+}11] Lin Hua, Ping Zhou, Hong Liu, Lin Li, Zheng Yang, and Zhi cheng Liu. Mining susceptibility gene modules and disease risk genes from SNP data by combining network topological properties with support vector regression. Journal of Theoretical Biology, 289(??):225–236, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004437.

Hua:2013:PBC

[HZL^{+}13] Lin Hua, Ping Zhou, Lin Li, Hong Liu, and Zheng Yang. Prioritizing breast cancer subtype related miRNAs using miRNA-mRNA dysregulated relationships extracted from their dual expression profiling. Journal of Theoretical Biology, 331(??):1–11, August 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001616.

Huang:2011:AFC

[HZLX11] Guohua Huang, Houqing Zhou, Yongfan Li, and Lixin Xu. Alignment-free comparison of genome sequences by a new numerical characterization. Journal of Theoretical Biology, 281(1):107–112, July 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001937.

Hu:2014:TPR

[HZW^{+}14] Jie Hu, Xiaomei Zhu, Xinan Wang, Ruoshi Yuan, Wei Zheng, Minjuan Xu, and Ping Ao. Two programmed replicative lifespans of Saccharomyces cerevisiae formed by the endogenous molecular-cellular network. Journal of Theoretical Biology, 362(??):69–74, December 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000162.

Huang:2010:SMA

[HZWH10] Wei Huang, Jianmin Zhang, Yurong Wang, and Dan Huang. A simple method to analyze the similarity of biological sequences based on the fuzzy theory. Journal of Theoretical Biology, 265(3):323–328, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002377.

Ibrahim:2017:EFP

[IA17] Wisam Ibrahim and Mohammad Saniee Abadeh. Extracting features from protein sequences to improve deep extreme learning machine for protein fold recognition. Journal of Theoretical Biology, 421(??):1–15, May 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730142X.

Iarosz:2015:MMB

[IBB^{+}15] Kelly C. Iarosz, Fernando S. Borges, Antonio M. Batista, Murilo S. Baptista, Regiane A. N. Siqueira, Ricardo L. Viana, and Sergio R. Lopes. Mathematical model of brain tumour with glia-neuron interactions and chemotherapy treatment. Journal of Theoretical Biology, 368(??):113–121, March 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000077.

Ishii:2011:HIC

[IC11] Reiichiro Ishii and Michael J. Crawley. Herbivore-induced coexistence of competing plant species. Journal of Theoretical Biology, 268(1):50–61, January 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004005.

Ishimoto:2016:SSS

[ICG16] Kenta Ishimoto, Jacky Cosson, and Eamonn A. Gaffney. A simulation study of sperm motility hydrodynamics near fish eggs and spheres. Journal of Theoretical Biology, 389 (??):187–197, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519315005068.

Ito:2014:EBU

[ID14] Hiroshi C. Ito and Ulf Dieckmann. Evolutionary branching under slow directional evolution. Journal of Theoretical Biology, 360(??):290–314, November 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004025.

Ito:2019:JEC

[ID19] Koichi Ito and Michael Doebeli. The joint evolution of cooperation and competition. Journal of Theoretical Biology, 480(??):1–12, November 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302875.

Ikemoto:2011:SRE

[IDI11] Shuhei Ikemoto, Fabio DallaLibera, and Hiroshi Ishiguro. Stochastic resonance emergence from a minimalistic behavioral rule. Journal of Theoretical Biology, 273(1):179–187, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000063.

Ingr:2015:EDM

[IDM15] Marek Ingr, Jirí Dostál, and Tatána Majerová. Enzymological description of multitemplate PCR-shrinking amplification bias by optimizing the polymerase-template ratio. Journal of Theoretical Biology, 382(??):178–186, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315003276.

Ikegawa:2015:EGS

[IEN15] Yusuke Ikegawa, Hideo Ezoe, and Toshiyuki Namba. Effects of generalized and specialized adaptive defense by shared prey on intra-guild predation. Journal of Theoretical Biology, 364(??):231–241, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005414.

Ishimoto:2014:SSS

[IG14] Kenta Ishimoto and Eamonn A. Gaffney. A study of spermatozoan swimming stability near a surface. Journal of Theoretical Biology, 360(??):187–199, November 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003774.

Ishimoto:2018:HSS

[IGG^{+}18] Kenta Ishimoto, Hermes Gadêlha, Eamonn A. Gaffney, David J. Smith, and Jackson Kirkman-Brown. Human sperm swimming in a high viscosity mucus analogue. Journal of Theoretical Biology, 446(??):1–10, June 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300651.

Immonen:2012:HSD

[IGL^{+}12] Taina Immonen, Richard Gibson, Thomas Leitner, Melanie A. Miller, Eric J. Arts, Erkki Somersalo, and Daniela Calvetti. A hybrid stochastic-deterministic computational model accurately describes spatial dynamics and virus diffusion in HIV-1 growth competition assay. Journal of Theoretical Biology, 312(??):120–132, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003396.

Ikegawa:2017:LMT

[IH17] Yusuke Ikegawa and Chihiro Himuro. Limited mobility of target pests crucially lowers controllability when sterile insect releases are spatiotemporally biased. Journal of Theoretical Biology, 421(??):93–100, May 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301455.

Hironaka:2012:MFL

[iHIM12] Ken ichi Hironaka, Yoh Iwasa, and Yoshihiro Morishita. Multiple feedback loops achieve robust localization of wingless expression in Drosophila notum development. Journal of Theoretical Biology, 292(??):18–29, January 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004863.

Hironaka:2017:ASC

[iHM17] Ken ichi Hironaka and Yoshihiro Morishita. Adaptive significance of critical weight for metamorphosis in holometabolous insects. Journal of Theoretical Biology, 417(??):68–83, March 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300140.

Ilamathi:2016:IPT

[IHNS16] M. Ilamathi, R. Hemanth, S. Nishanth, and V. Sivaramakrishnan. Identification of potential transmembrane protease serine 4 inhibitors as anti-cancer agents by integrated computational approach. Journal of Theoretical Biology, 389(??):253–262, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005202.

Inouye:2017:DSG

[III17] Masayori Inouye, Yojiro Ishida, and Keiko Inouye. Designing of a single gene encoding four functional proteins. Journal of Theoretical Biology, 419(??):266–268, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300516.

Iijima:2012:DDE

[Iij12] Ryota Iijima. On delayed discrete evolutionary dynamics. Journal of Theoretical Biology, 300(??):1–6, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000021.

Islam:2016:BSS

[IIKH16] Md. Nasrul Islam, Sumaiya Iqbal, Ataur R. Katebi, and Md. Tamjidul Hoque. A balanced secondary structure predictor. Journal of Theoretical Biology, 389(??):60–71, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005147.

Ichinose:2013:MCM

[IIKT13] Motoshi Ichinose, Masaru Iizuka, Junko Kusumi, and Masasuke Takefu. Models of compensatory molecular evolution: Effects of back mutation. Journal of Theoretical Biology, 323(??):1–10, April 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000362.

Ishihara:2015:MSP

[iITIM15] Jun ichi Ishihara, Masashi Tachikawa, Hideo Iwasaki, and Atsushi Mochizuki. Mathematical study of pattern formation accompanied by heterocyst differentiation in multicellular cyanobacterium. Journal of Theoretical Biology, 371(??):9–23, April 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931500051X.

Izquierdo:2010:OLC

[IIVR10] Segismundo S. Izquierdo, Luis R. Izquierdo, and Fernando Vega-Redondo. The option to leave: Conditional dissociation in the evolution of cooperation. Journal of Theoretical Biology, 267(1):76–84, November 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003966.

Ioannidis:2015:CSS

[IK15] Nikolaos E. Ioannidis and Kiriakos Kotzabasis. Could structural similarity of specific domains between animal globins and plant antenna proteins provide hints important for the photoprotection mechanism? Journal of Theoretical Biology, 364(??):71–79, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005360.

Iyer:2016:EDS

[IK16] Swami Iyer and Timothy Killingback. Evolutionary dynamics of a smoothed war of attrition game. Journal of Theoretical Biology, 396(??):25–41, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001041.

Ingr:2016:EOP

[IKHL16] Marek Ingr, Eva Kutálková, Josef Hrncirík, and Reinhard Lange. Equilibria of oligomeric proteins under high pressure — a theoretical description. Journal of Theoretical Biology, 411(??):16–26, December 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630323X.

Inarra:2012:ADR

[IL12] Elena Inarra and Annick Laruelle. Artificial distinction and real discrimination. Journal of Theoretical Biology, 305(??):110–117, July 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931200166X.

Iwasa:2013:GPE

[IL13] Yoh Iwasa and Joung-Hun Lee. Graduated punishment is efficient in resource management if people are heterogeneous. Journal of Theoretical Biology, 333(??):117–125, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002221.

Iwagami:2010:URH

[IM10] Akio Iwagami and Naoki Masuda. Upstream reciprocity in heterogeneous networks. Journal of Theoretical Biology, 265(3):297–305, August 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002390.

Ichinose:2018:ZDS

[IM18] Genki Ichinose and Naoki Masuda. Zero-determinant strategies in finitely repeated games. Journal of Theoretical Biology, 438(??):61–77, February 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305039.

Ibanez-Marcelo:2014:TRE

[IMA14] Esther Ibáñez-Marcelo and Tomás Alarcón. The topology of robustness and evolvability in evolutionary systems with genotype-phenotype map. Journal of Theoretical Biology, 356(??):144–162, September 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002343.

Ibanez-Marcelo:2016:EEC

[IMA16] Esther Ibáñez-Marcelo and Tomás Alarcón. Evolutionary escape on complex genotype-phenotype networks. Journal of Theoretical Biology, 394(??):18–31, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000412.

Ispolatov:2016:IBM

[IMD16] Iaroslav Ispolatov, Vaibhav Madhok, and Michael Doebeli. Individual-based models for adaptive diversification in high-dimensional phenotype spaces. Journal of Theoretical Biology, 390(??):97–105, February 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005020.

Iqbal:2015:IPA

[IMH15] Sumaiya Iqbal, Avdesh Mishra, and Md Tamjidul Hoque. Improved prediction of accessible surface area results in efficient energy function application. Journal of Theoretical Biology, 380(??):380–391, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500291X.

Iacobelli:2016:LEG

[IMM16] G. Iacobelli, D. Madeo, and C. Mocenni. Lumping evolutionary game dynamics on networks. Journal of Theoretical Biology, 407(??):328–338, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630220X.

Issarow:2015:MRA

[IMW15] Chacha M. Issarow, Nicola Mulder, and Robin Wood. Modelling the risk of airborne infectious disease using exhaled air. Journal of Theoretical Biology, 372(??):100–106, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000697.

Ingber:2016:SMN

[Ing16] Lester Ingber. Statistical mechanics of neocortical interactions: Large-scale EEG influences on molecular processes. Journal of Theoretical Biology, 395(??):144–152, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931600093X.

Iida:2019:QHS

[IOK19] Keita Iida, Nobuaki Obata, and Yoshitaka Kimura. Quantifying heterogeneity of stochastic gene expression. Journal of Theoretical Biology, 465(??):56–62, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300025.

Iosilevskii:2016:LNB

[Ios16] Gil Iosilevskii. Locomotion of neutrally buoyant fish with flexible caudal fin. Journal of Theoretical Biology, 399(??):159– 165, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300200.

Ito:2015:RBA

[IOY15] Koichi Ito, Hisashi Ohtsuki, and Atsushi Yamauchi. Relationship between aggregation of rewards and the possibility of polymorphism in continuous snowdrift games. Journal of Theoretical Biology, 372(??):47–53, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000740.

Iosilevskii:2012:EYY

[IPMH12] Gil Iosilevskii, Yannis P. Papastamatiou, Carl G. Meyer, and Kim N. Holland. Energetics of the yo-yo dives of predatory sharks. Journal of Theoretical Biology, 294(??):172–181, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005704.

Ingber:2014:EFI

[IPS14] Lester Ingber, Marco Pappalepore, and Ronald R. Stesiak. Electroencephalographic field influence on calcium momentum waves. Journal of Theoretical Biology, 343(??):138–153, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005274.

Iranzo:2011:SUG

[IRS11] Jaime Iranzo, Javier Román, and Angel Sánchez. The spatial ultimatum game revisited. Journal of Theoretical Biology, 278(1):1–10, June 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001226.

Ingram:2010:MUF

[IS10] Travis Ingram and Mike Steel. Modelling the unpredictability of future biodiversity in ecological networks. Journal of Theoretical Biology, 264(3):1047–1056, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001232.

Ishihara:2012:BIF

[IS12] Shuji Ishihara and Kaoru Sugimura. Bayesian inference of force dynamics during morphogenesis. Journal of Theoretical Biology, 313(??):201–211, November 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004316.

Ito:2016:EBU

[IS16] Hiroshi Ito and Akira Sasaki. Evolutionary branching under multi-dimensional evolutionary constraints. Journal of Theoretical Biology, 407(??):409–428, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301941.

Ishimoto:2017:MMD

[IS17] Yukitaka Ishimoto and Kaoru Sugimura. A mechanical model for diversified insect wing margin shapes. Journal of Theoretical Biology, 427(??):17–27, August 1, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302485.

Isella:2011:WCA

[ISB^{+}11] Lorenzo Isella, Juliette Stehlé, Alain Barrat, Ciro Cattuto, Jean-François Pinton, and Wouter Van den Broeck. What’s in a crowd? Analysis of face-to-face behavioral networks. Journal of Theoretical Biology, 271(1):166–180, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006284.

Ishimoto:2016:HES

[Ish16] Kenta Ishimoto. Hydrodynamic evolution of sperm swimming: Optimal flagella by a genetic algorithm. Journal of Theoretical Biology, 399(??):166–174, June 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300170.

Ishikawa:2011:TPM

[ISM^{+}11] T. Ishikawa, T. Sato, G. Mohit, Y. Imai, and T. Yamaguchi. Transport phenomena of microbial flora in the small intestine with peristalsis. Journal of Theoretical Biology, 279(1):63–73, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100172X.

Iseki:2011:ARB

[IST11] Naoyuki Iseki, Akira Sasaki, and Hirokazu Toju. Arms race between weevil rostrum length and camellia pericarp thickness: Geographical cline and theory. Journal of Theoretical Biology, 285(1):1–9, September 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002852.

Ilies:2018:GAS

[ISZ18] Iulian Ilies, Rifat Sipahi, and Günther K. H. Zupanc. Growth of adult spinal cord in knifefish: Development and parametrization of a distributed model. Journal of Theoretical Biology, 437(??):101–114, January 21, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304757.

Inoue:2011:MMD

[ITN^{+}11] Yasuhiro Inoue, Shunsuke Tsuda, Koji Nakagawa, Masaki Hojo, and Taiji Adachi. Modeling myosin-dependent rearrangement and force generation in an actomyosin network. Journal of Theoretical Biology, 281(1):65–73, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001949.

Inaba:2016:RLS

[ITO16] Misato Inaba, Nobuyuki Takahashi, and Hisashi Ohtsuki. Robustness of linkage strategy that leads to large-scale cooperation. Journal of Theoretical Biology, 409(??):97–107, November 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302788.

Ito:2018:DHC

[ITR^{+}18] Yusuke Ito, Alexandra Tauzin, Azaria Remion, Keisuke Ejima, Fabrizio Mammano, and Shingo Iwami. Dynamics of HIV-1 coinfection in different susceptible target cell populations during cell-free infection. Journal of Theoretical Biology, 455(??):39–46, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303254.

Ito:2010:FIN

[IU10] Yohei Ito and Kenko Uchida. Formulas for intrinsic noise evaluation in oscillatory genetic networks. Journal of Theoretical Biology, 267(2):223–234, November 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004480.

Iwamoto:2018:BSO

[IU18] Mayuko Iwamoto and Daishin Ueyama. Basis of self-organized proportion regulation resulting from local contacts. Journal of Theoretical Biology, 440(??):112–120, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305751.

Iwamoto:2014:AMA

[IUK14] Mayuko Iwamoto, Daishin Ueyama, and Ryo Kobayashi. The advantage of mucus for adhesive locomotion in gastropods. Journal of Theoretical Biology, 353(??):133–141, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400099X.

Ivanov:2016:SSS

[IvdSW16] Oleksandr Ivanov, Arjan van der Schaft, and Franz J. Weissing. Steady states and stability in metabolic networks without regulation. Journal of Theoretical Biology, 401(??):78–93, July 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001272.

Ilie:2012:MMT

[IvLP12] Olga Ilie, Mark C. M. van Loosdrecht, and Cristian Picioreanu. Mathematical modelling of tooth demineralisation and pH profiles in dental plaque. Journal of Theoretical Biology, 309(??):159–175, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002664.

Johnson:2013:CBC

[JA13] Christopher A. Johnson and Priyanga Amarasekare. Competition for benefits can promote the persistence of mutualistic interactions. Journal of Theoretical Biology, 328(??):54–64, July 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001203.

Jabot:2010:SDL

[Jab10] Franck Jabot. A stochastic dispersal-limited trait-based model of community dynamics. Journal of Theoretical Biology, 262(4):650–661, February 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900530X.

Jabot:2015:WPP

[Jab15] Franck Jabot. Why preferring parametric forecasting to nonparametric methods? Journal of Theoretical Biology, 372(??):205–210, May 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001034.

Jentsch:2018:SCE

[JAB18] Peter C. Jentsch, Madhur Anand, and Chris T. Bauch. Spatial correlation as an early warning signal of regime shifts in a multiplex disease-behaviour network. Journal of Theoretical Biology, 448(??):17–25, July 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301401.

Jinha:2012:RLE

[JAHKH12] Azim Jinha, Rachid Ait-Haddou, Motoshi Kaya, and Walter Herzog. Response to letter to the Editor regarding Jinha et al. (2009) “A task-specific validation of homogeneous non-linear optimization approaches”. Journal of Theoretical Biology, 306(??):145, August 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001701.

J:2019:CCR

[JAK19] Murray P. J., Carrieri F. A., and Dale J. K. Cell cycle regulation of oscillations yields coupling of growth and form in a computational model of the presomitic mesoderm. Journal of Theoretical Biology, 481(??):75–83, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301912.

James:2011:FSH

[Jam11] William H. James. Further support for a hypothesis on the reported offspring sex ratios of hepatitis B carriers. Journal of Theoretical Biology, 274(1):183–184, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000294.

James:2012:HSV

[Jam12] William H. James. Hypotheses on the stability and variation of human sex ratios at birth. Journal of Theoretical Biology, 310(??):183–186, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003311.

James:2013:EVM

[Jam13] William H. James. Evolution and the variation of mammalian sex ratios at birth: Reflections on Trivers and Willard (1973). Journal of Theoretical Biology, 334(??):141–148, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002956.

James:2014:UHO

[Jam14] William H. James. An update on the hypothesis that one cause of autism is high intrauterine levels of testosterone of maternal origin. Journal of Theoretical Biology, 355(??):33–39, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001921.

James:2015:HHL

[Jam15] William H. James. Hypothesis: High levels of maternal adrenal androgens are a major cause of miscarriage and other forms of reproductive suboptimality. Journal of Theoretical Biology, 364(??):316–320, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005669.

James:2016:REF

[Jam16a] Ananthu James. Role of epistasis on the fixation probability of a non-mutator in an adapted asexual population. Journal of Theoretical Biology, 407(??):225–237, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301837.

James:2016:SRS

[Jam16b] William H. James. The sex ratio of the sibs of probands diagnosed with autism. Journal of Theoretical Biology, 400 (??):154–157, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316001612.

Jalilvand:2018:FSB

[JAM18] A. Jalilvand, B. Akbari, and F. Zare Mirakabad. S-FLN: a sequence-based hierarchical approach for functional linkage network construction. Journal of Theoretical Biology, 437(??):149–162, January 21, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931730485X.

Jansson:2015:WGS

[Jan15] Fredrik Jansson. What games support the evolution of an ingroup bias? Journal of Theoretical Biology, 373(??):100–110, May 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001113.

Johnson:2011:PII

[JB11] Leah R. Johnson and Cheryl J. Briggs. Parameter inference for an individual based model of chytridiomycosis in frogs. Journal of Theoretical Biology, 277(1):90–98, May 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001044.

Jager:2018:LPR

[JB18a] Christoph G. Jäger and Dietrich Borchardt. Longitudinal patterns and response lengths of algae in riverine ecosystems: a model analysis emphasising benthic-pelagic interactions. Journal of Theoretical Biology, 442(??):66–78, April 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300171.

Janouskova:2018:HPD

[JB18b] Eva Janousková and Ludek Berec. Host-pathogen dynamics under sterilizing pathogens and fecundity-longevity trade-off in hosts. Journal of Theoretical Biology, 450(??):76–85, August 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301759.

Jain:2019:CAD

[JB19a] Sahil Jain and Manoj Baranwal. Computational analysis in designing T cell epitopes enriched peptides of Ebola glycoprotein exhibiting strong binding interaction with HLA molecules. Journal of Theoretical Biology, 465(??):34–44, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300165.

Jaworski:2019:SBP

[JB19b] Jacek Jaworski and Jason H. T. Bates. Sources of breathing pattern variability in the respiratory feedback control loop. Journal of Theoretical Biology, 469(??):148–162, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300943.

Jensen:2012:ASU

[JBSFB12] Kaare H. Jensen, Kirstine Berg-Sørensen, Søren M. M. Friis, and Tomas Bohr. Analytic solutions and universal properties of sugar loading models in Münch phloem flow. Journal of Theoretical Biology, 304(??):286–296, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001294.

Jian-Cheng:2010:STA

[JC10] Shi Jian-Cheng. Signal transduction and amplification in a circadian oscillator: Interaction between two colored noises. Journal of Theoretical Biology, 265(4):565–571, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310002833.

Jin:2016:TCS

[JC16] Weifeng Jin and Fangyue Chen. Topological chaos of the spatial prisoner’s dilemma game on regular networks. Journal of Theoretical Biology, 391(??):43–50, February 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005664.

Ju:2015:ITL

[JCG15] Zhe Ju, Jun-Zhe Cao, and Hong Gu. iLM-2L: a twolevel predictor for identifying protein lysine methylation sites and their methylation degrees by incorporating k-gap amino acid pairs into Chou’s general PseAAC. Journal of Theoretical Biology, 385(??):50–57, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003756.

Ju:2016:PLP

[JCG16] Zhe Ju, Jun-Zhe Cao, and Hong Gu. Predicting lysine phosphoglycerylation with fuzzy SVM by incorporating k-spaced amino acid pairs into Chou’s general PseAAC. Journal of Theoretical Biology, 397(??):145–150, May 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001168.

Jordan-Cooley:2011:BDE

[JCLS^{+}11] William C. Jordan-Cooley, Romuald N. Lipcius, Leah B. Shaw, Jian Shen, and Junping Shi. Bistability in a differential equation model of oyster reef height and sediment accumulation. Journal of Theoretical Biology, 289(??):1–11, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004085. See correc-

tion [JCLS^{+}21].

Jordan-Cooley:2021:CBD

[JCLS^{+}21] William C. Jordan-Cooley, Romuald N. Lipcius, Leah B. Shaw, Jian Shen, and Junping Shi. Corrigendum to “Bistability in a differential equation model of oyster reef height and sediment accumulation” [J. Theor. Biol. 289 (2011) 1– 11]. Journal of Theoretical Biology, 525(??):??, September 21, 2021. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519321001570. See [JCLS^{+}11].

Jover:2013:MMS

[JCW13] Luis F. Jover, Michael H. Cortez, and Joshua S. Weitz. Mechanisms of multi-strain coexistence in host-phage systems with nested infection networks. Journal of Theoretical Biology, 332(??):65–77, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001641.

Jaffar:2013:BRP

[JD13] M. Z. A. M. Jaffar and F. A. Davidson. Basic rules for polarised cell growth. Journal of Theoretical Biology, 336(??):44–51, November 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313003226.

Jiao:2016:PGRa

[JD16a] Ya-Sen Jiao and Pu-Feng Du. Predicting Golgi-resident protein types using pseudo amino acid compositions: Approaches with positional specific physicochemical properties. Journal of Theoretical Biology, 391(??):35–42, February 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315005585.

Jiao:2016:PGRb

[JD16b] Ya-Sen Jiao and Pu-Feng Du. Prediction of Golgi-resident protein types using general form of Chou’s pseudo-amino acid compositions: Approaches with minimal redundancy maximal relevance feature selection. Journal of Theoretical Biology, 402(??):38–44, August 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300601.

Jiao:2017:PPS

[JD17] Ya-Sen Jiao and Pu-Feng Du. Predicting protein submitochondrial locations by incorporating the positional-specific physicochemical properties into Chou’s general pseudo-amino acid compositions. Journal of Theoretical Biology, 416(??):81– 87, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316304362.

Jaberi-Douraki:2015:CMC

[JDPK15] Majid Jaberi-Douraki, Massimo Pietropaolo, and Anmar Khadra. Continuum model of T-cell avidity: Understanding autoreactive and regulatory T-cell responses in type 1 diabetes. Journal of Theoretical Biology, 383(??):93–105, October 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003781.

Jaberi-Douraki:2015:UCP

[JDSPK15] Majid Jaberi-Douraki, Santiago Schnell, Massimo Pietropaolo, and Anmar Khadra. Unraveling the contribution of pancreatic beta-cell suicide in autoimmune type 1 diabetes. Journal of Theoretical Biology, 375(??):77–87, June 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002707.

Joshi:2018:NAN

[JEA18] Tejas Joshi, Bret D. Elderd, and Karen C. Abbott. No appendix necessary: Fecal transplants and antibiotics can resolve Clostridium difficile infection. Journal of Theoretical Biology, 442(??):139–148, April 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300213.

Jensen:2010:LDO

[Jen10] Kenneth Kragh Jensen. Light-dependent orientation responses in animals can be explained by a model of compass cue integration. Journal of Theoretical Biology, 262(1):129–141, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004251.

Joshi:2014:ACE

[JG14] Saumitra V. Joshi and Sat. Ghosh. On the air cleansing efficiency of an extended green wall: a CFD analysis of mechanistic details of transport processes. Journal of Theoretical Biology, 361(??):101–110, November 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004159.

Jesse:2011:DCP

[JH11] Marieke Jesse and Hans Heesterbeek. Divide and conquer? Persistence of infectious agents in spatial metapopulations of hosts. Journal of Theoretical Biology, 275(1):12–20, April 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000440.

Johnston:2012:OVE

[JH12] Thomas R. Johnston and Mont Hubbard. Optimization of the visco-elastic parameters describing the heel-region of a prosthesis. Journal of Theoretical Biology, 311(??):1–7, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003104.

Janzen:2015:SFEa

[JHE15a] Thijs Janzen, Bart Haegeman, and Rampal S. Etienne. A sampling formula for ecological communities with multiple dispersal syndromes. Journal of Theoretical Biology, 374(??):94–106, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001277.

Janzen:2015:SFEb

[JHE15b] Thijs Janzen, Bart Haegeman, and Rampal S. Etienne. A sampling formula for ecological communities with multiple dispersal syndromes. Journal of Theoretical Biology, 387(??):258–261, December 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004774.

Jia:2016:SIM

[Jia16] Chen Jia. Simplification of irreversible Markov chains by removal of states with fast leaving rates. Journal of Theoretical Biology, 400(??):129–137, July 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300224.

John:2015:EDS

[JJ15] Sona John and Kavita Jain. Effect of drift, selection and recombination on the equilibrium frequency of deleterious mutations. Journal of Theoretical Biology, 365(??):238–246, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006134.

Jain:2017:FPN

[JJ17] Kavita Jain and Ananthu James. Fixation probability of a nonmutator in a large population of asexual mutators. Journal of Theoretical Biology, 433(??):85–93, November 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304010.

Jung:2012:HMP

[JK12] Inkyung Jung and Dongsup Kim. Histone modification profiles characterize function-specific gene regulation. Journal of Theoretical Biology, 310(??):132–142, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200286X.

Jenner:2019:OVT

[JKF19] Adrianne L. Jenner, Peter S. Kim, and Federico Frascoli. Oncolytic virotherapy for tumours following a Gompertz growth law. Journal of Theoretical Biology, 480(??):129–140, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931930308X.

Juher:2015:AEM

[JKS15] David Juher, Istvan Z. Kiss, and Joan Saldaña. Analysis of an epidemic model with awareness decay on regular random networks. Journal of Theoretical Biology, 365(??):457–468, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400602X.

Jeon:2011:SST

[JL11] Wonju Jeon and Sang-Hee Lee. Simulation study of territory size distributions in subterranean termites. Journal of Theoretical Biology, 279(1):1–8, June 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001561.

Jin:2016:HMM

[JL16] Huiqin Jin and Jinzhi Lei. A hybrid model of molecular regulation and population dynamics for yeast autophagy. Journal of Theoretical Biology, 402(??):45–53, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300467.

Jamieson-Lane:2015:FPS

[JLH15a] Alastair Jamieson-Lane and Christoph Hauert. Fixation probabilities on superstars, revisited and revised. Journal of Theoretical Biology, 382(??):44–56, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003082.

Jones:2015:LVD

[JLH^{+}15b] S. K. Jones, R. Laurenza, T. L. Hedrick, B. E. Griffith, and L. A. Miller. Lift vs. drag based mechanisms for vertical force production in the smallest flying insects. Journal of Theoretical Biology, 384(??):105–120, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003963.

Jia:2012:KBG

[JLQ^{+}12] Chen Jia, Xu-Feng Liu, Min-Ping Qian, Da-Quan Jiang, and Yu-Ping Zhang. Kinetic behavior of the general modifier mechanism of botts and Morales with non-equilibrium binding. Journal of Theoretical Biology, 296(??):13–20, March 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311005686.

Jia:2019:IPC

[JLQ^{+}19] Jianhua Jia, Xiaoyan Li, Wangren Qiu, Xuan Xiao, and Kuo-Chen Chou. iPPI-PseAAC(CGR): Identify protein–protein interactions by incorporating chaos game representation into PseAAC. Journal of Theoretical Biology, 460(??):195–203, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304971.

Jia:2015:IEE

[JLX^{+}15] Jianhua Jia, Zi Liu, Xuan Xiao, Bingxiang Liu, and Kuo-Chen Chou. iPPI-Esml: An ensemble classifier for identifying the interactions of proteins by incorporating their physicochemical properties and wavelet transforms into PseAAC. Journal of Theoretical Biology, 377(??):47–56, July 21, 2015.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001733.

Jia:2016:PLP

[JLX^{+}16] Jianhua Jia, Zi Liu, Xuan Xiao, Bingxiang Liu, and Kuo-Chen Chou. pSuc-lys: Predict lysine succinylation sites in proteins with PseAAC and ensemble random forest approach. Journal of Theoretical Biology, 394(??):223–230, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000539.

Jahandideh:2012:RSB

[JM12] Samad Jahandideh and Abbas Mahdavi. RFCRYS: Sequencebased protein crystallization propensity prediction by means of random forest. Journal of Theoretical Biology, 306(??): 115–119, August 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519312002093.

Johnston:2010:PCS

[JMC^{+}10] Matthew D. Johnston, Philip K. Maini, S. Jonathan Chapman, Carina M. Edwards, and Walter F. Bodmer. On the proportion of cancer stem cells in a tumour. Journal of Theoretical Biology, 266(4):708–711, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003887.

Jayathilake:2019:MMU

[JMH^{+}19] Chathranee Jayathilake, Philip K. Maini, Harriet W. Hopf, D. L. Sean McElwain, Helen M. Byrne, Mark B. Flegg, and Jennifer A. Flegg. A mathematical model of the use of supplemental oxygen to combat surgical site infection. Journal of Theoretical Biology, 466(??):11–23, April 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931930027X.

Jones:2017:DSS

[JMK^{+}17] Marjorie T. Jones, William R. Milligan, Lee B. Kats, Thomas L. Vandergon, Rodney L. Honeycutt, Robert N.

Fisher, Courtney L. Davis, and Timothy A. Lucas. A discrete stage-structured model of California newt population dynamics during a period of drought. Journal of Theoretical Biology, 414(??):245–253, February 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303769.

Jimenez-Munoz:2012:CEI

[JMS12] Juan C. Jiménez-Muñoz and José A. Sobrino. Comment on “ecological importance of the thermal emissivity of avian eggshells”. Journal of Theoretical Biology, 304(??):304–307, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001610.

Jin:2018:ELG

[JMS18] Wang Jin, Scott W. McCue, and Matthew J. Simpson. Extended logistic growth model for heterogeneous populations. Journal of Theoretical Biology, 445(??):51–61, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300870.

Jager:2013:DQS

[JMZ13] Tjalling Jager, Benjamin T. Martin, and Elke I. Zimmer. DE-Bkiss or the quest for the simplest generic model of animal life history. Journal of Theoretical Biology, 328(??):9–18, July 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931300115X.

Jain:2014:LER

[JN14] Rakesh K. Jain and Paolo A. Netti. Letter to the editor regarding recent article by Wu et al. [J. Theor. Biol. 355 (2014) pp. 194–207]. Journal of Theoretical Biology, 363(??):453, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005463. See corri-

gendum [WFC^{+}14].

Joly:2013:MIB

[JO13] Marcel Joly and Darci Odloak. Modeling interleukin-2-based immunotherapy in AIDS pathogenesis. Journal of Theoretical Biology, 335(??):57–78, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002919.

Jeong:2014:OGC

[JOAN14] Hyeong-Chai Jeong, Seung-Yoon Oh, Benjamin Allen, and Martin A. Nowak. Optional games on cycles and complete graphs. Journal of Theoretical Biology, 356(??):98–112, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002458.

Jedrak:2016:IGC

[JOM16] Jakub Jedrak and Anna Ochab-Marcinek. Influence of gene copy number on self-regulated gene expression. Journal of Theoretical Biology, 408(??):222–236, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302521.

Joshi:2013:ESP

[JP13] Badal Joshi and Mainak Patel. Encoding with synchrony: Phase-delayed inhibition allows for reliable and specific stimulus detection. Journal of Theoretical Biology, 328(??):26–32, July 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001161.

Jansen:2015:SGE

[JP15] Martin Jansen and Peter Pfaffelhuber. Stochastic gene expression with delay. Journal of Theoretical Biology, 364 (??):355–363, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519314005700.

Johnston:2019:ISL

[JP19] S. T. Johnston and K. J. Painter. The impact of shortand long-range perception on population movements. Journal of Theoretical Biology, 460(??):227–242, January 7, 2019.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305071.

Jones:2017:FPE

[JPBM17] Candace Jones, Brad Pearce, Ingrid Barrera, and Amanda Mummert. Fetal programming and eating disorder risk. Journal of Theoretical Biology, 428(??):26–33, September 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302503.

Jombart:2010:PPA

[JPDP10] Thibaut Jombart, Sandrine Pavoine, Sébastien Devillard, and Dominique Pontier. Putting phylogeny into the analysis of biological traits: a methodological approach. Journal of Theoretical Biology, 264(3):693–701, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001736.

Jiao:2017:PIR

[JR17a] Xiong Jiao and Shoba Ranganathan. Prediction of interface residue based on the features of residue interaction network. Journal of Theoretical Biology, 432(??):49–54, November 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303806.

Jordan:2017:TPP

[JR17b] Jillian J. Jordan and David G. Rand. Third-party punishment as a costly signal of high continuation probabilities in repeated games. Journal of Theoretical Biology, 421(??):189– 202, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301522.

Johnston:2016:QEE

[JRB^{+}16] Stuart T. Johnston, Joshua V. Ross, Benjamin J. Binder, D. L. Sean McElwain, Parvathi Haridas, and Matthew J. Simpson. Quantifying the effect of experimental design choices for in vitro scratch assays. Journal of Theoretical Biology, 400(??):19–31, July 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300406.

Jithinraj:2014:CZO

[JRG14a] P. K. Jithinraj, Ushasi Roy, and Manoj Gopalakrishnan. Corrigendum to “Zero-order ultrasensitivity: a study of criticality and fluctuations under the total quasi-steady state approximation in the linear noise regime” [J. Theor. Biol. 344 (2014) 1–11]. Journal of Theoretical Biology, 349(??):170, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000940. See [JRG14c].

Jithinraj:2014:EZO

[JRG14b] P. K. Jithinraj, Ushasi Roy, and Manoj Gopalakrishnan. Erratum to “Zero-order ultrasensitivity: a study of criticality and fluctuations under the total quasi-steady state approximation in the linear noise regime” [J. Theor. Biol. 344 (2014) 1–11]. Journal of Theoretical Biology, 350(??):111, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000939. See [JRG14c].

Jithinraj:2014:ZOU

[JRG14c] P. K. Jithinraj, Ushasi Roy, and Manoj Gopalakrishnan. Zeroorder ultrasensitivity: a study of criticality and fluctuations under the total quasi-steady state approximation in the linear noise regime. Journal of Theoretical Biology, 344(??):1–11, March 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005419. See corri-

gendum [JRG14a] and erratum [JRG14b].

Johansson:2012:TPF

[JRMS12] Anders Johansson, Kai Ramsch, Martin Middendorf, and David J. T. Sumpter. Tuning positive feedback for signal detection in noisy dynamic environments. Journal of Theoretical Biology, 309(??):88–95, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002652.

Jahangiri:2012:PDM

[JRR^{+}12] Abolfazl Jahangiri, Iraj Rasooli, Mohammad Reza Rahbar, Saeed Khalili, Jafar Amani, and Kobra Ahmadi Zanoos. Precise detection of L. monocytogenes hitting its highly conserved region possessing several specific antibody binding sites. Journal of Theoretical Biology, 305(??):15–23, July 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312001919.

Jiu-sheng:2012:EDF

[Js12] Ren Jiu-sheng. Effects of dispersion of fiber orientation on the mechanical property of the arterial wall. Journal of Theoretical Biology, 301(??):153–160, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000926.

Jimenez-Sanchez:2015:CRS

[JS15] Alfonso Jiménez-Sánchez. Chromosome replication status and DNA content at any cell age in a bacterial cell cycle. Journal of Theoretical Biology, 380(??):585–589, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002878.

Jimenez-Sanchez:2017:BCC

[JS17] Alfonso Jiménez-Sánchez. Bacterial cell cycle classification. Application to DNA synthesis and DNA content at any cell age. Journal of Theoretical Biology, 419(??):8–12, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300541.

Jimenez-Sanchez:2018:CBC

[JS18] Alfonso Jiménez-Sánchez. Corrigendum to “Bacterial cell cycle classification. Application to DNA synthesis and DNA content at any cell age” [Journal of Theoretical Biology 419 (2017) 8–12]. Journal of Theoretical Biology, 446(??):260, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300018.

Johnston:2015:MMI

[JSB15] Stuart T. Johnston, Matthew J. Simpson, and Ruth E. Baker. Modelling the movement of interacting cell populations: a moment dynamics approach. Journal of Theoretical Biology, 370(??):81–92, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500034X.

Jain:2016:SHA

[JSC^{+}16] Ravi Jain, Subash C. Sonkar, Uma Chaudhry, Manju Bala, and Daman Saluja. In-silico hierarchical approach for the identification of Potential Universal Vaccine Candidates (PUVCs) from Neisseria gonorrhoeae. Journal of Theoretical Biology, 410(??):36–43, December 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302867.

Jones:2014:ADM

[JSDEK14] Leslie Braziel Jones, Timothy W. Secomb, Mark W. Dewhirst, and Ardith W. El-Kareh. The additive damage model: a mathematical model for cellular responses to drug combinations. Journal of Theoretical Biology, 357(??):10–20, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002628.

Jingbo:2011:UCP

[JSF^{+}11] Xia Jingbo, Zhang Silan, Shi Feng, Xiong Huijuan, Hu Xuehai, Niu Xiaohui, and Li Zhi. Using the concept of pseudo amino acid composition to predict resistance gene against Xanthomonas oryzae pv. oryzae in rice: An approach from chaos games representation. Journal of Theoretical Biology, 284(1):16–23, September 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003018.

Jesionek:2016:LDL

[JSK16] Katarzyna Jesionek, Aleksandra Slapik, and Marcin Kostur. Low-density lipoprotein transport through an arterial wall under hypertension — a model with time and pressure dependent fraction of leaky junction consistent with experiments. Journal of Theoretical Biology, 411(??):81–91, December 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303174.

Jin:2016:RSA

[JSP^{+}16] Wang Jin, Esha T. Shah, Catherine J. Penington, Scott W. McCue, Lisa K. Chopin, and Matthew J. Simpson. Reproducibility of scratch assays is affected by the initial degree of confluence: Experiments, modelling and model selection. Journal of Theoretical Biology, 390(??):136–145, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005676.

Jiang:2012:ABM

[JSSZ12] Wen Jiang, Adam M. Sullivan, Chunlei Su, and Xiaopeng Zhao. An agent-based model for the transmission dynamics of Toxoplasma gondii. Journal of Theoretical Biology, 293(??):15–26, January 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005133.

Jiang:2019:RPU

[JSWY19] Jie Jiang, Anglu Shen, Hao Wang, and Sanling Yuan. Regulation of phosphate uptake kinetics in the bloomforming dinoflagellates Prorocentrum donghaiense with emphasis on two-stage dynamic process. Journal of Theoretical Biology, 463(??):12–21, February 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306015.

Jedrzejewska-Szmek:2010:DSD

[JSZ10] Joanna Jedrzejewska-Szmek and JarosLaw Zygierewicz. Depressing synapse as a detector of frequency change. Journal of Theoretical Biology, 266(3):380–390, October 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003188.

Jahandideh:2012:CCA

[JSZ12] Samad Jahandideh, Vinodh Srinivasasainagendra, and Degui Zhi. Comprehensive comparative analysis and identification of RNA-binding protein domains: Multi-class classification and feature selection. Journal of Theoretical Biology, 312(??):65–75, November 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003475.

Jade:2014:FEA

[JTSG14] Sameer Jade, Kelli H. Tamvada, David S. Strait, and Ian R. Grosse. Finite element analysis of a femur to deconstruct the paradox of bone curvature. Journal of Theoretical Biology, 341(??):53–63, January 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004402.

Juska:2011:MMG

[Jus11] Alfonsas Juska. Minimal models of growth and decline of microbial populations. Journal of Theoretical Biology, 269(1):195–200, January 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005643.

James:2014:FNR

[JV14] William H. James and John Valentine. A further note on the rises in sex ratio at birth during and just after the two world wars. Journal of Theoretical Biology, 363(??):404–411, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004524.

Ju:2018:PSS

[JW18] Zhe Ju and Shi-Yun Wang. Prediction of s-sulfenylation sites using mRMR feature selection and fuzzy support vector machine algorithm. Journal of Theoretical Biology, 457(??):6–13, November 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304004.

Wang:2019:EGD

[jWlGQ19] Xian jia Wang, Cui ling Gu, and Ji Quan. Evolutionary game dynamics of the Wright–Fisher process with different selection intensities. Journal of Theoretical Biology, 465(??):17–26, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300062.

Ji:2010:CBP

[JWS^{+}10] Guoli Ji, Xiaohui Wu, Yingjia Shen, Jiangyin Huang, and Qingshun Quinn Li. A classification-based prediction model of messenger RNA polyadenylation sites. Journal of Theoretical Biology, 265(3):287–296, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000250X.

Jing:2017:DIA

[JWTW17] Yi Jing, Xia Wang, Sanyi Tang, and Jianhong Wu. Data informed analysis of 2014 dengue fever outbreak in Guangzhou: Impact of multiple environmental factors and vector control. Journal of Theoretical Biology, 416(??):161–179, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316304234.

Jenner:2018:MHV

[JYY^{+}18] Adrianne Jenner, Chae-Ok Yun, Arum Yoon, Peter S. Kim, and Adelle C. F. Coster. Modelling heterogeneity in viraltumour dynamics: The effects of gene-attenuation on viral characteristics. Journal of Theoretical Biology, 454(??):41–52, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302777.

Jia:2018:NPS

[JYZ18] Cangzhi Jia, Qing Yang, and Quan Zou. NucPosPred: Predicting species-specific genomic nucleosome positioning via four different modes of general PseKNC. Journal of Theoretical Biology, 450(??):15–21, August 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301929.

Jia:2017:SSP

[JZ17] Cangzhi Jia and Yun Zuo. S-SulfPred: a sensitive predictor to capture s-sulfenylation sites based on a resampling onesided selection undersampling-synthetic minority oversampling technique. Journal of Theoretical Biology, 422(??):84– 89, June 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301364.

Jaruszewicz:2013:TND

[JZL13] Joanna Jaruszewicz, Paweł J. Zuk, and Tomasz Lipniacki. Type of noise defines global attractors in bistable molecular regulatory systems. Journal of Theoretical Biology, 317(??):140–151, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005292.

Jiang:2011:PTT

[JZZK11] Bo Jiang, Xiao Zhang, Yijun Zuo, and Guolian Kang. A powerful truncated tail strength method for testing multiple null hypotheses in one dataset. Journal of Theoretical Biology, 277(1):67–73, May 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000415.

Kang:2011:DED

[KA11] Yun Kang and Dieter Armbruster. Dispersal effects on a discrete two-patch model for plant-insect interactions. Journal of Theoretical Biology, 268(1):84–97, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005102.

Krick:2015:AEP

[KA15] Julian Krick and Josef Daniel Ackerman. Adding ecology to particle capture models: Numerical simulations of capture on a moving cylinder in crossflow. Journal of Theoretical Biology, 368(??):13–26, March 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006924.

Kotini:2016:ADT

[KA16] A. Kotini and P. Anninos. Alpha, delta and theta rhythms in a neural net model. Comparison with MEG data. Journal of Theoretical Biology, 388(??):11–14, January 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004981.

Kuruoglu:2017:ICG

[KA17] Ercan E. Kuruoğlu and Peter F. Arndt. The information capacity of the genetic code: Is the natural code optimal? Journal of Theoretical Biology, 419(??):227–237, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300553.

Kaehler:2017:FRN

[Kae17] Benjamin D. Kaehler. Full reconstruction of non-stationary strand-symmetric models on rooted phylogenies. Journal of Theoretical Biology, 420(??):144–151, May 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301194.

Koyama:2019:CSI

[KAKK19] Kento Koyama, Hiroki Abe, Shuso Kawamura, and Shige Koseki. Calculating stochastic inactivation of individual cells in a bacterial population using variability in individual cell inactivation time and initial cell number. Journal of Theoretical Biology, 469(??):172–179, May 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300293.

Kamimura:2011:PSR

[Kam11] Yoshitaka Kamimura. Population stochasticity, random determination of handedness, and the genetic basis of antisymmetry. Journal of Theoretical Biology, 290(??):73–80, December 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004243.

Krishnan:2011:IDP

[KAN11] J. Krishnan and Aiman Alam-Nazki. An investigation of design principles underlying repulsive and attractive gradient sensing and their switching. Journal of Theoretical Biology, 273(1):80–99, March 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310006417.

Kedziora:2012:PBQ

[KAPR12] D. J. Kedziora, R. G. Abeysuriya, A. J. K. Phillips, and P. A. Robinson. Physiologically based quantitative modeling of unihemispheric sleep. Journal of Theoretical Biology, 314(??):109–119, December 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004535.

Karenlampi:2011:ADT

[Kär11] Petri P. Kärenlampi. Age distribution of trees in stationary forest system. Journal of Theoretical Biology, 270(1):13–18, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006041.

Karonen:2012:STC

[Kar12] Ilmari Karonen. Stable trimorphic coexistence in a lattice model of spatial competition with two site types. Journal of Theoretical Biology, 295(??):77–85, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005376.

Kareva:2015:IET

[Kar15] Irina Kareva. Immune evasion through competitive inhibition: The shielding effect of cancer non-stem cells. Journal of Theoretical Biology, 364(??):40–48, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004962.

Kareva:2016:ETD

[Kar16] Irina Kareva. Escape from tumor dormancy and time to angiogenic switch as mitigated by tumor-induced stimulation of stroma. Journal of Theoretical Biology, 395(??):11–22, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000576.

Kember:2011:NCH

[KAZ11] G. Kember, J. A. Armour, and M. Zamir. Neural control of heart rate: The role of neuronal networking. Journal of Theoretical Biology, 277(1):41–47, May 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100107X.

Kember:2013:DNN

[KAZ13] G. Kember, J. A. Armour, and M. Zamir. Dynamic neural networking as a basis for plasticity in the control of heart rate. Journal of Theoretical Biology, 317(??):39–46, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005024.

Khrustalev:2010:LCU

[KB10a] Vladislav Victorovich Khrustalev and Eugene Victorovich Barkovsky. The level of cytosine is usually much higher than the level of guanine in two-fold degenerated sites from third codon positions of genes from simplex- and varicelloviruses with G+C higher than 50%. Journal of Theoretical Biology, 266(1):88–98, September 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003139.

Kononenko:2010:MSA

[KB10b] Nikolai I. Kononenko and Nataliya M. Berezetskaya. Modeling the spontaneous activity in suprachiasmatic nucleus neurons: Role of cation single channels. Journal of Theoretical Biology, 265(2):115–125, July 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001748.

Khrustalev:2011:PHI

[KB11] Vladislav V. Khrustalev and Eugene V. Barkovsky. Percent of highly immunogenic amino acid residues forming b-cell epitopes is higher in homologous proteins encoded by GC-rich genes. Journal of Theoretical Biology, 282(1):71–79, August 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002608.

Kareva:2015:CIP

[KB15] Irina Kareva and Faina Berezovskaya. Cancer immunoediting: a process driven by metabolic competition as a predator– prey-shared resource type model. Journal of Theoretical Biology, 380(??):463–472, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002866.

Khetan:2019:AFN

[KB19] Jawahar Khetan and Dipak Barua. Analysis of Fn14-NF-κB signaling response dynamics using a mechanistic model. Journal of Theoretical Biology, 480(??):34–42, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303017.

Kesseler:2013:PMM

[KBE^{+}13] Kevin J. Kesseler, Michael L. Blinov, Timothy C. Elston, William K. Kaufmann, and Dennis A. Simpson. A predictive mathematical model of the DNA damage G2 checkpoint. Journal of Theoretical Biology, 320(??):159–169, March 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200639X.

Kursawe:2018:ABC

[KBF18] Jochen Kursawe, Ruth E. Baker, and Alexander G. Fletcher. Approximate Bayesian computation reveals the importance of repeated measurements for parameterising cell-based models of growing tissues. Journal of Theoretical Biology, 443(??):66– 81, April 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300365.

Khrustalev:2016:MMB

[KBK16] Vladislav Victorovich Khrustalev, Eugene Victorovich Barkovsky, and Tatyana Aleksandrovna Khrustaleva. Magnesium and manganese binding sites on proteins have the same predominant motif of secondary structure. Journal of Theoretical Biology, 395(??):174–185, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000965.

Kilian:2014:MGP

[KBKK14] H. G. Kilian, D. Bartkowiak, M. Kazda, and D. Kaufmann. Modelling the growth of plants with a uniform growth logistics. Journal of Theoretical Biology, 349(??):57–65, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000368.

Khayyeri:2015:PCM

[KBL15] Hanifeh Khayyeri, Sara Barreto, and Damien Lacroix. Primary cilia mechanics affects cell mechanosensation: a computational study. Journal of Theoretical Biology, 379(??):38–46, August 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002167.

King:2019:MRD

[KBS^{+}19] Clarence C. King, Amelia Ann Brown, Irmak Sargin, K. M. Bratlie, and S. P. Beckman. Modeling of reaction–diffusion transport into a core-shell geometry. Journal of Theoretical Biology, 460(??):204–208, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304624. See [KBS^{+}20].

King:2020:CMR

[KBS^{+}20] Clarence C. King, Amelia Ann Brown, Irmak Sargin, K. M. Bratlie, and S. P. Beckman. Corrigendum to: Modeling of reaction–diffusion transport into a core-shell geometry. Journal of Theoretical Biology, 507(??):Article 110439, December 21, 2020. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519320302940. See [KBS^{+}19].

Keinj:2011:MMB

[KBV11] R. Keinj, T. Bastogne, and P. Vallois. Multinomial modelbased formulations of TCP and NTCP for radiotherapy treatment planning. Journal of Theoretical Biology, 279(1):55–62, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001718.

Keinj:2012:TGM

[KBV12] R. Keinj, T. Bastogne, and P. Vallois. Tumor growth modeling based on cell and tumor lifespans. Journal of Theoretical Biology, 312(??):76–86, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003426.

Kurbatova:2015:MMC

[KBV^{+}15] P. Kurbatova, N. Bessonov, V. Volpert, H. A. W. M. Tiddens, C. Cornu, P. Nony, and D. Caudri. Model of mucociliary clearance in cystic fibrosis lungs. Journal of Theoretical Biology, 372(??):81–88, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000946.

Krepkin:2011:DRC

[KC11a] Konstantin Krepkin and José Costa. Defining the role of cooperation in early tumor progression. Journal of Theoretical Biology, 285(1):36–45, September 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003407.

Kurochkina:2011:HHI

[KC11b] Natalya Kurochkina and Tsering Choekyi. Helix-helix interfaces and ligand binding. Journal of Theoretical Biology, 283(1):92–102, August 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002566.

Kucken:2013:MCI

[KC13] Michael Kücken and Christophe Champod. Merkel cells and the individuality of friction ridge skin. Journal of Theoretical Biology, 317(??):229–237, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005346.

Kwon:2014:GHM

[KC14] H. W. Kwon and M. Y. Choi. Generalized hydromechanical model for stomatal responses to hydraulic perturbations. Journal of Theoretical Biology, 340(??):119–130, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931300444X.

Krivan:2017:ITC

[KC17] Vlastimil Křivan and Ross Cressman. Interaction times change evolutionary outcomes: Two-player matrix games. Journal of Theoretical Biology, 416(??):199–207, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300103.

Korell:2011:ERB

[KCD11a] Julia Korell, Carolyn V. Coulter, and Stephen B. Duffull. Evaluation of red blood cell labelling methods based on a statistical model for red blood cell survival. Journal of Theoretical Biology, 291(??):88–98, December 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004723.

Korell:2011:SMR

[KCD11b] Julia Korell, Carolyn V. Coulter, and Stephen B. Duffull. A statistical model for red blood cell survival. Journal of Theoretical Biology, 268(1):39–49, January 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005382.

Krakauer:2011:CST

[KCE^{+}11] David C. Krakauer, James P. Collins, Douglas Erwin, Jessica C. Flack, Walter Fontana, Manfred D. Laubichler, Sonja J. Prohaska, Geoffrey B. West, and Peter F. Stadler. The challenges and scope of theoretical biology. Journal of Theoretical Biology, 276(1):269–276, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000774.

Kochova:2011:HAV

[KCJ^{+}11] Petra Kochová, Robert Cimrman, Jirí Janácek, Kirsti Witter, and Zbynek Tonar. How to assess, visualize and compare the anisotropy of linear structures reconstructed from optical sections — a study based on histopathological quantification of human brain microvessels. Journal of Theoretical Biology, 286(??):67–78, October 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003481.

Kandaswamy:2011:APR

[KCM^{+}11] Krishna Kumar Kandaswamy, Kuo-Chen Chou, Thomas Martinetz, Steffen Möller, P. N. Suganthan, S. Sridharan, and Ganesan Pugalenthi. AFP-pred: a random forest approach for predicting antifreeze proteins from sequence-derived properties. Journal of Theoretical Biology, 270(1):56–62, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005849.

Kang:2011:MMO

[KCMF11] Yun Kang, Rebecca Clark, Michael Makiyama, and Jennifer Fewell. Mathematical modeling on obligate mutualism: Interactions between leaf-cutter ants and their fungus garden. Journal of Theoretical Biology, 289(??):116–127, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004309.

Kochova:2015:MMC

[KCS^{+}15] Petra Kochová, Robert Cimrman, Milan Stengl, Bohuslav Ostádal, and Zbynek Tonar. A mathematical model of the carp heart ventricle during the cardiac cycle. Journal of Theoretical Biology, 373(??):12–25, May 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500123X.

Kaiping:2016:CPC

[KCS16] G. A. Kaiping, S. J. Cox, and T. J. Sluckin. Cooperation and punishment in community-structured populations with migration. Journal of Theoretical Biology, 405(??):116–126, September 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316000199.

Kissler:2014:DPD

[KCSB14] S. M. Kissler, C. Cichowitz, S. Sankaranarayanan, and D. M. Bortz. Determination of personalized diabetes treatment plans using a two-delay model. Journal of Theoretical Biology, 359(??):101–111, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003427.

Karami:2019:ARC

[KCZ^{+}19] Mina Karami, Begoña Calvo, Hassan Zohoor, Keikhosrow Firoozbakhsh, and Jorge Grasa. Assessing the role of Ca^{2+} in skeletal muscle fatigue using a multi-scale continuum model. Journal of Theoretical Biology, 461(??):76–83, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305101.

Keegan:2016:EFP

[KD16] Lindsay T. Keegan and Jonathan Dushoff. Estimating finitepopulation reproductive numbers in heterogeneous populations. Journal of Theoretical Biology, 397(??):1–12, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000552.

Kettle:2014:PFP

[KDFM14] Helen Kettle, Ruairi Donnelly, Harry J. Flint, and Glenn Marion. pH feedback and phenotypic diversity within bacterial functional groups of the human gut. Journal of Theoretical Biology, 342(??):62–69, February 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005134.

Kamran-Disfani:2013:LCI

[KDG13] Ahmad R. Kamran-Disfani and Antonio J. Golubski. Lateral cascade of indirect effects in food webs with different types of adaptive behavior. Journal of Theoretical Biology, 339(??):58–69, December 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002944.

Kowald:2014:EMM

[KDK14a] Axel Kowald, Marcus Dawson, and Thomas B. L. Kirkwood. Erratum to “Mitochondrial mutations and ageing: Can mitochondrial deletion mutants accumulate via a size based replication advantage?” [J. Theor. Biol. 340 (2014) 111– 118]. Journal of Theoretical Biology, 349(??):171, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000836. See corrigen-

dum [KDK14b].

Kowald:2014:MMA

[KDK14b] Axel Kowald, Marcus Dawson, and Thomas B. L. Kirkwood. Mitochondrial mutations and ageing: Can mitochondrial deletion mutants accumulate via a size based replication advantage? Journal of Theoretical Biology, 340(??):111–118, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004372. See erra-

tum [KDK14a].

Kurian:2016:HQE

[KDL16] P. Kurian, G. Dunston, and J. Lindesay. How quantum entanglement in DNA synchronizes double-strand breakage by type II restriction endonucleases. Journal of Theoretical Biology, 391(??):102–112, February 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500569X.

Kim:2018:MMI

[KdlRJ18] Soyoung Kim, Aurelio A. de los Reyes, and Eunok Jung. Mathematical model and intervention strategies for mitigating tuberculosis in the Philippines. Journal of Theoretical Biology, 443(??):100–112, April 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300420.

Karimi:2013:NCI

[KDMD13] Safoora Karimi, Mitra Dadvar, Hamid Modarress, and Bahram Dabir. A new correlation for inclusion of leaky junctions in macroscopic modeling of atherosclerotic lesion initiation. Journal of Theoretical Biology, 329(??):94–100, July 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000921.

Koffel:2016:GEE

[KDMK16] Thomas Koffel, Tanguy Daufresne, François Massol, and Christopher A. Klausmeier. Geometrical envelopes: Extending graphical contemporary niche theory to communities and eco-evolutionary dynamics. Journal of Theoretical Biology, 407(??):271–289, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302119.

Kodikara:2018:IAE

[KDS18] Saritha Kodikara, Haydar Demirhan, and Lewi Stone. Inferring about the extinction of a species using certain and uncertain sightings. Journal of Theoretical Biology, 442(??):98–109, April 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300237.

Kaito:2015:BPD

[KDST15] Chiho Kaito, Ulf Dieckmann, Akira Sasaki, and Fugo Takasu. Beyond pairs: Definition and interpretation of third-order structure in spatial point patterns. Journal of Theoretical Biology, 372(??):22–38, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000636.

Kritz:2010:OMN

[KdSUS10] Maurício Vieira Kritz, Marcelo Trindade dos Santos, Sebastián Urrutia, and Jean-Marc Schwartz. Organising metabolic networks: Cycles in flux distributions. Journal of Theoretical Biology, 265(3):250–260, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002171.

Keane:2016:CCH

[Kea16] Christopher Keane. Chaos in collective health: Fractal dynamics of social learning. Journal of Theoretical Biology, 409(??):47–59, November 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931630282X.

Keener:2010:MRM

[Kee10] J. P. Keener. A molecular ruler mechanism for length control of extended protein structures in bacteria. Journal of Theoretical Biology, 263(4):481–489, April 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005827.

Koch:2017:CBN

[KEG17] Daniel Koch, Robert S. Eisinger, and Alexander Gebharter. A causal Bayesian network model of disease progression mechanisms in chronic myeloid leukemia. Journal of Theoretical Biology, 433(??):94–105, November 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303983.

Khrustalev:2017:TAM

[KEHK17] Vladislav Victorovich Khrustalev, Marina Anatolyevna Ermalovich, Judith M. Hübschen, and Tatyana Aleksandrovna Khrustaleva. Transcription-associated mutational pressure in the parvovirus b19 genome: Reactivated genomes contribute to the variability of viral populations. Journal of Theoretical Biology, 435(??):199–207, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730437X.

Khamis:2018:EEU

[KEKB18] Doran Khamis, Claire El Mouden, Klodeta Kura, and Michael B. Bonsall. Ecological effects on underdominance threshold drives for vector control. Journal of Theoretical Biology, 456(??):1–15, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303461.

Kenakin:2019:GPR

[Ken19] Terry Kenakin. G-protein/receptor inhibitors as blockers of receptor signaling. Journal of Theoretical Biology, 480(??):23–33, November 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302991.

Kwansa:2010:EES

[KF10] Albert L. Kwansa and Joseph W. Freeman. Elastic energy storage in an unmineralized collagen type I molecular model with explicit solvation and water infiltration. Journal of Theoretical Biology, 262(4):691–697, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005050.

Kulich:2011:PEM

[KF11] Tomás Kulich and Jaroslav Flegr. Positive effects of multiple gene control on the spread of altruism by group selection. Journal of Theoretical Biology, 284(1):1–6, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002591.

Kremling:2012:SAC

[KF12] Andreas Kremling and Dietrich Flockerzi. Structural analysis of a core model for carbohydrate uptake in Escherichia coli. Journal of Theoretical Biology, 303(??):62–74, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200121X.

Kou:2015:IFK

[KF15] Gaoshan Kou and Yonge Feng. Identify five kinds of simple super-secondary structures with quadratic discriminant algorithm based on the chemical shifts. Journal of Theoretical Biology, 380(??):392–398, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002854.

Khatri:2012:OMD

[KFA12] Bhavin S. Khatri, Andrew Free, and Rosalind J. Allen. Oscillating microbial dynamics driven by small populations, limited nutrient supply and high death rates. Journal of Theoretical Biology, 314(??):120–129, December 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004171.

Kebir:2018:GSM

[KFB18] Amira Kebir, Nina H. Fefferman, and Slimane Ben Miled. A general structured model of a hermaphrodite population. Journal of Theoretical Biology, 449(??):53–59, July 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301711.

Knipfer:2014:RUP

[KFG^{+}14] T. Knipfer, J. Fei, G. A. Gambetta, K. A. Shackel, and M. A. Matthews. A revised and unified pressure-clamp/ relaxation theory for studying plant cell water relations with pressure probes: In-situ determination of cell volume for calculation of volumetric elastic modulus and hydraulic conductivity. Journal of Theoretical Biology, 359(??):80–91, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400318X.

Kershenbaum:2015:EVR

[KFG15] Arik Kershenbaum, Todd M. Freeberg, and David E. Gammon. Estimating vocal repertoire size is like collecting coupons: a theoretical framework with heterogeneity in signal abundance. Journal of Theoretical Biology, 373(??):1–11, May 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001125.

Kocieniewski:2012:IDP

[KFL12] PaweL Kocieniewski, James R. Faeder, and Tomasz Lipniacki. The interplay of double phosphorylation and scaffolding in MAPK pathways. Journal of Theoretical Biology, 295(??):116–124, February 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005844.

Konrad:2013:AMS

[KFS^{+}13] Wilfried Konrad, Felix Flues, Florian Schmich, Thomas Speck, and Olga Speck. An analytic model of the self-sealing mechanism of the succulent plant Delosperma cooperi. Journal of Theoretical Biology, 336(??):96–109, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003366.

Knudsen:2012:EAI

[KFW12] Michael Knudsen, Elisenda Feliu, and Carsten Wiuf. Exact analysis of intrinsic qualitative features of phosphorelays using mathematical models. Journal of Theoretical Biology, 300(??): 7–18, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000082.

Kozlovsky:2012:RCP

[KG12] Pavel Kozlovsky and Amit Gefen. The relative contributions of propulsive forces and receptor-ligand binding forces during early contact between spermatozoa and zona pellucida of oocytes. Journal of Theoretical Biology, 294(??):139–143, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005649.

Kar:2013:SEC

[KG13] T. K. Kar and Bapan Ghosh. Sustainability and economic consequences of creating marine protected areas in multispecies multiactivity context. Journal of Theoretical Biology, 318(??):81–90, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005796.

Khatri:2015:CGB

[KG15] Bhavin S. Khatri and Richard A. Goldstein. A coarsegrained biophysical model of sequence evolution and the population size dependence of the speciation rate. Journal of Theoretical Biology, 378(??):56–64, August 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002039.

Kisdi:2013:EPV

[KGB13] Eva Kisdi, Stefan A. H. Geritz, and Barbara Boldin. Evolution of pathogen virulence under selective predation: a construction method to find eco-evolutionary cycles. Journal of Theoretical Biology, 339(??):140–150, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931300252X.

Krivan:2018:BRD

[KGC18] Vlastimil Křivan, Theodore E. Galanthay, and Ross Cressman. Beyond replicator dynamics: From frequency to density dependent models of evolutionary games. Journal of Theoretical Biology, 455(??):232–248, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303163.

Kar:2012:DPP

[KGJ12] T. K. Kar, Abhijit Ghorai, and Soovoojeet Jana. Dynamics of pest and its predator model with disease in the pest and optimal use of pesticide. Journal of Theoretical Biology, 310(??):187–198, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003256.

Kim:2010:MTD

[KGK10] Byul Nim Kim, Luis F. Gordillo, and Yongkuk Kim. A model for the transmission dynamics of Orientia tsutsugamushi among its natural reservoirs. Journal of Theoretical Biology, 266(1):154–161, September 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002997.

Korvasova:2015:ITC

[KGM^{+}15a] K. Korvasová, E. A. Gaffney, P. K. Maini, M. A. Ferreira, and V. Klika. Investigating the Turing conditions for diffusiondriven instability in the presence of a binding immobile substrate. Journal of Theoretical Biology, 367(??):286–295, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006766.

Kumar:2015:SRD

[KGM15b] C. Selvaa Kumar, Nikhil Gadewal, and Sudheer MM Mohammed. Seminal role of deletion of amino acid residues in H1-S2 and S-loop regions in eukaryotic β-tubulin investigated from docking and dynamics perspective. Journal of Theoretical Biology, 378(??):79–88, August 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002179.

Khan:2015:TCI

[KGP^{+}15] Faez Iqbal Khan, Algasan Govender, Kugen Permaul, Suren Singh, and Krishna Bisetty. Thermostable chitinase II from Thermomyces lanuginosus SSBP: Cloning, structure prediction and molecular dynamics simulations. Journal of Theoretical Biology, 374(??):107–114, June 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001502.

Kwang-Hua:2019:TDV

[KH19] Chu W. Kwang-Hua. Temperature-dependent viscosity dominated transport control through AQP1 water channel. Journal of Theoretical Biology, 480(??):92–98, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303182.

Kharabian:2010:ECM

[Kha10] Ardashir Kharabian. An efficient computational method for screening functional SNPs in plants. Journal of Theoretical Biology, 265(1):55–62, July 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002006.

Klein:2018:FEM

[KHB^{+}18] Hartmut Klein, Linnea Hesse, Matthias Boljen, Tim Kampowski, Irina Butschek, Thomas Speck, and Olga Speck. Finite element modelling of complex movements during selfsealing of ring incisions in leaves of Delosperma cooperi. Journal of Theoretical Biology, 458(??):184–206, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303990.

Kaunas:2010:KMC

[KHH10] Roland Kaunas, Zuyi Huang, and Juergen Hahn. A kinematic model coupling stress fiber dynamics with JNK activation in response to matrix stretching. Journal of Theoretical Biology, 264(2):593–603, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000950.

Krishnamoorthy:2017:HMH

[KHH^{+}17] Ezhilarasi Krishnamoorthy, Sameer Hassan, Luke Elizabeth Hanna, Indira Padmalayam, Rama Rajaram, and Vijay Viswanathan. Homology modeling of Homo sapiens lipoic acid synthase: Substrate docking and insights on its binding mode. Journal of Theoretical Biology, 420(??):259–266, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302879.

Khan:2015:DAA

[KHK15] Zaheer Ullah Khan, Maqsood Hayat, and Muazzam Ali Khan. Discrimination of acidic and alkaline enzyme using Chou’s pseudo amino acid composition in conjunction with probabilistic neural network model. Journal of Theoretical Biology, 365(??):197–203, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006031.

Khan:2017:BPP

[KHK^{+}17] Muslim Khan, Maqsood Hayat, Sher Afzal Khan, Saeed Ahmad, and Nadeem Iqbal. Bi-PSSM: Position specific scoring matrix based intelligent computational model for identification of mycobacterial membrane proteins. Journal of Theoretical Biology, 435(??):116–124, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304307.

Khan:2017:UDI

[KHKI17] Muslim Khan, Maqsood Hayat, Sher Afzal Khan, and Nadeem Iqbal. Unb-DPC: Identify mycobacterial membrane protein types by incorporating un-biased dipeptide composition into Chou’s general PseAAC. Journal of Theoretical Biology, 415(??):13–19, February 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304143.

Kimori:2016:QMF

[KHNM16] Yoshitaka Kimori, Kazumi Hikino, Mikio Nishimura, and Shoji Mano. Quantifying morphological features of actin cytoskeletal filaments in plant cells based on mathematical morphology. Journal of Theoretical Biology, 389(??):123–131, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005305.

Kazmi:2012:ATG

[KHP^{+}12] N. Kazmi, M. A. Hossain, R. M. Phillips, M. A. Al-Mamun, and R. Bass. Avascular tumour growth dynamics and the constraints of protein binding for drug transportation. Journal of Theoretical Biology, 313(??):142–152, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312003785.

Keil:2010:PTP

[KHRS10] Petr Keil, Tomás Herben, James Rosindell, and David Storch. Predictions of Taylor’s power law, density dependence and pink noise from a neutrally modeled time series. Journal of Theoretical Biology, 265(1):78–86, July 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001979.

Ke:2013:MAA

[KHS13] Ruian Ke, Ken Haynes, and Jaroslav Stark. Modelling the activation of alkaline pH response transcription factor PacC in Aspergillus nidulans: Involvement of a negative feedback loop. Journal of Theoretical Biology, 326(??):11–20, June 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931300074X.

Kang:2019:FST

[KHX^{+}19] Chuanze Kang, Yanhao Huo, Lihui Xin, Baoguang Tian, and Bin Yu. Feature selection and tumor classification for microarray data using relaxed lasso and generalized multi-class support vector machine. Journal of Theoretical Biology, 463(??):77–91, February 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306003.

Kurochkina:2015:HAS

[KI15] Natalya A. Kurochkina and Michael J. Iadarola. Helical assemblies: Structure determinants. Journal of Theoretical Biology, 369(??):80–84, March 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000211.

Kurokawa:2017:EGW

[KI17] Shun Kurokawa and Yasuo Ihara. Evolution of group-wise cooperation: Is direct reciprocity insufficient? Journal of Theoretical Biology, 415(??):20–31, February 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630412X.

Kheirabadi:2016:CKM

[KIH16] Ramesh Kheirabadi, Mohammad Izadyar, and Mohammad Reza Housiandokht. Computational kinetic modeling of the selenol catalytic activity as the glutathione peroxidase nanomimic. Journal of Theoretical Biology, 409(??): 108–114, November 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316302831.

Kusumi:2019:EGD

[KII19] Junko Kusumi, Motoshi Ichinose, and Masaru Iizuka. Effects of gene duplication, epistasis, recombination and gene conversion on the fixation time of compensatory mutations. Journal of Theoretical Biology, 467(??):134–141, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300578.

Kinjo:2018:CFT

[Kin18] Akira R. Kinjo. Cooperative “folding transition” in the sequence space facilitates function-driven evolution of protein families. Journal of Theoretical Biology, 443(??):18–27, April 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300353.

Kirschner:2011:FYJ

[Kir11] Denise Kirschner. Fifty years of JTB: Past, present and future: a letter from the Editors-in-Chief. Journal of Theoretical Biology, 268(1):iii–iv, January 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004704.

Kisdi:2010:CDC

[Kis10] Éva Kisdi. Costly dispersal can destabilize the homogeneous equilibrium of a metapopulation. Journal of Theoretical Biology, 262(2):279–283, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004706.

Kisdi:2016:DPS

[Kis16] Eva Kisdi. Dispersal polymorphism in stable habitats. Journal of Theoretical Biology, 392(??):69–82, March 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005925.

Kittas:2010:ERB

[Kit10] Aristotelis Kittas. Evolution of the rate of biological aging using a phenotype based computational model. Journal of Theoretical Biology, 266(3):401–407, October 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003619.

Kusumi:2016:MCM

[KIT^{+}16] Junko Kusumi, Motoshi Ichinose, Masasuke Takefu, Robert Piskol, Wolfgang Stephan, and Masaru Iizuka. A model of compensatory molecular evolution involving multiple sites in RNA molecules. Journal of Theoretical Biology, 388(??):96–107, January 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005019.

Kleinhans:2011:IDA

[KJ11a] David Kleinhans and Per R. Jonsson. On the impact of dispersal asymmetry on metapopulation persistence. Journal of Theoretical Biology, 290(??):37–45, December 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004589.

Kumar:2011:RMN

[KJ11b] Ajit Kumar and Kresimir Josić. Reduced models of networks of coupled enzymatic reactions. Journal of Theoretical Biology, 278(1):87–106, June 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001275.

Krivan:2015:EAD

[KJ15] Vlastimil Křivan and Debaldev Jana. Effects of animal dispersal on harvesting with protected areas. Journal of Theoretical Biology, 364(??):131–138, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005499.

Kim:2019:PVS

[KJ19] Soyoung Kim and Eunok Jung. Prioritization of vaccine strategy using an age-dependent mathematical model for 2009 A/H1N1 influenza in the Republic of Korea. Journal of Theoretical Biology, 479(??):97–105, October 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302887.

Khan:2019:PPP

[KJH^{+}19] Yaser Daanial Khan, Mehreen Jamil, Waqar Hussain, Nouman Rasool, Sher Afzal Khan, and Kuo-Chen Chou. pSSbond-PseAAC: Prediction of disulfide bonding sites by integration of PseAAC and statistical moments. Journal of Theoretical Biology, 463(??):47–55, February 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306040.

Karami:2017:CVS

[KJM17] Masoumeh Karami, Chiya Jalali, and Sako Mirzaie. Combined virtual screening, MMPBSA, molecular docking and dynamics studies against deadly anthrax: An in silico effort to inhibit bacillus anthracis nucleoside hydrolase. Journal of Theoretical Biology, 420(??):180–189, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301224.

Koschutzki:2010:SAM

[KJSS10] Dirk Koschützki, Björn H. Junker, Jörg Schwender, and Falk Schreiber. Structural analysis of metabolic networks based on flux centrality. Journal of Theoretical Biology, 265(3):261–269, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002389.

Kim:2012:HTS

[KK12a] Jong-Hoon Kim and James S. Koopman. HIV transmissions by stage in dynamic sexual partnerships. Journal of Theoretical Biology, 298(??):147–153, April 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100645X.

Krizek:2012:WNI

[KK12b] Michał Krízek and Pavel Krízek. Why has nature invented three stop codons of DNA and only one start codon? Journal of Theoretical Biology, 304(??):183–187, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001580.

Kremer:2013:CVE

[KK13] Colin T. Kremer and Christopher A. Klausmeier. Coexistence in a variable environment: Eco-evolutionary perspectives. Journal of Theoretical Biology, 339(??):14–25, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002208.

Kohn:2015:USL

[KK15] Richard A. Kohn and Seon-Woo Kim. Using the second law of thermodynamics for enrichment and isolation of microorganisms to produce fuel alcohols or hydrocarbons. Journal of Theoretical Biology, 382(??):356–362, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003562.

Karisto:2017:EDU

[KK17a] Petteri Karisto and Éva Kisdi. Evolution of dispersal under variable connectivity. Journal of Theoretical Biology, 419(??):52–65, April 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303678.

Kuznetsov:2017:UBM

[KK17b] I. A. Kuznetsov and A. V. Kuznetsov. Utilization of the bootstrap method for determining confidence intervals of parameters for a model of MAP1B protein transport in axons. Journal of Theoretical Biology, 419(??):350–361, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931730070X.

Kuznetsov:2018:TAT

[KK18] Maxim B. Kuznetsov and Andrey V. Kolobov. Transient alleviation of tumor hypoxia during first days of antiangiogenic therapy as a result of therapy-induced alterations in nutrient supply and tumor metabolism — analysis by mathematical modeling. Journal of Theoretical Biology, 451(??):86–100, August 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302182.

Kakade:2018:FSU

[KKD18] Aishwarya Kakade, Baby Kumari, and Pankaj Singh Dholaniya. Feature selection using logistic regression in case-control DNA methylation data of Parkinson’s disease: a comparative study. Journal of Theoretical Biology, 457(??):14–18, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303977.

Koke:2014:CMN

[KKG^{+}14] Christoph Koke, Takuma Kanesaki, Jörg Grosshans, Ulrich S. Schwarz, and Carina M. Dunlop. A computational model of nuclear self-organisation in syncytial embryos. Journal of Theoretical Biology, 359(??):92–100, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003385.

Khaji:2016:PSP

[KKGN16] Erfan Khaji, Masoumeh Karami, and Zahra Garkani-Nejad. 3D protein structure prediction using imperialist competitive algorithm and half sphere exposure prediction. Journal of Theoretical Biology, 391(??):81–87, February 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005779.

Kori:2012:SCN

[KKM12] Hiroshi Kori, Yoji Kawamura, and Naoki Masuda. Structure of cell networks critically determines oscillation regularity. Journal of Theoretical Biology, 297(??):61–72, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311006163.

Khalo:2018:NAV

[KKO^{+}18] Irina V. Khalo, Anastasiya I. Konokhova, Darya Y. Orlova, Konstantin V. Trusov, Maxim A. Yurkin, Eva Bartova, Stanislav Kozubek, Valeri P. Maltsev, and Andrei V. Chernyshev. Nuclear apoptotic volume decrease in individual cells: Confocal microscopy imaging and kinetic modeling. Journal of Theoretical Biology, 454(??):60–69, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302819.

Kawasumi-Kita:2018:MSO

[KKOM18] Aiko Kawasumi-Kita, Daisuke Ohtsuka, and Yoshihiro Morishita. Morphometric staging of organ development based on cross sectional images. Journal of Theoretical Biology, 440(??):80–87, March 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305593.

Kumar:2014:PAD

[KKS^{+}14] Chundi Vinay Kumar, K. M. Kumar, Rayapadi Swetha, Sudha Ramaiah, and Anand Anbarasu. Protein aggregation due to nsSNP resulting in p56s VABP protein is associated with amyotrophic lateral sclerosis. Journal of Theoretical Biology, 354(??):72–80, August 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001660.

King:2010:PMI

[KKU^{+}10] Julian King, Helin Koc, Karl Unterkofler, Paweł Mochalski, Alexander Kupferthaler, Gerald Teschl, Susanne Teschl, Hartmann Hinterhuber, and Anton Amann. Physiological modeling of isoprene dynamics in exhaled breath. Journal of Theoretical Biology, 267(4):626–637, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005023.

Kato:2010:ORM

[KKUM10] Naoto Kato, Koji Kotani, Seiya Ueno, and Hiroyuki Matsuda. Optimal risk management of human alveolar echinococcosis with vermifuge. Journal of Theoretical Biology, 267(3):265–271, December 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003942.

Kazimierski:2018:WSP

[KKWA18] Laila D. Kazimierski, Marcelo N. Kuperman, Horacio S. Wio, and Guillermo Abramson. Waves of seed propagation induced by delayed animal dispersion. Journal of Theoretical Biology, 436(??):1–7, January 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304484.

Krause:2018:SEM

[KKYV18] Andrew L. Krause, Lawrence Kurowski, Kamran Yawar, and Robert A. Van Gorder. Stochastic epidemic metapopulation models on networks: SIS dynamics and control strategies. Journal of Theoretical Biology, 449(??):35–52, July 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301917.

Komarova:2010:ELD

[KL10] Natalia L. Komarova and Simon A. Levin. Eavesdropping and language dynamics. Journal of Theoretical Biology, 264(1):104–118, May 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000147.

Kim:2011:CST

[KL11] Peter S. Kim and Peter P. Lee. T cell state transition produces an emergent change detector. Journal of Theoretical Biology, 275(1):59–69, April 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000439.

Kandler:2013:TBS

[KL13] Anne Kandler and Kevin N. Laland. Tradeoffs between the strength of conformity and number of conformists in variable environments. Journal of Theoretical Biology, 332(??):191–202, September 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001914.

Konstorum:2018:AHC

[KL18] Anna Konstorum and John S. Lowengrub. Activation of the HGF/c-Met axis in the tumor microenvironment: a multispecies model. Journal of Theoretical Biology, 439(??):86–99, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305325.

Klausmeier:2010:SSD

[Kla10] C. A. Klausmeier. Successional state dynamics: a novel approach to modeling nonequilibrium foodweb dynamics. Journal of Theoretical Biology, 262(4):584–595, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004937.

Krivan:2016:DMB

[KLB^{+}16] Vlastimil Křivan, Mark Lewis, Barbara J. Bentz, Sharon Bewick, Suzanne M. Lenhart, and Andrew Liebhold. A dynamical model for bark beetle outbreaks. Journal of Theoretical Biology, 407(??):25–37, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301928.

Klein:2014:IHT

[Kle14] Eili Y. Klein. The impact of heterogeneous transmission on the establishment and spread of antimalarial drug resistance. Journal of Theoretical Biology, 340(??):177–185, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004529.

Kulasiri:2017:GSA

[KLHS17] Don Kulasiri, Jingyi Liang, Yao He, and Sandhya Samarasinghe. Global sensitivity analysis of a model related to memory formation in synapses: Model reduction based on epistemic parameter uncertainties and related issues. Journal of Theoretical Biology, 419(??):116–136, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300590.

Kim:2017:MMT

[KLJ17] Soyoung Kim, Jonggul Lee, and Eunok Jung. Mathematical model of transmission dynamics and optimal control strategies for 2009 A/H1N1 influenza in the Republic of Korea. Journal of Theoretical Biology, 412(??):74–85, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316303228.

Kempe:2014:CTC

[KLM14] Marius Kempe, Stephen J. Lycett, and Alex Mesoudi. From cultural traditions to cumulative culture: parameterizing the differences between human and nonhuman culture. Journal of Theoretical Biology, 359(??):29–36, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400335X.

Kaiser:2012:DCF

[KLN^{+}12] J. Kaiser, T. Lemaire, S. Naili, V. Sansalone, and S. V. Komarova. Do calcium fluxes within cortical bone affect osteocyte mechanosensitivity? Journal of Theoretical Biology, 303(??):75–86, June 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931200118X.

Konrad:2012:ESG

[KM12] Kai A. Konrad and Florian Morath. Evolutionarily stable in-group favoritism and out-group spite in intergroup conflict. Journal of Theoretical Biology, 306(??):61–67, August 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312001944.

Konini:2015:MIV

[KM15] Angjelina Konini and Seyed M. Moghadas. Modelling the impact of vaccination on curtailing haemophilus influenzae serotype ‘a’. Journal of Theoretical Biology, 387(??):101– 110, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519315004750.

Keeling:2017:EFM

[KM17] Patrick J. Keeling and John P. McCutcheon. Endosymbiosis: The feeling is not mutual. Journal of Theoretical Biology, 434(??):75–79, December 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302795.

Karimnezhad:2018:AFE

[KM18] Ali Karimnezhad and Fahimeh Moradi. Allelic frequency estimation in presence of uncertain priors. Journal of Theoretical Biology, 459(??):119–129, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830465X.

Kohnke:2019:WPC

[KM19] M. C. Köhnke and H. Malchow. Wave pinning in competitiondiffusion models in variable environments. Journal of Theoretical Biology, 461(??):204–214, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305307.

Khalili:2010:DSM

[KMA10] Samira Khalili, James M. Monaco, and Antonios Armaou. Development of a stochastic model for the efficacy of NR-TIs using known mechanisms of action. Journal of Theoretical Biology, 265(4):704–717, August 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002353.

Kim:2014:GDE

[KMCH14] Peter S. Kim, John S. McQueen, James E. Coxworth, and Kristen Hawkes. Grandmothering drives the evolution of longevity in a probabilistic model. Journal of Theoretical Biology, 353(??):84–94, July 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001465.

Kim:2017:AVB

[KMCJ17] Sungchan Kim, M. A. Masud, Giphil Cho, and Il Hyo Jung. Analysis of a vector-bias effect in the spread of malaria between two different incidence areas. Journal of Theoretical Biology, 419(??):66–76, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300589.

Karp:2012:CLI

[KMD^{+}12] Robert L. Karp, Mercedes Pérez Millán, Tathagata Dasgupta, Alicia Dickenstein, and Jeremy Gunawardena. Complexlinear invariants of biochemical networks. Journal of Theoretical Biology, 311(??):130–138, October 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003384.

Kim:2019:WDW

[KMH19] Peter S. Kim, John S. McQueen, and Kristen Hawkes. Why does women’s fertility end in mid-life? Grandmothering and age at last birth. Journal of Theoretical Biology, 461(??):84–91, January 14, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305113.

Kebir:2010:EDD

[KMHdlP10] A. Kebir, S. Ben Miled, M. L. Hbid, and R. Bravo de la Parra. Effects of density dependent sex allocation on the dynamics of a simultaneous hermaphroditic population: Modelling and analysis. Journal of Theoretical Biology, 263(4):521–529, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005803.

Kumari:2014:CMM

[KMLT14] Sweta Kumari, Arumugam MohanaPriya, Sajitha Lulu, and Mohammad Tauqueer. Corrigendum to “Molecular modeling, simulation and virtual screening of ribosomal phosphoprotein P1 from Plasmodium falciparum” [J. Theor. Biol. 343 (2014) 113–119]. Journal of Theoretical Biology, 354(??):146, August 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002100.

Kiflawi:2016:HPC

[KMM16] Moshe Kiflawi, Ofri Mann, and Mark G. Meekan. Heterogeneous ‘proportionality constants’ — a challenge to Taylor’s Power Law for temporal fluctuations in abundance. Journal of Theoretical Biology, 407(??):155–160, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301977.

Korayem:2018:IDB

[KMM18] M. H. Korayem, Z. Mahmoodi, and M. Mohammadi. 3D investigation of dynamic behavior and sensitivity analysis of the parameters of spherical biological particles in the first phase of AFM-based manipulations with the consideration of humidity effect. Journal of Theoretical Biology, 436(??):105–119, January 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304344.

Kharratzadeh:2017:SHL

[KMMS17] Milad Kharratzadeh, Marcel Montrey, Alex Metz, and Thomas R. Shultz. Specialized hybrid learners resolve Rogers’ paradox about the adaptive value of social learning. Journal of Theoretical Biology, 414(??):8–16, February 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303812.

Kakizoe:2015:CLV

[KMN^{+}15] Yusuke Kakizoe, Satoru Morita, Shinji Nakaoka, Yasuhiro Takeuchi, Kei Sato, Tomoyuki Miura, Catherine A. A. Beauchemin, and Shingo Iwami. A conservation law for virus infection kinetics in vitro. Journal of Theoretical Biology, 376(??):39–47, July 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001496.

Klanjscek:2016:FTP

[KMN16] Tin Klanjscek, Erik B. Muller, and Roger M. Nisbet. Feedbacks and tipping points in organismal response to oxidative stress. Journal of Theoretical Biology, 404(??):361–374, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301242.

Karimi:2018:MMD

[KMZR18] A. H. Karimi, H. R. Mirdamadi, and S. Ziaei-Rad. Mathematical modeling of dynamic behavior of fluid bilayer membranes under the effect of density asymmetry. Journal of Theoretical Biology, 454(??):330–344, October 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303205.

Klinkenberg:2011:CBI

[KN11] Don Klinkenberg and Hiroshi Nishiura. The correlation between infectivity and incubation period of measles, estimated from households with two cases. Journal of Theoretical Biology, 284(1):52–60, September 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003146.

Kitagawa:2018:PMM

[KNA^{+}18] Kosaku Kitagawa, Shinji Nakaoka, Yusuke Asai, Koichi Watashi, and Shingo Iwami. A PDE multiscale model of hepatitis C virus infection can be transformed to a system of ODEs. Journal of Theoretical Biology, 448(??):80–85, July 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301668.

Koike:2010:ECR

[KNT10] Shimpei Koike, Mayuko Nakamaru, and Masahiro Tsujimoto. Evolution of cooperation in rotating indivisible goods game. Journal of Theoretical Biology, 264(1):143–153, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000020.

Kay:2018:ERM

[KO18] Taryn M. Kay and Paul R. Ohmann. Effects of random motion in traveling and grazing herds. Journal of Theoretical Biology, 456(??):168–174, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303850.

Koenig:2015:PPP

[Koe15] Michel Koenig. Phosphoribosylphosphate and phosphoribosylnicotinate pairing with phosphoribosylamine at the origin of the RNA world. Journal of Theoretical Biology, 379(??):94–97, August 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002416.

Kawamura:2014:KMP

[KOF^{+}14] Koji Kawamura, Laurence Hibrand-Saint Oyant, Fabrice Foucher, Tatiana Thouroude, and Sébastien Loustau. Kernel methods for phenotyping complex plant architecture. Journal of Theoretical Biology, 342(??):83–92, February 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005158.

Konkoli:2011:MRN

[Kon11a] Zoran Konkoli. Multiparticle reaction noise characteristics. Journal of Theoretical Biology, 271(1):78–86, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006247.

Konkoli:2011:SNR

[Kon11b] Zoran Konkoli. Spontaneous noise reduction in a strongly cooperative reaction model. Journal of Theoretical Biology, 285(1):96–102, September 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003419.

Konno:2011:CCG

[Kon11c] Tomohiko Konno. A condition for cooperation in a game on complex networks. Journal of Theoretical Biology, 269(1):224–233, January 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005783.

Konkoli:2012:MRN

[Kon12] Zoran Konkoli. Modeling reaction noise with a desired accuracy by using the X level approach reaction noise estimator (XARNES) method. Journal of Theoretical Biology, 305(??):1–14, July 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001786.

Kondo:2017:UKB

[Kon17] Shigeru Kondo. An updated kernel-based Turing model for studying the mechanisms of biological pattern formation. Journal of Theoretical Biology, 414(??):120–127, February 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316303630.

Korobeinikov:2018:IRW

[Kor18] Andrei Korobeinikov. Immune response and within-host viral evolution: Immune response can accelerate evolution. Journal of Theoretical Biology, 456(??):74–83, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303734.

Kashima:2013:FFB

[KOS13] Kazutaka Kashima, Hisashi Ohtsuki, and Akiko Satake. Fission-fusion bat behavior as a strategy for balancing the conflicting needs of maximizing information accuracy and minimizing infection risk. Journal of Theoretical Biology, 318(??):101–109, February 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005747.

Karsai:2012:RTD

[KP12] Istvan Karsai and Michael D. Phillips. Regulation of task differentiation in wasp societies: a bottom-up model of the “common stomach”. Journal of Theoretical Biology, 294(??):98–113, February 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005601.

Krivan:2015:SPI

[KP15] Vlastimil Křivan and Anupam Priyadarshi. L-shaped prey isocline in the Gause predator–prey experiments with a prey refuge. Journal of Theoretical Biology, 370(??):21–26, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000302.

Kozlowska:2016:ABT

[KP16] Emilia Kozlowska and Krzysztof Puszynski. Application of bifurcation theory and siRNA-based control signal to restore the proper response of cancer cells to DNA damage. Journal of Theoretical Biology, 408(??):213–221, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931630251X.

Kenny:2018:MSM

[KPD18] Allanah Kenny, Michael J. Plank, and Tim David. Macro scale modelling of cortical spreading depression and the role of astrocytic gap junctions. Journal of Theoretical Biology, 458(??):78–91, December 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304363.

Kenny:2019:ECC

[KPD19] Allanah Kenny, Michael J. Plank, and Tim David. The effects of cerebral curvature on cortical spreading depression. Journal of Theoretical Biology, 472(??):11–26, July 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930147X.

Knutsdottir:2014:MMM

[KPEK14] Hildur Knútsdóttir, Eirikur Pálsson, and Leah Edelstein-Keshet. Mathematical model of macrophage-facilitated breast cancer cells invasion. Journal of Theoretical Biology, 357(??):184–199, September 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002616.

Kar:2019:MYR

[KPG19] Tapan Kumar Kar, Debprasad Pal, and Bapan Ghosh. Managing yield and resilience in a harvested tri-trophic food chain model. Journal of Theoretical Biology, 469(??):35–46, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930061X.

Kandaswamy:2013:EPE

[KPK^{+}13] Krishna Kumar Kandaswamy, Ganesan Pugalenthi, Kai-Uwe Kalies, Enno Hartmann, and Thomas Martinetz. EcmPred: Prediction of extracellular matrix proteins based on random forest with maximum relevance minimum redundancy feature selection. Journal of Theoretical Biology, 317(??):377–383, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005486.

Kumari:2014:MMS

[KPLT14] Sweta Kumari, Arumugam Mohana Priya, Sajitha Lulu, and Mohammad Tauqueer. Molecular modeling, simulation and virtual screening of ribosomal phosphoprotein P1 from Plasmodium falciparum. Journal of Theoretical Biology, 343(??):113–119, February 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519313005122. See corrigendum [NO14b].

Kovalenko:2017:SDM

[KPS17] Tatiana A. Kovalenko, Mikhail A. Panteleev, and Anastasia N. Sveshnikova. Substrate delivery mechanism and the role of membrane curvature in factor X activation by extrinsic tenase. Journal of Theoretical Biology, 435(??):125–133, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304320.

Kim:2015:TDO

[KR15] Jong-Hoon Kim and Seong-Hwan Rho. Transmission dynamics of oral polio vaccine viruses and vaccine-derived polioviruses on networks. Journal of Theoretical Biology, 364(??):266–274, January 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005657.

Krivan:2019:PCM

[KR19] Vlastimil Krivan and Tomás A. Revilla. Plant coexistence mediated by adaptive foraging preferences of exploiters or mutualists. Journal of Theoretical Biology, 480(??):112–128, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303091.

Khalili:2015:SAW

[KRDJ15] Saeed Khalili, Mohammad Reza Rahbar, Mohammad Haj Dezfulian, and Abolfazl Jahangiri. In silico analyses of Wilms’ tumor protein to designing a novel multi-epitope DNA vaccine against cancer. Journal of Theoretical Biology, 379(??):66–78, August 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002027.

Krivan:2010:ESO

[Kři10] Vlastimil Křivan. Evolutionary stability of optimal foraging: Partial preferences in the diet and patch models. Journal of Theoretical Biology, 267(4):486–494, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004881.

Krivan:2011:GPP

[Kři11] Vlastimil Křivan. On the Gause predator–prey model with a refuge: a fresh look at the history. Journal of Theoretical Biology, 274(1):67–73, April 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000282.

Krivan:2013:BRP

[Kři13] Vlastimil Křivan. Behavioral refuges and predator–prey coexistence. Journal of Theoretical Biology, 339(??):112–121, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312006522.

Krivan:2014:CTT

[Kři14] Vlastimil Křivan. Competition in di- and tri-trophic food web modules. Journal of Theoretical Biology, 343(??):127–137, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931300547X.

Krishnan:2016:CVH

[Kri16] S. Muthu Krishnan. Classify vertebrate hemoglobin proteins by incorporating the evolutionary information into the general PseAAC with the hybrid approach. Journal of Theoretical Biology, 409(??):27–37, November 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302612.

Krishnan:2018:UCG

[Kri18] S. Muthu Krishnan. Using Chou’s general PseAAC to analyze the evolutionary relationship of receptor associated proteins (RAP) with various folding patterns of protein domains. Journal of Theoretical Biology, 445(??):62–74, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300602.

Kroon:2010:CBC

[Kro10] Martin Kroon. On the correlation between continuum mechanics entities and cell activity in biological soft tissues: Assessment of three possible criteria for cell-controlled fibre reorientation in collagen gels and collagenous tissues. Journal of Theoretical Biology, 264(1):66–76, May 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251930900602X.

Kroon:2011:IDM

[Kro11] Martin Kroon. Influence of dispersion in myosin filament orientation and anisotropic filament contractions in smooth muscle. Journal of Theoretical Biology, 272(1):72–82, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006326.

Kim:2014:FTT

[KRR14] MunJu Kim, Damon Reed, and Katarzyna A. Rejniak. The formation of tight tumor clusters affects the efficacy of cell cycle inhibitors: a hybrid model study. Journal of Theoretical Biology, 352(??):31–50, July 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001027.

Kucken:2011:NRP

[KRSD11] Michael Kücken, Baruch Rinkevich, Lee Shaish, and Andreas Deutsch. Nutritional resources as positional information for morphogenesis in the stony coral Stylophora pistillata. Journal of Theoretical Biology, 275(1):70–77, April 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000300.

Kushwaha:2010:PIN

[KS10] Sandeep K. Kushwaha and Madhvi Shakya. Protein interaction network analysis — approach for potential drug target identification in Mycobacterium tuberculosis. Journal of Theoretical Biology, 262(2):284–294, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004652.

Kandler:2013:NEN

[KS13] Anne Kandler and Stephen Shennan. A non-equilibrium neutral model for analysing cultural change. Journal of Theoretical Biology, 330(??):18–25, August 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000982.

Kessler:2014:NLA

[KS14] David A. Kessler and Nadav M. Shnerb. Neutral-like abundance distributions in the presence of selection in a continuous fitness landscape. Journal of Theoretical Biology, 345(??):1–11, March 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005328.

Koch:2015:DTC

[KS15] Gilbert Koch and Johannes Schropp. Distributed transit compartments for arbitrary lifespan distributions in aging populations. Journal of Theoretical Biology, 380(??):550–558, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002970.

Koenders:2016:MMC

[KS16] M. A. Koenders and R. Saso. A mathematical model of cell equilibrium and joint cell formation in multiple myeloma. Journal of Theoretical Biology, 390(??):73–79, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315005639.

Kaufman:2019:MCR

[KS19] Marcelle Kaufman and Christophe Soulé. On the multistationarity of chemical reaction networks. Journal of Theoretical Biology, 465(??):126–133, March 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300098.

Karagoz:2016:IMB

[KSA16] Kubra Karagoz, Tuba Sevimoğlu, and Kazim Yalcin Arga. Integration of multiple biological features yields high confidence human protein interactome. Journal of Theoretical Biology, 403(??):85–96, August 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301047.

Kemp:2012:QFA

[KSBvL^{+}12] Michael U. Kemp, Judy Shamoun-Baranes, E. Emiel van Loon, James D. McLaren, Adriaan M. Dokter, and Willem Bouten. Quantifying flow-assistance and implications for movement research. Journal of Theoretical Biology, 308(??):56–67, September 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002688.

Kang:2011:RAF

[KSK^{+}11] John Kang, Robert L. Steward, YongTae Kim, Russell S. Schwartz, Philip R. LeDuc, and Kathleen M. Puskar. Response of an actin filament network model under cyclic stretching through a coarse grained Monte Carlo approach. Journal of Theoretical Biology, 274(1):109–119, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000233.

Kobayashi:2016:MMC

[KSK^{+}16] Yasuaki Kobayashi, Yusuke Sawabu, Hiroyuki Kitahata, Mitsuhiro Denda, and Masaharu Nagayama. Mathematical model for calcium-assisted epidermal homeostasis. Journal of Theoretical Biology, 397(??):52–60, May 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001284.

Kumar:2017:MDP

[KSK^{+}17] P. Narendra Kumar, T. H. Swapna, Mohamed Yahya Khan, Jayasimha Rayalu Daddam, and Bee Hameeda. Molecular dynamics and protein interaction studies of lipopeptide (Iturin A) on α- amylase of Spodoptera litura. Journal of Theoretical Biology, 415(??):41–47, February 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304131.

Kumar:2015:PLC

[KSKK15] Ravindra Kumar, Abhishikha Srivastava, Bandana Kumari, and Manish Kumar. Prediction of β-lactamase and its class by Chou’s pseudo-amino acid composition and support vector machine. Journal of Theoretical Biology, 365(??):96–103, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005979.

Koyano:2013:FPA

[KSKS13] Hitoshi Koyano, Dimitar Serbezov, Hirohisa Kishino, and Tore Schweder. Fractional parentage analysis and a scalefree reproductive network of brown trout. Journal of Theoretical Biology, 336(??):18–35, November 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003214.

Kupczik:2015:RMF

[KSM^{+}15] Kornelius Kupczik, Heiko Stark, Roger Mundry, Fabian T. Neininger, Thomas Heidlauf, and Oliver Röhrle. Reconstruction of muscle fascicle architecture from iodine-enhanced microCT images: a combined texture mapping and streamline approach. Journal of Theoretical Biology, 382(??):34–43, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003136.

Kerimoglu:2012:RPC

[KSP12] Onur Kerimoğlu, Dietmar Straile, and Frank Peeters. Role of phytoplankton cell size on the competition for nutrients and light in incompletely mixed systems. Journal of Theoretical Biology, 300(??):330–343, May 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000641.

Kawka:2014:RRS

[KSP^{+}14] Joanna Kawka, Dominik Sturm, Sabrina Pleier, Stefan Pfister, and Anna Marciniak-Czochra. Revealing the role of SGK1 in the dynamics of medulloblastoma using a mathematical model. Journal of Theoretical Biology, 354(??):105–112, August 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001672.

Khan:2018:ARF

[KSP18] Abhinandan Khan, Goutam Saha, and Rajat Kumar Pal. An approach for reduction of false predictions in reverse engineering of gene regulatory networks. Journal of Theoretical Biology, 445(??):9–30, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300754.

Kuritz:2017:RBC

[KSPA17] K. Kuritz, D. Stöhr, N. Pollak, and F. Allgöwer. On the relationship between cell cycle analysis with ergodic principles and age-structured cell population models. Journal of Theoretical Biology, 414(??):91–102, February 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304040.

Konrad:2019:MEF

[KSRN19] Wilfried Konrad, Rena Schott, and Anita Roth-Nebelsick. A model for extracellular freezing based on observations on Equisetum hyemale. Journal of Theoretical Biology, 478(??):161–168, October 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302589.

Katsnelson:2011:RMV

[KSSM11] Leonid B. Katsnelson, Tatiana Sulman, Olga Solovyova, and Vladimir S. Markhasin. Role of myocardial viscoelasticity in disturbances of electrical and mechanical activity in calcium overloaded cardiomyocytes: Mathematical modeling. Journal of Theoretical Biology, 272(1):83–95, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006302.

Kumar:2017:MRI

[KST17] Anuj Kumar, Prashant K. Srivastava, and Yasuhiro Takeuchi. Modeling the role of information and limited optimal treatment on disease prevalence. Journal of Theoretical Biology, 414(??):103–119, February 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303824.

Kadota:2018:MMT

[KSU^{+}18] Rei Kadota, Kazuma Sugita, Kenko Uchida, Hitoshi Yamada, Masashi Yamashita, and Hidenori Kimura. A mathematical model of type 1 diabetes involving leptin effects on glucose metabolism. Journal of Theoretical Biology, 456(??):213–223, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303813.

Khan:2018:ORT

[KSvdH18] Mohammad Farhan Khan, Sarah Spurgeon, and Tobias von der Haar. Origins of robustness in translational control via eukaryotic translation initiation factor (eIF) 2. Journal of Theoretical Biology, 445(??):92–102, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300808.

Kingsbury:2016:IEN

[KTH16] Nathaniel J. Kingsbury, Stephanie R. Taylor, and Michael A. Henson. Inhibitory and excitatory networks balance cell coupling in the suprachiasmatic nucleus: a modeling approach. Journal of Theoretical Biology, 397(??):135–144, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316001454.

Kameyama:2018:SSR

[KTI18] Masashi Kameyama, Sousuke Takeuchi, and Shinya Ishii. Steady-state relationship between average glucose, HbA1c and RBC lifespan. Journal of Theoretical Biology, 447(??):111– 117, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301371.

Kuga:2019:VVC

[KTJ19] Kazuki Kuga, Jun Tanimoto, and Marko Jusup. To vaccinate or not to vaccinate: a comprehensive study of vaccinationsubsidizing policies with multi-agent simulations and meanfield modeling. Journal of Theoretical Biology, 469(??):107– 126, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300827.

Kobayashi:2010:SMN

[KTN^{+}10] Taichiro Kobayashi, Kenshi Terajima, Motohiro Nozumi, Michihiro Igarashi, and Kouhei Akazawa. A stochastic model of neuronal growth cone guidance regulated by multiple sensors. Journal of Theoretical Biology, 266(4):712–722, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003930.

Kasimova:2014:AMO

[KTO^{+}14] R. G. Kasimova, D. Tishin, Yu.V. V. Obnosov, G. M. Dlussky, F. B. Baksht, and A. R. Kacimov. Ant mound as an optimal shape in constructal design: Solar irradiation and circadian brood/fungi-warming sorties. Journal of Theoretical Biology, 355(??):21–32, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000575.

Kumar:2019:CFF

[KTT^{+}19] Rakesh Kumar, Abhishek Kumar Tiwari, Dharmendra Tripathi, Nikhil Vivek Shrivas, and Fahad Nizam. Canalicular fluid flow induced by loading waveforms: a comparative analysis. Journal of Theoretical Biology, 471(??):59–73, June 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301304.

Kuetche:2016:IBD

[Kue16] Victor K. Kuetche. Ab initio bubble-driven denaturation of double-stranded DNA: Self-mechanical theory. Journal of Theoretical Biology, 401(??):15–29, July 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300546.

Kundu:2016:SMS

[Kun16] Siddhartha Kundu. Stochastic modelling suggests that an elevated superoxide anion–hydrogen peroxide ratio can drive extravascular phagocyte transmigration by lamellipodium formation. Journal of Theoretical Biology, 407(??):143–154, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301734.

Kurochkina:2010:HHI

[Kur10] Natalya Kurochkina. Helix-helix interactions and their impact on protein motifs and assemblies. Journal of Theoretical Biology, 264(2):585–592, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001013.

Kurbel:2014:HHE

[Kur14] Sven Kurbel. Hypothesis of homeothermy evolution on isolated South China Craton that moved from equator to cold north latitudes 250–200 Myr ago. Journal of Theoretical Biology, 340(??):232–237, January 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004463.

Kurokawa:2017:ERS

[Kur17] Shun Kurokawa. The extended reciprocity: Strong belief outperforms persistence. Journal of Theoretical Biology, 421 (??):16–27, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519317301418.

Kurokawa:2019:TPR

[Kur19] Shun Kurokawa. Three-player repeated games with an optout option. Journal of Theoretical Biology, 480(??):13–22, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302899.

Koseska:2010:CDT

[KUV^{+}10] A. Koseska, E. Ullner, E. Volkov, J. Kurths, and J. García-Ojalvo. Cooperative differentiation through clustering in multicellular populations. Journal of Theoretical Biology, 263(2):189–202, March 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005335.

Kuzma:2019:EIC

[Kuz19] Strelnikov Kuzma. Energy-information coupling during integrative cognitive processes. Journal of Theoretical Biology, 469(??):180–186, May 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300980.

Kaul:2015:DRR

[KV15] Himanshu Kaul and Yiannis Ventikos. Dynamic reciprocity revisited. Journal of Theoretical Biology, 370(??):205–208, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315000259.

Kaveh:2016:GM

[KVN16] Kamran Kaveh, Carl Veller, and Martin A. Nowak. Games of multicellularity. Journal of Theoretical Biology, 403(??): 143–158, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316300716.

Komarova:2010:OMO

[KW10] Natalia L. Komarova and Dominik Wodarz. ODE models for oncolytic virus dynamics. Journal of Theoretical Biology, 263(4):530–543, April 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000135.

Koykka:2015:EGD

[KW15] Cody Koykka and Geoff Wild. The evolution of group dispersal with leaders and followers. Journal of Theoretical Biology, 371(??):117–126, April 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000673.

Kobayashi:2015:PCC

[KWO15] Yutaka Kobayashi, Joe Yuichiro Wakano, and Hisashi Ohtsuki. A paradox of cumulative culture. Journal of Theoretical Biology, 379(??):79–88, August 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002301.

Kobayashi:2019:ECC

[KWO19] Yutaka Kobayashi, Joe Yuichiro Wakano, and Hisashi Ohtsuki. Evolution of cumulative culture for niche construction. Journal of Theoretical Biology, 472(??):67–76, July 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301547.

Kribs-Zaleta:2014:GAE

[KZ14] Christopher M. Kribs-Zaleta. Graphical analysis of evolutionary trade-off in sylvatic Trypanosoma cruzi transmission modes. Journal of Theoretical Biology, 353(??):34–43, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314001374.

Kong:2014:APP

[KZL14] Liang Kong, Lichao Zhang, and Jinfeng Lv. Accurate prediction of protein structural classes by incorporating predicted secondary structure information into the general form of Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 344(??):12–18, March 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005481.

Lashari:2012:FMR

[LA12] Abid Ali Lashari and Faiz Ahmad. False mathematical reasoning in biology. Journal of Theoretical Biology, 307(??):211, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002317.

Lin:2015:MPE

[LA15] Yuan Lin and Nicole Abaid. Modeling perspectives on echolocation strategies inspired by bats flying in groups. Journal of Theoretical Biology, 387(??):46–53, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004427.

Labowsky:2016:AES

[Lab16] M. Labowsky. An analysis of early-stage IL-2 capture times in populations of T cells diffusively interacting in a confined environment. Journal of Theoretical Biology, 411(??):37–47, December 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302983.

Leander:2014:DEC

[LAG^{+}14] R. Leander, E. J. Allen, S. P. Garbett, D. R. Tyson, and V. Quaranta. Derivation and experimental comparison of cell-division probability densities. Journal of Theoretical Biology, 359(??):129–135, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003415.

Leber:2016:BAC

[LAH^{+}16] Andrew Leber, Vida Abedi, Raquel Hontecillas, Monica Viladomiu, Stefan Hoops, Stanca Ciupe, John Caughman, Tricity Andrew, and Josep Bassaganya-Riera. Bistability analyses of CD4+ T follicular helper and regulatory cells during helicobacter pylori infection. Journal of Theoretical Biology, 398(??):74–84, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001387.

Laird:2011:GBE

[Lai11] Robert A. Laird. Green-beard effect predicts the evolution of traitorousness in the two-tag Prisoner’s dilemma. Journal of Theoretical Biology, 288(??):84–91, November 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003730.

Laird:2018:SIW

[Lai18] Robert A. Laird. Sequential interactions-in which one player plays first and another responds-promote cooperation in evolutionary-dynamical simulations of single-shot Prisoner’s dilemma and snowdrift games. Journal of Theoretical Biology, 452(??):69–80, September 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302339.

Lopez-Alamilla:2019:ASM

[LAJC19] N. J. López-Alamilla, M. W. Jack, and K. J. Challis. Analysing single-molecule trajectories to reconstruct freeenergy landscapes of cyclic motor proteins. Journal of Theoretical Biology, 462(??):321–328, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305691.

Lange:2016:RDT

[Lan16] Alexander Lange. Reconstruction of disease transmission rates: Applications to measles, dengue, and influenza. Journal of Theoretical Biology, 400(??):138–153, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300455.

Lane:2017:SES

[Lan17] Nick Lane. Serial endosymbiosis or singular event at the origin of eukaryotes? Journal of Theoretical Biology, 434(??):58–67, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302011.

Lika:2014:BDP

[LAPK14] Konstadia Lika, Starrlight Augustine, Laure Pecquerie, and Sebastiaan A. L. M. Kooijman. The bijection from data to parameter space with the standard DEB model quantifies the supply-demand spectrum. Journal of Theoretical Biology, 354(??):35–47, August 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001647.

Larjavaara:2010:MCT

[Lar10] Markku Larjavaara. Maintenance cost, toppling risk and size of trees in a self-thinning stand. Journal of Theoretical Biology, 265(1):63–67, July 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002043.

Larson:2016:EAL

[Lar16] Jennifer M. Larson. The evolutionary advantage of limited network knowledge. Journal of Theoretical Biology, 398 (??):43–51, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316001582.

Lambert:2014:PAA

[LAS14] Amaury Lambert, Helen K. Alexander, and Tanja Stadler. Phylogenetic analysis accounting for age-dependent death and sampling with applications to epidemics. Journal of Theoretical Biology, 352(??):60–70, July 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001064.

Lazar:2013:SSS

[Laz13] Dusan Lazár. Simulations show that a small part of variable chlorophyll a fluorescence originates in photosystem I and contributes to overall fluorescence rise. Journal of Theoretical Biology, 335(??):249–264, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003007.

Lutkenhoner:2011:AMV

[LB11] Bernd Lütkenhöner and Türker Basel. An analytical model of the vestibular evoked myogenic potential. Journal of Theoretical Biology, 286(??):41–49, October 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003468.

Lutkenhoner:2012:DVE

[LB12] Bernd Lütkenhöner and Türker Basel. Deconvolution of the vestibular evoked myogenic potential. Journal of Theoretical Biology, 294(??):87–97, February 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100556X.

Lawn:2013:IBE

[LBB^{+}13] Brian R. Lawn, Mark B. Bush, Amir Barani, Paul J. Constantino, and Stephen Wroe. Inferring biological evolution from fracture patterns in teeth. Journal of Theoretical Biology, 338(??):59–65, December 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004037.

Liao:2013:RCC

[LBF13] Kang-Ling Liao, Xue-Feng Bai, and Avner Friedman. The role of CD200-CD200R in tumor immune evasion. Journal of Theoretical Biology, 328(??):65–76, July 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001215.

Lee:2013:MAA

[LBGW13] S. Seirin Lee, R. E. Baker, E. A. Gaffney, and S. M. White. Modelling aedes aegypti mosquito control via transgenic and sterile insect techniques: Endemics and emerging outbreaks. Journal of Theoretical Biology, 331(??):78–90, August 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313001823.

Llaurens:2013:EDP

[LBJ13] V. Llaurens, S. Billiard, and M. Joron. The effect of dominance on polymorphism in Müllerian mimicry. Journal of Theoretical Biology, 337(??):101–110, November 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003780.

Lalande:2016:MMS

[LBMG16] Laure Lalande, Laurent Bourguignon, Pascal Maire, and Sylvain Goutelle. Mathematical modeling and systems pharmacology of tuberculosis: Isoniazid as a case study. Journal of Theoretical Biology, 399(??):43–52, June 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300145.

Le:2010:NIA

[LBS^{+}10] Tuyen Quang Le, Doyoung Byun, Saputra, Jin Hwan Ko, Hoon Choel Park, and Minjun Kim. Numerical investigation of the aerodynamic characteristics of a hovering coleopteran insect. Journal of Theoretical Biology, 266(4):485–495, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003620. See corri-

gendum [LBS^{+}11].

Le:2011:CNI

[LBS^{+}11] Tuyen Quang Le, Doyoung Byun, Saputra, Jin Hwan Ko, Hoon Choel Park, and Minjun Kim. Corrigendum to “Numerical investigation of the aerodynamic characteristics of a hovering Coleopteran insect” [J. Theor. Biol. 266 (2010) 485– 495]. Journal of Theoretical Biology, 270(1):196, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006260. See [LBS^{+}10].

Lyons:2014:PFR

[LBS^{+}14] James Lyons, Neela Biswas, Alok Sharma, Abdollah Dehzangi, and Kuldip K. Paliwal. Protein fold recognition by alignment of amino acid residues using kernelized dynamic time warping. Journal of Theoretical Biology, 354(??):137–145, August 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001805.

Lignet:2013:TIE

[LBW^{+}13] F. Lignet, S. Benzekry, S. Wilson, F. Billy, O. Saut, M. Tod, B. You, A. Adda Berkane, S. Kassour, M. X. Wei, E. Grenier, and B. Ribba. Theoretical investigation of the efficacy of antiangiogenic drugs combined to chemotherapy in xenografted mice. Journal of Theoretical Biology, 320(??):86–99, March 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312006418.

Liu:2018:PTP

[LBZ18] Junli Liu, Zhenguo Bai, and Tailei Zhang. A periodic twopatch SIS model with time delay and transport-related infection. Journal of Theoretical Biology, 437(??):36–44, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304745.

Lorthois:2010:FAV

[LC10] Sylvie Lorthois and Francis Cassot. Fractal analysis of vascular networks: Insights from morphogenesis. Journal of Theoretical Biology, 262(4):614–633, February 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005268.

Li:2011:MNH

[LC11] Yongfeng Li and Francis A. Cucinotta. Modeling nonhomologous end joining. Journal of Theoretical Biology, 283(1):122–135, August 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002578.

Lou:2016:SMH

[LC16] Yuting Lou and Yu Chen. Simulating the multicellular homeostasis with a cell-based discrete receptor dynamics model: The non-mutational origin of cancer and aging. Journal of Theoretical Biology, 404(??):15–29, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300698.

Lee:2017:EOC

[LC17] Sunmi Lee and Gerardo Chowell. Exploring optimal control strategies in seasonally varying flu-like epidemics. Journal of Theoretical Biology, 412(??):36–47, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303204.

Lee:2015:QSL

[LCA^{+}15] Chung-Hao Lee, Christopher A. Carruthers, Salma Ayoub, Robert C. Gorman, Joseph H. Gorman, and Michael S. Sacks. Quantification and simulation of layer-specific mitral valve interstitial cells deformation under physiological loading. Journal of Theoretical Biology, 373(??):26–39, May 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315001071.

Lorente:2012:CLP

[LCBO^{+}12] S. Lorente, E. Cetkin, T. Bello-Ochende, J. P. Meyer, and A. Bejan. The constructal-law physics of why swimmers must spread their fingers and toes. Journal of Theoretical Biology, 308(??):141–146, September 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002755.

Lee:2015:RRT

[LCC15] Sunmi Lee and Carlos Castillo-Chavez. The role of residence times in two-patch dengue transmission dynamics and optimal strategies. Journal of Theoretical Biology, 374(??):152–164, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001083.

Letelier:2011:LMM

[LCCB11] Juan-Carlos Letelier, María Luz Cárdenas, and Athel Cornish-Bowden. From L’Homme Machine to metabolic closure: Steps towards understanding life. Journal of Theoretical Biology, 286(??):100–113, October 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311003389.

Lee:2010:OCP

[LCCC10] Sunmi Lee, Gerardo Chowell, and Carlos Castillo-Chávez. Optimal control for pandemic influenza: The role of limited antiviral treatment and isolation. Journal of Theoretical Biology, 265(2):136–150, July 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001864.

Lee:2012:EPH

[LCCK12] Jung Min Lee, Donghoon Choi, Giphil Cho, and Yongkuk Kim. The effect of public health interventions on the spread of influenza among cities. Journal of Theoretical Biology, 293(??):131–142, January 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005157.

Lorenzi:2015:DDE

[LCDH15] Tommaso Lorenzi, Rebecca H. Chisholm, Laurent Desvillettes, and Barry D. Hughes. Dissecting the dynamics of epigenetic changes in phenotype-structured populations exposed to fluctuating environments. Journal of Theoretical Biology, 386(??):166–176, December 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500435X.

Laurie:2017:PL

[LCF17] Jason Laurie, Amit K. Chattopadhyay, and Darren R. Flower. Protein lipograms. Journal of Theoretical Biology, 430(??):109–116, October 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303405.

Lim:2012:HSM

[LCG12] Valery I. Lim, James F. Curran, and Maria B. Garber. Hydration shells of molecules in molecular association: a mechanism for biomolecular recognition. Journal of Theoretical Biology, 301(??):42–48, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000811.

Li:2015:SPE

[LCG^{+}15] Duo-Fang Li, Tian-Guang Cao, Jin-Peng Geng, Jian-Zhong Gu, Hai-Long An, and Yong Zhan. Statistical properties and error threshold of quasispecies on single-peak Gaussian-distributed fitness landscapes. Journal of Theoretical Biology, 380(??):53–59, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002428.

Lopez-Caamal:2012:PFA

[LCGMH12] Fernando López-Caamal, Míriam R. García, Richard H. Middleton, and Heinrich J. Huber. Positive feedback in the Akt/ mTOR pathway and its implications for growth signal progression in skeletal muscle cells: An analytical study. Journal of Theoretical Biology, 301(??):15–27, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000367.

Lythe:2016:HMT

[LCHMP16] Grant Lythe, Robin E. Callard, Rollo L. Hoare, and Carmen Molina-París. How many TCR clonotypes does a body maintain? Journal of Theoretical Biology, 389(??):214–224, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005159.

Lee:2016:DCM

[LCJ16] Jonggul Lee, Gerardo Chowell, and Eunok Jung. A dynamic compartmental model for the middle East respiratory syndrome outbreak in the Republic of Korea: a retrospective analysis on control interventions and superspreading events. Journal of Theoretical Biology, 408(??):118–126, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302430.

Lee:2018:MMU

[LCJ18] Jonggul Lee, Bo Youl Choi, and Eunok Jung. Metapopulation model using commuting flow for national spread of the 2009 H1N1 influenza virus in the Republic of Korea. Journal of Theoretical Biology, 454(??):320–329, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303114.

Liu:2014:PCP

[LCL14] Qian Liu, Yi-Ping Phoebe Chen, and Jinyan Li. k-partite cliques of protein interactions: a novel subgraph topology for functional coherence analysis on PPI networks. Journal of Theoretical Biology, 340(??):146–154, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004414.

Larsen:2014:MFE

[LCMC14] Peter E. Larsen, Leland J. Cseke, R. Michael Miller, and Frank R. Collart. Modeling forest ecosystem responses to elevated carbon dioxide and ozone using artificial neural networks. Journal of Theoretical Biology, 359(??):61–71, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003361.

Liu:2018:EAS

[LCQ^{+}18] Yong Liu, Jingan Chen, Jieying Qian, Hao Lin, Ning Sun, and Zunnan Huang. Evolutionary analysis and structural characterization of Aquilaria sinensis sesquiterpene synthase in agarwood formation: a computational study. Journal of Theoretical Biology, 456(??):249–260, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303795.

Lamberton:2014:LMP

[LCSH14] Timothy O. Lamberton, Nicholas D. Condon, Jennifer L. Stow, and Nicholas A. Hamilton. On linear models and parameter identifiability in experimental biological systems. Journal of Theoretical Biology, 358(??):102–121, October 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314003014.

Liu:2011:DSE

[LCT11] Xianning Liu, Xiaoping Chen, and Yasuhiro Takeuchi. Dynamics of an SIQS epidemic model with transport-related infection and exit-entry screenings. Journal of Theoretical Biology, 285(1):25–35, September 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003249.

Lade:2015:FLE

[LCTG15] Steven J. Lade, Miguel Coelho, Iva M. Tolić, and Thilo Gross. Fusion leads to effective segregation of damage during cell division: An analytical treatment. Journal of Theoretical Biology, 378(??):47–55, August 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001988.

Lin:2011:PIC

[LD11] Hao Lin and Hui Ding. Predicting ion channels and their types by the dipeptide mode of pseudo amino acid composition. Journal of Theoretical Biology, 269(1):64–69, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310005539.

Lucia:2018:IIF

[LD18] Umberto Lucia and Thomas S. Deisboeck. The importance of ion fluxes for cancer proliferation and metastasis: a thermodynamic analysis. Journal of Theoretical Biology, 445(??):1–8, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300791.

Letellier:2013:WCL

[LDA13] C. Letellier, F. Denis, and L. A. Aguirre. What can be learned from a chaotic cancer model? Journal of Theoretical Biology, 322(??):7–16, April 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000222.

Lorthois:2014:KMC

[LDB^{+}14] Sylvie Lorthois, Paul Duru, Ian Billanou, Michel Quintard, and Pierre Celsis. Kinetic modeling in the context of cerebral blood flow quantification by H_{2}^{1}5 O positron emission tomography : The meaning of the permeability coefficient in R e n k i n - - C r o n e’s model revisited at capillary scale. Journal of Theoretical Biology, 353(??):157–169, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001398.

Longhi:2013:APA

[LDdA^{+}13] Daniel Angelo Longhi, Francieli Dalcanton, Gláucia Maria Falcão de Aragão, Bruno Augusto Mattar Carciofi, and João Borges Laurindo. Assessing the prediction ability of different mathematical models for the growth of Lactobacillus plantarum under non-isothermal conditions. Journal of Theoretical Biology, 335(??):88–96, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313003020.

Laise:2011:DSA

[LDF^{+}11a] Pasquale Laise, Francesca Di Patti, Duccio Fanelli, Marika Masselli, and Annarosa Arcangeli. Deterministic and stochastic aspects of VEGF-A production and the cooperative behavior of tumoral cell colony. Journal of Theoretical Biology, 272(1):55–63, March 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000665X.

Lozada:2011:BEH

[LDF11b] Mariana Lozada, Paola D’Adamo, and Miguel Angel Fuentes. Beneficial effects of human altruism. Journal of Theoretical Biology, 289(??):12–16, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004115.

Li:2012:QDB

[LDH^{+}12] Yao Li, Gaurav Dwivedi, Wen Huang, Melissa L. Kemp, and Yingfei Yi. Quantification of degeneracy in biological systems for characterization of functional interactions between modules. Journal of Theoretical Biology, 302(??):29–38, June 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000938.

Lu:2014:CEM

[LDH^{+}14] Xinguo Lu, Yong Deng, Lei Huang, Bingtao Feng, and Bo Liao. A co-expression modules based gene selection for cancer recognition. Journal of Theoretical Biology, 362(??):75–82, December 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000149.

Langebrake:2014:TWR

[LDHD14] Jessica B. Langebrake, Gabriel E. Dilanji, Stephen J. Hagen, and Patrick De Leenheer. Traveling waves in response to a diffusing quorum sensing signal in spatially-extended bacterial colonies. Journal of Theoretical Biology, 363(??):53–61, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004457.

Liu:2016:BII

[LDJW16] Yang Liu, Yong Deng, Marko Jusup, and Zhen Wang. A biologically inspired immunization strategy for network epidemiology. Journal of Theoretical Biology, 400(??):92–102, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300479.

Lemay:2011:ESC

[LdLK11] Mathieu Lemay, Enno de Lange, and Jan P. Kucera. Effects of stochastic channel gating and distribution on the cardiac action potential. Journal of Theoretical Biology, 281(1):84–96, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311002190.

Lu:2016:OSB

[LDM16] Yaojie Lu, Remko A. Duursma, and Belinda E. Medlyn. Optimal stomatal behaviour under stochastic rainfall. Journal of Theoretical Biology, 394(??):160–171, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000047.

Ding:2015:CMP

[lDsWM^{+}15] Wei long Ding, Shui sheng Wu, Nelson Max, Fu li Wu, and Li feng Xu. A calculation method of plant similarity giving consideration to different plant features. Journal of Theoretical Biology, 387(??):136–143, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315004580.

Liu:2014:CDS

[LDWW14] Xiaowu Liu, Riqiang Deng, Jinwen Wang, and Xunzhang Wang. COStar: a d-star Lite-based dynamic search algorithm for codon optimization. Journal of Theoretical Biology, 344(??):19–30, March 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005493.

Lessard:2019:LHS

[LE19] Sabin Lessard and Warren J. Ewens. The left-hand side of the fundamental theorem of natural selection: a reply. Journal of Theoretical Biology, 472(??):77–83, July 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301559.

Lee:2016:LII

[Lee16a] Sungrim Seirin Lee. Lateral inhibition-induced pattern formation controlled by the size and geometry of the cell. Journal of Theoretical Biology, 404(??):51–65, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301175.

Lee:2016:PPF

[Lee16b] Sungrim Seirin Lee. Positioning of polarity formation by extracellular signaling during asymmetric cell division. Journal of Theoretical Biology, 400(??):52–64, July 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300236.

Lehtinen:2018:UVF

[Leh18] Sami Lehtinen. Understanding the Venus flytrap through mathematical modelling. Journal of Theoretical Biology, 444 (??):1–10, May 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519318300560.

Lei:2009:SSG

[Lei09] Jinzhi Lei. Stochasticity in single gene expression with both intrinsic noise and fluctuation in kinetic parameters. Journal of Theoretical Biology, 256(4):485–492, February 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519308005778. See corrigendum [Lei10].

Lei:2010:CSS

[Lei10] Jinzhi Lei. Corrigendum to: “Stochasticity in single gene expression with both intrinsic noise and fluctuation in kinetic parameters” [J. Theor. Biol. 256 (2009) 485–492]. Journal of Theoretical Biology, 262(2):381, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004718. See [Lei09].

Lenstra:2014:EGC

[Len14] Reijer Lenstra. Evolution of the genetic code through progressive symmetry breaking. Journal of Theoretical Biology, 347(??):95–108, April 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000113.

Leviyang:2014:CLB

[Lev14] Sivan Leviyang. Constructing lower-bounds for CTL escape rates in early SIV infection. Journal of Theoretical Biology, 352(??):82–91, July 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000952.

Li:2010:DDM

[LF10] Xu-Qing Li and Ping Fan. A duplex DNA model with regular inter-base-pair hydrogen bonds. Journal of Theoretical Biology, 266(3):374–379, October 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003450.

Lucas:2013:CDT

[LF13] Eric R. Lucas and Jeremy Field. Caste determination through mating in primitively eusocial societies. Journal of Theoretical Biology, 335(??):31–39, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002634.

Lawson:2016:MMI

[LF16] Brodie A. J. Lawson and Mark B. Flegg. A mathematical model for the induction of the mammalian ureteric bud. Journal of Theoretical Biology, 394(??):43–56, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000321.

Lai:2017:EMC

[LF17] Xiulan Lai and Avner Friedman. Exosomal microRNA concentrations in colorectal cancer: a mathematical model. Journal of Theoretical Biology, 415(??):70–83, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304180.

Lai:2019:MMS

[LF19] Xiulan Lai and Avner Friedman. Mathematical modeling in scheduling cancer treatment with combination of VEGF inhibitor and chemotherapy drugs. Journal of Theoretical Biology, 462(??):490–498, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830571X.

Levin:2016:SME

[LFB^{+}16] Drew Levin, Stephanie Forrest, Soumya Banerjee, Candice Clay, Judy Cannon, Melanie Moses, and Frederick Koster. A spatial model of the efficiency of T cell search in the influenzainfected lung. Journal of Theoretical Biology, 398(??):52–63, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001181.

Liu:2011:EEG

[LFD11] Xiao Liu, Yubo Fan, and Xiaoyan Deng. Effect of the endothelial glycocalyx layer on arterial LDL transport under normal and high pressure. Journal of Theoretical Biology, 283(1):71–81, August 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002827.

Li:2011:EAP

[LFM11] Jianquan Li, Paolo Fergola, and Zhien Ma. Effects of allelochemicals produced by one population in a chemostat-like environment. Journal of Theoretical Biology, 284(1):99–105, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003006.

Lee:2014:CBV

[LFM14] Sang-Woo Lee, Francois Gabriel Feugier, and Yoshihiro Morishita. Canalization-based vein formation in a growing leaf. Journal of Theoretical Biology, 353(??):104–120, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314001404.

Liu:2015:IMP

[LFW^{+}15] Bin Liu, Longyun Fang, Shanyi Wang, Xiaolong Wang, Hongtao Li, and Kuo-Chen Chou. Identification of microRNA precursor with the degenerate k-tuple or kmer strategy. Journal of Theoretical Biology, 385(??):153–159, November 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315004269.

Lascano:2018:EAM

[LFW^{+}18] Elena C. Lascano, Juan I. Felice, Sandra Wray, Sara Kosta, Pierre C. Dauby, Edmundo I. Cabrera-Fischer, and Jorge A. Negroni. Experimental assessment of a myocyte-based multiscale model of cardiac contractile dysfunction. Journal of Theoretical Biology, 456(??):16–28, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303692.

Landman:2011:BSC

[LFZN11] Kerry A. Landman, Anthony E. Fernando, Dongcheng Zhang, and Donald F. Newgreen. Building stable chains with motile agents: Insights into the morphology of enteric neural crest cell migration. Journal of Theoretical Biology, 276(1):250– 268, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000695.

Lavric:2010:BGD

[LG10a] Vasile Lavric and David W. Graham. Birth, growth and death as structuring operators in bacterial population dynamics. Journal of Theoretical Biology, 264(1):45–54, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310000329.

Liu:2010:PCN

[LG10b] Xinsheng Liu and Wanlin Guo. Persistent cooperators in nature. Journal of Theoretical Biology, 267(4):647–652, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005114.

Louzoun:2012:EVL

[LG12] Yoram Louzoun and Vitaly V. Ganusov. Evolution of viral life-cycle in response to cytotoxic T lymphocyte-mediated immunity. Journal of Theoretical Biology, 310(??):3–13, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003074.

Leloup:2013:CPS

[LG13] Jean-Christophe Leloup and Albert Goldbeter. Critical phase shifts slow down circadian clock recovery: Implications for jet lag. Journal of Theoretical Biology, 333(??):47–57, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002142.

Lebre:2017:COG

[LG17] Sophie Lèbre and Olivier Gascuel. The combinatorics of overlapping genes. Journal of Theoretical Biology, 415(??):90–101, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303150.

Levy:2018:EMP

[LG18] Orr Levy and Isak Nethanël Gath. Effects of mating patterns on genealogical trees: Assessment of the high carrier rate of familial Mediterranean fever in rural Israeli districts. Journal of Theoretical Biology, 443(??):92–99, April 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300444.

Lehtinen:2019:CCP

[LG19a] Sami O. Lehtinen and Stefan A. H. Geritz. Coevolution of cannibalistic predators and timid prey: evolutionary cycling and branching. Journal of Theoretical Biology, 483(??):Article 110001, December 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303704.

Lehtinen:2019:CPE

[LG19b] Sami O. Lehtinen and Stefan A. H. Geritz. Cyclic prey evolution with cannibalistic predators. Journal of Theoretical Biology, 479(??):1–13, October 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302668.

Li:2015:MEG

[LGBT15] Xiang-Yi Li, Stefano Giaimo, Annette Baudisch, and Arne Traulsen. Modeling evolutionary games in populations with demographic structure. Journal of Theoretical Biology, 380(??):506–515, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002714.

Lopez-Garcia:2017:SEE

[LGEM17] Purificación López-García, Laura Eme, and David Moreira. Symbiosis in eukaryotic evolution. Journal of Theoretical Biology, 434(??):20–33, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300954.

Lee:2009:MSE

[LGK^{+}09] Ha Youn Lee, Elena E. Giorgi, Brandon F. Keele, Brian Gaschen, Gayathri S. Athreya, Jesus F. Salazar-Gonzalez, Kimmy T. Pham, Paul A. Goepfert, J. Michael Kilby, Michael S. Saag, Eric L. Delwart, Michael P. Busch, Beatrice H. Hahn, George M. Shaw, Bette T. Korber, Tanmoy Bhattacharya, and Alan S. Perelson. Modeling sequence evolution in acute HIV-1 infection. Journal of Theoretical Biology, 261(2):341–360, November 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519309003506. See corrigendum [LGK^{+}12].

Lee:2012:CMS

[LGK^{+}12] Ha Youn Lee, Elena E. Giorgi, Brandon F. Keele, Brian Gaschen, Gayathri S. Athreya, Jesus F. Salazar-Gonzalez, Kimmy T. Pham, Paul A. Goepfert, J. Michael Kilby, Michael S. Saag, Eric L. Delwart, Michael P. Busch, Beatrice H. Hahn, George M. Shaw, Bette T. Korber, Tanmoy Bhattacharya, and Alan S. Perelson. Corrigendum to “Modeling sequence evolution in acute HIV-1 infection” [J. Theor. Biol. 261 (2009) 341–360]. Journal of Theoretical Biology, 297(??):187, March 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006199. See [LGK^{+}09].

Gu:2014:MCS

[lGLLL14] Jian lei Gu, Yao Lu, Cong Liu, and Hui Lu. Multiclass classification of sarcomas using pathway based feature selection method. Journal of Theoretical Biology, 362(??):3–8, December 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003816.

Lemmer:2014:MIT

[LGMV14] Yolandy Lemmer, Anne Grobler, Clint Moody, and Hendrik Viljoen. A model of isoniazid treatment of tuberculosis. Journal of Theoretical Biology, 363(??):367–373, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004366.

Lucia:2017:ETA

[LGPS17] Umberto Lucia, Giulia Grisolia, Antonio Ponzetto, and Francesca Silvagno. An engineering thermodynamic approach to select the electromagnetic wave effective on cell growth. Journal of Theoretical Biology, 429(??):181–189, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303077.

Lipfert:2012:MEC

[LGR^{+}12] Susanne W. Lipfert, Michael Günther, Daniel Renjewski, Sten Grimmer, and Andre Seyfarth. A model-experiment comparison of system dynamics for human walking and running. Journal of Theoretical Biology, 292(??):11–17, January 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004851.

Liu:2013:MBA

[LH13] Qin Liu and Hai-Chao Han. Mechanical buckling of arterioles in collateral development. Journal of Theoretical Biology, 316(??):42–48, January 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005073.

Li:2014:MID

[LH14] Yan Li and Andreas Handel. Modeling inoculum dose dependent patterns of acute virus infections. Journal of Theoretical Biology, 347(??):63–73, April 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000174.

Loeuille:2018:MPD

[LH18] Nicolas Loeuille and Céline Hauzy. Multidimensionality of plant defenses and herbivore niches: Implications for eco-evolutionary dynamics. Journal of Theoretical Biology, 445(??):110–119, May 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300584.

Landi:2015:FID

[LHD15] Pietro Landi, Cang Hui, and Ulf Dieckmann. Fisheriesinduced disruptive selection. Journal of Theoretical Biology, 365(??):204–216, January 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006067.

Liu:2017:LDI

[LHD^{+}17] Hui Liu, Dawei Hu, Chen Dong, Yuming Fu, Guanghui Liu, Youcai Qin, Yi Sun, Dianlei Liu, Lei Li, and Hong Liu. Low-dose ionizing radiation limitations to seed germination: Results from a model linking physiological characteristics and developmental-dynamics simulation strategy. Journal of Theoretical Biology, 427(??):10–16, August 1, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302382.

Leung:2016:PSS

[LHFM16] Tiffany Leung, Barry D. Hughes, Federico Frascoli, and James M. McCaw. Periodic solutions in an SIRWS model with immune boosting and cross-immunity. Journal of Theoretical Biology, 410(??):55–64, December 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302776.

Lu:2012:NHF

[LHH12] Jin-Long Lu, Xue-Hai Hu, and Dong-Gang Hu. A new hybrid fractal algorithm for predicting thermophilic nucleotide sequences. Journal of Theoretical Biology, 293(??):74–81, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004929.

Lv:2014:ICM

[LHL^{+}14] Hongqiang Lv, Jiuqiang Han, Jun Liu, Jiguang Zheng, Dexing Zhong, and Ruiling Liu. ISDTool 2.0: a computational model for predicting immunosuppressive domain of retroviruses. Journal of Theoretical Biology, 360(??):78–82, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003762.

Lopez:2018:CEM

[LHPF18] J. M. López, J. Herrero, D. Puigjaner, and G. Fortuny. Corrigendum to “A 3D electro-mechanical continuum model for simulating skeletal muscle contraction”, J. Theor. Biol. 335 (2013) 108–118. Journal of Theoretical Biology, 446(??):259, June 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301115. See [HGGCR13].

Levy:2016:RSP

[LHV16] Dorian Levy, Heather A. Harrington, and Robert A. Van Gorder. Role of seasonality on predator–prey-subsidy population dynamics. Journal of Theoretical Biology, 396(??):163– 181, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316000953.

Lettmann:2014:ILM

[LHW14] Karsten Alexander Lettmann and Matthias Hardtke-Wolenski. The importance of liver microcirculation in promoting autoimmune hepatitis via maintaining an inflammatory cytokine milieu — a mathematical model study. Journal of Theoretical Biology, 348(??):33–46, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000332.

Li:2011:PHG

[Li11] Wentian Li. On parameters of the human genome. Journal of Theoretical Biology, 288(??):92–104, November 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003717.

Lopez:2017:DDP

[LIB^{+}17] Álvaro G. López, Kelly C. Iarosz, Antonio M. Batista, Jesús M. Seoane, Ricardo L. Viana, and Miguel A. F. Sanjuán. The dose-dense principle in chemotherapy. Journal of Theoretical Biology, 430(??):169–176, October 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303284.

Lion:2016:MES

[Lio16] Sébastien Lion. Moment equations in spatial evolutionary ecology. Journal of Theoretical Biology, 405(??):46– 57, September 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251931500507X.

Lion:2018:PES

[Lio18] Sébastien Lion. From the Price equation to the selection gradient in class-structured populations: a quasi-equilibrium route. Journal of Theoretical Biology, 447(??):178–189, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301474. See corrigen-

dum [Lio24].

Lion:2024:CPE

[Lio24] Sébastien Lion. Corrigendum to “From the Price equation to the selection gradient in class-structured populations: a quasi-equilibrium route” [J. Theor. Biol. 447 (2018) 178– 189]. Journal of Theoretical Biology, 591(??):??, August 21, 2024. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519324001577. See [Lio18].

Louca:2012:SNP

[LIPD12] Stilianos Louca, Sébastien Ibanez, Didier Piau, and Laurence Després. Specialized nursery pollination mutualisms as evolutionary traps stabilized by antagonistic traits. Journal of Theoretical Biology, 296(??):65–83, March 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005984.

Liu:2010:HAP

[LJ10] Tian Liu and Cangzhi Jia. A high-accuracy protein structural class prediction algorithm using predicted secondary structural information. Journal of Theoretical Biology, 267(3):272–275, December 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000473X.

Lee:2015:STT

[LJ15] Jonggul Lee and Eunok Jung. A spatial-temporal transmission model and early intervention policies of 2009 A/H1N1 influenza in South Korea. Journal of Theoretical Biology, 380(??):60–73, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002362.

Liu:2018:BDG

[LJ18] Wei Liu and Yaolin Jiang. Bifurcation of a delayed Gause predator–prey model with Michaelis–Menten type harvesting. Journal of Theoretical Biology, 438(??):116–132, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305088.

Lee:2017:GCP

[LJI17] Joung-Hun Lee, Marko Jusup, and Yoh Iwasa. Games of corruption in preventing the overuse of common-pool resources. Journal of Theoretical Biology, 428(??):76–86, September 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931730262X.

Lakizadeh:2015:PGP

[LJM15] Amir Lakizadeh, Saeed Jalili, and Sayed-Amir Marashi. PCD-GED: Protein complex detection considering PPI dynamics based on time series gene expression data. Journal of Theoretical Biology, 378(??):31–38, August 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001964.

Liao:2017:SDP

[LJM^{+}17] Kang-Ling Liao, Roger D. Jones, Patrick McCarter, Meral Tunc-Ozdemir, James A. Draper, Timothy C. Elston, David Kramer, and Alan M. Jones. A shadow detector for photosynthesis efficiency. Journal of Theoretical Biology, 414(??):231–244, February 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316304052.

Llensa:2014:EED

[LJS14] Carlos Llensa, David Juher, and Joan Saldaña. On the early epidemic dynamics for pairwise models. Journal of Theoretical Biology, 352(??):71–81, July 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001210.

Lohier:2016:PSC

[LJW^{+}16] Théophile Lohier, Franck Jabot, Alexandra Weigelt, Bernhard Schmid, and Guillaume Deffuant. Predicting stochastic community dynamics in grasslands under the assumption of competitive symmetry. Journal of Theoretical Biology, 399(??):53– 61, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300194.

Leung:2015:MSW

[LK15] M.-R. Leung and M. Kot. Models for the spread of white pine blister rust. Journal of Theoretical Biology, 382(??):328–336, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003550.

Liu:2019:AIA

[LK19] Benjamin R. Liu and Mark Kot. Accelerating invasions and the asymptotics of fat-tailed dispersal. Journal of Theoretical Biology, 471(??):22–41, June 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301237.

Lone:2018:EMT

[LKAJ18] Mohsin Yousuf Lone, Sivakumar Prasanth Kumar, Mohd Athar, and Prakash Chandra Jha. Exploration of mycobacterium tuberculosis structural proteome: An in-silico approach. Journal of Theoretical Biology, 439(??):14–23, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305283.

Lee:2013:OCI

[LKK13] Jeehyun Lee, Jungeun Kim, and Hee-Dae Kwon. Optimal control of an influenza model with seasonal forcing and age-dependent transmission rates. Journal of Theoretical Biology, 317(??):310–320, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005723.

Liang:2013:NSF

[LKL13] Chen-Wei Liang, Chien-Kuo Ku, and Jeng-Jong Liang. A new scale-free network model for simulating and predicting epidemics. Journal of Theoretical Biology, 317(??):11–19, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004985.

Lee:2017:MDI

[LKO^{+}17] Elizabeth C. Lee, Michael R. Kelly, Brad M. Ochocki, Segun M. Akinwumi, Karen E. S. Hamre, Joseph H. Tien, and Marisa C. Eisenberg. Model distinguishability and inference robustness in mechanisms of cholera transmission and loss of immunity. Journal of Theoretical Biology, 420(??):68–81, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300322.

Li:2012:SBO

[LKP^{+}12] Weihan Li, Sandeep Krishna, Simone Pigolotti, Namiko Mitarai, and Mogens H. Jensen. Switching between oscillations and homeostasis in competing negative and positive feedback motifs. Journal of Theoretical Biology, 307(??):205–210, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001920.

Laga:2014:LFS

[LKSM14] Hamid Laga, Sebastian Kurtek, Anuj Srivastava, and Stanley J. Miklavcic. Landmark-free statistical analysis of the shape of plant leaves. Journal of Theoretical Biology, 363(??):41–52, December 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004500.

Lisowski:2015:CMO

[LKZB15] Bartosz Lisowski, Lukasz Kuśmierz, Michał Zabicki, and Martin Bier. “cargo-mooring” as an operating principle for molecular motors. Journal of Theoretical Biology, 374(??):26–34, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001101.

Li:2012:APR

[LL12] Tao Li and Qian-Zhong Li. Annotating the protein-RNA interaction sites in proteins using evolutionary information and protein backbone structure. Journal of Theoretical Biology, 312(??):55–64, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003724.

Li:2013:CSA

[LL13a] Yao-Wang Li and Bo Li. Characterization of structureantioxidant activity relationship of peptides in free radical systems using QSAR models: Key sequence positions and their amino acid properties. Journal of Theoretical Biology, 318(??):29–43, February 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005681.

Liu:2013:SRF

[LL13b] Ying Liu and Chunguang Li. Stochastic resonance in feedforward-loop neuronal network motifs in astrocyte field. Journal of Theoretical Biology, 335(??):265–275, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003287.

Li:2014:ECC

[LL14] Jonathan F. Li and John Lowengrub. The effects of cell compressibility, motility and contact inhibition on the growth of tumor cell clusters using the cellular Potts model. Journal of Theoretical Biology, 343(??):79–91, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005006.

Lejeune:2017:QRB

[LL17a] Emma Lejeune and Christian Linder. Quantifying the relationship between cell division angle and morphogenesis through computational modeling. Journal of Theoretical Biology, 418(??):1–7, April 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300267.

Lyu:2017:MLI

[LL17b] Hui Lyu and Dusan Lazár. Modeling the light-induced electric potential difference (ΔΨ), the pH difference (Δ pH) and the proton motive force across the thylakoid membrane in C_{3} leaves. Journal of Theoretical Biology, 413(??):11–23, January 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303563.

Li:2018:SSM

[LL18] Yazhi Li and Xianning Liu. A sex-structured model with birth pulse and release strategy for the spread of Wolbachia in mosquito population. Journal of Theoretical Biology, 448(??):53–65, July 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301619.

Lombarte:2018:VMR

[LLB^{+}18] Mercedes Lombarte, Maela Lupo, L. Fina Brenda, German Campetelli, A. R. Buzalaf Marilia, Marta Basualdo, and Alfredo Rigalli. In vivo measurement of the rate constant of liver handling of glucose and glucose uptake by insulin-dependent tissues, using a mathematical model for glucose homeostasis in diabetic rats. Journal of Theoretical Biology, 439(??):205–215, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305350.

Li:2015:NTG

[LLC15] Xiong Li, Bo Liao, and Haowen Chen. A new technique for generating pathogenic barcodes in breast cancer susceptibility analysis. Journal of Theoretical Biology, 366(??):84–90, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006432.

Liu:2012:SDP

[LLCC12] Guoqing Liu, Jia Liu, Xiangjun Cui, and Lu Cai. Sequencedependent prediction of recombination hotspots in Saccharomyces cerevisiae. Journal of Theoretical Biology, 293(??): 49–54, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251931100511X.

Li:2017:SPR

[LLD^{+}17] Qiang Li, Furong Lu, Chenxi Dai, Minjun Fan, Weiming Wang, and Kaifa Wang. Simulating the potential role of media coverage and infected bats in the 2014 Ebola outbreak. Journal of Theoretical Biology, 412(??):123–129, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303526.

Li:2014:MEC

[LLDW14] Qiang Li, Furong Lu, Guohong Deng, and Kaifa Wang. Modeling the effects of covalently closed circular DNA and dendritic cells in chronic HBV infection. Journal of Theoretical Biology, 357(??):1–9, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002719.

Lehmann:2016:FAA

[LLG16] Jean Lehmann, Albert Libchaber, and Benjamin D. Greenbaum. Fundamental amino acid mass distributions and entropy costs in proteomes. Journal of Theoretical Biology, 410(??):119–124, December 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931630248X.

Liao:2013:SPL

[LLH^{+}13] Jinbao Liao, Zhenqing Li, David E. Hiebeler, Yoh Iwasa, Jan Bogaert, and Ivan Nijs. Species persistence in landscapes with spatial variation in habitat quality: a pair approximation model. Journal of Theoretical Biology, 335(??):22–30, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002877.

Liu:2018:DAE

[LLJ18] Guirong Liu, Zhimei Liu, and Zhen Jin. Dynamics analysis of epidemic and information spreading in overlay networks. Journal of Theoretical Biology, 444(??):28–37, May 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300626.

Lee:2014:LIF

[LLK14] Sang Joon Lee, Seung Chul Lee, and Bo Heum Kim. Liquidintake flow around the tip of butterfly proboscis. Journal of Theoretical Biology, 348(??):113–121, May 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000617.

Liu:2013:DBS

[LLL13] Yiliang Liu, Qun Liu, and Zhenhai Liu. Dynamical behaviors of a stochastic delay logistic system with impulsive toxicant input in a polluted environment. Journal of Theoretical Biology, 329(??):1–5, July 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000970.

Li:2017:SCE

[LLL^{+}17] Yushuang Li, Yanfen Lv, Xiaonan Li, Wenli Xiao, and Chun Li. Sequence comparison and essential gene identification with new inter-nucleotide distance sequences. Journal of Theoretical Biology, 418(??):84–93, April 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300310.

Lander:2011:SPS

[LLLV11] Michelle E. Lander, Miles L. Logsdon, Thomas R. Loughlin, and Glenn R. Van Blaricom. Spatial patterns and scaling behaviors of Steller sea lion (Eumetopias jubatus) distributions and their environment. Journal of Theoretical Biology, 274(1):74–83, April 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000270.

Liu:2013:UPG

[LlLYB13] Zhi-Xin Liu, Song lei Liu, Hong-Qiang Yang, and Li-Hua Bao. Using protein granularity to extract the protein sequence features. Journal of Theoretical Biology, 331(??):48–53, August 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001872.

Laager:2019:MMD

[LLN^{+}19] Mirjam Laager, Monique Léchenne, Kemdongarti Naissengar, Rolande Mindekem, Assandi Oussiguere, Jakob Zinsstag, and Nakul Chitnis. A metapopulation model of dog rabies transmission in N’Djamena, Chad. Journal of Theoretical Biology, 462(??):408–417, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305800.

Li:2019:NEP

[LLP^{+}19] Gaoshi Li, Min Li, Wei Peng, Yaohang Li, Yi Pan, and Jianxin Wang. A novel extended Pareto Optimality Consensus model for predicting essential proteins. Journal of Theoretical Biology, 480(??):141–149, November 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303170.

Lamberton:2015:CDS

[LLS^{+}15] Timothy O. Lamberton, James Lefevre, Kieran M. Short, Ian M. Smyth, and Nicholas A. Hamilton. Comparing and distinguishing the structure of biological branching. Journal of Theoretical Biology, 365(??):226–237, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005906.

Laidre:2013:MSI

[LLSO13] Mark E. Laidre, Alex Lamb, Susanne Shultz, and Megan Olsen. Making sense of information in noisy networks: Human communication, gossip, and distortion. Journal of Theoretical Biology, 317(??):152–160, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004870.

Labavic:2019:SLC

[LLTP19] Darka Labavić, Mohamed Tahar Ladjimi, Quentin Thommen, and Benjamin Pfeuty. Scaling laws of cell-fate responses to transient stress. Journal of Theoretical Biology, 478(??):14–25, October 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302413.

Lau:2015:CHG

[LLW15] John Wei Lau, David N. Levy, and Dominik Wodarz. Contribution of HIV-1 genomes that do not integrate to the basic reproductive ratio of the virus. Journal of Theoretical Biology, 367(??):222–229, February 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006936.

Li:2018:IEP

[LLW^{+}18] Min Li, Wenkai Li, Fang-Xiang Wu, Yi Pan, and Jianxin Wang. Identifying essential proteins based on sub-network partition and prioritization by integrating subcellular localization information. Journal of Theoretical Biology, 447(??):65– 73, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301437.

Liu:2015:IPC

[LLY15] Yu Liu, Xue Liang, and Rongcun Yang. Identification of peculiar and common effects of histone modifications on transcription. Journal of Theoretical Biology, 380(??):24–39, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002222.

Li:2013:EGB

[LLZ13] Xiaobo Li, Scott C. Lenaghan, and Mingjun Zhang. Evolutionary game based control for biological systems with applications in drug delivery. Journal of Theoretical Biology, 326(??):58–69, June 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000040.

Lei:2011:MAS

[LM11] Jinzhi Lei and Michael C. Mackey. Multistability in an age-structured model of hematopoiesis: Cyclical neutropenia. Journal of Theoretical Biology, 270(1):143–153, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006090.

Lim:2014:HID

[LM14] Aaron G. Lim and Philip K. Maini. HTLV-i infection: a dynamic struggle between viral persistence and host immunity. Journal of Theoretical Biology, 352(??):92–108, July 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000976.

Lee:2015:ATC

[LM15] Robert H. Lee and Cassie S. Mitchell. Axonal transport cargo motor count versus average transport velocity: Is fast versus slow transport really single versus multiple motor transport? Journal of Theoretical Biology, 370(??):39–44, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000193.

Lefevre:2013:MCT

[LMC^{+}13] J. Lefevre, D. J. Marshall, A. N. Combes, A. L. Ju, M. H. Little, and N. A. Hamilton. Modelling cell turnover in a complex tissue during development. Journal of Theoretical Biology, 338(??):66–79, December 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004074.

Lindberg:2018:WHD

[LMCW18] Heather M. Lindberg, Kurt A. McKean, T. Caraco, and Ing-Nang Wang. Within-host dynamics and random duration of pathogen infection: Implications for between-host transmission. Journal of Theoretical Biology, 446(??):137–148, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300468.

Lopez:2011:FIP

[LMdL11] Diego Lopez, Sébastien Michelin, and Emmanuel de Langre. Flow-induced pruning of branched systems and brittle reconfiguration. Journal of Theoretical Biology, 284(1):117–124, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003328.

Lo:2013:MMC

[LMHF13] Wing-Cheong Lo, Edward W. Martin, Charles L. Hitchcock, and Avner Friedman. Mathematical model of colitis-associated colon cancer. Journal of Theoretical Biology, 317(??):20–29, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005036.

Lin:2016:MMP

[LMJ^{+}16] Mabelle Lin, Benjamin Mauroy, Joanna L. James, Merryn H. Tawhai, and Alys R. Clark. A multiscale model of placental oxygen exchange: The effect of villous tree structure on exchange efficiency. Journal of Theoretical Biology, 408(??):1–12, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301710.

Liu:2018:NHB

[LMJ^{+}18] Wei Liu, Liangyu Ma, Byeungwoo Jeon, Ling Chen, and Bolun Chen. A network hierarchy-based method for functional module detection in protein–protein interaction networks. Journal of Theoretical Biology, 455(??):26–38, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303266.

Lee:2013:AHD

[LMM^{+}13] Eric F. Lee, Mark A. Matthews, Andrew J. McElrone, Ronald J. Phillips, Kenneth A. Shackel, and Craig R. Brodersen. Analysis of HRCT-derived xylem network reveals reverse flow in some vessels. Journal of Theoretical Biology, 333(??):146–155, September 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002506.

Lopez-Marin:2018:SMA

[LMM18] N. López-Marín and R. Mulet. In silico modelling of apoptosis induced by photodynamic therapy. Journal of Theoretical Biology, 436(??):8–17, January 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304460.

Li:2010:DMD

[LMW^{+}10] Yao Li, Changxing Ma, Zhong Wang, Gang Chen, Kwangmi Ahn, Philip Lazarus, and Rongling Wu. A disequilibrium model for detecting genetic mutations for cancer. Journal of Theoretical Biology, 265(3):218–224, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000202X.

Li:2015:DAC

[LMYL15] Jianquan Li, Ke Men, Yali Yang, and Duan Li. Dynamical analysis on a chronic hepatitis C virus infection model with immune response. Journal of Theoretical Biology, 365(??):337–346, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006298.

Li:2013:RQA

[LN13] Jun Li and Fahima Nekka. A rational quantitative approach to determine the best dosing regimen for a target therapeutic effect: a unified formalism for antibiotic evaluation. Journal of Theoretical Biology, 319(??):88–95, February 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005942.

Lee:2019:STP

[LN19] Hyojung Lee and Hiroshi Nishiura. Sexual transmission and the probability of an end of the Ebola virus disease epidemic. Journal of Theoretical Biology, 471(??):1–12, June 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301298.

Lintott:2013:IID

[LNH13] R. A. Lintott, R. A. Norman, and A. S. Hoyle. The impact of increased dispersal in response to disease control in patchy environments. Journal of Theoretical Biology, 323(??):57–68, April 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313000520.

Li:2011:DMP

[LNL^{+}11] Zhiyuan Li, Ming Ni, Jikun Li, Yuping Zhang, Qi Ouyang, and Chao Tang. Decision making of the p53 network: Death by integration. Journal of Theoretical Biology, 271(1):205–211, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006363.

Lutianov:2011:MMC

[LNRK11] Michael Lutianov, Shailesh Naire, Sally Roberts, and Jan-Herman Kuiper. A mathematical model of cartilage regeneration after cell therapy. Journal of Theoretical Biology, 289(??):136–150, November 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004024.

Liu:2015:CCM

[LO15] Weifan Liu and Sarah D. Olson. Compartment calcium model of frog skeletal muscle during activation. Journal of Theoretical Biology, 364(??):139–153, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005372.

LeCunff:2012:PGR

[LP12a] Yann Le Cunff and Khashayar Pakdaman. Phenotypegenotype relation in Wagner’s canalization model. Journal of Theoretical Biology, 314(??):69–83, December 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004420.

Liz:2012:GDS

[LP12b] Eduardo Liz and Paweł Pilarczyk. Global dynamics in a stage-structured discrete-time population model with harvesting. Journal of Theoretical Biology, 297(??):148–165, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311006217.

LeCunff:2014:RCR

[LP14] Yann Le Cunff and Khashayar Pakdaman. Reproduction cost reduces demographic stochasticity and enhances inter-individual compatibility. Journal of Theoretical Biology, 360(??):263–270, November 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004019.

Lazcano:2017:OMC

[LP17] Antonio Lazcano and Juli Peretó. On the origin of mitosing cells: a historical appraisal of Lynn Margulis endosymbiotic theory. Journal of Theoretical Biology, 434(??):80–87, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303223.

Lee:2019:IKF

[LP19] Tamsin E. Lee and Melissa A. Penny. Identifying key factors of the transmission dynamics of drug-resistant malaria. Journal of Theoretical Biology, 462(??):210–220, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305320.

Landmann:2017:STA

[LPB17] Stefan Landmann, Nicolas Preuss, and Ulrich Behn. Selftolerance and autoimmunity in a minimal model of the idiotypic network. Journal of Theoretical Biology, 426(??):17–39, August 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302047.

Landmann:2018:CST

[LPB18] Stefan Landmann, Nicolas Preuss, and Ulrich Behn. Corrigendum to ‘Self-tolerance and autoimmunity in a minimal model of the idiotypic network’ [Journal of Theoretical Biology 426 (2017) 17–39]. Journal of Theoretical Biology, 447(??):199, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301450.

Lyons:2016:PFR

[LPD^{+}16] James Lyons, Kuldip K. Paliwal, Abdollah Dehzangi, Rhys Heffernan, Tatsuhiko Tsunoda, and Alok Sharma. Protein fold recognition using HMM-HMM alignment and dynamic programming. Journal of Theoretical Biology, 393(??):67–74, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316000229.

Livnat:2011:ACB

[LPF11] Adi Livnat, Christos Papadimitriou, and Marcus W. Feldman. An analytical contrast between fitness maximization and selection for mixability. Journal of Theoretical Biology, 273(1):232–234, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000634X.

Lema-Perez:2019:PBM

[LPGTBMA19] Laura Lema-Perez, Jose Garcia-Tirado, Carlos Builes-Montaño, and Hernan Alvarez. Phenomenological-based model of human stomach and its role in glucose metabolism. Journal of Theoretical Biology, 460(??):88–100, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305009.

Liu:2018:PCE

[LPH18] Xuesong Liu, Qiuhui Pan, and Mingfeng He. Promotion of cooperation in evolutionary game dynamics with local information. Journal of Theoretical Biology, 437(??):1–8, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304769.

Liu:2019:SDF

[LPH19] Xuesong Liu, Qiuhui Pan, and Mingfeng He. Stochastic dynamics in the fitness-based process which can be on behalf of the standard Moran, local and Wright–Fisher processes. Journal of Theoretical Biology, 460(??):79–87, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305046.

Liu:2015:FPE

[LPKH15a] Xuesong Liu, Qiuhui Pan, Yibin Kang, and Mingfeng He. Fixation probabilities in evolutionary games with the Moran and Fermi processes. Journal of Theoretical Biology, 364(??):242–248, January 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005347.

Liu:2015:FTE

[LPKH15b] Xuesong Liu, Qiuhui Pan, Yibin Kang, and Mingfeng He. Fixation times in evolutionary games with the Moran and Fermi processes. Journal of Theoretical Biology, 387(??):214–220, December 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004592.

Labarthe:2019:MMI

[LPP^{+}19] Simon Labarthe, Bastien Polizzi, Thuy Phan, Thierry Goudon, Magali Ribot, and Beatrice Laroche. A mathematical model to investigate the key drivers of the biogeography of the colon microbiota. Journal of Theoretical Biology, 462(??):552–581, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830599X.

Lahiri:2014:IAS

[LPS^{+}14] Chandrajit Lahiri, Shrikant Pawar, Radhakrishnan Sabarinathan, Md. Izhar Ashraf, Yamini Chand, and Dipshikha Chakravortty. Interactome analyses of Salmonella pathogenicity islands reveal SicA indispensable for virulence. Journal of Theoretical Biology, 363(??):188–197, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004615.

Lloyd-Price:2014:DSG

[LPTR14] Jason Lloyd-Price, Huy Tran, and Andre S. Ribeiro. Dynamics of small genetic circuits subject to stochastic partitioning in cell division. Journal of Theoretical Biology, 356(??):11–19, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002380.

Lyon:2011:MSP

[LPvSP11] James E. Lyon, Sagar A. Pandit, Carel P. van Schaik, and Gauri R. Pradhan. Mating strategies in primates: a game theoretical approach to infanticide. Journal of Theoretical Biology, 274(1):103–108, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519311000099. See corrigendum [LPvSP12].

Lyon:2012:CMS

[LPvSP12] James E. Lyon, Sagar A. Pandit, Carel P. van Schaik, and Gauri R. Pradhan. Corrigendum to “Mating strategies in primates: a game theoretical approach to infanticide” [J. Theor. Biol. 294 (2011) 103–108]. Journal of Theoretical Biology, 301(??):174, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000860. See [LPvSP11].

Laruson:2016:SUG

[LR16] Áki J. Láruson and Floyd A. Reed. Stability of underdominant genetic polymorphisms in population networks. Journal of Theoretical Biology, 390(??):156–163, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005743.

Long:2013:CSM

[LRA^{+}13] Byron L. Long, Rahul Rekhi, Amada Abrego, Jiwon Jung, and Amina A. Qutub. Cells as state machines: Cell behavior patterns arise during capillary formation as a function of BDNF and VEGF. Journal of Theoretical Biology, 326(??):43–57, June 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312006194.

Lutz:2013:ICF

[LRGA13] Alessandra F. Lütz, Sebastián Risau-Gusman, and Jeferson J. Arenzon. Intransitivity and coexistence in four species cyclic games. Journal of Theoretical Biology, 317(??):286–292, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005632.

Lyles:2015:PRE

[LRH15] Danielle Lyles, Todd S. Rosenstock, and Alan Hastings. Plant reproduction and environmental noise: How do plants do it? Journal of Theoretical Biology, 371(??):137–144, April 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000685.

Larbat:2016:MDP

[LRL^{+}16] Romain Larbat, Christophe Robin, Cathrine Lillo, Tormod Drengstig, and Peter Ruoff. Modeling the diversion of primary carbon flux into secondary metabolism under variable nitrate and light/dark conditions. Journal of Theoretical Biology, 402(??):144–157, August 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300868.

Lee:2010:EMD

[LS10a] Thomas J. Lee and Michael P. Speed. The effect of metapopulation dynamics on the survival and spread of a novel, conspicuous prey. Journal of Theoretical Biology, 267(3):319–329, December 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004509.

Levy:2010:PSS

[LS10b] David-Elie Levy and Avraham Seifert. Parameter study of simplified dragonfly airfoil geometry at Reynolds number of 6000. Journal of Theoretical Biology, 266(4):691–702, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003656.

Lloyd-Smith:2010:MDD

[LS10c] James O. Lloyd-Smith. Modeling density dependence in heterogeneous landscapes: Dispersal as a case study. Journal of Theoretical Biology, 265(2):160–166, July 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001657.

Lou:2011:MEA

[LS11a] Jie Lou and Robert J. Smith. Modelling the effects of adherence to the HIV fusion inhibitor enfuvirtide. Journal of Theoretical Biology, 268(1):1–13, January 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005163.

Lutscher:2011:ETV

[LS11b] Frithjof Lutscher and Gunog Seo. The effect of temporal variability on persistence conditions in rivers. Journal of Theoretical Biology, 283(1):53–59, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002840.

Lambert:2013:PLP

[LS13a] Amaury Lambert and Mike Steel. Predicting the loss of phylogenetic diversity under non-stationary diversification models. Journal of Theoretical Biology, 337(??):111–124, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003834.

Lehmann:2013:GRN

[LS13b]  Malte Lehmann and Kim Sneppen.  Genetic regula-

tory networks that count to 3. Journal of Theoretical Biology, 329(??):15–19, July 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001495.

Liang:2014:DFS

[LS14a] Bin Liang and Mao Sun. Dynamic flight stability of a hovering model dragonfly. Journal of Theoretical Biology, 348(??):100– 112, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000435.

Lin:2014:IME

[LS14b] Yangchen Lin and William J. Sutherland. Interaction modification effects on ecological networks are affected by ratio dependence and network topology. Journal of Theoretical Biology, 363(??):151–157, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400486X.

Laird:2015:CIP

[LS15a] Robert A. Laird and Brandon S. Schamp. Competitive intransitivity, population interaction structure, and strategy coexistence. Journal of Theoretical Biology, 365(??):149–158, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005992.

Lee:2015:SOA

[LS15b] Sungrim Seirin Lee and Tatsuo Shibata. Self-organization and advective transport in the cell polarity formation for asymmetric cell division. Journal of Theoretical Biology, 382(??):1–14, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003112.

Liu:2015:IHC

[LS15c] L. J. Liu and M. Schlesinger. Interstitial hydraulic conductivity and interstitial fluid pressure for avascular or poorly vascularized tumors. Journal of Theoretical Biology, 380(??):1–8, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002404.

Lapytsko:2016:RTD

[LS16a] Anastasiya Lapytsko and Jörg Schaber. The role of time delay in adaptive cellular negative feedback systems. Journal of Theoretical Biology, 398(??):64–73, June 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001491.

Lessard:2016:DFA

[LS16b] Sabin Lessard and Cintia Soares. Definitions of fitness in age-structured populations: Comparison in the haploid case. Journal of Theoretical Biology, 391(??):65–73, February 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315005688.

Liu:2016:MCA

[LS16c] L. J. Liu and M. Schlesinger. MRI contrast agent concentration and tumor interstitial fluid pressure. Journal of Theoretical Biology, 406(??):52–60, October 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301606.

Liu:2018:AMF

[LS18] Longgui Liu and Mao Sun. The added mass forces in insect flapping wings. Journal of Theoretical Biology, 437(??):45–50, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304770.

Liu:2019:DFS

[LS19] Longgui Liu and Mao Sun. Dynamic flight stability of hovering mosquitoes. Journal of Theoretical Biology, 464(??):149–158, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318306441.

Lutkenhoner:2010:MVE

[LSB10] Bernd Lütkenhöner, Wolfgang Stoll, and Türker Basel. Modeling the vestibular evoked myogenic potential. Journal of Theoretical Biology, 263(1):70–78, March 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005256.

Link:2018:MLE

[LSC^{+}18] Roman M. Link, Bernhard Schuldt, Brendan Choat, Steven Jansen, and Alexander R. Cobb. Maximum-likelihood estimation of xylem vessel length distributions. Journal of Theoretical Biology, 455(??):329–341, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303679.

Leimar:2013:LSS

[LSDD13] Olof Leimar, Akira Sasaki, Michael Doebeli, and Ulf Dieckmann. Limiting similarity, species packing, and the shape of competition kernels. Journal of Theoretical Biology, 339(??):3–13, December 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313003779.

Lee:2015:GCH

[LSDI15] Joung-Hun Lee, Karl Sigmund, Ulf Dieckmann, and Yoh Iwasa. Games of corruption: How to suppress illegal logging. Journal of Theoretical Biology, 367(??):1–13, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006274.

Luo:2013:MEP

[LSLL13] Jiawei Luo, Dan Song, Cheng Liang, and Guanghui Li. Model the evolution of protein interaction network assisted with protein age. Journal of Theoretical Biology, 333(??):10–17, September 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300221X.

Leslie:2017:GTF

[LSM17] Martin P. Leslie, Deborah E. Shelton, and Richard E. Michod. Generation time and fitness tradeoffs during the evolution of multicellularity. Journal of Theoretical Biology, 430(??):92–102, October 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303326.

Lakatos:2018:SVS

[LSMC18] Dóra Lakatos, Ellák Somfai, Előd Méhes, and András Czirók. Soluble VEGFR1 signaling guides vascular patterns into dense branching morphologies. Journal of Theoretical Biology, 456(??):261–278, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303758.

Leguebe:2014:CPS

[LSMP14] M. Leguèbe, A. Silve, L. M. Mir, and C. Poignard. Conducting and permeable states of cell membrane submitted to high voltage pulses: Mathematical and numerical studies validated by the experiments. Journal of Theoretical Biology, 360(??):83–94, November 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003701.

Lee:2013:CSV

[LSP13] Pilhwa Lee, Eric A. Sobie, and Charles S. Peskin. Computer simulation of voltage sensitive calcium ion channels in a dendritic spine. Journal of Theoretical Biology, 338(??):87–93, December 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003937.

Laranjeira:2017:MMC

[LSP^{+}17] Simão Laranjeira, Mkael Symmonds, Jacqueline Palace, Stephen J. Payne, and Piotr Orlowski. A mathematical model of cellular swelling in neuromyelitis optica. Journal of Theoretical Biology, 433(??):39–48, November 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303946.

Lopez:2014:AHC

[LSS^{+}14] Álvaro G. López, Juan Sabuco, Jesús M. Seoane, Jorge Duarte, Cristina Januário, and Miguel A. F. Sanjuán. Avoiding healthy cells extinction in a cancer model. Journal of Theoretical Biology, 349(??):74–81, May 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000599.

Lu:2010:SSM

[LSSG10] Zhao Lu, Ynte H. Schukken, Rebecca L. Smith, and Yrjo T. Grohn. Stochastic simulations of a multi-group compartmental model for Johne’s disease on US dairy herds with test-based culling intervention. Journal of Theoretical Biology, 264(4):1190–1201, June 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001670.

Li:2010:MMC

[LSY^{+}10] Bo Li, Bin Shao, Chenlu Yu, Qi Ouyang, and Hongli Wang. A mathematical model for cell size control in fission yeast. Journal of Theoretical Biology, 264(3):771–781, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001463.

Luo:2011:FMD

[LT11] Jamie X. Luo and Matthew S. Turner. Functionality and metagraph disintegration in Boolean networks. Journal of Theoretical Biology, 282(1):65–70, August 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002487.

Lin:2014:CLQ

[LT14] Thy-Hou Lin and Tsung-Lin Tsai. Constructing a linear QSAR for some metabolizable drugs by human or pig flavincontaining monooxygenases using some molecular features selected by a genetic algorithm trained SVM. Journal of Theoretical Biology, 356(??):85–97, September 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002410.

Lague:2012:EFC

[LTHEK12] Marysa Laguë, Nessy Tania, Joel Heath, and Leah Edelstein-Keshet. The effects of facilitation and competition on group foraging in patches. Journal of Theoretical Biology, 310(??):88–96, October 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003116.

Liu:2015:PSL

[LTL^{+}15] Taigang Liu, Peiying Tao, Xiaowei Li, Yufang Qin, and Chunhua Wang. Prediction of subcellular location of apoptosis proteins combining tri-gram encoding based on PSSM and recursive feature elimination. Journal of Theoretical Biology, 366(??):8–12, February 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006535.

Li:2019:QLR

[LTP19] Wentian Li, Dimitrios Thanos, and Astero Provata. Quantifying local randomness in human DNA and RNA sequences using Erdös motifs. Journal of Theoretical Biology, 461(??):41–50, January 14, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304673.

Lin:2016:BMS

[LTS^{+}16] Xiao Lin, Gabriel Terejanu, Sajan Shrestha, Sourav Banerjee, and Anindya Chanda. Bayesian model selection framework for identifying growth patterns in filamentous fungi. Journal of Theoretical Biology, 398(??):85–95, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001624.

Laomettachit:2015:DMH

[LTSTD15] Teeraphan Laomettachit, Teerasit Termsaithong, Anuwat Sae-Tang, and Orawan Duangphakdee. Decision-making in honeybee swarms based on quality and distance information of candidate nest sites. Journal of Theoretical Biology, 364(??):21–30, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005438.

Li:2017:IRD

[LTZ17] Lingling Li, Tianhai Tian, and Xinan Zhang. The impact of radiation on the development of lung cancer. Journal of Theoretical Biology, 428(??):147–152, September 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302904.

Li:2019:SMM

[LTZ19] Lingling Li, Tianhai Tian, and Xinan Zhang. Stochastic modelling of multistage carcinogenesis and progression of human lung cancer. Journal of Theoretical Biology, 479(??):81–89, October 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302838.

Lugli:2015:TBS

[Lug15] Marco Lugli. The tradeoff between signal detection and recognition rules auditory sensitivity under variable background noise conditions. Journal of Theoretical Biology, 386(??):1–6, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004385.

Lukeman:2014:ODC

[Luk14] Ryan Lukeman. Ordering dynamics in collectively swimming surf scoters. Journal of Theoretical Biology, 355(??):151–159, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001490.

Lundberg:2013:ESP

[Lun13] Per Lundberg. On the evolutionary stability of partial migration. Journal of Theoretical Biology, 321(??):36–39, March 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312006534.

Luo:2014:UFR

[Luo14] Shishi Luo. A unifying framework reveals key properties of multilevel selection. Journal of Theoretical Biology, 341(??):41–52, January 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313004542.

Lutkenhoner:2016:ETR

[Lüt16] Bernd Lütkenhöner. Estimation of a transient response from steady-state responses by deconvolution with built-in constraints. Journal of Theoretical Biology, 404(??):143–159, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301229.

Lepage:2017:RBC

[LV17] Kyle Q. Lepage and Sujith Vijayan. The relationship between coherence and the phase-locking value. Journal of Theoretical Biology, 435(??):106–109, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304095.

Lee:2016:SPE

[LvBJ16] William Lee, Minus van Baalen, and Vincent A. A. Jansen. Siderophore production and the evolution of investment in a public good: An adaptive dynamics approach to kin selection. Journal of Theoretical Biology, 388(??):61–71, January 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004889.

LoIacono:2012:EPP

[LvdBP12] Giovanni Lo Iacono, Frank van den Bosch, and Neil Paveley. The evolution of plant pathogens in response to host resistance: Factors affecting the gain from deployment of qualitative and quantitative resistance. Journal of Theoretical Biology, 304(??):152–163, July 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001658.

Lorenzi:2018:RSV

[LVLC18] Tommaso Lorenzi, Chandrasekhar Venkataraman, Alexander Lorz, and Mark A. J. Chaplain. The role of spatial variations of abiotic factors in mediating intratumour phenotypic heterogeneity. Journal of Theoretical Biology, 451(??):101–110, August 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302273.

Liu:2010:PESa

[LW10a] Meng Liu and Ke Wang. Persistence and extinction of a stochastic single-specie model under regime switching in a polluted environment. Journal of Theoretical Biology, 264(3):934–944, June 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001311.

Liu:2010:PESb

[LW10b] Meng Liu and Ke Wang. Persistence and extinction of a stochastic single-specie model under regime switching in a polluted environment II. Journal of Theoretical Biology, 267(3):283–291, December 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004534.

Liu:2012:MMR

[LW12] Kaihui Liu and Ruiqi Wang. MicroRNA-mediated regulation in the mammalian circadian rhythm. Journal of Theoretical Biology, 304(??):103–110, July 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001695.

Li:2015:EDS

[LW15] Aming Li and Long Wang. Evolutionary dynamics of synergistic and discounted group interactions in structured populations. Journal of Theoretical Biology, 377(??):57–65, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315001708.

Leung:2017:MSE

[LW17] Chung Yin (Joey) Leung and Joshua S. Weitz. Modeling the synergistic elimination of bacteria by phage and the innate immune system. Journal of Theoretical Biology, 429(??):241–252, September 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303235.

Ly:2018:AHC

[LW18] Cheng Ly and Seth H. Weinberg. Analysis of heterogeneous cardiac pacemaker tissue models and traveling wave dynamics. Journal of Theoretical Biology, 459(??):18–35, December 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304594.

Lenk:2015:IAF

[LWB^{+}15] Claudia Lenk, Frank M. Weber, Martin Bauer, Mario Einax, Philipp Maass, and Gunnar Seeman. Initiation of atrial fibrillation by interaction of pacemakers with geometrical constraints. Journal of Theoretical Biology, 366(??):13–23, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006201.

Larsson:2011:MCL

[LWH^{+}11] Jan-Åke Larsson, Niclas Wadströmer, Ola Hermanson, Urban Lendahl, and Robert Forchheimer. Modelling cell lineage using a meta-Boolean tree model with a relation to gene regulatory networks. Journal of Theoretical Biology, 268(1):62–76, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931000531X.

Liu:2011:NPS

[LWL^{+}11] Xuemei Liu, Lin Wan, Jing Li, Gesine Reinert, Michael S. Waterman, and Fengzhu Sun. New powerful statistics for alignment-free sequence comparison under a pattern transfer model. Journal of Theoretical Biology, 284(1):106–116, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003195.

Lempidakis:2018:WCK

[LWLM18] Emmanouil Lempidakis, Rory P. Wilson, Adrian Luckman, and Richard S. Metcalfe. What can knowledge of the energy landscape tell us about animal movement trajectories and space use? A case study with humans. Journal of Theoretical Biology, 457(??):101–111, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830403X.

Latif:2014:MIR

[LWRE14] Nurul S. Abdul Latif, Graeme C. Wake, Tony Reglinski, and Philip A. G. Elmer. Modelling induced resistance to plant diseases. Journal of Theoretical Biology, 347(??):144–150, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005742.

Lynch:2019:MTF

[LWvB^{+}19] T. A. Lynch, Y. Wang, B. van Brunt, D. Pacheco, and P. H. Janssen. Modelling thermodynamic feedback on the metabolism of hydrogenotrophic methanogens. Journal of Theoretical Biology, 477(??):14–23, September 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302164.

Li:2014:RFG

[LWW14] Xiaomei Li, Xindong Wu, and Gongqing Wu. Robust feature generation for protein subchloroplast location prediction with a weighted GO transfer model. Journal of Theoretical Biology, 347(??):84–94, April 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000125.

Li:2016:DRE

[LWY^{+}16] Chu-Chu Li, Jia-Ning Wu, Yun-Qiang Yang, Ren-Gao Zhu, and Shao-Ze Yan. Drag reduction effects facilitated by microridges inside the mouthparts of honeybee workers and drones. Journal of Theoretical Biology, 389(??):1–10, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005032.

Li:2010:SHE

[LX10] Xiaohai Li and Jizhong Xiao. Swarming in homogeneous environments: a social interaction based framework. Journal of Theoretical Biology, 264(3):747–759, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000913.

Li:2016:MPM

[LX16] Xuefang Li and Jian-Xin Xu. A mathematical prognosis model for pancreatic cancer patients receiving immunotherapy. Journal of Theoretical Biology, 406(??):42–51, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301552.

Liu:2015:UWF

[LXC15] Guoqing Liu, Yongqiang Xing, and Lu Cai. Using weighted features to predict recombination hotspots in Saccharomyces cerevisiae. Journal of Theoretical Biology, 382(??):15–22, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003094.

Liu:2015:TEM

[LXJ15] Xiaofeng Liu, Ning Xu, and Aimin Jiang. Tortuosity entropy: a measure of spatial complexity of behavioral changes in animal movement. Journal of Theoretical Biology, 364(??):197–205, January 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005645.

Louzoun:2014:MMP

[LXLF14] Yoram Louzoun, Chuan Xue, Gregory B. Lesinski, and Avner Friedman. A mathematical model for pancreatic cancer growth and treatments. Journal of Theoretical Biology, 351(??):74–82, June 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001039.

Li:2015:STG

[LXS15] Rui Li, Liufang Xu, and Hualin Shi. Strategy of tuning gene expression ratio in prokaryotic cell from perspective of noise and correlation. Journal of Theoretical Biology, 365(??):377–389, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006341.

Li:2014:MPI

[LXZQ14] Long-Fei Li, Cheng Xiang, Ya-Bing Zhu, and Kai-Rong Qin. Modeling of progesterone-induced intracellular calcium signaling in human spermatozoa. Journal of Theoretical Biology, 351(??):58–66, June 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001015.

Li:2011:SSE

[LY11] Jianquan Li and Yali Yang. SIR-SVS epidemic models with continuous and impulsive vaccination strategies. Journal of Theoretical Biology, 280(1):108–116, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001536.

Li:2014:SAS

[LY14] Sen Li and Jie Yang. System analysis of synonymous codon usage biases in archaeal virus genomes. Journal of Theoretical Biology, 355(??):128–139, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001611.

Li:2015:PSP

[LYF^{+}15] Chun Li, Yan Yang, Wenchao Fei, Ping an He, Xiaoqing Yu, Defu Zhang, Shumin Yi, Xuepeng Li, Jin Zhu, Changzhong Wang, and Zhifu Wang. Prediction of success for polymerase chain reactions using the Markov maximal order model and support vector machine. Journal of Theoretical Biology, 369(??):51–58, March 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000260.

Lev-Yadun:2012:MIF

[LYK12] Simcha Lev-Yadun and Gadi Katzir. Mirror images: Fish versus terrestrial animals. Journal of Theoretical Biology, 294(??):182–184, February 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005388.

Li:2019:SBP

[LYM^{+}19] Zhongyan Li, Fugang Yan, Qingqing Miao, Yang Meng, Li Wen, Qianhu Jiang, and Peng Zhou. Self-binding peptides: Binding-upon-folding versus folding-upon-binding. Journal of Theoretical Biology, 469(??):25–34, May 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300839.

Liu:2011:INH

[LYW11] Bo Liu, Shiwei Yan, and Qi Wang. Intrinsic noise and Hill dynamics in the p53 system. Journal of Theoretical Biology, 269(1):104–108, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005370.

Liu:2018:IPI

[LYZ^{+}18] Xiaoxia Liu, Zhihao Yang, Ziwei Zhou, Yuanyuan Sun, Hongfei Lin, Jian Wang, and Bo Xu. The impact of protein interaction networks’ characteristics on computational complex detection methods. Journal of Theoretical Biology, 439(??):141–151, February 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305362.

Liang:2018:IGN

[LZ18] Yunyun Liang and Shengli Zhang. Identify Gram-negative bacterial secreted protein types by incorporating different modes of PSSM into Chou’s general PseAAC via Kullback–Leibler divergence. Journal of Theoretical Biology, 454(??):22–29, October 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302820.

Larson:2019:DDS

[LZA^{+}19] Karen Larson, Loukas Zagkos, Mark Mc Auley, Jason Roberts, Nikos I. Kavallaris, and Anastasios Matzavinos. Data-driven selection and parameter estimation for DNA methylation mathematical models. Journal of Theoretical Biology, 467(??):87–99, April 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300128.

Li:2019:EAC

[LZG^{+}19] Weidong Li, Qingniao Zhou, Yong Gao, Yonghua Jiang, Yuanjie Huang, Zengnan Mo, Yiming Zou, and Yanling Hu. eQTL analysis from co-localization of 2739 GWAS loci detects associated genes across 14 human cancers. Journal of Theoretical Biology, 462(??):240–246, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305472.

Li:2017:NMS

[LZL^{+}17] Rong Li, Dexing Zhong, Ruiling Liu, Hongqiang Lv, Xinman Zhang, Jun Liu, and Jiuqiang Han. A novel method for in silico identification of regulatory SNPs in human genome. Journal of Theoretical Biology, 415(??):84–89, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316304039.

Lu:2019:PEE

[LZL^{+}19] Fuhua Lu, Maoshu Zhu, Ying Lin, Hongbin Zhong, Lei Cai, Lin He, and Kuo-Chen Chou. The preliminary efficacy evaluation of the CTLA-4-Ig treatment against Lupus nephritis through in-silico analyses. Journal of Theoretical Biology, 471(??):74–81, June 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301249.

Lotfi:2015:RSS

[LZMM15] Mohadeseh Lotfi, Fatemeh Zare-Mirakabad, and Soheila Montaseri. RNA secondary structure prediction based on SHAPE data in helix regions. Journal of Theoretical Biology, 380(??):178–182, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002611.

Liu:2012:SIM

[LZT12] Zijian Liu, Shouming Zhong, and Zhidong Teng. n species impulsive migration model with Markovian switching. Journal of Theoretical Biology, 307(??):62–69, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002263. See erratum [LZT12].

Liu:2013:ESI

[LZT13] Zijian Liu, Shouming Zhong, and Zhidong Teng. Erratum to “n Species impulsive migration model with Markovian switching”: [J. Theor. Biol. 307 (2012) 62–69]. Journal of Theoretical Biology, 318(??):226, February 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006169. See [LZT12].

Luo:2018:PMP

[LZTD18] Yantao Luo, Long Zhang, Zhidong Teng, and Donald L. DeAngelis. A parasitism-mutualism-predation model consisting of crows, cuckoos and cats with stage-structure and maturation delays on crows and cuckoos. Journal of Theoretical Biology, 446(??):212–228, June 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300882.

Liu:2011:LIC

[LZWK11] Y. Liu, W. Zhang, C. Wang, and G. S. Kassab. A linearized and incompressible constitutive model for arteries. Journal of Theoretical Biology, 286(??):85–91, October 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002396.

Liang:2015:IIN

[LZY^{+}15] Peifeng Liang, Jian Zu, Juan Yin, Hao Li, Longfei Gao, Fuqiang Cui, Fuzhen Wang, Xiaofeng Liang, and Guihua Zhuang. The independent impact of newborn hepatitis B vaccination on reducing HBV prevalence in China, 1992– 2006: a mathematical model analysis. Journal of Theoretical Biology, 386(??):115–121, December 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004348.

Morozov:2010:TCD

[MA10] A. Yu. Morozov and E. G. Arashkevich. Towards a correct description of zooplankton feeding in models: Taking into account food-mediated unsynchronized vertical migration. Journal of Theoretical Biology, 262(2):346–360, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004536.

Mehrabian:2011:GSP

[MA11a] Amin Mehrabian and Younane Abousleiman. General solutions to poroviscoelastic model of hydrocephalic human brain tissue. Journal of Theoretical Biology, 291(??):105–118, December 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100467X.

Morozov:2011:EVD

[MA11b] A. Yu. Morozov and M. W. Adamson. Evolution of virulence driven by predator–prey interaction: Possible consequences for population dynamics. Journal of Theoretical Biology, 276(1):181–191, May 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001019.

Matsuda:2013:FCE

[MA13] Hiroyuki Matsuda and Peter A. Abrams. Is feedback control effective for ecosystem-based fisheries management? Journal of Theoretical Biology, 339(??):122–128, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002774.

Mondal:2017:CSA

[MAA^{+}17] Abhijit Mondal, Daniel A. Appadurai, Nazem W. Akoum, Frank B. Sachse, and Alonso P. Moreno. Computational simulations of asymmetric fluxes of large molecules through gap junction channel pores. Journal of Theoretical Biology, 412(??):61–73, January 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302818.

Moore:2018:HDR

[MAA18] James Moore, Hasan Ahmed, and Rustom Antia. High dimensional random walks can appear low dimensional: Application to influenza H3N2 evolution. Journal of Theoretical Biology, 447(??):56–64, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830136X.

Maciejewski:2010:AOO

[Mac10] Wes Maciejewski. An analysis of the orientation of an orbweb spider. Journal of Theoretical Biology, 265(4):604–608, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002687.

MacKay:2011:MSC

[Mac11] Niall J. MacKay. Mass scale and curvature in metabolic scaling: Comment on: T. Kolokotrones et al., Curvature in metabolic scaling, Nature 464 (2010) 753–756. Journal of Theoretical Biology, 280(1):194–196, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001056.

Maciejewski:2014:RVG

[Mac14] Wes Maciejewski. Reproductive value in graph-structured populations. Journal of Theoretical Biology, 340(??):285–293, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004785.

Muolo:2019:PNN

[MAF^{+}19] Riccardo Muolo, Malbor Asllani, Duccio Fanelli, Philip K. Maini, and Timoteo Carletti. Patterns of non-normality in networked systems. Journal of Theoretical Biology, 480 (??):81–91, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519319302814.

Munir:2018:EWB

[MAFB18] Anum Munir, Shumaila Azam, Sahar Fazal, and A. I. Bhatti. Evaluation of the whole body physiologically based pharmacokinetic (WB-PBPK) modeling of drugs. Journal of Theoretical Biology, 451(??):1–9, August 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302091.

McKinley:2012:AAM

[MAFK12] Scott A. McKinley, Avanti Athreya, John Fricks, and Peter R. Kramer. Asymptotic analysis of microtubule-based transport by multiple identical molecular motors. Journal of Theoretical Biology, 305(??):54–69, July 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001671.

Murtada:2012:EMM

[MAH12] Saeil C. Murtada, Anders Arner, and Gerhard A. Holzapfel. Experiments and mechanochemical modeling of smooth muscle contraction: Significance of filament overlap. Journal of Theoretical Biology, 297(??):176–186, March 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005741.

Mallik:2010:HTM

[Mal10] Mrinmay Kumar Mallik. A hypothesis and theoretical model speculating the possible role of therapy mediated neoplastic cell loss in promoting the process of glioblastoma relapse. Journal of Theoretical Biology, 266(4):496–503, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003589.

Muneepeerakul:2011:EDE

[MAL^{+}11] Rachata Muneepeerakul, Sandro Azaele, Simon A. Levin, Andrea Rinaldo, and Ignacio Rodriguez-Iturbe. Evolution of dispersal in explicitly spatial metacommunities. Journal of Theoretical Biology, 269(1):256–265, January 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005886.

Malanson:2015:DDA

[Mal15] George P. Malanson. Diversity differs among three variations of the stress gradients hypothesis in two representations of niche space. Journal of Theoretical Biology, 384(??):121–130, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004051.

Malanson:2017:ICM

[Mal17] George P. Malanson. Interactions and constraints in model species response to environmental heteroscedasticity. Journal of Theoretical Biology, 419(??):343–349, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300784.

Mallik:2018:AUG

[Mal18] Mrinmay Kumar Mallik. An attempt to understand glioma stem cell biology through centrality analysis of a protein interaction network. Journal of Theoretical Biology, 438(??):78–91, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305027.

Mier:2017:PPI

[MALAN17] Pablo Mier, Gregorio Alanis-Lobato, and Miguel A. Andrade-Navarro. Protein-protein interactions can be predicted using coiled coil co-evolution patterns. Journal of Theoretical Biology, 412(??):198–203, January 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303599.

Mazzuca:2016:CMS

[MAM16] Enrico Mazzuca, Andrea Aliverti, and Giuseppe Miserocchi. Computational micro-scale model of control of extravascular water and capillary perfusion in the air blood barrier. Journal of Theoretical Biology, 400(??):42–51, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300121.

Mehrabian:2015:DPP

[MAMEA15] Amin Mehrabian, Younane N. Abousleiman, Timothy B. Mapstone, and Christian A. El-Amm. Dual-porosity poroviscoelasticity and quantitative hydromechanical characterization of the brain tissue with experimental hydrocephalus data. Journal of Theoretical Biology, 384(??):19–32, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003793.

Marshall:2011:QRO

[Mar11] James A. R. Marshall. Queller’s rule ok: Comment on van Veelen ‘When inclusive fitness is right and when it can be wrong’. Journal of Theoretical Biology, 270(1):185–188, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310003644.

Martineau:2012:STM

[Mar12] Louis C. Martineau. Simple thermodynamic model of unassisted proton shuttle uncoupling and prediction of activity from calculated speciation, lipophilicity, and molecular geometry. Journal of Theoretical Biology, 303(??):33–61, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312001178.

Martin:2017:PAO

[Mar17a] William F. Martin. Physiology, anaerobes, and the origin of mitosing cells 50 years on. Journal of Theoretical Biology, 434(??):2–10, December 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300048.

Martini:2017:MQD

[Mar17b] Johannes W. R. Martini. A measure to quantify the degree of cooperativity in overall titration curves. Journal of Theoretical Biology, 432(??):33–37, November 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303764.

Mokhtar:2017:EDS

[MAR^{+}17c] N. Hafizah Mokhtar, Aizat Abas, N. A. Razak, Muhammad Najib Abdul Hamid, and Soon Lay Teong. Effect of different stent configurations using lattice Boltzmann method and particles image velocimetry on artery bifurcation aneurysm problem. Journal of Theoretical Biology, 433(??):73–84, November 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730382X.

Marzen:2019:NDI

[Mar19] Sarah E. Marzen. Novelty detection improves performance of reinforcement learners in fluctuating, partially observable environments. Journal of Theoretical Biology, 477(??):44–50, September 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302346.

Masuda:2012:IFI

[Mas12] Naoki Masuda. Ingroup favoritism and intergroup cooperation under indirect reciprocity based on group reputation. Journal of Theoretical Biology, 311(??):8–18, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312003360.

Masuda:2014:EIA

[Mas14] Naoki Masuda. Evolution via imitation among like-minded individuals. Journal of Theoretical Biology, 349(??):100–108, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000642.

Morse:2015:PMM

[MAS15a] R. P. Morse, D. Allingham, and N. G. Stocks. A phenomenological model of myelinated nerve with a dynamic threshold. Journal of Theoretical Biology, 382(??):386–396, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003148.

Morse:2015:SDR

[MAS15b] R. P. Morse, D. Allingham, and N. G. Stocks. Stimulusdependent refractoriness in the Frankenhaeuser–Huxley model. Journal of Theoretical Biology, 382(??):397–404, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003318.

Massey:2016:NEE

[Mas16] Steven E. Massey. The neutral emergence of error minimized genetic codes superior to the standard genetic code. Journal of Theoretical Biology, 408(??):237–242, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302569.

Masso:2018:AAF

[Mas18] Majid Masso. All-atom four-body knowledge-based statistical potential to distinguish native tertiary RNA structures from nonnative folds. Journal of Theoretical Biology, 453(??):58–67, September 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302601.

Maury:2015:OLP

[Mau15] Carl Peter J. Maury. Origin of life. Primordial genetics: Information transfer in a pre-RNA world based on selfreplicating beta-sheet amyloid conformers. Journal of Theoretical Biology, 382(??):292–297, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003379.

McMeeking:2012:CDA

[MAW12] Robert M. McMeeking, Eduard Arzt, and Rüdiger Wehner. Cataglyphis desert ants improve their mobility by raising the gaster. Journal of Theoretical Biology, 297(??):17–25, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311006047.

Maxin:2010:TSD

[MB10] D. Maxin and L. Berec. A two-sex demographic model with single-dependent divorce rate. Journal of Theoretical Biology, 265(4):647–656, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002973.

Marcu:2012:NCH

[MB12a] Loredana G. Marcu and Eva Bezak. Neoadjuvant cisplatin for head and neck cancer: Simulation of a novel schedule for improved therapeutic ratio. Journal of Theoretical Biology, 297(??):41–47, March 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006011.

Morozov:2012:PIH

[MB12b] Andrew Morozov and Alex Best. Predation on infected host promotes evolutionary branching of virulence and pathogens’ biodiversity. Journal of Theoretical Biology, 307(??):29–36, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002044.

Marechal:2014:MRL

[MB14] Eric Maréchal and Olivier Bastien. Modeling of regulatory loops controlling galactolipid biosynthesis in the inner envelope membrane of chloroplasts. Journal of Theoretical Biology, 361(??):1–13, November 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400410X.

McQuaid:2015:PSR

[MB15] C. Finn McQuaid and Nicholas F. Britton. Parasite species richness and its effect on persistence in food webs. Journal of Theoretical Biology, 364(??):377–382, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005827.

Morsky:2016:TSP

[MB16] B. Morsky and C. T. Bauch. Truncation selection and payoff distributions applied to the replicator equation. Journal of Theoretical Biology, 404(??):383–390, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301540.

Maleki:2017:OLA

[MB17] Farzaneh Maleki and Attila Becskei. An open-loop approach to calculate noise-induced transitions. Journal of Theoretical Biology, 415(??):145–157, February 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304222.

Marrec:2018:QIP

[MB18a] Loïc Marrec and Anne-Florence Bitbol. Quantifying the impact of a periodic presence of antimicrobial on resistance evolution in a homogeneous microbial population of fixed size. Journal of Theoretical Biology, 457(??):190–198, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304284.

McKenna:2018:FSA

[MB18b] Joseph P. McKenna and Richard Bertram. Fast-slow analysis of the integrated oscillator model for pancreatic β-cells. Journal of Theoretical Biology, 457(??):152–162, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304089.

Murase:2018:SRA

[MB18c] Yohsuke Murase and Seung Ki Baek. Seven rules to avoid the tragedy of the commons. Journal of Theoretical Biology, 449(??):94–102, July 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301954.

Morsky:2019:ITS

[MB19] Bryce Morsky and Chris T. Bauch. The impact of truncation selection and diffusion on cooperation in spatial games. Journal of Theoretical Biology, 466(??):64–83, April 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300359.

McAvity:2013:PSM

[MBBD13] David M. McAvity, Tristen Bristow, Eric Bunker, and Alex Dreyer. Perception without self-matching in conditional tag based cooperation. Journal of Theoretical Biology, 333(??):58–67, September 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001951.

Mathew:2014:GSA

[MBBV14] Shibin Mathew, John Bartels, Ipsita Banerjee, and Yoram Vodovotz. Global sensitivity analysis of a mathematical model of acute inflammation identifies nonlinear dependence of cumulative tissue damage on host interleukin-6 responses. Journal of Theoretical Biology, 358(??):132–148, October 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003191.

Maxin:2012:ISA

[MBC^{+}12a] D. Maxin, L. Berec, M. Covello, J. Jessee, and M. Zimmer. The impact of sexually abstaining groups on persistence of sexually transmitted infections in populations with ephemeral pair bonds. Journal of Theoretical Biology, 292(??):1–10, January 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004875.

Miller:2012:DPA

[MBC^{+}12b] James K. Miller, Hope T. Badawy, Curtis Clemons, K. L. Kreider, Pat Wilber, Amy Milsted, and Gerald Young. Development of the Pseudomonas aeruginosa mushroom morphology and cavity formation by iron-starvation: a mathematical modeling study. Journal of Theoretical Biology, 308(??):68–78, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002718.

Moller:2012:DRS

[MBCM12] Jan Kloppenborg Møller, Kirsten Riber Bergmann, Lasse Engbo Christiansen, and Henrik Madsen. Development of a restricted state space stochastic differential equation model for bacterial growth in rich media. Journal of Theoretical Biology, 305(??):78–87, July 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001968.

Malkoc:2010:ICB

[MBE10] Berkin Malkoç, Duygu Balcan, and Ayse Erzan. Information content based model for the topological properties of the gene regulatory network of Escherichia coli. Journal of Theoretical Biology, 263(3):281–294, April 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005591.

Mohabatkar:2011:PGR

[MBE11] Hassan Mohabatkar, Majid Mohammad Beigi, and Abolghasem Esmaeili. Prediction of GABA a receptor proteins using the concept of Chou’s pseudo-amino acid composition and support vector machine. Journal of Theoretical Biology, 281(1):18–23, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002177.

Menshikov:2015:INR

[MBF^{+}15] Igor Menshikov, Liubov Beduleva, Maksim Frolov, Nadezhda Abisheva, Tatyana Khramova, Elena Stolyarova, and Kseniya Fomina. The idiotypic network in the regulation of autoimmunity: Theoretical and experimental studies. Journal of Theoretical Biology, 375(??):32–39, June 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400592X.

Muraro:2011:ICA

[MBK^{+}11] Daniele Muraro, Helen Byrne, John King, Ute Voß, Joseph Kieber, and Malcolm Bennett. The influence of cytokininauxin cross-regulation on cell-fate determination in Arabidopsis thaliana root development. Journal of Theoretical Biology, 283(1):152–167, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002530.

Muraro:2013:RAC

[MBKB13] Daniele Muraro, Helen Byrne, John King, and Malcolm Bennett. The role of auxin and cytokinin signalling in specifying the root architecture of Arabidopsis thaliana. Journal of Theoretical Biology, 317(??):71–86, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004547.

Mesgaran:2017:HGN

[MBLC17] Mohsen B. Mesgaran, Juliette Bouhours, Mark A. Lewis, and Roger D. Cousens. How to be a good neighbour: Facilitation and competition between two co-flowering species. Journal of Theoretical Biology, 422(??):72–83, June 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301662.

Markert:2010:HOB

[MBLV10] Elke K. Markert, Nils Baas, Arnold J. Levine, and Alexei Vazquez. Higher order Boolean networks as models of cell state dynamics. Journal of Theoretical Biology, 264(3):945– 951, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001384.

MacNamara:2011:DGU

[MBM11] Shev MacNamara, Ruth E. Baker, and Philip K. Maini. Distinguishing graded and ultrasensitive signalling cascade kinetics by the shape of morphogen gradients in Drosophila. Journal of Theoretical Biology, 285(1):136–146, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311003110.

Morozov:2016:LTT

[MBP16] A. Yu. Morozov, M. Banerjee, and S. V. Petrovskii. Longterm transients and complex dynamics of a stage-structured population with time delay and the Allee effect. Journal of Theoretical Biology, 396(??):116–124, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001120.

Mascheroni:2017:EIM

[MBPS17] Pietro Mascheroni, Daniela Boso, Luigi Preziosi, and Bernhard A. Schrefler. Evaluating the influence of mechanical stress on anticancer treatments through a multiphase porous media model. Journal of Theoretical Biology, 421(??):179–188, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301467.

Muneepeerakul:2019:EBP

[MBRRI19] Rachata Muneepeerakul, Enrico Bertuzzo, Andrea Rinaldo, and Ignacio Rodriguez-Iturbe. Evolving biodiversity patterns in changing river networks. Journal of Theoretical Biology, 462(??):418–424, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305757.

Matsiaka:2019:CDS

[MBS19] Oleksii M. Matsiaka, Ruth E. Baker, and Matthew J. Simpson. Continuum descriptions of spatial spreading for heterogeneous cell populations: Theory and experiment. Journal of Theoretical Biology, 482(??):Article 109997, December 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303479.

Massahi:2015:SDP

[MÇ15a] Aslan Massahi and Pinar Çalik. In-silico determination of Pichia pastoris signal peptides for extracellular recombinant protein production. Journal of Theoretical Biology, 364(??):179–188, January 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314005359.

Mulder:2015:DIC

[MC15b] Willem H. Mulder and Forrest W. Crawford. On the distribution of interspecies correlation for Markov models of character evolution on Yule trees. Journal of Theoretical Biology, 364(??):275–283, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005554.

Macnamara:2016:DDO

[MC16a] Cicely K. Macnamara and Mark A. J. Chaplain. Diffusion driven oscillations in gene regulatory networks. Journal of Theoretical Biology, 407(??):51–70, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302041.

Massahi:2016:ESP

[MÇ16b] Aslan Massahi and Pinar Çalik. Endogenous signal peptides in recombinant protein production by Pichia pastoris: From in-silico analysis to fermentation. Journal of Theoretical Biology, 408(??):22–33, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302223.

Mason:2019:ERK

[MC19] John C. Mason and Markus W. Covert. An energetic reformulation of kinetic rate laws enables scalable parameter estimation for biochemical networks. Journal of Theoretical Biology, 461(??):145–156, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830523X.

Mari:2017:HST

[MCC^{+}17] Lorenzo Mari, Manuela Ciddio, Renato Casagrandi, Javier Perez-Saez, Enrico Bertuzzo, Andrea Rinaldo, Susanne H. Sokolow, Giulio A. De Leo, and Marino Gatto. Heterogeneity in schistosomiasis transmission dynamics. Journal of Theoretical Biology, 432(??):87–99, November 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303818.

Mubayi:2010:TDU

[MCCC^{+}10] Anuj Mubayi, Carlos Castillo-Chavez, Gerardo Chowell, Christopher Kribs-Zaleta, Niyamat Ali Siddiqui, Narendra Kumar, and Pradeep Das. Transmission dynamics and underreporting of Kala-azar in the Indian state of Bihar. Journal of Theoretical Biology, 262(1):177–185, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004329.

Mettikolla:2011:CBK

[MCL^{+}11] P. Mettikolla, N. Calander, R. Luchowski, I. Gryczynski, Z. Gryczynski, J. Zhao, D. Szczesna-Cordary, and J. Borejdo. Cross-bridge kinetics in myofibrils containing familial hypertrophic cardiomyopathy r58q mutation in the regulatory light chain of myosin. Journal of Theoretical Biology, 284(1):71–81, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003134.

Macfarlane:2019:SIB

[MCL19] FR Macfarlane, MAJ Chaplain, and T. Lorenzi. A stochastic individual-based model to explore the role of spatial interactions and antigen recognition in the immune response against solid tumours. Journal of Theoretical Biology, 480(??):43–55, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303042.

Manapat:2010:BRR

[MCN10] Michael L. Manapat, Irene A. Chen, and Martin A. Nowak. The basic reproductive ratio of life. Journal of Theoretical Biology, 263(3):317–327, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005931.

Mendez:2012:DDD

[MCPF12] Vicenç Méndez, Daniel Campos, Ignacio Pagonabarraga, and Sergei Fedotov. Density-dependent dispersal and population aggregation patterns. Journal of Theoretical Biology, 309(??):113–120, September 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003025.

Mari:2018:ETW

[MCRG18] Lorenzo Mari, Renato Casagrandi, Andrea Rinaldo, and Marino Gatto. Epidemicity thresholds for water-borne and water-related diseases. Journal of Theoretical Biology, 447(??):126–138, June 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301383.

McGuinness:2014:MSO

[MCWF14] M. J. McGuinness, L. B. Cribbin, H. F. Winstanley, and A. C. Fowler. Modelling spatial oscillations in soil borehole bacteria. Journal of Theoretical Biology, 363(??):74–79, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400472X.

Monteiro:2016:IBP

[MD16] Angelo B. Monteiro and Lucas Del Bianco Faria. The interplay between population stability and food-web topology predicts the occurrence of motifs in complex food-webs. Journal of Theoretical Biology, 409(??):165–171, November 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302946.

Marashi:2012:FCA

[MDB12a] Sayed-Amir Marashi, Laszlo David, and Alexander Bockmayr. On flux coupling analysis of metabolic subsystems. Journal of Theoretical Biology, 302(??):62–69, June 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000963.

Molacek:2012:FAS

[MDB12b] Jan Molácek, Mark Denny, and John W. M. Bush. The fine art of surfacing: Its efficacy in broadcast spawning. Journal of Theoretical Biology, 294(??):40–47, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005285.

Mousavi:2013:CMC

[MDD13] Seyed Jamaleddin Mousavi, Mohamed Hamdy Doweidar, and Manuel Doblaré. 3d computational modelling of cell migration: a mechano-chemo-thermo-electrotaxis approach. Journal of Theoretical Biology, 329(??):64–73, July 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001379.

Ma:2011:ASM

[MDE11] Junling Ma, Jonathan Dushoff, and David J. D. Earn. Agespecific mortality risk from pandemic influenza. Journal of Theoretical Biology, 288(??):29–34, November 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003894.

Mata:2013:MIA

[MDEKH13] May Anne Mata, Meghan Dutot, Leah Edelstein-Keshet, and William R. Holmes. A model for intracellular actin waves explored by nonlinear local perturbation analysis. Journal of Theoretical Biology, 334(??):149–161, October 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002920.

Marques:2010:CCD

[MdFDM10] José Rui Ferreira Marques, Rute R. da Fonseca, Brett Drury, and André Melo. Conformational characterization of disulfide bonds: a tool for protein classification. Journal of Theoretical Biology, 267(3):388–395, December 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004789.

Miller:2010:EGG

[MDM^{+}10] Joel C. Miller, Bahman Davoudi, Rafael Meza, Anja C. Slim, and Babak Pourbohloul. Epidemics with general generation interval distributions. Journal of Theoretical Biology, 262(1):107–115, January 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309003749.

Mazzucco:2016:EEE

[MDM16] Rupert Mazzucco, Ulf Dieckmann, and Johan A. J. Metz. Epidemiological, evolutionary, and economic determinants of eradication tails. Journal of Theoretical Biology, 405 (??):58–65, September 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316001600.

Mbah:2014:CAS

[MDMG14] Martial L. Ndeffo Mbah, David P. Durham, Jan Medlock, and Alison P. Galvani. Country- and age-specific optimal allocation of dengue vaccines. Journal of Theoretical Biology, 342(??):15–22, February 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004980.

Mangalam:2016:SOL

[MDS16] Madhur Mangalam, Nisarg Desai, and Mewa Singh. Selforganization of laterally asymmetrical movements as a consequence of space-time optimization. Journal of Theoretical Biology, 390(??):50–60, February 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005482.

Miranda:2016:TKO

[MdSPBL16] Pedro J. Miranda, Sandro E. de S. Pinto, Murilo S. Baptista, and Giuliano G. La Guardia. Theoretical knockouts on biological networks. Journal of Theoretical Biology, 403(??):38–44, August 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301059.

Marchettini:2010:WMW

[MDVT10] Nadia Marchettini, Emilio Del Giudice, Vladimir Voeikov, and Enzo Tiezzi. Water: a medium where dissipative structures are produced by a coherent dynamics. Journal of Theoretical Biology, 265(4):511–516, August 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002560.

Macnamara:2015:MVE

[ME15] Cicely Macnamara and Raluca Eftimie. Memory versus effector immune responses in oncolytic virotherapies. Journal of Theoretical Biology, 377(??):1–9, July 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001587.

Miller:2012:RMA

[MED12] Callie J. Miller, G. Bard Ermentrout, and Lance A. Davidson. Rotational model for actin filament alignment by myosin. Journal of Theoretical Biology, 300(??):344–359, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000562.

Mehri:2017:NED

[Meh17] Ali Mehri. Non-extensive distribution of human eye photoreceptors. Journal of Theoretical Biology, 419(??):305–309, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300930.

Mei:2012:MKT

[Mei12a] Suyu Mei. Multi-kernel transfer learning based on Chou’s PseAAC formulation for protein submitochondria localization. Journal of Theoretical Biology, 293(??):121–130, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005303. See corri-

gendum [Mei13a].

Mei:2012:PPP

[Mei12b] Suyu Mei. Predicting plant protein subcellular multilocalization by Chou’s PseAAC formulation based multilabel homolog knowledge transfer learning. Journal of Theoretical Biology, 310(??):80–87, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519312003153. See corrigendum [Mei13b].

Mei:2013:CMK

[Mei13a] Suyu Mei. Corrigendum to “Multi-kernel transfer learning based on Chou’s PseAAC formulation for protein submitochondria localization” [J. Theor. Biol. 293 (2012) 121– 130]. Journal of Theoretical Biology, 338(??):112, December 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313004438. See [Mei12a].

Mei:2013:CPP

[Mei13b] Suyu Mei. Corrigendum to “Predicting plant protein subcellular multi-localization by Chou’s PseAAC formulation based multi-label homolog knowledge transfer learning” [J. Theor. Biol. 310 (2012) 80–87]. Journal of Theoretical Biology, 338(??):111, December 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004426. See [Mei12b].

Mei:2014:SEB

[Mei14] Suyu Mei. SVM ensemble based transfer learning for largescale membrane proteins discrimination. Journal of Theoretical Biology, 340(??):105–110, January 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004281.

Moreh:2018:APP

[MEJ18] Ágnes Móréh, Anett Endrédi, and Ferenc Jordán. Additivity of pairwise perturbations in food webs: Topological effects. Journal of Theoretical Biology, 448(??):112–121, July 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301693.

Marino:2011:HMC

[MEKK11] Simeone Marino, Mohammed El-Kebir, and Denise Kirschner. A hybrid multi-compartment model of granuloma formation and T cell priming in tuberculosis. Journal of Theoretical Biology, 280(1):50–62, July 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001688.

Mengel:2012:ECC

[Men12] Friederike Mengel. On the evolution of coarse categories. Journal of Theoretical Biology, 307(??):117–124, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312002585.

Macklin:2012:PCA

[METC12] Paul Macklin, Mary E. Edgerton, Alastair M. Thompson, and Vittorio Cristini. Patient-calibrated agent-based modelling of ductal carcinoma in situ (DCIS): From microscopic measurements to macroscopic predictions of clinical progression. Journal of Theoretical Biology, 301(??):122–140, May 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000665.

Mochizuki:2015:SCR

[MF15] Atsushi Mochizuki and Bernold Fiedler. Sensitivity of chemical reaction networks: a structural approach. 1. Examples and the carbon metabolic network. Journal of Theoretical Biology, 367(??):189–202, February 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006158.

Moise:2019:RAM

[MF19] Nicolae Moise and Avner Friedman. Rheumatoid arthritis — a mathematical model. Journal of Theoretical Biology, 461(??):17–33, January 14, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305216.

McCaig:2013:UPA

[MFG^{+}13] Chris McCaig, Andy Fenton, Andrea Graham, Carron Shankland, and Rachel Norman. Using process algebra to develop predator–prey models of within-host parasite dynamics. Journal of Theoretical Biology, 329(??):74–81, July 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000933.

Middleton:2014:CAL

[MFG14] Alistair M. Middleton, Christian Fleck, and Ramon Grima. A continuum approximation to an off-lattice individual-cell based model of cell migration and adhesion. Journal of Theoretical Biology, 359(??):220–232, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003543.

Mochizuki:2013:DCF

[MFKS13] Atsushi Mochizuki, Bernold Fiedler, Gen Kurosawa, and Daisuke Saito. Dynamics and control at feedback vertex sets. II: a faithful monitor to determine the diversity of molecular activities in regulatory networks. Journal of Theoretical Biology, 335(??):130–146, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002816.

Michel:2018:MMP

[MFL^{+}18] T. Michel, J. Fehrenbach, V. Lobjois, J. Laurent, A. Gomes, T. Colin, and C. Poignard. Mathematical modeling of the proliferation gradient in multicellular tumor spheroids. Journal of Theoretical Biology, 458(??):133–147, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304107.

Mansouri:2016:AIA

[MFM16] Ali Mansouri, Ali Fadavi, and Seyed Mohammad Mahdi Mortazavian. An artificial intelligence approach for modeling volume and fresh weight of callus — a case study of cumin (Cuminum cyminum L.). Journal of Theoretical Biology, 397(??):199–205, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001508.

Mirams:2012:TIE

[MFMB12] Gary R. Mirams, Alexander G. Fletcher, Philip K. Maini, and Helen M. Byrne. A theoretical investigation of the effect of proliferation and adhesion on monoclonal conversion in the colonic crypt. Journal of Theoretical Biology, 312(??):143–156, November 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003876.

Magnuson-Ford:2010:CSP

[MFMPS10] Karen Magnuson-Ford, Arne Mooers, Sébastien Rioux Paquette, and Mike Steel. Comparing strategies to preserve evolutionary diversity. Journal of Theoretical Biology, 266(1):107–116, September 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002882.

Marquez-Florez:2018:CMJ

[MFMS^{+}18] Kalenia M. Márquez-Flórez, James R. Monaghan, Sandra J. Shefelbine, Angélica Ramirez-Martínez, and Diego A. Garzón-Alvarado. A computational model for the joint onset and development. Journal of Theoretical Biology, 454(??):345–356, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301760.

Maderazo:2018:PFL

[MFN^{+}18] Dominic L. Maderazo, Jennifer A. Flegg, Melanie R. Neeland, Michael J. de Veer, and Mark B. Flegg. Physiological factors leading to a successful vaccination: a computational approach. Journal of Theoretical Biology, 454(??):215–230, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303011.

Mitchell:2014:WMW

[MFP^{+}14] K. M. Mitchell, A. M. Foss, H. J. Prudden, Z. Mukandavire, M. Pickles, J. R. Williams, H. C. Johnson, B. M. Ramesh, R. Washington, S. Isac, S. Rajaram, A. E. Phillips, J. Bradley, M. Alary, S. Moses, C. M. Lowndes, C. H. Watts, M.-C. Boily, and P. Vickerman. Who mixes with whom among men who have sex with men? Implications for modelling the HIV epidemic in southern India. Journal of Theoretical Biology, 355(??):140–150, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002148.

Mei:2018:API

[MFZ18] Juan Mei, Yi Fu, and Ji Zhao. Analysis and prediction of ion channel inhibitors by using feature selection and Chou’s general pseudo amino acid composition. Journal of Theoretical Biology, 456(??):41–48, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303709.

Mbah:2010:OCS

[MG10a] Martial L. Ndeffo Mbah and Christopher A. Gilligan. Optimization of control strategies for epidemics in heterogeneous populations with symmetric and asymmetric transmission. Journal of Theoretical Biology, 262(4):757–763, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900527X.

Mizera:2010:SME

[MG10b] Andrzej Mizera and Barbara Gambin. Stochastic modelling of the eukaryotic heat shock response. Journal of Theoretical Biology, 265(3):455–466, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002286.

Melak:2014:PNH

[MG14a] Tilahun Melak and Sunita Gakkhar. Potential non homologous protein targets of mycobacterium tuberculosis h37rv identified from protein–protein interaction network. Journal of Theoretical Biology, 361(??):152–158, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004433.

MGonigle:2014:AES

[MG14b] Leithen K. M’Gonigle and Philip B. Greenspoon. Allee effects and species co-existence in an environment where resource abundance varies. Journal of Theoretical Biology, 361(??):61–68, November 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004111.

Murray:2015:SSC

[MG15] John M. Murray and Ashish Goyal. In silico single cell dynamics of hepatitis B virus infection and clearance. Journal of Theoretical Biology, 366(??):91–102, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006729.

Meacham:2017:DLD

[MG17] Frazer Meacham and Thomas Getty. Deadbeats or losers: Discretionary male parental investment can make females less choosy. Journal of Theoretical Biology, 422(??):50–58, June 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301728.

Mohseni:2017:RDN

[MGB17] Ali Mohseni, Shahriar Gharibzadeh, and Fatemeh Bakouie. The role of driver nodes in managing epileptic seizures: Application of Kuramoto model. Journal of Theoretical Biology, 419(??):108–115, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730067X.

Moulton:2012:MGM

[MGC12] D. E. Moulton, A. Goriely, and R. Chirat. Mechanical growth and morphogenesis of seashells. Journal of Theoretical Biology, 311(??):69–79, October 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003438.

Mills:2013:PSC

[MGC13] H. L. Mills, A. Ganesh, and C. Colijn. Pathogen spread on coupled networks: Effect of host and network properties on transmission thresholds. Journal of Theoretical Biology, 320(??):47–57, March 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312006303.

Moulton:2015:MMB

[MGC15] D. E. Moulton, A. Goriely, and R. Chirat. The morphomechanical basis of ammonite form. Journal of Theoretical Biology, 364(??):220–230, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005608.

Martinez-Garcia:2013:SPM

[MGCL13] Ricardo Martínez-García, Justin M. Calabrese, and Cristóbal López. Spatial patterns in mesic savannas: The local facilitation limit and the role of demographic stochasticity. Journal of Theoretical Biology, 333(??):156–165, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002531.

Mesterton-Gibbons:2011:MCA

[MGGGA11] Mike Mesterton-Gibbons, Sergey Gavrilets, Janko Gravner, and Erol Akçay. Models of coalition or alliance formation. Journal of Theoretical Biology, 274(1):187–204, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006867.

Martin:2010:LVP

[MGGM10a] Natasha K. Martin, Eamonn A. Gaffney, Robert A. Gatenby, and Philip K. Maini. Leaky vessels as a potential source of stromal acidification in tumours. Journal of Theoretical Biology, 267(3):454–460, December 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000398X.

Martin:2010:TSI

[MGGM10b] Natasha K. Martin, Eamonn A. Gaffney, Robert A. Gatenby, and Philip K. Maini. Tumour-stromal interactions in acidmediated invasion: a mathematical model. Journal of Theoretical Biology, 267(3):461–470, December 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004510.

Moghtadaei:2013:PCD

[MGM13a] Motahareh Moghtadaei, Mohammad Reza Hashemi Golpayegani, and Reza Malekzadeh. Periodic and chaotic dynamics in a map-based model of tumor-immune interaction. Journal of Theoretical Biology, 334(??):130–140, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002622.

Moghtadaei:2013:VSF

[MGM13b] Motahareh Moghtadaei, Mohammad Reza Hashemi Golpayegani, and Reza Malekzadeh. A variable structure fuzzy neural network model of squamous dysplasia and esophageal squamous cell carcinoma based on a global chaotic optimization algorithm. Journal of Theoretical Biology, 318(??):164–172, February 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005887.

Marleau:2010:NFB

[MGML10] Justin N. Marleau, Frédéric Guichard, François Mallard, and Michel Loreau. Nutrient flows between ecosystems can destabilize simple food chains. Journal of Theoretical Biology, 266(1):162–174, September 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003127.

Martins:2015:URB

[MGO^{+}15] F. Martins, R. Gonçalves, J. Oliveira, M. Cruz-Monteagudo, J. M. Nieto-Villar, C. Paz y Miño, I. Rebelo, and E. Tejera. Unravelling the relationship between protein sequence and low-complexity regions entropies: Interactome implications. Journal of Theoretical Biology, 382(??):320–327, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315003288.

Mesterton-Gibbons:2016:HRD

[MGS16] Mike Mesterton-Gibbons and Tom N. Sherratt. How residency duration affects the outcome of a territorial contest: Complementary game-theoretic models. Journal of Theoretical Biology, 394(??):137–148, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000497.

Martinez-Garcia:2017:SCI

[MGT17] Ricardo Martínez-García and Corina E. Tarnita. Seasonality can induce coexistence of multiple bet-hedging strategies in Dictyostelium discoideum via storage effect. Journal of Theoretical Biology, 426(??):104–116, August 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302278.

Malherbe:2011:SMG

[MH11] G. Malherbe and David Holcman. Stochastic modeling of gene activation and applications to cell regulation. Journal of Theoretical Biology, 271(1):51–63, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006272.

Marshall:2012:CGD

[MH12] John M. Marshall and Bruce A. Hay. Confinement of gene drive systems to local populations: a comparative analysis. Journal of Theoretical Biology, 294(??):153–171, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005558.

Morin:2013:MDB

[MH13] Claire Morin and Christian Hellmich. Mineralization-driven bone tissue evolution follows from fluid-to-solid phase transformations in closed thermodynamic systems. Journal of Theoretical Biology, 335(??):185–197, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002907.

Murtada:2014:IRS

[MH14] Sae-Il Murtada and Gerhard A. Holzapfel. Investigating the role of smooth muscle cells in large elastic arteries: a finite element analysis. Journal of Theoretical Biology, 358(??):1–10, October 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002586.

Mayer:2017:ERP

[MH17] Christine Mayer and Thomas F. Hansen. Evolvability and robustness: a paradox restored. Journal of Theoretical Biology, 430(??):78–85, October 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303296.

Minoarivelo:2018:AAP

[MH18] Henintsoa O. Minoarivelo and Cang Hui. Alternative assembly processes from trait-mediated co-evolution in mutualistic communities. Journal of Theoretical Biology, 454(??):146–153, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302960.

Malerba:2016:IDP

[MHC16] Martino E. Malerba, Kirsten Heimann, and Sean R. Connolly. Improving dynamic phytoplankton reserve-utilization models with an indirect proxy for internal nitrogen. Journal of Theoretical Biology, 404(??):1–9, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301072.

Munir:2018:SNC

[MHD18] Anum Munir, Shahid Hussain, and Erum Dilshad. Silver nanoparticles conjugated with neurotrophin 3 upregulate myelin gene transcription pathway. Journal of Theoretical Biology, 459(??):111–118, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304697.

Muller:2013:BHS

[MHF^{+}13] J. Müller, B. A. Hense, T. M. Fuchs, M. Utz, and Ch. Pötzsche. Bet-hedging in stochastically switching environments. Journal of Theoretical Biology, 336(??):144–157, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003408.

Morin:2013:FSE

[MHH13] Claire Morin, Christian Hellmich, and Peter Henits. Fibrillar structure and elasticity of hydrating collagen: a quantitative multiscale approach. Journal of Theoretical Biology, 317(??):384–393, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005048.

Mikhchi:2016:ACD

[MHKA16] Abbas Mikhchi, Mahmood Honarvar, Nasser Emam Jomeh Kashan, and Mehdi Aminafshar. Assessing and comparison of different machine learning methods in parentoffspring trios for genotype imputation. Journal of Theoretical Biology, 399(??):148–158, June 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630011X.

Murphy:2011:TCM

[MHMM11] Kelly E. Murphy, Cameron L. Hall, Scott W. McCue, and D. L. Sean McElwain. A two-compartment mechanochemical model of the roles of transforming growth factor β and tissue tension in dermal wound healing. Journal of Theoretical Biology, 272(1):145–159, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006661.

Malunguza:2018:PIA

[MHMM18] Noble. J. Malunguza, Senelani D. Hove-Musekwa, and Zindoga Mukandavire. Projecting the impact of anal intercourse on HIV transmission among heterosexuals in high HIV prevalence settings. Journal of Theoretical Biology, 437(??):163–178, January 21, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304836.

Muijres:2012:AFP

[MHSH12] Florian T. Muijres, Per Henningsson, Melanie Stuiver, and Anders Hedenström. Aerodynamic flight performance in flap-gliding birds and bats. Journal of Theoretical Biology, 306(??):120–128, August 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001956.

Manore:2014:CDC

[MHX^{+}14] Carrie A. Manore, Kyle S. Hickmann, Sen Xu, Helen J. Wearing, and James M. Hyman. Comparing dengue and chikungunya emergence and endemic transmission in A. aegypti and A. albopictus. Journal of Theoretical Biology, 356(??):174–191, September 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931400263X.

Mougi:2011:UCD

[MI11a] Akihiko Mougi and Yoh Iwasa. Unique coevolutionary dynamics in a predator–prey system. Journal of Theoretical Biology, 277(1):83–89, May 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001093.

Muko:2011:LTE

[MI11b] Soyoka Muko and Yoh Iwasa. Long-term effect of coral transplantation: Restoration goals and the choice of species. Journal of Theoretical Biology, 280(1):127–138, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002128.

Muko:2011:OCS

[MI11c] Soyoka Muko and Yoh Iwasa. Optimal choice of species and size class for transplanting coral community. Journal of Theoretical Biology, 273(1):130–137, March 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006879.

Mamiya:2019:SEL

[MI19] Azumi Mamiya and Genki Ichinose. Strategies that enforce linear payoff relationships under observation errors in Repeated Prisoner’s Dilemma game. Journal of Theoretical Biology, 477(??):63–76, September 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931930236X.

Michel:2011:HCM

[Mic11] Denis Michel. Hierarchical cooperativity mediated by chromatin remodeling; the model of the MMTV transcription regulation. Journal of Theoretical Biology, 287(??):74–81, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003705.

Michel:2013:KAL

[Mic13] Denis Michel. Kinetic approaches to lactose operon induction and bimodality. Journal of Theoretical Biology, 325(??):62– 75, May 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313000659.

Michel:2014:GSR

[Mic14] Christian J. Michel. A genetic scale of reading frame coding. Journal of Theoretical Biology, 355(??):83–94, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314001684.

Michel:2015:EGS

[Mic15a] Christian J. Michel. An extended genetic scale of reading frame coding. Journal of Theoretical Biology, 365(??):164–174, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005876.

Michel:2015:MCS

[Mic15b] Christian J. Michel. The maximal C^{3} self-complementary trinucleotide circular code X in genes of bacteria, eukaryotes, plasmids and viruses. Journal of Theoretical Biology, 380(??):156–177, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500171X.

Miguel:2016:TOD

[Mig16] Antonio F. Miguel. Toward an optimal design principle in symmetric and asymmetric tree flow networks. Journal of Theoretical Biology, 389(??):101–109, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005263.

Mishra:2016:DPD

[MIH16] Avdesh Mishra, Sumaiya Iqbal, and Md Tamjidul Hoque. Discriminate protein decoys from native by using a scoring function based on ubiquitous phi and psi angles computed for all atom. Journal of Theoretical Biology, 398(??):112–121, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001764.

Malik:2016:OCM

[MIJ16] Tufail Malik, Mudassar Imran, and Raja Jayaraman. Optimal control with multiple human papillomavirus vaccines. Journal of Theoretical Biology, 393(??):179–193, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000175.

Milgram:2011:SBP

[Mil11] A. Milgram. The stability of the Boubaker polynomials expansion scheme (BPES)-based solution to Lotka–Volterra problem. Journal of Theoretical Biology, 271(1):157–158, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006442.

Milne:2016:CBA

[Mil16] Nick Milne. Curved bones: An adaptation to habitual loading. Journal of Theoretical Biology, 407(??):18–24, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302028.

Mitrokhin:2014:TFE

[Mit14] Yuriy Mitrokhin. Two faces of entropy and information in biological systems. Journal of Theoretical Biology, 359(??):192–198, October 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003610.

Mitchison:2016:CGA

[Mit16] Graeme Mitchison. Conformal growth of Arabidopsis leaves. Journal of Theoretical Biology, 408(??):155–166, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302570.

Mityushev:2017:OSR

[Mit17] Vladimir Mityushev. Optimal and stable random pattern formations. Journal of Theoretical Biology, 422(??):12–17, June 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301649.

Miyazawa:2016:SMP

[Miy16] Sanzo Miyazawa. Selection maintaining protein stability at equilibrium. Journal of Theoretical Biology, 391(??):21–34, February 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005767.

Miyazawa:2017:SOP

[Miy17] Sanzo Miyazawa. Selection originating from protein stability/foldability: Relationships between protein folding free energy, sequence ensemble, and fitness. Journal of Theoretical Biology, 433(??):21–38, November 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303922.

Mahdavi:2011:ADS

[MJ11] Abbas Mahdavi and Samad Jahandideh. Application of density similarities to predict membrane protein types based on pseudo-amino acid composition. Journal of Theoretical Biology, 276(1):132–137, May 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000749.

Min:2011:PKR

[MJC^{+}11] Yong Min, Xiaogang Jin, Ming Chen, Zhengzheng Pan, Ying Ge, and Jie Chang. Pathway knockout and redundancy in metabolic networks. Journal of Theoretical Biology, 270(1):63–69, February 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005977.

Matsuo:2014:CSN

[MJI14] Tadasu Matsuo, Marko Jusup, and Yoh Iwasa. The conflict of social norms may cause the collapse of cooperation: Indirect reciprocity with opposing attitudes towards in-group favoritism. Journal of Theoretical Biology, 346(??):34–46, April 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005699.

Ma:2011:FPD

[MJS11] Qi Ma, Anders Johansson, and David J. T. Sumpter. A first principles derivation of animal group size distributions. Journal of Theoretical Biology, 283(1):35–43, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002323.

Murphy:2016:MAE

[MJV16] James T. Murphy, Mark P. Johnson, and Frédérique Viard. A modelling approach to explore the critical environmental parameters influencing the growth and establishment of the invasive seaweed Undaria pinnatifida in Europe. Journal of Theoretical Biology, 396(??):105–115, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000783.

Maeda:2011:QMS

[MK11] Kazuhiro Maeda and Hiroyuki Kurata. Quasi-multiparameter sensitivity measure for robustness analysis of complex biochemical networks. Journal of Theoretical Biology, 272(1): 174–186, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519310006673.

Ma:2014:HMS

[MK14a] Qianqian Ma and Caner Kazanci. How much of the storage in the ecosystem is due to cycling? Journal of Theoretical Biology, 357(??):134–142, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002859.

Mann:2014:SFP

[MK14b] Ofri Mann and Moshe Kiflawi. Social foraging with partial (public) information. Journal of Theoretical Biology, 359(??):112–119, October 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003336.

Mironchenko:2014:OAP

[MK14c] Andrii Mironchenko and Jan Kozlowski. Optimal allocation patterns and optimal seed mass of a perennial plant. Journal of Theoretical Biology, 354(??):12–24, August 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001623.

Mougi:2014:SCA

[MK14d] Akihiko Mougi and Michio Kondoh. Stability of competitionantagonism-mutualism hybrid community and the role of community network structure. Journal of Theoretical Biology, 360(??):54–58, November 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003737.

Matsumoto:2016:IRB

[MK16a] Tomotaka Matsumoto and Jun Kitano. The intricate relationship between sexually antagonistic selection and the evolution of sex chromosome fusions. Journal of Theoretical Biology, 404(??):97–108, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301266.

Morozov:2016:TDG

[MK16b] Andrew Yu. Morozov and Oleg. A. Kuzenkov. Towards developing a general framework for modelling vertical migration in zooplankton. Journal of Theoretical Biology, 405(??):17–28, September 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000448.

Miles:2017:BCT

[MK17] Christopher E. Miles and James P. Keener. Bidirectionality from cargo thermal fluctuations in motor-mediated transport. Journal of Theoretical Biology, 424(??):37–48, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317302023.

Maeda:2018:LNF

[MK18a] Kazuhiro Maeda and Hiroyuki Kurata. Long negative feedback loop enhances period tunability of biological oscillators. Journal of Theoretical Biology, 440(??):21–31, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317305544.

Melnick:2018:TAT

[MK18b] Corey Melnick and Massoud Kaviany. Thermal actuation in TRPV1: Role of embedded lipids and intracellular domains. Journal of Theoretical Biology, 444(??):38–49, May 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300535.

Meakin:2019:CBS

[MK19] Sophie R. Meakin and Matt J. Keeling. Correlations between stochastic endemic infection in multiple interacting subpopulations. Journal of Theoretical Biology, 483 (??):Article 109991, December 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931930342X.

Martinez:2017:JGWa

[MKBE17a] Carlos Alberto Martínez, Kshitij Khare, Arunava Banerjee, and Mauricio A. Elzo. Joint genome-wide prediction in several populations accounting for randomness of genotypes: a hierarchical Bayes approach. I: Multivariate Gaussian priors for marker effects and derivation of the joint probability mass function of genotypes. Journal of Theoretical Biology, 417(??):8–19, March 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304325.

Martinez:2017:JGWb

[MKBE17b] Carlos Alberto Martínez, Kshitij Khare, Arunava Banerjee, and Mauricio A. Elzo. Joint genome-wide prediction in several populations accounting for randomness of genotypes: a hierarchical Bayes approach. II: Multivariate spike and slab priors for marker effects and derivation of approximate Bayes and fractional Bayes factors for the complete family of models. Journal of Theoretical Biology, 417(??):131–141, March 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316304313.

Martinez:2015:BMA

[MKE15] Carlos Alberto Martínez, Kshitij Khare, and Mauricio A. Elzo. On the Bayesness, minimaxity and admissibility of point estimators of allelic frequencies. Journal of Theoretical Biology, 383(??):106–115, October 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003768.

Mousavi:2014:MLM

[MKF^{+}14] Sayyed R. Mousavi, Ilnaz Khodadadi, Hossein Falsafain, Reza Nadimi, and Nasser Ghadiri. Maximum likelihood model based on minor allele frequencies and weighted Max-SAT formulation for haplotype assembly. Journal of Theoretical Biology, 350(??):49–56, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000551.

Montovan:2013:LBR

[MKJS13] Kathryn J. Montovan, Nathaniel Karst, Laura E. Jones, and Thomas D. Seeley. Local behavioral rules sustain the cell allocation pattern in the combs of honey bee colonies (Apis mellifera). Journal of Theoretical Biology, 336(??):75–86, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003317.

Manem:2014:SID

[MKKS14] V. S. K. Manem, M. Kohandel, N. L. Komarova, and S. Sivaloganathan. Spatial invasion dynamics on random and unstructured meshes: Implications for heterogeneous tumor populations. Journal of Theoretical Biology, 349(??):66–73, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000186.

McGillen:2014:GLM

[MKMG^{+}14] Jessica B. McGillen, Catherine J. Kelly, Alicia Martínez-González, Natasha K. Martin, Eamonn A. Gaffney, Philip K. Maini, and Víctor M. Pérez-García. Glucose-lactate metabolic cooperation in cancer: Insights from a spatial mathematical model and implications for targeted therapy. Journal of Theoretical Biology, 361(??):190–203, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005578.

McIntyre:2014:HRV

[MKMT14] Stephen D. McIntyre, Virendra Kakade, Yoichiro Mori, and Elena G. Tolkacheva. Heart rate variability and alternans formation in the heart: The role of feedback in cardiac dynamics. Journal of Theoretical Biology, 350(??):90–97, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000903.

Mund:2016:ADM

[MKPVH16] A. Mund, C. Kuttler, J. Pérez-Velázquez, and B. A. Hense. An age-dependent model to analyse the evolutionary stability of bacterial quorum sensing. Journal of Theoretical Biology, 405(??):104–115, September 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000266.

Martinez:2018:MCM

[MKRE18] Carlos Alberto Martínez, Kshitij Khare, Syed Rahman, and Mauricio A. Elzo. Modeling correlated marker effects in genome-wide prediction via Gaussian concentration graph models. Journal of Theoretical Biology, 437(??):67–78, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304800.

Mai:2009:BNB

[ML09] Zhongxing Mai and Haiyan Liu. Boolean network-based analysis of the apoptosis network: Irreversible apoptosis and stable surviving. Journal of Theoretical Biology, 259(4):760–769, August 21, 2009. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309001878. See corri-

gendum [ML12a].

Meng:2010:DPD

[ML10] Xinzhu Meng and Zhenqing Li. The dynamics of plant disease models with continuous and impulsive cultural control strategies. Journal of Theoretical Biology, 266(1):29–40, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002766.

Mitchell:2011:SGS

[ML11a] Cassie S. Mitchell and Robert H. Lee. Synaptic glutamate spillover increases NMDA receptor reliability at the cerebellar glomerulus. Journal of Theoretical Biology, 289(??):217–224, November 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004139.

Murrugarra:2011:RPM

[ML11b] David Murrugarra and Reinhard Laubenbacher. Regulatory patterns in molecular interaction networks. Journal of Theoretical Biology, 288(??):66–72, November 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004103.

Mai:2012:CBN

[ML12a] Zhongxing Mai and Haiyan Liu. Corrigendum to “Boolean network-based analysis of the apoptosis network: Irreversible apoptosis and stable surviving” [J. Theor. Biol. 259 (4) (2009) 760–769]. Journal of Theoretical Biology, 313(??):217, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200416X. See [ML09].

Mitchell:2012:CDD

[ML12b] Cassie S. Mitchell and Robert H. Lee. Cargo distributions differentiate pathological axonal transport impairments. Journal of Theoretical Biology, 300(??):277–291, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000276.

Marcoux:2013:NMP

[ML13] Marianne Marcoux and David Lusseau. Network modularity promotes cooperation. Journal of Theoretical Biology, 324(??):103–108, May 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312006406.

Markovitch:2014:MPD

[ML14] Omer Markovitch and Doron Lancet. Multispecies population dynamics of prebiotic compositional assemblies. Journal of Theoretical Biology, 357(??):26–34, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002756.

Moslonka-Lefebvre:2012:WSA

[MLBA12] Mathieu Moslonka-Lefebvre, Sebastian Bonhoeffer, and Samuel Alizon. Weighting for sex acts to understand the spread of STI on networks. Journal of Theoretical Biology, 311(??):46–53, October 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519312003244. See corrigendum [MLBA13].

Moslonka-Lefebvre:2013:CWS

[MLBA13] Mathieu Moslonka-Lefebvre, Sebastian Bonhoeffer, and Samuel Alizon. Corrigendum to “Weighting for sex acts to understand the spread of STI on networks” [J. Theor. Biol. 311 (2012) 46–53]. Journal of Theoretical Biology, 317(??):429, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931200584X. See [MLBA12].

Mendez:2011:EEF

[MLCH11] Vicenç Méndez, Isaac Llopis, Daniel Campos, and Werner Horsthemke. Effect of environmental fluctuations on invasion fronts. Journal of Theoretical Biology, 281(1):31–38, July 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002256.

Muraro:2016:MSM

[MLL^{+}16] D. Muraro, A. Larrieu, M. Lucas, J. Chopard, H. Byrne, C. Godin, and J. King. A multi-scale model of the interplay between cell signalling and hormone transport in specifying the root meristem of Arabidopsis thaliana. Journal of Theoretical Biology, 404(??):182–205, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300704.

Moslonka-Lefebvre:2015:EMT

[MLMG^{+}15] Mathieu Moslonka-Lefebvre, Hervé Monod, Christopher A. Gilligan, Elisabeta Vergu, and João A. N. Filipe. Epidemics in markets with trade friction and imperfect transactions. Journal of Theoretical Biology, 374(??):165–178, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931500096X.

McLean:2010:TOR

[MLT10] Alison K. McLean, Fabio Luciani, and Mark M. Tanaka. Trade-offs in resource allocation in the intracellular lifecycle of hepatitis C virus. Journal of Theoretical Biology, 267(4):565–572, December 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005084.

Magombedze:2012:MRD

[MM12] Gesham Magombedze and Nicola Mulder. A mathematical representation of the development of mycobacterium tuberculosis active, latent and dormant stages. Journal of Theoretical Biology, 292(??):44–59, January 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004899.

Merrill:2015:TAW

[MM15a] Stephen J. Merrill and Ying Mu. Thyroid autoimmunity as a window to autoimmunity: An explanation for sex differences in the prevalence of thyroid autoimmunity. Journal of Theoretical Biology, 375(??):95–100, June 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314007127.

Montevil:2015:BOC

[MM15b] Maël Montévil and Matteo Mossio. Biological organisation as closure of constraints. Journal of Theoretical Biology, 372(??):179–191, May 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001009.

Muniz:2018:MSI

[MM18] Danilo G. Muniz and Glauco Machado. Mate sampling influences the intensity of sexual selection and the evolution of costly sexual ornaments. Journal of Theoretical Biology, 447(??):74–83, June 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301413.

Marcu:2019:RHR

[MM19] Loredana G. Marcu and David Marcu. The role of hypofractionated radiotherapy in the management of head and neck cancer — a modelling approach. Journal of Theoretical Biology, 482(??):Article 109998, December 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303480.

MartinezMolina:2013:SDO

[MMAS13] Mario Martínez Molina, Marco A. Moreno-Armendáriz, and Juan Carlos Seck Tuoh Mora. On the spatial dynamics and oscillatory behavior of a predator–prey model based on cellular automata and local particle swarm optimization. Journal of Theoretical Biology, 336(??):173–184, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300338X.

Murray:2011:CWM

[MMB11] Philip J. Murray, Philip K. Maini, and Ruth E. Baker. The clock and wavefront model revisited. Journal of Theoretical Biology, 283(1):227–238, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002384.

Murall:2014:REA

[MMB14] Carmen Lía Murall, Kevin S. McCann, and Chris T. Bauch. Revising ecological assumptions about human papillomavirus interactions and type replacement. Journal of Theoretical Biology, 350(??):98–109, June 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000022.

Martinez:2018:TTM

[MMB18] Carlos Martínez, Francis Mairet, and Olivier Bernard. Theory of turbid microalgae cultures. Journal of Theoretical Biology, 456(??):190–200, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303400.

Macnamara:2019:SSM

[MMC19a] Cicely K. Macnamara, Elaine I. Mitchell, and Mark A. J. Chaplain. Spatial-stochastic modelling of synthetic gene regulatory networks. Journal of Theoretical Biology, 468(??):27– 44, May 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931930058X.

Moravec:2019:WIP

[MMC19b] Jirí C. Moravec, Stephen Marsland, and Murray P. Cox. Warfare induces post-marital residence change. Journal of Theoretical Biology, 474(??):52–62, August 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301869.

Montaseri:2012:HAR

[MMCZM12] Soheila Montaseri, Nasrollah Moghadam-Charkari, and Fatemeh Zare-Mirakabad. A heuristic approach to RNA-RNA interaction prediction. Journal of Theoretical Biology, 300(??):206–211, May 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000355.

Marino:2010:TMF

[MMFK10] Simeone Marino, Amy Myers, JoAnne L. Flynn, and Denise E. Kirschner. TNF and IL-10 are major factors in modulation of the phagocytic cell environment in lung and lymph node in tuberculosis: a next-generation two-compartmental model. Journal of Theoretical Biology, 265(4):586–598, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002419.

Mesk:2016:IST

[MMG^{+}16] Mohammed Mesk, Tewfik Mahdjoub, Sébastien Gourbière, Jorge E. Rabinovich, and Frédéric Menu. Invasion speeds of Triatoma dimidiata, vector of Chagas disease: An application of orthogonal polynomials method. Journal of Theoretical Biology, 395(??):126–143, April 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000503.

Mark:2010:NSV

[MMH10] Justin T. Mark, Brian B. Marion, and Donald D. Hoffman. Natural selection and veridical perceptions. Journal of Theoretical Biology, 266(4):504–515, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003772.

Martins:2012:DTC

[MMH^{+}12] Leonardo Martins, Jarno Mäkelä, Antti Häkkinen, Meenakshisundaram Kandhavelu, Olli Yli-Harja, José M. Fonseca, and Andre S. Ribeiro. Dynamics of transcription of closely spaced promoters in Escherichia coli, one event at a time. Journal of Theoretical Biology, 301(??):83–94, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000884.

Maaloy:2017:DSO

[MMJB17] Marthe Måløy, Frode Måløy, Per Jakobsen, and Bjørn Olav Brandsdal. Dynamic self-organisation of haematopoiesis and (a)symmetric cell division. Journal of Theoretical Biology, 414(??):147–164, February 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316304106.

Mitre:2016:MMG

[MMK16] Tina M. Mitre, Michael C. Mackey, and Anmar Khadra. Mathematical model of galactose regulation and metabolic consumption in yeast. Journal of Theoretical Biology, 407 (??):238–258, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316301813.

Mirsky:2011:SBA

[MMLK11] Henry P. Mirsky, Mark J. Miller, Jennifer J. Linderman, and Denise E. Kirschner. Systems biology approaches for understanding cellular mechanisms of immunity in lymph nodes during infection. Journal of Theoretical Biology, 287(??):160–170, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003420.

Murillo:2013:TMM

[MMP13] Lisa N. Murillo, Michael S. Murillo, and Alan S. Perelson. Towards multiscale modeling of influenza infection. Journal of Theoretical Biology, 332(??):267–290, September 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313001501.

Morin:2010:SBE

[MMRCCC10] Benjamin R. Morin, Liana Medina-Rios, Erika T. Camacho, and Carlos Castillo-Chavez. Static behavioral effects on gonorrhea transmission dynamics in a MSM population. Journal of Theoretical Biology, 267(1):35–40, November 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931000384X.

Mahdavi:2012:CAR

[MMY^{+}12] Sedigheh Mahdavi, Ali Mohades, Ali Salehzadeh Yazdi, Samad Jahandideh, and Ali Masoudi-Nejad. Computational analysis of RNA-protein interaction interfaces via the Voronoi diagram. Journal of Theoretical Biology, 293(??):55–64, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005042.

Masuda:2011:NAR

[MN11] Naoki Masuda and Mitsuhiro Nakamura. Numerical analysis of a reinforcement learning model with the dynamic aspiration level in the iterated Prisoner’s dilemma. Journal of Theoretical Biology, 278(1):55–62, June 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001342.

McNair:2012:TPT

[MN12] James N. McNair and J. Denis Newbold. Turbulent particle transport in streams: Can exponential settling be reconciled with fluid mechanics? Journal of Theoretical Biology, 300(??):62–80, May 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000240.

Minakuchi:2014:STF

[MN14] Mototaka Minakuchi and Seido Nagano. Stable territory formation in ecology and its potential generality in pattern formations. Journal of Theoretical Biology, 350(??):17–23, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005912.

McPherson:2012:SMP

[MNH^{+}12] N. J. McPherson, R. A. Norman, A. S. Hoyle, J. E. Bron, and N. G. H. Taylor. Stocking methods and parasite-induced reductions in capture: Modelling argulus foliaceus in trout fisheries. Journal of Theoretical Biology, 312(??):22–33, November 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200361X.

Mohammadi:2016:IAT

[MNSZ16] Mozafar Mohammadi, Foroogh Nejatollahi, Amirhossein Sakhteman, and Neda Zarei. Insilico analysis of three different tag polypeptides with dual roles in scFv antibodies. Journal of Theoretical Biology, 402(??):100–106, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300443.

Marriage:2012:MSB

[MO12] Tara N. Marriage and Maria E. Orive. Mutation-selection balance and mixed mating with asexual reproduction. Journal of Theoretical Biology, 308(??):25–35, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002147.

Mobilia:2010:ODR

[Mob10] Mauro Mobilia. Oscillatory dynamics in rock-paper-scissors games with mutations. Journal of Theoretical Biology, 264(1): 1–10, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000123.

Mahasa:2016:MMT

[MOEdP16] Khaphetsi Joseph Mahasa, Rachid Ouifki, Amina Eladdadi, and Lisette de Pillis. Mathematical model of tumor-immune surveillance. Journal of Theoretical Biology, 404(??):312–330, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301473.

Moghimbeigi:2015:TPZ

[Mog15] Abbas Moghimbeigi. Two-part zero-inflated negative binomial regression model for quantitative trait loci mapping with count trait. Journal of Theoretical Biology, 372(??):74–80, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315000752.

Moller:2012:MAN

[Möl12] Ralf Möller. A model of ant navigation based on visual prediction. Journal of Theoretical Biology, 305(??):118–130, July 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312002032.

Molski:2018:TMP

[Mol18] Marcin Molski. Theoretical modeling of pre and postnatal growth. Journal of Theoretical Biology, 458(??):58–67, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304326.

Monk:2018:MFP

[Mon18] Travis Monk. Martingales and the fixation probability of high-dimensional evolutionary graphs. Journal of Theoretical Biology, 451(??):10–18, August 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302236.

Moore:2014:BDH

[Moo14] James R. Moore. The benefits of diversity: Heterogeneous DC populations allow for both immunity and tolerance. Journal of Theoretical Biology, 357(??):86–102, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002641.

Moran:2010:MCA

[Mor10a] Oscar Moran. Model of the cAMP activation of chloride transport by CFTR channel and the mechanism of potentiators. Journal of Theoretical Biology, 262(1):73–79, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004068.

Morozov:2010:EHT

[Mor10b] Andrew Yu. Morozov. Emergence of Holling type III zooplankton functional response: Bringing together field evidence and mathematical modelling. Journal of Theoretical Biology, 265(1):45–54, July 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001992.

Morford:2011:TAS

[Mor11] Livingstone Morford. A theoretical application of selectable markers in bacterial episomes for a DNA cryptosystem. Journal of Theoretical Biology, 273(1):100–102, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310007010.

Morbey:2013:PSS

[Mor13] Yolanda E. Morbey. Protandry, sexual size dimorphism, and adaptive growth. Journal of Theoretical Biology, 339 (??):93–99, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519313002245.

Morozov:2016:MBE

[Mor16] Andrew Morozov. Modelling biological evolution: Linking mathematical theories with empirical realities. Journal of Theoretical Biology, 405(??):1–4, September 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301849.

Morimoto:2019:FDM

[Mor19] Juliano Morimoto. Foraging decisions as multi-armed bandit problems: Applying reinforcement learning algorithms to foraging data. Journal of Theoretical Biology, 467(??):48–56, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300566.

Manhart:2015:EFB

[MOSS15] Angelika Manhart, Dietmar Oelz, Christian Schmeiser, and Nikolaos Sfakianakis. An extended Filament Based Lamellipodium Model produces various moving cell shapes in the presence of chemotactic signals. Journal of Theoretical Biology, 382(??):244–258, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003239.

Mougi:2012:UPP

[Mou12a] Akihiko Mougi. Unusual predator–prey dynamics under reciprocal phenotypic plasticity. Journal of Theoretical Biology, 305(??):96–102, July 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001932.

Mousavi:2012:IHA

[Mou12b] Sayyed R. Mousavi. Improved haplotype assembly using xor genotypes. Journal of Theoretical Biology, 298(??):122–130, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000045.

Mishra:2012:ERB

[MP12] Pooja Mishra and Paras Nath Pandey. Elman RNN based classification of proteins sequences on account of their mutual information. Journal of Theoretical Biology, 311(??):40–45, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003323.

Maury:2013:IPC

[MP13a] Olivier Maury and Jean-Christophe Poggiale. From individuals to populations to communities: a dynamic energy budget model of marine ecosystem size-spectrum including life history diversity. Journal of Theoretical Biology, 324(??):52–71, May 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931300043X.

Michel:2013:PST

[MP13b] Christian J. Michel and Giuseppe Pirillo. A permuted set of a trinucleotide circular code coding the 20 amino acids in variant nuclear codes. Journal of Theoretical Biology, 319(??):116–121, February 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312006121.

Moulay:2013:MMC

[MP13c] Djamila Moulay and Yoann Pigné. A metapopulation model for chikungunya including populations mobility on a largescale network. Journal of Theoretical Biology, 318(??):129– 139, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519312005838.

Maciejewski:2014:EEG

[MP14a] Wes Maciejewski and Gregory J. Puleo. Environmental evolutionary graph theory. Journal of Theoretical Biology, 360(??):117–128, November 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003919.

Mondal:2014:CPA

[MP14b] Sukanta Mondal and Priyadarshini P. Pai. Chou’s pseudo amino acid composition improves sequence-based antifreeze protein prediction. Journal of Theoretical Biology, 356(??): 30–35, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251931400215X.

McMorris:2016:SAL

[MP16] F. R. McMorris and Robert C. Powers. Some axiomatic limitations for consensus and supertree functions on hierarchies. Journal of Theoretical Biology, 404(??):342–347, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301503.

Murallidharan:2019:TGM

[MP19] Janani Srree Murallidharan and Stephen J. Payne. Thrombus growth modelling and stenosis prediction in the cerebral microvasculature. Journal of Theoretical Biology, 478(??):1–13, October 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302401.

Matsiaka:2017:CAL

[MPBS17] Oleksii M. Matsiaka, Catherine J. Penington, Ruth E. Baker, and Matthew J. Simpson. Continuum approximations for lattice-free multi-species models of collective cell migration. Journal of Theoretical Biology, 422(??):1–11, June 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301650.

McGovern:2012:DMC

[MPC12] Andrew P. McGovern, Benjamin E. Powell, and Timothy J. T. Chevassut. A dynamic multi-compartmental model of DNA methylation with demonstrable predictive value in hematological malignancies. Journal of Theoretical Biology, 310(??):14–20, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003050.

Marrero-Ponce:2015:NBM

[MPCTGJ^{+}15] Yovani Marrero-Ponce, Ernesto Contreras-Torres, César R. García-Jacas, Stephen J. Barigye, Néstor Cubillán, and Ysaías J. Alvarado. Novel 3D bio-macromolecular bilinear descriptors for protein science: Predicting protein structural classes. Journal of Theoretical Biology, 374(??):125–137, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315001411.

Maeng:2013:MAE

[MPJH13] Joo-Sung Maeng, Jae-Hyung Park, Seong-Min Jang, and Seog-Young Han. A modeling approach to energy savings of flying Canada geese using computational fluid dynamics. Journal of Theoretical Biology, 320(??):76–85, March 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312006212.

Mercker:2012:MAC

[MPK^{+}12] M. Mercker, M. Ptashnyk, J. Kühnle, D. Hartmann, M. Weiss, and W. Jäger. A multiscale approach to curvature modulated sorting in biological membranes. Journal of Theoretical Biology, 301(??):67–82, May 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000598.

Malkemper:2016:SMA

[MPL16] E. Pascal Malkemper, Michael S. Painter, and Lukas Landler. Shifted magnetic alignment in vertebrates: Evidence for neural lateralization? Journal of Theoretical Biology, 399(??):141–147, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300169.

Morin:2014:DRM

[MPLK14] Benjamin R. Morin, Charles Perrings, Simon Levin, and Ann Kinzig. Disease risk mitigation: The equivalence of two selective mixing strategies on aggregate contact patterns and resulting epidemic spread. Journal of Theoretical Biology, 363(??):262–270, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004718.

Marks:2010:EIS

[MPM10] Alison J. Marks, Deenan Pillay, and Angela R. McLean. The effect of intrinsic stochasticity on transmitted HIV drug resistance patterns. Journal of Theoretical Biology, 262(1):1–13, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004378.

Moreira:2012:EDC

[MPNP12] João A. Moreira, Flávio L. Pinheiro, Ana Nunes, and Jorge M. Pacheco. Evolutionary dynamics of collective action when individual fitness derives from group decisions taken in the past. Journal of Theoretical Biology, 298(??):8–15, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311006515.

Mellor:2015:MAL

[MPP^{+}15] Nathan Mellor, Benjamin Péret, Silvana Porco, Ilkka Sairanen, Karin Ljung, Malcolm Bennett, and John King. Modelling of Arabidopsis LAX3 expression suggests auxin homeostasis. Journal of Theoretical Biology, 366(??):57–70, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006353.

Marcolino:2016:SIA

[MPP^{+}16a] Antonio C. S. Marcolino, William F. Porto, Állan S. Pires, Octavio L. Franco, and Sérgio A. Alencar. Structural impact analysis of missense SNPs present in the uroguanylin gene by long-term molecular dynamics simulations. Journal of Theoretical Biology, 410(??):9–17, December 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630296X.

Michel:2016:MDT

[MPP16b] Christian J. Michel, Marco Pellegrini, and Giuseppe Pirillo. Maximal dinucleotide and trinucleotide circular codes. Journal of Theoretical Biology, 389(??):40–46, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315004324.

Melchionda:2018:SBI

[MPP^{+}18] D. Melchionda, E. Pastacaldi, C. Perri, M. Banerjee, and E. Venturino. Social behavior-induced multistability in minimal competitive ecosystems. Journal of Theoretical Biology, 439(??):24–38, February 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305234.

Menconi:2011:TLA

[MPS^{+}11] Giulia Menconi, Aldamaria Puliti, Isabella Sbrana, Valerio Conti, and Roberto Marangoni. A top-down linguistic approach to the analysis of genomic sequences: The metabotropic glutamate receptors 1 and 5 in human and in mouse as a case study. Journal of Theoretical Biology, 270(1):134–142, February 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006053.

McVinish:2016:LSD

[MPS16] R. McVinish, P. K. Pollett, and A. Shausan. Limiting the spread of disease through altered migration patterns. Journal of Theoretical Biology, 393(??):60–66, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000242.

Menini:2019:BNA

[MPT19] Laura Menini, Corrado Possieri, and Antonio Tornambè. Boolean network analysis through the joint use of linear algebra and algebraic geometry. Journal of Theoretical Biology, 472(??):46–53, July 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301535.

Mahalingam:2014:PFB

[MPY14] Rajasekaran Mahalingam, Hung-Pin Peng, and An-Suei Yang. Prediction of FMN-binding residues with three-dimensional probability distributions of interacting atoms on protein surfaces. Journal of Theoretical Biology, 343(??):154–161, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005195.

Mohammadpour:2016:KRD

[MPZK16] Hemn Mohammadpour, Ali Akbar Pourfathollah, Mahin Nikougoftar Zarif, and Saeed Khalili. Key role of dkk3 protein in inhibition of cancer cell proliferation: An in silico identification. Journal of Theoretical Biology, 393(??):98–104, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000369.

Majumdar:2010:TIL

[MR10] Priyadarshi Majumdar and Sudipto Roy. A theoretical interpretation of length-biomass allometry of predominantly bidimensional seaweeds. Journal of Theoretical Biology, 266(2):226–230, September 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003322.

Magnus:2011:ROM

[MR11a] Carsten Magnus and Roland R. Regoes. Restricted occupancy models for neutralization of HIV virions and populations. Journal of Theoretical Biology, 283(1):192–202, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100302X.

Mann:2011:MEH

[MR11b] Joanne Mann and Mick Roberts. Modelling the epidemiology of hepatitis B in new Zealand. Journal of Theoretical Biology, 269(1):266–272, January 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000562X.

Milinski:2012:ISC

[MR12] Manfred Milinski and Bettina Rockenbach. On the interaction of the stick and the carrot in social dilemmas. Journal of Theoretical Biology, 299(??):139–143, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001548.

Malanson:2015:NFA

[MR15] George P. Malanson and Lynn M. Resler. Neighborhood functions alter unbalanced facilitation on a stress gradient. Journal of Theoretical Biology, 365(??):76–83, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314005943.

Muthukumaran:2018:QBM

[MR18] Panchaksaram Muthukumaran and Muniyan Rajiniraja. MIA-QSAR based model for bioactivity prediction of flavonoid derivatives as acetylcholinesterase inhibitors. Journal of Theoretical Biology, 459(??):103–110, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304661.

Muthukumaran:2019:AQB

[MR19] Panchaksaram Muthukumaran and Muniyan Rajiniraja. Aug– MIA–QSAR based strategy in bioactivity prediction of a series of flavonoid derivatives as HIV-1 inhibitors. Journal of Theoretical Biology, 469(??):18–24, May 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300888.

Mahlbacher:2019:MMT

[MRF19] Grace E. Mahlbacher, Kara C. Reihmer, and Hermann B. Frieboes. Mathematical modeling of tumor-immune cell interactions. Journal of Theoretical Biology, 469(??):47–60, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300955.

Morales:2015:NMM

[MROS15] M. A. Morales, J. F. Rojas, J. Oliveros, and A. A. Hernández S. A new mechanochemical model: Coupled Ginzburg– Landau and Swift–Hohenberg equations in biological patterns of marine animals. Journal of Theoretical Biology, 368(??):37– 54, March 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006948.

Marini:2017:EVB

[MRPH17] Giovanni Marini, Roberto Rosá, Andrea Pugliese, and Hans Heesterbeek. Exploring vector-borne infection ecology in multi-host communities: a case study of West Nile virus. Journal of Theoretical Biology, 415(??):58–69, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316304179.

Mercado-Reyes:2015:OPT

[MRPLAS15] Agustín Mercado-Reyes, Pablo Padilla-Longoria, and Alfonso Arroyo-Santos. Objects and processes: Two notions for understanding biological information. Journal of Theoretical Biology, 380(??):115–122, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002490.

Manapat:2012:SED

[MRPN12] Michael L. Manapat, David G. Rand, Christina Pawlowitsch, and Martin A. Nowak. Stochastic evolutionary dynamics resolve the Traveler’s dilemma. Journal of Theoretical Biology, 303(??):119–127, June 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001312.

Muller:2014:EAP

[MRS14] Stefan Müller, Georg Regensburger, and Ralf Steuer. Enzyme allocation problems in kinetic metabolic networks: Optimal solutions are elementary flux modes. Journal of Theoretical Biology, 347(??):182–190, April 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005420.

Means:2010:STC

[MS10a] Shawn A. Means and James Sneyd. Spatio-temporal calcium dynamics in pacemaking units of the interstitial cells of Cajal. Journal of Theoretical Biology, 267(2):137–152, November 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004170.

Mochizuki:2010:ASS

[MS10b] Atsushi Mochizuki and Daisuke Saito. Analyzing steady states of dynamics of bio-molecules from the structure of regulatory networks. Journal of Theoretical Biology, 266(2):323–335, September 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002912.

Martyn:2012:IIL

[MS12a] Iain Martyn and Mike Steel. The impact and interplay of long and short branches on phylogenetic information content. Journal of Theoretical Biology, 314(??):157–163, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004626.

Misra:2012:DMM

[MS12b] A. K. Misra and Vishal Singh. A delay mathematical model for the spread and control of water borne diseases. Journal of Theoretical Biology, 301(??):49–56, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200077X.

Molina:2012:MSD

[MS12c] Chai Molina and Lewi Stone. Modelling the spread of diseases in clustered networks. Journal of Theoretical Biology, 315(??):110–118, December 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004584.

Margolskee:2013:LMF

[MS13] Alison Margolskee and James F. Selgrade. A lifelong model for the female reproductive cycle with an antimüllerian hormone treatment to delay menopause. Journal of Theoretical Biology, 326(??):21–35, June 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000751.

Ma:2014:SBP

[MS14a] Xin Ma and Xiao Sun. Sequence-based predictor of ATPbinding residues using random forest and mRMR-IFS feature selection. Journal of Theoretical Biology, 360(??):59–66, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003804.

Maykranz:2014:CAF

[MS14b] Daniel Maykranz and Andre Seyfarth. Compliant ankle function results in landing-take off asymmetry in legged locomotion. Journal of Theoretical Biology, 349(??):44–49, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000460.

Morishita:2014:BIW

[MS14c] Yoshihiro Morishita and Takayuki Suzuki. Bayesian inference of whole-organ deformation dynamics from limited space-time point data. Journal of Theoretical Biology, 357(??):74–85, September 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002471.

Mossel:2014:MRT

[MS14d] Elchanan Mossel and Mike Steel. Majority rule has transition ratio 4 on Yule trees under a 2-state symmetric model. Journal of Theoretical Biology, 360(??):315–318, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400441X.

Means:2016:IHV

[MS16] Shawn A. Means and Robert J. Smith. The impact of human and vector distributions on the spatial prevalence of malaria in sub-Saharan Africa. Journal of Theoretical Biology, 409(??):70–85, November 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302727.

Mehrafrooz:2018:MDB

[MS18] Behzad Mehrafrooz and Amir Shamloo. Mechanical differences between ATP and ADP actin states: a molecular dynamics study. Journal of Theoretical Biology, 448(??):94–103, July 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931830170X.

Moulton:2016:MMS

[MSA^{+}16] D. E. Moulton, V. Sulzer, G. Apodaca, H. M. Byrne, and S. L. Waters. Mathematical modelling of stretch-induced membrane traffic in bladder umbrella cells. Journal of Theoretical Biology, 409(??):115–132, November 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302740.

Meyer:2016:PBC

[MSB16] A. D. Meyer, D. F. Stevens, and J. C. Blackwood. Predicting bat colony survival under controls targeting multiple transmission routes of white-nose syndrome. Journal of Theoretical Biology, 409(??):60–69, November 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302752.

McLaren:2013:SET

[MSBB13] James D. McLaren, Judy Shamoun-Baranes, and Willem Bouten. Stop early to travel fast: modelling risk-averse scheduling among nocturnally migrating birds. Journal of Theoretical Biology, 316(??):90–98, January 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004997.

Melkikh:2010:AMI

[MSC10] Alexey V. Melkikh, Vladimir D. Seleznev, and Oksana I. Chesnokova. Analytical model of ion transport and conversion of light energy in chloroplasts. Journal of Theoretical Biology, 264(3):702–710, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001852.

Mochan:2014:MMI

[MSE^{+}14] Ericka Mochan, David Swigon, G. Bard Ermentrout, Sarah Lukens, and Gilles Clermont. A mathematical model of intrahost pneumococcal pneumonia infection dynamics in murine strains. Journal of Theoretical Biology, 353(??):44–54, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000964.

Murase:2010:RWG

[MSIR10] Yohsuke Murase, Takashi Shimada, Nobuyasu Ito, and Per Arne Rikvold. Random walk in genome space: a key ingredient of intermittent dynamics of community assembly on evolutionary time scales. Journal of Theoretical Biology, 264(3):663–672, June 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001785.

Maruvka:2010:UFS

[MSK10] Yosi E. Maruvka, Nadav M. Shnerb, and David A. Kessler. Universal features of surname distribution in a subsample of a growing population. Journal of Theoretical Biology, 262(2):245–256, January 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004421.

Murray:2016:VLT

[MSL^{+}16] John M. Murray, Ognjen Stancevic, Marc Lütgehetmann, Karsten Wursthorn, Joerg Petersen, and Maura Dandri. Variability in long-term hepatitis B virus dynamics under antiviral therapy. Journal of Theoretical Biology, 391(??):74–80, February 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005895.

Markham:2014:CMM

[MSM^{+}14] Deborah C. Markham, Matthew J. Simpson, Philip K. Maini, Eamonn A. Gaffney, and Ruth E. Baker. Comparing methods for modelling spreading cell fronts. Journal of Theoretical Biology, 353(??):95–103, July 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000988.

Markvoort:2014:CSC

[MSN14] Albert J. Markvoort, Sam Sinai, and Martin A. Nowak. Computer simulations of cellular group selection reveal mechanism for sustaining cooperation. Journal of Theoretical Biology, 357(??):123–133, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002598.

Mahdevar:2012:TFB

[MSND12] Ghasem Mahdevar, Mehdi Sadeghi, and Abbas Nowzari-Dalini. Transcription factor binding sites detection by using alignment-based approach. Journal of Theoretical Biology, 304(??):96–102, July 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001713.

Martinkova:2019:MIP

[MSP19] Natália Martínková, Pavel Skrabánek, and Jiri Pikula. Modelling invasive pathogen load from non-destructive sampling data. Journal of Theoretical Biology, 464(??):98–103, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318306258.

Meher:2016:CAP

[MSRW16] Prabina Kumar Meher, Tanmaya Kumar Sahu, A. R. Rao, and S. D. Wahi. A computational approach for prediction of donor splice sites with improved accuracy. Journal of Theoretical Biology, 404(??):285–294, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301461.

Montgomery-Smith:2010:SME

[MSS10] Stephen J. Montgomery-Smith and Francis J. Schmidt. Statistical methods for estimating complexity from competition experiments between two populations. Journal of Theoretical Biology, 264(3):1043–1046, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001219.

Michiko:2018:ERS

[MSSM18] Nosaka Michiko, Budi Setyawan, and Syunya Sunaba-Mitsumasu. Enzymatic reaction sites on a plane formed with four exon-junctions. Journal of Theoretical Biology, 457(??):51–56, November 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303849.

Mai:2019:JID

[MSZW19] Ali Mai, Guowei Sun, Fengqin Zhang, and Lin Wang. The joint impacts of dispersal delay and dispersal patterns on the stability of predator–prey metacommunities. Journal of Theoretical Biology, 462(??):455–465, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305885.

Marashi:2014:MAE

[MT14] Sayed-Amir Marashi and Mojtaba Tefagh. A mathematical approach to emergent properties of metabolic networks: Partial coupling relations, hyperarcs and flux ratios. Journal of Theoretical Biology, 355(??):185–193, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002318.

Murakawa:2015:CMC

[MT15] Hideki Murakawa and Hideru Togashi. Continuous models for cell–cell adhesion. Journal of Theoretical Biology, 374(??):1– 12, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001058.

Malaspinas:2016:STM

[MTdS^{+}16] O. Malaspinas, A. Turjman, D. Ribeiro de Sousa, G. Garcia-Cardena, M. Raes, P.-T. T. Nguyen, Y. Zhang, G. Courbebaisse, C. Lelubre, K. Zouaoui Boudjeltia, and B. Chopard. A spatio-temporal model for spontaneous thrombus formation in cerebral aneurysms. Journal of Theoretical Biology, 394(??):68–76, April 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931600028X.

Matsuoka:2015:SAE

[MTE15] Takeshi Matsuoka, Shigenori Tanaka, and Kuniyoshi Ebina. Systems approach to excitation-energy and electron transfer reaction networks in photosystem II complex: model studies for chlorophyll a fluorescence induction kinetics. Journal of Theoretical Biology, 380(??):220–237, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002349.

Mackey:2011:MDG

[MTKY11] Michael C. Mackey, Marta Tyran-Kamińska, and Romain Yvinec. Molecular distributions in gene regulatory dynamics. Journal of Theoretical Biology, 274(1):84–96, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000324.

Munoz-Tamayo:2010:MMC

[MTLW^{+}10] Rafael Muñoz-Tamayo, Béatrice Laroche, Éric Walter, Joël Doré, and Marion Leclerc. Mathematical modelling of carbohydrate degradation by human colonic microbiota. Journal of Theoretical Biology, 266(1):189–201, September 7, 2010.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002845.

Marin:2012:CTQ

[MTNM12] Arturo Marín, Héctor Tejero, Juan Carlos Nuño, and Francisco Montero. Characteristic time in quasispecies evolution. Journal of Theoretical Biology, 303(??):25–32, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312001142.

Mulder:2011:PDA

[Mul11] Willem H. Mulder. Probability distributions of ancestries and genealogical distances on stochastically generated rooted binary trees. Journal of Theoretical Biology, 280(1):139–145, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001998. See corri-

gendum [Mul12].

Mulder:2012:CPD

[Mul12] Willem H. Mulder. Corrigendum to “Probability distributions of ancestries and genealogical distances on stochastically generated rooted binary trees” [J. Theor. Biol. 280 (2011) 139– 145]. Journal of Theoretical Biology, 314(??):216–217, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004638. See [Mul11].

Munk:2012:MIP

[Mun12] Ole Lajord Munk. Model-independent plot of dynamic PET data facilitates data interpretation and model selection. Journal of Theoretical Biology, 295(??):1–8, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005662.

Masso:2010:KBC

[MV10] Majid Masso and Iosif I. Vaisman. Knowledge-based computational mutagenesis for predicting the disease potential of human non-synonymous single nucleotide polymorphisms. Journal of Theoretical Biology, 266(4):560–568, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310003838.

Morsky:2018:CAS

[MV18] Bryce Morsky and Dervis Can Vural. Cheater-altruist synergy in public goods games. Journal of Theoretical Biology, 454(??):231–239, October 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303060.

Muller:2017:UNF

[MvAKR17] E. J. Müller, S. J. van Albada, J. W. Kim, and P. A. Robinson. Unified neural field theory of brain dynamics underlying oscillations in Parkinson’s disease and generalized epilepsies. Journal of Theoretical Biology, 428(??):132–146, September 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317302886.

Ma:2013:ICN

[MvdDW13] Junling Ma, P. van den Driessche, and Frederick H. Willeboordse. The importance of contact network topology for the success of vaccination strategies. Journal of Theoretical Biology, 325(??):12–21, May 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000258.

Marinelli:2018:TBB

[MVGGB18] Isabella Marinelli, Theodore Vo, Luca Gerardo-Giorda, and Richard Bertram. Transitions between bursting modes in the integrated oscillator model for pancreatic β-cells. Journal of Theoretical Biology, 454(??):310–319, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303126.

Martin:2011:MMH

[MVH11] Natasha K. Martin, Peter Vickerman, and Matthew Hickman. Mathematical modelling of hepatitis C treatment for injecting drug users. Journal of Theoretical Biology, 274(1):58–66, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000038.

Munther:2013:ESF

[MW13] Daniel Munther and Jianhong Wu. Enhanced surveillance on food-borne disease outbreaks: Dynamics of crosscontamination in biocidal wash procedure. Journal of Theoretical Biology, 321(??):28–35, March 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000064.

McLeod:2014:RBE

[MW14] David V. McLeod and Geoff Wild. The relationship between ecology and the optimal helping strategy in cooperative breeders. Journal of Theoretical Biology, 354(??):25–34, August 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314001386.

Manzourolajdad:2013:ITU

[MWSM13] Amirhossein Manzourolajdad, Yingfeng Wang, Timothy I. Shaw, and Russell L. Malmberg. Information-theoretic uncertainty of SCFG-modeled folding space of the non-coding RNA. Journal of Theoretical Biology, 318(??):140–163, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005620.

Martin:2012:MCA

[MYC12] Geoffrey V. Martin, Yeoheung Yun, and Laura Conforti. Modulation of T cell activation by localized K^{+} accumulation at the immunological synapse — a mathematical model. Journal of Theoretical Biology, 300(??):173–182, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000264.

Mak:2011:DRD

[MYK^{+}11] Arthur F. T. Mak, Yanyan Yu, Linda P. C. Kwan, Lei Sun, and Eric W. C. Tam. Deformation and reperfusion damages and their accumulation in subcutaneous tissues during loading and unloading: a theoretical modeling of deep tissue injuries. Journal of Theoretical Biology, 289(??):65–73, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004255.

Matsumoto:2017:CGF

[MYK17] Tomotaka Matsumoto, Kohta Yoshida, and Jun Kitano. Contribution of gene flow to the evolution of recombination suppression in sex chromosomes. Journal of Theoretical Biology, 431(??):25–31, October 27, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303648.

Mellard:2011:VDP

[MYLK11] Jarad P. Mellard, Kohei Yoshiyama, Elena Litchman, and Christopher A. Klausmeier. The vertical distribution of phytoplankton in stratified water columns. Journal of Theoretical Biology, 269(1):16–30, January 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005242.

Matsumoto:2013:DFT

[MYN^{+}13] Tomotaka Matsumoto, Akiko A. Yasumoto, Kozue Nitta, Tetsukazu Yahara, and Hidenori Tachida. Difference in flowering time as an isolating barrier. Journal of Theoretical Biology, 317(??):161–167, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005267.

Matsumoto:2015:DFT

[MYN^{+}15] Tomotaka Matsumoto, Akiko A. Yasumoto, Kozue Nitta, Shun K. Hirota, Tetsukazu Yahara, and Hidenori Tachida. Difference in flowering time can initiate speciation of nocturnally flowering species. Journal of Theoretical Biology, 370(??):61– 71, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315000533.

Mpolya:2014:EDV

[MYOS14] Emmanuel A. Mpolya, Kenta Yashima, Hisashi Ohtsuki, and Akira Sasaki. Epidemic dynamics of a vector-borne disease on a villages-and-city star network with commuters. Journal of Theoretical Biology, 343(??):120–126, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005511.

Magal:2017:CLF

[MZ17] Pierre Magal and Zhengyang Zhang. Competition for light in forest population dynamics: From computer simulator to mathematical model. Journal of Theoretical Biology, 419(??):290–304, April 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300802.

Mei:2018:APP

[MZ18] Juan Mei and Ji Zhao. Analysis and prediction of presynaptic and postsynaptic neurotoxins by Chou’s general pseudo amino acid composition and motif features. Journal of Theoretical Biology, 447(??):147–153, June 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301486.

Mehra:2019:OBA

[MZAI19] Amir Hossein Amiri Mehra, Iman Zamani, Zohreh Abbasi, and Asier Ibeas. Observer-based adaptive PI sliding mode control of developed uncertain SEIAR influenza epidemic model considering dynamic population. Journal of Theoretical Biology, 482(??):Article 109984, December 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303273.

Man:2016:QDH

[MZM^{+}16] Menghua Man, Ya Zhang, Guilei Ma, Karl Friston, and Shanghe Liu. Quantification of degeneracy in Hodgkin–Huxley neurons on Newman–Watts small world network. Journal of Theoretical Biology, 402(??):62–74, August 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300789.

Moleriu:2014:IMT

[MZMM^{+}14] Radu Dumitru Moleriu, Daniela Zaharie, Lavinia Cristina Moatar-Moleriu, Alexandra Teodora Gruia, Ani Aurora Mic, and Felix Aurel Mic. Insights into the mechanisms of thymus involution and regeneration by modeling the glucocorticoidinduced perturbation of thymocyte populations dynamics. Journal of Theoretical Biology, 348(??):80–99, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400037X.

Mackiewicz:2010:GAM

[MZWC10] Dorota Mackiewicz, Marta Zawierta, Wojciech Waga, and StanisLaw Cebrat. Genome analyses and modelling the relationships between coding density, recombination rate and chromosome length. Journal of Theoretical Biology, 267(2):186–192, November 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004315.

Nahalka:2012:GTF

[Nah12] Jozef Nahalka. Glycocodon theory-the first table of glycocodons. Journal of Theoretical Biology, 307(??):193–204, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002366.

Nahalka:2014:PRR

[Nah14] Jozef Nahalka. Protein–RNA recognition: Cracking the code. Journal of Theoretical Biology, 343(??):9–15, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005316.

Nakamura:2011:MCH

[NAK^{+}11] Yoshihiro Nakamura, Shoichi Awa, Hitoshi Kato, Yoichi M. Ito, Akira Kamiya, and Takashi Igarashi. Model combining hydrodynamics and fractal theory for analysis of in vivo peripheral pulmonary and systemic resistance of shunt cardiac defects. Journal of Theoretical Biology, 287(??):64–73, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003663.

Nakabayashi:2012:CMH

[Nak12] Jun Nakabayashi. A compartmentalization model of hepatitis C virus replication: An appropriate distribution of HCV RNA for the effective replication. Journal of Theoretical Biology, 300(??):110–117, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000331.

Nakahashi:2014:ECI

[Nak14] Wataru Nakahashi. The effect of cultural interaction on cumulative cultural evolution. Journal of Theoretical Biology, 352(??):6–15, July 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001076.

Nakabayashi:2016:IDH

[Nak16a] Jun Nakabayashi. The intracellular dynamics of hepatitis B virus (HBV) replication with reproduced virion “recycling”. Journal of Theoretical Biology, 396(??):154–162, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316000989.

Nakahashi:2016:CFO

[Nak16b] Wataru Nakahashi. Coevolution of female ovulatory signals and male-male competition in primates. Journal of Theoretical Biology, 392(??):12–22, March 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005986.

Nakamaru:2016:ECE

[Nak16c] Mayuko Nakamaru. Evolution of costly explicit memory and cumulative culture. Journal of Theoretical Biology, 399 (??):71–83, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316001557.

Nakahashi:2019:CSL

[Nak19] Wataru Nakahashi. Cultural skill and language: How structuration affects cultural evolution. Journal of Theoretical Biology, 471(??):13–21, June 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301262.

Naasell:2012:RCC

[Nås12] Ingemar Nåsell. Recurrence conditions for childhood infections. Journal of Theoretical Biology, 313(??):212–216, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519312003840.

Nguyen:2018:CCT

[NBA^{+}18] Trong Hieu Nguyen, Timothée Brochier, Pierre Auger, Viet Duoc Trinh, and Patrice Brehmer. Competition or cooperation in transboundary fish stocks management: Insight from a dynamical model. Journal of Theoretical Biology, 447(??):1–11, June 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301310.

Nikin-Beers:2018:UWH

[NBBCC18] Ryan Nikin-Beers, Julie C. Blackwood, Lauren M. Childs, and Stanca M. Ciupe. Unraveling within-host signatures of dengue infection at the population level. Journal of Theoretical Biology, 446(??):79–86, June 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301188.

Niu:2016:CMB

[NBC16a] Yuanling Niu, Kevin Burrage, and Luonan Chen. Corrigendum for “modelling biochemical reaction systems by stochastic differential equations with reflection”. Journal of Theoretical Biology, 396(??):207–209, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001648.

Niu:2016:MBR

[NBC16b] Yuanling Niu, Kevin Burrage, and Luonan Chen. Modelling biochemical reaction systems by stochastic differential equations with reflection. Journal of Theoretical Biology, 396(??):90–104, May 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001004.

Nanni:2010:HPS

[NBL10] Loris Nanni, Sheryl Brahnam, and Alessandra Lumini. High performance set of PseAAC and sequence based descriptors for protein classification. Journal of Theoretical Biology, 266(1):1–10, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002900.

Nanni:2014:PPS

[NBL14] Loris Nanni, Sheryl Brahnam, and Alessandra Lumini. Prediction of protein structure classes by incorporating different protein descriptors into general Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 360(??):109–116, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003968.

Nandi:2013:ECS

[NBS^{+}13] Anjan K. Nandi, Anindita Bhadra, Annagiri Sumana, Sujata A. Deshpande, and Raghavendra Gadagkar. The evolution of complexity in social organization — a model using dominance-subordinate behavior in two social wasp species. Journal of Theoretical Biology, 327(??):34–44, June 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000350.

Nouvellet:2010:TLA

[NBW10] P. Nouvellet, J. P. Bacon, and D. Waxman. Testing the level of ant activity associated with quorum sensing: An empirical approach leading to the establishment and test of a nullmodel. Journal of Theoretical Biology, 266(4):573–583, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003590.

Nouvellet:2011:TLA

[NBW11] P. Nouvellet, J. P. Bacon, and D. Waxman. Testing the level of ant activity associated with quorum sensing: An empirical approach leading to the establishment and test of a null-model (response to the comment of Richardson et al.). Journal of Theoretical Biology, 280(1):191–193, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001810.

Novak:2015:FDD

[NC15] Sebastian Novak and Sylvia Cremer. Fungal disease dynamics in insect societies: Optimal killing rates and the ambivalent effect of high social interaction rates. Journal of Theoretical Biology, 372(??):54–64, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000776.

Nabawy:2016:OHW

[NC16] Mostafa R. A. Nabawy and William J. Crowther. Optimum hovering wing planform. Journal of Theoretical Biology, 406(??):187–191, October 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301588.

Nardini:2016:MKW

[NCLB16] John T. Nardini, Douglas A. Chapnick, Xuedong Liu, and David M. Bortz. Modeling keratinocyte wound healing dynamics: cell–cell adhesion promotes sustained collective migration. Journal of Theoretical Biology, 400(??):103–117, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300431.

Novak:2013:DG

[NCN13] Sebastian Novak, Krishnendu Chatterjee, and Martin A. Nowak. Density games. Journal of Theoretical Biology, 334(??):26–34, October 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002609.

Nguyen:2013:REC

[NCS^{+}13] Phuong T. T. Nguyen, Alan J. Conley, James Sneyd, Rita S. F. Lee, Tanya K. Soboleva, and Paul R. Shorten. The role of enzyme compartmentalization on the regulation of steroid synthesis. Journal of Theoretical Biology, 332(??):52–64, September 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001896.

Nguyen:2014:IRP

[NDG14] H. Nguyen, P. Dayan, and G. J. Goodhill. The influence of receptor positioning on chemotactic information. Journal of Theoretical Biology, 360(??):95–101, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003658.

Ngoc:2010:CSC

[NdlPZA10] Doanh Nguyen Ngoc, Rafael Bravo de la Parra, Miguel A. Zavala, and Pierre Auger. Competition and species coexistence in a metapopulation model: Can fast asymmetric migration reverse the outcome of competition in a homogeneous environment? Journal of Theoretical Biology, 266(2):256–263, September 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003103.

Nakamura:2013:SIE

[NdMLLBLC13] Miguel Nakamura, Pablo del Monte-Luna, Daniel Lluch-Belda, and Salvador E. Lluch-Cota. Statistical inference for extinction rates based on last sightings. Journal of Theoretical Biology, 333(??):166–173, September 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002440.

Neto:2011:ESC

[NdORC11] Elias D. C. Neto, Viviane M. de Oliveira, Alexandre Rosas, and Paulo R. A. Campos. The effect of spatially correlated environments on genetic diversity-area relationships. Journal of Theoretical Biology, 288(??):57–65, November 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004140.

Ngonghala:2014:QID

[NDZMA14] Calistus N. Ngonghala, Sara Y. Del Valle, Ruijun Zhao, and Jemal Mohammed-Awel. Quantifying the impact of decay in bed-net efficacy on malaria transmission. Journal of Theoretical Biology, 363(??):247–261, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004731.

Neill:2014:EL

[Nei14] David Neill. Evolution of lifespan. Journal of Theoretical Biology, 358(??):232–245, October 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003579.

Nemzer:2017:SIE

[Nem17] Louis R. Nemzer. Shannon information entropy in the canonical genetic code. Journal of Theoretical Biology, 415(??):158–170, February 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316304209.

Newey:2011:OOT

[New11] Philip Newey. Not one odour but two: a new model for nestmate recognition. Journal of Theoretical Biology, 270(1):7–12, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005631.

Nabaei:2014:MMC

[NF14a] Malikeh Nabaei and Nasser Fatouraee. Microstructural modelling of cerebral aneurysm evolution through effective stress mediated destructive remodelling. Journal of Theoretical Biology, 354(??):60–71, August 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001556.

Nakahashi:2014:EDL

[NF14b] Wataru Nakahashi and Marcus W. Feldman. Evolution of division of labor: Emergence of different activities among group members. Journal of Theoretical Biology, 348(??):65–79, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000447.

Neto:2016:SMS

[NF16] Manoel Figueiredo Neto and Marxa L. Figueiredo. Skeletal muscle signal peptide optimization for enhancing propeptide or cytokine secretion. Journal of Theoretical Biology, 409(??):11–17, November 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302764.

Ng:2017:MVD

[NF17] Chin F. Ng and Hermann B. Frieboes. Model of vascular desmoplastic multispecies tumor growth. Journal of Theoretical Biology, 430(??):245–282, October 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302199.

Nelson:2011:SRD

[NG11] Ronald M. Nelson and Jaco M. Greeff. Sex ratio dependent dispersal when sex ratios vary between patches. Journal of Theoretical Biology, 290(??):81–87, December 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004334.

North:2017:DDM

[NG17] Ace R. North and H. Charles J. Godfray. The dynamics of disease in a metapopulation: The role of dispersal range. Journal of Theoretical Biology, 418(??):57–65, April 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300371.

Nezafat:2014:NME

[NGJ^{+}14] Navid Nezafat, Younes Ghasemi, Gholamreza Javadi, Mohammad Javad Khoshnoud, and Eskandar Omidinia. A novel multi-epitope peptide vaccine against cancer: An in silico approach. Journal of Theoretical Biology, 349(??):121–134, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000356.

Nie:2010:CMC

[NGL^{+}10] Jingxin Nie, Lei Guo, Gang Li, Carlos Faraco, L. Stephen Miller, and Tianming Liu. A computational model of cerebral cortex folding. Journal of Theoretical Biology, 264(2):467–478, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000615.

Neigenfind:2013:RBR

[NGN13] Jost Neigenfind, Sergio Grimbs, and Zoran Nikoloski. On the relation between reactions and complexes of (bio)chemical reaction networks. Journal of Theoretical Biology, 317(??): 359–365, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519312005498.

Namboodiri:2012:LSB

[NGND12] Saritha Namboodiri, Alessandro Giuliani, Achuthsankar S. Nair, and Pawan K. Dhar. Looking for a sequence based allostery definition: a statistical journey at different resolution scales. Journal of Theoretical Biology, 304(??):211–218, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001221.

Numminen:2016:IHM

[NGS^{+}16] Elina Numminen, Michael Gutmann, Mikhail Shubin, Pekka Marttinen, Guillaume Méric, Willem van Schaik, Teresa M. Coque, Fernando Baquero, Rob J. L. Willems, Samuel K. Sheppard, Edward J. Feil, William P. Hanage, and Jukka Corander. The impact of host metapopulation structure on the population genetics of colonizing bacteria. Journal of Theoretical Biology, 396(??):53–62, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001156.

Nakahashi:2012:EAH

[NH12] Wataru Nakahashi and Shiro Horiuchi. Evolution of ape and human mating systems. Journal of Theoretical Biology, 296(??):56–64, March 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005960.

Nakanishi:2019:PSH

[NH19] Ayako Nakanishi and Yoshito Hirata. Practically scheduling hormone therapy for prostate cancer using a mathematical model. Journal of Theoretical Biology, 478(??):48–57, October 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302395.

Nessler:2016:IHP

[NHE^{+}16] K. H. L. Neßler, J. R. Henstock, A. J. El Haj, S. L. Waters, J. P. Whiteley, and J. M. Osborne. The influence of hydrostatic pressure on tissue engineered bone development. Journal of Theoretical Biology, 394(??):149–159, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000254.

Nitzany:2010:QBV

[NHM10] Eyal Nitzany, Ilan Hammel, and Isaac Meilijson. Quantal basis of vesicle growth and information content, a unified approach. Journal of Theoretical Biology, 266(1):202–209, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003097.

Nikolaou:2016:DCP

[NHS^{+}16] Andreas Nikolaou, Philipp Hartmann, Antoine Sciandra, Benoît Chachuat, and Olivier Bernard. Dynamic coupling of photoacclimation and photoinhibition in a model of microalgae growth. Journal of Theoretical Biology, 390(??):61–72, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005408.

Niu:2014:PDB

[NHSX14] Xiao-Hui Niu, Xue-Hai Hu, Feng Shi, and Jing-Bo Xia. Predicting DNA binding proteins using support vector machine with hybrid fractal features. Journal of Theoretical Biology, 343(??):186–192, February 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005018.

Nosaka:2014:PFF

[NHTS14] Michiko Nosaka, Katsuki Hirata, Ryotarou Tsuji, and Syunya Sunaba. Planes formed with four intron-positions in tertiary structures of retinol binding protein and calpain domain VI. Journal of Theoretical Biology, 340(??):139–145, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004177.

Nishiura:2011:EIP

[NI11] Hiroshi Nishiura and Hisashi Inaba. Estimation of the incubation period of influenza A (H1N1-2009) among imported cases: Addressing censoring using outbreak data at the origin of importation. Journal of Theoretical Biology, 272(1):123–130, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006727.

Nojo:2019:ESS

[NI19] Saori Nojo and Yasuo Ihara. The effect of sexual selection on phenotypic diversification among human populations: a simulation study. Journal of Theoretical Biology, 462(??):1–11, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305460.

Nicholson:2019:CRM

[Nic19] Daniel J. Nicholson. Is the cell really a machine? Journal of Theoretical Biology, 477(??):108–126, September 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319302292.

Nagatani:2018:ERW

[NIiT18a] Takashi Nagatani, Genki Ichinose, and Kei ichi Tainaka. Epidemics of random walkers in metapopulation model for complete, cycle, and star graphs. Journal of Theoretical Biology, 450(??):66–75, August 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302078.

Nagatani:2018:MMRa

[NIiT18b] Takashi Nagatani, Genki Ichinose, and Kei ichi Tainaka. Metapopulation model for rock-paper-scissors game: Mutation affects paradoxical impacts. Journal of Theoretical Biology, 450(??):22–29, August 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301656.

Nagatani:2019:MDR

[NIiT19] Takashi Nagatani, Genki Ichinose, and Kei ichi Tainaka. Metapopulation dynamics in the rock-paper-scissors game with mutation: Effects of time-varying migration paths. Journal of Theoretical Biology, 462(??):425–431, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305836.

Nijhout:2011:DMA

[Nij11] H. Frederik Nijhout. Dependence of morphometric allometries on the growth kinetics of body parts. Journal of Theoretical Biology, 288(??):35–43, November 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004036.

Nikolic:2015:PHL

[Nik15] Danko Nikolić. Practopoiesis: Or how life fosters a mind. Journal of Theoretical Biology, 373(??):40–61, May 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500106X.

Nagatani:2018:MMRb

[NiTI18] Takashi Nagatani, Kei ichi Tainaka, and Genki Ichinose. Metapopulation model of rock-scissors-paper game with subpopulation-specific victory rates stabilized by heterogeneity. Journal of Theoretical Biology, 458(??):103–110, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304399.

Norton:2018:MTN

[NJP18] Kerri-Ann Norton, Kideok Jin, and Aleksander S. Popel. Modeling triple-negative breast cancer heterogeneity: Effects of stromal macrophages, fibroblasts and tumor vasculature. Journal of Theoretical Biology, 452(??):56–68, September 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318302297.

Nath:2018:EPR

[NK18] Abhigyan Nath and S. Karthikeyan. Enhanced prediction of recombination hotspots using input features extracted by class specific autoencoders. Journal of Theoretical Biology, 444(??):73–82, May 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300766.

Neofytou:2016:MMP

[NKB16a] G. Neofytou, Y. N. Kyrychko, and K. B. Blyuss. Mathematical model of plant-virus interactions mediated by RNA interference. Journal of Theoretical Biology, 403(??):129–142, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300960.

Neofytou:2016:TDM

[NKB16b] G. Neofytou, Y. N. Kyrychko, and K. B. Blyuss. Timedelayed model of immune response in plants. Journal of Theoretical Biology, 389(??):28–39, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005196.

Nah:2010:DED

[NKL10] Kyeongah Nah, Yongkuk Kim, and Jung Min Lee. The dilution effect of the domestic animal population on the transmission of P. vivax malaria. Journal of Theoretical Biology, 266(2):299–306, September 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003231.

Netzer:2012:NBF

[NKM^{+}12] Michael Netzer, Karl G. Kugler, Laurin A. J. Müller, Klaus M. Weinberger, Armin Graber, Christian Baumgartner, and Matthias Dehmer. A network-based feature selection approach to identify metabolic signatures in disease. Journal of Theoretical Biology, 310(??):216–222, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002809.

Nordon:2011:MTB

[NKOS11] Robert E. Nordon, Kap-Hyoun Ko, Ross Odell, and Timm Schroeder. Multi-type branching models to describe cell differentiation programs. Journal of Theoretical Biology, 277(1):7– 18, May 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001007.

Nanni:2014:SDI

[NLB14] Loris Nanni, Alessandra Lumini, and Sheryl Brahnam. A set of descriptors for identifying the protein-drug interaction in cellular networking. Journal of Theoretical Biology, 359(??):120–128, October 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003452.

Nyabadza:2011:MHA

[NM11] F. Nyabadza and Z. Mukandavire. Modelling HIV/AIDS in the presence of an HIV testing and screening campaign. Journal of Theoretical Biology, 280(1):167–179, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002219.

Ngonghala:2016:IBI

[NMAZP16] Calistus N. Ngonghala, Jemal Mohammed-Awel, Ruijun Zhao, and Olivia Prosper. Interplay between insecticide-treated bednets and mosquito demography: implications for malaria control. Journal of Theoretical Biology, 397(??):179–192, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316001430.

Nene:2018:EGD

[NMI18] Nuno R. Nené, Ville Mustonen, and Christopher J. R. Illingworth. Evaluating genetic drift in time-series evolutionary analysis. Journal of Theoretical Biology, 437(??):51–57, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304332.

Nikoloski:2010:ACM

[NMS10] Zoran Nikoloski, Patrick May, and Joachim Selbig. Algebraic connectivity may explain the evolution of gene regulatory networks. Journal of Theoretical Biology, 267(1):7–14, November 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310003851.

Ning:2019:DKP

[NMZ19] Qiao Ning, Zhiqiang Ma, and Xiaowei Zhao. dForml(KNN)- PseAAC: Detecting formylation sites from protein sequences using K-nearest neighbor algorithm via Chou’s 5step rule and pseudo components. Journal of Theoretical Biology, 470(??):43–49, June 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301122.

Nicolis:2018:FNN

[NN18] G. Nicolis and S. C. Nicolis. Feedbacks, nonlinearities and nonequilibria: a thermodynamic perspective. Journal of Theoretical Biology, 458(??):1–9, December 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304272.

Nierop:2016:MIL

[NNH^{+}16] Andreas F. M. Nierop, Aimon Niklasson, Anton Holmgren, Lars Gelander, Sten Rosberg, and Kerstin Albertsson-Wikland. Modelling individual longitudinal human growth from fetal to adult life — QEPS I. Journal of Theoretical Biology, 406(??):143–165, October 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301424.

Noonburg:2010:ELH

[NNK10] Erik G. Noonburg, Roger M. Nisbet, and Tin Klanjscek. Effects of life history variation on vertical transfer of toxicants in marine mammals. Journal of Theoretical Biology, 264(2):479– 489, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000925.

Nasiri:2015:PAS

[NNK^{+}15] Jaber Nasiri, Mohammad Reza Naghavi, Amir Hossein Kayvanjoo, Mojtaba Nasiri, and Mansour Ebrahimi. Precision assessment of some supervised and unsupervised algorithms for genotype discrimination in the genus Pisum using SSR molecular data. Journal of Theoretical Biology, 368(??):122–132, March 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315000028.

Nev:2017:OMN

[NNvdB17] Olga A. Nev, Oleg A. Nev, and Hugo A. van den Berg. Optimal management of nutrient reserves in microorganisms under time-varying environmental conditions. Journal of Theoretical Biology, 429(??):124–141, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302862.

Nakamura:2014:EIR

[NO14a] Mitsuhiro Nakamura and Hisashi Ohtsuki. Erratum to “Indirect reciprocity in three types of social dilemmas” [J. Theor. Biol. 355 (2014) 117–127]. Journal of Theoretical Biology, 361(??):206, November 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004482. See [NO14b].

Nakamura:2014:IRT

[NO14b] Mitsuhiro Nakamura and Hisashi Ohtsuki. Indirect reciprocity in three types of social dilemmas. Journal of Theoretical Biology, 355(??):117–127, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400191X. See erratum [NO14a].

Nakahashi:2015:WEC

[NO15] Wataru Nakahashi and Hisashi Ohtsuki. When is emotional contagion adaptive? Journal of Theoretical Biology, 380(??):480–488, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002933.

Nakahashi:2018:EEC

[NO18] Wataru Nakahashi and Hisashi Ohtsuki. Evolution of emotional contagion in group-living animals. Journal of Theoretical Biology, 440(??):12–20, March 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305520.

Nonacs:2010:BTW

[Non10] Peter Nonacs. Bordered tug-of-war models are neither general nor predictive of reproductive skew. Journal of Theoretical Biology, 266(4):739–741, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003863.

Noshita:2014:QGA

[Nos14] Koji Noshita. Quantification and geometric analysis of coiling patterns in gastropod shells based on 3D and 2D image data. Journal of Theoretical Biology, 363(??):93–104, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004585.

Nishimura:2017:EFC

[NOS17] Takeshi Nishimura, Akira Okada, and Yasuhiro Shirata. Evolution of fairness and coalition formation in three-person ultimatum games. Journal of Theoretical Biology, 420(??):53– 67, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300978.

Nowak:2012:EC

[Now12] Martin A. Nowak. Evolving cooperation. Journal of Theoretical Biology, 299(??):1–8, April 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200015X.

Nurmi:2011:JES

[NP11] Tuomas Nurmi and Kalle Parvinen. Joint evolution of specialization and dispersal in structured metapopulations. Journal of Theoretical Biology, 275(1):78–92, April 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100035X.

Nurmi:2013:ESU

[NP13] Tuomas Nurmi and Kalle Parvinen. Evolution of specialization under non-equilibrium population dynamics. Journal of Theoretical Biology, 321(??):63–77, March 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006546.

Neupane:2015:ISA

[NP15] Ram C. Neupane and James A. Powell. Invasion speeds with active dispersers in highly variable landscapes: Multiple scales, homogenization, and the migration of trees. Journal of Theoretical Biology, 387(??):111–119, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315004798.

Nonaka:2013:ESC

[NPB13] Etsuko Nonaka, Kalle Parvinen, and Åke Brännström. Evolutionary suicide as a consequence of runaway selection for greater aggregation tendency. Journal of Theoretical Biology, 317(??):96–104, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005115.

Niccum:2012:UCE

[NPH^{+}12] Brittany A. Niccum, Roby Poteau, Glen E. Hamman, Jan C. Varada, Jewgeni H. Dshalalow, and Richard R. Sinden. On an unbiased and consistent estimator for mutation rates. Journal of Theoretical Biology, 300(??):360–367, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000392.

Nicolas:2015:MMF

[NPMM15] M. Nicolás, E. Peña, M. Malvè, and M. A. Martínez. Mathematical modeling of the fibrosis process in the implantation of inferior vena cava filters. Journal of Theoretical Biology, 387(??):228–240, December 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004786.

Nevai:2010:SCH

[NPS10] Andrew L. Nevai, Kevin M. Passino, and Parthasarathy Srinivasan. Stability of choice in the honey bee nest-site selection process. Journal of Theoretical Biology, 263(1):93–107, March 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005323.

Nurmi:2017:ESS

[NPS17] Tuomas Nurmi, Kalle Parvinen, and Vesa Selonen. The evolution of site-selection strategy during dispersal. Journal of Theoretical Biology, 425(??):11–22, July 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302060.

Nurmi:2018:JED

[NPS18] Tuomas Nurmi, Kalle Parvinen, and Vesa Selonen. Joint evolution of dispersal propensity and site selection in structured metapopulation models. Journal of Theoretical Biology, 444(??):50–72, May 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300638.

Neilan:2014:SEF

[NR14] Rachael Miller Neilan and Kenneth Rose. Simulating the effects of fluctuating dissolved oxygen on growth, reproduction, and survival of fish and shrimp. Journal of Theoretical Biology, 343(??):54–68, February 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005298.

Nax:2016:AES

[NR16] Heinrich H. Nax and Alexandros Rigos. Assortativity evolving from social dilemmas. Journal of Theoretical Biology, 395(??):194–203, April 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000710.

Nath:2018:DWC

[NRKE18] Madhurima Nath, Yihui Ren, Yasamin Khorramzadeh, and Stephen Eubank. Determining whether a class of random graphs is consistent with an observed contact network. Journal of Theoretical Biology, 440(??):121–132, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730560X.

Naushad:2016:SAI

[NRS^{+}16] Shaik Mohammad Naushad, M. Janaki Ramaiah, Balraj Alex Stanley, S. Prasanna Lakshmi, J. Vishnu Priya, Tajamul Hussain, Salman A. Alrokayan, and Vijay Kumar Kutala. In silico approaches to identify the potential inhibitors of glutamate carboxypeptidase II (GCPII) for neuroprotection. Journal of Theoretical Biology, 406(??):137–142, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301990.

Nakabayashi:2011:MMI

[NS11] Jun Nakabayashi and Akira Sasaki. A mathematical model of the intracellular replication and within host evolution of hepatitis type B virus: Understanding the long time course of chronic hepatitis. Journal of Theoretical Biology, 269(1):318–329, January 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005588.

Nov:2013:OCR

[NS13] Yuval Nov and Danny Segev. Optimal codon randomization via mathematical programming. Journal of Theoretical Biology, 335(??):147–152, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002658.

Nath:2016:IMB

[NS16] Abhigyan Nath and Karthikeyan Subbiah. Insights into the molecular basis of piezophilic adaptation: Extraction of piezophilic signatures. Journal of Theoretical Biology, 390(??):117–126, February 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500572X.

Namba:2018:PRF

[NS18] Toshinori Namba and Tatsuo Shibata. Propagation of regulatory fluctuations induces coordinated switching of flagellar motors in chemotaxis signaling pathway of single bacteria. Journal of Theoretical Biology, 454(??):367–375, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303217.

Nath:2019:EEL

[NS19] Abhigyan Nath and Gopal Krishna Sahu. Exploiting ensemble learning to improve prediction of phospholipidosis inducing potential. Journal of Theoretical Biology, 479(??):37–47, October 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302863.

Nanni:2010:PCU

[NSBL10] Loris Nanni, Jian-Yu Shi, Sheryl Brahnam, and Alessandra Lumini. Protein classification using texture descriptors extracted from the protein backbone image. Journal of Theoretical Biology, 264(3):1024–1032, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001438.

Noble:2010:UAM

[NSH^{+}10] Sarah L. Noble, Eric Sherer, Robert E. Hannemann, Doraiswami Ramkrishna, Terry Vik, and Ann E. Rundell. Using adaptive model predictive control to customize maintenance therapy chemotherapeutic dosing for childhood acute lymphoblastic leukemia. Journal of Theoretical Biology, 264(3):990–1002, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000597.

Neidhart:2013:ERA

[NSK13] Johannes Neidhart, Ivan G. Szendro, and Joachim Krug. Exact results for amplitude spectra of fitness landscapes. Journal of Theoretical Biology, 332(??):218–227, September 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002178.

Nakamaru:2018:ESE

[NSKO18] Mayuko Nakamaru, Hayato Shimura, Yoko Kitakaji, and Susumu Ohnuma. The effect of sanctions on the evolution of cooperation in linear division of labor. Journal of Theoretical Biology, 437(??):79–91, January 21, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304708.

Neeman:2015:SPB

[NSO15] Noga Neeman, James R. Spotila, and Michael P. O’Connor. A simple, physiologically-based model of sea turtle remigration intervals and nesting population dynamics: Effects of temperature. Journal of Theoretical Biology, 380(??):516–523, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003021.

Napoletani:2011:HCP

[NSS^{+}11] D. Napoletani, M. Signore, T. Sauer, L. Liotta, and E. Petricoin. Homologous control of protein signaling networks. Journal of Theoretical Biology, 279(1):29–43, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001664.

Noshita:2016:GAE

[NSS16] Koji Noshita, Keisuke Shimizu, and Takenori Sasaki. Geometric analysis and estimation of the growth rate gradient on gastropod shells. Journal of Theoretical Biology, 389(??):11–19, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005044.

Nemeth:2010:PCB

[NT10] A. Németh and K. Takács. The paradox of cooperation benefits. Journal of Theoretical Biology, 264(2):301–311, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931000072X.

Neal:2014:LMC

[NT14] Allison T. Neal and Peter D. Taylor. Local mate competition and transmission bottlenecks: a new model for understanding malaria parasite and other sex ratios. Journal of Theoretical Biology, 363(??):381–389, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004986.

Nakorn:2011:VSB

[NTC^{+}11] Poonsap Na Nakorn, Witcha Treesuwan, Kiattawee Choowongkomon, Supa Hannongbua, and Nonlawat Boonyalai. In vitro and in silico binding study of the peptide derived from HIV-1 CA-CTD and LysRS as a potential HIV-1 blocking site. Journal of Theoretical Biology, 270(1):88–97, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005953.

Ngwa:2019:TSS

[NTED^{+}19] Gideon A. Ngwa, Miranda I. Teboh-Ewungkem, Yves Dumont, Rachid Ouifki, and Jacek Banasiak. On a threestage structured model for the dynamics of malaria transmission with human treatment, adult vector demographics and one aquatic stage. Journal of Theoretical Biology, 481(??):202–222, November 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318306490.

Noble:2011:STA

[NTFK11] Andrew E. Noble, Nico M. Temme, William F. Fagan, and Timothy H. Keitt. A sampling theory for asymmetric communities. Journal of Theoretical Biology, 273(1):1–14, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006764.

Nishikawa:2016:MMC

[NTOI16] Seiya Nishikawa, Atsuko Takamatsu, Shizue Ohsawa, and Tatsushi Igaki. Mathematical model for cell competition: predator–prey interactions at the interface between two groups of cells in monolayer tissue. Journal of Theoretical Biology, 404(??):40–50, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301217.

Nguyen:2016:OAP

[NTY16] Linh Thi Hoai Nguyen, Vi{ě \atop .}t Tôn Ta., and Atsushi Yagi. Obstacle avoiding patterns and cohesiveness of fish school. Journal of Theoretical Biology, 406(??):116–123, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302004.

Nakauma:2017:RGF

[NvD17] Alberto Nakauma and G. Sander van Doorn. Reconstructing the genotype-to-fitness map for the bacterial chemotaxis network and its emergent behavioural phenotypes. Journal of Theoretical Biology, 420(??):200–212, May 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301340.

Nathan:2013:SIU

[NvHM13] Jonathan Nathan, Jost von Hardenberg, and Ehud Meron. Spatial instabilities untie the exclusion-principle constraint on species coexistence. Journal of Theoretical Biology, 335 (??):198–204, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519313002981.

Nian:2010:EIS

[NW10] Fuzhong Nian and Xingyuan Wang. Efficient immunization strategies on complex networks. Journal of Theoretical Biology, 264(1):77–83, May 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000111.

Norton:2010:MMB

[NWB^{+}10] Kerri-Ann Norton, Michael Wininger, Gyan Bhanot, Shridar Ganesan, Nicola Barnard, and Troy Shinbrot. A 2d mechanistic model of breast ductal carcinoma in situ (DCIS) morphology and progression. Journal of Theoretical Biology, 263(4):393–406, April 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005669.

Nicholson:2017:MSE

[NWWL17] Arwen E. Nicholson, David M. Wilkinson, Hywel T. P. Williams, and Timothy M. Lenton. Multiple states of environmental regulation in well-mixed model biospheres. Journal of Theoretical Biology, 414(??):17–34, February 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303848.

Nicholson:2018:AMG

[NWWL18] Arwen E. Nicholson, David M. Wilkinson, Hywel T. P. Williams, and Timothy M. Lenton. Alternative mechanisms for Gaia. Journal of Theoretical Biology, 457(??): 249–257, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519318304132.

Niu:2015:PEC

[NWZ15] Yuanling Niu, Yue Wang, and Da Zhou. The phenotypic equilibrium of cancer cells: From average-level stability to path-wise convergence. Journal of Theoretical Biology, 386(??):7–17, December 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004300.

Nishiura:2012:ETP

[NYSM12] Hiroshi Nishiura, Ping Yan, Candace K. Sleeman, and Charles J. Mode. Estimating the transmission potential of supercritical processes based on the final size distribution of minor outbreaks. Journal of Theoretical Biology, 294(??):48–55, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005625.

Niu:2019:MDS

[NZZ19] Rui-Juan Niu, Qing-Chuan Zheng, and Hong-Xing Zhang. Molecular dynamics simulations study of influence of tyr422ala mutation on transcriptional enhancer activation domain 4 (TEAD4) and transcription co-activators complexes. Journal of Theoretical Biology, 472(??):27–35, July 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931930150X.

Omori:2011:DSC

[OA11] Ryosuke Omori and Ben Adams. Disrupting seasonality to control disease outbreaks: The case of koi herpes virus. Journal of Theoretical Biology, 271(1):159–165, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006466.

Okamoto:2012:BCD

[OA12] Kenichi W. Okamoto and Priyanga Amarasekare. The biological control of disease vectors. Journal of Theoretical Biology, 309(??):47–57, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002627.

Ogino:2015:MCC

[OA15] Yoshiyuki Ogino and Toru Asahi. Model of complex chiral drug metabolic systems and numerical simulation of the remaining chirality toward analysis of dynamical pharmacological activity. Journal of Theoretical Biology, 373(??):117–131, May 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001149.

Oprisan:2018:STM

[OABB18] Sorinel A. Oprisan, Tristan Aft, Mona Buhusi, and Catalin V. Buhusi. Scalar timing in memory: a temporal map in the hippocampus. Journal of Theoretical Biology, 438(??):133–142, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305192.

Orman:2012:MNA

[OABI12] Mehmet A. Orman, Ioannis P. Androulakis, Francois Berthiaume, and Marianthi G. Ierapetritou. Metabolic network analysis of perfused livers under fed and fasted states: Incorporating thermodynamic and futile-cycle-associated regulatory constraints. Journal of Theoretical Biology, 293(??):101–110, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005340.

Ouattara:2010:SDF

[OAJK10] Djomangan A. Ouattara, Wassim Abou-Jaoudé, and Marcelle Kaufman. From structure to dynamics: Frequency tuning in the p53-Mdm2 network. II: Differential and stochastic approaches. Journal of Theoretical Biology, 264(4):1177–1189, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001645.

Okamoto:2014:MOV

[OAP14] Kenichi W. Okamoto, Priyanga Amarasekare, and Ian T. D. Petty. Modeling oncolytic virotherapy: Is complete tumortropism too much of a good thing? Journal of Theoretical Biology, 358(??):166–178, October 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002604.

Omori:2010:CCS

[OAS10] Ryosuke Omori, Ben Adams, and Akira Sasaki. Coexistence conditions for strains of influenza with immune crossreaction. Journal of Theoretical Biology, 262(1):48–57, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004354.

Olejarz:2016:EWP

[OAV^{+}16] Jason W. Olejarz, Benjamin Allen, Carl Veller, Raghavendra Gadagkar, and Martin A. Nowak. Evolution of worker policing. Journal of Theoretical Biology, 399(??):103–116, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316001417.

Olofsson:2010:MGT

[OB10] Peter Olofsson and Alison A. Bertuch. Modeling growth and telomere dynamics in Saccharomyces cerevisiae. Journal of Theoretical Biology, 263(3):353–359, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005712.

OBrien:2015:PCF

[OB15] Haley D. O’Brien and Jason Bourke. Physical and computational fluid dynamics models for the hemodynamics of the artiodactyl carotid rete. Journal of Theoretical Biology, 386(??):122–131, December 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004518.

Orman:2011:PAL

[OBAI11] Mehmet A. Orman, Francois Berthiaume, Ioannis P. Androulakis, and Marianthi G. Ierapetritou. Pathway analysis of liver metabolism under stressed condition. Journal of Theoretical Biology, 272(1):131–140, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006375.

Oborny:2017:PLA

[OBE^{+}17] Beáta Oborny, Veronika Benedek, Péter Englert, Máté Gulyás, and András G. Hubai. The plant in the labyrinth: Adaptive growth and branching in heterogeneous environments. Journal of Theoretical Biology, 412(??):146–153, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303484.

Overton:2019:MAS

[OBHS19] Christopher E. Overton, Mark Broom, Christoforos Hadjichrysanthou, and Kieran J. Sharkey. Methods for approximating stochastic evolutionary dynamics on graphs. Journal of Theoretical Biology, 468(??):45–59, May 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300724.

Orlova:2011:DFA

[OBK^{+}11] Darya Yu. Orlova, Vyacheslav I. Borisov, Vladimir S. Kozhevnikov, Valeri P. Maltsev, and Andrei V. Chernyshev. Distribution function approach to the study of the kinetics of IgM antibody binding to FcγRIIIb (CD16b) receptors on neutrophils by flow cytometry. Journal of Theoretical Biology, 290(??):1–6, December 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004292.

Oyarzun:2012:MOG

[OCHHZ12] Diego A. Oyarzún, Madalena Chaves, and Marit Hoff-Hoffmeyer-Zlotnik. Multistability and oscillations in genetic control of metabolism. Journal of Theoretical Biology, 295(??):139–153, February 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931100587X.

Oka:2010:CCK

[OCN10] Chiaki Oka, Chae Young Cha, and Akinori Noma. Characterization of the cardiac Na^{+} /K^{+} pump by development of a comprehensive and mechanistic model. Journal of Theoretical Biology, 265(1):68–77, July 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002274.

Obermayer:2010:ETS

[OF10] Benedikt Obermayer and Erwin Frey. Error thresholds for self- and cross-specific enzymatic replication. Journal of Theoretical Biology, 267(4):653–662, December 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000490X.

OFallon:2011:TOM

[O’F11] Brendan O’Fallon. Two optimal mutation rates in obligate pathogens subject to deleterious mutation. Journal of Theoretical Biology, 276(1):150–158, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000452.

OBrien:2010:MGA

[OFB10] Ricky O’Brien, Nev Fowkes, and Andrew P. Bassom. Models for gibberellic acid transport and enzyme production and transport in the aleurone layer of barley. Journal of Theoretical Biology, 267(1):15–21, November 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003875.

Ohara:2015:EMM

[OFT15] Takayuki Ohara, Hirokazu Fukuda, and Isao T. Tokuda. An extended mathematical model for reproducing the phase response of Arabidopsis thaliana under various light conditions. Journal of Theoretical Biology, 382(??):337–344, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315003537.

Olsson:2016:DDD

[OGA16] Karin H. Olsson, Henrik Gislason, and Ken Haste Andersen. Differences in density-dependence drive dual offspring size strategies in fish. Journal of Theoretical Biology, 407(??):118–127, October 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302090.

Ogunnaike:2010:PFM

[OGE10] Babatunde A. Ogunnaike, Claudio A. Gelmi, and Jeremy S. Edwards. A probabilistic framework for microarray data analysis: Fundamental probability models and statistical inference. Journal of Theoretical Biology, 264(2):211–222, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310000962.

Oteo-Garcia:2019:EDC

[OGO19] Gonzalo Oteo-García and José Angel Oteo. Evolutionary distances corrected for purifying selection and ancestral polymorphisms. Journal of Theoretical Biology, 483 (??):Article 110004, December 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319303741.

Ohtsuki:2010:SED

[Oht10] Hisashi Ohtsuki. Stochastic evolutionary dynamics of bimatrix games. Journal of Theoretical Biology, 264(1):136–142, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000287.

Ohtsuki:2012:DSR

[Oht12] Hisashi Ohtsuki. Does synergy rescue the evolution of cooperation? An analysis for homogeneous populations with non-overlapping generations. Journal of Theoretical Biology, 307(??):20–28, August 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002111.

Ohara:2018:GRN

[OHWS18] Takayuki Ohara, Timothy J. Hearn, Alex A. R. Webb, and Akiko Satake. Gene regulatory network models in response to sugars in the plant circadian system. Journal of Theoretical Biology, 457(??):137–151, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304016.

Okabe:2011:PPP

[Oka11] Takuya Okabe. Physical phenomenology of phyllotaxis. Journal of Theoretical Biology, 280(1):63–75, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001895.

Okabe:2012:SVD

[Oka12] Takuya Okabe. Systematic variations in divergence angle. Journal of Theoretical Biology, 313(??):20–41, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312003931.

Okamoto:2015:DSA

[Oka15] Kenichi W. Okamoto. The dynamics of strangling among forest trees. Journal of Theoretical Biology, 384(??):95–104, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003513.

Okonjo:2015:BEH

[Oko15] Kehinde O. Okonjo. Bohr effect of hemoglobins: Accounting for differences in magnitude. Journal of Theoretical Biology, 380(??):436–443, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003008.

Okonjo:2016:BEA

[Oko16] Kehinde O. Okonjo. Bohr effect of avian hemoglobins: Quantitative analyses based on the Wyman equation. Journal of Theoretical Biology, 410(??):25–35, December 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302855.

Okuyama:2015:DSA

[Oku15] Toshinori Okuyama. Demographic stochasticity alters the outcome of exploitation competition. Journal of Theoretical Biology, 365(??):347–351, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006304.

Olejarz:2018:SSC

[OKVN18] Jason Olejarz, Kamran Kaveh, Carl Veller, and Martin A. Nowak. Selection for synchronized cell division in simple multicellular organisms. Journal of Theoretical Biology, 457(??):170–179, November 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304260.

Osman:2019:AVB

[OL19] Mojeeb Al-Rahman El-Nor Osman and Jinhui Li. Analysis of a vector-bias malaria transmission model with application to Mexico, Sudan and democratic Republic of the Congo. Journal of Theoretical Biology, 464(??):72–84, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318306337.

Ochab-Marcinek:2010:ENP

[OM10] Anna Ochab-Marcinek. Extrinsic noise passing through a Michaelis–Menten reaction: a universal response of a genetic switch. Journal of Theoretical Biology, 263(4):510–520, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309006018.

OMalley:2017:EHE

[O’M17] Maureen A. O’Malley. From endosymbiosis to holobionts: Evaluating a conceptual legacy. Journal of Theoretical Biology, 434(??):34–41, December 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301200.

Onana:2013:EHP

[OMO13] Denis Fils Nkoa Onana, Boulchard Mewoli, and Djomangan Adama Ouattara. Excitability in the host-pathogen interactions of HIV infection and emergence of viral load blips. Journal of Theoretical Biology, 317(??):407–417, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005516.

Olejarz:2014:EST

[ON14] Jason W. Olejarz and Martin A. Nowak. Evolution of staying together in the context of diffusible public goods. Journal of Theoretical Biology, 360(??):1–12, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400366X.

Olah:2015:MSP

[ONO15] Judit Oláh, Vic Norris, and Judit Ovádi. Modeling of sensing potency of cytoskeletal systems decorated with metabolic enzymes. Journal of Theoretical Biology, 365(??):190–196, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006079.

Okuonghae:2011:AMM

[OO11] D. Okuonghae and S. E. Omosigho. Analysis of a mathematical model for tuberculosis: What could be done to increase case detection. Journal of Theoretical Biology, 269(1):31–45, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005278.

Ose:2017:SHN

[OO17a] Nicholas J. Ose and Paul R. Ohmann. The selfish herd: Noise effects in local crowded horizon and Voronoi models. Journal of Theoretical Biology, 424(??):84–90, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301947.

Oya:2017:SPC

[OO17b] Gaku Oya and Hisashi Ohtsuki. Stable polymorphism of cooperators and punishers in a public goods game. Journal of Theoretical Biology, 419(??):243–253, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303770.

OConnor:2018:IIC

[OO18] Michael P. O’Connor and Sean O’Donnell. Implications of iterative communication for biological system performance.

Journal of Theoretical Biology, 436(??):93–104, January 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304423.

Oroji:2016:ISD

[OOY16] Amin Oroji, Mohd Omar, and Shantia Yarahmadian. An Îto stochastic differential equations model for the dynamics of the MCF-7 breast cancer cell line treated by radiotherapy. Journal of Theoretical Biology, 407(??):128–137, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931630217X.

Oprisan:2010:ESC

[Opr10] Sorinel Adrian Oprisan. Existence and stability criteria for phase-locked modes in ring neural networks based on the spike time resetting curve method. Journal of Theoretical Biology, 262(2):232–244, January 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004743.

Ottesen:2019:BBC

[OPS^{+}19] Johnny T. Ottesen, Rasmus K. Pedersen, Zamra Sajid, Johanne Gudmand-Hoeyer, Katrine O. Bangsgaard, Vibe Skov, Lasse Kjær, Trine A. Knudsen, Niels Pallisgaard, Hans C. Hasselbalch, and Morten Andersen. Bridging blood cancers and inflammation: The reduced Cancitis model. Journal of Theoretical Biology, 465(??):90–108, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300013.

Orr:2019:MEN

[Orr19] H. Allen Orr. Measures of the efficiency of natural selection during gene substitution. Journal of Theoretical Biology, 476(??):1–4, September 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302012.

Ooyama:2011:HON

[OS11] Shunsuke Ooyama and Tatsuo Shibata. Hierarchical organization of noise generates spontaneous signal in Paramecium cell. Journal of Theoretical Biology, 283(1):1–9, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100258X.

Omori:2013:TEN

[OS13] Ryosuke Omori and Akira Sasaki. Timing of the emergence of new successful viral strains in seasonal influenza. Journal of Theoretical Biology, 329(??):32–38, July 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001537.

Olson:2011:CBH

[OSF11] Sarah D. Olson, Susan S. Suarez, and Lisa J. Fauci. Coupling biochemistry and hydrodynamics captures hyperactivated sperm motility in a simple flagellar model. Journal of Theoretical Biology, 283(1):203–216, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002888.

Okada:2018:SPA

[OSN18] Isamu Okada, Tatsuya Sasaki, and Yutaka Nakai. A solution for private assessment in indirect reciprocity using solitary observation. Journal of Theoretical Biology, 455(??):7–15, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303138.

Oommen:2016:GWD

[OSV^{+}16] Anup Mammen Oommen, Neeti Somaiya, Jisha Vijayan, Satheesh Kumar, Suri Venkatachalam, and Lokesh Joshi. GlycoGAIT: a web database to browse glycogenes and lectins under gastric inflammatory diseases. Journal of Theoretical Biology, 406(??):93–98, October 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630203X.

Oizumi:2013:OLS

[OT13] Ryo Oizumi and Takenori Takada. Optimal life schedule with stochastic growth in age-size structured models: Theory and an application. Journal of Theoretical Biology, 323(??):76–89, April 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313000453.

Owens:2010:ECS

[OTGT10] Nick D. L. Owens, Jon Timmis, Andrew Greensted, and Andy Tyrrell. Elucidation of T cell signalling models. Journal of Theoretical Biology, 262(3):452–470, February 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004925.

Otsuka:2011:LSE

[Ots11] Jinya Otsuka. The large-scale evolution by generating new genes from gene duplication; similarity and difference between monoploid and diploid organisms. Journal of Theoretical Biology, 278(1):120–126, June 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001354.

Oster:2011:LCM

[OTTF11] Andrew M. Oster, Balbir Thomas, David Terman, and Christopher P. Fall. The low conductance mitochondrial permeability transition pore confers excitability and CICR wave propagation in a computational model. Journal of Theoretical Biology, 273(1):216–231, March 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006788.

Ou:2019:ANF

[Ou19] Zhonghui Ou. Approximate nutrient flux and concentration solutions of the Nye–Tinker–Barber model by the perturbation expansion method. Journal of Theoretical Biology, 476(??):19–29, September 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302097.

Olivera:2010:RDS

[OUMA10] Beatriz Carely Luna Olivera, Edgardo Ugalde, and Agustino Martínez-Antonio. Regulatory dynamics of standard twocomponent systems in bacteria. Journal of Theoretical Biology, 264(2):560–569, May 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000755.

Oraby:2014:MSB

[OVKL14] Tamer Oraby, Olga Vasilyeva, Daniel Krewski, and Frithjof Lutscher. Modeling seasonal behavior changes and disease transmission with application to chronic wasting disease. Journal of Theoretical Biology, 340(??):50–59, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313004244.

Oien:2016:EEH

[ØW16] Alf H. Øien and Helge Wiig. Electrostatic, elastic and hydration-dependent interactions in dermis influencing volume exclusion and macromolecular transport. Journal of Theoretical Biology, 400(??):80–91, July 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630025X.

Okasha:2014:IFM

[OWB14] Samir Okasha, John A. Weymark, and Walter Bossert. Inclusive fitness maximization: An axiomatic approach. Journal of Theoretical Biology, 350(??):24–31, June 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000630.

Ohnishi:2014:ASP

[OYA14] Shuhei Ohnishi, Takashi Yamakawa, and Tatsuro Akamine. On the analytical solution for the Pütter–Bertalanffy growth equation. Journal of Theoretical Biology, 343(??):174–177, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931300516X.

Ohkubo:2010:LTB

[OYIM10] Jun Ohkubo, Kazushi Yoshida, Yuichi Iino, and Naoki Masuda. Long-tail behavior in locomotion of Caenorhabditis elegans. Journal of Theoretical Biology, 267(2):213–222, November 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004297.

Olyaee:2016:PPS

[OYY16] Mohammad H. Olyaee, Ali Yaghoubi, and Mahdi Yaghoobi. Predicting protein structural classes based on complex networks and recurrence analysis. Journal of Theoretical Biology, 404(??):375–382, September 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301527.

Penta:2015:RMT

[PA15] R. Penta and D. Ambrosi. The role of the microvascular tortuosity in tumor transport phenomena. Journal of Theoretical Biology, 364(??):80–97, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400455X.

Pagnutti:2011:EGL

[PAA11] Christopher Pagnutti, Madhur Anand, and Mohamed Azzouz. Estimating gap lifetime and memory from a simple model of forest canopy dynamics. Journal of Theoretical Biology, 274(1):154–160, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000336.

Pang:2010:DBH

[PaCZ10] Jianhua Pang, Jing an Cui, and Xueyong Zhou. Dynamical behavior of a hepatitis B virus transmission model with vaccination. Journal of Theoretical Biology, 265(4):572–578, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931000281X.

Page:2019:OWM

[Pag19] Karen M. Page. Oscillations in well-mixed, deterministic feedback systems: Beyond ring oscillators. Journal of Theoretical Biology, 481(??):44–53, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301882.

Painter:2019:MMC

[Pai19] Kevin J. Painter. Mathematical models for chemotaxis and their applications in self-organisation phenomena. Journal of Theoretical Biology, 481(??):162–182, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830314X.

Pahlajani:2011:SRM

[PAK11] Chetan D. Pahlajani, Paul J. Atzberger, and Mustafa Khammash. Stochastic reduction method for biological chemical kinetics using time-scale separation. Journal of Theoretical Biology, 272(1):96–112, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006089.

Panchanathan:2011:TWD

[Pan11] Karthik Panchanathan. Two wrongs don’t make a right: The initial viability of different assessment rules in the evolution of indirect reciprocity. Journal of Theoretical Biology, 277(1):48– 54, May 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001032.

Pantokratoras:2019:CPP

[Pan19] Asterios Pantokratoras. Comment on the paper “Peristaltic transport of multilayered power-law fluids with distinct viscosities: a mathematical model for intestinal flows, S. K. Pandey, M. K. Chaube, Dharmendra Tripathi, Journal of Theoretical Biology 278 (2011) 11–19”. Journal of Theoretical Biology, 467(??):174, April 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931930044X.

Pajaro:2017:SMN

[PAOMV17] Manuel Pájaro, Antonio A. Alonso, Irene Otero-Muras, and Carlos Vázquez. Stochastic modeling and numerical simulation of gene regulatory networks with protein bursting. Journal of Theoretical Biology, 421(??):51–70, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301352.

Paradis:2015:RPD

[Par15] Emmanuel Paradis. Random phylogenies and the distribution of branching times. Journal of Theoretical Biology, 387(??):39–45, December 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004403.

Parag:2019:SEL

[Par19] Kris V. Parag. On signalling and estimation limits for molecular birth-processes. Journal of Theoretical Biology, 480(??):262–273, November 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931930284X.

Painter:2010:IAC

[PAS10] Kevin J. Painter, Nicola J. Armstrong, and Jonathan A. Sherratt. The impact of adhesion on cellular invasion processes in cancer and development. Journal of Theoretical Biology, 264(3):1057–1067, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001669.

Patel:2016:SMM

[Pat16] Mainak Patel. A simplified model of mutually inhibitory sleep-active and wake-active neuronal populations employing a noise-based switching mechanism. Journal of Theoretical Biology, 394(??):127–136, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000461.

Patel:2019:AFM

[Pat19] Mainak Patel. Analysis of feedforward mechanisms of multiwhisker receptive field generation in a model of the rat barrel cortex. Journal of Theoretical Biology, 477(??):51–62, September 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302358.

Preziosi:2010:EVP

[PAV10] L. Preziosi, D. Ambrosi, and C. Verdier. An elastovisco-plastic model of cell aggregates. Journal of Theoretical Biology, 262(1):35–47, January 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309003841.

Pavesi:2014:PDT

[Pav14] Angelo Pavesi. Prediction of the determinants of thermal stability by linear discriminant analysis: The case of the glutamate dehydrogenase protein family. Journal of Theoretical Biology, 357(??):160–168, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002847.

Pavoine:2019:OAE

[Pav19] Sandrine Pavoine. An ordination approach to explore similarities among communities. Journal of Theoretical Biology, 462(??):85–96, February 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318305496.

Pigolotti:2016:CBF

[PB16] Simone Pigolotti and Roberto Benzi. Competition between fast- and slow-diffusing species in non-homogeneous environments. Journal of Theoretical Biology, 395(??):204–210, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000722.

Pharaon:2018:ISB

[PB18] Joe Pharaon and Chris T. Bauch. The influence of social behaviour on competition between virulent pathogen strains. Journal of Theoretical Biology, 455(??):47–53, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303102.

Purcell:2012:CEB

[PBA12] Jessica Purcell, Alan Brelsford, and Leticia Avilés. Coevolution between sociality and dispersal: The role of synergistic cooperative benefits. Journal of Theoretical Biology, 312(??):44–54, November 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003608.

Pavlova:2010:CBT

[PBB10] Viola Pavlová, Ludĕk Berec, and David S. Boukal. Caught between two Allee effects: Trade-off between reproduction and predation risk. Journal of Theoretical Biology, 264(3):787– 798, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001323.

Paixao:2015:TUF

[PBB^{+}15a] Tiago Paixão, Golnaz Badkobeh, Nick Barton, Doğan Çörüş, Duc-Cuong Dang, Tobias Friedrich, Per Kristian Lehre, Dirk Sudholt, Andrew M. Sutton, and Barbora Trubenová. Toward a unifying framework for evolutionary processes. Journal of Theoretical Biology, 383(??):28–43, October 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003409.

Petrie:2015:QRW

[PBB^{+}15b] Stephen M. Petrie, Jeff Butler, Ian G. Barr, Jodie McVernon, Aeron C. Hurt, and James M. McCaw. Quantifying relative within-host replication fitness in influenza virus competition experiments. Journal of Theoretical Biology, 382(??):259–271, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931500332X.

Pietra:2010:ECA

[PBBB10] Daniele Pietra, Alice Borghini, Maria Cristina Breschi, and Anna Maria Bianucci. Enhancer and competitive allosteric modulation model for g-protein-coupled receptors. Journal of Theoretical Biology, 267(4):663–675, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004479.

Poos:2011:HIM

[PBD11] J. J. Poos, Å. Brännström, and U. Dieckmann. Harvestinduced maturation evolution under different life-history trade-offs and harvesting regimes. Journal of Theoretical Biology, 279(1):102–112, June 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001305.

Postlethwaite:2013:NMS

[PBD13] Claire M. Postlethwaite, Pieta Brown, and Todd E. Dennis. A new multi-scale measure for analysing animal movement data. Journal of Theoretical Biology, 317(??):175–185, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005322.

Porter:2012:ISN

[PBEI12] Joshua R. Porter, John S. Burg, Peter J. Espenshade, and Pablo A. Iglesias. Identifying a static nonlinear structure in a biological system using noisy, sparse data. Journal of Theoretical Biology, 300(??):232–241, May 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000574.

Piwowar:2013:SRE

[PBKR13] Monika Piwowar, Mateusz Banach, Leszek Konieczny, and Irena Roterman. Structural role of exon-coded fragment of polypeptide chains in selected enzymes. Journal of Theoretical Biology, 337(??):15–23, November 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003391.

Pendergraft:2015:ACC

[PBP15] William F. Pendergraft, Anshul K. Badhwar, and Gloria A. Preston. Autoantigen complementarity and its contributions to hallmarks of autoimmune disease. Journal of Theoretical Biology, 375(??):88–94, June 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400695X.

Pattni:2017:EDE

[PBR17a] Karan Pattni, Mark Broom, and Jan Rychtár. Evolutionary dynamics and the evolution of multiplayer cooperation in a subdivided population. Journal of Theoretical Biology, 429(??):105–115, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303119.

Polizzi:2017:TSM

[PBR17b] B. Polizzi, O. Bernard, and M. Ribot. A time-space model for the growth of microalgae biofilms for biofuel production. Journal of Theoretical Biology, 432(??):55–79, November 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303892.

Petchey:2011:FEB

[PBRW11] Owen L. Petchey, Andrew P. Beckerman, Jens O. Riede, and Philip H. Warren. Fit, efficiency, and biology: Some thoughts on judging food web models. Journal of Theoretical Biology, 279(1):169–171, June 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001652.

Poh:2012:QGS

[PBS^{+}12] Yong Cheng Poh, Arthur Beyder, Peter R. Strege, Gianrico Farrugia, and Martin L. Buist. Quantification of gastrointestinal sodium channelopathy. Journal of Theoretical Biology, 293(??):41–48, January 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100470X.

Pitcher:2019:MMH

[PBSM19] Ashley B. Pitcher, Annick Borquez, Britt Skaathun, and Natasha K. Martin. Mathematical modeling of hepatitis c virus (HCV) prevention among people who inject drugs: a review of the literature and insights for elimination strategies. Journal of Theoretical Biology, 481(??):194–201, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305666.

Peletier:2010:IPB

[PBvdG10] Lambertus A. Peletier, Neil Benson, and Piet H. van der Graaf. Impact of protein binding on receptor occupancy: a two-compartment model. Journal of Theoretical Biology, 265(4):657–671, August 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000278X.

Pal:2018:EMS

[PBYB18] Arijit Pal, Amiya Ranjan Bhowmick, Farhana Yeasmin, and Sabyasachi Bhattacharya. Evolution of model specific relative growth rate: Its genesis and performance over Fisher’s growth rates. Journal of Theoretical Biology, 444(??):11–27, May 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931830064X.

Pigolotti:2010:CIT

[PC10a] Simone Pigolotti and Massimo Cencini. Coexistence and invasibility in a two-species competition model with habitatpreference. Journal of Theoretical Biology, 265(4):609–617, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002870.

Posta:2010:MMI

[PC10b] Filippo Posta and Tom Chou. A mathematical model of intercellular signaling during epithelial wound healing. Journal of Theoretical Biology, 266(1):70–78, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002729.

Pigolotti:2013:SAL

[PC13] Simone Pigolotti and Massimo Cencini. Species abundances and lifetimes: From neutral to niche-stabilized communities. Journal of Theoretical Biology, 338(??):1–8, December 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003986.

Persons:2017:FOD

[PC17] W. Scott Persons and Philip J. Currie. The functional origin of dinosaur bipedalism: Cumulative evidence from bipedally inclined reptiles and disinclined mammals. Journal of Theoretical Biology, 420(??):1–7, May 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300942.

Peruzzi:2014:NIQ

[PCB14] Lorenzo Peruzzi, Katia Francesca Caparelli, and Gianni Bedini. A new index for the quantification of chromosome number variation: An application to selected animal and plant groups. Journal of Theoretical Biology, 353(??):55–60, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314001477.

Piedrafita:2012:SMI

[PCBMM12] Gabriel Piedrafita, Athel Cornish-Bowden, Federico Morán, and Francisco Montero. Size matters: Influence of stochasticity on the self-maintenance of a simple model of metabolic closure. Journal of Theoretical Biology, 300(??):143–151, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000148.

Plazzotta:2015:MSB

[PCC15] Giacomo Plazzotta, Ted Cohen, and Caroline Colijn. Magnitude and sources of bias in the detection of mixed strain M. tuberculosis infection. Journal of Theoretical Biology, 368(??):67–73, March 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314007061.

Pearcy:2016:CRH

[PCC16] Nicole Pearcy, Nadia Chuzhanova, and Jonathan J. Crofts. Complexity and robustness in hypernetwork models of metabolism. Journal of Theoretical Biology, 406(??):99–104, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301667.

Philipsen:2010:MCS

[PCHM10] K. R. Philipsen, L. E. Christiansen, H. Hasman, and H. Madsen. Modelling conjugation with stochastic differential equations. Journal of Theoretical Biology, 263(1):134–142, March 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309005530.

Pienaar:2015:CTI

[PCL^{+}15] Elsje Pienaar, Nicholas A. Cilfone, Philana Ling Lin, Véronique Dartois, Joshua T. Mattila, J. Russell Butler, JoAnne L. Flynn, Denise E. Kirschner, and Jennifer J. Linderman. A computational tool integrating host immunity with antibiotic dynamics to study tuberculosis treatment. Journal of Theoretical Biology, 367(??):166–179, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006730.

Priklopil:2017:OIS

[PCN17] Tadeas Priklopil, Krishnendu Chatterjee, and Martin Nowak. Optional interactions and suspicious behaviour facilitates trustful cooperation in prisoners dilemma. Journal of Theoretical Biology, 433(??):64–72, November 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303995.

Pool:2018:MMM

[PCS^{+}18] Frances Pool, Richard Currie, Peter K. Sweby, José Domingo Salazar, and Marcus J. Tindall. A mathematical model of the mevalonate cholesterol biosynthesis pathway. Journal of Theoretical Biology, 443(??):157–176, April 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305623.

Pandey:2011:PTM

[PCT11] S. K. Pandey, M. K. Chaube, and Dharmendra Tripathi. Peristaltic transport of multilayered power-law fluids with distinct viscosities: a mathematical model for intestinal flows. Journal of Theoretical Biology, 278(1):11–19, June 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001299.

Pandey:2019:CPT

[PCT19] S. K. Pandey, M. K. Chaube, and Dharmendra Tripathi. Corrigendum to “Peristaltic transport of multilayered power-law fluids with distinct viscosities: a mathematical model for intestinal flows” [Journal of Theoretical Biology 278 (2011) 11–19]. Journal of Theoretical Biology, 470(??):119, June 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930092X.

Parvinen:2013:SET

[PD13] Kalle Parvinen and Ulf Dieckmann. Self-extinction through optimizing selection. Journal of Theoretical Biology, 333 (??):1–9, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519313001513.

Prodanov:2016:MSF

[PD16] Dimiter Prodanov and Jean Delbeke. A model of spacefractional-order diffusion in the glial scar. Journal of Theoretical Biology, 403(??):97–109, August 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300595.

Pantic:2015:DAF

[PDB^{+}15] Igor Pantic, Sanja Dacic, Predrag Brkic, Irena Lavrnja, Tomislav Jovanovic, Senka Pantic, and Sanja Pekovic. Discriminatory ability of fractal and grey level co-occurrence matrix methods in structural analysis of hippocampus layers. Journal of Theoretical Biology, 370(??):151–156, April 7, 2015.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000521.

Pane:2017:APC

[PDC^{+}17] Katia Pane, Lorenzo Durante, Orlando Crescenzi, Valeria Cafaro, Elio Pizzo, Mario Varcamonti, Anna Zanfardino, Viviana Izzo, Alberto Di Donato, and Eugenio Notomista. Antimicrobial potency of cationic antimicrobial peptides can be predicted from their amino acid composition: Application to the detection of “cryptic” antimicrobial peptides. Journal of Theoretical Biology, 419(??):254–265, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300681.

Piretto:2018:CTI

[PDF18] Elena Piretto, Marcello Delitala, and Mario Ferraro. Combination therapies and intra-tumoral competition: Insights from mathematical modeling. Journal of Theoretical Biology, 446(??):149–159, June 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301280.

Plitzko:2012:CSR

[PDG12] Sebastian J. Plitzko, Barbara Drossel, and Christian Guill. Complexity-stability relations in generalized food-web models with realistic parameters. Journal of Theoretical Biology, 306(??):7–14, August 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001816.

Pal:2017:NRV

[PDG17] Krishnendu Pal, Biswajit Das, and Gautam Gangopadhyay. Nonequilibrium response of a voltage gated sodium ion channel and biophysical characterization of dynamic hysteresis. Journal of Theoretical Biology, 415(??):113–124, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316304155.

Peurichard:2017:SMC

[PDL^{+}17] D. Peurichard, F. Delebecque, A. Lorsignol, C. Barreau, J. Rouquette, X. Descombes, L. Casteilla, and P. Degond. Simple mechanical cues could explain adipose tissue morphology. Journal of Theoretical Biology, 429(??):61–81, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303065.

Pai:2017:SBD

[PDM17] Priyadarshini P. Pai, Tirtharaj Dash, and Sukanta Mondal. Sequence-based discrimination of protein-RNA interacting residues using a probabilistic approach. Journal of Theoretical Biology, 418(??):77–83, April 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300498.

Pantic:2016:GLC

[PDNP16] Igor Pantic, Draga Dimitrijevic, Dejan Nesic, and Danica Petrovic. Gray level co-occurrence matrix algorithm as pattern recognition biosensor for oxidopamine-induced changes in lymphocyte chromatin architecture. Journal of Theoretical Biology, 406(??):124–128, October 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302016.

Pedruzzi:2016:UPL

[PDRC16] Gabriele Pedruzzi, Phonindra Nath Das, Kanury VS. Rao, and Samrat Chatterjee. Understanding PGE2, LXA4 and LTB4 balance during mycobacterium tuberculosis infection through mathematical model. Journal of Theoretical Biology, 389(??):159–170, January 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500524X.

Palamara:2013:PEM

[PDSP13] Gian Marco Palamara, Gustav W. Delius, Matthew J. Smith, and Owen L. Petchey. Predation effects on mean time to extinction under demographic stochasticity. Journal of Theoretical Biology, 334(??):61–70, October 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002798.

Politi:2010:MSD

[PDT^{+}10] Antonio Z. Politi, Graham M. Donovan, Merryn H. Tawhai, Michael J. Sanderson, Anne-Marie Lauzon, Jason H. T. Bates, and James Sneyd. A multiscale, spatially distributed model of asthmatic airway hyper-responsiveness. Journal of Theoretical Biology, 266(4):614–624, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003899.

Pressley:2010:ECC

[PDW10] Joanna Pressley, Erika M. C. D’Agata, and Glenn F. Webb. The effect of co-colonization with communityacquired and hospital-acquired methicillin-resistant Staphylococcus aureus strains on competitive exclusion. Journal of Theoretical Biology, 264(3):645–656, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001712.

Pretorius:2011:ISP

[PDW11] Carel Pretorius, Peter Dodd, and Robin Wood. An investigation into the statistical properties of TB episodes in a south African community with high HIV prevalence. Journal of Theoretical Biology, 270(1):154–163, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005400.

Peedicayil:2012:CEE

[Pee12] Jacob Peedicayil. A comparison of Erikson’s epigenetic principle with epigenetics. Journal of Theoretical Biology, 315(??):144–145, December 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004882.

Pelce:2018:CEB

[Pel18] Pierre Pelcé. Competition of energy between active transport and vesicle fusion at the origin of intracellular gradient fields. Journal of Theoretical Biology, 438(??):165–173, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305258.

Paarporn:2018:ISC

[PEW18] Keith Paarporn, Ceyhun Eksin, and Joshua S. Weitz. Information sharing for a coordination game in fluctuating environments. Journal of Theoretical Biology, 454(??):376–385, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303229.

Price:2013:VVS

[PEZ^{+}13] Ian Price, Bard Ermentrout, Ruben Zamora, Bo Wang, Nabil Azhar, Qi Mi, Gregory Constantine, James R. Faeder, Shirley Luckhart, and Yoram Vodovotz. In vivo, in vitro, and in silico studies suggest a conserved immune module that regulates malaria parasite transmission from mammals to mosquitoes. Journal of Theoretical Biology, 334(??):173–186, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002580.

Pontier:2010:CCP

[PFB10] Dominique Pontier, David Fouchet, and Joël Bried. Can cat predation help competitors coexist in seabird communities? Journal of Theoretical Biology, 262(1):90–96, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309004263.

Pavlova:2015:NVS

[PFJS15] J. Pavlova, A. Fasano, J. Janela, and A. Sequeira. Numerical validation of a synthetic cell-based model of blood coagulation. Journal of Theoretical Biology, 380(??):367–379, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002830.

Puckeridge:2011:ICQ

[PFPR11] M. Puckeridge, B. D. Fulcher, A. J. K. Phillips, and P. A. Robinson. Incorporation of caffeine into a quantitative model of fatigue and sleep. Journal of Theoretical Biology, 273(1):44–54, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310006739.

Peedicayil:2018:EBO

[PG18a] Jacob Peedicayil and Dennis R. Grayson. An epigenetic basis for an omnigenic model of psychiatric disorders. Journal of Theoretical Biology, 443(??):52–55, April 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300432.

Peedicayil:2018:SIE

[PG18b] Jacob Peedicayil and Dennis R. Grayson. Some implications of an epigenetic-based omnigenic model of psychiatric disorders. Journal of Theoretical Biology, 452(??):81–84, September 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302443.

Prisman:2011:SAR

[PGF11] Eliezer Z. Prisman, Amiram Gafni, and Antonio Finelli. A stochastic approach to risk management for prostate cancer patients on active surveillance. Journal of Theoretical Biology, 284(1):61–70, September 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003122.

Pichugin:2015:MMM

[PGG^{+}15] Yuriy Pichugin, Chaitanya S. Gokhale, Julián Garcia, Arne Traulsen, and Paul B. Rainey. Modes of migration and multilevel selection in evolutionary multiplayer games. Journal of Theoretical Biology, 387(??):144–153, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315004762.

Parnell:2012:EIE

[PGGvdB12] S. Parnell, T. R. Gottwald, W. R. Gilks, and F. van den Bosch. Estimating the incidence of an epidemic when it is first discovered and the design of early detection monitoring. Journal of Theoretical Biology, 305(??):30–36, July 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312001269.

Panovska-Griffiths:2019:WPN

[PGGvL^{+}19] Jasmina Panovska-Griffiths, Luca Grieco, Edwin van Leeuwen, Marc Baguelin, Richard Pebody, and Martin Utley. Are we prepared for the next influenza pandemic? Lessons from modelling different preparedness policies against four pandemic scenarios. Journal of Theoretical Biology, 481(??): 223–232, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519319301870.

Poethke:2011:AIA

[PGH11] Hans Joachim Poethke, Andreas Gros, and Thomas Hovestadt. The ability of individuals to assess population density influences the evolution of emigration propensity and dispersal distance. Journal of Theoretical Biology, 282(1):93–99, August 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311002542.

Powathil:2012:MEC

[PGHC12] Gibin G. Powathil, Kirsty E. Gordon, Lydia A. Hill, and Mark A. J. Chaplain. Modelling the effects of cell-cycle heterogeneity on the response of a solid tumour to chemotherapy: Biological insights from a hybrid multiscale cellular automaton model. Journal of Theoretical Biology, 308(??):1–19, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002573.

Pilosof:2017:ADD

[PGKZ17] Shai Pilosof, Gili Greenbaum, Boris R. Krasnov, and Yuval R. Zelnik. Asymmetric disease dynamics in multihost interconnected networks. Journal of Theoretical Biology, 430(??):237–244, October 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931730351X.

Pei:2014:ILC

[PGLZ14] Yonggang Pei, Qinghui Gao, Juntao Li, and Xiting Zhao. Identifying local co-regulation relationships in gene expression data. Journal of Theoretical Biology, 360(??):200–207, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003750.

Pinky:2019:SCR

[PGPD19] Lubna Pinky, Gilberto González-Parra, and Hana M. Dobrovolny. Superinfection and cell regeneration can lead to chronic viral coinfections. Journal of Theoretical Biology, 466(??):24– 38, April 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300116.

Potvin:2010:SLF

[PGS10] J. Potvin, J. A. Goldbogen, and R. E. Shadwick. Scaling of lunge feeding in rorqual whales: An integrated model of engulfment duration. Journal of Theoretical Biology, 267(3):437–453, December 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004492.

Pan:2019:TFM

[PGT^{+}19] Michael Pan, Peter J. Gawthrop, Kenneth Tran, Joseph Cursons, and Edmund J. Crampin. A thermodynamic framework for modelling membrane transporters. Journal of Theoretical Biology, 481(??):10–23, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304703.

Painter:2013:MMG

[PH13] K. J. Painter and T. Hillen. Mathematical modelling of glioma growth: The use of diffusion tensor imaging (DTI) data to predict the anisotropic pathways of cancer invasion. Journal of Theoretical Biology, 323(??):25–39, April 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000398.

Perfahl:2017:HMV

[PHA^{+}17] Holger Perfahl, Barry D. Hughes, Tomás Alarcón, Philip K. Maini, Mark C. Lloyd, Matthias Reuss, and Helen M. Byrne. 3D hybrid modelling of vascular network formation. Journal of Theoretical Biology, 414(??):254–268, February 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303782.

Pantoja-Hernandez:2016:RED

[PHÁBAI^{+}16] Libertad Pantoja-Hernández, Elena Álvarez-Buylla, Carlos F. Aguilar-Ibáñez, Adriana Garay-Arroyo, Alberto Soria-López, and Juan Carlos Martínez-García. Retroactivity effects dependency on the transcription factors binding mechanisms. Journal of Theoretical Biology, 410(??):77–106, December 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302454.

Pruett:2012:AHH

[PHC12] William A. Pruett, Robert L. Hester, and Thomas G. Coleman. The apparent hysteresis in hormone-agonist relationships. Journal of Theoretical Biology, 296(??):1–5, March 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311005959.

Painter:2018:CMF

[PHH18] Kevin J. Painter, William Ho, and Denis J. Headon. A chemotaxis model of feather primordia pattern formation during avian development. Journal of Theoretical Biology, 437(??):225–238, January 21, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304952.

Pellis:2015:EAM

[PHK15] Lorenzo Pellis, Thomas House, and Matt J. Keeling. Exact and approximate moment closures for non-Markovian network epidemics. Journal of Theoretical Biology, 382(??):160–177, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002210.

Piltz:2018:IPP

[PHPM18] Sofia H. Piltz, Lauri Harhanen, Mason A. Porter, and Philip K. Maini. Inferring parameters of prey switching in a 1 predator-2 prey plankton system with a linear preference tradeoff. Journal of Theoretical Biology, 456(??):108–122, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303291.

Pantic:2012:CFD

[PHTP^{+}12] Igor Pantic, Ljubica Harhaji-Trajkovic, Aleksandar Pantovic, Nebojsa T. Milosevic, and Vladimir Trajkovic. Changes in fractal dimension and lacunarity as early markers of UVinduced apoptosis. Journal of Theoretical Biology, 303(??):87– 92, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001300.

Piast:2019:SIB

[Pia19] Radoslaw W. Piast. Shannon’s information, Bernal’s biopoiesis and Bernoulli distribution as pillars for building a definition of life. Journal of Theoretical Biology, 470(??):101– 107, June 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301109.

Pietruszka:2010:EAS

[Pie10] Mariusz Pietruszka. Exact analytic solutions for a global equation of plant cell growth. Journal of Theoretical Biology, 264(2):457–466, May 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000792.

Pietruszka:2012:BIM

[Pie12] Mariusz Pietruszka. A biosynthesis/inactivation model for enzymatic WLFs or non-enzymatically mediated cell evolution. Journal of Theoretical Biology, 315(??):119–127, December 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004948.

Parker:2010:SCG

[PIPB10] G. A. Parker, S. Immler, S. Pitnick, and T. R. Birkhead. Sperm competition games: Sperm size (mass) and number under raffle and displacement, and the evolution of P_{2}. Journal of Theoretical Biology, 264(3):1003–1023, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001268.

Park:2012:IER

[PIT12] Sang Woo Park, Marcos Intaglietta, and Daniel M. Tartakovsky. Impact of endothelium roughness on blood flow. Journal of Theoretical Biology, 300(??):152–160, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000252.

Patel:2013:DSO

[PJ13] Mainak Patel and Badal Joshi. Decoding synchronized oscillations within the brain: Phase-delayed inhibition provides a robust mechanism for creating a sharp synchrony filter. Journal of Theoretical Biology, 334(??):13–25, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002518.

Patel:2018:DSR

[PJ18] Mainak Patel and Badal Joshi. Deterministic stability regimes and noise-induced quasistable behavior in a pair of reciprocally inhibitory neurons. Journal of Theoretical Biology, 441(??):68–83, March 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305507.

Podder:2014:HDR

[PJL14] Avijit Podder, Nidhi Jatana, and N. Latha. Human dopamine receptors interaction network (DRIN): a systems biology perspective on topology, stability and functionality of the network. Journal of Theoretical Biology, 357(??):169–183, September 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002823.

Powell:2011:GTS

[PKH11] Eric N. Powell, John M. Klinck, and Eileen E. Hofmann. Generation time and the stability of sex-determining alleles in oyster populations as deduced using a gene-based population dynamics model. Journal of Theoretical Biology, 271(1):27–43, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005916.

Papadakis:2012:MDM

[PKL12] Ioannis A. Papadakis, Kiriakos Kotzabasis, and Konstadia Lika. Modeling the dynamic modulation of light energy in photosynthetic algae. Journal of Theoretical Biology, 300(??):254– 264, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000604.

Pelosse:2012:RRV

[PKZ12] Perrine Pelosse and Christopher M. Kribs-Zaleta. The role of the ratio of vector and host densities in the evolution of transmission modes in vector-borne diseases. The example of sylvatic Trypanosoma cruzi. Journal of Theoretical Biology, 312(??):133–142, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003803.

Pienaar:2014:EMM

[PL14a] Elsje Pienaar and Maria Lerm. Erratum to “A mathematical model of the initial interaction between Mycobacterium tuberculosis and macrophages” [J. Theoret. Biol. 342 (2014) 23–32]. Journal of Theoretical Biology, 349(??):172, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000824.

Pienaar:2014:MMI

[PL14b] Elsje Pienaar and Maria Lerm. A mathematical model of the initial interaction between mycobacterium tuberculosis and macrophages. Journal of Theoretical Biology, 342(??):23–32, February 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004657.

Platkowski:2010:CTP

[Pła10] Tadeusz Płatkowski. Cooperation in two-person evolutionary games with complex personality profiles. Journal of Theoretical Biology, 266(4):522–528, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003759.

Powell:2017:MDI

[PLD^{+}17] Christopher D. Powell, Secundino López, André Dumas, Dominique P. Bureau, Sarah E. Hook, and James France. Mathematical descriptions of indeterminate growth. Journal of Theoretical Biology, 425(??):88–96, July 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301960.

Polo:2018:BRN

[PLF18] Gina Polo, Marcelo B. Labruna, and Fernando Ferreira. Basic reproduction number for the Brazilian spotted fever. Journal of Theoretical Biology, 458(??):119–124, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304405.

Powell:2019:EFM

[PLF19] Christopher D. Powell, Secundino López, and James France. Elementary functions modified for seasonal effects to describe growth in freshwater fish. Journal of Theoretical Biology, 461(??):133–144, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305186.

Pourziaei:2019:SMD

[PLHL19] Babak Pourziaei, Gregory M. Lewis, Huaxiong Huang, and John E. Lewis. Spatiotemporal model for depth perception in electric sensing. Journal of Theoretical Biology, 461(??):157–169, January 14, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304995.

Pena:2014:GSE

[PLN14] Jorge Peña, Laurent Lehmann, and Georg Nöldeke. Gains from switching and evolutionary stability in multi-player matrix games. Journal of Theoretical Biology, 346(??):23–33, April 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005675.

Pereira:2010:PPD

[PM10a] Marcelo Alves Pereira and Alexandre Souto Martinez. Pavlovian Prisoner’s Dilemma — analytical results, the quasiregular phase and spatio-temporal patterns. Journal of Theoretical Biology, 265(3):346–358, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002572.

Preece:2010:EGL

[PM10b] T. Preece and Y. Mao. The evolution of gynodioecy on a lattice. Journal of Theoretical Biology, 266(2):219–225, September 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003152.

Parham:2011:OPE

[PM11a] Paul E. Parham and Edwin Michael. Outbreak properties of epidemic models: The roles of temporal forcing and stochasticity on pathogen invasion dynamics. Journal of Theoretical Biology, 271(1):1–9, February 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006004.

Peris:2011:NGA

[PM11b] Guillermo Peris and Andrés Marzal. Normalized global alignment for protein sequences. Journal of Theoretical Biology, 291(??):22–28, December 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004735.

Psiuk-Maksymowicz:2013:MMD

[PM13] K. Psiuk-Maksymowicz. Multiphase modelling of desmoplastic tumour growth. Journal of Theoretical Biology, 329(??):52–63, July 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313000994.

Patel:2014:KBN

[PM14] Maulika S. Patel and Himanshu S. Mazumdar. Knowledge base and neural network approach for protein secondary structure prediction. Journal of Theoretical Biology, 361(??):182–189, November 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004536.

Perretti:2015:ERE

[PM15] Charles T. Perretti and Stephan B. Munch. On estimating the reliability of ecological forecasts. Journal of Theoretical Biology, 372(??):211–216, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001046.

Porfiri:2017:SDA

[PM17] Maurizio Porfiri and Manuel Ruiz Marín. Symbolic dynamics of animal interaction. Journal of Theoretical Biology, 435(??):145–156, December 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304162.

Powathil:2016:BET

[PMCS16] Gibin G. Powathil, Alastair J. Munro, Mark A. J. Chaplain, and Maciej Swat. Bystander effects and their implications for clinical radiation therapy: Insights from multiscale in silico experiments. Journal of Theoretical Biology, 401(??):1–14, July 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300315.

Payne:2011:ECD

[PMD11] Joshua L. Payne, Rupert Mazzucco, and Ulf Dieckmann. The evolution of conditional dispersal and reproductive isolation along environmental gradients. Journal of Theoretical Biology, 273(1):147–155, March 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006995.

Pond:2018:MBC

[PMD^{+}18] D. Pond, A. T. McBride, L. M. Davids, B. D. Reddy, and G. Limbert. Microstructurally-based constitutive modelling of the skin — linking intrinsic ageing to microstructural parameters. Journal of Theoretical Biology, 444(??):108–123, May 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300225.

Phipps:2015:MMM

[PMK15] Colin Phipps, Hamid Molavian, and Mohammad Kohandel. A microscale mathematical model for metabolic symbiosis: Investigating the effects of metabolic inhibition on ATP turnover in tumors. Journal of Theoretical Biology, 366(??):103–114, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006687.

Prakash:2010:NNA

[PMKM10] Om Prakash, Shakti Mehrotra, Aneesh Krishna, and Bhartendu N. Mishra. A neural network approach for the prediction of in vitro culture parameters for maximum biomass yields in hairy root cultures. Journal of Theoretical Biology, 265(4):579–585, August 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002559.

Price:2015:IRI

[PMKS^{+}15] Ian Price, Ericka D. Mochan-Keef, David Swigon, G. Bard Ermentrout, Sarah Lukens, Franklin R. Toapanta, Ted M. Ross, and Gilles Clermont. The inflammatory response to influenza A virus (H1N1): An experimental and mathematical study. Journal of Theoretical Biology, 374(??):83–93, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315001265.

Prosnier:2018:PEC

[PML18] Loïc Prosnier, Vincent Médoc, and Nicolas Loeuille. Parasitism effects on coexistence and stability within simple trophic modules. Journal of Theoretical Biology, 458(??): 68–77, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251931830434X.

Pechenick:2013:IAR

[PMP13] Dov A. Pechenick, Jason H. Moore, and Joshua L. Payne. The influence of assortativity on the robustness and evolvability of gene regulatory networks upon gene birth. Journal of Theoretical Biology, 330(??):26–36, August 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001239.

Pawlowitsch:2011:NSD

[PMR11] Christina Pawlowitsch, Panayotis Mertikopoulos, and Nikolaus Ritt. Neutral stability, drift, and the diversification of languages. Journal of Theoretical Biology, 287(??):1–12, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003493.

Priyadarshi:2017:MSV

[PMSY17] Anupam Priyadarshi, Sandip Mandal, S. Lan Smith, and Hidekatsu Yamazaki. Micro-scale variability enhances trophic transfer and potentially sustains biodiversity in plankton ecosystems. Journal of Theoretical Biology, 412(??):86–93, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303368.

Pierre:2012:PLR

[PMW12] Jean-Sebastien Pierre, Jean-Pierre Masson, and Eric Wajnberg. Patch leaving rules: a stochastic version of a wellknown deterministic motivational model. Journal of Theoretical Biology, 313(??):1–11, November 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003888.

Potapov:2012:ECS

[PMYHR12] Ilya Potapov, Jarno Mäkelä, Olli Yli-Harja, and Andre S. Ribeiro. Effects of codon sequence on the dynamics of genetic networks. Journal of Theoretical Biology, 315(??):17–25, December 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004511.

Pena:2016:VGS

[PN16] Jorge Peña and Georg Nöldeke. Variability in group size and the evolution of collective action. Journal of Theoretical Biology, 389(??):72–82, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005226.

Pena:2018:GSE

[PN18] Jorge Peña and Georg Nöldeke. Group size effects in social evolution. Journal of Theoretical Biology, 457(??):211– 220, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519318303746.

Park:2016:GAW

[PNK16] Su-Chan Park, Johannes Neidhart, and Joachim Krug. Greedy adaptive walks on a correlated fitness landscape. Journal of Theoretical Biology, 397(??):89–102, May 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001375.

Pena:2015:EDC

[PNL15] Jorge Peña, Georg Nöldeke, and Laurent Lehmann. Evolutionary dynamics of collective action in spatially structured populations. Journal of Theoretical Biology, 382(??):122–136, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003185.

Nguyen:2010:DAR

[pNMK^{+}10] Nam phuong Nguyen, Chris Myers, Hiroyuki Kuwahara, Chris Winstead, and James Keener. Design and analysis of a robust genetic Muller C-element. Journal of Theoretical Biology, 264(2):174–187, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005153.

Pantic:2016:FAG

[PNP^{+}16] Igor Pantic, Zorica Nesic, Jovana Paunovic Pantic, Sanja Radojević-Škodrić, Mila Cetkovic, and Gordana Basta Jovanovic. Fractal analysis and Gray level co-occurrence matrix method for evaluation of reperfusion injury in kidney medulla. Journal of Theoretical Biology, 397(??):61–67, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316001405.

Pandey:2018:MIB

[PNVN18] Surabhi Pandey, Seema Nanda, Amit Vutha, and Ram Naresh. Modeling the impact of biolarvicides on malaria transmission. Journal of Theoretical Biology, 454(??):396–409, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302947.

Pring:2012:MMB

[POK^{+}12] Stephen R. Pring, Markus Owen, John R. King, Kevin D. Sinclair, Robert Webb, Anthony P. F. Flint, and Philip C. Garnsworthy. A mathematical model of the bovine oestrous cycle: Simulating outcomes of dietary and pharmacological interventions. Journal of Theoretical Biology, 313(??):115–126, November 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003967.

Polev:2012:TND

[Pol12] Dmitrii Polev. Transcriptional noise as a driver of gene evolution. Journal of Theoretical Biology, 293(??):27–33, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100508X.

Presbitero:2012:DML

[POLT12] Gerardo Presbitero, Fergal J. O’Brien, T. Clive Lee, and David Taylor. Distribution of microcrack lengths in bone in vivo and in vitro. Journal of Theoretical Biology, 304(??):164–171, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001592.

Pontzer:2012:RRE

[Pon12] Herman Pontzer. Relating ranging ecology, limb length, and locomotor economy in terrestrial animals. Journal of Theoretical Biology, 296(??):6–12, March 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005881.

Payne:2012:VDI

[POP12] S. J. Payne, C. N. J. Oakes, and C. S. Park. Vasomotion does inhibit mass exchange between axisymmetric blood vessels and tissue. Journal of Theoretical Biology, 302(??):1–5, June 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001063.

Peurichard:2019:ECM

[POP^{+}19] D. Peurichard, M. Ousset, J. Paupert, B. Aymard, A. Lorsignol, L. Casteilla, and P. Degond. Extra-cellular matrix rigidity may dictate the fate of injury outcome. Journal of Theoretical Biology, 469(??):127–136, May 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300864.

Portet:2013:DVI

[Por13] Stéphanie Portet. Dynamics of in vitro intermediate filament length distributions. Journal of Theoretical Biology, 332(??):20–29, September 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001574.

Parvinen:2017:EFD

[POW17] Kalle Parvinen, Hisashi Ohtsuki, and Joe Yuichiro Wakano. The effect of fecundity derivatives on the condition of evolutionary branching in spatial models. Journal of Theoretical Biology, 416(??):129–143, March 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304295.

Parvinen:2018:SHE

[POW18] Kalle Parvinen, Hisashi Ohtsuki, and Joe Yuichiro Wakano. Spatial heterogeneity and evolution of fecundity-affecting traits. Journal of Theoretical Biology, 454(??):190–204, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302972.

Prescott:2012:GEB

[PP12] Thomas P. Prescott and Antonis Papachristodoulou. Guaranteed error bounds for structured complexity reduction of biochemical networks. Journal of Theoretical Biology, 304(??):172–182, July 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001750.

Prescott:2014:LDM

[PP14] Thomas P. Prescott and Antonis Papachristodoulou. Layered decomposition for the model order reduction of timescale separated biochemical reaction networks. Journal of Theoretical Biology, 356(??):113–122, September 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002161.

Parag:2017:OPP

[PP17a] Kris V. Parag and Oliver G. Pybus. Optimal point process filtering and estimation of the coalescent process. Journal of Theoretical Biology, 421(??):153–167, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301492.

Potts:2017:FFB

[PP17b] Jonathan R. Potts and Sergei V. Petrovskii. Fortune favours the brave: Movement responses shape demographic dynamics in strongly competing populations. Journal of Theoretical Biology, 420(??):190–199, May 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730125X.

Pozzobon:2018:HMP

[PP18] Victor Pozzobon and Patrick Perre. Han’s model parameters for microalgae grown under intermittent illumination: Determined using particle swarm optimization. Journal of Theoretical Biology, 437(??):29–35, January 21, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304733. See corrigendum [PP19].

Pozzobon:2019:CHM

[PP19] Victor Pozzobon and Patrick Perre. Corrigendum to “Han’s model parameters for microalgae grown under intermittent illumination: Determined using particle swarm optimization” [Journal of Theoretical Biology 437 (2018) 29–35]. Journal of Theoretical Biology, 469(??):201, May 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300906. See [PP18].

Pakka:2010:CFC

[PPBD10] Vijayanarasimha H. Pakka, Adam Prügel-Bennett, and Srinandan Dasmahapatra. Correlated fluctuations carry signatures of gene regulatory network dynamics. Journal of Theoretical Biology, 266(3):343–357, October 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000336X.

Pena:2011:PCC

[PPBT11] Jorge Peña, Enea Pestelacci, André Berchtold, and Marco Tomassini. Participation costs can suppress the evolution of upstream reciprocity. Journal of Theoretical Biology, 273(1):197–206, March 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000051.

Pires:2017:TSC

[PPC^{+}17] Állan S. Pires, William F. Porto, Pryscilla O. Castro, Octavio L. Franco, and Sérgio A. Alencar. Theoretical structural characterization of lymphoguanylin: a potential candidate for the development of drugs to treat gastrointestinal disorders. Journal of Theoretical Biology, 419(??):193–200, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300711.

Porto:2017:AAP

[PPF17] William F. Porto, Állan S. Pires, and Octavio L. Franco. Antimicrobial activity predictors benchmarking analysis using shuffled and designed synthetic peptides. Journal of Theoretical Biology, 426(??):96–103, August 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302205.

Pechenick:2012:IAR

[PPM12] Dov A. Pechenick, Joshua L. Payne, and Jason H. Moore. The influence of assortativity on the robustness of signalintegration logic in gene regulatory networks. Journal of Theoretical Biology, 296(??):21–32, March 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005996.

Proia:2016:CFC

[PPT^{+}16] Frédéric Proïa, Alix Pernet, Tatiana Thouroude, Gilles Michel, and Jérémy Clotault. On the characterization of flowering curves using Gaussian mixture models. Journal of Theoretical Biology, 402(??):75–88, August 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300509.

Proia:2019:PRG

[PPTC19] Frédéric Proïa, Fabien Panloup, Chiraz Trabelsi, and Jérémy Clotault. Probabilistic reconstruction of genealogies for polyploid plant species. Journal of Theoretical Biology, 462(??):537–551, February 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830585X.

Peace:2016:SGD

[PPW16] Angela Peace, Monica D. Poteat, and Hao Wang. Somatic growth dilution of a toxicant in a predator–prey model under stoichiometric constraints. Journal of Theoretical Biology, 407(??):198–211, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302193.

Potapov:2013:ATS

[PR13] Alex Potapov and Harshana Rajakaruna. Allee threshold and stochasticity in biological invasions: Colonization time at low propagule pressure. Journal of Theoretical Biology, 337(??):1–14, November 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003743.

Patel:2017:RLC

[PR17] Mainak Patel and Aaditya Rangan. Role of the locus coeruleus in the emergence of power law wake bouts in a model of the brainstem sleep-wake system through early infancy. Journal of Theoretical Biology, 426(??):82–95, August 7, 2017.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302497.

Paul:2018:RSS

[PR18] Debdas Paul and Nicole Radde. The role of stochastic sequestration dynamics for intrinsic noise filtering in signaling network motifs. Journal of Theoretical Biology, 455(??):86–96, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303370.

Pedruzzi:2015:MMM

[PRC15] Gabriele Pedruzzi, Kanury V. S. Rao, and Samrat Chatterjee. Mathematical model of mycobacterium-host interaction describes physiology of persistence. Journal of Theoretical Biology, 376(??):105–117, July 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001460.

Pokhilko:2011:ULS

[PRH^{+}11] Alexandra Pokhilko, Jason A. Ramos, Hans Holtan, Don R. Maszle, Rajnish Khanna, and Andrew J. Millar. Ubiquitin ligase switch in plant photomorphogenesis: a hypothesis. Journal of Theoretical Biology, 270(1):31–41, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006065.

Phillips:2013:ASF

[PRK13] A. J. K. Phillips, P. A. Robinson, and E. B. Klerman. Arousal state feedback as a potential physiological generator of the ultradian REM/NREM sleep cycle. Journal of Theoretical Biology, 319(??):75–87, February 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006182.

Palmer:2011:KMG

[PRM11] K. J. Palmer, M. S. Ridout, and B. J. T. Morgan. Kinetic models of guanidine hydrochloride-induced curing of the yeast [ PSI + ] prion. Journal of Theoretical Biology, 274(1):1–11, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006818.

Prosper:2012:ARS

[PRM12] Olivia Prosper, Nick Ruktanonchai, and Maia Martcheva. Assessing the role of spatial heterogeneity and human movement in malaria dynamics and control. Journal of Theoretical Biology, 303(??):1–14, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000835.

Prosper:2014:OVB

[PRM14] Olivia Prosper, Nick Ruktanonchai, and Maia Martcheva. Optimal vaccination and bednet maintenance for the control of malaria in a region with naturally acquired immunity. Journal of Theoretical Biology, 353(??):142–156, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314001489.

Plowes:2014:EBS

[PRMH14] Nicola J. R. Plowes, Kai Ramsch, Martin Middendorf, and Bert Hölldobler. An empirically based simulation of group foraging in the harvesting ant, Messor pergandei. Journal of Theoretical Biology, 340(??):186–198, January 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003378.

Perrin:2010:CTD

[PRN10] Dimitri Perrin, Heather J. Ruskin, and Tohru Niwa. Cell typedependent, infection-induced, aberrant DNA methylation in gastric cancer. Journal of Theoretical Biology, 264(2):570–577, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001153.

Parr:2011:ASV

[PRSC11] W. C. H. Parr, A. Ruto, C. Soligo, and H. J. Chatterjee. Allometric shape vector projection: a new method for the identification of allometric shape characters and trajectories applied to the human astragalus (talus). Journal of Theoretical Biology, 272(1):64–71, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006259.

Pires:2014:MFN

[PRV^{+}14] Marcelo Pires, Frank Raischel, Sandra H. Vaz, Andreia Cruz-Silva, Ana M. Sebastião, and Pedro G. Lind. Modeling the functional network of primary intercellular Ca^{2+} wave propagation in astrocytes and its application to study drug effects. Journal of Theoretical Biology, 356(??):201–212, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002446.

Prum:2012:HDG

[PS12] Richard O. Prum and Larry Samuelson. The Hairy–Downy Game: a model of interspecific social dominance mimicry. Journal of Theoretical Biology, 313(??):42–60, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312003633.

Panigrahi:2013:CSA

[PS13] Priya P. Panigrahi and Tiratha Raj Singh. Computational studies on Alzheimer’s disease associated pathways and regulatory patterns using microarray gene expression and network data: Revealed association with aging and other diseases. Journal of Theoretical Biology, 334(??):109–121, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002853.

Priya:2013:MBA

[PSAA13] S. Binny Priya, Subhojit Saha, Ramesh Anishetty, and Sharmila Anishetty. A matrix based algorithm for protein– protein interaction prediction using domain-domain associations. Journal of Theoretical Biology, 326(??):36–42, June 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000908.

Peng:2018:MBE

[PSD^{+}18] Valery Peng, Natalka Suchowerska, Ana Dos Santos Esteves, Linda Rogers, Elizabeth Claridge Mackonis, Joanne Toohey, and David R. McKenzie. Models for the bystander effect in gradient radiation fields: Range and signalling type. Journal of Theoretical Biology, 455(??):16–25, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303278.

Poorebrahim:2017:SES

[PSG^{+}17] Mansour Poorebrahim, Solmaz Sadeghi, Raziyeh Ghorbani, Matin Asghari, Mohammad Foad Abazari, Hourieh Kalhor, and Hamzeh Rahimi. In silico enhancement of the stability and activity of keratinocyte growth factor. Journal of Theoretical Biology, 418(??):111–121, April 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300085.

Pan:2015:HDD

[PSJ15] Wei Pan, Gui-Quan Sun, and Zhen Jin. How demographydriven evolving networks impact epidemic transmission between communities. Journal of Theoretical Biology, 382 (??):309–319, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519315003380.

Prohaska:2010:IGR

[PSK10] Sonja J. Prohaska, Peter F. Stadler, and David C. Krakauer. Innovation in gene regulation: The case of chromatin computation. Journal of Theoretical Biology, 265(1):27–44, July 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001347.

Preedy:2010:MCS

[PSL^{+}10] Katharine F. Preedy, Pietà G. Schofield, Sijia Liu, Anastasios Matzavinos, Mark A. J. Chaplain, and Stephen F. Hubbard. Modelling contact spread of infection in host-parasitoid systems: Vertical transmission of pathogens can cause chaos. Journal of Theoretical Biology, 262(3):441–451, February 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309004780.

Palk:2012:MEC

[PSP^{+}12] Laurence Palk, James Sneyd, Kate Patterson, Trevor J. Shuttleworth, David I. Yule, Oliver Maclaren, and Edmund J. Crampin. Modelling the effects of calcium waves and oscillations on saliva secretion. Journal of Theoretical Biology, 305(??):45–53, July 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001828.

Palk:2010:DMS

[PSS^{+}10] Laurence Palk, James Sneyd, Trevor J. Shuttleworth, David I. Yule, and Edmund J. Crampin. A dynamic model of saliva secretion. Journal of Theoretical Biology, 266(4):625–640, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003176. See corri-

gendum [PSS^{+}13].

Palk:2013:CDM

[PSS^{+}13] Laurence Palk, James Sneyd, Trevor J. Shuttleworth, David I. Yule, and Edmund J. Crampin. Corrigendum to ‘A dynamic model of saliva secretion’ [J. Theor. Biol. 266, 2010, 625– 640]. Journal of Theoretical Biology, 317(??):428, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005668. See [PSS^{+}10].

Petrovskii:2017:RSE

[PSV17] Sergei Petrovskii, Yadigar Sekerci, and Ezio Venturino. Regime shifts and ecological catastrophes in a model of plankton-oxygen dynamics under the climate change. Journal of Theoretical Biology, 424(??):91–109, July 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301820.

Pedro:2010:HAD

[PT10] S. A. Pedro and J. M. Tchuenche. HIV/AIDS dynamics: Impact of economic classes with transmission from poor clinical settings. Journal of Theoretical Biology, 267(4):471–485, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004935.

Powers:2012:PCP

[PTB12] Simon T. Powers, Daniel J. Taylor, and Joanna J. Bryson. Punishment can promote defection in group-structured populations. Journal of Theoretical Biology, 311(??):107–116, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200344X.

Poletto:2013:HMT

[PTC13] Chiara Poletto, Michele Tizzoni, and Vittoria Colizza. Human mobility and time spent at destination: Impact on spatial epidemic spreading. Journal of Theoretical Biology, 338(??):41–58, December 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313004062.

Poelwijk:2011:RSE

[PTNKT11] Frank J. Poelwijk, Sorin Tanase-Nicola, Daniel J. Kiviet, and Sander J. Tans. Reciprocal sign epistasis is a necessary condition for multi-peaked fitness landscapes. Journal of Theoretical Biology, 272(1):141–144, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006703.

Pepin:2010:PDV

[PVB^{+}10] K. M. Pepin, I. Volkov, J. R. Banavar, C. O. Wilke, and B. T. Grenfell. Phenotypic differences in viral immune escape explained by linking within-host dynamics to host-population immunity. Journal of Theoretical Biology, 265(4):501–510, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002791.

Pliego:2018:IES

[PVCEC18] Emilene Pliego Pliego, Jorge Velázquez-Castro, Markus P. Eichhorn, and Andrés Fraguela Collar. Increased efficiency in the second-hand tire trade provides opportunity for dengue control. Journal of Theoretical Biology, 437(??):126–136, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304848.

Prokharau:2012:MDB

[PVGA12] Pavel A. Prokharau, Fred J. Vermolen, and José Manuel García-Aznar. Model for direct bone apposition on preexisting surfaces, during peri-implant osseointegration. Journal of Theoretical Biology, 304(??):131–142, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200149X.

Portillo:2019:MMS

[PVGV19] A. M. Portillo, E. Varela, and J. A. García-Velasco. Mathematical model to study the aging of the human follicle according to the telomerase activity. Journal of Theoretical Biology, 462(??):446–454, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305897.

Postlethwaite:2011:GMI

[PW11] Claire M. Postlethwaite and Michael M. Walker. A geometric model for initial orientation errors in pigeon navigation. Journal of Theoretical Biology, 269(1):273–279, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005837.

Postlethwaite:2014:MNE

[PW14] Claire M. Postlethwaite and Michael M. Walker. A model for navigational errors in complex environmental fields. Journal of Theoretical Biology, 363(??):134–144, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004627.

Peck:2018:WAH

[PW18a] Joel R. Peck and David Waxman. What is adaptation and how should it be measured? Journal of Theoretical Biology, 447(??):190–198, June 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301103.

Phadumdeo:2018:HRV

[PW18b] Vrishti M. Phadumdeo and Seth H. Weinberg. Heart rate variability alters cardiac repolarization and electromechanical dynamics. Journal of Theoretical Biology, 442(??):31–43, April 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300158.

Picco:2019:TCY

[PW19] Noemi Picco and Thomas E. Woolley. Time to change your mind? Modelling transient properties of cortex formation highlights the importance of evolving cell division strategies. Journal of Theoretical Biology, 481(??):110–118, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303989.

Parr:2012:TIG

[PWC^{+}12] W. C. H. Parr, S. Wroe, U. Chamoli, H. S. Richards, M. R. McCurry, P. D. Clausen, and C. McHenry. Toward integration of geometric morphometrics and computational biomechanics: New methods for 3d virtual reconstruction and quantitative analysis of finite element models. Journal of Theoretical Biology, 301(??):1–14, May 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000409.

Panitchob:2015:CMA

[PWC^{+}15] N. Panitchob, K. L. Widdows, I. P. Crocker, M. A. Hanson, E. D. Johnstone, C. P. Please, C. P. Sibley, J. D. Glazier, R. M. Lewis, and B. G. Sengers. Computational modelling of amino acid exchange and facilitated transport in placental membrane vesicles. Journal of Theoretical Biology, 365(??):352–364, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006328.

Pineda:2015:MCM

[PWE15] M. Pineda, C. J. Weijer, and R. Eftimie. Modelling cell movement, cell differentiation, cell sorting and proportion regulation in Dictyostelium discoideum aggregations. Journal of Theoretical Biology, 370(??):135–150, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000594.

Palmer:2016:RSX

[PWG16] Michael L. Palmer, Richard A. Williams, and Derek Gatherer. Rosen’s (M, R) system as an X-machine. Journal of Theoretical Biology, 408(??):97–104, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302417.

Pan:2013:ACU

[PWH^{+}13] Lu-Lu Pan, Yu Wang, Jian-Hui Hu, Zhao-Tang Ding, and Chen Li. Analysis of codon use features of stearoyl–acyl carrier protein desaturase gene in Camellia sinensis. Journal of Theoretical Biology, 334(??):80–86, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002786.

Peng:2016:DLC

[PWHW16] Xiaoqing Peng, Jianxin Wang, Jun Huan, and Fang-Xiang Wu. Double-layer clustering method to predict protein complexes based on power-law distribution and protein sublocalization. Journal of Theoretical Biology, 395(??):186–193, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316000849.

Petroff:2011:RDM

[PWL^{+}11] Alexander P. Petroff, Ting-Di Wu, Biqing Liang, Jeannie Mui, Jean-Luc Guerquin-Kern, Hojatollah Vali, Daniel H. Rothman, and Tanja Bosak. Reaction-diffusion model of nutrient uptake in a biofilm: Theory and experiment. Journal of Theoretical Biology, 289(??):90–95, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003900.

Pan:2019:APA

[PWZ^{+}19] Yi Pan, Shiyuan Wang, Qi Zhang, Qianzi Lu, Dongqing Su, Yongchun Zuo, and Lei Yang. Analysis and prediction of animal toxins by various Chou’s pseudo components and reduced amino acid compositions. Journal of Theoretical Biology, 462(??):221–229, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305630.

Pandey:2019:MDA

[PYG^{+}19] Shrish Kumar Pandey, Saveg Yadav, Yugal Goel, Mithlesh Kumar Temre, Vinay Kumar Singh, and Sukh Mahendra Singh. Molecular docking of anti-inflammatory drug diclofenac with metabolic targets: Potential applications in cancer therapeutics. Journal of Theoretical Biology, 465(??):117–125, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300207.

Ponce:2019:AEM

[PZLF19] Joan Ponce, Yiqiang Zheng, Guang Lin, and Zhilan Feng. Assessing the effects of modeling the spectrum of clinical symptoms on the dynamics and control of Ebola. Journal of Theoretical Biology, 467(??):111–122, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931930013X.

Pivonka:2010:TIR

[PZS^{+}10] Peter Pivonka, Jan Zimak, David W. Smith, Bruce S. Gardiner, Colin R. Dunstan, Natalie A. Sims, T. John Martin, and Gregory R. Mundy. Theoretical investigation of the role of the RANK-RANKL-OPG system in bone remodeling. Journal of Theoretical Biology, 262(2):306–316, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900441X.

Zhang:2016:GPP

[pZWZ^{+}16] Yan ping Zhang, Wuyunqiqige, Wei Zheng, Shuyi Liu, and Chunguang Zhao. gDNA-prot: Predict DNA-binding proteins by employing support vector machine and a novel numerical characterization of protein sequence. Journal of Theoretical Biology, 406(??):8–16, October 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301308.

Quinn:2015:EBR

[QA15] Noel P. Quinn and Josef D. Ackerman. The effect of bottom roughness on scalar transport in aquatic ecosystems: Implications for reproduction and recruitment in the benthos. Journal of Theoretical Biology, 369(??):59–66, March 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000089.

Quiroz:2010:TBG

[QF10] G. Quiroz and R. Femat. Theoretical blood glucose control in hyper- and hypoglycemic and exercise scenarios by means of an H_{∞} algorithm. Journal of Theoretical Biology, 263(1):154–160, March 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005578.

Quijano:2016:STA

[QJR^{+}16] Alex John Quijano, Michele L. Joyner, Chelsea Ross, J. Colton Watts, Edith Seier, and Thomas C. Jones. Spatio-temporal analysis of foraging behaviors of Anelosimus studiosus utilizing mathematical modeling of multiple spider interaction on a cooperative web. Journal of Theoretical Biology, 408(??):243–259, November 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302557.

Qiu:2018:PPS

[QLC^{+}18] Wenying Qiu, Shan Li, Xiaowen Cui, Zhaomin Yu, Minghui Wang, Junwei Du, Yanjun Peng, and Bin Yu. Predicting protein submitochondrial locations by incorporating the pseudo-position specific scoring matrix into the general Chou’s pseudo-amino acid composition. Journal of Theoretical Biology, 450(??):86–103, August 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301942.

Qu:2019:NMB

[QLWC19] Rui Qu, Shu-Shen Liu, Ze-Jun Wang, and Fu Chen. A novel method based on similarity and triangulation for predicting the toxicities of various binary mixtures. Journal of Theoretical Biology, 480(??):56–64, November 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303030.

Qi:2011:SSA

[QLZQ11] Zhao-Hui Qi, Ling Li, Zhi-Meng Zhang, and Xiao-Qin Qi. Selfsimilarity analysis of eubacteria genome based on weighted graph. Journal of Theoretical Biology, 280(1):10–18, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001792.

Qi:2015:EIS

[QMJW15] Hang Qi, Shoufeng Ma, Ning Jia, and Guangchao Wang. Experiments on individual strategy updating in iterated snowdrift game under random rematching. Journal of Theoretical Biology, 368(??):1–12, March 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400705X.

Qiu:2013:SIM

[QQJW13] Zhijun Qiu, Cuili Qin, Min Jiu, and Xicheng Wang. A simple iterative method to optimize protein-ligand-binding residue prediction. Journal of Theoretical Biology, 317(??):219–223, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200567X.

Qi:2010:NMG

[QQL10] Zhao-Hui Qi, Xiao-Qin Qi, and Chen-Chen Liu. New method for global alignment of 2 DNA sequences by the tree data structure. Journal of Theoretical Biology, 263(2):227–236, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005797.

Qi:2011:CDR

[QW11] Zhao-Hui Qi and Ruo-Yan Wei. A combination dimensionality reduction approach to codon position patterns of eubacteria based on their complete genomes. Journal of Theoretical Biology, 272(1):26–34, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006697.

Qiu:2012:PPP

[QW12] Zhijun Qiu and Xicheng Wang. Prediction of protein–protein interaction sites using patch-based residue characterization. Journal of Theoretical Biology, 293(??):143–150, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311005364.

Qu:2010:ASR

[QYO10] Kun Qu, Anastasia M. Yesnik, and Peter J. Ortoleva. Alternative splicing regulatory network reconstruction from exon array data. Journal of Theoretical Biology, 263(4):471–480, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005980.

Qin:2014:SCA

[QZ14] Guimin Qin and Xing-Ming Zhao. A survey on computational approaches to identifying disease biomarkers based on molecular networks. Journal of Theoretical Biology, 362(??):9–16, December 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003440.

Qiu:2017:PPI

[QZY17] Zhijun Qiu, Bo Zhou, and Jiangfeng Yuan. Protein-protein interaction site predictions with minimum covariance determinant and Mahalanobis distance. Journal of Theoretical Biology, 433(??):57–63, November 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304009.

Rakshit:2010:NAL

[RA10] Sourav Rakshit and G. K. Ananthasuresh. A novel approach for large-scale polypeptide folding based on elastic networks using continuous optimization. Journal of Theoretical Biology, 262(3):488–497, February 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004858.

Radulescu:2013:HCI

[RA13] Anca Rǎdulescu and Paul Adams. Hebbian crosstalk and input segregation. Journal of Theoretical Biology, 337(??): 133–149, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519313003767.

Rimer:2017:KOD

[RA17] Oren Rimer and Gil Ariel. Kinetic order-disorder transitions in a pause-and-go swarming model with memory. Journal of Theoretical Biology, 419(??):90–99, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300644.

Rozum:2018:SSP

[RA18] Jordan C. Rozum and Réka Albert. Self-sustaining positive feedback loops in discrete and continuous systems. Journal of Theoretical Biology, 459(??):36–44, December 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304533.

Rathore:2016:SAG

[RAA^{+}16] Poonam K. Rathore, V. Arathy, Vijaylaxmi S. Attimarad, Pramod Kumar, and Subarna Roy. In-silico analysis of gymnemagenin from Gymnema sylvestre (Retz.) R.Br. with targets related to diabetes. Journal of Theoretical Biology, 391(??):95–101, February 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005792.

Rabhi:2015:NPB

[RAAS15] Mokded Rabhi, Abdallah Atia, Chedly Abdelly, and Abderrazak Smaoui. New parameters for a better evaluation of vegetative bioremediation, leaching, and phytodesalination. Journal of Theoretical Biology, 383(??):7–11, October 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315003720.

Rooman:2014:SNR

[RAD14] Marianne Rooman, Jaroslav Albert, and Mitia Duerinckx. Stochastic noise reduction upon complexification: Positively correlated birth-death type systems. Journal of Theoretical Biology, 354(??):113–123, August 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001428.

Radzvilavicius:2016:EDC

[Rad16] Arunas L. Radzvilavicius. Evolutionary dynamics of cytoplasmic segregation and fusion: Mitochondrial mixing facilitated the evolution of sex at the origin of eukaryotes. Journal of Theoretical Biology, 404(??):160–168, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301278.

Reynolds:2014:MPC

[RAF^{+}14] Joseph Reynolds, Inês F. Amado, Antonio A. Freitas, Grant Lythe, and Carmen Molina-París. A mathematical perspective on CD4 + T cell quorum-sensing. Journal of Theoretical Biology, 347(??):160–175, April 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005705.

Rayhan:2019:CCF

[RAF^{+}19] Farshid Rayhan, Sajid Ahmed, Dewan Md Farid, Abdollah Dehzangi, and Swakkhar Shatabda. CFSBoost: Cumulative feature subspace boosting for drug-target interaction prediction. Journal of Theoretical Biology, 464(??):1–8, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318306210.

Ramsey:2010:LPP

[Ram10] David M. Ramsey. A large population parental care game: Polymorphisms and feedback between patterns of care and the operational sex ratio. Journal of Theoretical Biology, 266(4):675–690, October 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003814.

Ruokolainen:2011:RSH

[RAMS11] Lasse Ruokolainen, Peter A. Abrams, Kevin S. McCann, and Brian J. Shuter. The roles of spatial heterogeneity and adaptive movement in stabilizing (or destabilizing) simple metacommunities. Journal of Theoretical Biology, 291(??):76–87, December 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004607.

Rand:2012:PMT

[Ran12] David G. Rand. The promise of Mechanical Turk: How online labor markets can help theorists run behavioral experiments. Journal of Theoretical Biology, 299(??):172–179, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001330.

Rand:2010:ASP

[RANO10] David G. Rand, Joseph J. Armao IV, Mayuko Nakamaru, and Hisashi Ohtsuki. Anti-social punishment can prevent the co-evolution of punishment and cooperation. Journal of Theoretical Biology, 265(4):624–632, August 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002948.

Rao:2012:UTI

[Rao12] Arni S. R. Srinivasa Rao. Understanding theoretically the impact of reporting of disease cases in epidemiology. Journal of Theoretical Biology, 302(??):89–95, June 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001105.

Roy:2019:VCI

[RAR19] Amit Kumar Roy, Fahad Al Basir, and Priti Kumar Roy. A vivid cytokines interaction model on psoriasis with the effect of impulse biologic (TNF-α inhibitor) therapy. Journal of Theoretical Biology, 474(??):63–77, August 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301481.

Reynolds:2013:WCS

[RB13] Sara A. Reynolds and Chad E. Brassil. When can a single-species, density-dependent model capture the dynamics of a consumer-resource system? Journal of Theoretical Biology, 339(??):70–83, December 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003998.

Ringa:2014:DCF

[RB14] N. Ringa and C. T. Bauch. Dynamics and control of footand-mouth disease in endemic countries: a pair approximation model. Journal of Theoretical Biology, 357(??):150– 159, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519314002811.

Root-Bernstein:2015:TIM

[RB15] Robert Root-Bernstein. Towards an integration of mathematical models, theories and observations concerning autoimmune diseases. Journal of Theoretical Biology, 375(??):1–3, June 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001575.

Roemer:2012:MMC

[RBB^{+}12] R. B. Roemer, D. Booth, A. A. Bhavsar, G. H. Walter, and L. I. Terry. Mathematical model of cycad cones’ thermogenic temperature responses: Inverse calorimetry to estimate metabolic heating rates. Journal of Theoretical Biology, 315(??):87–96, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004717.

Ritterskamp:2016:EEC

[RBB16a] Daniel Ritterskamp, Daniel Bearup, and Bernd Blasius. Emergence of evolutionary cycles in size-structured food webs. Journal of Theoretical Biology, 408(??):187–197, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302582.

Ritterskamp:2016:NDE

[RBB16b] Daniel Ritterskamp, Daniel Bearup, and Bernd Blasius. A new dimension: Evolutionary food web dynamics in two dimensional trait space. Journal of Theoretical Biology, 405(??):66–81, September 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300182.

Root-Bernstein:2015:UIT

[RBF15] Robert Root-Bernstein and DeLisa Fairweather. Unresolved issues in theories of autoimmune disease using myocarditis as a framework. Journal of Theoretical Biology, 375(??):101–123, June 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006742.

Rudrauf:2017:MME

[RBG^{+}17] David Rudrauf, Daniel Bennequin, Isabela Granic, Gregory Landini, Karl Friston, and Kenneth Williford. A mathematical model of embodied consciousness. Journal of Theoretical Biology, 428(??):106–131, September 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302540.

Ritchie:2014:HOS

[RBHK14] Martin Ritchie, Luc Berthouze, Thomas House, and Istvan Z. Kiss. Higher-order structure and epidemic dynamics in clustered networks. Journal of Theoretical Biology, 348(??):21– 32, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000423.

Rodriguez-Brenes:2019:RTS

[RBKW19] Ignacio A. Rodriguez-Brenes, Natalia L. Komarova, and Dominik Wodarz. The role of telomere shortening in carcinogenesis: a hybrid stochastic-deterministic approach. Journal of Theoretical Biology, 460(??):144–152, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304338.

Ruiz-Blanco:2013:GSP

[RBMPP^{+}13] Yasser B. Ruiz-Blanco, Yovani Marrero-Ponce, Waldo Paz, Yamila García, and Jesús Salgado. Global stability of protein folding from an empirical free energy function. Journal of Theoretical Biology, 321(??):44–53, March 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000052.

Ruiz-Blanco:2015:HLB

[RBMPP^{+}15] Yasser B. Ruiz-Blanco, Yovani Marrero-Ponce, Pablo J. Prieto, Jesús Salgado, Yamila García, and Clivia M. Sotomayor-Torres. A Hooke’s law-based approach to protein folding rate. Journal of Theoretical Biology, 364(??):407–417, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005402.

Rahimi:2017:OSB

[RBMS17] Maryam Rahimi, Mohammad Reza Bakhtiarizadeh, and Abdollah Mohammadi-Sangcheshmeh. OOgenesis Pred: a sequence-based method for predicting oogenesis proteins by six different modes of Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 414(??):128–136, February 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316304088.

Rezazadegan:2018:MPR

[RBR18] Reza Rezazadegan, Chris Barrett, and Christian Reidys. Multiplicity of phenotypes and RNA evolution. Journal of Theoretical Biology, 447(??):139–146, June 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301425.

Root-Bernstein:2015:RML

[RBRB15] Meredith Root-Bernstein and Robert Root-Bernstein. The ribosome as a missing link in the evolution of life. Journal of Theoretical Biology, 367(??):130–158, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006778.

Root-Bernstein:2016:RML

[RBRB16] Robert Root-Bernstein and Meredith Root-Bernstein. The ribosome as a missing link in prebiotic evolution II: Ribosomes encode ribosomal proteins that bind to common regions of their own mRNAs and rRNAs. Journal of Theoretical Biology, 397(??):115–127, May 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001260.

Roy:2010:GIF

[RBSD10] Manas K. Roy, Paromita Banerjee, Tapas K. Sengupta, and Sushanta Dattagupta. Glucose induced fractal colony pattern of Bacillus thuringiensis. Journal of Theoretical Biology, 265(3):389–395, August 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002511.

Requejo:2011:ECM

[RC11] Rubén J. Requejo and Juan Camacho. Evolution of cooperation mediated by limiting resources: Connecting resource based models and evolutionary game theory. Journal of Theoretical Biology, 272(1):35–41, March 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006478.

Rodriguez-Caso:2013:CCM

[RC13] Carlos Rodríguez-Caso. Can cell mortality determine division of labor in tissue organization? Journal of Theoretical Biology, 332(??):161–170, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002129.

Robertson:2016:HSS

[RC16] Suzanne L. Robertson and Kevin A. Caillouët. A host stagestructured model of enzootic West Nile virus transmission to explore the effect of avian stage-dependent exposure to vectors. Journal of Theoretical Biology, 399(??):33–42, June 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300017.

Royer-Carenzi:2016:CAS

[RCD16] Manuela Royer-Carenzi and Gilles Didier. A comparison of ancestral state reconstruction methods for quantitative characters. Journal of Theoretical Biology, 404(??):126–142, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301205.

Royer-Carenzi:2019:TCB

[RCD19] Manuela Royer-Carenzi and Gilles Didier. Testing for correlation between traits under directional evolution. Journal of Theoretical Biology, 482(??):Article 109982, December 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930325X.

Radulescu:2014:SSM

[RCH14] I. R. Rădulescu, D. Cândea, and A. Halanay. A study on stability and medical implications for a complex delay model for CML with cell competition and treatment. Journal of Theoretical Biology, 363(??):30–40, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004573.

Ren:2010:QPT

[RCL^{+}10] Yanrong Ren, Xiaolin Chen, Xiaobo Li, Han Lai, Qiang Wang, Peng Zhou, and Guoping Chen. Quantitative prediction of the thermal motion and intrinsic disorder of protein cofactors in crystalline state: a case study on halide anions. Journal of Theoretical Biology, 266(2):291–298, September 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003358.

Rozada:2016:CEH

[RCL16] Ignacio Rozada, Daniel Coombs, and Viviane D. Lima. Conditions for eradicating hepatitis C in people who inject drugs: a fibrosis aware model of hepatitis C virus transmission. Journal of Theoretical Biology, 395(??):31–39, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000692.

Rizvi:2012:RMA

[RD12] Mohd Suhail Rizvi and Sovan Lal Das. Role of membrane addition in animal cell cytokinesis. Journal of Theoretical Biology, 315(??):139–143, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004924.

Ray:2014:PCC

[RD14] Sanhita Ray and Anjan Kr. Dasgupta. Probiotics as cheater cells: Parameter space clustering for individualized prescription. Journal of Theoretical Biology, 361(??):165–174, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004160.

Rubin:2017:RES

[RD17] Ilan N. Rubin and Michael Doebeli. Rethinking the evolution of specialization: a model for the evolution of phenotypic heterogeneity. Journal of Theoretical Biology, 435(??):248–264, December 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304381.

Robert:2014:ATA

[RDD14] Marianne Robert, Laurent Dagorn, and Jean Louis Deneubourg. The aggregation of tuna around floating objects: What could be the underlying social mechanisms? Journal of Theoretical Biology, 359(??):161–170, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314003531.

Renton:2011:DCH

[RDMP11] Michael Renton, Art Diggle, Sudheesh Manalil, and Stephen Powles. Does cutting herbicide rates threaten the sustainability of weed management in cropping systems? Journal of Theoretical Biology, 283(1):14–27, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002529.

Roman:2016:TAU

[RDP16] Ahmed Roman, Debanjan Dasgupta, and Michel Pleimling. A theoretical approach to understand spatial organization in complex ecologies. Journal of Theoretical Biology, 403(??):10–16, August 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931630087X.

Rolls:2012:MHC

[RDSD^{+}12] D. A. Rolls, G. Daraganova, R. Sacks-Davis, M. Hellard, R. Jenkinson, E. McBryde, P. E. Pattison, and G. L.

Robins. Modelling hepatitis C transmission over a social network of injecting drug users. Journal of Theoretical Biology, 297(??):73–87, March 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006175.

Ruiz-Diaz:2013:IRP

[RDTHSM13] Claudia Patricia Ruiz-Diaz, Carlos Toledo-Hernández, Alberto M. Sabat, and Mariano Marcano. Immune response to a pathogen in corals. Journal of Theoretical Biology, 332(??):141–148, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001963.

Reynolds:2010:MMP

[REC10] Angela Reynolds, G. Bard Ermentrout, and Gilles Clermont. A mathematical model of pulmonary gas exchange under inflammatory stress. Journal of Theoretical Biology, 264(2):161– 173, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000159.

Reick:2012:SSS

[Rei12] Christian H. Reick. Self-similarity and scaling in two models of phyllotaxis and the selection of asymptotic divergence angles. Journal of Theoretical Biology, 313(??):181–200, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003955.

Remien:2015:MDS

[Rem15] Christopher H. Remien. Modeling the dynamics of stable isotope tissue-diet enrichment. Journal of Theoretical Biology, 367(??):14–20, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006705.

Ren:2013:GRS

[Ren13] Jiu-Sheng Ren. Growth and residual stresses of arterial walls. Journal of Theoretical Biology, 337(??):80–88, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003822.

Revilla:2015:NRR

[Rev15] Tomás A. Revilla. Numerical responses in resource-based mutualisms: a time scale approach. Journal of Theoretical Biology, 378(??):39–46, August 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001745.

Reynolds:2013:SPG

[Rey13] A. M. Reynolds. Selection pressures give composite correlated random walks Lévy walk characteristics. Journal of Theoretical Biology, 332(??):117–122, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002038.

Rodriguez-Fernandez:2015:PMA

[RFdL15] José Luis Rodríguez-Fernández and Mario García de Lacoba. Plasma membrane-associated superstructure: Have we overlooked a new type of organelle in eukaryotic cells? Journal of Theoretical Biology, 380(??):346–358, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002647.

Richard:2012:ODR

[RFM^{+}12] Morgiane Richard, Matthew Fryett, Samantha Miller, Ian Booth, Celso Grebogi, and Alessandro Moura. Optimality in DNA repair. Journal of Theoretical Biology, 292(??):39–43, January 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004279.

Riera-Fernandez:2012:NMS

[RFME^{+}12] Pablo Riera-Fernández, Cristian R. Munteanu, Manuel Escobar, Francisco Prado-Prado, Raquel Martín-Romalde, David Pereira, Karen Villalba, Aliuska Duardo-Sánchez, and Humberto González-Díaz. New Markov–Shannon entropy models to assess connectivity quality in complex networks:

From molecular to cellular pathway, parasite–host, neural, industry, and legal–social networks. Journal of Theoretical Biology, 293(??):174–188, January 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005315.

Rangarajan:2015:DOD

[RFS^{+}15] Nivedita Rangarajan, Zach Fox, Abhyudai Singh, Prakash Kulkarni, and Govindan Rangarajan. Disorder, oscillatory dynamics and state switching: the role of c-Myc. Journal of Theoretical Biology, 386(??):105–114, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004567.

Reneaux:2010:TRC

[RG10] Melissa Reneaux and Manoj Gopalakrishnan. Theoretical results for chemotactic response and drift of E. coli in a weak attractant gradient. Journal of Theoretical Biology, 266(1):99–106, September 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002961.

Rybinski:2012:MBS

[RG12] Mikolaj Rybiński and Anna Gambin. Model-based selection of the robust JAK-STAT activation mechanism. Journal of Theoretical Biology, 309(??):34–46, September 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002123.

Rockenfeller:2017:HEH

[RG17] R. Rockenfeller and M. Günther. Hill equation and Hatze’s muscle activation dynamics complement each other: enhanced pharmacological and physiological interpretability of modelled activity-pCa curves. Journal of Theoretical Biology, 431(??):11–24, October 27, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303545.

Rockenfeller:2018:IFS

[RG18] Robert Rockenfeller and Michael Günther. Inter-filament spacing mediates calcium binding to troponin: a simple geometric-mechanistic model explains the shift of forcelength maxima with muscle activation. Journal of Theoretical Biology, 454(??):240–252, October 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303035.

Ramirez:2010:ARS

[RGA^{+}10] A. Ramírez, J. Grasa, A. Alonso, F. Soteras, R. Osta, M. J. Muñoz, and B. Calvo. Active response of skeletal muscle: In vivo experimental results and model formulation. Journal of Theoretical Biology, 267(4):546–553, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004923.

Rezaei-Ghaleh:2017:RAB

[RGBR17] Nasrollah Rezaei-Ghaleh, Davood Bakhtiari, and Armin Rashidi. Reverse allostasis in biological systems: Minimal conditions and implications. Journal of Theoretical Biology, 426(??):134–139, August 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302473.

Rodriguez-Girones:2010:SEP

[RGCML10] Miguel A. Rodríguez-Gironés, Guadalupe Corcobado, and Jordi Moya-Laraño. Silk elasticity as a potential constraint on spider body size. Journal of Theoretical Biology, 266(3):430–435, October 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003218.

Risau-Gusman:2012:MEP

[RGG12] Sebastián Risau-Gusman and Pablo M. Gleiser. Modelling the effect of phosphorylation on the circadian clock of Drosophila. Journal of Theoretical Biology, 307(??):53–61, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312002287.

Roberts:2017:MMR

[RGL^{+}17] Paul A. Roberts, Eamonn A. Gaffney, Philip J. Luthert, Alexander J. E. Foss, and Helen M. Byrne. Mathematical models of retinitis pigmentosa: The oxygen toxicity hypothesis. Journal of Theoretical Biology, 425(??):53–71, July 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931730214X.

Rosandic:2013:SSC

[RGP13] Marija Rosandić, Matko Gluncić, and Vladimir Paar. Start/ stop codon like trinucleotides extensions in primate alpha satellites. Journal of Theoretical Biology, 317(??):301–309, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005000.

Rhodes:2019:MMI

[RH19] Adam Rhodes and Thomas Hillen. A mathematical model for the immune-mediated theory of metastasis. Journal of Theoretical Biology, 482(??):Article 109999, December 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303492.

Rieger:2011:MBD

[RHJ11] Romain Rieger, Ridha Hambli, and Rachid Jennane. Modeling of biological doses and mechanical effects on bone transduction. Journal of Theoretical Biology, 274(1):36–42, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311000075.

Rockwell:2014:LHS

[RHS14a] Fulton E. Rockwell, N. Michele Holbrook, and Abraham D. Stroock. Leaf hydraulics i: Scaling transport properties from single cells to tissues. Journal of Theoretical Biology, 340(??):251–266, January 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313004827.

Rockwell:2014:LHI

[RHS14b] Fulton E. Rockwell, N. Michele Holbrook, and Abraham D. Stroock. Leaf hydraulics II: Vascularized tissues. Journal of Theoretical Biology, 340(??):267–284, January 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004013.

Ruiz-Herrera:2018:EDT

[RHT18] Alfonso Ruiz-Herrera and Pedro J. Torres. Effects of diffusion on total biomass in simple metacommunities. Journal of Theoretical Biology, 447(??):12–24, June 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301322.

Richards:2017:VIO

[Ric17] Kelvin J. Richards. Viral infections of oceanic plankton blooms. Journal of Theoretical Biology, 412(??):27–35, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303186.

Richard:2019:PNC

[Ric19] Adrien Richard. Positive and negative cycles in Boolean networks. Journal of Theoretical Biology, 463(??):67–76, February 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305812.

Renardy:2018:CCF

[RJSC18] Marissa Renardy, Alexandra Jilkine, Leili Shahriyari, and Ching-Shan Chou. Control of cell fraction and population recovery during tissue regeneration in stem cell lineages. Journal of Theoretical Biology, 445(??):33–50, May 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300778.

Revilla:2018:CTM

[RK18] Tomás A. Revilla and Vlastimil Křivan. Competition, traitmediated facilitation, and the structure of plant-pollinator communities. Journal of Theoretical Biology, 440(??):42–57, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305581.

Renardy:2019:EVS

[RK19] Marissa Renardy and Denise E. Kirschner. Evaluating vaccination strategies for tuberculosis in endemic and non-endemic settings. Journal of Theoretical Biology, 469(??):1–11, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930089X.

Roterman:2011:TIM

[RKJ^{+}11] I. Roterman, L. Konieczny, W. Jurkowski, K. Prymula, and M. Banach. Two-intermediate model to characterize the structure of fast-folding proteins. Journal of Theoretical Biology, 283(1):60–70, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002694.

Rhodes:2019:DFN

[RKK^{+}19] Sophie J. Rhodes, Gwenan M. Knight, Denise E. Kirschner, Richard G. White, and Thomas G. Evans. Dose finding for new vaccines: The role for immunostimulation/immunodynamic modelling. Journal of Theoretical Biology, 465(??):51–55, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300177.

Ristanovic:2012:MMT

[RKMG12] Dusan Ristanović, Bojana Krstonosić, Nebojsa T. Milosević, and Radmila Gudović. Mathematical modelling of transformations of asymmetrically distributed biological data: An application to a quantitative classification of spiny neurons of the human putamen. Journal of Theoretical Biology, 302(??):81– 88, June 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001117.

Riley:2015:SAS

[RL15] Emily E. Riley and Eric Lauga. Small-amplitude swimmers can self-propel faster in viscoelastic fluids. Journal of Theoretical Biology, 382(??):345–355, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003240.

Rajakaruna:2017:TCA

[RL17] Harshana Rajakaruna and Mark Lewis. Temperature cycles affect colonization potential of calanoid copepods. Journal of Theoretical Biology, 419(??):77–89, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730053X.

Riotte-Lambert:2013:PAM

[RLBCJ13] Louise Riotte-Lambert, Simon Benhamou, and Simon Chamaillé-Jammes. Periodicity analysis of movement recursions. Journal of Theoretical Biology, 317(??):238–243, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005656.

Roca-Lacostena:2018:EKM

[RLFS18] Jordi Roca-Lacostena and Jesús Fernández-Sánchez. Embeddability of Kimura 3ST Markov matrices. Journal of Theoretical Biology, 445(??):128–135, May 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300572.

Raghib:2010:MAC

[RLK10] M. Raghib, S. A. Levin, and I. G. Kevrekidis. Multiscale analysis of collective motion and decision-making in swarms: An advection-diffusion equation with memory approach. Journal of Theoretical Biology, 264(3):893–913, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001050.

Roselius:2014:MAG

[RLM^{+}14] Louisa Roselius, Dirk Langemann, Johannes Müller, Burkhard A. Hense, Stefan Filges, Dieter Jahn, and Richard Münch. Modelling and analysis of a gene-regulatory feed-forward loop with basal expression of the second regulator. Journal of Theoretical Biology, 363(??):290–299, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005281.

Richter:2017:ASW

[RLSM17] Yvonne Richter, Pedro G. Lind, Gunnar Seemann, and Philipp Maass. Anatomical and spiral wave reentry in a simplified model for atrial electrophysiology. Journal of Theoretical Biology, 419(??):100–107, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300620.

Rubin:2014:MEB

[RLSN14] Katy J. Rubin, Katherine Lawler, Peter Sollich, and Tony Ng. Memory effects in biochemical networks as the natural counterpart of extrinsic noise. Journal of Theoretical Biology, 357(??):245–267, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003397.

Ramakrishnan:2014:DDM

[RLW^{+}14] Sridhar Ramakrishnan, Srinivas Laxminarayan, Nancy J. Wesensten, Gary H. Kamimori, Thomas J. Balkin, and Jaques Reifman. Dose-dependent model of caffeine effects on human vigilance during total sleep deprivation. Journal of Theoretical Biology, 358(??):11–24, October 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002884.

Rhodes:2010:IVI

[RM10] C. J. Rhodes and A. P. Martin. The influence of viral infection on a plankton ecosystem undergoing nutrient enrichment. Journal of Theoretical Biology, 265(3):225–237, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002055.

Ruokolainen:2013:EWT

[RM13] Lasse Ruokolainen and Kevin McCann. Environmental weakening of trophic interactions drives stability in stochastic food webs. Journal of Theoretical Biology, 339(??):36–46, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003950.

Radulescu:2017:MMR

[RM17a] Anca Rădulescu and Rachel Marra. A mathematical model of reward and executive circuitry in obsessive compulsive disorder. Journal of Theoretical Biology, 414(??):165–175, February 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316304076.

Ramirez:2017:CNS

[RM17b] Juan Camilo Ramírez and James A. R. Marshall. Can natural selection encode Bayesian priors? Journal of Theoretical Biology, 426(??):57–66, August 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302254.

Rusinko:2017:STE

[RM17c] Joseph Rusinko and Matthew McPartlon. Species tree estimation using neighbor joining. Journal of Theoretical Biology, 414(??):5–7, February 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303587.

Ryser:2017:BRS

[RM17d] Marc D. Ryser and Kevin A. Murgas. Bone remodeling as a spatial evolutionary game. Journal of Theoretical Biology, 418(??):16–26, April 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300218.

Rout:2019:SAP

[RM19] Subhashree Rout and Rajani Kanta Mahapatra. In silico analysis of Plasmodium falciparum CDPK5 protein through molecular modeling, docking and dynamics. Journal of Theoretical Biology, 461(??):254–267, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305277.

Rauwolf:2015:VHB

[RMB15] Paul Rauwolf, Dominic Mitchell, and Joanna J. Bryson. Value homophily benefits cooperation but motivates employing incorrect social information. Journal of Theoretical Biology, 367(??):246–261, February 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006754.

Rahman:2018:HDM

[RMF^{+}18] Ashrafur Rahman, Daniel Munther, Aamir Fazil, Ben Smith, and Jianhong Wu. With-in host dynamics of L. monocytogenes and thresholds for distinct infection scenarios. Journal of Theoretical Biology, 454(??):80–90, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302741.

Rajkovic:2016:MMN

[RMM^{+}16a] Katarina Rajković, Dušica L. Marić, Nebojša T. Milošević, Sanja Jeremic, Valentina Arsić Arsenijević, and Nemanja Rajković. Mathematical modeling of the neuron morphology using two dimensional images. Journal of Theoretical Biology, 390(??):80–85, February 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005706.

Riahi:2016:MCH

[RMM^{+}16b] Aouatef Riahi, Abdelmonem Messaoudi, Ridha Mrad, Asma Fourati, Habiba Chabouni-Bouhamed, and Maher Kharrat. Molecular characterization, homology modeling and docking studies of the r2787h missense variation in BRCA2 gene: Association with breast cancer. Journal of Theoretical Biology, 403(??):188–196, August 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300911.

Ramirez-Malule:2016:ISC

[RMRC^{+}16] Howard Ramirez-Malule, Albeiro Restrepo, Wilson Cardona, Stefan Junne, Peter Neubauer, and Rigoberto Rios-Estepa. Inversion of the stereochemical configuration (3s, 5s)-clavaminic acid into (3r, 5r)-clavulanic acid: a computationally-assisted approach based on experimental evidence. Journal of Theoretical Biology, 395(??):40–50, April 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000679.

Rahimi:2013:EFS

[RMSTG13] Amir Rahimi, Armin Madadkar-Sobhani, Rouzbeh Touserkani, and Bahram Goliaei. Efficacy of function specific 3Dmotifs in enzyme classification according to their EC-numbers. Journal of Theoretical Biology, 336(??):36–43, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931300324X.

Rand:2012:EDF

[RN12] David G. Rand and Martin A. Nowak. Evolutionary dynamics in finite populations can explain the full range of cooperative behaviors observed in the centipede game. Journal of Theoretical Biology, 300(??):212–221, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000124.

Rydzewski:2015:IPS

[RNN15] Jakub Rydzewski, Wieslaw Nowak, and Giuseppe Nicosia. Inferring pathological states in cortical neuron microcircuits. Journal of Theoretical Biology, 386(??):34–43, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315004361.

Randic:2010:SPM

[RNVP10] Milan Randić, Marjana Novic, Marjan Vracko, and Dejan Plavsić. Study of proteome maps using partial ordering. Journal of Theoretical Biology, 266(1):21–28, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310002924.

Robinson:2011:NFT

[Rob11] P. A. Robinson. Neural field theory of synaptic plasticity. Journal of Theoretical Biology, 285(1):156–163, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311003225.

Reeves:2017:APG

[ROF17] T. Reeves, H. Ohtsuki, and S. Fukui. Asymmetric public goods game cooperation through pest control. Journal of Theoretical Biology, 435(??):238–247, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304174.

Rogers:2019:EGC

[Rog19] Scott O. Rogers. Evolution of the genetic code based on conservative changes of codons, amino acids, and aminoacyl tRNA synthetases. Journal of Theoretical Biology, 466(??):1– 10, April 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300281.

Rosas:2010:EGT

[Ros10a] Alejandro Rosas. Evolutionary game theory meets social science: Is there a unifying rule for human cooperation? Journal of Theoretical Biology, 264(2):450–456, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000901.

Ross:2010:CEM

[Ros10b] J. V. Ross. Computationally exact methods for stochastic periodic dynamics: Spatiotemporal dispersal and temporally forced transmission. Journal of Theoretical Biology, 262(1):14–22, January 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004342.

Ross:2011:IID

[Ros11] J. V. Ross. Invasion of infectious diseases in finite homogeneous populations. Journal of Theoretical Biology, 289(??):83–89, November 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004383.

Rosenstrom:2013:BMD

[Ros13] Tom Rosenström. Bargaining models of depression and evolution of cooperation. Journal of Theoretical Biology, 331(??):54–65, August 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001653.

Rose:2015:MMC

[Ros15] Noel R. Rose. Molecular mimicry and clonal deletion: a fresh look. Journal of Theoretical Biology, 375(??):71–76, June 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004950.

Rouch:2014:EFG

[Rou14] Duncan A. Rouch. Evolution of the first genetic cells and the universal genetic code: a hypothesis based on macromolecular coevolution of RNA and proteins. Journal of Theoretical Biology, 357(??):220–244, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003403.

Richards:2018:NKA

[RP18] Christopher T. Richards and Laura B. Porro. A novel kinematics analysis method using quaternion interpolation — a case study in frog jumping. Journal of Theoretical Biology, 454(??):410–424, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303047.

Ross:2018:GMF

[RPCW18] Cody Ross, Luis Pacheco-Cobos, and Bruce Winterhalder. A general model of forager search: Adaptive encounterconditional heuristics outperform Lévy flights in the search for patchily distributed prey. Journal of Theoretical Biology, 455(??):357–369, October 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303552.

Ristl:2014:CRF

[RPD14] Konstantin Ristl, Sebastian J. Plitzko, and Barbara Drossel. Complex response of a food-web module to symmetric and asymmetric migration between several patches. Journal of Theoretical Biology, 354(??):54–59, August 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001441.

Rangel:2019:BEP

[RPGG^{+}19] G. Capo Rangel, J. Prezioso, L. Gerardo-Giorda, E. Somersalo, and D. Calvetti. Brain energetics plays a key role in the coordination of electrophysiology, metabolism and hemodynamics: Evidence from an integrated computational model. Journal of Theoretical Biology, 478(??):26–39, October 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319302309.

Rizzo:2016:NME

[RPP16] Alessandro Rizzo, Biagio Pedalino, and Maurizio Porfiri. A network model for Ebola spreading. Journal of Theoretical Biology, 394(??):212–222, April 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000485.

Rowland:2017:MCM

[RQB17] J. Mark Rowland, Clifford R. Qualls, and Bruno A. Buzatto. A model for conditional male trimorphisms. Journal of Theoretical Biology, 419(??):184–192, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300619.

Ruokolainen:2012:SSI

[RR12] Lasse Ruokolainen and Jörgen Ripa. The strength of species interactions modifies population responses to environmental variation in competitive communities. Journal of Theoretical Biology, 310(??):199–205, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003281.

Reimers:2016:SSA

[RR16] Alexandra-M. Reimers and Arne C. Reimers. The steadystate assumption in oscillating and growing systems. Journal of Theoretical Biology, 406(??):176–186, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301655.

Ryabov:2010:VDC

[RRB10] Alexei B. Ryabov, Lars Rudolf, and Bernd Blasius. Vertical distribution and composition of phytoplankton under the influence of an upper mixed layer. Journal of Theoretical Biology, 263(1):120–133, March 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005232.

Ramsch:2012:MMF

[RRBM12] Kai Ramsch, Chris R. Reid, Madeleine Beekman, and Martin Middendorf. A mathematical model of foraging in a dynamic environment by trail-laying Argentine ants. Journal of Theoretical Biology, 306(??):32–45, August 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001762.

Richardson:2011:CPN

[RRC^{+}11] T. O. Richardson, E. J. H. Robinson, K. Christensen, J. H. Jensen, K. Christensen, H. J. Jensen, N. R. Franks, and A. B. Sendova-Franks. Comment on P. Nouvellet, J. P. Bacon, D. Waxman, “Testing the level of ant activity associated with quorum sensing: An empirical approach leading to the establishment and test of a null-model”. Journal of Theoretical Biology, 269(1):356–358, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005898.

Rahbar:2010:SAA

[RRG^{+}10] Mohammad Reza Rahbar, Iraj Rasooli, Seyed Latif Mousavi Gargari, Jafar Amani, and Yaser Fattahian. In silico analysis of antibody triggering biofilm associated protein in Acinetobacter baumannii. Journal of Theoretical Biology, 266(2):275–290, September 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310002985.

Rahbar:2012:PSA

[RRG^{+}12] Mohammad Reza Rahbar, Iraj Rasooli, Seyed Latif Mousavi Gargari, Gunnar Sandstrom, Jafar Amani, Yaser Fattahian, Abolfazl Jahangiri, and Marziyeh Jalali. A potential in silico antibody-antigen based diagnostic test for precise identification of Acinetobacter baumannii. Journal of Theoretical Biology, 294(??):29–39, February 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005418.

Ramakrishnan:2013:BMR

[RRL^{+}13] Sridhar Ramakrishnan, Srinivasan Rajaraman, Srinivas Laxminarayan, Nancy J. Wesensten, Gary H. Kamimori, Thomas J. Balkin, and Jaques Reifman. A biomathematical model of the restoring effects of caffeine on cognitive performance during sleep deprivation. Journal of Theoretical Biology, 319(??):23–33, February 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005905.

Ross:2015:SPM

[RRR15] Caitlin Ross, Jan Rychtár, and Olav Rueppell. A structured population model suggests that long life and post-reproductive lifespan promote the evolution of cooperation. Journal of Theoretical Biology, 369(??):85–94, March 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000296.

Roman-Roman:2016:MTG

[RRRRSPTR16] Patricia Román-Román, Sergio Román-Román, Juan José Serrano-Pérez, and Francisco Torres-Ruiz. Modeling tumor growth in the presence of a therapy with an effect on rate growth and variability by means of a modified Gompertz diffusion process. Journal of Theoretical Biology, 407(??):1–17, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302065.

Roman-Roman:2010:DPM

[RRRTR10] Patricia Román-Román, Desirée Romero, and Francisco Torres-Ruiz. A diffusion process to model generalized von Bertalanffy growth patterns: Fitting to real data. Journal of Theoretical Biology, 263(1):59–69, March 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005761.

Roman-Roman:2012:IET

[RRTR12] Patricia Román-Román and Francisco Torres-Ruiz. Inferring the effect of therapies on tumor growth by using diffusion processes. Journal of Theoretical Biology, 314(??):34–56, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200392X.

Russell:2019:SCQ

[RRÚJ19] Timothy W. Russell, Matthew J. Russell, Francisco Úbeda, and Vincent A. A. Jansen. Stable cycling in quasilinkage equilibrium: Fluctuating dynamics under gene conversion and selection. Journal of Theoretical Biology, 477(??):84–95, September 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302371.

Reznik:2010:SMC

[RS10] Ed Reznik and Daniel Segrè. On the stability of metabolic cycles. Journal of Theoretical Biology, 266(4):536–549, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003802.

Richards:2012:ERC

[RS12] Christopher T. Richards and Gregory S. Sawicki. Elastic recoil can either amplify or attenuate muscle-tendon power, depending on inertial vs. fluid dynamic loading. Journal of Theoretical Biology, 313(??):68–78, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003864.

Robertson:2014:EEM

[RS14a] Benjamin D. Robertson and Gregory S. Sawicki. Exploiting elasticity: Modeling the influence of neural control on mechanics and energetics of ankle muscle-tendons during human hopping. Journal of Theoretical Biology, 353(??):121–132, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001453.

Roy:2014:TAD

[RS14b] Tuhin K. Roy and Timothy W. Secomb. Theoretical analysis of the determinants of lung oxygen diffusing capacity. Journal of Theoretical Biology, 351(??):1–8, June 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000848.

Royero:2014:CWH

[RS14c] S. P. Vega Royero and G. J. Sibona. Can we heal Chagas infection? Journal of Theoretical Biology, 340(??):23–29, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004232.

Rocha:2019:ETB

[RS19] Mariana Rocha and Roger Springett. Electron transfer between cytochrome c and the binuclear center of cytochrome oxidase. Journal of Theoretical Biology, 460(??):134–141, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304983.

Rossetti:2010:EPT

[RSBB10] Valentina Rossetti, Bettina E. Schirrmeister, Marco V. Bernasconi, and Homayoun C. Bagheri. The evolutionary path to terminal differentiation and division of labor in cyanobacteria. Journal of Theoretical Biology, 262(1):23–34, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309004299.

Reynolds:2014:DAD

[RSC14] Andy M. Reynolds, Patrick Schultheiss, and Ken Cheng. Does the Australian desert ant Melophorus bagoti approximate a Lévy search by an intrinsic bi-modal walk? Journal of Theoretical Biology, 340(??):17–22, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300427X.

Robinson:2012:SDA

[RSD12] Marguerite Robinson, Nikolaos I. Stilianakis, and Yannis Drossinos. Spatial dynamics of airborne infectious diseases. Journal of Theoretical Biology, 297(??):116–126, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311006321.

Roblitz:2013:MMH

[RSD^{+}13] Susanna Röblitz, Claudia Stötzel, Peter Deuflhard, Hannah M. Jones, David-Olivier Azulay, Piet H. van der Graaf, and Steven W. Martin. A mathematical model of the human menstrual cycle for the administration of GnRH analogues. Journal of Theoretical Biology, 321(??):8–27, March 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005954.

Raicar:2016:IPF

[RSD^{+}16] Gaurav Raicar, Harsh Saini, Abdollah Dehzangi, Sunil Lal, and Alok Sharma. Improving protein fold recognition and structural class prediction accuracies using physicochemical properties of amino acids. Journal of Theoretical Biology, 402(??):117–128, August 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300741.

Reluga:2019:DGT

[RSH19] Timothy C. Reluga, Rachel A. Smith, and David P. Hughes. Dynamic and game theory of infectious disease stigmas. Journal of Theoretical Biology, 476(??):95–107, September 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302188.

Rempala:2011:MCA

[RSI11] Grzegorz A. Rempala, MichaL Seweryn, and Leszek Ignatowicz. Model for comparative analysis of antigen receptor repertoires. Journal of Theoretical Biology, 269(1):1–15, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005291.

Roques:2011:SRD

[RSR11] Lionel Roques, Samuel Soubeyrand, and Jérôme Rousselet. A statistical-reaction-diffusion approach for analyzing expansion processes. Journal of Theoretical Biology, 274(1):43–51, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000105.

Rahman:2018:DPD

[RSS^{+}18] M. Saifur Rahman, Swakkhar Shatabda, Sanjay Saha, M. Kaykobad, and M. Sohel Rahman. DPP-PseAAC: a DNA-binding protein prediction model using Chou’s general PseAAC. Journal of Theoretical Biology, 452(??):22–34, September 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302327.

Roper:2010:CSI

[RSV10] Ryan T. Roper, Maria Pia Saccomani, and Paolo Vicini. Cellular signaling identifiability analysis: a case study. Journal of Theoretical Biology, 264(2):528–537, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001049.

Rudnicki:2015:SGE

[RT15] Ryszard Rudnicki and Andrzej Tomski. On a stochastic gene expression with pre-mRNA, mRNA and protein contribution. Journal of Theoretical Biology, 387(??):54–67, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004555.

Reed:2019:DLP

[RT19] Daniel T. Reed and Colin R. Tosh. Diversity loss is predicted to increase extinction risk of specialist animals by constraining their ability to expand niche. Journal of Theoretical Biology, 476(??):44–50, September 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302140.

Robertson-Tessi:2012:MMT

[RTEKG12] Mark Robertson-Tessi, Ardith El-Kareh, and Alain Goriely. A mathematical model of tumor-immune interactions. Journal of Theoretical Biology, 294(??):56–73, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100542X.

Robertson-Tessi:2015:MEA

[RTEKG15] Mark Robertson-Tessi, Ardith El-Kareh, and Alain Goriely. A model for effects of adaptive immunity on tumor response to chemotherapy and chemoimmunotherapy. Journal of Theoretical Biology, 380(??):569–584, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500288X.

Robertson-Tessi:2019:CMM

[RTEKG19a] Mark Robertson-Tessi, Ardith El-Kareh, and Alain Goriely. Corrigendum to “A mathematical model of tumor-immune interactions” [Journal of Theoretical Biology 294 (2012) 56– 73]. Journal of Theoretical Biology, 464(??):180, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318306477.

Robertson-Tessi:2019:CME

[RTEKG19b] Mark Robertson-Tessi, Ardith El-Kareh, and Alain Goriely. Corrigendum to “A model for effects of adaptive immunity on tumor response to chemotherapy and chemoimmunotherapy” [Journal of Theoretical Biology 380 (2015) 569–584]. Journal of Theoretical Biology, 464(??):181, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306465.

Raybaud:2017:CCC

[RTFP^{+}17] Virginie Raybaud, Sylvie Tambutté, Christine Ferrier-Pagès, Stéphanie Reynaud, Alexander A. Venn, Éric Tambutté, Paul Nival, and Denis Allemand. Computing the carbonate chemistry of the coral calcifying medium and its response to ocean acidification. Journal of Theoretical Biology, 424(??):26–36, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301984.

Rajdev:2013:UMM

[RTR^{+}13] Pooja Rajdev, David Thorsley, Srinivasan Rajaraman, Tracy L. Rupp, Nancy J. Wesensten, Thomas J. Balkin, and Jaques Reifman. A unified mathematical model to quantify performance impairment for both chronic sleep restriction and total sleep deprivation. Journal of Theoretical Biology, 331(??):66–77, August 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001811.

Ramirez-Torres:2017:RMT

[RTRRS^{+}17] Ariel Ramírez-Torres, Reinaldo Rodríguez-Ramos, Federico J. Sabina, Catherine García-Reimbert, Raimondo Penta, José Merodio, Raúl Guinovart-Díaz, Julián Bravo-Castillero, Aura Conci, and Luigi Preziosi. The role of malignant tissue on the thermal distribution of cancerous breast. Journal of Theoretical Biology, 426(??):152–161, August 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302552.

Roa-Ureta:2016:SDM

[RU16] Ruben H. Roa-Ureta. Size-dependent mortality rate profiles. Journal of Theoretical Biology, 402(??):107–116, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931630090X.

Rasolonjanahary:2016:SMP

[RV16] Manan’Iarivo Rasolonjanahary and Bakhtier Vasiev. Scaling of morphogenetic patterns in reaction-diffusion systems. Journal of Theoretical Biology, 404(??):109–119, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301254.

Roos:2011:ESD

[RVA^{+}11] Gert Roos, Sam Van Wassenbergh, Peter Aerts, Anthony Herrel, and Dominique Adriaens. Effects of snout dimensions on the hydrodynamics of suction feeding in juvenile and adult seahorses. Journal of Theoretical Biology, 269(1):307–317, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005576.

Rosandic:2019:NLD

[RVP19] Marija Rosandić, Ines Vlahović, and Vladimir Paar. Novel look at DNA and life-symmetry as evolutionary forcing. Journal of Theoretical Biology, 483(??):Article 109985, December 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319303285.

Rahman:2014:ITG

[RVZ14] S. M. Ashrafur Rahman, Naveen K. Vaidya, and Xingfu Zou. Impact of Tenofovir gel as a PrEP on HIV infection: a mathematical model. Journal of Theoretical Biology, 347(??):151–159, April 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005729.

Rateitschak:2010:TTD

[RW10] Katja Rateitschak and Olaf Wolkenhauer. Thresholds in transient dynamics of signal transduction pathways. Journal of Theoretical Biology, 264(2):334–346, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000603.

Roy:2012:IHS

[RW12] Sarah M. Roy and Dominik Wodarz. Infection of HIV-specific CD4 T helper cells and the clonal composition of the response. Journal of Theoretical Biology, 304(??):143–151, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312001324.

Roy:2014:TAF

[RW14a] Sarah M. Roy and Dominik Wodarz. Tissue architecture, feedback regulation, and resilience to viral infection. Journal of Theoretical Biology, 340(??):131–138, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004396.

Ruxton:2014:EAP

[RW14b] Graeme D. Ruxton and David M. Wilkinson. Energetic arguments predict larger-bodied animals will be increasingly confined to flat environments. Journal of Theoretical Biology, 355(??):236–238, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002355.

Randhawa:2015:MMP

[RW15a] Avleen Randhawa and James M. Wakeling. Multidimensional models for predicting muscle structure and fascicle pennation. Journal of Theoretical Biology, 382(??):57–63, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002805.

Ross:2015:SWV

[RW15b] Cody T. Ross and Bruce Winterhalder. Sit-and-wait versus active-search hunting: a behavioral ecological model of optimal search mode. Journal of Theoretical Biology, 387(??):76–87, December 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004658.

Rabbani:2016:DMP

[RW16] Mahnaz Rabbani and Lindi M. Wahl. The dynamics of mobile promoters: Enhanced stability in promoter regions. Journal of Theoretical Biology, 407(??):401–408, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302132.

Rosenstrom:2016:SUT

[RWH16] Tom Rosenström, Karoline Wiesner, and Alasdair I. Houston. Scalar utility theory and proportional processing: What does it actually imply? Journal of Theoretical Biology, 404(??):222–235, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630131X.

Reynolds:2011:PDC

[RWSB11] Jennifer J. H. Reynolds, Andrew White, Jonathan A. Sherratt, and Mike Boots. The population dynamical consequences of density-dependent prophylaxis. Journal of Theoretical Biology, 288(??):1–8, November 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003857.

Reddyhoff:2015:TAM

[RWW^{+}15] Dennis Reddyhoff, John Ward, Dominic Williams, Sophie Regan, and Steven Webb. Timescale analysis of a mathematical model of acetaminophen metabolism and toxicity.

Journal of Theoretical Biology, 386(??):132–146, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315004142.

Rich:2016:HTR

[RZ16] Patricia Rich and Kevin J. S. Zollman. Honesty through repeated interactions. Journal of Theoretical Biology, 395 (??):238–244, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519316000801.

Rossi:2016:TIH

[RZIR16] Franca Rossi, Teresa Zotta, Lucilla Iacumin, and Anna Reale. Theoretical insight into the heat shock response (HSR) regulation in Lactobacillus casei and L. rhamnosus. Journal of Theoretical Biology, 402(??):21–37, August 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300571.

Ramirez-Zuniga:2019:MME

[RZRSC19] Ivan Ramirez-Zuniga, Jonathan E. Rubin, David Swigon, and Gilles Clermont. Mathematical modeling of energy consumption in the acute inflammatory response. Journal of Theoretical Biology, 460(??):101–114, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304120.

Suzuki:2013:SCM

[SA13] Reiji Suzuki and Takaya Arita. A simple computational model of the evolution of a communicative trait and its phenotypic plasticity. Journal of Theoretical Biology, 330(??):37–44, August 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001598.

Suzuki:2014:EDR

[SA14] Reiji Suzuki and Takaya Arita. Emergence of a dynamic resource partitioning based on the coevolution of phenotypic plasticity in sympatric species. Journal of Theoretical Biology, 352(??):51–59, July 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001106.

Sahmani:2017:SDA

[SA17] S. Sahmani and M. M. Aghdam. Size-dependent axial instability of microtubules surrounded by cytoplasm of a living cell based on nonlocal strain gradient elasticity theory. Journal of Theoretical Biology, 422(??):59–71, June 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301686.

Saadatpour:2010:AAA

[SAA10] Assieh Saadatpour, István Albert, and Réka Albert. Attractor analysis of asynchronous Boolean models of signal transduction networks. Journal of Theoretical Biology, 266(4):641–656, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003796.

Siegmund:2010:CDM

[SAB10] Thomas Siegmund, Matthew R. Allen, and David B. Burr. Can deterministic mechanical size effects contribute to fracture and microdamage accumulation in trabecular bone? Journal of Theoretical Biology, 265(2):202–210, July 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000192X.

Sigdel:2017:CBI

[SAB17] Ram P. Sigdel, Madhur Anand, and Chris T. Bauch. Competition between injunctive social norms and conservation priorities gives rise to complex dynamics in a model of forest growth and opinion dynamics. Journal of Theoretical Biology, 432(??):132–140, November 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303600.

Schmidt:2019:RDM

[SAB^{+}19] Barbara Schmidt, Dalia Alansary, Ivan Bogeski, Barbara A. Niemeyer, and Heiko Rieger. Reaction-diffusion model for STIM-ORAI interaction: The role of ROS and mutations. Journal of Theoretical Biology, 470(??):64–75, June 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300736.

Sircar:2015:DEO

[SABB15] S. Sircar, E. Aisenbrey, S. J. Bryant, and D. M. Bortz. Determining equilibrium osmolarity in poly(ethylene glycol)/ chondrotin sulfate gels mimicking articular cartilage. Journal of Theoretical Biology, 364(??):397–406, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005840.

Sachs:2013:WEB

[Sac13] Gottfried Sachs. Wind effects on bounding flight. Journal of Theoretical Biology, 316(??):35–41, January 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004614.

Sachs:2015:NMF

[Sac15] Gottfried Sachs. New model of flap-gliding flight. Journal of Theoretical Biology, 377(??):110–116, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001319.

Spyrou:2017:HMV

[SAD17] L. A. Spyrou, M. Agoras, and K. Danas. A homogenization model of the Voigt type for skeletal muscle. Journal of Theoretical Biology, 414(??):50–61, February 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303836.

Safarzynska:2013:CCE

[Saf13] Karolina Safarzynska. The coevolution of culture and environment. Journal of Theoretical Biology, 322(??):46–57, April 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000234.

Saunders:2019:TRT

[SAG19] Amanda M. Saunders, Daniel Ashlock, and Steffen P. Graether. Testing the rogue taxa hypothesis for clustering instability. Journal of Theoretical Biology, 472(??):36–45, July 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301432.

Serrano-Alcalde:2017:RNM

[SAGAGB17] Francisco Serrano-Alcalde, José Manuel García-Aznar, and María José Gómez-Benito. The role of nuclear mechanics in cell deformation under creeping flows. Journal of Theoretical Biology, 432(??):25–32, November 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303594.

Steckmann:2012:KPS

[SAGC12] Timothy Steckmann, Zubair Awan, Bernard S. Gerstman, and Prem P. Chapagain. Kinetics of peptide secondary structure conversion during amyloid β-protein fibrillogenesis. Journal of Theoretical Biology, 301(??):95–102, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000859.

Strilka:2014:QAS

[SAI14] Richard J. Strilka, Scott B. Armen, and Matthew C. Indeck. Qualitative analysis of subcutaneous Lispro and regular insulin injections for stress hyperglycemia: a pilot numerical study. Journal of Theoretical Biology, 356(??):192–200, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002434.

Sakai:2010:WGC

[Sak10] Satoki Sakai. With whom is the gene in conflict in offspring production?: Synthesis of the theories of intragenomic and parent-offspring conflict. Journal of Theoretical Biology, 266(3):367–373, October 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003401.

Stamatelos:2013:SMI

[SAKG13] Spyros K. Stamatelos, Ioannis P. Androulakis, Ah-Ng Tony Kong, and Panos G. Georgopoulos. A semi-mechanistic integrated toxicokinetic-toxicodynamic (TK/TD) model for arsenic(III) in hepatocytes. Journal of Theoretical Biology, 317(??):244–256, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004973.

Salomonsson:2010:GSQ

[Sal10] Marcus Salomonsson. Group selection: The quest for social preferences. Journal of Theoretical Biology, 264(3):737–746, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001104.

Salgado:2015:EPC

[Sal15] Mauricio Salgado. The evolution of paternal care can lead to population growth in artificial societies. Journal of Theoretical Biology, 380(??):192–202, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002696.

Sofonea:2017:EDM

[SAM17] Mircea T. Sofonea, Samuel Alizon, and Yannis Michalakis. Exposing the diversity of multiple infection patterns. Journal of Theoretical Biology, 419(??):278–289, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300668.

Sarvestani:2010:CBA

[Sar10] Alireza S. Sarvestani. Compliance of bio-adhesive substrates controls the kinetics of membrane-substrate association. Journal of Theoretical Biology, 266(4):516–521, October 21, 2010.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003784.

Satake:2010:DPL

[Sat10] Akiko Satake. Diversity of plant life cycles is generated by dynamic epigenetic regulation in response to vernalization. Journal of Theoretical Biology, 266(4):595–605, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003760.

Sato:2014:MDC

[Sat14] Kohji Sato. Minimum dynamic core in consciousness. Journal of Theoretical Biology, 343(??):208–210, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004803.

Sato:2017:RTB

[Sat17] Naoki Sato. Revisiting the theoretical basis of the endosymbiotic origin of plastids in the original context of Lynn Margulis on the origin of mitosing, eukaryotic cells. Journal of Theoretical Biology, 434(??):104–113, December 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304022.

Sato:2018:WDH

[Sat18] Kohji Sato. Why does hyperglycinemia exhibit so grave brain anomalies and so severe neurological symptoms? Journal of Theoretical Biology, 455(??):97–100, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303345.

Sato:2019:WGC

[Sat19] Kohji Sato. Why is glycine cleavage system segmentally expressed in radial glia? Journal of Theoretical Biology, 470(??):17–19, June 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301080.

Sun:2014:EPP

[SAZ^{+}14] Jin-Tu Sun, Bin Ao, Sheng Zhang, Zhitong Bing, and Lei Yang. Evolving protein–protein interaction networks: a model based on duplication and mutation at different rates. Journal of Theoretical Biology, 350(??):32–36, June 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000538.

Sagitov:2012:ICN

[SB12] Serik Sagitov and Krzysztof Bartoszek. Interspecies correlation for neutrally evolving traits. Journal of Theoretical Biology, 309(??):11–19, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002858.

Szarka:2015:SAT

[SB15] András Szarka and Tibor Balogh. In silico aided thoughts on mitochondrial vitamin C transport. Journal of Theoretical Biology, 365(??):181–189, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006043.

Su:2017:IRQ

[SB17] Min Su and Mike Boots. The impact of resource quality on the evolution of virulence in spatially heterogeneous environments. Journal of Theoretical Biology, 416(??):1–7, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316304271.

Stopka:2019:SMA

[SB19] Anna Stopka and Marcelo Boareto. A stochastic model of adult neurogenesis coupling cell cycle progression and differentiation. Journal of Theoretical Biology, 475(??):60–72, August 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302115.

Smolen:2014:SSP

[SBB14] Paul Smolen, Douglas A. Baxter, and John H. Byrne. Simulations suggest pharmacological methods for rescuing long-term potentiation. Journal of Theoretical Biology, 360(??):243–250, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003993.

Smolen:2018:PLM

[SBB18] Paul Smolen, Douglas A. Baxter, and John H. Byrne. Paradoxical LTP maintenance with inhibition of protein synthesis and the proteasome suggests a novel protein synthesis requirement for early LTP reversal. Journal of Theoretical Biology, 457(??):79–87, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304053.

Sergeev:2010:IEV

[SBCR10] R. A. Sergeev, R. E. Batorsky, J. M. Coffin, and I. M. Rouzine. Interpreting the effect of vaccination on steady state infection in animals challenged with simian immunodeficiency virus. Journal of Theoretical Biology, 263(3):385–392, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309005852.

Schindler:2013:PFM

[SBJ13] Susanne Schindler, Olaf Breidbach, and Jürgen Jost. Preferring the fittest mates: An analytically tractable model. Journal of Theoretical Biology, 317(??):30–38, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004961.

Stritzler:2018:UIE

[SBJ^{+}18] Margarita Stritzler, Carolina Berini, Cintia Jozefkowicz, Gabriela Soto, and Nicolás Ayub. Understanding the intracellular-to-extracellular localization switch of polyhydroxybutyrate polymerase in pseudomonas backgrounds as a microevolutionary process. Journal of Theoretical Biology, 456(??):29–33, November 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303680.

Sherborne:2016:CPM

[SBK16] N. Sherborne, K. B. Blyuss, and I. Z. Kiss. Compact pairwise models for epidemics with multiple infectious stages on degree heterogeneous and clustered networks. Journal of Theoretical Biology, 407(??):387–400, October 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301989.

Suweis:2012:SPT

[SBM^{+}12] S. Suweis, E. Bertuzzo, L. Mari, I. Rodriguez-Iturbe, A. Maritan, and A. Rinaldo. On species persistence-time distributions. Journal of Theoretical Biology, 303(??):15–24, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000951.

Sameen:2016:MMD

[SBM^{+}16] Sheema Sameen, Roberto Barbuti, Paolo Milazzo, Antonio Cerone, Marzia Del Re, and Romano Danesi. Mathematical modeling of drug resistance due to KRAS mutation in colorectal cancer. Journal of Theoretical Biology, 389(??):263–273, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005184.

Sharp:2019:OCA

[SBM^{+}19] Jesse A. Sharp, Alexander P. Browning, Tarunendu Mapder, Kevin Burrage, and Matthew J. Simpson. Optimal control of acute myeloid leukaemia. Journal of Theoretical Biology, 470(??):30–42, June 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301079.

Sheikh-Bahaei:2010:CER

[SBMH10] Shahab Sheikh-Bahaei, Jacquelyn J. Maher, and C. Anthony Hunt. Computational experiments reveal plausible mechanisms for changing patterns of hepatic zonation of xenobiotic clearance and hepatotoxicity. Journal of Theoretical Biology, 265(4):718–733, August 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000295X.

Scianna:2013:RMM

[SBP13] M. Scianna, C. G. Bell, and L. Preziosi. A review of mathematical models for the formation of vascular networks. Journal of Theoretical Biology, 333(??):174–209, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002117.

Sergeev:2010:MTT

[SBR10] R. A. Sergeev, R. E. Batorsky, and I. M. Rouzine. Model with two types of CTL regulation and experiments on CTL dynamics. Journal of Theoretical Biology, 263(3):369–384, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005293.

Shafique:2016:SIP

[SBR16] Shagufta Shafique, Nousheen Bibi, and Sajid Rashid. In silico identification of putative bifunctional Plk1 inhibitors by integrative virtual screening and structural dynamics approach. Journal of Theoretical Biology, 388(??):72–84, January 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004993.

Szilagyi:2017:ASC

[SBS17] András Szilágyi, Gergely Boza, and István Scheuring. Analysis of stability to cheaters in models of antibiotic degrading microbial communities. Journal of Theoretical Biology, 423(??):53– 62, June 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301959.

Sjodin:2014:PLC

[SBSE14] Henrik Sjödin, Åke Brännström, Mårten Söderquist, and Göran Englund. Population-level consequences of heterospecific density-dependent movements in predator–prey systems.

Journal of Theoretical Biology, 342(??):93–106, February 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313004475.

Schokker:2013:MMR

[SBSR13] D. Schokker, A. Bannink, M. A. Smits, and J. M. J. Rebel. A mathematical model representing cellular immune development and response to Salmonella of chicken intestinal tissue. Journal of Theoretical Biology, 330(??):75–87, August 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313001586.

Sharma:2018:MPC

[SBT^{+}18] Ronesh Sharma, Maitsetseg Bayarjargal, Tatsuhiko Tsunoda, Ashwini Patil, and Alok Sharma. MoRFPred-plus: Computational identification of MoRFs in protein sequences using physicochemical properties and HMM profiles. Journal of Theoretical Biology, 437(??):9–16, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304782.

Schiffner:2011:MAN

[SBW11] Ingo Schiffner, Johann Baumeister, and Roswitha Wiltschko. Mathematical analysis of the navigational process in homing pigeons. Journal of Theoretical Biology, 291(??):42–46, December 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004656.

Salazar-Ciudad:2010:TLM

[SC10a] Isaac Salazar-Ciudad. A two level mutation-selection model of cultural evolution and diversity. Journal of Theoretical Biology, 267(2):171–185, November 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004303.

Shen:2010:GMT

[SC10b] Hong-Bin Shen and Kuo-Chen Chou. Gneg-mPLoc: a topdown strategy to enhance the quality of predicting subcellular localization of Gram-negative bacterial proteins. Journal of Theoretical Biology, 264(2):326–333, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000305.

Sumner:2010:PEP

[SC10c] J. G. Sumner and M. A. Charleston. Phylogenetic estimation with partial likelihood tensors. Journal of Theoretical Biology, 262(3):413–424, February 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004755.

Sharma:2011:STP

[SC11a] Ajeet K. Sharma and Debashish Chowdhury. Stochastic theory of protein synthesis and polysome: Ribosome profile on a single mRNA transcript. Journal of Theoretical Biology, 289(??):36–46, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004267.

Sibani:2011:NSA

[SC11b] Paolo Sibani and Simon Christiansen. Non-stationary aging dynamics in ant societies. Journal of Theoretical Biology, 282(1):36–40, August 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311002347.

Sibly:2012:EDP

[SC12] Richard M. Sibly and Robert N. Curnow. Evolution of discrimination in populations at equilibrium between selfishness and altruism. Journal of Theoretical Biology, 313(??):162–171, November 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004183.

Straube:2013:ERE

[SC13a] Ronny Straube and Carsten Conradi. Erratum to “Reciprocal enzyme regulation as a source of bistability in covalent modification cycles” [J. Theor. Biol. 330 (2013) 56– 74]. Journal of Theoretical Biology, 333(??):210, September 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002592. See corri-

gendum [SC13b].

Straube:2013:RER

[SC13b] Ronny Straube and Carsten Conradi. Reciprocal enzyme regulation as a source of bistability in covalent modification cycles. Journal of Theoretical Biology, 330(??):56–74, August 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001550. See erra-

tum [SC13a].

Suzuki:2016:DER

[SC16] Takanori Mizuno Suzuki and Satoshi Chiba. Dynamics of evolutionary radiation under ecological neutrality. Journal of Theoretical Biology, 406(??):1–7, October 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301400.

Scheff:2013:PCT

[SCA13] Jeremy D. Scheff, Steve E. Calvano, and Ioannis P. Androulakis. Predicting critical transitions in a model of systemic inflammation. Journal of Theoretical Biology, 338(??):9–15, December 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003858.

Sanchez-Corrales:2010:ATF

[SCÁBM10] Yara-Elena Sánchez-Corrales, Elena R. Álvarez-Buylla, and Luis Mendoza. The Arabidopsis thaliana flower organ specification gene regulatory network determines a robust differentiation process. Journal of Theoretical Biology, 264(3):971–983, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001293. See corri-

gendum [SCÁBM11].

Sanchez-Corrales:2011:CAT

[SCÁBM11] Yara-Elena Sánchez-Corrales, Elena R. Álvarez-Buylla, and Luis Mendoza. Corrigendum to “The Arabidopsis thaliana flower organ specification gene regulatory network determines a robust differentiation process” [J. Theor. Biol. 264 (2010) 971–983]. Journal of Theoretical Biology, 274(1):185, April 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000397. See [SCÁBM10].

Sergeev:2011:MUP

[SCC11] Rinat Sergeev, Caroline Colijn, and Ted Cohen. Models to understand the population-level impact of mixed strain M. tuberculosis infections. Journal of Theoretical Biology, 280(1):88– 100, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311002013.

Shi:2012:DMM

[SCF^{+}12] Junfen Shi, Neil Curtis, Laura C. Fitton, Paul O’Higgins, and Michael J. Fagan. Developing a musculoskeletal model of the primate skull: Predicting muscle activations, bite force, and joint reaction forces using multibody dynamics analysis and advanced optimisation methods. Journal of Theoretical Biology, 310(??):21–30, October 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002834.

Shanafelt:2018:SDB

[SCF^{+}18] David W. Shanafelt, Jean Clobert, Eli P. Fenichel, Michael E. Hochberg, Ann Kinzig, Michel Loreau, Pablo A. Marquet, and Charles Perrings. Species dispersal and biodiversity in human-dominated metacommunities. Journal of Theoretical Biology, 457(??):199–210, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304296.

Scheuring:2010:EMP

[Sch10] István Scheuring. Egalitarian motive in punishing defectors. Journal of Theoretical Biology, 264(4):1293–1295, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001220.

Schrago:2014:LDE

[Sch14] Carlos G. Schrago. The limiting distribution of the effective population size of the ancestor of humans and chimpanzees. Journal of Theoretical Biology, 357(??):55–61, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400280X.

Schulman:2017:BRA

[Sch17] Lawrence S. Schulman. Bacterial resistance to antibodies: a model evolutionary study. Journal of Theoretical Biology, 417(??):61–67, March 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S002251931730022X. See corrigendum [Sch23].

Schlomann:2018:SMD

[Sch18] Brandon H. Schlomann. Stationary moments, diffusion limits, and extinction times for logistic growth with random catastrophes. Journal of Theoretical Biology, 454(??):154–163, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302996. See corri-

gendum [Sch21].

Schepens:2019:RRA

[SCH^{+}19a] Diana Schepens, Ross P. Carlson, Jeff Heys, Ashley E. Beck, and Tomás Gedeon. Role of resource allocation and transport in emergence of cross-feeding in microbial consortia. Journal of Theoretical Biology, 467(??):150–163, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300475.

Schinazi:2019:CEP

[Sch19b] Rinaldo B. Schinazi. Can evolution paths be explained by chance alone? Journal of Theoretical Biology, 467(??):63–65, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300591.

Schumacher:2019:NCM

[Sch19c] Linus J. Schumacher. Neural crest migration with continuous cell states. Journal of Theoretical Biology, 481(??): 84–90, November 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519319300463.

Schlomann:2021:CSM

[Sch21] Brandon H. Schlomann. Corrigendum to “Stationary moments, diffusion limits, and extinction times for logistic growth with random catastrophes” [J. Theor. Biol. 454 (2018) 154–163]. Journal of Theoretical Biology, 514(??):Article 110595, April 7, 2021. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519321000175. See [Sch18].

Schulman:2023:CBR

[Sch23] Lawrence S. Schulman. Corrigendum to “Bacterial resistance to antibodies: a model evolutionary study” [J. Theor. Biol. 417 (2017) 61–67]. Journal of Theoretical Biology, 565(??):??, May 21, 2023. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519323000607. See [Sch17].

Song:2015:ILC

[SCKL15] Renhua Song, Daniel R. Catchpoole, Paul J. Kennedy, and Jinyan Li. Identification of lung cancer miRNA-miRNA coregulation networks through a progressive data refining approach. Journal of Theoretical Biology, 380(??):271–279, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931500260X.

Scheff:2010:MIC

[SCLA10] Jeremy D. Scheff, Steve E. Calvano, Stephen F. Lowry, and Ioannis P. Androulakis. Modeling the influence of circadian rhythms on the acute inflammatory response. Journal of Theoretical Biology, 264(3):1068–1076, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001591.

Stan:2013:ISL

[SCLC13] Cristina Stan, Monica Teodora Cristescu, Buimaga Iarinca Luiza, and C. P. Cristescu. Investigation on series of length of coding and non-coding DNA sequences of bacteria using multifractal detrended cross-correlation analysis. Journal of Theoretical Biology, 321(??):54–62, March 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300009X.

Stan:2010:SAD

[SCS10] Cristina Stan, Constantin P. Cristescu, and Eugen I. Scarlat. Similarity analysis for DNA sequences based on chaos game representation. Case study: The albumin. Journal of Theoretical Biology, 267(4):513–518, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005011.

Sun:2014:RLP

[SCS14] Mengyang Sun, Xianrui Cheng, and Joshua E. S. Socolar. Regulatory logic and pattern formation in the early sea urchin embryo. Journal of Theoretical Biology, 363(??):80–92, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004354.

Sharp:2015:SMF

[SCS15] Katie Sharp, Edmund Crampin, and James Sneyd. A spatial model of fluid recycling in the airways of the lung. Journal of Theoretical Biology, 382(??):198–215, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931500329X.

Shan:2018:MPT

[SCsMzX18] Guo Shan, Hong Chuan-shan, Sun Ming-zhu, and Zhao Xin. Meshwork pattern transformed from branching pattern in spherical shell domain. Journal of Theoretical Biology, 455(??):293–302, October 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303667.

Swan:2012:CMT

[SCT12] Annalisa J. Swan, Alys R. Clark, and Merryn H. Tawhai. A computational model of the topographic distribution of ventilation in healthy human lungs. Journal of Theoretical Biology, 300(??):222–231, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000628.

Sindi:2016:CDP

[SD16] Suzanne S. Sindi and Rick Dale. Culturomics as a data playground for tests of selection: Mathematical approaches to detecting selection in word use. Journal of Theoretical Biology, 405(??):140–149, September 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000138.

Saeki:2015:OCC

[SDD15] Koichi Saeki, Hilje M. Doekes, and Rob J. De Boer. Optimal T cell cross-reactivity and the role of regulatory T cells. Journal of Theoretical Biology, 375(??):4–12, June 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400650X.

Shanafelt:2015:BPS

[SDJ^{+}15] David W. Shanafelt, Ulf Dieckmann, Matthias Jonas, Oskar Franklin, Michel Loreau, and Charles Perrings. Biodiversity, productivity, and the spatial insurance hypothesis revisited. Journal of Theoretical Biology, 380(??):426–435, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002969.

Spagnuolo:2011:MVC

[SDK11] Anna Maria Spagnuolo, Victor DiRita, and Denise Kirschner. A model for Vibrio cholerae colonization of the human intestine. Journal of Theoretical Biology, 289(??):247–258, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004310.

Sopakayang:2012:EVP

[SDKF12] Ratchada Sopakayang, Raffaella De Vita, Albert Kwansa, and Joseph W. Freeman. Elastic and viscoelastic properties of a type I collagen fiber. Journal of Theoretical Biology, 293(??):197–205, January 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005339.

Sueur:2011:GSG

[SDPC11] Cédric Sueur, Jean-Louis Deneubourg, Odile Petit, and Iain D. Couzin. Group size, grooming and fission in primates: a modeling approach based on group structure. Journal of Theoretical Biology, 273(1):156–166, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006909.

Shamloo-Dashtpagerdi:2015:NPC

[SDRA^{+}15] Roohollah Shamloo-Dashtpagerdi, Hooman Razi, Massumeh Aliakbari, Angelica Lindlöf, Mahdi Ebrahimi, and Esmaeil Ebrahimie. A novel pairwise comparison method for in silico discovery of statistically significant cis-regulatory elements in eukaryotic promoter regions: Application to Arabidopsis. Journal of Theoretical Biology, 364(??):364–376, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005852.

Sasmal:2017:MMC

[SDT17] Sourav Kumar Sasmal, Yueping Dong, and Yasuhiro Takeuchi. Mathematical modeling on T-cell mediated adaptive immunity in primary dengue infections. Journal of Theoretical Biology, 429(??):229–240, September 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303211.

Sardanyes:2010:ETR

[SE10] Josep Sardanyés and Santiago F. Elena. Error threshold in RNA quasispecies models with complementation. Journal of Theoretical Biology, 265(3):278–286, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002535.

Sands:2010:MII

[SEK^{+}10] Scott A. Sands, Bradley A. Edwards, Vanessa J. Kelly, Malcolm R. Davidson, Malcolm H. Wilkinson, and Philip J. Berger. A model investigation of the impact of ventilationperfusion mismatch on oxygenation during apnea in preterm infants. Journal of Theoretical Biology, 264(3):657–662, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001761.

Seki:2012:IIC

[Sek12] Motohide Seki. Intra-individual conflicts between autosomal and X-linked altruistic genes: Evolutionary perspectives of sex-specific grandmothering. Journal of Theoretical Biology, 304(??):273–285, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001646.

Seligmann:2010:PCB

[Sel10] Hervé Seligmann. Positive correlations between molecular and morphological rates of evolution. Journal of Theoretical Biology, 264(3):799–807, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001426.

Seligmann:2011:PMA

[Sel11] Hervé Seligmann. Pathogenic mutations in antisense mitochondrial tRNAs. Journal of Theoretical Biology, 269 (1):287–296, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519310005928.

Seligmann:2012:OGCb

[Sel12a] Hervé Seligmann. Overlapping genes coded in the 3’-to-5’- direction in mitochondrial genes and 3’-to-5’ polymerization of non-complementary RNA by an ‘invertase’. Journal of Theoretical Biology, 315(??):38–52, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004663.

Seligmann:2012:OGCa

[Sel12b] Hervé Seligmann. An overlapping genetic code for frameshifted overlapping genes in Drosophila mitochondria: Antisense antitermination tRNAs UAR insert serine. Journal of Theoretical Biology, 298(??):51–76, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006503.

Seligmann:2013:SAN

[Sel13] Hervé Seligmann. Systematic asymmetric nucleotide exchanges produce human mitochondrial RNAs cryptically encoding for overlapping protein coding genes. Journal of Theoretical Biology, 324(??):1–20, May 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000490.

Seligmann:2014:PAM

[Sel14a] Hervé Seligmann. Putative anticodons in mitochondrial tRNA sidearm loops: Pocketknife tRNAs? Journal of Theoretical Biology, 340(??):155–163, January 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004049.

Seligmann:2014:SRD

[Sel14b] Hervé Seligmann. Species radiation by DNA replication that systematically exchanges nucleotides? Journal of Theoretical Biology, 363(??):216–222, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004974.

Seligmann:2015:CES

[Sel15a] Hervé Seligmann. Codon expansion and systematic transcriptional deletions produce tetra-, pentacoded mitochondrial peptides. Journal of Theoretical Biology, 387(??):154–165, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004804.

Seligmann:2015:SEB

[Sel15b] Hervé Seligmann. Systematic exchanges between nucleotides: Genomic swinger repeats and swinger transcription in human mitochondria. Journal of Theoretical Biology, 384(??):70–77, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003975.

Seligmann:2016:SRS

[Sel16] Hervé Seligmann. Swinger RNA self-hybridization and mitochondrial non-canonical swinger transcription, transcription systematically exchanging nucleotides. Journal of Theoretical Biology, 399(??):84–91, June 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300273.

Seligmann:2017:NMP

[Sel17] Hervé Seligmann. Natural mitochondrial proteolysis confirms transcription systematically exchanging/deleting nucleotides, peptides coded by expanded codons. Journal of Theoretical Biology, 414(??):76–90, February 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303861.

Serdoz:2017:MLE

[SENS^{+}17] Stuart Serdoz, Attila Egri-Nagy, Jeremy Sumner, Barbara R. Holland, Peter D. Jarvis, Mark M. Tanaka, and Andrew R. Francis. Maximum likelihood estimates of pairwise rearrangement distances. Journal of Theoretical Biology, 423(??):31–40, June 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301790.

Szomolay:2012:MIB

[SER^{+}12] Barbara Szomolay, Tim D. Eubank, Ryan D. Roberts, Clay B. Marsh, and Avner Friedman. Modeling the inhibition of breast cancer growth by GM-CSF. Journal of Theoretical Biology, 303(??):141–151, June 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001488.

Serafino:2016:ATC

[Ser16] Loris Serafino. Abiogenesis as a theoretical challenge: Some reflections. Journal of Theoretical Biology, 402(??):18–20, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300674.

Straka:2015:TES

[SF15] P. Straka and S. Fedotov. Transport equations for subdiffusion with nonlinear particle interaction. Journal of Theoretical Biology, 366(??):71–83, February 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006559.

Simon:2012:HRM

[SFD12] Burton Simon, Jeffrey A. Fletcher, and Michael Doebeli. Hamilton’s rule in multi-level selection models. Journal of Theoretical Biology, 299(??):55–63, April 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100364X.

Sedaghat:2016:IFC

[SFMS16] Nafiseh Sedaghat, Mahmood Fathy, Mohammad Hossein Modarressi, and Ali Shojaie. Identifying functional cancerspecific miRNA-mRNA interactions in testicular germ cell tumor. Journal of Theoretical Biology, 404(??):82–96, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301151.

Shipman:2013:TCP

[SFS13] Patrick D. Shipman, Sérgio H. Faria, and Christopher Strickland. Towards a continuous population model for natural language vowel shift. Journal of Theoretical Biology, 332(??):123–135, September 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001859.

Sumner:2012:LMM

[SFSJ12] J. G. Sumner, J. Fernández-Sánchez, and P. D. Jarvis. Lie Markov models. Journal of Theoretical Biology, 298(??):16– 31, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311006345.

Sekimura:2014:MPD

[SFT14] Toshio Sekimura, Yuta Fujihashi, and Yasuhiro Takeuchi. A model for population dynamics of the mimetic butterfly Papilio polytes in the Sakishima Islands, Japan. Journal of Theoretical Biology, 361(??):133–140, November 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314003725.

Santoni:2016:NVR

[SFV16] Daniele Santoni, Giovanni Felici, and Davide Vergni. Natural vs. random protein sequences: Discovering combinatorics properties on amino acid words. Journal of Theoretical Biology, 391(??):13–20, February 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005731.

Singh:2015:CSS

[SG15a] Surya Pratap Singh and Dwijendra K. Gupta. A comparative study of structural and conformational properties of casein kinase-1 isoforms: Insights from molecular dynamics and principal component analysis. Journal of Theoretical Biology, 371(??):59–68, April 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000491.

Song:2015:DSL

[SG15b] Tao Song and Hong Gu. Discovering short linear protein motif based on selective training of profile hidden Markov models. Journal of Theoretical Biology, 377(??):75–84, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315001137.

Subramanian:2018:EAD

[SG18] Hemachander Subramanian and Robert A. Gatenby. Evolutionary advantage of directional symmetry breaking in selfreplicating polymers. Journal of Theoretical Biology, 446 (??):128–136, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519318301267.

Shim:2012:GDM

[SGA^{+}12] Eunha Shim, John J. Grefenstette, Steven M. Albert, Brigid E. Cakouros, and Donald S. Burke. A game dynamic model for vaccine skeptics and vaccine believers: Measles as an example. Journal of Theoretical Biology, 295(??):194–203, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005674.

Sandoval:2012:MMP

[SGAM12] Rosy Paola Cárdenas Sandoval, Diego Alexander Garzón-Alvarado, and Angélica Maria Ramírez Martínez. A mathematical model of the process of ligament repair: Effect of cold therapy and mechanical stress. Journal of Theoretical Biology, 302(??):53–61, June 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000550.

Siebert:2012:GMD

[SGB12] Tobias Siebert, Michael Günther, and Reinhard Blickhan. A 3d-geometric model for the deformation of a transversally loaded muscle. Journal of Theoretical Biology, 298(??):116– 121, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000100.

Santana:2019:QBD

[SGB19] Leonardo M. Santana, Shridar Ganesan, and Gyan Bhanot. A quasi birth-and-death model for tumor recurrence. Journal of Theoretical Biology, 480(??):175–191, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303029.

Schmitt:2014:EIC

[SGCD14] Christoph K. Schmitt, Christian Guill, Eddy Carmack, and Barbara Drossel. Effect of introducing a competitor on cyclic dominance of sockeye salmon. Journal of Theoretical Biology, 360(??):13–20, November 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003646.

Schmitt:2012:RCD

[SGD12] Christoph K. Schmitt, Christian Guill, and Barbara Drossel. The robustness of cyclic dominance under random fluctuations. Journal of Theoretical Biology, 308(??):79–87, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002706.

Srivastava:2016:SAU

[SGD^{+}16] Ambuj Srivastava, Prerana Gogoi, Bhagyashree Deka, Shrayanti Goswami, and Shankar Prasad Kanaujia. In silico analysis of 5’-UTRs highlights the prevalence of Shine– Dalgarno and leaderless-dependent mechanisms of translation initiation in bacteria and archaea, respectively. Journal of Theoretical Biology, 402(??):54–61, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316300777.

Song:2012:TFS

[SGDL12] Liyan Song, Yunbo Guo, Qingkai Deng, and Jinming Li. TH17 functional study in severe asthma using agent based model. Journal of Theoretical Biology, 309(??):29–33, September 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002469.

Sanchez-Gomez:2019:PSR

[SGGHS19] Laura Sánchez-Gómez, Agustín Guerrero-Hernández, and Moisés Santillán. Polymerization of sarcoplasmic-reticulum calcium-binding proteins might explain observed reticulum kinetics-on-demand behavior. Journal of Theoretical Biology, 482(??):Article 109986, December 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303297.

Schofield:2011:TAG

[SGGM11] J. W. Schofield, E. A. Gaffney, R. A. Gatenby, and P. K. Maini. Tumour angiogenesis: The gap between theory and experiments. Journal of Theoretical Biology, 274(1):97–102, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000245.

Silva:2010:QTM

[SGGY10] Ariosto S. Silva, Robert A. Gatenby, Robert J. Gillies, and Jose Andres Yunes. A quantitative theoretical model for the development of malignancy in ductal carcinoma in situ. Journal of Theoretical Biology, 262(4):601–613, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309005207.

Sanchez-Garduno:2019:THP

[SGKCP19] Faustino Sánchez-Garduño, Andrew L. Krause, Jorge A. Castillo, and Pablo Padilla. Turing–Hopf patterns on growing domains: the torus and the sphere. Journal of Theoretical Biology, 481(??):136–150, November 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304648.

Scholz:2010:BMH

[SGL10] Markus Scholz, Arnd Gross, and Markus Loeffler. A biomathematical model of human thrombopoiesis under chemotherapy. Journal of Theoretical Biology, 264(2):287–300, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310000056.

Soula:2015:MAT

[SGS15] H. A. Soula, A. Géloën, and C. O. Soulage. Model of adipose tissue cellularity dynamics during food restriction. Journal of Theoretical Biology, 364(??):189–196, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314005335.

Shafiey:2017:ECT

[SGW17] Hassan Shafiey, Toni I. Gossmann, and David Waxman. Evolutionary control: Targeted change of allele frequencies in natural populations using externally directed evolution. Journal of Theoretical Biology, 419(??):362–374, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300231.

Stadler:2018:FBD

[SGW^{+}18] Tanja Stadler, Alexandra Gavryushkina, Rachel C. M. Warnock, Alexei J. Drummond, and Tracy A. Heath. The fossilized birth-death model for the analysis of stratigraphic range data under different speciation modes. Journal of Theoretical Biology, 447(??):41–55, June 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830119X.

Suryawanshi:2015:MSM

[SH15] Gajendra W. Suryawanshi and Alexander Hoffmann. A multiscale mathematical modeling framework to investigate antiviral therapeutic opportunities in targeting HIV-1 accessory proteins. Journal of Theoretical Biology, 386(??):89–104, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004373.

Staalhand:2016:LAS

[SH16] Jonas Stålhand and Gerhard A. Holzapfel. Length adaptation of smooth muscle contractile filaments in response to sustained activation. Journal of Theoretical Biology, 397(??):13– 21, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001247.

Steingass:2017:IBE

[SH17] Sheanna Steingass and Markus Horning. Individual-based energetic model suggests bottom up mechanisms for the impact of coastal hypoxia on Pacific harbor seal (Phoca vitulina richardii) foraging behavior. Journal of Theoretical Biology, 416(??):190–198, March 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730005X.

Sharma:2014:BAR

[Sha14] Abhay Sharma. Bioinformatic analysis revealing association of exosomal mRNAs and proteins in epigenetic inheritance. Journal of Theoretical Biology, 357(??):143–149, September 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002902.

Schmutz:2017:CAB

[SHB^{+}17] Christian F. Schmutz, Trey D. Hatton, Michael P. Blomfield, Benjamin D. Chandler, Trent C. Benedick, and Marc D. H. Hansen. A cellular automaton based on plasma membrane turnover accurately recapitulates cell mechanics during epithelial scattering. Journal of Theoretical Biology, 430(??):141–156, October 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302928.

Schoenmakers:2014:SHS

[SHBT14] Sarah Schoenmakers, Christian Hilbe, Bernd Blasius, and Arne Traulsen. Sanctions as honest signals — the evolution of pool punishment by public sanctioning institutions. Journal of Theoretical Biology, 356(??):36–46, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002392.

Shen:2010:BPR

[She10] Hui-Shen Shen. Buckling and postbuckling of radially loaded microtubules by nonlocal shear deformable shell model. Journal of Theoretical Biology, 264(2):386–394, May 21, 2010.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000810.

Shen:2011:NAL

[She11a] Hui-Shen Shen. Nonlinear analysis of lipid tubules by nonlocal beam model. Journal of Theoretical Biology, 276(1):50– 56, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000816.

Shen:2011:CSE

[She11b] Xiaojuan Shen. Conformation and sequence evidence for two-fold symmetry in left-handed beta-helix fold. Journal of Theoretical Biology, 285(1):77–83, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003109.

Segura:2016:ATH

[SHF16] Juan Segura, Frank M. Hilker, and Daniel Franco. Adaptive threshold harvesting and the suppression of transients. Journal of Theoretical Biology, 395(??):103–114, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000795.

Safarbali:2016:MMN

[SHG16] Bahareh Safarbali, Fatemeh Hadaeghi, and Shahriar Gharibzadeh. Mesoscopic model of neuronal system deficits in multiple sclerosis. Journal of Theoretical Biology, 411(??):6–15, December 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301965.

Shay:2015:OSC

[SHH15] Julie A. Shay, Christopher Huynh, and Paul G. Higgs. The origin and spread of a cooperative replicase in a prebiotic chemical system. Journal of Theoretical Biology, 364(??):249–259, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400558X.

Shibasaki:2019:EGI

[Shi19] Shota Shibasaki. The evolutionary game of interspecific mutualism in the multi-species model. Journal of Theoretical Biology, 471(??):51–58, June 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931930133X.

Shavlyugin:2014:DPC

[SHK14] Evgeny A. Shavlyugin, Leonid G. Hanin, and Mikhail A. Khanin. Dynamics of pathologic clot formation: a mathematical model. Journal of Theoretical Biology, 340(??):96–104, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003792.

Shi:2011:PTT

[SHLL11] Xiaobo Shi, Xiuzhen Hu, Shaobo Li, and Xingxing Liu. Prediction of β-turn types in protein by using composite vector. Journal of Theoretical Biology, 286(??):24–30, October 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311003456.

Singh:2018:FRB

[SHM^{+}18] Soibam Shyamchand Singh, Dineshchandra Haobijam, Md. Zubbair Malik, Romana Ishrat, and R. K. Brojen Singh. Fractal rules in brain networks: Signatures of self-organization. Journal of Theoretical Biology, 437(??):58–66, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304319.

Sajitz-Hermstein:2012:RCG

[SHN12] Max Sajitz-Hermstein and Zoran Nikoloski. Restricted cooperative games on metabolic networks reveal functionally important reactions. Journal of Theoretical Biology, 314(??):192–203, December 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312004328.

Steel:2013:MAN

[SHS13] Mike Steel, Wim Hordijk, and Joshua Smith. Minimal autocatalytic networks. Journal of Theoretical Biology, 332(??):96–107, September 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002002.

Sturrock:2015:MMP

[SHSR15] Marc Sturrock, Wenrui Hao, Judith Schwartzbaum, and Grzegorz A. Rempala. A mathematical model of prediagnostic glioma growth. Journal of Theoretical Biology, 380(??):299–308, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002829.

Shtylla:2017:MMS

[Sht17] Blerta Shtylla. Mathematical modeling of spatiotemporal protein localization patterns in C. crescentus bacteria: a mechanism for asymmetric FtsZ ring positioning. Journal of Theoretical Biology, 433(??):8–20, November 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303776.

Shutters:2013:TRF

[Shu13] Shade T. Shutters. Towards a rigorous framework for studying 2-player continuous games. Journal of Theoretical Biology, 321(??):40–43, March 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000088.

Sewalt:2016:IAE

[SHvHB16] Lotte Sewalt, Kristen Harley, Peter van Heijster, and Sanjeeva Balasuriya. Influences of Allee effects in the spreading of malignant tumours. Journal of Theoretical Biology, 394(??):77– 92, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931600031X.

Sun:2016:AAT

[SHW16] Shiqi Sun, Yan Huang, and Qining Wang. Adding adaptable toe stiffness affects energetic efficiency and dynamic behaviors of bipedal walking. Journal of Theoretical Biology, 388(??):108–118, January 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004890.

Saeki:2010:ONR

[SI10a] Koichi Saeki and Yoh Iwasa. Optimal number of regulatory T cells. Journal of Theoretical Biology, 263(2):210–218, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309005542.

Suzuki:2010:ACR

[SI10b] Kenta Suzuki and Takashi Ikegami. Allometry and catastrophic regime shifts in food chains. Journal of Theoretical Biology, 267(1):121–128, November 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004017.

Saeki:2011:CAS

[SI11a] Koichi Saeki and Yoh Iwasa. T cell anergy as a strategy to reduce the risk of autoimmunity. Journal of Theoretical Biology, 277(1):74–82, May 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001081.

Seto:2011:RSR

[SI11b] Mayumi Seto and Yoh Iwasa. Regime shift and robustness of organism-created environments: a model for microbial ecosystems. Journal of Theoretical Biology, 269(1):297–306, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005825.

Satake:2012:SMC

[SI12] Akiko Satake and Yoh Iwasa. A stochastic model of chromatin modification: Cell population coding of winter memory in plants. Journal of Theoretical Biology, 302(??):6–17, June 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000823.

Seto:2019:PDC

[SI19] Mayumi Seto and Yoh Iwasa. Population dynamics of chemotrophs in anaerobic conditions where the metabolic energy acquisition per redox reaction is limited. Journal of Theoretical Biology, 467(??):164–173, April 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300542.

Siero:2018:NGP

[Sie18] E. Siero. Nonlocal grazing in patterned ecosystems. Journal of Theoretical Biology, 436(??):64–71, January 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304563.

Sigmund:2012:MAI

[Sig12] Karl Sigmund. Moral assessment in indirect reciprocity. Journal of Theoretical Biology, 299(??):25–30, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001706.

Sunada:2015:SAP

[SIHO15] S. Sunada, A. Ide, Y. Hoshino, and M. Okamoto. A study of autorotating plant seeds. Journal of Theoretical Biology, 386(??):55–61, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004282.

Sabooh:2018:IMS

[SIK^{+}18] M. Fazli Sabooh, Nadeem Iqbal, Mukhtaj Khan, Muslim Khan, and H. F. Maqbool. Identifying 5-methylcytosine sites in RNA sequence using composite encoding feature into Chou’s PseKNC. Journal of Theoretical Biology, 452(??):1–9, September 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302224.

Silanikove:2016:TRM

[Sil16] Nissim Silanikove. Transcellular route as the most probable explanation for the presence of plasminogen in mammal’s milk. Journal of Theoretical Biology, 395(??):221–226, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000588.

Simon:2014:CTM

[Sim14] Burton Simon. Continuous-time models of group selection, and the dynamical insufficiency of kin selection models. Journal of Theoretical Biology, 349(??):22–31, May 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000472.

Sinclair:2015:NRN

[Sin15] Robert Sinclair. Necessary relations for nucleotide frequencies. Journal of Theoretical Biology, 374(??):179–182, June 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931500140X.

Sgardeli:2017:FED

[SIVH17] Vasiliki Sgardeli, Yoh Iwasa, Harry Varvoglis, and John M. Halley. A forecast for extinction debt in the presence of speciation. Journal of Theoretical Biology, 415(??):48–52, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303654.

Stump:2018:HLO

[SJK18] Simon Maccracken Stump, Evan Curtis Johnson, and Christopher A. Klausmeier. How leaking and overproducing resources affect the evolutionary robustness of cooperative crossfeeding. Journal of Theoretical Biology, 454(??):278–291, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303072.

Soula:2013:MAS

[SJSG13] H. A. Soula, H. Julienne, C. O. Soulage, and A. Géloën. Modelling adipocytes size distribution. Journal of Theoretical Biology, 332(??):89–95, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001938.

Stump:2018:HSS

[SJSK18] Simon Maccracken Stump, Evan Curtis Johnson, Zepeng Sun, and Christopher A. Klausmeier. How spatial structure and neighbor uncertainty promote mutualists and weaken black queen effects. Journal of Theoretical Biology, 446(??):33–60, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300894.

Shtylla:2010:MMM

[SK10] Blerta Shtylla and James P. Keener. A mechanomolecular model for the movement of chromosomes during mitosis driven by a minimal kinetochore bicyclic cascade. Journal of Theoretical Biology, 263(4):455–470, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005967.

Shiino:2011:TAF

[SK11a] Yuta Shiino and Osamu Kuwazuru. Theoretical approach to the functional optimisation of spiriferide brachiopod shell: Optimum morphology of sulcus. Journal of Theoretical Biology, 276(1):192–198, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000981.

Suzuki:2011:ODC

[SK11b] Shinsuke Suzuki and Hiromichi Kimura. Oscillatory dynamics in the coevolution of cooperation and mobility. Journal of Theoretical Biology, 287(??):42–47, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003699.

Shtylla:2012:MMP

[SK12a] Blerta Shtylla and James P. Keener. A mathematical model of ParA filament-mediated chromosome movement in caulobacter crescentus. Journal of Theoretical Biology, 307(??):82–95, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002299.

Sorace:2012:ANM

[SK12b] Ron Sorace and Natalia L. Komarova. Accumulation of neutral mutations in growing cell colonies with competition. Journal of Theoretical Biology, 314(??):84–94, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004298.

Saraee:2015:DSM

[SK15] Mahdieh B. Saraee and Moharam H. Korayem. Dynamic simulation and modeling of the motion modes produced during the 3D controlled manipulation of biological micro/nanoparticles based on the AFM. Journal of Theoretical Biology, 378(??):65–78, August 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001976.

Schmickl:2016:HRB

[SK16a] Thomas Schmickl and Istvan Karsai. How regulation based on a common stomach leads to economic optimization of honeybee foraging. Journal of Theoretical Biology, 389(??):274–286, January 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005378.

Stump:2016:CCB

[SK16b] Simon M. Stump and Christopher A. Klausmeier. Competition and coexistence between a syntrophic consortium and a metabolic generalist, and its effect on productivity. Journal of Theoretical Biology, 404(??):348–360, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301539.

Steel:2018:NRC

[SK18] Mike Steel and Stuart Kauffman. A note on random catalytic branching processes. Journal of Theoretical Biology, 437(??):222–224, January 21, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304885.

Saravanan:2019:ABC

[SK19] Kandasamy Saravanan and Poomani Kumaradhas. Acylguanidine-BACE1 complex: Insights of intermolecular interactions and dynamics. Journal of Theoretical Biology, 464(??):33–49, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318306167.

Sanches:2018:PAE

[SKAG18] Vitor H. Sanches, Dhyan V. H. Kuraoka, Pedro R. Almeida, and Carla Goldman. A phenomenological analysis of ecoevolutionary coupling under dilution. Journal of Theoretical Biology, 438(??):156–164, February 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730526X.

Stamatakos:2010:ADS

[SKD^{+}10] G. S. Stamatakos, E. A. Kolokotroni, D. D. Dionysiou, E. Ch. Georgiadi, and C. Desmedt. An advanced discrete statediscrete event multiscale simulation model of the response of a solid tumor to chemotherapy: Mimicking a clinical study. Journal of Theoretical Biology, 266(1):124–139, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310002547.

Sassaroli:2016:NEH

[SKF16] Angelo Sassaroli, Jana M. Kainerstorfer, and Sergio Fantini. Nonlinear extension of a hemodynamic linear model for coherent hemodynamics spectroscopy. Journal of Theoretical Biology, 389(??):132–145, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005354.

Sarkar:2019:ECS

[SKGM19] Anyesha Sarkar, Brian M. Kobylkevich, David M. Graham, and Mark A. Messerli. Electromigration of cell surface macromolecules in DC electric fields during cell polarization and galvanotaxis. Journal of Theoretical Biology, 478(??):58–73, October 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302504.

Staalhand:2011:MCM

[SKH11] J. Stålhand, A. Klarbring, and G. A. Holzapfel. A mechanochemical 3d continuum model for smooth muscle contraction under finite strains. Journal of Theoretical Biology, 268(1):120–130, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005369.

Schmidt:2017:NSS

[SKH17] Markus Schmidt, Arne Kutzner, and Klaus Heese. A novel specialized single-linkage clustering algorithm for taxonomically ordered data. Journal of Theoretical Biology, 427(??):1– 7, August 1, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302163.

Spasic:2011:SDM

[SKK^{+}11] Sladjana Spasic, Aleksandar Kalauzi, Srdjan Kesic, Milica Obradovic, and Jasna Saponjic. Surrogate data modeling the relationship between high frequency amplitudes and Higuchi fractal dimension of EEG signals in anesthetized rats. Journal of Theoretical Biology, 289(??):160–166, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004401.

Schraiber:2012:CUL

[SKK^{+}12] Joshua G. Schraiber, Angela N. Kaczmarczyk, Ricky Kwok, Miran Park, Rachel Silverstein, Florentine U. Rutaganira, Taruna Aggarwal, Michael A. Schwemmer, Carole L. Hom, Richard K. Grosberg, and Sebastian J. Schreiber. Constraints on the use of lifespan-shortening Wolbachia to control dengue fever. Journal of Theoretical Biology, 297(??):26–32, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311006151.

Srivastava:2018:BPC

[SKK18] Abhishikha Srivastava, Ravindra Kumar, and Manish Kumar. BlaPred: Predicting and classifying β-lactamase using a 3-tier prediction system via Chou’s general PseAAC. Journal of Theoretical Biology, 457(??):29–36, November 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304090.

Shin:2017:EOD

[SKPK17] Nina Shin, Taewoo Kwag, Sangwook Park, and Yon Hui Kim. Effects of operational decisions on the diffusion of epidemic disease: a system dynamics modeling of the MERS-CoV outbreak in south Korea. Journal of Theoretical Biology, 421(??):39–50, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931730139X.

Shtilerman:2015:ESC

[SKS15] Elad Shtilerman, David A. Kessler, and Nadav M. Shnerb. Emergence of structured communities through evolutionary dynamics. Journal of Theoretical Biology, 383(??):138–144, October 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003574.

Sun:2019:MCF

[SKS^{+}19] Zepeng Sun, Thomas Koffel, Simon M. Stump, Ghjuvan M. Grimaud, and Christopher A. Klausmeier. Microbial crossfeeding promotes multiple stable states and species coexistence, but also susceptibility to cheaters. Journal of Theoretical Biology, 465(??):63–77, March 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300104.

Shiino:2012:SCR

[SKSO12] Yuta Shiino, Osamu Kuwazuru, Yutaro Suzuki, and Satoshi Ono. Swimming capability of the remopleuridid trilobite Hypodicranotus striatus: Hydrodynamic functions of the exoskeleton and the long, forked hypostome. Journal of Theoretical Biology, 300(??):29–38, May 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000136.

Swierniak:2016:DPC

[ŚKSRW16] Andrzej Świerniak, Michał Krześlak, Sebastian Student, and Joanna Rzeszowska-Wolny. Development of a population of cancer cells: Observation and modeling by a mixed spatial evolutionary games approach. Journal of Theoretical Biology, 405(??):94–103, September 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301187.

Sapijanskas:2010:CEF

[SL10] Jurgis Sapijanskas and Michel Loreau. Cascading extinctions, functional complementarity, and selection in two-trophic-level model communities: a trait-based mechanistic approach. Journal of Theoretical Biology, 267(3):375–387, December 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310004546.

Su:2012:AAG

[SL12] Linlin Su and Roger Lui. Advance of advantageous genes for a multiple-allele population genetics model. Journal of Theoretical Biology, 315(??):1–8, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004730.

Steel:2017:ORR

[SL17] Mike Steel and Christoph Leuenberger. The optimal rate for resolving a near-polytomy in a phylogeny. Journal of Theoretical Biology, 420(??):174–179, May 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301121.

Syed:2019:SFG

[SL19] Christopher K. Syed and Robert H. Lee. Slow and fast grouping of cargo velocities in axonal transport due to single versus multi-motor transport. Journal of Theoretical Biology, 480(??):65–70, November 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305745.

Steinacher:2012:CMA

[SLC12] Arno Steinacher, Ottoline Leyser, and Richard H. Clayton. A computational model of auxin and pH dynamics in a single plant cell. Journal of Theoretical Biology, 296(??):84–94, March 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100590X.

Stephanou:2017:HTI

[SLD^{+}17] A. Stéphanou, A. C. Lesart, J. Deverchère, A. Juhem, A. Popov, and F. Estève. How tumour-induced vascular changes alter angiogenesis: Insights from a computational model. Journal of Theoretical Biology, 419(??):211–226, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300735.

Sharma:2013:FET

[SLDP13] Alok Sharma, James Lyons, Abdollah Dehzangi, and Kuldip K. Paliwal. A feature extraction technique using bi-gram probabilities of position specific scoring matrix for protein fold recognition. Journal of Theoretical Biology, 320(??):41–46, March 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312006327.

Steel:2013:IST

[SLHS13] Mike Steel, Simone Linz, Daniel H. Huson, and Michael J. Sanderson. Identifying a species tree subject to random lateral gene transfer. Journal of Theoretical Biology, 322(??):81–93, April 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313000283.

Sun:2010:IIR

[SLJ^{+}10] Gui-Quan Sun, Quan-Xing Liu, Zhen Jin, Amit Chakraborty, and Bai-Lian Li. Influence of infection rate and migration on extinction of disease in spatial epidemics. Journal of Theoretical Biology, 264(1):95–103, May 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000010X.

Smadi:2018:LRM

[SLL18] Charline Smadi, Hélène Leman, and Violaine Llaurens. Looking for the right mate in diploid species: How does genetic dominance affect the spatial differentiation of a sexual trait? Journal of Theoretical Biology, 447(??):154–170, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301462.

Shelton:2017:MCD

[SLM17] Deborah E. Shelton, Martin P. Leslie, and Richard E. Michod. Models of cell division initiation in chlamydomonas: a challenge to the consensus view. Journal of Theoretical Biology, 412(??):186–197, January 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303502.

Strelioff:2010:EDE

[SLO10] Christopher C. Strelioff, Richard E. Lenski, and Charles Ofria. Evolutionary dynamics, epistatic interactions, and biological information. Journal of Theoretical Biology, 266(4):584–594, October 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003826.

Sanz-Leon:2017:MCS

[SLR17] Paula Sanz-Leon and P. A. Robinson. Multistability in the corticothalamic system. Journal of Theoretical Biology, 432(??):141–156, November 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303478.

Sturrock:2017:INC

[SLS17] Marc Sturrock, Shiyu Li, and Vahid Shahrezaei. The influence of nuclear compartmentalisation on stochastic dynamics of self-repressing gene expression. Journal of Theoretical Biology, 424(??):55–72, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302059.

Song:2018:PIA

[SLT^{+}18] Jiangning Song, Fuyi Li, Kazuhiro Takemoto, Gholamreza Haffari, Tatsuya Akutsu, Kuo-Chen Chou, and Geoffrey I. Webb. PREvaIL, an integrative approach for inferring catalytic residues using sequence, structural, and network features in a machine-learning framework. Journal of Theoretical Biology, 443(??):125–137, April 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300390.

Stirk:2010:SCE

[SLvdBMP10] Emily R. Stirk, Grant Lythe, Hugo A. van den Berg, and Carmen Molina-París. Stochastic competitive exclusion in the maintenance of the naïve T cell repertoire. Journal of Theoretical Biology, 265(3):396–410, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000233X.

Song:2018:MWS

[SLW^{+}18a] Wei Song, Huaping Liu, Jiajia Wang, Yan Kong, Xia Yin, and Weidong Zang. MATHT: a web server for comprehensive transcriptome data analysis. Journal of Theoretical Biology, 455(??):140–146, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830345X.

Su:2018:ECI

[SLW18b] Qi Su, Aming Li, and Long Wang. Evolution of cooperation with interactive identity and diversity. Journal of Theoretical Biology, 442(??):149–157, April 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300377.

Santos:2010:SGC

[SM10a] José Santos and Ángel Monteagudo. Study of the genetic code adaptability by means of a genetic algorithm. Journal of Theoretical Biology, 264(3):854–865, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001165.

Szilagyi:2010:CFE

[SM10b] András Szilágyi and Géza Meszéna. Coexistence in a fluctuating environment by the effect of relative nonlinearity: a minimal model. Journal of Theoretical Biology, 267(4):502–512, December 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004947.

Stich:2011:MFE

[SM11a] Michael Stich and Susanna C. Manrubia. Motif frequency and evolutionary search times in RNA populations. Journal of Theoretical Biology, 280(1):117–126, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001500.

Subirana:2011:DAS

[SM11b] Juan A. Subirana and Xavier Messeguer. The distribution of alternating AT sequences in eukaryotic genomes suggests a role in homologous chromosome recognition in meiosis. Journal of Theoretical Biology, 283(1):28–34, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311002670.

Sakaguchi:2014:CAD

[SM14] Hidetsugu Sakaguchi and Satomi Maeyama. Competitive aggregation dynamics using phase wave signals. Journal of Theoretical Biology, 359(??):155–160, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003609.

Sherratt:2016:HDT

[SM16] Jonathan A. Sherratt and Julia J. Mackenzie. How does tidal flow affect pattern formation in mussel beds? Journal of Theoretical Biology, 406(??):83–92, October 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301618.

Sankari:2017:PMP

[SM17a] E. Siva Sankari and D. Manimegalai. Predicting membrane protein types using various decision tree classifiers based on various modes of general PseAAC for imbalanced datasets. Journal of Theoretical Biology, 435(??):208–217, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304368.

Sasaki:2017:PLS

[SM17b] Akira Sasaki and Akiko N. Mizuno. Partitioning light spectra: Adaptive stratification of phytobenthic communities in Antarctic lakes. Journal of Theoretical Biology, 424(??):1– 10, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301868.

Sankari:2018:PMP

[SM18] E. Siva Sankari and D. Manimegalai. Predicting membrane protein types by incorporating a novel feature set into Chou’s general PseAAC. Journal of Theoretical Biology, 455(??):319–328, October 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318303564.

Smith:2011:MMT

[SMA11] Amber M. Smith, Jonathan A. McCullers, and Frederick R. Adler. Mathematical model of a three-stage innate immune response to a pneumococcal lung infection. Journal of Theoretical Biology, 276(1):106–116, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000786.

Starnini:2013:ISE

[SMC^{+}13] Michele Starnini, Anna Machens, Ciro Cattuto, Alain Barrat, and Romualdo Pastor-Satorras. Immunization strategies for epidemic processes in time-varying contact networks. Journal of Theoretical Biology, 337(??):89–100, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003251.

Scornavacca:2017:FAR

[SMC17] Celine Scornavacca, Joan Carles Pons Mayol, and Gabriel Cardona. Fast algorithm for the reconciliation of gene trees and LGT networks. Journal of Theoretical Biology, 418(??):129– 137, April 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300243.

Schrempf:2016:RPA

[SMD^{+}16] Dominik Schrempf, Bui Quang Minh, Nicola De Maio, Arndt von Haeseler, and Carolin Kosiol. Reversible polymorphismaware phylogenetic models and their application to tree inference. Journal of Theoretical Biology, 407(??):362–370, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302259.

Schneider:2016:MDB

[SMdA16] David M. Schneider, Ayana B. Martins, and Marcus A. M. de Aguiar. The mutation-drift balance in spatially structured populations. Journal of Theoretical Biology, 402(??):9–17, August 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300522.

Scott:2018:CSF

[SMG18] Ryan Scott, Brian MacPherson, and Robin Gras. A comparison of stable and fluctuating resources with respect to evolutionary adaptation and life-history traits using individualbased modeling and machine learning. Journal of Theoretical Biology, 459(??):52–66, December 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304557.

Samozino:2010:JAT

[SMHB10] Pierre Samozino, Jean-Benoît Morin, Frédérique Hintzy, and Alain Belli. Jumping ability: a theoretical integrative approach. Journal of Theoretical Biology, 264(1):11–18, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310000330.

Smith:2011:IEC

[Smi11] D. A. Smith. The interaction energy of charged filaments in an electrolyte: Results for all filament spacings. Journal of Theoretical Biology, 276(1):8–15, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000725.

Smith:2014:EFS

[Smi14] David A. Smith. Electrostatic forces or structural scaffolding: What stabilizes the lattice spacing of relaxed skinned muscle fibers? Journal of Theoretical Biology, 355(??):53–60, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001933.

Shinde:2014:MMM

[SMJS14] Sonali Shinde, Milsee Mol, Virashree Jamdar, and Shailza Singh. Molecular modeling and molecular dynamics simulations of GPI 14 in Leishmania major: Insight into the catalytic site for active site directed drug design. Journal of Theoretical Biology, 351(??):37–46, June 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000927.

Soler:2013:WWP

[SMM^{+}13] L. Soler, A. Molenaar, N. Merola, P. D. Eckersall, A. Gutiérrez, J. J. Cerón, V. Mulero, and T. A. Niewold. Why working with porcine circulating serum amyloid a is a pig of a job. Journal of Theoretical Biology, 317(??):119–125, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931200536X.

Saez:2015:TME

[SMM15] P. Sáez, M. Malvè, and M. A. Martínez. A theoretical model of the endothelial cell morphology due to different waveforms. Journal of Theoretical Biology, 379(??):16–23, August 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002209.

Sartakhti:2017:MTI

[SMS17] Javad Salimi Sartakhti, Mohammad Hossein Manshaei, and Mehdi Sadeghi. MMP-TIMP interactions in cancer invasion: An evolutionary game-theoretical framework. Journal of Theoretical Biology, 412(??):17–26, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303162.

Sheraton:2019:PQB

[SMS19] M. V. Sheraton, V. R. Melnikov, and P. M. A. Sloot. Prediction and quantification of bacterial biofilm detachment using Glazier–Graner–Hogeweg method based model simulations. Journal of Theoretical Biology, 482(??):Article 109994, December 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303443.

Sneyd:2017:MCW

[SMZ^{+}17] James Sneyd, Shawn Means, Di Zhu, John Rugis, Jong Hak Won, and David I. Yule. Modeling calcium waves in an anatomically accurate three-dimensional parotid acinar cell. Journal of Theoretical Biology, 419(??):383–393, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300583.

Sekiguchi:2011:HIB

[SN11] Takuya Sekiguchi and Mayuko Nakamaru. How inconsistency between attitude and behavior persists through cultural transmission. Journal of Theoretical Biology, 271(1):124–135, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006399.

Sakurai:2012:MNU

[SN12] Shunsuke Sakurai and Seido Nagano. A molecular network underlying spontaneous cAMP oscillation and synchronization in Dictyostelium. Journal of Theoretical Biology, 307(??):37–41, August 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002275.

Sekiguchi:2014:HII

[SN14] Takuya Sekiguchi and Mayuko Nakamaru. How intergenerational interaction affects attitude-behavior inconsistency. Journal of Theoretical Biology, 346(??):54–66, April 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005730.

Seki:2016:MGM

[SN16] Motohide Seki and Mayuko Nakamaru. A model for gossipmediated evolution of altruism with various types of false information by speakers and assessment by listeners. Journal of Theoretical Biology, 407(??):90–105, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301746.

Szatmary:2017:DWO

[SN17] A. C. Szatmary and R. Nossal. Determining whether observed eukaryotic cell migration indicates chemotactic responsiveness or random chemokinetic motion. Journal of Theoretical Biology, 425(??):103–112, July 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302230.

Shimura:2018:LGS

[SN18] Hayato Shimura and Mayuko Nakamaru. Large group size promotes the evolution of cooperation in the mutual-aid game. Journal of Theoretical Biology, 451(??):46–56, August 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301796.

Shayeghi:2012:DRA

[SNC12] N. Shayeghi, T. Ng, and A. C. C. Coolen. Direct response analysis in cellular signalling networks. Journal of Theoretical Biology, 304(??):219–225, July 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001749.

Suzuki:2012:ICA

[SNCM12] Yasuyuki Suzuki, Taishin Nomura, Maura Casadio, and Pietro Morasso. Intermittent control with ankle, hip, and mixed strategies during quiet standing: a theoretical proposal based on a double inverted pendulum model. Journal of Theoretical Biology, 310(??):55–79, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003062.

Sukhov:2011:SAP

[SNOV11] Vladimir Sukhov, Vladimir Nerush, Lyubov Orlova, and Vladimir Vodeneev. Simulation of action potential propagation in plants. Journal of Theoretical Biology, 291(??): 47–55, December 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519311004759.

Sabzekar:2017:EDP

[SNS17] Mostafa Sabzekar, Mahmoud Naghibzadeh, and Javad Sadri. Efficient dynamic programming algorithm with prior knowledge for protein β-strand alignment. Journal of Theoretical Biology, 417(??):43–50, March 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300188.

Shoji:2013:NIB

[SNY^{+}13] Hiroto Shoji, Yasuhiro Nakatomi, Chihiro Yokoyama, Kenji Fukui, and Kazumitsu Hanai. New index based on the physical separation of motion into three categories for characterizing the effect of cocaine in mice. Journal of Theoretical Biology, 333(??):68–77, September 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002233.

Sasaki:2017:MMM

[SNY^{+}17] Daiki Sasaki, Hitomi Nakajima, Yoshimi Yamaguchi, Ryuji Yokokawa, Shin-Ichiro Ei, and Takashi Miura. Mathematical modeling for meshwork formation of endothelial cells in fibrin gels. Journal of Theoretical Biology, 429(??):95–104, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302837.

Somersalo:2012:RDM

[SOBC12] Erkki Somersalo, Rossana Occhipinti, Walter F. Boron, and Daniela Calvetti. A reaction-diffusion model of CO 2 influx into an oocyte. Journal of Theoretical Biology, 309(??):185–203, September 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003037.

Simons:2014:DSD

[SOCF14] Julie Simons, Sarah Olson, Ricardo Cortez, and Lisa Fauci. The dynamics of sperm detachment from epithelium in a coupled fluid-biochemical model of hyperactivated motility. Journal of Theoretical Biology, 354(??):81–94, August 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001635.

Soto-Ortiz:2016:CTB

[SOF16] Luis Soto-Ortiz and Stacey D. Finley. A cancer treatment based on synergy between anti-angiogenic and immune cell therapies. Journal of Theoretical Biology, 394(??):197–211, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931600059X.

Suzuki-Ohno:2010:AGP

[SOIO10] Yukari Suzuki-Ohno, Maki N. Inoue, and Kazunori Ohno. Applying geographic profiling used in the field of criminology for predicting the nest locations of bumble bees. Journal of Theoretical Biology, 265(2):211–217, July 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001931.

Soslau:2018:CRC

[Sos18] Gerald Soslau. Circular RNA (circRNA) was an important bridge in the switch from the RNA world to the DNA world. Journal of Theoretical Biology, 447(??):32–40, June 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318301358.

Sozou:2013:KSL

[Soz13] Peter D. Sozou. Kin selection and local competition in a heterogeneous community. Journal of Theoretical Biology, 332(??):261–266, September 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300218X.

Sirot:2011:DPR

[SP11] Etienne Sirot and Olivier Pays. On the dynamics of predation risk perception for a vigilant forager. Journal of Theoretical Biology, 276(1):1–7, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000713.

Scianna:2013:MIN

[SP13a] Marco Scianna and Luigi Preziosi. Modeling the influence of nucleus elasticity on cell invasion in fiber networks and microchannels. Journal of Theoretical Biology, 317(??):394–406, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005784.

Szolnoki:2013:ECP

[SP13b] Attila Szolnoki and Matjaz Perc. Effectiveness of conditional punishment for the evolution of public cooperation. Journal of Theoretical Biology, 325(??):34–41, May 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000763.

Sutradhar:2014:TWB

[SP14] S. Sutradhar and R. Paul. Tug-of-war between opposing molecular motors explains chromosomal oscillation during mitosis. Journal of Theoretical Biology, 344(??):56–69, March 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931300550X.

Simon:2016:GLE

[SP16] Burton Simon and Michael Pilosov. Group-level events are catalysts in the evolution of cooperation. Journal of Theoretical Biology, 410(??):125–136, December 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302545.

Susree:2018:CPI

[SPA18] M. Susree, Mikhail A. Panteleev, and M. Anand. Coated platelets introduce significant delay in onset of peak thrombin production: Theoretical predictions. Journal of Theoretical Biology, 453(??):108–116, September 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302613.

Steel:2018:EIP

[SPCM18] Mike Steel, Vahab Pourfaraj, Abhishek Chaudhary, and Arne Mooers. Evolutionary isolation and phylogenetic diversity loss under random extinction events. Journal of Theoretical Biology, 438(??):151–155, February 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305064.

Spencer:2015:SCS

[Spe15] Matthew Spencer. Size change, shape change, and the growth space of a community. Journal of Theoretical Biology, 369(??):23–41, March 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500003X.

Sneha:2018:SDH

[SPG^{+}18] Patil Sneha, Pritam Kumar Panda, Fatemeh Rahimi Gharemirshamlu, Kourosh Bamdad, and Seetharaman Balaji. Structural discordance in HIV-1 Vpu from brain isolate alarms amyloid fibril forming behavior- a computational perspective. Journal of Theoretical Biology, 451(??):35–45, August 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302108.

Staaterman:2012:OBF

[SPH12] Erica Staaterman, Claire B. Paris, and Judith Helgers. Orientation behavior in fish larvae: a missing piece to Hjort’s critical period hypothesis. Journal of Theoretical Biology, 304(??):188–196, July 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001336.

Szymanska:2014:MMI

[SPL14] Zuzanna Szymańska, Martin Parisot, and Miroslaw Lachowicz. Mathematical modeling of the intracellular protein dynamics: The importance of active transport along microtubules. Journal of Theoretical Biology, 363(??):118–128, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004238.

Santos:2012:RDE

[SPLP12] Francisco C. Santos, Flávio L. Pinheiro, Tom Lenaerts, and Jorge M. Pacheco. The role of diversity in the evolution of cooperation. Journal of Theoretical Biology, 299(??):88–96, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004590.

Serebrennikova:2010:QAM

[SPMGR10] Yulia M. Serebrennikova, Janus Patel, Wilbur K. Milhous, and Luis H. García-Rubio. Quantitative analysis of morphological alterations in Plasmodium falciparum infected red blood cells through theoretical interpretation of spectral measurements. Journal of Theoretical Biology, 265(4):493–500, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002808.

Shorten:2010:MMP

[SPO^{+}10] P. R. Shorten, A. J. Peterson, A. R. O’Connell, J. L. Juengel, K. P. McNatty, and T. K. Soboleva. A mathematical model of pregnancy recognition in mammals. Journal of Theoretical Biology, 266(1):62–69, September 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002894.

Simon:2013:MBC

[SPRF13] Carl P. Simon, Bethany Percha, Rick Riolo, and Betsy Foxman. Modeling bacterial colonization and infection routes in health care settings: Analytic and numerical approaches. Journal of Theoretical Biology, 334(??):187–199, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002452.

Santos:2011:CEP

[SPS11] Francisco C. Santos, Jorge M. Pacheco, and Brian Skyrms. Co-evolution of pre-play signaling and cooperation. Journal of Theoretical Biology, 274(1):30–35, April 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000087.

Samal:2015:CGS

[SPSM15] Himanshu Bhusan Samal, Jyoti Prava, Mrutyunjay Suar, and Rajani Kanta Mahapatra. Comparative genomics study of Salmonella Typhimurium LT2 for the identification of putative therapeutic candidates. Journal of Theoretical Biology, 369(??):67–79, March 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000314.

Santos:2012:DPS

[SPSP12] Marta D. Santos, Flávio L. Pinheiro, Francisco C. Santos, and Jorge M. Pacheco. Dynamics of N-person snowdrift games in structured populations. Journal of Theoretical Biology, 315(??):81–86, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004699.

Shi:2012:MDB

[SQS^{+}12] Shao-Ping Shi, Jian-Ding Qiu, Xing-Yu Sun, Sheng-Bao Suo, Shu-Yun Huang, and Ru-Ping Liang. A method to distinguish between lysine acetylation and lysine methylation from protein sequences. Journal of Theoretical Biology, 310(??):223–230, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003232.

Sun:2016:NCB

[SQZ^{+}16] Pingping Sun, Jialiang Qi, Yizhu Zhao, Yanxin Huang, Guifu Yang, Zhiqiang Ma, and Yuxin Li. A novel conformational bcell epitope prediction method based on mimotope and patch analysis. Journal of Theoretical Biology, 394(??):102–108, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316000540.

Sheng:2017:JPE

[SQZZ17] Xiaona Sheng, Yihong Qiu, Ying Zhou, and Wensheng Zhu. Joint parameter estimation in the QTL mapping of ordinal traits. Journal of Theoretical Biology, 432(??):100–108, November 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303739.

Shen:2010:RST

[SR10] Sheng-Feng Shen and H. Kern Reeve. Reproductive skew theory unified: The general bordered tug-of-war model. Journal of Theoretical Biology, 263(1):1–12, March 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005359.

Simpson:2012:MHH

[SR12] Joanne Simpson and Mick Roberts. Modelling heterogeneous host immune response in a multi-strain system. Journal of Theoretical Biology, 304(??):60–65, July 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001233.

Simkin:2014:SMS

[SR14] M. V. Simkin and V. P. Roychowdhury. Stochastic modeling of a serial killer. Journal of Theoretical Biology, 355(??):111–116, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001957.

Sumen:2012:CAB

[SRAL12] Cenk Sumen, Timothy E. Reddy, Mark S. Alber, and Peter P. Lee. Corrigendum to: “Agent-based modeling of the context dependency in T cell recognition” [J. Theor. Biol. 236 (2005) 376–391]. Journal of Theoretical Biology, 303(??):152–153, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001427. See [CSR^{+}05].

Soubeyrand:2011:PPD

[SRCF11] Samuel Soubeyrand, Lionel Roques, Jérôme Coville, and Julien Fayard. Patchy patterns due to group dispersal. Journal of Theoretical Biology, 271(1):87–99, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006429.

Saini:2015:SLG

[SRD^{+}15] Harsh Saini, Gaurav Raicar, Abdollah Dehzangi, Sunil Lal, and Alok Sharma. Subcellular localization for Gram positive and Gram negative bacterial proteins using linear interpolation smoothing model. Journal of Theoretical Biology, 386(??):25–33, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004130.

Spoljaric:2019:MMC

[SRH19] Ana Martincić Spoljarić, Ivica Rubelj, and Miljenko Huzak. Mathematical model and computer simulations of telomere loss. Journal of Theoretical Biology, 465(??):78–89, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300074.

Saitou:2012:MMI

[SRK^{+}12] Takashi Saitou, Mahemuti Rouzimaimaiti, Naohiko Koshikawa, Motoharu Seiki, Kazuhisa Ichikawa, and Takashi Suzuki. Mathematical modeling of invadopodia formation. Journal of Theoretical Biology, 298(??):138–146, April 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006424.

Sefid:2016:HMC

[SRP16] Fateme Sefid, Iraj Rasooli, and Zahra Payandeh. Homology modeling of a Camelid antibody fragment against a conserved region of Acinetobacter baumannii biofilm associated protein (Bap). Journal of Theoretical Biology, 397(??):43–51, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001053.

Saini:2015:PES

[SRS^{+}15] Harsh Saini, Gaurav Raicar, Alok Sharma, Sunil Lal, Abdollah Dehzangi, James Lyons, Kuldip K. Paliwal, Seiya Imoto, and Satoru Miyano. Probabilistic expression of spatially varied amino acid dimers into general form of Chou’s pseudo amino acid composition for protein fold recognition. Journal of Theoretical Biology, 380(??):291–298, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002659.

Song:2018:GSF

[SRS18] Chuliang Song, Rudolf P. Rohr, and Serguei Saavedra. A guideline to study the feasibility domain of multi-trophic and changing ecological communities. Journal of Theoretical Biology, 450(??):30–36, August 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302042.

Shen:2011:PCC

[SRV11] Sheng-Feng Shen, H. Kern Reeve, and Sandra L. Vehrencamp. Parental care, cost of reproduction and reproductive skew: a general costly young model. Journal of Theoretical Biology, 284(1):24–31, September 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002803.

Suzuki:2011:HDR

[SS11] Sayaki U. Suzuki and Akira Sasaki. How does the resistance threshold in spatially explicit epidemic dynamics depend on the basic reproductive ratio and spatial correlation of crop genotypes? Journal of Theoretical Biology, 276(1):117–125, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000828.

Stadler:2012:DBL

[SS12a] Tanja Stadler and Mike Steel. Distribution of branch lengths and phylogenetic diversity under homogeneous speciation models. Journal of Theoretical Biology, 297(??):33–40, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311005893.

Szabo:2012:SFO

[SS12b] György Szabó and Attila Szolnoki. Selfishness, fraternity, and other-regarding preference in spatial evolutionary games. Journal of Theoretical Biology, 299(??):81–87, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931100155X.

Sharifimajd:2014:CME

[SS14] Babak Sharifimajd and Jonas Stålhand. A continuum model for excitation-contraction of smooth muscle under finite deformations. Journal of Theoretical Biology, 355(??):1–9, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001519.

Seri:2015:SPT

[SS15a] Efrat Seri and Nadav Shnerb. Spatial patterns in the tropical forest reveal connections between negative feedback, aggregation and abundance. Journal of Theoretical Biology, 380(??):247–255, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002702.

Shen:2015:APL

[SS15b] ZhuoHua Shen and Justin Seipel. Animals prefer leg stiffness values that may reduce the energetic cost of locomotion. Journal of Theoretical Biology, 364(??):433–438, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005475.

Shen:2015:LSA

[SS15c] ZhuoHua Shen and Justin Seipel. The leg stiffnesses animals use may improve the stability of locomotion. Journal of Theoretical Biology, 377(??):66–74, July 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001721.

Soares:2015:ITS

[SS15d] André E. R. Soares and Carlos G. Schrago. The influence of taxon sampling on Bayesian divergence time inference under scenarios of rate heterogeneity among lineages. Journal of Theoretical Biology, 364(??):31–39, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005426.

Sober:2015:HPC

[SS15e] Elliott Sober and Mike Steel. How probable is common ancestry according to different evolutionary processes? Journal of Theoretical Biology, 373(??):111–116, May 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001216.

Sumathy:2015:MBM

[SS15f] S. Sumathy and S. V. M. Satyanarayana. Model for bidirectional movement of cytoplasmic dynein. Journal of Theoretical Biology, 380(??):48–52, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002118.

Saghapour:2017:PMA

[SS17a] Ehsan Saghapour and Mohammadreza Sehhati. Prediction of metastasis in advanced colorectal carcinomas using CGH data. Journal of Theoretical Biology, 429(??):116–123, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303016.

Stroberg:2017:ONM

[SS17b] Wylie Stroberg and Santiago Schnell. On the origin of nonmembrane-bound organelles, and their physiological function. Journal of Theoretical Biology, 434(??):42–49, December 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301546.

Shen:2018:ELS

[SS18] ZhuoHua Shen and Justin Seipel. Effective leg stiffness of animal running and the co-optimization of energetic cost and stability. Journal of Theoretical Biology, 451(??):57–66, August 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301802.

Sato:2019:SAW

[SS19] Naoki Sato and Kaoru Sato. Statistical analysis of word usage in biological publications since 1965: Historical delineation highlighting an emergence of function-oriented discourses in contemporary molecular and cellular biology. Journal of Theoretical Biology, 462(??):293–303, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305708.

Serov:2015:OVD

[SSB^{+}15] A. S. Serov, C. M. Salafia, P. Brownbill, D. S. Grebenkov, and M. Filoche. Optimal villi density for maximal oxygen uptake in the human placenta. Journal of Theoretical Biology, 364(??):383–396, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400561X.

Szabo:2013:CFE

[SSC13] György Szabó, Attila Szolnoki, and Lilla Czakó. Coexistence of fraternity and egoism for spatial social dilemmas. Journal of Theoretical Biology, 317(??):126–132, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005395.

Serrano-Solis:2016:GSV

[SSCJ16] Víctor Serrano-Solís, Germinal Cocho, and Marco V. José. Genomic signatures in viral sequences by in-frame and out-frame mutual information. Journal of Theoretical Biology, 403(??):1–9, August 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300923.

Safaeian:2011:CMH

[SSD11a] Navid Safaeian, Mathieu Sellier, and Tim David. A computational model of hemodynamic parameters in cortical capillary networks. Journal of Theoretical Biology, 271(1):145–156, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006338.

Song:2011:GPC

[SSD11b] Yixian Song, Stefan Scheu, and Barbara Drossel. Geographic parthenogenesis in a consumer-resource model for sexual reproduction. Journal of Theoretical Biology, 273(1):55–62, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006752.

Song:2012:LHC

[SSD12] Yixian Song, Stefan Scheu, and Barbara Drossel. Life histories and Cope’s rule from an explicit resource-consumer model based on metabolic theory. Journal of Theoretical Biology, 310(??):175–182, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312003141.

Sah:2013:SBP

[SSD13] Pratha Sah, Joseph Paul Salve, and Sutirth Dey. Stabilizing biological populations and metapopulations through adaptive limiter control. Journal of Theoretical Biology, 320(??):113–123, March 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931200642X.

Santos:2015:PPB

[SSF15] Mauro Santos, Eörs Szathmáry, and José F. Fontanari. Phenotypic plasticity, the Baldwin effect, and the speeding up of evolution: The computational roots of an illusion. Journal of Theoretical Biology, 371(??):127–136, April 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000715.

Serov:2015:ATO

[SSFG15] A. S. Serov, C. M. Salafia, M. Filoche, and D. S. Grebenkov. Analytical theory of oxygen transport in the human placenta. Journal of Theoretical Biology, 368(??):133–144, March 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314007139.

Senapati:2019:IAM

[SSG^{+}19] Abhishek Senapati, Tridip Sardar, Krishnendra Sankar Ganguly, Krishna Sankar Ganguly, Asis Kumar Chattopadhyay, and Joydev Chattopadhyay. Impact of adult mosquito control on dengue prevalence in a multi-patch setting: a case study in Kolkata (2014–2015). Journal of Theoretical Biology, 478(??):139–152, October 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302565.

Sharma:2019:OCB

[SSH^{+}19] Saurabh Kumar Sharma, Soibam Shyamchand Singh, Dineshchandra Haobijam, Md. Zubbair Malik, and R. K. Brojen Singh. Organization in complex brain networks: Energy distributions and phase shift. Journal of Theoretical Biology, 476(??):30–35, September 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319302139.

Satake:2016:FDD

[SSI^{+}16] Akiko Satake, Motohide Seki, Makoto Iima, Takashi Teramoto, and Yasumasa Nishiura. Florigen distribution determined by a source–sink balance explains the diversity of inflorescence structures in Arabidopsis. Journal of Theoretical Biology, 395(??):227–237, April 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000746.

Scheiblauer:2018:FSC

[SSJK18] Johannes Scheiblauer, Stefan Scheiner, Martin Joksch, and Barbara Kavsek. Fermentation of Saccharomyces cerevisiae — combining kinetic modeling and optimization techniques points out avenues to effective process design. Journal of Theoretical Biology, 453(??):125–135, September 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302558.

Strang:2017:SLB

[SSKK17] Jan Eric Sträng, Rainer Schuler, Michael Kühl, and Hans A. Kestler. Switch-like behavior enables Wnt11 concentration specific response during dorso-ventral axis formation in Xenopus laevis. Journal of Theoretical Biology, 429(??):82–94, September 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303041.

Stadler:2018:SAS

[SSKS18] Tanja Stadler, Stavroula Skylaki, Konstantinos D. Kokkaliaris, and Timm Schroeder. On the statistical analysis of single cell lineage trees. Journal of Theoretical Biology, 439(??):160–165, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305301. See erra-

tum [SSKS20].

Stadler:2020:ESA

[SSKS20] Tanja Stadler, Stavroula Skylaki, Konstantinos D. Kokkaliaris, and Timm Schroeder. Erratum to “On the statistical analysis of single cell lineage trees” [J. Theor. Biol. 439 (2018) 160–165]. Journal of Theoretical Biology, 506 (??):Article 110424, December 7, 2020. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519320302794. See [SSKS18].

Sikic:2015:SME

[SSLB15] Hrvoje Sikić, Yanrong Shi, Snjezana Lubura, and Steven Bassnett. A stochastic model of eye lens growth. Journal of Theoretical Biology, 376(??):15–31, July 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001307.

Schank:2015:EFD

[SSM15] Jeffrey C. Schank, Paul E. Smaldino, and Matt L. Miller. Evolution of fairness in the dictator game by multilevel selection. Journal of Theoretical Biology, 382(??):64–73, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003100.

Shariatpanahi:2018:MMT

[SSM^{+}18] Seyed Peyman Shariatpanahi, Seyed Pooya Shariatpanahi, Keivan Madjidzadeh, Moustapha Hassan, and Manuchehr Abedi-Valugerdi. Mathematical modeling of tumor-induced immunosuppression by myeloid-derived suppressor cells: Implications for therapeutic targeting strategies. Journal of Theoretical Biology, 442(??):1–10, April 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300067.

Sole:2010:SMR

[SSME10] Ricard V. Solé, Joan Saldaña, Jose M. Montoya, and Douglas H. Erwin. Simple model of recovery dynamics after mass extinction. Journal of Theoretical Biology, 267(2):193–200, November 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004248.

Suzuki:2014:NPI

[SSN^{+}14] Yasuyuki Suzuki, Daichi Sakai, Taishin Nomura, Yoshito Hirata, and Kazuyuki Aihara. A new protocol for intermittent androgen suppression therapy of prostate cancer with unstable saddle-point dynamics. Journal of Theoretical Biology, 350(??):1–16, June 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000654.

Shreif:2015:JPS

[SSP15] Zeina Shreif, Deborah A. Striegel, and Vipul Periwal. The jigsaw puzzle of sequence phenotype inference: Piecing together Shannon entropy, importance sampling, and empirical Bayes. Journal of Theoretical Biology, 380(??):399–413, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002891.

Santos:2015:EDG

[SSPP15] Fernando P. Santos, Francisco C. Santos, Ana Paiva, and Jorge M. Pacheco. Evolutionary dynamics of group fairness. Journal of Theoretical Biology, 378(??):96–102, August 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002015.

Sorrell:2019:MML

[SSR^{+}19] Ian Sorrell, Rebecca J Shipley, Sophie Regan, Iain Gardner, Michael P. Storm, Marianne Ellis, John Ward, Dominic Williams, Pratibha Mistry, José Domingo Salazar, Andrew Scott, and Steven Webb. Mathematical modelling of a liver hollow fibre bioreactor. Journal of Theoretical Biology, 475(??):25–33, August 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302000.

Swetha:2016:ICC

[SSRA16] Rayapadi G. Swetha, Madangopal Sandhya, Sudha Ramaiah, and Anand Anbarasu. Identification of CD4+ T-cell epitope and investigation of HLA distribution for the immunogenic proteins of Burkholderia pseudomallei using in silico approaches — a key vaccine development strategy for melioidosis. Journal of Theoretical Biology, 400(??):11–18, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300303.

Sand:2013:SLG

[SSS13] Andreas Sand, Andreas Sand, and Mike Steel. The standard lateral gene transfer model is statistically consistent for pectinate four-taxon trees. Journal of Theoretical Biology, 335(??):295–298, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003238.

Shokry:2015:DSD

[SSS15] Abdallah Shokry, Per Ståhle, and Ingrid Svensson. Determination of spatially dependent diffusion parameters in bovine bone using Kalman filter. Journal of Theoretical Biology, 384(??):10–18, November 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003707.

Shibasaki:2017:CIO

[SSS17] Shota Shibasaki, Yuka Shirokawa, and Masakazu Shimada. Cooperation induces other cooperation: Fruiting bodies promote the evolution of macrocysts in Dictyostelium discoideum. Journal of Theoretical Biology, 421(??):136–145, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301509.

Shatabda:2017:IIB

[SSSD17] Swakkhar Shatabda, Sanjay Saha, Alok Sharma, and Abdollah Dehzangi. iPHLoc-ES: Identification of bacteriophage protein locations using evolutionary and structural features. Journal of Theoretical Biology, 435(??):229–237, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317304393.

Shah:2019:OCH

[SSST19] Nita H. Shah, Nehal Shukla, Moksha H. Satia, and Foram A. Thakkar. Optimal control of HCV transmission under liquoring. Journal of Theoretical Biology, 465(??):27–33, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300049.

Szalisznyo:2019:NDC

[SST19] Krisztina Szalisznyó, David N. Silverstein, and János Tóth. Neural dynamics in co-morbid schizophrenia and OCD: a computational approach. Journal of Theoretical Biology, 473(??):80–94, July 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300554.

Shokrollahzade:2015:PCA

[SSV^{+}15] Soheila Shokrollahzade, Fatemeh Sharifi, Akbar Vaseghi, Maryam Faridounnia, and Samad Jahandideh. Protein cold adaptation: Role of physico-chemical parameters in adaptation of proteins to low temperatures. Journal of Theoretical Biology, 383(??):130–137, October 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003501.

Schumm:2010:NMS

[SSvdM10] Phillip Schumm, Caterina M. Scoglio, and Deon van der Merwe. A network model of successive partitioning-limited solute diffusion through the stratum corneum. Journal of Theoretical Biology, 262(3):471–477, February 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004913.

Schumm:2015:GEI

[SSZB15] Phillip Schumm, Caterina Scoglio, Qian Zhang, and Duygu Balcan. Global epidemic invasion thresholds in directed cattle subpopulation networks having source, sink, and transit nodes. Journal of Theoretical Biology, 367(??):203–221, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006961.

Sataric:2017:BMH

[SSZR17] Miljko V. Sataric, Dalibor L. Sekulic, Slobodan Zdravkovic, and Nebojsa M. Ralevic. A biophysical model of how α-tubulin carboxy-terminal tails tune kinesin-1 processivity along microtubule. Journal of Theoretical Biology, 420(??):152–157, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301236.

Sanchez-Taltavull:2016:OAD

[ST16] Daniel Sánchez-Taltavull. Optimal architecture of differentiation cascades with asymmetric and symmetric stem cell division. Journal of Theoretical Biology, 407(??):106–117, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302120.

Sakai:2017:AES

[ST17a] Yuma Sakai and Takenori Takada. The analysis of an effect of seed propagation on defense strategy against pathogen transmission within clonal plant population using lattice model. Journal of Theoretical Biology, 427(??):65–76, August 1, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302035.

Syomin:2017:SMC

[ST17b] Fyodor A. Syomin and Andrey K. Tsaturyan. A simple model of cardiac muscle for multiscale simulation: Passive mechanics, crossbridge kinetics and calcium regulation. Journal of Theoretical Biology, 420(??):105–116, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300760.

Stadler:2010:STT

[Sta10] Tanja Stadler. Sampling-through-time in birth-death trees. Journal of Theoretical Biology, 267(3):396–404, December 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310004765.

Sanchez-Taltavull:2015:SMV

[STA15] Daniel Sánchez-Taltavull and Tomás Alarcón. Stochastic modelling of viral blips in HIV-1-infected patients: Effects of inhomogeneous density fluctuations. Journal of Theoretical Biology, 371(??):79–89, April 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000600.

Stepanskiy:2012:SIU

[Ste12] Leonard G. Stepanskiy. Sonication-induced unfolding proteins. Journal of Theoretical Biology, 298(??):77–81, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000094.

Steel:2013:CBI

[Ste13] Mike Steel. Consistency of Bayesian inference of resolved phylogenetic trees. Journal of Theoretical Biology, 336(??):246–249, November 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300386X.

Schenk:2017:CPK

[STG17] Hanna Schenk, Arne Traulsen, and Chaitanya S. Gokhale. Chaotic provinces in the kingdom of the Red Queen. Journal of Theoretical Biology, 431(??):1–10, October 27, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303582.

Shen:2019:IPS

[STG19] Yinan Shen, Jijun Tang, and Fei Guo. Identification of protein subcellular localization via integrating evolutionary and physicochemical information into Chou’s general PseAAC. Journal of Theoretical Biology, 462(??):230–239, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305654.

Schuss:2018:DCS

[STH18] Z. Schuss, K. Tor, and D. Holcman. Do cells sense time by number of divisions? Journal of Theoretical Biology, 452(??):10–16, September 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302261.

Seto:2013:ICS

[STI13] Mayumi Seto, Noriko Takamura, and Yoh Iwasa. Individual and combined suppressive effects of submerged and floating-leaved macrophytes on algal blooms. Journal of Theoretical Biology, 319(??):122–133, February 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005917.

Sarbaz:2012:DCF

[STJG12] Yashar Sarbaz, Farzad Towhidkhah, Ayyoob Jafari, and Shahriar Gharibzadeh. Do the chaotic features of gait change in Parkinson’s disease? Journal of Theoretical Biology, 307(??):160–167, August 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002135.

Sudarsan:2012:FCV

[STKE12] Rangarajan Sudarsan, Cody Thompson, Peter G. Kevan, and Hermann J. Eberl. Flow currents and ventilation in Langstroth beehives due to brood thermoregulation efforts of honeybees. Journal of Theoretical Biology, 295(??):168–193, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005698.

Synodinos:2018:IIA

[STLJ18] Alexis D. Synodinos, Britta Tietjen, Dirk Lohmann, and Florian Jeltsch. The impact of inter-annual rainfall variability on African savannas changes with mean rainfall. Journal of Theoretical Biology, 437(??):92–100, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304824.

Sasmal:2019:CMA

[STN19a] Sourav Kumar Sasmal, Yasuhiro Takeuchi, and Shinji Nakaoka. T-cell mediated adaptive immunity and antibodydependent enhancement in secondary dengue infection. Journal of Theoretical Biology, 470(??):50–63, June 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301110.

Shiraishi:2019:DSL

[STN^{+}19b] Masashi Shiraishi, Rito Takeuchi, Hiroyuki Nakagawa, Shin I. Nishimura, Akinori Awazu, and Hiraku Nishimori. Diverse stochasticity leads a colony of ants to optimal foraging. Journal of Theoretical Biology, 465(??):7–16, March 21, 2019.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300037.

Sano:2017:DCB

[STNT17] Hitomi I. Sano, Tamami Toki, Yasuhiro Naito, and Masaru Tomita. Developmental changes in the balance of glycolytic ATP production and oxidative phosphorylation in ventricular cells: a simulation study. Journal of Theoretical Biology, 419(??):269–277, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300759.

Strombom:2011:CML

[Str11] Daniel Strömbom. Collective motion from local attraction. Journal of Theoretical Biology, 283(1):145–151, August 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931100261X.

Strother:2013:CMF

[Str13] James A. Strother. A computational model of flow between the microscale respiratory structures of fish gills. Journal of Theoretical Biology, 338(??):23–40, December 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003895.

Straube:2015:ASC

[Str15] Ronny Straube. Analysis of substrate competition in regulatory network motifs: Stimulus-response curves, thresholds and ultrasensitivity. Journal of Theoretical Biology, 380(??):74–82, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315002465.

Straube:2017:ORC

[Str17] Ronny Straube. Operating regimes of covalent modification cycles at high enzyme concentrations. Journal of Theoretical Biology, 431(??):39–48, October 27, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303661.

Sturrock:2011:STM

[STX^{+}11] Marc Sturrock, Alan J. Terry, Dimitris P. Xirodimas, Alastair M. Thompson, and Mark A. J. Chaplain. Spatio-temporal modelling of the Hes1 and p53-Mdm2 intracellular signalling pathways. Journal of Theoretical Biology, 273(1):15–31, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006715.

Sasaki:2011:RDP

[SU11] Tatsuya Sasaki and Tatsuo Unemi. Replicator dynamics in public goods games with reward funds. Journal of Theoretical Biology, 287(??):109–114, October 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003766.

Su:2016:AFP

[Su16] Qiang Su. Analyzing fractal property of species abundance distribution and diversity indexes. Journal of Theoretical Biology, 392(??):107–112, March 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315006013.

Sueur:2012:VDM

[Sue12] Cédric Sueur. Viability of decision-making systems in human and animal groups. Journal of Theoretical Biology, 306(??):93–103, August 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312002019.

Sumner:2013:LGM

[Sum13] Jeremy G. Sumner. Lie geometry of 2×2 Markov matrices. Journal of Theoretical Biology, 327(??):88–90, June 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313000519.

Sen:2014:PFU

[SV14] Siddhartha Sen and H. Paul Voorheis. Protein folding: Understanding the role of water and the low Reynolds number environment as the peptide chain emerges from the ribosome and folds. Journal of Theoretical Biology, 363(??):169–187, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314004378.

Shen:2017:PPS

[SV17] Lulan Shen and Robert A. Van Gorder. Predator–preysubsidy population dynamics on stepping-stone domains. Journal of Theoretical Biology, 420(??):241–258, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301248.

Srinivasu:2018:AFS

[SVA18] P. D. N. Srinivasu, D. K. K. Vamsi, and V. S. Ananth. Additional food supplements as a tool for biological conservation of predator–prey systems involving type III functional response: a qualitative and quantitative investigation. Journal of Theoretical Biology, 455(??):303–318, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318303424.

Serra:2010:DRB

[SVB^{+}10] R. Serra, M. Villani, A. Barbieri, S. A. Kauffman, and A. Colacci. On the dynamics of random Boolean networks subject to noise: Attractors, ergodic sets and cell types. Journal of Theoretical Biology, 265(2):185–193, July 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001955.

Szabo:2010:ISE

[SVCS10] László Zsolt Szabó, János Vincze, László Csernoch, and Péter Szentesi. Improved spark and ember detection using stationary wavelet transforms. Journal of Theoretical Biology, 264(4):1279–1292, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001888.

Schellekens:2012:CTS

[SvK12] Tim Schellekens and Tobias van Kooten. Coexistence of two stage-structured intraguild predators. Journal of Theoretical Biology, 308(??):36–44, September 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002597.

Stefanoni:2011:NMD

[SVMN11] Filippo Stefanoni, Maurizio Ventre, Francesco Mollica, and Paolo A. Netti. A numerical model for durotaxis. Journal of Theoretical Biology, 280(1):150–158, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001913.

Skutkova:2015:PAG

[SVSP15] Helena Skutkova, Martin Vitek, Karel Sedlar, and Ivo Provaznik. Progressive alignment of genomic signals by multiple dynamic time warping. Journal of Theoretical Biology, 385(??):20–30, November 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003951.

Shiri:2011:MIA

[SW11] Tinevimbo Shiri and Alex Welte. Modelling the impact of acute infection dynamics on the accumulation of HIV-1 mutations. Journal of Theoretical Biology, 279(1):44–54, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001512.

Salehipour:2013:CKE

[SW13a] Hesam Salehipour and David J. Willis. A coupled kinematicsenergetics model for predicting energy efficient flapping flight. Journal of Theoretical Biology, 318(??):173–196, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005334.

Stamper:2013:MMI

[SW13b] I. Johanna Stamper and Xujing Wang. Mathematical modeling of insulin secretion and the role of glucose-dependent mobilization, docking, priming and fusion of insulin granules. Journal of Theoretical Biology, 318(??):210–225, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005772.

Schmeiser:2015:FLE

[SW15] Christian Schmeiser and Christoph Winkler. The flatness of Lamellipodia explained by the interaction between actin dynamics and membrane deformation. Journal of Theoretical Biology, 380(??):144–155, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002386.

Sainudiin:2016:TPC

[SW16] Raazesh Sainudiin and David Welch. The transmission process: a combinatorial stochastic process for the evolution of transmission trees over networks. Journal of Theoretical Biology, 410(??):137–170, December 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302211.

Shafiey:2017:ERP

[SW17] Hassan Shafiey and David Waxman. Exact results for the probability and stochastic dynamics of fixation in the Wright– Fisher model. Journal of Theoretical Biology, 430(??):64–77, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730303X.

Sachs:2018:FFI

[SW18] Gottfried Sachs and Henri Weimerskirch. Flight of frigatebirds inside clouds — energy gain, stability and control. Journal of Theoretical Biology, 448(??):9–16, July 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301292.

Stamper:2019:IMM

[SW19] I. Johanna Stamper and Xujing Wang. Integrated multiscale mathematical modeling of insulin secretion reveals the role of islet network integrity for proper oscillatory glucose-dose response. Journal of Theoretical Biology, 475(??):1–24, August 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301924.

Schmitt:2016:NPP

[SWD16] Christoph K. Schmitt, Christian Wildner, and Barbara Drossel. Noisy predator–prey model explains oscillation patterns in sockeye salmon data. Journal of Theoretical Biology, 389(??):198–205, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005287.

Schwen:2015:AGR

[SWG^{+}15] Lars Ole Schwen, Weiwei Wei, Felix Gremse, Josef Ehling, Lei Wang, Uta Dahmen, and Tobias Preusser. Algorithmically generated rodent hepatic vascular trees in arbitrary detail. Journal of Theoretical Biology, 365(??):289–300, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400616X.

Spronck:2016:CSA

[SWPC^{+}16] Bart Spronck, John Walmsley, Georgina Palau-Caballero, Koen D. Reesink, and Tammo Delhaas. Can significance of aortic sinus vortices be assessed using the assumptions in the model of Aboelkassem et al.? Journal of Theoretical Biology, 389(??):304–305, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002672.

Schmitz:2011:NCS

[SWR11] Yvonne Schmitz, Olaf Wolkenhauer, and Katja Rateitschak. Nucleo-cytoplasmic shuttling of APC can maximize βcatenin/TCF concentration. Journal of Theoretical Biology, 279(1):132–142, June 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001640.

Solowiej-Wedderburn:2019:WSE

[SWSLMJ19] J. Solowiej-Wedderburn, D. J. Smith, S. S. Lopes, and T. D. Montenegro-Johnson. Wall stress enhanced exocytosis of extracellular vesicles as a possible mechanism of left-right symmetry-breaking in vertebrate development. Journal of Theoretical Biology, 460(??):220–226, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304855.

Shao:2015:MNC

[SWTO15] Bin Shao, Jiayi Wu, Binghui Tian, and Qi Ouyang. Minimum network constraint on reverse engineering to develop biological regulatory networks. Journal of Theoretical Biology, 380(??):9–15, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002337.

Song:2012:SSM

[SX12] Zigen Song and Jian Xu. Stability switches and multistability coexistence in a delay-coupled neural oscillators system. Journal of Theoretical Biology, 313(??):98–114, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003979.

Shi:2012:PFI

[SXZY12] Jichen Shi, Jian Xu, Xiaorong Zhang, and Ling Yang. Positive feedback induced memory effect in ischemic preconditioning. Journal of Theoretical Biology, 300(??):317–323, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312000434.

Simon:2011:MMD

[SY11] Cory Simon and Niño Yosinao. A mathematical model to distinguish sociological and biological susceptibility factors in disease transmission in the context of H1N1/09 influenza. Journal of Theoretical Biology, 286(??):50–56, October 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311003584.

Suzuki:2012:NRS

[SY12] Kenta Suzuki and Takehito Yoshida. Non-random spatial coupling induces desynchronization, chaos and multistability in a predator–prey-resource system. Journal of Theoretical Biology, 300(??):81–90, May 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006205.

Suzuki:2015:ERP

[SY15] Kenta Suzuki and Takehito Yoshida. Ecological resilience of population cycles: a dynamic perspective of regime shift. Journal of Theoretical Biology, 370(??):103–115, April 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000351.

Smerlak:2017:LFD

[SY17] Matteo Smerlak and Ahmed Youssef. Limiting fitness distributions in evolutionary dynamics. Journal of Theoretical Biology, 416(??):68–80, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300061.

Syi:2014:HPR

[Syi14] Syi. Hidden patterns of reciprocity. Journal of Theoretical Biology, 345(??):92–98, March 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005626.

Sawada:2017:GLT

[SYI17] Kota Sawada, Sachi Yamaguchi, and Yoh Iwasa. Be a good loser: a theoretical model for subordinate decision-making on bi-directional sex change in haremic fishes. Journal of Theoretical Biology, 421(??):127–135, May 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301479.

Smith:2011:EIM

[SYR11] M. L. Smith, G. M. Yanega, and A. Ruina. Elastic instability model of rapid beak closure in hummingbirds. Journal of Theoretical Biology, 282(1):41–51, August 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002499.

Siewe:2017:GFL

[SYSF17] Nourridine Siewe, Abdul-Aziz Yakubu, Abhay R. Satoskar, and Avner Friedman. Granuloma formation in leishmaniasis: a mathematical model. Journal of Theoretical Biology, 412(??):48–60, January 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303356.

Sheffer:2011:MVR

[SYSM11] Efrat Sheffer, Hezi Yizhaq, Moshe Shachak, and Ehud Meron. Mechanisms of vegetation-ring formation in water-limited systems. Journal of Theoretical Biology, 273(1):138–146, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310006831.

Suzuki:2017:CIB

[SYY17a] Kenta Suzuki, Yuji Yamauchi, and Takehito Yoshida. Corrigendum to “Interplay between microbial trait dynamics and population dynamics revealed by the combination of laboratory experiment and computational approaches” [Journal of Theoretical Biology 419 (2017) 201–210]. Journal of Theoretical Biology, 426(??):180, August 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302175.

Suzuki:2017:IBM

[SYY17b] Kenta Suzuki, Yuji Yamauchi, and Takehito Yoshida. Interplay between microbial trait dynamics and population dynamics revealed by the combination of laboratory experiment and computational approaches. Journal of Theoretical Biology, 419(??):201–210, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300693.

Stamatakis:2010:MCA

[SZ10] Michail Stamatakis and Kyriacos Zygourakis. A mathematical and computational approach for integrating the major sources of cell population heterogeneity. Journal of Theoretical Biology, 266(1):41–61, September 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002857.

Shoele:2015:DMB

[SZ15a] Kourosh Shoele and Qiang Zhu. Drafting mechanisms between a dolphin mother and calf. Journal of Theoretical Biology, 382(??):363–377, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003549. See addendum [SZ16].

Si:2015:CEG

[SZ15b] Wen Si and Weinian Zhang. Control exponential growth of tumor cells with slow spread of oncolytic virus. Journal of Theoretical Biology, 367(??):111–129, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006675.

Shoele:2016:ADM

[SZ16] Kourosh Shoele and Qiang Zhu. Addendum to: “Drafting mechanisms between a dolphin mother and calf” [J. Theor. Biol. 382 (2015) 363–377]. Journal of Theoretical Biology, 389(??):310, January 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519315005391. See corrigendum [SZ15a].

Sipahi:2018:SCA

[SZ18] Rifat Sipahi and Günther K. H. Zupanc. Stochastic cellular automata model of neurosphere growth: Roles of proliferative potential, contact inhibition, cell death, and phagocytosis. Journal of Theoretical Biology, 445(??):151–165, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300857.

Szathmary:2015:E

[Sza15a] Eörs Szathmáry. Editorial. Journal of Theoretical Biology, 381(??):1, September 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003653.

Szathmary:2015:FSC

[Sza15b] Eörs Szathmáry. Founder of systems chemistry and foundational theoretical biologist: Tibor Gánti (1933–2009). Journal of Theoretical Biology, 381(??):2–5, September 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315002192.

Takemoto:2010:NSA

[TA10] Kazuhiro Takemoto and Masanori Arita. Nested structure acquired through simple evolutionary process. Journal of Theoretical Biology, 264(3):782–786, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001621.

Tanaka:2013:ESV

[TA13] Gouhei Tanaka and Kazuyuki Aihara. Effects of seasonal variation patterns on recurrent outbreaks in epidemic models. Journal of Theoretical Biology, 317(??):87–95, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005164.

Tomassini:2015:LFC

[TA15] Marco Tomassini and Alberto Antonioni. Lévy flights and cooperation among mobile individuals. Journal of Theoretical Biology, 364(??):154–161, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005529.

Tripuraneni:2016:PMQ

[TA16] Naga Srinivas Tripuraneni and Mohammed Afzal Azam. Pharmacophore modeling, 3D-QSAR and docking study of 2phenylpyrimidine analogues as selective PDE4B inhibitors. Journal of Theoretical Biology, 394(??):117–126, April 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000230.

Tosenberger:2013:MTG

[TAB^{+}13] A. Tosenberger, F. Ataullakhanov, N. Bessonov, M. Panteleev, A. Tokarev, and V. Volpert. Modelling of thrombus growth in flow with a DPD-PDE method. Journal of Theoretical Biology, 337(??):30–41, November 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003469.

Tachikawa:2010:MMP

[Tac10] Masashi Tachikawa. A mathematical model for periodmemorizing behavior in Physarum plasmodium. Journal of Theoretical Biology, 263(4):449–454, April 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000093.

Tarafder:2018:RMP

[TAI^{+}18] Sumit Tarafder, Md. Toukir Ahmed, Sumaiya Iqbal, Md Tamjidul Hoque, and M. Sohel Rahman. RBSURFpred: Modeling protein accessible surface area in real and binary space using regularized and optimized regression. Journal of Theoretical Biology, 441(??):44–57, March 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305763.

Takashina:2016:SRE

[Tak16] Nao Takashina. Simple rules for establishment of effective marine protected areas in an age-structured metapopulation. Journal of Theoretical Biology, 391(??):88–94, February 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315005901.

Takai:2017:TRC

[Tak17a] Kazuyuki Takai. Translational resistivity/conductivity of coding sequences during exponential growth of Escherichia coli. Journal of Theoretical Biology, 413(??):66–71, January 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316303800.

Takesue:2017:PSE

[Tak17b] Hirofumi Takesue. Partner selection and emergence of the merit-based equity norm. Journal of Theoretical Biology, 416(??):45–51, March 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316304374.

Tal:2012:CEG

[Tal12] Omri Tal. The cumulative effect of genetic markers on classification performance: Insights from simple models. Journal of Theoretical Biology, 293(??):206–218, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005121.

Tokuda:2019:RCM

[TAL19] Isao T. Tokuda, Ozgur E. Akman, and James C. W. Locke. Reducing the complexity of mathematical models for the plant circadian clock by distributed delays. Journal of Theoretical Biology, 463(??):155–166, February 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306039.

Tomezak:2016:BMC

[TALC16] M. Tomezak, C. Abbadie, E. Lartigau, and F. Cleri. A biophysical model of cell evolution after cytotoxic treatments: Damage, repair and cell response. Journal of Theoretical Biology, 389(??):146–158, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005160.

Tsaneva-Atanasova:2010:FSB

[TAORS10] Krasimira Tsaneva-Atanasova, Hinke M. Osinga, Thorsten Rieß, and Arthur Sherman. Full system bifurcation analysis of endocrine bursting models. Journal of Theoretical Biology, 264(4):1133–1146, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001633.

Thillainayagam:2016:CMF

[TAR16] Mahalakshmi Thillainayagam, Anand Anbarasu, and Sudha Ramaiah. Comparative molecular field analysis and molecular docking studies on novel aryl chalcone derivatives against an important drug target cysteine protease in Plasmodium falciparum. Journal of Theoretical Biology, 403(??):110–128, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300972.

Tauber:2015:RAO

[Tau15] Alfred I. Tauber. Reconceiving autoimmunity: An overview. Journal of Theoretical Biology, 375(??):52–60, June 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003026.

Taylor:2013:IPF

[Tay13] Peter Taylor. Inclusive and personal fitness in synergistic evolutionary games on graphs. Journal of Theoretical Biology, 325(??):76–82, May 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931300060X.

Taylor:2016:VMD

[Tay16a] Brian K. Taylor. Validating a model for detecting magnetic field intensity using dynamic neural fields. Journal of Theoretical Biology, 408(??):53–65, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302442.

Taylor:2016:HRF

[Tay16b]  Peter Taylor.  Hamilton’s Rule in finite populations

with synergistic interactions. Journal of Theoretical Biology, 397(??):151–157, May 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001302.

Tang:2011:DBC

[TB11] Bin Tang and M. J. Brennan. On the dynamic behaviour of the “click” mechanism in dipteran flight. Journal of Theoretical Biology, 289(??):173–180, November 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004425.

Toor:2015:EHR

[TB15] Jaspreet Toor and Alex Best. The evolution of host resistance to disease in the presence of predators. Journal of Theoretical Biology, 365(??):104–111, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005980.

Takashina:2016:EES

[TB16] Nao Takashina and Marissa L. Baskett. Exploring the effect of the spatial scale of fishery management. Journal of Theoretical Biology, 390(??):14–22, February 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005470.

Theuer:2018:IIA

[TB18] Michał Theuer and Ludek Berec. Impacts of infection avoidance for populations affected by sexually transmitted infections. Journal of Theoretical Biology, 455(??):64–74, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303242.

Teichmann:2014:ATD

[TBA14] Jan Teichmann, Mark Broom, and Eduardo Alonso. The application of temporal difference learning in optimal diet models. Journal of Theoretical Biology, 340(??):11–16, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313004189.

Taylor:2014:FTI

[TBB14] Duncan Taylor, Jo-Anne Bright, and John Buckleton. The ‘factor of two’ issue in mixed DNA profiles. Journal of Theoretical Biology, 363(??):300–306, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004767.

tenBrink:2018:LAC

[tBdR18] Hanna ten Brink and André M. de Roos. Large-amplitude consumer-resource cycles allow for the evolution of ontogenetic niche shifts in consumer life history. Journal of Theoretical Biology, 457(??):237–248, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304156.

Takahashi:2013:ACT

[TBM^{+}13] Daisuke Takahashi, Åke Brännström, Rupert Mazzucco, Atsushi Yamauchi, and Ulf Dieckmann. Abrupt community transitions and cyclic evolutionary dynamics in complex food webs. Journal of Theoretical Biology, 337(??):181–189, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003755.

Thew:2019:MOD

[TBMM19] Johnny Thew, Pamela Burrage, Natalie Medlicott, and Dann Mallet. Modelling optimal delivery of bFGF to chronic wounds using ODEs. Journal of Theoretical Biology, 465(??):109–116, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318306283.

Tani:2014:OCS

[TBQG14] Noriyuki P. Tani, Alan Blatt, David A. Quint, and Ajay Gopinathan. Optimal cooperative searching using purely repulsive interactions. Journal of Theoretical Biology, 361(??):159–164, November 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314004391.

Terry:2011:STM

[TC11] Alan J. Terry and Mark A. J. Chaplain. Spatio-temporal modelling of the NF-κB intracellular signalling pathway: The roles of diffusion, active transport, and cell geometry. Journal of Theoretical Biology, 290(??):7–26, December 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311004395.

Towers:2012:IWS

[TC12] S. Towers and G. Chowell. Impact of weekday social contact patterns on the modeling of influenza transmission, and determination of the influenza latent period. Journal of Theoretical Biology, 312(??):87–95, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200375X.

Tully:2013:CRP

[TCB13] Stephen Tully, Monica Cojocaru, and Chris T. Bauch. Coevolution of risk perception, sexual behaviour, and HIV transmission in an agent-based model. Journal of Theoretical Biology, 337(??):125–132, November 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003883.

Taylor:2014:VRC

[TCH14] Stephanie R. Taylor, Allyson Cheever, and Sarah M. Harmon. Velocity response curves demonstrate the complexity of modeling entrainable clocks. Journal of Theoretical Biology, 363(??):307–317, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005311.

Tealdi:2013:ISC

[TCR13] Stefano Tealdi, Carlo Camporeale, and Luca Ridolfi. Interspecies competition-facilitation in stochastic riparian vegetation dynamics. Journal of Theoretical Biology, 318(??):13–21, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005814.

Taylor:2014:VMV

[TCW14] Bradford P. Taylor, Michael H. Cortez, and Joshua S. Weitz. The virus of my virus is my friend: Ecological effects of virophage with alternative modes of coinfection. Journal of Theoretical Biology, 354(??):124–136, August 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400143X.

Tsourkas:2012:FBO

[TCYY^{+}12] Philippos K. Tsourkas, Das Somkanya C., Paul Yu-Yang, Wanli Liu, Susan K. Pierce, and Subhadip Raychaudhuri. Formation of BCR oligomers provides a mechanism for B cell affinity discrimination. Journal of Theoretical Biology, 307(??):174–182, August 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002342.

Teseo:2017:RFI

[TD17] Serafino Teseo and Francesco Delloro. Reduced foraging investment as an adaptation to patchy food sources: a phasic army ant simulation. Journal of Theoretical Biology, 428(??):48–55, September 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317302813.

Thiel:2018:ISM

[TD18] Tatjana Thiel and Barbara Drossel. Impact of stochastic migration on species diversity in meta-food webs consisting of several patches. Journal of Theoretical Biology, 443(??):147–156, April 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300547.

Tracht:2012:EAU

[TDE12] Samantha M. Tracht, Sara Y. Del Valle, and Brian K. Edwards. Economic analysis of the use of facemasks during pandemic (H1N1) 2009. Journal of Theoretical Biology, 300(??):161–172, May 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000422.

Tanaka:2018:BAM

[TDHC^{+}18] Gouhei Tanaka, Elisa Domínguez-Hüttinger, Panayiotis Christodoulides, Kazuyuki Aihara, and Reiko J. Tanaka. Bifurcation analysis of a mathematical model of atopic dermatitis to determine patient-specific effects of treatments on dynamic phenotypes. Journal of Theoretical Biology, 448(??):66– 79, July 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301620.

Ternent:2015:BFS

[TDKJ15] Lucy Ternent, Rosemary J. Dyson, Anne-Marie Krachler, and Sara Jabbari. Bacterial fitness shapes the population dynamics of antibiotic-resistant and -susceptible bacteria in a model of combined antibiotic and anti-virulence treatment. Journal of Theoretical Biology, 372(??):1–11, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000703.

Trauer:2014:CMM

[TDM14] James M. Trauer, Justin T. Denholm, and Emma S. McBryde. Construction of a mathematical model for tuberculosis transmission in highly endemic regions of the Asia-pacific. Journal of Theoretical Biology, 358(??):74–84, October 7, 2014.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002963.

Tindall:2012:MAC

[TDSM12] Marcus J. Tindall, Louise Dyson, Kieran Smallbone, and Philip K. Maini. Modelling acidosis and the cell cycle in multicellular tumour spheroids. Journal of Theoretical Biology, 298(??):107–115, April 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005716.

Tadrist:2016:LOD

[TDT16] Loïc Tadrist and Baptiste Darbois-Texier. Are leaves optimally designed for self-support? An investigation on giant monocots. Journal of Theoretical Biology, 396(??):125–131, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001144.

Taylor:2016:SLC

[TDW16] Bradford P. Taylor, Jonathan Dushoff, and Joshua S. Weitz. Stochasticity and the limits to confidence when estimating R_{0} of Ebola and other emerging infectious diseases. Journal of Theoretical Biology, 408(??):145–154, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302466.

Teng:2018:GWH

[TDZ^{+}18] Yanbo Teng, Yanjun Ding, Mingming Zhang, Xinren Chen, Xizi Wang, Hang Yu, Chonghui Liu, Hongchao Lv, and Ruijie Zhang. Genome-wide haplotype association study identifies risk genes for non-small cell lung cancer. Journal of Theoretical Biology, 456(??):84–90, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303801.

Tabone:2010:BOF

[TED10] Michaelangelo Tabone, Bard Ermentrout, and Brent Doiron. Balancing organization and flexibility in foraging dynamics.

Journal of Theoretical Biology, 266(3):391–400, October 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310003474.

Teboh-Ewungkem:2012:MFO

[TEW12] Miranda I. Teboh-Ewungkem and Miao Wang. Male fecundity and optimal gametocyte sex ratios for Plasmodium falciparum during incomplete fertilization. Journal of Theoretical Biology, 307(??):183–192, August 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002639.

Teboh-Ewungkem:2010:WVM

[TEY10] Miranda I. Teboh-Ewungkem and Thomas Yuster. A withinvector mathematical model of Plasmodium falciparum and implications of incomplete fertilization on optimal gametocyte sex ratio. Journal of Theoretical Biology, 264(2):273–286, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005840.

Teboh-Ewungkem:2016:EID

[TEY16] Miranda I. Teboh-Ewungkem and Thomas Yuster. Evolutionary implications for the determination of gametocyte sex ratios under fecundity variation for the malaria parasite. Journal of Theoretical Biology, 408(??):260–273, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302491.

Toyoizumi:2014:DSQ

[TF14] Hiroshi Toyoizumi and Jeremy Field. Dynamics of social queues. Journal of Theoretical Biology, 346(??):16–22, April 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005687.

Tseng:2015:CAF

[TF15] Zhijie Jack Tseng and John J. Flynn. Convergence analysis of a finite element skull model of Herpestes javanicus (Carnivora, Mammalia): Implications for robust comparative inferences of biomechanical function. Journal of Theoretical Biology, 365(??):112–148, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005918.

Thornley:2016:BTM

[TF16] John H. M. Thornley and James France. Blue tongue — a modelling examination of fundamentals — seasonality and chaos. Journal of Theoretical Biology, 403(??):17–29, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300753.

Toman:2017:SBS

[TF17] Jan Toman and Jaroslav Flegr. Stability-based sorting: The forgotten process behind (not only) biological evolution. Journal of Theoretical Biology, 435(??):29–41, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304137.

Thornley:2018:EBK

[TF18] John H. M. Thornley and James France. Exploration of bimodal kinetics in marker digesta outflows using compartmental models. Journal of Theoretical Biology, 439(??):226–231, February 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305490.

Taylor:2010:RDI

[TG10] Peter D. Taylor and A. Grafen. Relatedness with different interaction configurations. Journal of Theoretical Biology, 262(3):391–397, February 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004871.

Tam:2018:NLG

[TGB^{+}18] Alexander Tam, J. Edward F. Green, Sanjeeva Balasuriya, Ee Lin Tek, Jennifer M. Gardner, Joanna F. Sundstrom, Vladimir Jiranek, and Benjamin J. Binder. Nutrientlimited growth with non-linear cell diffusion as a mechanism for floral pattern formation in yeast biofilms. Journal of Theoretical Biology, 448(??):122–141, July 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301644.

Telschow:2017:IWM

[TGC^{+}17] Arndt Telschow, Florian Grziwotz, Philip Crain, Takeshi Miki, James W. Mains, George Sugihara, Stephen L. Dobson, and Chih hao Hsieh. Infections of Wolbachia may destabilize mosquito population dynamics. Journal of Theoretical Biology, 428(??):98–105, September 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302242.

Tekwa:2015:LDC

[TGL15] Edward W. Tekwa, Andrew Gonzalez, and Michel Loreau. Local densities connect spatial ecology to game, multilevel selection and inclusive fitness theories of cooperation. Journal of Theoretical Biology, 380(??):414–425, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002957.

Tahir:2019:MTD

[TGLK19] Daniah Tahir, Sylvain Glémin, Martin Lascoux, and Ingemar Kaj. Modeling a trait-dependent diversification process coupled with molecular evolution on a random species tree. Journal of Theoretical Biology, 461(??):189–203, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305083.

Towers:2011:ATP

[TGZF11] S. Towers, K. Vogt Geisse, Y. Zheng, and Z. Feng. Antiviral treatment for pandemic influenza: Assessing potential repercussions using a seasonally forced SIR model. Journal of Theoretical Biology, 289(??):259–268, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004061.

Tupper:2017:ETR

[TH17] Andrew S. Tupper and Paul G. Higgs. Error thresholds for RNA replication in the presence of both point mutations and premature termination errors. Journal of Theoretical Biology, 428(??):34–42, September 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302618.

Tounsi:2010:VLD

[THBM10] Abdelouahed Tounsi, Houari Heireche, Hachemi Benhassaini, and Miloud Missouri. Vibration and length-dependent flexural rigidity of protein microtubules using higher order shear deformation theory. Journal of Theoretical Biology, 266(2):250–255, September 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310003346.

Tektonidis:2011:IIV

[THC^{+}11] Marco Tektonidis, Haralambos Hatzikirou, Arnaud Chauvière, Matthias Simon, Karl Schaller, and Andreas Deutsch. Identification of intrinsic in vitro cellular mechanisms for glioma invasion. Journal of Theoretical Biology, 287(??):131–147, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003626.

Townsend:2010:GSC

[THM10] Sunny E. Townsend, Daniel T. Haydon, and Louise Matthews. On the generality of stability-complexity relationships in Lotka–Volterra ecosystems. Journal of Theoretical Biology, 267(2):243–251, November 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004273.

Thuraisingham:2015:DNS

[Thu15] R. A. Thuraisingham. Dense neuron system interacting with the gravitational potential. Journal of Theoretical Biology, 382(??):279–283, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003331.

Tanaka:2012:CEB

[TI12a] Cinthia Marie Tanaka and Yoh Iwasa. Cultural evolution of a belief controlling human mate choice: Dynamic modeling of the hinoeuma superstition in Japan. Journal of Theoretical Biology, 309(??):20–28, September 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002780.

Thomas:2012:EEE

[TI12b] Neil Thomas and Yasuhiro Imafuku. Effect of elastic energy on the folding of an RNA hairpin. Journal of Theoretical Biology, 312(??):96–104, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003736.

Tachiki:2013:CMS

[TI13] Yuuya Tachiki and Yoh Iwasa. Coevolution of mast seeding in trees and extended diapause of seed predators. Journal of Theoretical Biology, 339(??):129–139, December 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313002555.

Takigawa-Imamura:2015:TGI

[TIMI^{+}15] Hisako Takigawa-Imamura, Ritsuko Morita, Takafumi Iwaki, Takashi Tsuji, and Kenichi Yoshikawa. Tooth germ invagination from cell–cell interaction: Working hypothesis on mechanical instability. Journal of Theoretical Biology, 382(??):284–291, October 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003355.

Tachiki:2010:PCC

[TIS10] Yuuya Tachiki, Yoh Iwasa, and Akiko Satake. Pollinator coupling can induce synchronized flowering in different plant species. Journal of Theoretical Biology, 267(2):153– 163, November 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519310004467.

Tanaka:2011:HET

[TIY^{+}11] Yumi Tanaka, Kei’ichiro Iguchi, Jin Yoshimura, Nariyuki Nakagiri, and Kei ichi Tainaka. Historical effect in the territoriality of ayu fish. Journal of Theoretical Biology, 268(1):98–104, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005230.

Tjorve:2010:HRS

[Tjø10] Even Tjørve. How to resolve the SLOSS debate: Lessons from species-diversity models. Journal of Theoretical Biology, 264(2):604–612, May 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310000767.

Tania:2010:CMC

[TK10a] Nessy Tania and James P. Keener. Calsequestrin mediates changes in spontaneous calcium release profiles. Journal of Theoretical Biology, 265(3):359–376, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002717.

Tugrul:2010:ATR

[TK10b] M. Tugrul and A. Kabakçioğlu. Anomalies in the transcriptional regulatory network of the yeast Saccharomyces cerevisiae. Journal of Theoretical Biology, 263(3):328–336, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930900575X.

Turner:2010:ILB

[TK10c] C. Turner and M. Kohandel. Investigating the link between epithelial-mesenchymal transition and the cancer stem cell phenotype: a mathematical approach. Journal of Theoretical Biology, 265(3):329–335, August 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002675.

Thieffry:2019:PSI

[TK19] Denis Thieffry and Marcelle Kaufman. Prologue to the special issue of JTB dedicated to the memory of René Thomas (1928–2017): a journey through biological circuits, logical puzzles and complex dynamics. Journal of Theoretical Biology, 474(??):42–47, August 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301705.

Takashina:2018:SEA

[TKB^{+}18] Nao Takashina, Buntarou Kusumoto, Maria Beger, Suren Rathnayake, and Hugh P. Possingham. Spatially explicit approach to estimation of total population abundance in field surveys. Journal of Theoretical Biology, 453(??):88–95, September 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318302480.

Tahir:2014:PSL

[TKK14] Muhammad Tahir, Asifullah Khan, and Hüseyin Kaya. Protein subcellular localization in human and hamster cell lines: Employing local ternary patterns of fluorescence microscopy images. Journal of Theoretical Biology, 340(??):85–95, January 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003913.

Takashina:2019:GAS

[TKKE19] Nao Takashina, Buntarou Kusumoto, Yasuhiro Kubota, and Evan P. Economo. A geometric approach to scaling individual distributions to macroecological patterns. Journal of Theoretical Biology, 461(??):170–188, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830506X.

Tiwari:2018:SMB

[TKTB18] Abhishek Kumar Tiwari, Rakesh Kumar, Dharmendra Tripathi, and Subham Badhyal. In silico modeling of bone adaptation to rest-inserted loading: Strain energy density versus fluid flow as stimulus. Journal of Theoretical Biology, 446(??):110–127, June 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301231.

Takaguchi:2015:SCE

[TL15] Taro Takaguchi and Renaud Lambiotte. Sufficient conditions of endemic threshold on metapopulation networks. Journal of Theoretical Biology, 380(??):134–143, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002593.

Tang:2012:LED

[TLCZ12] Changbing Tang, Xiang Li, Lang Cao, and Jingyuan Zhan. The σ law of evolutionary dynamics in community-structured population. Journal of Theoretical Biology, 306(??):1–6, August 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002056.

Tyson:2019:MDB

[TLK19] John J. Tyson, Teeraphan Laomettachit, and Pavel Kraikivski. Modeling the dynamic behavior of biochemical regulatory networks. Journal of Theoretical Biology, 462(??):514–527, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305873.

Taghipoor:2012:MMT

[TLL^{+}12] Masoomeh Taghipoor, Philippe Lescoat, Jean-René Licois, Christine Georgelin, and Guy Barles. Mathematical modeling of transport and degradation of feedstuffs in the small intestine. Journal of Theoretical Biology, 294(??):114–121, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100539X.

Thompson:2014:SDA

[TLM14] C. J. Thompson, T. E. Lee, and M. A. McCarthy. Species distributions and area relationships. Journal of Theoretical Biology, 363(??):129–133, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004597.

Tsuda:2011:WCM

[TLPH11] Akira Tsuda, Fiona E. Laine-Pearson, and Peter E. Hydon. Why chaotic mixing of particles is inevitable in the deep lung. Journal of Theoretical Biology, 286(??):57–66, October 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003432.

Tron:2014:EWT

[TLR14] Stefania Tron, Francesco Laio, and Luca Ridolfi. Effect of water table fluctuations on phreatophytic root distribution. Journal of Theoretical Biology, 360(??):102–108, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314003786.

Thompson:2013:IER

[TLS^{+}13] C. J. Thompson, T. E. Lee, L. Stone, M. A. McCarthy, and M. A. Burgman. Inferring extinction risks from sighting records. Journal of Theoretical Biology, 338(??):16–22, December 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313003974.

Thai:2015:CAZ

[TLT^{+}15] Khac-Minh Thai, Duy-Phong Le, Nguyen-Viet-Khoa Tran, Thi-Thu-Ha Nguyen, Thanh-Dao Tran, and Minh-Tri Le. Computational assay of Zanamivir binding affinity with original and mutant influenza neuraminidase 9 using molecular docking. Journal of Theoretical Biology, 385(??):31– 39, November 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519315004129.

Tilman:2018:RSC

[TLW18] Andrew R. Tilman, Simon Levin, and James R. Watson. Revenue-sharing clubs provide economic insurance and incentives for sustainability in common-pool resource systems. Journal of Theoretical Biology, 454(??):205–214, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318302959.

Tao:2019:EGS

[TLY^{+}19] Jin Tao, Xiaoqing Liu, Siqian Yang, Chaohui Bao, Pingan He, and Qi Dai. An efficient genomic signature ranking method for genomic island prediction from a single genome. Journal of Theoretical Biology, 467(??):142–149, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300712.

Tanabe:2012:ECF

[TM12] Shoma Tanabe and Naoki Masuda. Evolution of cooperation facilitated by reinforcement learning with adaptive aspiration levels. Journal of Theoretical Biology, 293(??):151–160, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005352.

Takashina:2014:EMP

[TM14] Nao Takashina and Akihiko Mougi. Effects of marine protected areas on overfished fishing stocks with multiple stable states. Journal of Theoretical Biology, 341(??):64–70, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004578.

Tachikawa:2015:NCV

[TM15a] Masashi Tachikawa and Atsushi Mochizuki. Nonlinearity in cytoplasm viscosity can generate an essential symmetry breaking in cellular behaviors. Journal of Theoretical Biology, 364(??):260–265, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005621.

Takashina:2015:MSY

[TM15b] Nao Takashina and Akihiko Mougi. Maximum sustainable yields from a spatially-explicit harvest model. Journal of Theoretical Biology, 383(??):87–92, October 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003732.

Tang:2016:DPA

[TM16] Zhe F. Tang and David R. McMillen. Design principles for the analysis and construction of robustly homeostatic biological networks. Journal of Theoretical Biology, 408(??):274–289, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301709.

Tung:2014:CSM

[TMD14] Sudipta Tung, Abhishek Mishra, and Sutirth Dey. A comparison of six methods for stabilizing population dynamics. Journal of Theoretical Biology, 356(??):163–173, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002665.

Tarkhov:2017:SMC

[TMF17] Andrei E. Tarkhov, Leonid I. Menshikov, and Peter O. Fedichev. Strehler–Mildvan correlation is a degenerate manifold of Gompertz fit. Journal of Theoretical Biology, 416(??):180–189, March 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300176.

Trimmer:2012:DNS

[TMHM12] Pete C. Trimmer, John M. McNamara, Alasdair I. Houston, and James A. R. Marshall. Does natural selection favour the Rescorla–Wagner rule? Journal of Theoretical Biology, 302(??):39–52, June 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000872.

Tejero:2010:ELE

[TMM10] Hector Tejero, Arturo Marín, and Francisco Montero. Effect of lethality on the extinction and on the error threshold of quasispecies. Journal of Theoretical Biology, 262(4):733–741, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900486X.

Tripathi:2013:EPD

[TMS13] Sunil Kumar Tripathi, Ravikumar Muttineni, and Sanjeev Kumar Singh. Extra precision docking, free energy calculation and molecular dynamics simulation studies of CDK2 inhibitors. Journal of Theoretical Biology, 334(??):87–100, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002439.

Tosh:2011:WCP

[Tos11] Colin R. Tosh. Which conditions promote negative density dependent selection on prey aggregations? Journal of Theoretical Biology, 281(1):24–30, July 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002141.

Tourigny:2014:GPS

[Tou14] David S. Tourigny. Geometric phase shifts in biological oscillators. Journal of Theoretical Biology, 355(??):239–242, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002379.

Toxvaerd:2017:RPO

[Tox17] Søren Toxvaerd. The role of the peptides at the origin of life. Journal of Theoretical Biology, 429(??):164–169, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303004.

Toxvaerd:2018:SAP

[Tox18] Søren Toxvaerd. The start of the abiogenesis: Preservation of homochirality in proteins as a necessary and sufficient condition for the establishment of the metabolism.

Journal of Theoretical Biology, 451(??):117–121, August 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318302285.

Toxvaerd:2019:PL

[Tox19] Søren Toxvaerd. A prerequisite for life. Journal of Theoretical Biology, 474(??):48–51, August 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301833.

Tozzi:2015:HTO

[Toz15] Arturo Tozzi. How to turn an oscillation in a pink one. Journal of Theoretical Biology, 377(??):117–118, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001940.

Takezawa:2010:RER

[TP10a] Masanori Takezawa and Michael E. Price. Revisiting “The Evolution of Reciprocity in Sizable Groups”: Continuous reciprocity in the repeated n-person prisoner’s dilemma. Journal of Theoretical Biology, 264(2):188–196, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000482.

Tamir:2010:REC

[TP10b] Boaz Tamir and Avner Priel. Reversibility and efficiency in coding protein information. Journal of Theoretical Biology, 267(4):519–525, December 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004996.

Tanaka:2010:REB

[TP10c] Natsuki Tanaka and Garegin A. Papoian. Reverse-engineering of biochemical reaction networks from spatio-temporal correlations of fluorescence fluctuations. Journal of Theoretical Biology, 264(2):490–500, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000974.

Thornley:2014:ANG

[TP14] John H. M. Thornley and Anthony J. Parsons. Allocation of new growth between shoot, root and mycorrhiza in relation to carbon, nitrogen and phosphate supply: Teleonomy with maximum growth rate. Journal of Theoretical Biology, 342(??):1–14, February 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004955.

Tripathi:2017:NAF

[TP17] Pooja Tripathi and Paras N. Pandey. A novel alignment-free method to classify protein folding types by combining spectral graph clustering with Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 424(??):49–54, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301972.

Traulsen:2012:GGC

[TR12] Arne Traulsen and Floyd A. Reed. From genes to games: Cooperation and cyclic dominance in meiotic drive. Journal of Theoretical Biology, 299(??):120–125, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002335.

Thouzeau:2017:EMM

[TR17] Valentin Thouzeau and Michel Raymond. Emergence and maintenance of menopause in humans: a game theory model. Journal of Theoretical Biology, 430(??):229–236, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317303508.

Trassinelli:2016:ECO

[Tra16] M. Trassinelli. Energy cost and optimisation in breath-hold diving. Journal of Theoretical Biology, 396(??):42–52, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000990.

Trenchard:2019:CPM

[Tre19] Hugh Trenchard. Cell pelotons: a model of early evolutionary cell sorting, with application to slime mold Dictyostelium discoideum. Journal of Theoretical Biology, 469(??):75–95, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300748.

Tromeur:2016:IDS

[TRG16] Eric Tromeur, Lars Rudolf, and Thilo Gross. Impact of dispersal on the stability of metapopulations. Journal of Theoretical Biology, 392(??):1–11, March 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005949.

Trifonov:2010:BPS

[Tri10] Edward N. Trifonov. Base pair stacking in nucleosome DNA and bendability sequence pattern. Journal of Theoretical Biology, 263(3):337–339, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005621.

Tung:2019:CIR

[TRJD19] Sudipta Tung, M. Rajamani, Amitabh Joshi, and Sutirth Dey. Complex interaction of resource availability, lifehistory and demography determines the dynamics and stability of stage-structured populations. Journal of Theoretical Biology, 460(??):1–12, January 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304880.

Terranova:2014:MMG

[TRM^{+}14] N. Terranova, P. Rebuzzini, G. Mazzini, E. Borella, C. A. Redi, M. Zuccotti, S. Garagna, and P. Magni. Mathematical modeling of growth and death dynamics of mouse embryonic stem cells irradiated with γ-rays. Journal of Theoretical Biology, 363(??):374–380, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400527X.

Tehrani:2017:FDG

[TS17a] Mohammad Tehrani and Alireza S. Sarvestani. Forcedriven growth of intercellular junctions. Journal of Theoretical Biology, 421(??):101–111, May 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301480.

Tulub:2017:HSD

[TS17b] Alexander A. Tulub and Vassily E. Stefanov. Hidden symmetries of DNA molecule. Journal of Theoretical Biology, 416(??):144–148, March 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300012.

Till:2010:MAI

[TSB10] Olaf Till, Tobias Siebert, and Reinhard Blickhan. A mechanism accounting for independence on starting length of tension increase in ramp stretches of active skeletal muscle at short half-sarcomere lengths. Journal of Theoretical Biology, 266(1):117–123, September 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003115.

Tiedemann:2017:MCO

[TSB^{+}17] Hendrik B. Tiedemann, Elida Schneltzer, Johannes Beckers, Gerhard K. H. Przemeck, and Martin Hrabe de Angelis. Modeling coexistence of oscillation and delta/notchmediated lateral inhibition in pancreas development and neurogenesis. Journal of Theoretical Biology, 430(??):32–44, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302771.

Tadrist:2014:WGM

[TSdL14] Loïc Tadrist, Marc Saudreau, and Emmanuel de Langre. Wind and gravity mechanical effects on leaf inclination angles. Journal of Theoretical Biology, 341(??):9–16, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313004554.

Trichilo:2019:CMD

[TSF^{+}19] Silvia Trichilo, Stefan Scheiner, Mark Forwood, David M. L. Cooper, and Peter Pivonka. Computational model of the dual action of PTH — application to a rat model of osteoporosis. Journal of Theoretical Biology, 473(??):67–79, July 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319301699.

Tavoni:2012:SCS

[TSL12] Alessandro Tavoni, Maja Schlüter, and Simon Levin. The survival of the conformist: Social pressure and renewable resource management. Journal of Theoretical Biology, 299(??):152–161, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100347X.

Tanabe:2013:CIR

[TSM13a] Shoma Tanabe, Hideyuki Suzuki, and Naoki Masuda. Corrigendum to “Indirect Reciprocity with Trinary Reputations” [J. Theor. Biol. 317 (2013) 338–347]. Journal of Theoretical Biology, 328(??):100–101, July 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000830. See [TSM13b].

Tanabe:2013:IRT

[TSM13b] Shoma Tanabe, Hideyuki Suzuki, and Naoki Masuda. Indirect reciprocity with trinary reputations. Journal of Theoretical Biology, 317(??):338–347, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519312005711. See corrigendum [TSM13a].

Treloar:2014:VEC

[TSMB14] Katrina K. Treloar, Matthew J. Simpson, D. L. Sean McElwain, and Ruth E. Baker. Are in vitro estimates of cell diffusivity and cell proliferation rate sensitive to assay geometry? Journal of Theoretical Biology, 356(??):71–84, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400246X.

Terentyeva:2015:KMS

[TSP15] Vladislava A. Terentyeva, Anastasia N. Sveshnikova, and Mikhail A. Panteleev. Kinetics and mechanisms of surfacedependent coagulation factor XII activation. Journal of Theoretical Biology, 382(??):235–243, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003306.

Thompson:2016:DMM

[TSS16] Colin J. Thompson, Elad Shtilerman, and Lewi Stone. A discrete Markov metapopulation model for persistence and extinction of species. Journal of Theoretical Biology, 404(??):391–397, September 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301448.

Thomas:2013:SNS

[TST^{+}13] Philipp Thomas, Arthur V. Straube, Jens Timmer, Christian Fleck, and Ramon Grima. Signatures of nonlinearity in single cell noise-induced oscillations. Journal of Theoretical Biology, 335(??):222–234, October 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313002932.

Tsuruyama:2019:NET

[Tsu19] Tatsuaki Tsuruyama. Non-equilibrium thermodynamics of biological signal transduction predicts conservation of entropy production rate. Journal of Theoretical Biology, 472(??):84– 87, July 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301493.

Tjorve:2010:UAR

[TT10a] Even Tjørve and Kathleen M. C. Tjørve. A unified approach to the Richards-model family for use in growth analyses: Why we need only two model forms. Journal of Theoretical Biology, 267(3):417–425, December 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004741.

Tyvand:2010:WFM

[TT10b] Peder A. Tyvand and Steinar Thorvaldsen. Wright–Fisher model of social insects with haploid males and diploid females. Journal of Theoretical Biology, 266(3):470–478, October 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003486.

Toropova:2017:NQC

[TT17] Alla P. Toropova and Andrey A. Toropov. Nano-QSAR in cell biology: Model of cell viability as a mathematical function of available eclectic data. Journal of Theoretical Biology, 416(??):113–118, March 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300139.

Tamura:2019:MEE

[TT19] Kohei Tamura and Hiroki Takikawa. Modelling the emergence of an egalitarian society in the n-player game framework. Journal of Theoretical Biology, 461(??):1–7, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305198.

Terranova:2018:MTG

[TTB^{+}18] N. Terranova, E. M. Tosca, E. Borella, E. Pesenti, M. Rocchetti, and P. Magni. Modeling tumor growth inhibition and toxicity outcome after administration of anticancer agents in xenograft mice: a Dynamic Energy Budget (DEB) approach. Journal of Theoretical Biology, 450(??):1–14, August 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301723.

Tang:2010:OTI

[TTC10] Sanyi Tang, Guangyao Tang, and Robert A. Cheke. Optimum timing for integrated pest management: Modelling rates of pesticide application and natural enemy releases. Journal of Theoretical Biology, 264(2):623–638, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001098.

Tahir:2019:IPI

[TTC19] Muhammad Tahir, Hilal Tayara, and Kil To Chong. iRNA-PseKNC(2methyl): Identify RNA 2’-o-methylation sites by convolution neural network and Chou’s pseudo components. Journal of Theoretical Biology, 465(??):1–6, March 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318306349.

Taheriyan:2019:PNM

[TTG19] Fatemeh Taheriyan, Mohammad Teshnehlab, and Shahriar Gharibzadeh. Presenting a neuroid model of wind-up based on dynamic synapse. Journal of Theoretical Biology, 465(??):45–50, March 21, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319300189.

Tarnita:2013:ECS

[TTN13] Corina E. Tarnita, Clifford H. Taubes, and Martin A. Nowak. Evolutionary construction by staying together and coming together. Journal of Theoretical Biology, 320(??):10–22, March 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005978.

Tal:2017:TST

[TTP17] Omri Tal, Tat Dat Tran, and Jacobus Portegies. From typical sequences to typical genotypes. Journal of Theoretical Biology, 419(??):159–183, April 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300656.

Takagi:2013:HES

[TTW13] Tsutomu Takagi, Yumiko Tamura, and Daniel Weihs. Hydrodynamics and energy-saving swimming techniques of Pacific bluefin tuna. Journal of Theoretical Biology, 336(??):158–172, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931300341X.

Tung:2013:PPS

[Tun13] Chun-Wei Tung. Prediction of pupylation sites using the composition of k-spaced amino acid pairs. Journal of Theoretical Biology, 336(??):11–17, November 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003305.

Tutone:2018:RSI

[TVA18] Marco Tutone, Alessia Virzì, and Anna Maria Almerico. Reverse screening on indicaxanthin from Opuntia ficus-indica as natural chemoactive and chemopreventive agent. Journal of Theoretical Biology, 455(??):147–160, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303412.

Taghipoor:2016:SAE

[TvMG16] Masoomeh Taghipoor, Jaap van Milgen, and Florence Gondret. A systemic approach to explore the flexibility of energy stores at the cellular scale: Examples from muscle cells. Journal of Theoretical Biology, 404(??):331–341, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316301485.

Thurber:2012:MCM

[TW12] Greg M. Thurber and K. Dane Wittrup. A mechanistic compartmental model for total antibody uptake in tumors. Journal of Theoretical Biology, 314(??):57–68, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004560.

Taskinen:2013:RTO

[TW13] S. Taskinen and D. I. Warton. Robust tests for one or more allometric lines. Journal of Theoretical Biology, 333(??):38–46, September 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002257.

Tudge:2016:GTT

[TWB16] S. J. Tudge, R. A. Watson, and M. Brede. Game theoretic treatments for the differentiation of functional roles in the transition to multicellularity. Journal of Theoretical Biology, 395(??):161–173, April 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000825.

Tian:2019:PPP

[TWC^{+}19] Baoguang Tian, Xue Wu, Cheng Chen, Wenying Qiu, Qin Ma, and Bin Yu. Predicting protein–protein interactions by fusing various Chou’s pseudo components and using wavelet denoising approach. Journal of Theoretical Biology, 462(??):329–346, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305642.

Thomaseth:2013:MSS

[TWH^{+}13] Caterina Thomaseth, Patrick Weber, Thomas Hamm, Kenji Kashima, and Nicole Radde. Modeling sphingomyelin synthase 1 driven reaction at the Golgi apparatus can explain data by inclusion of a positive feedback mechanism. Journal of Theoretical Biology, 337(??):174–180, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003962.

Turner:2016:LGM

[TWP16] Rebecca M. Turner, Michael M. Walker, and Claire M. Postlethwaite. Literal grid map models for animal navigation: Assumptions and predictions. Journal of Theoretical Biology, 404(??):169–181, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630128X.

Tahir:2018:IMD

[TWR^{+}18] Rana Adnan Tahir, Hao Wu, Muhammad Ahmad Rizwan, Tassadaq Hussain Jafar, Shahzad Saleem, and Sheikh Arslan Sehgal. Immunoinformatics and molecular docking studies reveal potential epitope-based peptide vaccine against DENV-NS3 protein. Journal of Theoretical Biology, 459(??):162–170, December 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304739.

Taylor:2015:SFH

[TWS15] Rachel A. Taylor, Andrew White, and Jonathan A. Sherratt. Seasonal forcing in a host-macroparasite system. Journal of Theoretical Biology, 365(??):55–66, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005967.

Tait:2018:CCA

[TWTA^{+}18] Luke Tait, Kyle Wedgwood, Krasimira Tsaneva-Atanasova, Jon T. Brown, and Marc Goodfellow. Control of clustered action potential firing in a mathematical model of entorhinal cortex stellate cells. Journal of Theoretical Biology, 449(??):23– 34, July 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301735.

Tan:2019:PAT

[TWW19] Chengguan Tan, Yuanshi Wang, and Hong Wu. Population abundance of a two-patch chemostat system with asymmetric diffusion. Journal of Theoretical Biology, 474(??):1–13, August 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301766.

Tang:2015:PVI

[TXCW15] Biao Tang, Yanni Xiao, Robert A. Cheke, and Ning Wang. Piecewise virus-immune dynamic model with HIV-1 RNAguided therapy. Journal of Theoretical Biology, 377(??):36– 46, July 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001770.

Tang:2016:MWV

[TXTW16] Biao Tang, Yanni Xiao, Sanyi Tang, and Jianhong Wu. Modelling weekly vector control against dengue in the Guangdong Province of China. Journal of Theoretical Biology, 410(??):65–76, December 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316303009.

Tang:2012:PHV

[TXWW12] Sanyi Tang, Yanni Xiao, Ning Wang, and Hulin Wu. Piecewise HIV virus dynamic model with CD4 + T cell count-guided therapy: I. Journal of Theoretical Biology, 308(??):123–134, September 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002640.

Tang:2012:CQF

[TXY^{+}12] Sanyi Tang, Yanni Xiao, Lin Yuan, Robert A. Cheke, and Jianhong Wu. Campus quarantine (Fengxiao) for curbing emergent infectious diseases: Lessons from mitigating A/H1N1 in Xi’an, China. Journal of Theoretical Biology, 295(??):47–58, February 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311005583.

Tang:2011:IND

[TZ11] Junfeng Tang and Shurong Zhou. The importance of niche differentiation for coexistence on large scales. Journal of Theoretical Biology, 273(1):32–36, March 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006806.

Tang:2013:HNN

[TZ13] Junfeng Tang and Shurong Zhou. Hybrid niche-neutral models outperform an otherwise equivalent neutral model for fitting coral reef data. Journal of Theoretical Biology, 317(??):212–218, January 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312005528.

Tian:2018:CHA

[TZY18] Kun Tian, Xin Zhao, and Stephen S.-T. Yau. Convex hull analysis of evolutionary and phylogenetic relationships between biological groups. Journal of Theoretical Biology, 456(??):34–40, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303655.

Uz:2010:MLE

[UAL10] Bilge Uz, Erdem Arslan, and Ian J. Laurenzi. Maximum likelihood estimation of the kinetics of receptor-mediated adhesion. Journal of Theoretical Biology, 262(3):478–487, February 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004901.

Ullah:2012:SRM

[UBP12] Ghanim Ullah, William J. Bruno, and John E. Pearson. Simplification of reversible Markov chains by removal of states with low equilibrium occupancy. Journal of Theoretical Biology, 311(??):117–129, October 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003414.

Ud-Dean:2010:SEE

[UD10] S. M. Minhaz Ud-Dean. Structural explanation for the effect of humidity on persistence of airborne virus: Seasonality of influenza. Journal of Theoretical Biology, 264(3):822–829, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001360.

Uehara:2010:GML

[UI10] Takashi Uehara and Yoh Iwasa. Global mutations and local mutations have very different effects on evolution, illustrated by mixed strategies of asymmetric binary games. Journal of Theoretical Biology, 262(2):223–231, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251930900455X.

Uchinomiya:2013:ESS

[UI13] Kouki Uchinomiya and Yoh Iwasa. Evolution of stalk/spore ratio in a social amoeba: Cell-to-cell interaction via a signaling chemical shaped by cheating risk. Journal of Theoretical Biology, 336(??):110–118, November 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313003470.

Uitdehaag:2011:BHB

[Uit11] Joost C. M. Uitdehaag. Bet hedging based cooperation can limit kin selection and form a basis for mutualism. Journal of Theoretical Biology, 280(1):76–87, July 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001925.

Utz:2014:PPI

[UJLG14] Margarete Utz, Jonathan M. Jeschke, Volker Loeschcke, and Wilfried Gabriel. Phenotypic plasticity with instantaneous but delayed switches. Journal of Theoretical Biology, 340(??):60–72, January 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313004207.

UlHasan:2017:MCM

[UK17] A. B. M. Shamim Ul Hasan and Hiroyuki Kurata. Mathematical comparison of memory functions between mutual activation and repression networks in a stochastic environment. Journal of Theoretical Biology, 427(??):28–40, August 1, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317302588.

Unckless:2011:PRX

[Unc11] Robert L. Unckless. The potential role of the X chromosome in the emergence of male-killing from mutualistic endosymbionts. Journal of Theoretical Biology, 291(??):99–104, December 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004632.

Udiani:2015:IRR

[UPWK15] Oyita Udiani, Noa Pinter-Wollman, and Yun Kang. Identifying robustness in the regulation of collective foraging of ant colonies using an interaction-based model with backward bifurcation. Journal of Theoretical Biology, 367(??):61–75, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400678X.

Ursino:2014:EHD

[Urs14] Nadia Ursino. Eco-hydrology driven fire regime in savanna. Journal of Theoretical Biology, 355(??):68–76, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314002112.

Uchida:2010:CAR

[US10] Satoshi Uchida and Karl Sigmund. The competition of assessment rules for indirect reciprocity. Journal of Theoretical Biology, 263(1):13–19, March 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005554.

Uziel:2012:DPS

[US12] Asher Uziel and Lewi Stone. Determinants of periodicity in seasonally driven epidemics. Journal of Theoretical Biology, 305(??):88–95, July 21, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312001166.

Uddin:2018:ESA

[USF^{+}18] Md. Raihan Uddin, Alok Sharma, Dewan Md Farid, Md. Mahmudur Rahman, Abdollah Dehzangi, and Swakkhar Shatabda.

EvoStruct-sub: An accurate Gram-positive protein subcellular localization predictor using evolutionary and structural features. Journal of Theoretical Biology, 443(??):138–146, April 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300559.

Usherwood:2016:PAG

[Ush16] James Richard Usherwood. Physiological, aerodynamic and geometric constraints of flapping account for bird gaits, and bounding and flap-gliding flight strategies. Journal of Theoretical Biology, 408(??):42–52, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301722.

Uriu:2017:FLI

[UT17] Koichiro Uriu and Hajime Tei. Feedback loops interlocked at competitive binding sites amplify and facilitate genetic oscillations. Journal of Theoretical Biology, 428(??):56–64, September 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302758.

Verdy:2010:ASS

[VA10] Ariane Verdy and Priyanga Amarasekare. Alternative stable states in communities with intraguild predation. Journal of Theoretical Biology, 262(1):116–128, January 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004317.

Voinson:2018:SDE

[VABS18] Marina Voinson, Alexandra Alvergne, Sylvain Billiard, and Charline Smadi. Stochastic dynamics of an epidemic with recurrent spillovers from an endemic reservoir. Journal of Theoretical Biology, 457(??):37–50, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303965.

Valencia:2017:AMM

[VACGF17] Pedro L. Valencia, Carolina Astudillo-Castro, Diego Gajardo, and Sebastián Flores. Application of the median method to estimate the kinetic constants of the substrate uncompetitive inhibition equation. Journal of Theoretical Biology, 418(??):122– 128, April 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300346.

Vermolen:2016:MIS

[VAG16] F. J. Vermolen, E. C. M. M. Arkesteijn, and A. Gefen. Modelling the immune system response to epithelial wound infections. Journal of Theoretical Biology, 393(??):158–169, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316000370.

Vandiver:2015:BIA

[Van15] Rebecca M. Vandiver. Buckling instability in arteries. Journal of Theoretical Biology, 371(??):1–8, April 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000569.

VanCleve:2016:CCC

[Van16a] Jeremy Van Cleve. Cooperation, conformity, and the coevolutionary problem of trait associations. Journal of Theoretical Biology, 396(??):13–24, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001028.

VanOosterwyck:2016:COC

[Van16b] Hans Van Oosterwyck. Corrigendum to: “Oxygen as a critical determinant of bone fracture healing — a multiscale model” [J. Theor. Biol. 365 (2015) 247–264]. Journal of Theoretical Biology, 392(??):124, March 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005494. See [CGvG^{+}15].

VanDyken:2017:NSR

[Van17a] J. David Van Dyken. Noise slows the rate of Michaelis–Menten reactions. Journal of Theoretical Biology, 430(??):21–31, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317303259.

VanDyken:2017:PCG

[Van17b] J. David Van Dyken. Propagation and control of gene expression noise with non-linear translation kinetics. Journal of Theoretical Biology, 430(??):185–194, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303314.

Veprauskas:2018:EER

[VAT18] Amy Veprauskas, Azmy S. Ackleh, and Tingting Tang. Examining the effect of reoccurring disturbances on population persistence with application to marine mammals. Journal of Theoretical Biology, 455(??):109–117, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303357.

Vazquez:2010:OCD

[Vaz10] Alexei Vazquez. Optimal cytoplasmatic density and flux balance model under macromolecular crowding effects. Journal of Theoretical Biology, 264(2):356–359, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000998.

Viertel:2019:CMM

[VB19a] Ryan Viertel and Alla Borisyuk. A computational model of the mammalian external tufted cell. Journal of Theoretical Biology, 462(??):109–121, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304752.

Vitale:2019:PTP

[VB19b] Caterina Vitale and Alex Best. The paradox of tolerance: Parasite extinction due to the evolution of host defence. Journal of Theoretical Biology, 474(??):78–87, August 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301742.

Voss-Bohme:2010:CBC

[VBD10] A. Voß-Böhme and A. Deutsch. The cellular basis of cell sorting kinetics. Journal of Theoretical Biology, 263(4):419–436, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005785.

Velasco:2013:LCR

[VBHM^{+}13] Ana María Velasco, Arturo Becerra, Ricardo Hernández-Morales, Luis Delaye, María Eugenia Jiménez-Corona, Samuel Ponce de Leon, and Antonio Lazcano. Low complexity regions (LCRs) contribute to the hypervariability of the HIV-1 gp120 protein. Journal of Theoretical Biology, 338(??):80–86, December 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004219.

VanDerlinden:2010:SVS

[VBV10] E. Van Derlinden, K. Bernaerts, and J. F. Van Impe. Simultaneous versus sequential optimal experiment design for the identification of multi-parameter microbial growth kinetics as a function of temperature. Journal of Theoretical Biology, 264(2):347–355, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000007X.

Vidal-DiezdeUlzurrun:2017:MTD

[VBVD17] G. Vidal-Diez de Ulzurrun, J. M. Baetens, J. Van den Bulcke, and B. De Baets. Modelling three-dimensional fungal growth in response to environmental stimuli. Journal of Theoretical Biology, 414(??):35–49, February 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630385X.

Vickerstaff:2010:WCS

[VC10] Robert J. Vickerstaff and Allen Cheung. Which coordinate system for modelling path integration? Journal of Theoretical Biology, 263(2):242–261, March 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005633.

Veliz-Cuba:2011:RBN

[VC11] Alan Veliz-Cuba. Reduction of Boolean network models. Journal of Theoretical Biology, 289(??):167–172, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311004450.

Vaz:2015:OI

[VC15] Nelson M. Vaz and Claudia R. Carvalho. On the origin of immunopathology. Journal of Theoretical Biology, 375(??):61– 70, June 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003439.

Venturino:2019:EMS

[VCF^{+}19] Ezio Venturino, Simona Collino, Luca Ferreri, Luigi Bertolotti, Sergio Rosati, and Mario Giacobini. An effective management strategy for the control of two lentiviruses in goat breedings. Journal of Theoretical Biology, 469(??):96–106, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300876.

vandenBerg:2010:ERD

[vdBD10] Hugo A. van den Berg and Zoe A. Duncombe. Eradicationresolution dynamics with stochastic flare-ups. Journal of Theoretical Biology, 264(3):962–970, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001335.

Vaughan:2012:WHD

[VDD12] T. G. Vaughan, P. D. Drummond, and A. J. Drummond. Within-host demographic fluctuations and correlations in early retroviral infection. Journal of Theoretical Biology, 295(??):86–99, February 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005868.

Viger:2017:SAG

[VDD^{+}17] Louise Viger, Fabrice Denis, Clément Draghi, Thibault Ménard, and Christophe Letellier. Spatial avascular growth of tumor in a homogeneous environment. Journal of Theoretical Biology, 416(??):99–112, March 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316304210.

VanWassenbergh:2015:SPO

[VDH^{+}15] Sam Van Wassenbergh, Steven W. Day, L. Patricia Hernández, Timothy E. Higham, and Tyler Skorczewski. Suction power output and the inertial cost of rotating the neurocranium to generate suction in fish. Journal of Theoretical Biology, 372(??):159–167, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315001022.

Viger:2014:CMA

[VDRL14] Louise Viger, Fabrice Denis, Martin Rosalie, and Christophe Letellier. A cancer model for the angiogenic switch. Journal of Theoretical Biology, 360(??):21–33, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003634.

vanDoorn:2014:CVR

[vDRT14] G. Sander van Doorn, Thomas Riebli, and Michael Taborsky. Coaction versus reciprocity in continuous-time models of cooperation. Journal of Theoretical Biology, 356(??):1–10, September 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001544.

vanderSman:2012:PPF

[vdSS12] R. G. M. van der Sman and M. Sanders. Prediction of postharvest firmness of apple using biological switch model. Journal of Theoretical Biology, 310(??):239–248, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200330X.

Vagne:2015:RSE

[VDTF15] Constance Vagne, Jacques David, Muriel Tavaud, and Bénédicte Fontez. Reciprocal sign epistasis and truncation selection: When is recombination favorable in a prebreeding program with a selfing species? Journal of Theoretical Biology, 386(??):44–54, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004063.

vanDrongelen:2018:MET

[vDVFH^{+}18] Ruben van Drongelen, Tania Vazquez-Faci, Teun A. P. M. Huijben, Maurijn van der Zee, and Timon Idema. Mechanics of epithelial tissue formation. Journal of Theoretical Biology, 454(??):182–189, October 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302935.

vanElburg:2011:SCT

[vE11] Ronald A. J. van Elburg. Stochastic continuous time neurite branching models with tree and segment dependent rates. Journal of Theoretical Biology, 276(1):159–173, May 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000658.

Veglio:2010:SAS

[Veg10] Andrea Veglio. The shade avoidance syndrome: a non-Markovian stochastic growth model. Journal of Theoretical Biology, 264(3):722–728, June 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001141.

Veloso:2017:DSR

[Vel17] Felipe A. Veloso. On the developmental self-regulatory dynamics and evolution of individuated multicellular organisms. Journal of Theoretical Biology, 417(??):84–99, March 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316304349.

Vahidi:2012:BSU

[VF12] Bahman Vahidi and Nasser Fatouraee. A biomechanical simulation of ureteral flow during peristalsis using intraluminal morphometric data. Journal of Theoretical Biology, 298(??):42–50, April 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006436.

Vasas:2015:PEI

[VFS^{+}15] Vera Vasas, Chrisantha Fernando, András Szilágyi, István Zachár, Mauro Santos, and Eörs Szathmáry. Primordial evolvability: Impasses and challenges. Journal of Theoretical Biology, 381(??):29–38, September 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003264.

Vermolen:2013:SSC

[VG13] F. J. Vermolen and A. Gefen. A semi-stochastic cellbased model for in vitro infected ‘wound’ healing through motility reduction: a simulation study. Journal of Theoretical Biology, 318(??):68–80, February 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005826.

Vessman:2016:EPC

[VGL16] Björn Vessman, Philip Gerlee, and Torbjörn Lundh. Estimating the probability of coexistence in cross-feeding communities. Journal of Theoretical Biology, 408(??):13–21, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302260.

Valderrama-Gomez:2018:PCM

[VGPS18] Miguel A. Valderrama-Gómez, Rebecca E. Parales, and Michael A. Savageau. Phenotype-centric modeling for elucidation of biological design principles. Journal of Theoretical Biology, 455(??):281–292, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303321.

Vargas-Garcia:2018:STM

[VGZS18] Cesar Vargas-Garcia, Ryan Zurakowski, and Abhyudai Singh. Synaptic transmission may provide an evolutionary benefit to HIV through modulation of latency. Journal of Theoretical Biology, 455(??):261–268, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303540.

Vasalou:2011:MMD

[VH11a] Christina Vasalou and Michael A. Henson. A multicellular model for differential regulation of circadian signals in the core and shell regions of the suprachiasmatic nucleus. Journal of Theoretical Biology, 288(??):44–56, November 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931100405X.

Vuong:2011:BFM

[VH11b] Jenny Vuong and Christian Hellmich. Bone fibrillogenesis and mineralization: Quantitative analysis and implications for tissue elasticity. Journal of Theoretical Biology, 287(??):115–130, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100378X.

vanHeijster:2014:CMB

[vHHKB14] Peter van Heijster, Heather Hardway, Tasso J. Kaper, and Cynthia A. Bradham. A computational model for BMP movement in sea urchin embryos. Journal of Theoretical Biology, 363(??):277–289, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004871.

Vargason:2017:MMM

[VHM^{+}17] Troy Vargason, Daniel P. Howsmon, Stepan Melnyk, S. Jill James, and Juergen Hahn. Mathematical modeling of the methionine cycle and transsulfuration pathway in individuals with autism spectrum disorder. Journal of Theoretical Biology, 416(??):28–37, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316304301.

Venkatesh:2017:TTR

[VHS17] Thejaswini Venkatesh, Shaharabhanu A. Hussain, and Padmanaban S. Suresh. A tale of three RNAs in mitochondria: tRNA, tRNA derived fragments and mitomiRs. Journal of Theoretical Biology, 435(??):42–49, December 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304149.

Vergni:2012:IHH

[VIBC12] Davide Vergni, Sandro Iannaccone, Stefano Berti, and Massimo Cencini. Invasions in heterogeneous habitats in the presence of advection. Journal of Theoretical Biology, 301(??):141– 152, May 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000914.

vanIersel:2011:WTT

[vIK11] Leo van Iersel and Steven Kelk. When two trees go to war. Journal of Theoretical Biology, 269(1):245–255, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005771.

Vandermeer:2010:CCR

[VK10] John Vandermeer and Aaron King. Consequential classes of resources: Subtle global bifurcation with dramatic ecological consequences in a simple population model. Journal of Theoretical Biology, 263(2):237–241, March 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005736.

Vainstein:2012:SCS

[VKKA12] Vladimir Vainstein, Oleg U. Kirnasovsky, Yuri Kogan, and Zvia Agur. Strategies for cancer stem cell elimination: Insights from mathematical modeling. Journal of Theoretical Biology, 298(??):32–41, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006333.

Vallina:2011:SCF

[VL11] Sergio M. Vallina and Corinne Le Quéré. Stability of complex food webs: Resilience, resistance and the average interaction strength. Journal of Theoretical Biology, 272(1):160–173, March 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006387.

vanLeeuwen:2013:GFR

[vLBJ^{+}13] E. van Leeuwen, Å. Brännström, V. A. A. Jansen, U. Dieckmann, and A. G. Rossberg. A generalized functional response for predators that switch between multiple prey species. Journal of Theoretical Biology, 328(??):89–98, July 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000611.

Vazquez:2010:DPC

[VLCM10] Federico Vazquez, Cristóbal López, Justin M. Calabrese, and Miguel A. Muñoz. Dynamical phase coexistence: a simple solution to the “savanna problem”. Journal of Theoretical Biology, 264(2):360–366, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000780.

Veasy:2019:CPS

[VLCT19] Joshua Veasy, Yi Ming Lai, Stephen Coombes, and Rüdiger Thul. Complex patterns of subcellular cardiac alternans. Journal of Theoretical Biology, 478(??):102–114, October 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319302516.

Vazquez:2012:ENL

[VLFF12] José Antonio Vázquez, José M. Lorenzo, Pablo Fuciños, and Daniel Franco. Evaluation of non-linear equations to model different animal growths with mono and bisigmoid profiles. Journal of Theoretical Biology, 314(??):95–105, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004493.

vanLingen:2019:BMM

[vLFM^{+}19] Henk J. van Lingen, James G. Fadel, Luis E. Moraes, André Bannink, and Jan Dijkstra. Bayesian mechanistic modeling of thermodynamically controlled volatile fatty acid, hydrogen and methane production in the bovine rumen. Journal of Theoretical Biology, 480(??):150–165, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303200.

Veltman:2017:STA

[VLPH17] Denise Veltman, Thessa Laeremans, Egle Passante, and Heinrich J. Huber. Signal transduction analysis of the NLRP3inflammasome pathway after cellular damage and its paracrine regulation. Journal of Theoretical Biology, 415(??):125–136, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931630426X.

vanLoon:2011:USB

[vLSBBD11] E. E. van Loon, J. Shamoun-Baranes, W. Bouten, and S. L. Davis. Understanding soaring bird migration through interactions and decisions at the individual level. Journal of Theoretical Biology, 270(1):112–126, February 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005850.

Vickerstaff:2012:PIM

[VM12] Robert J. Vickerstaff and Tobias Merkle. Path integration mediated systematic search: a Bayesian model. Journal of Theoretical Biology, 307(??):1–19, August 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312002159.

Voets:2016:SCT

[VM16] Philip J. G. M. Voets and Roderick P. P. W. M. Maas. Serum cardiac troponin I analysis to determine the excessiveness of exercise intensity: a novel equation. Journal of Theoretical Biology, 392(??):48–52, March 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315006001.

Vig-Milkovics:2019:MSB

[VMZK^{+}19] Zsuzsanna Vig-Milkovics, István Zachar, Ádám Kun, András Szilágyi, and Eörs Szathmáry. Moderate sex between protocells can balance between a decrease in assortment load and an increase in parasite spread. Journal of Theoretical Biology, 462(??):304–310, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305733.

vanNoort:2012:RWR

[vNÁBG12] Sander P. van Noort, Ricardo Águas, Sébastien Ballesteros, and M. Gabriela M. Gomes. The role of weather on the relation between influenza and influenza-like illness. Journal of Theoretical Biology, 298(??):131–137, April 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006448.

Verma:2018:BGI

[VNS18] Prateek Verma, Anjan K. Nandi, and Supratim Sengupta. Bribery games on interdependent complex networks. Journal of Theoretical Biology, 450(??):43–52, August 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302054.

Voets:2018:CLA

[Voe18] Philip J. G. M. Voets. Central line-associated bloodstream infections and catheter dwell-time: a theoretical foundation for a rule of thumb. Journal of Theoretical Biology, 445(??):31–32, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300845.

Vohradsky:2017:LPG

[Voh17] Jiri Vohradsky. Lambda phage genetic switch as a system with critical behaviour. Journal of Theoretical Biology, 431(??):32–38, October 27, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303557.

Vanderelst:2018:MBP

[VP18] Dieter Vanderelst and Herbert Peremans. Modeling bat prey capture in echolocating bats: The feasibility of reactive pursuit. Journal of Theoretical Biology, 456(??):305–314, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303515.

Vela-Perez:2013:AFG

[VPFV13] M. Vela-Pérez, M. A. Fontelos, and J. J. L. Velázquez. Ant foraging and geodesic paths in labyrinths: Analytical and computational results. Journal of Theoretical Biology, 320(??):100–112, March 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312006145.

Vazquez-Prieto:2014:QLM

[VPGDP^{+}14] Severo Vázquez-Prieto, Humberto González-Díaz, Esperanza Paniagua, Román Vilas, and Florencio M. Ubeira. A QSPR-like model for multilocus genotype networks of Fasciola hepatica in Northwest Spain. Journal of Theoretical Biology, 343(??):16–24, February 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005304.

Vesipa:2017:ISF

[VR17] Riccardo Vesipa and Luca Ridolfi. Impact of seasonal forcing on reactive ecological systems. Journal of Theoretical Biology, 419(??):23–35, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931730036X.

Valdez:2017:ERC

[VSD^{+}17] L. D. Valdez, G. J. Sibona, L. A. Diaz, M. S. Contigiani, and C. A. Condat. Effects of rainfall on Culex mosquito population dynamics. Journal of Theoretical Biology, 421(??):28–38, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301431.

VandePaer:2015:JEM

[VSLVB15] Céline Van de Paer, Pierre Saumitou-Laprade, Philippe Vernet, and Sylvain Billiard. The joint evolution and maintenance of self-incompatibility with gynodioecy or androdioecy. Journal of Theoretical Biology, 371(??):90–101, April 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000624.

Vukov:2011:ETC

[VSP11] Jeromos Vukov, Francisco C. Santos, and Jorge M. Pacheco. Escaping the tragedy of the commons via directed investments. Journal of Theoretical Biology, 287(??):37–41, October 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003729.

Vincent:2010:EEG

[VSW10] P. E. Vincent, S. J. Sherwin, and P. D. Weinberg. The effect of the endothelial glycocalyx layer on concentration polarisation of low density lipoprotein in arteries. Journal of Theoretical Biology, 265(1):1–17, July 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001980.

vonThienen:2015:MSP

[vTMW15] Wolfhard von Thienen, Dirk Metzler, and Volker Witte. Modeling shortest path selection of the ant linepithema humile using psychophysical theory and realistic parameter values. Journal of Theoretical Biology, 372(??):168–178, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315001010.

vanVeelen:2011:RDP

[vV11a] Matthijs van Veelen. The replicator dynamics with n players and population structure. Journal of Theoretical Biology, 276(1):78–85, May 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311000701.

vanVeelen:2011:RRI

[vV11b] Matthijs van Veelen. A rule is not a rule if it changes from case to case (a reply to Marshall’s comment). Journal of Theoretical Biology, 270(1):189–195, February 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005965.

vanVeelen:2018:FFT

[vV18] Matthijs van Veelen. Fisher’s fundamental theorem of natural selection, dynamic sufficiency, and the necessity of higher moments. Journal of Theoretical Biology, 437(??):299–304, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304691.

vanVeelen:2017:HR

[vVAH^{+}17] Matthijs van Veelen, Benjamin Allen, Moshe Hoffman, Burton Simon, and Carl Veller. Hamilton’s rule. Journal of Theoretical Biology, 414(??):176–230, February 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302533.

vanVelzen:2015:IEC

[vVE15] Ellen van Velzen and Rampal S. Etienne. The importance of ecological costs for the evolution of plant defense against herbivory. Journal of Theoretical Biology, 372(??):89–99, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315000983.

vanVeelen:2010:ITG

[vVGA10] Matthijs van Veelen, Julián García, and Leticia Avilés. It takes grouping and cooperation to get sociality. Journal of Theoretical Biology, 264(4):1240–1253, June 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001189.

vanVeelen:2012:GSI

[vVGSE12] Matthijs van Veelen, Julián García, Maurice W. Sabelis, and Martijn Egas. Group selection and inclusive fitness are not equivalent; the price equation vs. models and statistics. Journal of Theoretical Biology, 299(??):64–80, April 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311003754.

vanVeelen:2014:SMG

[vVLS14] Matthijs van Veelen, Shishi Luo, and Burton Simon. A simple model of group selection that cannot be analyzed with inclusive fitness. Journal of Theoretical Biology, 360(??):279–289, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931400397X.

vanVeelen:2012:MPG

[vVN12] Matthijs van Veelen and Martin A. Nowak. Multiplayer games on the cycle. Journal of Theoretical Biology, 292(??):116–128, January 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311004346.

Villarreal:2010:VEA

[VW10] Luis P. Villarreal and Guenther Witzany. Viruses are essential agents within the roots and stem of the tree of life. Journal of Theoretical Biology, 262(4):698–710, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309004895.

Vanselow:2019:WVS

[VWF19] Anna Vanselow, Sebastian Wieczorek, and Ulrike Feudel. When very slow is too fast — collapse of a predator–prey system. Journal of Theoretical Biology, 479(??):64–72, October 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302851.

Vandermeer:2012:SOS

[VY12] John Vandermeer and Senay Yitbarek. Self-organized spatial pattern determines biodiversity in spatial competition. Journal of Theoretical Biology, 300(??):48–56, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000069.

VanDyken:2019:CCS

[VZ19] J. David Van Dyken and Bo Zhang. Carrying capacity of a spatially-structured population: Disentangling the effects of dispersal, growth parameters, habitat heterogeneity and habitat clustering. Journal of Theoretical Biology, 460(??):115–124, January 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304454.

Vinogradova:2015:PLT

[VZB^{+}15] Svetlana V. Vinogradova, Kirill V. Zhudenkov, Neil Benson, Piet H. Van Der Graaf, Oleg V. Demin, and Tatiana A. Karelina. Prediction of long-term treatment outcome in HCV following 24 day PEG-IFN alpha-2b therapy using population pharmacokinetic–pharmacodynamic mixture modeling and classification analysis. Journal of Theoretical Biology, 382(??):91–98, October 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003203.

Woodall:2014:PCE

[WA14] Hannah Woodall and Ben Adams. Partial cross-enhancement in models for dengue epidemiology. Journal of Theoretical Biology, 351(??):67–73, June 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000915.

Wu:2014:PNA

[WAC14] Xiaogang Wu, Mohammad Al Hasan, and Jake Yue Chen. Pathway and network analysis in proteomics. Journal of Theoretical Biology, 362(??):44–52, December 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400304X.

Wallace:2010:PFD

[Wal10] Rodrick Wallace. Protein folding disorders: Toward a basic biological paradigm. Journal of Theoretical Biology, 267(4):582–594, December 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310004984.

Walters:2012:PEI

[Wal12] Kevin Walters. Parameter estimation for an immortal model of colonic stem cell division using approximate Bayesian computation. Journal of Theoretical Biology, 306(??):104–114, August 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312002020.

Wallace:2016:HMD

[Wal16] Rodrick Wallace. High metabolic demand in neural tissues: Information and control theory perspectives on the synergism between rate and stability. Journal of Theoretical Biology, 409(??):86–96, November 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302739.

Wallace:2018:NSM

[Wal18] Rodrick Wallace. New statistical models of nonergodic cognitive systems and their pathologies. Journal of Theoretical Biology, 436(??):72–78, January 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317304411.

Wang:2019:PMT

[Wan19] Yuanshi Wang. Pollination-mutualisms in a two-patch system with dispersal. Journal of Theoretical Biology, 476(??):51–61, September 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302310.

Waxman:2011:CRT

[Wax11a] D. Waxman. A compact result for the time-dependent probability of fixation at a neutral locus. Journal of Theoretical Biology, 274(1):131–135, April 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000312.

Waxman:2011:CCW

[Wax11b] D. Waxman. Comparison and content of the Wright– Fisher model of random genetic drift, the diffusion approximation, and an intermediate model. Journal of Theoretical Biology, 269(1):79–87, January 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005424.

Whitacre:2010:DDP

[WB10] James Whitacre and Axel Bender. Degeneracy: a design principle for achieving robustness and evolvability. Journal of Theoretical Biology, 263(1):143–153, March 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005347.

Waqar:2015:SAB

[WB15] Maleeha Waqar and Sidra Batool. In silico analysis of binding of neurotoxic venom ligands with acetylcholinesterase for therapeutic use in treatment of Alzheimer’s disease. Journal of Theoretical Biology, 372(??):107–117, May 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000995.

Wechsler:2019:TRM

[WB19] Daniel Wechsler and Jordi Bascompte. Thresholds in the resilience of modular social networks to invasion by defectors. Journal of Theoretical Biology, 460(??):56–63, January 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318304892.

Wu:2018:IME

[WBB^{+}18] Dan Wu, Richard C. Boucher, Brian Button, Timothy Elston, and Ching-Long Lin. An integrated mathematical epithelial cell model for airway surface liquid regulation by mechanical forces. Journal of Theoretical Biology, 438(??):34–45, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305179.

Wiff:2015:MPP

[WBM^{+}15] Rodrigo Wiff, Mauricio A. Barrientos, Andrés C. Milessi, J. C. Quiroz, and John Harwood. Modelling production per unit of food consumed in fish populations. Journal of Theoretical Biology, 365(??):67–75, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005931.

White:2012:MWH

[WBMH12] S. M. White, J. P. Burden, P. K. Maini, and R. S. Hails. Modelling the within-host growth of viral infections in insects. Journal of Theoretical Biology, 312(??):34–43, November 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003748.

Watson:2018:TPM

[WBMM18] Michael G. Watson, Helen M. Byrne, Charlie Macaskill, and Mary R. Myerscough. A two-phase model of early fibrous cap formation in atherosclerosis. Journal of Theoretical Biology, 456(??):123–136, November 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303837.

Wendelsdorf:2010:MCI

[WBRHE10] Katherine Wendelsdorf, Josep Bassaganya-Riera, Raquel Hontecillas, and Stephen Eubank. Model of colonic inflammation: Immune modulatory mechanisms in inflammatory bowel disease. Journal of Theoretical Biology, 264(4):1225–1239, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001608.

Winter:2010:ESM

[WC10] Samantha L. Winter and John H. Challis. The expression of the skeletal muscle force-length relationship in vivo: a simulation study. Journal of Theoretical Biology, 262(4):634–643, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005177.

Wang:2013:FSL

[WCC13] Hengtong Wang, Yueling Chen, and Yong Chen. Firstspike latency in Hodgkin’s three classes of neurons. Journal of Theoretical Biology, 328(??):19–25, July 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000957.

Wong:2014:CSN

[WCC14] Yung-Hao Wong, Ru-Hong Chen, and Bor-Sen Chen. Core and specific network markers of carcinogenesis from multiple cancer samples. Journal of Theoretical Biology, 362(??):17–34, December 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003348.

Westoby:2012:EAM

[WCF12] Mark Westoby, William K. Cornwell, and Daniel S. Falster. An evolutionary attractor model for sapwood cross section in relation to leaf area. Journal of Theoretical Biology, 303(??):98– 109, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001257.

Wu:2011:PAO

[WCHC11] Zimei Wu, C.-K. Chui, G.-S. Hong, and Stephen Chang. Physiological analysis on oscillatory behavior of glucoseinsulin regulation by model with delays. Journal of Theoretical Biology, 280(1):1–9, July 7, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001780.

Whang:2011:DMT

[WCJ11] Sungim Whang, Sunhwa Choi, and Eunok Jung. A dynamic model for tuberculosis transmission and optimal treatment strategies in south Korea. Journal of Theoretical Biology, 279(1):120–131, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001494.

Wirkus:2015:MMF

[WCM15] Stephen Wirkus, Erika T. Camacho, and Pamela A. Marshall. Mathematical modeling of fungal infection in immune compromised individuals: The effect of back mutation on drug treatment. Journal of Theoretical Biology, 385(??):66–76, November 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004105.

Wang:2015:DDS

[WCPF15] Shaopeng Wang, Anping Chen, Stephen W. Pacala, and Jingyun Fang. Density-dependent speciation alters the structure and dynamics of neutral communities. Journal of Theoretical Biology, 372(??):128–134, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000661.

Wang:2019:EVS

[WCS19] Lingzi Wang, Stephen J. Cornell, and Michael P. Speed. The evolution of variance in sequential defences. Journal of Theoretical Biology, 462(??):194–209, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830479X.

Wang:2012:ICI

[WCSA12] Bing Wang, Lang Cao, Hideyuki Suzuki, and Kazuyuki Aihara. Impacts of clustering on interacting epidemics. Journal of Theoretical Biology, 304(??):121–130, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001464.

Wang:2014:PEM

[WcW14] Hong Wang and Xi cheng Wang. Parameter estimation for metabolic networks with two stage Bregman regularization homotopy inversion algorithm. Journal of Theoretical Biology, 343(??):199–207, February 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004487.

Wang:2015:EDF

[WCW15] Xiaofeng Wang, Xiaojie Chen, and Long Wang. Evolutionary dynamics of fairness on graphs with migration. Journal of Theoretical Biology, 380(??):103–114, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002489.

Wang:2011:TIO

[WD11] Yuanshi Wang and Donald L. DeAngelis. Transitions of interaction outcomes in a uni-directional consumer-resource system. Journal of Theoretical Biology, 280(1):43–49, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311001901.

Weaver:2012:ITE

[WD12] Iain S. Weaver and James G. Dyke. The importance of timescales for the emergence of environmental self-regulation. Journal of Theoretical Biology, 313(??):172–180, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200389X.

Wendelsdorf:2011:HIR

[WDH^{+}11] K. Wendelsdorf, G. Dean, Shuhua Hu, S. Nordone, and H. T. Banks. Host immune responses that promote initial HIV spread. Journal of Theoretical Biology, 289(??):17–35, November 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311004073.

Wang:2016:IWR

[WDH^{+}16] Chuan Wang, Xiaobao Dong, Lei Han, Xiao-Dong Su, Ziding Zhang, Jinyan Li, and Jiangning Song. Identification of WD40 repeats by secondary structure-aided profile-profile alignment. Journal of Theoretical Biology, 398(??):122–129, June 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001661.

Wu:2013:DTD

[WDL^{+}13] Xiaotian Wu, Venkata R. Duvvuri, Yijun Lou, Nicholas H. Ogden, Yann Pelcat, and Jianhong Wu. Developing a temperature-driven map of the basic reproductive number of the emerging tick vector of Lyme disease ixodes scapularis in Canada. Journal of Theoretical Biology, 319(??):50–61, February 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005899.

Warsinske:2017:IMD

[WDL^{+}17] Hayley C. Warsinske, Robert M. DiFazio, Jennifer J. Linderman, JoAnne L. Flynn, and Denise E. Kirschner. Identifying mechanisms driving formation of granuloma-associated fibrosis during mycobacterium tuberculosis infection. Journal of Theoretical Biology, 429(??):1–17, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302874.

Wosniack:2017:PEE

[WdLS17] M. E. Wosniack, M. G. E. da Luz, and L. S. Schulman. Punctuated equilibrium as an emergent process and its modified thermodynamic characterization. Journal of Theoretical Biology, 412(??):113–122, January 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303460.

Wery:2019:FEP

[WDN^{+}19] Méline Wery, Olivier Dameron, Jacques Nicolas, Elisabeth Remy, and Anne Siegel. Formalizing and enriching phenotype signatures using Boolean networks. Journal of Theoretical Biology, 467(??):66–79, April 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300153.

White:2015:MPP

[WdVMD15] D. White, G. de Vries, J. Martin, and A. Dawes. Microtubule patterning in the presence of moving motor proteins. Journal of Theoretical Biology, 382(??):81–90, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003197.

Weiss:2012:MTT

[Wei12] Michael Weiss. A model for transit time distributions through organs that accounts for fractal heterogeneity. Journal of Theoretical Biology, 301(??):57–61, May 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312000768.

Weigang:2017:CPT

[Wei17] Helene C. Weigang. Coevolution of patch-type dependent emigration and patch-type dependent immigration. Journal of Theoretical Biology, 426(??):140–151, August 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730228X.

Wicke:2017:CRO

[WF17] Kristina Wicke and Mareike Fischer. Comparing the rankings obtained from two biodiversity indices: the fair proportion index and the Shapley value. Journal of Theoretical Biology, 430(??):207–214, October 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303430.

Whitehead:2018:CCD

[WF18] Hal Whitehead and John K. B. Ford. Consequences of culturally-driven ecological specialization: Killer whales and beyond. Journal of Theoretical Biology, 456(??):279–294, November 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303941.

Wu:2014:EIP

[WFC^{+}14] Min Wu, Hermann B. Frieboes, Mark A. J. Chaplain, Steven R. McDougall, Vittorio Cristini, and John S. Lowengrub. The effect of interstitial pressure on therapeutic agent transport: Coupling with the tumor blood and lymphatic vascular systems. Journal of Theoretical Biology, 355(??):194–207, August 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931400232X. See comments [GS13a].

Wu:2013:EIP

[WFM^{+}13] Min Wu, Hermann B. Frieboes, Steven R. McDougall, Mark A. J. Chaplain, Vittorio Cristini, and John Lowengrub. The effect of interstitial pressure on tumor growth: Coupling with the blood and lymphatic vascular systems. Journal of Theoretical Biology, 320(??):131–151, March 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006200.

Wu:2019:PAM

[WFW19] Te Wu, Feng Fu, and Long Wang. Phenotype affinity mediated interactions can facilitate the evolution of cooperation. Journal of Theoretical Biology, 462(??):361–369, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305769.

Wu:2013:ATS

[WFZW13] Te Wu, Feng Fu, Yanling Zhang, and Long Wang. Adaptive tag switching reinforces the coevolution of contingent cooperation and tag diversity. Journal of Theoretical Biology, 330(??):45–55, August 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313001604.

Wilson:2013:MST

[WGC13] Natasha P. Wilson, Bradford Gates, and Mariajosé Castellanos. Modeling the short time-scale dynamics of βamyloid-neuron interactions. Journal of Theoretical Biology, 331(??):28–37, August 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000805.

Wu:2018:NMP

[WGDZ18] Chuanyan Wu, Rui Gao, Yang De Marinis, and Yusen Zhang. A novel model for protein sequence similarity analysis based on spectral radius. Journal of Theoretical Biology, 446(??):61– 70, June 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318301097.

Wu:2014:SSF

[WGH^{+}14] Tai-Hsien Wu, Rurng-Sheng Guo, Guo-Wei He, Ying-Ming Liu, and Dewei Qi. Simulation of swimming of a flexible filament using the generalized lattice-spring lattice-Boltzmann method. Journal of Theoretical Biology, 349(??):1–11, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314000381.

Woolley:2015:GCL

[WGO^{+}15] Thomas E. Woolley, Eamonn A. Gaffney, James M. Oliver, Sarah L. Waters, Ruth E. Baker, and Alain Goriely. Global contraction or local growth, bleb shape depends on more than just cell structure. Journal of Theoretical Biology, 380(??):83–97, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931500199X.

Walker:2010:ASC

[WGS10] D. C. Walker, N. T. Georgopoulos, and J. Southgate. Antisocial cells: Predicting the influence of E-cadherin loss on the growth of epithelial cell populations. Journal of Theoretical Biology, 262(3):425–440, February 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309004779.

Woodhouse:2015:STC

[WGS15] Francis G. Woodhouse, Bruce S. Gardiner, and David W. Smith. Short-term consolidation of articular cartilage in the long-term context of osteoarthritis. Journal of Theoretical Biology, 368(??):102–112, March 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000041.

Wakano:2011:PFC

[WH11] Joe Yuichiro Wakano and Christoph Hauert. Pattern formation and chaos in spatial ecological public goodsgames. Journal of Theoretical Biology, 268(1):30–38, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005138.

Wang:2014:RSM

[WH14] Y. X. Rachel Wang and Haiyan Huang. Review on statistical methods for gene network reconstruction using expression data. Journal of Theoretical Biology, 362(??):53–61, December 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001969.

White:2017:EEE

[WHH17] Diana White, Stéphane Honoré, and Florence Hubert. Exploring the effect of end-binding proteins and microtubule targeting chemotherapy drugs on microtubule dynamic instability. Journal of Theoretical Biology, 429(??):18–34, September 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302850.

Willener:2015:EWS

[WHHS15] Astrid S. T. Willener, Yves Handrich, Lewis G. Halsey, and Siobhán Strike. Effect of walking speed on the gait of king penguins: An accelerometric approach. Journal of Theoretical Biology, 387(??):166–173, December 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004622.

Whittlesey:2011:WBU

[Whi11] Robert W. Whittlesey. Wake-based unsteady modeling of the aquatic beetle Dytiscus marginalis. Journal of Theoretical Biology, 291(??):14–21, December 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003912.

Wu:2013:UFC

[WHL13] Gang Wu, Keqiang Han, and Fenglin Lv. Use of fast conformational sampling to improve the characterization of VEGF a-peptide interactions. Journal of Theoretical Biology, 317(??):293–300, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005607.

Wu:2013:AAD

[WHS^{+}13] Hsien-Tsai Wu, Po-Chun Hsu, Cheuk-Kwan Sun, Hou-Jun Wang, Cyuan-Cin Liu, Hong-Ruei Chen, An-Bang Liu, Chieh-Ju Tang, and Men-Tzung Lo. Assessment of autonomic dysfunction in patients with type 2 diabetes using reactive hyperemia. Journal of Theoretical Biology, 330(??):9–17, August 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001483.

Wang:2017:IBE

[WHT17] Bing Wang, Yuexing Han, and Gouhei Tanaka. Interplay between epidemic spread and information propagation on metapopulation networks. Journal of Theoretical Biology, 420(??):18–25, May 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300747.

Wang:2018:MBT

[WHWZ18] Dayang Wang, Chaobo He, Chengqing Wu, and Yongshan Zhang. Mechanical behaviors of tension and relaxation of tongue and soft palate: Experimental and analytical modeling. Journal of Theoretical Biology, 459(??):142–153, December 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304715.

Wyse:2017:ICE

[WHYMG17] J. Max Wyse, Ian C. W. Hardy, Lisa Yon, and Mike Mesterton-Gibbons. The impact of competition on elephant musth strategies: a game-theoretic model. Journal of Theoretical Biology, 417(??):109–130, March 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300255.

Wu:2017:CMP

[WHZ^{+}17] Sijia Wu, Jiuqiang Han, Xinman Zhang, Dexing Zhong, and Ruiling Liu. A computational model for predicting integrase catalytic domain of retrovirus. Journal of Theoretical Biology, 423(??):63–70, June 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301844.

Williams:2011:CAE

[Wil11] R. J. P. Williams. Chemical advances in evolution by and changes in use of space during time. Journal of Theoretical Biology, 268(1):146–159, January 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004959.

Wills:2013:FPP

[Wil13] Peter R. Wills. Frameshifted prion proteins as pathological agents: Quantitative considerations. Journal of Theoretical Biology, 325(??):52–61, May 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000799.

Wang:2016:MMH

[WJL^{+}16] Yuancheng Wang, Peter H. Janssen, Tammy A. Lynch, Bruce van Brunt, and David Pacheco. A mechanistic model of hydrogen-methanogen dynamics in the rumen. Journal of Theoretical Biology, 393(??):75–81, March 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000333.

Witthoft:2012:BMC

[WK12] Alexandra Witthoft and George Em Karniadakis. A bidirectional model for communication in the neurovascular unit. Journal of Theoretical Biology, 311(??):80–93, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312003487.

Weigang:2015:EDU

[WK15] Helene C. Weigang and Éva Kisdi. Evolution of dispersal under a fecundity-dispersal trade-off. Journal of Theoretical Biology, 371(??):145–153, April 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000727.

Wild:2017:EDD

[WK17] Geoff Wild and Judith Korb. Evolution of delayed dispersal and subsequent emergence of helping, with implications for cooperative breeding. Journal of Theoretical Biology, 427(??):53–64, August 1, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931730259X.

Williams:2018:EPV

[WK18a] Paul David Williams and Stephanie Jill Kamel. The evolution of pathogen virulence: Effects of transitions between host types. Journal of Theoretical Biology, 438(??):1–8, February 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317305155.

Wood:2018:CBB

[WK18b] Karen E. Wood and Natalia L. Komarova. Cooperation-based branching as a mechanism of evolutionary speciation. Journal of Theoretical Biology, 445(??):166–186, May 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318300936.

Webb:2013:RSP

[WKB13] Steven D. Webb, Matt J. Keeling, and Mike Boots. The role of spatial population structure on the evolution of parasites with acquired immunity and demography. Journal of Theoretical Biology, 324(??):21–31, May 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000404.

Wilkinson:2016:DNB

[WKH16] Richard D. Wilkinson, Apostolos Kapranas, and Ian C. W. Hardy. Detecting non-binomial sex allocation when developmental mortality operates. Journal of Theoretical Biology, 408(??):167–178, November 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316302429.

Wannige:2014:NDS

[WKS14] C. T. Wannige, D. Kulasiri, and S. Samarasinghe. A nutrient dependant switch explains mutually exclusive existence of meiosis and mitosis initiation in budding yeast. Journal of Theoretical Biology, 341(??):88–101, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004669.

Waxman:2010:SMS

[WL10] D. Waxman and L. Loewe. A stochastic model for a single click of Muller’s ratchet. Journal of Theoretical Biology, 264(4):1120–1132, June 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310001372.

Wakano:2012:CEC

[WL12a] Joe Yuichiro Wakano and Laurent Lehmann. Corrigendum to “Evolutionary and convergence stability for continuous phenotypes in finite populations derived from two-allele models” [J. Theor. Biol. 310 (2012) 206–215]. Journal of Theoretical Biology, 312(??):157, November 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003815. See [WL12b].

Wakano:2012:ECS

[WL12b] Joe Yuichiro Wakano and Laurent Lehmann. Evolutionary and convergence stability for continuous phenotypes in finite populations derived from two-allele models. Journal of Theoretical Biology, 310(??):206–215, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312003293. See corrigen-

dum [WL12a].

Whitehead:2012:ASN

[WL12c] Hal Whitehead and David Lusseau. Animal social networks as substrate for cultural behavioural diversity. Journal of Theoretical Biology, 294(??):19–28, February 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005406.

Wagner:2013:CSF

[WL13] Gregory L. Wagner and Eric Lauga. Crawling scallop: Friction-based locomotion with one degree of freedom. Journal of Theoretical Biology, 324(??):42–51, May 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000465.

Wakano:2014:EBD

[WL14] Joe Yuichiro Wakano and Laurent Lehmann. Evolutionary branching in deme-structured populations. Journal of Theoretical Biology, 351(??):83–95, June 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001209.

Whalen:2015:CBT

[WL15] Andrew Whalen and Kevin Laland. Conformity biased transmission in social networks. Journal of Theoretical Biology, 380(??):542–549, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003070.

Wang:2019:PGG

[WL19] Xianjia Wang and Shaojie Lv. The public goods game with shared punishment cost in well-mixed and structured populations. Journal of Theoretical Biology, 476(??):36–43, September 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302176.

Wu:2011:HDM

[WLD^{+}11] Sangwook Wu, Shubin Liu, Charles H. Davis, Darrel W. Stafford, John D. Kulman, and Lee G. Pedersen. A heterodimer model for concerted action of vitamin K carboxylase and vitamin K reductase in vitamin K cycle. Journal of Theoretical Biology, 279(1):143–149, June 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001767.

Wu:2014:NAR

[WLL^{+}14] Duojiao Wu, Xiaoping Liu, Chen Liu, Zhiping Liu, Ming Xu, Ruiming Rong, Mengjia Qian, Luonan Chen, and Tongyu Zhu. Network analysis reveals roles of inflammatory factors in different phenotypes of kidney transplant patients. Journal of Theoretical Biology, 362(??):62–68, December 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001416.

Luo:2010:MDU

[wLW10] Jia wei Luo and Ting Wang. Motif discovery using an immune genetic algorithm. Journal of Theoretical Biology, 264(2):319– 325, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000779.

Wei:2019:VDE

[WLW19] Yuting Wei, Yaosen Lin, and Bin Wu. Vaccination dilemma on an evolving social network. Journal of Theoretical Biology, 483(??):Article 109978, December 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519319303212. See corrigendum [WLW20].

Wei:2020:CVD

[WLW20] Yuting Wei, Yaosen Lin, and Bin Wu. Corrigendum to “Vaccination dilemma on an evolving social network” [Journal of Theoretical Biology 483 (2019) 109978]. Journal of Theoretical Biology, 490(??):Article 110154, April 7, 2020. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519320300102.

Wen:2017:PPP

[WLY^{+}17] Yu-Ting Wen, Hai-Jun Lei, Zhu-Hong You, Bai-Ying Lei, Xing Chen, and Li-Ping Li. Prediction of protein–protein interactions by label propagation with protein evolutionary and chemical information derived from heterogeneous network. Journal of Theoretical Biology, 430(??):9–20, October 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317302643.

Warrell:2017:SSP

[WM17] Jonathan Warrell and Musa Mhlanga. Stability and structural properties of gene regulation networks with coregulation rules. Journal of Theoretical Biology, 420(??):304–317, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303381.

Wang:2018:EBE

[WMCL18] Yi Wang, Junling Ma, Jinde Cao, and Li Li. Edge-based epidemic spreading in degree-correlated complex networks. Journal of Theoretical Biology, 454(??):164–181, October 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318302984.

Wang:2011:CTN

[WMD^{+}11] Zhen Wang, Aleksandra Murks, Wen-Bo Du, Zhi-Hai Rong, and Matjaz Perc. Coveting thy neighbors fitness as a means to resolve social dilemmas. Journal of Theoretical Biology, 277(1):19–26, May 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311001184.

Wcislo:2016:PPA

[WMD16] Rafal Wcislo, S. Shea Miller, and Witold Dzwinel. PAM: Particle automata model in simulation of Fusarium graminearum pathogen expansion. Journal of Theoretical Biology, 389(??):110–122, January 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315005172.

Wellard:2010:ECP

[WMHH10] C. Wellard, J. Markham, E. D. Hawkins, and P. D. Hodgkin. The effect of correlations on the population dynamics of lymphocytes. Journal of Theoretical Biology, 264(2):443–449, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310000949.

Wan:2013:GSL

[WMK13] Shibiao Wan, Man-Wai Mak, and Sun-Yuan Kung. GOASVM: a subcellular location predictor by incorporating termfrequency gene ontology into the general form of Chou’s pseudo-amino acid composition. Journal of Theoretical Biology, 323(??):40–48, April 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313000374.

Wan:2014:RLC

[WMK14] Shibiao Wan, Man-Wai Mak, and Sun-Yuan Kung. R3P-Loc: a compact multi-label predictor using ridge regression and random projection for protein subcellular localization. Journal of Theoretical Biology, 360(??):34–45, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314003749.

Wan:2015:MHL

[WMK15] Shibiao Wan, Man-Wai Mak, and Sun-Yuan Kung. mLASSOhum: a LASSO-based interpretable human-protein subcellular localization predictor. Journal of Theoretical Biology, 382(??):223–234, October 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003215.

Wan:2016:MAT

[WMK16] Shibiao Wan, Man-Wai Mak, and Sun-Yuan Kung. Mem-ADSVM: a two-layer multi-label predictor for identifying multi-functional types of membrane proteins. Journal of Theoretical Biology, 398(??):32–42, June 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001545.

Wang:2016:REP

[WML16] Mao-Xiang Wang, Yu-Qiang Ma, and Pik-Yin Lai. Regulatory effects on the population dynamics and wave propagation in a cell lineage model. Journal of Theoretical Biology, 393(??):105–117, March 21, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000436.

Wang:2017:IRI

[WML^{+}17] Xia Wang, Gregory Mink, Daniel Lin, Xinyu Song, and Libin Rong. Influence of raltegravir intensification on viral load and 2-LTR dynamics in HIV patients on suppressive antiretroviral therapy. Journal of Theoretical Biology, 416(??):16–27, March 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316304258.

West:2018:CCE

[WMN18] Jeffrey West, Yongqian Ma, and Paul K. Newton. Capitalizing on competition: An evolutionary model of competitive release in metastatic castration resistant prostate cancer treatment. Journal of Theoretical Biology, 455(??):249–260, October 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318303527.

Walmsley:2015:ASP

[WMPF^{+}15] John Walmsley, Gary R. Mirams, Joe Pitt-Francis, Blanca Rodriguez, and Kevin Burrage. Application of stochastic phenomenological modelling to cell-to-cell and beat-to-beat electrophysiological variability in cardiac tissue. Journal of Theoretical Biology, 365(??):325–336, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006195.

Wittmann:2010:BMU

[WMT10] Dominik M. Wittmann, Carsten Marr, and Fabian J. Theis. Biologically meaningful update rules increase the critical connectivity of generalized Kauffman networks. Journal of Theoretical Biology, 266(3):436–448, October 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003565.

Wei:2015:RST

[WMT15] Ning Wei, Yoichiro Mori, and Elena G. Tolkacheva. The role of short term memory and conduction velocity restitution in alternans formation. Journal of Theoretical Biology, 367(??):21–28, February 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006663.

Wei:2016:DEE

[WMT16] Ning Wei, Yoichiro Mori, and Elena G. Tolkacheva. The dual effect of ephaptic coupling on cardiac conduction with heterogeneous expression of connexin 43. Journal of Theoretical Biology, 397(??):103–114, May 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001259.

Waxman:2019:SST

[WN19] David Waxman and Pierre Nouvellet. Sub- or supercritical transmissibilities in a finite disease outbreak: Symmetry in outbreak properties of a disease conditioned on extinction. Journal of Theoretical Biology, 467(??):80–86, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300505.

Wang:2011:FMP

[WO11] Peng Wang and William L. Olbricht. Fluid mechanics in the perivascular space. Journal of Theoretical Biology, 274(1):52– 57, April 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311000269.

Waldherr:2015:DOM

[WOB15] Steffen Waldherr, Diego A. Oyarzún, and Alexander Bockmayr. Dynamic optimization of metabolic networks coupled with gene expression. Journal of Theoretical Biology, 365(??):469–485, January 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006250.

Wodarz:2018:ECD

[Wod18] Dominik Wodarz. Effect of cellular de-differentiation on the dynamics and evolution of tissue and tumor cells in mathematical models with feedback regulation. Journal of Theoretical Biology, 448(??):86–93, July 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301504.

Wohlin:2015:NSH

[Woh15] Åsa Wohlin. Numeral series hidden in the distribution of atomic mass of amino acids to codon domains in the genetic code. Journal of Theoretical Biology, 369(??):95–109, March 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315000223.

Wolpert:2011:PIP

[Wol11] Lewis Wolpert. Positional information and patterning revisited. Journal of Theoretical Biology, 269(1):359–365, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005795.

Wolff:2012:RRP

[Wol12] Irenaeus Wolff. Retaliation and the role for punishment in the evolution of cooperation. Journal of Theoretical Biology, 315(??):128–138, December 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312004900.

Wolpert:2014:RMR

[Wol14] Lewis Wolpert. A relaxed modification of the Rappaport model for cytokinesis. Journal of Theoretical Biology, 345(??):109, March 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005596.

Wood:2010:SSU

[Woo10] A. Jamie Wood. Strategy selection under predation; evolutionary analysis of the emergence of cohesive aggregations. Journal of Theoretical Biology, 264(4):1102–1110, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001682.

Wang:2017:PDE

[WP17] Yi-Shan Wang and Jonathan R. Potts. Partial differential equation techniques for analysing animal movement: a comparison of different methods. Journal of Theoretical Biology, 416(??):52–67, March 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300024.

Watkinson-Powell:2017:RGI

[WPA17] Benjamin Watkinson-Powell and Nina Alphey. Resistance to genetic insect control: Modelling the effects of space. Journal of Theoretical Biology, 413(??):72–85, January 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316303575.

Wang:2012:SSR

[WPH^{+}12] Yunjiao Wang, Paweł Paszek, Caroline A. Horton, Hong Yue, Michael R. H. White, Douglas B. Kell, Mark R. Muldoon, and David S. Broomhead. A systematic survey of the response of a model NF-κB signalling pathway to TNFα stimulation. Journal of Theoretical Biology, 297(??):137–147, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931100631X.

Wade:2016:EBC

[WPPD16] M. J. Wade, R. W. Pattinson, N. G. Parker, and J. Dolfing. Emergent behaviour in a chlorophenol-mineralising three-tiered microbial ‘food web’. Journal of Theoretical Biology, 389(??):171–186, January 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005317.

Wang:2014:SPS

[WR14] Sunpeng Wang and Libin Rong. Stochastic population switch may explain the latent reservoir stability and intermittent viral blips in HIV patients on suppressive therapy. Journal of Theoretical Biology, 360(??):137–148, November 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003932.

Wang:2019:TIA

[WRC^{+}19] Lingzi Wang, Graeme D. Ruxton, Stephen J. Cornell, Michael P. Speed, and Mark Broom. A theory for investment across defences triggered at different stages of a predator– prey encounter. Journal of Theoretical Biology, 473(??):9–19, July 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301572.

Walsh:2016:CMP

[WRH^{+}16] Darragh M. Walsh, Philipp T. Röth, William R. Holmes, Kerry A. Landman, Tobias D. Merson, and Barry D. Hughes. Is cell migration or proliferation dominant in the formation of linear arrays of oligodendrocytes? Journal of Theoretical Biology, 406(??):17–30, October 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630162X.

Wu:2019:MRG

[WRWX19] Chengyuan Wu, Shiquan Ren, Jie Wu, and Kelin Xia. Magnus representation of genome sequences. Journal of Theoretical Biology, 480(??):104–111, November 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319303169.

Wang:2012:INP

[WRZ^{+}12] Yongcui Wang, Xianwen Ren, Chunhua Zhang, Naiyang Deng, and Xiangsun Zhang. Interrogating noise in protein sequences from the perspective of protein–protein interactions prediction. Journal of Theoretical Biology, 315(??):64–70, December 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312004754.

Wu:2010:RVN

[WS10] Jihua Wu and Edward Susko. Rate-variation need not defeat phylogenetic inference through pairwise sequence comparisons. Journal of Theoretical Biology, 263(4):587–589, April 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519309005955.

Weissmann:2016:PCS

[WS16] Haim Weissmann and Nadav M. Shnerb. Predicting catastrophic shifts. Journal of Theoretical Biology, 397(??):128– 134, May 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001296.

Wang:2010:NSD

[WSH^{+}10] Minghao Wang, Xiaofeng Song, Ping Han, Wei Li, and Bin Jiang. New syntax to describe local continuous structuresequence information for recognizing new pre-miRNAs. Journal of Theoretical Biology, 264(2):578–584, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001128.

Wasik:2019:DLS

[WSK^{+}19] Szymon Wasik, Natalia Szostak, Mateusz Kudla, Michał Wachowiak, Krzysztof Krawiec, and Jacek Blazewicz. Detecting life signatures with RNA sequence similarity measures. Journal of Theoretical Biology, 463(??):110–120, February 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318306155.

Wang:2014:IAM

[WSLC14] Hengtong Wang, Yongjuan Sun, Yichen Li, and Yong Chen. Influence of autapse on mode-locking structure of a Hodgkin– Huxley neuron under sinusoidal stimulus. Journal of Theoretical Biology, 358(??):25–30, October 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002999.

Wolkenhauer:2012:RTP

[WSM12] Olaf Wolkenhauer, Darryl Shibata, and Mihajlo D. Mesarović. The role of theorem proving in systems biology. Journal of Theoretical Biology, 300(??):57–61, May 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311006473.

Wang:2014:REF

[WSP14] Zhen Wang, Attila Szolnoki, and Matjaz Perc. Rewarding evolutionary fitness with links between populations promotes cooperation. Journal of Theoretical Biology, 349(??):50–56, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000563.

Werner:2018:EIF

[WSRG18] Jan Werner, Nikolaos Sfakianakis, Alan D. Rendall, and Eva Maria Griebeler. Energy intake functions and energy budgets of ectotherms and endotherms derived from their ontogenetic growth in body mass and timing of sexual maturation. Journal of Theoretical Biology, 444(??):83–92, May 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300596.

Watteaux:2015:SRN

[WST15] Romain Watteaux, Roman Stocker, and John R. Taylor. Sensitivity of the rate of nutrient uptake by chemotactic bacteria to physical and biological parameters in a turbulent environment. Journal of Theoretical Biology, 387(??):120–135, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931500394X.

Wang:2016:IMC

[WSTL16] Shengqin Wang, Beili Sun, Jing Tu, and Zuhong Lu. Improving the microbial community reconstruction at the genus level by multiple 16s rRNA regions. Journal of Theoretical Biology, 398(??):1–8, June 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316001570.

Wu:2014:DER

[WSVT14] Yue Wu, Clare A. Saddler, Frank Valckenborgh, and Mark M. Tanaka. Dynamics of evolutionary rescue in changing environments and the emergence of antibiotic resistance. Journal of Theoretical Biology, 340(??):222–231, January 7, 2014.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313004566.

Wiegmann:2010:MCO

[WSW10] Daniel D. Wiegmann, Steven M. Seubert, and Gordon A. Wade. Mate choice and optimal search behavior: Fitness returns under the fixed sample and sequential search strategies. Journal of Theoretical Biology, 262(4):596–600, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251930900410X.

Warren:2010:EIC

[WTC10] N. J. Warren, M. H. Tawhai, and E. J. Crampin. The effect of intracellular calcium oscillations on fluid secretion in airway epithelium. Journal of Theoretical Biology, 265(3):270–277, August 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002365.

Wang:2016:SSM

[WTC16] Xia Wang, Sanyi Tang, and Robert A. Cheke. A stage structured mosquito model incorporating effects of precipitation and daily temperature fluctuations. Journal of Theoretical Biology, 411(??):27–36, December 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303046.

Wang:2010:BSF

[WTQL10] Shiyuan Wang, Fengchun Tian, Yu Qiu, and Xiao Liu. Bilateral similarity function: a novel and universal method for similarity analysis of biological sequences. Journal of Theoretical Biology, 265(2):194–201, July 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001967.

Watts:2011:SVD

[WTZ^{+}11] Margaret Watts, Joel Tabak, Charles Zimliki, Arthur Sherman, and Richard Bertram. Slow variable dominance and phase resetting in phantom bursting. Journal of Theoretical Biology, 276(1):218–228, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000683.

Wu:2014:NMS

[Wu14] Bo Wu. A novel method to study aging in bacteria that reproduce by morphologically symmetric fission. Journal of Theoretical Biology, 347(??):1–6, April 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005523.

Wentz:2015:MMR

[WVO^{+}15] J. M. Wentz, V. Vainstein, D. Oldson, Z. Gluzman-Poltorak, L. A. Basile, and D. Stricklin. Mathematical model of radiation effects on thrombopoiesis in rhesus macaques and humans. Journal of Theoretical Biology, 383(??):44–60, October 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003495.

Wang:2012:ITM

[WW12a] Lin Wang and Xiao Wang. Influence of temporary migration on the transmission of infectious diseases in a migrants’ home village. Journal of Theoretical Biology, 300(??):100– 109, May 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312000057.

Wang:2012:HIM

[WW12b] Xiunan Wang and Wendi Wang. An HIV infection model based on a vectored immunoprophylaxis experiment. Journal of Theoretical Biology, 313(??):127–135, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312004456.

Wang:2019:GDP

[WW19] Yuanshi Wang and Hong Wu. Global dynamics of parasitismcompetition systems with one host and multiple parasites. Journal of Theoretical Biology, 461(??):268–275, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318305344.

Wensink:2014:NSD

[WWB14] Maarten J. Wensink, Tomasz F. Wrycza, and Annette Baudisch. No senescence despite declining selection pressure: Hamilton’s result in broader perspective. Journal of Theoretical Biology, 347(??):176–181, April 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005432.

Walker:2019:BBL

[WWIG19] Benjamin J. Walker, Richard J. Wheeler, Kenta Ishimoto, and Eamonn A. Gaffney. Boundary behaviours of Leishmania mexicana: a hydrodynamic simulation study. Journal of Theoretical Biology, 462(??):311–320, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830568X.

Wang:2012:RMR

[WWY12] Xiang-Sheng Wang, Jianhong Wu, and Yong Yang. Richards model revisited: Validation by and application to infection dynamics. Journal of Theoretical Biology, 313(??):12–19, November 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003761.

Wu:2010:MWS

[WXC10] Zhi-Cheng Wu, Xuan Xiao, and Kuo-Chen Chou. 2d-MH: a web-server for generating graphic representation of protein sequences based on the physicochemical properties of their constituent amino acids. Journal of Theoretical Biology, 267(1):29–34, November 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004169.

Wang:2010:GPP

[WXmH10] Tong Wang, Tian Xia, and Xiao ming Hu. Geometry preserving projections algorithm for predicting membrane protein types. Journal of Theoretical Biology, 262(2):208–213, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004573.

Wang:2012:MME

[WXWL12] Xia Wang, Yanni Xiao, Junrui Wang, and Xinxin Lu. A mathematical model of effects of environmental contamination and presence of volunteers on hospital infections in China. Journal of Theoretical Biology, 293(??):161–173, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311005248.

Wang:2013:GSB

[WY13] Doris Z. Wang and Hong Yan. A graph spectrum based geometric biclustering algorithm. Journal of Theoretical Biology, 317(??):200–211, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005371.

Wang:2017:IDT

[WYKT17] Shi-Chang Wang, Jie-Ru Yu, Shun Kurokawa, and Yi Tao. Imitation dynamics with time delay. Journal of Theoretical Biology, 420(??):8–11, May 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317300966.

Wu:2018:VSL

[WYL^{+}18] Jiayu Wu, Hong Yuan, Longyuan Li, Kunjie Fan, Shanguang Qian, and Bing Li. Viscoelastic shear lag model to predict the micromechanical behavior of tendon under dynamic tensile loading. Journal of Theoretical Biology, 437(??):202–213, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317304812.

Wang:2019:PRP

[WYL^{+}19] Lei Wang, Xin Yan, Meng-Lin Liu, Ke-Jian Song, Xiao-Fei Sun, and Wen-Wen Pan. Prediction of RNA-protein interactions by combining deep convolutional neural network with feature selection ensemble method. Journal of Theoretical Biology, 461(??):230–238, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305058.

Wang:2017:APA

[WYX^{+}17] Lei Wang, Zhu-Hong You, Shi-Xiong Xia, Feng Liu, Xing Chen, Xin Yan, and Yong Zhou. Advancing the prediction accuracy of protein–protein interactions by utilizing evolutionary information from position-specific scoring matrix and ensemble classifier. Journal of Theoretical Biology, 418(??):105– 110, April 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317300036.

Weiss:2012:IPA

[WZ12] Meital Weiss and Amotz Zahavi. The information provided by the adrenal cortical steroids: a hypothesis. Journal of Theoretical Biology, 311(??):66–68, October 21, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312003335.

Wu:2015:FML

[WZ15] Xiaowei Wu and Hongxiao Zhu. Fast maximum likelihood estimation of mutation rates using a birth-death process. Journal of Theoretical Biology, 366(??):1–7, February 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006523.

Wei:2017:DBG

[WZ17] Ze-Gang Wei and Shao-Wu Zhang. DBH: a de Bruijn graph-based heuristic method for clustering large-scale 16s rRNA sequences into OTUs. Journal of Theoretical Biology, 425(??):80–87, July 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301832.

Wang:2018:VPF

[WZ18] Xiaoli Wang and Guohong Zhang. Vegetation pattern formation in seminal systems due to internal competition reaction between plants. Journal of Theoretical Biology, 458(??):10–14, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304247.

Wang:2013:MAT

[WZJ^{+}13] Kai Wang, Xueliang Zhang, Zhen Jin, Haimei Ma, Zhidong Teng, and Lei Wang. Modeling and analysis of the transmission of Echinococcosis with application to Xinjiang Uygur Autonomous Region of China. Journal of Theoretical Biology, 333(??):78–90, September 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001884.

Wei:2015:PSR

[WZLW15] Xinzhu Wei, Lei Zhao, Martin Lascoux, and David Waxman. Population structure and the rate of evolution. Journal of Theoretical Biology, 365(??):486–495, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006316.

Wang:2019:FSI

[WZM19] Lidong Wang, Ruijun Zhang, and Yashuang Mu. Fu-SulfPred: Identification of protein S-sulfenylation sites by fusing forests via Chou’s general PseAAC. Journal of Theoretical Biology, 461(??):51–58, January 14, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305289.

Wu:2011:CBF

[WZWY11] Qinqin Wu, Weiqiang Zhou, Jiajun Wang, and Hong Yan. Correlation between the flexibility and periodic dinucleotide patterns in yeast nucleosomal DNA sequences. Journal of Theoretical Biology, 284(1):92–98, September 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003250.

Wen:2014:MSM

[WZY14] Jia Wen, YuYan Zhang, and Stephen S. T. Yau. kmer sparse matrix model for genetic sequence and its applications in sequence comparison. Journal of Theoretical Biology, 363(??):145–150, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004895.

Xia:2018:MCB

[XCW^{+}18] Yun Xia, Zhang Cheng, Shaobing Wang, Di Guan, and Feng Liu. Modeling the crosstalk between the circadian clock and ROS in neurospora crassa. Journal of Theoretical Biology, 458(??):125–132, December 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304442.

Xu:2011:MAS

[XDC^{+}11] Aiping Xu, Jan Dolfing, Thomas P. Curtis, Gary Montague, and Elaine Martin. Maintenance affects the stability of a two-tiered microbial ‘food chain’? Journal of Theoretical Biology, 276(1):35–41, May 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311000385.

Xu:2015:PPP

[XDD^{+}15] Yan Xu, Ya-Xin Ding, Jun Ding, Ling-Yun Wu, and Nai-Yang Deng. Phogly-PseAAC: Prediction of lysine phosphoglycerylation in proteins incorporating with position-specific propensity. Journal of Theoretical Biology, 379(??):10–15, August 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001794.

Xiao:2013:AMM

[XF13] Xiao Xiao and Gregor F. Fussmann. Armstrong–McGehee mechanism revisited: Competitive exclusion and coexistence of nonlinear consumers. Journal of Theoretical Biology, 339(??):26–35, December 21, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313002543.

Xu:2012:REP

[XG12] Yangqing Xu and Jeremy Gunawardena. Realistic enzymology for post-translational modification: Zero-order ultrasensitivity revisited. Journal of Theoretical Biology, 311(??):139–152, October 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003463.

Xu:2017:EAF

[XGZ17] Bingxiang Xu, Hao Ge, and Zhihua Zhang. An efficient and assumption-free method to approximate protein level distribution in the two-states gene expression model. Journal of Theoretical Biology, 433(??):1–7, November 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303934.

Xie:2011:MDM

[Xie11] Ping Xie. A model for the dynamics of mammalian family X DNA polymerases. Journal of Theoretical Biology, 277(1):111– 122, May 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001202.

Xie:2013:DMP

[Xie13a] Ping Xie. A dynamical model of programmed −1 ribosomal frameshifting. Journal of Theoretical Biology, 336 (??):119–131, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519313003457.

Xie:2013:DTO

[Xie13b] Ping Xie. Dynamics of tRNA occupancy and dissociation during translation by the ribosome. Journal of Theoretical Biology, 316(??):49–60, January 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005012.

Xie:2015:UMN

[Xie15] Ping Xie. A unified model of nucleic acid unwinding by the ribosome and the hexameric and monomeric DNA helicases. Journal of Theoretical Biology, 380(??):359–366, September 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519315002908.

Xie:2018:MDU

[Xie18] Ping Xie. A model of DNA unwinding dynamics by the RecBCD complex and its regulation by chi recognition. Journal of Theoretical Biology, 448(??):142–156, July 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301747.

Xu:2019:BCR

[XJ19] Jin Xu and Junghyo Jo. Broad cross-reactivity of the Tcell repertoire achieves specific and sufficiently rapid target searching. Journal of Theoretical Biology, 466(??):119–127, April 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300372.

Xue:2015:FDM

[XKCG15] Julian Z. Xue, Artem Kaznatcheev, Andre Costopoulos, and Frederic Guichard. Fidelity drive: a mechanism for chaperone proteins to maintain stable mutation rates in prokaryotes over evolutionary time. Journal of Theoretical Biology, 364(??):162–167, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005566.

Xue:2015:TSB

[XLG^{+}15] Fudong Xue, Junjie Liu, Longfang Guo, Lirong Zhang, and Qianzhong Li. Theoretical study on the bactericidal nature of nanopatterned surfaces. Journal of Theoretical Biology, 385(??):1–7, November 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931500404X.

Xi:2010:GLP

[XLL^{+}10] Lili Xi, Shuyan Li, Huanxiang Liu, Jiazhong Li, Beilei Lei, and Xiaojun Yao. Global and local prediction of protein folding rates based on sequence autocorrelation information. Journal of Theoretical Biology, 264(4):1159–1168, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001773.

Xue:2017:NPT

[XLLF17] Shi-Lei Xue, Shao-Zhen Lin, Bo Li, and Xi-Qiao Feng. A nonlinear poroelastic theory of solid tumors with glycosaminoglycan swelling. Journal of Theoretical Biology, 433(??):49–56, November 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303958.

Xu:2019:IBE

[XLSF19] Zhongpu Xu, Kezan Li, Mengfeng Sun, and Xinchu Fu. Interaction between epidemic spread and collective behavior in scale-free networks with community structure. Journal of Theoretical Biology, 462(??):122–133, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305502.

Xie:2011:TGR

[XM11] Guosen Xie and Zhongxi Mo. Three 3d graphical representations of DNA primary sequences based on the classifications of DNA bases and their applications. Journal of Theoretical Biology, 269(1):123–130, January 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005527.

Xiao:2013:IPI

[XMWC13] Xuan Xiao, Jian-Liang Min, Pu Wang, and Kuo-Chen Chou. iCDI-PseFpt: Identify the channel-drug interaction in cellular networking with PseAAC and molecular fingerprints. Journal of Theoretical Biology, 337(??):71–79, November 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003871.

Xiaohui:2013:UCC

[XNJ^{+}13] Niu Xiaohui, Li Nana, Xia Jingbo, Chen Dingyan, Peng Yuehua, Xiao Yang, Wei Weiquan, Wang Dongming, and Wang Zengzhen. Using the concept of Chou’s pseudo amino acid composition to predict protein solubility: An approach with entropies in information theory. Journal of Theoretical Biology, 332(??):211–217, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313001148.

Xu:2017:STS

[XNW17] Xuying Xu, Li Ni, and Rubin Wang. Synchronous transitions of up and down states in a network model based on stimulations. Journal of Theoretical Biology, 412(??):130–137, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316303538.

Xu:2013:LDF

[XS13] Na Xu and Mao Sun. Lateral dynamic flight stability of a model bumblebee in hovering and forward flight. Journal of Theoretical Biology, 319(??):102–115, February 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006224.

Xue:2012:NBM

[XSCS12] Ling Xue, H. Morgan Scott, Lee W. Cohnstaedt, and Caterina Scoglio. A network-based meta-population approach to model Rift valley fever epidemics. Journal of Theoretical Biology, 306(??):129–144, August 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200210X.

Xenitidis:2017:EAT

[XSKA17] P. Xenitidis, I. Seimenis, S. Kakolyris, and A. Adamopoulos. Evaluation of artificial time series microarray data for dynamic gene regulatory network inference. Journal of Theoretical Biology, 426(??):1–16, August 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317302187.

Xu:2016:PSA

[XSLZ16] Chunrui Xu, Dandan Sun, Shenghui Liu, and Yusen Zhang. Protein sequence analysis by incorporating modified chaos game and physicochemical properties into Chou’s general pseudo amino acid composition. Journal of Theoretical Biology, 406(??):105–115, October 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316301680.

Xu:2012:IAV

[XSMF12] Dashun Xu, Gregory J. Sandland, Dennis J. Minchella, and Zhilan Feng. Interactions among virulence, coinfection and drug resistance in a complex life-cycle parasite. Journal of Theoretical Biology, 304(??):197–210, July 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001725.

Xiao:2014:TPN

[XSTW14] Yanni Xiao, Xiaodan Sun, Sanyi Tang, and Jianhong Wu. Transmission potential of the novel avian influenza A (H7N9) infection in mainland China. Journal of Theoretical Biology, 352(??):1–5, July 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314001222.

Xiao:2013:PHA

[XTZ^{+}13] Yanni Xiao, Sanyi Tang, Yicang Zhou, Robert J. Smith, Jianhong Wu, and Ning Wang. Predicting the HIV/AIDS epidemic and measuring the effect of mobility in mainland China. Journal of Theoretical Biology, 317(??):271–285, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005152.

Xu:2017:QPF

[XW17] Li Xu and Jin Wang. Quantifying the potential and flux landscapes of multi-locus evolution. Journal of Theoretical Biology, 422(??):31–49, June 7, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301716.

Xiao:2011:IVM

[XWC11] Xuan Xiao, Zhi-Cheng Wu, and Kuo-Chen Chou. iLoc-virus: a multi-label learning classifier for identifying the subcellular localization of virus proteins with both single and multiple sites. Journal of Theoretical Biology, 284(1):42–51, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003043.

Xu:2010:LAS

[XWD^{+}10] Yan Xu, Xiao-Bo Wang, Jun Ding, Ling-Yun Wu, and Nai-Yang Deng. Lysine acetylation sites prediction using an ensemble of support vector machine classifiers. Journal of Theoretical Biology, 264(1):130–135, May 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000172.

Xu:2014:PPM

[XWW^{+}14] Yan Xu, Xiaobo Wang, Yongcui Wang, Yingjie Tian, Xiaojian Shao, Ling-Yun Wu, and Naiyang Deng. Prediction of posttranslational modification sites from amino acid sequences with kernel methods. Journal of Theoretical Biology, 344(??):78–87, March 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005390.

Xu:2019:LOW

[XWW^{+}19] Xiaoyan Xu, Jianing Wu, Kai Wang, Yunqiang Yang, and Shaoze Yan. Locking of the operculum in a water snail: Theoretical modeling and applications for mechanical sealing. Journal of Theoretical Biology, 464(??):104–111, March 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318306362.

Xu:2010:BRV

[XWZ10] Zhaojin Xu, Zhen Wang, and Lianzhong Zhang. Bounded rationality in volunteering public goods games. Journal of Theoretical Biology, 264(1):19–23, May 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000378.

Xu:2017:PSP

[XXD^{+}17] Qian Xu, Yi Xiong, Hao Dai, Kotni Meena Kumari, Qin Xu, Hong-Yu Ou, and Dong-Qing Wei. PDC-SGB: Prediction of effective drug combinations using a stochastic gradient boosting algorithm. Journal of Theoretical Biology, 417(??):1–7, March 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730019X.

Xu:2014:CPE

[XY14] Ying Xu and Jiang-Nan Yang. Cooperation promotes the evolution of separate sexes from hermaphrodites with unitary growth. Journal of Theoretical Biology, 341(??):102–110, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004797.

Xi:2015:MMS

[XYN15] Man Xi, Guohong Yun, and B. Narsu. A mathematical model on stress-strain of the epimysium of skeletal muscles. Journal of Theoretical Biology, 365(??):175–180, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004937.

Xue:2015:FSI

[XYZ15] Xu Xue, Wang Yongjun, and Li Zhihong. Folding of SAM-II riboswitch explored by replica-exchange molecular dynamics simulation. Journal of Theoretical Biology, 365(??):265–269, January 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314006122.

Yonenoh:2014:SOP

[YA14] Hiroto Yonenoh and Eizo Akiyama. Selection of opponents in the prisoner’s dilemma in dynamic networks: An experimental approach. Journal of Theoretical Biology, 351(??):25–36, June 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000678.

Yang:2017:DCN

[YAK17] Jienian Yang, David E. Axelrod, and Natalia L. Komarova. Determining the control networks regulating stem cell lineages in colonic crypts. Journal of Theoretical Biology, 429(??):190–203, September 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317303107.

Yamaguchi:2016:TRS

[Yam16a] Sachi Yamaguchi. Time required for sex change in teleost fishes: Hormonal dynamics shaped by selection. Journal of Theoretical Biology, 407(??):339–348, October 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301953.

Yamazaki:2016:CMP

[Yam16b] Kazuo Yamazaki. Caterpillar mimicry by plant galls as a visual defense against herbivores. Journal of Theoretical Biology, 404(??):10–14, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301163.

Yan:2010:ISE

[Yan10a] Ping Yan. Impulsive SUI epidemic model for HIV/AIDS with chronological age and infection age. Journal of Theoretical Biology, 265(2):177–184, July 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001943.

Yang:2010:CEA

[Yan10b] Jiang-Nan Yang. Cooperation and the evolution of anisogamy. Journal of Theoretical Biology, 264(1):24–36, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000317.

Yates:2014:DCM

[Yat14] Christian A. Yates. Discrete and continuous models for tissue growth and shrinkage. Journal of Theoretical Biology, 350(??):37–48, June 7, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314000605.

Yi:2017:CAT

[YBC17] Su Do Yi, Seung Ki Baek, and Jung-Kyoo Choi. Combination with anti-tit-for-tat remedies problems of tit-for-tat. Journal of Theoretical Biology, 412(??):1–7, January 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316303149.

Yao:2019:NAQ

[YBH^{+}19] Yu-Xiang Yao, Zhi-Tong Bing, Liang Huang, Zi-Gang Huang, and Ying-Cheng Lai. A network approach to quantifying radiotherapy effect on cancer: Radiosensitive gene group centrality. Journal of Theoretical Biology, 462(??):528–536, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318305915.

You:2017:SVN

[YBT^{+}17] Li You, Joel S. Brown, Frank Thuijsman, Jessica J. Cunningham, Robert A. Gatenby, Jingsong Zhang, and Kateřina Staňková. Spatial vs. non-spatial eco-evolutionary dynamics in a tumor growth model. Journal of Theoretical Biology, 435(??):78–97, December 21, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303971.

Yousef:2013:NMB

[YC13] Abdulaziz Yousef and Nasrollah Moghadam Charkari. A novel method based on new adaptive LVQ neural network for predicting protein–protein interactions from protein sequences. Journal of Theoretical Biology, 336(??):231–239, November 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003172.

Yousef:2015:SNS

[YC15] Abdulaziz Yousef and Nasrollah Moghadam Charkari. SFM: a novel sequence-based fusion method for disease genes identification and prioritization. Journal of Theoretical Biology, 383(??):12–19, October 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315003392.

Yeoh:2016:MMP

[YCB16] Jing Wui Yeoh, Alberto Corrias, and Martin L. Buist. A mechanistic model of a PDGFR α + cell. Journal of Theoretical Biology, 408(??):127–136, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302387.

Yan:2017:CMC

[YCH^{+}17a] Ada W. C. Yan, Pengxing Cao, Jane M. Heffernan, Jodie McVernon, Kylie M. Quinn, Nicole L. La Gruta, Karen L. Laurie, and James M. McCaw. Corrigendum to “Modelling cross-reactivity and memory in the cellular adaptive immune response to influenza infection in the host” [ J.Theor. Biol. 413 (2017) 34–49]. Journal of Theoretical Biology, 419(??):394, April 21, 2017. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317301273. See [YCH^{+}17b].

Yan:2017:MCR

[YCH^{+}17b] Ada W. C. Yan, Pengxing Cao, Jane M. Heffernan, Jodie McVernon, Kylie M. Quinn, Nicole L. La Gruta, Karen L. Laurie, and James M. McCaw. Modelling cross-reactivity and memory in the cellular adaptive immune response to influenza infection in the host. Journal of Theoretical Biology, 413(??):34–49, January 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519316303629. See corrigendum [YCH^{+}17a].

Yoon:2017:MMM

[YCL^{+}17] Dennis J. Yoon, Kevin Y. Chen, André M. Lopes, April A. Pan, Joseph Shiloach, Anne B. Mason, and Daniel T. Kamei. Mathematical modeling of mutant transferrin-CRM107 molecular conjugates for cancer therapy. Journal of Theoretical Biology, 416(??):88–98, March 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300097.

Yu:2015:AMC

[YCR^{+}15] Jia-Feng Yu, Qing-Li Chen, Jing Ren, Yan-Ling Yang, Ji-Hua Wang, and Xiao Sun. Analysis of the multi-copied genes and the impact of the redundant protein coding sequences on gene annotation in prokaryotic genomes. Journal of Theoretical Biology, 376(??):8–14, July 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001563.

Yin:2014:MDS

[YCY14] Changchuan Yin, Ying Chen, and Stephen S.-T. Yau. A measure of DNA sequence similarity by Fourier transform with applications on hierarchical clustering. Journal of Theoretical Biology, 359(??):18–28, October 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314003324.

Yenkie:2014:OCP

[YD14] Kirti M. Yenkie and Urmila Diwekar. Optimal control for predicting customized drug dosage for superovulation stage of in vitro fertilization. Journal of Theoretical Biology, 355(??):219–228, August 21, 2014. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314002331.

Yenkie:2015:UCD

[YD15] Kirti M. Yenkie and Urmila Diwekar. Uncertainty in clinical data and stochastic model for in vitro fertilization. Journal of Theoretical Biology, 367(??):76–85, February 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006365.

Yu:2013:PSN

[YDC^{+}13] Chenglong Yu, Mo Deng, Shiu-Yuen Cheng, Shek-Chung Yau, Rong L. He, and Stephen S.-T. Yau. Protein space: a natural method for realizing the nature of protein universe. Journal of Theoretical Biology, 318(??):197–204, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005802.

Young:2012:CCC

[YFB^{+}12] Todd R. Young, Bastien Fernandez, Richard Buckalew, Gregory Moses, and Erik M. Boczko. Clustering in cell cycle dynamics with general response/signaling feedback. Journal of Theoretical Biology, 292(??):103–115, January 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311005091.

Yeh:2012:GPM

[YG12] Angela Yi-Chen Yeh and Andy Gardner. A general ploidy model for the evolution of helping in viscous populations. Journal of Theoretical Biology, 304(??):297–303, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001774.

Yang:2018:NPM

[YG18] Yongxiao Yang and Xinqi Gong. A new probability method to understand protein–protein interface formation mechanism at amino acid level. Journal of Theoretical Biology, 436(??):18–25, January 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304435.

Yu:2010:SIB

[YGL^{+}10] Lezheng Yu, Yanzhi Guo, Yizhou Li, Gongbing Li, Menglong Li, Jiesi Luo, Wenjia Xiong, and Wenli Qin. SecretP: Identifying bacterial secreted proteins by fusing new features into Chou’s pseudo-amino acid composition. Journal of Theoretical Biology, 267(1):1–6, November 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003991.

Yang:2012:CPN

[YGMT12] Jinliang Yang, Rui Gao, Max Q.-H. Meng, and Tzyh-Jong Tarn. Colored Petri nets to model gene mutation and amino acids classification. Journal of Theoretical Biology, 300(??):183–192, May 7, 2012. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519312000379.

Yamaguchi:2014:ESD

[YHI14] Sachi Yamaguchi, Jens T. Høeg, and Yoh Iwasa. Evolution of sex determination and sexually dimorphic larval sizes in parasitic barnacles. Journal of Theoretical Biology, 347(??):7–16, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000204.

Yu:2014:VGP

[YHY14] Chenglong Yu, Rong Lucy He, and Stephen S.-T. Yau. Viral genome phylogeny based on Lempel–Ziv complexity and Hausdorff distance. Journal of Theoretical Biology, 348(??):12–20, May 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000393.

Yang:2014:OPS

[YHZ14] Guoli Yang, Jincai Huang, and Weiming Zhang. Older partner selection promotes the prevalence of cooperation in evolutionary games. Journal of Theoretical Biology, 359(??):171–183, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003567.

Yamaguchi:2016:SNI

[YI16] Ryo Yamaguchi and Yoh Iwasa. Smallness of the number of incompatibility loci can facilitate parapatric speciation. Journal of Theoretical Biology, 405(??):36–45, September 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005238.

Yamaguchi:2017:TPP

[YI17] Ryo Yamaguchi and Yoh Iwasa. A tipping point in parapatric speciation. Journal of Theoretical Biology, 421(??):81– 92, May 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301376.

Yamaguchi:2018:TDS

[YI18a] Sachi Yamaguchi and Yoh Iwasa. Temperature-dependent sex determination, realized by hormonal dynamics with enzymatic reactions sensitive to ambient temperature. Journal of Theoretical Biology, 453(??):146–155, September 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318302637.

Yamaguchi:2018:WBS

[YI18b] Sachi Yamaguchi and Yoh Iwasa. Why is bidirectional sex change rare? Journal of Theoretical Biology, 453(??):136– 145, September 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519318302625.

Yin:2017:IRD

[Yin17] Changchuan Yin. Identification of repeats in DNA sequences using nucleotide distribution uniformity. Journal of Theoretical Biology, 412(??):138–145, January 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931630354X.

Yamada:2017:IKO

[YIS^{+}17] Masatsugu Yamada, Masashi Iwanaga, Kansuporn Sriyudthsak, Masami Y. Hirai, and Fumihide Shiraishi. Investigation of kinetic-order sensitivities in metabolic reaction networks. Journal of Theoretical Biology, 415(??):32–40, February 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316304118.

Yakimov:2014:MAN

[YISG14] Basil N. Yakimov, Dmitry I. Iudin, Leonid A. Solntsev, and David B. Gelashvili. Multifractal analysis of neutral community spatial structure. Journal of Theoretical Biology, 343(??):44–53, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519313005092.

Yokoi:2019:MMP

[YiTS19] Hiroki Yokoi, Kei ichi Tainaka, and Kazunori Sato. Metapopulation model for a prey-predator system: Nonlinear migration due to the finite capacities of patches. Journal of Theoretical Biology, 477(??):24–35, September 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302255.

Yamauchi:2011:JES

[YK11] Atsushi Yamauchi and Yutaka Kobayashi. Joint evolution of sex ratio and reproductive group size under local mate competition with inbreeding depression. Journal of Theoretical Biology, 270(1):127–133, February 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931000593X.

Yoshimura:2016:NMM

[YKO^{+}16] Kenji Yoshimura, Ryo Kobayashi, Tomohisa Ohmura, Yoshinaga Kajimoto, and Takashi Miura. A new mathematical model for pattern formation by cranial sutures. Journal of Theoretical Biology, 408(??):66–74, November 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302375.

Yang:2016:TAT

[YLC16] Cheng-Hong Yang, Yu-Da Lin, and Li-Yeh Chuang. tR-NAfeature: An algorithm for tRNA features to identify tRNA genes in DNA sequences. Journal of Theoretical Biology, 404(??):251–261, September 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316301412.

Yu:2017:CSG

[YLF^{+}17] Yang Yu, Jie Liu, Nuan Feng, Bo Song, and Zeyu Zheng. Combining sequence and gene ontology for protein module detection in the weighted network. Journal of Theoretical Biology, 412(??):107–112, January 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316303514.

Yong:2012:HCM

[YLH12] Kamuela E. Yong, Yi Li, and Stephen D. Hendrix. Habitat choice of multiple pollinators in almond trees and its potential effect on pollen movement and productivity: a theoretical approach using the Shigesada–Kawasaki–Teramoto model. Journal of Theoretical Biology, 305(??):103–109, July 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931200197X.

Yang:2014:HPC

[YLL^{+}14] Lei Yang, Yingli Lv, Tao Li, Yongchun Zuo, and Wei Jiang. Human proteins characterization with subcellular localizations. Journal of Theoretical Biology, 358(??):61–73, October 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314002793.

Youssefpour:2012:MMC

[YLLL12] H. Youssefpour, X. Li, A. D. Lander, and J. S. Lowengrub. Multispecies model of cell lineages and feedback control in solid tumors. Journal of Theoretical Biology, 304(??):39–59, July 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001154.

Ying:2014:SCL

[YLW^{+}14] Zhixia Ying, Jinbao Liao, Shichang Wang, Hui Lu, Yongjie Liu, Liang Ma, and Zhenqing Li. Species coexistence in a lattice-structured habitat: Effects of species dispersal and interactions. Journal of Theoretical Biology, 359(??):184–191, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003373.

Yang:2010:MMB

[YLWZ10] Jie Yang, Jia-Huang Li, Jin Wang, and Chen-Yu Zhang. Molecular modeling of BAD complex resided in a mitochondrion integrating glycolysis and apoptosis. Journal of Theoretical Biology, 266(2):231–241, September 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310002936.

Yamazaki:2015:DWT

[YLY15] Kazuo Yamazaki and Simcha Lev-Yadun. Dense white trichome production by plants as possible mimicry of arthropod silk or fungal hyphae that deter herbivory. Journal of Theoretical Biology, 364(??):1–6, January 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314005323.

Yamashita:2014:QAC

[YM14] Satoshi Yamashita and Tatsuo Michiue. Quantitative analysis of cell arrangement indicates early differentiation of the neural region during xenopus gastrulation. Journal of Theoretical Biology, 346(??):1–7, April 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314000034.

Yamashita:2016:BPP

[YM16] Satoshi Yamashita and Tatsuo Michiue. Boundary propagation of planar cell polarity is robust against cell packing pattern. Journal of Theoretical Biology, 410(??):44–54, December 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316302995.

Yang:2012:MFC

[YMW12] Xiaofeng Yang, Vladimir Mironov, and Qi Wang. Modeling fusion of cellular aggregates in biofabrication using phase field theories. Journal of Theoretical Biology, 303(??):110–118, June 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312001208.

Yu:2018:TDP

[YMZ18] Don Yu, Neal Madras, and Huaiping Zhu. Temperaturedriven population abundance model for Culex pipiens and Culex restuans (Diptera: Culicidae). Journal of Theoretical Biology, 443(??):28–38, April 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300407.

Yang:2010:SED

[YNY^{+}10] Zhiwei Yang, Yingkun Nie, Gang Yang, Yuangang Zu, Yujie Fu, and Lijun Zhou. Synergistic effects in the designs of neuraminidase ligands: Analysis from docking and molecular dynamics studies. Journal of Theoretical Biology, 267(3):363–374, December 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004522.

Yang:2017:DNS

[YRD17] Qian Yang, Tim Rogers, and Jonathan H. P. Dawes. Demographic noise slows down cycles of dominance. Journal of Theoretical Biology, 432(??):157–168, November 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303569.

Youssef:2011:IBA

[YS11] Mina Youssef and Caterina Scoglio. An individual-based approach to SIR epidemics in contact networks. Journal of Theoretical Biology, 283(1):136–144, August 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002815.

Yakovlev:2014:MRC

[YS14] A. G. Yakovlev and V. A. Shuvalov. Modeling of reversible charge separation in reaction centers of photosynthesis: An incoherent approach. Journal of Theoretical Biology, 343(??):92–101, February 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313005365.

Yan:2014:AAC

[YSH^{+}14] Wenying Yan, Maomin Sun, Guang Hu, Jianhong Zhou, Wenyu Zhang, Jiajia Chen, Biao Chen, and Bairong Shen. Amino acid contact energy networks impact protein structure and evolution. Journal of Theoretical Biology, 355(??):95–104, August 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001799.

Yang:2017:MPF

[YSI17] Jinyou Yang, Yuji Shimogonya, and Takuji Ishikawa. Mixing and pumping functions of the intestine of zebrafish larvae. Journal of Theoretical Biology, 419(??):152–158, April 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317300607.

Yang:2018:WCS

[YSI18] Jinyou Yang, Yuji Shimogonya, and Takuji Ishikawa. What causes the spatial heterogeneity of bacterial flora in the intestine of zebrafish larvae? Journal of Theoretical Biology, 446(??):101–109, June 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318301218.

Yamaguchi:2013:SPF

[YSST13] Sachi Yamaguchi, Satoko Seki, Kota Sawada, and Satoshi Takahashi. Small and poor females change sex: a theoretical and empirical study on protogynous sex change in a triggerfish under varying resource abundance. Journal of Theoretical Biology, 317(??):186–191, January 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312005590.

Yu:2014:NDF

[YST14] Jianshe Yu, Qiwen Sun, and Moxun Tang. The nonlinear dynamics and fluctuations of mRNA levels in crosstalking pathway activated transcription. Journal of Theoretical Biology, 363(??):223–234, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004792.

Yamaguchi:2013:DMH

[YSYI13] Sachi Yamaguchi, Kota Sawada, Yoichi Yusa, and Yoh Iwasa. Dwarf males and hermaphrodites can coexist in marine sedentary species if the opportunity to become a dwarf male is limited. Journal of Theoretical Biology, 334(??):101–108, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002567.

Yamauchi:2012:BEE

[YT12] Atsushi Yamauchi and Arndt Telschow. Bistability of endosymbiont evolution of genome size and host sex control. Journal of Theoretical Biology, 309(??):58–66, September 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312002561.

Yan:2016:MCH

[YTGW16] Qinling Yan, Sanyi Tang, Sandra Gabriele, and Jianhong Wu. Media coverage and hospital notifications: Correlation analysis and optimal media impact duration to manage a pandemic. Journal of Theoretical Biology, 390(??):1–13, February 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315005366.

Yamauchi:2010:ECS

[YTK10] Atsushi Yamauchi, Arndt Telschow, and Yutaka Kobayashi. Evolution of cytoplasmic sex ratio distorters: Effect of paternal transmission. Journal of Theoretical Biology, 266(1):79–87, September 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310003085.

Mier-y-Teran-Romero:2013:BSI

[yTRSC13] Luis Mier y Teran-Romero, Ira B. Schwartz, and Derek A. T. Cummings. Breaking the symmetry: Immune enhancement increases persistence of dengue viruses in the presence of asymmetric transmission rates. Journal of Theoretical Biology, 332(??):203–210, September 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300204X.

Yamauchi:2018:SPG

[YvBS18] Atsushi Yamauchi, Minus van Baalen, and Maurice W. Sabelis. Spatial patterns generated by simultaneous cooperation and exploitation favour the evolution of altruism. Journal of Theoretical Biology, 441(??):58–67, March 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931730574X.

Yang:2013:LRM

[YW13a] Xiwu Yang and Tianming Wang. Linear regression model of short k-word: a similarity distance suitable for biological sequences with various lengths. Journal of Theoretical Biology, 337(??):61–70, November 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931300369X.

Yang:2013:NSM

[YW13b] Xiwu Yang and Tianming Wang. A novel statistical measure for sequence comparison on the basis of k-word counts. Journal of Theoretical Biology, 318(??):91–100, February 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005759.

Yang:2013:BMW

[YWP13] Le Yang, Tarynn M. Witten, and Ramana M. Pidaparti. A biomechanical model of wound contraction and scar formation. Journal of Theoretical Biology, 332(??):228–248, September 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313001173.

Yang:2014:AIT

[YWW^{+}14] Lei Yang, Jizhe Wang, Huiping Wang, Yingli Lv, Yongchun Zuo, and Wei Jiang. Analysis and identification of toxin targets by topological properties in protein–protein interaction network. Journal of Theoretical Biology, 349(??):82–91, May 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000629.

Yang:2016:CBN

[YWZ^{+}16] Lei Yang, Shiyuan Wang, Meng Zhou, Xiaowen Chen, Yongchun Zuo, and Yingli Lv. Characterization of Bio-Plex network by topological properties. Journal of Theoretical Biology, 409(??):148–154, November 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316302703.

Yuan:2015:RSF

[YX15] Jia Yuan and Bin Xue. Role of structural flexibility in the evolution of emerin. Journal of Theoretical Biology, 385(??):102–111, November 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315004026.

Yan:2014:TWD

[YXH^{+}14] Jing Yan, Yan Xu, Brian Hu, Sammy Alnajm, Lina Liu, Yin Lu, Zhiguang Sun, and Feng Cheng. TIBS: a web database to browse gene expression in irritable bowel syndrome. Journal of Theoretical Biology, 354(??):48–53, August 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001659.

Yin:2015:IMW

[YY15] Changchuan Yin and Stephen S.-T. Yau. An improved model for whole genome phylogenetic analysis by Fourier transform. Journal of Theoretical Biology, 382(??):99–110, October 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315003124.

Yasui:2018:BHA

[YY18] Yukio Yasui and Jin Yoshimura. Bet-hedging against male-caused reproductive failures may explain ubiquitous cuckoldry in female birds. Journal of Theoretical Biology, 437(??):214–221, January 21, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317304988.

Yamauchi:2019:TCC

[YYB^{+}19] Atsushi Yamauchi, Takahiro Yamagishi, Ross Booton, Arndt Telschow, and Gaku Kudo. Theory of coevolution of cytoplasmic male-sterility, nuclear restorer and selfing. Journal of Theoretical Biology, 477(??):96–107, September 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319302383.

Yang:2019:PPS

[YYC19] Cheng-Hong Yang, Huai-Shuo Yang, and Li-Yeh Chuang. PB-MDR: a particle swarm optimization-based multifactor dimensionality reduction for the detection of multilocus interactions. Journal of Theoretical Biology, 461(??):68–75, January 14, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931830482X.

Yao:2014:NDP

[YYH^{+}14] Yuhua Yao, Shoujiang Yan, Jianning Han, Qi Dai, and Ping an He. A novel descriptor of protein sequences and its application. Journal of Theoretical Biology, 347(??):109–117, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000101.

Yamaguchi:2013:SSD

[YYST13] Sachi Yamaguchi, Yoichi Yusa, Kota Sawada, and Satoshi Takahashi. Sexual systems and dwarf males in barnacles: Integrating life history and sex allocation theories. Journal of Theoretical Biology, 320(??):1–9, March 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931200625X.

Yu:2019:PCI

[YZ19] Yang Yu and Zeyu Zheng. Protein complex identification based on weighted PPI network with multi-source information. Journal of Theoretical Biology, 477(??):77–83, September 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319302322.

Yang:2016:ELA

[YZFY16] Lianping Yang, Xiangde Zhang, Haoyue Fu, and Chenhui Yang. An estimator for local analysis of genome based on the minimal absent word. Journal of Theoretical Biology, 395(??):23–30, April 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316000564.

Yan:2018:EBS

[YZMY18] Shuixian Yan, Yu Zhang, Junling Ma, and Sanling Yuan. An edge-based SIR model for sexually transmitted diseases on the contact network. Journal of Theoretical Biology, 439(??):216–225, February 14, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519317305349.

Yang:2010:BWS

[YZW10] Lianping Yang, Xiangde Zhang, and Tianming Wang. The Burrows–Wheeler similarity distribution between biological sequences based on Burrows–Wheeler transform. Journal of Theoretical Biology, 262(4):742–749, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519309005220.

Yang:2015:NRE

[YZX15] Guoli Yang, Weiming Zhang, and Baoxin Xiu. Neighbourhood reaction in the evolution of cooperation. Journal of Theoretical Biology, 372(??):118–127, May 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000934.

Yang:2016:NCM

[YZY^{+}16] Jing Yang, Tong-Jun Zhao, Chang-Qing Yuan, Jing-Hui Xie, and Fang-Fang Hao. A nonlinear competitive model of the prostate tumor growth under intermittent androgen suppression. Journal of Theoretical Biology, 404(??):66–72, September 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316301230.

Yang:2012:AFC

[YZZ12] Lianping Yang, Xiangde Zhang, and Hegui Zhu. Alignment free comparison: Similarity distribution between the DNA primary sequences based on the shortest absent word. Journal of Theoretical Biology, 295(??):125–131, February 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311005911.

Yang:2013:AFC

[YZZ13] Lianping Yang, Xiangde Zhang, and Hegui Zhu. Alignment free comparison: k word voting model and its applications. Journal of Theoretical Biology, 335(??):276–282, October 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313003202.

Zabet:2011:NFP

[Zab11] Nicolae Radu Zabet. Negative feedback and physical limits of genes. Journal of Theoretical Biology, 284(1):82–91, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311003201.

Zadeh:2011:IPD

[Zad11] Kouroush Sadegh Zadeh. An integro-partial differential equation for modeling biofluids flow in fractured biomaterials. Journal of Theoretical Biology, 273(1):72–79, March 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310007022.

Zhang:2015:FTI

[ZADB15] Lai Zhang, Ken H. Andersen, Ulf Dieckmann, and Åke Brännström. Four types of interference competition and their impacts on the ecology and evolution of size-structured populations and communities. Journal of Theoretical Biology, 380(??):280–290, September 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519315002581.

Zupanc:2019:GMM

[ZAL^{+}19] Günther K. H. Zupanc, Stephanie M. Amaro, Dávid Lehotzky, Frederick B. Zupanc, and Nicholas Y. Leung. Glia-mediated modulation of extracellular potassium concentration determines the sexually dimorphic output frequency of a model brainstem oscillator. Journal of Theoretical Biology, 471(??):117–124, June 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519319301146.

Zagkos:2019:MMD

[ZARK19] Loukas Zagkos, Mark Mc Auley, Jason Roberts, and Nikos I. Kavallaris. Mathematical models of DNA methylation dynamics: Implications for health and ageing. Journal of Theoretical Biology, 462(??):184–193, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305538.

Zambrano:2014:SMN

[ZBA14] Samuel Zambrano, Marco E. Bianchi, and Alessandra Agresti. A simple model of NF-κB dynamics reproduces experimental observations. Journal of Theoretical Biology, 347(??):44–53, April 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314000241.

Zinn-Bjorkman:2018:MFR

[ZBA18] Leif Zinn-Björkman and Frederick R. Adler. Modeling factors that regulate cell cooperativity in the zebrafish posterior lateral line primordium. Journal of Theoretical Biology, 444(??):93–99, May 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318300614.

Zhao:2014:NBB

[ZC14] Xing-Ming Zhao and Luonan Chen. Network-based biomarkers for complex diseases. Journal of Theoretical Biology, 362 (??):1–2, December 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519314004007.

Zhang:2014:NBM

[ZCA^{+}14] Suhua Zhang, Yafei Chen, Hailong An, Hui Liu, Junwei Li, Chunli Pang, Qing Ji, and Yong Zhan. A novel biophysical model on calcium and voltage dual dependent gating of calcium-activated chloride channel. Journal of Theoretical Biology, 355(??):229–235, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002136.

Zhai:2017:HAP

[ZCAB17] Jing-Xuan Zhai, Tian-Jie Cao, Ji-Yong An, and Yong-Tao Bian. Highly accurate prediction of protein selfinteractions by incorporating the average block and PSSM information into the general PseAAC. Journal of Theoretical Biology, 432(??):80–86, November 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317303752.

Zubkov:2015:SAM

[ZCS^{+}15] V. S. Zubkov, A. N. Combes, K. M. Short, J. Lefevre, N. A. Hamilton, I. M. Smyth, M. H. Little, and H. M. Byrne. A spatially-averaged mathematical model of kidney branching morphogenesis. Journal of Theoretical Biology, 379(??):24–37, August 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315001782.

Zamora-Chimal:2012:IFL

[ZCSRG12] Criseida Zamora-Chimal, Moisés Santillán, and Jesús Rodríguez-González. Influence of the feedback loops in the trp operon of B. subtilis on the system dynamic response and noise amplitude. Journal of Theoretical Biology, 310(??):119–131, October 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312003013.

Zenati:2018:GSA

[ZCT18] Abdelhafid Zenati, Messaoud Chakir, and Mohamed Tadjine. Global stability analysis and optimal control therapy of blood cell production process (hematopoiesis) in acute myeloid leukemia. Journal of Theoretical Biology, 458(??):15–30, December 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318304314.

Zhang:2013:DIA

[ZCW13] Jianlei Zhang, Tianguang Chu, and Franz J. Weissing. Does insurance against punishment undermine cooperation in the evolution of public goods games? Journal of Theoretical Biology, 321(??):78–82, March 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006558.

Zhu:2013:ESC

[ZCXF13] Guanghu Zhu, Guanrong Chen, Xin-Jian Xu, and Xinchu Fu. Epidemic spreading on contact networks with adaptive weights. Journal of Theoretical Biology, 317(??):133–139, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519312005140.

Zhang:2018:PPS

[ZD18] Shengli Zhang and Xin Duan. Prediction of protein subcellular localization with oversampling approach and Chou’s general PseAAC. Journal of Theoretical Biology, 437(??):239–250, January 21, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931730499X.

Zhong:2012:MEP

[ZDF^{+}12] Xue Zhong, Jason Droesch, Randal Fox, Eva M. Top, and Stephen M. Krone. On the meaning and estimation of plasmid transfer rates for surface-associated and well-mixed bacterial populations. Journal of Theoretical Biology, 294(??):144–152, February 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311005571.

Zhang:2010:PPA

[ZDG^{+}10] Ning Zhang, Guangyou Duan, Shan Gao, Jishou Ruan, and Tao Zhang. Prediction of the parallel/antiparallel orientation of beta-strands using amino acid pairing preferences and support vector machines. Journal of Theoretical Biology, 263(3):360–368, April 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251930900592X.

Zhao:2011:NCM

[ZDY11] Bo Zhao, Victor Duan, and Stephen S.-T. Yau. A novel clustering method via nucleotide-based Fourier power spectrum analysis. Journal of Theoretical Biology, 279(1):83–89, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311001755.

Zook:2011:FWO

[ZEJA11] Alexander E. Zook, Anna Eklof, Ute Jacob, and Stefano Allesina. Food webs: Ordering species according to body size yields high degree of intervality. Journal of Theoretical Biology, 271(1):106–113, February 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310006405.

Zhang:2017:DRI

[ZFWK17] Jie Zhang, Angela G. Fleischman, Dominik Wodarz, and Natalia L. Komarova. Determining the role of inflammation in the selection of JAK2 mutant cells in myeloproliferative neoplasms. Journal of Theoretical Biology, 425(??):43–52, July 21, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317302229.

Zhang:2010:CSB

[ZG10a] Jiang Zhang and Liangpeng Guo. Corrigendum to “Scaling behaviors of weighted food webs as energy transportation networks” [J. Theor. Biol. 264 (2010) 760–770]. Journal of Theoretical Biology, 266(3):483, October 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310003164. See [ZG10b].

Zhang:2010:SBW

[ZG10b] Jiang Zhang and Liangpeng Guo. Scaling behaviors of weighted food webs as energy transportation networks. Journal of Theoretical Biology, 264(3):760–770, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001475. See corrigen-

dum [ZG10a].

Zumsande:2010:BCM

[ZG10c] Martin Zumsande and Thilo Gross. Bifurcations and chaos in the MAPK signaling cascade. Journal of Theoretical Biology, 265(3):481–491, August 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931000216X.

Zaman:2019:MBF

[ZG19] Salma Zaman and Mahwish Gul. Magnetohydrodynamic bioconvective flow of Williamson nanofluid containing gyrotactic microorganisms subjected to thermal radiation and Newtonian conditions. Journal of Theoretical Biology, 479(??):22–28, October 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300840.

Zaoli:2019:PNS

[ZGG^{+}19] Silvia Zaoli, Andrea Giometto, Jonathan Giezendanner, Amos Maritan, and Andrea Rinaldo. On the probabilistic nature of the species-area relation. Journal of Theoretical Biology, 462(??):391–407, February 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305848.

Zhang:2010:RDB

[ZGS^{+}10] Lihai Zhang, Bruce S. Gardiner, David W. Smith, Peter Pivonka, and Alan J. Grodzinsky. On the role of diffusible binding partners in modulating the transport and concentration of proteins in tissues. Journal of Theoretical Biology, 263(1):20–29, March 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005657.

Zhao:2016:MWF

[ZGW16] Lei Zhao, Toni I. Gossmann, and David Waxman. A modified Wright–Fisher model that incorporates N_{e}: a variant of the standard model with increased biological realism and reduced computational complexity. Journal of Theoretical Biology, 393(??):218–228, March 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316000035.

Zhang:2014:EPD

[ZGWL14] Hui Zhang, Meng Gao, Wenting Wang, and Zhiguang Liu. Evolutionary prisoner’s dilemma game on graphs and social networks with external constraint. Journal of Theoretical Biology, 358(??):122–131, October 7, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931400321X.

Zhang:2011:NCD

[ZGY11] Jiapu Zhang, David Y. Gao, and John Yearwood. A novel canonical dual computational approach for prion AGAAAAGA amyloid fibril molecular modeling. Journal of Theoretical Biology, 284(1):149–157, September 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003237.

Zhang:2010:FFS

[ZGZ^{+}10] Xi Zhang, Chen Guo, Wen Zhang, Huai Cao, Huazhen Xie, Kan Wang, and Ciquan Liu. A folding “framework structure” of Tetrahymena group I intron. Journal of Theoretical Biology, 267(4):495–501, December 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004674.

Zhang:2010:SSS

[Zha10a] Jiapu Zhang. Studies on the structural stability of rabbit prion probed by molecular dynamics simulations of its wildtype and mutants. Journal of Theoretical Biology, 264(1):119– 122, May 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000366.

Zhao:2010:POC

[Zha10b] Gang Zhao. Phase organization of circadian oscillators in extended gate and oscillator models. Journal of Theoretical Biology, 264(2):367–376, May 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310000627.

Zhang:2011:CSS

[Zha11] Jiapu Zhang. Comparison studies of the structural stability of rabbit prion protein with human and mouse prion proteins. Journal of Theoretical Biology, 269(1):88–95, January 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310005540.

Zhang:2015:PBS

[Zha15] Shang-Hong Zhang. Persistence and breakdown of strand symmetry in the human genome. Journal of Theoretical Biology, 370(??):202–204, April 7, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314007115.

Zelik:2014:RSA

[ZHAK14] Karl E. Zelik, Tzu-Wei P. Huang, Peter G. Adamczyk, and Arthur D. Kuo. The role of series ankle elasticity in bipedal walking. Journal of Theoretical Biology, 346(??):75–85, April 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519313005651.

Zhu:2010:SLP

[ZHC10] Wei Zhu, Jingyu Hou, and Yi-Ping Phoebe Chen. Semantic and layered protein function prediction from PPI networks. Journal of Theoretical Biology, 267(2):129–136, November 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310004042.

Zheng:2015:LDP

[Zhe15] Qi Zheng. The Luria–Delbrück protocol is still the most practical. Journal of Theoretical Biology, 386(??):188–190, December 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004336.

Zheng:2016:SLF

[Zhe16] Qi Zheng. A second look at the final number of cells in a fluctuation experiment. Journal of Theoretical Biology, 401(??):54– 63, July 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300558.

Zhou:2011:DLP

[Zho11] Guo-Ping Zhou. The disposition of the LZCC protein residues in wenxiang diagram provides new insights into the protein–protein interaction mechanism. Journal of Theoretical Biology, 284(1):142–148, September 7, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311003055.

Zhu:2011:SMV

[Zhu11] Hao Zhu. Spatiotemporally modulated vestigial gradient by wingless signaling adaptively regulates cell division for precise wing size control. Journal of Theoretical Biology, 268(1):131–140, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005266.

Zaric:2011:NCI

[ZJS11] Bozidarka L. Zarić, Vesna B. Jovanović, and Srdan D. Stojanović. Non-covalent interactions across subunit interfaces in Sm proteins. Journal of Theoretical Biology, 271(1):18–26, February 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310006107.

Zhang:2016:PHP

[ZJSC16] Xiaopan Zhang, Xiong Jiao, Jie Song, and Shan Chang. Prediction of human protein–protein interaction by a domainbased approach. Journal of Theoretical Biology, 396(??):144– 153, May 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316001223.

Zhang:2019:ASF

[ZJY19] Juan Zhang, Zhen Jin, and Yuan Yuan. Assessing the spread of foot and mouth disease in mainland China by dynamical switching model. Journal of Theoretical Biology, 460(??):209–219, January 7, 2019. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519318304636.

Zuo:2010:MTS

[ZK10] Yijun Zuo and Guolian Kang. A mixed two-stage method for detecting interactions in genomewide association studies. Journal of Theoretical Biology, 262(4):576–583, February 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309005189.

Zhang:2018:IAI

[ZK18] Lichao Zhang and Liang Kong. iRSpot-ADPM: Identify recombination spots by incorporating the associated dinucleotide product model into Chou’s pseudo components. Journal of Theoretical Biology, 441(??):1–8, March 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317305702.

Zhang:2016:SCP

[ZKHL16] Lichao Zhang, Liang Kong, Xiaodong Han, and Jinfeng Lv. Structural class prediction of protein using novel feature extraction method from chaos game representation of predicted secondary structure. Journal of Theoretical Biology, 400(??): 1–10, July 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300297.

Zhukov:2019:CAS

[ZKMB19] Oleg A. Zhukov, Tatiana A. Kazakova, Georgy V. Maksimov, and Alexey R. Brazhe. Cost of auditory sharpness: Modelbased estimate of energy use by auditory brainstem “octopus” neurons. Journal of Theoretical Biology, 469(??):137– 147, May 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300311.

Zavodszky:2015:EFP

[ZKP15] Gábor Závodszky, György Károlyi, and György Paál. Emerging fractal patterns in a real 3D cerebral aneurysm. Journal of Theoretical Biology, 368(??):95–101, March 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000053.

Zhao:2015:SWO

[ZKR15] X. Zhao, J. W. Kim, and P. A. Robinson. Slow-wave oscillations in a corticothalamic model of sleep and wake. Journal of Theoretical Biology, 370(??):93–102, April 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000375.

Zhong:2010:AMP

[ZKTK10] Xue Zhong, Jarosław E. Kro·l, Eva M. Top, and Stephen M. Krone. Accounting for mating pair formation in plasmid population dynamics. Journal of Theoretical Biology, 262(4):711–719, February 21, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309004883.

Zhang:2018:PAP

[ZL18a] Shengli Zhang and Yunyun Liang. Predicting apoptosis protein subcellular localization by integrating auto-cross correlation and PSSM into Chou’s PseAAC. Journal of Theoretical Biology, 457(??):163–169, November 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318304302.

Zhou:2018:WFS

[ZL18b] Jian Zhou and Guanghui Lin. Will forest size structure follow the −2 power-law distribution under ideal demographic equilibrium state? Journal of Theoretical Biology, 452(??):17–21, September 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318302364.

Zhao:2013:AIB

[ZLB13] Xiaoqing Zhao, Hong Li, and Tonglaga Bao. Analysis on the interaction between post-spliced introns and corresponding protein coding sequences in ribosomal protein genes. Journal of Theoretical Biology, 328(??):33–42, July 7, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519313000945.

Zienkiewicz:2018:DDM

[ZLB^{+}18] Adam K. Zienkiewicz, Fabrizio Ladu, David A. W. Barton, Maurizio Porfiri, and Mario Di Bernardo. Data-driven modelling of social forces and collective behaviour in zebrafish. Journal of Theoretical Biology, 443(??):39–51, April 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519318300195.

Zhao:2018:MSR

[ZLCH18] Shi Zhao, Yijun Lou, Alice P. Y. Chiu, and Daihai He. Modelling the skip-and-resurgence of Japanese encephalitis epidemics in Hong Kong. Journal of Theoretical Biology, 454(??):1–10, October 7, 2018. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931830256X.

Zhou:2013:PPP

[ZLDZ13] Xuan Zhou, Zhanchao Li, Zong Dai, and Xiaoyong Zou. Predicting promoters by pseudo-trinucleotide compositions based on discrete wavelets transform. Journal of Theoretical Biology, 319(??):1–7, February 21, 2013. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312006133.

Zhou:2012:PNE

[ZLL^{+}12] Junhong Zhou, Dongdong Liu, Xin Li, Jing Ma, Jue Zhang, and Jing Fang. Pink noise: Effect on complexity synchronization of brain activity and sleep consolidation. Journal of Theoretical Biology, 306(??):68–72, August 7, 2012. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519312001798.

Zhang:2017:ECT

[ZLL17] Fengqin Zhang, Jianquan Li, and Jia Li. Epidemic characteristics of two classic SIS models with disease-induced death. Journal of Theoretical Biology, 424(??):73–83, July 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519317301996.

Zhuge:2012:NDR

[ZLM12] Changjing Zhuge, Jinzhi Lei, and Michael C. Mackey. Neutrophil dynamics in response to chemotherapy and g-CSF. Journal of Theoretical Biology, 293(??):111–120, January 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519311005327.

Zheng:2019:UAS

[ZLT^{+}19] Bo Zheng, Xiaoping Liu, Moxun Tang, Zhiyong Xi, and Jianshe Yu. Use of age-stage structural models to seek optimal Wolbachia-infected male mosquito releases for mosquito-borne disease control. Journal of Theoretical Biology, 472(??):95– 109, July 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319301511.

Zhao:2014:ESC

[ZLW14] Lei Zhao, Martin Lascoux, and David Waxman. Exact simulation of conditioned Wright–Fisher models. Journal of Theoretical Biology, 363(??):419–426, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004883.

Zhao:2016:ITC

[ZLW16] Lei Zhao, Martin Lascoux, and David Waxman. An informational transition in conditioned Markov chains: Applied to genetics and evolution. Journal of Theoretical Biology, 402(??):158–170, August 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519316300492.

Zhou:2018:CNG

[ZLW18] Lei Zhou, Aming Li, and Long Wang. Coevolution of nonlinear group interactions and strategies in well-mixed and structured populations. Journal of Theoretical Biology, 440(??):32–41, March 7, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317305374.

Zhao:2019:PPS

[ZLW^{+}19] Wei Zhao, Guang-Ping Li, Jun Wang, Yuan-Ke Zhou, Yang Gao, and Pu-Feng Du. Predicting protein sub-Golgi locations by combining functional domain enrichment scores with pseudo-amino acid compositions. Journal of Theoretical Biology, 473(??):38–43, July 21, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319301754.

Zhang:2013:BET

[ZLY^{+}13] Yijia Zhang, Hongfei Lin, Zhihao Yang, Jian Wang, and Yanpeng Li. Biomolecular event trigger detection using neighborhood hash features. Journal of Theoretical Biology, 318(??):22–28, February 7, 2013. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002251931200570X.

Zhang:2014:IPA

[ZLY14] Shengli Zhang, Yunyun Liang, and Xiguo Yuan. Improving the prediction accuracy of protein structural class: Approached with alternating word frequency and normalized Lempel–Ziv complexity. Journal of Theoretical Biology, 341(??):71–77, January 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004943.

Zheng:2017:SRD

[ZLY^{+}17] Xiu-Deng Zheng, Cong Li, Jie-Ru Yu, Shi-Chang Wang, Song-Jia Fan, Bo-Yu Zhang, and Yi Tao. A simple rule of direct reciprocity leads to the stable coexistence of cooperation and defection in the Prisoner’s dilemma game. Journal of Theoretical Biology, 420(??):12–17, May 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300991.

Zhao:2011:NAC

[ZLZ^{+}11] Xuefeng Zhao, Yi Liu, Wei Zhang, Chong Wang, and Ghassan S. Kassab. A novel arterial constitutive model in a commercial finite element package: Application to balloon angioplasty. Journal of Theoretical Biology, 286(??):92–99, October 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100289X.

Zhang:2015:MMT

[ZLZ15a] Jinhui Zhang, Yong Li, and Xinan Zhang. Mathematical modeling of tuberculosis data of China. Journal of Theoretical Biology, 365(??):159–163, January 21, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314006080.

Zhang:2015:DBS

[ZLZ^{+}15b] Qiang Zhang, Hong Li, Xiaoqing Zhao, Yan Zheng, and Deliang Zhou. Distribution bias of the sequence matching between exons and introns in exon joint and EJC binding region in C. elegans. Journal of Theoretical Biology, 364(??):295–304, January 7, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314005487.

Zhang:2016:APS

[ZLZ^{+}16] Qiang Zhang, Hong Li, Xiaoqing Zhao, Yan Zheng, Hu Meng, Yun Jia, Hui Xue, and Sulin Bo. Analysis on the preference for sequence matching between mRNA sequences and the corresponding introns in ribosomal protein genes. Journal of Theoretical Biology, 392(??):113–121, March 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005780.

Zadeh:2010:CES

[ZM10] Kouroush Sadegh Zadeh and Hubert J. Montas. A class of exact solutions for biomacromolecule diffusion-reaction in live cells. Journal of Theoretical Biology, 264(3):914–933, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931000161X.

Zareh:2019:NAI

[ZMAM19] Masoumeh Zareh, Mohammad Hossein Manshaei, Mehdi Adibi, and Mohammad Ali Montazeri. Neurons and astrocytes interaction in neuronal network: a game-theoretic approach. Journal of Theoretical Biology, 470(??):76–89, June 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519319300931.

Zhou:2018:BSA

[ZMC^{+}18] Da Zhou, Shanjun Mao, Jing Cheng, Kaiyi Chen, Xiaofang Cao, and Jie Hu. A Bayesian statistical analysis of stochastic phenotypic plasticity model of cancer cells. Journal of Theoretical Biology, 454(??):70–79, October 7, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318302789.

Zhuge:2019:OOP

[ZML19] Changjing Zhuge, Michael C. Mackey, and Jinzhi Lei. Origins of oscillation patterns in cyclical thrombocytopenia. Journal of Theoretical Biology, 462(??):432–445, February 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318305770.

Zahiri:2010:NED

[ZMNd^{+}10] J. Zahiri, G. Mahdevar, A. Nowzari-dalini, H. Ahrabian, and M. Sadeghi. A novel efficient dynamic programming algorithm for haplotype block partitioning. Journal of Theoretical Biology, 267(2):164–170, November 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004285.

Zakeri:2011:PPS

[ZMS11] Pooya Zakeri, Behzad Moshiri, and Mehdi Sadeghi. Prediction of protein submitochondria locations based on data fusion of various features of sequences. Journal of Theoretical Biology, 269(1):208–216, January 21, 2011. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005606.

Zhao:2017:NMM

[ZMS17] Lei Zhao, Dong Mi, and Yeqing Sun. A novel multitarget model of radiation-induced cell killing based on the Gaussian distribution. Journal of Theoretical Biology, 420(??):135–143, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317301145.

Zu:2011:AEF

[ZMT11] Jian Zu, Masayasu Mimura, and Yasuhiro Takeuchi. Adaptive evolution of foraging-related traits in a predator–prey community. Journal of Theoretical Biology, 268(1):14–29, January 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310005151.

Zu:2010:EPT

[ZMW10] Jian Zu, Masayasu Mimura, and Joe Yuichiro Wakano. The evolution of phenotypic traits in a predator–prey system subject to Allee effect. Journal of Theoretical Biology, 262(3):528–543, February 7, 2010. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519309005037.

Zheng:2018:MTD

[ZN18] Ting-Ting Zheng and Lin-Fei Nie. Modelling the transmission dynamics of two-strain dengue in the presence awareness and vector control. Journal of Theoretical Biology, 443(??):82–91, April 14, 2018. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318300316.

Zhao:2016:IGS

[ZNA^{+}16] Xiaowei Zhao, Qiao Ning, Meiyue Ai, Haiting Chai, and Guifu Yang. Identification of S-glutathionylation sites in speciesspecific proteins by incorporating five sequence-derived features into the general pseudo-amino acid composition. Journal of Theoretical Biology, 398(??):96–102, June 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001776.

Zhao:2015:PPP

[ZNC^{+}15] Xiaowei Zhao, Qiao Ning, Haiting Chai, Meiyue Ai, and Zhiqiang Ma. PGlcS: Prediction of protein o-GlcNAcylation sites with multiple features and analysis. Journal of Theoretical Biology, 380(??):524–529, September 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315003057.

Zhao:2015:ASI

[ZNCM15] Xiaowei Zhao, Qiao Ning, Haiting Chai, and Zhiqiang Ma. Accurate in silico identification of protein succinylation sites using an iterative semi-supervised learning technique. Journal of Theoretical Biology, 374(??):60–65, June 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315001447.

Zhang:2014:RNR

[ZOG14] Yunjun Zhang, Qi Ouyang, and Zhi Geng. Refining network reconstruction based on functional reliability. Journal of Theoretical Biology, 353(??):170–178, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314001258.

Zoli:2014:TVN

[Zol14] Marco Zoli. Twist versus nonlinear stacking in short DNA molecules. Journal of Theoretical Biology, 354(??):95–104, August 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314001787.

Zahavi:2011:IES

[ZP11] Amotz Zahavi and Marina Perel. The information encoded by the sex steroid hormones testosterone and estrogen: a hypothesis. Journal of Theoretical Biology, 280(1):146–149, July 7, 2011. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002251931100213X.

Zamudio:2019:NET

[ZPdFJ19] Gabriel S. Zamudio, Francisco Prosdocimi, Sávio Torres de Farias, and Marco V. José. A neutral evolution test derived from a theoretical amino acid substitution model. Journal of Theoretical Biology, 467(??):31–38, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300451.

Zhang:2011:ACL

[ZPHS11] Ya-Nan Zhang, Xiao-Yong Pan, Yan Huang, and Hong-Bin Shen. Adaptive compressive learning for prediction of protein–protein interactions from primary sequence. Journal of Theoretical Biology, 283(1):44–52, August 21, 2011.

CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311002657.

Zollner:2015:HHS

[ZPM^{+}15] Alexander M. Zöllner, Jacquelynn M. Pok, Emily J. McWalter, Garry E. Gold, and Ellen Kuhl. On high heels and short muscles: a multiscale model for sarcomere loss in the gastrocnemius muscle. Journal of Theoretical Biology, 365(??):301–310, January 21, 2015. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519314006262.

Zheng:2016:ETD

[ZR16] Xianbing Zheng and Jiusheng Ren. Effects of the threedimensional residual stresses on the mechanical properties of arterial walls. Journal of Theoretical Biology, 393(??):118–126, March 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316000163.

Zhang:2019:PEP

[ZRGW19] Zhaopeng Zhang, Jishou Ruan, Jianzhao Gao, and Fang-Xiang Wu. Predicting essential proteins from proteinprotein interactions using order statistics. Journal of Theoretical Biology, 480(??):274–283, November 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319302577.

Zou:2015:STD

[ZRZ15] Lan Zou, Shigui Ruan, and Weinian Zhang. On the sexual transmission dynamics of hepatitis B virus in China. Journal of Theoretical Biology, 369(??):1–12, March 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315000065.

Zeron:2010:DNF

[ZS10] Eduardo S. Zeron and Moisés Santillán. Distributions for negative-feedback-regulated stochastic gene expression: Dimension reduction and numerical solution of the chemical master equation. Journal of Theoretical Biology, 264(2):377–385, May 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310000639.

Zhang:2012:CCR

[ZS12] Qing Zhang and Hualin Shi. Coupling chromosomal replication to cell growth by the initiator protein DnaA in Escherichia coli. Journal of Theoretical Biology, 314(??):164– 172, December 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519312004675.

Zhuge:2016:CPI

[ZSCL16a] Changjing Zhuge, Xiaojuan Sun, Yingyu Chen, and Jinzhi Lei. Corrigendum to “PDCD5 interacts with p53 and functions as a regulator of p53 dynamics in the DNA damage response” [J. Theor. Biol., 388 (2016) 1–10]. Journal of Theoretical Biology, 396(??):210, May 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519316001636. See [ZSCL16b].

Zhuge:2016:PFR

[ZSCL16b] Changjing Zhuge, Xiaojuan Sun, Yingyu Chen, and Jinzhi Lei. PDCD5 functions as a regulator of p53 dynamics in the DNA damage response. Journal of Theoretical Biology, 388(??):1–10, January 7, 2016. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0022519315004683. See corrigendum [ZSCL16a].

Zhao:2016:NSA

[ZSH^{+}16] Jia Zhao, Ya Shen, Markus Haapasalo, Zhejun Wang, and Qi Wang. A 3D numerical study of antimicrobial persistence in heterogeneous multi-species biofilms. Journal of Theoretical Biology, 392(??):83–98, March 7, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005597.

Zheng:2016:INT

[ZSO16] M. Mocarlo Zheng, Bin Shao, and Qi Ouyang. Identifying network topologies that can generate Turing pattern. Journal of Theoretical Biology, 408(??):88–96, November 7, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316302405.

Zintzaras:2010:SVF

[ZSS10] Elias Zintzaras, Mauro Santos, and Eörs Szathmáry. Selfishness versus functional cooperation in a stochastic protocell model. Journal of Theoretical Biology, 267(4):605–613, December 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310004777.

Zhao:2014:AIT

[ZSSM14] Junsong Zhao, Kimberly D. Siegmund, Darryl Shibata, and Paul Marjoram. Ancestral inference in tumors: How much can we know? Journal of Theoretical Biology, 359(??):136–145, October 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003002.

Zhang:2011:OAR

[ZSW11] Jiapu Zhang, Jie Sun, and Changzhi Wu. Optimal atomicresolution structures of prion AGAAAAGA amyloid fibrils. Journal of Theoretical Biology, 279(1):17–28, June 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519311001068.

Zhou:2014:MCG

[ZSW^{+}14] Linghua Zhou, Yong Shen, Weimiao Wu, Zuoheng Wang, Wei Hou, Sheng Zhu, and Rongling Wu. A model for computing genes governing marital dissolution through sentimental dynamics. Journal of Theoretical Biology, 353(??):24–33, July 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931400085X.

Zhang:2014:PPE

[ZSZM14] Jian Zhang, Pingping Sun, Xiaowei Zhao, and Zhiqiang Ma. PECM: Prediction of extracellular matrix proteins using the concept of Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 363(??):412–418, December 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519314004494.

Zhang:2016:PPI

[ZT16] Shu-Bo Zhang and Qiang-Rong Tang. Protein-protein interaction inference based on semantic similarity of gene ontology terms. Journal of Theoretical Biology, 401(??):30–37, July 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300480.

Zollner:2012:BMG

[ZTK12] Alexander M. Zöllner, Adrian Buganza Tepole, and Ellen Kuhl. On the biomechanics and mechanobiology of growing skin. Journal of Theoretical Biology, 297(??):166–175, March 21, 2012. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519311006461.

Zhang:2018:VTS

[ZTT18] Qianqian Zhang, Biao Tang, and Sanyi Tang. Vaccination threshold size and backward bifurcation of SIR model with state-dependent pulse control. Journal of Theoretical Biology, 455(??):75–85, October 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318303333.

Zhang:2012:NCM

[ZTWL12] Lei Zhang, Fengchun Tian, Shiyuan Wang, and Xiao Liu. A novel coding method for gene mutation correction during protein translation process. Journal of Theoretical Biology, 296(??):33–40, March 7, 2012. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519311006035.

Zucca:2014:PSB

[Zuc14] Fabio Zucca. Persistent and susceptible bacteria with individual deaths. Journal of Theoretical Biology, 343(??):69–78, February 21, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313005353.

Zelnik:2017:DFP

[ZUFM17] Yuval R. Zelnik, Hannes Uecker, Ulrike Feudel, and Ehud Meron. Desertification by front propagation? Journal of Theoretical Biology, 418(??):27–35, April 7, 2017. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317300280.

Zongo:2010:STM

[ZVMB10] Pascal Zongo, Anne-France Viet, Pierre Magal, and Catherine Beaumont. A spatio-temporal model to describe the spread of Salmonella within a laying flock. Journal of Theoretical Biology, 267(4):595–604, December 21, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310005047.

Zhao:2016:IGD

[ZW16] Lei Zhao and David Waxman. The influence of genetic drift on the formation and stability of polymorphisms arising from negative frequency-dependent selection. Journal of Theoretical Biology, 391(??):51–64, February 21, 2016. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315005615.

Zhang:2013:MSU

[ZWC^{+}13] Yanling Zhang, Te Wu, Xiaojie Chen, Guangming Xie, and Long Wang. Mixed strategy under generalized public goods games. Journal of Theoretical Biology, 334(??):52–60, October 7, 2013. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313002269.

Zhou:2010:ESQ

[ZWG10] Da Zhou, Bin Wu, and Hao Ge. Evolutionary stability and quasi-stationary strategy in stochastic evolutionary game dynamics. Journal of Theoretical Biology, 264(3):874–881, June 7, 2010. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519310001414.

Zanghellini:2010:PDB

[ZWvG10] J. Zanghellini, F. Wodlei, and H. H. von Grünberg. Phospholipid demixing and the birth of a lipid droplet. Journal of Theoretical Biology, 264(3):952–961, June 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310001001.

Zhou:2014:MPC

[ZWW14] Da Zhou, Yue Wang, and Bin Wu. A multi-phenotypic cancer model with cell plasticity. Journal of Theoretical Biology, 357(??):35–45, September 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002732.

Zhao:2015:EMG

[ZWY15] Jieliang Zhao, Jianing Wu, and Shaoze Yan. Erection mechanism of glossal hairs during honeybee feeding. Journal of Theoretical Biology, 386(??):62–68, December 7, 2015. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519315004312.

Zou:2010:UIV

[ZXC^{+}10] Xiufen Zou, Xueshuang Xiang, Yan Chen, Tao Peng, Xuelian Luo, and Zishu Pan. Understanding inhibition of viral proteins on type I IFN signaling pathways with modeling and optimization. Journal of Theoretical Biology, 265(4):691–703, August 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310002304.

Zhou:2019:TTP

[ZXH19] Weike Zhou, Yanni Xiao, and Jane M. Heffernan. A twothresholds policy to interrupt transmission of West Nile virus to birds. Journal of Theoretical Biology, 463(??):22–46, February 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519318306052.

Zou:2019:IMA

[ZXS^{+}19] Ren-Chao Zou, Shu-Feng Xiao, Zhi-Tian Shi, Yang Ke, Hao-Ran Tang, Tian-Gen Wu, Zhi-Tang Guo, Fan Ni, Wen-Xing Li, and Lin Wang. Identification of metabolism-associated pathways and genes involved in male and female liver cancer patients. Journal of Theoretical Biology, 480(??):218–228, November 7, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319303236.

Zu:2015:TPI

[ZYD15] Jian Zu, Bo Yuan, and Jianqiang Du. Top predators induce the evolutionary diversification of intermediate predator species. Journal of Theoretical Biology, 387(??):1–12, December 21, 2015. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519315004671.

Zhang:2018:DAS

[ZYJL18] Juping Zhang, Chan Yang, Zhen Jin, and Jia Li. Dynamics analysis of SIR epidemic model with correlation coefficients and clustering coefficient in networks. Journal of Theoretical Biology, 449(??):1–13, July 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002251931830167X.

Zhang:2010:SSC

[ZZ10] Luping Zhang and Gang Zhao. Superiority of single covalent modification in specificity: From deterministic to stochastic viewpoint. Journal of Theoretical Biology, 264(4):1111–1119, June 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519310001840.

Zhang:2014:MDS

[ZZ14a] Jiapu Zhang and Yuanli Zhang. Molecular dynamics studies on the NMR and X-ray structures of rabbit prion proteins. Journal of Theoretical Biology, 342(??):70–82, February 7, 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519313004979.

Zhang:2014:BOH

[ZZ14b] Xuebing Zhang and Hongyong Zhao. Bifurcation and optimal harvesting of a diffusive predator–prey system with delays and interval biological parameters. Journal of Theoretical Biology, 363(??):390–403, December 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314004925.

Zhang:2018:EIE

[ZZ18] Yue Zhang and Qing-Chuan Zheng. Exploring the interactions of EGFR with phosphorylated mig6 by molecular dynamics simulations and MM-PBSA calculations. Journal of Theoretical Biology, 447(??):118–125, June 14, 2018. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519318301449.

Zhang:2010:CES

[ZZC10] Jianlei Zhang, Chunyan Zhang, and Tianguang Chu. Cooperation enhanced by the ‘survival of the fittest’ rule in prisoner’s dilemma games on complex networks. Journal of Theoretical Biology, 267(1):41–47, November 7, 2010. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519310004200.

Zhang:2014:EID

[ZZC14] Wanwei Zhang, Tao Zeng, and Luonan Chen. EdgeMarker: Identifying differentially correlated molecule pairs as edgebiomarkers. Journal of Theoretical Biology, 362(??):35–43, December 2014. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519314003245.

Zhang:2017:PSP

[ZZC^{+}17] Shanxin Zhang, Zhiping Zhou, Xinmeng Chen, Yong Hu, and Lindong Yang. pDHS-SVM: a prediction method for plant DNase I hypersensitive sites based on support vector machine. Journal of Theoretical Biology, 426(??):126–133, August 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519317302527.

Zhang:2017:PFS

[ZZCZ17] Chunyan Zhang, Yuying Zhu, Zengqiang Chen, and Jianlei Zhang. Punishment in the form of shared cost promotes altruism in the cooperative dilemma games. Journal of Theoretical Biology, 420(??):128–134, May 7, 2017. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519317301182.

Zhang:2016:UST

[ZZG^{+}16] Lina Zhang, Chengjin Zhang, Rui Gao, Runtao Yang, and Qing Song. Using the SMOTE technique and hybrid features to predict the types of ion channel-targeted conotoxins. Journal of Theoretical Biology, 403(??):75–84, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519316300686.

Zhang:2014:PPS

[ZZK14] Lichao Zhang, Xiqiang Zhao, and Liang Kong. Predict protein structural class for low-similarity sequences by evolutionary difference information into the general form of Chou’s pseudo amino acid composition. Journal of Theoretical Biology, 355(??):105–110, August 21, 2014. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519314002173.

Zhu:2011:CSC

[ZZL^{+}11] Xiuwei Zhu, Xiaobo Zhou, Michael T. Lewis, Ling Xia, and Stephen Wong. Cancer stem cell, niche and EGFR decide tumor development and treatment response: a biocomputational simulation study. Journal of Theoretical Biology, 269(1):138–149, January 21, 2011. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022519310005448.

Zhang:2019:IIS

[ZZL19] Nan Zhang, Pengcheng Zhao, and Yuguo Li. Increased infection severity in downstream cities in infectious disease transmission and tourists surveillance analysis. Journal of Theoretical Biology, 470(??):20–29, June 7, 2019. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519319300979.

Zhao:2019:IMS

[ZZN^{+}19] Xiaowei Zhao, Ye Zhang, Qiao Ning, Hongrui Zhang, Jinchao Ji, and Minghao Yin. Identifying N^{6}-methyladenosine sites using extreme gradient boosting system optimized by particle swarm optimizer. Journal of Theoretical Biology, 467(??):39– 47, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519319300529.

Zou:2010:MTD

[ZZR10] Lan Zou, Weinian Zhang, and Shigui Ruan. Modeling the transmission dynamics and control of hepatitis B virus in China. Journal of Theoretical Biology, 262(2):330–338, January 21, 2010. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519309004731. See corri-

gendum [ZZR13].

Zou:2013:CMT

[ZZR13] Lan Zou, Weinian Zhang, and Shigui Ruan. Corrigendum to “Modeling the transmission dynamics and control of Hepatitis B virus in China” [J. Theor. Biol. 262 (2010) 330– 338]. Journal of Theoretical Biology, 317(??):425–427, January 21, 2013. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519312005383. See [ZZR10].

Zhang:2011:SRS

[ZZRZ11] Lei Zhang, Ning Zhang, Ji-Shou Ruan, and Tao Zhang. Studies on the rules of β-strand alignment in a protein βsheet structure. Journal of Theoretical Biology, 285(1):69– 76, September 21, 2011. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519311003353.

Zupanc:2019:SCA

[ZZS19] Günther K. H. Zupanc, Frederick B. Zupanc, and Rifat Sipahi. Stochastic cellular automata model of tumorous neurosphere growth: Roles of developmental maturity and cell death. Journal of Theoretical Biology, 467(??):100–110, April 21, 2019. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251931930030X.

Zheng:2016:GRG

[ZZX^{+}16] Gang Zheng, Wei Zhang, Jinfeng Xu, Ao Yuan, Qizhai Li, and Joseph L. Gastwirth. Genetic risks and genetic model specification. Journal of Theoretical Biology, 403(??):68–74, August 21, 2016. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022519316300947.